A Complete Bibliography of Publications in SIAM Review

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

31 May 2024 Version 4.45

Title word cross-reference

#14107 [AR80].

(0, 1) [Sch72a, Sch73b, Ven72, VN73]. (1, −1) [Sch72b, Sch73c]. 1 [SBH^{+}20, SRK99]. 1(mod4) [Sch72b, Sch73c]. 2 [IWMW21]. $25 [BL06]. 3 [Ano79e, LH19, MMH04, TH96]. 30^{∘} [Rei66]. $44.95 [BC96b]. 4^{k} [Sch72a, Sch73b]. 5 [PPR11]. $57.00 [Hig05]. $59.50 [Bor06]. $59.95 [Tre13]. 6 [Wil76b]. 622 [Wil76b]. 0F_{3} [Lar69a]. _{4}F_{3} [Dam86b, Dam87]. A [Zie70]. A = B [PWZK97]. A = tridiag[−12 − 1] [Koh98a]. A_{n} [EH98]. α [BFM^{+}05]. AX = A [Mey69b]. Ax = ΛBx [Mcc78]. C_{0} [Her78b]. C_{1} [Law78]. χ^{2} [JS71, Tre70]. δ [Hol70b]. ΔU = e^{U} [BF96]. ΔU = U^{\frac{n+2}{n−2}} [BF96]. e^{At} [Zie70]. ℓ_{1} [KKBG09]. F [GPP88, Phi87, Goo65a]. f′(x) = F (x, f(g(x))) [And66]. f(x) = x [Hua86]. G [Def76]. Γ [PSV12]. Γ(N) [Kon62]. H [Ask80, Kha88, BS94b]. h^{2} [Pru75]. H^{∞} [Tan93b, Tan93a]. H_{∞} [Bal93]. hp [MMT^{+}21]. ∫_{0}^{∞} f(x) log |\frac{x+1x−1}{|}dx

[Sch59]. ∫_{0}^{∞} f(x) log |\frac{x+1}{x−1} | dx [Sch62a]. J [Hig03]. K [GPP88, GP88, Kei89b, Phi87, Sku05b, Ste59]. L [Han92]. l^{p} [Gev82]. L_{1} [CBS76, Kam78]. L_{2} [CP76, Deb76]. L_{∞} [Deb76]. L_{m} [Def76]. L_{p} [WC77]. L_{p}(1 ≤ p) [Ahm75]. LR [Par64]. M [BH76b, PB74, Sch78a, Min70]. M_{3} [Isb68]. R^{N} [Sim16b]. N [CN05, BL84, Eas93, Ebe75, EK59, Fol64, GPP88, GG89, Gre88, KW76, PRM^{+}66, Sch60b, Xu96c, Xu97, And77, Ban88c, Fin76a, Gla89b, Grz83, JHSB22, Kla76a, Luc71, Mei68, OSS82, Phi87, Ren80a, Sch75b, Sin79a, Smi86, Tij81]. n + 1 [Mei68]. n: F [GP88]. n × n [Bus76b, Jes77b]. O(N^{2}) [Dor75, Dor73]. P [Sch04a, BS94b, Har67b]. Q [CD75, Ask84, Ism85, Wil85c]. QL [PH78]. QR [DD02, Wat82, Wat08, BG96]. QZ [War75]. R^{n} [Kap85a, Sik92]. S [PT07, Gao18]. sℓ(3, R) [GL98]. S ⟶ I ⟶ R ⟶ I [NW97]. SO(3) [Ric85b, Eic84, Ric85a]. S → I → R → I [MY97]. ∑_{ρ=1}^{r} f_{ρ}(A)Xg_{ρ}(B) = C [WZ72]. t [Bla87, Gao18, LS22, Mal88b]. V [Nie76]. ε [Hol70b]. εy′′ − xy′ + y = 0 [Tre20]. x [Cat21]. X = −U(T)U(T)′X [MN76]. Y A = A [Mey69b]. Z [WM63, Wel89]. ||.||_{1} [GHU22]. ||.||_{∞} [GHU22].

* [Cat21].

-1 [Coh75a]. -Approximation [Deb76]. -Beads [CN05]. -Body [Eas93, Gre88]. -Center [Min70]. -City [BL84]. -Coloring [Ano79e]. -curve [Han92]. -D [IWMW21]. -Dimensional [Fin76a, Ren80a, Sch60b, Sch75b, Sin79a, EK59, KW76]. -Distribution [Tre70, JS71]. -Elements [Ebe75]. -Extension [Ask84, Wil85c]. -Extrapolation [Pru75]. -Finger [Goo65a]. -Fold [OSS82, Grz83, BH76b]. -Function [Ask80]. -Hypergeometric [Ism85]. -Laplacian [SBH^{+}20]. -Lemma [PT07]. -Limit [PSV12]. -Matrices [PB74, Sch78a]. -Matrix [Sch04a]. -Optimal [Tan93b, Tan93a]. -Order [Sch72a, Sch73b]. -Orthogonal [Hig03]. -out-of- [GP88, Phi87, GPP88]. -Person [Luc71, Tij81]. -Plane [Kei89b]. -Problems [Kap85a]. -SNE [LS22]. -Solutions [CP76]. -Space [Ban88c, Gla89b, GG89, Xu96c, Xu97]. -Spaces [Gev82]. -Splines [Nie76]. -Subinfinity [Kha88]. -th [And77, Kla76a]. -Torus [JHSB22]. -Transforms [WM63]. -Widths [Smi86]. -wise [PPR11].

0 [BC96b, Bor06, Hig05]. 0-12-558630-2 [BC96b]. 0-19-850763-1 [Bor06]. 0-89871-561-X [Hig05].

1 [Bor06, PN99, Arb78, Cha70a, Cur64b, EK91a, Fle77a, Kap65, Kni94b, KZ93, Rei77b, Tut94, Wei69, Win62, Rad73]. 100-Digit [Hig05]. 12 [Mul64a]. 15th [Jun64]. 16 [Sal63c]. 1724120 [Bel10]. 1750 [Bar91a]. 17th [Sag82]. 1809-1882 [Kaz91b]. 1862 [Jon92c]. 19 [Gud78b]. 1900 [Bar89]. 1927 [Han86a]. 1927-1984 [Han86a]. 1962 [Rei64]. 1969 [Ano70j]. 1970 [Ano71o, Ano71p]. 1971 [Ano72n]. 1972 [Sus79]. 1972-1977 [Sus79]. 1975 [Ano76p]. 1977 [Sus79]. 1978 [AR80]. 1979 [SW80]. 1981 [Bei85]. 1983 [Pip86, Por88]. 1984 [Ano85k, Han86a, Por86a]. 1985 [Mor86c]. 1986 [Ano87l]. 1987 [Ano88j]. 1988 [Ano89m, Coh93b]. 1989 [Ano90m, Dea90b, PC90b]. 1990 [Ano91l]. 1991 [Ano92i]. 1992 [Ano93j]. 1993 [Ano94g]. 1994 [Ano95q]. 1995 [Ano96r]. 19th [Sag82, Van94d]. 19th-Century [Van94d]. 1st [DA78, Han76, Lan76b, MR78, Ric76b].

2 [BC96b, Fle77b, Gue69, Pin72a, Rei78c, Rus95, Sie69, Sto66, Wax91a, Zlo91]. 2-Dimensional [Gas76a, YF78]. 2-Inverses [Zlo91]. 2. [Kap90b]. 20 [AR80, Ano87b]. 2003 [ADPR04]. 21 [SW80]. 21st [HF05, Sch93d]. 27 [Mor86c]. 2Nd [BE78, Bus76a, GC78, HW76, Bir64, Che86, Elc76, HK92a, Hoc79a, Kup75, Lee77, Ser80a, Sop82, Wei77, Whi80, Yan94]. 2nd-Order [HK92a].

3-Dimensional [Kos79a]. 30 [Coh93b]. 31 [Dea90b, PC90b]. 34 [Rob94b]. 35 [Win96b]. 36 [For96]. 370 [Zwe16]. 39 [Mac98a]. 3rd [FGD^{+}92, Hoc80]. 3rd-Order [FGD^{+}92].

4 [Ock92a]. 4An [MG95]. 4F3 [Rah87a].

4th [Man76].

5 [Noh62]. 58 [AR80, Fir97].

6 [Car08]. 63-2 [New65b].

76-17 [Hol78].

80h [SW80]. 82-17 [Hen82]. 85 [Ano87b]. 85-20 [Ano87b]. 86f [Mor86c]. 89g [Coh93b].

9 [Tre13]. 90a [Lax89]. 90g [PC90b]. 90k [Dea90b]. 93-17 [Wol94b]. 93i [Rob94b]. 94g [Win96b]. 95-18 [Auc97b]. 96-1 [Ada97b]. 96-2 [RT97]. 96-3 [BPV97b]. 96-4 [Gla97b]. 97-1 [AB97]. 97-2 [Kri97]. 97-3 [Con97]. 97-4 [Wan97]. 97-5 [GV97]. 978 [Tre13]. 978-3-0348-0179-9 [Tre13].

= [Boo94].

A. [Eva84a, Fin84a, Her84a, Joh83b, Sle84a, Tis83a]. A.M.S [Jun64]. Åke [Jac76, Kor85]. Aarts [Ham90, Ker91a, Ker91b]. Abaffy [Ari91a, Ari91b]. Abarbanel [Por86a]. Abbott [Car87]. Abdelkader [For92, Lew91b, Loc84b, Osb85]. Abdul [Coc86]. Abel [JB88, Boh88b, Deu87]. Abelian [Mil81a]. Ablation [Mun61, Kla63a]. Ablowitz [Sco83]. Abraham [Car81, Dav91a, Lee84b, Man81, Rob81, WP88, You71]. ABS [Ari91a, Ari91b]. Abscissa [JS71, Tre70]. Absolute [Asn75, BS67, CS70, Ham69, Sch61, Wal66, Kal70]. Absorbed [Por64]. Absorbers [ABJK10]. Absorbing [HS70, IWMW21]. Absorption [Epp86, HS70, Spa62b]. Absorptivity [Fet66, Wil64b]. Abstract [Aiz94a, Bir86, Kah66, Lie97, Per93a, Per93b, Tod84, Aiz94b, Pin85a]. Abstracts [Ano64i]. Abtastsysteme [Hig62a].

Academic [BC96b, Lin76]. Accelerated [FR16, FP96, Ger94, HYS76, JGM95, Jor95, Kla94b, Kla95b, RJ96, Jor96]. Accelerating [AW76, GL04]. Acceleration [BRZS18]. Acceptable [Bra76b]. Acceptance [Flu90]. Acceptance-Rejection [Flu90]. Access [Erd92b, Erd92a, Jev91, Jev92a, Jev92b, Jev93, Jev95a]. Accident [Dic75, TT76]. Accompany [Mil89b]. According [Man61c]. Accounting [Tod74]. Accumulation [VCKJ16]. Accuracy [For21, GAT75, HM78, Hig05, Ike76, Kab93, Ste97b]. Accurate [Ole94]. Accurately [NF16]. Achenbach [Whe74]. Acid [DBS^{+}07]. Acids [Rei83b]. Ackoff [Sal62]. Acoustic [Arn70, Col96, Col84, CCM00, CCM01, KK75, MWWC96, NY90, Sle94]. Acoustics [CB76, Maj91, Nor90]. Action [BLS92, BBLS92, Cen81, Grc18, New91b, Pro84a, Ral93a, Ral93b, Bry94, O’R92, Pro84b]. Activator [Ada91]. Active [Kel75b, Tur04]. Activities [Lin76]. Activity [CSW^{+}08]. Acton [Dan71, Mon60]. ACTORS [Kow88]. Actuarial [Nas87a]. Aczél [Cam77]. Adam [Pit90b, Pit90a]. Adamek [Gha94]. Adams [Har78]. Adaptation [Sam76a]. Adapted [WC96, Chu96]. Adaptive [Aok70, BH03, Bau88, Ber07e, BFM^{+}05, GJRW24, KW76, McC95, MNS02, PvySF23, Pug66a, RFS03, RM95, Wei91, Bau91, Dup93, Ham62a, Hüb91, Lai78, LeV93]. Addendum [Dea90b, SW80]. Adder [Jev93, Jev95a]. Adding [GAK17]. Addition [Dea90a, JRZ97, Sal80]. additions [AR80]. Additive [McC92a]. Adelmeyer [Van94c, Van94b]. Adel’son [GG91]. Adel’son-Vel’ski [GG91]. Adi [Eck82, Mey76c]. Adiabatic [AvK^{+}08, Coo91]. Adjacency [Har62, Mur71]. Adjoint [CHT21, HK92a, MS96a, NS24, PG00, SS16, HK92b, Svo96]. Adjustment [Lew83c]. Admit [Ben70]. Adrian [Ber97a]. Adrianus [Shi97].

Adsorbing [Tre78]. Advanced [And90, Gut97, Mor97]. Advances [AMO91, CKCH23, Hu68, LaS64, Lef65, Sal63b, VBVM22, Kau86, McG67a]. Advection [FL18, GSYY22, LW13, Mic88]. Advertising [Set77]. Aerodynamics [Hel75, LRJ78, Sul93b, Che89, Hol66, Kan62b, Sul93a]. Aeroelasticity [Ash89]. Aerofoil [Sie61]. Affairs [Sel76]. Affine [Ber92a, JP89]. Affirmative [BLS92, BBLS92, New91b]. Affirmative-Action [BLS92]. Afternotes [Kin97]. AG [Tre13]. Agarwal [Haj96]. Age [KA89, WR96b, Ran96b]. Aggarwal [Bou76a]. Aggregate [Dau90]. Aggregation [BT13, BFM^{+}05, Dau90, Saa95]. Agha [Kow88]. AGM [TBDS92, Wim88, BDS92, CV91]. Agricultural [PM87]. Agriculture [HHSS75]. Aharon [Eck82]. Ahlberg [Bir69a]. Ahlfors [Hoc80]. Ahmed [Ann95b, Don90, Gro93a, Neu94, Sac83b, Sac83a]. Ahuja [Bal95, Hun84]. Aid [Gol62]. Aided [BP76b, Bos95, DeR91a, DeR91b, Duf86, HLB95]. Aids [Tud92]. Aigner [Can91a, Can91b]. Aiken [Jel87]. Aim [BH76a, SS61]. Aimo [AKP93a, AKP93b]. Air [Mul79, WM61a, WM61b]. Airborne [BCG77]. Aircraft [DeS88, Mil98, Sed91, Wol94c, For96]. Airfoil [Che92b, GM66, Mac68a, Che92c]. Airlines [Zwe16]. Airplane [Som79]. Airy [Gib73, GdMGR91, Eva90, FO99, Gro91b, Smi72]. Aisermann [Har72b]. Aizerman [Kal70]. Ajit [Spu89]. Akhiezer [Fle63a, Ral62b]. Akilov [Sop69]. Akoev [Hol89a]. Alai [Mat94b]. Alain [Ber85b, Fle89, Raz87, Mat94a]. Alamos [Buz78]. Alan [Als92a, Amo64, Bla83, Bla93, DRWK64, Duf84, Hey88, Hig06, Hop79, Liu88, Lus87, Man90a, Man90b, Man89, MS65, Mur94a, Ros87, Saa96, Sha92a, Spi84b, Als92b, Mur94b, Sha92b]. Alarm [Bur78b]. Albada [FZ63]. Albert [Cot94, Fan82, Mei89, Whi80, Zlo91]. Albeverio [Bro90]. Albrecht [Ork95]. Alcuin [CHW12]. Aldo [Jer81, Tau80]. Aldous [Sie90]. Alefeld [Moo85]. Aleksandr [Nas77b, Pin70]. Aleksandrov [Ham65b]. Alekseev [Hol89a]. Aleksei [Lev88b]. Aleksey [Har72b]. Alexander [Baz73, Bir64, Bla88d, Bla89a, Bol91b, Bur94a, Fir92b, Fir92a, HPF85, Lev63a, Maw96, Mor82, Spä97, Tis83b, Bol91d]. Alfred [Cer96]. Alfredo [Stu90]. Algebra [Bak66, BDO95, Bjø90, Bre62b, BL06, Cha83b, Che86, Cia92a, Cia92b, DD02, DGK84, DW95, For67, GPS90, GG99, Gus76, Hei95, Hel78, Hig99, Hig00, Joh81, Kla84j, Kla84c, KZ93, Leh65, LHFE97, LW17, Mal80, May60, MW95, Mit91a, Nas90, NdA14, Pin85a, Ral93b, RLGBRMC06, Ros82, Sch88, Sto76, Stu94a, Wat86c, Wil81, Wit22, Bir86, Bre62a, Bre64a, Kee88a, Mit91b, Mul67, Par72, Ral93a, Row79, Rus60, Som93, Stu94b, Win95]. Algebra-Symbolic [Che86]. Algebraic [Ale98, BKST15, Cam91b, Cam91c, Cha83b, Che67a, Che86, Chi90, Dwa64, GM87, Gra66b, Gre97, Hal86, HA71, Joh68, Kap70a, KJ87, KJH87, LRM96, Ost61b, Par66, Rud61, Sag86, SS98, Som93, Str77, TTM05, Zie78, Boo79, Bou77, Gos84, Meh96, Mul69, Wol93, You71, Zak85]. Algebraic-Geometric [Wol93]. Algebraically [CD75]. Algebraische [Ham85]. Algebras [BKS66, BKS68, BK69, Gru74, Gru88, Gui92, Tu92, Vog93]. Algèbre [Ost61a]. ALGOL [Din72, Car66c]. Algorithm [AW76, Aro78b, Ars97, BCF95, BS77, BL84, BJK11, BH90, CG61, CC76, Cop07, Cor77, CD75, Fra75, Fra95a, FW62, GL89, GMS05, GGK96, Haw95, Hor78, Hou98b, Hou74a, KW76, KB76b, Kli78, LeV93, Mac98a, MS75, MR18, MM95, McA87, Min78, Naz76, OHBG00, PR91, Pap85, PH78, Poh78, PGS75, RW84, She09, Shr76, Sou76, Ste61a, SPR76, Swa77, Van78, Vis01, VDC22, War75, Wat82, Wat08, Wit75, Wol65, WPTZ01, BG96, DGH^{+}18, Hou98a, Ort96a, Ort96b]. Algorithmen [Alf93]. Algorithmes [Dan72]. Algorithmic [Ant92c, Ant92d, CJR18, Hig01a, Mil83b, MW95, Nas77b, Pal94, PLC96, Man89, Mil83c, Win95]. Algorithmiques [Ham87]. Algorithms [AK90, Ari91a, Ari91b, Asp84a, Aur18, BH03, Bar74, BG76, BB96, BPE96, Bla88a, Bla89a, Cho87, CKCH23, Dev90, DGK84, DFG99, Duf69, Dup93, EGJK04, EMUP98, FG04, Fat93, Fla97, FLGM23, Fre82, GPS90, GG80, Gar94a, GG91, GO89, Gra72a, GDR78, HMT11, HNP91, Hel78, Hil75, Iba73, Joh83e, JS20, Kel92a, Law18, Len93b, Len93c, LT13, MS76, Mat85, McG67b, Pac88c, Par71, Pow86, Ral93a, Ral93b, RW80, Sho99, Som93, ST91b, Tar78, TK82, TB76, Van93, Wan83, Whi80, Wic94a, Wic94b, WS13, Won83b, Won83a, Ye94, ST91a, Ste97b, Asp84b, Bak91a, Bur81b, Cha86, Dob87, Ebe96, Fra89, Fro80, GB89, Har72b, Hüb88, Joh83f, Kv93, Kri91, Mac69, Mac70, Pac86, Pei83, Whi81, Whi83, vdV94, Dav77, Fro87a, Kel92b, Tha79]. Algoritmij [GG91]. Ali [Col76a, Hol86a, How82, Ran96a]. Alignment [ZZ09]. Aliprantis [Art92a, Art92b]. Alistair [Ran89, Wei88]. All-Pairs [Wil21]. All-Pass [Mak86]. All-Terminal [Kar01]. Allan [Sch78c, Sie89, Sle89a, Smi86]. Allele [Cop07]. Allen [Ban88b, Fia73, Les76, Rob74, Wou67, Bra81]. Alley [Hig01b]. Allgower [Sai91a, Sai91b]. Allina [Dei92]. Allocation [Bel65, Ber60, Cof83, Coo64, FSV06, McC91b, She72, Saa88, Wet75]. Allocations [Par80]. Alloys [AR63]. Almeida [Stu90]. Almost [FG04, Fin72, Lax76, Mas92, PE76, Sei65, Sam76c]. Almost-Sure [PE76]. Alois [Han81a]. Along [Mor76]. Alpert [Jos95]. Also [Ste59]. Alston [Fra72, Gre72, Var66]. Alternate [Dec65a, Hey63]. Alternating [BD93, BdD93, Den76, Hen92]. Alternative [Ang92a, Bel63d, Edw00, McC92b]. Alternatives [Dub76, Pru78]. Altitude [DeS88, Sed91]. Altman [Pic87a]. Altruism [Lud82]. Alvin [Shu82]. Amann [Eas91, Lie97]. Amateurs [Bra91b, Bra91a]. Ambartzumian [Mec92]. AMCD [Sin76]. Amedeo [MN83]. Amer [Lax89]. America [Lax89]. American [Saw59, VP09]. Ames [Bar70, Blu91, Fer79]. Amir [Din94a, Din94b]. Amos [SLF66]. Amount [Str76a]. Ampere [Pin84b, Urb93a, Pin84a, Urb93b]. Amperometric [BM03]. Amplification [CM89]. Amplitude [CM89, Har66b, LM90, Mak86, MB76]. Amplitudes [Bon67]. AMS [GM96, Mey96]. AMS-LATEX [GM96, Mey96]. Amson [Ker92]. Analogous [Bel69]. Analogs [SS62]. Analogue [WA63]. Analyse [BI68b]. Analyses [HFKG06, JS94, Lig86, Lin86, WMZ94, Rog60]. Analysing [Sav18]. Analysis [Ahl61, AA78, AW02, And78, Aro78a, Art91b, Ask72, Asp84b, Asp84a, BK95, BJ75, Beh76, Ber62, Ber65a, Ber91a, Ber85b, Bes92, Bla88a, Boa87, Boy90a, Boy90b, Bra82, Bry90b, But78, Cab75, Car87, CCA08, CE89, Che75, CH68, Chu89, Chu96, CLH97, Cla78c, Cle63, Cru86, Den76, Dim91a, Dim91b, DZH^{+}04, DBS^{+}07, Dob87, Dru64, DGP92, DGLZ12, Duf86, Dwy63, Ell85a, Ell92a, Ell92b, Els89, Els91, Enr89, Epp86, EKW96, Far85, FG19, Feu76, Fla97, Fle76, FGA97, Fre75, Gei89, Ger96, Gho87, Gil90a, Gil90b, Gon06, Gor75, Gra72a, Gra66a, Gra93b, Grc11, GWZ16, Gue79, Gue83c, Hab83a, HH16, Hag82, HHSS75, Ham73b, Ham73c, Han92, HT78, Hau86, Hau88, Hig06, Hin87, Hin91, HU07, Hoc80, Hop75, HG96, Hut89, IL97, JHSB22, Jor87b, Jor89b]. Analysis [Jow72, Joy71, JS20, KP87, KR98, Kv93, Kaz91a, Kel88a, Ken86a, Ker92, Kir10, KR78, KSW96, Koh88, Kra67, Kre78, KP06, Kui06, Lae83a, LLSN96, LM90, Lee84b, Lem85, Lin92, Luc87, Luc10, LRV76, Mac91b, Mac91d, Mal87, Man87, Man88b, Man63, MS96a, Maw91, McC92b, MG96, Mei19, Mer89, Mir71, Mit65, Moo77b, Mor81, Nas77b, Net20, Nic96, Nie89a, Nie94b, O’M10e, Par71, Par78, Par76, PN99, Pin84b, Pin84a, Pit90a, Rab68a, RCC78, Ral62a, Raz92a, Raz92b, Rei77a, Rei87, Ric67a, Ric88, Ros76a, Sae73a, SZ96, Sal63d, Sch89a, Sch70a, Sch78d, SHL^{+}79, Sco91a, Sco91b, Sco94, She09, Sik92, Sin88, Sin93a, Sku04, Sle89b, Sle67, Sma90, Smi75, Smo91a, Smo91b, Sno76, Sop69, Sop79, Sop82, Spi84b, Spi84a, Spi74, Sta75, Ste01, Ste62, Str89b, Svo96]. Analysis [Swa77, Tad87, TF96, Ter82, Tod75, Ton88, Ton89, Tre84b, Tre84a, Tri78, Tul90, Tur04, Ula64, Wah03, Wal97, Wet74a, WC96, Wil85a, Wil86a, Wil71, Wil77b, Wim96, WS23, Yak94, Yau68, You73, ZM20, Zwe16, Hen79, Sle83b, Daa97, Hig02b, Kin97, Neu97b, Aga94, Ans64c, Ans80, Arb77, Art91a, Bau88, Ber85a, Ber92b, Bla88c, Bla86b, Blu65, Boa76, Boa78, Boa81, Boa82b, Boa84, Boa85, Boa88, Bry90a, Cab74, Eas91, Eps59a, Far96, Fie93, Fis76, Fri59, Gol65b, Gra93a, Gri88, Gup96, Hab83b, Hag86, Ham65a, Han67, Han81b, Her77, Kel86b, Kel90b, Kin82, Kle71, Kol63, Kul60, Lae83b, Lee82b, Lei66, Man85, McK81, Mei77, Moo63, Ner70, Olk59, Par69, Par67a, Ple88, Rhe60, Riv60a, Sch82a, Sch86a, Sle89a, Sop72, Ste81b, Ste93b, Thi63b, Var66, Wat63]. Analysis [Wat65, Whi76b, Wil60a, Wil65b, Wou67, Zha96, vV65, Gea74, Hag83, Jow76, Kor95b, Moc85, Mon60, Moo69, Old62, Pin85b, Pit90b, Pol59]. Analysis-Cauchy [Boa87]. Analysis-Correspondence [Ber85b]. Analytic [BP94, DA79a, DeM65, Dev94b, Haw95, Hoc80, McA94c, McA94a, McA94b, Ort96a, PTK11, Rou97b, Sha64b, Shi68, Sim96, TK97, Van75, Wal70, Wim88, Rou97a, Coh82, Ort96b, Wal63a, Whi67]. Analytical [Col91, Cra89, DF98, Ghi76, Kla84j, Kla84c, McL86a, NC76, NBR96, Ran96a, RMT^{+}03, Sar10, Sha89b, Sto76, Str77, TeB97b, XC11, TeB97a, Pol69]. Analyzing [GN22, Zie70]. Anastassiou [Ath94]. Anatol [Coz85]. Anatole [Swa96]. Anatoliy [Hoc83b]. anatomy [DGH^{+}18]. Ancient [RR97b, RR97a]. Anderberg [Che75]. Anders [Bar91a, Bar91b]. Andersen [McK95]. Anderson [Abl80, Bau88, Feu76, Moo77a, Rob94a, vL86a, BRZS18, SBR08]. André [Koe95]. Andreas [Fat93]. Andrew [Hur93a, Par92b, Ran91a, Ran91b, You72b, Hur93b, Ran91b]. Andrews [Bre81, Bur79, Jor84b]. Andronov [Mey75b]. Andrzej [Gre88, Guc87]. Andy [Ren93d]. Angel [Dor74]. Angeles [Fre90]. Angelo [Hin89]. Angle [Ano90a, EH16, Hof71, Kla90b, MV93]. Angular [KA76, Sim71a, Ant72]. Angus [Sop82]. Anile [Rau95]. Animal [Tud90]. Animated [Dev92]. Animation [Odg86]. Anisotropic [Joh67, Rog69b]. Ann [Sha83]. Annealing [Ham90, Ker91a, Ker91b, Ham90]. Annie [GM85]. Announcement [Ano75a, Ano75b, Ano75c, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano80a, Ano80b, Ano80c, Ano81a, Ano82a]. Annulus [DEP01, Quo75]. Anomaly [CD93]. Anosov [Sib95]. Ansélone [Moo76]. Anshel [Ull88]. Antanas [AKP93a, AKP93b]. Antenna [Str75]. Antes [Shi93]. Anthony [Boj93, Hul61b, Jag72a, Jag87b, Ken93, Lem85, Lin79c, Man70, Mil81b, Smi83]. Anthropology [Sei85]. Anti [HW77, Hor95a]. Anti-Ballistic [HW77]. Anti-plane [Hor95a]. Antiderivative [Str71, Jag72a]. Antimirov [Mey94a]. Antiviral [GAM93]. Antman [Ren95]. Anton [Mor86a]. Antoni [Var87]. António [Ell87]. Anything [Gea81]. Aoki [Lai91, Mor71, Tze91]. Aperture [Che01, Che18]. Aplevich [Hei93a]. Appearing [Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d]. Appell [Lan76b]. Appendix [Lev65]. Appledorn [EB88]. Application [AN83, Beh76, Bel65, BM93, Bir65, BD63, CH66, Dec65b, Dri75, EFMM13, Flo69, Goo91b, Gre59, Ham77, Hau86, Hel75, HSŠ16, HM01, Hir75, Hof71, Joh84b, Joh84a, Ken86a, Kri76, Lan76b, LB10, LW13, Mac68a, MTGF78, Mer64, Mil76, MM10, Mor78, Mou67, New65a, Orb18, PC61, PR70, Pro65a, RCC78, RLGBRMC06, San65, Sch75a, Sle76, Sou76, Spa62b, Tuc73, Wou64b, YS76, BI68a, Dri65, Fat93, Kel88a, Mor89c, O’M86, Tek89, Zlo91, Spi87]. Applications [Aga93, Aiz94a, Ame86a, And74c, Art87, Ask79a, Ask80, Ask92b, BS91, BCF95, BBN98, BT69, Bar91a, Bar91b, Bei78, BKS66, BKS68, BK69, BK17, BBR13, Ber65b, BBC11, Bla66, BGH^{+}04, Boe88a, Boj93, BCZ21, BT20, Car87, CCCM06, CCA08, Cen81, Cha70a, Cha81a, Cha81b, CFG95, Cla83a, Cla83b, Col94b, Coo91, DM67, Din94b, DPS73, DFG99, Dum96, Dys66, Eas89, Eck80, Edw90, Elc76, Ell87, Enr84, FLM22, FT76, FP97a, Fir97, Fle77a, Fle77b, Fle82, Fle84a, FGA97, Fos90, Fra86c, Fre86c, FSY65, Gar94a, Geo71, Gil90b, GF78, Gol78, Gon06, GAK17, Gra00, Gra86b, Gre60c, Gru74, Gru88, Hab83a, Hac75, Han81b, Hau90, Hel97, HK93, Hoc79a, Hor95b, JSF95, Joh87, Joh91a, Joh91b, Jor87b, Jor89b, KS95, Kaz82, Kei89a, Kei89b, Kel60b]. Applications [Kel84b, Kel84a, Kes69, KN61b, Kok84, KB09, Kre78, Kup75, Kus19, Kus76, Lak80, Lan86, Law95, Leh63a, Lev97, LLMM14, Lon77, Luc10, MYDF08, Mac00, Mac76b, Mac98b, Mag86, MV68, Mar60, Mat94a, McG67b, Mey76c, Mey82a, MXPW11, MPW23, MD96, MBK97, Moo78, Mul75, Neu94, O’M10f, Obe64a, Ode89, PSSW06, Pay67, Pic87a, PFM76, Pin72a, PD97, PC95, Rab68a, RM89, Ric67a, Ros76b, RM96, Rub77a, SS79, Sel94, Sha95a, SST95, SAH96, Sie69, SS95, Sul93a, Sul93b, Syn93b, Thr94b, TD20, TFL23, Tot91, Van94b, Var67, Wan91, Web85, WL95, Whi63b, Whi92b, WP96, WM15, Win62, Xu21, Bel10, Ham90, Hey97, Wic94b, Abl80, Aga95a, Aiz94b, AK90, Alp85, Ars62a, Ask78, Ask92a, Aur18, BR67, BR76, Bir69a, But89, Cha94c, Che75, Chu88b]. Applications [Coc86, Cor95, Den78, Din94a, Dor89, Eva81, Fan82, Fin95, Fle84b, Fra95a, Fro74, Gil90a, Gog85, Goo72, Gos79, GM85, Guc79, Gue83b, Gue85, Gue88, Haj96, Ham79a, Hoc92, Hom86, Hun81, Ism85, Joh88, Kar78, Koh86, Lau74, Law18, Lee84a, Lee84b, Lor91, Luk82, Mal88a, Man88a, Man81, Mar87, Mat94b, McA96b, McC91a, McG61, Mey79, Muk86, O’M88b, Orl93a, Pal96, Par74, Pic90, Pin87b, Pol65, Red83, Ros82, Rox84, Sat78, Sav18, Sch89a, Sha89b, Shu85b, Sie89, Sle81, Sop79, Sto81, Stu91, Syn93a, Sys63, Tak83, Tap86, Thr94a, Ton95, Van94c, Wal79, Wal93a, WP88, Wel89, Whi92a, Wic94a, Wim65, Yau68, Arb77, Bir86, Hab83b, Har64a, Kap68, Kel86b, Kes65, Kum88, Mar64]. Applications-Oriented [Geo71]. Applications. [Red90]. Applied [Ano72o, Ano73n, Ano75d, Ano75o, Ano77q, Ano78e, Ano78f, Are89, Ask79b, Bla81, CCR^{+}62, Car67e, CO07, Det85, Dwa85, FKL^{+}06, Fis76, FM18, GBF78, Gor59, Gue79, Hai85, Ham76, Hen86, Hig00, Hig02c, HH19, Hoc79b, Hoc82, HG96, Jun64, Kap71, Kap76, Kaz91b, Lae83a, Lat69, Lem83a, LCK^{+}17, Lin67, Lin76, Lin78a, Luc86, Mac68a, Man82, Mey82b, NBR96, O’M10e, Pan93a, Pan93b, Pap62b, Pin92a, RFS03, Ros67a, Sag83, Sal63c, Sch80a, Spa77a, SKK20, Tho77a, Wei69, Wig85, Wol97, Lax89, Ram88, Tre15, Ula67, Dur97, Boa76, Boa78, Boa87, Bro77, Cha64a, Dav91b, Fel72, Gue83b, Gup96, Har73, Hau88, Hut89, Jer81, Lae83b, Lee86a, Lem83b, Lin92, Liu88, Mey94a, Mul67, Pin92b, Por83, Ran96a, Sch82a, Seg80, Sle74, Sle89a]. Applying [Gue83a, Pen95, Pon83, Gue83d]. Appointed [Ano80e, Ano80f, Ano81b, Ano82g, Ano83e, Ano84e, Ano85a, Ano86a, Ano86b, Ano86c, Ano87c, Ano88a, Ano89c, Ano90b, Ano91c, Ano92c, Ano93e, Ano94a, Ano95e, Ano96f]. Apportionment [BY78]. Appraisal [HMH00]. Approach [Ars97, BT69, BEKS17, Bis78, Bom02, BEG23, BGL01, CM76, CN15, Cum92a, Cum92b, Don85, EW75, EKW96, Faz98, Gar00, GHU22, HFS01, Hei95, HUS10, HS15, Ike76, KL76, KL97b, Ker91a, Ker91b, KM85, KD76, KM86, LW13, Mac98a, MG76, McA92, McM61a, MTM96, MMB^{+}07b, MT96, Mor75b, Mor85b, Mro91, Nas77b, Olv78, Pal94, PBL98, Pic87a, Red90, Sch78a, Sch78b, Ser80b, Shi93, Spa62b, Sta92c, Ste62, Stu94b, Stu94a, TTY88, VSM14, VC06, VDC22, WPL98, Wit90, KL97a, MMB^{+}07a, Dur97, Sha97a, Ber88a, Cor92b, Cum91, Dav91b, Fle76, Ham88, Hey89, Kal63, Kap64, Kha86, Kha87, Mar88, McA96a, Mor83, Mul66b, Pin85a, Pro85, Sta92b, Was94a, Wim96]. Approaches [Bel63d, Cul68, MC21, Naz80, Rei69, Smi74, Zlo83]. Approval [Hof85]. Approximants [GM85, McC93b, McC93a]. Approximate [BE78, FC99, Gon06, GM66, HMT11, HLS09, LL75, Lew78a, Nas77b, War98, Riv60a, Ste73b, Wou68a]. Approximated [Zua05]. Approximating [AG80, Ang92b, GDR78, KM85, LSY16, McG82, Mey75a, Pit74]. Approximation [AH19, Ahm75, All87, Ans85, ADR76, Ask74, Ask76, Ath94, Ban88a, BBKV96, Boj97, Byr65, CS06, CBG75, Cha81a, Cha68, CS63, Chu81, Cod70, Deb76, DM67, Deu90, DGP92, DGLZ12, Eub88, Ewi67, Fab80, FG19, Fet68b, Fra75, GK91, GRL98, GGT13, Hau78, Hol81, Hol70a, Hua86, Jer95, Kam78, KL75, Kar01, KF06, Las07, Law72, Lew73, Man76, Mer64, Moo76, Mou67, New65a, New62b, New62c, New62d, PTK11, Pro83b, Pro83a, Pru75, Ray69, Roo10, Rot62, Rub78, RBZ23, Sch78c, Ser78, Smi86, Ste90a, Tre17, WM63, Wim81, ZWGB16, Cor16, FK06, Cor16, Ask72, Ban88b, Cav90, FR63, Gre65, Jor77a, Kin76, Kor85, Moo60, Por85, Row86, Seg06, Tot91, Abb92, Ask81, Che65, Che67b, Chu82, Fix74, Gra86b, Jet91, Sho87, Vai96]. Approximations [DdSG^{+}22, DG20, Eck77, FLM22, Fle79, Fro87a, FH76, Ful76, GT10, GNZ20, Her91, Her78b, JR96, Jok66b, Jok66a, KTSB21, LT13, Mor76, Olv80, PG00, Red76a, Sie90, Tad87, Tre84b, Tre84a, Cod78, Dup93, Olv72, Rei96, Wah92]. April [Ano73d]. Apt [Mit93, Orl93b]. Aragnol [Mar85c]. Arbib [Kap70b, Mor75b]. Arbitrage [Ell84]. Arbitrary [All78, BP76b, Dec65a, Fat76, Kat76, Kat78, Ung77, Wil59, Zlo70]. Arc [Yos60]. Archimedean [CS75]. Architecture [Sco92]. Architectures [And90]. Area [Fee00, Gia63, PD89]. Areas [Gra18, Mel64, BI70a]. Argument [Gau62, Sch76a, Seb75, Noh77]. Arguments [Dri75, Wou68b]. Aris [Coh76, Gun63, Seg79]. Arise [ZWY14]. Arising [Ada97b, Ano92a, ABG96, BG95a, Bil93b, Bil94b, Boe91, BdD94, Boe96b, Bro60a, Car64, CR96, CZ92, Geo76, Gla93b, Gla93d, Gla94c, Gla95d, Gla95e, GM66, Hel75, Hoc66, Ism96b, JH92, Lar94, LRJ78, Ric89, Sle62, Ver90, VBdD91, Ada97a, AH98, Boe98a, Ism96a]. Aristid [Ant92c, Ant92d]. Aristide [Con96]. Arithmetic [BB83a, BB84b, BB84a, Car70c, Erd77, Fat76, Flo69, GBF78, Hig02c, Jac69, Jun64, KM86, LS63, MW76a, MNC71, Neu96, New76d, Olv78, Ral83b, Ral83a, SB72, Zie76, Sei97, Hoc82]. Arithmetic-Geometric [BB83a, BB84b, BB84a]. Arithmetics [Yoh79]. Arkin [Car67b]. Armand [Coo67a]. Armbruster [Kee88a]. Armed [McC91b]. Arms [Kre66]. Arnaud [Mey79]. Arndt [Hey85]. Arne [Sal85]. Arnljot [Mee96]. Arnol’d [Kaz91a, Ber85e, BR76, Con85, Hig02b, Hor93, Jor77a, Jow86, Mey79, Mey89b, KL08]. Arnoldi [BNT21]. Arrangement [FS62, Fri64]. Arrangements [MYDF08]. Arrays [Str75]. Arrival [Giv63]. Arrow [Bas88, Man61d, Ric67c]. Arrowsmith [Mey91a]. Arsac [Kap67]. Arscott [Hoc65]. Arsenin [Wil79]. Art [Dre89, Enr87, Isa88, Isa89a, Mac69, Mac70, Nie94b, Raz92a, Raz92b, Sey93a, Sey93b, Sha75, Ste94a, Ste94b, Yan94, SWD76, Fra89]. Arthur [Bla81, Bro77, Ham80, Ork95, Ran89, Sop65, Was94a, Wei88, Wei87, Yau68]. Arthurs [SW81]. Article [Ano99b, The99a]. Articles [Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d]. Artificial [Ars97, CL72, Gra86a, Mac98a, Sam76a, Sco92, Sam76b]. Artificial-Free [Ars97, Mac98a]. Arto [Har75]. Ascent [Gri72, Wol69, Wol71]. Asch [Law18]. Ascher [Her93a, Isa89b, Her93b]. Ash [Mos93]. Ashish [Ber92b]. Ashwani [Buc85]. Askey [Car70a, Gas71, Hol75, Jag74b, Jag74a, Wil85c]. Aslam [Mil95b]. Aspect [LS85]. Aspects [Bom02, BK83, CBM98, DA79a, DA79b, DF98, Fie93, Gau67, Hal90a, Hal90b, Ham90, Har60, Hea76, Kaz80a, Kni94b, Kni94a, LS96, MQ95, MW06, Olk59, Olv93, Pip81, Pra89a, Pug86, Qua95, Ros76a, RS75, Ruo91a, SW93b, Str76a, Str77, Wat86c, WL95, PI88, Boa82b, Bou91b, Law95, Ruo91b]. Assaults [Bei80]. Assem [Nie89a]. Assembly [Giu18, HS03]. Assessing [DP76].

Assessment [DH09, LSS98]. Assignment [BR74, HK72, MO78]. Assignments [NT78]. Assimilation [Law18, RBC^{+}07]. Assistance [Hol87, McG93]. Assisted [KSW96]. Associated [AW76, BG77, EG98, Kla66, Kla78b, KG79, Köl83, Ste73c, Tra66, Koh98a]. Associations [Gri77]. Associative [Hun85]. Assumption [SS89]. Astap [Hac75]. Asteroid [Kla59]. Astor [RCC71]. Astrom [Kus72]. Astronomy [Gol71, Ung85]. Astrophysical [Lan76b]. Astrophysics [Kaz86]. Asymmetric [Bla84, Gel79]. Asymptotes [BS63]. Asymptotic [Ang92a, Ari92, Bar71, Bar85, Ben74, Boj65, Bos96, Boy05, CS06, Cho76a, Cla67, Daa97, DO98, De 84, Dri68, Dwa80, Ede90a, Ede90b, ER79, EKW96, FO99, FT76, FH76, GM81a, GM81b, GPP88, Gla89a, GS63, GJ90, Hop75, Jon72, Jor72, KP87, KCI^{+}80, Lan67a, Lin86, Lóp08, Lud70, Luk68, MS72, Mcn66, Mil66, MN76, NR90, O’M74, OSS82, Olv80, Phi87, Pin92c, PS64, Rei71d, Ren93a, Rub78, Rue72c, Rue89, Sle76, Tre65, TS90, Van86, Van66, VCG^{+}86, VM95, Wei63a, Wol77, Won80, dB64, dB65a, BY98, Ben76b, Sha97b, Buc85, Fok84, GP88, Gro90, Grz83, Har70, Kaz62, Knu80, Lan67b, Mie95, New90, O’M88b, Ren93c, Rhe60, Sco91a, Sco91b, Ser86, Was74, Wim96, Yan91, dB65b]. Asymptotically [Hag83, TD20]. Asymptotics [BC98, BK99, CO07, HH75, KM96, Mey80a, O’B11, PW08, Ver15, Ame81, Maw96, Wim75]. Asynchronous [NT78]. Athanasios [Kul65, Pol65, Por86b]. Atkin [Hor83a, Sel76]. Atkinson [Mar73, Not94, Hei69]. ATLAST [LHFE97]. Atmosphere [HH16]. Atmospheric [Dre75, Jin91, Sto11, Yan91]. Atomic [CDS01]. Attack [CR67]. Attenuated [Boh88b, Deu87, JB88]. Attenuation [Haw95, Ort96b, Ort96a]. Attitude [Sin76, Wah65, FSW^{+}66]. Attraction [Agg74, Bou76b, EG98, Kla92a, Sey93c, Sla64, Gro96a, Bou65b, Bou76a, Gro94a, Gro96e]. Attractor [VŞ10]. Attractors [ADW23, EFN^{+}96, Jol96, Mac91b, Mac91d, Hol85, Gol93]. Aubin [Fix74, Pin84b, Sch82a, Pin84a]. Audrey [Jor89b]. Augmentation [Car93]. Augmented [CG76, WE19]. August [LW64]. Augusti [Yao86]. Augustin [Gue92]. Augustin-Louis [Gue92]. Aulbach [Has86]. Aunxey [Jor87b]. Author [Cia92a, Gar94b, Gar94a, Bur89, Cas88, Pan86, Pen88, Ram90b, Cia92b, New87, Owe85]. Author-s [Gar94b]. authored [Gra18]. Authorities [DZH^{+}04]. Autocorrelation [Dum65, Fri63, LCE76, Zad67]. Autocorrelations [BD63]. Automata [Boo79, Gos84, Har72b, Hur93b, Man81, Mul64a, Paz76, Zie83a, Gra70, Hur93a, You71, Zie83c]. Automata-Theoretic [Boo79]. Automatic [Hig62c, LW64, Lyn83, McG62b, Nei10, Par72, SS16, Car65b, Fat93]. Automation [Gre60a, Kar60, Sch60a]. Automaton [Nie66]. Automotive [Gor75]. Autonomous [De 84, Her65]. Autowave [Erm89]. aux [Fal70a]. Avascular [RCM07]. Average [ADR76, CCG^{+}02, Dal76, DV76, Dob87, Jor77a, KA89, Lei78a, Lei79, LFK78, LL69, Ose75, Rul98, Sen89, WGG^{+}90, Wat90, Wil67d, WC77, vdM68, New84]. Average-Case [CCG^{+}02]. Averaged [Tot91]. Averages [Lew78a, Thy73, Thy74]. Averaging [Bau88, Coo91, Mor66, Mul76, OT11, O’M87b]. Avery [Che67b]. Avis [Hey82]. Avner [Cum96, Eva84b, Fle77a, Fle77b, Kap90b, Ock92a, Kap90d, Ock92b]. Avoid [Owe06, SKP19]. Avoidable [BS00]. Avoiding [Bur94b, Bur94c, Lon78]. Avondo [Leh63a]. Avondo-Bodino [Leh63a]. Avraham [Kha86]. Avriel [Bor90, Kan78]. Avtar [Kno83]. Aware [BCH14]. Axeisson [Han95]. Axelsson [Gun88]. Axes [Kla74a]. Axial [Min78]. Axially [SPR76]. Axially-Symmetric [SPR76]. Axiomatic [Vog87, Shu82]. Axisymmetric [Cru86, LH19, Núñ96, Por64]. Azencott [Gut88]. Aziz [Kev85].

B

[Aga93, AK90, Ari96, AF69, Ask79b, Ask92a, Ask92b, Bal95, Bau88, Bla85b, Boa81, BdD92, Bou91a, Bur81b, CL62, Cas72a, Cha95b, Che86, Chu81, Cra84, Dri75, Dur87, Ell90, Ell81, FS68, Gei89, Gib95, GM87, GK81, Goo84, Gre63a, GdMGR91, Gut82, Hag85a, Hag89a, Hag80, Ham77, Han86b, Her70, Hig62b, Hin89, Hoc86, Hol87, Hol89a, Hol90a, Hol90b, HR85, How63, Jes77b, Jd93, Joh68, Jol96, JG70, Jos95, Jun64, Kod96, Koh86, Kor65, Kri82, Lah94, Lan67b, Lem84, Lev77b, Lig86, Liu91, Luc86, Mac61, Man61a, May90, McI71, McL86b, MNC71, Mer64, MT62, Miu76a, Mor88b, Mos93, New94a, Par81a, Pin70, Pin87b, Pol59, Pra85, Pro84c, Ral70, Rei62a, Rei69, Ren94b, Rhe70, Rhe88, Ric85b, Rob62b, Rog60, Ros63, Ros78, Rox92, Rüd97]. B [Run91, Rya63, Sam86, Sau62, Sch63a, Sch86a, SR93b, Sed88, Ser94, She87, SS70, SW81, Sto81, Str89b, Syl96, Tak77, TBDS92, Tre84b, Wan91, Was74, Wei80b, Whi63b, Whi94, Wim72, Wim88, Win95, Yak94, Yan77, Yan94, Ben95, Lew76]. B-Spline [Lew76]. B. [Tre84a]. Babbage [Smi83]. Babic [Ram93]. Babister [Wim67]. Babskii [Dij90]. Baby [LKGK07]. Baby-Bathwater [LKGK07]. Back [CK18]. Backboard [Pom67, Ste61a]. Backlogging [Mor69]. Bäcklund [Abl80, Lee84a]. Backorder [Had65, Guh66]. Bad [Bur82, ES08, Erd83, Pat19]. Bade [Ano63a, ZN62]. Badii [Ask75b]. Badminton [Mar89]. Bahgat [Reh89]. Bahnasawi [Fel90]. Bai [Hol86b]. Bai-Lin [Hol86b]. Bailey [Lev63a, Ral70]. Baines [Hua96]. Bainov [Sei93]. Baiocchi [Ell87]. Bakel’Man [Ham77]. Baker [Bry91a, Bry91b, Pre81]. Bakonyi [Fra95a]. Bakr [Cru86]. Balaban [Bei78]. Balachandran [Ran96a]. Balakotaiah [Can86b]. Balakrishnan [Chu85, Har93, Hor93, Lae83b, Yak95, Lae83a]. Balakumar [Ran96a]. Balance [MM10]. Balanced [LCE76]. Balancing [Jon71, VBVM22]. Balas [Geo67]. Balasko [Ric90b, Ric90c]. Baley [Bla86b]. Balinski [May84c, May84d, May84b]. Ball [Ame85, BZ05, Coz87, Fre14, REG19, Xu96c, Xu97]. Ball-Dropping [REG19]. Ballistic [AV65, HW77]. Ballistics [Sha64a]. Ballot [JK59]. Bamberg [Wax91a, Wax91b]. Banach [Ama77b, Ber78, Dat80, Don90, Ort74, Ama76, DM67, Her78b]. Banach-Spaces [Her78b]. Band [BS92, Hue92a, RKM76]. Band-Limited [BS92, Hue92a]. Bandit [McC91b]. Bands [Ran76]. Bandwidth [PGS75, WKW75]. Bang [Art80, Gla89b]. Bang-Bang [Art80]. Bange [JA80]. Banichuk [Sym92]. Bank [Cra84, Mit94]. Banks [De 74, Lud91]. Baratta [Yao86]. Barbara [Cha87a]. Barbeau [Kla90a]. Barbra [Bus95]. Barbu [Dat80, Eva85, Kor95b]. Bard [Rei75]. Barenblatt [Ame81]. Bargaining [Shu82]. Barkauskas [JA80]. Barker [Gun88]. Barlow [Ros65]. Barnes [Aga93]. Barnett [Sin79b]. Barnsley [Goo90a, Goo90b]. Barratt [Che92a, Gua92]. Barraucand [Jor79]. Barrett [Mag86]. Barrie [Ros68a, Sim82]. Barrier [HS70]. Barriers [HH16, Her69]. Barrow [Kaz91a]. Barrucand [BMMO76, Jor72, Rue72c, Van86]. Barry [Hor93, Jag73, Sul93a, Sul93b]. Bart [Lan81]. Bary [Boa65]. Barycentric [BT04]. Basar [Art87, Seg85, Tan93b]. Base [Kli78, SG73]. Base-30 [Kli78]. Baseball [Gou10]. Based [ABG^{+}95, BH76a, BGW15, Bor62, CS63, CN15, CN19, Dra83, FTM95, FG19, FLGM23, Goo86, HH16, Hei95, JS94, KTSB21, Mar95, McA94c, McA94a, McA94b, McK95, NGTB15, RB23, Ste81a, TTY88, VDC22, Was94a, ZKP17, Nov94]. Basel [LW64, Tre13]. Baser [Bin84, Tan93a]. Bases [Lew76, LS08, Stu94b, Stu94a]. Basheleishvili [Kno81]. Basic [And74c, Ask85, Ask86a, BS94a, BS94b, BKS68, BR67, Ble81, Car87, Gor63, Gos64, Gun63, HLS09, Kla84h, Kla84a, LC72, Mal85, Muk86, Wim91a, Row79, Wim91b]. Basics [Pit89]. Basins [EG98]. Basis [CM78a, CMSM79, CDS01, Cle63, Jer76, Koh98a, Nas87b, RFS03, SM79, Vas97, Vog87, WS13, Zar68, Fix78]. Basketball [GL05, Gou10]. Basset [MT62]. Batchelor [Sop59]. Bateman [RLGBRMC06]. Bathwater [LKGK07]. Battery [Bel86]. Bauer [Car66c, Pro65a, Kob10]. Baumann [Car66c]. Bauwens [Ric88]. Baxley [Cas87b]. Bayard [Coo67a]. Bayes [CPSV18]. Bayesian [Har66a, Rei87, ABL^{+}13, AB23, Ber97a, COSG19, DMP22, Lat23, LSS98, OSS15, Ruk93a, Ruk93b, Ric88, Tay68]. Bazaraa [Fro80]. be [OW95b]. Bead [War70, War71]. Beads [CN05]. Beam [Dum96, MD96]. Bean [Bra82]. Bean-San [Bra82]. Beard [Nas87b]. Beardon [Bla93]. Bearing [SB98a]. Beauchamp [Sta89]. Beaumont [Dwa85]. Beauty [Ant92c, Ant92d]. Beavis [Art91a, Art91b]. Bebernes [Dol91a, Dol91b]. Becker [Sal64, Stu94b, Stu94a]. Beckmann [Gas76b, Hes86]. Bedford [Red89]. Bedienungstheorie [Tak77]. Beekman [Gin85]. Beem [Her83b, Her83a]. Beetle [GZ11]. Before [Bar89, Bar91b, Boy08, Bar91a]. Begehr [Tut94]. Beggs [Duf85]. Behandlung [Che65]. Behavior [Aga95a, AR78, Ari92, BLMT18, CCG^{+}02, Coh75a, DD04, ER79, FT76, Gla89a, Gre67b, GJ90, Gro90, Grz83, Jon72, JC88a, Jor88a, KCI^{+}80, Knu80, KLA95c, MSO96, Mil66, NR90, New90, Ode82, OSS82, Ott96, Pip81, Rob92a, Rob92b, Rue89, SR75, Sle76, Str76a, AR80]. Behavioral [Car87, Coz85, MHC01]. Behaviour [Cha87b, KS90, Mie95, VM95]. behind [BL06]. Beilby [Sch94]. Beker [Bla86a]. Bela [Zlo77]. Belhoste [Gue92]. Belinfante [Gru74]. Bell [Hop79, Sag77, Tru94, CM91, Mil93]. Belleni [Jer81]. Belleni-Morante [Jer81]. Bellman [Bar64, Bar65, Bec85, Car65a, Cha64a, Coh82, Dor74, Fle64, Fra86a, Ham62a, KK78, Kaw65, Mar88, Ral66b, Sal63a, Wil60a]. Bellomo [Log97]. Belt [Bra81]. Beltrametti [Bro83a]. Beltrami [Coo89, McK81, Nas77b, Shi68]. BEM [Say13]. Ben [Bla88f, Eck82, Kap68, Mey76c, Pan93a, Pan93b, Bog75]. Ben-Haim [Pan93a, Pan93b]. Ben-Israel [Eck82]. Ben-Zion [Bla88f]. Benchmark [AB23]. Bender [But79]. Bending [Cla64a]. Bendixson [MM98]. Benedict [Bro65]. Benjamin [Reh89]. Bennett [McC85]. Benoît [Dev92, Goo84]. Bensonssan [Rus95]. Bensoussan [Fle89]. Bent [Cha94d]. Benveniste [Dup93]. Benvenuto [Spi91a, Spi91b]. Benzi [Aur18]. Berberian [GA63]. Beresford [Saa83]. Berg [Gra72b, Lan86]. Berg-Man [Lan86]. Berger [Key94, Rei87]. Bergman [Kra24]. Bergs [BI68c]. Berkeley [Rei63]. Berkovitz [Rox76]. Berlekamp [YY96]. Berman [Car81, Dav91a, Gib77, Knu77a, Sil78, Tho74, Wil77c]. Bernadette [Nas90]. Bernard [Bri64, Gol65b, Gor59, Gru74, Lil67, Loe92a, Sam79, Sch89a, Sch89b, Tar81, Zah80, Loe92b]. Bernardo [Ber97a]. Bernd [Has86]. Berndt [HR85, Lev77b]. Bernhard [Tan93b, Tan93a]. Bernoulli [KLG69, Los85, NS77, Sab68, SM84, Ser78, Sim96, Tak79]. Bernstein [Der86, Wat91b, Wer89, Cha85, DCL^{+}86]. Bert [Bla81]. Bertelè [Lar74]. Berthier [Sch84b, Sch84a]. Berthold [Nag88]. Bertil [Sül97]. Bertram [BdD92, Bus95].

Bertrand [Boy90a, Boy90b, DMS23, Lev65]. Bertsekas [Beu89, Ell80, Lar78, Zlo85]. Bessel [Ask75b, Bra66, Bre89, COW96, Fet65, Gla78a, Gla86c, Gro96g, Han85b, Jag87b, Jag88a, Jor88b, LG80, Lev77b, Los86a, Mil94a, Ber76c, Cha96b, Dam85b, Dam86a, Eks60, Fet68c, Fet84b, GN78, Gla79, Gla82a, Gla87c, Gla93c, Gro96b, Han85c, Hoc67, Hui95, Ike76, JR87, JGA^{+}88, JC88b, Lar69a, Lar80, LB77, Los86b, Nås77a, OW95a, Pet85, Raw86, Rie87, RPP88, Sha63a, Sha63b, Sol67, Ste85, Vel64, WG76, WG77a, WG77b, Wim80]. Bessel-Function [JR87]. Bessel-Functions [Los86b]. Besseling [Hav95]. Best [Bor62, CU03, Ful76, JS76, MM81, Ram89]. Beta [Ask84, Wil85c]. Betchov [Kan68]. Beth [Mil69a]. Bethe [BSS05]. Better [Tre08]. Bettis [Bar78]. Between [Bla81, BGML96, FS78a, GG80, HK92a, HK92b, KP87, Lar69a, Naz76, Wil78c, BGH^{+}04, Boe96a, Dix91, HS70, Her88, Lam04, LS95, TW69, YA18]. Beyond [CMSZ24, Gle15, PCLS23]. Bezalel [Luc87]. Bezdek [Pei83, Wan83]. Bezoutiants [Hou70a]. BFGS [Naz76]. Bharucha [Daw74, Sys63, Wei69]. Bharucha-Reid [Daw74, Sys63, Wei69]. Bhatia [Dav89, Sac74]. Bhimsen [Elu88, Wei90b]. Bias [DHM07]. Bibliographic [PR86]. Bicycle [Ols03]. Bidomain [WiOH10]. Bierman [Ber62]. Bifurcation [Abe88, Ado89, Ant86a, BKR75, BP94, Bry90a, Bry90b, Chi94a, Chi94b, CLWM95, Dan91, Kaz90a, Kaz90b, Kee88a, Kir75a, Kir75b, Mey86b, Mey89c, Mey95a, Olm76, PEGK03, Rhe85, Rhe88, Sat78, Van94c, Van94b, Mey82a, Mey88, Ant78, Mar85c]. Bifurcations [Cas94a, Cas94b, Cra89, GAK17, Hol85, JH11, Kaz82, Kaz84b, Kaz84a, Mey75b, PTK96, PMR76, Sim16b, dBC^{+}08, Cho94, Kra96a]. Big [Sav18]. Biggs [Bou77]. Bigradients [Hou68a, Hou74a, Hou68b, Fra72, Gre72]. Biharmonic [GP79, Kel60b, Lig86, Red76a, ST94]. Bilateral [RVBE21]. Bilinear [GSYY22, Hin87, JS20, Abl86, Bro76a]. Bill [dM94]. Billiard [DG98, HH00]. Billingham [BdD94, Dav94a]. Billingsley [BR62, Pin69, Tul68]. Bimatrix [Mil68]. Bin [CCG^{+}02]. Binary [Blo94, Gok70, Inn69, Ken93, LCE76, Los95, McG62a, Ost61b, Sch68, Sch69b, Sch71b, Sch93c, Sch65c, Sta92a, Boe98b]. Bingni [Luc18]. Bini [Ebe96]. Binomial [BPK84, Bro60a, Car64, Car67a, Cha82a, CS83, FSP93b, FSP^{+}93a, Gro92a, Gro96f, Hig02a, Jag73, MM06, Rop83, Sib91, Ski95, Sle62, Wil92, WGLS92, Wol72b, AL98, Han98, Knu98]. Bioassay [JSR82b, Rae81, JSR82a]. Bioattenuation [WPL98]. Biochemical [Gou16, But78]. Bioeconomics [RV78]. Biofilms [KD10]. Biofluiddynamics [His77]. Biograph [Mil89b]. Biography [Ano89a, Gue92]. Biological [Bél88, Bra82, Cro77, EK91a, EK91b, FS78a, JC12, Kaz80b, KR78, Mil90a, Saw85, She91, Mil90b]. Biology [Boj93, But74a, But74b, But74c, Erm89, Man90d, MHC01, Mer81, San83a, Sha95a, SS95, Bel90b, Bra85b, But78, Chi83, Gla86a, Man90c, Mil89b, Sán83b, Seg88, Sel94]. Biomechanics [Iva04, ML80]. Biomolecular [XC11]. Biorthogonal [Cha72, Car73a]. Biostatistics [Dur97]. Bipartite [San76]. Biquadrates [SR70]. Birger [Jun70a]. Birkhäuser [Nic96, Tre13]. Birkhäuser/Springer [Tre13]. Birkhoff [Ant63, Wei87]. Birthday [Mat91, Pac88a]. Bisecting [Kon62]. Bisection [Cor77, FK06, KF06]. Bisectors [Par71]. Bishop [Moo69]. Bispectral [Hin87]. Biswas [Dur97]. Bit [Bla88f, Kar96]. Bitmead [Bau88]. Bitsadze [Pro65c]. Bittanti [Die92, Dis92]. Bivariate [Ger62, KHS63, Kri67, Str62a].

Björck [Jac76]. Bjorner [Whi94]. Bjørstad [Mit91b]. Blachman [Mal88b]. Black [DeW85]. Bladel [Pic92]. Blagovest [Tot91]. Blaine [Mil92]. Blake [Ass77, Ran91a, Ran91b, Zie78]. Blasius [Boy08, Hen76]. Blended [CBG75]. Blending [Dub76, Wix76]. Blending-Function [Wix76]. Blips [ZWY14]. Block [AT17, CD75, DU10, FG04, HSŠ16, Kel60b, RP96, Hor97b, Pea97]. Blocked [EGJK04]. Blom [Los95]. Blomberg [Zak85]. Blood [Kel75b, TTY88]. Bloom [Rog94b, Rog94a]. Bloomfield [Wil77b]. Blotto [Bel69]. Blow [NE76, SGK^{+}96, Lev96]. Blow-Up [NE76, SGK^{+}96, Lev96]. Blowout [Pac88a]. Blowup [Lev90b, PS00]. Blum [Ham73b, Ham73c, Joh91a, Joh91b]. Bluman [Ber86c, Olm77, Olv90]. Board [AN78]. Boas [Bel63a, DeM65, Lev88b]. Bobrow [Sch87]. Bobylev [Vai96]. Boccara [Dim91a, Dim91b]. Bocquet [Law18]. Bodies [Poh63]. Bodino [Leh63a]. Body [AMR04, Eas93, Eri66, Gre88, Hag85b, Sac67, Sim71a, SPR76, Ste00, Wu75, Ant72, Joh67]. Boehm [Pet95, Ren94c]. Boerner [Cur64a]. Bogart [Als85]. Bogdanoff [Tay86]. Bogoyavlensky [Kaz86]. Böhm [WBGP91]. Bohr [Sch77c, Amo78a]. Bois [Pap62a]. Bolibruch [Sib95]. Boltyanski [Fir92b, Fir92a]. Boltyanskii [Dat73]. Boltzmann [Dor89, Ker91a, GO09, Ker91b, PI88, Man61c]. Bondy [Wal79]. Bong [BKS96, Car80]. Book [Aal71, Abb92, Abl77a, Abl80, Abl84b, Abl86, Abr94, Ack74, Aeb94b, Aga93, Aga94, Aga95a, Aga95b, Ahl82, Ahl93, Aiz94b, AK90, AKP91a, AKP93a, AS85, ASF86, Ale90, Alf93, All78, All92b, Alm87, AA82, Alp85, Als85, Als92a, Ama82, Ame81, Ame84, Ame85, Ame86a, Ame86b, Ame88, Ame91a, Ame95, Amo64, Amo78a, And75a, And95, AG67, And73, And74a, And77, Ann95b, Ano61a, Ano63a, Ano68a, Ano71a, Ano73a, Ano80d, Ano82f, Ano87a, Ano87b, Ano89b, Ano91b, Ano92a, Ano92b, Ano93c, Ano93d, Ano93b, Ano93a, Ano97a, Ano98a, Ano99a, Ans64a, Ans64b, Ans64c, Ans64d, Ans66, Ans68, Ans80, Ans85, Ant72, Ant78, Ant82, Ant83b, Ant86a, Ant87, Ant90, Ant92a, Ant91, Ant92d, Ant63, Ant86b, Aok70, Apl95, Arb76, Arb77, Arb78, Are86, Are89]. Book [Are91a, Ari96, Ari91b, Arm94, Arm90a, Arn89, Arn92, Aro63, Aro69, Ars62a, Ars62b, Art83b, Art85a, Art85b, Art88, Art87, Art91a, Art92a, AGR83, Ash89, AF69, Ask72, Ask75b, Ask78, Ask79a, Ask79b, Ask80, Ask81, AD87, Ask92a, Asn75, Asn78, Asp84b, Ass77, Ast87, ABG96, Ath94, Ath87, Ath93b, Atk66, Aur18, Axf65, Bai87, Bai74a, Bai74b, Bai76, Bai93a, Bak66, Bak91a, Bak91b, Bak92, Bal88a, Bal88b, Bal93, Bal95, Ban84b, Ban85, Ban87, Ban88b, Bar86, Bar95, Bar64, Bar65, Bar67, Bar70, Bar94, Bar89, Bar91a, Bar83b, Bar79, Bar59, BS67, Bas88, Bau88, Bau89, Bau91, Bax85, Baz73, Bea90a, Beb68, Beb79, Bec85, Bec94, Bei78, Bei80, Bei81, Bei85, Bei96, Bél88, Bel90a, Bel90b, Bel62, Bel63b, Bel63a, BI67, BI68a, BI68b, BI68c, BI70a]. Book [BI70b, Ben87, Ber93, Ber95, Ber62, Ber65a, Ber83a, Ber84b, Ber85a, Ber91a, Ber94, Ber85b, Ber91b, Ber92b, Ber92c, Ber76a, Ber85c, Ber97a, Ber81, Ber86a, Ber91d, Ber85d, Ber78, Ber83b, Ber84e, Ber85e, Ber88a, Ber90b, BW65, Ber85f, Ber85g, Ber86c, Bes92, Beu83b, Beu85, Beu87, Beu89, Beu92, Bha93, BR62, BR65, BR67, BR69, BR76, BR79, BR85, Bic94, Bin84, Bir64, Bir69a, Bir86, Bjö61, Bjö85, Bla70, Bla82, Bla85a, Bla85b, Bla88a, Bla88b, Bla88c, Bla88d, Bla89a, Bla89b, Bla90, Bla83, Bla81, Bla84, Bla86a, Bla88e, Bla88f, Bla93, Bla66, Bla86b, Ble81, Blo89, Blo91, Blo76, Blo77, Blo95, Blu65, Blu86, Blu91, Blu94a, Blu94b, Blu78, Boa82a, Boa65, Boa76, Boa78, Boa81, Boa82b, Boa83, Boa84, Boa85, Boa87, Boa88, Boe88a]. Book [Boe89, BGML96, Boe76, BdDJ83, BdD92, BdD93, BdD94, BG95b, Boe95b, BC96a, Boe96b, Böh88c, Boj93, Boj65, BGL^{+}67, Boj97, Bol91a, Bol91b, BW73, Bor73, Bor79, Bor90, BC96b, Bos95, Bos73a, Bos73b, BS86, Bou91a, Bou91b, Bou77, Bou65b, Bou65a, Bou66a, Bou66b, Bou76a, Bou86, Boy70, Boy88, Boy90a, Boy97, Bra94b, Bra85a, Bra96, Bra66, Bra65, Bra82, Bra85b, Bra91a, Bra63, Bre70, Bre74, BMMO76, Bre77, BPK84, Bre89, BKS96, Bre62a, Bre62b, Bre64a, Bre68, Bre94a, Bre81, Bre93, Bri64, Bri96, BS80, Bro76a, BBLS92, Bro77, Bro65, Bro93, Bro59b, Bro76b, Bro83a, Bro90, Bro83b, Bru72, Bru73, Bru95, Bru83, Bru86, BOS73c, Bry90a, Bry91a, Bry93, Bry94, Bry96, Buc69, Buc91a, Buc85, Bue66, Bue71, Bun83b, Bur94a, Bur94b, Bur81a, BW96b]. Book [Bur79, Bur81b, BG64, Bur93, Bus86, Bus95, Bus70, But89, But92, But61, But74a, But74b, But74c, But78, But79, Byr65, Cab74, Cab75, Caf92, Caf94, Cam77, Cam91a, Cam91b, Cam93, Can91a, Can86b, Car93, Car66a, Car67b, Car70a, CS70, Car71a, Car71b, Car72a, Car73a, Car80, Car76b, Car81, Car66c, Car65b, Car65a, Car87, Cas87b, Cas95, CL62, Cas72a, Cas94a, Cav90, Cer96, Ces75, Cha94a, Cha94b, Cha87a, Cha95b, Cha67a, Cha70a, Cha78, Cha66, CL72, Cha94c, Cha94d, Cha82b, Cha83b, Cha86, Cha95d, Cha96c, COW96, Cha64a, Cha64b, Cha81b, Cha68, Che91, Che92a, Che96b, Che65, Che67b, Che18, Che89, Che92c, Che75, Che86, CS64, Chi90, Chi79, Chi83, Chi71, Chi94a, Chi95, Cho94, Cho66, Cho70, Cho87, Chu81, Chu82, Chu85, Chu88b, Chu89, Chu90, Chu96, Chu78]. Book [Chu92, Chu94b, Cia92b, CMR73, Çin93a, Cla62, Cla64a, Cla64b, Cla78c, Cla83b, Cle69, Coc84, Coc86, Cod78, Coh76, Coh90a, Coh75b, Coh82, Coh84b, Col76a, Col77, Col87a, Col91, Col73, Col87b, CZ89, Col92, Col96, Col94b, Col86, Col89, Col94c, Con85, Con96, Coo66, Coo67a, Coo73, Coo89, Coo91, Coo65, Cop86, CK86, Cor92a, Cor95, Cor16, Cor79, Cor92b, Cot73, Cot94, Cou97, Cox65, Coz85, Coz87, Cra84, Cra86, Cra89, Cre97, Cro73, Cro85, Cru86, Cul77, Cul94b, Cul66, Cul68, Cum91, Cum92a, Cum96, Cur64a, Cur77, Cur64b, Cus96, Cuv90, Cyb93, Daa97, Daf89, DC86, DV76, Dam84b, Dan93, Dan91, Dan66, Dan71, Dan72, Dat67, Dat69, Dat73, Dat78, Dat80, Dat86, Dat90, Dau90, Dav75a, Dav75c, Dav77, Dav89, Dav91b, Dav91a, Dav94a, Daw74]. Book [Day87, Day89, De 74, DeM65, DG96, DK97, DeR91a, DeW85, Dei92, Del95, Dem86, Dem90, Den78, Den82, Der86, Des68, Det85, Deu63, DRWK64, Deu81c, Deu90, Dev90, Dev91, Dev92, Dev94a, Dev94b, Dia59, Dia64, Dic67, Die92, Dij90, Dim91a, Din72, Din94a, Dis92, Dix91, DG68, Djo68, Dob87, Dod73, Dol91a, DD74, Dol62, Dol72, Don85, Don90, Doo63, Dor89, Dor74, Dra61a, Dra61b, Dra62, Dre89, Dri63, Dri65, Dri75, Dri81, Dru64, Duf84, Duf85, Duf86, Duf91a, Duf91b, Dum96, Dum65, Dun94, Dup93, Dur93a, Dur84b, Dur87, Dur90a, Dur92a, Dur92b, Dur93b, Dur97, Dwa64, Dwa66, Dwa70, Dwa72, Dwa80, Dwa85, Dwy63, Ear81, Eas89, Eas91, Eas93, Eat96, Eba82, Ebe96, Eck82, EK91a, Ede90a, Edw90, Edw63, Eic67, Eis86a, Eis86b, Elk86, Ell85a, Ell92a]. Book [Ell90, Ell80, Ell81, Ell85b, Ell87, Ell88, Ell94, Els89, Els91, Elu88, Eng78, EMUP98, Eng94, ELRS66, Enr84, Enr87, Enr89, Eps59a, Eps59b, Erd77, Erd83, Erd92b, Erm89, Esc70, EF65, Eva81, Eva84b, Eva85, EB88, Fab80, Fab87, Fad62, Fal70a, Fal68, Fal70b, Fal73, Fal93a, Fan59, Fan82, Far94, Far85, Far93a, Far65, Far69, FSW^{+}66, Far96, Fat93, FS73, Fei96, Fej88, FGB81, Fel90, Fel72, Fen94, Fer79, Fet65, Fet66, Fet68b, Fet68a, FS68, Fet72a, Fet80, FLR^{+}81b, FW82, Feu76, Fia73, Fie93, Fif79, Fif86, Fig93a, Fin84b, Fin95, Fin76a, Fin76b, Fin93, Fir92b, Fir97, Fis93, Fis71b, Fis71a, Fis76, Fis88d, Fix74, Fix78, Fix93, FSP93b, Fla97, Fla66, Fla67, Fle62, Fle73, Fle76, Fle63a, Fle63b, Fle64, Fle72, Fle77a, Fle77b, Fle79]. Book [Fle82, Fle83, Fle84b, Fle85, Fle88, Fle89, Fle90, Flo69, Flu93a, Fly89b, Fok84, FB78, Foo63, For92, For75, FARE69, Fos90, Fox71, Fra86a, Fra66b, Fra86b, Fra72, Fra86c, Fra88, Fra89, FR63, Fra95a, Fra95b, Fre80, Fre78, Fre79, Fre83, Fre86a, Fre95, Fre90, Fre86c, Fri78a, Fri82, Fri86, Fri64, Fri71, Fri59, Fro60, Fro73, Fro74, Fro80, Fro83, Fro85, Fro86, Fro87a, Fro87b, FZ63, Gal95, Gar93, Gar61, Gar91, Gar67, Gas63, Gas71, Gas76b, Gas85b, Gas70, Gas95, Gat88, Gea74, Gea80, Gea85, Gei89, GR95, Gen86, Gen76, Gen96, GK86a, GP88, Ger96, Get96, Gha89, Gha94, Gho87, Gib85, Gib95, Gib73, Gib90, Gib77, Gib87b, Gib97, GA63, Gil65b, Gil78, Gil82, Gil90a, Gil95, Gin85, Giu18, Giv78a, Giv78b]. Book [GC79, Gla95a, Gla86a, Gla94a, Gla78a, GR81, Gla85a, Gla86b, Gla86c, Gla87a, GD88, Gla88a, Gla89b, Gla93b, Gla95b, Gle83, Gli91a, GC69, Göb70, God87, Gog85, GM87, Gol65b, Gol65a, Gol66b, Gol69a, GGG69, Gol71, Gol73a, GK81, GG91, GH86, Gol93, Goo72, Goo82, Goo84, Goo86, Goo90a, Goo91a, Goo91b, Goo68, Gor59, Gos63, Gos64, Gos79, Gos83a, Gos83b, Gos84, Gou68, Gra18, Gra67, Gra72b, Gra86a, GRL87, Gra66a, Gra83a, Gra86b, Gra66b, Gra70, GM85, Gra79, Gra93a, Gra94a, Grc18, Gre60a, Gre63a, Gre97, Gre62a, Gre88, GB89, Gre76, Gre65, Gre87, Gre89, Gre72, Gre63b, Gri61, Gri88, Gro87, Gro94a, GRYL89, Gro90, Gro91a, GdMGR91, Gro93d, Gro93a, GdM95, Gro96e, Gro96f, Gro96g, Gro96h, Gro88b, Gru74, Gru88, Gru93, Grz83, Gua92]. Book [Guc87, Guc79, Gue69, Gue72, Gue79, Gue82, Gue83d, Gue83b, Gue83c, Gue84, Gue85, Gue88, Gue89, Gue92, Gue95, Guh66, Guh68, Gui92, GWK95, Gun63, Gun88, Gun95, Gup82, Gup89, Gup91a, Gup96, Gur62a, Gur62b, Gur66, Gur63, Gus76, Gus90a, Gut97, Gut82, Gut88, Gut90, Haa93, Hab83b, Hag85b, Hag85a, Hag89a, Hag80, Hag82, Hag83, Hag85c, Hag86, Hai85, Hai77, Haj96, Hal66, Hal95, Hal90a, HPF85, Ham85, Ham87, Ham88, Ham90, Ham61, Ham62a, Ham62b, Ham65a, Ham65b, Ham70, Ham71, Ham73a, Ham76, Ham77, Ham79a, Ham79b, Ham80, Ham73b, Ham73c, Han95, Han94b, Han75b, Han79, Han81a, Han85a, Han87, Han91, Han64, Han67, Han81b, Han83b, Han85b, Han86b, Har64a, Har73, Har74, Har83, Har96, Har61b, Har61a, Har64b, Har66a, Har67d, Har70]. Book [Har72a, Har76b, Har76a, Har78, Har93, Har63, Har72b, Har75, Har82, Has86, Has92a, Hau85, Hau86, Hau88, Hau89, Hav95, Hea91, Hea74, Hea81, Hei93a, Hei93b, Hei89, Hei97, Hei96, Hel97, HBO66, Hen86, Hep67, Her77, Her78a, Her80a, Her80b, Her83b, Her84b, Her97, Her65, Her67b, Her67a, Her69, Her71, Her88, Her93a, HF05, Her70, Hes83, Hes86, Hes72, Hey82, Hey85, Hey86, Hey88, Hey89, Hey93, Hey97, HR62, Hig62a, Hig62c, Hig62d, Hig62b, Hig65, Hig92, Hig99, Hig05, Hil76, Hil96, Him78, HS72, Hin87, Hin91, Hin89, Hin94, Hir71, Hir73, His77, Hoc65, Hoc79a, Hoc79b, Hoc80, Hoc82, Hoc83b, Hoc86, Hoc92, Hoc94, Hof85, Hof95, Hol91b, Hol70a, Hol85, Hol86b, Hol86a, Hol87, Hol89a, Hol89b, Hol95, Hol66, Hol75, Hol90a, Hom86]. Book [Hop79, Hop83, Hor83b, Hor83a, Hor89, Hor97a, Hor93, HR85, Hou67, Hou70b, Hou71a, HA71, HT81, How63, How82, How83, How94, Hua96, Hüb88, Hüb91, Hud95, Hug93a, Hul61a, Hul61b, Hul71, Hun81, Hun84, Hun85, Hur65, Hur93a, Hut89, Hut91, Hye73, Isa88, Isa89a, Isa89b, Ism76, Ism78, Ism83b, Ism85, Ism96b, Isr92b, Ivi65, Jac76, Jac97, JS71, Jag64, Jag71, Jag72a, Jag72c, JBL72, Jag72b, Jag72d, Jag73, Jag74b, Jag74a, Jag76a, Jag79a, Jag79b, Jag80, Jag82, Jag83, JGF87, Jag87a, Jag87b, Jag88b, Jag88a, Jag90, Jag92, Jam59, JW83, JRR71, Jan70, Jar76, Jef95, Jel87, Jer81, Jer91a, Jes77b, Jes93b, Jet91, Jin91, Jd93, Jód96, Joh67, Joh81, Joh83f, Joh83c, Joh83d, Joh84b, Joh87, Joh88, Joh91a, Joh68, Joh73, Jol96, JSR82a, Jon92a]. Book [Jon92c, JG70, JF71, Jor72, Jor73, Jor77a, Jor77b, wBJ77, Jor78a, Jor78b, Jor79, Jor80a, JA80, Jor80c, Jor81, Jor82, Jor83a, Jor84a, Jor84b, Jor84c, Jor86a, Jor86b, JB88, Jor88a, Jor88b, Jor89a, Jor90a, Jor90b, Jor90c, Jor93b, JG93, JGM95, Jor87a, Jor87b, Jor89c, Jor89b, Jos95, Jow72, Jow76, Jow86, Jun64, Jun65, Jun70b, Jun70a, Jur95, Kal63, Kal93a, Kal65, Kal70, Kv93, Kan78, KH85, Kan62c, Kan62a, Kan62b, Kan66, Kan68, Kap85a, Kap85b, Kap90a, Kap90b, Kap64, Kap65, Kap66, Kap67, Kap68, Kap70b, Kap71, Kap72, Kap76, Kap77, Kap78, Kap81, Kar90, Kar60, Kar63, Kar78, Kar82, Kar83, Kar84b, Kar64, KJ87, Kar87, Kat65, Kau86, Kau88, Kaw63, Kaw65, Kaw66, Kaw67, Kaw68, Kaz62, Kaz67, Kaz75, Kaz80a, Kaz80b, Kaz82, Kaz83, Kaz84b]. Book [Kaz86, Kaz89, Kaz90a, Kaz91a, Kaz91b, Kaz92, Kea96b, Kea80, Kea93a, Kee95, Kee88a, Kee61, Kee87, Kee88b, Kee90, Keh89, Kei64, Kei65, Kei71, Kel60a, Kel67, Kel84b, Kel86a, Kel83, Kel84d, Kel88a, Kel96, Kel97a, Kel97b, Kel97c, Kel97d, Kel98a, Kel98b, Ke98, Kel99d, Kel99a, Kel99b, Kel99c, Kel00, Kel75b, Kel86b, Kel90b, Kel92b, Kel86c, Kel90c, Kem67, Kem85, Kem92a, Ken86a, Ken86b, Ken91a, Ken93, Ken74, KS94, Ker91a, Ker92, Kes65, Kes69, Kes75, Kes85, Kev85, Key94, Kha86, Kha87, Kha88, Kha71, Kie94, KLG69, Kin94b, Kin97, Kin76, Kin82, Kin94c, Kir93a, Kir90a, Kir76, Kla63a, Kla80, Kla83a, Kla84h, Kla84i, Kla84j, Kla84g, Kla86a, Kla88a, KG89, Kla90a, Kla92b, Kla97, Kle71, Kle72, Kle76, Kle75, Klo93, Kni84b, Kni86]. Book [Kni94b, Kno81, Kno83, Knu77a, Knu80, KL87, Kob10, Koç89, Koç94, Kod96, Koe95, Kog66, Koh85, Koh86, Koh88, Köl83, Köl87, Kol63, Kol65, Kol76, KF62, KS73, Koo77, Kor65, Kor68, Kor73, Kor95b, Kor95a, Kor85, Kov96, Kow83, Kow86, Kow88, KZ93, Kra96a, Kra68, Kra67, Kre96, Kri85, Kri86, Kri91, Kri82, Kul60, Kul65, Kul69, Kum82, Kum88, Kun93, Kup94, Kup62, Kup75, Kur69, Kus72, Kus74, Kus93, LW64, Lae76, Lae82, Lae83b, Lag89, Lag97, Lah94, Lai91, Lai78, Lak80, Lam86, Lan81, Lan86, LF66, LS66, Lan67b, LG80, Lar94, Lar74, Lar78, Lat69, Lat90a, Lau74, Law18, Law95, Law65, Law66, Law72, Lay24a, Lay24b, Laz82, Laz91a, Le 92, LeV77a, LeV93, Lea94, Lee77, Lee82a, Lee82b, Lee84a, Lee84b, Lee86a, Leh63a]. Book [Leh63b, Leh82, Lei66, Lem88, Lem84, Lem85, Lem83b, Len80, Len95, Len93b, Ler64, Les76, Lev62, Lev67a, Lev63a, Lev88b, Lev96, Lev84, Lev87, Lev63b, Lev77b, LG89, Lew83a, Lew86, Lew91b, Lew91a, Lie97, Lie95, Lig86, Lil67, Lil68, Lin91a, Lin80, Lin90, Lin79c, Lin92, Lit63, Liu88, Liu91, LM81, Loc84b, Loe95, Loe92a, Log97, Lon87, Lor91, LJW73, Los73, Los74a, Los74b, Los75, Los76b, Los76a, LFR80, Los81, LJ83, Los83a, Los83b, Los84a, Los85, Los86a, Los89, Los95, Lou92, Lu92, Luc87, Luc86, Luc18, Lud77, Lud82, Lud87, Lud91, Luk75, LSB67, Luk68, Luk82, Lum85, Lus87, LS84, Mac91b, Mac91a, Mac68b, Mac69, Mac70, Mac92, Mac61, Mac76a, Mag86, Mah94, MR71, Mal80, Mal85, Mal87, Mal88a, Mal90, Mal93a, Mal77, Mal88b, Man90a]. Book [Man89, Man88a, Man61a, Man61b, Man61d, Man61c, Man81, Man70, Man87, Man88b, Man90c, Man85, Man63, Mar66, Mar85a, Mar64, Mar73, Mar75a, Mar83b, Mar81, Mar86a, Mar88, Mar76b, Mar85c, Mar86b, Mar65, Mar87, Mar95, Mas95, Mas87, Mat89a, Mat80, Mat88, Mat94b, Maw86, Maw91, Maw93a, Maw95, Maw96, May91, May84c, May84d, May90, McA88, McA92, McA96a, McA96b, McA94c, McC91a, McC93a, McC92a, McC91b, McC79, McC83, McC85, McC71, McC95, McC86, McG62b, McG67a, McG67b, McG64, McG65a, McG65b, McG93, McG61, McI71, McK95, MP62, McK81, MN83, ML85, MBCG83a, McL79, McL86b, McM61c, McM87, McN59, McN89a, McN89b, MNMF63, Mea86, Mea88a, Mea88b, MNC71, Mec92, Med78, Mee96, Meg88, Meh73, Meh96, Mei78a, Mei83a, Mei77, Mei89]. Book [Mel92b, Men94a, Mer64, Mer81, Mer85, Mer84b, Mey72, Mey75b, Mey76c, Mey77b, Mey78, Mey79, Mey82b, Mey82a, Mey83, Mey84, Mey86b, Mey88, Mey89b, Mey89c, Mey90, Mey91a, Mey91b, Mey93b, Mey94a, Mey95a, Mey95b, Mey96, Mic95b, Mic92b, Mie95, Mik89, Mil69a, Mil82a, Mil83a, Mil80, Mil82b, Mil61a, Mil82c, Mil92, Mil94a, Mil95b, Mil81a, Mil81b, Mil83c, MT61, MT62, MT64, Mil89b, Mil90b, Mil91a, Mil94b, Min66, Mir85, Mir88, Mit65, Mit93, Mit87, Mit94, Mit91b, Miu76a, Mla97, Mla94, Moc85, Mod92, Moi89, Mon60, Moo74b, Moo74a, Moo75, Moo60, Moo63, Moo69, Moo76, Moo77b, Moo77a, Moo85, MS65, Mor74b, Mor86a, Mor88b, Mor97, Mor93, Mor96, Mor68, Mor71, Mor73, Mor75b, Mor82, Mor83, Mor85b, Mor88d, Mor89c, Mor91a, Mos93, Mot88, Mro91, Muk86, Mul64a]. Book [Mul66b, Mul67, Mul69, MR65b, Mur94a, Mur97, MG95, Nag96, Nag88, Nas87a, Nas87b, Nas90, Nas65, Nas77b, Nas86, Nas95, Nau86, Nel74, NPW95, Ner70, Neu94, Neu97b, Nev82, New87, New90, New94a, New76c, NW81, New82a, New74, New76e, Nic96, Nic65, Nic94, Nie89a, Nit85, Nit90, Noe72, Noe81, Noe82, Noh62, Noh67, Noh71, Noh76, Noh77, Nor90, Not94, Nov94, O’B95, O’M74, O’M86, O’M87b, O’M87a, O’M88b, O’M92, O’M96, O’M00a, O’M00b, O’M00c, O’M01a, O’M01b, O’M01c, O’M01d, O’M02a, O’M02b, O’M02c, O’M02d, O’M03a, O’M03b, O’M03c, O’M03d, O’M04a, O’M04b, O’M04c, O’M04d, O’M05d, O’M05a, O’M05b, O’M05c, O’M06a, O’M06b, O’M06c, O’M06d, O’M07a, O’M07b, O’M07d, O’M07e, O’M08a, O’M08b, O’M08c, O’M08d, O’M09a, O’M09b, O’M09c, O’M09d, O’M10a, O’M10b, O’M10c]. Book [O’M10d, O’M11a, O’M11b, O’M11c, O’M11d, O’M12a, O’M12b, O’M12c, O’M12d, O’M13a, O’M13b, O’M13c, O’M13d, O’M14a, O’M14b, O’M14c, O’M14d, O’R92, Obe64a, Obe64b, Obe77, Ock91, Ock92a, Oco87, Ode86, Ode89, Odg86, Old62, Ole68, Olk59, Olk72a, Olk72b, Olk77, Olk78, Olm77, Olv72, Olv86, Olv90, Olv91a, Olv93, Olv97, Orb18, Ork95, Orl93a, Orl93b, OP88, Ort74, OW89b, OG90, Ort96b, Osb85, Ost61a, Ost61b, Oth82, Ott96, OR67, Owe85, Pac86, Pac87, Pac88b, Pac10, Pal84b, Pal95, Pal96, Pal94, Pan93a, Pan95, Pao88, Pao90a, Pao93, Pap62a, Pap62b, Pap63a, Pap63b, Par94, Par88, Par80, Par87, Par90, Par66, Par74, Par81a, PT81, Par10, Par69, Par67a, Par72, Par67b, Par76, Par92b, Par92a, Pas95, Pat64, Pat87, Pav94, Pay68]. Book [Paz76, Pei83, Pel97, Per89, Per65, Per68, Per76, Per93a, Pet95, Pet68, Pfl92, Pic87a, Pic90, Pic92, Pic87b, Pil87, Pim87, Pim93, Pin88, Pin85a, Pin85b, Pin87a, Pin92b, Pin69, Pin70, Pin72b, Pin72a, Pin77, Pin84b, Pin87b, Pip81, Pip86, Pip92, Pit89, Pit90b, Pit91, Pit93a, Pit74, Pit77, Pit79, Pit81, Ple88, Poh63, Pol59, Pol65, Pol69, Pol76, Pon90a, PRM^{+}66, Por72, Por80, Por83, Por85, Por86a, Por88, Por81, Por86b, Por89, Por92, Pow75, Pow80, Pra68, Pra85, Pra86, Pra89a, Pra89b, Pré95b, Pre81, Pro65a, Pro65b, Pro83b, Pro84b, Pro65c, Pro84c, Pro85, Pro87a, Pro91, Pro95, Pro96, Pug86, Pul64, Pul91, Qua95, Qui92b, Rab94, Rad90b, Rad73, RV78, RCC71, Ral62a, Ral62b, RA63, Ral66a, Ral66b, Ral70, Ral74, Ral80, Ral83b]. Book [Ral83a, Ral86, Ral93a, Ram90a, Ram93, Ran89, Ran96a, Ran96b, Ran91a, Rao94, Rat87, Rat92a, Rat93, Rau95, Raz87, Raz92b, Red89, Red90, Red93a, Red83, Reh89, Rei71a, Rei77a, Rei96, Rei62a, Rei62b, Rei63, Rei69, Rei71c, Rei75, Rei77b, Rei77c, Rei78c, Rei79, Rei83a, Rei87, Rei88, Rei89, Rei64, Ren84b, Ren89, RJ90, Ren90, Ren92b, Ren93c, Ren93d, Ren94a, Ren94b, Ren94d, Ren94c, Ren95, RJ96, Ren97b, Ren80a, RL62, Rhe60, Rhe70, Rhe85, Rhe88, Ric67a, Ric74, Ric88, Ric85b, RO66, RR89, Ric90b, Rie72, Riv60a, Riz94, Rob81, Rob85, Rob95b, Rob94a, Rob62a, Rob62b, Rob74, Rob76, Rob91, Rob92a, Roc92a, Rog60, RB90, Rog94b, Rog96, RFC72, Ron95, Rop83, Ros82, Ros71, Ros63, Ros65, Ros68a, Ros72a, Ros78, Ros73, Ros68b]. Book [Ros80a, Ros85, RSS65, Ros72b, Ros87, RD65, Rou72, RAG72, Rou73, Rou91, Rou88, Row64, Row76, Row79, Row86, Rox76, Rox77, Rox84, Rox90a, Rox92, Roy91, Roy93a, Rub69, Rub77a, Rüd97, Rud61, Rue72c, Rue72a, Rue72d, Rue73a, Rue73b, RSTW73, Rue74a, Rue76b, Rue78, Rue79, RLG81, Rue81a, Rue82b, RAK88, Ruk93b, Run91, Ruo90a, Ruo91b, Ruo92a, Rus95, Rus60, Rus90, Rus85, Rya63, Saa83, Saa73, Saa88, Saa96, Saa68, Sac83b, Sac74, Sae63, Saf96, Sag77, Sag78, Sag81, Sag82, Sag83, Sag86, Sai91a, Sal85, Sal64, Sal66, Sal62, Sal63a, Sal63b, Sal63c, Sal63d, Sam86, Sam76a, Sam76b, Sam79, Sam76c, Sam80, San82, Sán83b, San95, Sar84b, Sat78, Sau62, Sau78, Sau79, Sav18, Saw82, Saw85, Say79, Sca59, Sch77a, Sch93a, Sch63a, Sch71a, Sch73a, Sch82a]. Book [Sch82b, Sch84b, Sch86a, Sch86b, Sch89a, Sch65a, SNS66, Sch95, Sch89b, Sch80a, Sch87, Sch93c, Sch94, Sch81a, Sch60a, Sch60b, Sch65b, Sch65c, Sch93d, Sch76c, Sch18c, Sch18a, Sch18b, Sch18d, Sch19a, Sch19b, Sch19c, Sch19d, Sch20a, Sch20b, Sch20c, Sch20d, Sch21a, Sch21b, Sch21c, Sch21d, Sch22a, Sch22b, Sch22c, Sch22d, Sch23d, Sch23a, Sch23b, Sch23c, Sch78c, Sch85a, Sch91a, Sch60c, SR93b, Sch81c, Sch85c, Sco83, Sco84b, Sco85, Sco88, Sco90a, Sco90b, Sco91a, Sco91b, Sco92, Sco94, Seb70, Sed88, Seg79, Seg80, Seg85, Seg88, Sei62, Sei85, Sei93, Sei86a, SK80, Sel94, Sel76, Sen94, Ser80a, Ser80b, Ser84b, Ser86, Ser94, Ser96, Ser64, Set62a, Set62b, Set63, Sey93b, Sez90, Sha74, Sha82a, Sha83, Sha95a, Sha96, Sha97a, Sha65, Sha92a, Sha97b, Sha85]. Book [Sha95c, SM79, Sha89b, Sha82b, She77b, She63a, She63b, She66b, She87, She91, Shi94, Shi68, Shi69, Shi71, Shi93, Shi97, Sho69, SR79, Sho66a, Sho87, Shu87, Shu80, Shu82, Shu85a, Shu85b, Shu90a, Sib95, Sie69, Sie89, Sie90, Sik92, Sil78, Sim89, Sim76, Sim93a, Sim82, Sin59, Sin69, Sin93a, Sin79b, Sin93b, Sit92, Siv94d, Siv94a, SS70, Ske95, Sle74, Sle89a, Sle92, Sle94, Sle79a, Sle81, Sle83a, Sle85, Sle86, Sle87, Sle89b, Sle62, Sle67, Sle84b, Slo93, Slo97, Sma76, Smi74, Smi83, Smi80b, SW81, Smi82, Smi86, Smo91b, Sno76, Som93, Sop59, Sop60, Sop62, Sop65, Sop69, Sop72, Sop79, Sop82, Spä97, Spi84b, Spi87, Spi91a, SLF66, Spi74, Spi80, Spi81a, Spi81b, Spi82, Spo85, Spo88, Spu89, St.82, Sta66, Sta85, Sta88, Sta89]. Book [Sta90, Sta92b, Ste87, Ste67b, Ste68a, Ste71, Ste72, Ste73b, Ste81b, Ste88, Ste90a, Ste93a, SKGM67, Ste68b, Ste70a, Ste93b, Ste94a, Ste97b, Sto96, Sto61, Sto62a, Sto62b, Sto66, Sto81, Str86a, Str89b, SFH^{+}66, Str68, Str70, Str76b, Str82, Str86c, Str62a, Str62c, Str62b, Stu90, Stu91, Stu94b, Sub90, Sül97, Sul93a, Sup77, Sus79, Sus88, Svo96, Swa96, Syl96, Sym92, Syn83, Syn93a, Sys63, Szy87, Ta’90a, Tab95, Tab97, Tad89, Tad93, Tak77, Tak83, Tal81a, Tal81b, Tan93b, Tap84b, Tap86, Tar81, Tau80, Tay68, TT76, Tay86, Tay89, Tay96, Tay97, Tek89, Ter93a, TA90b, Ter92, Tha79, Tha82, Tha96, Thi63a, Thi63b, Thi84b, Tho74, Tho76a, Tho87, Tho60a, Tho59, Tho75, Thr79, Thr94a, Thy74, Tis83b, TBDS92, Ton95, Ton89, Tot91, Tre84b]. Book [Tre13, Tre65, TS65, Tri95, Tru94, Tu92, Tul68, Tul90, Tur60, Tut94, Uhl73a, Uhl73b, Ull88, Ung85, Urb93a, Uri92, Vai96, Vak93, Van86, Van96, Van88, Van66, Van68, Van76, VJ83, Van63, Van93, Van94c, Var87, Var85, Var66, Var67, VK62, Vel71, VN73, VBdD91, Via85, Vog87, Vog91a, Vog93, Wag92, Wah92, Wal95, Wal87, Wal79, Wal88, Wal93a, Wal97, Wal63a, Wal93c, Wal75b, Wan91, WP88, War92, War64, Was94a, Was74, Was94b, Wat91b, WJ94, Wat63, Wat65, Wat86a, Wat10, Wat15a, Wat15b, Wat15c, Wat15d, Wat16a, Wat16b, Wat16c, Wat16d, Wat17a, Wat17b, Wat17c, Wat17d, WS71, Wat86b, Wat90, Wax91a, Web85, Wei80b, Wei84b, Wei88, Wei87, Wei91, Wei69, Wei77, Wei79, Wei67b, Wei67a, Wei90b, Wel88, Wel89, Wen68, Wen78]. Book [Wer89, Wer63, Wes92, Wes84b, Wet74a, Wet75, Wet93, Whe74, Whe76, Whi93, Whi63a, Whi63b, Whi66a, Whi66b, Whi67, Whi69, Whi76a, Whi76b, Whi92a, Whi80, Whi81, Whi83, Whi84b, Whi88, Whi94, Wib87, Wic94a, Wic94b, Wie87, Wig85, Wig90, Wig91, Wil62a, Wil62b, Wil62c, Wil67a, Wil70a, Wil84b, Wil85a, Wil86a, Wil88, Wil66, WLP^{+}61, Wil65a, Wil69, Wil61, WJV79, Wil91a, Wil60a, Wil60b, Wil64c, Wil65b, Wil72, Wil73, Wil77b, Wil78b, Wil79, Wil81, Wil86b, Wil93, Wil77c, Wil82b, Wil85c, WM85, WBGP91, Wil92, WR94b, Wim65, Wim67, Wim72, Wim75, Wim80, Wim81, Wim86, Wim88, Wim91b, Wim96, Win74, Win95, Win60, Win62, Wis60, Wit90, Wol86, Wol97, Wol72a, Wol94a, Wol93, Won83b, WGM93, Woo92, WA63, Wou64a, Wou64b, Wou67, Wou68a, Wou68b, Woź87]. Book [Wri60, Xan87, XL94, Xu96a, Xu96b, Yak94, Yak95, Yan77, Yan91, Yan92, Yan94, Yan65, Yan76, Yao86, Yat84b, Yau68, Yau69, Ye94, Yon95, YY96, You72a, You71, You72b, You74, Zad67, Zah80, Zak84b, Zak85, Zal83, Zar65, Zas65, Zei72, Zem96, ZN62, Zha96, Zha97, Zie76, Zie78, Zie83c, Zie83b, Zie84b, Zim90, Zit63, Zlo77, Zlo83, Zlo85, Zlo91, Zuo96, Zwe84b, dB65b, dM94, vH87, vK79, vL86b, vL86a, vV65, vWHE91, vdM68, vdV94, Bel10, Bor09, Tre15, Ver09, Che67b]. Books [Ano62e, Ano63f, Ano63g, Ano63h, Ano63i, Ano64e, Ano64f, Ano64h, Ano65e, Ano65f, Ano65g, Ano65h, Ano66g, Ano66h, Ano66i, Ano66j, Ano67e, Ano67f, Ano67g, Ano67h, Ano68g, Ano68h, Ano68i, Ano68j, Ano69f, Ano69g, Ano69h, Ano69i, Ano70f, Ano70g, Ano70h, Ano70i, Ano71h, Ano71i, Ano71j, Ano71k, Ano72g, Ano72h, Ano72i, Ano72j, Ano73g, Ano73h, Ano73i, Ano73j, Ano74e, Ano74f, Ano74g, Ano74h, Ano75k, Ano75l, Ano75m, Ano76j, Ano76k, Ano76l, Ano76m, Ano77j, Ano77k, Ano77l, Ano77m, Ano78l, Ano78m, Ano78n, Ano78o, Ano79i, Ano79j, Ano79k, Ano79l, Ano80l, Ano80m, Ano80n, Ano80o, Ano81g, Ano81j, Ano81h, Ano81i, Ano82m, Ano82n, Ano82o, Ano82p, Ano83j, Ano83k, Ano83l, Ano83m, Ano84l]. Books [Ano84m, Ano84n, Ano84o, Ano85g, Ano85h, Ano85i, Ano85j, Ano86i, Ano86j, Ano86h, Ano86k, Ano87h, Ano87i, Ano87j, Ano87k, Ano88f, Ano88g, Ano88h, Ano88i, Ano89i, Ano89j, Ano89k, Ano89l, Ano90g, Ano90h, Ano90i, Ano90j, Hig99, O’M05d, Ano64g]. Boole [Har61a]. Boolean [Mul67, Leh65, Mar60, BI70a, Whe76]. Boor [Ron95, Smi80b]. Boorman [Lud82]. Booth [Sch60a]. Bootstrap [Siv94b, Siv94d]. Border [Fra86c, Sim16b]. Border-Collision [Sim16b]. Bordered [PR70]. Borel [Daa97]. Borgan [McK95]. Borgwardt [Bla88c]. Boring [GZ11]. Boris [Dei92, Par80, Swa96, Daa97]. Borisov [Kre96]. Borkar [Fle90, Gut97]. Borkowski [Pit90b, Pit90a]. Bornemann [Hig05]. Boron [Ral62a]. Borovkov [How94, Ser86].

Borowiecki [Bei85]. Borrelli [Koç94]. Borwein [Kri82, Van96, Wim88]. Boss [Chu88b, Sch95]. Both [HT78]. Bottema [BW73, Los76a, Wei80b]. Bottiger [Gla87a]. Botvinnik [Ber85c]. Bouncing [BZ05]. Bound [Deb76, DJLQ21, Esl93, Esl96, GG73, Gre62b, HE91, JS71, Kah66, Kor91, Los74a, Rez62, Sch72a, Sch73b, Tre70, Van98, WS23, vW90, vWHE91]. Bound-Preserving [WS23]. Boundaries [CS90, O’M92]. Boundary [Ale95, AMR04, Ann95b, Ann95a, Bai74a, Bai74b, Ban87, Ber84b, Ber84a, Bes92, BG77, Ble81, Blu86, Blu91, Boy05, BK78, Bus76a, Cla67, CGRG^{+}23, Con97, Con66, Cru86, Dam84b, Dam84a, Eck72, Ell87, Ell88, Els91, Eva84b, Eva84a, Eva90, Faz96, Faz98, Fel76, Fix74, Fos90, Fri79b, GS23a, GHC97, Gei89, GdMGR91, Gro91b, Gro93a, Gud75, Gun88, Her70, Het72, Hig86, HMO97, HM73, HA75, Isa89b, IA75, Joh75, Joh83d, Joh83b, Jub70, Kap81, Kel75a, Kel79b, Kla62b, Kla64a, Kla65, KR74, Lau74, Let69, Lev88b, Lor91, Man76, Mat75, Mat76b, MG76, Mat85, Mey77a, MCS69, Muk86, Na67, Na68, NC76, O’M71, O’M08e, Obe64b, Pon90b, Pro87a, Ral66b, Ral70, Red80, RT97, Ril75, Roo10, Ros76b, Rus90, Sar84b, Sar84a, Sch76b, She60b, Shi68]. Boundary [Shi93, Sle76, SPR76, Sun76, Tho59, Van94a, Van88, Van78, Van94d, VH78, Wat96, WS75a, Wil88, Wil67b, Wil67c, Zay92, ZKP17, ZC74, AR81, Boe98a, Con98, Ren97a, Ans68, Day89, Els89, Gue69, Gur62a, Lig86, Pon90a, Riz94, Sag86]. Boundary-Field [GHC97]. Boundary-Layer [Ble81, CGRG^{+}23, Kel75a, MG76, Van94a, Van94d, VH78]. Boundary-Value [Ale95, Ann95b, Ann95a, Ber84a, Bus76a, Fel76, Fix74, Gei89, Gro93a, HA75, IA75, Joh75, NC76, Ral66b, Ros76b, Sar84a, Sch76b, Van88, Van78, WS75a, Zay92]. Bounded [Cha87b, Don23, Har66b, JC88a, Kar22, Rog69b, Jor88a]. Boundedness [She77a, She78, Won64b]. Bounding [MW76b, Moo78, dB74]. Bounds [Ahl62, Ang92b, BK88, BT19, Bic94, Bus76a, Dav89, Deh89, Far60, FR66, Gol66a, Hog74, Hog75, Jac69, Jag87c, JK88, JS76, KS68, Lan64, Mor69, New63, New65b, Olv80, Ram89, Sch69a, Sch71c, Sch72b, Sch73c, Sha64b, Sig67, Sta76, Ste73c, Tin76, VBC07, Won80, And75a, Jag88b]. Bouwkamp [Goo68, Wol72a]. Bowden [Fab87]. Bowers [Ste93a]. Bowles [Tre84b, Tre84a]. Bowman [Ars62a, McC91a, Pol59, Rog60]. Box [Gre89, dBHRR95, Ron95]. Boyarintsev [Cam93]. Boyce [Koç94]. Boyd [Che92a, Gua92, Rue79, RLG81, Yak95]. Boyer [Tal81a]. Boynton [Jag83]. Bracewell [Wel88]. Bradley [Fre83, Siv94d, Siv94b]. Bradshaw [Pra89a]. Brady [Gra86a]. Braess [Deu90]. Brailsford [LSB67]. Brain [Seg85]. Bram [Ral66a]. Bramble [Kin94c]. Brams [Hof85, ML85]. Branch [Olm76, PR91]. Branch-and-Cut [PR91]. Branching [BR65, DSL04, Goo72, Lae76, Sta71, Sta72]. Brands [JRA22, Ser94]. Brandstatter [Row64]. Brannan [Boa82b]. Brassard [Buc91a, Buc91b]. Bratley [McC85]. Brauer [Cha67a, Her71]. Braun [Gue84, Hoc79a]. Breakdown [GP02]. Breaking [Hea91, Mul79, PB75, YFS76]. Brebbia [Ban87, Fos90, Lor91, Muk86]. Brechtken [Ahl93]. Brechtken-Mandersheid [Ahl93]. Brekhovskikh [Mil82b]. Brenner [Ano68a, BG64, Jag72b, Kat65, MR65b, Pas95, Sei86a]. Bressan [Tau80]. Brewer [Wie87]. Brezinski [McC93a, Wim81, McC93b]. Brezzi [Fal93a, Fal93b, Por88, Sim93a, Sim93b]. Brian [Art91a, Boy97, Isa88, Isa89a, Art91b, Auc97b]. Bridge [CGP21, SKK20, GK12]. Bridges [LM90, Mey95b]. Brief [Sal63a, Str89a]. Brigadients [Hou71b]. Brioschi [Lar74]. Brittleness [OSS15]. Britton [Rus85]. Brock [SKGM67]. Brøndsted [Sal85]. Brook [ELRS66, JF71, Sle62]. Brooks [Car65b]. Brown [Art92a, Art92b, Chu92, Gre63a, Lu92, Pan95, TS65, Gay76]. Brownian [Kni94b, Pit93a, BRHC20, Kni84a, Kni94a, MV68, Pit93b, Kni84b, Kni86, Pin72b, Spi80]. Broyden [Gay78]. Brualdi [Gib97]. Bruce [Dat69, Hun84, Kha88]. Bruijn [Rhe60]. Bruno [Gue92, Rei77b, Rei78c]. Brunton [Luc18]. Bruter [Mar85c]. Bryan [Dra62, Gib97]. Brzoska [Tad93]. Bubble [Kel80b]. Bubbles [BFF^{+}99]. Buchanan [Pro85]. Buchberger [Che86, Cop07]. Buck [DeM65]. Buckled [AR78, AR80]. Buckley [CP78]. Buckmaster [Kap85b]. Budden [Fej88]. Budgeting [Ber65a]. Bug [Ede97]. Bugrov [Kla84j, Kla84c]. Building [CT95, Gre89, Mla94, Tab95, Gut88]. Bulakh [Sed88]. Buldyrev [Ram93]. Bulirsch [Kin82]. Bulk [Beu85]. Bump [ADW23]. Bunce [Wie87]. Bunch [Par81a]. Bundling [GD78]. Bundy [Lus87]. Buniakowskiĭ [EV91]. Bunyan [Lew78b, LCG^{+}79]. Buoyancy [Reh89]. Buoyancy-Induced [Reh89]. Buoyant [FS76]. Burchuladze [Kno81]. Burckel [Boa81]. Bureau [Tod75]. Burger [Cha64b]. Burgers [BW11]. Burghes [Gue83d, Mor82, Gue83a]. Büringer [Noe82]. Burke [Lee86a]. Burkinshaw [Art92a, Art92b]. Burman [Vai96]. Burn [Per93a, Per93b]. Burnashev [Erd83]. Burnside [Kee61, Ost61b]. Burr [Sle84b, Sle84a]. Burrage [Jac97]. Burrows [She77b]. Bursters [Izh01]. Burton [Han85a, Han87, Par67a]. Bus [FR83, San87]. Busacker [Min66]. Bush [Mur94a, Mur94b]. Bushard [Jes77b]. Business [Ber62, Tab97, Wri60]. Butcher [Enr89]. Butkovskiy [Hoc83b, Jef95, Hoc83a]. Buttigieg [Nas90].

Butzer [Ask72]. By-Products [For76]. Byron [Kel86c].

C [All92b, Ano82f, Ant63, Arb77, Are89, AF69, Ask79b, Ass77, Ban87, Bar95, Bar64, Bar65, Bau88, Bel90a, BI68a, Ber94, Ber81, Ber85g, Beu85, Bla82, Bla81, Bla86a, Blo76, Blo77, Blu91, Blu94b, Boa83, Boy88, Bro83b, Bur93, Bus86, Cas87b, Che91, Col87a, Col96, Coo91, Cox65, Cyb93, Dam84b, Dam84a, Dan66, Day87, Day89, DeM65, Dev92, Die92, Din72, Dis92, Enr89, EB88, Far85, FSP93b, For75, Fos90, Fra86c, Fro80, Gar91, Gas63, Ger96, GC69, GK81, Goo68, Gra72b, Guc87, Gus90a, Gus90b, HPF85, Han64, Har83, Har67d, Hau89, Her80a, Hey89, Hig00, Hof85, Hol87, Hop83, Hor93, HR85, Hul61a, Isa89a, Jag90, Jel87, Joh83f, Jol96, JG70, JG93, Jor87a, Jos95, Jun64, Jun65, Kan62b, Kap85a, Kap76, Kar90, Kaz80a, Kee88b]. C [Ker92, Köl87, Kor95a, Kup75, Lan86, Law65, Lee84a, Leh63b, Lem88, Lem84, Lev77b, Liu91, Lor91, Los83b, Los89, Lud87, Luk75, Lum85, Mal85, Man81, Mar85c, McC91b, McG62b, McG65b, MN83, McL86b, McM87, MNC71, Meh73, Mey91a, Mic95b, MT61, Mil94b, Muk86, Nas86, Noh76, Not94, Par81a, Par72, Par92b, Par92a, Pas95, Pat87, Pei83, Pin85a, Pip81, Pow75, Pre81, Pul64, Rad73, Ral74, Ran96b, Rei71a, Rei79, Rei64, Riz94, Ron95, Ros78, Ros68b, Ros80a, Roy91, Saa73, San82, Sch65c, Sco84b, Sco84a, SM79, Shu85b, Sin93a, Sle86, Sma76, Smi74, Sop82, Spi80, Spu89, Ste71, Str62c, Sus79, TBDS92, Ton89, Wag92, Whi93, Whi69, Wil60b, Wil73, Wil81, Wol72a, Won83a, Yan92, Zie78, Zlo91, Zwe84b, Zwe84a, vL86a, vdV94, Han86a, Kin94a]. C [OH95, EMUP98, Hof95, Kin94b]. C. [Joh83e, Naa01, Rob94b, Rüd97, Sco84a, Wan83]. CA [BC96b]. Cable [Bel90a]. CAD [Far85]. Cading [Wyn69]. Cage [CHT15]. Cahit [Chu78]. Caines [Lin90]. Cake [BJK11]. Cake-Cutting [BJK11].

Calcul [BI68b, Ost61a]. Calculated [Ham74]. Calculating [KB76b]. Calculation [Bar74, For98b, For98a, GM76a, Kij93, Koh98b, NE76, Sie65, Spa62b, Ste73b, Van75, VdLBF17, WM61a, WM61b, Köl87]. Calculations [Gef75, Ric63, SCS10, Hul61a]. Calculator [Fre78]. Calculators [Whi80]. Calculus [Ahl93, Bla97b, Bou91a, CCA08, DC86, DY11, DF98, Eic67, Fle63a, Fle63b, Fle84b, Fle84a, Hal66, Ham62b, Har61a, Her84b, Her84a, Kel60a, Kla84a, Loe95, Mal83, Mec92, MP72, MW66, Mor92, MT96, Mos92, Net20, Pin77, RB23, Rei71a, Sag82, Sop65, Str77, Tis83a, TL95, TTY88, Vet73, Bla97a, Ber86c, Fre78, Maw86, Tis83b, Bel63b, Kla84h]. Calculus-Based [RB23]. Caldvvell [McN59]. Calibration [BP95, Pan95]. Calin [Chu90]. Calvin [Mar86a]. Calvo [Ske95]. Cam [Koh98b]. Cam-Form [Koh98b]. Cambanis [RR89, RAK88]. Cambridge [Hig02b]. Camera [Tur04]. Camion [Mil81a]. Campaigns [Tud92]. Campbell [Ber81]. Can [CEH12, Pan84, Pro87b, ZWY14]. Canards [BBK20, DGP^{+}16]. Cancer [Mer81, RMT^{+}03, SFOM99, SKK20]. Candles [Pac88a]. Canonical [GW76, GGK96, HKD20, KL97b, LR05, OHBG00, KL97a, Ken86a]. Canopies [Gan89, Gan90]. Canuto [Boy88]. Capacitor [MA04]. Capelo [Ell87]. Capillary [Alm87, CF76, CS90, Dij90, YFS76]. Capital [Ber65a]. Capiz [Wil91b]. Capobianco [Har83]. Capriciousness [FR11]. Capriz [Wil91a]. Capturing [Gue61]. Car [MMT^{+}21, Mic95a, Moo94, Pin88, Sal87, Mic95b]. Carbonization [For76]. Carcinogenesis [WK78]. Carcinogens [BK83]. Card [BK76, Knu77a, Knu77b]. Cardiac [CS06]. Cardinal [Mer84b, Mer84a, Ste81a]. Care [PGH^{+}88]. Carl [Cho70, Gup96, Rei89, Sae63, Smi80b]. Carleman [EK87]. Carleson [Dev94a]. Carlitz [AG67, And73, And74a, Bre77, Giv78b, Pet68, RAG72]. Carlo [Bru72, Dor89, Ham77, Jun65, Jun70b, Slo93, DKST19, DM07, DMP22, Fos68, Hal70, LW17, NA06, Pug66a, Spa62b, YKS78, Zar68, Ric88]. Carlson [Ask79b, TBDS92]. Carmona [Dur87, Fig93a, Fig93b]. Carnegie [Ral86]. Carnegie-Mellon [Ral86]. Caroline [Mal90]. Carré [Tar81]. Carrées [BI68b]. Carrier [Cho70, Dan93]. Carroll [All92b, Her78a]. Cartesian [Str62a, Hen75]. Cascade [Gel79]. Cascades [PH13]. Casciati [Yao86]. Case [Ame86b, AG14, AW19, BTKNZ00, CGT^{+}97a, CCG^{+}02, Col76b, Cum91, CWW99, Dob87, Esl96, Fee00, FP21, Fly76, JRZ97, KSW96, Lau03, Lei76, LS22, Nag96, SS89, SN96, SS61, Wey62, WJV79, Yan97b, Cul77, Ell80, Fre83]. Cases [BDvD^{+}98]. Cash [Enr84]. Cassels [Joh83c, Joh83a]. Cassinelli [Bro83a]. Cast [Chi95]. Cast-Rigiano [Chi95]. Castelle [Gut88]. caster [Yan97a, VS95, Yan97b]. Casti [Eas89, Lak80, Rue73b, Sus88]. Catalan [RC95, Cas95]. Catalysts [Coh76]. Catastrophe [Ban84a, CHC95, Gol78, Gra83b, Kaz91a, Som79, Tho77a, TH96, Ban84b, Ber85e, Chi95, Guc79, Kaz82, Sus79]. Catastrophes [Ber83b]. Catastrophic [Lon78]. Categories [Ano74d]. Category [AM74a, CH61]. Catenary [Ves95]. Catherine [Ros85]. Catlin [CC90]. Cattermole [Art88, Beu87]. Cauchy [Boa87, Gue92, Her78a, Tru94, Wil85c, Ask84, EV91, Fus68a, GI97, Lev65, Naz94, Ric61, Sha04, Wan79]. Causality [Sae70, New74]. Caustics [Gra83b]. Cavendish [Pow75, SM79]. Cayley [Dec65b]. Céa [Dan72]. Cebeci [Pra89a]. Cecil [Jes93b]. Cees [Mey88]. Celestial [Gar67, Pol69, Saa73, Wol88, Pol76]. Cell [Mer81, NRS19, WHMN16, VB75, Ton89]. Cell-to-Cell [Ton89]. Cells [DGS05, GDS05, Kel75b]. Cellular [But74b, Chi83, Gla86a, Hur93a, Hur93b, Mah94, Mil89b]. Cemal [Hut91]. Centennial [BS77]. Center [Abb92, Abe88, Eis62, GK91, Kla60, Min70, NRS19, Qui92b, Qui92a, RR97b, Rou97b, RR97a, Rou97a]. Centered [JK72]. Centers [BP94, Mas92]. Central [Har78]. Centre [PLC96]. Centroidal [DFG99]. Centrosymmetric [And98a, And98b]. century [Lax89, Van94a, Van94d, Sag82, Sch93d, HF05]. Cercignani [Bar95, Dor89]. Certain [AW76, Dic67, Fen68, Gil62b, Gla91, Gra72a, GM66, GG68, Han64, Hoc60, Jub70, Max60, Mul66a, NW67b, Rub78, Sch63c, Sco59, Ste61b, Ste73c, Ter92, War68, Wen61, Win67, Won64b, dO69]. Cesari [Noh62, Rox84]. Ceschino [Gra67]. CGLS [CPSV18]. Chadan [Roy91]. Chai [Car73a]. Chain [AW19, Bou21, BDX04, DKST19, DMP22, Pit81, She76, Rob62b]. Chains [BR69, Dur93b, Kar78, Kij93, KTSB21, Mey75c, Mey89a, Ros95, Sch70b, Sch71e, Yon06, Dwa72, Jag71, Pit74, Rob94a, Tay68]. Chakravarti [Ste71]. Chalker [Ada97b]. Challenge [Gli91b, Hig05]. Challenges [CAiJS23, DGS05, GDS05, HFKG06]. Chamberland [Cha96c]. Champ [HH75]. Chance [BCH14, GLY^{+}18]. Chance-Constrained [BCH14]. Chances [Ham74, HH75]. Chandler [Spo88]. Chandra [Mor85b]. Chandrasekaran [Esc70]. Chandrasekhar [DeW85]. Chandy [Bru83]. Chang [Ame91a, Ame91b, Ber94, Blu94b, BPK84, Cro85, Der86, Mor74b, O’M86, Ren84b, Rop83]. Change [DS79, FC76, Joh84b, Joh84a, Mac76b]. Changepoint [VC06]. Changes [Ano96a, Ano97b, ST94]. Changing [DP76, VCKJ16]. Channels [Jev92a, Jev93, Jev95a]. Chao [Kel92a, Kel92b]. Chaos [Ant92b, Are93, Bry90a, Bry90b, Bry93, BZ05, CC23, Cra89, Hol85, Hol87, Kea93b, Koç89, LH95, Mac91c, RW95, Sed02, Sey93a, Sey93b, Sha95a, Ste94a, Ste94b, SS95, Stu90, Thu01, Stu90, Abr94, Ant92a, Bar94, Hol95, Kea93a, Pal94, Wal95, Are86, Hol86b, Mac91a, Pal94]. Chaotic [Bry91b, Dev90, Hol87, KS90, Kra96a, Mac91d, MSO96, PTK96, Rob92b, Saa95, Stu90, Are89, Bry91a, Mac91b, Ott96, Rob91, Rob92a]. Char [Mor74b]. Char-Tung [Mor74b]. Character [Bri64]. Characteristic [Alh96, Far60, Han91, Hou98b, Ike75, Kes65, Red76b, Hou98a, Rob62a]. Characteristics [BPJ95, TKMP11, Jef95]. Characterization [CH66, Dri68, JGM95, Jor95, Kla64b, Kla94b, Kla95b, MO68, Par88, RJ96, Sto73, Jor96, Luk75]. Characterizations [Kla88b, LM00, Man68, MS89, Cam77, Mat89a]. Characterizing [Pat19]. Charalambos [Art92a, Art92b]. Charge [WZCX12]. Charges [Hni89, JH90, Jor90c]. Charles [Bai93a, Bre93, Bru83, Cha81b, Col94c, Dem86, Dem90, Gen76, Goo72, Hei97, Kap64, Luc86, Pin85a, Ple88, Qua95, Sch78c, Spi80, Sta88, Sta90, Ste93b, Stu91, Wal93a, WJV79, Bai93b, Col94d, Smi83]. Charnes [Win62]. Charpit [Del97b, Del97a]. Chart [AW19]. Chatelin [Ans85]. Chattering [Kre96, ZBK96]. Chaudhry [Beu85, Mil95b]. Chaudry [Gla93b]. Chawla [Jor88a, Kau88]. Chebotarev [Pol69]. Chebyshev [Boy13, Cha67b, Dun68, Fra75, Kra80b, Mei68, Mou67, VBC07, Che67b]. Checkerboard [New71, New72b]. Chemical [Hig08, Jon71, PSV12, Bar94, Bei78]. Chemically [Buc85]. Chemistry [Erm89, Sha95a, SS95, Rou88]. Chemostat [Cus96, SWC96]. Chemotherapy [Mer81]. Chen [Dur93b, Har73, Hau89]. Cheney [Cha68]. Cheng [WLP^{+}61]. Chengzhi [Mey95a]. Chernavskii [Erm89]. Chernikov [Bry93]. Chernov [Wil62c]. Chervonenkis [Kog66]. Chess [Ber85c, HH75, Ber76a, Mar86b]. Chess-Solving [Ber85c]. Chi [KHS63, Kri67, Rei71c]. Chi-Squared [Rei71c]. Chiang [Ger96]. Chicatelli [Wan91]. Chih [Car80, Lee82a]. Chih-Bong [Car80]. Child [MG23]. Childress [Yat84b, Yat84a]. Chilling [BWYY96, YY96]. Chin [Mor74b]. Chin-Liang [Mor74b]. chiral [Gro98d, Ric98]. Chistyakov [Par80]. Chladni [GK12]. Choi [Koh88]. Choice [Bas88, CP78, Fis74, Ser04, Lea94]. Cholesky [KF76]. Choosing [MC91, Pon83]. Chor [Bla88f]. Chorafas [Cul66, Cul68]. Chosen [BGML96, LS95, Boe96a]. Chossat [Sch95]. Chow [Jag72c, Mey95a, Rue73a, Sam80]. Chris [Nit85]. Christensen [Moi89, Spi87]. Christian [Cam91b, Cam91c, Dav75c]. Christoffel [DEP01]. Christoph [Jon92a, Kle72, Jon92b]. Christopolous [Zem96]. Christos [Won83b, Zem96]. Christy [Mei89]. Chronicle [Ano68b, Ano68c, Ano68d, Ano69a, Ano70a, Ano70b, Ano73b, Ano73c, Ano76e, Ano79d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96b, Ano96c, Ano96d, Ano96e, Ano97c, Ano97d, Ano97e, Ano98b]. Chua [Dev90]. Chuan [Lee82a]. Chuan-Chih [Lee82a]. Chuen [BPV97b]. Chui [Bre93, Hau89, Wal93a]. Chun [Arm90a, Arm90b]. Chun-Hung [Arm90a, Arm90b]. Chung [BR69, Wol97]. Church [BGL^{+}67]. Churchill [Obe64a, Obe64b]. Chvatal [Bla85a]. Ciarlet [Bar79, Ber83a, Mac92, Nas90, Sco94, Sik92, Sit92, Nas92]. cinlar [Pit77]. Ciprian [Bea90a, Bea90b]. Circle [Ano91a, FK62, Kla89, KMH^{+}90, Lee97b, Rog78, Sku05b, SS61, Ste96, Lee97a, Sku05a, Bor79, Mey90, Mey91b]. Circling [Hni89, Jor90c, JH90]. Circuit [Coh60, Kra68, PBL98, Row66, Wil21, Bos73b]. Circuits [McN59, MA04, Yau68, Hig62b]. Circulant [Cha82b]. Circular [Amo64, CHL08, DG98, Dub97b, Las76, LFK78, MN65, MS65, RR97b, Sku05b, Sta92a, Tri62a, Zit76, Dub97a, RR97a, Sku05a, Sch93c]. Circular-Loop [MN65]. Circumference [Kol91, Kol92]. Cities [HW77]. City [BL84, Zit76]. Civil [Spi84a, Spi84b]. Claerbou [Fri86]. Clair [Fro60]. Claremont [Spa77b, Spa77a]. Clarification [For96]. Clark [Fab80, Kee90, Ken91a, Mic92b, RV78, Sha83, Wal95, Jef94, Ken91b, Mic92a]. Clarke [Man85]. Class [Ber75, Ber65b, Car73a, Cha72, CM78b, Cho69, Cle59, CB76, Cot76, DJLQ21, Fai71, Fol73, GZ19, GHC97, GF78, Hea74, Hoc66, HMO97, Ken93, Kla62b, LCE76, LW59, MR00, Ric61, San65, Sco91a, Sco91b, SXY19, Str86b, WM78]. Classes [Mil66]. Classic [Bjö61, CGRG^{+}23]. Classical [Ahl61, Bak91b, Bak92, Bar74, Boa81, Bue71, Cha70b, CH68, Coo91, CBM98, Duf69, Dum96, Gud84, Gue95, Gui92, Hir71, JM92, Jef92, Kir90b, Kli78, KLT03, KLT04, Kov96, LK65, Mac77, MD96, Mit65, Olv97, Pal84b, Pal84a, Pro83a, Rog94b, Rog94a, RRG95, SC81, SR93b, SR93a, Spi80, TtK96, Tu92, Whi66b, Rob94b, Aga94, Bax85, Kir90a, Lee77, Pap63a, Pro83b, Sch89a]. Classification [Beh76, BF16, Gil70, Hil07, New76f, RMT^{+}03, RBZ23]. CLASSLIB [Pin85a]. Classroom [And98b, Ang97b, Ano75d, Ano78e, Ano78f, Ars97, Bat21, Bla97b, BCD98b, Dar97, Dax97b, Del97b, Dra83, Dub97b, Fer97b, For98b, Gan98b, Gup98b, Hou98b, KL97b, Kwa98, Lee97b, Liz98b, Mel97, MBR98b, MY97, NJN98b, NW97, Pet97b, Rei97b, RR97b, Rou97b, Smi97, Sny97b, TeB97b, VBU97a]. Claude [BI68c, Har64a, McC93a, Wim81, McC93b]. Claudette [McA92]. Claudio [Ell87]. Clements [Cum91]. Clenshaw [FW66, Tre08]. Clerk [Jon92c]. Cleve [Gup91a, Gup91b, Joh68]. Cleveland [Fra86b]. Client [TH96]. Client-Server [TH96]. Clifford [Car87, Har93]. climate [Dut82]. Climbing [BCD98b, BCD98a]. Clinic [Spa77b, Spa77a]. Cloak [BL10]. Cloaking [GKLU09]. Closed [Sma83, Var80]. Closeness [Far94]. Clouaire [McA92]. Cloud [Boe91, Ver90, VBdD91]. Clumping [Sie90]. Clunie [Boa82b]. Cluster [Beh76, Che75]. clustering [ADPR04]. Clusters [FLGM23]. Coagulation [MP06]. Coal [For76]. Coarse [LKGK07]. Coaster [Fly89b, Fly89a, Lew88]. Coating [TS90]. Cochlear [Ins78]. Cochran [Gue83b]. Coconuts [FP04]. Code [Jef90, Mic92a, SG76a, WS75a, Mic92b]. Codebooks [Jev92a]. Codes [EK13, Las76, Mat94b, Mat94a, Sav97, SH95, Mat80, Wol93, Gha89, Hal95]. Codes-Theory [Mat94b, Mat94a]. Codimension [HMO97]. Codimension-Two [HMO97]. Coding [AM74b, Erd92a, Jev91, Mat88, McG62b, Zie76, Zie78, Boy97, Ass77, Erd92b, Gha94, Kem67]. Cody [Tha82]. Coefficient [Cha82a, Enr78, Roa78, Rop83, Sch77a]. Coefficients [AB23, FSP^{+}93a, Gil62b, GG73, Gro92a, Gup98b, Han64, Hil76, Kev87, LW81, Los74a, McL86a, MM06, Moh66, NS63, Owh17, Por64, Ram89, Sch65b, SD74, Sha82a, Sib91, Ski95, Str62b, WGLS92, Woj06, Gup98a, FSP93b, Gro96f, Mey77b, Str62c, Wil92]. Coercivity [Elc76, Mor74a]. Coffman [Kv93]. Cognition [Cyb93]. Cohen [Aal71, Ano71a, Beu83b, Bos73b, Har93, Kei71, Vog91a, Beu83a, Vog91b]. Coherence [Jos95, WA94]. Cohoon [Gla87a]. Coin [DHM07, FB78, MM89, NW61, New77, PF62, Pul64, REG19, She63a]. Coin-Flipping [REG19]. Coke [For76].

Colbourn [Hei97]. Cold [Kri76]. Cole [Che89, Lat69, Mey82b, Olm77]. Coleman [Gue84, Koç94]. Colin [Hol85, Kee90, Kir90a, RV78, Kir90b]. collaboration [Fal68, McA92, Mei89]. Collatz [Gra66a]. Collection [Ask75b, Ham70, Koç94]. Collections [Ano93k, Ano93l, Ano94h, Ano94i, Ano94j, Ano94k, Ano95r, Ano95s, Ano95t, Ano95u, Ano96s, Ano96t, Ano96u, Ano96v, Ano97l, Ano97m, Ano97n, Ano98e]. Collective [SB98a]. Collectively [Moo76]. Collector [MW06]. Collegiate [TKMP11]. Collet [Kir93a, Kir93b]. Collins [Che86]. Collision [Sim16b]. Collisions [Don23]. Collocation [Arc75, BNT10, DA78, IA75, Kel17]. Collocation-Galerkin [DA78]. Colloid [Wil61]. Colonel [Bel69]. Colonius [Rox90a, Rox90b]. Colony [Sco92]. Colony-Style [Sco92]. Color [Bei80, GMP05]. Coloring [Ano79e, Erd72, FS73, GMP05, Meg78, Meg79]. Colton [Gil78, Sle94, Wil84b, Wil84a]. Column [AAJ76, Cha87c, Kau88, Kel90a, WKW75]. Combat [Abd90, CT76b, Lew91b, Sed02]. Combination [Jor80c, Jor80b, Per79, War75]. Combinations [Whi88]. Combinatoire [Ham87]. Combinatorial [And75b, And76, Ark66, Bar76a, Bar75, Bla89a, Bla90, Bre70, BMMO76, BKS96, Bre64b, Bru72, Can91b, Car67b, Cof83, Fir92b, Fir92a, Fis88a, Göb70, Hay68, Hei97, Hul89, HJ90, Ier96, Jag80, Joh83e, Kar76, Ker91a, Ker91b, Kul60, Mor88a, Mor89a, Mor89b, Mor78, MR65b, NÖ79, NdA14, PW08, RS02, Roh70, Roh71, RFC72, Ros76a, SS79, Sav97, Sho78, SR79, SF97, Tar78, Ula64, Whi69, Whi84a, Won83a, XRTM95, Zie78, Boe96c, Bre96, IR97, Kra96b, Len93a, Sch96, Can91a, Dwa70, Hei89, Joh83f, Sha85, Whi80, Whi81, Whi84b, Won83b]. Combinatorics [Als92b, AM74b, Liu88, Luc86, Pil87, Als85, Als92a, Gra72b, How94, Moo75, Fis93]. Combustion [Dol91a, Dol91b]. Come [Lar69b]. Comment [Bur89, Gar94b, Gar94a, Ram90b, Swi60, Cas88, Pen88]. Commentaries [CGO07]. Commentary [Jef94, Man90a, Man90b]. Comments [Hil75, Hue91, Jef94, Mac98a, Mac68a, New87, Owe85, Pan86, Rob94b, Rub78, Wat22, Sle83b]. Committees [Luc87]. Communicability [EH16]. Communicating [FKL^{+}06]. Communication [Beu87, SS65, Deu63, HR62, Kul69, Jar76]. Communications [Bla84, Gre63b, Sil61, Sin61, Art88, Bla86a]. Community [TKMP11]. Commutation [Ros68b]. Commutative [Bre63, Kat65, Som93, Stu94b, Stu94a, Wie87]. Commuting [Ack73, Ber88b, Lei76, WBLQ89, Ack74, Wer89]. Como [Por88]. Compact [CL75, Mar73, Moo76, Wil81]. Companion [Boy13, Kam76, Mel12, Bor09, Tre15]. Comparative [Thi84b, Thi84a]. Comparing [TKMP11]. Comparison [Aro78a, BG76, Far94, Gol76b, Jac69, Kre64, Kru83, Lam04, LS63, MS76, MC76, Mor66, RAB02, SK78, Spu89, Tre78, Tri78, WS75b, Wol77, Le 92]. Comparisons [CG76, Ebe75]. Compartment [AEH^{+}21, GS23b]. Compartmental [JS93, VLPR20]. Compatibility [CS16]. Compatible [TD20]. Competition [Ken93, SB98a, SWC96, SAT22, Cus96]. Competitive [Art92a, Art92b]. Compilation [Per93b, Per93a]. Complement [LM00, Mil89a, OG90, OGGR90]. Complementarily [Bla89b, Cot73, Mai72, Mor93]. Complementarity [FP97a, Mor74a, Sch04a, Wil70c]. Complementary [Kla74a, Kle83, Fre86c]. Complementation [Mey89a]. Complete [Alh96, ABL^{+}13, Fin71, Hoc70, Kat78, Lee86b, SR98, TS86]. Completed [Sch66]. Completely [Kam78, Qui87]. Completeness [Cle63, YT95, Har82, Tri95]. Completion [CESV15, Mat69a]. Complex [BI70c, Boa76, Boa78, Boa81, Boa82b, Boa87, Bre61, Buc69, Bur94a, Cav90, CLOT99, Cho70, Cor78, Dec65a, Dev94b, EERG20, Fis71b, Gil78, Gil90a, Gil90b, GKP75, Haa93, Har96, Hoc80, Jow76, Kap77, KM80, Luc18, MQ18, MS96a, MH96, Mer84a, Mil69b, Mil73, Mit65, Mon70, Mor64, New03, Rue81a, Seb75, Sma76, ST98, Svo96, Thi63a, Thi63b, TBBER21, VŞ10, Whi63b, Whi66a, Whi66b, Zlo70, Hen79, Tre13, Dev94a, Hoc80, Mer84b, Mot88, Obe64a, Sto61, Sto66, Whi63a, Whi76a]. Complexes [SBH^{+}20, Whi81]. Complexity [ABGJ21, Bla85b, Bla88b, Boo94, BB88, Bra17, HLH03, Lev89, MPT^{+}02, Pan92, Tar78, TA90b, TLM90, Was94a, Wil21, Won83a, Nov94, Shu87, Ste88, Wim88, Won83b, Ric74, Wol86]. Component [Dre75]. Components [Flu93b, Kim65, Flu93a]. Componentwise [Rum99]. Composing [BKST15]. Composite [Wol65]. Composites [Gra18]. Composition [Con97, Con98, CEH12]. Compositions [Mye75]. Compound [Cho76a]. Comprehensive [Bei96, HRB96, DMS23, May91, NWF97, ZMPK24]. Compressed [BCT11, BL13]. Compressible [Mar75b, RFS03, Vel21, Sle85]. Compressor [GH01]. Computability [Zie83c, Zie83a, Sho69, Smi82]. Computable [Bal20, Kah66]. Computation [AvK^{+}08, Ame73, AB73, BB84b, BCZ21, CGO07, Cha81a, Che86, Cor72, Cul78, CS90, EH98, GW76, GMP66, Gre60a, Ham73b, Ham73c, Ike75, JHSB22, JM78, Kar60, Kel75a, Kin94b, Kin94a, Koc92, Kow88, Kri10, Küç92, Leh82, LMDLS23, Mir97, O’R10, Orb18, Par72, Quo75, RK78, Sch93d, Seb75, Seg06, SB72, Tha79, Urq68, War76, Boj97, Log97, Don85, Gen86, Gun88, Jel87, Kow86, McG67a, Row76, Tho87, Cyb93, Hul71, Yau69]. Computation-Steeples [Ame73]. Computational [Bai93b, BK69, Boa76, Boa78, Boa87, CG24, CAiJS23, ERS14, Els89, FPM97, Fie93, Fla97, For67, FR66, Gar89, Gau67, Gre73, Gru74, Ham90, HLH03, Hol90b, Luc10, McC95, Nas77b, NBR96, Naz94, OH95, Pai76, Pin85a, Pol73, Por86a, Pra89a, Ran96a, Roy93b, RM95, RWMD18, SIA01, Sha97a, Sim16a, Som93, Ste98, Stu94b, Stu94a, Tal81a, Tu76, TPCC^{+}11, Was94a, Wim88, YL03, Hen79, PAB^{+}01, Ver09, Bai93a, Fra88, Gro87, Hag86, Hof95, Hol90a, Rhe70, Rhe85, Row86, Roy93a, Sav18, Sch85a, Ste88, vL86b, vL86a]. Computationally [Bus76a, PdHM16]. Computations [Alh96, Aur18, BES05, BPE96, Boy08, CL23, DKW^{+}09, Dem90, DW95, Ebe96, Eck80, FLM22, Fix93, GPS90, Hen76, Hug93b, Mil82c, Ng94, OLHB02, Pan92, Tew70, TTM05, Cho87, Dem86, Gar93, Hug93a, Moo85, Ric74, Rub77a]. Compute [MV78a, MV78b, MV03, MBK08]. Computed [Fin84a, KMS22, Fin84b]. Computer [AKP91a, Ask79a, BP76b, Bla88a, Boe88a, Bre89, Bun83a, DeR91a, DeR91b, Den78, Efr88, Flo69, GG99, GBF78, Gor63, Gos83a, Han83b, Han83a, HB95, HLB95, Kee88a, Kei89a, Kei89b, KPSP96, KSW96, KM86, LHFE97, Lud87, Lus87, Mac69, Mac70, Mar60, McG67b, Mil89b, PFM76, Pra89a, Pro96, Ral83b, Ral86, Ran96b, Raz92a, Raz92b, Ric74, RLGBRMC06, RPP88, Sch88, Sho99, Sin59, Ste99, Sto76, Tal81b, TB76, Van93, WS75a, Wes92, WR96b, Zie84b, Zie84a, Grc18, Koç94, Pin87a, Smi83, Ber95, Ber76a, Beu92, Bos95, Bru83, Bru86, Bun83b, Che86, Duf84, Duf86, Enr87, Fla66, Joh68, Kel83, Odg86, Ral83a, Smi74]. Computer-Aided [BP76b].

Computer-Assisted [KSW96]. Computer-Intensive [Efr88]. Computerized [McA92, Min78]. Computers [Boy08, DW95, Efr79, Hul61a, Kar60, LS63, OV85, Whi80, Wil81, Hig00, Fla66, Hul61b, Kar63, Lin79c, McG62b, RA63, WA63, Wri60, Ber85c, Har82]. Computing [AMSV02, Ale85, BEKS17, Bla88a, Buz78, CHT21, CD75, FK06, Fix75, Gal95, GW12, GMP05, Gla95e, Hig02c, Hig05, Ike76, Ips97, Ism96b, Jun64, KB76b, KN98, Kel59, Ker91b, Kog66, LDG95, Lax69, Leh63b, Leo78, Mcc78, Mil91b, NF16, Nan20, O’M10f, PD89, Sal63c, SPR76, Vis01, Woj06, Yan94, Ism96a, Cha83b, Isa88, Isa89a, Ker91a, Mil91a, Sch60a, Zie83b]. Comstock [Bla81]. Concave [BI70b, PR86]. Concavity [Bor90]. Concentrated [SG76b]. Concentrating [Mor78]. Concentration [BF96, SDW06, Tho77b, Kes75]. Concept [Far85, Nie83, Sin76]. Concepts [BPP95, Dre89, Eck94, Kla84h, Kla84a, Kus19, LC72, NW82, Pac87, Pet95, Rou88, Fre86a]. Concerning [ABG96, Bac60, BG95a, Raw70, Sie61, TF65, AH98]. Concise [FGP^{+}22, New76b]. Concrete [AABT15, Cum92a, Cum92b]. Concurrent [Kow88, Mit93, Orl93b]. Condensation [Ock91]. Condition [Alp96, Hig87, Sei78, She77a, She78, SPR76]. Conditional [BW96b, LT13, Roe95, Zha88, BW96a]. Conditioned [Bar22, Gol76c, GW76, MCS69, Neu98, OT13, PW79, Rum99, Tre20, WE76]. Conditions [BTKNZ00, BG77, Car73a, CCCM06, CD94, Cha72, Ewi77, HS88, KK98, MW76b, Meh74c, Mor74a, Na67, Nas77b, PLC96, Sun76, TT94, Tre73, Urq68, Wol69, Wol71, ZZ18]. Conducting [Chi79, JM78]. Conduction [BK88, CZ92, Gla93b, Hui95, Hui96, Jag87c, JK88, Let69, RK75, Ros87, Wil59, ZC74, Gro96b, Hui97, Gro96g, Jag88b].

Conductor [Rai59]. Conductors [Kla70a]. Cone [BI70c, GE63, Law65, New67a, PD89, Sha88]. Cone-Monotonicity [BI70c]. Conference [Wey67]. Confidence [Gho87, Ang90]. Configural [Tuk87]. Configuration [LM99]. Configuring [FLNU18]. Confined [Gia63, Kri76, VN76a]. Conflict [Bau91]. Conflux [Bau91]. Conformal [Boa87, Cer70, Cer71, ET99, PFM76, Ric89, Tru94, Tur04, Whi63a]. Conforming [Kat76, Kat78]. Congestion [Des68]. Congress [LW64]. Congruence [LR05]. Conic [Han90]. Conical [Sed88]. Conjecture [ABG96, BG95a, FG76, GC79, LM78, LMGC79, Lau82, Lau83, Lau84, MS72, New76d, RC95, Wan84, AH98, Cas95, Erd77, FGB81]. Conjectured [Bou86, Cah77, CM83, Chu78, CMR73, Flo77, Gau87, GS88, Gib87b, Gib87a, Gla77a, GRLR87, GWK95, GW95, HL86, Jag90, JBG^{+}90, Jor80a, Kaw95, Kla94a, Mak86, NW81, NY90, Nut85, RS79, Sch72c, SV89, Usm72, Vei86, Dut82, Rob80, GRL87, Jor89a, LJW73, Rue78]. Conjectures [HU07, Fix93, Her88]. Conjugacy [Naz86]. Conjugate [CN96, Ehr76, GO89, HB96, Luc10, Man09, Naz76, Naz86, Bra96, Rus85]. Connected [Ari92, ET99]. Connectedness [CH61]. Connecting [Sil66, Sil67]. Connection [Lau98, SBXD06]. Connections [CHW12, FS78a, Her88, LM90]. Connell [Sco92]. Conquer [BJK11]. Conquest [Bei80]. Conrad [Kee90]. Conroy [Flu78]. Consecutive [GPP88, GP88, Phi87]. Consensus [MT14, OT11]. Conservation [Bal97, ES80, EHS17, GRL98, HLvL83, Jef92, Pit91, Sle85, WPL98, Rob94b]. Conservative [Zha97]. Conserving [CN15]. Considerations [BL09, Sny97b, Sny97a]. Considering [BS00]. Consistency [Dri91]. Consisting [Deu95]. Consolidation [Sch89b]. Conspiracy [SH95, Hal95].

Constant [BC86, Enr78, Eub88, Gup98b, Jev92a, Meh74b, Ram89, Gup98a, Gla87a]. Constant-Weight [Jev92a]. Constantin [Bea90a, Bea90b, Liu91]. Constantinescu [Fra95a]. Constrained [ASPL03, BCH14, CM76, Fab80, GMS05, Jag93, Jag94, Kre64, Moo74c, PR70, Rus76, Tap76, Tap84b, Tap84a, Wol81, Wol82, AK90, Sch76c, Zlo85]. Constraint [Fel76, JLM^{+}17, LT13, Pec62, Pet73]. Constraints [Bla97b, Che91, DGLZ12, HSV95, JU76, Lew73, PdHM16, Ric67c, Sau76, Sch70a, Sez90, Bla97a, Bla82, Tad89]. Constructing [FH76, GGK96, GNZ20, HMT11, Kor76, Lam85, MKS86, Ang90, McC86]. Construction [Ath93b, Ath93a, Boa87, Cod74, Hue96, MW76b, Mey69b, Ven72, Sim89, VN73, Wil78b]. Constructive [BT69, Buc69, HA75, Paz76, Wou67]. Contact [BTKNZ00, Bes92, Kel80b, Shi93, Kee87]. Contain [Gil65a]. Containing [DB70, Eks60, JF71, MM06]. Contaminants [WPL98]. Contemporary [Ahl61, Boa82b, Dav91b]. Content [AKP^{+}91b]. Contests [GLY^{+}18]. Context [Aro78b]. Conti [Fin76b, Noh67]. Contiguous [Bar22]. Continua [HHOT93, Wil91b, Hut89, Hut91, Pra68, Wil91a]. Continuation [AG80, KP87, Sai91a, Sai91b, Was73]. Continued [Bla64, HT12, McC93a, McC93b, Mer64, Phi68, Phi71b, Sin69, Thr94b, McC91a, Thr94a]. Continuity [Dun68, Fin71]. Continuous [Abl75, Abl77b, Ahm75, Ang92b, Ars60, Art80, Bak91b, Bak92, Chu88a, Dun94, Ede90a, Ede90b, Fle69, GJRW24, HW89, Hol81, KH98, Kel86b, Kel90b, KTSB21, Mea88b, MR00, Mor75b, Mor85b, Pit93b, Ray69, Sta88, Und76, Vak93, VS95, Yan97b, Zit76, Yan97a, Has86, Pit93a, Rob94a, Rus90]. Continuous-Parameter [Fle69].

Continuous-Time

[Ang92b, KTSB21, Rob94a]. Continuously [BLT08]. Continuum [Ant83b, Ant83a, Edw90, Hau90, Hoc79b, Hor95b, Red89, Sch80a, Wol97, Hor83b, Obe77]. Continuum-Mechanics [Ant83a]. Contour [BEG23, Gol62, Mer84b, Mer84a, ZD78]. Contouring [Gra00]. Contrôle [Fal70a]. Contraction [Den67, San65, Wal76a]. Contraction-Mapping [San65]. Contribution [Hol87]. Contributions [Par92a, Har63, Man61a, Noh62]. Control [Ada91, Aok70, Arm76, Ath87, Bal78, BK95, Bau91, BDD^{+}95, BCH14, Bis78, Blo89, Blo91, BvJ09, BGF^{+}95, Bro76a, CJR18, Cas87b, Cas87a, Cas80, Ces75, Cha66, Che91, CGP21, CM78b, Chu90, CC90, Cla78b, Cla79b, Coh90b, Dat69, Dau90, Dun94, Ell80, Ewi77, Fle69, Fle79, Fle88, Fle90, FS78b, Fre75, Fri78b, Fri79a, GQJ96, GSYY22, Goo91b, GZ11, Gue88, Hag80, Hag82, Ham62a, HSV95, Har66b, Her67a, Her69, Hig62c, HM01, Hüb88, Hüb91, Joh91a, Joh91b, JS68, JH98, Kal63, Kal65, Kar60, Kha88, Kok84, Kor95b, KKY95, Kre66, Kre78, KG98, Kus72, Kus76, LLSN96, Lew83c, Lew91a, LM99, LmYS96, Lim23, MPRW02, Man87, Man88b, MO78, MT76, McG64, McG65a, McG65b, MR76, Med78, Mil83b, MB86, Mor82, Mor88d, Mor89c, Nic96, Nie66, Old62]. Control [Pai76, Pfl92, Pol73, Rei71a, Rei78b, Rox76, Rox90b, RBZ23, Rus95, Sac83b, Sac83a, Sag77, Sag81, SS16, Sch81a, Ser84a, Set77, Sin76, Sta75, Sto96, Tan93b, Tan93a, War68, Yak95, Yan90, YZZ04, You76, ZBK96, ZZ18, Zua05, Bal93, Bau89, Bus70, Coh90a, Col89, Dat67, Dat73, Ell85b, Fle89, Goo91b, Gre60a, Hau89, Jód96, KS94, Kre96, LW64, Lar78, Lin80, McC79, Mil83c, Mor68, Mor82, Neu94, Rox90a, Sag83, Ser84b, Sno76, Bal88a, Cul68, Fel90, Kha87, Yon95]. Controllability [FC99, Háj79, JU76, Lag97, LM99, Lio88, PAT^{+}80, Rus78, Und78, Chu89, Pow80]. Controllable [Qui87]. Controlled [Far93a, Far93b, Pal84a, Pal84b, Ell81, Fle83, Men94a, Oco87]. Controller [Che92a, Gua92]. Controlling [GMSW91]. Controls [Hau78, Hau81, Lef65, Sle89b, Win75]. Conundrums [Bol91c, Bol91a]. Convection [BJK24, Gar60, MK97, RS75, Shu09, SR75, Gar61, May90]. Convection-Diffusion [MK97]. Convective [Ber90a, Pra89a, WB76]. Convenient [MB64]. Convergence [AW76, Bag95, BES05, Bog75, Bre67, BRP88, Dax90, DG14, FF78, FP96, Fra75, Gay78, Ger94, Gil62a, Giv78a, GK78, Gra86b, Han94a, Hoc67, Jor86b, Jor86c, Kri77, Kui06, Mel97, MP06, MNS02, OT11, OW95b, PH78, Pow86, Ros95, SW79, Sed02, Ser80a, SS62, SS05, Ste70b, WS13, Wil85b, Wol69, Wol71, SW80, Par88, Sam76c, Ano68a, Han94b, Pin69]. Convergent [GBC^{+}90, Gib90, Kel84d, Sha89a, TW14, Wat86c]. Conversion [All75, All76, Les76]. Conversions [Fre14]. Converting [Dar97]. Convex [Ame86a, BB96, BLT08, BI70b, BGL01, CPW12, CTW21, CKCH23, Ell85a, Ewi77, FGP^{+}22, Fle65, JP89, KL75, Kot65, Lew73, Luc10, Mar93, Mar94, Mur71, Pan93b, Ren94a, SW79, Sal85, Wal23, Wil67d, Ye94, ZKP17, vdM68, SW80, Cla78c, Kle71, Kle76, Pan93a, Rya63]. Convexity [Gud77, Gud78b, JvD90, JvD91, JW83, Kol91, Kol92, Pon67, Rat92b, Dat80, Kle72, Maw86, Rat92a]. Convolution [Aga93, Boe68, Böh88c, Boh88a, Bus64, CW94, Fen68, Gol76a, JS20, Kel92a, Koh85, Kuc87, Lóp08, MNMF63, Reu67, Tur68, Wei60a, Kel92b]. Convolutions [Moh66, AL98, Han98, Knu98]. Convolved [NS77]. Conway [GC69, GGG69, Sch86a, Bos99]. Cook [Che89]. Cooking [Kla61]. Coolers [Sta75].

Coolidge [Bra91a, Bra91b]. Cooling

[Hey63, Way82]. Coonce [Cas72a]. Cooper [New94a, Win62, Kro66]. Cooperative [Dri91, Mie17]. Coordinate [FR16, KL75, MT76, SG73]. Coordinates [KN98, SB98a]. Coordinating [Dau90]. Coordination [CB09]. Copositive [HUS10]. Coppersmith [Pan86]. Cordes [Tay96]. Corduneanu [Noh76]. Core [MT84, RPFM17, SES75]. Core-Periphery [RPFM17]. Corliss [Fat93]. Cornelis [Hol90a, Hol90b]. Cornwell [Mit87]. Correcting [Mat80, Mat94b, Mat94a, Ste76]. Correction [Ama77a, Bal77, Ben76a, Gud78a, Haw95, Hni92, Kus77, Mat76a, May84a, Miu77a, Mor86b, Ort96b, Sei87, Smi88a, SFS88, Sto91, Sul93a, Sul93b, Wil80a, Wil80b, Xu92, Ort96a]. Corrections [AR80, Wol71]. Corrector [Tho76b]. Correlation [HSŠ16, SK20]. Correlations [AMS^{+}23]. Correspondence [Ber85b]. Corresponding [CD75, Cul78]. Corrigenda [Ano78g]. Corrington [Far69]. Cosine [Str99]. Cosines [Gar69, Fis71a]. Cosmological [Ano93a, Gro93c]. Cosmology [Gla94a]. COSRIMS [Ham70]. Cost [GH22, Ham88, Zit76]. Costa [Wol94a]. Costs [CD94, Guh66, Had65, Mor69]. Cottle [Mor93, Wet75]. Couette [Sch95, CIS95]. Coulomb [Ike76]. Council [Ano75e]. Countable [Pit74]. Counter [Mey86a]. Counter-Examples [Mey86a]. Counterexamples [Har83, Gol65b]. Counterpart [Bax85]. Counterparts [PdHM16]. Counting [ABG^{+}95, Kra68, McK95, Row66, Tak83]. Coupled [BT20, EFMM13, Kno83, NRS19, PSSW06, Pea81, Rog69a, Wil64a, WE19]. Coupling [Ede90a, Ede90b, Kal93a, Kal93b, Say13, SBM22]. Coupon [MW06]. Courant [Dia64]. Course [Beu85, Boa83, Bur93, CAiJS23, Cum96, Ell94, FLC96, Gue83a, Gue83d, HS95, Hor93, IL97, Kar83, Kaw67, Kes85, Kha88, Lag89, Lee82a, Mat88, Mir85, NN96, Owe85, Par69, Por92, San95, Sch86a, Sch75a, Spi82, SKK20, Wax91a, Wax91b, Wei77, Gut97]. Courses [AW19]. Courtney [Gue84, Koç94]. Courtois [Den78]. Covariance [BGL01, DMGL22]. Covariances [Cha95b, DEPC95]. Covariant [SG73]. Covariates [EW19]. Cover [CHW12, Kie94]. Coverage [SS61]. Covering [FK62, GE63, Law65, Los84a, Mei83a, Mei83b, New82b, SA75, Sch85c, Vij83, Cox65]. Cowling [Núñ96]. Cox [Cha95d, Pin85b, Rei77c, Som93, Tab95, Tab97]. Coxeter [Zas65]. CR [SM22]. Crack [Bue71]. Cracking [BK78]. Craft [Hig01b]. Crafton [Sto62b]. Craig [Bla81, Che92a, Gua92]. Craik [Ben87]. Cramér [Kaw68]. Crank [Ell88]. Crapo [Bru72]. Craven [Hag80]. Crawford [For75]. CRC [Hei97, ZK96, Kel96]. Creature [Sco92]. Credibility [BAH^{+}80, SK79, BS80]. Creep [Cho78]. Creeping [Sac67]. Criminale [Kan68]. Crispin [Grc18]. Criteria [Asn75, Ike76, Mro91, SWM^{+}07, Mas87]. Criterion [Kra66, MM98]. Critical [BSS05, Cel78, GM95, Laz91b, Lev90b, Maw95, Rei77a, Rei78c, Laz91a]. Croft [Kla92b]. Cronin [Lae82, vV65]. Crop [Ant91]. Cross [Cer70, Cer71, Dre89, Fre83, GHW76, Mel64, Mor76]. Cross-Section [Mel64]. Crossings [LC65, Wit90, Yak94]. Crouch [Blo89]. Crowder [Sin93a]. Crowds [BD11]. Cruse [Els89]. Cryptography [Bla88f, DDN03, Bla88e, Gha89, Lag89, Bla83]. Cryptology [Buc91a, Buc91b, Kob10]. Cryptosystems [KM04, Bla84]. Crystal [Gla93d, LG74, LG75, Gla94c, Lar94]. Crystalline [Sch93b, Sch93a]. Crystallography [Mul75, Wu72]. Crystals [Gla95e, Ism96a, Ism96b]. Csaki [Med78].

Cubature [Ahl62, Tre17, Str82]. Cube [Lei78a, Lei79, MW61, RKDD14, Sin79a, Xu96c, Xu97, Ren80a, She63b]. Cubic [Dav75a, Dav75b, Jes93b, Jes93a, Kar22, MR92, PS74, Sei78, Sil66, Sil67]. Cubics [Dub97b, Dub97a]. Cuboctahedral [Pow93, Pow94]. Cullum [Cho87]. Cultural [FS78a]. Cumulative [Amo78b, Sch61, Tay86]. Currency [TE96]. Current [Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Som78]. Curriculum [Kir12, Lax67, Mon67, Ral86]. Curse [SKP19]. Curtain [Gue79, Mor96]. Curtis [Kee88b, Loe95, FW66, Tre08]. Curvature [BT85]. Curve [Ber92a, CS06, Jag87a, LC65, Nas95, Nie76, Pec62, PMK96, RM86, Sto73, TLN95, Win93, Zha13, Han92, Win96b, Win96a, Pra89b]. Curved [Bou91b, EMB76, Kap90a, Kap90c, Lud77]. Curves [DeR91b, Dub97b, LF73, Ray69, Roc92a, Roc92b, Dub97a, DeR91a]. Curvilinear [MT76]. Cushion [Mul79, Pet78]. Cut [Amo64, Bol91b, Bol91d, MS65, PR91, Tri62a]. Cut-Out [MS65, Tri62a, Amo64]. Cutland [Smi82]. Cuts [Ste04, VBK22]. Cutting [BJK11, Gol61, JRZ97]. Cuyt [GM85]. Cvetković [Bei81]. Cybernetics [Ano64i]. Cycle [AW19, Fre82, New65a]. Cycles [BP94]. Cyclic [Ano90a, DGS05, DG96, GDS05, Kla90b, Kla95a, SAT22, Swa77, Gro96c]. Cylinder [Tri62a, Amo64, MS65]. Cylinders [Han89, Mor76]. Cylindrical [BK88, Hol89b, MT84]. Cylindrically [Kri76].

D

[Abl84b, Aga94, Ame86a, Ame86b, Ame95, And95, Ano73a, Arb76, Aro63, Art83a, Ask81, Asn78, Asp84a, Bai87, Bak91a, Bak91b, Bak92, Bal93, Ban84b, Ban84a, Bar86, Bau88, Bei81, Bel90b, Ben87, Ber81, Bla82, Bla85b, Boa82b, Bor73, Bor79, Bou66a, Boy70, Bre64a, BBLS92, BOS73c, Bry93, Bur81a, But89, Caf92, Car71b, Car87, Cha83b, Cha81b, Che89, Cho66, Chu82, Cot94, Cra84, Dan91, Dev92, Dia64, Dij90, Dix91, Dum65, Ell94, Eps59b, Erd77, Erm89, Eva81, Fif79, FR63, Fre80, Fre83, Gil78, Gil95, Gla86b, Gla86c, Gli91a, Gou68, Gra70, Gre60a, Gro93d, Gru88, Grz83, Gue83a, Gue83d, Gut90, Hag89a, Hag80, Hai77, Ham65b, Han79, Har70, Har72a, Hea81, Hei93a, Hei89, Hel97, Her92a, Hey82, Hir73, HT81, Hüb91, Hud95, Hut89, Isa89b, Jag76a, Jag79a]. D [Jag87b, Joh81, Joh83b, JSR82a, JA80, Jor84c, Jun65, Jur95, Kap85b, Kap72, Kap78, Kar60, Kea80, Kha71, Kla84f, Kla97, Kno81, KL87, KF62, Koo77, Kow86, Lan86, Lat69, Lin92, LFR80, LSB67, Lum85, Mal80, Mal85, Mal87, Man87, McG93, McK95, ML85, MBCG83a, Mei83a, Mer85, Mey82b, Mey82a, Mey91a, Mot88, Naa01, Nas95, Nel74, New76c, New82a, Noh77, Olm77, Pal84a, Pap63a, Par92a, Per68, Pin85b, Pip92, Por72, Por80, Rei77c, Rei64, Ren80a, Rob95b, Rob74, Rob76, Ros78, Rox76, Rud61, Saa96, Sag77, Sal63d, Sau62, Sau79, Sca59, SNS66, Sch94, Sch60a, Sch65b, Sco84b, Sco84a, Sei93, Ser80b, Set63, Sha95c, She63a, She91, Shi93, Sib95, Sie90, Sil78, Sin59, SS70, Sle94, Sle87, Som93, Spi82, Sta66, Ste70a, Sto62a, Str70, Str62c]. D [Tab97, Var85, Vog93, War92, Wat63, Wen68, Whe74, Whi84a, Wil84a, Wil65a, Wil69, Wil77c, Zad67, Zit63, vL86b, Her92b, IWMW21, LH19, MMH04, TH96]. D. [Abl84a, Cha63, Kni84a]. D.B [Erd92b]. Dacorogna [Bou91a]. Dacunha [Gut88]. Dacunha-Castelle [Gut88]. DAEs [SRK99]. Dagpunar [Kel90c, Kel90d]. Dahlard [Pal84b]. Dahlquist [Jac76]. Dai [Lew91a]. Daiuto [Wan91]. Dale [Kle76, vL86a]. Dalembert [TP78, Ste81c]. Daley [Gut90]. Damage [Tay86].

Damjanović [Ano87a]. Damme [Shu90a, Shu90b]. Damped [EG98, Lan64, RLT65, Str64]. Damper [Pet97b, Pet97a]. Damping [Sle76, Tew77]. Dan [Hes83, Kri91]. Daners [Aiz94a, Aiz94b]. Daniel [Aiz94b, Art83b, Art85b, Asp84b, BW65, Bur94a, Caf92, Dan91, Fif79, Hin94, Hul71, Kel96, Klo93, Lat90a, Nas77b, New87, Odg86, Par87, Pit93a, Sch87, Wis60, Wri60, Aiz94a, Lat90b, Pit93b]. Danny [Fis88d, FB78, Hig00]. Danskin [Saa68]. DaPrato [Eva85]. Darboux [CS63]. Dario [Ebe96]. DARMA [Ram89]. Daróczy [Cam77]. Darrell [Fab87]. Data [BNT10, BP76b, BGL17, BF16, Ber64, Bor62, Car93, Car65b, Che92b, Cho78, CSN09, Dix91, Ell92b, EGJK04, FP21, Gra93a, Gra93b, Han76, Hor78, Kar60, Kli78, Lin92, Luc18, MYDF08, MS75, McL86a, NO60, NGTB15, RBC^{+}07, Sav18, Sch85a, Sch91b, Sei78, Sto60, VB05, War98, ZKP17, Che92c, Ell92a, Fra86b, Mon60, Orb18, Pin85b, Sch91a, Sin93a, Sin79b, Str89b, You74, Law18]. Data-Driven [Luc18]. Databases [TH96]. Daubechies [Hei93b]. Daut [Tul90]. Dautray [Ker92]. Dave [Fri64]. Davenport [SAH96, Haj96]. David [Byr65, Caf92, Cor92b, Dav91b, Dol91a, Dol91b, Dwa72, Fal70b, Fet68b, Gar93, Gho87, Gla87a, Gup91a, Har82, Hoc92, Joh73, Kap72, Kaz90a, Kes85, Kni84b, Knu77a, Koc92, Lee77, Len80, Mac91b, Mey86b, Mey89c, Mor82, OP88, Pin72b, Pit74, Por83, Por81, Ran96b, Roc92a, RB90, Sag81, Sop82, Spo88, Vog91a, Whi84b, Wil84b, YY96, Gup91b, Kaz90b, Küç92, Mac91d, Roc92b, Vog91b]. Davidon [Dav75b, PS74, Dav75a]. Davidson [Gra86a, SV00]. Davies [Her67b, Koh86, Sop59]. Davis [Cha82b, Dat90, Gre65, Jag64, Len95, Lin80, Mea88b, Ste68a, Str76b, Str86c, Sub90]. Davydov [Sco88]. Dawe [Mal87]. Day [Man90a, Ros87, Sto62a, WJV79, Man90b].

Dayanand [Gei89]. Daykin [MBCG83a]. Deak [Mod92]. Dean [Fab80, Kar78]. DeAngelis [Kee88b]. Death [Ano77e, Gia63]. Deaths [Ano69b, Ano70c, Ano70d, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano74a, Ano75f, Ano76f, Ano76g]. Debnath [Gib85, Kau86, Syn93a, Syn93b]. Dec [Ano87b]. Decay [BBR13, HW76, Núñ96]. Decaying [DL76]. December [Ano70j, Ano71p, Ano85k, Ano86l, Ano87l, Ano88j, Ano89m, Ano90m, Ano91l, Ano92i, Ano93j, Ano94g, Ano95q, Ano96r]. Decentralization [Sez90]. Deciding [LKGK07]. Decision [Aok70, Fei96, Fox71, Fre86c, KM12, Mal81, Mal83, Man61b, Ngu76, PF96, Rei87, Ser80b, Tay68, Bas88, Har67d]. Decision-Making [Bas88]. Decisions [Ber62, Sca59]. Deck [Kel95]. Decomposability. [Den78]. Decomposition [CDS01, DD02, Gol75c, HKD20, KB76b, KF76, Luc18, MMH04, Owh17, Pap85, Ros69, Shr76, SBGP98, Ste93c, Val67, Xu92, XZ98, DGH^{+}18, Mur97, Kun93]. Decompositions [HMT11, HL76, KB09, NF16, Neu97b]. Deconvolution [CW94, McG93, Mor86a]. Decoupling [Mat85, Rus90, Smi87]. Decrypted [Kob10]. Dedication [Dav94b]. Deep [HH19, LMMK21, YL75]. Deep-Water [YL75]. DeepXDE [LMMK21]. Defect [Sul93a, Sul93b]. Defective [Gre60b]. Defense [CR67, HW77, KS77]. Definite [Amo78a, Ano63a, Ano71a, Ano91m, Ano92a, Bad61, Ber84c, BG95b, Bou86, Bro63, Byr65, CQRZ96, CMR73, CW70, DB70, Duf84, Far65, Fet68b, Fet68a, FH70, Gla67, Gla77b, Gla78b, Gla78c, Gla84a, Gla84c, Gla85a, Gla85d, Gla87b, Gla87c, GG88, Gla89a, Gla93d, GBdD94, Gla94c, Gla95b, Gra79, GSL79, GJ90, Gro90, Gro91a, Gro88b, GAB^{+}88, Hen82, Her90, HR85, Jag88a, JGA^{+}88, JH92, Jer82, Jin84, Jin85, Joh83g, JG70, JF71, Kle63, Lar94, LMS68, LJ83, Los83c, Mor84, Nut85, Ort94, Ort95, Ost91, RSS65, Rot62, Sch72c, Sch63c, Sch77c, Sho64, Sho66a, Sti87, Str65a, TS65, VG78, VJ83, Wag92, WGG^{+}92, ZM22, Gla97a, Gla98]. Definiteness [HSŠ16]. Definition [Fel76]. Deformable [MT64]. Deformation [BvdGH95, Sch89b, Hav95]. Deformations [Hor95a]. Deforming [GHU22]. Degeneracy [Bom02, Ste76]. Degenerate [Her78a, Kus76]. Degree [FG19, FGA97, FLNU18, GL89, Gib87b, Gib87a, Hol81, Kel88b, LCK^{+}17, Vei86, Zel77, vV65]. Degree-Based [FG19]. Deif [Nie89a]. Deimling [Bre94a, Bre94b]. DeKerf [Sco91a, Sco91b]. Delay [Asn78, Bul67, Dri81, HY71, HT78, JvD90, JvD91, Liu94, Liu95, MC21, Mil90a, Mil90b, NH78, PEGK03, Sei93, Sei93, Cor92a]. Delay-Differential [Bul67]. Delay-Induced [MC21]. Delayed [ZM22]. Delays [Dri68]. Delfour [Rus95]. Delivery [SFOM99]. Delta [Lam69, Lam70]. Delves [Gro87, Hag83]. Demagnetizing [Jor78a, Jos76]. Demand [Ell85a, Mor69]. Dembo [Din94a, Din94b]. Demography [Lud87, Pug86, Gin85]. Demonstration [GL98]. Demyanov [Nas77b]. Dendritic [Bel90a]. Deng [Hol91b]. Denn [Ram90a]. Dennis [Hal95, Han64, Han86b, Wat91b, Wer89]. Denoising [BCM10]. Dense [DGK84, EGJK04, GPS90, MS76]. Densities [LSY16, RW84]. Density [BL09, Goo82, GO67, Kha71, NW70, RO66, Sch64]. Denumerable [Ath93b, Ath93a]. Departed [Jef18]. Dependent [Ban87, Bar65, Del95, Den76, Fri79b, GM81a, GM81b, Giv63, Kel82, KW18, LL76a, Naz86, Smi97, Sül97, Tad87, VB75, ZD95]. Depletion [CD94, Ngu76]. Depth [Mil91b, She75, Mil91a]. Derald [BS67]. Derby [Zei72]. Derek [Lig86, Rob85].

Derivation [Ang92a, AR63, Cho78, GG99, JS20, PLC96, SV98, Vog87, Was74]. Derivations [SK20]. Derivative [Gla78a, GN78, Nås77a, Pea65, Rez62, Rub78, SK76b, Str62b, WS13]. Derivative-Expansion [Rub78]. Derivative-Free [WS13]. Derivatives [GMP05, Gol66a, Gri88, HF78, Ike76, Jok66b, LOT76a, Lev90a, Lev91, LOT76b, Mat70, ST98, Ste59]. Derived [PW08, Sta76]. Dérivées [Fal70a]. Deriving [HS81, MA04, SB98b]. Derogatory [Gre60b]. Dertouzos [Mul66b]. Descent [CF80, FR16, Lem88, She09]. Descents [Olv70]. Description [KD10, McL86b]. Descriptions [FGP^{+}22]. Descriptive [Ber85b]. Design [BP76b, BTKNZ00, BPP95, But78, CFHP10, Che92a, Che92b, Che92c, Dan66, DeR91a, DeR91b, Edw90, Gel79, GP92, Gor75, Gra91, Gua92, GAM93, Hau90, HLB95, Jow72, Kal65, Kla87, KDJM88, Laf93, LL94, McG65b, MMB^{+}07b, Mro91, NH78, New60, New64c, Nie76, Old62, Pet78, Pet97b, Pol87, Raz92a, Raz92b, Ruk93a, Ruk93b, Spi84a, Str86a, Tad89, Tan93b, Tan93a, TWW65, VS95, Wil76a, WS23, MMB^{+}07a, Pet97a, Bos95, Kla88a, Lin92, Man88b, McN59, Pet95, Rei69, Sag81, Spi84b, Koh88]. Designing [De 00, GRK81, Lau03, Ren80b, GK81]. Designs [Mil81b, Hei97, Hei89, Jow76, Not94]. Desoer [Kap64]. deSwart [Jin91]. Detailed [Dau90]. d’Été [Dur87]. Detecting [AMR04, Ste76]. Detection [Arn70, BW76, Gri77, LS63, VC06, ZWGB16]. Detector [Bur78b]. Determinant [BB79, Bha98, Bos73a, Boy80, Boy82, Boy83, CL71, Deu81a, Far84, GG73, GM92, HL76, Har62, Kel89, KSS80b, Lew83b, Los74a, Olk72a, Pra69, Ren90, RAH^{+}90, RP96, RLG81, RL81, Sch71d, Sch72a, Sch73b, Sch73f, Sch74, She65, Sle78, Sle79b, Ven74, Ven75, Ven72, WGM93, WG93, Hor97b, Pea97, Hou70b, Hou71a, Jag72c, Jag79a, Jag83, JA80, Los75, She66b, VN73, Deu81c]. Determinantal [Gib87b, Gib87a, Vei86]. Determinants [Ars62b, Got60, LM78, LMGC79, Sch69a, Sch71c, Sch72b, Sch73c, Sch73d, Sha78, SJ79, GC79, Jag79b]. Determination [EC76, Nar73, Tor62, WPTZ01, Bre74]. Determine [Mos92]. Determining [FR66, FR83, LRV76, Qui87, RRW07, Yan97b, Zie70, Yan97a]. Determinism [Sey93a, Sey93b]. Deterministic [Ant92a, Ant92b, BvdD78, Beu89, BdCPT12, Guc87, Kap76, Ken93, Mac91b, Mac91d, MSO96, Ott96, Pal94, Rob92a, Rob92b, Tud90, Wet74b, ZWY14, Koç89, Pug86]. Deterrence [KS77]. Dettman [Ans64b]. Deutsch [JB88, Kei64, Nev82]. Devaney [Hol87, Stu90, Are93]. Developing [Mil76]. Development [Che76, Col94a, Geo71, Kel88b, Par64, Vas94, Wan79, Ypm95]. Developments [Boo94, Cas82, Che78, CCM00, CCM01, FGN13, Jür86, KP06, Luc71, Mos86, ML80, Pon83, Lum85]. Deviating [Dri75, Wou68b]. Deviations [Din94a, Din94b, Eis86a, Var85]. Device [Pim87]. Devices [FT11, GKLU09, Moc85]. Devroye [Ste87]. Dezin [Lev88b]. Dhaliwal [Kno83]. Dhall [Gal95]. Dhatt [Mea86]. Dhir [JF71]. Diaconis [Ano99b, The99a]. Diagonal [Cle59, FG04, Hag83, Hea81, HS88, PC89, PC90b, Ros80b]. Diagonalization [SBR08]. Diagonals [Kla70b, Kla71c, MKR76]. Diagonizable [Nik78]. Diagram [Are93]. Diagrams [Ric67c]. Diameter [RKDD14]. Diamond [Ham73a]. Dice [KL80]. Dichotomy [Swa78]. Dickey [Ant78, Kup94]. Dickinson [Gup82, Noe72]. Dickson [TT76]. Dictated [LL76a]. Dictionary [How94]. Didier [Gog85, Gut88, McA92]. Diederich [CS64]. Diego [Bec94, BC96b]. Dieter [Bos95, Cav90, Hey85, Kee88a, O’B95]. Dieting [MBR98b, MBR98a]. Dietmar [Goo91a]. Dietrich [Deu90]. Dieudonné [Her77]. Diewert [Bor90]. Difference [Aga95a, Aga95b, Aro69, Ber65b, Bro59b, But89, Car74, Car75b, DL92, For98b, Gev82, GAT75, Gup98b, Har61b, HM78, Kel60a, Kel60b, KM85, KLA95c, Kus76, Lei78b, Mea88b, MA95, MM06, Por80, Sch61, Sch80c, Sch81c, Sik92, SM14, Stu89, Sül97, Tad87, Tho69, Tre82, Wak86, WJ75, Wen61, Wil64c, Zha97, Zua05, For98a, Gro98d, Gup98a, Ric98, But92, Mea88a, Mil81a, Wen68]. Difference-Differential [Bro59b, Car74, Car75b]. Differences [BPV97b, Ham69, Kel60a, Ng96, BPV97a, CS70, Har61a, Str62b]. Differencing [DJLQ21, HLvL83, Hir75]. Different [Bom02, Gev82]. Differentiable [CBS76, Kaz90a, Kaz90b]. Differential [Aar07, Abd65, Abd66, Abd84, Abd91, Abl84b, Abl84a, Aga95b, Ale90, All78, Ame85, Ame86b, Ame95, Amo65, And77, Ano93d, Ans66, Ant82, Ant63, Ari96, AB23, Asn78, Ast87, BNT10, Bai87, Bar78, Bar70, Ber91a, Ber78, BR76, Bob66, Boj93, BR18, Bra65, Bra63, Bre94a, Bre94b, Bro59b, Bro76b, Bul67, But92, Cam91b, Cam91c, Cam93, Car74, Car75b, Car77, CCCM06, CSW^{+}08, Ces75, Cha94a, Cha94b, Cha67a, Cha84, Cha81a, Cha81b, Che96b, Che76, Che78, Chi90, Cho66, Chu94b, Chu94a, Çin93a, Cin93b, Col91, Col94b, Con85, CK86, Cor92a, Cro85, DO98, Dan93, Dat78, De 84, DTV14, Del95, Det85, Deu85, DL92, Dor74, Dri68, Dri75, Eas91, Enr89, EKZA96, EW75, EC76, FGN13, Fer79, FW82, Fin76b, Fle62, Fle77a, Fle77b, Fle82]. Differential [For92, FGD^{+}92, Fre80, Fri80, GVL97, Gat64, Gat88, Gea74, Gea80, Gea81, GR95, Geo84a, Geo84b, Gil78, Gil90a, Gil90b, Giv78b, GC78, Gla95a, Gli91b, Gos83a, Gos83b, GS22, GL98, Gra94a, Gra94b, Gre62a, Gru83, Gue85, Gue95, Gup89, Hac75, HY71, HL63, Han75b, Han79, Han81a, Han85a, Han87, HT78, Har70, Has92b, Her84b, Her84a, Her67b, Her71, Her73, Hig01a, Hin89, Hin94, Hoc65, Hor97a, How83, Hul67, Hul71, IM88, Isa89b, IL97, Jac97, Jag64, JU76, JR08, Kah66, Kan88, Kap85a, Kap72, Kap77, Kaz83, Kea80, KSS80a, Kel83, KMH66, Kla62b, Kla66, KN67, Kla76a, Kla97, KPGR95, Kol65, Kor95a, Kor95c, KKS82, Kra80b, Kra80a, Kum88, Kus74, Kus93, Lam86, Lat90a, Lat90b, Lax69, Laz82, LP75, Lee82a, Lee86a, Lei76, Ler63, Lev62, Lev67a]. Differential [Lev88b, LG89, Lew89, LR80, Lew86, LBD^{+}86, Lie95, Liu91, LM81, LLR^{+}81, LA95, Lov79a, LMMK21, Mac76b, Mah75, Mat69b, MG76, MMS98, MS72, McG82, McL86a, Mel91, Mel92b, Mer89, Mey72, Mey77b, Mey78, MR78, Mey93b, MA95, Mie17, Mit94, Na67, NH78, New76f, New78, NS24, Noh67, OL98, O’M10f, Olm77, Olv90, Olv91b, Olv93, OV85, Osb85, PY98, Pal84a, Pap86, Pay68, PD76, Per68, Pic87a, Pin87b, Por72, Por80, Por85, Ral65b, Rao94, Red76b, RRK95, Ric76b, Rod81, RL95, Rog83b, RSTK97, Ros79, Rou73, Rox92, Rub69, Rub78, Rub77b, RRG95, Rue72b, Rue73c, Rue74b, Rus78, Sch77a, Sch63a, Sch73a, Sch82b, Sch80b, Sch60c, Sch88, Sch85b, Sco59, Sed96, Sei65, Sei93, Sha75, SG79, SG95, Sha97a, SS62, SS98, SR98, SA78, Sle92]. Differential [Sle79a, Smi88b, SBGP98, Smo91a, Smo91b, SLF66, Ste61b, Ste93a, SY69, Syn83, TK97, Tho59, Tri95, TS90, Und78, Ung77, VC06, Vel63, War68, Was94a, Was94b, Web85, Wei91, Wil85a, Wil86a, Wil72, Wil73, Wim67, Won64b, Wou68a, Wou68b, YT95, Yan65, ZD95, Zie79, ZMPK24, Ano97j, Che96a, IM89, SWD76, Beb68, Ber90b, Blu94a, Cha70a, Coo67a, Dia59, Dri63, Fle72, Gue84, Gue89, Gur62b, Gus76, Has92a, Her80a, Her80b, Hoc79a, Hoc86, Koç94, Kor65, Lae82, Ler64, Olv91a, Pal84b, Uri92, Wil67a].

Differential- [Wil67a]. Differential-Algebraic [Cam91b, Cam91c]. Differential-Delay [HY71, HT78]. Differential-Equation [FGD^{+}92, KSS80a, Kra80a, Lew89, LBD^{+}86, MR78]. Differential-Geometric [Pap86, VC06]. Differentialgleichungen [Sch71a]. Differentialoperatoren [Gil65b]. Differentiation [Cos93, Fat93, Nei10, SS16, SG72, Vil74, Wol94a]. Diffraction [Ble81, Fel70, Fel72, Kel67, Lud70, Ram93]. Diffuse [BF16]. Diffusing [Kaz80a]. Diffusion [Aeb94b, Aeb94a, Bal78, BJK24, Ber90a, Bil93b, Bil94b, BdD94, CT75, Coh76, DGLZ12, EERG20, Fif86, Fle83, Fle90, Gar60, Gau62, GSYY22, GNP95, Jer91b, Jer95, KL08, Kaz89, Kni84a, KW18, LW13, LW17, MK97, NPW95, Nie98, OH76, OK94, PSV12, PS00, Pin72b, Var67, WMZ94, ZC74, Gro98b, Rul98, Gar61, Kni84b, Chu92, Jer91a, Lu92]. Diffusion-Convection [Gar60, Gar61]. Diffusional [BRP88]. Diffusions [Fle88, Por81]. Diffusive [Ame88, EHS17, RS75, SR75]. Digit [Hig05, TL02]. Digital [And78, Arm76, Cam91a, Fla66, Hul61b, Kal65, Kar63, KM86, Lin79c, McG62b, Mul76, RA63, Chu88b, Gos83a, Hul61a, Sin59]. Digits [GF76, GBF78]. Digrams [Gil65a]. Digraph [CH61]. Dijk [Oco87]. dilation [Str89a]. Dilemma [Mar89, GO09]. Dilute [Bar95, CIPB95]. Dimension [Fis93, GS04, TFL23, Kle72]. Dimensional [AG14, BK95, Bec74, BDD^{+}95, Ber94, Ber90a, BF16, Blu94b, Boe96b, BB83b, BK78, Cel78, Coo67b, Cul77, Cur75, CZM96, EK59, Fin76a, Gas76a, Gla95d, GP79, HM78, KW76, KR74, Kor95b, Kos79a, Kos79b, Kra67, LmYS96, Mac92, Mas92, MTGF78, Mei19, Mey76a, Mey77a, MP20, MM10, Mor85a, Mor96, Mor91b, Mun61, New76f, New78, Pet73, PZ81, Quo75, RB23, Ren80a, Rus95, Sag81, Sch60b, Sch75b, Sei76, Sin79a, Sit92, Sto96, Stu64, Thu01, VB05, VN76b, VDC22, Win90, YF78, ZD78, ZWGB16, ZKP17, Mor86c, dMvSB95, Gas76b, Hol66, Kla63a, Kno81, Mor91a, Bru95]. Dimensionality [LS22, SKP19]. Dimensions [Arc75, FR16, Fis88b, Gue61, KS84, Mar21, Pre95a, Roa78, SW79, San60, Urb93b, VS95, Yan97b, ZW94, SW80, Yan97a, Pré95b, Urb93a]. Dimitri [Beu89, Cha87a, Ell80, Lar78, Zlo85]. Dimitris [Cul66, Cul68]. Dinca [Dav75c]. Dines [Mor85b]. Dingjun [Hoc94]. Dingle [Was74]. Dinits [GG91]. Dinitz [Hei97, Poh78]. Dinklebach [Gas70]. Dio [Fro60]. Diophantine [Los74b, Kla73]. Dipole [JM78]. DiPrima [Han86a]. Diprotic [Rei83b]. Dirac [Lam69, Lam70]. Direct [Bre68, CM76, Dor70, Dor73, Dor75, Hir78, JP89, Kap65, Kel17, KLT03, KLT04, Roa78, Sem59, SES75, VC06, Wou64b, dEJL07, Bou91a, CZ89, Dei92, Eis86b]. Directed [Che67a, De 20]. Direction [Den76]. Directional [You74]. Directionality [AMS^{+}23]. Directions [Zie79]. Directly [CC75]. Directors [AN78]. Dirichlet [FOTG96, Lei78b, McC03, Siv69, Get96]. Dirk [Hig05]. Disc [LFK78, Lew78a]. Discharge [Ber60]. Discipline [Gor63]. Disciplines [Ask80]. Discontinuities [Kel64]. Discontinuity [ZWGB16]. Discontinuous [Ber90b, RFS03, Bus70]. Discovery [Par86]. Discret [DD74]. Discrete [Abl75, Abl77b, ADPR04, Art80, Bar67, BJK24, Bec74, Ber84e, Ber84d, Boa87, BvJ09, BK87, Cho76a, Chu90, CC90, CTW21, Con96, DY11, DN59, Dor70, Dor73, Dor75, Ell80, Fei96, Fel76, FSY65, Gas76b, Gas76a, Ger62, GYS95, GJRW24, HIC96, Has86, HW89, HS95, Kel86b, Kel90b, Kel92a, Kel92b, Kla84b, Lai91, LT98, Man87, McC83, Mor78, Mor75b, PI88, PF96, Rin74, RRW07, San95, Saw85, Sed02, Sha95c, SS62, Sho99, Sta88, Str99, Swa77, TFL23, Tze91, Vak93, Var76, Var79, Zas65, Ham87, Han92, Gre62a, Kla84i, Luc86, Per76, Thr79, Zim90]. Discrete-Event [GYS95, Sha95c]. Discrete-Time [Lai91, McC83, Mor78, Tze91]. Discretization [AG14, Gea74, GST01, TD20, Oco87, Gea85]. Discretizations [AT17]. Discretized [Tin76]. Discriminant [Gup96, HG96]. discrimination [Gro98d, Ric98]. Discussion [Dev92, JK72, Tuk67]. Diseases [Het00, Yor75]. Disjoint [Jev92a]. Disjunctive [Bla82, Cre97]. Disk [DL76, Gee78, GK91, IWMW21, Kol91, Kol92, Men94b, Mos92, Wil78c, Abb92, Blo95]. Disks [Mel12, WC77]. Dislocation [HHOT93]. Disney [Hey89]. Dispersal [LPL05, PLL05]. Dispersion [Fel70, New74, Pro65b, Sto11]. Dispersive [EHS17, JK72]. Displacement [BT19, KL76, KS95, Mal87]. Displayed [Mea86]. Dissipation [BRMR08, JH92, OH76, PS00, Ano92a]. Dissipation-Induced [BRMR08]. Dissipative [EFN^{+}96, Jol96, Liu91, MSO96, Ott96, PvySF23, Rhe85, Rob92a, Rob92b, Stu89]. Distance [Dal76, Ham77, Lei78a, Lei79, LLMM14, Ose75, Sen89, She67, WGG^{+}90, DV76, Wat90]. Distances [CT76a, LFK78, WC77]. Distant [SR94]. Distinct [Jev95b, Sch81b]. Distortion [Gil74]. Distortionless [Mor88c]. Distributed [BPJ95, CB09, Jef95, Kow88, Lio88, Lud91, MR76, OT11, Sac83b, Sac83a, Sle89b, TH96, Wal93b]. Distribution [Amo78b, App87, BBN98, Bla87, Boe88b, CS84, Cla67, Coh75a, EF63, EF65, Eva88, Geo95, Hey63, KHS63, Kri67, LeV77a, MK78, Mil69a, Nag96, Sch61, SN96, Smi88c, Spi86, SS78, Ste67a, Tre70, VB05, WW77, WW78, Wei63a, dB65b, EB88, JS71, Mal88b, Mei78a, Rei71c, Sau62, Sau78, Lei66]. Distributional [EKW96, Wal92, Wal93c, Wim96]. Distributionally [PdHM16]. Distributions [Ars60, Ask75b, Bec74, Bel63d, Cho76a, CSN09, Col94b, DMGL22, Gas76a, Hey97, Hud95, JMH95, Kim65, KTSB21, Mar83b, Mar83a, Moh66, OW95b, Udw89, WP78, Wal70, WHZG24, Gas76b, Gur66, Kap67, Ros78, Vog93, Shi68]. Disturbances [Har66b]. Ditch [LS78]. Dive [DeS88]. Diverge [Boy05]. Divergence [JLM^{+}17]. Diverse [Bor62]. Divide [BJK11]. Divide-and-Conquer [BJK11]. Divided [WJ75]. Divided-Difference [WJ75]. Divisible [Jow86]. Division [Ede97, FK62]. Dixon [Eps59a]. Do [BM03, Bro61, BL13, Col76b, Gea81, Kel17, Pac88c]. Dobbs [Art91a, Art91b]. Doberkat [Sta85]. Doctoral [Gom67]. Documentation [Sal63b]. Dodd [Sco84b, Sco84a]. Dodge [Dru64]. Dodson [Jor77a]. Doelder [BG95b, Fet80]. Does [Ben70, Bol91b, Bol91d, Cor77, GS23a]. Doherty [Jor81]. Doing [Leh82, Wes92]. Dolezal [Kap66]. Dollard [Cha81b]. Domain [Asn75, BC76, GM81a, GM81b, Hir78, Mur97, SBGP98, XZ98, Gat88, Gil78, Kap77]. Domains [CHL08, DN59, Don23, ET99, Hen75, Hin89, IWMW21, Leb82, Wix76, Mit65, Sch86b]. Domenico [Chi95]. Dominance [DGS05, GDS05]. Dominant [LW59]. Dominated [Shu09]. Dominating [Yan97b, Yan97a]. Dominic [Gha89]. Domino [EJ07]. Dominoes [Ber86b, Day71, MBCG83b, MBCG83a]. Don [Pan86]. Donald [Art92a, Art92b, Asp84b, Chu90, Dav91a, Fle62, Gue84, Hoc82, Hom86, Kee88b, Kin76, Kla84g, Lee82b, Luk68, Mil91a, Par74, Per76, RR89, RAK88, Say79, SLF66, Str89b, CC90, Mil91b, Pul64]. Donating [Zha13]. Donev [Not94]. Dongarra [Hig00, Par81a]. Doob [Bax85, Bei81]. Doolan [Joh83d, Joh83b]. Doorn [Rue82b]. Dorin [Hor89]. Dormand [Sha97a]. Dose [JSR82a, JSR82b, Rae81, WM61a, WM61b]. Double [Abr84, Aga93, ASF86, Als86, Ber73b, Ber74, Bou71, Car67c, Car68a, Fis88c, Hep66, Hep67, Jor78a, Jor88c, Jos76, LF66, Lop94, MR65a, New67a, Nic65, OW88, Ove66, OR67, Phi71a, Rie72, Rob95b, Rob95a, RPP88, RS75, Sch63b, SR75, Siv68, SS70, Sri85, Sum87, Wil61, Wol72a, Bre89, Fis88d]. Double-Cone [New67a]. Double-Diffusive [RS75, SR75]. Double-Sum [Phi71a, Rie72]. Doubling [PH13]. Doubly [SG76b]. Dougherty [Sta88, Sta90]. Douglas [Ast87, Boy97, Noe81]. Dowell [Ash89]. Down [Gou68, NW67a]. Drábek [Cas94a, Cas94b]. Draft [Pen95]. Drag [New64a, Wil65a]. Draining [TS90]. Draper [Gre89]. Drawdown [Lew79]. Drazin [De 74, Gru93, Mil82a, Sco85, Sco90a, Sco90b, Cli78]. Dream [Gra94a, Gra94b]. Dream-Group [Gra94a, Gra94b]. Drew [Gue84]. Dreyfus [Cha64a, Hal66]. Drift [Jer95, PB75]. Drift-Diffusion [Jer95]. Drive [Mos92]. Driven [Kon03, Luc18]. Driver [Asn78]. Driving [Ben79, WHMN16]. Droplet [WB76]. Dropping [REG19]. Drug [Tho77b]. Drum [CU03, FN73, Pro87b]. Drums [Dri97]. Dual [PL78, Pea81, Pet76a, Wal76b, Wes65, Wit22, YZZ04]. Duality [Abd65, Abd66, BT69, Bis78, Geo71, GM95, HJ00, Kla66, Kor76, Rus85, Sha64a, Maw95]. Duals [Ben95]. Dubin [Vog93]. Dubious [MV78a, MV78b, MV03]. DuBois [Hil76, Gog85, McA92]. Duct [EMB76]. Dudkin [Dau90]. Due [Cha63, GP02]. Duff [CZ89, Hig00]. Duffin [BI68a]. Duffing [McC92b]. Dugas [Man61c]. Duistermaat [Tay97]. Dumas [Coo91]. Dunn [Cha95b]. Duo [Mey95a]. Dupuis [Dun94]. Dust [BCG77]. Dwork [Loe92a, Loe92b]. Dyke [Col77]. Dym [Gol73a]. Dynamic [Bel65, Ber91a, Ber86a, Cha64a, Cof83, Coo89, Dav91a, Den67, Ell85b, Fei96, Fly76, FS78b, Gla86a, GH13, Gue82, Kle83, LLSN96, Lar74, Lin91b, Mey75b, ML98, Mil89b, MMB^{+}07b, Nic96, PF96, Qui87, Ser80b, Ser84b, Ser84a, Set77, Sey93a, Sey93b, Shi93, SR75, Wou62, MMB^{+}07a, Art87, Beu89, Bin84, Dor74, Hal66, Kal63, Kno83, Lar78, Lin91a, Luc18]. Dynamical [AE23, Beb79, BGW15, Blo91, BT20, BBKK22, Chi94a, Chi94b, CB09, Dev94b, DHM07, Eas89, Eas93, Fel90, FG19, FM18, GT10, Gus76, Hal90a, Hal90b, Han91, Has92a, Has92b, HMK96, Hoc94, Hol87, HS95, HPS07, Jef90, Kan62c, KHS96, Kaz84b, Kaz84a, Kaz86, KS90, Kno98, Kod96, LS22, Lud75b, MQ18, Mei89, MP06, Mer85, Mey86a, Mey91a, MWL97, Mor73, Mor85b, Mor91a, Mor91b, O’M87b, Pal95, Rat87, RW95, Roo64, Sac74, San95, Sch70a, Sta92b, Sta92c, Stu86, Stu90, Swa96, WP95, YAKM20, dBC^{+}08, Gre97, Lak80, Mey89b, Sel94, Sle81, Wal95]. Dynamically [FL18]. Dynamics [Aar07, Ame91b, Ant87, Ari92, Bir83, Bom02, Boy97, Bry91a, Bry91b, Bur94a, Chr68, CN05, Cus96, DMT^{+}18, Dum96, EW75, FS96, FPM97, Fre79, Fre86a, Gue88, Has86, Hig62c, Hol87, HFKG06, Hur93a, Hur93b, Ins78, JMFB15, Kap90c, Kaz90a, Kaz90b, KT92, KCD^{+}97, Kra96a, Lan76a, Lin69, Mah94, Mar75b, Mea88b, MPT^{+}02, MXPW11, MPW23, MD96, MT14, NBR96, PTK96, Pal96, Pal94, PN99, Pip92, Por86a, Por88, Pro84c, Ram78, Ran96a, RVBE21, Rob91, RW95, Rog73, Saf96, SHL16, Sha95a, SK20, SAB19, SS24, SWC96, Ste00, SS95, Stu90, Thu01, Tze91, Van88, Wal95, WP96, Yor75, ZD78, dMvSB95, ES80, Ame91a, Are86, Boy88, Bru95, Caf94, Cro73, Dev94a, Dri81, Duf86, Kaz86, Lai91, Mir85, Rob85, Ant90, Ber88a, Cho94, Kap90a, Kap66, Poh63]. Dynamics-Como [Por88]. Dynamics-Optimal [Gue88].

Dynkin [Ell81, Pin70]. Dyson [Wig90]. Dzhuraev [XL94].

E. [Ant83a, Asp84a, Her83a, Mil83b, Pro84a]. Each [CH61]. Earl [Ash89, Cho66, Sam76b, Tab95, Tab97]. Earll [Por86a]. Early [Kaz92, PC89, Ste93c, PC90b]. Earth [Fri86, Wol88, Fri82]. Earthquake [Run91]. Eastwood [Bun83b, Bun83a]. Easy [Flu90, Shi17, WHZG24]. EasyUQ [WHZG24]. Eaves [Wil60b]. Eberhard [Bor06]. Eberle [Ork95]. Eberly [Dol91a, Dol91b]. Eccentric [ST94]. Eckart [Gea85]. Eckhard [GR95]. Eckhart [Saw85]. Eckhaus [O’M74]. Eckmann [Kir93a, Kir93b]. École [Dur87]. Ecology [CN15, Dix91, Ken91a, Ken91b, Ken93, Ken86a, Wal75b]. Econometric [Ell92a, Ell92b, Fis76]. Econometrics [Joh81]. Economic [Art91a, Art91b, Bur63, FP97a, Joh84b, Joh84a, Nel74, Pet76a, Bur64, Leh63a]. Economics [Cla79a, Fra86c, Gue88, HM01, Joh83a, Kee90, KB59, Nie83, Sel94, Wet74a, Joh83c]. Economics. [Art87]. Economies [Ell85a, vLMZ19]. Economies-of-Scale [vLMZ19]. Economy [DP76, Hyd86]. Ecosystem [WJV79]. Ecosystems [CN15]. ed [Ame85, Ame86b, Art87, Ban87, Bei78, Bla84, Blu65, Bra85b, Bre62b, Bro77, Bru86, But74a, But74b, But74c, Che86, Chi83, Cla62, Cyb93, Dol62, Duf86, Dur93a, Dur87, Fin84b, Fle64, Foo63, Fos90, Gib85, Gra86a, Gur63, Hag85a, Ham76, Ham77, Ham80, Jel87, Jon92c, Kap90a, Kau86, Kel96, Ken86b, Kow86, Kur69, Lor91, Mal90, Man61b, Mar86b, May90, Mey89b, Mil80, Moo60, Muk86, Mul64a, Nit90, Olv91a, Orl93a, Par74, Pip86, Por88, Pul91, Ral86, Rei63, Rei89, Sal62, Sal63b, Sch93d, Sen94, Stu91, Tek89, Tot91, Wal93a, Wei69, Wig85, Zie84b]. Ed. [Lee77, Yan94, Hoc80]. Eddy [MMT^{+}21]. Edelstein [Seg88]. Edelstein-Keshet [Seg88]. Eden [Jol96]. Edgar [Aal71, Dev91, God87, Moo74b]. Edge [Ano79e, Meg78, Meg79, REG19]. Edged [GK81, GRK81, Ren80b]. Edited [Bor06, Aur18, Gra18]. Editing [Hig01b]. Edition [Eas89, Hin94, Kob10, Koh86, Ler64, Orb18, May90, Ver09, Pac10, Wat10, Bir64, Che86, DeM65, Dem90, Gol65a, Joh87, Koç89, Kup75, Lam86, Liu88, Mar86b, Obe64a, Por86b, Rei87, Sch94, Ser80a, Sle86, Str86c, Wes92, Whi80, Har64b, Hoc79a, Kem67, Nit90, Pao93, Sop82, Zas65]. Edition. [Nas87b]. Editions [Ano95g, Ano95h, Ano95i, Ano96h, Ano96i, Ano97g, Ano97h, Ano97i, Ano98c]. Editor [Hun81]. Edits [Kru83]. Edmond [Bro59b, Pro65a]. Edna [Ham71]. Edoardo [Spi91a, Spi91b]. eds [Are86, Bai87, Bei85, Beu87, Bjö61, Bla88b, Bla81, Boa82b, Bri64, Caf92, Chu92, Des68, Die92, Don85, Fat93, Fre83, Goo91a, GB89, Gue84, Ham65b, Hop79, Hul61b, Kar60, Kaw65, Kla84i, Leh63b, Lin79c, Lu92, Man61a, Man61d, Mar85c, Mas87, McG64, ML85, Noh62, Pap62b, Par92a, Por86a, Pra85, Rei64, Ric74, Ros78, Sch87, Sco94, Seg85, Sik92, Sle87, Szy87, Tho87, Ull88, War92, Wei87, Wet75, WJV79, Yan92]. Eduard [Coo73]. Education [Ber05a, Ber05b, Ber05c, Ber05d, Ber06a, Ber06b, Ber06c, Ber06d, Ber07a, Ber07b, Ber07c, Ber07d, Ber08c, Ber08d, Ber08a, Ber08b, Ber09d, Ber09a, Ber09b, Ber09c, Ber10a, Ber10b, Ber10c, Ber10d, CCR^{+}62, DA79a, DA79b, Den17a, Den17b, Den17c, Den17d, Den18a, Den18b, Den18c, Den18d, Den19a, Den19b, Den19c, Den19d, Den20a, Den20b, Den20c, Den20d, Den21a, Den21b, Den21c, Den21d, Den22a, Den22b, Den22c, Den22d, Fra23d, Fra23a, Fra23b, Fra23c, Fra24a, Fra24b, LCK^{+}17, Lin67, Lin78a, Ros11a, Ros11b, Ros11c, Ros11d, Ros12a, Ros12b, Ros12c, Ros12d, Ros13a, Ros13b, Ros13c, Ros14a, Ros14b, Ros14c, Ros14d, Ros15a, Ros15b, Ros15c, Ros15d, Ros16a, Ros16b, Ros16c, Ros16d, RWMD18, SIA01, Sch01a, Sch99e, Sch99a, Sch99b, Sch99c, Sch99d, Sch00a, Sch00b, Sch00c, Sch00d, Sch00e, Sch00f, Sch01c, Sch01b]. Education [Sch01d, Sch01e, Sch02a, Sch02c, Sch02d, Sch02e, Sch02f, Sch03a, Sch03b, Sch03c, Sch03d, Sch03e, Sch03f, Sch03g, Sch04b, Sch04c, Sch04d, Sch04e, Sch05, Spa77a, TPCC^{+}11, YL03, PAB^{+}01, Ula67]. Edward [Ban88b, Coo89, Dev92, Dor74, Flu93a, Koh88, McK81, Mor88b, Nas77b, Rei69, Rog60, Sta88, Sta90, Flu93b]. Edwards [Coh75b]. Effect [BS00, DL76, Dre76, GS23a, GS04, GF76, LPL05, Mat89b, PLL05, Pet98, Tud92]. Effective [GBS08, JSR82a, JSR82b, Rae81, Sho69, Sle87]. Effects [BdCPT12, Mir97, OLHB02]. Efficiency [Bla85b, EH16, Sha97b, Yan90]. Efficient [Bet78, CC75, Enr78, FL18, MR00, MMB^{+}07b, Sei76, SXY19, SS78, MMB^{+}07a]. Efficiently [Hor78]. Efron [Siv94d, Siv94b]. Efroymson [PT81, SK80]. Ehrig [GM87]. Ehrlich [Her83b, Her83a]. Eichinger [Ric85b]. Eigen [PL78]. Eigen-Structures [PL78]. Eigenanalysis [EV10]. Eigencurves [BV96]. Eigenelements [Cha81a]. Eigenfaces [MMH04]. Eigenfunction [Sco59]. Eigenfunctions [Bra76b, Cla67, GN13, Koh72, NE76]. Eigenmode [CMZ94]. Eigenmodes [Dri97]. Eigenproblem [Cha95c, Cox94, Cha95d]. Eigenproblems [BG77]. Eigenspace [CD75]. Eigenstructure [McC03]. Eigensystem [Gol75c]. Eigensystems [GW76, TW69, Wei68]. Eigenvalue [Ale98, Ama76, Ama77b, BES05, BK99, BH15, Bou67b, Bou68, BEG23, Cho87, Chu98, CEH12, Cul94b, Cul94a, ESS79, GK12, Gol73b, Goo65b, Gur62a, Los76b, Mar73, Nie94a, Par66, PT81, Rei97b, Saa83, SK80, SKG^{+}80, Sle92, SV00, Ste73c, TM01, TB76, Was75, Wat93, Wat05, Par10, Rei97a]. Eigenvalues [Ann95a, BKR75, Baz73, Bir65, Bus76b, Cho69, Cho72, Cla67, CD75, Cul78, Eva93, Eva94, Fai71, FARE69, FR66, GLO20, Hoc66, Hul89, HJ90, Ike78, Jes77b, Jes77a, Kat68, KS68, KS84, New63, New65b, Ode85, Phi68, Phi71b, Ros80b, RS71, Sae63, Sea61, WC86, Zay92, Ann95b, Dav89, Gro93a, Ren94b, Sin69, Hea81, Rue73a, Wat86a]. Eigenvector [BL06, Ips97, LM05, TE96]. Eigenvectors [Cul78, Fai71, GLO20, Hul89, HJ90, LW59, Ode85, Wat86a, WC86, Koh98a]. Eight [Tre20]. Eikelder [Rat87]. Eilbeck [Sco84b, Sco84a]. Einar [Kap77]. Einführung [Hig62a, Tak77]. Eisen [EK91a, Mer81, EK91b]. Eisenstein [Sch60b]. Ekeland [Cla78c, Rat92a, Rat92b]. El-Hodiri [Keh89]. Elastic [AMR04, AR78, Cla62, Coo67b, Dat86, Hut89, JHSB22, Joh67, Kel64, LLSN96, LS96, Mor76, Nic96, Pit90a, Pit90b, Ren97b, Spi91a, Spi91b, Whe74, AR80, Bes92]. Elastic-Plastic [Pit90a, Pit90b]. Elasticity [AR95, HA75, Hut89, Joh67, Kno81, Mac92, Nie83, Sit92, Wu75, Ant78, Ant87, Bue71, Gos64, Hor83a, Mac92, Ren95, Sit92]. Elasticoviscous [Pip81]. Elastics [Han89]. Elbert [Mey83]. Eldon [Los86a]. Eleanor [Her67b]. Elected [Ano80e, Ano80f, Ano81b, Ano82g, Ano83e, Ano84e, Ano85a, Ano86a, Ano86b, Ano86c, Ano87c, Ano88a, Ano89c, Ano90b, Ano91c, Ano92c, Ano93e, Ano94a, Ano95e, Ano96f]. Elections [VLPR20]. Electric [Ber60, CM91, JM78, Mil93, Wil64a]. Electrical [BG18, Che73, CIN99, Mac61, SW11, Wil61]. Electrically [Chi79, Kel75b]. Electro [Jer91b, Jer91a]. Electro-Diffusion [Jer91b, Jer91a]. Electrodynamics [Jef92, Lee77, Rob94b, Pap63a, Wig91, Hut91]. Electromagnetic [GKLU09, Kri76, Lor91, MP93, Pic92, Rog94b, Rog94a, Rog96, Sle94, SPR76, WR96a, Dol62, Kel67, Mar66, Mil80]. Electromagnetics [CB76]. Electromagnetism [Sta66]. Electron [Gla96a, GJ97]. Electronic [BBR13, Pim93, SCS10, Sto62a, WA63]. Electronics [Art88]. Électronique [Mar85a]. Electrostatic [Gee78, XC11]. Electrostatics [JM92, Kan88, Pap86, SV98, SR93a, SR93b]. Element [AD76, BS94a, BS94b, Bak91a, Ban87, Bar79, BJK24, BG98, BL05, BSP95, Cab75, CL23, CLH97, Cuv90, Els89, Fal93a, Fal93b, Fel78, Fix76, Fix78, Fos90, GS23a, Gun88, Hen75, JLM^{+}17, Kel83, LeV93, LH75, Lor91, Mal85, Mal87, Mal88a, Mal90, Man76, Mar75b, Mea86, MMT^{+}21, MNS02, Muk86, Ode86, Ode89, Pas95, Red76a, Sch89b, Sch69a, Sch71c, Sco94, Sim93a, Sim93b, Tha96, VB75, Wah92, aYT96]. Elementary [BB84b, BCV98, Cop71, Dax97b, De 20, Dea89, Eng17, GL98, How63, KL97b, Kea95, Kea96b, Kog66, MPRW02, McM61a, Mil81a, Nie86, Pen95, Rei62a, Sch67, Ser78, ST94, Tha82, Zie83a, Dax97a, Dea90b, KL97a, Kea96a, Car87, Dan91, DG68, Eng78, Rus60, Zie83c]. Elements [BH96, BI68b, Bre61, CD93, Den90, Ebe75, Fix75, Fra86b, GK98, Hol91a, HS88, Kat76, Kat78, Kaz90b, Kri10, Leo78, Mal85, Roo10, Sal64, SX96, Wei84b, Wei84a, YL03, Cuv90, Bes92, Els91, Hua96, Tad93, Wen78, Xu96a, Cha83b, Hol91b, Kaz90a, Kei65, Kel86c, Kie94, Kol63, Lil68, Mil83a, Ral62a, Sys63, Thi63a, Wil65b]. Elevator [New72a, BOS73c]. Elias [Tre13]. Eliminating [Pet98]. Elimination [Grc11, Hou70a, Par61, SM22]. Elishakoff [Pan93a, Pan93b]. Elizabeth [Bri64]. Ellen [Ros85]. Elliott [Col89, Dam84b, Fle84b, Dam84a, Fle84a]. Ellipse [EV10, ST94]. Ellipsoid [Mur93, Woo96, Ano93c]. Ellipsoidal [Jac69, JS76]. Ellipsoids [Ale85, DKS96, UW13]. Elliptic [Ars62a, AG14, BNT10, Bar79, Boy78, BAD^{+}79, BE78, CG76, CE89, Cla67, DX98, Eck72, EKS97, Elc76, Fet72a, Fix74, Fle79, Gil69, Gla71, Gla81b, Gla82b, Gro96g, GS67, Hil76, HW76, Hui95, Izh01, Jor80c, Jor80b, Jor82, Jor90a, JW90, Kel83, Kus76, LA75, Lio82, Mit94, Mur97, Par86, Per79, Pro85, Ric76b, Roa78, Roo10, Rue79, SD74, Sha82b, SG82, SS62, SBGP98, Str86a, Urb93a, Urb93b, Var67, Wac89, Wah92, Gro96b, Gil82, Rao94, Sch86b, Wei87]. Elliptical [CMZ94, IWMW21, Jan70]. Ellis [Eis86a]. Elna [Wim72]. Elongation [VK84]. Elongational [Pip81]. Elsewhere [Vas94]. Elsgolc [Bel63b, Ham65a]. El’sgol’ts [Dri63, Dri75, Wou68b]. Elwyn [YY96]. Emanuel [Boa84, McG61]. Embedded [Bet78]. Embedding [Xu21]. Emden [Won75]. Emergence [WHMN16]. Emerson [Kaz92]. Emil [Wim80]. Emile [Ker91a, Ker91b]. Emilio [Ari91a, Ari91b]. Emission [CFHP10, HS15, Por64]. Emmer [Sen94]. Empirical [CSN09, Jok66b, Kli78, Gre89]. Empty [Jev92b]. Encounter [Abr94]. Encyclopedia [BC96b, Fir97, SPBC96]. End [HT78, Sei78]. End-Points [HT78]. Ended [Ano87l, Ano88j, Ano89m, Ano90m, Ano91l, Ano92i, Ano93j, Ano94g, Ano95q, Ano96r]. Ending [Ano85k, Ano86l]. Endo [EERG20]. Enemy [Ame86a]. Energy [Ant87, BF96, BL09, DP76, Gla95e, Ism96b, Kel88b, LT98, Pro84c, SXY19, Wei80a, Ism96a]. Engelbrecht [Par90]. Engels [Str82]. Engineering [Bau73, Bir65, Blu86, Dan66, Edw90, FP97a, Hac75, Hau90, Kri86, Lev63a, Maw93b, McA96b, McG64, MG96, Mor97, Pap62b, Pol87, RM96, RWMD18, Ruo92b, SIA01, She91, Spi84b, Spi84a, SS95, Tho87, TPCC^{+}11, Wil76a, YL03, PAB^{+}01, Ans64b, Ber83b, Beu87, Blu91, Boj93, Car65a, Ger96, Gos83b, Ham61, Har74, Hau88, Kap68, Kaw65, Maw93a, Per76, Ral80, Raz87, Sha95a, Wer63, Yao86, Ruo92a, Tap86]. Engineers [Hol87, Mal85, Mur94a, Mur94b, Sta89, Hig01b, Hig06, Are89, Mal85, Sto61, Sto62a]. Engines [Día78]. English [Cuv90, How94, May90]. Engstrom [Ahl93]. Enhance [CGT^{+}97a, CGT^{+}97b]. Enhancement [FGTW16]. Enhances [MT14]. ENIAC [Grc18]. Enigmas [Bol91a, Bol91c]. Enrichment [Gel79]. Enrique [Ede90a, Ede90b]. Ensemble [SBM22, Thy73, Thy74]. Enslein [Lin79c]. Enss [Hag85a]. Enthusiasts [Gou10]. Entire [Kaz62, Lau98]. Entropy [App87, Bal93, Boe88b, EB88, Eva88, Jea78, KR78, LRV76, Min78, PF13, Smi88c, Spo85, Udw89, WPTZ01, Eis86a]. Enumerating [HC16]. Enumeration [Ben74, Whi84a, Geo67, Ben76b, Moo74b, Whi84b]. Enumerative [Whi88]. Environmental [BK83, Ken86b]. Environments [Kem92b, VCKJ16, Kem92a]. Envy [BJK11]. Ephemerides [New76b]. Epidemic [AMS^{+}23, Yan92]. Epidemics [CK76]. Epidemiological [Rou97b, Rou97a]. Epidemiology [Hop83]. Epochs [Gre67b]. Eppler [Che92c, Che92b]. Epstein [Sch63a, Wei79]. EQS [DEPC95, Cha95b]. Equal [Fee00, Gla78a, GN78, Jag79b, Nås77a, Sha78, SJ79]. Equality [KA89, Sha74]. Equation [Abd84, Abd91, Amo65, Asm98, BS00, Bar78, BW11, Bob66, Bou67a, Boy13, CC78, CH68, CP89, CPRG90, CP78, Coo71, Cru86, Dat82, Dor89, Dor70, Dor73, Dor75, DG14, EH78, Elh78, EL85, EL86, EC76, FGD^{+}92, Fus68b, Fus69, GQJ96, GZ19, GM81a, GM81b, GHC97, Geo84a, Geo84b, Gie63, Gla75b, GP79, Gol76c, Gru83, Gue84, HT78, Her87, Hun76, Ike75, IM88, JR87, JK72, Jód96, Joh75, Jor86d, KSS80a, KM85, Kla73, Kla76a, Kle79, KR74, Kor95c, Kra80b, Kra80a, Lax76, Lei76, LG74, LG75, Lew89, LBD^{+}86, LKGK07, Lig86, Lin10, Los76c, Lov79a, Lud77, Mah75, MB76, MS72, McC92b, McG82, Mel91, Mel92a, Mel93, MR78, Mic88, Miu75, MS14, Moo65, MC76, MN76, Mor69, Pan97, Par81b, Pin92c, PI88, Raw70, Raw86, Red76a, RGJ90]. Equation [Ren93a, Ric88, RK75, RLGBRMC06, Rod81, Rog73, RBB^{+}90, Ros79, Rub77b, Rue72b, Rue73c, Rue74b, Rue75, Sch77a, Sch84a, Sch85b, Sco59, She71, Sim96, Siv72, SC70, Str65b, SP82, Swa77, Tre20, Tur68, Vel63, Vel64, VN76b, Wak86, Wal92, Wal93b, Wal63b, WiOH10, Wil84b, Wil84a, Wil89, WZ72, Won75, YFS76, YF78, Ano97j, Che96a, IM89, Miu76b, And77, Ano93d, Bou66b, Bra66, Car76b, Cas72a, Che96b, Die92, Dis92, Eba82, FW82, For92, Fre80, Goo68, Hai77, HS72, Jag87b, Kea80, Kir76, Kol65, LS66, LG89, Lew86, Los74b, Mel92b, Mey78, Osb85, Ren93c, RB90, Rou73, Sch84b, SLF66, St.82, Wal93c]. Equations [Abd65, Abd66, Abl75, Abl77b, Abl84b, Abl84a, Aeb94b, Aeb94a, Aga95a, Aiz94a, Aiz94b, Ale90, All78, AW76, AG80, Ama76, Ame78, Ame86b, Ant82, AB23, Ari91a, AG14, Ast70, AC78, BNT10, Bai76, BF96, Bar70, BJK24, BG76, Bea90b, Beb68, BI69, Ber84b, Ber84a, Ber65b, Ber91c, Ber78, Ber90b, BR76, BV96, Blu94a, Boj93, BR18, Bre94b, BDE09, Bur83, Bus64, But89, But92, Car68b, Car69, Car72b, Car74, Car75b, Car77, CW94, Cas75, Cas72b, CCCM06, Cas94b, CSW^{+}08, Cel78, CG76, Cha84, Che67a, Che76, Che78, CT75, Chu94a, CL75, Cin93b, Coc86, Col91, Cor92a, Cor95, CS90, DO98, Dar97, Dax90, De 84, DTV14, Det85, Deu85, Dia59, DA78, Don90, Dri68, Dri75, Dri81, DJLQ21, Eas91, ERHO95, EFN^{+}96, Ell61, Enr89]. Equations [EKZA96, Eri66, EW75, EK87, Eva85, FLM22, Far96, FGN13, FC99, Fic60, Fin76b, Fle77a, Fle77b, Fri80, FEY61, GVL97, Gar75, Gat64, Gat88, Gea81, Gil78, Gil90a, Gil90b, GC78, Gli91b, GM87, Gol76a, GF78, Goo65b, Gos83b, GS22, GL98, Gra94b, Gre59, Gri85, Gue85, Gue95, Gun63, Gup98b, GS67, Gus76, Hac75, HY71, HL63, Han94a, Han76, Har96, Has92a, Has92b, Her73, Hig01a, Hin89, Hoc66, Hoc79a, Hoc92, Hol85, HW76, Hor97a, Hou67, HM73, Hüb88, Hul67, Ike72, IL97, Jen75, Jes93a, Jd93, JU76, Jon71, JR08, Kah66, KL76, Kap85a, Kap72, Kap77, Kel60b, Kel83, Kel88a, KMH66, Ker92, Kla62b, Kla66, KR74, KPGR95, Koç94, KLA95c, KG95, Kor65, KKS82, KM96, Kup94, Kus76, Kus93, Lae82, Lam86, Lan70, LRM96]. Equations [Lat90b, Lax69, LP75, Lei78b, Ler63, Ler64, LA75, Lev88b, Lev65, LR80, LW04, Lio82, Liu94, Liu95, LM81, LLR^{+}81, LA95, Lon77, LRJ78, LMMK21, LS02, LH19, Mac76b, MS96a, MH96, Mat69b, MG76, MMS98, McG93, McL86a, Mei68, MP06, MR92, Mey69b, Mey77b, MA95, Mie17, MBK97, Miu77b, Moo78, Na67, Nas86, NH78, Noh76, OL98, O’M10f, OT12, Ole94, Olv91a, Olv91b, Ork95, Ort74, OV85, Ost61b, PY98, Pal84b, Pal84a, Par74, PZ64, PSSW06, Pea81, PD76, Per65, PW79, Pic87a, Pic90, Pin84a, Pin87b, Pon90b, Por80, PC95, Ral65a, Ral69b, Ral69a, Ral74, Ral65b, RFP10, RRK95, Ric76b, Riv63, Roa78, Rog83b, RSTK97, Ros61, Rub78, RRG95, RS71, Rus78, SGK^{+}96, SB98a, San65, SS16, Sch63a, Sch80b, Sch88, Sch90, Sco84a, Sed96, Sei65, Sei93]. Equations [Sem59, Sha75, SG79, SG95, Sha97a, Sha04, SS62, SES75, SS98, SR98, SA78, Sig67, Sik92, Sim16a, Ske78, Sle79a, Sle81, Slo97, Smi88b, SBGP98, Sta71, Sta76, Ste62, Sto76, Str88, SY69, Sul93a, Sul93b, Svo96, Syn83, TK97, TF96, Tew77, TP78, Tre84a, TS90, Und78, Ung77, Urb93a, Urb93b, Var67, VH78, Vel21, VM95, Wal76b, Was94a, Wat22, Wei60a, Wes65, Win90, Win67, WMZ94, Won64b, Wou68b, Yan65, You76, ZD95, Zie79, ZMPK24, dO69, Gup98a, Ram88, SWD76, Str89a, Jac97, Kla97, Sta72, Ber83a, Fal70a, Aga95b, Ame85, Ame95, And73, Ans66, Ant63, Ari96, Ari91b, Arn89, Asn78, Ast87, Bai87, Bea90a, Ber91a, Ber91d, Bra65, Bra63, Bre94a, Bro59b, Bue66, Bur93, But92, Cam93, CL62, Cas94a, Ces75, Cha94a, Cha94b, Cha67a]. Equations [Cha81b, Chi90, Cho66, Chu92, Chu94b, Çin93a, Col94b, Con85, CK86, Cro85, Dan93, Dat78, Daw74, Del95, Dor74, Dri63, Fer79, Fle62, Fle79, Fle82, Fle85, Gar91, Gea74, Gea80, GR95, Gib95, Giv78b, Gla95a, Gol93, Gos83a, Gra94a, Gre62a, Gro87, Gue72, Gup89, Gur62b, Han94b, Han75b, Han79, Han81a, Han85a, Han87, Han86b, Har61b, Har70, Her65, Her67b, Her71, Hin94, Hoc65, Hoc86, Hou67, Hul71, Hur65, Isa89b, Jag64, Jer81, Jes93b, Jol96, Kaz83, Kel60a, Kor95a, Kum88, Kus74, Lae76, Lat90a, Laz82, LeV93, Lev62, Lev67a, Lev96, Liu91, Maw96, MNMF63, Mea88a, Meh96, Mey72, Mey93b, Mie95, Mit94, Noh67, Olm77, Olv90, Olv93, Par90, Pay68, Per68, Pic87a, Pin84b, Pon90a, Por72, Pre81, RCC71, Rao94, Rhe70, Rox92, Rub69, Rue73b]. Equations [Sch73a, Sch60c, Sco84b, Sei93, Sle92, Ste93a, Ta’90a, Tho59, Tre84b, Was94b, Web85, Wei91, Whe76, Wil66, Wil72, Wil73, Wim67, Wou68a, XL94, Yan76, Dol72, Noh77, Pro65c]. Equations-Introduction [Lae82]. Equations-Second [Lam86]. Equations. [Lu92, Pim87]. Equilateral [McC03]. Equilibria [Art92b, BT13, Fin93, Pet76a, Shu90b, Tij81, Art92a, Has86, Lem84, Shu90a]. Equilibrium [Bot72, Bry90a, Bry90b, HHSS75, HHOT93, Kir90a, Kir90b, Mil68, Ric90c, SK20, Str88, Per89, Ric90b, Alm87, BW73, Cur64b, Spo85]. Equispaced [PTK11]. Equivalence [GG68, Kel84d, LR05]. Equivalent [AvK^{+}08, Wet74b]. Equivariant [BFM95, Mey95b]. Eradication [GZ11]. Erbe [Ari96]. Erdélyi [Van96, Sop65]. Erdos [Nic94, FS73, Moo75]. Ergodic [Kim65, Shi94, Tul68]. Erhan [Pit77]. Eric [Kaz91a, Shu90a, Shu90b]. Erich [Det85, Sch82b]. Erick [Che18]. Ericksen [Pip86, Sle87]. Erik [Le 92]. Eringen [Hut91]. Erisman [CZ89]. Erlang [Jor78b, New94a, Coo93, Coo94, JvD90, JvD91, Kau77]. Erlebach [Knu80]. Ernest [Fet66, Har74, Spi84b, Vel71]. Ernst [Cam91b, Mit93, Ork95, Orl93b, Cam91c]. Ernst-Rudiger [Mit93, Orl93b]. Errata [Ama77b, Ano66a, Ano92d, Ano94b, Ano95f, Ano96g, Ben76b, Bri63a, Bur64, Fra66a, Gro94b, Mat76b, Miu77b, Mor86c, Sta72, WM61b, Ano93d]. Erratum [Ano66b, Ano68e, Ano73d, Ano78h, Coh93b, Gla76, Gre67c, Gud78b, Hou68b, Mac98a, MPW23, PC90b, ST91a, Wil80c, Win96b, Ano80d, May84c]. Error [Ahl62, All87, Ang92b, Ban88b, Ban88a, Beu66, CL23, HPS07, Jac69, Kah66, Köl87, LRV76, Mat80, Mat94a, Olv78, Olv80, Spi86, Ste73c, Tin76, Wil71, Won80, dB74, Leh82, Mat94b, Wil64c, Zie76]. Error-Bound [Kah66]. Error-Correcting [Mat80]. Error-Free [Beu66, Leh82]. Error-Like [Köl87]. Errors [For59, Gra66b, Jok66b, Koh98b, Rei62b, Sto60]. Erwin [Gil90a, Gil90b, Ruo90a, Sop79, Ruo90b]. Escape [HS14, Kla59, New64a, Wil65a]. Especially [O’M94]. Essay [Har66a]. Essays [Man61a, Ham70, Ham80, Ung85]. Essence [Hol95, LH95]. Essentially [Shu09]. Establishing [Kre91]. Estimate [Bor62, FSW^{+}66, Fre69, ST98, Wah65]. Estimates [AABT15, GBF78, HW76, Hor97a, Kaz62, Lev90a, Lev91, Pri64, Pro61, WP78]. Estimating [ABL^{+}13, HP76, KB76a, Law78, MT84, Mel64, SM94, WPL98]. Estimation [ASPL03, BE78, BK83, CL23, CC90, CR78, Dix91, Har66a, Har93, Har63, Hea76, Hig87, HPS07, JSR82b, Lin80, MYDF08, MS75, Mor88d, Pai76, Pap85, Rae81, RW64, Ruk93a, Ruk93b, Rus76, Ste99, TLN95, WE76, Wit75, Bic94, Bjö85, Bro77, Col92, Goo82, Nas95, Rei75, Sam86, Chu90, JSR82a, Lud91]. Estimators [AB69, Cho76a, Far94]. Estrada [Wim96]. Etching [MPRW02]. Ethics [IGWL90, Loc89, LIW90]. Ethier [Par88]. Ethnomathematics [Her93b, Her93a]. Etudes [Fir92b, Fir92a]. Euclid [Hou74a]. Euclidean [Ort74, Kli78, LLMM14, Ter82]. Eugene [Fix78, Moo69, Mor73, Sai91a, Sco88, Ser80a, Spä97, Var67, Sai91b]. Euler [LeV93, Ork95, Ta’90a, CJ16, ERHO95, GW12, Gau08, LH19, Mel97, Vel21]. Eusanio [Law65]. Evaluated [BR94, Bus95]. Evaluation [Amo78b, AD76, Bla64, Cor68, Far69, Gas85a, GBC^{+}90, Gib90, Gol62, GM66, Hab70, Kam76, Lin78b, Lin79b, Mac67, MN65, NS77, Raz87, Rue84, Ruo92a, Ruo92b, Sch76a, Sch63c, Sha89a, Sle78, Sle79b, Sne67, Beu92, Bla88a, Gas85b, Jag79a]. Evangelos [Bla88e]. Evans [GdMGR91, Mal85, McG62b, Ren94b, KM14]. Evar [Cot94]. Evasion [GK15]. Even [Isr92a, Nie94a, SS91, Wat96, Isr92b]. Evenly [Ber07e]. Event [GYS95, Sha95c]. Events [FHRO95, O’B95]. Everitt [Cha95b, Dur93a]. Every [Día78]. Everywhere [Goo90b, Goo90a]. Evgenii [Pin70]. Evgrafov [Kaz62, Whi67]. Evidence [VCKJ16]. Evolution [Abl75, Abl77b, Aiz94a, Aiz94b, Ame78, Don90, EFN^{+}96, EFMM13, FS78a, GL89, God87, Gol93, Hal90a, Hal90b, Ins78, Jer81, JMFB15, Jol96, Mac91b, Mac91d, MB76, MBK97, Par90, Pic87a, Pic90, Sho87, Sle81, Wah92]. Evolutionary [Ber88a, Mie95, PC95, VM95, WHMN16, Cor95]. Evolvents [Kaz91a]. Evtushenko [Col87b]. Ewald [Cha64b]. Ewens [Saw82]. Exact [CC76, IM88, IM89, Kla97, Lag97, Lew89, Lio88, Mic88, MV93, SS05, TP78, LG89]. Exact-Solutions [Lew89]. Exactness [Tre22]. Examination [Var79]. Example [AABT15, Ale95, Ale98, ABL^{+}13, CS20, Edw00, Isa89a, JS68, Mac77, VS95]. Examples [Ano99b, Col76b, CS20, The99a, FKL^{+}06, FH76, Hau81, Lou70, Mey86a, PW08, Har83]. Exchange [Log98, TE96, Wol86]. Excitable [KT92, WM15, Win90]. Excitation [Dat86]. Excited [Ell90]. Exciting [CJR18]. Exciton [Gla95c, BC96a]. Excursion [Hou74c]. Exercise [HH75]. Exercises [LHFE97, Mil89b, Hig02c, Bak66]. Exhibiting [Mat75, Mat76b, SHL16]. Existence [And66, Ann95a, Art92b, Dun68, Faz96, Fox67, GL98, Het72, Inn69, Lio82, Mac68a, Mas92, Ren97a, RT97, RS71, Vei67, Ver98, Wan97, Win75, Zay92, dBL75, Ann95b, Art92a, Gro93a]. Exo [EERG20]. Exp [Kam76]. Expanded [Bre61]. Expansion [Bar71, Bru65, FO99, Fet66, Gil62b, Hag83, Hau75, Hau76, Hop75, Jag73, Lan67a, Lan67b, Mcn66, Mor66, PS64, Rub78, Sha71a, Wil64b, Wol72b, Han64, Ism76, Jor72, Luk68, Rue72c, Van66, Rue72d]. Expansions [Ask75b, Bos96, Car70b, Car71d, DeM65, FH76, Jor83b, Kea95, Kea96a, Leb82, Lóp08, Los84b, Lud70, O’M74, Rei71d, Rot81, Vet73, Was74, Wol77, Won80, Tre65, Kea96b]. Expectation [FZ63, KA89, Zha88, vA61]. Expectations [Mal68, Mal69, NW67b]. Expected [AGR83, AG83, Boe96a, Gro93b, Kla71a, LS95, New72a, New92, New94b, Rao82, Rog78, Rue72a, She61, SS61, Van63, BGML96, Bor79, BOS73c, Gro93d]. Experience [Dra83, Ehr76, Gom67]. experiences [New84]. Experiment [HF05, YL75]. Experimental [Dan66, Jow76, Jun64, Lin92, NBR96, Not94, Ran96a, Raz92a, Raz92b, Rei62b, Ruk93a, Ruk93b]. Experiments [Arc75, CEH12, Gou16, KPSP96, Koç94, Pin87a, Jow72, Jow86, Le 92, Pro96]. Explanation [Mey89d, Lev88a]. Explication [FH76]. Explicit [Fre69, Hol70b, Lan70, Lei76, LS02, Lut66, Zlo70]. Exploiting [Aur18]. Exploration [Saa95]. Explorations [Koç94]. Exploratory [Gra93b, Gra93a]. Exploring [De 00]. Exponential [BBN98, Ber90c, Cav74, DJLQ21, EFN^{+}96, Hig09, HW76, Köl87, Kra66, Leo96, Liz98b, Mac98b, Mey80a, MV78a, MV78b, MV03, Nag96, PJ75, Pru78, Rus85, SN96, War76, Wat91a, Liz98a, Pow75, Wat91b, Jol96]. Exponentiality [Pit81]. Exponentially [Bar22, Lev90a, Lev91, TW14, Tre20]. Exponentials [Ber88b, Gla91, HFS01, Kam78, Ter96, WBLQ89, Ste97a, Wer89, Ter92]. Exponents [Lev90b]. Exposed [CS84]. Expository [AM74a, Ips10a, Ips10b, Ips10c, Ips10d, Ips11a, Ips11b, Ips11c, Ips11d]. Expression [Ier96, Lam69, Lam70, IR97]. Expressions [BSFW82, BSW82, Fos81, PR70]. Extended [FH76, Kir93a, Kir93b, KN67]. Extending [WM15, Gra18]. Extension [Alt76, Ask84, Esl96, Gud84, KK78, Kro66, Na68, OHBG00, Pon65, SK85, TT94, Wil85c, Jor86a]. Extensions [Ask85, Ask86a, Bra76a, Hor78, Rei77a, Zuo96]. Extensive [LMDLS23]. Extent [Olm76]. Exterior [Her84a, HA75, Sha64a, Her84b]. Exton [Ask78, Ask79a, Ism85]. Extraction [BGH^{+}04, CD94, Zit63]. Extrapolated [Hen75]. Extrapolation [Deu85, Joy71, Pru75, Rüd97, SFS87, SFS88]. Extrema [Mal68, Mal69]. Extremal [Erd92c, Erd94, HT81, Hue96, LM00, Mcc78, NN67, Nic94, Sin80, Wil69, Ano97f]. Extremals [Rei77a]. Extreme [Art80, Bus76b, Dwa80, Gro96a, Jes77b, Jes77a, Kla76b, Kla92a, Mey80b, VD68, DGH^{+}18, Gro96e, Los81]. Extremes [FHRO95, O’B95]. Extremum [Bal85, Can86b, Can86a, KS73, Nas77b, Tap72].

F [Ano71a, Arn89, Ars62a, Ask75b, Ass77, Bar70, Ber97a, Ber90b, Bla93, Blu91, Böh88c, Bro90, BG64, Car66c, Cho70, Cou97, Dan93, Dem86, Dem90, Dur92a, Ell90, Ell80, Eps59a, Fat93, Fel72, Fer79, Fin93, Fis76, Fre86c, Fri86, Gar61, Gla95a, Gos83b, Gra67, Gra86a, Gra83a, Gru74, Gur62b, Har61b, Har70, Hav95, Hea91, Hea74, Hey82, Hey86, Hey88, Hoc65, JG70, Jor81, Kal70, Kaz91a, Kla84i, Kla84b, Kla97, Kle75, Kra96a, Kre96, Lan86, Lee84a, Lud77, Mar64, Mar73, Mat80, May90, McA92, ML85, Med78, MT61, Mil94b, Mit87, Mor96, Mor89c, Nas77b, New87, Nov94, O’M86, O’M87b, Pac86, Pac88b, Pap62b, Par80, Pat87, Pay68, PRM^{+}66, Por88, Ral62a, Ral70, Ren94d, Row86, Sam76c, San82, Sán83b, San83a, Sco91a, Sco91b, Sha92a, Sha92b, She63b]. F [Shu87, Spi81b, Tay86, Wei91, Wet93, Wie87, Wil66, Wim75, Woź87, Yao86, Zar65, Zie78, Dod73, MNMF63]. F. [Jef94, Lem83a, San83a]. F.-A [Woź87]. Fa [Dur93b]. Faber [Saf96]. Fabian [Fro87b]. Face [Kar22]. Facial [Art80]. Facility [Spi76]. FACR [Swa77]. Factorable [Sed96]. Factorial [Fra83, LS84, Lux84]. Factorials [KM85]. Factorization [Ber73a, Bos87, CL71, Gle81, Gle85, HL76, Hoc60, HS71, HMH00, Kha87, Lan81, Mul64b, Nie94a, Sho99, Tho74, Fra66b, Gle83, Jag72c, Mec92]. Factorizations [CFG95, DG20, GZ19]. Factors [DM07]. Facts [Wat10]. Faddeev [Hou98b, Hou98a]. Fahrmeir [Mar95]. Faierman [Sle92]. Failure [Bur78b, Kel80a, Sku04]. Fair [May84c, May84b, May84d, Saa88]. Fairer [BI18, VSM14]. Fairweather [Ano82f]. Falb [Kap70b]. Falconer [Kla92b]. Falk [O’B95]. Fall [Ano69k, Ano72l, Ano73l, Ano74l, Ano75n, Ano76p, Ano77o, Ano78p, Ano71m].

Falling [Ber86b, Day71, MBCG83b, MBCG83a]. False [Bur78b]. Familiar [BB88]. Families [DS61, SN96, Ste64, Nag96]. Family [Cas72b, Cle69, Jag92, Jes93b, Jes93a, MR92, Mey75a, PH76, Rue73b, Sei91, Wah68, Gro98a, SGD98]. Fantasy [GLY^{+}18]. Far [BC76]. Faraday [CHT15]. Farin [DeR91a, DeR91b]. Farkas [Fre95, Dax97b, Dax97a]. Farokh [Wit90]. Farrar [Smi74]. Farreny [McA92]. Fast [Bai93a, BB84b, CU03, DGP^{+}16, Eds05, GL04, GJRW24, Hir78, JS20, Kel84d, LW17, OT13, PTK11, Set99, SP01, War98, WPTZ01, XC11, Bai93b, Hen79]. Faster [Wil21, Cop86]. Fastest [BDX04, SBXD06]. Fathi [Mei89]. Fatigue [Mat89b]. Fatou [Bur94a]. Favorable [GO67]. Fazio [RT97]. Feasibility [BB96]. Feature [Sch75a]. Featured [Bor06, Car08, Hig99]. Federal [GH22]. Federgruen [Hüb88]. Fedoriuk [Pro83b, Pro83a]. Feed [KW18]. Feedback [Bal78, BH76a, BvJ09, Kee88b, MO78, McG65b, Noh76, Old62, Yan90]. Feinerman [Kin76]. Feintuch [Kha86]. Felberg [Sku04]. Feliciano [Car66c]. Feller [Kes69, Pin72a, Sie69]. Fellman [Mal93a]. FEM [BS00, Say13]. Fenyo [Fle73]. Feron [Yak95]. Ferromagnetism [KP06]. Ferziger [Uhl73a, Uhl73b]. Feshchenko [Har70]. Fetter [Pol59, Rog60]. Fettis [Boa82a, Han85b, RSTW73]. Few [Mcc78, SM78]. Feynman [MW66]. FFT [BH90, Eng17]. FHMA [Jev92b, Jev94]. Fiacco [Lem85, Man70]. Fiber [All76, Lew79, Wu72]. Fibers [WiOH10, Wil76b]. Fibonacci [Mye75]. Fibonaccian [Kro66]. Fibre [All75, Les76]. Fichera [Ste81b]. Field [BC76, Bou91b, CS17, Chi79, GHC97, Gee78, Her91, KN61a, Kla82, Kla83a, MN65, MP93, Pro65b, SPR76, Ste94a, Ste94b, Str75, WM61a, WM61b, WB76, Pul91, VK62]. Fields [CLWM95, DS73, EK13, GP02, JM78, Jos76, Kaz84a, LH75, Núñ96, PL78, Pic92, SG72, Vil74, Xan87, Sch92, Jor78a, Kaz84b, Mey95a, Pic87b]. Fienberg [Par92a]. Fife [Kaz80a]. Fifth [LK65, Lut66]. Fifth-Order [LK65, Lut66]. Fifty [Dur93a]. Figures [GK12]. Fike [Rad73]. Filippov [Ber90b]. Film [Lin86]. Films [Bas86, Kon03, Mye98]. Filter [Che73, CC90, HRW12, LRV76, RBC^{+}07, Chu90, Par92b]. Filtering [Bjö85, BdCPT12, BvJ09, Chu85, Cur75, KKBG09, Mer89, Mul76, SBM22, VC06, Bel10]. Filters [Bri65, BGL01, Chu88b, Ram89, SKP19]. Finance [O’M05d]. Financial [HS15, Hyd86, Mal81, Mal83, Ste01]. Find [Aro78b, Cor77, Ebe75, LB78, Hig62d]. Finding [Boy13, HMT11, Max60, PH76, PO76, RR97b, RR97a, Ste70b, Ypm83]. Finetti [Rei77b, Rei78c]. Finger [Goo65a]. Finis [Cuv90]. Finite [Aga95b, AD76, AW19, BS94a, BS94b, BH96, Bar64, Bar79, BJK24, Ben70, BR91, BG98, BL05, BdD92, BD92, BGL^{+}67, BSP95, BGL01, Cab75, CHL08, CL23, Chu66, Chu67, CLH97, Cul77, Der71, EK13, Fal93a, Fal93b, FG76, FGB81, Fel78, Fix75, Fix76, Fix78, FSP93b, FSP^{+}93a, For98b, Gia63, GAT75, Gro92a, Har61a, Hen75, Hua96, JLM^{+}17, Kat76, Kel60a, Kel83, Kle72, Kus76, LeV93, LH75, LH19, Mal85, Mal87, Mal88a, MO78, Man76, Mar75b, MK78, Mea86, Mey75a, Mey75c, MA95, MNS02, Ode86, Ode89, Pas95, Pet73, Por80, Red76a, Rei59a, Roo10, Sag81, SW79, Sch89b, Sch80c, Sch81c, Sco94, Ser80b, SX96, Sik92, Sim93a, Sim93b, Sin93b, Sin93c, Stu89, Tad87, Tre82, Udw89, VB75, Vel21, Ves95, Wah92, Wen61, Wen78, Xu96a, Yau69, Zei72]. Finite [Zha97, Zua05, For98a, SW80, Cuv90, Neu97b, Bak91a, Gun88, Har61b, Hir71, Mal90, Mea88b, Min66, Ser80b, Wig91]. Finite-Difference [GAT75, Zha97]. Finite-Dimensional [Pet73, Sag81]. Finite-Element [BS94a, Fel78, Hen75, Kel83, LH75, Man76, Mar75b, Red76a, VB75]. Finite-Time [LH19]. Finite-Volume [Vel21]. Fink [Aga94]. Finkbeiner [Wis60]. Finn [Alm87]. Firby [Pat87]. Fire [ADW23]. First [AW19, Ber86c, Beu85, Bla97b, Blo73, Bob66, CD20, GP79, GLO20, Hil76, HS95, Hor93, HM73, IL97, IWMW21, Jor78b, Kau77, Kaw67, Kes85, KA95, Lee82a, Lev90a, Lev91, Man61d, Mat88, McG93, Mey78, Mir85, Nas86, O’M87a, Owe85, Rub77b, San95, SD74, Wei77, Bla97a, Apl95, Pro85]. First-Order [Bob66, CD20, GLO20, KA95, Apl95]. First-Year [Bla97b, Bla97a]. Fischer [Are86, Boa84, Bri64]. Fish [MPT^{+}02]. Fishburn [Hof85, Kar90, McM87]. Fisher [Mot88]. Fit [Wal66, BS67]. Fitted [MOSK97]. Fitting [Ber64, Ell02, Gou16, Han76, Pec62, Pra89b, Sei78, Sto60, Sch85a]. Five [MV03]. Fix [Cab75]. Fixed [AG80, Ama76, BH76a, DG14, FLNU18, HSŠ16, Lem84, LP71, New65a, San65, Sez90, Wet74b, Zha13, Ama77b, Fra86c, vV65]. Fixed-Cycle [New65a]. Fixed-Point [BH76a]. Flagellar [Lig76]. Flame [Squ61, War64]. Flament [Har64a]. Flames [Kap85b]. Flannery [Isa88, Isa89a, Yan94]. Flat [Jac07]. Flatto [RJ90]. Flavin [Hor97a]. Fleeting [Men94a]. Fletcher [Den82, Tap84b, Tap84a]. Flexible [MT96]. Flight [Edw63, KN61a, Kla76b, Kla82, Lan76a, Zar65, Zwe16, Edw63, Kla83a, VK62]. Flipping [REG19]. Floating [Esp95, Fat76, Gol76b, GF76, Hig02c, Row76]. Floating-Point [Esp95, Row76]. Flohr [Seg85]. Flor [Rue78]. Florea [Dav75c]. Flores [Fla66]. Florian [Ant92a, Ant92b]. Flour [Dur87]. Flow [Ber90a, BCH14, Cel78, Dea89, DL76, ERHO95, ERS14, FLGM23, GG91, Gos63, GJ19, GWK95, GW95, Ham88, Hen76, JHSB22, Kaw95, Kel75b, Kla94a, Kni86, Kon03, Maj91, MTGF78, Ork95, Pit89, Poh78, PDS78, RFS03, SHL16, Sei76, SR94, Sim96, Sle85, Smi88b, Smi97, Stu86, Sun76, Tu76, VB75, Vel21, Wil78c, Dea90b, Axf65, Cor79, Fre79, Pat64, Riz94, Sed88, Ser64, Shi71, vL86b]. Flow-Based [FLGM23]. flowering [Lax89]. Flows [AMO91, AMOB95, Bak91a, Bal20, Ber85f, Caf92, CS75, Çin93a, Cin93b, EMB76, Hu68, JLM^{+}17, Kel75a, LR00, MB76, PL78, Pip81, Reh89, SXY19, TKG95, TS90, Wat84, ZD78, Bal95, Ben87, Gun95, Kan68, Kra67, Ock91, Yan91]. Fluctuation [Tak79]. Flügge [Bus70]. Fluid [Aar07, Ame73, Ant87, Bat21, Ben87, Bir83, Boy88, CL08, Col77, Cro73, Dij90, FS96, FPM97, Fif79, FS76, Gli91a, Gun63, Hir75, Kon03, LB10, Leo78, Lin86, Maj91, MTGF78, Mir85, Mir88, Moi89, Por86a, Por88, Pro84c, Ram78, Ren94d, SHL16, She75, Sim96, Sle85, Sun76, Van88, VB75, Wei90b, vL86b, vL86a, ES80, Pip92, Saf96]. Fluid-Dynamics [Ram78]. Fluid-Flow [MTGF78, Sun76, VB75]. Fluid-Plasma [Wei90b]. Fluid-Structure [Bat21, LB10]. Fluides [Cuv90]. Fluidization [Dre76]. Fluids [AHP^{+}95, BL09, Moi89, NC76, Rau95, Red93b, Cuv90, Chi79, Dat90, Mar86a, Pip81, Red93a, Sub90]. Flux [Zha13]. Flying [Yat84a, Yat84b]. Fock [Gol65a, Kel67]. Foias [Bea90a, Jol96, Liu91, Bea90b]. Fokas [DTV14]. Fold [BH76b, OSS82, Grz83]. Folded [DGP^{+}16]. Folding [MK99]. Foley [Hea74, PRM^{+}66]. Folger [McA88]. Folkmar [Hig05]. Following [GG80, Gon92, PGH^{+}88, Sal87, Pin88]. Fomin [Ham77, Ral62a]. Football [Gou10, Kee93]. Force [Dre76, HJ00, Sie65, Wu75]. Forced [Dre75, Fle65, Wil24c]. Forecasting [BDC98, Fin95, Gut88, JSF95, Sey93a, Sey93b, TE96, VLPR20, Par92b]. Form [BCV98, Cli76, Dec65a, DU10, Far85, GW76, Gup75, Koh98b, Lev89, Mar76a, Mey86a, TA90b, TLM90, Tew66, Tew67, Tij81, Var80, Zlo70, AR81, Mar76b, McL86b]. Formal [Boo79, Eck77, McN89a, McN89b, Zie83c, Zie83a, Ber85d, Gra70, Har75]. Formalism [Len80]. Formalisms [BT20]. Forman [Dan71, Mon60]. Formation [GN22, KMH66, KW18, LH19]. Forms [CLWM95, FOTG96, Get96, Her80a, KL97b, Kaz83, LR05, Mey95a, Mil81b, Nei19, Pat19, KL97a, Ost61b]. Formula [Boa70, Che98, CS63, CFG95, KK78, Lev65, Lin59, NS24]. Formulae [Con97, Wah68, Con98, ZK96, Kel96, Ste72]. Formulas [Ask75b, Boa70, For98b, Gar00, HF78, Hue92a, Lan76b, LK65, Lut66, SK76b, Tre22, Wil78b, For98a, Cle69, Wat10, Kaz67]. Formulation [Far85, Gef75, Joh75, Kel88a, Sun76, Vie15, dEJL07]. Formulations [BH76a, Pat03]. Forsyth [Ham62b]. Forsythe [Joh68]. Fortin [Fal93a, Sim93a, Fal93b, Sim93b]. FORTRAN [Din72, Yan94, Nik78, EMUP98, Kri86]. FORTRAN-Scientific [Kri86]. FORTRAN-The [Yan94]. Forum [Bro61]. Fossard [McC79]. Fosset [Hey82]. Foster [Mac68b, Sey93b, BSW82, Sey93a]. Foucault [Moe17]. Foulds [Sha85]. Foundational [Sin88]. Foundations [Bak91b, Bak92, Bru72, Coo65, Eic67, Ham73a, Hol79, Kes85, Ner70, Owe85, Pol87, Ric90c, Sma90, Arb76, Por81, But74a, But74b, But74c, Cha87a, Dat69, Dod73, Gha94, Haa93, Kaz75, Pra68, Ric90b, Rob81]. Four [Bei80, Ell02, FKL^{+}06, Gla93c, Mil94a, TL02, Par87]. Four-Digit [TL02]. Four-Manifolds [Par87]. Fouraker [Ber62]. Fourier [Ano73a, Ask75b, Bai93a, Bai93b, Boa87, GL04, JBL72, JGF87, Kel86b, Kel90b, Kel92b, Mar85a, Pol65, Rus85, Wat63, Wil86a, BS91, Bar22, CE89, GF86, Hen79, Hoc67, Jor80b, Kel79b, Kel92a, Mac67, Per79, RCC78, Rao72, Sae73a, SD74, Sha71b, Sha72a, Sle83b, Smi71, Squ67, Str68, Swa77, Tad87, Tay97, Tre84a, Tur04, Wal70, War98, Boa88, Ask72, Gol73a, Hil76, Jor80c, Kap67, Obe64b, Spi74, Tre84b, Wil77b]. Fourth [GAT75, Hir75, Kob10, Rei63]. Fowler [Won75]. Fox [Bak66, Hor83a, McC85, McG67a, Mul64a, Tho59]. Fracking [CS17]. Fractal [Dev91, EG98, GS04, Goo84, Goo91a]. Fractals [Are86, Are91a, Are91b, Goo90b, Pal94, Dev92, Goo90a]. Fraction [Bru65, Phi68, Phi71b, Sin69, Val67]. Fractional [AS85, AEH^{+}21, BS91, LOT76a, Lew76, LOT76b, MV68, Sri84, UB00]. Fractions [Bla64, HT12, McC91a, McC93a, McC93b, Mer64, Thr94a, Thr94b]. Fracture [Els89, Gos63]. Frames [AH19]. Framework [AEH^{+}21, CN15, CN19, FLGM23, Hug93a, Hug93b, Str88, WS23]. Frameworks [Bai93a, Bai93b, FG19]. Francesco [Lar74]. Franciosi [Hol87]. Francis [Are89, Kha88, Pfl92, Seg80, Thi84b, Zlo91, Thi84a]. Franco [Fal93a, Fra88, Sim93a, Fal93b, Sim93b]. Francoise [Ans85]. Frank [Bau89, Boj97, Bre64a, Gue92, Isr92b, Jar76, Lem83b, Man85, McA94c, Moo74b, Mor88d, O’M92, Pat64, Ros65, Sán83b, Sei85, Tho75, Wol86, You72a, McA94a, McA94b]. Franken [Hey85, Ser94]. Franklin [Bir64, CS64]. Fransén [LS84]. Franz [Fif86, Mul67, Red93a, Rus60, Red93b]. Fraser [Eps59b, Har72a]. Frauenthal [Hop83]. Fred [Bla86a, Cha67a, Dan66, Her71, Kla84i, Luc86, Thr79]. Frederick [Car65b, Hey82, Par92a, Rog94a]. Fredholm [Nas86, Rue73b, Cas72b, Ike72, KK78, Win67]. Fredrick [Rog94b]. Free [Aro78b, Ars97, BTKNZ00, Beu66, BK78, CS90, Ell87, Eva84b, Eva84a, Faz98, Fri79b, GL05, GWG13, HMO97, Kij93, KS68, Mac98a, Mey76a, Mey77a, She60b, TLN95, VB75, WS13, Ell88, Fri82, Leh82, Nas95]. Free-Boundary [BK78, Ell87, Eva84a].

Free-Surface [Mey76a]. Freed [Par87]. Freedman [Dwa72, Fri64, Kni84b, Pin72b, Pit74, Ano99b, The99a, Kni84a]. Freedom [Kel88b, Kel88c]. Freeman [Hag83, Wig90, vdV94]. Frehland [Saw85]. Freiberger [Ham76]. Freidlin [Mer85]. French [Man61c]. Frenzen [Jor87a]. Frequency [AW02, AKP^{+}91b, CL08, CM89, DK89, Deh89, Kel88c, LW95, Lin59, Mor76, Asn75]. Fresh [BEKS17, HRW12]. Frey [Fle73, Mar86b]. Friction [Ste00]. Fridman [Par90]. Fried [Gea80]. Friedlander [Bjö61, Lud77]. Friedman [Cha81b, Cum96, Dum65, Eva84b, Fle72, Fle77a, Fle77b, Gor59, Kap90b, Ock92a, Pul64, Shu80, Sop60, Zad67, Eva84a, Kap90d, Ock92b]. Friedmar [Urb93a, Urb93b]. Friedrich [Kob10]. Friedrichs [Coo66]. Friend [BCD98b, BCD98a]. Friendly [Fra95b, KF95]. fries [Mic92b]. Frisch [NGTB15]. Frish [Jar76]. Fritz [Ask75b, Kaz67, Kea93a, Rox90a, Ano78i, Kea93b, Rox90b]. Frobenius [Cop71, GTH23, Kee93]. Frogs [Mil91a, Mil91b]. Front [Xin00]. Fronts [Kap90c, Kir93a, Kir93b, Kap90a]. Fruits [FS78b]. Fryer [Bou66a]. Fu [Sup77]. Fuchs [Whi63b]. Fucik [Han81a, Ber84b, Ber84a]. Fukunaga [Pao93, YH93]. Fukushima [Get96]. Full [Ben70, Fre82, MBOV17, Ric88]. Fulling [Bou91b]. Fully [FGN13, Kar01, LT98, PSSW06]. Fun [RKM76]. Function [And74a, And85, And87, Ask80, BLT08, Ber92a, Ber76c, BJ76, Boe96b, Boy78, BAD^{+}79, Boy08, Bru65, But78, Car73b, CZ94, CZ96a, Che91, Dam85a, Dam85b, Dam86a, Dam86b, Dam87, EF63, Fel78, Fet68c, Fin71, Fri63, Gas85a, GR85, Ger62, Gla78a, GN78, Gla81b, Gla82a, Gla82b, Gla86b, Gla95d, GS63, HT72, Hil76, Hir78, Hol70b, Ism78, Jag87b, JR87, Jag88c, JvD90, JvD91, Jor83b, Jor90a, JW90, Kar77, Kau77, Kla74a, Köl87, KKS82, Lam69, Lam70, LB77, Lev77b, Los84b, Med62, Meh73, Meh74a, Mil95a, Moo78, Mor75a, Nås77a, NF16, NW70, New03, NJN98b, Pea65, Pei83, Pro85, Rah87a, Raw86, Ric67b, Ros75, Rue84, Sch64, Sch81b, SD74, ST94, SG82, Sim70, Sim71b, Smi82, Sol67, Ste85, Wac89, Wal63a, Wan83, WG76, WG77a, WG77b, WS13]. Function [Wil81, Wil67d, Wix76, Ang90, CZ97, Gro98d, NJN98a, Ric98, Ano89b, Boj65, Bus86, Coh75b, Dum65, EF65, Gas85b, Guh68, Jag76a, Jor78b, Jor82, Kha71, Mil95b, RO66, RD65, Rue79, Sha82b, Zad67, vdM68]. Functional [Ari96, Ber91d, Ber91c, Ces75, CK86, Dat78, Dim91b, Eba82, EKZA96, Far96, Fle76, Hag86, Han87, Hau86, Hau88, HS72, Hol89a, Hüb88, Kel70, Kir10, Kla66, Kle79, Kol63, Lae83a, Lae83b, Ler63, Moo74c, Mor85b, Par81b, Per93a, Per93b, Ral62a, Rog73, RB90, RBB^{+}90, Rox92, Sam86, San65, Sch82a, Sch86a, Sch90, She71, Sop72, Sop79, Sop82, St.82, SC70, SP82, TK97, TF96, Tul90, Wag69, Wil89, JRR71, Nas65, Sch76c, Cor92a, Dim91a, Gei89, Gue79, Hag82, Jd93, Pav94, Sno76, Sop69, Wet74a]. Functionals [BK17, Ewi77, Nas77b, PG00, Pru75]. Functions [Ahm75, Amo78b, And74c, Ano99b, Ars62a, AS80, Ask74, Ask76, AC76, Ask78, Ask79b, Ask86b, Ask92a, Ask92b, BM93, BR94, BB84b, BB88, Bre89, BCZ21, BAH^{+}80, BD63, BS92, BS70, Car75a, Car70d, Cle63, Cod70, Cod78, CBS76, Coo93, Coo94, Cul68, DeM65, DF99, Dub76, DB78, The99a, Eks60, ET99, EH10, Eva90, FO99, Fet84b, FJT19, Fin72, Fis74, FSY65, Gau62, Gla79, Gla87c, Gla93c, Gol66a, GO67, GM66, Gro91b, Gro96g, Han85c, Hoc80, Hoc83b, Hoc83a, Hol81, Hue92a, Hui95, Hye73, Ike76, Ism83b, Ism83a, Ism85, JGA^{+}88, JvD90, JvD91, Jok66b, Jor80b, JC88b, Kam78, KL75, Kap65, Kap81, Kaz67, Kea95, Kla62a, KB68, Köl87, KM14, Kot65, Kre91, Lar69a, Lar80, LOT76a, Law72, Lin78b, Lin79b, Loe92b, Los86b, Lou70, LOT76b, Maz84, McA94a]. Functions [MW76b, Mer84a, Mit65, NF16, Olv72, Olv97, OW95a, Pap62b, PW08, Per79, Pet85, PTK11, Pom67, Rah87b, RS79, Ral62a, Ral69b, Ray69, Rez62, Rie87, RPP88, Ros72c, Rub69, Rue76a, Rue81b, RK78, Rya63, SK79, Sch78c, Seb75, Sha63a, Sha63b, SK20, Sha64b, Sha04, She62, She64, Sho69, Sim96, Smi72, Spa62a, ST98, Sri75, Ste59, Ste81a, Sto66, SM14, Tri95, Tut94, Ung77, VBK22, Wah03, Whe76, Whi66b, WM78, Wil82a, Wim65, Wim67, YT95, Zem68, Cha96b, Gro96b, Kea96a, Lau98, Rue98, Tre13, Ahl82, All78, Ame86a, Ask81, AD87, Bel63a, Bla93, Boa83, Boa87, Bor73, BS80, Bus95, COW96, Chu81, Dav77, Fet65, Gib73, Gla86c, GdMGR91, Han91, Han85b, Hol70a, Jag88a, Jet91, Jor80a, Jor80c, Jor88b, Kar87, Kaz62, Kea96b, Kes65, Kes75]. Functions [Kle76, Kog66, Köl83, Köl87, Kur69, Lan81, LG80, Loe92a, Los86a, McA94c, McA94b, Mil94a, New94a, Rob62a, Rue76b, Sal63a, Shi68, SW81, Sop65, Ste90a, Tha79, Tha82, Var87, Whi67, Wil82b, Wim72, Wim75, Wim86, Buc69, Cho70, Whi63b, Whi66a]. Fundamental [Aar07, BPE96, Dut82, Ebe96, Eck94, Mac69, NF16, SY69]. Fundamentals [Chi94b, Kla84c, KSM95, Pao90b, Ros71, Ber91b, Chi94a, Dob87, Gar93, Gas63, GM87, Kla84j, Mas95, Pao90a, Say79]. Funktional [Gra66a]. Funktionen [Che65]. Furnace [Lew79]. Furniture [Gol69a, Mos66b, Seb70]. Fursikov [Mik89]. Furstenberg [Ano61a]. Furter [Mey95b]. Further [Gol75c, Gol76b, Lut66, Na68, TW69]. Fusion [Gar89]. Future [Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Cas82, Ham76]. Fuzzy [Arb77, Arb78, CT95, MTM96, Pei83, RM96, Tab95, Wan83, Tab97, Gog85, Man81, McA88, McA96a, McA96b, WP88].

G [Aga95a, Ahl93, Ale90, Ame81, Ano82f, Ans64a, Ans66, Ant72, Ant63, Ari96, Bar79, Ber83a, Ber91b, Beu85, Beu92, Bla85a, Bla89b, Bla83, Bla86b, Boa82b, BdD92, BS86, Bra66, Bre81, Bro83a, Bro83b, Bry93, Bus95, Caf92, Caf94, Cha95b, Che86, Cia92b, Col87a, Col87b, Cot73, Cou97, Dan93, Dat73, Day89, De 74, Dei92, Der86, Dij90, Din72, Dra62, Dru64, Dun94, Dwa85, Eas93, Erm89, Eva81, Eva85, Fad62, Fal70b, Fej88, Fin76b, Fir92b, Fir92a, Fle88, Foo63, Fre86a, Fro60, Gar61, Gas85b, Gib95, Giv78a, Gla88a, GG91, GRYL89, Gru74, Gru93, Gue79, Guh66, Gui92, Gur62b, Gur63, Ham61, Han91, Har63, Hau89, Hey86, Him78, Hoc83b, Hoc86, Hoc92, Hol89a, How83, Hun81, Ism76, Jag87a, Jef95, Joh67, Joh81, Jor73, Jor84b, Kv93, Kap85a]. G [Kap85b, Kap65, Kar64, Kaz82, Ken93, Kir76, Kno81, Köl87, Kol65, Kul69, Lan86, Law95, Leh63a, Lev67a, LM81, Los95, Lud77, Mac92, Man61a, May91, May90, McG62b, McG93, McN89a, Meh73, Mey75b, Mey86b, Mey89c, Mil82a, Mil83a, Mil92, Min66, Miu76a, Moo77a, Nas90, Nas65, New87, New90, Nic96, Noh67, Nov94, O’M87a, Ock91, Old62, Olm77, Olv90, Ort96b, Oth82, Pap62a, Par88, Par81a, PT81, Per68, Pol69, Por80, Rau95, Rei62b, Rei88, Rhe60, Ros68a, Ros80a, Rou73, Rus95, Sac74, Sau62, Sch87, Sco85, Sco90a, Sco90b, Sco94, Sei93, Sei86a, SK80, Sha83, She63b, Shu87, Sik92, Sim76, Sim82, Sit92, Sop59, Sop69, Spo85, Sta89, Ste81b, Ste88, Str62a, Tap86, Tay89, Tot91, Tu92, Uhl73a, Uhl73b, Was94a, Wet74a, Whi94, Wig91]. G [Wil82b, WR94b, Wol93, WA63, Wou68a, Yao86, Zie83b, Zwe84b, Zwe84a]. G. [Hoc83a, Kin97, Wei84a]. G.C.D. [Bar74]. Gaal [Moo77b]. Gabeskiria [Sim89].

Gabler [Sin93b, Sin93c]. Gabor [CW94]. Gabr [Hin87]. Gabriel [Bur81b, Yan92]. Gaier [Cav90]. Gajewski [Tad89]. Gajic [Jód96]. Gal [AA82]. Galaktionov [Lev96]. Galambos [Dwa80]. Galaxies [Lin69]. Galcit [Col94a]. Galdi [Ant87]. Galerkin [DA78, GW12, Ike72, PvySF23, RFS03, SHL16, Tad87, Ode86]. gales [Por81]. Gallager [Kul69]. Gallavotti [Wei84b, Wei84a]. Galler [McI71]. Gallery [Fra89]. Galoi [Gra94a]. Galois [Gra94b]. Gambler [Jev94, Tro71]. Gamblers [Gou10]. Gambling [Alb10, Wei79]. Game [Arm90a, Arm90b, Ber76b, Bin84, Bos99, CC23, Dri91, Fra86c, GWZ16, Kon62, Liu92a, Luc71, Mar85b, New66, New68b, Ren93d, Ren93b, Sei62, Sil78, SW93b, SS96, TL02, TW69, TF65, Wei68, Wil77c, Luc87, Pac87, Saa96, Shu85a]. Game-Tree [Arm90a, Arm90b]. Gamelin [Dev94a]. Games [Alb10, Alp85, Art87, Bel69, BI68c, Bol91c, Bom02, BHKL09, CS17, Gil62a, LL69, Mil68, O’R10, Owh17, PCH98, SS24, Tij81, Und76, AA82, Cha64b, Fin93, Fle72, Gri61, Ros72b, Rox77, Shu80, Bol91a, Sca59, Shu85b]. Gamma [Jag76a, Jor83b, Köl87, Los84b, Mei78a, Ros75, Sau78, WW77, WW78]. Gandz [Gol71]. gang [Pin92b]. Ganti [Mor85b]. Gantmacher [Kal70, Mar64, Zar65]. Ganzzahligen [Ham85]. Gaoyuan [Rou91]. Gaponov [Bry94]. Gaponov-Grekhov [Bry94]. GARCH [ASPL03]. Garcia [Lem84]. Garden [DSL04]. Gardiner [Pat87]. Gardner [Fis71a, Hin91, Fir97]. Gareth [Fin95]. Garey [Har82]. Garrett [Mil89b, Wei87]. Garth [Man70, Via85]. Gary [Rei89, Van88, Zim90]. Gas [Ame91a, Ame91b, CS84, Gla96a, KB76a, Kaz86, Mar75b, Ock91, Rog73, Wil64b, GJ97, Fet66]. Gases [CIPB95, Bar95, Uhl73a, Uhl73b]. Gaspard [CGP21]. Gasper [Wim91a, Wim91b].

Gasson [Far85]. Gastinel [Whi76b]. Gatlinburgs [Hou74b]. Gauge [Gue89, Jor89c]. Gauss [Bag95, GJRW24, Koh98a, Par61, Squ67, Tre08, Str68]. Gaussian [Gla93b, Gur66, Ism78, Nag96, Pel97, Ahl62, CZ92, FJT19, Fra65c, Grc11, Kar77, Mil69b, SN96, VDC22]. Gautschi [Seg06]. Gaviano [Jor93b]. Gavin [Col92]. GCD [Kli78]. Ge [Sav18]. Gear [Cyb93, Wil73]. Gebhart [Reh89]. Geert [Bin84]. Gegelia [Kno81]. Gegenbauer [Sne67]. Geisser [Siv94a, Siv94c]. Gekeler [Gea85]. Gelbaum [Gol65b]. Gel’fand [Bre62a]. Gelfond [Rud61]. Geller [Jor77b]. Gels [Bro77]. Gene [Dem86, Dem90, CGO07]. General [AEH^{+}21, Ath76, Bus63, Chu81, CT76b, Enr89, Fly76, GMSW91, GK95, Lev88b, Lou92, McG82, Mor76, Pac86, Ric90b, Ric90c, Sim89, Sto81, Str86b, VBK22, Wim81, Xan87, Zha97, Par10, Mor92, Arb76, Kan62a, Kee95, Pul91]. Generalization [Ang90, Bla87, CR78, Fet68c, Los75, Mal88b, Sha70, Sha71c, Sha71d, Sha73a, Ven74, Ven75]. Generalizations [Com72, JvD90, JvD91, Mer64, SC81, Sch78a, Sch78b]. Generalized [Boa70, BB83a, BB84a, Bus64, CZ92, CH68, Cli64, CH66, Coh88, Coh93a, Dec65b, Dec65a, FTM95, Gar75, Gil70, Gla93b, GMP66, Gre66, Gro88a, HKD20, Kor76, Loe92b, Los89, Lou70, Mar95, Mat91, Mey75c, MP70, MO68, NS24, PR70, Rob74, Rob76, Roh66, Roo64, Ros69, Siv69, Sop65, Tew69, UFW17, Urq68, Urq69, VBC07, Wit22, Won75, YH76, Zem68, Zie73, Coh93b, Gre67c, Ber81, Bor90, Fle85, Kap65, Kor73, Loe92a, McC92a, Mey76c, Rub69, Wat86b]. Generate [Smi88b, Ste64]. Generating [Ano89b, Ber60, BM93, Bus86, FSY65, GR85, Ger62, GS22, Guh68, Jag88c, LS08, MW76a, MB64, Mor98, PW08, REG19, Ric67b, Tu76, WHZG24, Wim72, Wim86]. Generation [Ber92a, Chi79, Hig03, Kel90d, KSM95, LS85, MT76, Slo93, YFS76, Kel90c, Mas95, Ste87]. Generator [PS04, SS78]. Generators [And90, HD62, Mod92, Jun70a, Zas65]. Generic [JH11]. Generics [JRA22]. Genetics [Lud82, Saw82]. Gengzhe [BPK84, Cro85, Ren84b, Rop83]. Genin [Kap71]. Gennady [Pel97]. Genotypes [Ken93]. Geo [Rog96, WR96a]. Geo-Electromagnetic [Rog96, WR96a]. Geodesic [HH16]. Geodesics [Woo96]. Geodetic [Ale85, KN98]. Geodynamics [Pit89]. Geoffrey [Dur90a, Fro85, Per93a, Dur90b, Per93b]. Geomechanics [Fos90]. Geometric [BP76b, BPP95, BB83a, BB84b, BB84a, Bos95, BBK20, Car70c, Con94, Cor92b, CB09, DeR91a, DeR91b, Dia59, Duf70, Eck80, EW75, Far93b, Fir92a, Haj96, Het72, HLB95, Hug93a, Hug93b, Jon92b, JH98, Kan88, Ler64, MNC71, Mec92, MS96b, MQ95, Neu96, Pap86, Pet95, Pet76b, Pro84a, Qua95, Sed96, SAH96, Sta88, Sug76, Sug77, Tur04, VC06, WS23, Sei97, Fir97, Alp85, BI68a, Bla89a, Far93a, Fir92b, Gra83a, Jon92a, Pet95, Pro84b, Syl96, Wol93]. Geometrical [Cha63, GN13, Hol87, Kri76, Lud75a, Mar66, Wil62a, Ber85a, Con85, Olv93]. Geometries [Bru72]. Geometry [ABGJ21, Bél88, Boe88a, BM95, Cha95a, Duf91b, Fra88, Goo84, GH01, Gru88, Gue89, Hei95, Her80b, Her83a, Hof95, Hug93a, Hug93b, Iva04, JHSB22, JH11, KL08, Kla84c, LLMM14, Mag86, Mel97, Mul75, OH95, PF13, SA75, Som93, Wey62, Wit22, Cha96a, Ham77, Ban85, Cha96c, Dev91, Her83b, Kla84j, Kla92b, Lee82a, Lee86a, Far85, JW83]. Geometry-with [Mag86]. Geophysics [Car70e]. Georg [Koç89, Sai91a, Sai91b]. George [Ath94, Bir86, Bla90, Bri64, Bur79, Bus86, Cab75, Cho70, Col86, Dat67, Duf84, Fat93, Fre78, Giu18, Gos83b, Gre89, Ham79b, Har61a, Hop79, Joh68, Kv93, Lin91a, Maw93a, McA88, Mey96, Mro91, Rei64, Ton95, Whi76a, Wim91b, Lin91b, Maw93b, Wim91a]. Georgi [Rub69]. Geospatial [FP21]. Geostatistics [Cre97, DK97]. Gerald [Ber93, DeR91a, Dev91, Fla67, Fra95b, DeR91b]. Geramita [Mil81b]. Gerard [Dan91, Hut91, Sch95, Van94c, Van94b]. Gerasimenko [Bak91b, Bak92]. Gerasoulis [Jor89a]. Gerhard [Bei96, Eas91, Jes93b, Mar95]. Gerhardt [Gra86a]. Germund [Jac76]. Gerschgorin [Far60]. Gershgorin [Mel12]. Gesztesy [Bro90]. Gets [BWYY96, YY96]. Getson [Zlo91]. Getting [TFL23]. Gevirtz [Gla88a]. Gevrey [Lie95, RL95]. Gewirtz [Ull88]. Ghaffari [Hol87]. Ghaoui [Yak95]. Gheorghe [Pic90, Wib87]. Ghez [Kaz89]. Ghizetti [Ste72]. Ghosts [Jef18]. Ghouila [BI68c]. Ghoussoub [Maw95]. Ghurye [Man61a]. Gian [Bru72]. Gianfranco [Wil91a, Wil91b]. Gianni [Bro83a]. Giardina [Sta88, Sta90]. Gibb [Pet98]. Gibbon [Sco84b, Sco84a]. Gibbons [Cum92a, Gil95, Gup82, Man89, Noe72, Cum92b]. Gibbs [CS97, GS97]. Giessen [Hav95]. Gihman [Kus74]. Gilbert [Bha93, Cab75, Hen86, Mea86, Pro85, Ros82, Tut94, WM85]. Giles [Coo65]. Gill [Kow83, KZ93, McK95, Yau68]. Gilles [Buc91a, Buc91b]. Gillespie [Klo93, MR18]. Gillies [Jan70]. Gilliland [Han64]. Gilman [Che18]. Gino [Whi66a]. Ginsburg [Boo79, You71]. Ginzburg [DGP92, GP02]. Giovanni [Ant87, Wei84b]. Girault [Par67b]. Giri [Rei79]. Girson [Fre83]. Gitterman [Gue83c]. Gittins [Ken86a, McC91b]. Glaciology [Pit89]. Gladwell [WR94b]. Glashoff [Kor85]. Glass [Lew79]. Glassar [Ism96b]. Glasser [Abb92, Boe76, Boe95b, BC96a, Boe96b, Car76b, Chi71, Fet68a, Fet72a, FLR^{+}81b, GK86a, GH86, Gra79, Gro90, Jag82, Jag88a, Jor82, Jor84a, Jor90b, LG80, Lar94, Mil94a, OW89b, Ren89, Sha82b, Ter92].

Glasserman [Sha95c]. Glassey

[Kin94b, Kin94a]. Glazman [Ral62b]. Glenn [Gla97b]. Glick [Ort96b]. Glimpse [Sch92]. Gliner [Kor65]. Global [AK90, AKP91a, AKP93b, AR78, AR80, BW11, CBM98, Ewi77, Fic60, FH76, GG80, Hag83, Her83a, Hur93a, Hur93b, KH98, Ken74, KLA95c, KG98, Lee82b, MW76b, MY97, MW97, NJN98b, NW97, PR86, PH78, TtK96, Ton89, WP96, WS13, NJN98a, Aga95a, AKP93a, Cra89, Fif86, Her83b, Kov96, Mei89, Pal96, Sch77a]. Globalization [PSSW06]. Globally [Wat86c]. Glover [Bal93]. GMRES [Emb03]. Gnedenko [How63, Rei62a, Tak77]. Go [BWYY96, YY96]. Goal [KD76, LW13]. Goal-Oriented [LW13]. Gockeler [Gue89]. Godunov [Aro69, HLvL83]. Godunov-Type [HLvL83]. Goh [Bra82]. Gohberg [Bar83b, Joh88, Lan81, Bar83a]. Goldberg [Lew83a]. Golden [Day89]. Gol’denblat [MT64]. Goldstein [Los74a, Mar87, Rei64, Wou67]. Goldstine [Sag82]. Golf [Sch75a]. Gollub [Bry91a, Bry91b]. Golub [Dem86, Dem90, CGO07]. Golubitsky [Mey86b, Mey89c, Ste94a, Ste94b]. Gompertzian [WHMN16]. Gondran [Cha86, Raz87]. Gonzalez [Ros80a]. Good [BCD98b, Har66a, BCD98a]. Goodwin [Sal63c]. Google [BL06]. Goose [GWZ16]. Gopal [Tak83]. Gorbunov [Ock91]. Gordan [SS79]. Gordon [Bur94b, Bur94c, Mey75b, Rei64, Whi76a]. Gossel [Ber84e, Ber84d]. Gotskalk [Hin89]. Gottfried [Lil68]. Götz [Moo85]. Gould [Fre86c]. Goulden [Whi84b, Whi84a]. Gouri [Mea86]. Gouvernés [Fal70a]. Governed [Ber91a, CS90, Don90, Pea81]. Government [Gas60]. Govindarajulu [Mal77]. GR [Par10]. Grabbe [Gre60a, Kar60]. Grace [Cle69, FSW^{+}66, Sch91a, Sch91b]. Gradient [Bra96, CN96, CMSZ24, Ehr76, GO89, HB96, LT13, Naz76, Naz86, O’B11, Pug66a, She09, SXY19, Und76, ZM22]. Grading [VSM14]. Graduate [Kir12, PAB^{+}01, SIA01]. Graeme [Gra18]. Graffi [Abl84b, Abl84a]. Graham [Boj65, Cor92b, Day89, Los74a, Mor82, Tis83b, Bou65a, Tis83a]. Grand [Gli91b]. Grandell [Ruo91b, Ruo91a]. Granular [GWG13]. Granville [Gup89]. Graph [Ada97b, Bei78, Bei96, BGH^{+}04, Bou77, BDX04, BLS92, CR96, CPW12, Che67a, CHW12, Deo64, Erd72, FS73, FLNU18, GMP05, GBS08, Har60, Har62, Har64a, Har73, Har83, HRB96, Hul76, LMSM83, LMSM84, Man89, May91, New91b, Pil87, PO76, SBXD06, Tew67, Wol86, Xu21, Ada97a, BBLS92, Bei85, Har74, Wal79, You72a]. Graphes [Ham87]. Graphical [Ano99b, The99a, KM14, Mil89b, Ros76b, God87, Moo74b]. Graphics [Ber85d, PFM76, WB09, Boe88a]. Graphing [Fra86b]. Graphs [BR97, BF16, CS16, Chu89, CT76a, De 20, DMGL22, Har67a, Kar22, KK99, Lim20, Mac61, Mar60, McM87, Min66, Par61, REG19, San76, Sou76, Tuc73, Bei81, God87, Hey82, Leh63a, Cha86, Tar81, Whi83]. Grasman [O’M88b]. Grass [REG19]. Grass-Hopping [REG19]. Grassmann [AMSV02]. Graves [Wet74a, WGM93]. Gravitating [Lin69]. Gravitation [Gol65a]. Gravitational [Agg74, Auc97b, Bou76b, Gro94a, Gro96e, Kan62a, Kla92a, NS02, Pul91, Sey93c, Sla64, Sut95, Auc97a, Gro96a, Bou65b, Bou76a]. Gravity [Bro93, Gro88b, GAB^{+}88, Kla60, Kon03, Qui92b, Qui92a, Sti87, YFS76, Gue89, Dij90]. Gray [Fri71, Ran89, Shi94, Wei88, GN22, Sav97]. Graybill [Bir64]. Grazing [MV93]. Gräztzer [Mey96]. Great [Bra91a, Bra91b, SS88, Dra61a]. Greek [RR97b, RR97a]. Green [Sma76, ET99, Fin71, Hoc83a, Kre91, Lou70, ST94, Kap81]. Greene [Asp84a]. Greens [Asp84b, Hoc83b]. GreenSpan [Per76, Fle62, Hoc82, Por72]. Greg [Sag83]. Gregory [Bry91a, Dur92a, Fro85, Kaz83, Leh82, Bry91b]. Grekhov [Bry94]. Grenander [Dwa64, Kee95, Pin87a]. Greville [Mey76c]. Grid [Bra76a, KSM95, Mas95, VH78, ZD95, CD93, Str62a, Del95, Man88a]. Grid-Point [VH78]. Grids [CHL08, Zha97]. Griewank [Fat93]. Griffiths [Her84b, Mal85, Her84a]. Grimmett [Dur90a, Dur90b]. Grimshaw [Cha94a, Cha94b]. Grindrod [Ter93a, Ter93b]. Grisvard [Sch86b]. Grit [BCG77]. Gröbner [Stu94b]. Groebner [Stu94a]. Groeneboom [Bic94]. Groenewod [Law65]. Groetsch [Col94c, Nas86, Col94d]. Grosjean [FSP93b, Los89]. Grosswald [Wim80]. Grotschel [Bla89a]. Grouchko [Sau79]. Groun [Sau79]. Ground [Pap63b, WPL98]. Groundwater [GG95, Gil95, Hil96, SH96]. Group [Blu94a, Bre62b, Den90, Der71, Dys66, Gra94a, Gra94b, Har64a, Hol91a, Isr92a, KM12, Kir12, Mey75c, Mla97, RKDD14, SS91, Stu64, Tre82, Hol91b, Isr92b, Zei72, Blu86, Jor89c, Mit87, Pro65a, Var87]. Grouped [Har63]. Grouping [Eub88]. Groups [BKS66, BKS68, BK69, Gru74, Gru88, Her80b, Hof71, Joh84a, Kum88, Mey86b, Mey89c, Sch92, Zas65, Cur64a, Joh84b, Mil81a, Tur60]. Growth [Ada91, CT75, CK76, DM07, Ham71, Jea78, JC12, LG74, LG75, Lew83c, Ngu76, RCM07, Sma83, WHMN16, Lau98, Bél88]. Grusa [Ber91a]. Gruver [Mil83c, Mil83b]. Gruyter [Bre94a, Bre94b]. Guaranteed [RFP10]. Guard [GF76]. Guckenheimer [Kaz84b, Kaz84a]. Guddat [Cra84, Wet93]. Guerra [Wet93]. Guerrilla [Abd90, Lew91b]. Guessing [BHKL09]. Guiasu [Gos79]. Guidance [Gur63]. Guide [Bak91a, Ber93, Bor06, CT95, CWW99, DeR91b, Flu93a, Flu93b, Fra95b, Har82, Hig01b, KF95, Kri85, McC85, Smi80b, Sta89, Tab95, DeR91a, Par94, Par81a]. Guidebook [Kri85]. Guided [BS98, Ham62a, LH19]. Guiding [Dre60]. Guillemin [Wig85]. Guillermo [Shu85a]. Gul [Kow88]. Gummed [BFF^{+}99]. Gundersen [Kra67]. Gunter [Bai74a, Che65, Gil65b, McC83]. Gunther [Wil93, Jet91, Vog87]. Gunzburger [Bak91a]. Gupta [Gut82, Mar76b]. Guri [Svo96]. Gurtin [Ant83b, Ant83a]. Gusev [Har72b]. Gusfield [Kri91]. Gushing [Dri81]. Gustafson [Ant82, Kor85, Syn83]. Gustafsson [Sül97]. Gustavus [Bla84]. Gut [Sie89]. Gutman [Rou88]. Guttag [Orb18]. Guy [Kla92b, Roc92a, Roc92b]. Györi [Sei93].

H [Abl80, AKP91a, AS85, ASF86, Alf93, Alp85, Ant91, Ari96, Ars62b, Ash89, Asn75, Asp84b, Asp84a, Bai87, Bai74a, Bai76, Bar65, Bél88, Ber62, Ber65a, Ber91b, Ber86a, Ber78, BW65, Bir69a, Blo91, Boa82a, Boa83, Bou76a, Bou86, Bre68, Bro90, Bru72, Bru83, Cab74, Cam91a, Cas72a, CGO07, Che92a, Chi79, Cla64a, Coh75b, Coh84a, Coo91, DV76, De 74, Dem86, Dev92, Din72, Dra61b, Duf91b, Dur87, Fel72, Fox71, Fro87a, Gar61, Gen86, Gen96, GC79, GM87, Gol65b, GGG69, Gol73a, Gra86a, Gra66b, Gru88, Gua92, Gue82, Gur62a, Ham90, Han83a, Har61b, Hei97, Hig62d, Hol87, Hop79, Hor93, Isa88, Isa89a, Ism83b, Ism83a, Ism85, JBL72, JGF87, Jag90, Jag92, Jef94, Jin91, Joh83d, Joh83b, Jun64, Kan62c, KJ87, Kaz92, Kee88a, Kel75b, Kel86b, Kel90b]. H [Kin94c, Kla92b, Kor73, Kow83, KZ93, Kra68, Lan81, Law95, LeV77a, Lea94, Leh63b, Lei66, Len95, Lin80, Lud91, Mal85, Mal90, Man85, Mar81, Mar86a, May84c, May84d, May84b, McG64, McN59, Men94a, Mey82a, Mil82a, MT61, Mla94, Mor96, Nas87a, NW81, New74, Noe81, Noe82, Nov94, O’M96, Pac86, Pac88b, Par66, Par69, Par72, Par76, Par92a, Pay68, Pet95, Pim93, Pol59, Pre81, Pro96, Rab94, RCC71, Ral62a, Ran96a, Rat87, Rat93, Rei71c, Rob62b, Rog60, Ros78, Row76, RSTW73, Rue74a, Rue76b, Saa88, Sag77, Sag82, Sam76a, Sca59, Sch73a, Sch94, Sch76c, Sco84b, Sco84a, Sco92, Seg85, Sel76, Sha82a, Sha95a, She63b, Shi93, Shu87, Slo93, Ste73b, Str82, Str62c, Str62b, Tay96, Tre65, Tut94, Uhl73a, Uhl73b, Van66, War92, Wat65, Wet93]. H [Whi81, Wil72, Wim86, Won83b, Won83a, Yan94, Zak85, Zas65, Zwe84b, Zwe84a, vL86a, BS94a]. H-P [BS94a]. H. [Dur84a]. H.-J [Jag92]. H.-O [Dev92]. Haack [Sch71a]. Haag [Fre86a]. Haake [Kea93a, Kea93b]. Haakon [Thr94a, Thr94b]. Haberman [Fre79]. Hackbusch [Man88a, Rao94, Slo97]. Hadamard [Los74a, Mil81b, Cop71, GG73, Sch73d, Sch73e, Sch73f, Sch74]. Hadamard-Type [Los74a, GG73]. Hadley [Guh66, Kar64]. Hagedorn [Nit90, Pro65b]. Hageman [Por83]. Hags [Sei85]. Hahn [Ber93, Col73, Wou64b, DM67, Lev67b]. Haigh [Grc18]. Haight [Lem83b, Pat64, Lem83a]. Haim [Pan93a, Pan93b]. Hainsworth [May90]. Hairer [Ale90, Cam91b, Cam91c]. Hajek [Rox77, Pow80]. Haken [Seg85]. Hal [Cus96]. Halanay [Beb68, Con96]. Hald [Bar91a, Bar91b]. Hale [Cho94, Dat78, Mey72]. Halfin [vLMZ19]. Halfman [Poh63]. Halftoning [She09]. Hall [Eas93, Tur60, WJV79]. Halley [Mel97]. Hallway [Mos66b, Gol69a, Seb70]. Halpern [Gue83c]. Halton [KJ87, Owe06]. Halvorsen [Hin89]. Hamdy [Rub77a]. Hamilton [Fle85, Dec65b, Eva85, Red89]. Hamiltonian [Blo89, Eas93, Kup94, Laz91a, Mey88, Rat87, Rat92a, Ske95, Cla78b, Cla79b, EW75, Jür86, Laz91b, MA04, Mos86, Rat92b, SSCS95, Stu90].

Hamiltonians [PY00]. Hamiltonicity [Kar22]. Hammer [BI70a]. Hammerline [Wil93]. Hammersley [CS70, Jun65]. Hanan [Are91a, Are91b]. Handbook [Bra94a, CLH97, CT95, Hei97, Hoc83a, Kla97, Lat90b, Moo74a, PD97, Spä97, Tab95, Ver09, Shi97, Bru86, Hoc83b, Ask79a, Bra94b, Gre60a, Hin94, Kar60, Kog66, Lat90a, Par72, Ros78, Sco94, Sik92]. Handicapping [Sch75a]. Handing [McA92]. Handler [Bur81b]. Hands [Lon87]. Handscomb [Jun65]. Hanging [VK84]. Hanif [Bla82]. Hanke [Bra96]. Hankel [Cor72, SM14]. Hannan [Fro87b, Van76]. Hans [Chu94b, DK97, Din72, Fal68, Gea74, Han83b, Olv91a, Pay68, Pul91, Chu94a, Olv91b]. Hans-Paul [Han83b]. Hansen [Los86a]. Hanson [Gen76]. Hao [Hol86b]. Harald [Kaw68]. Harary [Moo74b, Sei85, You72a]. Hard [Shi17]. hardcover [BC96b, Hig05]. Harding [Sch94]. Hardle [Pin92b, Pin92a]. Hardy [Dur93a]. Hare [Emb03]. Harman [Deu63, Jon92c]. Harmonic [BF96, Col84, Fin76a, Kla62a, MTGF78, Maz84, Mei77, Sch75b, Ter82, Rue98, Jor87b, Jor89b, Mit65]. Harmonics [Sah95, Sha70, Sha71c, Sha71d, Sha73a]. Harold [Ask78, Ask79a, Coh84b, Dun94, Fab80, Fle79, Gra86b, Jag64, Jow76, Man61a, Mor68, Wri60]. Harriet [Bai76, Gri88]. Harris [BR65, Lil67, MT61, San82]. Harrison [Kni86]. Harry [Ano61a, Dur84b, Gar67, Hye73, Nas87a, Roy93a, Sle74, Yan76, BL10, Roy93b]. Hartley [Nau86, Sho69, Wan91, Wel88, WA63]. Hartman [Bra65, Han75b]. Hartmanis [Mul69]. Hartmut [Jow86]. Hartog [Set62a, Set62b]. Hartsfield [Bei96, May91]. Harvested [FS78b]. Harvesting [Sán77]. Harvey [Par92b, Sco83]. Hasan [Col76a, Hol86a, How82]. Hassan [Man88b].

Hassard [Mey82a]. Hasselblatt [Swa96]. Hastie [McC92a]. Hauer [Meh73]. Haug [Duf86, Koh88, Pim93]. Haugen [Rei69]. Hauptmanns [Ruo92a, Ruo92b]. Haussmann [Fle88, Ism76]. Having [Hop75, RKM76]. Hawkes [Lim23]. Hawkins [Ort96b]. Hayes [Bre70, Chi95, Göb70, Sal64]. Hazard [Ben79]. Heading [Lev63b]. Health [GAM93, Tud92]. Hear [CEH12, Pro87b]. Heart [WiOH10]. Hearts [Bat21]. Heat [BK88, CZ92, CH68, Día78, Gla93b, Gro96g, GWK95, GW95, Hai77, Hui95, Hui96, Jag87c, JK88, Jag88b, Kaw95, Kev85, Kla94a, Let69, Log98, MM10, Pra85, Pra86, Pra89a, Vel71, Wil59, Wil70b, ZC74, vL86a, Gro96b, Hui97, May90, Ros87]. Heat-Flow [GW95, Kaw95]. Heating [Hey63, Lew69, Lew71, VN76a]. Heavy [Hey86, New87, vLMZ19]. Heavy-Traffic [vLMZ19]. Hebrew [Gol71]. Heck [Koe95]. Hedayat [Olk72b]. Hedging [CD03]. Height [CF76]. Heijmans [Ser96]. Heinrich [Ork95]. Heins [Whi66b]. Heinz [Bla88c, Goo91a, Koç89, Pon90a, Sül97, Wil93, Pon90b]. Heinz-Otto [Goo91a, Pon90a, Sül97, Pon90b]. Hejtmanek [Zwe84b, Zwe84a]. Helffer [Sch89a]. Helgason [Her80b, Zal83]. Helices [Nie96]. Hellwig [Ans66, Gil65b]. Helly [Mac76a]. Helmholtz [BS00, DGK16, GS23a, GZ19, KR74, Lin10, MS14, OT12]. Helmke [Kod96]. Helmut [Jow86]. Helps [SW22]. Hemisphere [CS84]. Hengartner [Kes75]. Henk [Hig00, Ser96]. Henkel [BdD93, GRL87]. Henley [Har74]. Hennequin [Dur87]. Hennings [Vog93]. Henri [Gog85, McA92]. Henrici [Boa76, Boa78, Boa87, Gre62a, VJ83, Wil64c, Wil65b]. Henry [Bel90a, Bla86a, Bru72, Dum65, Gus90a, Han85b, Man61a, Mey79, Pul64, Sam80, Sno76, Tek89, Zad67, Dwy63, Gus90b]. Henry-Labordère [Mey79]. Herbert [Bos73b, Eas91, Her70, Hir71, Hul61b, Lin79c, Moo69, Rhe88, Str62a, Whi80, Lie97]. Herglotz [GGK96, OHBG00]. Herman [Hun81, Sag82]. Hermann [Cur64a]. Hermans [Col87a]. Hermes [Sno76]. Hermite [SM79, CM78a, CMSM79, Cha97, Gau87, Lam04, VG96]. Hermitian [Bos87, Gol75c, HS88, LR05, Ros69]. Hernandez [Hüb91]. Hernandez-Lerma [Hüb91]. Herpes [Tud90]. Herrera [Sag86]. Herron [Gro91a]. Herve [Bas88]. Herzberger [Moo85, Wil62a]. Herzenberg [Pap62b]. Hesselink [Lea94]. Hessenberg [Cho69, Fai71]. Hestenes [Cul77]. Heteroclinic [BRMR08]. Heterogeneity [AMS^{+}23, WHMN16]. Heterogeneous [Xin00]. Heterophilious [MT14]. Hettmansperger [Goo86]. Heuristic [Arm90a, Arm90b, Coo64, Sie90]. Hexagonal [Wil76b]. Heyman [Art83b, Art85b, New87, Art83a]. HHD [SAT22]. Hidden [Aur18, Car70d]. Hierarchical [Ber88a, Kat78, Owh17, Vas97]. Hierarchy [Cha00, Giu18]. Higgins [McG64]. High [BS00, BF16, Bru65, Bur78a, CL08, CM89, DW95, DG14, FR16, GST01, Her91, Hig00, Hig05, Jun64, Kla84h, Kla84a, LW95, MMT^{+}21, Mor76, Mye98, Ock91, Shu09, Sun76, Tur78, VDC22, ZWGB16, ZKP17]. High-Accuracy [Hig05]. High-Dimensional [VDC22, ZWGB16, ZKP17]. High-Frequency [CL08, CM89, Mor76]. High-Order [Bur78a, DG14, GST01, Her91, Tur78]. High-Performance [Hig00, MMT^{+}21]. High-Speed [Ock91]. Higham [Bra94b, Bra94a, Ste97b]. Higher [Aga95a, BGL17, BGHS23, CL75, Fin76b, GS67, Her78b, Riv60a, Yak94]. Higher-Order [BGL17, BGHS23, CL75, Her78b]. Highly [AC78, RLT65]. Highly-Damped [RLT65].

Hilbert

[Ant82, Coo73, Dia64, Erd83, Eva85, GA63, Hey97, Kha86, Ral62a, Ral62b, Ros68b, Sib95, Syn83, Syn93a, Ter92, Vog87, Vog91a, ABS95, Bur82, FR66, Fri80, Gla91, HL76, Lev65, Sae70, Sae73a, Sae73b, SR98, Syn93b, TS66, Vog91b, Fen94]. Hildreth [Ros85]. Hill [Blu94a, Mat88]. Hillier [Hey82]. Hills [Kap77, Wal63a]. Hinkley [Rei77c]. Hino [Cor92a]. Hinton [Spi84b, Spi84a]. Hiroshi [Çin93a, Cin93b]. Hirotaka [Cra86]. Hirsch [Gus76]. Hirschel [Ork95]. Hislop [Hel97]. Historical [BCZ21, BS92, Har60, Har76a, Hou74c, HMH00, Pol73, Rei78a, Sto81, Ypm95]. Histories [WJV79]. History [Bar89, Bar91a, Bar91b, Bur94a, GO89, McC93b, Pan97, Ros71, Sag82, Spi91a, Spi91b, Ste93c, Ung85, McC93a, Ros73]. History-Selected [Ung85]. Hit [Dun64, Wol60]. Hitiris [She77b]. HIV [GAM93, PN99]. HIV-1 [PN99]. Hjorth [Ber95]. Hnizdo [Jor90c]. Hoaglin [Par92a]. Hochberg [Spu89]. Hochstadt [Hye73, Yan76]. Hocking [Fad62]. Hockney [Bun83b, Bun83a]. Hodge [Lim20, SBH^{+}20]. Hodgkin [Mah75]. Hodiri [Keh89]. Hoeffding [Man61a]. Hoegh [Bro90]. Hoegh-Krohn [Bro90]. Hofbauer [Hal90a, Hal90b]. Hoffman [Jon92a, Wil93, Jon92b, Koc92, Küç92, Lee97b, Lee97a]. Hofri [Bla88a, Fla97]. Hoggatt [And75a]. Hohn [Mul67, Rus60]. Holcombe [Gos84]. Hold [CU03]. Holden [Bro90, Fis76]. Hölder [Wan79]. Holes [DeW85]. Holl [Fro60]. Holland [Chu81, Sam76a]. Hollig [Ron95]. Holloway [Rab94]. Holmes [Kaz84b, Kaz84a, O’M96]. Holmgren [San95]. Holomorphic [Sha04]. Holschneider [Wal97]. Holstein [Bai87]. Holtzman [Fle76]. Home [Nie89b]. Home-Made [Nie89b]. Homing [GK15]. Homoclinic [Kra96a, PTK96, Pal96, WP96]. homogénéisation [Lan86]. Homogeneous [Día78, Gle81, Gle83, Gle85, GT10, KK75, New76f, New78, Wal92, Wal93c, Wal93b, Lau74]. Homogenization [Sle87]. Homology [FP21]. Homotopic [Bha98]. Homotopy [Wat86c]. Hong [Gro96h, Pré95b]. Honor [Man61a]. Hooke [Kaz91a]. Hoop [Shi97]. Hoorn [Fro87a]. Hopcroft [Gra70]. Hopf [Mey82a, Mey88, Sat78, Abe88, Asm98, PEGK03]. Hoppensteadt [Mil94b, Sán83b, San83a]. Hopping [REG19]. Horizons [Wet74a]. Horizontal [Dub97b, Dub97a]. Hörmander [Smo91b, Wil85a, Wil86a, Smo91a]. Horn [Nag88, Ple88, Ste93b]. Horseshoe [BD94]. Horst [Bei81]. Hoschek [Bos95]. Hospital [MM95]. Host [ZWY14]. Hot [CS84, SR94]. Hot-Water [SR94]. Hotelling [Man61a]. House [Gri77, Mil91a, Mil91b]. Householder [Fra72, Gre72, Var66]. Houwen [Wil78b]. Howard [Art85a, Jar76, Kar83, Wei77]. Howell [Sta92b, vdM68, Sta92c]. Howes [O’M86]. Howie [McN89b]. Howlett [Pao88]. Høyland [Mee96]. Hrusa [Daf89]. Hsiao [Ame91a, Col87a, Ame91b]. Hsiung [Lee82a]. Hsu [Ton89]. Hsuan [Zlo91]. Hu [Joh83f, Jor73, Joh83e]. Hua [Hab83b, Mit65, Hab83a]. Huber [Rei83a]. Hubert [Coo66, Ede90a, Ede90b]. Huberty [Gup96]. Hubs [DZH^{+}04]. Hudson [Dat86]. Huffel [Fie93]. Huge [Cen81]. Hughston [Lou92]. Huguet [Mat94b, Mat94a]. Hui [Gro96g]. Huilgol [Obe77]. Hull [ZKP17]. Human [BK83, Iva04, Sel76, She09, Tud90, WiOH10]. Hundred [Hig02c]. Hung [Arm90a, Arm90b]. Hunt [Duf91b, Goo91b, Sam76b]. Hunter [Hai85, Rau95]. Huntley [Gue83a, Gue83d]. Hurley [Hig62b]. Hurst [GS04]. Hurt [Pro84b, Pro84a]. Hurwitz [Hou68a, Hou68b]. Huseyin [Cho94, Gol66b]. Husler [O’B95]. Hussaini [Boy88]. Hutton [Ran89, Wei88]. Huxley [Mah75]. Huygens [Kaz91a]. Hwa [Wri60]. Hya [Dau90]. Hybrid [CG24, EGJK04, Fal93a, Fix76, GT10, MMB^{+}07b, Red76a, SWM^{+}07, Sim93a, Sim93b, MMB^{+}07a]. Hydraulic [Pet78]. Hydraulics [Hol90a, Hol90b]. Hydro [Ber60]. Hydro-Electric [Ber60]. Hydrocarbon [Ode82]. Hydrodynamic [FEY61, HM78, Kir75b, Kir75a, Mil82a]. Hydrodynamics [Col87a, Hol66, Lig76, Pru00, Cha87a, MT62]. Hydromagnetic [Elu88]. Hydromechanics [Kan66, Mik89]. Hyman [Smi83]. Hyper [Dob84, Mac85, MW61, She63b, Sta85]. Hyper-Cube [MW61, She63b]. Hyper-Volume [Dob84, Sta85]. Hyperasymptotics [Boy05]. Hyperbolic [Ame78, BF96, Car80, CFL^{+}80, Cha78, CL75, DD04, DA78, GRL98, GAT75, HLvL83, Hig86, Kov96, Kre91, LA75, Lin79a, Pal95, Sar84a, Sco59, TtK96, Tre84b, Tre84a, Wan91, WP95, YF78, Sar84b]. Hyperbolic-Elliptic [LA75]. Hyperbolicity [PTK96, Kra96a]. Hypercube [Tre17]. Hyperfunctions [DC86]. Hypergeometric [ASF86, Als86, And74c, AC76, Ask78, Ask79a, Ask85, Ask86a, Ask86b, AD87, Ask92b, Dam85a, Dam86b, Dam87, FL68, FL70, FL84, Gas85b, Gas85a, Gla86b, Gro88a, Ism78, Ism85, Kar77, Lan76b, Loe92a, Loe92b, Los89, Rah87a, Rah87b, Rue76b, Rue76a, Rue81b, Rue84, Sau62, Sri75, Sri85, Wil82b, Wil82a, Wim91a, Wim91b, Ask92a]. Hypergeometric-Type [Gro88a, Los89]. Hypergraph [UA07, VBK22]. Hyperspherical [ZWGB16]. Hypersurfaces [SG73]. Hypervolume [Bog73]. Hypotheses [Kup62]. Hypothesis [LMDLS23]. Hysteresis [CMR23, MNZ93]. Hysteretic [Ell90].

I- [MW97]. I. [Thi84a, Art83b]. I.U.T.A.M [Cla62]. I/II [Kau86]. Iain [Hig00]. Ian [Ant91, Art91a, Art91b, Ass77, Bol91a, Fra88, Guc79, Mey89c, Rob85, Tul90, Whi84b, Zie78, Bol91c]. IBM [Kaz92]. Ibragimov [Abl80]. Ice [Mul79]. Idea [Van94a, Van94d]. Ideal [Gel79, Sim96]. Idealized [All75, All76, Les76]. Ideals [Dem74, Som93]. Ideas [Dra61a, Ham79a, Her93b, Her93a]. Idempotent [KF76]. Identifiability [MXPW11, MPW23, NW82, Lah94]. Identification [Bau91, DH09, Don90, Gil62a, HKS75, MS91, MM89, NS75, YS76, Wib87, You72b, Mor85b]. Identifying [EBTZ94, ETZ94]. Identities [And74b, Ben64, Ber96, Bre64b, Car66b, Car73b, Cos93, GRLR87, HL86, Kap70a, Mor88a, Mor89a, Mor89b, NdA14, NY90, Raw83, XRTM95, Ber97b, Boe96c, Bre96, Kra96b, Sch96, AG67, And74a, BKS96, Bre81, Bro65, GRL87, HA71, Jor84c, MR65b, Wol94a]. Identity [ASF86, Als86, And85, And87, BD73, Bar76a, Bar75, Bha98, BSD^{+}86, Boy80, Bro60a, Car64, Car67a, CFL^{+}80, CR82, CR83, Dam86b, Dam87, Dam88, Fet84a, GLL^{+}88, Gla85c, Gla88b, GK86b, Gon06, GM92, Gro88a, GV97, JW90, Kar77, KM80, KR85, Lin79a, Lov79b, Moo87, NÖ79, Phi71a, Rah87a, Ric89, Roh70, Roh71, Rue81a, RL81, Rue81b, Rue82a, Sch71d, Sho78, Sib91, Sle62, Sri73, Sri85, VG96, Wac89, Wil82a, Wil92, WGLS92, Wol72b, WG93, vD81, Cha97, Len93a, Wil98b, Bos73a, BS86, BMMO76, Car80, De 74, GD88, GH86, Ism78, Jag73, Jag80, Jor90a, LFR80, Los89, Olk72a, Rie72, RFC72, Rue74a, RLG81, Rue82b, Sha83, SR79, WGM93, Wil82b]. Identric [Neu96, Sei97]. IEEE [Hig02c]. IFAC [LW64]. IFISS [ERS14]. Igor [Ott96, Rob92a, Rob92b]. ihre [Che65]. II [Art85b, But74b, Ell85b, Haa93, Kla92b, Lax89, McG62b, Olv72, Par72, Poh63, Ral62a, Rei63, Sco94, Spi91a, Vak93, BKS68, Cha00, Kla82, Lam70, Spi91b, Vil74, Wol71, Blo77, Dia64, Fif79, Jor89b, Kau86, Kla83a, Law72, Mey89c, Mit87, Pin72b, Wei87, Whi63b, Wil85a, Mil61a]. II. [Lax67]. III [But74c, McG65b, Sal63b, Wil86a, BK69, Mey89b, Sto66]. III-2 [Sto66]. III. [Mon67]. Ikonić [LG89]. Ilgamov [Ash89]. Ilia [Jor78b]. Ill [Bar22, Bec94, Bra96, Eng94, Gol76c, GW76, HB96, Hea76, Kah76, MCS69, Neu98, PW79, Rum99, Rus76, Tre20, Var76, Var79, WE76, Wil79, Han92]. Ill-Conditioned [Bar22, Gol76c, GW76, MCS69, Neu98, PW79, Rum99, Tre20, WE76]. Ill-Posed [Bec94, Bra96, Eng94, HB96, Hea76, Kah76, Rus76, Var76, Var79, Wil79, Han92]. Illner [Bar95]. Illumination [De 00]. Illusions [Hof71]. Illustrating [Mac77]. Illustration [Ano99b, The99a]. Ilya [Dau90]. Image [And78, Ang78, BCM10, Hei95, HS96, MSS96, Ser96, Sha96, Sta90, Hun81, Sta88, Tek89]. Images [BDE09, Goo91a]. Imaginary [Sha04]. Imaging [ABJK10, BdCPT12, BL10, DMP22, PCLS23, Che18, Fri86]. Imbedding [Bai76, Bar64, Bar65, Mar88, Van78, Win74, Wou62]. Imitating [SK76b]. Immiscible [LR00]. Immunization [Elv75, EFA76]. Immunology [Hop79]. Impact [CN05, DP76, Kac62, Pet97b, Ste00, Pet97a]. Impagliazzo [Pug86]. Impedance [BL10, CIN99, SW11, SPR76]. Implement [Hig02a]. Implementation [Epp86, Fat76, Fat93, Hun81, Nik78, Ser04, TB76, Lud87]. Implementations [Van93]. Implementing [DGK84, Gay76, Shr76, Yoh79]. Implication [DA79a, DA79b]. Implications [LC65, MM98, MV15, TH96]. Implicit [CL75, Hol70b, Med62, MMT^{+}21, NJN98b, RD65, Geo67, NJN98a, Hei93a]. Importance [SKP19]. Imposed [Sch70a]. Impossibilities [Mey76b]. Impossibility [Han90, PTK11]. Impossible [Tok95, Tok94]. Improbability [Don23].

Improved [GBF78]. Improvement [Koh72]. Improvements [Poh78]. Improving [Fel78, For21, FLGM23, Yan90]. Impulsive [CSW^{+}08]. In-Host [ZWY14]. Incidence [PO76, Wal85, Wal87]. Include [AW19]. Including [Hig02c, Mor82]. Inclusion [AMR04]. Inclusions [DL92]. Incomplete [Sho66b, Ste67b]. Incompressible [Bak91a, Ber90a, ERS14, GJ19, JLM^{+}17, Sei76, Ser64, Shi71, ZD78]. Inconsistent [Mei68]. Incorporating [CAiJS23]. Incorrectly [Fro86]. Increasing [CD94, Gla77a, MR76]. Incrementing [GAK17]. Increments [PC89, PC90b]. Indefinite [BG76, BS66, DGK16, MS14, Pap62a]. Independence [Man63, PPR11, Sam80]. Independent [Bry96, Har76c, PB96, Smi80a, Smi81, ZM20]. Index [SRK99]. Index- [SRK99]. Indexes [Har76c]. Indicators [EBTZ94, ETZ94]. Indices [McC91b]. Indritz [Ans64c]. Induced [BRMR08, FM18, MC21, MBK08, Reh89]. Induction [WR96a, Woź87, Rog96]. Inductor [MA04]. Inductor-Capacitor [MA04]. Industrial [Ano72o, Ano73n, Ano75o, Ano77q, FLC96, Hig00, Hig02c, Kap90b, Kap90d, Ock92a, Ock92b, Rog60, Win62, Cum96]. Industry [CGT^{+}97a, Gas60, JRZ97, MMT^{+}21, VS95, Yan97b, Tab97, Bla81]. Industry-Relevant [MMT^{+}21]. Inelastic [BvdGH95, Hav95]. Inequalities [Aga94, BJ79, Bar77, Ber75, BSFW82, BSW82, BEV96, BGF^{+}95, Cha70a, Ell87, Eva81, Flo77, Fos81, GS67, Her88, Hor95b, Hou65, JP89, Kel89, Kot65, LM00, Neu96, Pay67, Red76b, RAH^{+}90, Ril68, Ros72c, Ros75, Ton88, Wan79, Yak95, tR83, Sei97, Ama82, Ame84, Hir71, How83, Jor79, Lev84, Ren90, Rue78, Ste70a, Str70, Van96, Bor73, Dur93a, Fan82, Jag76a]. Inequality [AB97, AT76, Ang97b, Ano91d, Ask73b, Ask73a, Ask75a, Auc97b, Bjø90, BDS92, Bre71, BNR85, Car70c, CV91, Cha69, Elc76, EK86, FL84, Fra83, Gau87, JBG^{+}90, Jag92, Kla71b, Kla77, Kla78a, Kla78b, Kla79, KG79, Kla83b, Kla84d, Kla85, Kla86b, Kla90c, LMSM83, LMSM84, Leh60, Lem70, Lux79, Lux84, Lyn67, MG79, MNC71, Mei78b, Mit91b, Mit91a, MATL77, New67b, New68a, New69, Nor89, RM86, Roe95, Rue82a, Sch77b, Sei86b, Sei91, Sha73b, SO78, Sim70, Sim71b, SV89, Ste74, Ste75, SY75, Sut95, Sza88, Sza91, Tre73, TC68, VKK79, Wal75a, Wil77a, Wil78a, Wil96, YKF96, vD81, Ang97a, Auc97a, BW96a, CBW98, EV91, Ram97, Ren98, Wil98a, Wil80c, Ano80d, AF69, Blu78, Boe89, Boy70, BW96b, Car71a, Elk86, Esc70, GC69, Hol75, Jag72b, Jag74b, Jag74a, Jag87a]. Inequality [Jag90, Kla80, Kle75, Kol76, LS84, McL79, Moo77a, RL62, Rou72, Sei86a, Ste68b, Yan77, vK79, Ame86a, Kla86a, Rue82b, TBDS92]. Inertia [Ano93c, GMSW91, KB76b, Mur93]. Inertia-Controlling [GMSW91]. Inertial [Liu91, Gur63]. Inexact [Ber85c, SS05]. Infection [VLPR20]. Infections [Tud90]. Infectious [Het00]. Inference [BR62, Car93, DMP22, Goo86, Har93, Kup75, LW13, Lim23, OSS15, PWG18, Rad90a, Sha92b, Siv94a, Siv94c, Har72a, Hey88, Noe72, Rad90b, Rei79, Rei88, Rei89, Ros72a, Sha92a]. Infestation [GZ11]. Infestations [FS78b]. Infinite [Ano93b, Atk66, BK95, BDD^{+}95, Bil93a, Bil94a, Car68b, Car69, Car72a, CZ96a, Chi71, Coo93, Coo94, Cor92a, Cur75, CZM96, Dav94a, De 79, Eks60, FG76, FGB81, Fet80, FLR^{+}81b, FLR81a, Gel76, GK86a, GKG^{+}86, Gla70, Gla77a, GJ80, Gla80, GR81, Gla81a, Gla85b, Gla93c, Gro92b, GNP95, Her87, HK93, Jor77b, JG93, Jor93a, Kla63b, Kor95b, Kor76, LG95, LmYS96, Lie77, Lie78, Mil94a, MP20, Mor96, Mor91b, NPW95, New91a, New94a, Olk72b, OH76, OK94, PD76, Pet85, PZ81, Rog83b, Rus95, Sar71, Sha65, Ski95, Sto96, TT94, Ula64, VBU97a, Wet74a, Yas71, Yau69, aYT96, CZ97, VBU97b, Gas95, Ber94, Blu94b, Gla86c, Gro96f, Mor91a, Tha96]. Infinite-Dimensional [Cur75, CZM96, MP20, Mor96, Mor91b, PZ81, Mor91a]. Infinitely [Kaw63, Jow86]. Infinitesimals [Kap78]. Infinity [Ahm75, Hey86, New87]. Inflation [Her81]. Inflationary [Hyd86]. Influence [EERG20, JMFB15]. Informal [Ham80, Mir88, Sch93a, Sch93b]. Information [AE23, Arm90a, Arm90b, BW76, BDO95, BDJ99, Bra17, Cam77, Gor63, Kem67, Kie94, Kot65, KKY95, Lai78, LM05, LL69, McG67b, McM61a, Mor73, Owh17, Ric88, Ull88, Was94a, Wyn69, McA88, Tul68, Yon95, Bic94, Bla81, Gos79, Kul69, Man61b, Mos93, Nov94, Sal64, Sal66, Sal63b, Shu87, Tho60a]. Information-Based [Was94a, Nov94]. Information. [Lev87]. Infrared [Ran76]. Ingham [Lig86]. Ingram [Fan82, Gup82, Man61a]. Ingrid [Hei93b]. Inhibitor [Ada91]. Inhomogeneous [Coo67b, Kel64, Sah95, SB98a]. Initial [Fus68a, Gea85, GM87, Gra67, Hig86, Joh83d, Joh83b, Kla62b, LP75, Na67, Na68, O’M71, O’M88a, Pon90b, Rog75, Sco91a, Sco91b, SH94, Van88, Van75, Wen68, Wil73, Wil78b, Win67, ZC74, Bai74a, Pon90a, Wil88]. Initial-Boundary [Hig86, Pon90b, Pon90a]. Initial-Value [Wen68]. Initialization [Ars97, Ghi76, Mac98a]. Inner [Kel88b, MC91]. Innovation [BR79]. Innovations [KVM78]. Input [BNT10, MP20, Qui87]. Input-to-State [MP20]. Inputs [Bon67]. Insitu [Geo76]. Inspection [Ang97b, Ang97a]. Inspired [CPSV18]. Instabilities [CS75, Kir93b, Kon03, Ber83b, Kir93a]. Instability [MC21, Mie17, RK75, Stu89, Pro84c]. Instance [Sku04]. Instant [BS70]. Instantons [Par87]. Institute [Ano73m]. Instrumental [Fab87, Wib87].

Instruments [AW02, HS15]. Insulated [Wil59]. Insurance [Nas87b, Ruo90a, Ruo90b]. Integer [Alh96, BJ76, Bla88d, BC96b, Bos99, CC76, Geo67, Kli78, Lev89, Mey76b, Moo74a, Pat03, SPBC96, TLM90, Vie15, TA90b, Bla90, Mey79, Rub77a]. Integrability [Stu90]. Integrable [CBM98, Gib95, KG95, Gui92, Tu92]. Integral [Ada97b, AW76, AG76, Amo78a, Ang94, Ano80d, Ano92a, Ano93a, Ask69, Ask73b, Ask75a, Ask84, AG14, ABG96, Bad61, Bai87, Ban88c, BD73, Bar73, BG95a, BK17, Ber84c, Ber96, BR91, Bil93a, Bil94a, Bla91, BF82, Boe68, BD92, BM98, Bou66b, Bou67a, Bou71, Boy70, Boy78, BAD^{+}79, BEG23, Bro63, Bur93, Bus63, Car67d, Car71c, Car76b, Cas72b, CR96, Cha95a, Cha70a, Cha96c, Cha81a, CQRZ96, CZ96a, Coc84, Coc86, CW70, Cor95, Cru86, Dam85a, Dam85b, Dam86a, Daw74, De 74, DB70, Dol93, Eic84, EK87, Eva93, Eva94, FLM22, FS68, Fet68c, Fet81, Fus68b, Fus69, Gar75, GR87, Gie63, Gil69, Gla67, Gla70, Gla75a, Gla75b, Gla77b, Gla78b, Gla78c, Gla84a, Gla84c, Gla85c, Gla85d, Gla87b, Gla87c, GG88, GGR^{+}88, Gla89a, Gla89c, GG89, Gla93c, GBdD94, Gla95e]. Integral [Gla96a, Gla97b, Gol76a, GF78, GH86, GK86b, Goo65b, Goo68, GSL79, Gro87, GJ90, Gro93c, Gro88b, GAB^{+}88, Gue72, Gut74, Han85a, Han76, Hen82, Hoc66, Hol75, HM73, Hun76, Ike72, Ism96b, Jag64, Jag74b, Jag88a, JGA^{+}88, JH92, Jer76, Jer82, Jin84, Jin85, Joh75, Joh83g, JF71, Jor78a, JG90, Jos76, Jos95, KL76, Ker92, Kir76, Kla90c, Kle63, KR74, Kol76, Kor73, KKS82, Kri97, LS66, Led96, Lev65, Lev63b, LG74, LG75, Lig86, LMS68, LS71, LS72, Lon77, Los76c, Los83c, Low61, Lux69, Mal78, McG93, McN67, Mei78b, Mer84b, Mer84a, Mey80b, Mey94a, Mil89a, MJ61, Mil94a, MM10, Moo65, Moo74c, Mor84, MR65a, MR72, Mur90, New91a, New67b, New68a, New69, Nor96, Nut85, OSS82, OW88, OGGR90, Ort94, Ort95, Ost91, Ove66, Pol65, Pre81]. Integral [PC95, Ral62a, Ral65a, Ral69b, Rei71b, Ren94b, Ric85b, Ric85a, Ric76b, Riz94, Rot62, Rue73b, Rue75, Rue79, RS71, San72, Sch76c, Sch77c, Sha63a, Sha71a, Sha73b, Sha74, Shi93, Sho64, Sie61, Siv72, Siv69, Slo97, Sne67, Sol67, Sri84, Sta76, Ste85, Sti87, Sto76, Str65b, Tay97, Tur68, VKK79, VD68, VBU97a, WGG^{+}92, Wal75a, Wal76b, WJ75, Was94a, Wei90a, Wes65, Wil67a, Wil84a, Wil86a, Wil77a, Wil78a, Wil80c, Wil96, Win67, WA94, Wou68a, XL94, Yas71, Zie73, vK79, Ada97a, AH98, Ber97b, BT85, Boe98a, Cha96a, CZ97, FP97b, Gla97a, Gla98, GJ97, Ism96a, Rob97, VBU97b, Gla76, AS85, Ano63a, Ano71a, Ano93b, Bla70, Boa82a, Boe76, BdD92, BG95b, Bou86, Byr65, Car70a, Chi71, Dav94a, Far65, Fet68b, Fet68a, Fet72a, Gas71, Gla85a]. Integral [Gla86b, Gla88a, Gla89b, Gla95b, Gra79, Gro90, Grz83, HR85, Jan70, JG70, Jor90b, LF66, Los73, Los81, LJ83, Mar75a, Mil92, Olk72b, OG90, OR67, Pet68, RSS65, RAG72, Rou91, Rue72d, TS65, VJ83, Wag92, Wil85c, Wol72a, Bai76, Hoc92, Koh86, Liu91, Noh76, Rob74, Rob76, Wil84b, Yan65, Yan76]. Integral- [Wil67a]. Integral-Equation [Joh75]. Integrals [Abr84, Aga93, AD76, Ano91m, Ask79a, Ask85, Ask86a, BR94, Bil93b, Bil94b, Bjö85, Boa87, Boe91, BdD94, Boh88a, BK66, Bus95, CZ92, Cor68, Eks60, Fet65, Gla82a, Gla93b, Gla93d, Gla94c, Gla96b, GRS97, Gro91c, Gro96b, Hab70, Har67b, Her90, Hil76, Hui95, Hui96, Hui97, Jon72, Köl87, Kuc87, KWP95, Lar94, Lev65, Lop94, Lóp08, MN65, MR76, Meh74d, Mil66, Pro95, Ric61, Rob95a, Roc92a, Roc92b, Sch59, Sch63c, SD74, Sha63b, Sha64b, Siv68, SM94, Str65a, Ver90, VBdD91, War69, WM61a, WM61b, Won80, Gro98a, Gro98e, Sch62a, Ano91b, Böh88c, Far69, Gol73a, Gro91a, LSB67, MR71, Pap62a, Rob95b, Sho66a, SS70, Ste73b, Gro96g, Ste71]. Integrals-Construction [Boa87]. Integrands [FH70]. Integrate [ADW23]. Integrate-and-Fire [ADW23]. Integrated [SZ96, Zha96]. Integration [BS66, Cha81b, Cha87c, Cla83a, CH66, DL66, Dra83, Flu78, FH70, Gol62, HLS09, Hul67, Kel84a, Lam04, NE76, OR21, PC61, Ral65a, Ric88, RLT65, SJG96, Str86c, Str62c, Var80, Wil78b, YKS78, Cla83b, Gen96, Kar82, Kau88, Kel84b, Ste68a, Str76b, Per65]. Integration-Second [Str86c]. Integro [Ano93d, Mel91, Mel92b]. Integro-Differential [Ano93d, Mel91, Mel92b]. Integrodifferential [Mel92a, Mel93, Dri81]. Intellectual [Seg80]. Intelligence [CL72, Gra86a, Sam76b]. Intelligent [BDO95, Rad90a, Rad90b]. Intensive [Ber95, Efr88, HB95]. Interacting [DBS^{+}07, Fre86a]. Interaction [Ame91a, Ame91b, Bal97, Bat21, CLG14, Che60, LB10, NC09, Str86b, YA18, WLP^{+}61]. Interactions [Ben87, CN15, Hen76]. Interactive [Ber85d, Bra17, Nie76, Ros76b, SZ99]. Intercept [Gan98b, NN96, Gan98a]. Interchangeability [Sam80]. Interface [BF16, NR90, Rue89, New90]. Interfaces [San60, Say13]. Interfacial [Whi93]. Interference [Bas86, Mir97]. Interim [TH96]. Interior [DdSG^{+}22, EBTZ94, ETZ94, FGW02, Fri86, Ye94]. Interior-Point [EBTZ94, ETZ94, Ye94]. Intermediate [Ame81, Baz73, MD96, RM75, Boa84, Cho66, Dum96]. Intermittency [BC98]. Intermittent [Gia63]. Internal [AG14]. International [LW64, Mil61a]. Interplants [CBG75]. Interplay [SN04]. Interpolating [Zel77]. Interpolation [Alf93, BP76b, Bat21, BT04, Ber07e, Boy13, CS20, Dav75b, DB78, Gau87, Hue92a, KL75, Law78, Mat70, Nei19, PS74, Riv59, Riv60b, Sha72b, Str62a, Wix76, dB74, Dav75a, Gre65]. Interpretation [Per93a, Per93b, Was74]. Interpreted [OW95b]. Interpretive [Lan76a]. Intersection [Sha88, Xu96c, Xu97]. Intersections [Nie96, Dur92a]. Interval [CR78, FK62, Han81b, Hea76, Jac69, MW76a, McM87, Moo85, Rus76, Nit85, Han67, McM87]. Intervals [Ber93, Her87, Udw89, Ang90, Gho87]. Intractability [Har82]. Intriligator [Nel74]. Intrinsic [Bél88]. Introducing [Ste98]. Introduction [Als92a, Als92b, Ame95, Ano99e, Ano99f, Ano00d, Ano05d, Ano07a, Ano08a, Ano08b, Ano09, Ans64a, Ant83b, Ant83a, AM74a, Are89, Art85a, Ast87, Bak66, BKS66, BKS68, BK69, Ber91a, Ber05a, Ber08c, Ber08d, Ber09d, Bjö85, Boa81, Boy90a, Boy90b, Boy97, Bre68, Bre93, BDC98, BL13, Bry91b, But79, Cam91a, Cha94c, Cia92b, Cof83, Coh84a, Cou97, Cre97, CZM96, Daa97, DTG98, Dat67, Dim91a, Dim91b, Eas91, Edw90, Ell84, EK13, Eva81, Fis76, Fla99a, Fla99e, Fla00d, Fla02a, Fla03a, Fla03b, Fla03f, Fla04a, Fla05c, Fle63b, Gar67, Gat88, Gla94a, God87, Gol78, GO09, Gri61, GM96, Gut90, HRB96, Hau90, HK95, Hig01a, HH19, Hin87, Hir73, Hoc80, Hol70a, Hol87, HO96, Hol90b, Hom86, Hor83a, How63, Ips08c, Ips08d, Ips09d, Jon92b, KHS96, Kaw63]. Introduction [Kel60a, Kel99d, Kel17, Kes69, Key94, KG97, Klo93, Koç89, KG95, Kor95a, Kul60, Lae82, LK98, LeV03b, LeV03c, LeV03d, LeV04a, LeV05c, LB10, Lev88b, Lev63b, LA95, Lou92, Mal85, Man90c, Man90d, Mey83, Mey91a, Mey96, Mir88, Mor96, Mor88d, MG95, Nan20, Nau86, Nei10, NWF97, O’M07c, Ode89, Olv97, Ost61b, PMG95, Pin72a, Pin87b, Pul91, Rei62a, RRK95, Row64, Row86, Sag83, Sai91b, Sal85, Sal63a, San08c, San08d, San09d, Sch93b, Sch99e, Sch99a, Sch00a, Sch00b, Sch02a, Sch03a, Sch03b, Sch03c, Sel94, Sie69, Siv94d, Siv94b, Siv94c, Smi82, Som93, Spi91a, Spi91b, Stu90, Syn93b, Tre99a, Tre99b, Tre00c, Tre00d, Tre02a, Vak93, Var85, Vog91a, Vog91b, Wat63, Wen78, WJV79, WA63, Zem68, Zie78, Str89a, Tre13, Cia92a, Hel97]. Introduction [Her97, Hig02b, Hor97a, Stu90, Aro69, Bei96, Bry91a, Coh82, Col87a, Enr87, Fra88, Hol90a, Jon92a, May91, Pra89b, Ros80a, Sai91a, Sch93a, Sch63a, Sco90a, Sco90b, Siv94a, Sle74, Sle79a, Wil77b, Ahl93, Ant82, Arb78, Ars62a, Bel90a, Bir64, Bre64a, Car66c, Cha64b, Cha68, Chi90, Coc86, Coh84b, Cro73, Dra62, Dri75, Duf91a, Eas93, Eps59a, Fis71b, Fle62, Fro60, Fro87b, Gib95, GA63, Gol66b, Gre87, Gue85, Har76b, Hen86, Hig62c, How82, Jag64, Kap71, Kap78, Kin82, Koe95, Kus72, Lee82b, Lem85, McN89a, Moo85, Mor82, Nas90, Noe81, O’M96, Orb18, Pit77, Por89, Pro65b, Rei89, Rob85, Sop82, Spo88, Sto62a, Swa96, Sym92, Syn83, Syn93a, Thi63b, Wei90b, Whi69, Wil60a, Wis60, Wou68b]. Introductory [Bis78, CL23, Gei89, Kri85, Rei78c, Als85, Dwa85, Gue88, Sop79]. Intuitive [Har60, Kla92b]. Invariance [Blu86, Joh84b, Joh84a, Lew78a, Mor75a]. Invariant [AMSV02, AG01, BW11, Dem74, Mar88, Pal95, Ral69b, Van78, WP95, Wou62, Bar64, Bar65, Joh88, Win74]. Invariants [CPW12, Chi90, Ars62b]. Invented [Fin72]. Inventories [Bar59]. Inventory [Guh66, Had65, Mor69]. Inverse [BC76, BK87, BL10, BB83b, CZ92, Che68, Che73, Cho76b, Chu98, CG24, Cli64, Cli78, CB76, Col84, Col94d, CCM00, CCM01, CK18, CEH12, Dec65a, Dei92, Den68, Den69, DMP22, Fin71, Fis71a, FN73, Fre69, Gar69, Gla81c, Gla93b, Gol62, Gon06, GMP66, GDR78, Gre66, Gre67c, GRYL89, Hag89b, HS81, Hue92b, Hue96, Ips97, Jag82, JGK^{+}82, KG97, Kla88c, KG89, Lat23, Mey75c, Nag96, New70b, Nie98, PW79, Ric76a, Rob74, Rob76, Ros69, RY88, Sco83, SN96, SS78, SH96, TDS18, Ten01, UFW17, Urq68, Urq69, Zie73, Zlo70, Bar86, Col94c, Hil96, Roy91, Sle94]. Inverses [Ask75b, Ber81, Cho69, Gil70, KVM78, MO68, PR70, Roh66, Ste77, Tew66, Tew67, Tew69, Wan79, Mey76c, Zlo91]. Inversion [ABL^{+}13, AB23, Bus63, DK89, Dec65b, Kei89a, Kei89b, MBOV17, Nie86, Sta76, Swi60, TS66]. Inversions [Ham77, Reu67]. Inverted [LW95, Lev88a]. Invertible [Chu94a, Chu94b]. Inverting [NS02, PC90b, PC90a, Wit22]. Investigating [ERS14]. Investment [Car71a, Lem70, Ngu76]. Inviscid [ZD78]. Invitation [BvJ09, Lam04, Ban85]. Involutions [Isr92b, Isr92a, SS91]. Involving [All78, BSFW82, BSW82, CZ94, CZ96a, Fos81, Han85c, Hil76, KB59, Los86a, Los86b, Mil95b, Mil95a, SD74, Sne67, Ung77, CZ97]. Ioan [Pil87]. Ion [BM03, Tre78]. Ionescu [Con96, Pav94]. Ionization [Rei83b]. Ionosphere [Fej88]. Ions [Jer91b, Jer91a]. Iooss [Van94b]. Ioss [Van94c]. Irena [Har72b]. Iris [DH09]. Irmgard [Bus70]. Irreducible [Sch69b]. Irregular [Gut88, Hir78, KK99]. Irregularly [Ber64, Beu66, War98]. Irreversibility [Spo85]. Irrotational [KN61a, Kla82, Kla83a, VK62]. Irvin [Mar66]. Irving [Kri91]. Isaac [Gea80, Pan93a, Pan93b]. Isaacson [Kar78, Moo69]. Isaak [Jer91a, Jer91b]. ISBN [BC96b, Bor06, Hig05, Tre13]. Iserles [Hig92]. Isidore [Gre63b]. Islam [Gla94a, Xan87]. Ismael [Sag86]. Isometry [BCT11]. Isoparametric [AD76]. Isoperimetric [Elk86, EK86, Her88, Kla83b, Kla84d, Pay67, Her88]. Isospectral [Dri97, Wat84]. Isothermal [Lin86]. Isotope [Gel79]. Isotropy [Tre17]. Israel [Bor90, Eck82, Hag85b, Mey76c, Por86a, Bog75]. Issues [Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Bla88f, Sin88, Wol86]. Istvan [Mod92]. Itô [Gre87]. Itô’s [Mal83]. Iterated [Ano99b, Ber92a, CN05, DF99, The99a]. Iterates [Nit85]. Iteration [AMSV02, Bog75, Bre67, Dev94b, Ful84, GKP75, GH13, Ips97, Jos70, LW17, Liu83, Nik78, PW79, SV00, Sma76, TDS18, Ano68a, Bla93]. Iterations [DTT98, Han94a, Han94b, Kel60b, Kui06, NF16]. Iterative [AH95b, CE89, Chu88a, Dax90, Fel78, GZ19, GDR78, Kir10, KD76, Max60, Par86, PH76, PW15, Por83, WP78, Woź87, Xu92, Dau90, Gre76, Han95, Joh73, Ral74, Var67, Wil66]. Iteratively [ST91b, ST91a]. Iuanović [Gla86b]. IV [Her77, Wil86a, CLH97, Yan65]. IV. [Car67e]. Ivan [Fla66, Jar76, Rou88]. Ivar [Ans68, Cla78c, Gue69, Kap81, Rat92a, Rat92b]. Iverson [Car65b]. Ivor [Thi84b]. Iyengar [Tho87]. Izaak [Ham77].

J [Abb92, Abr94, Aga94, AK90, AKP91a, Ame85, Ame95, Amo78a, Ano71a, Ano80d, Ans64b, Ant86b, Arm94, Art83b, Art85b, Art88, AF69, Ask72, Ask81, Bai87, Bal88a, Bal88b, Bal95, Ban88b, Bas88, Bau91, Bax85, Beb79, Bei85, Bel90b, BI68a, Ber93, Ber95, Ber92c, Ber83b, Ber88a, Beu83b, Beu83a, Beu85, Beu87, Bir69a, Bla70, Bla84, Bla86a, Blo89, Blo91, Blu78, Boa82b, BG95b, BGL^{+}67, BC96b, Bou66b, Boy70, Bra94b, Bra94a, Bre89, BS80, BBLS92, BOS73c, Bun83b, Bun83a, BG64, But92, Cab75, Cam77, Car71a, Car71b, Car81, Cas87b, Cer96, Ces75, CL72, Cha82b, Che89, Chu82, Chu92, Coc86, Col87a, Col91, CZ89, Col92, Col89, Cor95, Cor92b, Cra84, Cre97, Daf89, DV76, Dam84b, Dam84a, Dat86, Dat90, Dav91b, Dav91a, Dav94a, Day89, Den78, Die92, Dis92]. J [DG68, Dra62, Dre89, Dri81, Duf86, Duf91a, Dun94, Dur87, Enr84, Enr89, Eps59a, Erd77, Fab87, Fad62, Fal70a, Fet66, Fet80, FLR^{+}81b, Fie93, Fin76a, Fir97, Fle79, Fle84b, For96, FR63, Fre80, Fre86c, Fro83, Gas63, Gea74, Gen86, Gen76, Gha94, Gla94a, Gla88a, GC69, GGG69, Goo72, Goo91b, Goo68, Gou68, Gra67, GRL87, Gra86b, Gra66b, GB89, Gre76, Gre65, Gro87, Gro93d, Gue79, Gue83d, Gup96, Gut90, Hag82, Hai85, HPF85, Ham88, Ham90, Ham73a, Han86b, Har74, Har66a, Har78, Hau85, Hau88, Hav95, Hea74, Hei93a, Hei97, Her77, Her83a, Hes86, Hey93, Hey97, Hig62a, Hig00, Hig06, Hof85, Hof95, Hol70a, Hol87, Hop83, Hor83b, Hor83a, Hou70b, Hou71a, Hua96, Hud95, Hun84, Hur65, Hut89, Jag71, Jag72a, Jag83, Jag87a, Jag92, JW83]. J [Joh83c, Joh83d, Joh83a, Joh83b, JF71, Jor86b, Jun65, Kan62a, Kap85b, Kap67, Kap78, Kaz84a, Kaz91a, Kei71, Kel86c, Kem67, Kha71, Kin76, Kin82, Kir90a, Kir90b, Kla84g, Kla90a, Kla92b, Kni86, Koh88, Kol63, Kol76, KF62, Kow86, Kra96a, Kum88, Kus72, Lan86, LF66, Lau74, Lem84, Lev63a, Lev63b, Lin91a, Lin91b, Lin92, Lon87, Los73, Los74a, LFR80, LJ83, Los83b, Los84a, Lu92, Luc18, Lud87, Mac68b, Mal87, Mal90, Mal93a, Man61d, Man90c, Mar85a, Mar75a, Mat80, McA88, McC92a, McC91b, McC79, McC86, McG64, McG65a, McI71, McK81, ML85, McN89b, Meg88, Meh73, Mei83a, Mey82b, Mey84, Mey95b, Mik89, Mit87, Moo74a, Mor93, Mor96, Mor68, Mor83, Mul69, Mur97, Nas77b, NPW95, New82a, Nic96, Noh62, O’M87b, Old62, Olm77, OG90, Pac86]. J [Pal94, Pan95, Pao88, Pap63a, Par94, Par90, Par66, Par74, Par81a, Par72, Par92a, Pfl92, Pic92, Pin77, Pip81, Pip86, Pit91, Pow75, Pro91, Pul64, Ral74, Rau95, Red90, Rei63, Rei83a, Rei88, Ren94d, Ric67a, Rob94a, Rob62b, Rog96, RSS65, Row64, Rue73b, Saa73, Sag77, San82, Sat78, Sau62, Saw82, SNS66, Sch65b, Sch78c, SR93b, Sch81c, Sch85c, Sco83, Sco84b, Sco84a, Sco94, Seg80, Seg85, Sei86a, Ser80b, Ser96, Set62a, Set62b, SM79, She63a, Shi68, Sho66a, Shu87, Sik92, Sim76, Sin93a, Siv94d, Siv94b, Ske95, Sle81, Sle87, Sle89b, Sma76, Smi74, Smi82, Som93, Spi80, Ste67b, Ste68a, Ste93a, Ste68b, Ste97b, Str89b, Str76b, Str86c, Sub90, Szy87, Tad93, Tal81a, Tay68, TT76, Tay86, Tay97, Tha82, Ton95, Tre84b, Tre84a]. J [TS65, Uhl73a, Uhl73b, Van76, Vel71, VN73, Wal79, Wal75b, Wan91, Wan83, War92, Wat86b, Wen78, Wet93, Whe74, Wig90, Wil66, Wil65a, Wil69, Wil60b, Wil78b, Wil81, Wim75, Win74, Wol97, Wol72a, Wol94a, Yan77, Yan92, Zas65, Zlo83, Zlo91, Zwe84b, vH87, Man90d]. J. [Art83a, Fle84a, Gue83a, Kla84f, Tay97, Zwe84a]. J.-L [Sle87]. J.-P [Yan92]. J.-S [Mor93]. Jack [Ans64c, Blo76, Blo77, Cho94, Dat78, Fle76, Hig00, Mey72, Rei89]. Jackknifing [FL81]. Jackson [Flu93a, Lee77, Pap63a, Whi84b, Whi84a, Flu93b]. Jacob [Shi69, Wet75]. Jacobi [Eva85, Fle85, Rei64, Bag95, Gau87, SV00]. Jacobian [Jor80c, GMP05, Jor80b, Per79]. Jacobson [Sag77]. Jacoby [Gre76]. Jacod [Pro91]. Jacqueline [Ede90a, Ede90b]. Jacques [DD74, Eic67, Joh91a, Ker92, Mey95b, Tul90, Joh91b]. Jacques-Louis [Ker92, Tul90]. Jaedicke [Ber62]. Jagdish [Olk78, Ros78]. Jagers [Ano89b, Olk77]. Jagjit [Dra61a]. Jai [Sle89b]. Jain [Lam86, Raz92b, Tho87, Raz92a]. Jaluria [Reh89]. Jamal [Xan87]. Jambu [Gra93a, Gra93b]. James [Ant72, Are91a, Are91b, Ask92a, Asn75, Blu94a, Fro87b, Goo82, Gue83b, Her67b, His77, Hor97a, Hul71, Jon92c, Kin94c, Mar75a, Mir88, Nas77b, O’M87a, Pei83, PRM^{+}66, Pro85, Rei87, Ric74, Sav18, Shu80, Ste70a, Str70, Wie87, Wig90, You72b, Ask92b]. Jan [Bin84, Ker91a, Ker91b, Mey88, Ruo91b, Ruo91a]. Jan-Cees [Mey88]. Janacek [Fin95]. Jane [Cho87, Lae82, vV65]. Janos [Dwa80]. Janssen [Jow86, Jur95, SR93b]. Jansson [Jun70a]. January [Ano71o, Ano72n]. Jaromir [Neu97b]. Jaroslav [Gat88]. Jati [Fro74]. Javed [Jód96]. Jayme [Hau86]. Jean [Dan72, Fig93a, Fig93b, Fix74, Gup82, Kir93a, Kir93b, Laz91a, Nas90, Noe72, Ros71, Sch82a, Sha96, Laz91b]. Jean-Marie [Nas90]. Jean-Michel [Sha96]. Jean-Pierre [Fix74, Kir93a, Kir93b, Sch82a]. Jeff [Mic92b]. Jefferson [BY78]. Jeffery [Jam59]. Jeffrey [Bla66, Cha78, Fok84, Gra70, GdM95, Hai85, Har67d, Hei97, Lai78]. Jeffries [Ken91a, Rob94b, Jef94, Ken91b, Mic92a, Rob94b]. Jennifer [Mil81b]. Jens [Pon90a, Pon90b, Rei96]. Jensen [Rei96]. Jerold [Gla87a]. Jerome [Bai74b, Far94, Mar87, Mul64a, Sho87]. Jerri [Coc86, FLR^{+}81b]. Jerrold [Dol91a, Lee84b, Rob81, Dol91b, Her97]. Jerry [Bry91a, Bry91b]. Jesper [Kaz91b]. Jessop [MT61]. Jet [AAJ76, Luk82]. Jevtić [Erd92b]. Jian [Coh90a, Coh90b]. Jiang [Els91]. Jill [Sch93d]. Jin [Gla89b]. Jingyuan [Coh90a, Coh90b]. Jiongmin [Sto96]. Jo [Sha83]. Joachim [Ano78i, Ham85]. Jobson [Lin92]. Joe [Gen96, Len93b, Len93c]. Joel [Ano68a, Jag72b, Kat65, Moo75, MR65b]. Johan [Gru74, O’M88b]. Johann [Ama82, Ame84, Eat96]. John [And95, Ano93d, Ant91, Boa87, BdD94, Bro93, Cam91a, CS70, Cha67a, Cha83b, Cha81b, Daf89, Eas89, Ell88, Fly89b, Gin85, GD88, Gol65b, Gra70, Har83, Hea91, Her83b, Her71, Jow72, Jun64, Kap71, KJ87, Kaz83, Kaz84b, Kaz92, Kel90c, Ker92, Kod96, Kra68, Kul60, Lak80, Law72, Lee77, Man88b, McG93, Mey72, Mil82c, Orb18, Pit79, Pug86, Pul91, Ral93a, Saa68, Sam76a, Sch86a, Sha97a, Sus88, Whi93, Wie87, WJV79, Wil72, Zie83b, vL86a, Grc11, Kel90d, Lig76, Ral93b, Tod74]. Johnson [Ant91, Bau88, Dan66, Har82, Kaz92, Kee87, Ple88, Sco90a, Sco90b, Ste93b, Stu91, Say13]. Joining [Kos79a, Kos79b]. Joint [ML80]. Jon [Bic94, Fri86, Kee90]. Jonathan [Sco92, Wim88]. Jones [Gut90, Ran89, Sta66, Wei88]. Jonge [Dra61b]. Jongen [Wet93]. Joppich [Xu96b]. Jordan [Han79, BCV98, GW76]. Jörg [Hig05]. Jorge [Fre90, Hug93a, Par94, Hug93b]. Jörgens [Coc84]. Jorgensen [Cha94d]. Jose [Ber97a]. Josef [Bos95, Hal90a, Kle72, Hal90b]. Joseph [Caf92, Car67b, Dan91, Duf84, Fif79, Fra89, Jor78a, Kap71, Kel86b, Kel90b, Mei89, Nov94, Owe85, Pip92, Ral66a, Sal64, Sho87, Sno76, Sül97, Kaz91b]. Joshua [Luc18]. Journal [ADPR04, Ano80h, Ano82h, FK06]. Journals [Ano77f, Ano77n, Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d]. Journey [Bro93]. Joy [Kie94]. Jozsef [Ari91b, Ari91a]. Jr [Aal71, And75a, Ano80d, Bau88, Ber81, Ber78, Bri64, Car65b, DeM65, Ell90, Eps59a, Fet66, Fis71a, Gur63, Han86b, Har76b, Hop79, Kv93, Kan68, Lon87, ML85, Mil83a, Moo77a, O’M87a, Rie72, Sho69, Sin69, Tha82, TS65, Tur60, Vel71, VN73, Was94b, WJV79]. Jr. [Wig90]. Judea [Rad90b, Rad90a]. Julia [BEKS17, Bur94a]. Julian [Bra91a, Lat69, Mea88b, Pit81, Bra91b]. Julio [Cha94c]. Julius [Kal65, Row64]. July [Ano70j, Ano71p]. JuMP [DHL17]. Jumping [LS78]. Jumps [Wet93]. Junctions [DC76]. June [Ano71o, Ano72n]. Jupiter [HH16]. Jurek [Hud95]. Jurg [O’B95]. Jürgen [Moo85, Ruk93b, Ruk93a]. Jurgens [Dev92].

k-Way [KK99]. K. [Ban84a, Won83a]. K.-H [Noe82]. K.-U [Ber91a]. Ka-Sing [Jd93]. Kaashoek [Lan81]. Kac [Man63, GO09]. Kachanov [Dod73]. Kaczkowski [Tad93]. Kagan [Luk75]. Kagiwada [Bai76, Bar65, Gri88]. Kahaner [Gup91a, Gup91b]. Kahn [Lee82b]. Kai [BR69]. Kailath [Zak84b, Zak84a]. Kainen [Bei80]. Kaiser [Fra95b, Fri59]. Kakeya [Pla15]. Kakutani [She66a, Wei67a]. Kalaba [Bai76, Bar64, Bar65, Gri88, Ral66b]. Kalbfleisch [Rei88]. Kall [Bri96]. Kalman [Chu90, Kap70b, Par92b, Bel10, CC90, HRW12, RBC^{+}07, Chu85]. Kamel [Ral62a]. Kandel [Man81, WP88]. Kantorovich [Riv60a, Sop69, Hou65]. Kanwal [Gue72, Wim96]. Kapadia [Ros78]. Kaper [Uhl73a, Uhl73b, Zwe84b, Zwe84a]. Kapila [Buc85]. Kapitaniak [Mac91a, Mac91c]. Kaplan [HA71, Sch60c]. Kappa [Kei89a]. Kappa-Plane [Kei89a]. Kapteyn [BY98]. Karel [Ral80]. Karen [Par87]. Karl [Ant82, Bla88c, Cam91a, Gri88, Hal90a, Hal90b, Kus72, Syn83, Wil93]. Karl-Heinz [Wil93]. Karlin [Art85a, Kar83, Kaw67, Man61d, Ric67a, Sch78c, Wei77]. Karlsson [Ism78]. Kármán [Tay73]. Kármán. [Ber83a]. Karni [Sle67]. Karplus [Gos83a]. Karr [Hey88, Sha92a, Sha92b]. Karush [TT94]. Karzanov [GG91, Poh78]. Kashya [Ism83b]. Kashyap [Ism83a]. Kato [Wil70a]. Katok [Swa96]. Katta [Bla85a]. Katz [FARE69, Zah80]. Kaufman [Jor78b, Sau79]. Kaufmann [Arb78, Mey79]. Kawahara [Fok84]. Kawata [Spi74]. Kay [Mar66]. Kazarinoff [Her92b, Mey82a, Her92a]. Kazimierz [Kel86a]. Kazuo [Chu89]. KdV [Lax76]. Keating [Far94]. Kecs [Koh85]. Kedem [Yak94]. Keiding [McK95]. Keilson [Pit81]. Keinosuke [Pao93, YH93]. Keller [Her70, Moo69]. Kelley [Rhe88]. Kellner [JRR71]. Kelly [JW83, Spi81b]. Kelmanson [Lig86]. Kelsey [Ren90]. Kelvin [DK89]. Kemp [Dob87]. Kemperman [Rob62b]. Keng [Hab83b]. Kennan [Boa85]. Kennedy [Bei85]. Kenneth [Als85, Bas88, Bla88b, Car65b, Gur66, Kel60a, Lon87, Luc86, Man61d, Won83b]. Kennett [She87]. Kent [Sal63b]. Ker [Bla88b]. Ker-I [Bla88b]. Kernel [Reu67, TF65, Tut94]. Kernels [Bus64, KL76, Mor64, WJ75]. Kersch [Cer96]. Keshet [Seg88]. Kesten [Dur84b, Dur87, Dur84a]. Kestin [Owe85]. Kestutis [Pra89b]. Kevin [Jac97]. Kevorkian [Col91, Mey82b]. Key [Bla88f, KM04]. Keyfitz [Gin85]. Keys [VdLBF17]. Khalil [Man88b]. Khinchin [How63, Rei62a]. Khokhrayakov [Ock91]. Khovanskii [Mer64]. Kibel [Kan66]. Kidney [HSSZ98]. Kiefer [Rei89]. Kiessler [Hey89]. Kilian [Len93b, Len93c]. Killingbeck [Ral93a, Ral93b]. Kim [Fra86c, NPW95, Pac88b, BPV97b]. Kimber [Sin93a]. Kimbro [Str62a]. Kind [Cha67b, Han76, HM73, Ike72, Lan76b, Man68, McG93, Ric76b, SD74, Hil76, Nas86]. Kinderlehrer [Eva81, Sle87]. Kinds [Pon67]. Kinematic [Duf91b]. Kinematics [Duf91a, Fre90, Wei80b, Duf85]. Kinetic [Erm89, GI97, Gou16, MS75, MBK97]. Kinetics [For76, WHMN16]. King [Wig90]. Kinnmark [Kel88a]. Kipiniak [Kal63]. Kirchhoff [LS96]. Kirmser [Jor73]. Kirpicnikova [Ble81]. Kites [Ado89]. Kiwiel [Lem88]. Kiyosi [Gre87]. Klamkin [And77, Atk66, BS86, DG96, Deu81c, Elk86, FARE69, Gro96e, GWK95, HBO66, Jor86a, JGM95, KH85, Knu77a, KF62, Los74b, LSB67, Mat89a, Mey90, Mey91b, Ole68, Qui92b, Rab94, RJ96, Rou72, Rue72a, Rue81a, Sch60b, Sha65, Zuo96, Dav94b]. Klatte [Cra84]. Klaus [Bre94a, Kor85, Nag88, Bre94b, Gre97]. Kleijnen [Blo76, Blo77]. Klein [Far65, Gun95]. Kleinrock [Tho76a]. Kline [Mar66]. Klir [McA88]. Kloeden [GR95, Pro96]. Klotzler [Hes72]. Knapsack [Bla88f, MR00, Van93]. Knockout [Knu86, KL87, Knu87]. Knots [Sch92]. Know [Bar78]. Knowles [Sag83]. Known [Cho69, Fel90, She76, Udw89]. Knupp [Mas95]. Knuth [Asp84b, Asp84a, Mac69, Mac70]. Ko [Bla88b, Sha82a]. Ko-Wei [Sha82a]. Koblitz [Lag89]. Kocak [Cho94]. Koch [Aiz94a, Aiz94b, Pim93]. Kochen [Sal66]. Kochin [Kan66]. Kochina [Pap63b]. Kocic [Aga95a]. Koene [Ham88]. Kogan [Ham77]. Kohn [Sle87]. Kohonen [Hun85]. Koiter [Cla62]. Kokotovic [Bau88, Man88b]. Kolchin [Par80]. Kolen [Meg88]. Kolman [Gru74]. Kolmanovskii [Rox92]. Kolmogorov [Ham65b, Ral62a]. Kolyshin [Mey94a]. Komkov [Edw90, Koh88, Hau90]. Komm [Ral62a]. Komornik [Lag97]. Kong [Ari96]. Konheim [Bla83]. König [Hey85]. Konopelchenko [Gib95]. Koopman [Sto81, BBKK22]. Koopmans [Par76]. Kopachevskii [Dij90]. Koplowitz [Abb92]. Koren [Sul93a, Sul93b]. Korevaar [Shi69]. Korfhage [McG67b]. Korganoff [BI68b, Ost61a]. Korman [Che96b]. Korn [Hor95b]. Körner [Boa88]. Korovin [Vai96]. Korsak [Sma76]. Korst [Ker91a, Ker91b]. Korteweg [Miu77b, Sco91a, Sco91b, JK72, Miu75, Miu76b]. Koshlyakov [Kor65]. Kostrov [Run91]. Kosut [Bau88]. Kowalenko [Tak77]. Kowalewski [Fus68a]. Kowalik [Gre76]. Kowalski [Boj97]. Kozin [Tay86]. Kracht [Gil90a, Gil90b]. Krall [Sle89a]. Kramer [Ham71]. Kranakis [Bla88e]. Krarup [Wet75]. Krasnoesl’skii [Rya63]. Krasnosel’skii [Ber85a, Maw96, Bue66]. Krasovskii [Wou64a, Yon95]. Krein [KK78, KM14]. Kreiss [Pon90a, Pon90b, Sül97]. Krelle [Fal68]. Kress [Sle94, Wil84b, Wil84a]. Kreyszig [Gil90a, Sop79, Gil90b]. Krickeberg [Kaw66]. Kriging [Cre97]. Krishnan [Bjö85]. Kristiansen [Giv78a]. Krohn [Bro90]. Krolak [Kro66]. Kronecker [Boa70, RM89, Tis83a, Tis83b]. Krook [Cho70]. Krushkal [Ear81]. Kruskal [Par92a]. Kryachko [Sco88]. Krylov [Fle83, Hol89a, Par10, Riv60a, BES05, CPSV18, Kui06, LS08, PSSW06, SS05].

Krzysztof [Boj97, Lem88, Mit93, Orl93b]. Kubícek [Rhe85]. Kuchuk [Böh88c]. Kufner [Han81a]. Kuga [Gra94a, Gra94b]. Kuhn [TT94, JP89, MW76a]. Kuijper [Apl95]. Kuipers [LeV77a]. Kulikovskii [Ren97b]. Kulisch [Don85, Ral83b, Ral83a]. Kullback [Tho60a]. Kulldorf [Har63]. Kumar [BS80]. Kumei [Olv90]. Kummer [Cra84]. Kumpati [Asn75]. Kunisch [Lud91]. Kunita [Çin93a, Cin93b]. Kuntzmann [Gra67]. Kunz [Fri59]. Kunzi [Din72, Fal68]. Kupradze [Kno81]. Kurdyumov [Lev96]. Kurt [Lin79c, Sai91a, Sai91b]. Kurtz [Par88]. Kurzweil [Gat88]. Kushner [Dun94, Fab80, Fle79, Gra86b, Mor68]. Kutta [Cam91b, Enr89, Sku04, BH76b, Bet78, Cam91c, Cas75, Hor78, LT98, LK65, Lut66, Ros67b, SS16]. Kutz [Luc18]. Kwapien [Pro95]. Kyung [Koh88].

L [Abb92, Abl80, Aga95a, Ame88, Amo64, AG67, And73, And74a, Ano63a, Ano68a, Are93, Ari96, Arn92, Ask72, Ask75b, Bak66, Bal95, Bar83b, Bau88, Bau89, Bax85, Bel63b, BI68a, BI70a, Ber62, Ber81, Beu85, Bha93, Bir69a, Bjö61, Bla90, Bla81, Boe76, BG95b, Boe95b, BC96a, Boe96b, Boj65, Bou65b, Bou65a, Bra63, Bre77, Bri64, Bru73, BG64, Car66a, Car76b, Car66c, Che89, Chi71, Chi95, Cla83a, Cod78, Cus96, Dan66, Dat90, Des68, DRWK64, Dev90, Dij90, Dod73, Don85, Dri63, Dri75, Dur87, Ear81, Eas89, Fab80, Fal70a, Far94, Fei96, Fet72a, FLR^{+}81b, Fin84a, Fis71a, Fix78, Fle63b, Fle85, Fro60, FZ63, Gen76, GK86a, Gib97, Gil78, Gil82, Giv78b, GC79, Gla95a, GR81, Gla85a, GH86, Gos84, Gra86a, Gra79, GB89, Gre76, Gro87]. L [Gro90, Gro96f, Gru88, Grz83, Gue95, Gun95, Hab83a, Hag83, Ham90, Ham65a, Har70, Har76a, Hey82, Hey89, Hig65, HS72, Hoc94, Hol87, Hol66, Ism96b, Jag82, JGF87, Jag88a, Jam59, JW83, Jef95, Jes77b, Joh88, JG70, Jor77a, wBJ77, Jor82, Jor84a, Jor90a, Jor90b, Jor87a, Jor89c, Kap85a, Kap70b, Kap76, Kar78, Kee87, Kee88b, Kei65, Kel84a, Kel83, Kla84h, Kla84a, Knu80, Kob10, Koç94, Kog66, Kol63, KF62, Kup94, Kup62, Lah94, Lak80, Lan86, LG80, Lar94, Lau74, Law65, LeV77a, Lee86a, Lew91a, Lit63, Los74a, Lou92, Luc18, Luk68, May84c, May84d, May84b, McC91a, McC85, McC71, McG62b, McG67a, McL79, Mea88b, Mil83a, Mil80, Mil82b, Mil94a, Min66, Mit65, MS65, Mor88d, Mul66b, MR65b, Noh62, O’M92, Olv72, OW89b, Pal84b, Pal94]. L [Par76, Pas95, Pet68, Pip86, Pit89, Poh63, Pra68, Pro85, Ral62a, Ral66a, Ral83b, Ral83a, Reh89, Rei71a, Rei62b, Ren89, Rhe70, Riv60a, RAG72, Rox76, Rue73b, Saa73, Sac83b, Sag81, Sal62, Sam86, SNS66, Sco94, Sei86a, Sez90, Sha85, Sha82b, She66b, She87, Shu85b, Sik92, Sin93a, Sle87, Sop60, Sop69, Ste93a, Stu90, Sub90, Sus88, TT76, Tay86, Ter92, Tha79, Thi63a, Tho59, Tre65, TS65, Van66, Van63, Var67, Wat65, Wei67a, Wet74a, Whi69, Wil61, WM85, Wim86, Wou68a, Wou68b, Yak95, Yan77, Yau69, You72b, Zar65, ZN62, Zim90, Zit63, vWHE91, vdV94, OW95b]. L. [Bar83a, BSW82, Cor16, Pal84a, Sac83a]. Laan [Köl87]. Laarhoven [Ham90]. Labels [BFF^{+}99]. Laboratory [ERS14, Koç94, Sal63c]. Labordère [Mey79]. Lacey [Chu92, Lu92]. Lack [KKY95, Urq69, Yon95]. Lacki [GRL87]. Lacroix [Fig93a, Fig93b]. Ladas [Aga95a, Sei93]. Ladde [How83]. Ladder [Bou66a, Fry64]. Ladyzhenskaya [Gol93, Pro87a, Ser64, Shi71]. Lag [Her67a]. Lagarias [Los83b]. Lagged [Mor69]. Lagnese [Nic96]. Lagów [Bei85]. Lagrange [Zlo85, BT04, Del97b, Del97a, Fer97b, Fer97a, Kno98, Pon83, Roc93, Ves95]. Lagrangean [HM78]. Lagrangian [Bal20, CG76, FL18, LS96, MA04, Nor89, Zha13]. Lags [Kaz80b, Beb68]. Laguerre [Bus64, GV97, McC60, Wil98b]. Laha [Kes65]. Lai [BR69]. Lajos [Dwa70]. Lakin [Kap72]. Lakshmikantham [But89, Cha70a, How83, Maw91]. Lakshmivarahan [Gal95]. Lamb [Kan62c, Mil83a]. Lambert [McC86]. Lamberto [Rox84]. Laminar [Kap85b, Ril75]. Lamp [Sil65, Law66]. Lamperti [Pit79, Van68]. Lancaster [Bar83b, Bar83a, Ell92a, Joh87, Joh88, Meh96, Pra89b, Rei71c, Ell92b]. Lanchester [CT76b]. Lanczos [CG61, CD75, GO89, Cho87]. Land [TFL23]. Landahl [Moi89]. Landau [DGP92, GP02]. Landkof [Kaz75]. Landsberg [Kum82]. Landscapes [RS02]. Langer [Dol62, Moo60]. Langevin [Mei89, DMP22, SK20]. Langhaar [GC79]. Langlois [Axf65]. Language [Bur78a, DHL17]. Languages [Aro78b, Gra70, Per93a, Per93b, Ros71, Zie83c, Zie83a, Ber85d, Boo79, Har75, McI71, McN89a, McN89b]. Lannes [Mar85a]. Lansing [Luk68]. Lapidus [Gos83b, Wil72]. Laplace [Ask75b, Daa97, Fro60, Gas63, GRYL89, Jag82, Lan67b, OW89b, Ste67b, Sto61, Sto62a, WS71, CZ94, Che68, Che73, ES08, Gla81c, Gla88c, Gol62, Gro98b, JGK^{+}82, Kel60b, Mcn66, Mil95a, OW89a, Ric76a, RY88, Sho66b, Smi72, WY70, Woj06, Zem68, Gib73, Mil95b]. Laplacian [Boe98a, Fin71, GN13, KS84, RRW07, SBH^{+}20]. Laplacians [Lim20]. Lapwood [Fri82]. Large [Ber85b, BCN18, Buz78, Cho87, CG24, Cot76, CD75, Cul78, Cul94b, Cul94a, Dau90, Din94b, Duf84, Eis86a, Ell85a, Esl96, Fal73, GMS05, Gri72, HYS76, Jen78, Joh73, LM90, LSY16, Lon78, LRJ78, MMT^{+}21, MBK08, Naz80, Ode85, OR21, PR91, PL78, PEGK03, PO76, RLT65, SBR08, Sch93d, Som78, Ste62, Tor62, Tur78, Var85, Wat86a, WC86, Wes84b, Wes84a, Cha94c, Din94a, Par74, Sez90]. Large-Amplitude [LM90]. Large-Eddy [MMT^{+}21]. Large-Parameter [Esl96].

Large-Scale [BCN18, Buz78, CG24, GMS05, PR91, PL78, SBR08, Som78, Tur78]. Largest [CD75, Cul78]. Larry [Ahl82, MP62, Mil92, RJ90, Wil93]. Lars [Hoc80, Smo91b, Wil85a, Wil86a, Smo91a]. Larson [Ber85g, MN83]. LaSalle [Noh62, Sno76, Beb79]. Lasdon [Fal73]. Laser [All92a, All92b, Car91, Hoc66, PEGK03, VN76a]. Lasers [Kut06]. Lasota [Guc87]. Lasser [Bos95]. Last [BWYY96, YY96]. Laszló [Bla89a]. Late [Ano80g, Ano81c, Ano83f, Ano84f, Ano84g, Ano85b, Ano91e, Coh81, LT90, MO66, SM65, Tri86, Ano82f, Ano87a]. Latent [Cha95b, DEPC95, Lan64]. Later [Ano95g, Ano95h, Ano95i, Ano96h, Ano96i, Ano97g, Ano97h, Ano97i, Ano98c, MV03]. LATEX [GM96]. Latin [Gil65a]. Latitudes [Woo96]. Latta [Whi76a]. Lattice [Ben70, BSS05, Lin10, Maz84, NC09, Sil66, Sil67, SJG96, Wei63a, dB65b, Gen96, Tak83]. Lau [Jd93]. Laub [Die92, Dis92, Hig06]. Launder [May90]. Laurence [Bla90]. Laurent [Kor68, Pao88, Car70b, Car71d]. Laurie [Ame86b, Hig05]. Lavenberg [Bru86]. Lavine [Gas95]. Lavrent’ev [Ham65b]. Law [CSN09, GLY^{+}18, Jef92, OR21, Rei78b, Rob94b]. Lawler [Dur92a]. Lawrence [Cla83b, Dat69, Fet68a, Fin84b, Has92a, Kel84b, Koo77, Ral70, Has92b]. Laws [Ant92a, Ant92b, Bal97, BC98, EHS17, FHRO95, GRL98, HJ00, HLvL83, Sle85, ES80, Pit91, O’B95]. Lawson [Gen76]. Lay [Sop82, TFL23]. Layer [Ble81, BdCPT12, Boy05, CGRG^{+}23, GZ19, Kel75a, Mat75, Mat76b, MG76, O’M71, OT12, Van94a, Van94d, VH78, Wil61]. Layered [Mil82b]. Layers [AG14, DGK16, Eck72, Joh83b, Let69, Ril75, Joh83d]. Layouts [Wol86]. Lazy [Per93b, Per93a]. Lead [Ano90a, BKR75, Kla90b]. Leadbetter [Kaw68]. Leah [Seg88]. Leakage [LH75]. Leaky [ADW23]. Learning [AE23, BCN18, Gut82, HH19, LMMK21, Nor74, SG81, Tho75, ZZ09, Dre89]. Least [BS67, Ber64, BG98, Bog75, CS64, Cho76a, DS79, Die62, Ewi67, FSW^{+}66, Fie93, FR63, Gen76, Gla88d, GO67, GNZ20, Jok66b, Jok66a, Kel79b, Mei68, Mil73, Naz80, New62b, Nie94b, NS75, Pro61, Ren89, RJ89, RW80, She09, Ste77, ST91b, Wah65, Wal66, Wat22, ST91a]. Least-Squares [BG98, Bog75, CS64, Cho76a, Die62, GNZ20, NS75, She09]. Leavitt [Lud82]. Lebart [Ber85b]. Lebedev [Buc69, Wim65]. Lebesgue [Ral62a, Van98]. Lecture [O’M10e, Lig76]. Lectures [Ham77, Hei93b, Hes86, Kal93b, Arm94, Bre62a, Cav90, Day87, Kal93a, Por72, Rei71a, Rhe88, Saa73, Sei62, Tal81b]. Ledder [Gla97b]. Ledet [Rei96]. Ledley [RA63]. Lee [Chi83, Dat69, Gla86a, HR62, Hoc79b, Man81, Mil89b, Mor74b, Sch80a, Wri60]. Leela [Cha70a, Maw91]. Lefschetz [Dia59, Ler64, Noh62]. Left [Cat21, Gea81]. Leftward [Olm76]. Léfvre [Yan92]. Legendre [Jor80a, Jor84a, Jor84c, Bus63, Gla83, RS79, Raw83, Sim70, Sim71b, Zem68]. Legged [HFKG06]. Lehmann [Kup62]. Lehnigk [Bre68]. Leiba [Meh96]. Leibniz [Ros73, Kre91]. Leibovich [Abl77a]. Leigh [Hag82]. Leighton [Wol86]. Leis [Wil88]. Leitmann [Dat67, Foo63]. Lekhnitskii [Joh67]. Lekkerkerker [Zwe84b, Zwe84a]. Lemeshow [Whi92a, Whi92b]. Lemma [Bel68, Dax97b, Mor78, PT07, Wag69, Dax97a]. Lempel [Car71a]. Length [Abd83, BGML96, Fre82, Kla76b, LS95, Loc84a, LRV76, MT84, Rei59b, SB98a, Boe96a, Loc84b]. Length-Scale [SB98a]. Lennart [Dev94a]. Lens [LW17]. Lenses [NS02]. Leo [Dum96, Gol69a, Seb70]. Leon [Fal73, Gos83b, Wil72]. Leonard [Fan59, Köl83, Tho76a]. Leone [Dan66]. Leonhard [Gau08]. Leonid [Har72b]. Leontovich [Mey75b]. Leopold [RJ90].

LePage [Sto61]. Lerma [Hüb91]. lesan [Hor89]. Lesieur [Moi89]. Leslie [Lum85]. Less [BL13, Naz86]. Lesser [Hur93a, Hur93b]. Lessman [Har61b]. Lesson [O’B11]. Lester [Wet74a]. Letters [Jon92c]. Leugeuring [Nic96]. Leung [Boj93]. Lev [Dau90, Har72b, Wil62c]. LeVeque [Pit91]. Lever [TA90b]. Leverett [CP78]. Leverrier [Hou98b, Hou98a]. Levin [Bra85b, Whi63b, Zar65]. Levine [Mar81]. Levinson [Fis71b, Pap85]. Levy [Gas63, Har61b, Whi92a, Whi92b]. Lew [Enr87, Fly89b]. Lewin [Köl83]. Lewis [Bau89, BdDJ83, LM81, Mor88d, Pra85, Sch89b, Sin79b, Spi84b, Spi84a, WS71]. Li [Kov96, Mey95a, Sto96]. Liaisons [CGP21]. Liang [Mor74b]. Liapunov [Bre68, Wou64b, LaS64]. Libang [Hoc94]. Liberman [Sau62]. Libraries [Cod74, DW95]. Library [Bur78a, EGJK04, LMMK21, Sal63b, ZMPK24, Kri86]. Lichnerowicz [Mar85c]. Lichtenberg [Mey84]. Lick [But92]. Lie [Gru74, Gui92, Her80b, Joh84b, Kum88, Tu92, BKS66, BKS68, BK69, Hof71, Joh84a, Sch88, TTM05, Udw89, Abl80, Gru74, Gru88]. Lieb [FLM22]. Lieberman [Mey84]. Lieberstein [Kel75b, Par69, Sch73a]. Liem [Rüd97]. Life [Bel86, Coz87, Gau08, Har78, KA89, Mil94b, Ruo90b, VdLBF17, Bos99, Ruo90a]. Lift [Dre76]. Lifting [FGP^{+}22]. Light [De 00, Gla93d, Gla94c, Lar94, New65a]. Lighthill [His77, Mir88, Com72]. Lights [Lag82, Los83b]. Lih [Hol91b, Sha82a]. Lih-Yuan [Hol91b]. Like [Köl87, Win90, Ypm83]. Likelihood [AB69, Bic94, LMDLS23, MS75, RW84, WP78]. Limit [AG76, BP94, Boe93, Boe94, Boe95a, Cha83a, Cha85, Der86, DCL^{+}86, Gav82, Gla75a, Gol75a, Gol75b, Har78, JG82, JMH95, KR98, Liu94, Liu95, Lux69, MR72, NS66, NW70, New81, Ost91, PSV12, PB96, Ren84a, Rob80, SNS66, Sie89, WGG^{+}92, Gla76, NW81, Ren94c, Bla70, Boe76, Bry96, Dru64, Hud95, Jor93b, Kha71, Los73, New82a, Pro91, Ren84b, Wag92]. Limited [Bon67, BS92, Hue92a, SES75]. Limiting [Moo65, Van75, LS66]. Limits [Che92a, Fer97b, Gua92, Fer97a]. Lin [Hol86b, Kap76, O’B11]. Lind [Boy97]. Lindberger [Jag72c]. Lindblad [Spo85]. Lindenmayer [Ant92d, Ant92c, Are91b, Are91a]. Lindgren [Ros63]. Lindman [Jow76]. Lindstedt [Vis01]. Lindvall [Kal93a, Kal93b]. Line [BK87, CZ96b, Deu95, Ell02, MT76, Mon60, NO60, PBL98, RR97b, Sou76, Zem96, RR97a]. Linéaire [Ham87, Ost61a]. Linear [ABGJ21, All78, Ame78, AB73, And77, Ano93d, Ans85, Ant86b, Apl95, Ari91a, Ari91b, Arm76, Aro78b, Ars60, Ast70, AC78, Bak66, Bal78, Bal88a, BT69, Bar78, Bel90a, Ber81, Ber90a, BDO95, BR85, Bjø90, Bla85a, Bla88d, Bla89b, Bon67, Bor62, BGF^{+}95, Boy05, Bre62a, Bre64a, Bre68, Bri63a, Bri63b, BL06, CL62, Cha94d, Cha81a, Che67a, Che76, Che78, CS64, Cia92a, Cia92b, CP76, CBS76, Con96, Con66, Cot73, Cou97, Cre97, CZM96, DO98, DTG98, Dar97, Dax90, Day89, DTV14, DGP^{+}16, Die62, DGK84, DW95, Dri68, DG14, Eas89, Eck72, Ell61, Enr78, Enr89, EC76, FTM95, FG04, Fic60, FM68, Fle65, For67, Fro83, GPS90, Geo84a, Geo84b, Gib97, Gol76c, Gon92, Goo59, GL98, Gre63a, Gre59, Gri61, Gri72, Gru83, GS23c]. Linear [Gup98b, Gus76, Hag82, HIC96, Han94a, Han94b, Har96, Har70, Har66b, HV75, HYS76, HNP91, Hei93a, Hel78, Hew76, Hig86, Hig99, Hig00, Hor95a, Hue92b, Hue96, Hur65, Joh68, Joh73, Jos70, JR08, KL76, KW71, Kap66, KB76b, KSS80a, Kel79a, Kel88c, KMH66, Kla76a, Kla84j, Kla84c, Kor76, KZ93, KA95, Lae76, Lak80, Lan70, Lau10, LHFE97, Lev88b, Lie97, LW13, LW17, LLR^{+}81, LRJ78, LL76b, Mac76b, Mai72, MS76, MO78, Mar83b, Mar83a, Mar87, Mar95, MH96, Mas92, Maw93b, Mei68, Mel91, Mel92b, Mel92a, Mel93, MR78, Mie17, Mil73, Mil90a, Mil90b, Mit91b, Mit91a, MN76, Mor93, Mor96, Nas90, Nau86, Neu98, NW82, Nie89a, NGTB15, NS75, O’M75, O’M08e, Pac88c, PPR11, Par74, Par61, PE76, Par72, PZ64, Por72, Pra68, PR70, Pru75]. Linear [Rab68a, RCC71, Ral62b, Ral69b, RFP10, Ren92a, Riv63, RL95, Ros61, Ros79, Ros87, Ruk93a, Ruk93b, Rus78, Rus90, SZ96, SK76a, San82, Sau76, Sch77a, Sch04a, Sem59, SES75, SS98, SR98, SA78, Sim16a, Ske78, SV00, Smo91a, Smo91b, Sta92c, Ste61b, Ste77, Sto62b, Str75, The91, Tra66, Tre97, Ung77, UFW17, VG78, VP09, Var76, Var79, Vie15, Wan91, Wat86b, Wat86c, Web85, Wil70a, Wil85a, Wil86a, Wil70c, Wil81, Wim67, Win60, Win62, Wis60, Wit22, Won64a, Wu75, YZZ04, You76, YH76, YF78, Zak84a, Zak85, Zie79, ZM22, Zwe84b, Zwe84a, Cas80, Gup98a, MCS69, Lie97, BI67, Bla85a, Bla89b, Bla66, Bre62b, Che92a, Coc84, Cot94, Cra84, Eas89, Fia73, Fin76b, Gua92, Gue72, Hau89, Kap64, Kar64, Kor85, Kup75, Lie95, Lin80, Lin90, Mac61, Maw93a]. Linear [Mey77b, Miu76a, Moo77b, Nau86, Nit90, Olv86, Ren92b, Ros82, Whi76b, Wie87, Yak95, Yau68, Zak84b, Zha96, Sta92b]. Linear-Differential [Geo84b]. Linear-Programming [Sau76]. Linear-Quadratic [YZZ04, Cas80]. Linear-System [CBS76, LL76b, NS75, Ren92a, SK76a, YH76]. Linearity [HL63]. Linearized [LW95, Riz94, Kra67]. Linearizing [Duf70]. Linearly [Smi80a, Smi81]. Lines [CH61, Nie96, Rai59, Van75, Wal92, Wal93c, Wal93b, Wil64a]. Ling [Ame91a, Ame91b, Car80]. Link [DZH^{+}04, Dix91, KP87, LLSN96, Nic96]. Linkage [Bla81]. Linnik [Luk75]. LINPACK [Par81a]. Linus [McC85]. Linz [Ans80]. Lions [Fal70a, Fle85, Ker92, Lan86, Lau74, Sco94, Sik92, Sle87, Tul90]. Liouville [And95, Kaz91b, BF96, BV96, CA75, DS61, PA95, Ste64]. Lipka [Tad93]. Lipow [Aro63]. Lipson [Cha83b]. Liquid [Kel90a]. Liquids [LR00, Fri71, Obe77, Pip92]. Liron [wBJ77]. Lisa [Thr94a, Thr94b]. Lisek [Ser94]. List [Mac68b]. Literature [BD64]. Lithotripsy [HSSZ98]. Little [Som93, Str89b]. Littlewood [Dur93a]. Littman [Cum96]. Liu [Duf84, Ren93d, Sav18, Whi69]. Liusternik [Kol63]. Liverman [Kap65]. Livesley [Mal85]. Lj [Ano87a, Gla86b]. Lliboutry [Pit89]. Llorenc [Mat94b, Mat94a]. Lloyd [Aro63]. lntegral [Gla71]. Load [VBVM22]. Loaded [Lin91a, Lin91b]. Loading [SSW99]. Loads [Hea91]. Loan [Bai93a, Dem86, Dem90, Bai93b]. Loans [GH22]. Local [Fel78, Fic60, FGTW16, GG80, Por64, Van12, Ant86a]. Localization [BH15, HT12, Hou70a, SBR08]. Locally [CBG75, Mey76a]. Location [Abb92, AB69, Bel65, Bor62, Coo64, Deu81b, Eis62, GK91, Mak86, Meg88, ND82, Spi76, Bur81b, Nev82]. Location-Allocation [Bel65, Coo64]. Locations [Rog83a]. Lochak [Coo91]. Lock [Nas95, VdLBF17]. Locked [Kut06]. Locomotion [HFKG06]. Locus [BS63, Cop07, Kra70]. Loeb [Wil61]. Loeve [Har64b]. Log [BK88]. Logan [Ame95, Dav91b, JG70]. Logarithmic [BDS92, CV91, TBDS92]. Logarithms [MM06, Sho99]. Logic [Bro83a, Har67d, Hig62b, KSW96, McA96a, McA96b, MTM96, Mor74b, Mul66b, Ros73, RM96, Tab97, Tal81b, Vog91b, Tal81a, Vog91a, Wei79, Har72b, McG67b]. Logical [McN59]. Logically [CHL08]. Logistic [Geo95, Mic88]. Logit [Wit75]. Lokenath [Gib85, Kau86, Syn93a, Syn93b]. Long [New76a]. Loo [Hab83b]. Look [Art80, Asm98, Bar74, HRW12, O’M87a, Way82]. Looking [CK18]. Loon [Rus90]. Loop [MN65]. Loos [Che86, Dan91]. Loper [Rob95b]. Lorch [Ans64d]. Lord [McL86b].

Lorden [Rei89]. Loren [Ber85g]. Lorentzen [Thr94a, Thr94b]. Lorentzian [Her83b, Her83a]. Lorenz [Dev92, Hol85, Hol95, Paz76, Pon90a, Pon90b, VŞ10]. Loskutov [Haa93]. Loss [Coo93, Coo94, DeS88, JvD90, JvD91, New94a, Rai59]. Lot [New60, New64c]. Lotfi [Kap64]. Lothar [Gra66a]. Lotka [Ole94]. Lottery [Pen95]. Lotto [Liu92a, Ren93b, Ren93d]. Lotz [Bus70]. Loud [Her65]. Louis [Ker92, Mor75b, Por83, Thi63a, Tul90, Wer63, Gue92]. Louise [Fin95]. Lovász [Bla89a]. Love [Fre80, LFR80, FLM22]. Low [Bri65, DK89, DG20, Fej88, Jin91, LR00, LW04, Dij90]. Low-Frequency [DK89]. Low-Gravity [Dij90]. Low-Order [Jin91]. Low-Pass [Bri65]. Low-Rank [DG20, LW04]. Lowell [Bra91a, Bra91b]. Lowengrub [Bue71]. Lower [BK17, FR66, KS68, MW76b]. Lowest [AA78]. LS [She09]. LS-MGD [She09]. Lu [Kel92b, Rüd97, SM22, Kel92a]. Lubich [Cam91b, Cam91c]. Lubrication [Lin86]. Luc [Ric88, Ste87]. Lucas [Kla84i, ML85, Kla84b]. Luce [Sca59]. Lucien [Sag78]. Luck [GLY^{+}18]. Ludford [Kap85b]. Ludovic [Ber85b]. Ludwig [BR76, Mar95, Vog87]. Ludyk [McC83]. Lueker [Kv93]. Luenberger [Fal70b]. Luigi [Lie95, Log97]. Lukacs [Kes65, Rob62a, Ser80a]. Luke [Cod78, Olv72, Tha79]. Lukes [Pal84b, Pal84a]. Lund [Ste93a]. Luneburg [Ivi65]. Lung [Tha96]. Lung-an [Tha96]. Lunneborg [Car87]. Lützen [Kaz91b]. Luxemburg [Bla70, Kap78]. Lyapunov [Dri65, Bus69, Dat82, GQJ96, LW04, Jód96]. Lyle [Kaz92]. Lyness [AF69, Ste68b]. Lynn [Ham80]. Lyusternik [Kog66].

M [Abb92, Aga94, AK90, AS85, ASF86, Ame85, And77, Ano71a, Ans64a, Ant83a, Ant91, Arb76, Aro63, Art83a, Art85a, Ask80, ABG96, Atk66, Bai74b, Bar95, Bar64, Bar65, Bar89, Bar59, Bau88, Bei81, Bei85, Bel62, BI68b, Ber85a, Ber85b, Ber85c, Ber97a, Ber86a, Ber83b, Ber84e, Beu85, Bir64, Bla82, Ble81, Blo95, Blu94a, Boa83, BGML96, Boe76, BdDJ83, BG95b, Boe95b, BC96a, Boe96b, BS86, Boy88, Bra85a, Bre62a, Bre64a, Bro76b, Bry93, Bry94, Bue66, Bue71, CS70, Car71a, Car76b, Car66c, Cha94c, Cha86, Chi71, Chi94a, Chu82, Coh75b, CZ89, Col94b, Col86, Coz87, Dam84b, Dam84a, Dau90, Day89, DG96, Deu81c, Dod73, Dri65, Dri81, Duf91a, Dup93, Elk86, Erd83, Erm89, Fan59, Fel90, Fet68a, FS68, Fet72a, FLR^{+}81b, Fia73, Fla67, Fle76, For92]. M [FARE69, Fra95a, Fre83, Fro80, Fro87a, GK86a, GR81, Gla85a, Gla86c, Gla93b, Gle83, God87, Gol65b, GG91, GH86, Gos84, Gra86a, GRL87, Gra79, Gre60a, Gre63a, Gro87, Gro94a, Gro90, Gro96e, Gro88b, Grz83, Gue83c, Gue89, Gui92, GWK95, Hag83, Ham90, Ham65b, Ham77, Har72b, Hel97, Her67b, Her67a, Hey82, Hey86, Hin87, Hoc65, Hol87, Hor83b, Hou67, Hua96, Hur65, Isa89b, Ism83b, Ism83a, Ism96b, Jag80, Jag82, JGF87, Jag88a, Jag90, Jef95, Joh73, Jon92a, Jon92c, Jon92b, Jor81, Jor82, Jor84a, Jor88a, Jor90b, Jor93b, JGM95, Jow72, Jun65, Jur95, Kal70, KH85, Kar60, Kar82, Kar83, Kar84b, Kar87, Kau88, Kaw68, Kaz62, Kee88b, Kee90, Key94, Kha87, Kie94, Kla80, Kla83a, Kla84i, Kla84j, Kla86a, Kla88a, KG89, Kno81, Knu77a, Köl87]. M [KF62, Kor65, Kor95a, Kra67, Kre96, Lae76, Lam86, Lan81, Lan67b, LG80, Lar94, Len95, Lev63a, Lev87, Lew83a, Lew91b, Lin80, Loc84b, Los74b, Luk75, LSB67, Mac68b, Mac61, Mah94, Mal80, Mal90, Mal88b, Man63, Mar76b, Mas87, Mat89a, Maw96, May84b, McA96a, McL79, Mei78a, Mer85, Mey77b, Mey84, Mey91a, Mey93b, Mey94a, Mie95, Mik89, Mil82b, Mil94a, Mil95b, Mil89b, Moi89, Moo74b, Moo76, Nas77b, Nas95, Neu97b, New87, New74, O’M96, Oco87, Old62, Olv97, Ort74, OW89b, Osb85, Par67b, Par92a, Pav94, Por83, Por86a, Por92, Pra86, Pro83b, Pro83a, Qui92b, Rab94, Ral62b, Ral74, Ram90a, Ram93, Rat87, Rau95, Red90, Ren89, Rhe85, RD65, Rou72, Rüd97, Rue72a, Rue74a, Rue76b, Rue81a, Rus95, Rus90, Rya63, Saa68, Sal64, Sam76b, Sat78, Sau78, Sch84b]. M [Sch94, Sch81a, Sch60b, SR93b, Sed88, Sen94, Ser96, Sha65, Sha82b, Shi94, Shi68, Sin93a, Ske95, Sle89a, Sle92, Sle83a, Sma76, Smi74, Sop59, Sop60, SFH^{+}66, Str62a, Stu90, Tad89, Tad93, Tap86, Ter92, Tho87, Tho75, TBDS92, Tot91, Tu92, Vai96, VK62, Vog93, Wei77, Whi67, Whi83, Whi84b, Whi94, Wil62a, WR94b, Wim86, Wim88, WA63, Zar65, Zas65, vWHE91, Chi94b, Sch78b]. M-Matrices [Sch78b]. M. [Ber84d, Kar84a, Kla84b, Kla84c, Sch84a, Wal97, Whi84a, Wol93]. M/M/Infinity [Hey86, New87]. Ma [BKS96]. MacCallum [Olv91a]. MacDonald [Kaz80b, Mil90b, Mil95b, Mil90a, CZ94, CZ96a, CZ97, Mil95a]. Machielsen [Che91]. Machine [BM95, BCN18, DGK84, GP92, Mar86b, Sal63b, Tod74]. Machine-Second [Mar86b]. Machine-Tool [GP92]. Machines [AM74a, Ker91a, Ker91b, NT78, Mul69, Yau69]. Machining [Far93b, Far93a]. Machol [Eng78, Man61b]. Mackey [Guc87]. Mackiw [Bir86]. Maclaurin [Jor87a, Fre86b, JF87]. Macmillan [McG64]. Macon [Moo63]. Macroeconomic [She77b]. Macromolecules [Kru83]. Macroscopic [CHO92, XC11]. Macsyma [Fat76]. MacWilliams [Mat80]. Made [Flu90, Nie89b]. Madow [Man61a]. Madras [Bur94b, Bur94c]. Madsen [Kar78]. Maëlle [Law18]. Magenes [Lau74]. Magnanti [Bal95]. Magnenat [Odg86]. Magnenat-Thalmann [Odg86]. Magnetic [LH75, Núñ96, SW11, Chi79]. Magnetically [VN76a]. Magnetohydrodynamic [Kra67]. Magnetohydrodynamics [Bla66]. Magnetosphere [Fej88]. Magnitude [Hue91, LS63]. Magnus [Cul77, Kaz67]. Mahajan [Reh89]. Mahaney [Bla88b]. Mahmoud [Fel90]. Mahr [GM87]. Maier [Cot73, Mey75b]. Maindonald [Gen86]. Mainframe [TH96]. Maintaining [CT95, Tab95]. Maintenance [Bar59]. Maistrov [Har76a]. Majda [Sle85]. Majorization [Fan82, Ton88]. Makes [BCD98b, FC76, BCD98a]. Making [BI18, BL13, Grc18, KM12, Mal81, Mal83, Pfl92, Bas88]. Malaysia [Zwe16]. Mal’cev [Dav77]. Malcolm [Olv91a]. Malevolent [Lag82, Los83b]. Malhotra [Los85]. Maller [Ken93]. Mallgren [Ber85d]. Mallows [JG70]. Malte [BdD93]. Malyshev [Bak91b, Bak92]. Man [Gau08, Lan86, Mar86b]. Management [Aro63, Cla78a, Gue88, MM95, PM87, RV78, Fro85, Pol59, Bra82, Win62]. Managers [Gou10, Hig01b]. Managing [TL02]. Mandelbrot [Are93, Goo84, Dev92]. Mandell [Seg85]. Mandersheid [Ahl93]. Manfred [Gil90a, Sal66, Gil90b]. Mangasarian [McC71]. Mani [Bru83]. Manifold [Abe88, CMSZ24, Rou97b, ZZ09, Rou97a]. Manifolds [BW11, Has86, Liu91, NRS19, O’M92, Pal95, Pin84b, Pin84a, WP95, Hag85c, Par87, Rat93, Lee84b, Ran89, Wei88]. Manipulation [Lon87]. Manipulators [MT96]. Mann [Leh63b, Man61a, O’M87a]. Manna [Tal81b]. Manning [Sch94]. Manocha [Wim86]. Manometer [Kel90a]. Manoukian [Mor88b]. Manpower [TK82]. Mansfield [Cla64a]. Manual [Pin85a, Pra89a, Tha82]. Many [Cat21, Hag85b, Kaw63, Kel95, Mac00, vLMZ19]. Many-Body [Hag85b]. Many-Server [vLMZ19]. Mao [Kus93]. Map [BD94, Wan91]. Maple [Fro60, NWF97, HK95, Koe95]. Mapping [Cer70, Cer71, DEP01, ET99, Pea82, PFM76, Ric89, San65, Ton89, Tru94, Whi63a].

Mappings [Den67, Ear81, Hes83, MWL97]. Maps [Boa87, BFM95, Fee00, Gil74, Hog73, Nit85, Mey95b]. Maps-Univalent [Boa87]. Marat [Ash89]. Marc [Cha96c, Kni94b, Law18, Pit93b, Wal88, Kni94a]. Marcel [Gra83a]. Marching [Mur97, Set99]. Marchuk [Svo96]. Marcia [Her93a, Her93b]. Marciniak [Far93a, Gre88, Far93b]. Marcus [Ren94a, Boy97]. Mardia [You74]. Mardo [Jag87a]. Mare [Pit93a]. Mareels [Bau88]. Marek [Boj97, Ott96, Rhe85, Rob92a, Rob92b]. Margaret [Kow83, KZ93]. Marginal [CD94, LMDLS23]. Marginally [WJ75]. Margreet [Apl95]. Marie [Nas90]. Marjan [Sch85a]. Marjorie [Sch93a, Sch93b]. Mark [Grc18, Gue85, Har72b, Law18, Lig86, Ren93c, Sch93c, Sco83]. Markcov [Rob62b]. Marker [VB75]. Market [Car71a, CFHP10, Lem70, Ell85a]. Markets [HS15]. Markov [Ath93b, BR62, Dur93b, Ell81, Get96, Hey89, Hüb91, Jag71, Men94a, Oco87, Pit74, Por81, Rob94a, Ser80b, Sys63, Tay68, Ang92b, Ath93a, AW19, BDX04, DL95, DKST19, DMP22, Fox67, FOTG96, Kij93, KTSB21, Mey75c, Mey89a, PF96, Ros95, Sch70b, Sch71e, She09, She76, SBXD06, Par88, BR69, Dwa72, Fei96, Kar78, Klo93, Len95, Pin70, Pit81, Tho75]. Markov-Chain [She76]. Markovian [MR18, Nor74, Fox71, Hüb88]. Markowich [Gar91, Pim87]. Markowitz [Ste01]. Markus [Dat69, Hol89b]. Markusuevich [Whi63a]. Marquardts [Shr76]. Marriage [Kri91]. Marriages [BR97]. Marsden [Arm94, Her97, Lee84b, Rob81, Sat78]. Marshall [Fan82, Rei64, Tur60]. Marston [Rei77a]. Martello [Van93]. Martin [Ber65a, Ber78, Bla89a, Bra96, Cab74, Can91a, Car93, Che67b, Coo67a, EK91a, Fei96, Gue84, Hes86, Hoc79a, McA92, Mer81, Mey86b, Mey89c, Noe82, Pac87, Por81, Pul91, Sle79a, Sop72, Ste94a, Ste94b, Tay68, Can91b, EK91b]. Martin-Clouaire [McA92]. Martingale [Mal81]. Martingales [Bjö85, Pit93a, Pit93b, Sam80, Kus93, DD74]. Martos [Zlo77]. Martynyuk [Maw91]. Marvasti [Wit90]. Marvin [Ken74, Mee96, Ren94a, Yau69]. Mary [Bri64, Mil69a, Ral86]. Masaaki [Wil92]. Masanao [Lai91, Mor71, Tze91]. Masao [Aeb94b, Aeb94a]. Masatake [Mal88a]. Masatoshi [Get96]. Masayoshi [Get96]. Maslov [Har96, Pro83b, Pro83a]. Mason [Far94, Hau86, Hud95, Jur95, Lon87]. Mass [Ars60, BS09b, CN15, Pet97b, PS00, WPL98, Pet97a]. Mass-Conserving [CN15]. Massey [Eps59a]. Master [Kaz91b]. Matched [DGK16, Hop75, Wol77, O’M74]. Matching [Eck94, O’M94]. Matchings [Bal70]. Material [BTKNZ00, MT84]. Materials [JC12, SCS10, Sle87, Set62b, Tau80]. Math [Bol91a, Bol91c, GM96, Lax89, SZ99, Ula67, Mey96]. Mathai [Ask80]. Mathematica [Car04, Car08]. Mathematical [AK77, ABJK10, Ano75g, Ans68, Arb76, Ari91a, Ari91b, AR63, Ass77, Bar95, BNT76, Ber91b, BWYY96, BvdGH95, BD64, Bra94b, Bra94a, Bra85a, Bra81, BSP95, Bur78a, But74a, But74b, But74c, But79, Car81, CIPB95, Che01, Cla79a, Coh76, Col94c, Col94d, Cro73, Cum92a, Cum92b, Dau90, DeW85, DC76, Dia64, Día78, Dij90, Din72, Dol91b, Dol72, Dre89, DHL17, EK91b, Ell84, Enr87, Fix93, Fla97, Fle73, For76, FL81, Fre83, Fro87b, Geo76, Gla94a, Gos64, GZ11, Gre63a, GK95, Gue69, Gue83a, Gue83d, Gue95, GAM93, Gur62a, Hal90a, Hal90b, Ham65a, Ham65b, Ham70, HCH^{+}75, Hei95, Her80a, Her93a, Her93b, Hog73, Hol79, Hye73, Ins78, JW83, JRA22, Kap70b, Kaw65, Kaz67, Kaz91a, Kee95, Kel96, Ken91b, Ken74, Kes85, KG97, KD10, Kno81, KB59]. Mathematical [Kor65, Kur69, Lev97, Log97, Luk75, Lus87, MNZ93, Mac77, Man90c, Man90d, McL86b, MG96, MQ95, Mil82c, Mla94, Mul64a, Nau86, Neu97a, Neu97b, O’B11, Ode82, Olv97, Owe85, Pas95, PN99, Pit79, PI88, Pol87, PCH98, Pri93, Pro87a, Pug86, Rei63, Rog94a, RCM07, Ros73, RS75, RRG95, RM95, San83a, Sch93a, Sch93b, Sch80b, Ser64, Shi71, Smi74, Tha79, Thr79, Ulk76, WP88, WR96a, Wen78, Wet74a, Whi81, Wil88, Wil76a, Wol75, Wu72, Zar68, ZK96, Ans64b, Art88, Bak91b, Bak92, Bel90b, Ber65a, Ber91a, Beu87, Chi83, Cod78, Coo65, Coz85, Cum91, Daf89, Dat73, Dol91a, Dwa66, EK91a, Eic67, Fle64, Fre79, Fro85, Gar67, Ger96, Hag80, Hai85, Ham79a, Ham79b, Hav95, Him78, His77, Hop83, Hul61b, Ivi65, Kaz80a, Kel75b]. Mathematical [Ken91a, Ken93, Ker92, Lit63, Mac92, Man61d, McC95, Mer81, Mik89, Mor88b, Nel74, Ner70, Pat64, Pin87a, Qua95, RV78, Ren94d, Rog94b, Rog96, Ros65, Rou88, Sán83b, Sau79, Saw82, Seg79, Seg88, Sel76, She91, Shi69, Sit92, Sop62, Tul90, Uhl73a, Uhl73b, YY96]. Mathematician [Ano75g, HCH^{+}75]. Mathematicians [Col76b, Dur84a, Gas60, HH19, Joh83a, Lin78a, Man82, McM62, Dur84b, Joh83c]. Mathematics [Alb10, Ano72o, Ano73n, Ano73m, Ano75d, Ano75o, Ano77q, Ano78e, Ano78f, Asp84a, Bir69b, Bla81, BCD98b, Bra91a, Bra91b, CCR^{+}62, Car67e, Car65a, CHT15, Cro77, Cum96, DA79a, DA79b, Dav91b, Dra61a, Ede97, FKL^{+}06, Fir92a, Fir97, FLC96, GVL97, Gou10, Gue83a, Gue83d, Gue83b, Her81, Het00, Hig00, Hig02c, Jaf84, Jun64, Kap90b, Kap90d, Kap68, Key73, Kla92b, Kre66, Lax67, LCK^{+}17, Lin67, Lin76, Mac68a, Nie89b, O’M05d, Ock92b, PS07, Pet91, Ral80, Rap73, Ros67a, Row79, Row86, Ruo90b, Saa96, Sam73, Saw59, Seg80, Sle74, Sle84b, Sle84a, Spa77b, Spa77a, SKK20, Ste94a, Ste94b, Sto11, TS96, Tho77a, War92, Wat10, Wax91a, Wax91b, Wei80a, Wei69, Wes92, BCD98a, Lax89, Tre15, ADPR04, Aro63, Ask79b, Bor06, Det85, Eng78, Fir92b, Gol71, Gor59, Gue79]. Mathematics [Ham71, Ham73a, Ham76, Ham77, Hen86, Hoc82, Jun64, Kap71, Kaz91b, Ken86b, Lat69, Lev63a, Luc86, Mey82b, Nas87a, Ral66a, Ruo90a, Whi69, Wig85, Wil93, Bor09, Lax89, Asp84b, Coo89, Coz87, Ham80, HF05, Hoc79b, Kap90b, Kap76, Kor68, Mil94b, O’R92, Ock92a, Pac10, Sag81, Sch80a, Ull88, Wes92]. Mathematik [Gra66a]. Mathématiques [Ham87]. Mathematischen [Gil65b]. Mathieu [Alh96, BCZ21]. Mathieus [Ike75]. Mathletics [Gou10]. Matiur [Mor97]. Matiyasevich [Fen94]. MATLAB [Hig02a, Nei10, PS04, SRK99, Ver09]. Matri [Dec65a]. Matrices [Ack73, And98b, Ars62b, BT19, BI68b, BI69, BI70c, Ber73a, Ber86a, Ber90c, BDJ99, Bha98, BG77, Bos87, Bre61, Car81, CG61, Cho69, Cle59, Cli64, Cli78, Coh88, Coh93a, Cop86, Cul78, Dav91a, DGK84, DU10, Fai71, Fol73, Fou84, FW62, Gre60b, Har67a, HS81, Hig87, Hig03, HSŠ16, HUS10, HT12, Joh87, Joh88, KL97b, KB76b, KF76, LP71, LSY16, LW59, Lon78, Lou70, Man68, Mar73, Mar60, McA79, Mel12, Mon70, MP70, MO68, Mul66a, Nik78, Pan92, Par74, PC90a, PB74, REG19, RM89, Rei97b, RKM76, Ros69, Rum99, San76, Sch68, Sch69b, Sch71b, Sch73e, Sch78a, Sch78b, SB72, SM14, Tau67, Tew66, Tew67, Tew70, Urq68, Urq69, Var66, WKW75, Wat91b, Wat91a, Wig67, WM78, Wis60, ZZ09, And98a, Coh93b, KL97a, PC90b]. matrices [Rei97a, Gib97, Ack74, Ber85b, Cha82b, CZ89, Dic67, Djo68, Eis86b, Fan59, Hea74, Mar64, Mil81b, Tho74, Wil60b]. Matrix [Alp96, AT76, Aur18, Bar83a, Bar22, Ber88b, Ber90c, Ber96, BPE96, Bos87, BGF^{+}95, Boy13, BS62, Bro82, Bro83b, Bus76b, CESV15, Cav74, CGO07, Cha83a, Cha95d, Cha95c, Cho72, CFG95, Cos93, Cox94, Cra74, Cra75, CD75, Cur77, CV77, Dar97, Dav89, De 20, Dec65b, Den90, Den68, Den69, Deu81a, DTT98, DG20, Ebe96, EFMM13, ESS79, Ehr76, EGJK04, EH10, Fol64, For75, Fre69, FC76, GQJ96, Gar93, Gil70, Giv78a, GK78, Gol75a, Gol75b, Gol75c, Gol73b, GMP66, GDR78, Gre59, Gre60c, Gre66, Gre67c, GH13, Hag89b, HMT11, Har62, HFS01, Hig07, Hig09, Hol91a, Hol91b, HS88, HMH00, Hul89, HJ90, Hut89, Ike76, Ike78, Jag90, JBG^{+}90, Jes77a, Jód96, Jor81, Kam76, Kat68, KSS80b, Kri77, Kus19, Kwa98, Lan70, Lan81, LR05, Lei76]. Matrix [Leo96, Lew86, LBD^{+}86, Liu92b, Liz98b, LT13, Mac98b, MKR76, Mat69a, MQ69, Mey69b, MV78a, MV78b, MV03, Moo77a, MATL77, NF16, Net20, New63, New65b, Ode85, OLHB02, Pan84, PT81, PC89, PGS75, Pow75, PJ75, Pri64, PR70, PO76, RFP10, Ren84a, Ric63, Ros80b, Rot81, RP96, RK78, Rus60, SS79, Sch04a, Sch70a, Sch72c, Sch78d, SHL^{+}79, Sch85b, Sea61, SK80, SKG^{+}80, Ser76, Shi17, Sim16a, Sou80, SV89, Ste97a, Ste93b, SP01, Swi60, Ter96, Tis83b, Tis83a, TS66, UA07, Usm72, VB05, Ven76, Vet73, Vos75, Vos76, Wal85, Wal87, War76, Wat91b, Wat91a, WC86, Wer89, WBLQ89, Wil60a, Wil86b, WZ72, Wit22, Wol81, Wol82, Yak95, Zie73, Zlo70, dO69, Ber97b, Hor97b, Koh98a, Liz98a, Par10, Pea97, PC90b, Ram88, SGD98, Hig06]. Matrix [BG64, CMR73, Deu81c, FARE69, Hea81, Jes77b, LJW73, New76c, PRM^{+}66, Ren84b, Rob74, Rob76, Rue73a, Sae63, Wat86a, Bar83b, Dem86, Dem90, Gra83a, Mal87, Mil82c, Ple88, Stu91, Wat10, Wer63, Wol94a]. Matrix-Geometric [Gra83a]. Matrix-Vector [VB05]. Matroid [Kun93, Orl93a]. Matroids [Chu89, Whi94]. Matsuno [Abl86]. Matters [Sku04]. Mattheij [Isa89b]. Matthew [Art83b, Art85b, Lon87]. Maugin [Hut91]. Maurice [Sch60b, Sha89b, Whi66b].

Mawhin [Laz91a, Laz91b]. Max [Bak91a, Cho70, DD02, JW83, Pon65, Utz79, Saa68]. Max-Min [Saa68]. Max-Plus [DD02]. Maxima [Ano92b, Mur92]. Maximal [HK92a, HK92b, PO76, Sch69a, Sch71c, Sch72b, Sch72a, Sch73c, Sch73b, Ven72, VN73, OP88]. Maximization [Mou67, WPTZ01]. Maximizing [Bra76b, HS88, LS78, Rei78b]. Maxims [Kob10]. Maximum [AB69, AV65, BJK24, Bic94, Bla94, Den90, DPS73, DJLQ21, Fel92, FL82, Fle88, GS88, HSV95, Hau81, Hol91a, Hol91b, wBJ77, Jor77c, Jor89a, KR78, Lir76, Los83a, LRV76, Mak86, Mal93b, MS75, McA79, Min78, Ols03, PF13, RW84, Rei78a, San93, Sek82, Sen89, Sey93c, Sha90, Sin86, SBXD06, Udw89, WP78, WGG^{+}90, Wat90, dM94, Sha98, Gro94a, Mal93a, Pay68, Red90, Red83]. Maximum-Likelihood [MS75, WP78]. Maxwell [Jon92c]. May [ZWY14]. Maybes [Kap71]. Maynard [Wal75b]. Mazumdar [Man90c, Man90d]. Mcaloon [Bla88b]. McBride [Wim72]. McCarthy [Duf91a]. McCarty [Fre78]. McCauley [Pal94]. McClamroch [Gue82]. McComb [Gli91a]. McCormick [Man70, Via85, Wei91]. McCracken [Sat78, Sin59, Wri60]. Mccullagh [Wat86b]. McDonald [Gue83d, Gue83a]. McGee [Sch65c]. McGree [Sav18]. McKean [Gol73a]. McKelvey [Ken86b]. McLaughlin [Bar86, Cav90]. McNaughton [Zie83c, Zie83a]. McNeil [FS68, Lan67b]. McNeill [McA96a]. McShane [Pin77]. Mean [BK88, BT85, BB83a, BB84b, BB84a, BDS92, CV91, CS17, Har67b, HT72, Hoc67, IWMW21, Maz84, Ste01, TBDS92, Meh73]. Mean-Variance [Ste01]. Means [Car70c, Neu96, RAB02, Rei71b, Sei91, Han92, Sei97, Jag92, MNC71]. Measure [Bal20, BGH^{+}04, Far94, Gud84, Ral62a, Dev91]. Measurement [Bar89, BP95, Gri77, Mil76, PL78, Raz92a, Raz92b, Ros85, Pan95]. Measurements [AMR04, She67]. Measures [Cam77, CHT21, PdHM16, VN76b, Pin69]. Measuring [Sed91]. Mechanical [Duf86, Fre79, Hag85b, LM99, Mor74b, Sha89b, VdLBF17]. Mechanics [Ame73, Ame86a, Ant83a, Ant92b, Bak91b, Bak92, BL84, BK78, Coo73, Cur64b, Dij90, Edw90, Edw63, Fri71, GO09, Gui92, Gut74, Ham77, Hau90, Her97, Hir75, Hol89b, Hor95a, Hor95b, Jen78, Kel85, Kir90b, Kyn76, Lon87, Mac77, Mar85c, MT64, Moi89, Mor91a, Mor91b, Obe77, Pan93b, Pol76, Pra68, Pro83a, Rat92b, Red93b, SC81, Sar10, Spi91a, Spi91b, Sug98, Tu92, Vog87, Vog93, Wei84a, Wes84b, Wes84a, Yat84a, vL86a, Gro98a, Wol97, Ant83b, Arm94, Bal88b, Bro83a, Bro90, Col77, Eis86a, Els89, Gar67, Gru88, Gun63, Hau85, Hau86, Hoc79b, Hor83b, Kee87, Kir90a, Kum82, Mir88, Pan93a, Pol69, Pro65a, Pro83b, Ran89, Rat92a, Red89, Ren94d, Rob81, Run91, Saa73, Sch80a, Sop59, Spo88, Wei84b, Wei88, Ant92a, Red93a, Set62a, Yat84b]. Mechanisms [Duf91b]. Mechanoreceptors [Hol89a]. Medalists [Sch92]. Medhi [Hey93]. Media [BLMT18, Coo67b, JM78, KT92, Kel64, Rog69b, Tre78, WM15, Win90, Xin00, Ren97b, Mea88b, Mil82b, MT64, Row64, Sch89b, She87, Sle87]. Median [Ebe75, JSR82a, JSR82b, Rae81]. Medical [RCC78]. Medicine [Mil94b, Erm89]. Medick [RD65]. Medina [Aiz94b, Aiz94a]. Medium [Abd85, Wil62c]. Medved [Chi94a, Chi94b]. Meeker [Ber93]. Meer [Mey88]. Meeting [Ano68l, Ano69k, Ano69l, Ano71m, Ano71n, Ano72l, Ano72m, Ano73l, Ano74l, Ano74j, Ano74k, Ano75e, Ano76o, Ano76p, Ano77o, Ano77p, Ano78p, Ano78q]. Meetings [Ano75n]. Meets [CPSV18, CGP21, DMP22]. Megapixel [WB09]. Mehrdimensionale [Hes72]. Mei [Ger96]. Meijer [Pro65a]. Meinardus [Che65]. Meir [Rue81a, vK79]. Melda [Bre70, Göb70]. Melissen [SR93b]. Mellin [Aga93, Gro98c, Har67c, Lóp08]. Mellon [Ral86]. Melo [Bru95]. Melrose [Syl96]. Melsa [You72b]. Melting [NR90, New90, Rue89]. Melville [Ano93d]. Melvin [Kel75b, Par69, Sch73a]. Melzak [Ban85, Ham79a]. Members [Ano68f, Ano69c, Ano69d, Ano69e, Ano70e, Ano71e, Ano71f, Ano71g, Ano72c, Ano72d, Ano72e, Ano72f, Ano73e, Ano73f, Ano74b, Ano74c, Ano75h, Ano75i, Ano75j, Ano76h, Ano76i, Ano77g, Ano77h, Ano78j, Ano78k, Ano79f, Ano79g, Ano79h, Ano80i, Ano80j, Ano80k, Ano81d, Ano81e, Ano81f, Ano82i, Ano82j, Ano82k, Ano82l, Ano83g, Ano83h, Ano83i, Ano84h, Ano84i, Ano84j, Ano84k, Ano85c, Ano85d, Ano85e, Ano85f, Ano86d, Ano86e, Ano86f, Ano86g, Ano87d, Ano87e, Ano87f, Ano87g, Ano88b, Ano88c, Ano88d, Ano88e, Ano89f, Ano89g, Ano89h, Ano90c, Ano90d, Ano90e, Ano90f, Ano91f, Ano91g, Ano91h, Ano91i, Ano92e, Ano92f, Ano92g, Ano92h, Ano93f, Ano93g, Ano93h, Ano93i, Ano94c, Ano94d, Ano94e, Ano94f, Ano95j, Ano95k, Ano95l, Ano95m]. Members [Ano96j, Ano96k, Ano96l, Ano96m, Ano77i, Ano89e]. Membership [Ano74d]. Membrane [CMZ94, KS68, Sle76]. Memoriam [Ano78i, Her92b, Her92a]. Memories [Tay73]. Memory [GS23b, Hun85]. Meneguetle [Blo95]. Ment [Min78]. Mercier [Boy90a, Boy90b]. Merging [She76]. Merilees [Jor83a]. Merriam [McG65b]. Mesarović [Arb76]. Mesh [PS04]. Meshes [Ber07e]. Meshing [BH03]. Mesirov [Sch93d]. Mesterton [Cum92a, Cum92b]. Mesterton-Gibbons [Cum92a, Cum92b]. Metabolic [HKS75]. Metal [HS03]. Metaplex [EERG20]. Metastability [BW11]. Mete [Men94a]. Method [Alh96, Aro78a, AN83, BS94a, BS94b, BNT10, BY78, Bar79, Bec94, Ben95, Ber75, BT05, Bet78, Bla85b, Bla88c, Ble81, BR18, Bru65, CC75, CMSZ24, CS63, CZ96b, CGT^{+}97a, CP78, Cop71, Cru86, DTV14, Del97b, DA78, DMP22, Eds05, Edw00, EG98, Flu78, FP96, Fos68, GS23a, Gay76, Gay78, Ger94, Gof78, Gol76c, GS23c, Hal70, Har96, HYS76, Hel75, Hig02a, Hig09, Hir78, HA75, Ike72, IA75, Kal93b, Kel22, Kir12, Kno98, KKS82, Kra70, Kri76, Lee86b, LT98, Liu92b, Mal85, Mal88a, MB64, MH96, Mea86, MR00, MBOV17, MBK08, MA04, Mor66, Ode89, OT13, Per65, PW79, Por85, Pri76, RMT^{+}03, Rog69b, RH76, SB98b, SHL16, Sch89b, Sei76, Sem59, She75, SV00, Smi75, Ste70b, SP01, TDS18, Ton89, TE96, TS86, Und76, Val67, VB75]. Method [VB05, Ves95, Wal23, Was73, WM63, Wil84b, Wil84a, Woj06, Wol77, Ypm95, Zem96, CGT^{+}97b, Del97a, Geo67, Ham77, Lag97, Rüd97, Abl86, Bre68, Cab75, Hag85a, Jun70b, Kal93a, Kel83, Wou64b]. Methoden [Ham85]. Methodes [Lan86, Ham87, BI68b, Cuv90, Ost61a]. Methodologies [Bla88a]. Methodology [Gup82]. Methods [AKP91a, AK77, AG80, AHP^{+}95, Ans64b, Ant87, Aro63, Art88, AC78, AH95b, Bag95, Bai87, Bai74a, Bak91a, BBN98, Bal88b, Bar78, BJK24, BES05, BKS68, BK69, BI70a, Ben74, Ben76b, BGW15, Ber85a, BH76b, BG98, BS09a, BCN18, Bou91a, Boy88, Boy90b, Bra96, BEG23, BSP95, BCM10, But89, CHL08, Cam91c, CD20, Cen81, CG76, CE89, CN96, CL72, Cha94c, CS63, Che67b, Cho75, CG24, CLH97, COSG19, Col77, CZ89, Con85, Coo64, CR78, Cor92b, Cul94b, Cul94a, Cuv90, CS90, Dam84b, Dam84a, Dan71, Dat73, Dau90, DdSG^{+}22, Del95, Den76, DM77, DS79, Den82, Deu85, Din72, DL92, Dra61b, DM07, DG14, Dun94, Dwa70, Eck80, EK91a, EK91b, Ede90b, Efr88, Eis86b, EBTZ94, ETZ94, Eng94, Eps59b, Fab80, Fal93b, FGN13, Fer79, FPM97, Fix76]. Methods [Fla97, Fle73, Fle79, Fle89, Fok84, FGW02, For67, FR66, FR11, Fro74, Fro85, GG99, GZ19, GG80, GHC97, Gea74, Gea85, Gen96, GM95, GG95, Gil90b, GMSW91, Gil95, GP79, Gol76a, GF78, Gol76b, Gon92, GAT75, Got78, GST01, GS22, Gra86b, Gre62a, Gre76, Gro87, Gud75, Gup91a, Gup91b, Hag83, Ham61, Ham65a, Ham79b, HB96, Han86b, Har70, Hau85, Hau88, Her70, HK93, HB95, Hol87, HO96, HRM96, Hul61b, Hul67, Jac97, Jen75, JLM^{+}17, Joh83d, Joh83b, JMX96, KP87, Kap65, Kar76, Kee88a, Kel59, Kel83, Ker92, Kev85, Kir10, Kob10, KLT03, KLT04, KG98, Kui06, Kus76, LM05, Lat69, Law18, LP75, Leh82, Lin79c, Lit63, Loe95, Log97, LBL02, Mal81, Man88a, Man61d, Man88b, Mar75b, Maw86, Maw95, McA94c, McA94a, McA94b, McL86a]. Methods [Mea88b, Med78, Mel97, MMT^{+}21, Mey76a, Mey82b, MOSK97, Mil83b, Mil83c, MM10, Moo75, MSS96, MNS02, Mur94a, Mur94b, Mur97, Nas77b, Nau86, NBR96, Naz86, NO60, New65a, O’M71, O’M75, O’M87b, O’M88b, Ode86, Olk78, Olm77, Pao88, PR86, Par64, Par86, PH76, Pav94, PSSW06, PWG18, Per76, PW15, Pit91, Poh63, Pol73, Pol69, Por88, Pow70, Pra85, Pro84c, RAB02, Ral80, Ram78, Ran96b, Rat92a, Rat92b, Rau95, Red80, Rhe60, Rhe85, Rhe88, Riv60a, Riz94, Rog96, RSTK97, Roo10, Ros68a, RM95, SCS10, Sag86, Sai91a, Sai91b, Sal63c, Sav18, Sch76b, Sch85a, Sch80b, Sco94, Ser86, Set99, Sha96, SX96, Sha97a, Sha89b, SES75, Sim89, Sim93b, SS05, Sim16a, Sim82, Sle79a, SJG96, SFS87, SBGP98, Sny97b, Sop60, SM94, Sta90, Ste81a, Ste93a, Ste98, ST91b]. Methods [Sup77, Swa77, TK97, Tap76, Tap84b, Tap84a, Tap86, Tho87, Tho76b, Tin76, Tot91, Tra66, Tre78, TL95, Tul90, Tur78, Van88, Var76, Var79, Vas97, Vel21, VC06, Was94b, Wat86c, WR96a, Wei91, Wei69, Wen68, Wer63, Wet75, Whi92a, Whi92b, Wib87, WR96b, Wil66, Wil81, Wol69, Wol71, Wou62, Wou68a, Xu92, Xu96a, Xu96b, XZ98, XC11, Xu21, Yao86, aYT96, Ypm83, Zar68, ZD95, Zha97, ZM22, Zua05, Zwe16, Zwe84b, Zwe84a, vL86b, EV91, Hen79, Par10, SFS88, Sny97a, Spa62b, ST91a, Sül97, Ant82, Bai74b, Bai76, Ber95, Blu94a, Boy90a, Cam91b, Car93, Chu82, Col76a, Coz87, Cra89, Ede90a, Enr89, Fal70b, Fal93a, Gru74, Han95, Har66a, Har93, Him78, Jac76, JW83, Jun65, Kap85a, Kel84d, Kin94c, Lev63b, Mal90, McC95, Mee96, MT61, Moo69, O’M96, Pas95]. Methods [Por80, Por83, Ran96a, Shi69, Sik92, Sim93a, Slo93, Sta88, Syn83, Tha96, Tul90, Wil64c, Par10, Ans64c, Baz73, Cam93, Dia64, Dri65, Fro86, Gil90a, Han81b, Hau86, Kaz86, Lem88, Str76b, Str86c, XL94]. Métivier [Dup93]. Metric [Gao18, Kra24, Ral62a, TFL23]. Metrics [Dur92b]. Metropolis [Jun64, BL84]. Metzen [Eas91, Jes93b]. Meunier [Coo91]. Meyer [Bai74a, Ber81, Cro73, Dre89, Eas93, Los81, SM79, WJ94, Wic94a, Gar94a, Wic94b]. MGD [She09]. Miara [Nas90]. Micchelli [Qua95, Sch78c]. Micha [Bla88a, Fla97, Haj96]. Michael [Bei81, Boe88a, Bre74, But78, Che75, Cum92a, Daf89, Dav91a, Fra88, Goo90a, Guc87, Hag85b, Har83, Har82, Hig01b, Hig02c, Kap70b, Kni86, Mei89, Mor75b, Mul66b, Nel74, O’B95, Ros80a, Sch76c, Ste94a, Stu90, Ull88, Wig90, Cum92b, Goo90b, Ste94b]. Michailov [Kra66]. Michel [BI67, Cam91b, Cam91c, Fal93a, Fal93b, Gra93a, Ham87, Har64b, Laz91a, Laz91b, May84c, May84d, Raz87, Sha96, Sim93a, Sim93b, Gra93b]. Michele [Aur18]. Michelotti [Dum96]. Michio [Gra94a, Gra94b]. Mickens [Aga95b, Mea88a, Mey83]. Miclet [Pao88]. Microbial [Cus96, KD10, SWC96]. Microcomputer [Ral93b, Ral93a]. Microeconomic [Keh89]. Microelectronics [KMC96, Cer96]. Microprocessors [DKW^{+}09]. Microscopic [LKGK07]. Microscopie [Mar85a]. Microstructure [Wil91b, Wil91a].

Mieczyslaw [Pic87a]. Miele [Edw63]. Migration [Tre78]. Mihalas [Cha87a]. Mijalkovic [Xu96b]. Mikhail [Che18, Zha97]. Mikhailov [Haa93, Lev96]. Mikhlin [Nas65, Wou68a]. Miklós [Fre95]. Mikusiński [Syn93a, Syn93b]. Milanović [LG89]. Milbrodt [Jow86]. Mildly [CG76]. Miler [Law66]. Miles [Kan62b, OG90, San82]. Milestones [CGO07]. Miller [Ber92c, But61, GC79, Gur66, Joh83d, Joh83b, Kel60a, McM61c, Ric74]. Milner [Sil65]. Milos [Ott96, Rob92a, Rob92b]. Milton [Col77, Gup82, McG93, Gra18]. Min [Pon65, Utz79, Saa68]. Min-Max [Pon65, Utz79]. Minasian [Kar87]. Mind [Sen94]. Mine [Fox71]. Ming [BKS96]. Mingarelli [Hin89]. Mingle [Win74]. Minima [Ewi77, Han89]. Minimal [BJK11, HK92a, HK92b, Koc92, Küç92, Pla15, Pri64, SB98b, Zel77, Ham88, Lan81]. Minimal-Envy [BJK11]. Minimalist [Sco92]. Minimax [CC75, Har66b, Sin93c, Tan93b, Tan93a, Sin93b]. Minimization [CD20, CBS76, Gol81, GdM95, Jag71, Lew83a, Man70, Moo74c, Mou67, Nas77b, PR86, Pom67, RFP10, Rin74, Sch76c, Sch70b, Sch71e, Sha94, Sha95b, Spa62a, VH78, Wil81]. Minimize [Sto60]. Minimizing [GBS08, GH22, HS88, Mal78]. Minimum [Bad60a, But61, Cah77, Chu78, Ebe75, Fre75, GL89, Gib87b, Gib87a, HJ85, Jag93, Jag94, KH85, Kan85, Kla84e, Kla91, KF06, Lar86, Lee86c, Los83a, McM61b, McM61c, MW76b, Mil59, Mil60, Mur90, Nas65, New59, New62a, New64b, Ove66, OR67, Rai59, RFP10, Red90, Sek82, Ste73a, Vei86, VK84, ZN62, FK06, Sha98, Bal93]. Minimum-Loss [Rai59]. Minimum-Rank [RFP10]. Minneapolis [Pip86]. Minors [Sch73f, Sch74]. Minoux [Cha86]. Minsky [Yau69]. Minus [Isr92b, Isr92a, SS91, SG73]. Minus-One [SG73]. Miranker [Don85, Ral83a, Ral83b]. Mirchandani [Bur81b]. Miron [Joh87]. Mirsky [Bru73]. Mishev [Sei93]. Mishra [Win95]. Missile [Bor62, Rei78b]. Missiles [AV65, HW77]. Missing [Str89b]. Mitchell [She63b]. Mitigate [OO17]. Mitrinovic [Aga94]. Mitsuo [Mei77]. Mitter [Rus95]. Mix [Kel95]. Mixed [BJ76, Car70c, DGK^{+}12, DGP^{+}16, Fal93b, GJ19, JLM^{+}17, Joh75, Kel22, LA75, MYDF08, May90, MNC71, Mey79, Pro65c, Red76a, Sim93b, Vie15, Boe98a, Fal93a, Sim93a]. Mixed-Hybrid [Red76a]. Mixed-Mode [DGK^{+}12, DGP^{+}16]. Mixing [BDX04, FGTW16, SBXD06, Way82]. Mixture [CM89, RW84, WP78]. Mixtures [Gud77, Ken93, Gud78b]. Mizan [Wim91a, Wim91b]. Mladen [Chu96]. mm [Wil76b]. Mobile [Sco92]. Mode [All75, All76, DGK^{+}12, DGP^{+}16, Goo72, Kel88b, Kel88c, Kut06, Les76, Luc18, Lud87]. Mode-Locked [Kut06]. Model [Abd90, Ath76, BM03, BGW15, Boe91, BS09b, CZ92, Cop07, DeS88, DC76, DGP92, Epp86, For76, FL81, GN22, GAHK05, GP02, Gla93b, GZ11, Gre89, Gro93c, Gut88, Han75a, HS03, Hop92, Ins78, JRA22, Kow88, KD78, LMDLS23, LRJ78, MPRW02, Mic88, MBR98b, MY97, New76a, NW97, Par92a, Pip81, PH13, Rou97b, SS65, SG81, SBR08, Sed02, SH94, TLN95, TeB97b, Tho77b, Tud90, Ver90, WHZG24, WiOH10, WHMN16, WMZ94, YAKM20, YFS76, Zit76, Lew83c, MBR98a, MW97, Rou97a, TeB97a, Ano93a, Gut82, Lew91b, VBdD91, Mla94, Nas95]. Model-Building [Gre89]. Model-Free [TLN95, Nas95]. Modeling [ABJK10, AW19, BG18, BD11, BDE09, Bru86, BGL01, CZ96b, CGT^{+}97a, CL87, CN15, DKW^{+}09, Dre89, DHL17, FP04, GL05, Gou16, Gue83a, Hig08, Jon92a, Jon92b, JC12, Ken91a, Ken91b, Koç94, KP06, LLSN96, Lim23, Luc18, MYDF08, MQ95, Mil89b, MG95, Neu97a, O’B11, Ode82, PBL98, PMG95, Pit89, PM87, Raz92a, Sey93a, Sey93b, Sin89, Sto11, SH96, Tud92, WJV79, Yan91, Zem96, CGT^{+}97b, Sle83b, Art85a, Bru83, Coo89, Hil96, Qua95, Ram90a, Raz92b, Gla86a, Nic96]. Modelling [BG95a, BvdGH95, Cum91, Cum92b, DEPC95, FTM95, Far93a, Far93b, Ham79a, Hav95, Hop83, KCD^{+}97, Lus87, Mar95, MT96, PMK96, PCH98, Sam79, Seg79, She91, AH98, Log97, Ant91, ABG96, But79, Cum92a, Dre89, Fre83, Gue83d, Moi89, Wal88, Cha95b]. Models [Aga95b, ABL^{+}13, ASPL03, ABG^{+}95, Ang78, AEH^{+}21, Art83b, Art83a, Art85b, BvdD78, BD11, BT13, BF16, Bro90, BK87, CC96, CSW^{+}08, CHO92, Cha00, Chi83, CT75, Cla78a, Cla79a, Cof83, CT76b, Coz85, Coz87, DL95, Dri81, Dur99, EK91a, EK91b, Elv75, EFA76, FTM95, FLNU18, Fre79, Fre86a, Gar89, GGK^{+}22, GS23b, GS04, Gue82, GAM93, Han83a, Hey93, HFKG06, HPS07, HRM96, Hul76, Kar76, Ken93, KS94, KM12, Kla84b, Len95, MNZ93, MQ18, McK95, MHC01, Mea98, Mee96, Mer81, MXPW11, MPW23, MA95, MA04, MG23, NGTB15, O’M10f, Pan93a, Pan93b, Pit81, PI88, PR70, Pri93, Ric88, RCM07, Ruk93a, Sán83b, San83a, Sau79, Sch91a, Sch91b, Seg88, Sei85, Set77, Shu82, Ste01, SAT22, Str86b, Tay86, TD20, UA07, VCKJ16]. Models [VLPR20, WZCX12, Win62, Wit75, XC11, ZWY14, Beu89, Cha94d, Dix91, Dur92b, Gra83a, Gue84, Han83b, Hin87, Hun84, Jin91, Kaz80b, Kla84i, Lah94, Mac76a, Mar95, McC92a, ML85, Pin77, Ruk93b, Ser94, Tho75, Thr79, Wat86b, McK81, Wal75b]. Models-Rarity [Pit81]. Modern [Boa85, BBKK22, Buc91b, DKW^{+}09, Dra61a, GW12, Gol66b, Grc11, Grc18, Gue79, Ham71, Har74, Har66a, Hei95, KHS96, Kaz75, Kel59, KLT03, KLT04, Mac77, Mor97, Pon83, SZ96, Spo88, Swa96, Wil71, You73, Zha96, Buc91a, Fle73, McG65a, McG61, Sal63c, Sle79a, Wil62a].

Modes [BZ05, EFMM13, Phi85, Sez90]. Modified [Gla78a, GN78, Gol73b, Mel12, Mor66, Nås77a]. Modular [LS63, She67]. Modularity [EW19]. Modularization [Hil75]. Modulationally [BLMT18]. Moduli [Sle87, Tot91]. Modulus [Gla88d, RJ89, Ren89]. Moffatt [Chi79]. Mohamed [GB89, Gro87, Keh89]. Mohammed [CK86]. Mohanty [Tak83]. Mohler [Bro76a]. Mokhtar [Fro80]. Molding [CGT^{+}97a, CGT^{+}97b]. Molecular [But78, Chi83, Gla86a, Mil89b, Neu97a, RMT^{+}03, Sco88, WHMN16]. Moler [AF69, Gup91a, Gup91b, Joh68, Par81a]. Mollification [Bec94]. Mollifiers [Rei71b]. Mollo [Moi89]. Mollo-Christensen [Moi89]. Molluzzo [Har83]. Moment [Ant72, Bus86, CH66, GR85, Mor85b, Sim71a]. Moments [Deh89, HS70, Jag76b, Pri76, Olk77, Ath94]. Monastyrsky [Wig90]. Monge [Pin84b, Urb93a, CGP21, Pin84a, Urb93b]. Monique [Pao90a, Pao90b]. Monitoring [BCG77, GG95, Gil95]. Monographs [Ham65b]. Monoids [Jer76]. Monotone [BN76, BT20, GYS95, Hes83, Man68, BI70b, Gib77, Sha95c]. Monotonic [CM83, EF63, EF65, Kam78]. Monotonicity [BI70c, Her88, Jag76b, JC88b, KB68, LJW73, Meh74a, Mon70, Olk77, Rie87, Usm72, Jor88b]. Monotropic [Ber85f]. Monroe [Ber76a]. Monte [DKST19, Ham77, Jun70b, Ric88, Slo93, DM07, DMP22, Fos68, Hal70, LW17, NA06, Pug66a, Spa62b, YKS78, Zar68, Jun65]. Moody [LS66]. Moon [Are89, Bir64, GD88, CF80]. Moore [Hag86, Han67, Han81b, Hul71, Kod96, Mic95b, Rob74, Rob76, Tal81a, GM76a, Zie73, Zlo70]. Moorhouse [Cas87b]. Moors [Sch76c]. Morante [Jer81]. Mordecai [Bor90, Kan78]. Moreau [DMP22]. Morel [Sha96]. Moretti [Whi66a]. Morgan [Kel86c, Pra85]. Mori [Mal88a]. Morineau [Ber85b]. Moritz [Van94c, Van94b]. Morokoff [Cer96]. Morosanu [Pic90]. Morozov [Fro86]. Morozow [Eng94]. Morphological [HS96, Ser96, Sta90]. Morphology [Hei95]. Morra [Goo65a]. Morris [Gus76, Mar66, Sco84b, Sco84a, WGM93]. Morrison [Sey93b, Sey93a]. Morse [Bar59, Ber94, Blu94b, Rei77a]. Mortality [MV15]. Mortensen [Woo92]. Morton [Ant83b, Bou65b, Ram90a, Wen68]. Moscardini [Dre89]. Moscow [Vas94]. Moser [Gol69a, Saa73, Seb70, Zas65]. Moshe [JB88, Ton95]. Mosteller [Par92a]. Motility [Mah94]. Motion [Ale84, BRHC20, Dre75, Eri66, HJ00, HPF85, HP85, HLH03, Jen75, Jor95, Kla88b, Kla94b, Kla95b, Kni84a, Kni84b, Kni86, Kni94b, Kni94a, Lax75, MS89, Mey95c, Pin72b, Pit93b, Sch75b, Spi80, War70, War71, Jor96, Col73, Fin76a, JGM95, Mar81, Mat89a, Mey84, Pit93a, RJ96, Ros85, Wou64a]. Motions [Bre68, FF95, Fif79, Leo78, MV68, Mos66a, Fre95]. Motivating [GK12]. Motivation [DM77]. Motivational [Ale95, Ale98]. Mountain [Bis15]. Movement [Pap63b]. Moves [Gia63, LL76a]. Moving [Ano91a, BH96, Dam84b, Dam84a, Ell88, Kla74a, Kla89, KMH^{+}90, Mey90, Mey91b, Mos66b, ZD95, Del95, Gol69a, Hua96, Seb70]. Moving-Grid [ZD95, Del95]. MR [Lax89, Rob94b, Win96b, AR80, Bel10, Coh93b, Dea90b, Mor86c, PC90b, SW80]. MR2164298 [Par10]. MR2383888 [Par10]. MR2404036 [Ver09]. MR3380576 [Tre15]. MREIT [SW11]. Mu [Dur93b]. Muhammad [Jód96]. Muller [Sch82b]. Muller-Pfeiffer [Sch82b]. Mullin [Ass77, Dic67, Zie78]. Mullineux [Fra66b, LF66, Los73]. Multi [Bal85, Kel88b, LLSN96, Nic96, SK76b, Can86b, Man88a, McC91b]. Multi-Armed [McC91b]. Multi-Degree-of-Freedom [Kel88b]. Multi-Derivative [SK76b]. Multi-Grid [Man88a]. Multi-Link [LLSN96, Nic96]. Multi-Variable [Bal85, Can86b]. Multibody [Ant90]. Multicanonical [DM07]. Multichannel [Cle80]. Multiclass [RMT^{+}03]. Multicode [NT78]. Multicomponent [Tre78]. Multicriterion [Bas88]. Multicultural [Her93a, Her93b]. Multidegree [Kel88c]. Multidimensional [Ame78, BM93, Flu78, Gib95, HF78, KG95, Mey76a, Ram78, Rou91, San72, Wei90a, Wit90, Gur66]. Multidisciplinary [CAiJS23]. Multifidelity [PWG18]. Multifrontal [Liu92b]. Multigrid [BS09a, BFM^{+}05, BH90, JMX96, Owh17, SX96, Sul93b, Bai87, Kin94c, Sul93a, Ta’90a, Xu96a, Xu96b]. Multigroup [Spa62b]. Multilagged [BK88]. Multilevel [DKST19, KK99, McC95, RM95, SBGP98, Vas97, WiOH10, EV91, Wei91]. Multiload [BTKNZ00]. Multinomial [Ano89d, Bro66a, Moh66, RC87, RR89, RAK88, SKGM67]. Multiobjective [Cra86, KD78]. Multiparameter [Mar73]. Multiphase [BL09, Coo91]. Multiple [Ano91b, AN83, Bad60b, BKR75, Ber94, Blu94b, BD93, BdD93, Bre77, CC96, Car76a, DGK^{+}12, Dun64, Edw00, Erd92b, Erd92a, Esl96, Flu78, Gen96, Gro91c, Hab70, Hel75, Hen92, Jev91, Jev92a, Jev93, Jev95a, Kar87, KWP95, Meh74d, MJ61, Mil76, RAB02, Rub78, Sho66a, Siv69, SJG96, Ste73b, Str65a, Wol77, Ypm83, Pro95, Ask78, Spu89, Van76]. Multiple-Access [Jev92a, Jev93, Jev95a]. Multiple-Scattering [Flu78]. Multiplication [GF76, MKR76, McG62a, Pan84, VB05, Sch65c, Col94b]. Multiplicity [HP76, Jor77c, Lir76, wBJ77]. Multiplier [Fer97b, Lag97, Zlo85, Fer97a]. Multipliers [Pon83, Roc93]. Multiply [Cop86, ET99]. Multipole [SP01]. Multiresolution [JS94, Owh17]. Multiscale [WZCX12]. Multispecies [Ole94]. Multistage [Kla87, Kla88a, KDJM88]. Multistate [FG19]. Multistep [Bar78, LP75, SK76b]. Multitype [Goo72]. Multivalued [Bre94b, Bre94a]. Multivariable [Bou21, Can86a, Rei71d, Smi75, Zak85, Bla86b]. Multivariate [Car87, CBG75, DK97, FTM95, FSY65, Gra93a, Gra93b, Lin92, Nei19, Olk59, PW08, Ton88, Ber85b, Mar95, McC79, Rei79]. Muni [Ber92b]. Municipal [Tuc73]. Munk [Col96]. Munson [Kla63a]. Murad [Pel97]. Murakami [Cor92a]. Murase [Mah94]. Murat [Lan86]. Murio [Bec94]. Murlan [Far69]. Murman [Por86a]. Murota [Chu89]. Murotsu [Spi87]. Murray [Bel90b, Blu65, HBO66, Jor77b, Jor86a, Kow83, KZ93, Mey90, Mey91b, Ole68, Pay68, Pic87b, RJ96, Seg80, Zuo96, Dav94b]. Murthy [She91]. Murty [Bla85a, Bla89b, Wal79]. Music [RPP88, Bre89]. Musical [AW02]. Muskhelishvili [Gos64]. Mustafa [Bal93]. Mutual [KS77]. Mutualism [CN19]. Mutually [CJR18]. Myoung [Kel92a, Kel92b]. Myron [Zah80]. Myshkis [Dij90].

N [Amo78a, AF69, Bha93, BR67, Bir69a, Ble81, Boa65, BC96b, Buc69, Bue71, Cha81b, Chu81, Chu88b, Coh84b, Cor16, Cul66, Cul68, Dij90, Don90, Fle63a, Fle83, Fra66b, GP88, Gla94a, Gos64, Gue82, Gue83a, Gue83d, Gue88, Ham65b, Har70, Har93, Hau85, Hau88, Hey97, Hoc86, Hol89a, Hol95, Hor83a, Hor97a, Hor93, Jag72c, Jef94, wBJ77, Jor80c, Jun64, Kan66, Kaz75, Kem85, Köl87, Kor65, LF66, LM81, Los73, Mal88b, Mar83b, Mar83a, Mat80, Maw95, May91, McK95, Mey76c, Mey82a, Mil90b, Moo74a, Moo63, Mro91, Nas77b, Neu94, Not94, Ock91, Oco87, Olk59, Par88, Par90, Pro84a, Pro91, Ral62a, Ral62b, Rhe60, Ros73, Saa83, Sac83b, Sac83a, Sac74, Sel94, She87, She91, Sil78, Sim89, Smi82, Spi81a, Ste68b, SFH^{+}66, Sym92, Tak77, Tul90, Wat86a]. N [Wel88, Wen78, Whi83, Whi94, Wil79, Wil77c, WM85, Wim65, Wou64a, Yon95, Zie76, Mil90a]. N. [Coh84a, Mar83a, Sch77c]. Nachrichten [Hig62d]. Nachrichten-find [Hig62d]. Nadia [Odg86]. Nagaraja [Hor93]. Nagasawa [Aeb94b, Aeb94a]. Naidu [Man87]. Nail [Abl80]. Nakamizo [Aok70]. Nakayama [Cra86]. Name [JRA22]. Namik [Her67a]. Nancy [Jag83]. Nanospheres [YA18]. Narayan [Rei79]. Narayana [Jag80]. Narcowich [Bre74]. Narenda [Hun84]. Narendra [Asn75]. Narrow [HS14]. Nasell [Gla78a]. Nash [Gup91a, Shu90a, Wil81, Gup91b, Shu90b, Tij81]. Naslin [McG64]. Nathan [Gin85, Luc18]. National [Ano68l, Ano69l, Ano71n, Ano72o, Ano72m, Ano74j, Ano74k, Ano76o, Ano77p, Ano78q, Sal63c, Tod74, Tod75]. Natural [CD94, Cla78a, Giu18, Kap76, Kee88b, Kel88c, Ken86b, Koc92, Küç92, PW15, Sam76a, WPL98, Kee90]. Natural-Resource [Cla78a]. Nature [Dra61a, Ham71, LKGK07, Ste94b, Urq69, Goo84, Ste94a]. Naval [Man82]. Navier [Pon90a, Sul93a, Bea90b, CS90, FC99, PSSW06, Pon90b, Sul93b, TF96, TP78, Bea90a, Far96]. Navigation [GK15, Kab93, Nie84, RVBE21, Wei62, Wol88, Mar65]. Nayfeh [Col76a, Hol86a, How82, Ran96a]. Nazrul [Xan87]. NBA [Pen95]. Ne-Zheng [Hil96]. Neal [Bur94b, Bur94c]. Near [DGK16, Has86, IWMW21, MN65, Rum99, SR75, CL08]. Near-Disk [IWMW21]. Near-Optimal [DGK16]. Near-Zone [MN65]. Nearly [Ful76, Gri85, Mey89a]. Necessarily [BS92]. Necessary [JP89, Nas77b, ZZ18]. Necessity [TT94]. Nédélec [Say13]. Nee [Mey95a]. Need [Bro61, ZWY14]. Needed [CCR^{+}62, Leh82]. Negativity [Gil62b]. Negoita [Arb77]. Nehari [Thi63b]. Neil [Lag89, Orl93a, Rob91]. Nelder [Wat86b].

Nemhauser [Bla90]. Nemirovskii [Ye94]. Nemirovsky [Bla85b]. Neoclassical [Lew83c]. NEOS [CWW99]. Nering [Cot94]. Nerve [RK75]. Nerve-Conduction [RK75]. Nessel [Ask72]. Nesterov [Ye94, Wal23]. Net [OR21]. Nethercot [Mal85]. nets [Ani82, Per65]. Network [AMO91, AMOB95, BG18, BCH14, EW19, EH10, Fry64, GC69, GH13, Hu68, Jef90, Kar01, Leh60, LSS98, McM61b, Meg88, Mil69a, Mil59, Mil60, Ner70, Poh78, RL62, Ste67a, SAT22, Tu76, TC68, Bou66a, But61, McM61c, Bal95, Ber85f, Sle67]. Networked [VBVM22]. Networks [ADW23, BM93, Beu92, BGHS23, Bur81b, DK85, EH16, EERG20, FSV06, FG19, GGK^{+}22, Ham88, Hey89, Hop92, JHSB22, JMFB15, McA87, Mic92a, Mye75, New03, NRS19, PMG95, Rad90a, Rad90b, RPFM17, SAB19, Sle84a, Sou76, TKMP11, Whi83, BI68c, Jar76, Mac61, Mic92b, Min66, MG95, Sle84b, Spi81b, Tar81, Wol86]. Neugebauer [Ung85]. Neumaier [Hig02b]. Neumann [Dav91a, Lig76, AW76, BL05, Fri80, Grc11, Grz89, Man09, Seb75, Tod74]. Neural [Hop92, Jef90, Ker91a, Ker91b, Mic92a, Mic92b, Mil91a, Mil91b, MG95, PMG95, RBZ23, ZMPK24]. NeuralUQ [ZMPK24]. Neuringer [Wil69]. Neurobiology [Bel90a]. Neurons [ADW23]. Neurosciences [Gus90b, Gus90a]. Neustadt [Sag78]. Neutral [HS15, Sei93, Und78]. Neutron [Var67, VN76b]. Neutron-Transport [VN76b]. Neuts [Gra83a]. Nevanlinna [Han94b, Han94a]. Neveu [DD74, Eic67]. Newborn [Ber76a]. Newcomb [Sei86a]. Newell [Hey86]. Newman [Ask81, Boy70, BBLS92, BOS73c, Car71b, Erd77, FB78, FR63, Gou68, Gro93d, Kha71, Kin76, Kla84g, KF62, Mei83a, Sch65b, She63a, Wil65a, Wil69, Kla84f]. News [Ano60a, Ano60b, Ano60c, Ano60d, Ano61b, Ano61c, Ano61d, Ano61e, Ano62a, Ano62b, Ano62c, Ano62d, Ano63b, Ano63c, Ano63d, Ano63e, Ano64a, Ano64b, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano65d, Ano66c, Ano66d, Ano66e, Ano66f, Ano67a, Ano67b, Ano67c, Ano67d, Ano71l, Ano72k, Ano74i, Ano76n]. Newton [Ant92a, Kaz91a, Ant92b, DM77, DS79, DG14, EG98, FP96, GG80, Ger94, Kel22, MBOV17, Naz94, Nei19, PSSW06, PW79, RH76, SHL16, Tap76, TDS18, Tin76, Tod84, WPTZ01, Ypm83, Ypm95]. Newton-Like [Ypm83]. Neyman [Rei63, Wag69]. Ng [BPV97b]. Nguyen [Aga93, Ber91b]. Nicely [Ham74]. Niche [Giu18]. Nicholas [Bra94b, Flo69, Her92b, Bra94a, Her92a, Ste97b]. Nico [Olv97]. Nicola [Log97]. Nicolaenko [Jol96, Liu91]. Nicolis [Ber88a, Oth82]. Niederreiter [LeV77a, Slo93]. Nievergelt [Smi74]. Nijenhuis [Whi80]. Nijmeijer [Blo91]. Nikitin [Sha97b]. Nikolai [Vai96]. Nikolenko [Har70]. Nile [FS89, FSRJ90, RJ90]. Nilpotent [Her91]. Nils [CL72]. Nilson [Bir69a]. Nilsson [CL72]. Nima [Gri88]. Nine [Hig02a]. Nineteen [MV78b, MV03]. Nineteenth [Van94a]. Nineteenth-Century [Van94a]. Nino [Dim91a, Dim91b]. No [Gil65a, Mor69, Sal63c, AR80, Coh93b, Dea90b, For96, Gud78b, Mac98a, Mor86c, PC90b, Rob94b, SW80, Win96b]. Nobel [Daf89]. Noble [Kap68]. Node [AD76]. Nodes [DGP^{+}16, Iba73, Lam85, McC86, MKS86]. Nodet [Law18]. Noel [Whi76b]. Noether [Lil68, SC81]. Nohel [Cha67a, Her71]. Noise [Cho76b, FM18, Hig65, Mac91c, MBK08, Mac91a, Zit63]. Noise-Induced [FM18, MBK08]. Noises [MV68]. Noisy [NGTB15]. Nokolsky [Kla84j, Kla84c]. Nolnlinear [Lew86]. Non [Ame78, AC78, Cre97, DA79a, Ell85a, Fra65c, Gil62b, JMJ95, KL76, KSS80a, Kem92a, Kem92b, Kla95a, Lae76, LL76a, Lei76, LLR^{+}81, LRJ78, MR18, Maw93b, McG67a, MR78, Ost61a, Pel97, Pra68, Ros79, Ruo90b, Sha75, SA78, Sta92c, Wil64a, Wix76, YF78, Ter82, Bla66, Cra84, Cur64b, Fin76b, Hin89, Jur95, Lau74, Maw93a, Nit90, Ruo90a, Spo85, Sta92b, Ste87]. Non-Analytic [DA79a]. Non-Commuting [Lei76]. Non-Convex [Ell85a]. Non-Displacement [KL76]. Non-Equilibrium [Cur64b, Spo85]. Non-Gaussian [Pel97]. Non-Homogeneous [Lau74]. Non-Life [Ruo90b, Ruo90a]. Non-Linear [Ame78, AC78, Cre97, KSS80a, Lae76, LLR^{+}81, LRJ78, Maw93b, MR78, Pra68, Ros79, SA78, Sta92c, YF78, Bla66, Cra84, Fin76b, Maw93a, Nit90, Sta92b]. Non-Markovian [MR18]. Non-Negativity [Gil62b]. Non-Numerical [McG67a]. Non-Oscillation [Hin89]. Non-Paretian [LL76a]. Non-Random [Kem92a, Kem92b]. Non-Rectangular [Wix76]. Non-Standard [JMJ95, Jur95]. Non-Stationary [Fra65c]. Non-Stiff [Sha75]. Non-Symmetric [Kla95a]. Non-Uniform [Wil64a, Ste87]. Nonacademic [Ano75g, Col76b, HCH^{+}75]. Nonanalytic [DA79b]. Nonautonomous [JR08, MN76, OHBG00]. Noncentral [Kri67]. Noncomplete [Spi71, Jag72d]. Nonconvex [Ame86a, CD20, EHS17]. Noncooperative [Bin84, Tij81]. Nondifferentiable [Lem88, Pol87]. Nondiscrete [Woź87]. Nonequilibrium [Dur93b, Oth82, SK20, Ock91]. Nonexistence [Pro84c]. Nonharmonic [Boe95b, Gla94b, Gla95f]. Nonhomogeneous [AB73, CC78, Wim67]. Noniterative [Bru65]. Nonlinear [Abd84, Abl75, Abl77b, Abl84a, Aga95a, ASPL03, Ama76, Ama77b, Ame85, Ame95, And73, AHP^{+}95, Ant78, AR95, Ari91a, Ari91b, Aro78a, Ash89, Bal88a, Bal97, Bal88b, BK95, Bel90a, Ber84b, Ber84a, Ber85a, Ber91a, Ber84d, BH15, Bla89b, BBKV96, Bog75, Boj93, Bou67b, Bou68, BEG23, BKST15, Bus76a, Car72b, Cas82, Cas94a, Cas94b, CG76, Cha66, CM76, CT75, Chu94b, Chu94a, CC76, Cou97, CM91, DC76, Den76, DG14, Eas91, Eck82, Elc76, Far96, FGN13, FW82, Fic60, Fok84, FGW02, Fre82, GHC97, GTH23, Gef75, Geo71, Gli91b, Goo65b, Gre76, Gri85, Gri88, Han86b, Har96, HV75, Het72, Hil75, HU07, Hol87, Hor95a, Jac69, Jag64, Jag87c, JK88, Jag88b, JK72, Kau86, Kaz84a, Kea80, Kel90a, KK98, Kir75b, Kir76, KLA95c, Kor95b, Lax69]. Nonlinear [Lax75, LM90, LF73, LP75, Lef65, Lem85, LR80, LBD^{+}86, LM81, LA95, Los76c, Mac91b, Mac91d, MW76a, MH96, MB76, MM95, Maw91, MS72, MR00, Mey78, Mey83, Mey93b, MXPW11, MPW23, Miu76a, Moo78, MBK08, Mor74a, MC76, NC76, Nas77b, Nau86, NBR96, Naz80, Naz94, Noh62, O’M71, O’M87b, O’M88a, Ode89, Old62, OH76, Ono84, Ono85, Osb85, Pea81, PD76, Pic87a, Pin84a, Por72, Pow86, PMR76, Ral66b, Ral74, Ram88, Ran96a, Rau95, Raw70, Ren97b, Rhe70, RLT65, Riv65, Rob91, Rod81, RM75, Rog94b, Rog94a, Rou73, Rub77b, Rue72b, Rue73c, Rue74b, Rue75, RW80, SS65, Sac67, San65, Sau76, Sch75a, Sch80c, Sch81c, Sch85b, Sco84b, Sco84a, Sho87, Sle76, Spa62a, SBM22, Sta71, Sta72, SS95, Str64, Stu89, Stu90, TF96, Tew77, Ton89, Vai96]. Nonlinear [Van75, VS95, VC06, Wag69, Wak86, Wal63b, War92, WE76, Wei90b, Win90, WMZ94, Won68, YL75, ZC74, vV65, Kir75a, Abl77a, Abl84b, Ame88, Bar70, Ber78, Ber84e, Bjö85, Blo91, Blu91, Cha94a, Cha94b, Col92, Dav75c, Deu90, Fal68, Fle76, Fro80, Gib85, Gru93, Han79, Han81a, Her67b, Hes83, Kan78, Kar90, Kaz84b, Kei64, Kel84d, Lev62, Man70, McC71, Mor83, Noh67, O’M86, Par90, Pic90, Pin84b, Ral66a, Ral70, Rei75, Ren95, Sed88, Sha95a, Sus88, Via85, Wah92, Zlo77]. Nonlinear-Wave [Gef75, Lax75].

Nonlinearities [KM96, Maw96, Bry94]. Nonlinearity [HL63]. Nonlinearly [AR78, AR80]. Nonlocal [BT13, BCM10, DGLZ12, GS67, TD20, Noh71]. Nonmalleable [DDN03]. Nonmathematical [Ham73a]. Nonmonotonicity [KMH66]. Nonnegative [Ber73a, Far84, GTH23, GR87, Gla88a, GGR^{+}88, Gup75, Hue92b, Las07, Lew83b, Mar76b, Mar76a, Ros69, Shi17, Tho74, Ber86a, Car81, Dav91a]. Nonnegativity [Han64]. Nonobtuse [BKKŠ09]. Nonoscillation [Bur81a, Ros80c]. Nonoscillatory [Shu09]. Nonoverlapping [XZ98]. Nonparametric [Bic94, Lil68, Mor88b, Noe81, Pin92a, Pin92b, Sha97b, YS76, Eps59b, Goo82, Noe72, Noe82, Sam86]. Nonpoised [Sha72b]. Nonprobabilistic [Hin91]. Nonrenewable [CD94]. Nonserial [Lar74]. Nonsmooth [CMSZ24, CB09, Man85, Sch86b, dBC^{+}08]. Nonstandard [MA95, TL95, Yoh79, Aga95b, Loe95]. Nonstiff [Ale90, SWD76]. Nonsymmetric [Gro96c, San76, DG96]. Nonuniform [AG01, Wit90, GL04]. Nonuniqueness [BBK20]. Nonweilerz [Row86]. Nonzero [Gan98b, NN96, Gan98a]. Nora [Bei96]. Norbert [Coo67a, Dev94b, Gla95b]. Norkin [Dri75]. Norm [AG14, Cav74, Fre69, Pow75, PJ75, RFP10]. Norm-Preserving [AG14]. Normal [BCV98, CLWM95, Kel88b, LC65, MB64, Mey86a, Pat19, Sch61, She62, She64, SS78, Tij81, WP78, Wat22, Mey95a]. Normality [Ang92a]. Normalization [BC86, Gla87a, GF76, Meh74b]. Normalized [SBH^{+}20]. Normally [WP95, Pal95]. Normals [SG73, Ste73a]. Norman [Bou77, Dan66, Fis71b, Gre89, Kaz80b, Pro84b, Tho75]. Normed [Leh65, Ral62a, Sop69]. Norms [GHU22, RW64]. Norrie [Dur93a]. Nørsett [Ale90, Hig92]. Northover [Fel72]. Nosov [Rox92]. Nostrand [Bre64a]. Note [And98b, Ang97b, Ars97, Bac60, BI69, Bla97b, BCD98b, Bos87, Cli64, Dar97, Dax97b, De 84, Del97b, Dol93, Dub97b, Ell61, Eri66, Fer97b, For98b, For59, FSY65, Gan98b, Gan98a, GMP66, Gre66, Gup98b, Hou98b, Jos70, Kab93, KL97b, KHS63, Kwa98, Lee97b, Liz98b, Low61, Mat69b, Mel97, MBR98b, MY97, Mor64, Naa01, NJN98b, NW97, Pea65, Pet97b, Rei78a, Rei97b, Riv60b, Roo64, RR97b, Rou97b, San65, San87, Sie65, SS61, Smi97, Sny97b, SC70, Str64, TeB97b, VBU97a, Wal76a, Won64b, Won68, Wou66, Liz98a]. Notes [Ano75d, Ano78e, Ano78f, Kee90, Mac98b, O’M10e, Sal63c]. Notice [Jac69, Mac98a]. Notices [Ano60a, Ano60b, Ano60c, Ano60d, Ano61b, Ano61c, Ano61d, Ano61e, Ano62a, Ano62b, Ano62c, Ano62d, Ano63b, Ano63c, Ano63d, Ano63e, Ano64a, Ano64b, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano65d, Ano66c, Ano66d, Ano66e, Ano66f, Ano67a, Ano67b, Ano67c, Ano67d, Ano68k, Ano69j]. Nottrot [Hag85c]. Novel [MA04]. Novozhilov [Cla64b]. NP [Boo94, Har82]. NP-Completeness [Har82]. Nuances [Sch70a]. Nuclear [Hul61a, RFP10, Win74]. Nucleic [DBS^{+}07]. Null [FC99, Und78]. Nullity [Bro66b]. Number [Alp96, And90, Bra76b, BOS73c, CH61, Coh82, CP89, CPRG90, CHW12, Ebe75, Eng17, Hab83a, Hab83b, Hig87, HD62, Jev93, Jev95a, Jun70a, Kon62, Lag89, Lau82, Lau83, Lau84, LR00, Man63, MK78, McA79, Mod92, New72a, RGJ90, Slo93, SS78, Ste73a, TL02, Wan84, Wim88, New84, Ham77]. Numbers [Alh96, BS00, BD94, BB88, Coh84a, FHRO95, KM80, MK78, Mye75, NS77, O’B95, OR21, SA75, SK20, Sun76, Whi63a, Coh84b, Rud61, Rue81a]. Numerical [AH19, Ale95, Ale98, Amo78b, And78, Ans80, Arc75, Ast87, Bak66, BJ75, Bec94, BEKS17, Bir83, Bla64, Bos96, BR18, BRHC20, Boy90a, Boy90b, But89, Cam91b, Cam91c, Cha87c, Che67b, Che78, CM78b, Cia92a, Cia92b, CLH97, CH66, COSG19, Cod74, CP78, Con66, Cor78, Cor68, Cul94a, CS90, DO98, DL66, Dev90, ERHO95, Enr84, Enr89, Faz98, FGN13, FL18, Fra65c, FR11, FEY61, GG99, GVL97, Gat64, Gea81, Ghi76, GBC^{+}90, GP79, GRL98, Gol76a, Gor75, Got78, Gra72a, Grc11, GS23c, Gup89, Gup91b, Hab70, Hab83a, Hab83b, Han83a, HV75, Hau78, Hel78, Hig99, Hig01a, Hig02b, Hig05, HS71, Hul67, Ike72, Ike75, Isa89a, IL97, Jac69, Jen78, Joh83d, Joh83b, Joh91b, Joy71, Kam76, Kau88, Kel84d, KK98, Ker92, Kij93, Kin94a, Kin97, Kin82, KPGR95]. Numerical [KPSP96, LA75, LR00, LW17, LBL02, McA94a, McA94b, McG67a, MP72, MOSK97, Mir71, Moi89, Moo60, Moo69, MC76, MK97, Mou67, Nas90, NO60, New76b, Nie94b, PC61, Par78, Par69, Par67a, Pav94, Per76, Pol69, Por72, Por64, Pow70, Pre81, PA95, Rab68a, Ral65a, Ral65b, Ram78, Rhe88, Ric60, Ric63, RLT65, Ric76b, RSTK97, Roo10, Ros76a, Row79, Row86, RK78, SCS10, Sai91b, SSCS95, Say79, Sch76a, Sch76b, Sco94, Sei76, Sha89a, SG95, Sha97a, Sik92, Slo97, Sma90, SK78, Sny97b, Som78, Ste81a, Ste97b, Str76b, Str86c, SH94, Sun76, TeB97b, TeB97a, Tho59, Tod75, Tot91, Tre78, Tre84b, Tre84a, Tri78, TS90, Tul90, Van78, Var76, Var79, Var66, War76, Was73, Wat86c, Whi76b, Wil65b, Wil81, You73, vdV94, Sny97a, EMUP98, Hig00]. Numerical [Hig02c, Par10, Ame86b, And95, Arn89, Che91, Col87b, Cul94b, Dan71, Din72, Dun94, EMUP98, Far69, Fer79, Fri59, Gea80, GR95, Gib90, Gla95a, Gos83b, Gra67, Gre88, GB89, Gri88, Gup91a, Ham61, Ham73b, Ham73c, Han83b, Han86b, Her70, Isa88, Isa89a, Isa89b, Jac76, Joh91a, Kin94b, KZ93, Lam86, McA94c, Moo63, Ork95, Per68, Pfl92, Pit91, Por80, Por88, Pra85, Pra86, Pro96, Sai91a, Sal63d, Ske95, Ste68a, Ste81b, Str82, Tho87, Van88, Wil72, Wil73, Wil93, Yan94]. Numerically [Far93a, Far93b]. Numerics [Ale85, KP06, LH19, SK20]. Numérique [Ost61a, BI68b, BI68b]. Numérique-tome [BI68b]. Numerische [Gra66a, Che65]. Nürnberger [Jet91]. Nussenzveig [New74]. Nuttall [Bou86]. Nxn [Jes77a].

O [Abr94, Bau88, Ben95, BdD92, BW73, BS86, Coo66, Cuv90, Dev90, Dev92, Dix91, Dre89, Gla88a, GRYL89, Gru88, Gun88, Hil76, Kan68, Kir76, Kno81, Knu80, Kog66, LM81, Los76a, Mal93a, McC91a, New90, Pow80, Pro87a, Rei71c, Rei87, Rou73, Sei86a, Ser64, Sha83, Shi71, Sto81, Tay96, Tre65, Ung85, Van66, Wei80b, Wig90, Wil82b, Win74, Zas65, Dod73, MNMF63]. Oakes [Pin85b]. Oberguggenberger [Col94b]. Oberhettinger [Ask75b, Kaz67]. Object [Nei10]. Object-Oriented [Nei10]. Objective [Kar87, Pei83, Wan83]. Objectives [Lin67]. Objects [KK75]. Obligate [CN19]. Observation [Zua05]. Observational [RH96, Sch91a, Sch91b, Hei96]. Observations [Gut88]. Observed [Car93]. Observers [Ant86b]. Obstacles [Tay89]. Obtained [Jok66b]. Obtaining [Iba73, Str62b, WP78, Wim72]. Obtuse [HH00]. OC [Jag87a, RM86]. Occurrence [SS05]. Occurring [Gie63, Jos95, KMH66, Sie61, WA94, Gro98a]. Ocean [KCD^{+}97, LM79, MWWC96, Col96]. Oceanography [Ier96, RBC^{+}07, IR97]. Oceans [Nie84]. Ockendon [Dam84b, Dam84a]. Octic [Cas72a, Coo71]. Odd [Isr92b, Isr92a, Nie94a, SS91, Wat96]. Odd-Even [Nie94a]. ODE [Dar97, Mea98, MXPW11, MPW23, RBZ23, Sku04]. Odell [Mil89b]. Oden [Bal88b, Hau85, Wat86a, Wen78]. ODEs [Kel84d, Cor78, Enr78, FJT19, Wil24c]. Odlyzko [Grz83, McL79]. Odoni [MN83]. Oettli [Fal68]. Ofer [Din94a, Din94b]. Off [For59, Lon78, Kri76]. Officers [Ano72o]. Official [Ano89c]. Officials [Ano80e, Ano80f, Ano81b, Ano82g, Ano83e, Ano84e, Ano85a, Ano86a, Ano86b, Ano86c, Ano87c, Ano88a, Ano90b, Ano91c, Ano92c, Ano93e, Ano94a, Ano95e, Ano96f]. Offset [SS61]. Ogustoreli [Her67a]. Ohmic [Ano92a, JH92, Lew69, Lew71]. Oil [Geo76, KB76a]. Oksendal [Pin87b]. Olavi [Han94b, Han94a]. Old [Bra81]. Oldenbourg [Hig62a]. Olderog [Mit93, Orl93b]. Ole [Eis86b]. Oleg [Kaz86]. Olesiak [Tad93]. Olfactory [RVBE21]. Olga [Djo68, Gol93]. Oliger [Sül97]. Oligopoly [Shu80]. Oliver [Hey82]. Olivier [Str86a]. Olkin [Fan82, Gup82, Man61a]. Ollendorff [Pap62b]. Olmsted [Gol65b]. Olsder [Bin84]. Olver [Kum88, War92, Wim75]. Olvi [McC71]. O’Malley [Har76b, Was94b]. Omega [Net20]. One [Ask74, Ask76, AG14, BLT08, Bet78, Boe96b, Bol91b, Bol91d, BB83b, BK78, Cel78, CFG95, Coo67b, CEH12, Gla95d, Han76, HM78, Hig02c, Hoc80, Kin88, Kra67, LT13, Mac92, Mey76a, Mey77a, MM10, Mun61, Pro87b, RB23, SG73, Sit92, SS61, Thu01, VN76b, dMvSB95, Bru95, Kla63a]. One-Dimensional [AG14, Boe96b, BB83b, BK78, Cel78, Coo67b, Gla95d, HM78, Kra67, Mey76a, Mey77a, MM10, Mun61, RB23, Thu01, VN76b, dMvSB95, Kla63a, Bru95]. One-Sided [Ask74, Ask76]. One-Way [Kin88]. O’Neill [Mag86]. Online [CCG^{+}02, TKMP11]. Onset [SR75]. onto [Pea82]. Open [Boy08, HU07, MP20, Rus78, Ram88]. Operational [Kap65, Str77, DC86, Sop65]. Operations [Art83b, Art83a, Art85b, BI70a, Coh88, Coh93a, Dor73, Dor75, Lit63, MN83, Ng94, Sal62, Vet73, Coh93b, Rog60, Sop60]. Operator [Boe93, Boe94, Boe95a, Bue66, Dun68, Gok70, GF78, JMH95, KM96, Lan81, Maw96, MW66, Moo76, Moo78, Owh17, Pic87a, Raw70, Ren94c, Rhe70, Sch89a, Sch78a, Sch78b, Sch83, YH76, YF78, Ama82, Ame84, Hud95]. Operator-Limit [JMH95, Hud95]. Operators [Aar07, AT17, Ber78, BT20, Cha81a, CHT21, Con97, CT96, Dem74, EKS97, Fig93b, FR66, Gla91, GM85, HK92a, HK92b, HS96, KVM78, KN67, KKS82, Lie95, Man82, Mar87, Mey93b, Mul76, NS24, Ral62b, RL95, Ros68b, Sch84b, Sch84a, Smo91a, Smo91b, Ste61b, Tho69, Tre97, WJ75, Web85, Wil85a, Wil86a, Wou66, YT95, ZMPK24, Con98, Hel97, Tay97, Wil86a, Ans85, BR85, Bro76b, Coc84, Fig93a, Mar73, Sch82b, Ser96, Tay96, Ter92, Tri95, Uri92, Wil70a, Wil86a]. Opinion [JMFB15]. Opposed [FS12]. Optical [All75, All76, Les76, LBL02, Pim93, Bri64]. Optics [BC76, GKLU09, Kri76, Wil62b, Ivi65, Mar66, War92, Wil62a]. Optimal [Arm76, BW76, BH76a, Bec85, Bel63c, BH76b, BCH14, Bis78, Bro77, Bus70, BS09b, Cas87b, Cas87a, CD94, Che91, Cla78b, Cla79b, CCG^{+}02, Col86, Col89, CP18, Dat67, Dat69, Dat73, DGK16, Edw90, Ell80, Eub88, Fal70a, Fle69, Fle88, Fle89, FS78b, Fox67, Fra65a, Fra65b, Fra66a, Fra86a, Fre75, Fre14, Fri78b, Fri79a, FGTW16, Gan98b, GK15, GSYY22, GM76b, GM76c, Gol77, GBF78, Goo91b, Gro96d, Gue88, HSV95, Hau90, Hau78, Hau81, Hau89, HM01, Joh91a, Joh91b, JS68, Kla74b, Koo77, Kre78, KG98, Lam85, LCE76, LL75, Lew81a, Lew81b, Lew82a, Lew82b, Lew83c, LmYS96, LS08, McC86, MKS86, Mil83b, Mil83c, MB86, Mor82, Mor89c, Mor69, NN96, Ngu76, OSS^{+}13, Pac86, Pac88c, Pea82, Pol73, Pre95a, Rab68b, RV78, Rei71a, Rei78b, Rox90b, Sac83a, Sag77]. Optimal [Sag83, San87, SS16, Set77, Sno76, Spi86, Tan93b, Tan93a, TWW65, VN76a, Win75, Wol86, Yan97b, Yan97a, ZW93, ZW94, ZZ18, Zit76, Gan98a, Bau89, Ces75, Fle90, Gro96h, Hag85c, Her69, Mor88d, Pré95b, Rox76, Rox90a, Sac83b, Spi81a, Sto96, Str86a, Tad89]. Optimal-Aim [BH76a]. Optimal-Cost [Zit76]. Optimality [Art92a, Art92b, JP89, Mas87, MHC01, Mro91, Roc93, Eck82, Fin93]. Optimally [Tho76b]. Optimierung [Ham85]. Optimisation [Art88, Bal88a, Ham87, Cia92b, Nas90, Dan72, Ham87]. Optimization [AK90, AKP93b, Art91b, BBC11, Bla97b, Bom02, BS98, BS09a, BCN18, BGL01, CMSZ24, CM76, CTW21, Col87b, CKCH23, CB09, CWW99, Dat80, DKW^{+}09, DL95, Den82, Din72, Duf86, DHL17, DX98, Esl93, FR16, FT11, Fis88a, Fle64, FGW02, GMS05, Gof78, Gre76, Gue88, Hac75, HE91, Han83b, Han83a, Han86b, HMK96, Hil75, HU07, Hue91, IWMW21, Kal63, Kar76, Kel17, Ker91a, Ker91b, Kle72, KA76, KLT03, KLT04, Kor85, KZ93, KD78, Mas87, Mcc78, Nas77b, Naz94, Nel74, NBS12, PWG18, PR90, Pol87, Pow70, Pow86, Ros76b, RH76, SA75, Sau76, Sch81a, Ser84a, Sha85, SF97, Sym92, Tap76, Tap84b, Tap84a, Tap86, Tze91, WB09, Wal23, Wet74a, Wet93, WS13, Wil76a, Wol81, Wol82, Won83b, Won83a, Zlo85, vW90, Bla97a, Cia92a, AKP91a, AKP93a]. Optimization [Art85b, Ber85f, Bla85b, Bla89a, Bla90, Cra84, Cra86, Den82, Fia73, Ken74, Kow83, Lem88, Len95, Mro91, Nas77b, Tap84b, Zim90, Zlo83, Art91a, Bla82, Cul77, Don90, Ell85b, Fal70b, Fal73, Foo63, Kod96, Lai91, Man85, McG65b, Mor71, Par94, Rox84, Sag78, Ser84b, Wet75, vWHE91]. Optimized [GZ19]. Optimizing [Bel86, SFOM99, Tuc73, VS95, DGH^{+}18]. Optimum [Ber60, Bro59a, Bro60b, Eis62, Erd92a, GHW76, Hul67, Jev91, Jev92a, Jev92b, Kla87, Kla88a, KDJM88, Pet78, Pet97b, Pet97a, Quo75, SS65, Str75, Erd92b, Kal65, Not94]. Option [Hig02a]. Options [CD03, GF76, Kri97]. Optique [Mar85a]. O’Raifeartaigh [Jor89c]. Orbit [Are93]. Orbital [Kyn76]. Orbiting [Hey63]. Orbits [BRMR08, Dol93, Pal96, Vis01, WP96]. Order [Abd91, Aga95a, Amo65, And77, AEH^{+}21, BGL17, BGHS23, Bob66, BE78, Bru65, Bur78a, Bus76a, Car77, CD20, Cas75, CL75, Dar97, Den90, DA78, DG14, Dwa80, Elc76, FF78, FGN13, FW82, For92, FGD^{+}92, Giv78b, GC78, GAT75, GST01, GL98, GLO20, GS67, HK92a, HK92b, Har67b, Her65, Her91, Her78b, Hir75, Hol91a, Hol91b, HPS07, HW76, Hor93, Jen75, Jin91, Kat76, Kat78, KK98, Kla76a, Lew76, LK65, Lut66, Mac76b, MQ18, Mal68, Mal69, Man76, Mat69b, Mey75b, Mey78, MR78, NS24, Pro85, Rod81, RBC^{+}07, Rub77b, Sch77a, Sch72b, Sch72a, Sch73c, Sch73b, Seb75, SS62, SR98, Shu09, Smi88b, SLF66, TS90, Tur78, Won64b, Won68, Yak94, ZZ18, KA95, Smi87, Fin76b, Apl95, Har93, Hig92]. Order-Statistics [Mal68, Mal69]. Ordered [Ama76, Ama77b, Fis93, McM87]. Ordering [GL89, Gup82, Jaf84, OLHB02]. Orders [Alh96, Bet78, McM87, SST95, Ton95]. Ordinary [Ale90, Bar70, Cam93, CCCM06, Cha94a, Cha94b, Cha84, Con85, Cro85, DO98, Deu85, Enr89, EW75, Gat88, Gea74, Gea81, Gla95a, GS22, GL98, Gre62a, Gup89, Han79, Han87, Her71, Hul67, Hul71, Isa89b, Jac97, Kah66, Kap72, Kla62b, Kla97, Kra80b, Kra80a, Laz82, LP75, Mac76b, Mat69b, MG76, MMS98, Mie17, Na67, OL98, O’M10f, Ral65b, Sch77a, Sch82b, Sha75, SG79, SG95, Sle92, Smi88b, Tho59, TS90, War68, Was94b, Wil72, Wil73, SWD76, Ant63, Asn78, Bra65, Bra63, Bro59b, Cha67a, Dan93, Eas91, Gat88, Han75b, Kap85a, Kap77, Lev67a, Mey72, Sch60c]. O’Reilly [Ant86b, Art88, Beu87, Man88b]. Organic [Rou88]. Organization [Erm89, Giu18, Hes86, Hol89a, Hun85, Oth82]. Organizing [Rei64]. Orientation [Seg80]. Orientations [CT76a]. Oriented [Dur97, Geo71, Hug93b, LW13, Nei10, Hug93a, Whi94]. Orienting [Pip86]. Origami [KMS22]. Origin [Cha87b, JRA22, JC88a, Owe06, Gre97, Jor88a]. Origins [Grc11]. Orlicz [Rya63]. Orlin [Bal95]. O’Rourke [Fra89, Hof95]. Orr [Her87]. Ortega [Ral74]. Orthog [WS75a]. Orthogonal [Bac60, Ber90c, Cha70b, Coh75a, DN59, FL18, Gar00, Gon06, Hig03, KVM78, LS08, Meh74c, Rei78b, RFS03, SG76b, Ste59, Wah03, Wat91a, Wim81, Mil81b, Seg06, Wat91b]. Orthomodular [Ben70]. Orthonormalization [WS75a]. Ortner [Gla95b]. Osaki [Fox71]. Oscillation [BBKV96, EKZA96, Hei69, Vai96, Wal63b, Won68, Lau98, Sei93, Ari96, Col86, Hin89, Sei93]. Oscillations [Beb68, CL08, Cro77, DGK^{+}12, DGP^{+}16, DL76, Fle65, FS76, Fri82, Kaz84b, Kaz84a, LM90, ML98, Nit90, Noh62, O’M88b, Sed02, Str64, Mey83, Noh71]. Oscillator [BBK20, CM91, Kyn76]. Oscillators [WE19]. Oscillatory [Kev87]. Osculatory [Str62a, Str62c]. Oser [DV76]. O’Shea [Som93]. Oshima [Get96]. Oskar [Rou88]. Ossicini [Ste72]. Oster [Wag92]. østerby [Eis86b]. Ostrowski [Hou67, Ort74]. Otani [Keh89]. Other [And90, Ano62e, Ano63f, Ano63g, Ano63h, Ano63i, Ano64e, Ano64f, Ano64g, Ano64h, Ano65e, Ano65f, Ano65g, Ano65h, Ano66g, Ano66h, Ano66i, Ano66j, Ano67e, Ano67f, Ano67g, Ano67h, Ano68g, Ano68h, Ano68i, Ano68j, Ano69f, Ano69g, Ano69h, Ano69i, Ano70f, Ano70g, Ano70h, Ano70i, Ano71h, Ano71i, Ano71j, Ano71k, Ano72g, Ano72h, Ano72i, Ano72j, Ano73g, Ano73h, Ano73i, Ano73j, Ano74e, Ano74f, Ano74g, Ano74h, Ano75k, Ano75l, Ano75m, Ano76j, Ano76k, Ano76l, Ano76m, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano78l, Ano78m, Ano78n, Ano78o, Ano79i, Ano79j, Ano79k, Ano79l, Ano80l, Ano80m, Ano80n, Ano80o, Ano81g, Ano81j, Ano81h, Ano81i, Ano82b, Ano82c, Ano82d, Ano82e, Ano82m, Ano82n]. Other [Ano82o, Ano82p, Ano83a, Ano83b, Ano83c, Ano83d, Ano83j, Ano83k, Ano83l, Ano83m, Ano84a, Ano84b, Ano84c, Ano84d, Ano84l, Ano84m, Ano84n, Ano84o, Ano85g, Ano85h, Ano85i, Ano85j, Ano86i, Ano86j, Ano86h, Ano86k, Ano87h, Ano87i, Ano87j, Ano87k, Ano88f, Ano88g, Ano88h, Ano88i, Ano89i, Ano89j, Ano89k, Ano89l, Ano90g, Ano90h, Ano90i, Ano90j, Boo94, CG76, GLY^{+}18, Gra18, Wol86, Ask80]. Otomar [Rox77]. Otto [Bus95, Gas85b, Goo91a, Pon90a, Sül97, Pon90b]. Outliers [Sin79b]. Output [Hor78, WHZG24]. Outstanding [Cas80]. Overdamped [WS75b]. Overdetermined [CBS76, LL76b, Riv63]. Overflow [LS63]. Overhauser [NPW95]. Overlap [Kla64c, UW13, HBO66, Ole68]. Overlapping [All92a, All92b, Blo94, Car91, Los95]. Oversampled [Hue92a]. Overton [Hig02c]. Overview [BS92, Fat76, JS94, Kru83, Lim23, Ode82]. Owe [Han95]. Owen [Art92a, Art92b, Kes85, Ros78, Sau62, Shu85a]. Own [Kel17, Pon83]. Oxford [Bor06, Bor06]. Oxygen [Epp86]. Ozorio [Stu90]. Özsoyoglu [Jag80].

P [Ahl93, AK90, Ale90, Als85, Ame86a, Ame86b, Ame88, Amo78a, Ant87, Are86, Art83b, Art83a, Art85b, Bak91b, Bak92, Bar83b, Bar83a, Bau88, Beb79, Bel63a, Ber85a, Beu89, BR62, Blo89, Blo76, Blo77, BG95b, BGL^{+}67, BMMO76, Bre89, BS80, Caf94, Car65b, Cas95, Cas94a, Che91, Che92a, Che89, Chi95, Cia92b, Col96, Coo91, De 74, DeM65, Del95, Den78, Deu81c, Din72, Dra62, Dup93, Dur87, Dwa85, Ell80, Far94, FS73, Fet80, Fle85, Fra88, FZ63, Gar91, Gas76b, Gib73, Gib87b, Gol73a, Goo86, Gre62a, Gre89, Gru93, Gua92, Gue72, Guh68, Gut82, Ham90, Han79, Har96, Her83a, Her84a, Hig92, Hol89a, Isa88, Isa89a, Ism78, Jag87a, Joh83d, Joh88, Jon92c, Jor72, Jor73, Jor79, Jor83a, Jor89a, Kap65, Kar63, KLG69, Kin76, Kri85]. P [Kri82, Lar78, Lev88b, Lev96, Lew86, Liu91, Los85, Lou92, Lud82, Man70, Mar85c, Mas95, McC86, McG64, McK95, Mil82a, Mla94, MG95, New87, NW81, Nic96, Old62, Pap63b, Par80, Pat87, Pro83b, Pro83a, Pro85, Pro96, Ral93a, Ral93b, Rau95, RSS65, Roy91, Rue72c, Rus90, Sac74, Sal63d, Sam79, Sch86b, Sch93d, Sch81c, Sco85, Sco90a, Sco90b, Sco94, Sei93, Ser94, Set62a, Set62b, Sha95c, She91, Shi93, Sik92, Ske95, Sno76, Sop69, Spi81b, Ste67b, Ste71, Ste90a, Str62c, Str62b, Ta’90a, Tut94, Van96, Van93, Via85, Wan91, War92, Wat86b, Whi84b, Whi84a, Whi88, Wil78b, Wim96, WGM93, Yan92, Yan94, You72b, Zlo85, vWHE91, Ano87b, BS94a, Boo94, Cas94b]. P. [Han83a, Hel97, Kaz84a, May84b, Pea65, Ser84a]. P.-L [Fle85]. PA [Hig00, Hig02c]. Pablo [Aiz94a, Aiz94b]. Pacemaker [Bel86]. Package [Fat76, Hil75, Mil89b, Mit94, TTM05]. Packaging [BM95]. Packing [CCG^{+}02, Kv93, SA75, UW13, Cox65]. Packings [HC16]. Paddon [Bai87, Kow86]. Padé [All87, Ban88b, Ban88a, FH76, GGT13, Gra72a, McC93a, McC93b, WM63, GM85, Wim81]. Padé-Type [Wim81]. Padulo [Mor75b]. Page [She91]. PageRank [Gle15]. Pages [Raz87]. Pai [Sle83a]. Pair [AG67, And74a, Ben64, Bre64b, Bro65, Car66b, Car73b, Kla78b, KG79, MR65b]. Pairs [LR05, Wes65, Wil21]. Pairwise [CS16, Gri77]. Pakes [Ken93]. Pal [Kaz67, Kem92a, Kem92b]. Palamin [Bre89]. Palencia [Ede90a, Lan86, Ede90b]. Palindromic [SR70]. Palis [Kra96a]. Palls [Spi87]. Pallu [Her69]. palm [Win96b, Win96a, PMK96, Win93]. Palmer [God87, Kaz92, Moo74b]. Pan [Cop86, Ebe96]. Panagiotopoulos [Ame86a, Shi93]. Pandey [Hey97]. Pang [Mor93]. Panjer [Nas87a]. Pankaj [Haj96]. Pantano [Rau95]. Panton [Ost61b]. Papadimitriou [Won83b, Won83a]. Papanicolaou [Lan86]. Paper [Jef94, JRZ97, Sch62b, Rob94b]. Papers [Bjö61, Ips10a, Ips10b, Ips10c, Ips10d, Ips11a, Ips11b, Ips11c, Ips11d, Jon92c, Sus79]. Papoulis [Kul65, Pol65, Por86b]. Parabolic [Aro78a, DD04, DJLQ21, EC76, GS67, Kla88b, Kus76, Lev96, Lie97, Mat89a, MS89, Mey76a, Mey77a, Nie89b, Ode86, Per65, SGK^{+}96, Sig67]. Parachute [Mea98]. Paradigm [BH03]. Paradigms [OLHB02]. Paradox [Ang97b, Ste81c, Ang97a]. Paradoxes [Saa95]. Parallax [Ste98]. Parallel [BH03, FR16, GPS90, HNP91, Hel78, Jac97, Kan68, KK99, Kow86, LDG95, MKR76, MB76, Mel64, OV85, SK76a, SBGP98, VB05, WiOH10, Gal95, vdV94]. Parallelepipeds [DKS96, Tor62]. Parallelism [Mir71]. Parameter [AABT15, BV96, Bob66, BPJ95, Esl96, Fle69, Gou16, HKS75, Jef95, Lud91, Rei75, Sac83b, Sac83a, Sle92, Ste99, WE76, Wit75, Dix91]. Parameter-Estimation [WE76]. Parameters [ABL^{+}13, AB69, BM98, DB70, Goo65b, Har76c, NS75, Udw89, WP78, Hin89, JF71]. Parametric [BGW15, Cot73, Cra84, Mai72, PE96, Yan90, Boe98a, Eat96, Wet93]. Parametrische [Gas70]. Parametrization [Stu64]. Parametrized [TD20]. Pardalos [AK90]. Paretian [LL76a]. Parisi [Abr94]. Park [Dav75a]. Parker [Dev90, Wil60b, Zim90]. Parking [KNS60, KF62, New60, New64c]. Parlett [Saa83]. Parlon [Gla96b]. Parratt [Rei62b]. Pars [Fle63b]. Parsimonious [Sch85a]. Part [BR91, Blo76, Blo77, BdD92, BD92, CLH97, Kap90b, Kap90d, Kni94b, Kni94a, Lax89, Ock92a, Ock92b, PY00, Pol76, Sal63b, Sco94, Sik92, Spi91a, McC03, Spi91b].

Parthasarathy [Eba82, Ros72b, St.82].

Partial

[Aar07, Abl84b, Abl84a, All78, Ame85, Ame86b, Ame95, Ant82, AB23, BNT10, Ber91a, Bob66, Boj93, BR18, Bro76b, Bru65, Che76, Che78, Col94b, DTV14, Del95, Dor74, EC76, FGN13, Fer79, Fle62, Fri80, GVL97, Gil90a, Gil90b, Gli91b, Gos83a, Gos83b, Gue85, Gup89, HF78, Hor97a, Kel83, Kor95a, KKS82, Lax69, Ler63, Lie95, Liu91, LA95, Mah75, MG76, Mey93b, Mit94, Olv93, OV85, PY98, Per68, Pic87a, Por80, RRK95, RL95, Rub69, RRG95, Rus78, Sch73a, Sco59, SS62, Sle79a, SBGP98, Smo91a, Smo91b, Syn83, TK97, Tho69, TH96, Ung77, Val67, Web85, Wei91, Wil85a, Wil86a, Yan65, ZD95, vLMZ19, Ans66, Col91, Det85, Gil78, Gue95, Lev88b, Sch63a]. Partial-Differential [Abl84a, Che76, EC76, Mah75, MG76]. Partially [Fel90, Fis93, Har63, McM87, Sch66]. Partially-Known [Fel90]. Participants [Ano67i]. Particle [Dur93b, Dys66, NC09, SKP19, TWW65, Win90, Wil61]. Particle-Like [Win90]. Particles [Ano91a, Bun83a, CLG14, Kla89, KMH^{+}90, Mey90, Mey91b, Poh63, Bun83b]. Particular [BT05, CS63, Dob87, Gup98b, Gup98a]. Partielle [Sch71a]. Partielles [Fal70a]. Parties [YAKM20]. Partition [And85, And87, Swi60, Gro98d, Ric98]. Partitioned [BI69, Deu81c, Deu81a, KSS80b]. Partitioning [BP76a, Kv93, KK99, UA07]. Partitions [Beh76, BKKŠ09, Bur79]. Parts [Mil61a, Sha04]. Parvu [BI68b]. Parzen [Doo63, McG61]. Pascal [Sch95, Sch94, Ham77]. Pascali [Hes83]. Pascual [Hur65]. Pass [Bri65, Mak86]. Passage [IWMW21, Rob62b]. Passes [Bis15]. Passing [Sku05b, Sku05a]. Passivation [AR63]. Passivity [Bau88].

Pastel [Old62]. Pat [Row76]. Patel [Ros78]. Paterson [Mir85]. Path [Bad60a, Cla83a, Cla83b, FR83, Gao18, GG80, Gon92, Kel84b, Kel84a, Mic95a, Moo94, New59, New62a, New64b, Tak83, Zwe16, ZN62, Mic95b]. Path-Following [Gon92]. Pathfollowing [Wet93]. Pathria [New76e]. Paths [Iba73, Kla76b, Mor98, Rab68b, Wil21, Edw63]. Pathways [Lem84]. Patience [Mal62, Mal63]. Patients [SFOM99]. Patrick [Ber83a, Man61d, Mur97, Pin69, Tul68]. Pattern [Beh76, Fre86b, GN22, GK95, GK98, Jor87a, JF87, Kee95, KW18, LS85, Pao88, Pao90a, Pao90b, Pao93, Ros80a, Ste94a, Ste94b, Sup77, Wan83, YH93, Hun84, Pei83]. Patterns [LPL05, PLL05, SM84, Str75, Ter93b, ZM20, Bry93, Haa93, Los85, Ter93a]. Patterson [Boe88a]. Paul [Ask72, Bei80, Ber91d, BW96b, Cus96, DG68, Dun94, ELRS66, Fal68, Han83b, Hei96, Her83b, Jag71, JW83, Kaz80a, McC85, Mil81a, Moo75, Nag88, Ner70, Nic94, Pro65a, Ral70, She77b, Sle62, SKGM67, Sto62b, Wax91a, Whi92a, Lew78b, LCG^{+}79, Wax91b, Whi92b]. Pavel [BI68b, Pao90a, Pao90b]. Pavel-Parvu [BI68b]. Payoff [LL69]. Pazy [Web85]. PCA [dEJL07]. PDAEs [BG18]. PDE [BM03, BS09a, DD04, GK12, PG00, Swa78]. PDEs [Kel84d, AG14, CHL08]. Peaceman [Say79]. Peano [Ros73]. Pearl [Rad90b, Rad90a]. Pearls [HRB96, Bei96, May91]. Pearson [Cho70, Dan93, Hag89a, Wag69]. Pease [Sma76]. Pecaric [Aga94]. Pedestrian [GHU22]. Peeling [BK78]. Peerless [Kri86]. Peinke [Abr94]. Peitgen [Dev92, Goo91a]. Peixoto [RB23]. Pelce [Kap90a, Kap90c]. Peleg [Luc87]. Pelinovski [Par90]. Penalty [Fel78]. Pencil [Ser76]. Pendulum [Abd83, Abd85, Ant98, BS93, LW95, Loc84a, Moe17, Lev88a, Loc84b]. Penetrable [KK75]. Penna [Boe88a]. Pennisi [Thi63a]. Penrose [Rob74, Rob76, GM76a, Zie73, Zlo70]. Pentikainen [Nas87b]. Pentium [Ede97]. People [Alb10]. Percentiles [WW77, WW78, Mei78a, Sau78]. Perception [Mil91a, Mil91b]. Percival [Rob85]. Percolation [BSS05, Dur84a, Dur90b, Gla95d, Boe96b, Dur84b, Dur90a]. Pereira [Jor80c]. Perelomov [Gui92, Tu92]. Perelson [Hop79]. Peretto [MG95]. Pérez [Mar85a]. Perfect [Bal70, CCG^{+}02, Kor76, LL69, Tuc73]. Perfection [Shu90a, Shu90b]. Perfectly [DGK16]. Performance [BNT76, Beu92, Bla88a, Bru83, Bru86, DKW^{+}09, DW95, FLGM23, Har76c, Hig00, MMT^{+}21, Mil98, Mos92, Raz92a, Raz92b, SAT22, Wol94c, For96, Che92a, Gua92]. Period [GAK17, PH13]. Periodic [Aiz94a, Aiz94b, Ano97j, AN83, BG77, Che96a, De 84, Dol93, FF95, Fin72, Guh66, Had65, HH00, Han87, Kol65, Kor95c, Lax76, LM90, MK99, Mey77b, Rox90a, Rox90b, SHL16, Sei65, Vel63, Vis01, Che96b, Fre95, Her65, Hoc65]. Periodically [Lev90a, Lev91]. Periodicity [CK76, Thu01]. Periphery [RPFM17]. Perko [Has92a, Has92b]. Perky [McG62b]. Perlis [McI71]. Permeable [Coh76]. Permutation [Bro64, ELRS66, HT72, Kla63b, Atk66, Meh73, Sha65]. Permutations [Car75a, Gol61, NW67a, San76, WKW75, Gou68]. Perpetual [VP09]. Perplexing [Wey62]. Perron [GTH23, Kee93, Mac00]. Perry [Hag85a, Nic65, RO66]. Persico [Pul64]. Persistence [Lud75b]. Persistent [FP21]. Person [Luc71, Ros72b, Tij81]. Personal [Ano68k]. Perspective [BCZ21, NC09, Sin89]. Perspectives [BC98, BD11, Cas82, Eng03, KLT03, KLT04, Naz94, Tre20, Wat93]. Persymmetric [Gol75c, Rei97b, Rei97a]. PERT [Sou76].

PERT-Type [Sou76]. Perturbation [Bir65, BBK20, Car70e, Cho75, Col94a, Com72, DPS73, Eck72, Fre75, FGTW16, GM95, GS22, GLO20, Hig07, HO96, How82, Ike78, KP87, Kev87, Kir12, Kla60, Kok84, LC72, Leb82, Lev97, Man87, Man88b, Maw95, MOSK97, Mur94b, O’M86, O’M87a, O’M96, Qui92a, Ric90a, Sch80b, SS89, Smi75, Ste73c, Ste77, Ste90b, Vas94, Was94b, WR96b, Wol77, Hol86a, Qui92b, Car65a, Col76a, Col77, Coo66, Dav89, Fle89, Hom86, Kee88a, Kev85, Lat69, Mey82b, Mur94a, Ran96b, Wil70a]. Perturbations [BS98, Eck77, FR11, Hoc70, Kel88c, Lio88, Lud75b, Mer85, MBK08, O’M92, O’M94, dJFO98, Har76b, O’M74]. Perturbed [Her65, Leb82, MBK97, O’M75, O’M88a, O’M08e, RSTK97, Sco91a, Sco91b, WE19]. Pervozvanskii [Cha66]. Peskin [Mil94b]. Pesonal [Ano69j]. Pesonen [Nas87b]. Pest [PM87]. Petar [Man88b]. Peter [Ans80, Bea90a, BI70a, Bla81, Boa76, Boa78, Boa87, Bri96, Ell85b, Far85, Fri71, GR95, Hag85a, Hey85, Hey89, Hof85, JS71, Joh87, Jow72, Kap70b, Kar90, Kri85, Kum82, Kum88, Lin90, Mar86b, Maw86, McM87, Meh96, Nit90, Per89, Pim87, Pra89a, Pra89b, RCC71, Rei83a, Ren92b, Rue78, Ser84b, Ter93a, VJ83, Wat90, Whi81, Wil64c, Wil65b, Wil77b, Wim88, Wol94a, Bea90b, Ter93b]. Peterson [BI68a]. Petit [Pap62a]. Petković [Gla86c]. Petre [Wib87]. Petri [Ani82]. Petrie [Pip81]. Petrina [Bak91b, Bak92]. Petrov [Bry96]. Petrovski [Lev67a]. Petrushev [Ste90a]. Petter [Mit91b]. Petukhov [May90]. Peyret [vL86b]. Peyton [May84c, May84d, Saa88]. Pfaffsche [Sch71a]. Pfanzagl [Eat96]. Pfeiffer [Sch82b]. Pg [Boe95a, Coh93a, Mel93, New94b, ADPR04, Win96a]. Ph [Mar85a]. Pharmacokinetics [RLGBRMC06]. Pharmacology [Dra61b]. Phase [Caf92, CESV15, CS75, GKR20, JU76, Lev63b, New76f, Pal96, VN76b, WP96, WM15, WE19, WPTZ01, Tre13]. Phase-Integral [Lev63b]. Phases [Kim65]. Phenomena [Gla86a, Gli91b, Kel90a, Maw93a, Maw93b, Mil89b, O’M86, Dij90, Whi93]. Phenomenon [CS97, GS97, Hig07, Mey89d]. Philadelphia [Hig00, Hig02c, Hig05]. Philip [Bar59, Bél88, Bra65, Cha82b, Che96b, Dru64, Gre65, Han75b, Kaz84b, Kow83, KZ93, ML85, Pan95, Saa96, Ste68a, Str76b, Str86c]. Philippe [Ber83a, Nas90, Wal88]. Phillip [Her84b, Moo76]. Phillipe [Mac92, Sit92]. Phillippe [Bar79]. Phillips [Tot91, vdV94]. Philosophy [Bro61, Fat76, KN61b]. Phipps [Rie72, Sin69]. Photoacoustic [ABJK10]. Photoelasticity [MT61]. Photon [Gan89, Gan90]. Photons [Wal92, Wal93c, Wal93b]. Phylogenomics [PS07]. Physical [Abl84b, Abl84a, ABL^{+}13, BC76, Dan66, DP76, Gue83c, Kla64b, Klo93, Kor68, Man61c, Rap73, RBC^{+}07, Sal63c, Sch87, She91, Ula64, Pra89a]. Physicists [Dim91b, FS96, Dim91a, Saf96]. Physics [Aar07, AK77, Ans64b, Ans68, Bar78, Bla66, Car65a, Dia64, Dys66, Erm89, Gli91a, Gru88, Gue69, Joh91a, Joh91b, Kaw65, MD96, Mit87, Mor91b, Nan20, New70b, Per76, RRG95, Sch80b, Sha95a, SS95, TWW65, War92, Wax91a, Wax91b, Wig67, Mla97, Olv97, Dol72, Dum96, Gol66b, Gue95, Gur62a, Her80a, Hye73, Kaz67, Kor65, Mar85c, Mor91a, Pro87a, Var67, Wig90, Wil88]. Physik [Gil65b]. Physiology [Kel75b, Man90c, Man90d]. Physique [Man61c, Lan86]. Pi [Wim88]. Picard [Yan92]. Piccinini [Kap85a]. Pickles [Cha95b]. Pictograms [Ste96]. Picture [DP76]. Pie [MP62, She60a]. Piecewise [BBK20, CBS76, DGP^{+}16, Eub88, Fle65, MG23]. Piecewise-Linear [DGP^{+}16, Fle65]. Piehler [Ham85]. Pierce [Hea91]. Piergiorgio [Mil80]. Pierre [Fix74, Gue88, Kap90a, Kir93a, Kir93b, Sch82a, Sel94, Tan93b, Tan93a, Van86, Kap90c, Kir93b]. Piet [Bic94]. Piezoelectric [Wil76b]. Pig [Gri77]. Pilz [Ruk93b, Ruk93a]. Pimbley [Hop79]. Pin [McN89b]. Pinder [Gos83b]. Ping [Gro96g]. Pinkus [Sch78c, Smi86]. Pinney [Bro59b]. Pinsky [Gue85, Ren93c]. Pioneer [Smi83]. Pioneers [Kaz91a]. Piotr [Syn93a, Syn93b]. Pipe [BK88, Kos79a, Kos79b]. Pipeline [DGK84]. Piper [Bla86a]. Pipes [Wer63]. Pipkin [Bur93, Day87]. Pironneau [Cuv90, Str86a]. Pirumov [Ock91]. Pissanetsky [Wil86b]. Piston [LB10]. Pitfalls [Kah76, Ver15]. Pitman [Gur63, Far94]. Pitu [Bur81b]. pivot [Ang90]. Pivoting [Fre86c]. Pizzo [Gre76]. PL [Rad73, Fla67]. PL/1 [Rad73]. PL/I [Fla67]. Place [Mey91a]. Placement [Fre14, HK72, Pré95b, Pre95a, Ste61a, ZW93, ZW94, Gro96d, Gro96h]. Planar [BP94, CLWM95, Don23, Kar22, Meg88, Mey95a, ML98, MP93, PY00, Sed96, Sou76]. Plane [CL08, Cor78, DS73, Ell02, Kei89a, Kei89b, Kri76, Mey75b, Pea82, Sah95, Wu75, Hor95a, Gil82]. Plane-Wave [DS73]. Planes [Nie96]. Planetary [HJ00, New76b]. Planets [Mor92]. Plankton [MPT^{+}02]. Planned [Ano96a, Ano97b]. Planning [BS09b, HLH03, Hyd86, TH96]. Plant [Gan89, Gan90, Ken93, Ant91]. Plants [Ant92c, Are91b, Ant92d, Are91a]. Plasma [Joh91a, Joh91b, Row64, VN76a, Wei90b, Yos60]. Plasmas [AHP^{+}95, Kri76, Rau95]. Plasmonic [YA18]. Plastic [Pit90a, Pit90b, Gos63]. Plasticity [Dod73]. Plate [Kel80a, PMR76]. Plateaus [Hil07, MV15]. Platen [GR95, Pro96]. Plates [Dru64, GK12, PMR75, Cla64a]. Plausible [HF05, Rad90a, Rad90b]. Play [Alb10]. Playing [Isb68]. Plemelj [Ric61]. Plemmons [Car81]. Pletcher [vL86a]. Plis [Spä97]. Pliss [Noh71]. Plouffe [BC96b]. PLTMG [Mit94]. Plus [DD02, SM14].

PMF [Mor85b]. Pohlhasen [Pap62b]. Poincaré [Com72, Kra24, Lev65, Vis01]. Point [Ale95, Ama76, Ama77b, BH76a, BWYY96, Chu94b, Chu94a, DdSG^{+}22, Don23, DG14, EBTZ94, ETZ94, Ell94, Esp95, Eva90, Fat76, Fra86c, GM95, Gol76b, GF76, GdMGR91, Gro91b, Gud75, Gue61, Gut90, Her70, Hey85, Hig02c, Hog73, IA75, Kla77, Kla78a, Kla91, Laz91a, Laz91b, Maw95, MCS69, O’M08e, Olv86, PW15, Ral70, Red80, San65, Sch76b, Sha92b, Tho59, Van78, VH78, VDC22, WS75a, YY96, Hey88, Row76, Sha92a, Ye94, Bai74b]. Point-to-Set [Hog73]. Point-Vortices [Don23]. Points [AG80, Art80, Bis15, Bot72, Ell02, Han90, HT78, Lem84, Man09, Mil68, Sil66, Sil67, VD68, Wei63a, dB65b, vV65, BW73]. Pointwise [Sig67]. Poised [Sha72b]. Poiseuille [Smi97]. Poisson [EB88, Mor85b, Sie90, App87, Boe88b, Bus69, Che98, Dor70, Dor73, Dor75, Eva88, Hai78, HG79, Ser78, Smi88c, Ste85, Swa77]. Polak [Nas77b]. Polansky [Rou88]. Polarized [GZ19, YAKM20]. Polaron [Gro98e]. Polderman [Bau91]. Pole [Cor78, Shu62]. Poles [Bru65]. Poli [Mat94b, Mat94a]. Policies [Fly76, Fox67]. Policy [GAM93, LL76a, Wei80a, Par92a]. Political [ML85]. Politics [Pac10, Saa96, TS96]. Pollack [Zie84b, Zie84a]. Pollak [Tre65, Van66]. Pollard [Gar67, Sle74]. Pollock [Bri64, Joh81, Pul91]. Pollution [GS23a, BS00]. Polubarinova [Pap63b]. Polubarinova-Kochina [Pap63b]. Polya [Ham79b, Whi76a, Dur93a]. Polyadic [HKD20]. Polyakov [May90]. Polyanin [Kla97]. Polygon [EV10, Kla70b, Kla71c]. Polygonal [Say13]. Polyhedra [CG66, Mur71, Ulk76, GGG69]. Polyhedral [GHU22]. Polylogarithmic [KF06, FK06]. Polylogarithms [Köl83]. Polymers [Pip86]. Polynomial [Aal71, Bar76b, BES05, Ber07e, BPE96, Bor80, BEV96, Bus64, Cod70, Coh69, Deu81b, Fet84a, FL84, Fra83, Gib87a, Gle81, Gle85, GKP75, GV97, GNZ20, Hog74, Hog75, Hou98b, Hou78, HS71, Hua86, Jok66a, Jor84c, Kar01, Kat76, Kat78, Kla73, Kri82, LW81, Los74b, LS84, Lux84, Nei19, ND82, New62c, NS63, PY00, Pan97, Pec62, Pro61, Raw83, Rie80, RSS81, Sch77a, Sho99, Ste70b, SY75, Van96, VG96, Vei86, Yan77, Ye94, Cha97, Hou98a, Wil98b, And75a, Gib87b, Gle83, Nev82, Sch65b, Sha82a, Sma76, DeM65, Ebe96, Kin76]. Polynomial-Time [Sho99]. Polynomials [Bac60, Bar74, Bar83a, Bar73, Ber64, BEV96, CM78a, CMSM79, Cha85, Cha72, Cha70b, Cha67b, Cha87b, Coh75a, DN59, DCL^{+}86, Gau87, Gla83, GG73, Hoc60, HT12, JC88a, Jor88a, KVM78, Kra66, Kra80b, Kra80a, Las07, Lev67b, Max60, Meh74c, Mon82, Mul64b, Pan92, Rog83a, Sch67, Seg06, SG76b, Smi80a, Smi81, Sne67, Sof96, Swa65, Tra66, Zel77, Sof97, Bar83b, Car73a, Der86, DG68, Fra66b, Jor84a, Los74a, SM79, SFH^{+}66, Wim80, Wim81, Kla90a, Van96]. Polysampling [Tuk87]. Polytope [BR74]. Polytopes [BR97, Sal85]. Ponstein [Zlo83]. Pontecorvo [Dei92]. Pontryagin [Bra63]. Pool [GM76b, GM76c, Tok95, Tok94]. Poorly [Han76]. Popov [Ste90a, Tot91]. Popular [Ham77]. Population [Ari92, ABG96, BvdD78, BG95a, CT75, Coh90b, CK76, CN15, Dri81, Dur97, Fre79, Key73, KA89, KR78, KS77, Mic88, Sán83b, San83a, Saw82, Sma83, TeB97b, AH98, TeB97a, Bra85b, Coh90a]. Population-Growth [CT75]. Populations [Gup82, LPL05, PLL05, Sán77, Sin93c, Tud90, Whi92a, Whi92b, Bra82, Fre86a, Sin93b]. Porous [Sch89b]. Porschan [Ros65]. Port [Spi80]. Porter [Hoc92]. Portfolio [Ste01]. Portrait [New76f]. portraits [Tre13]. Posed [Bec94, Bra96, Eng94, Fro86, HB96, Han76, Hea76, Kah76, Lat23, Rus76, Var76, Var79, Wil79, Han92].

Posedness [BT20]. Position [Sto60]. Positioning [NJN98b, NJN98a]. Positive [Ask69, Bha98, Car70a, Car71c, CMR73, Dem74, Duf84, Gas71, Gol66a, JRR71, Kel70, Lio82, Ren92b, Ren92a, Rez62, RAG72, Sch72c, The91, VG78, Wou66, ZM22, Bue66, Kee88b]. Positivity [LW81, Meh74c, Sha82a]. Possibilities [Mey76b]. Possibility [McA92]. Possible [CLG14, Jac07, JGM95, Jor95, Kla94b, Lar69b, Sch66, TH96]. Post [Gom67, Haw95, Kee88b, Lew79, Ort96b, Ort96a]. Post-Doctoral [Gom67]. Post-Furnace [Lew79]. Posteriori [CL23]. Poston [Guc79]. Postori [BE78]. Postulates [Leh65]. Potential [Ant72, Bax85, DTT98, Fis88b, GJ19, Kaz75, Kla88c, KG89, Kui06, Lig86, PFM76, Sei76, Sim71a, Spi80, TE96, Tru94]. Potentials [Ars60, Dei92, GZ19]. Potokovije [GG91]. Potpourri [HU07, Kap72]. Potra [Woź87]. Potter [BL10]. pour [Cuv90]. Powell [Chu82]. Power [BCH14, Cha83a, CSN09, Fej88, Fol64, Gla81b, Gla82b, Hau75, Hau76, Ism76, JMFB15, Jor86b, Jor86c, KA76, Kwa98, NF16, Pac10, PRM^{+}66, Ren84b, Ren84a, SG82, Tu76, Van75, Wil85b, Jor82, Sha82b, Sle83a]. Power-Law [CSN09]. Power-Series [Hau76]. Powers [EL85, EL86, Fis71a, Gar69, Jor86d, NS77, Rot81]. Poynting [Jef94, Rob94b]. Pozrikidis [Riz94]. pp [BC96b, Bor06, Hig05, Tre13]. Prabhu [BR67]. Practical [Bry90a, Bry90b, Cod70, DeR91a, DeR91b, Hoc92, Ike76, McA96a, MTM96, Smi80b, Tap84a, Var79, Arn92, Den82, Dev90, Kow83, Tap84b]. Practice [Bak91a, Liu92b, MMS98, Ral83b, WJV79, Mor97, Ham73b, Ham73c, Odg86, Ral83a, Cor16]. Practitioner [CT95, Tab95]. Practitioners [Ber93]. Prade [Gog85, McA92]. Prager [Mas87, Mro91]. Prakasa [Lah94, Sam86]. Praly [Bau88]. Pranab [Cha94c, Far94]. Prasad [Hou70b, Hou71a]. Prasada [Mor85b]. Prato [Rus95]. Prautzsch [Pet95]. Preasymptotic [GT10]. Precedence [Ric67c]. Precipanu [Dat80]. Precise [Bou21]. Precision [Fat76, Kel22, MK78, Pro61]. Preconditioners [CPSV18, GZ19]. Preconditioning [Gon06, PW15, Wat22]. Predator [CC96, PH13]. Predator-Prey [CC96, PH13]. Predictability [Jin91]. Predicting [DeS88, LH75, MQ18, Pet76a]. Prediction [Ano61a, Ghi76, LRV76, Neu97a, PL78, Raz87, Som78, Coo67a, Sal63d]. Predictive [Siv94c, WHZG24, Siv94a]. Preface [Ano73k, Gre67a, Han84, Wri99]. Preference [Kar90, Wal88]. Preferences [Ell85a]. Prefix [Gal95, LDG95]. Preisendorfer [Bar67]. Preliminary [AW76]. Preparata [Fra88]. Prescribed [Aal71, Ann95b, Ann95a, Cha87b, Gro93a, Inn69, JC88a, Jor88a, Zay92]. Presence [EW19, PB75]. Presentation [Cha70b]. Preserve [WM78]. Preserving [AG14, GST01, MWL97, Smi88b, WS23]. President [Tho60b]. Press [BC96b, Bor06, Hig02b, Isa88, Isa89a, Yan94]. Pressure [Fis88b, Kel80a]. Pressures [Saw59]. Preston [Nit85]. Prestrud [Bar64, Bar65]. Previewing [TH96]. Previous [Sch62b]. Prey [CC96, PH13]. Preziosi [Log97]. Price [Bla86b, Sam73]. Pricing [CD03, HS15, VP09]. Priestly [Grc18]. Prigogine [Cur64b, Oth82]. Primal [Ham88, YZZ04]. Primal-Dual [YZZ04]. Primality [Bla88e]. Prime [Sho99]. Primer [Bel63a, BT13, FM18, Kaz89, McG93, Bla83, Fla67, Boa85]. Primrose [Kaz91a]. Princeton [Bor09, Tre15]. Principal [BF82, Fet81, Flu93a, Flu93b, Gla89c, Jor90b, JG90, Koh98a, Boa82a]. Principle [BJK24, BCM10, DPS73, Fle88, Hau81, Kla74a, Red89, Rei78a, Sch76a]. Principles [BS94a, BS94b, DJLQ21, Edw90, Eva84b, Eva84a, Gol66b, Gue83b, HSV95, Hau90, Kel90d, MT61, Muk86, Pay68, Pet76a, Red90, Red83, Sto81, Deu63, Gor59, Gra72b, Kel90c, Run91, Sop72]. Priouret [Dup93]. Pritchard [Gue79, Sle89b]. Probabilistic [Asm98, Bax85, Beh76, Bla88c, COSG19, DF98, GWZ16, HMT11, IGWL90, Kar76, Kle75, Lev84, Loc89, LIW90, McL79, MG79, Rad90a, SO78, Ste74, Ste75, Way82, tR83, Bla88a, Fro73, Guc87, Kv93, Moo75, Rad90b, Rei69, Tay86, Wei69, Yao86]. Probabilités [Dur87]. Probabilities [Ang92b, Dun64, Har66a, Kij93, Pow93, Pow94, REG19, She76, Wol60, BR69, Dwa64]. Probability [Ang94, Ano89b, Ath94, Bar91a, Bar91b, Bec74, Bla91, BD63, Bry96, Cha69, CQRZ96, Dic75, Dra62, Dwa66, Esc70, Fro87b, Gas76b, Gas76a, Gia63, Goo82, Gre87, Gre73, Gue61, HBO66, How94, Hud95, Jag88c, Jd93, JMH95, Kac62, Kel84c, Kes69, Key94, Kha71, Kla64c, Kur69, Kus76, LS78, Lem83a, LP71, Man61a, Man63, Mar83a, Mar85b, McG61, Mil92, Moh66, Mos92, NW70, Ole68, Pen95, PB96, Pin69, Pin72a, Rei63, Rei77b, Rei78c, RO66, Sau62, Sch64, Sie69, Spi74, Ste67a, SF97, Sug76, Sug77, Gla97a, Gla98, Ram88, Gut97, Dwa85, Eic67, Hai85, How63, Kee61, Lem83b, Lil67, Mec92, Rei62a, TT76, Dur92b, Fle79, Har64b, Har76a, Kaw66, Kem85, Kul65, Lev87, Mar83b, Mil69a, Por86b, Rei62b, Rei88, Sam80, Shi94, Sie90, Van68]. Probe [NO60]. Probing [PPR11]. Problem [Ada97b, AN78, Ale98, All92a, All92b, Alt76, AB97, Ame91a, Ame91b, And75b, And76, Ann95b, Ann95a, Ano73d, ABS95, Ark66, Arn70, Auc97b, Bal85, Bar85, Bei80, BPV97b, Bil93b, Bil94b, BL05, BDJ83, BL84, Boö86, Bor80, Bot74, Bou67b, Bou68, Boy05, Bro60a, Bro82, BB83b, BDvD^{+}98, Can86a, Car64, Car91, Cas87a, Cas80, Cha95a, Cha96c, CC75, Cha87c, CIS95, Cho76b, Chu81, CGRG^{+}23, Cli76, Col84, Con97, CEH12, Cra74, Cra75, CV77, DH09, Eds05, ESS79, Erd72, Erd92c, Erd94, Eva90, Faz96, Fie93, FL82, FN73, FS89, FSRJ90, Fra65a, Fra65b, Fra66a, Fre75, Ful84, Fus68a, GVL97, Gao18, Gie63, GS84, Giv63, GK78, GD78, Gla97b, GI97, GP79, Gol81, Gol76c, Gra63, Gri77, GdMGR91, Gro91b, Gro93a, GNP95, GV97, GW95, Hay68, Hen82]. Problem [Hil75, HS14, Hol78, Hou68a, Hou68b, Hue91, Jag93, Jag94, Jer95, Jev93, Jev95a, JK59, Joh75, JS76, JRZ97, Kai59, Kan88, Kar01, Kau88, Kaw95, KNS60, Kla64a, Kla65, KL80, Kla84g, Kla84f, Kla88c, Kla94a, Kri91, Kri97, Kri77, LDG95, Lan76b, Led96, Lee86b, Lee86c, LB10, Lev65, Lew82c, Liu83, Liu94, Liu95, Log98, LRJ78, LRV76, Mai72, MM89, Mat91, McC03, McG93, Mea98, Mei83b, MP72, Min70, MB86, NC76, Nas65, NN67, NW61, New65b, New77, New82b, NBS12, Nie98, NY90, O’B11, O’M75, Ono84, Ono85, OK94, Pac88a, Pap86, PC61, Pat03, Pea81, PF62, PR90, Pla15, Pul64, RT97, Ric89, Rob62b, Rog75, Rog69a, Row66, Sch04a, SSW99, SKG^{+}80, Sek82, Sha94, Sha95b, She60a, She72, She60b, She91, Shu62, Sil61]. Problem [Sim75, Sim76, SW93b, Sou80, Spi76, Spi86, Ste61a, SBXD06, Tap72, TM01, Tro71, Tu76, TTY88, TS86, Utz79, VCG^{+}86, VS95, Ven76, Vij83, Wag63, Wal85, Wal86, Wan97, Was75, Wat93, Wei87, Wei60b, Wil70b, Wil67c, Wol81, Wol82, Wol94b, Won64a, Zay92, vA61, vL60, vL61, Ano97f, Boe98a, Cha96a, Gro98b, Par10, Ren97a, Sha98, Ver98, MW06, BW65, Bha93, BdDJ83, BdD94, Bou65a, Bre70, Bro83b, Can86b, Car67b, Cas87b, Cot73, Cur77, Eas93, FS73, Fen94, FB78, For75, FZ63, Gal95, Giv78a, Göb70, Gre88, Gre63b, GdM95, GWK95, Hor89, Jor81, KG89, KF62, KS73, Kra68, Kri82, Lew83a, LM81, Los76b, Los76a, Los83a, Los84a, MP62, Mei83a, Mor93, NPW95, Nic94, Par66, PT81, Saa83, Sch95, Sch85c, SK80, She63a, Sib95]. Problem [Sle62, Van86, Vel71, Wal87, Wei67b, Wil69, WM85, Ano87b, Ber85g, Ber86c, Bla85b, CL72, RJ90]. Problem-Solving [Ber85g, CL72]. Problems [Abl84a, Aiz94a, Aiz94b, Ale95, AB97, Ama76, Ama77b, Ame86a, Ano95n, Ano95o, Ano95p, Ano96n, Ano96o, Ano96p, Ano96q, Ano97k, Ano98d, Ans68, Ant78, AR95, Aro78a, Bai74b, BB96, Baz73, Bel65, Ber84a, Beu87, BH15, Blu91, Bog75, BS98, Boo94, Boy08, BE78, BEG23, BL10, Bue71, BK78, Bus76a, Cas80, CE89, CZ92, Che91, CM78b, Chu98, CG24, Cla67, CTW21, Cle61, CH66, CB76, Col94c, Col94d, CA75, Con97, Con66, Coo64, CR67, CL87, Cul94a, Daf89, Dam84a, Dau90, Day89, Dei92, Den76, Dol91a, Dol91b, DPS73, DGK16, DGLZ12, Duf69, Dun94, DMP22, DX98, ERHO95, Eck72, ERS14, Elv75, EFA76, Eva84a, Faz98, Fel76, Fel70, FP97a, Fis88a, Fix93, Fla99a, Fla99e, Fla99b, Fla99c, Fla99d, Fla00d, Fla00a, Fla00b, Fla00c, Fla01a, Fla01b]. Problems [Fla01c, Fla01d, Fla02a, Fla02b, Fla02c, Fla02d, Fla02e, Fla03a, Fla03b, Fla03f, Fla03c, Fla03d, Fla03e, Fla04b, Fla04c, Fla04d, Fla04e, Fla05c, Fla05a, Fla05b, Flu78, Fri78b, Fri79a, Fri79b, FLC96, GK12, GHC97, Gei89, Geo76, Ger62, Gol69b, Gol73b, Gos64, GV97, Gud75, Gue69, Gun88, Gur62a, HY71, Hal86, HK72, HB96, Han76, HSV95, HV75, Hea76, Hel75, Het72, Hig86, Hil96, Hir75, Hol66, Hop75, HMO97, HM73, HA75, Hüb88, Hue92b, Hue96, Hul76, Ips05, Ips06a, Ips06b, Ips06c, Ips06d, Ips07a, Ips07b, Ips07c, Ips07d, Ips08c, Ips08d, Ips08a, Ips08b, Ips09d, Ips09a, Ips09b, Ips09c, Isa89b, IA75, Jac69, Jen78, Joh83d, Joh83b, Jub70, Kah76, Kap85a, Kap90b, Kap90d, Kap76, KJH87, KH98, Kel80b, Kel79b, KG97, Kla62b, Kla92b, Kno81, Kok84]. Problems [Kri97, Lan64, Lat23, Lau74, LP75, Lie97, Lud70, Luk75, MS91, Man09, Man76, Mar73, MTGF78, Mat75, Mat76b, MG76, Mat85, MQ69, MR76, MR00, Mer64, Mey75a, Mey76a, Mey77a, Mik89, MOSK97, MT64, MM10, Mor74a, MK97, MCS69, Mur92, Na68, Nas77b, Naz80, NW67b, New70b, Nie94a, Noh71, O’M71, O’M88a, O’M08e, Ock92a, Ock92b, Olm76, Pac88c, PR91, Par86, PF13, PC89, Pla15, Pon90a, Pon90b, Pri76, Pro87a, PA95, Ram78, Red80, Red76b, RFS03, Ren95, Ric90a, RM75, Rog94b, Rog94a, Rog69b, Roo10, Ros76b, RR95, RRE97a, RRE97b, RRE97c, RRE97d, RR98a, RR98b, RRe98, Roy91, RW80, Rus76, SSCS95, Sar84a, Sch86b, Sch81a, Sch76b, Sei76, Sha72b, Shi17, Shu09, Sle92, SV00, Ste73c, Ste77, SH94, Sül97, SH96, Sun76, Tad87]. Problems [Tak79, Tan93a, Tay68, Ten01, Tho59, TH96, TD20, TWW65, Tra66, Tur78, UFW17, Van93, VB75, Van78, Van75, Var76, Var79, VC06, Wah92, Wan97, War68, Wat96, Wat05, WS75a, WE76, Wil88, Wil76a, Wil67b, Wil73, Win67, Wu72, ZZ18, ZC74, AR81, Han92, PC90b, Ram88, SGD98, Par10, Abl84b, Ale90, And95, Ano92b, Bai74a, Bal88a, Bar86, Bar79, Bec94, Ber65a, Ber84b, Ber94, Ber85c, Ber85g, Ble81, Blu86, Blu94b, Bra96, Cla78c, Cot94, Cru86, Cul94b, Cum96, Dam84b, Ell87, Ell88, Eng94, Fix74, Fro86, Gea85, Gen76, Gin85, Gla93b, Gra67, Gre76, Gue83c, Her78a, Her70, Kap81, KJ87, Kaz89, Kee90, Kel67, Lev88b, Lig86, Obe64b, Ode86, Ode89, Ork95, Pap62b, Pin70, Ral66b, Ral70, Rhe88, Rus90, Sag77, Sar84b, Sco91a]. Problems [Sco91b, Shi93, Ske95, Smi74, Sop60, Tan93b, Van88, Wen68, Wil78b, Wil79, Hol86a, Kur69, Pil87]. Problems-Theory [Gun88]. Problems. [Ban87, Eva84b]. Procedure [BJ75, Bro59a, Bro60b, FH76, Mor64, Mor66, New76f, WP78, Mal77]. Procedures [BBKV96, Fel78, Gar75, Hen75, Vai96, Noe82, Spu89]. Proceedings [Man61d, Moo60, Pip86, Por86a, Blu65, Cla62, Des68, Jun64, Mil61a, Mul64a, Rei63].

Process [BGL17, CGT^{+}97a, Hai78, HG79, JMX96, Kaw65, MPRW02, SBXD06, Xu96b, Rul98, Ram90a]. Processes [ABG^{+}95, Ang92b, Ano61a, Ath93a, Bal78, BI70b, Ber91a, BR67, CJR18, Cha66, Chu88a, CAiJS23, Dax90, DK85, DSL04, Dur97, Erm89, Fei96, FG19, FJT19, Fra65c, FSY65, FOTG96, Ger62, Goo72, Gra86b, Gus90a, Gus90b, Hag85c, Ham62a, Hey88, Hey89, Joy71, Kar82, Kar84a, Klo93, LK98, LC65, Lim23, Lud87, Mar75a, MR18, Men94a, Mil69b, Mil61b, Moi89, Mor98, Mor73, Nor74, Oco87, Par88, Pin70, Pit79, Por81, Por86b, PF96, Sha92a, Sha92b, Shi94, SW93a, Ste85, Sys63, Tho75, Uhl73a, Uhl73b, Way82, Wei77, Yan92, Pel97, Ath93b, Bar65, Bar70, Bel62, BR62, BR65, BR79, Bro76a, Col86, Doo63, Dwa70, Ell81, Ell94, Fle83, Fle90, Fox71, Get96, Gra66b, Gut90, Hey85, Hüb91, Kar83, Kar84b, Kaw67, Kaw68, Kei64, Key94, Kul65, Man61b]. Processes [McK95, Par67b, Pit77, Por89, Pro91, Ser80b, Spi82, Woź87]. Processes-Second [Por86b]. Processing [And78, CGT^{+}97b, Ham88, Hei95, Hol79, Hop92, McA92, RM89, Sta88, WB09, Cam91a, Car65b, Kar60, Lai78, Mac68b, Sta89, Sta90, Ull88]. Processing-Continuous [Sta88]. Processor [RKM76]. Processors [GDR78, MKR76, Zie76]. Proctor [Luc18]. Product [CGT^{+}97a, Cli64, FG76, Gre66, Hen75, Kel88b, MC91, NS66, She61, Sne67, Tew66, Tew67, VS95, Wat05, Yan97b, CGT^{+}97b, Yan97a, Gre67c, FGB81, Koh85, SNS66, Van63, Cha81b]. Production [Gor75, Pol59]. Products [Coh88, Coh93a, Eks60, FSP93b, FSP^{+}93a, For76, GRS97, Gro92a, MM06, RM89, Tis83b, Tis83a, Coh93b]. Professional [Dre60, Gas60]. Professionals [Her81]. Profile [PGS75]. Profits [Ver15]. Program [AR63, Arm76, BJ76, Hac75, Ken74, Lan76a, Ros67a, VB75, Wet74b]. Programmation [Ham87]. Programmer [Sch94].

Programmers [Rad73]. Programmierung [Gas70]. Programming [ABGJ21, ASPL03, Bel65, BI68a, BTKNZ00, Ber65a, Ber85d, Beu89, Bla85a, Bos99, Che91, Che76, Che78, Cli76, CC76, Cul68, DTG98, Den67, Eck80, Ell85b, Fla66, Fla67, Fly76, FS78b, Fro74, Fro80, Geo71, GMSW91, Gon92, Gri61, Hag80, Hal66, HHSS75, HK93, Him78, Hog73, Hue96, KWB96, KW71, Kar87, KK98, Kle83, KSW96, Kor76, KD76, KW99, Lar78, Lev97, Mac69, Mac70, MW76a, Man70, Mar60, MM95, McG67a, McI71, Mey76b, Mey79, Nau86, Nei10, OLHB02, Orb18, Pat03, Pet76b, PF96, Rab68a, Rub77a, SS65, SZ96, Sau76, Ser80b, Ser84b, Ser84a, TT94, Ulk76, VB96, VBC07, VP09, Vie15, Wil70c, Win62, Wol75, Wou62, YZZ04, Zha96, dEJL07, Dor74, Geo67, BI67, Ber86a, Bla85a, Bla88d, Bla89b, Bri96, Cha64a, Eck82, Fal68, Fei96]. Programming [Fro73, Kan78, Kar64, Lar74, Lem85, McC71, Mla94, Nau86, Sin59, Sop62, Tal81b, Via85, Win60, Ye94, Zlo77, BI68c, McG62b, RA63, Ros71, Wri60]. Programming-Theory [BI68a]. Programs [BT69, Cot76, Cot94, Duf70, Fox67, Gri72, LCK^{+}17, Pat19, Wet74b, Ask79a, Din72, Fro83, Mit93, Orl93b, Pra89a, Lea94]. Progress [Deu85, Kel79a, Pan97, Par78, Rus78, Wol75, McG64, Por86a, Sal62, Wil62b]. Progressing [TH96]. Projected [PD89]. Projectile [Lar86, Sha90]. Projection [Ari91a, Ari91b, BB96, BGW15, CG24, DG20, GF78, Red80, Und76, YKF96, Ram97]. Projection-Based [BGW15]. Projections [Gon06, Hun81, SM78, Ste77, Zah80]. Projective [Hug93a, Hug93b, ST91b, ST91a, Boe88a]. Projectors [BBR13]. Projects [Boy08, Koç94]. Prolegomena [Día78]. Prony [WM63]. Proof [Abe88, AH95a, Bag95, Bel68, Cha63, Dax97b, De 20, EH78, Elh78, Hei69, Hou98b, JP89, Kno98, DMS23, Mor85a, PH78, RB23, ST94, BG96, Dax97a, Hil98, Hou98a, Mor86c, Pac10]. Proofs [Het72, KSW96, Mac00, Pat19]. Propagating [Mor76]. Propagation [AN83, Bla66, CO07, Cho75, Cho76b, Coo67b, Dat86, Dat90, Fej88, Kel64, Kel67, Kel79a, Lud70, Mar86a, Mar88, Mor88c, MP93, PWG18, Rog69b, She87, Squ61, Sub90, Whe74, Wil62c, Wil64a, Wil64c, Wil76b, Xin00, Zua05, Kel86a, War64]. Propelled [CLG14]. Properties [Ari92, BS94a, BS94b, BBR13, Boo79, Cho76a, Coh69, Dol93, Dri91, EH10, Fic60, Gay78, Guc87, GG68, Her88, Hig03, Inn69, Jin91, LCE76, Mah75, Pim93, Pow86, Rei97b, Ros68b, Sch83, Sed96, Wil76a, ZZ09, Rei97a, Aal71, Shi94]. Property [AB69, BCT11, Bra66, BS62, Bri65, BG64, Geo95, Gla81c, Har67c, HM78, Jag76b, Jag82, JGK^{+}82, Jor78b, Jor80a, Kau77, Kla62a, Kro66, Olk77, RS79, Vel64, Rue98]. Proportional [BJK11, GK15, Mor69]. Proposed [Mil76]. Proscribed [Coh69]. Prospective [Hal70]. Prospects [Kel79a]. Protein [Neu97a]. Proteins [Neu97a]. Protocols [Len93c, Len93b]. Prototype [WMZ94]. Protter [Pay68]. Prove [Mas92]. proved [Dut82]. Providence [Lax89]. Proving [Mor74b]. Proximal [CMSZ24, CKCH23, DMP22, FR16, VDC22]. Proximity [ZKP17]. Proxy [Boy13]. Prugovecki [Coo73]. Prusinkiewicz [Ant92c, Ant92d, Are91a, Are91b]. Pruss [Cor95]. Pryce [And95]. Przemyslaw [Ant92d, Are91a, Ant92c, Are91b]. Przeworska [Noh77]. Przeworska-Rolewicz [Noh77]. Pseudo [Ste77, Wil86a, Wit22, Uri92]. Pseudo-Differential [Wil86a, Uri92]. Pseudo-Inverses [Ste77]. Pseudodifferential [CT96, Tay96, Mey93b]. Pseudoinverse [GM76a, Gol75c, Gre59, Gre60c, Pri64, SB72]. Pseudoinverses [Mey69a]. Pseudolinear [KW99]. Pseudomanifold [Fre86c]. Pseudoplastic [NC76]. Pseudospectra [Tre97]. Pseudospectral [Tad87]. Pshenichnyi [Nas77b]. Psychology [Bir69b]. Ptak [Woź87]. Public [Bla88f, GAM93, KM04, McK81, Par92a, Tud92]. Public-Key [KM04]. Publication [BD64]. Pugh [Car66a, Kaz92]. Pull [DeS88]. Pull-Up [DeS88]. Pulse [CO07]. Pulsed [BM03]. Pulses [All92a, Car91, All92b]. Pulvierenti [Bar95]. Pumping [MG23]. Pure [BL05, Kaz91b]. Purkinje [WiOH10]. Purpose [Laf93, LL94]. Purposes [Cle59, Fin72, HMH00]. Pursuit [Ano90a, CDS01, DG96, Kla90b, Kla95a, Gro96c, Rox77]. Pustyl’nikov [Jef95]. Puterman [Fei96]. Putnam [Ros68b]. Putting [Bla97b, Bla97a]. Puzzles [Hal95, O’R10, SH95]. Python [Orb18].

Q [Ari96, Ber93, Bus86, Par86]. QED [vLMZ19]. Qn [SG76a]. Quadratic [CD20, Cas80, Che91, Cot76, GMSW91, HHSS75, HK72, Har76c, Lei78b, Lev90a, Lev91, Mil81b, Nas65, New76f, New78, PY00, Ral69a, TM01, Wan91, YZZ04, Kaz83]. Quadrative [Cli76]. Quadrature [Amo78b, Boa70, FW66, Gar75, HF78, KW76, Lam85, Lyn83, MKS86, Mon82, SK78, Squ67, Ste93a, Str82, Tre08, Tre22, McC86, Str68, Ste72]. Quadratures [Gat64]. Quadrilateral [CD93, KL75]. Quadruple [Car67d, Gla89c, Jor90b, JG90, Pet68]. Qualifications [Pet73]. Qualitative [Her71, Hor97a, JS93, Kaz86, Lae82, Mac76b, Mah75, MQ69, MWL97, Bal88b, Gue83c, Ham65a, Hoc94, Mey75b, Sch87]. Quality [CGT^{+}97a, CGT^{+}97b]. Quantification [MQ18, OSS^{+}13, WHZG24, ZMPK24]. Quantifying [GLY^{+}18, SAT22]. Quantiles [Ang92a]. Quantitative [Fre86a, Ste81b, WK78, Ber62, Dra61b]. Quantities [Jef18]. Quantization [Pro84b, Pro84a, Stu90]. Quantum [AvK^{+}08, Bou91b, BvJ09, Bro83a, Bro90, BZ05, Coo67a, Dei92, Fos68, Hag85b, Kea93b, Nan20, Pro65a, Pro83b, Pro83a, Roy91, Sch92, Sho99, Vog87, Vog91a, Vog91b, Wes84a, Wig91, Gro98a, Coo73, Hag89a, Kea93a, Pim93, Vog93, Wes84b]. Quarteroni [Boy88]. Quartic [HK69, HK70, HK94, Rab94]. Quasi [Bry93, DM77, DS79, Dri68, Mor98, SB98b, SS89, Slo93, SM94, Tap76, Tin76, Tod84, Zar68]. Quasi-Linear [Dri68]. Quasi-Minimal [SB98b]. Quasi-Monte [Slo93, Zar68]. Quasi-Newton [DM77, DS79, Tap76, Tin76, Tod84]. Quasi-Random [Mor98, SM94]. Quasi-Regular [Bry93]. Quasi-Steady-State [SS89]. Quasiconformal [Ear81]. Quasicrystals [Kaz91a]. Quasilinear [HW76, Kov96, Lev96, Lie97, SGK^{+}96, TtK96, Urb93a, Urb93b, Cha78]. Quasilinearization [WS23, Ral66b]. Quasiperiodic [Mos66a]. Quasivariational [Ell87]. Question [Naa01, Wal76b]. Questions [HU07, MP20, New66, New68b, Rus78, Zah80]. Queue [Beu83a, Gre67b, Beu83b, Kei71]. Queueing [BM93, Beu92, Den78, DK85, Fro87a, Hey89, Hey93, Ser86, vLMZ19, Hey82, Tho76a]. Queues [Kel82, Mor78, New65a, Beu85, Bar59, Hey85]. Queuing [Ism83b, Ism83a, Kei65]. Quick [Mas92]. Quickie [Ano90a, Ano90k, Ano90l, Ano91b, Ano91j, Ano91k, Ano92a, Ano92b, Ano93c, Ano93b, Ano93a, Kla90b]. Quinney [Ast87]. Quirk [Szy87]. Qun [Ath93b, Ath93a]. Quotient [AMSV02, Mcc78, TDS18]. Qureshi [Jód96].

R [Abr94, Ans64d, Ant78, Ant90, Ant91, Arb77, Are86, Ask80, Asn78, Bal95, Bar95, Bar64, Bar65, Bau88, Bel63a, BI68a, BI70b, Ber62, Ber85d, Ber85f, Bla88b, Blu86, Boe88a, BdD93, Boj65, Bou76a, Bre62b, Bro76a, Bro90, Bry93, Bun83b, Bun83a, Car66c, Cha94a, Cha94b, Cha95d, Che75, Che86, Col87a, Coo65, Cor92b, Cul77, Cum91, Dam84b, DeM65, Den82, Dia64, Dij90, Dol62, Dur87, Eas93, Eba82, Ell94, Enr84, Esc70, FGB81, Fin76b, Fis76, Fle84a, For75, Fre83, Fri82, Gib90, Gib87b, GC79, Gla87a, Gol65b, GGG69, Goo82, Gre89, GRYL89, Gun63, Gun95, Gur63, Gut82, Hag82, Ham61, Ham62a, Ham62b, Ham73a, Har74, Har82, Hei96, Her78a, Her69, Hes72, Hig62a, Hig62b, Hig65, Hol87, Hol75, Hom86, Hor83a, Ism83b, Ism83a, Ivi65, Jag74a]. R [Jag79b, Jam59, Jol96, Jor78a, Jor80c, Jor89a, Jos95, Jow76, Kal70, Kan62b, Kaw68, Kaz92, Kei64, Ken86a, Ken93, Kes65, Kes75, Kes85, Kin82, Kla84b, Kla92b, Kle71, Kog66, KS73, Kor95a, Kul69, Lai78, LF66, Law72, LeV93, Leh63b, Les76, Lev62, Liu91, LJW73, Los73, Los74a, Luc86, Lud82, Mal85, Mal90, Mar64, Mar88, May90, McC91a, McC92a, McG67b, McG64, McK95, MN83, Mec92, Mei78a, Mic95b, Mil82c, Mir85, Mit93, Moo60, MW97, Mor93, Mor96, Mul69, Nas87b, New76e, Noh67, Obe77, Ode89, Orl93b, Owe85, Pap62b, Par81a, Pav94, Pin85b, Ple88, Poh63, Pol59, Pra85, Pro65b, RA63, Ral66b, Rei77c, Ros68b, Row86, Rox92, Sal62, Sal63a, San95, Sau78, Sau79, Sca59, Sch94, Sco84b, Sco84a, Sco90a, Sco90b, Seg79, Sel76, Sha97a, Sha85]. R [She63b, Shi68, Sin93a, Sle94, Sle87, Sop59, Spi84a, St.82, Sta88, Sta90, Ste93b, Sto61, Stu91, Tap84b, Tho87, Tre84b, Tre84a, Tru94, Tut94, Uri92, VN73, Wal79, Was74, Wat63, WS71, Wig90, Wil88, Wim96, WGM93, Yan91, Zak85, Zar65, Zie76, Zim90, vH87]. R- [MW97]. R. [Dam84a, Joh84a, Pea65, Sar84a, Tap84a, Wil84a, Zie83a]. R.-D [Ell94]. Raa [Lev84]. Rabier [Ber83a]. Rabinovich [Bry94, Dau90]. Rabinowitz [Ste68a, Str76b, Str86c, Wol94a]. Rachev [Dur92b]. Racquetball [Kel84c, Mar85b]. Radack [Bri64]. Radar [Che01, New67a].

Rader [Eng17]. Radhakrishna

[Kup75, Luk75]. Radial [Mor88c, WS13]. Radiant [Bar64]. Radiated [KA76]. Radiation [Bra81, Cha87a, Fel70, Hni89, JH90, Ran76, Row64, SFOM99, Shi97, Jor90c]. Radiative [Eds05, Bar67]. Radio [Fej88, Sto62a]. Radionuclide [Tre78]. Radius [BF96, BK88, CLG14, Jor86c, New76c, Vos75, Vos76, Wil85b, Jor86b]. Radon [Zal83, Nie86]. Rafael [Ros80a]. Raghavan [Ros72b]. Ragland [Gue92]. Ragsdell [Tap86]. Rahman [Wim91b, Mor97, Wim91a]. Raiffa [Sca59]. Raina [Jor80a]. Rainer [Dob87, Wil84b]. Rainville [Cho66]. Raj [Kaz67, Raz92b, Raz92a]. Rajendra [Dav89]. Ralescu [Arb77]. Rall [Rhe70]. Ralph [Cho87, DeM65, Hey89, Lee84b, Lev88b, Par74, Ren90, Rob81]. Ralston [Hul61b, Lin79c]. Ram [Gue72]. Ramachandran [Jd93]. Ramanujan [Bre81, Jor84b, And74b, And83, BSS14]. Ramifications [Baz73]. Ramm [Tay89]. Ramo [Gre60a, Kar60]. Ramon [Hag86, Han67, Han81b, Hul71]. Ramsey [EH78, Elh78]. Rand [Kee88a]. Randall [Pit91]. Randles [Noe81]. Randolph [Mit94]. Random [Ada97b, And90, Ano99b, BNT10, BGL17, Ber92a, Bor79, BR18, BD63, Bry96, CR96, Coh88, Coh93a, CP78, Con71, CL87, Dav75c, DF99, DK85, Dra62, Dur92a, The99a, EV10, Fig93a, Fig93b, FJT19, FK62, FLNU18, Fra65c, God87, Gra86b, HS70, HT72, Hig65, HD62, Jev92b, Jok66b, Kei64, Kel90c, Kel90d, Kem92a, Kem92b, Knu86, KL87, Knu87, Kul65, Kur69, KWP95, Law95, Lud75b, Mal80, Meh73, MW61, Moi89, Mor98, Pel97, PB96, Phi85, Pic87b, Por86b, Por89, REG19, Rog78, SBH^{+}20, Sch61, SK20, She62, She63b, She64, Shi94, Sie89, Sku05b, SM94, SS78, Ste87, Ulk76, WL95, Wei63a, Wig67, Wil62c, dB65b, Ada97a, Coh93b, New84, Sku05a, Bel62, BR85, Cha66, Daw74, How83, Jun70a, Kem92a, Mer85, Mod92, Par80, Pro95, Slo93, Woo92]. Randomized [Con71, DG20, GGK^{+}22, Kar01, LW17]. Randomly [AKP^{+}91b, BGML96, Ell90, Gue61, LS95, Boe96a]. Randomness [Beu87, HMT11, Len93b, Len93c, Sey93a, Sey93b, SW22]. Range [AV65, CCCM06, Gan98b, NN96, Pit90a, Ros69, Sha90, Gan98a, Pit90b]. Ranjit [Kno83]. Rank [Ber73a, Bro66b, CFG95, DG20, FC76, JMJ95, Jur95, Kat68, LW04, LT13, RFP10, Shi17, Ste76, FARE69, Tho74]. Rank-One [CFG95, LT13]. Rank-Revealing [DG20]. Ranked [Hey86, New87]. Rankin [Coo67a]. Ranking [Kee93]. Ranks [Mey69a, SN04, Goo86]. Rao [Ano73a, Hin87, Kar82, Kup75, Lah94, Luk75, Man87, Mor85b, Sam86, Zie76]. Raphael [But61, McM61c]. Raphson [Ypm95]. Rapoport [Coz85]. Rardin [Zim90]. Rare [DM07, FHRO95, O’B95]. Rarity [Pit81]. Rasband [Rob91]. Rassias [Neu97b]. Raster [Lew81a, Lew82a]. Rate [KW18, WPL98]. Rates [Bur78b, Giv63, Ros95, Spa62b, TE96, WM61a, Lau98, WM61b]. Rating [Sch75a]. Ratio [Lop94, NBS12, Rob95b, Rob95a, Sho66a, Str65a]. Rational [Ask81, Bla93, Bru65, CS63, Cod70, Deu95, Fix78, Fra75, Ful76, KL75, Wah03, Dev94b, Fre90, Sle86, Ste90a]. Ratiu [Lee84b, Her97]. Ratschek [AKP91a]. Rausand [Mee96]. Ravindran [Tap86]. Rawlins [Jag87b, Jor84c]. Ray [AHP^{+}95, She75, Wu72, Tul90, Rau95]. Rayleigh [AMSV02, Koh72, Mcc78, Mor64, TDS18]. Raymond [Fis71b, Mat88, Ric74]. Raynaud [Bas88]. Rays [CLOT99, Fel70, Row64]. Reactance [Sha64b]. Reacting [Buc85, Kaz80a].

Reaction [Bil93b, Bil94b, BdD94, Coh76, Fif86, GSYY22, KW18, PS00, WMZ94, Rul98, Chu92, Lu92]. Reaction-Diffusion [Bil93b, Bil94b, BdD94, Fif86, KW18, PS00, WMZ94, Rul98, Chu92, Lu92]. Reactions [Hig08, PSV12]. Reactive [CM89]. Reactor [Hul61a, Var67]. Readings [Win60]. Real [Bel63a, Bha98, Boa83, Boa84, Bra66, CD75, Cum96, FLC96, Gat88, Gla93a, Lau82, Lau83, Lau84, Man68, Max60, Nik78, Rez62, Seb75, Sha04, Sof96, Sof97, ST98, Ste90a, Szy87, Tu76, Vel64, Wan84, WR94b, WR94a, Wou67, Bel10, Wat65]. Real-Time [Tu76, Bel10]. Real-World [Cum96, FLC96]. Realization [Chu88a]. Really [MS14]. Reamer [GRK81, Ren80b, GK81]. Rear [FR83]. Reasoning [HF05, KB59, Rad90a, Rad90b, Sch87, Lus87]. Received [Ano62e, Ano63f, Ano63g, Ano63h, Ano63i, Ano64e, Ano64f, Ano64g, Ano64h, Ano65e, Ano65f, Ano65g, Ano65h, Ano66g, Ano66h, Ano66i, Ano66j, Ano67e, Ano67f, Ano67g, Ano67h, Ano68g, Ano68h, Ano68i, Ano68j, Ano69f, Ano69g, Ano69h, Ano69i, Ano70f, Ano70g, Ano70h, Ano70i, Ano71h, Ano71i, Ano71j, Ano71k, Ano72g, Ano72h, Ano72i, Ano72j, Ano73g, Ano73h, Ano73i, Ano73j, Ano74e, Ano74f, Ano74g, Ano74h, Ano75k, Ano75l, Ano75m, Ano76j, Ano76k, Ano76l, Ano76m, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano78l, Ano78m, Ano78n, Ano78o, Ano79i, Ano79j, Ano79k, Ano79l, Ano80l, Ano80m, Ano80n, Ano80o, Ano81g, Ano81h, Ano81i, Ano82m, Ano82n, Ano82o, Ano82p, Ano83j, Ano83k, Ano83l, Ano83m]. Received [Ano84l, Ano84m, Ano84n, Ano84o, Ano85g, Ano85h, Ano85i, Ano85j, Ano86h, Ano87h, Ano87i, Ano87j, Ano87k, Ano88f, Ano88g, Ano88h, Ano88i, Ano89i, Ano89j, Ano89k, Ano89l, Ano90g, Ano90h, Ano90i, Ano90j]. Recieved [Ano86i, Ano86j]. Recipes [Isa88, Isa89a, Yan94]. Recived [Ano81j]. Recognition [Beh76, Jef90, Mic92a, Mic92b, Pao90b, Pao93, Pei83, Ros80a, Wan83, YH93, Bri64, Pao88, Pao90a, Sup77]. Reconstruct [Zwe16]. Reconstruction [AG01, Hun81, LBL02, Ran91b, SM78, Zah80, Ran91a]. Reconstructions [MMH04]. Recourse [Wet74b]. Recovering [McL86a, Sha04]. Recovery [Beu66, PG00, Tek89]. Rectangle [Dor70, Swa77, Wei63a, dB65b]. Rectangles [BI68b]. Rectangular [CHL08, Cli78, DKS96, Gla84b, Gre59, Hub83, Jos70, PMR75, Wix76]. Rectilinear [Meg88]. Recurrence [CM78a, CMSM79, FS68, Fet72b, Gau67, McN67, RSTW73, SM79]. Recurrences [FT76, FF78, LS08]. Recursion [AB73, Car66a, Lan76b, Lea94, Pug66b]. Recursions [Enr84]. Recursive [Dav77, EGJK04, FM68, Sho69, Smi82, Wou62]. Recycling [HM01]. Reddy [Gei89, Hau85, Hau88, Wen78]. Redheffer [Fis71b]. Reduced [CH66, HPS07, MQ18, RBC^{+}07]. Reduced-Order [HPS07, MQ18, RBC^{+}07]. Reducibility [KN67, Ste61b]. Reducible [Mey89a, Win67]. Reducing [NO60, OW88, PGS75]. Reduction [Ani82, BGW15, CFG95, Jok66a, Kel84d, LS22, PY00, Swa77, WKW75, War68, WM15, WE19, Smi87]. Redundancy [Fro83, Leh63b]. Reed [LF66, Los73, Mac61]. Reexamined [Grz89]. Referees [Ano70j, Ano71o, Ano71p, Ano72n]. Reference [CS63, GGK^{+}22]. Refined [BG18, MM10]. Refinement [Bra76a]. Reflected [Cho72, Rue73a]. Reflecting [HS70]. Reflection [BB83b, De 00, GS23c, MV93]. Reflections [Par71]. Reflector [Nie89b]. Reflectors [De 00, TWW65]. Reflexive [KW71]. Reformulated [Mah75]. Reformulation [AR81, Win96b]. regellose [Hig62d].

Regelungstechnik [Hig62d].

Regenerative [SW93a]. Regime [vLMZ19]. Regiment [Sim75, Sim76]. Region [BS09b, OH76, WS13]. Regional [Kaz82]. Regions [Bor79, MW76a, Rog78, Zha13]. Register [Fre82]. Regression [BP95, CR78, Efr88, GHW76, Lev90a, Lev91, Lin92, Mil76, Nie94b, Pan95, Pin92a, Ruk93a, Ruk93b, VSM14, YS76, Ber92c, Pin92b, Ber92b]. Regular [Ano93c, Bry93, GK95, Ike76, Kee95, Mur93, Tri95, YT95, Mey84]. Regularity [Bom02, Urb93b, Urb93a]. Regularization [BBK20, Nas86, Neu98, PCLS23, Ten01, UFW17, Eng94]. Regularized [CP18, Str76a]. Regulator [Arm76, Kal70, O’M75, Wan91]. Reid [CZ89, Daw74, Laz82, Mil82a, Par74, Sys63, Wei69]. Reiko [Sar84b]. Reiman [Hey82]. Reingold [Smi74]. Reinsch [Par72]. Reintjes [Pfl92]. Reiss [Ell94, O’B95]. Reissig [Fin76b]. Rejected [Hor78]. Rejection [Flu90, Grz89]. Reklaitis [Tap86]. Rektorys [Ral80]. Related [Ang94, Ano99b, BI70a, Ber85b, Cot94, The99a, Fet65, Hou65, Kaw68, ML85, Moh66, Mon82, OT12, Pro84c, Ros68b, Sha63a, Sha63b, Tan93b, Tan93a, Gro98d, Ric98]. Relating [Mor97]. Relation [Ber75, Eva93, Eva94, Fet72b, Gra72a, Gra70, Kla66, McN67, Pug66b, Ren94b, Sha64a, Car66a, FS68, RSTW73]. Relational [Zie83b]. Relations [AB73, CM78a, CMSM79, GG80, Gau67, Goo59, Lar69a, Ric61, SM79, TW69, Zas65, New74, Pro65b]. Relationship [Naz76]. Relative [ADR76, Jor77a, Kli78, Ral65b]. Relatively [SM94]. Relativistic [Tau80]. Relativity [Gol66b, Kan62a, Poh63, Pul91, Mor92, Rin61, Lou92, Mag86, New76e, Xan87]. Relativizations [Boo94]. Relaxation [Cho78, CTW21, Gof78, O’M88b]. Relay [Ath87]. Relevant [MMT^{+}21, TS90]. Reliability [Cle80, FSV06, GPP88, GP88, HRM96, Kar01, LC65, LSS98, Mee96, NW67b, Phi87, Raz87, Sau79, Sin88, Sin93a, Spi87, Wei63b, Ros65, Aro63, Ell90, Sch65a]. Reliable [Bou21, Kul69]. Remaining [KA89]. Remaking [Grc18]. Remark [Fen68, Sch62b]. Remarks [Alt76, Bra81, FW66, SS62, Sma90]. Remembering [Ben95]. Remi [Mei89, Mey94a]. Reminiscences [Han86a]. Remote [MS91]. Renardv [Kor95a]. Renardy [Daf89]. Renate [Cav90]. René [Fig93a, Man61c, Red83, Fig93b]. Renewable [Cla78a, Cla79a, RV78]. Renewables [CS17]. Renewal [Fox67, Hey89]. Rennie [GK81]. Renormalization [Kir12]. Renormalizing [BS93]. Renton [Hut89]. Repayment [GH22]. Repeated [Gat64, Gil65a, GS22, Wil24c]. Replacement [Bur63, Bur64]. Replenishable [Ngu76]. Reply [Nas92, Cia92a, Cia92b, Rob94b]. Report [AW76, Ano61f, Ano75p, Ano77r, Ano78r, Ano79m, Ano80p, Ano81k, Ano82q, Ano83n, Ano84p, Ano85k, Ano86l, Ano87l, Ano88j, Ano89m, Ano90m, Ano91l, Ano92i, Ano93j, Ano94g, Ano95q, Ano96r, LCK^{+}17, Tho60b]. Representable [MK78]. Representation [Bar73, BKS68, BK69, BDD^{+}95, CC23, GJ19, May84b, Mer84b, Mer84a, Moo77b, New76b, Phi68, Phi71b, Rob74, Rob76, Sah95, Sin69, Sou76, Wal70, YH76, Zie73, May84c, May84d, Rus95]. Representations [Apl95, DS73, KA95, Mei77, Tew69, TBBER21, Var87, Cur64a]. Representing [Aro78b]. Reprint [For59]. Reprinted [ADPR04, FK06, Lax89]. Required [Ebe75, Ler63]. Requirements [VH78]. Research [Arm90b, Art83b, Art83a, Art85b, Ban87, BI70a, Boy08, Bro61, CCR^{+}62, Car67e, Coh88, Coh93a, DA79a, DA79b, Fos90, Gar89, Ips10a, Ips10b, Ips10c, Ips10d, Ips11a, Ips11b, Ips11c, Ips11d, Ips12a, Kil18a, Kil18b, Kil18c, Kil18d, Kil19a, Kil19b, Kil19c, Kil19d, Kil20a, Kil20b, Kil20c, Kil20d, Kil21a, Kil21b, Kil21c, Kil21d, Kil22a, Kil22b, Kil22c, Kil22d, Lor91, Muk86, Ng94, Ros67a, RWMD18, Sal62, San14, San15a, San15b, San15c, San15d, San16a, San16b, San16c, San16d, San17a, San17b, San17c, San17d, Sop60, Tum12a, Tum12b, Tum12c, Tum13a, Tum13b, Tum13c, Tum14, Wil23d, Wil23a, Wil23b, Wil23c, Wil24a, Wil24b, Wri14a, Wri14b, Coh93b, Lit63, MN83]. Research-Methods [Sop60]. Researchi [Dre60]. Reservoir [Say79]. Reservoirs [Ode82]. Residual [Naz80, SB98b]. Residue [Jag72d, Spi71, SB72, Flo69]. Residues [BM93, BS66]. Resin [CGT^{+}97a, CGT^{+}97b]. Resistance [Bha93, Bou65a, Fry64, GBS08, GS84, Gra63, Kla70a, Pla15, Sha64b, Spi91a, Spi91b, Tri62a, WM85, Amo64, Bou66a, MS65]. Resistances [Sin79a, Ren80a]. Resistor [Leh60, Mye75, RL62]. Resolution [BBK20, CS97, GS97, PR91, Zah80]. Resolvents [KK78]. Resonance [Mat75, MG76, SW11, Mat76b]. Resonances [Pal96, WP96]. Resonant [Deu95, GS22, Hen76]. Resonantly [Wil24c]. Resource [CD94, Cla78a, CN19, Kee90, Ken86b, Ngu76, She72, Wet75]. Resource-Based [CN19]. Resources [Cla79a, RV78]. Respect [Kus93]. Respecting [BJK24]. Response [Bon67, Gre89, Jef94, Lin59, Nas92]. Responses [MQ18]. Restart [BES05, Emb03]. Restitution [CS06]. Restoration [Ang78]. Restoring [HSŠ16]. Restricted [BCT11]. Result [Boj65, GS63, JM92, MV93, Ros80c, SV98, SR93b, SR93a, Tri78, Bur81a]. Resultant [Bar74]. Resulting [LM00]. Results [Alt76, Ang94, BKS68, BE78, Bra81, Cas80, CA75, Ghi76, HY71, Her73, LM79, MP20, Miu75, Nie94a, Roh66, Sei65, Ser78, Smi88b, TF65, Udw89, Miu76b, Miu77b, Hag85b].

Resurgent [Daa97]. Retarded [MO78, Han91]. Retorting [Geo76]. Retrieval [BDO95, BDJ99, CESV15, GKR20, LM05, Sal64, Sal63b]. Retrospective [Hal70]. Rev [AR80, Coh93b, Dea90b, For96, Gud78b, Mor86c, PC90b, Rob94b, SW80, Win96b]. Revealed [EH10]. Revealing [DG20]. Reversal [Kar96]. Reverse [BK76, Knu77a, Knu77b]. Reversibility [Spi81b]. Reversion [FO99]. Revesz [Kem92a, McN89a, Kem92b]. Review [Aal71, Abb92, Abl77a, Abl80, Abl84b, Abl86, Abr94, Ack74, Aeb94b, Aga93, Aga94, Aga95a, Aga95b, Ahl82, Ahl93, Aiz94b, AK90, AKP91a, AKP93a, AS85, ASF86, Ale90, Alf93, All78, All92b, Alm87, AA82, Alp85, Als85, Als92a, Ama82, Ame81, Ame84, Ame85, Ame86a, Ame86b, Ame88, Ame91a, Ame95, Amo64, Amo78a, And75a, And95, AG67, And73, And74a, And77, Ann95b, Ano61a, Ano63a, Ano68a, Ano71a, Ano73a, Ano80d, Ano82f, Ano87a, Ano87b, Ano89b, Ano91b, Ano92a, Ano92b, Ano93c, Ano93d, Ano93b, Ano93a, Ans64a, Ans64b, Ans64c, Ans64d, Ans66, Ans68, Ans80, Ans85, Ant72, Ant78, Ant82, Ant83b, Ant86a, Ant87, Ant90, Ant92a, Ant91, Ant92d, Ant63, Ant86b, Aok70, Apl95, Arb76, Arb77, Arb78, Are86, Are89, Are91a, Are93, Ari96]. Review [Ari91b, Arm94, Arm90a, Arn89, Arn92, Aro63, Aro69, Ars62a, Ars62b, Art83b, Art85a, Art85b, Art88, Art87, Art91a, Art92a, AGR83, Ash89, AF69, Ask72, Ask75b, Ask78, Ask79a, Ask79b, Ask80, Ask81, AD87, Ask92a, Asn75, Asn78, Asp84b, Ass77, Ast87, ABG96, Ath94, Ath87, Ath93b, Atk66, Aur18, Axf65, Bai87, Bai74a, Bai74b, Bai76, Bai93a, Bak66, Bak91a, Bak91b, Bak92, Bal88a, Bal88b, Bal93, Bal95, Ban84b, Ban85, Ban87, Ban88b, Bar86, Bar95, Bar64, Bar65, Bar67, Bar70, Bar94, Bar89, Bar91a, Bar83b, Bar79, Bar59, BS67, Bas88, Bau88, Bau89, Bau91, Bax85, Baz73, Bea90a, Beb68, Beb79, Bec85, Bec94, Bei78, Bei80, Bei81, Bei85, Bei96, Bél88, Bel90a, Bel90b, Bel62, Bel63b, Bel63a, BI67, BI68a, BI68b, BI68c, BI70a, BI70b]. Review [Ben87, Ber93, Ber95, Ber62, Ber65a, Ber83a, Ber84b, Ber85a, Ber91a, Ber94, Ber85b, Ber91b, Ber92b, Ber92c, Ber76a, Ber85c, Ber97a, Ber81, Ber86a, Ber91d, Ber85d, Ber78, Ber83b, Ber84e, Ber85e, Ber88a, Ber90b, BW65, Ber85f, Ber85g, Ber86c, Bes92, Beu83b, Beu85, Beu87, Beu89, Beu92, Bha93, BR62, BR65, BR67, BR69, BR76, BR79, BR85, Bic94, Bin84, Bir64, Bir69a, Bir86, Bjö61, Bjö85, Bla70, Bla82, Bla85a, Bla85b, Bla88a, Bla88b, Bla88c, Bla88d, Bla89a, Bla89b, Bla90, Bla83, Bla81, Bla84, Bla86a, Bla88e, Bla88f, Bla93, Bla66, Bla86b, Ble81, Blo89, Blo91, Blo76, Blo77, Blo95, Blu65, Blu86, Blu91, Blu94a, Blu94b, Blu78, Boa82a, Boa65, Boa76, Boa78, Boa81, Boa82b, Boa83, Boa84, Boa85, Boa87, Boa88, Boe88a, Boe89]. Review [BGML96, Boe76, BdDJ83, BdD92, BdD93, BdD94, BG95b, Boe95b, BC96a, Boe96b, Böh88c, Boj93, Boj65, BGL^{+}67, Boj97, Bol91a, Bol91b, Boo79, BW73, Bor73, Bor79, Bor90, BC96b, Bor06, Bos95, Bos73a, Bos73b, BS86, Bou91a, Bou91b, Bou77, Bou65b, Bou65a, Bou66a, Bou66b, Bou76a, Bou86, Boy70, Boy88, Boy90a, Boy97, Bra94b, Bra85a, Bra96, Bra66, Bra65, Bra82, Bra85b, Bra91a, Bra63, Bre70, Bre74, BMMO76, Bre77, BPK84, Bre89, BKS96, Bre62a, Bre62b, Bre64a, Bre68, Bre94a, Bre81, Bre93, Bri64, Bri96, BS80, Bro76a, BBLS92, Bro77, Bro65, Bro93, Bro59b, Bro76b, Bro83a, Bro90, Bro83b, Bru72, Bru73, Bru95, Bru83, Bru86, BOS73c, Bry90a, Bry91a, Bry93, Bry94, Bry96, Buc69, Buc91a, Buc85, Bue66, Bue71, Bun83b, Bur94a, Bur94b, Bur81a]. Review [BW96b, Bur79, Bur81b, BG64, Bur93, Bus86, Bus95, Bus70, But89, But92, But61, But74a, But74b, But74c, But78, But79, Byr65, Cab74, Cab75, Caf92, Caf94, Cam77, Cam91a, Cam91b, Cam93, Can91a, Can86b, Car93, Car66a, Car67b, Car70a, CS70, Car71a, Car71b, Car72a, Car73a, Car80, Car76b, Car81, Car66c, Car65b, Car65a, Car87, Cas87b, Cas95, CL62, Cas72a, Cas94a, Cav90, Cel78, Cer96, Ces75, Cha94a, Cha94b, Cha87a, CL23, Cha95b, Cha67a, Cha70a, Cha78, Cha66, CL72, Cha94c, Cha94d, Cha82b, Cha83b, Cha86, Cha95d, Cha96c, COW96, Cha64a, Cha64b, Cha81b, Cha68, Che91, Che92a, Che96b, Che65, Che67b, Che18, Che89, Che92c, Che75, Che86, CS64, Chi90, Chi79, Chi83, Chi71, Chi94a, Chi95, Cho94, Cho66, Cho70, Cho87, Chu81, Chu82, Chu85, Chu88b]. Review [Chu89, Chu90, Chu96, Chu78, Chu92, Chu94b, Cia92b, CMR73, Çin93a, Cla62, Cla64a, Cla64b, Cla78c, Cla83b, Cle69, Coc84, Coc86, Cod78, Coh76, Coh90a, Coh75b, Coh82, Coh84b, Col76a, Col77, Col87a, Col91, Col73, Col87b, CZ89, Col92, Col96, Col94b, Col86, Col89, Col94c, Con85, Con96, Coo66, Coo67a, Coo73, Coo89, Coo91, Coo65, Cop86, CK86, Cor92a, Cor95, Cor79, Cor92b, Cot73, Cot94, Cou97, Cox65, Coz85, Coz87, Cra84, Cra86, Cra89, Cre97, Cro73, Cro85, Cru86, Cul77, Cul94b, Cul66, Cul68, Cum91, Cum92a, Cum96, Cur64a, Cur77, Cur64b, Cus96, Cuv90, Cyb93, Daa97, Daf89, DC86, DV76, Dam84b, Dan93, Dan91, Dan66, Dan71, Dan72, Dat67, Dat69, Dat73, Dat78, Dat80, Dat86, Dat90, Dau90, Dav75a, Dav75c, Dav77, Dav89, Dav91b]. Review [Dav91a, Dav94a, Daw74, Day87, Day89, De 74, DeM65, DG96, DK97, DeR91a, DeW85, Dei92, Del95, Dem86, Dem90, Den78, Den82, Der86, Des68, Det85, Deu63, DRWK64, Deu81c, Deu90, Dev90, Dev91, Dev92, Dev94a, Dev94b, Dia59, Dia64, Dic67, Die92, Dij90, Dim91a, Din72, Din94a, Dis92, Dix91, DG68, Djo68, Dob87, Dod73, Dol91a, DD74, Dol62, Dol72, Don85, Don90, Doo63, Dor89, Dor74, Dra61a, Dra61b, Dra62, Dre89, Dri63, Dri65, Dri75, Dri81, Dru64, Duf84, Duf85, Duf86, Duf91a, Duf91b, Dum96, Dum65, Dun94, Dup93, Dur93a, Dur84b, Dur87, Dur90a, Dur92a, Dur92b, Dur93b, Dur97, Dwa64, Dwa66, Dwa70, Dwa72, Dwa80, Dwa85, Dwy63, Ear81, Eas89, Eas91, Eas93, Eat96, Eba82, Ebe96, Eck82, EK91a, Ede90a, Edw90, Edw63, Eic67, Eis86a, Eis86b]. Review [Elk86, Ell85a, Ell92a, Ell90, Ell80, Ell81, Ell85b, Ell87, Ell88, Ell94, Els89, Els91, Elu88, Eng78, Eng94, ELRS66, Enr84, Enr87, Enr89, Eps59a, Eps59b, Erd77, Erd83, Erd92b, Erm89, Esc70, EF65, Eva81, Eva84b, Eva85, EB88, Fab80, Fab87, Fad62, Fal70a, Fal68, Fal70b, Fal73, Fal93a, Fan59, Fan82, Far94, Far85, Far93a, Far65, Far69, FSW^{+}66, Far96, Fat93, FS73, Fei96, Fej88, FGB81, Fel90, Fel72, Fen94, Fer79, Fet65, Fet66, Fet68b, Fet68a, FS68, Fet72a, Fet80, FLR^{+}81b, FW82, Feu76, Fia73, Fie93, Fif79, Fif86, Fig93a, Fin84b, Fin95, Fin76a, Fin76b, Fin93, Fir92b, Fir97, Fis93, Fis71b, Fis71a, Fis76, Fis88d, Fix74, Fix78, Fix93, FSP93b, Fla97, Fla66, Fla67, Fle62, Fle73, Fle76, Fle63a, Fle63b, Fle64, Fle72, Fle77a, Fle77b]. Review [Fle79, Fle82, Fle83, Fle84b, Fle85, Fle88, Fle89, Fle90, Flo69, Flu93a, Fly89b, Fok84, FB78, Foo63, For92, For75, FARE69, Fos90, Fox71, Fra86a, Fra66b, Fra86b, Fra72, Fra86c, Fra88, Fra89, FR63, Fra95a, Fra95b, Fre80, Fre78, Fre79, Fre83, Fre86a, Fre95, Fre90, Fre86c, Fri78a, Fri82, Fri86, Fri64, Fri71, Fri59, Fro60, Fro73, Fro74, Fro80, Fro83, Fro85, Fro86, Fro87a, Fro87b, FZ63, Gal95, Gar93, Gar61, Gar91, Gar67, Gas63, Gas71, Gas76b, Gas85b, Gas70, Gas95, Gat88, Gea74, Gea80, Gea85, Gei89, GR95, Gen86, Gen76, Gen96, GK86a, GP88, Ger96, Get96, Gha89, Gha94, Gho87, Gib85, Gib95, Gib73, Gib90, Gib77, Gib87b, Gib97, GA63, Gil65b, Gil78, Gil82, Gil90a, Gil95, Gin85, Giu18, Giv78a]. Review [Giv78b, GC79, Gla95a, Gla86a, Gla94a, Gla78a, GR81, Gla85a, Gla86b, Gla86c, Gla87a, GD88, Gla88a, Gla89b, Gla93b, Gla95b, Gle83, Gli91a, GC69, Göb70, God87, Gog85, GM87, Gol65b, Gol65a, Gol66b, Gol69a, GGG69, Gol71, Gol73a, GK81, GG91, GH86, Gol93, Goo72, Goo82, Goo84, Goo86, Goo90a, Goo91a, Goo91b, Goo68, Gor59, Gos63, Gos64, Gos79, Gos83a, Gos83b, Gos84, Gou68, Gra18, Gra67, Gra72b, Gra86a, GRL87, Gra66a, Gra83a, Gra86b, Gra66b, Gra70, GM85, Gra79, Gra93a, Gra94a, Grc18, Gre60a, Gre63a, Gre97, Gre62a, Gre88, GB89, Gre76, Gre65, Gre87, Gre89, Gre72, Gre63b, Gri61, Gri88, Gro87, Gro94a, GRYL89, Gro90, Gro91a, GdMGR91, Gro93d, Gro93a, GdM95, Gro96e, Gro96f, Gro96g, Gro96h, Gro88b, Gru74, Gru88, Gru93, Grz83]. Review [Gua92, Guc87, Guc79, Gue69, Gue72, Gue79, Gue82, Gue83d, Gue83b, Gue83c, Gue84, Gue85, Gue88, Gue89, Gue92, Gue95, Guh66, Guh68, Gui92, GWK95, Gun63, Gun88, Gun95, Gup82, Gup89, Gup91a, Gup96, Gur62a, Gur62b, Gur66, Gur63, Gus76, Gus90a, Gut97, Gut82, Gut88, Gut90, Haa93, Hab83b, Had65, Hag85b, Hag85a, Hag89a, Hag80, Hag82, Hag83, Hag85c, Hag86, Hai85, Hai77, Haj96, Hal66, Hal95, Hal90a, HPF85, Ham85, Ham87, Ham88, Ham90, Ham61, Ham62a, Ham62b, Ham65a, Ham65b, Ham70, Ham71, Ham73a, Ham76, Ham77, Ham79a, Ham79b, Ham80, Ham73b, Ham73c, HK72, Han95, Han94b, Han75b, Han79, Han81a, Han85a, Han87, Han91, Han64, Han67, Han81b, Han83b, Han85b, Han86b, Har64a, Har73, Har74, Har83, Har96, Har61b, Har61a, Har64b]. Review [Har66a, Har67d, Har70, Har72a, Har76b, Har76a, Har78, Har93, Har63, Har72b, Har75, Har82, Has86, Has92a, Hau85, Hau86, Hau88, Hau89, Hav95, Hea91, Hea74, Hea81, Hei93a, Hei93b, Hei89, Hei97, Hei96, Hel97, HBO66, Hen86, Hep67, Her77, Her78a, Her80a, Her80b, Her83b, Her84b, Her97, Her65, Her67b, Her67a, Her69, Her71, Her88, Her93a, HF05, Her70, Hes83, Hes86, Hes72, Hey82, Hey85, Hey86, Hey88, Hey89, Hey93, Hey97, HR62, Hig62a, Hig62c, Hig62d, Hig62b, Hig65, Hig92, Hig99, Hig00, Hig01b, Hig02b, Hig02c, Hig05, Hig06, Hig12a, Hig12b, Hig12c, Hig12d, Hig13a, Hig13b, Hig13c, Hig14a, Hig14b, Hig14c, Hig14d, Hig15a, Hig15b, Hig15c, Hig15d, Hig15e, Hig16a, Hig16b, Hig16c, Hig16d, Hil76, Hil96, Him78, HS72, Hin87, Hin91, Hin89, Hin94]. Review [Hir71, Hir73, His77, Hoc23d, Hoc23a, Hoc23b, Hoc23c, Hoc24a, Hoc24b, Hoc65, Hoc79a, Hoc79b, Hoc80, Hoc82, Hoc83b, Hoc86, Hoc92, Hoc94, Hof85, Hof95, Hol91b, Hol70a, Hol85, Hol86b, Hol86a, Hol87, Hol89a, Hol89b, Hol95, Hol66, Hol75, Hol90a, Hom86, Hop79, Hop83, Hor83b, Hor83a, Hor89, Hor97a, Hor93, HR85, Hou67, Hou70b, Hou71a, HA71, HT81, How63, How82, How83, How94, Hua96, Hüb88, Hüb91, Hud95, Hug93a, Hul61a, Hul61b, Hul71, Hun81, Hun84, Hun85, Hur65, Hur93a, Hut89, Hut91, Hye73, Isa88, Isa89a, Isa89b, Ism76, Ism78, Ism83b, Ism85, Ism96b, Isr92b, Ivi65, Jac76, Jac97, JS71, Jag64, Jag71, Jag72a, Jag72c, JBL72, Jag72b, Jag72d, Jag73, Jag74b, Jag74a, Jag76a, Jag79a, Jag79b, Jag80, Jag82, Jag83, JGF87, Jag87a]. Review [Jag87b, Jag88b, Jag88a, Jag90, Jag92, Jam59, JW83, JRR71, Jan70, Jar76, Jef95, Jel87, Jer81, Jer91a, Jes77b, Jes93b, Jet91, Jin91, Jd93, Jód96, Joh67, Joh81, Joh83f, Joh83c, Joh83d, Joh84b, Joh87, Joh88, Joh91a, Joh68, Joh73, Jol96, JSR82a, Jon92a, Jon92c, JG70, JF71, Jor72, Jor73, Jor77a, Jor77b, wBJ77, Jor78a, Jor78b, Jor79, Jor80a, JA80, Jor80c, Jor81, Jor82, Jor83a, Jor84a, Jor84b, Jor84c, Jor86a, Jor86b, JB88, Jor88a, Jor88b, Jor89a, Jor90a, Jor90b, Jor90c, Jor93b, JG93, JGM95, Jor87a, Jor87b, Jor89c, Jor89b, Jos95, Jow72, Jow76, Jow86, Jun64, Jun65, Jun70b, Jun70a, Jur95, Kal63, Kal93a, Kal65, Kal70, Kv93, Kan78, KH85, Kan62c, Kan62a, Kan62b, Kan66, Kan68, Kap85a, Kap85b, Kap90a, Kap90b, Kap64, Kap65, Kap66]. Review [Kap67, Kap68, Kap70b, Kap71, Kap72, Kap76, Kap77, Kap78, Kap81, Kar90, Kar60, Kar63, Kar78, Kar82, Kar83, Kar84b, Kar64, KJ87, Kar87, Kat65, Kau86, Kau88, Kaw63, Kaw65, Kaw66, Kaw67, Kaw68, Kaz62, Kaz67, Kaz75, Kaz80a, Kaz80b, Kaz82, Kaz83, Kaz84b, Kaz86, Kaz89, Kaz90a, Kaz91a, Kaz91b, Kaz92, Kea96b, Kea80, Kea93a, Kee95, Kee88a, Kee61, Kee87, Kee88b, Kee90, Keh89, Kei64, Kei65, Kei71, Kel60a, Kel67, Kel84b, Kel86a, Kel83, Kel84d, Kel88a, Kel96, Kel75b, Kel86b, Kel90b, Kel92b, Kel86c, Kel90c, Kem67, Kem85, Kem92a, Ken86a, Ken86b, Ken91a, Ken93, Ken74, KS94, Ker91a, Ker92, Kes65, Kes69, Kes75, Kes85, Kev85, Key94, Kha86, Kha87, Kha88, Kha71, Kie94, KLG69, Kin94b, Kin97, Kin76, Kin82, Kin94c]. Review [Kir93a, Kir90a, Kir76, Kla63a, Kla80, Kla83a, Kla84h, Kla84i, Kla84j, Kla84g, Kla86a, Kla88a, KG89, Kla90a, Kla92b, Kla97, Kle71, Kle72, Kle76, Kle75, Klo93, Kni84b, Kni86, Kni94b, Kno81, Kno83, Knu77a, Knu80, KL87, Kob10, Koç89, Koç94, Kod96, Koe95, Kog66, Koh85, Koh86, Koh88, Köl83, Köl87, Kol63, Kol65, Kol76, KF62, KS73, Koo77, Kor65, Kor68, Kor73, Kor95b, Kor95a, Kor85, Kov96, Kow83, Kow86, Kow88, KZ93, Kra96a, Kra68, Kra67, Kre96, Kri85, Kri86, Kri91, Kri82, Kul60, Kul65, Kul69, Kum82, Kum88, Kun93, KTSB21, Kup94, Kup62, Kup75, Kur69, Kus72, Kus74, Kus93, LW64, Lae76, Lae82, Lae83b, Lag89, Lag97, Lah94, Lai91, Lai78, Lak80, Lam86, Lan81, Lan86, LF66, LS66, Lan67b, LG80]. Review [Lar94, Lar74, Lar78, Lat69, Lat90a, Lau74, Law18, Law95, Law65, Law66, Law72, Laz82, Laz91a, Le 92, LeV77a, LeV93, LeV03a, LeV04b, LeV04c, LeV04d, LeV05a, LeV05b, Lea94, Lee77, Lee82a, Lee82b, Lee84a, Lee84b, Lee86a, Leh63a, Leh63b, Leh82, Lei66, Lem88, Lem84, Lem85, Lem83b, Len80, Len95, Len93b, Ler64, Les76, Lev62, Lev67a, Lev63a, Lev88b, Lev96, Lev84, Lev87, Lev63b, Lev77b, LG89, Lew83a, Lew86, Lew91b, Lew91a, Lie97, Lie95, Lig86, Lil67, Lil68, Lin91a, Lin80, Lin90, Lin79c, Lin92, Lit63, Liu88, Liu91, LMDLS23, LM81, Loc84b, Loe95, Loe92a, Log97, Lon87, Lor91, LJW73, Los73, Los74a, Los74b, Los75, Los76b, Los76a, LFR80, Los81, LJ83, Los83a, Los83b, Los84a, Los85, Los86a, Los89, Los95, Lou92, Lu92, Luc87, Luc86, Luc18, Lud77]. Review [Lud82, Lud87, Lud91, Luk75, LSB67, Luk68, Luk82, Lum85, Lus87, LS84, Mac91b, Mac91a, Mac68b, Mac69, Mac70, Mac92, Mac61, Mac76a, Mac98a, Mag86, Mah94, MR71, Mal80, Mal85, Mal87, Mal88a, Mal90, Mal93a, Mal77, Mal88b, Man90a, Man89, Man88a, Man61a, Man61b, Man61d, Man61c, Man81, Man70, Man87, Man88b, Man90c, Man85, Man63, Mar66, Mar85a, Mar64, Mar73, Mar75a, Mar83b, Mar81, Mar86a, Mar88, Mar76b, Mar85c, Mar86b, Mar65, Mar87, Mar95, Mas95, Mas87, Mat89a, Mat80, Mat88, Mat94b, Maw86, Maw91, Maw93a, Maw95, Maw96, May91, May84c, May84d, May90, McA88, McA92, McA96a, McA96b, McA94c, McA94b, McC91a, McC93a, McC92a, McC91b, McC79, McC83, McC85, McC71, McC95, McC86, McG62b, McG67a, McG67b, McG64, McG65a, McG65b]. Review [McG93, McG61, McI71, McK95, MP62, McK81, MN83, ML85, MBCG83a, McL79, McL86b, McM61c, McM87, McN59, McN89a, McN89b, MNMF63, Mea86, Mea88a, Mea88b, MNC71, Mec92, Med78, Mee96, Meg88, Meh73, Meh96, Mei78a, Mei83a, Mei77, Mei89, Mel92b, Men94a, Mer64, Mer81, Mer85, Mer84b, Mey72, Mey75b, Mey76c, Mey77b, Mey78, Mey79, Mey82b, Mey82a, Mey83, Mey84, Mey86b, Mey88, Mey89b, Mey89c, Mey90, Mey91a, Mey91b, Mey93b, Mey94a, Mey95a, Mey95b, Mey96, Mic95b, Mic92b, Mie95, Mik89, Mil69a, Mil82a, Mil83a, Mil80, Mil82b, Mil61a, Mil82c, Mil92, Mil94a, Mil95b, Mil81a, Mil81b, Mil83c, MT61, MT62, MT64, Mil89b, Mil90b, Mil91a, Mil94b, Min66, Mir85, Mir88, Mit65, Mit93, Mit87, Mit94, Mit91b, Miu76a, Mla97, Mla94, Moc85, Mod92, Moi89, Mon60, Moo74b]. Review [Moo74a, Moo75, Moo60, Moo63, Moo69, Moo76, Moo77b, Moo77a, Moo85, MS65, Mor74b, Mor86a, Mor88b, Mor97, Mor93, Mor96, Mor68, Mor71, Mor73, Mor75b, Mor82, Mor83, Mor85b, Mor88d, Mor89c, Mor91a, Mos93, Mot88, Mro91, Muk86, Mul64a, Mul66b, Mul67, Mul69, MR65b, Mur94a, Mur97, MG95, Nag96, Nag88, Nas87a, Nas87b, Nas90, Nas65, Nas77b, Nas86, Nas95, Nau86, Nel74, NPW95, Ner70, Neu94, Neu97b, Nev82, New87, New90, New94a, New76c, NW81, New82a, New74, New76e, NW82, Nic96, Nic65, Nic94, Nie89a, Nit85, Nit90, Noe72, Noe81, Noe82, Noh62, Noh67, Noh71, Noh76, Noh77, Nor90, Not94, Nov94, Núñ96, O’B95, O’M74, O’M86, O’M87b, O’M87a, O’M88b, O’M92, O’M96, O’R92, Obe64a, Obe64b, Obe77, Ock91, Ock92a, Oco87, Ode86, Ode89]. Review [Odg86, Old62, Ole68, Olk59, Olk72a, Olk72b, Olk77, Olk78, Olm77, Olv72, Olv86, Olv90, Olv91a, Olv93, Olv97, Orb18, Ork95, Orl93a, Orl93b, OP88, Ort74, OW89b, OG90, Ort96b, Osb85, Ost61a, Ost61b, Oth82, Ott96, OR67, Ove05, Owe85, Pac86, Pac87, Pac88b, Pac10, Pal84b, Pal95, Pal96, Pal94, Pan86, Pan93a, Pan95, Pao88, Pao90a, Pao93, Pap62a, Pap62b, Pap63a, Pap63b, Par94, Par88, Par80, Par87, Par90, Par66, Par74, Par81a, PT81, Par10, Par69, Par67a, Par72, Par67b, Par76, Par92b, Par92a, Pas95, Pat64, Pat87, Pav94, Pay68, Paz76, Pei83, Pel97, Per89, Per65, Per68, Per76, Per93a, Pet95, Pet73, Pet68, Pfl92, Pic87a, Pic90, Pic92, Pic87b, Pil87, Pim87, Pim93, Pin88, Pin85a, Pin85b, Pin87a, Pin92b, Pin69]. Review [Pin70, Pin72b, Pin72a, Pin77, Pin84b, Pin87b, Pip81, Pip86, Pip92, Pit89, Pit90b, Pit91, Pit93a, Pit74, Pit77, Pit79, Pit81, Ple88, Poh63, Pol59, Pol65, Pol69, Pol76, Pon90a, PRM^{+}66, Por72, Por80, Por83, Por85, Por86a, Por88, Por81, Por86b, Por89, Por92, Pow75, Pow80, Pra68, Pra85, Pra86, Pra89a, Pra89b, Pré95b, Pre81, Pro65a, Pro65b, Pro83b, Pro84b, Pro65c, Pro84c, Pro85, Pro87a, Pro91, Pro95, Pro96, Pug86, Pul64, Pul91, Qua95, Qui92b, Rab94, Rad90b, Rad73, RV78, RCC71, Ral62a, Ral62b, RA63, Ral66a, Ral66b, Ral70, Ral74, Ral80, Ral83b, Ral83a, Ral86, Ral93a, Ram90a, Ram93, Ran89, Ran96a, Ran96b, Ran91a, Rao94, Rat87, Rat92a, Rat93, Rau95, Raz87, Raz92b, Red89, Red90, Red93a, Red83, Reh89, Rei71a, Rei77a, Rei96, Rei62a, Rei62b]. Review [Rei63, Rei69, Rei71c, Rei75, Rei77b, Rei77c, Rei78c, Rei79, Rei83a, Rei87, Rei88, Rei89, Rei64, Ren84b, Ren89, RJ90, Ren90, Ren92b, Ren93c, Ren93d, Ren94a, Ren94b, Ren94d, Ren94c, Ren95, RJ96, Ren97b, Ren80a, RL62, Rhe60, Rhe70, Rhe85, Rhe88, Ric67a, Ric74, Ric88, Ric85b, RO66, RR89, Ric90b, Rie72, Riv60a, Riz94, Rob81, Rob85, Rob95b, Rob94a, Rob62a, Rob62b, Rob74, Rob76, Rob91, Rob92a, Roc92a, Rog60, RB90, Rog94b, Rog96, RFC72, Ron95, Rop83, Ros82, Ros71, Ros63, Ros65, Ros68a, Ros72a, Ros78, Ros73, Ros68b, Ros80a, Ros85, RSS65, Ros72b, Ros87, RD65, Rou72, RAG72, Rou73, Rou91, Rou88, Row64, Row76, Row79, Row86, Rox76, Rox77, Rox84, Rox90a, Rox92, Roy91, Roy93a, Rub69, Rub77a]. Review [Rüd97, Rud61, Rue72c, Rue72a, Rue72d, Rue73a, Rue73b, RSTW73, Rue74a, Rue76b, Rue78, Rue79, RLG81, Rue81a, Rue82b, RAK88, Ruk93b, Run91, Ruo90a, Ruo91b, Ruo92a, Rus95, Rus60, Rus90, Rus85, Rya63, Saa83, Saa73, Saa88, Saa96, Saa68, Sac83b, Sac74, Sae63, Saf96, Sag77, Sag78, Sag81, Sag82, Sag83, Sag86, Sai91a, Sal85, Sal64, Sal66, Sal62, Sal63a, Sal63b, Sal63c, Sal63d, Sam86, Sam76a, Sam76b, Sam79, Sam76c, Sam80, San82, Sán83b, San95, San06a, San06b, San06c, San06d, San07a, San07b, San07d, San07e, San08a, San08b, San09d, San09a, San09b, San09c, San10a, San10b, San10c, San10d, San11a, San11b, San11c, San11d, SS17a, SS17b, SS17c, SS17d, SS18a, SS18b, SS18c, SS18d, SS19a, SS19b, SS19c, SS19d, SS20a, SS20b, SS20c, SS20d]. Review [SS21a, SS21b, SS21c, SS21d, SS22a, SS22b, SS22c, SS22d, Sar84b, Sat78, Sau62, Sau78, Sau79, Sav18, Saw82, Saw85, Say79, Sca59, Sch77a, Sch93a, Sch63a, Sch71a, Sch73a, Sch82a, Sch82b, Sch84b, Sch86a, Sch86b, Sch89a, Sch65a, SNS66, Sch95, Sch89b, Sch80a, Sch87, Sch93c, Sch94, Sch81a, Sch60a, Sch60b, Sch65b, Sch65c, Sch93d, Sch76c, Sch78c, Sch85a, Sch91a, Sch60c, SR93b, Sch81c, Sch85c, Sco83, Sco84b, Sco85, Sco88, Sco90a, Sco90b, Sco91a, Sco91b, Sco92, Sco94, Seb70, Sed88, Seg79, Seg80, Seg85, Seg88, Seg06, Sei62, Sei85, Sei93, Sei86a, SK80, Sel94, Sel76, Sen94, Ser80a, Ser80b, Ser84b, Ser86, Ser94, Ser96, Ser64, Set62a, Set62b, Set63, Sey93b, Sez90, Sha74, Sha82a, Sha83, Sha95a, Sha96, Sha97a]. Review [Sha65, Sha92a, Sha97b, Sha85, Sha95c, SM79, Sha89b, Sha82b, She77b, She63a, She63b, She66b, She87, She91, Shi94, Shi68, Shi69, Shi71, Shi93, Shi97, Sho69, SR79, Sho66a, Sho87, Shu87, Shu80, Shu82, Shu85a, Shu85b, Shu90a, Sib95, Sie69, Sie89, Sie90, SW93a, Sik92, Sil78, Sim89, Sim76, Sim93a, Sim82, Sin59, Sin69, Sin93a, Sin79b, Sin93b, Sit92, Siv94d, Siv94a, SS70, Ske95, Sle74, Sle89a, Sle92, Sle94, Sle79a, Sle81, Sle83a, Sle85, Sle86, Sle87, Sle89b, Sle62, Sle67, Sle84b, Slo93, Slo97, Sma76, Smi74, Smi83, Smi80b, SW81, Smi82, Smi86, Smo91b, Sno76, Som93, Sop59, Sop60, Sop62, Sop65, Sop69, Sop72, Sop79, Sop82, Spa62a, Spä97, Spi84b, Spi87, Spi91a, SLF66, Spi74, Spi80, Spi81a, Spi81b, Spi82, Spo85, Spo88]. Review [Spu89, St.82, Sta66, Sta85, Sta88, Sta89, Sta90, Sta92b, Ste87, Ste67b, Ste68a, Ste71, Ste72, Ste73b, Ste81b, Ste88, Ste90a, Ste93a, SKGM67, Ste68b, Ste70a, Ste93b, Ste94a, Ste97b, Sto96, Sto61, Sto62a, Sto62b, Sto66, Sto81, Str86a, Str89b, SFH^{+}66, Str68, Str70, Str76b, Str82, Str86c, Str62a, Str62c, Str62b, Stu90, Stu91, Stu94b, Sub90, Sül97, Sul93a, Sup77, Sus79, Sus88, Svo96, Swa96, Syl96, Sym92, Syn83, Syn93a, Sys63, Szy87, Ta’90a, Tab95, Tab97, Tad89, Tad93, Tak77, Tak83, Tal81a, Tal81b, Tan93b, Tap84b, Tap86, Tar81, Tau80, Tay68, TT76, Tay86, Tay89, Tay96, Tay97, Tek89, Ter93a, TA90b, Ter92, Tha79, Tha82, Tha96, Thi63a, Thi63b, Thi84b, Thi84a, Tho74, Tho76a, Tho87, Tho60a, Tho59, Tho75, Thr79, Thr94a]. Review [Thy74, Tis83b, TBDS92, Ton95, Ton89, Tot91, Tre84b, Tre99c, Tre99d, Tre99e, Tre99f, Tre00a, Tre00b, Tre01a, Tre01b, Tre01c, Tre01d, Tre02b, Tre02c, Tre02d, Tre02e, Tre13, Tre65, TS65, Tri95, Tru94, Tu92, Tul68, Tul90, Tur60, Tut94, Uhl73a, Uhl73b, Ull88, Ung85, Urb93a, Uri92, Vai96, Vak93, Van86, Van96, Van88, Van66, Van68, Van76, VJ83, Van63, Van93, Van94c, Var87, Var85, Var66, Var67, VK62, Vel71, VN73, VBdD91, Via85, Vog87, Vog91a, Vog93, VDC22, Wag92, Wah92, Wal95, Wal87, Wal79, Wal88, Wal93a, Wal97, Wal63a, Wal93c, Wal75b, Wan91, WP88, War92, War64, Was94a, Was74, Was94b, Wat91b, WJ94, Wat63, Wat65, Wat86a, Wat10, WS71, Wat86b, Wat90, Wax91a, Web85, Wei80b, Wei84b, Wei88, Wei87]. Review [Wei91, Wei69, Wei77, Wei79, Wei67b, Wei67a, Wei90b, Wel88, Wel89, Wen68, Wen78, Wer89, Wer63, Wes92, Wes84b, Wet74a, Wet75, Wet93, Whe74, Whe76, Whi93, Whi63a, Whi63b, Whi66a, Whi66b, Whi67, Whi69, Whi76a, Whi76b, Whi92a, Whi80, Whi81, Whi83, Whi84b, Whi88, Whi94, Wib87, Wic94a, Wic94b, Wie87, Wig85, Wig90, Wig91, Wil62a, Wil62b, Wil62c, Wil67a, Wil70a, Wil84b, Wil85a, Wil86a, Wil88, Wil66, WLP^{+}61, Wil65a, Wil69, Wil61, WJV79, Wil91a, Wil60a, Wil60b, Wil64c, Wil65b, Wil72, Wil73, Wil77b, Wil78b, Wil79, Wil81, Wil86b, Wil93, Wil77c, Wil82b, Wil85c, WM85, WBGP91, Wil92, WR94b, Wim65, Wim67, Wim72, Wim75, Wim80, Wim81, Wim86, Wim88, Wim91b, Wim96, Win74, Win95, Win60, Win62, Wis60, Wit90, Wol86, Wol97, Wol72a]. Review [Wol94a, Wol93, Won83b, WGM93, Woo92, WA63, Wou64a, Wou64b, Wou67, Wou68a, Wou68b, Woź87, Wri60, Xan87, XL94, Xu96a, Xu96b, Yak94, Yak95, Yan77, Yan91, Yan92, Yan94, Yan65, Yan76, Yao86, Yat84b, Yau68, Yau69, Ye94, Yon95, YY96, You72a, You71, You72b, You74, Zad67, Zah80, Zak84b, Zak85, Zal83, Zar65, Zas65, Zei72, Zem96, ZN62, Zha96, Zha97, Zie76, Zie78, Zie83c, Zie83b, Zie84b, Zim90, Zit63, Zlo77, Zlo83, Zlo85, Zlo91, Zuo96, Zwe84b, dB65b, dM94, vH87, vK79, vL86b, vL86a, vLMZ19, vV65, vWHE91, vdM68, vdV94, Bel10, Bor09, Car08, Cia92a, Cor16, LeV03b, LeV03c, LeV03d, LeV05c, San07c, San08c, San08d, Tre99a, Tre99b, Tre00c, Tre00d, Tre02a, Tre15, Ver09, Ano96a, Ano97b, Thi84b]. Reviewer [Nas92]. Reviews [Ano86k, Ano97a, Ano98a, Ano99a, EMUP98, Kel97a, Kel97b, Kel97c, Kel97d, Kel98a, Kel98b, Ke98, Kel99d, Kel99a, Kel99b, Kel99c, Kel00, Lay24a, Lay24b, O’M00a, O’M00b, O’M00c, O’M01a, O’M01b, O’M01c, O’M01d, O’M02a, O’M02b, O’M02c, O’M02d, O’M03a, O’M03b, O’M03c, O’M03d, O’M04a, O’M04b, O’M04c, O’M04d, O’M05d, O’M05a, O’M05b, O’M05c, O’M06a, O’M06b, O’M06c, O’M06d, O’M07a, O’M07b, O’M07d, O’M07e, O’M08a, O’M08b, O’M08c, O’M08d, O’M09a, O’M09b, O’M09c, O’M09d, O’M10a, O’M10b, O’M10c, O’M10d, O’M11a, O’M11b, O’M11c, O’M11d, O’M12a, O’M12b, O’M12c, O’M12d, O’M13a, O’M13b, O’M13c, O’M13d, O’M14a, O’M14b, O’M14c, O’M14d, Sch18c, Sch18a, Sch18b, Sch18d, Sch19a, Sch19b, Sch19c, Sch19d, Sch20a, Sch20b, Sch20c, Sch20d, Sch21a, Sch21b, Sch21c, Sch21d, Sch22a, Sch22b]. Reviews [Sch22c, Sch22d, Sch23d, Sch23a, Sch23b, Sch23c, Wat15a, Wat15b, Wat15c, Wat15d, Wat16a, Wat16b, Wat16c, Wat16d, Wat17a, Wat17b, Wat17c, Wat17d]. Revised [Eas89, Koç89]. Revisited [Ben79, BSS05, Che98, DD02, Hig09, Man09, Mil93, Mil98, O’B11, Pro87b, RPFM17, Sha90, Sim75, Ste01, Wat08, Sim76]. Revisiting [UA07]. Reviving [BT05]. Revolution [BJ75]. Revuz [Pit93a, Pit93b]. Reweighted [ST91b, ST91a]. Reyman [Gui92, Tu92]. Reynolds [LR00, Sun76]. Rheinboldt [Ral74]. Rheometry [Fri78a]. RI [Lax89]. Ribaric [Sch85a]. Ribenboim [Cas95]. Ricardo [RT97]. Riccati [Die92, Dis92, Meh96, Dar97, LRM96, Lei76, MCS69, Smi87]. Rice [Fri71, Law72, Mil82c]. Richard [Abl77a, Bec85, Boe88a, Car70a, Car65a, Cha64a, Che92c, Coh82, Cro73, Dor74, Dwa66, Eis86a, Fix93, Fle64, Flo69, Fra86a, Fre79, Gas71, Gen76, Goo82, GdMGR91, Har67d, Jag74b, JRR71, Jel87, Kar78, Kaw65, Kaz89, Kee88a, Kel92b, Kha86, MN83, Mor74b, Ren94b, Ros65, Syl96, Wei79, Wet75, Whi88, Wil60a, Wil85c, Woo92, vL86a, CGP21, Che92b, Fir97, Gib97, Han86a, Kel92a]. Richards [Guh68, RR89, Rob85, RAK88]. Richter [Dix91]. Richtmyer [Wen68]. Ride [Sug98]. Ridge [CR78, GHW76]. Ridgway [Pas95]. Riedle [Bau88]. Riemann [Ame91a, Ear81, Sib95, Ame91b, ABS95, BK99, DL66, Dra83, Ral65a, Sha04, Van98, Wig90, Coh75b]. Riemannian [Nie84, Wol88, Mag86]. Riemenschneider [Ron95]. Rienstra [Jor88b]. Right [Amo64, Ant72, Deu95, MS65, Tri62a, Rei66]. Righthand [Ber90b]. Rightsizing [TH96]. Rigiano [Chi95]. Rigid [CL08, Eri66, Han89, HC16, Poh63, Sim71a, Ste00]. Rigid-Body [Ste00]. Rigorous [Hag85b, KH98]. Riley [Ste70a, Str70]. Ring [CN05, GO09, SSW99]. Ringel [Bei96, May91]. Ringhofer [Gar91]. Rings [Wie87]. Rionero [Ant87, Hor97a]. Riordan [Kul60]. Risk [AMS^{+}23, BCH14, BK83, HS15, Mor75a, PdHM16, Ruo91a, Ruo91b, Sin88, TL02, Nas87b]. Risk-Aware [BCH14]. Risk-Neutral [HS15]. Risks [Ruo92a, Ruo92b]. Ritz [Ber75, GW12, Koh72, Mor64]. Rivers [DC76]. Rivlin [Hol70a]. Rivoirard [Cre97]. Rizzi [Ork95]. Roach [Cou97]. Roache [Mur97]. Road [MMT^{+}21]. Roast [Kla61]. Roberson [Ant90, Str62b]. Robert [Abl80, Alm87, Are93, Bai76, Ber78, Boa81, But74a, But74b, But74c, Car81, Car87, Col89, Eng78, EB88, Fan59, Far94, Far65, Fet65, Fia73, Fle84b, Gil95, Gri88, Gut88, Han86b, Har76b, Hol87, Isa89b, Jag72d, Kan68, Kel83, Ken86b, Ker92, Kin94b, Kin76, Kla84i, Koç94, Kri91, Leh82, Man61b, McG67b, Min66, Mor89c, Mos93, New94a, Pro85, Ran89, Rat93, Rog60, Rue72d, Sal64, Shi94, Siv94d, Siv94b, Stu90, Tal81a, Tul90, Was94b, Wei88, Wel89, Wen68, Wet74a, Zie83c, Kin94a]. Robertazzi [Beu92]. Roberts [Bai74b, Kla84i, Kla84b, Kle76, Luc86, Thr79]. Robinson [Kaw63, NW81, Wal95, Jef94]. Robot [Lon87, Nag88]. Robotic [DGS05, GDS05]. Robotics [Sco92, Gra86a]. Robust [BBC11, EK13, GGT13, KS94, MMB^{+}07b, PdHM16, SXY19, Ste99, TD20, MMB^{+}07a, Rei83a]. Roche [Cam91b, Cam91c]. Rock [BCD98b, BCD98a]. Rockafellar [BI70b, Ber85f, Kle71]. Rocket [Kla87, Kla88a, KDJM88, Pru00]. Rockets [Zar65]. Rockey [Mal85]. Rod [FT11, Hea91, LS96]. Rod-Stirring [FT11]. Roderick [Str89b]. Rodica [Isr92b]. Rodin [FW82, She91]. Rodino [Lie95]. Rodman [Bar83b, Bar83a, Joh88, Meh96]. Rodrigues [Ren94d]. Rods [AR78, Kre64, AR80]. Roediger [BW96b, Jag87a]. Roel [Hag85c]. Roger [Cla78c, Fab87, Far96, McA92, Mor91a, Nau86, Ple88, Ste93b, vL86b, Mor91b]. Rogers [And83, Ber91b, Blu91, Bor79, Bre81, Cox65, Jor84b, Kor95a, Lee84a, Sho69, And74b]. Rogers-Ramanujan [Bre81, And83]. Rokne [AKP91a]. Roland [Law66, Sch89b, Spi84b]. Role [AMS^{+}23, Dre60, Gor75, Jaf84, Kac62, KS77, Lev90b, Mey75c, MWL97, PCLS23, EV91]. Rolewicz [Noh77]. Rolf [Ask72, O’B95]. Rolf-Dieter [O’B95]. Roller [Fly89b, Fly89a, Lew88]. Roller-Coaster [Fly89b, Fly89a, Lew88]. Romanovskii [Erm89]. Romesh [Zha96]. Ron [Bou65a]. Ronald [Aga95b, Ass77, Bla88b, Bro76a, Mea88a, Mey83, Noe81, Wel88, Zie78, Zim90]. Ronveaux [Bre89]. Rooij [Boa83]. Roop [Reh89]. Root [BS63, Coo71, Cor77, Far60, Gla93a, GS22, HP76, HT12, Kra70, PH76, SG76a, Tri95, WR94a, YT95, WR94b, Cas72a]. Root-Finding [PH76]. Rootfinders [Boy13]. Rootfinding [CS20]. Roots [CP89, CPRG90, EL85, EL86, FL84, GKP75, Jag87b, JR87, Jor86d, Lan64, Raw86, RGJ90, Rog83a, Sma76, Sof96, Van94a, Van94d, Wil24c, Wou66, Sof97]. Rope [VK84, Grc18]. Ropes [Yon06]. Rosakhoo [Gri88]. Rose [Par74]. Roseau [Sha89b]. Rosen [AK90, But74a, But74b, But74c, CL62, Jor77a]. Rosenbaum [Hei96].

Rosenblatt [Bel62, Blu65, Pic87b]. Rosinger [Kel84d]. Ross [Bor73, Bur81a, Col92, Hea81, Jag76a, Kar84b, Kea80, Luc86, McA96b, Por92, Kar84a]. Rössler [Abr94]. Rota [Ant63, Bru72]. Rotating [Kel90a, Wil78c, Xan87]. Rotation [GAK17, Stu64]. Rotational [Iva04]. Rotationally [Dru64]. Rotations [Bre63, Mar21, May60, Kat65]. Roth [Shu82, Wei80b]. Rothe [Fif86, Pap62b]. Rothman [Byr65, Fet68b]. Roubens [Wal88]. Rough [Owh17]. Round [Lon78]. Round-Off [Lon78]. Rounding [For59, Gra66b]. Rounding-Off [For59]. Roush [Pac88b]. Rousseau [Bus86, Jes93b]. Routes [Zit76]. Routh [Hou68b, Jor86a, Zuo96, BS77, Hou68a, SK85]. Routine [Lyn83]. Routing [Ols03]. Row [Cen81, WKW75]. Row-Action [Cen81]. Rowland [Kra68]. Rowlands [Maw93a, Maw93b]. Roy [Kra67, Olk59]. Royce [Zei72]. Royden [Wat65]. Rozanov [BR79, Por89]. Roze [Kan66]. Rozonoer [Har72b]. Rozvany [Mro91]. RSA [Bla88f]. Rubenfeld [wBJ77]. Rubik [RKDD14]. Rubin [Str89b]. Rubinstein [Gue95, Jer91a, Mey78, Jer91b]. Rubio [Bal88a]. Ruckle [Alp85]. Rüde [McC95]. Rudeanu [BI70a, Whe76]. Rudiger [Bry90a, Mit93, Orl93b, Bry90b]. Rudolf [Kap70b, WJ94]. Rudolph [Bar67]. Rueh [Sei86a]. Ruehr [BdD92, BS86, Bus95, Gas85b, Gla88a, Kir76, Knu80, LM81, New90, Rou73, Sha83, Wil82b]. Ruel [Obe64a, Obe64b]. Ruelle [Kaz90a, Len80, Mac91b, Kaz90b, Mac91d]. Rugby [Fre14]. Rugh [Mor83]. Ruin [Jev94, Kij93, Tro71]. Rule [Bou21, CN15, Fer97b, For21, Hig02c, Kre91, TW14, Fer97a]. Rule-Based [CN15]. Rules [Ben95, BI18, Gar75, Lam04, McA87, Mon82, Ser04, Spi81a, Zlo85]. Rumnov [Nas77b]. Runge [Cam91b, Enr89, BH76b, Bet78, Cam91c, Cas75, CS20, Hor78, LT98, LK65, Lut66, Ros67b, SS16, Sku04]. Running [GAHK05, Mat89b, Pri93]. Russak [Bla81]. Russel [TS86]. Russell [Isa89b, Sag81]. Russels [Lee86b]. Russian [How94, Ani82, How94]. Russo [Rau95]. Rustagi [Olk78]. Ruth [Gue79]. Rutherford [Coh76]. Rutickii [Rya63]. Ryabenki [Aro69]. Ryabov [Dri68]. Rybchinskaya [May90]. Ryuzo [Joh84b]. Ryzhov [Ock91].

S [Aga93, Aga94, Ale90, And77, Ano87a, Aro69, Asn75, ABG96, Atk66, Bai74b, Ber95, Ber91b, Ber81, Ber88a, Bes92, Bjö61, Bla85b, BS86, Bou91b, Bra63, Bro90, Bru95, Bru86, Buc69, Bur94a, Bus95, Car72a, Cha70a, Chi90, Chu81, Cla83b, CZ89, Col92, Coo91, CK86, Cor92a, Cra89, Cur77, Dan71, Dav75a, DG96, DeW85, Deu81c, Dev90, Die92, Dij90, Dis92, Dol72, Dri75, Eis86a, Elk86, Els91, Eps59b, Erm89, Fab80, Fad62, Fal73, Far69, Fel90, Fie93, Fix93, Fle73, Fle90, Flo69, Fly89b, For96, FARE69, Fox71, Fra86b, Fra72, Fri59, Fro73, Fro80, Gal95, Gar93, Gen96, GA63, Gla86b, Gla93b, Gre60a, Gre76, Gre72, Gri61, Gro96e, Gue84, GWK95, Gur66, Gut97, Haa93, Ham77, Har70, Har72a, Har82, HBO66, Her97, Her65, Hey82, Hig62d]. S [Hig92, Hig00, Hin89, Hir71, Hoc92, Hop79, How83, Hul61b, Jag72c, JBL72, Joh67, Joh81, Joh83c, JF71, Jor86a, Jor88b, JGM95, Kv93, KH85, Kap85b, Kap71, Kar60, Kar84a, Kaz75, Kel60a, Kel84b, Kla80, Kla83a, Kla84i, Kla84j, Kla84b, Kla84c, Kla86a, Kla88a, KG89, Kno83, Knu77a, KF62, Kor65, Lah94, Ler64, Lev96, Lin79c, Los74b, Los75, Luc86, Lud82, LSB67, Mac61, Man87, Mas95, Mat89a, Maw91, May90, Mei78a, Mey90, Mey91b, Mil95b, Mil94b, Mla97, Mon60, MW97, Mor93, Nas65, Noh62, O’M92, Ole68, Olk59, Olk78, Olv90, Ost61b, Par87, PT81, Par92a, Pel97, Pip81, Por86a, Qui92b, Rab94, Ral62a, RA63, Ram93, Ren95, RJ96, Rob91, Ron95, Ros72b, Rou72, Rue72a, Rue73a, Rue81a, Rus95, Sam86, Sau78, Sch60b, Sco88, Sco90a, Sco90b, Sei62]. S [SK80, Set63, Sha65, Sle84a, Sop62, Sta66, SFH^{+}66, Sup77, Ta’90a, Tal81a, Tho87, Thr79, Ton89, Tri95, Uri92, Van93, Var66, VK62, Wal87, Wal79, Wan91, Wat91b, Wer89, Whi80, Whi83, Wie87, WJV79, Win60, Wol93, Wou68a, Xu96b, Yak95, Yan94, Zas65, Zie84b, Zie84a, Zuo96, Gar94b, Ben95]. S- [MW97]. S-E [CK86]. S. [Ber84a, Cla83a, Joh83a, Kel84a, Yat84a]. SA [BFM^{+}05]. Saad [Cul94b, Cul94a]. Saaty [Bei80, Kei65, Lit63, Min66, Ral66a, TE96]. Sabatier [Roy91]. Sabharwal [BS80, Gut82, KLG69, Los85]. Sach [Mil83b, Mil83c]. Sachdev [Ame88]. Sachs [Bei81]. Saddle [Bis15, PW15]. Saddlepoint [JR96, Rei96]. Saeks [Kha86]. Safe [PGH^{+}88, Sal87, Pin88]. Saffman [Caf94]. Sagan [Gur62a]. Sagdeev [Bry93]. Sage [You72b, Bee09]. Saha [Mor85b]. Sahney [Chu81]. Saigal [Zha96]. Saint [Dur87, Hor89]. Saint-Flour [Dur87]. Saint-Venant [Hor89]. Sakamoto [Sar84b, Sar84a]. Sakarovitch [Ham87]. Salesman [BL84, BDvD^{+}98, Gao18, PR91, Pat03]. Salisbury [Lon87]. Salkauskas [Pra89b]. Salmon [Pin88]. Salomaa [Har75]. Salukvadze [Sch81a]. Salvatore [Ant87, Hor97a]. Salzer [Str62a, Str62c, Str62b]. Samarskii [Lev96]. Same [Reu67]. Samelson [Car66c]. Sammakia [Reh89]. Sammet [Ros71]. Samorodnitsky [Pel97]. Sample [Ang92a, Cha94c]. Sampled [Lev90a, Lev91]. Sampler [BRP88]. Samples [Beu66, PTK11, SM94, Har63]. Sampling [AG01, BS92, CW94, Flu90, Gar00, Sin93b, Sin93c, VDC22, Whi92b, Wit90, Whi92a]. Sampson [Lai78]. Samuel [Art85a, Kar83, Kaw67, Man61d, Man81, McN59, Ric67a, Sch78c, Wei77]. Samuil [Ear81]. Samulon [Lin59]. San [BC96b, Bra82]. Sanchez [Ede90a, Ede90b, Kap72, Lan86]. Sanchez-Palencia [Lan86]. Sanders [O’M87b]. Sanderson [Zie83b]. Sandra [Chi95]. Sands [dM94]. Sanford [Bai74b]. Sangren [Hul61a]. Sanjo [Eck82]. Sansone [Fin76b, Noh67]. Sanz [Ske95]. Sanz-Serna [Ske95]. Saran [Car72a]. Sasieni [Sop60]. Satellite [Ano61f, Cle61, FSW^{+}66, Gre63b, SS65, Sil61, Sin61, Wah65]. Satisficing [YAKM20]. Satisfy [Kra80b, Kra80a, Urq68]. Satisfying [Wim67]. Sato [Joh84b, Joh84a]. Sattinger [Gru88, War92]. Sauer [Bru83]. Saul [Isa88, Isa89a, Por86a]. Saul’yev [Per65]. Saupe [Dev92, Goo91a]. Savage [Naa01]. Savageau [But78]. Savant [Gas63]. Save [Mas87]. Saville [Cor92b]. Savings [EH78, Elh78]. Sawaragi [Aok70, Cra86]. Saxena [Ask80]. Sburlan [Hes83]. Scalability [SKP19]. Scalable [BKST15, VBVM22]. Scalar [DS73]. Scale [AB69, BCN18, Buz78, CG24, GMS05, Gri72, PR91, PL78, SB98a, SBR08, Sch93d, Sez90, Som78, Tur78, vLMZ19, DGH^{+}18]. Scales [AN83, CC96, DGK^{+}12, Edw00, Hel75, LL75, Pea81, Wol77]. Scaling [BC98, FW62, Hig09, Mei19, MP06, Rub78, Seg72, Ske78, ST91b, TE96, ST91a]. Scattered [AKP^{+}91b, SPR76, WM61a, WM61b]. Scattering [BC76, BK87, Col84, CCM00, CCM01, CK18, Cou97, DK89, Flu78, Hag89a, KK75, Kri76, MS96b, NY90, Roy91, Sco83, Shi97, Sle94, Syl96, Wil84b, Wil84a, Dei92, Mil80, Hag85b, Hag85a, Tay89]. Scattering-Theory [Wil84a]. Scavenger [PH13]. Schachter [Hun84]. Schaefer [Ken74]. Schaeffer [Caf92, Mey86b, Mey89c]. Schaft [Blo89, Blo91]. Schaible [Bor90]. Scharf [Wig91]. Schechter [Bro76b, Sle79a, Sop72].

Scheck [Ant92a, Ant92b]. Scheduling [DGS05, GDS05, TK82]. Scheffé [Dwy63]. Scheinok [Nic65, RO66]. Scheme [BP76b, FW66, Max60, NGTB15, Kar01, KK99, LKGK07, Sch80c, Sch81c]. Schemes [Aro69, CFHP10, CL75, DJLQ21, FL18, Gev82, HLvL83, HM78, LT98, PvySF23, SS16, SXY19, Shu09, Stu89, TD20, Tre82, Wen61, WS23, Arn92]. Schempp [Mer84b, Mer84a]. Schikhof [Boa83]. Schilders [Joh83d, Joh83b]. Schinzel [Haj96, SAH96]. Schlitt [Hig62d]. Schmeiser [Gar91]. Schmidt [Bos73a, CMR73, Hey85, Isr92b, Nic96, Olk72a, Gre97]. Schnabel [Han86b]. Schoenberg [Blu78, Fin76a, Sch78c]. Schoenstadt [Bla81, Wei87]. Schools [Kla84a, Kla84h]. Schoultz [Str62c]. Schrage [McC85]. Schrefler [Sch89b]. Schreiber [Ott96, Rob92a, Rob92b]. Schreifler [Pra85]. Schriever [Noe82]. Schrijver [Bla88d, Bla89a]. Schröder [Bur94a, Hou74c]. Schrödinger [Aeb94a, Fig93a, Fig93b, Hel97, Sch89a, CGP21, MBK08, Sch84b, Sch84a, YF78, Aeb94b]. Schrödinger-Equation [Sch84a]. Schroeder [Ama82, Ame84]. Schucker [Gue89]. Schulman [Cla83b, Kel84b, Cla83a, Kel84a]. Schultz [Cab74]. Schulz [Hil76, Urb93a, Urb93b]. Schulz-DuBois [Hil76]. Schumaker [Ahl82, Wil93]. Schur [BCF95, LM00, Fra95a]. Schurz [Pro96]. Schuss [Fle82]. Schuster [Koç89]. Schwartz [Hey97, Kor68, OW95b]. Schwarz [Mal90, DEP01, EV91, GZ19]. Schwefel [Han83b, Han83a]. Schweitzer [Amo78a, Ber91d, DG68, Jag71, Lew86]. Schwertassek [Ant90]. Science [Ber83b, Blu91, Bur78a, Enr87, Goo91a, Gos83b, Gue88, Ham61, Ker92, Maw93a, Maw93b, Mon67, Par92a, Ral80, RWMD18, SIA01, Sal63b, Sal63c, Ste98, Tul90, TPCC^{+}11, YL03, Zie84a, PAB^{+}01, Gra18, Ham79b, Ral86, Rei62b, Sal66, Zie84b]. Sciences [Bra94a, Car87, Col94d, Coz87, EK91a, EK91b, Gor63, Ham70, Man61d, Pac87, Rap73, Bra94b, Car81, Col94c, Coz85, Dan66, Fre86c, Kap76, Kor68, McG67b, Mil94b, She91]. Sciences-Applications [Car87]. Scientific [Buz78, Don85, Fix75, Isa88, Isa89a, Jon92c, Log97, Rad73, Tho87, Yan94, Fix93, Kri86, Sch94]. Scientists [Are89, Hig01b, Hig06, Hol87, Mur94b, Klo93, Mur94a]. Scott [Bar94, Pas95, GN22]. Scrambled [OR21]. Screening [Sto81]. Scroll [GH01, KT92]. sd. [Wei77]. SDE [Pro96, KPSP96]. Seaborn [Ask92a, Ask92b]. Sealander [Sae63]. Search [BW76, Bel63c, Ber85c, Can91b, Fra65a, Fra65b, Fra66a, Gia63, Hul67, KH98, KLT03, KLT04, Kro66, New67a, Rab68b, RVBE21, Ste94a, Ste94b, Won64a, Arm90a, Bec85, Can91a, Fra86a, Koo77, AA82, Sto81]. Searching [BGH^{+}04, DM07, Pet91]. Seasons [Ros67b]. Seating [FS62, Fri64]. Sebastião [Ste70b]. Seberry [Mil81b]. Secant [DS79]. Second [Boa83, Car77, Cha67b, Dar97, DeM65, Dem90, FGN13, FW82, Giv78b, Gol65a, GL98, HK92b, Her65, Hin94, Ike72, Jen75, Joh87, Kar83, KK98, Kem67, Koç89, Koh86, LW64, Lam86, Ler64, Lev90a, Lev91, Liu88, Mac76b, Mar86b, Mey75b, NS24, Nit90, Obe64a, Orb18, Pac10, Pao93, Pea65, Por86b, Rei87, Rod81, Sch77a, SS62, SR98, Sle86, Str86c, Str62b, Wat10, Wes92, Won64b, Won68, ZZ18, Zas65, Ver09]. Second-Order [Dar97, GL98, HK92b, Her65, Jen75, KK98, Mey75b, SS62, ZZ18]. Secondary [BKR75, DL76, PMR75, PMR76]. Secret [VdLBF17]. Secrets [Kob10]. Section [Ano71l, Ano72k, Ano73d, Ano74i, Ano76n, Mel64, Mor76]. Sections [Hir71]. Sector [KA76]. Secure [Bla84, Bla86a]. Security [Bla88f]. Sedov [Hol66, Pra68]. see [Lax89].

Seebass [Abl77a]. Segel [Chi83, Gla86a, Hoc79b, Kap76, Mil89b, Sch80a, O’B11]. Segmentation [MSS96, Sha96]. Segmenting [MYDF08]. Segur [Sco83]. Seidel [Bag95, Koh98a]. Seiffert [Jag92]. Seinfeld [Wil72]. Seismic [She87]. Seismology [BB83b]. Sekiguchi [Los83a]. Selberherr [Moc85]. Selected [Ano93k, Ano93l, Ano94h, Ano94i, Ano94j, Ano94k, Ano95r, Ano95s, Ano95t, Ano95u, Ano96s, Ano96t, Ano96u, Ano96v, Ano97l, Ano97m, Ano97n, Ano98e, Boj97, CGO07, Gue61, KB59, Sus79, TB76, Ung85, Hig99, Whi66b]. Selecting [Gup82]. Selection [Ber92c, Hal90b, LMDLS23, Mic92a, Mic92b, Noe82, Wit75, Hal90a]. Self [Ame81, Bur94b, Bur94c, CLG14, CHT21, Erm89, Goo91b, HK92a, HK92b, Kol91, Kol92, Lin69, NS24, Por64, Hun85, Oth82, Rei64]. Self-Absorbed [Por64]. Self-Adjoint [CHT21, HK92a, NS24, HK92b]. Self-Avoiding [Bur94b, Bur94c]. Self-Circumference [Kol91]. Self-Gravitating [Lin69]. Self-Organization [Erm89, Hun85, Oth82]. Self-Organizing [Rei64]. Self-Propelled [CLG14]. Self-Similarity [Ame81]. Self-Tuning [Goo91b]. Sellers [Whi81]. Seltzer [Nor90]. Semantics [GM87]. Semi [Aro78b, CZ96a, HK93, Kor76, OH76, Pro83a, VBU97a, CZ97, VBU97b, Kus93, Mag86, Pro83b, Sch89a]. Semi-Classical [Pro83a, Pro83b, Sch89a]. Semi-Infinite [CZ96a, HK93, Kor76, OH76, VBU97a, CZ97, VBU97b]. Semi-Linear [Aro78b]. Semi-martingales [Kus93]. Semi-Riemannian [Mag86]. Semialgebraic [HLS09]. Semiclassical [Kel85]. Semiconductor [MPRW02, Moc85, Pim87, Gar91]. Semiconductors [Pim93]. Semicontinuity [BK17]. Semidefinite [BTKNZ00, Pat19, VB96, VBC07, YZZ04, dEJL07]. Semidiscrete [LT98]. Semidual [CP18].

Semigroup [Neu94]. Semigroups [Ari92, Gol93, Her78b, Jer81, Sle76, Mar87, Web85]. Semilinear [DJLQ21, Lio82, Sig67]. Seminar [Kla84f, Kla84g]. Seminumerical [Mac70]. Semyon [Che18]. Sen [Ber92b, Cha94c, Far94]. Sendov [Tot91]. Senechal [Sch93a, Sch93b]. Sengupta [Fro74]. Senn [Wat90]. Sens [Man61c]. Sense [Ham74, OW95b]. Sensible [Ben95]. Sensing [BCT11, BL13, MS91]. Sensitive [Kra96a, PTK96]. Sensitivitätsanalysen [Gas70]. Sensitivity [Ath76, Bal20, CCA08, JS76, Koh88, Lem85, Nie89a]. Sensor [Gro96h, Pré95b, Pre95a, ZW93, ZW94, Gro96d]. Sensors [BM03, Sle89b]. Separable [RW80]. Separate [GS04]. Separates [MC91]. Separation [FP04, FS12, Pet98]. Separatrix [BP94]. September [LW64]. Sequeira [Ren94d]. Sequence [BRZS18, Bro64, CM83, ER79, Gav82, GS88, KLG69, KCI^{+}80, Kru83, Los85, Mak86, Nar73, Sab68, SM84, Ser78, Van75, Bre74, ELRS66, Jor89a, Jor93b, Knu80, Luk82]. Sequences [BPV97b, BPV97a, Blo94, Bry96, Car75a, FM68, FLNU18, Gok70, Haj96, Ham69, Inn69, LCE76, Ng96, Owe06, PB96, Pic87b, SW79, SAH96, SPBC96, SFS87, Sta92a, Vei67, Boe98b, SFS88, BC96b, LeV77a, Los95, Moo74a, Sch93c, CS70]. Sequential [Che91, Jac97, KM12, Man70, Mit93, Mul69, NT78, Noe82, Orl93b, Yau68, Gho87, Mal77, Ros68a, Sim82]. Sergey [Vai96]. Sergio [Sha96, Wil86b]. Sergiu [BI70a, Whe76]. Serieì [Wil85b]. Series [AA78, AW76, Als86, And83, AC76, Blu65, Boa70, Boy05, BDC98, Car80, CFL^{+}80, CC78, COW96, Chu66, Chu67, FLR81a, FL68, FL70, Fin95, Fre86b, GBC^{+}90, Gil62b, Gla77a, GJ80, Gla81b, Gla82b, GF86, Gla94b, Gla95f, Gol73a, GRS97, Gro96f, Han64, Han85c, Hau75, Hau76, Hin87, Hoc67, Ism76, JSF95, Jor80c, Jor80b, Jor82, Jor86c, JF87, Kel79b, Kim65, KR78, KWP95, Lan76b, Lei78b, Lev90a, Lev91, Lin79a, Los86b, Mac91b, Mac91d, Mal70, Mal71, Obe64b, OW95a, OW95b, Par92b, Per79, Pro95, Rue76b, Rue76a, Sco59, Sha71b, Sha72a, Sha89a, Sha82b, SG82, Ski95, Sri75, Sri85, Sta92b, Sta92c, Van75, Wal70, WG76, WG77a, WG77b, Wil77b, Wim91a, Yak94, dB64, dB65a, BY98, Cha96b, KK99, Len93a, ASF86, Boa65, Boe95b, BGL^{+}67, FLR^{+}81b, Feu76]. Series [Gib90, JGF87, Jam59, Jor84b, Jor86b, Jor87a, Par76, Van76, Wim91b, Gut88, Los86a]. Serious [Kri85]. Serna [Ske95]. Serve [Mar89]. Server [Beu83b, Beu83a, Kei71, TH96, vLMZ19]. Servers [Hey86, New87]. Service [Hey86, Kri86, New87]. Services [Tuc73]. Servo [Mil61b]. Seshadri [Blu86, Nag96]. Seshu [Mac61]. Session [Tuk67]. Set [And73, Ano92b, Ask86b, AD87, Ast70, BP76a, Bar73, Ben70, Bol91a, Bol91c, BT20, Car68b, Car69, Car72b, CL62, CL87, Fra66b, Hog73, Hur65, Mar93, Mar94, MNMF63, Mic92a, Mic92b, Mul64b, Mur92, PZ64, Rah87b, RCC71, Ril68, Ros61, Smi80a, Smi81, Ste70a, SFH^{+}66, Str70, Swa65, Wei60a, Are93, Ren94a]. Set-Valued [BT20]. Sets [Arb77, Aro78b, CH66, EW75, FGP^{+}22, Fis93, Gog85, HLS09, Hoc70, Jev95b, Kra80b, Kra80a, McA88, Mil81a, Moo78, Par74, RRW07, SW79, Ula64, Ang90, SW80, McM87]. Seung [Gla89b]. Seung-Jin [Gla89b]. Sevast’yanov [Par80]. Seven [Pon67]. Several [BCF95, Fra95a, GG80, Mit65, Sle85]. Sewell [Gup89, Red90]. Sextuple [Ano93a, Gro93c]. Seydel [Bry90a, Bry90b]. Seymour [Boo79, Gro94a, Siv94a, Siv94c]. Shabana [Vak93]. Shader [Gib97]. Shadwick [Lee84a]. Shafer [Fet65, Gib90, Rue72d]. Shaidurov [Xu96a]. Shaked [Ton95]. Shale [Geo76]. Shallit [GdM95]. Shallow [Arn89, Kel88a].

Shallow-Water [Arn89]. Shamita [Run91]. Shamos [Fra88]. Shampine [Gla95a, Ral70]. Shan [BPK84]. Shanks [BRZS18]. Shannon [CW94]. Shantaram [Jag79b]. Shanthikumar [Ton95]. Shape [Fly89a, Kel88c, Law66, Lew88, Luc10, McL86b, Pro87b, Sil65, Str86a, Fly89b]. Shapes [CMZ94, Lar69b]. Shapiro [Coh84b, LeV93, Coh84a]. Sharir [Haj96]. Sharma [Deu81c, Jor86a, Zuo96]. Sharp [BCT11, Dum65, Fri63, Zad67]. Shasha [Hal95]. Shashkov [Zha97]. Shatalov [Daa97]. Shaughan [Gas95]. Shaw [Ral86]. Shear [Cor79, Fel92, Hor95a, Hue91, Mal93a, Mal93b, MB76, PL78]. Shee [Mey95a]. Shee-Nee [Mey95a]. Sheet [HS03]. Sheets [MK99]. Sheldon [Kar84b, Por92]. Shelemyahu [Ros72a]. Shell [Wil59, Sim89]. Shells [BJ75, Cla62, Cla64b, Dru64, Hol89b]. Shen [Sav18]. Shenton [McC91a]. Shepp [Bha93, Boj65, Fin84b, Grz83, JG70, KF62, MP62, McL79, RJ90, Tre65, Van66, Van63, Wei67a, WM85, Fin84a]. Sherali [Bla82]. Shetty [Bla82, Fro80]. Shields [Kaz62, TWW65]. Shift [AG01, Fre82, War75]. Shift-Invariant [AG01]. Shifted [TDS18]. Shih [Sam80, Rüd97]. Shikin [Spä97]. Shilov [Rub69]. Shin [Pra86]. Ship [Col87a]. Shipman [Bai74b]. Shiryaev [Pro91]. Shiryayev [Kem85, Spi81a]. Shivamoggi [Elu88, Wei90b]. Shkil [Har70]. Shlomo [Pol76, Sle67, Wax91a, Wax91b]. Shmuel [AA82]. Shock [EHS17, HSSZ98, Ric60, Sto62b]. Shokin [Por85]. Shooting [Bai74b]. Shootout [SS88]. Short [Hei69, LS08, New76a, Pat19, RSS65, Ste67b, TE96, Ram93]. Short-Term [TE96]. Short-Wavelength [Ram93]. Shortest [Iba73, Wil21]. Shot [Tok95, Tok94]. Should [OW95b]. Shoulder [BM95]. Showalter [Her78a]. Shreider [Ham77, Jun70b]. Shreve [Ell80].

Shrinking [HSŠ16]. Shub [Mei89]. Shubik [Pac87]. Shuffle [BK76, Knu77b, Wol86, Knu77a]. Shuffle-Exchange [Wol86]. Shuffles [Kel95]. Shuffling [Gol61]. Shutout [Kel84c]. SIAM [For96, Gud78b, Hig05, Mac98a, Rob94b, Win96b, Ano68l, Ano69k, Ano69l, Ano70j, Ano71m, Ano71n, Ano71o, Ano71p, Ano72l, Ano72m, Ano72n, Ano73l, Ano73m, Ano74l, Ano74j, Ano74k, Ano75e, Ano75n, Ano76o, Ano76p, Ano77o, Ano77p, Ano78p, Ano78q, Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Ano96a, Ano97b, AR80, Blo73, Coh93b, Dea90b, FK06, LCK^{+}17, Mor86c, PC90b, SW80, ADPR04]. Sibirsky [Chi90]. Sibuya [Sch77a, Wil92]. Side [AE23]. Sided [Ask74, Ask76]. Sides [Ber90b]. Sidney [Abl77a]. Siegel [Coo67a, Kar63, Saa73]. Siegfried [Bor90, Bre68, Moc85, Sin93b, Sin93c]. Siegmund [Gho87]. Sigal [Hag85b, Hel97]. Sigest [Ano08a, Ano08b, Ano99e, Ano99f, Ano99c, Ano99d, Ano00d, Ano00a, Ano00b, Ano00c, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05d, Ano05a, Ano05b, Ano05c, Ano06, Ano07b, Ano07c, Ano09, Ano10, Ano11, Ano12a, Ano12b, Ano18, The99b, The99c, The00a, The00b, The00c, The00d, The01a, The01b, The02a, The02b, The02c, The03a, The03b, The03c, The03d, The04a, The04b, The04c, The05a, The05b, The05c, The05d, The06a, The06b, The06c, The07a, The07b, The08a, The08b, The08c, The08d, The09a, The09b, The09c, The09d, The10a, The10b, The10c, The10d, The11a, The11b, The11c, Edi11, The11d, Edi12a, Edi12b, Edi12c, The12, The13a, The13b, The13c, The14, Edi14a, Edi14b, Edi14c, Edi15a, Edi15b, Edi15c]. SIGEST [Edi15d, The16a, The16b, Edi16a, Edi16b, Edi17a, Edi17b, Edi17c, Edi17d, Edi18a, Edi18b, Edi18c, Edi19a, Edi19b, Edi19c, Edi19d, Edi20a, Edi20b, Edi20c, Edi20d, Edi21a, Edi21b, Edi21c, Edi21d, Edi22a, Edi22b, Edi22c, Edi22d, Edi23d, Edi23a, Edi23b, Edi23c, Edi24a, Edi24b]. Sigmund [Hal90a, Hal90b]. Sign [Ask86b, AD87, Fre86b, Gib97, JF87, MS14, NF16, Rah87b, ST94, Jor87a]. Sign-Indefinite [MS14]. Sign-Solvable [Gib97]. Signal [Bri63a, Bri63b, Cam91a, Cle63, Hol79, Hop92, RM89, Sta89, Sta90]. Signaling [Mor88c, SS24]. Signals [AKP^{+}91b, Beu66, BDE09, Wit90, Woo92, Zit63]. Signature [KM14]. Signatures [Kea93a, Kea93b]. Signed [SAB19]. Significance [Gev82]. Signomial [MW76b]. SIGNUM [Ano76p]. Sigurdur [Her80b, Zal83]. Sikorski [Boj97]. Silberman [Gre63b]. Silva [Ste70b]. Silver [Law66, AR63]. Silverman [Bre62b, AN78]. Silvester [Kri85]. Silviu [Gos79]. Silvu [Hes83]. Silvu-Sburlan [Hes83]. Simanca [Uri92]. Similar [Sed91]. Similarity [Ame81, BGH^{+}04, Faz98, Log98, PY98, Ame81, Olm77]. Simion [Isr92b]. Simitses [Lin91a, Lin91b]. Simmonds [Ant72, O’M87a]. Simmons [Bla84]. Simon [BC96b, Bra85b, Hag85a]. Simoncini [Aur18]. Simonnard [BI67]. Simple [Abe88, Ant98, BM03, BS93, Dra83, EH78, Elh78, Fin76a, Fri71, Gre67b, Gri85, Han75a, HS03, Hou98b, KS90, LM99, Mey89d, Mor85a, NS24, PS04, Rou97b, SV98, Sch75b, BG96, Hil98, Hou98a, Mor86c, Rou97a]. Simplest [MP72]. Simplex [Ars97, Bla88c, Kla85, Kla86b, Mac98a, Mur93, ST91b, Wol65, ST91a, Ano93c, Kla86a]. Simplexes [Kla78b, KG79]. Simplicial [AG80, BKKŠ09, SBH^{+}20]. Simplicity [FGP^{+}22]. Simplification [Quo75, Seg72]. Simplified [Geo71, Lev65]. Simply [GAHK05]. Simpson [Lam04]. Sims [Pin85a]. Simsa [Neu97b]. Simulated [Ham90, Ker91b, Ham90, Ker91a]. Simulating [Gou16, Hig08]. Simulation [Blo76, Blo77, Bun83b, Bun83a, Cer96, Elv75, EFA76, Fin95, Fra65c, Hig01a, JSF95, Joh91a, Joh91b, JMX96, KMC96, Lan76a, MMT^{+}21, Mil89b, Moc85, PM87, Raz92a, Raz92b, Tu76, Cul66, Kel86c, McC85, Mod92, Por92, Sam79, Say79, Xu96b]. Simulations [LKGK07, XC11]. Simulink [SRK99]. Simultaneous [Ber91c, Bor62, Car68b, Car69, Cul78, GKP75, Mcc78, Nik78, Ric88, Sch90, Sem59, Wes65, Ber91d, Sma76]. Sin [WLP^{+}61]. Sin-I [WLP^{+}61]. Sinc [McA94a, McA94b, Ste81a, Ste93a]. Sine [Ask73a, GF86, JBL72, Jag74a, JGF87, McA94c, Smi71]. Sing [Jd93]. Singer [Cha94c]. Singh [Dra61a, Jor80a, Kno83]. Single [Beu83b, Beu83a, Bob66, GZ19, Kei71, KWP95, Pro95, Ste04, WHZG24, Wit90]. Single-Valued [WHZG24]. Singmaster [HT81, OP88, Ren80a]. Singular [BT19, BBK20, Bus76a, Cam93, Car76b, Car70e, CA75, Dax90, DD02, DPS73, Eck72, Eck77, EK87, FH70, FR11, Fus68a, Gla75b, GF78, Gre59, Gri85, HK92a, HK92b, HT78, Har76b, Kok84, Kre78, LC72, Lev65, Lig86, MR71, MOSK97, MMH04, NC76, Neu98, O’M74, O’M86, O’M92, O’M94, Por72, Ric90a, Roc92a, Roc92b, Sch76b, Sch80b, Shr76, Smi75, Ste93c, Vas94, WJ75, War69, Wil67c, Wol77, XL94, dJFO98, DGH^{+}18, Her78a, Hom86, Lew91a, Man87, Man88b, O’M92, Pic92, Sag77, Was94b]. Singular-Perturbation [Hom86]. Singularities [Ame78, Gri85, LH19, VŞ10, Wet93, Mey86b, Mey89c]. Singularity [AK77, BFM95, Riz94, Rum99, Hea91, Mey95b]. Singularly [MBK97, O’M75, O’M88a, O’M08e, RSTK97]. Sinha [Ban84b, Ban84a]. Sinkhorn [CGP21]. Sinusoidal [Kim65]. Sinusoids [Nie96]. Six [DTT98, Han90, Mor88a, Mor89b]. Sixth [Cas75]. Size [GM76b, GM76c]. Sizes [Kar22]. Skeletal [Pit90a, Pit90b]. Sketch [Har76a]. Skill [Mar86b]. Skinner [Gro96f]. Skip [Kij93]. Skip-Free [Kij93]. Skiving [JRZ97]. Skorohod [BR85, Kus74]. Slab [MP93]. Slabs [Bar64]. Slackness [Kle83]. Slade [Bur94b, Bur94c]. Slash [Ber76b, Sil78, Wil77c]. Slater [Bou65b]. Slattery [Whi93]. Slepian [Grz83, Jag79a, Ner70, Cha63]. Sliding [JH11, War70, War71]. Slipping [BK78]. Slit [Cer70, Cer71]. Sloan [Gen96]. Sloane [BC96b, Mat80, Moo74a]. Slobozhanin [Dij90]. Slomson [Als92a, Als92b]. Slope [Els91]. Slopes [Bon67, Law78]. Sloshing [OO17]. Slow [DGP^{+}16, Pit89, Axf65]. Slow-Fast [DGP^{+}16]. Slowest [CF80]. Slowly [GBC^{+}90, Gib90, Kev87, Sha89a]. Smale [Gus76, BD94]. Small [ABJK10, BW11, Bob66, CM89, Cho78, Dri68, FHRO95, Gau62, Hig07, IWMW21, O’B95, SM94, Wal92, Wal93c, Wal93b]. Small-Amplitude [CM89]. Smallest [CLG14, Cul78, GE63, Hue92b, Kla59, Law65]. Smirnov [Bre62b, Buc69, Col86, Kor65, Sto66, Yan65, Alt76]. Smirnova [Har72b]. Smith [Are86, Boa85, BG64, Che18, Cus96, Des68, Deu81c, Gib73, Han79, Hom86, JBL72, Maw86, Per68, Por80, Sin93a, Spi84b, Wal75b, Ber97a, Spi84a]. Smolitskiy [Wou68a]. Smoluchowski [MP06]. Smooth [BBK20, DB78, FJT19, MG23, Riv59, Riv60b]. Smoothed [BFM^{+}05]. Smoothing [Bjö85, Mil61b, RW64, Str76a, Tew77]. Smoothness [Tot91]. SNE [LS22]. Sneddon [Bue71, Tul90]. SNOPT [GMS05]. Snyder [Che67b]. Sobczyk [Kel86a]. Sobel [Art83b, Art83a, Art85b, Gup82]. Sobol [Ham77]. Sobolev [Kol63, Coh75a, KK78]. Soc [Lax89]. Social [Coz85, Coz87, Fis74, JMFB15, Man61d, Pac87, Ser04, She91, TKMP11, Bas88].

Societies [Dre60]. Society [Ano72o, Ano73n, Ano75o, Ano77q, Hig00, Hig02c, Ano73m]. Sociology [Fre86a]. Sociology-The [Fre86a]. Sod [Van88]. Söderström [Wib87]. Sofonea [Pav94]. softcover [Bor06, Tre13]. Software [BNT76, Bau73, DW95, EGJK04, Fla66, FLGM23, Gup91b, Mit94, NH78, Par94, SK78, Thi84b, Thi84a, TTM05, WC96, Chu96, Gup91a, Mil82c, Szy87, Tha82]. Soifer [Bol91b, Fir92b, Bol91d, Fir92a]. Sojourn [AW19]. Solid [Hau86, Hor95a, Jon92a, Jon92b]. Solids [CLH97, Dat90, KR98, Let69, Pit89, Red93a, Red93b, Riv65, Spi91a, Spi91b, Sub90, Gos63, Nor90, Whe74]. Solimini [Sha96]. Soliton [Kut06, Mil83a, SB98a, Kup94]. Soliton-Bearing [SB98a]. Solitons [MBK08, Sco84a, Sco83, Sco84b, Sco85, Sco88, Sco90a, Sco90b, War92]. Solodovnikov [Hig62c]. Solomon [Dia59, Gol71, Tho60a]. Solution [Ale95, Ale98, Ame86b, And66, And95, Arn89, Ast87, AH95b, Bec94, Ber75, Ber94, Blu94b, BL05, BR18, BRHC20, Cam91b, Cam91c, Che91, Che76, Che78, CS64, CM78b, CH66, Col91, CP78, CBS76, Con66, Cop71, Cor78, DO98, Dat82, Die62, Dor70, Dor73, Dor75, Duf84, Eds05, Enr78, Enr84, Faz98, Fel76, FEY61, GM81a, GM81b, Gat64, Gea80, Gea81, GR95, Gla95a, Gos83b, Gra67, Gre59, GS23c, Gun88, Gup89, Hac75, Han95, HV75, Hir78, HM73, Ike72, Isa89b, Joh68, Joh73, Jos70, KL76, Kan88, KPGR95, KPSP96, Kol65, Lam86, LS66, LeV93, Lei76, LA75, Lew83c, LW04, Liu92b, Log98, Lon78, MS76, Man76, MTGF78, McG82, Mei68, Moo65, Mor69, MK97, MCS69, MO66, NC76, New70a, Nie98, O’M75, Ole94, Olm76, Olv91a, Olv91b, OV85]. Solution [Pap86, PSSW06, Per68, Pin92c, Por72, Por80, Por64, Pro96, PA95, Ral65a, Ral69b, Ral74, Ral65b, Raw70, Ren93a, Ren93c, Rhe70, Ric76b, Roa78, Sco59, Sei76, Sem59, Ser76, SG95, SES75, Sik92, Sto76, Str65b, Swa77, Ta’90a, Tew77, Tho59, Tra66, Tri86, Van78, Var76, Var67, VH78, Vel63, Wil66, Wil72, Wol94b, Wou68a, dBL75, Ram88, Car71b, Bou66b, Hou67, Ort74]. Solutions [Aar07, Ada97b, AG80, Ano80g, Ano81c, Ano83f, Ano84f, Ano84g, Ano85b, Ano91e, Ano95n, Ano95o, Ano95p, Ano96n, Ano96o, Ano96p, Ano96q, Ano97k, Ano98d, Ano98f, Auc97b, BT05, BPV97b, BSW82, BE78, Bra66, BDE09, Bue66, Bur83, Cas75, Che67a, Che96b, CT75, CP76, CB76, Coh81, Col89, DGS05, De 84, Dri68, ERHO95, Fel78, Fif86, Fle85, GVL97, GDS05, Gla97b, GS22, Gra83a, Gre88, Gri85, Gup98b, Han87, Her65, Hop75, Hue92b, Hue96, IM88, Jac69, JR08, Jub70, KMH66, Kel79b, Kla97, Kor95c, Kov96, Lae76, Lan70, LT90, Lax76, Lem84, LL75, LG89, Lew89, Lio82, LS02, Mat69b, MR76, Mcc78, Mey69b, Mic88, MP72, Mie95, OT12, Ork95, Pet98, Pra89a, Ral69a, RFP10, Ren92a, RT97, RK75, RRE97a, RRE97b, RRE97c, RRE97d, RR98a]. Solutions [RR98b, RRe98, Rub78, Sei65, SS98, Shi17, SA78, Sig67, Sin93b, Sin93c, SM65, Sta71, Sta72, Str76a, TtK96, TeB97b, The91, TP78, TH96, Ung77, UFW17, Van75, Vel64, VN76b, VM95, Wal92, Wal93c, Wal93b, Wal63b, Was73, Wat96, Wil67b, Win90, Ano97j, Che96a, Gup98a, IM89, RR95, TeB97a, Ano82f, Ano87a, Men94a, Pac87, Ren92b, All78, Wil79]. Solvability [Ber84a, Cas94b, Chu89, Mey75a, Ber84b, Cas94a]. Solvable [BDvD^{+}98, Cli76, Gib97, Hun76, New76a, Bro90]. Solve [CC75, HYS76]. Solved [Erd72, FS73]. Solvers [BKST15, GZ19, MC76, SK76a, Sku04, Wei87]. Solving [Bar78, BG76, BB96, Ber85c, Cam93, CCCM06, CG76, CC78, Cum96, Dax90, DTV14, DG14, DX98, FLC96, Fro86, GVL97, Gol76a, Gol76c, Gri85, KB76b, LP75, LRJ78, LMMK21, Mit94, Naz80, Neu98, NH78, Pan97, Pri76, Rog69b, Sha75, SG76a, SWD76, SG79, SRK99, She91, Und76, VG78, VB75, WiOH10, WZ72, Ale90, Ber85g, CL72, Gen76]. Some [Aar07, AMO91, Ame73, Amo78b, Ang92b, Ano99b, Arc75, Ari92, Ask85, Ask86a, Bad60b, BNT76, Bel63d, Ber96, Ber97b, BSFW82, BSW82, BE78, BK78, Cas80, CCA08, CT75, CA75, CR67, Cor68, Cro77, CS90, Den76, The99a, Ehr76, Far69, Fos81, Fox67, Gay78, Gol76a, GF78, Gol73b, GO89, Gre60c, Gud75, HY71, Ham74, Han86a, Har60, Hau81, Hea76, Her88, Hir71, HF78, Hop75, Hou65, Hue91, Jer76, Kap70a, Kel80b, Kla66, KN67, Kni94a, KLT03, KLT04, Kot65, Kus76, LM90, LM79, LP71, Lou70, Luc71, LK65, Mac68a, MR76, Moh66, MW06, Naz80, Pan97, Pon83, Rei97b, Rei97a, Ric90a, Roh66, RS75, Rub78, Sch70a, Sei65, Ser78, Sha72b, Sle83b, Sma90, Str76a, Tew69, TF65, TWW65, Ton88, Tri78, TS90, Udw89, Ula64, Var79, Wat96]. Some [Wat22, Wat93, Whi63b, Wol71, Wol75, Wu72, XZ98, Dut82, Gos64, HA71, Kni94b, MT64, Olk59, Ros72b]. Sommerfeld [Her87]. Söndgerath [Dix91]. Soner [Men94a]. Song [Coh90a, Coh90b]. Soni [Kaz67]. Sophus [Sch88]. Sorensen [Hig00]. Sorting [Bro59a, Bro60a, Bro60b, Car64, Kai59, Mal62, Mal63, Sle62, Wei60b, vL60, vL61, Wei67b]. Sound [Kan62c, Mar86a]. Source [EC76, GZ19, Gla84b, Hub83, MS91, Run91, SM78]. Sources [GJRW24, Lon77, Por64, Wil59, Pic92]. Sourisseau [Jor81]. Space [Ant82, Ban88c, Bur82, Coh75a, Coo67a, Fal70b, Far85, Fri80, Gla89b, GG89, Gla91, Gol65a, HL76, Kap64, Kha86, KW18, LS85, Mor75b, Pal96, PCLS23, Ral62b, Ros68b, Sae70, Sae73b, Sle85, Syn83, Ter92, TFL23, VN76b, Vog87, Vog91a, Vog91b, WP96, Xu92, Xu96c, Xu97, Bou91b, Coo66, Coo73, Don90, Erd83, GA63, Lud77, Ral62a, Med78]. Space-Dependent [KW18]. Space-Time [Bou91b, Lud77]. Spacecraft [Sin76]. Spaced [Ber07e, Ber64, Beu66, Str75, War98]. Spaces [Ahm75, AG01, Ama76, Ame78, Art80, BDJ99, Fan59, FR66, Gev82, Her78b, Jor87b, Jor89b, KW71, Mar83a, Moo78, O’M92, Pet73, Ral62a, RL95, Sae73a, SR98, Syn93a, Syn93b, Tod84, Ama77b, Bar67, Ber78, Che91, Dat80, Eva85, Her80b, Lie95, Mar83b, Ort74, Rya63, Sop69]. Spacetime [Bro93]. Spacey [BGL17]. Spacing [Eub88, MT76]. Spanning [Mye75]. Sparrow [Hol85]. Sparse [BDE09, Cen81, CZ89, CD75, DdSG^{+}22, Duf84, Eis86b, Hac75, HNP91, Liu92b, OLHB02, Par74, PGS75, San76, SES75, Tew66, Tew67, Tew70, UA07, VB05, WKW75, ZWGB16, dEJL07, Wil86b]. Sparsifiers [BSS14]. Sparsity [LT13]. Spath [Alf93]. Spatial [Dur99, Erm89, EH16, Far85, HHSS75, RK75, Wei62, YF78, Mar65]. Spatially [SB98a]. Spatiospectral [SDW06]. Spatiotemporal [MPT^{+}02]. Special [AS80, Ask74, Ask76, Bor73, BDvD^{+}98, Car75a, Car70d, CMZ94, Cot76, Der71, Fol73, Hea74, Ism83a, Kaz67, Kel60b, Köl87, LOT76a, LOT76b, Olv72, Olv97, Ros72c, SW81, Sto66, Var87, Wim75, Ask79b, Ism83b, Wim65, Zei72]. Species [Ken93]. Specific [Fin72]. Specification [Ber85d, GM87]. Specified [Iba73, Urq68]. Spectra [Coo66, RM75, Bei81, Bro76b]. Spectral [Ari92, AT17, BBR13, BR18, Boy90a, Boy90b, Cha81a, CHT21, EKS97, Fig93b, Gon06, Got78, Hag89a, Hel97, Hin91, Hoc92, Jin91, KR78, LSY16, LRV76, McL86a, MMT^{+}21, Mor81, New76c, OT13, Pap85, Par76, Sch84a, Sch78d, SHL^{+}79, Vos75, Vos76, ZZ09, Zwe84a, Ans64d, Ans85, Boy88, Fig93a, Sch82b, Sch84b, Zwe84b]. Spectral/ [MMT^{+}21]. Spectrum [Eng03, MO78, NS24, Rus85]. SPECTRWM [BR18]. Specular [MV93]. Speculations [BD11, Ham74]. Speculative [Sam73]. Spedicato [Ari91b, Ari91a].

Speech [Bla86a]. Speed [ADR76, Jor77a, Jun64, Ock91, OT11, Pan84, Rei78b, Sed91]. Spekreijse [Ta’90a]. Spencer [Ell90, Hor83b, Moo75]. Sperb [Red83]. Sphere [CL08, DL76, GE63, Hey63, HC16, Joh75, Kla64b, Kla95a, MV93, Pea82, Sha88, SDW06, Sin80, Gro96c, DG96, HT81, Law65]. Spherical [CHL08, Jor86a, Sah95, Sha70, Sha71c, Sha71d, Sha73a, SK85, Wil59, Wil61, Zuo96]. Spherically [Wil59]. Spikes [Hil07]. Spiking [CSW^{+}08]. Spingarn [Gri88]. Spira [Jag72d]. Spiral [PDS78]. Spivey [Fia73]. Spline [Chu81, Jet91, LF73, Lew76, Mer84a, Mul76, Pru78, Sch78c, Sch91b, Sei78, dB74, Alf93, Mer84b, Ahl82, Cab74, Sch91a]. Spline-Interpolation-Algorithmen [Alf93]. Splines [Bir69a, Deb76, Def76, Lew76, Nie76, Spä97, UB00, UFW17, dBHRR95, Ron95, Smi80b]. Splitting [CKCH23, Ehr76, Rüd97, VBK22]. Spneod [Ral74]. Sports [Alb10, BI18, GLY^{+}18, Gou10]. Spotlights [Ips12a, Kil18a, Kil18b, Kil18c, Kil18d, Kil19a, Kil19b, Kil19c, Kil19d, Kil20a, Kil20b, Kil20c, Kil20d, Kil21a, Kil21b, Kil21c, Kil21d, Kil22a, Kil22b, Kil22c, Kil22d, San14, San15a, San15b, San15c, San15d, San16a, San16b, San16c, San16d, San17a, San17b, San17c, San17d, Tum12a, Tum12b, Tum12c, Tum13a, Tum13b, Tum13c, Tum14, Wil23d, Wil23a, Wil23b, Wil23c, Wil24a, Wil24b, Wri14a, Wri14b]. Spotting [Shu62]. Spread [AMS^{+}23, CD03, Eng03, Kri97, Tud92, Yor75]. Spring [Pet97b, Pet97a]. Spring-Damper [Pet97b, Pet97a]. Springer [Hei96, Hig01b, Mal80, Tre13]. Springer-Verlag [Hei96, Hig01b]. SQP [GMS05]. Square [Bos87, Ewi67, FR63, New62b, SC70, Wou66]. Squared [Rei71c]. Squares [Ber64, BG98, Bog75, CS64, Cho76a, Die62, FSW^{+}66, Fie93, Gen76, Gil65a, GNZ20, Jok66b, Jok66a, Kel79b, Las07, Mei68, Mil73, Naz80, Nie94b, NS75, Pro61, RW80, She09, Ste77, ST91b, Wah65, Wat22, ST91a]. Squaring [Hig09]. Squire [Str68, War64]. Sri [Tak83]. Srivastav [Jor89a]. Srivastava [AS85, ASF86, Ber92b, Ism83b, Rue74a, Rue76b, Wim86, Ism83a]. SSOR [HYS76]. Stabilities [Bus69]. Stability [Abe88, AAJ76, Aro78a, Art91a, Art91b, BvdD78, Beb68, Beb79, Bot74, Bri63b, BT20, Bry90a, Bry90b, CMR23, CG66, Cur77, CV77, Dan91, De 84, Dre76, Duf69, Dur92b, Esl93, Esl96, GQJ96, Gev82, GT10, GST01, GKR20, GG68, HE91, Han75a, Han91, Jód96, Jor81, Kal70, Kir75b, Kus19, LaS64, Lef65, Lem85, Lev88a, LT98, Lin91a, Lin91b, Los76a, Mat85, MM98, Mei89, MC21, MWL97, Mil90a, Mil90b, MP20, MY97, MN76, Mor68, Noh76, NW97, PE76, PDS78, PZ81, RB23, Ral65b, Ram89, Ric60, RW95, SWM^{+}07, Shu90b, Sin86, Sny97b, Sny97a, Sou80, Ste97b, SH94, Tad89, Tad87, Tho77a, Tho69, Van78, Ven76, Wal95, WMZ94, ZM20, vW90, vWHE91, Kir75a, MW97, Bri63a, Asn75, Elu88, Mil82a, OP88, Sle83a, Bau88, Bre68]. Stability [Col73, Fif79, GGG69, Han87, Kan68, Kus93, Maw91, Rox92, Sac74, San82, Shu90a, Wou64a]. Stability-Preserving [GST01]. Stabilizability [PZ81, Rus78]. Stabilization [Hew76, Lag97, LW95, Lio88]. Stabilized [GAHK05]. Stabilizing [Vas97]. Stable [BR97, Eds05, Gea85, Jen75, Kri91, Pel97, PTK11, SK76a, SXY19, Tho76b, Win90, Enr84]. Stackgold [Kap81]. Stadium [RR97b, RR97a]. Stadler [Nit90]. Staircase [Fou84]. Stakgold [Ans68, Gue69]. Stamatis [RR89, RAK88]. Stamp [Laf93, LL94, Sch93c]. Stampacchia [Eva81, Kap85a]. Stan [Hig05, HA71, O’R92]. Stancu [Kar87]. Stancu-Minasian [Kar87]. Standard [AvK^{+}08, JMJ95, Kel96, Lev89, MM10, Nik78, SZ99, TA90b, TLM90, ZK96, AR81, Jur95]. Standards [Mil76, Tod75]. Standing [O’B11]. Stanford [Man61d]. Stanisić [Bra85a]. Stanislaw [Pro95]. Stanley [Whi92a, Whi92b, Whi88, Wol94a]. Stanton [Ham61]. Stark [Tek89]. Stars [Chr68, Hig92]. Starting [Gar75, RR97b, RR97a]. Starzhinskii [Mey77b]. State [BRP88, Che91, DRWK64, Fly76, FS76, Gar60, HSV95, Kap64, Kla63a, MP20, Mor75b, Mun61, NT78, Nie94b, O’M75, Sae73b, SS89, Sha75, She76, Tri62b, Vas94, Yos60, SWD76, Gar61, Gue82, Med78]. State-of-the-Art [Nie94b]. State-Space [Mor75b, Med78]. States [AR78, Ben70, PMR75, She76, AR80]. Static [Shi93]. Statics [Spi91a, Spi91b]. Station [Ber60]. Stationary [BR69, Dax90, Fox67, Fra65c, KA89, KTSB21, Pim87, Rob62b, Wil78c, Ano61a, Kaw68, Pic87b, Ser94]. Statische [Hig62d]. Statistical [Aar07, ABG^{+}95, Bak91b, Bak92, Ber95, Ber85b, Bjö61, BL84, BSFW82, BSW82, BK83, Car93, Cur64b, Deu63, Eat96, Efr88, Eis86a, Eps59a, FTM95, Far94, Feu76, Fos81, Fri71, GG95, GO09, Gup82, Gut74, Hey88, Hig62c, HB95, Jow86, Kir90a, Kir90b, Kum82, Kup62, Kup75, Le 92, Mac91b, Mac91d, MQ18, Mal77, Mar95, McC91a, Mik89, Noe72, Par92a, PE96, PC89, Rei79, Rei88, Rei89, Ros72a, Ros78, Sav18, Sha92a, Sha92b, Sin89, SW93b, Sin79b, Spo88, Sta88, Ten01, Thi84a, Ton88, Wei79, YH93, Zlo91, PC90b, Aok70, Ber93, BR62, Blo76, Blo77, Cor92b, Gen86, Gil95, Goo86, HR62, Hin91, Jow72, Lin79c, Man63, McK95, Pao93, Rei87, Ros63, Set63, Sin93a, Str89b, Thi84b]. Statistically [CPSV18]. Statistics [AABT15, AH95a, AG83, Ask80, Bar89, Bar91a, Bar91b, Bir64, Cha94c, Con94, Efr79, Fos68, Gre73, Har93, How94, HRM96, Kur69, Mal68, Mal69, Man61a, Mee96, MBK08, Par92a, Rao82, Rei63, Ver09, AGR83, Ber91b, Dwa66, Dwa80, Eps59b, Fro87b, Hor93, Lil68, Luk75, Mor88b, Noe81, Olk78, Rei77c, Rei83a, Ric67a, Ros68a, Sim82, Tho60a, Dan66, You74]. Status [Som78]. Steadily [Sac67]. Steady [BRP88, DRWK64, ERHO95, Fly76, FS76, Gar60, Kla63a, Mun61, Ork95, SS89, She76, Sul93a, Sul93b, Ta’90a, Tri62b, Yos60, Gar61]. Steady-State [DRWK64, Fly76, FS76, Gar60, Kla63a, Mun61, She76, Tri62b, Yos60, Gar61]. Stearns [Mil69a, Mul69]. Steeb [Chu94b, Jag90, Chu94a]. Steel [VS95]. Steen [Ham80]. Steep [YFS76]. Steepest [Gri72, Olv70]. Steeples [Ame73]. Stefan [Hol89b, Sle84b, Gla97b, Led96, MM10]. Steger [PT81, SK80]. Steiglitz [Won83b, Won83a]. Stein [Bri96]. Steinberg [Mas95, PT81, SK80]. Steinmetz [Dev94b]. Stekloff [KS68]. Stem [PMK96, Win93, Win96b, Win96a]. Stencil [DKW^{+}09]. Stengel [Mor89c]. Stenger [Baz73, Boj97, McA94c, McA94a, McA94b]. Step [Lin59]. Stépán [Han91]. Stephan [Pim93]. Stephani [Olv91a, Pul91, Olv91b]. Stephen [Bar94, Bar89, Bla88b, Bru86, Che92a, Gua92, Gup91a, Gup91b, Gus76, Mot88, Ort96b, Pal95, Pal96, Par94, Wei91, Wes92, Yat84b]. Stepping [Arn92]. Steps [Cat21, DTT98, Hor78]. Sterbenz [Row76]. Stern [Dav91a]. Sternberg [Pol76, Wax91a, Mla97, Wax91b]. Sternin [Daa97]. Stetter [Gea74]. Steutel [Kle75]. Steven [Ell80, Hof85, Luc18, ML85, Moo77b, Sha95a, Tru94]. Stevens [Bri64]. Stewart [Bol91a, Bre64a, Guc79, Hol87, Mey89c, Par88, Par81a, Pul91, Bol91c, Kin97]. Stiassnie [Gro88b]. Sticky [BRHC20]. Stiction [BBK20]. Stiefel [CMSZ24]. Stieltjes [DL66, Har67b, Hun76, Mon82, WM78, You76]. Stiemke [SS79]. Stiff [Enr84, Gud75, Sha75, SG79, Jel87]. Stiffly [Jen75]. Stigler [Bar89]. Stiles [Duf85]. Still [Cat21, Sku04]. Stimulating [Dre60]. Stirling [Hoc92]. Stirring [FT11]. Stochastic [Art83a, Art85a, Art85b, BNT10, Bal20, Bel90a, BGL17, Beu89, Bis78, Bjö85, BR18, CJR18, CGP21, Cho75, Çin93a, Cin93b, CAiJS23, DMT^{+}18, Dun94, Dup93, Dur92b, Dur97, Dur99, Dwa70, Ell85b, Elv75, EFA76, Fei96, FL18, Fin93, Fle69, Fle79, Fle82, Fle84a, Fle88, Fri78b, Fri79a, FSY65, GR95, Ger62, Gli91b, GS04, Gra83a, Gra86b, Gus90b, Hau78, Hau81, Hig01a, KWB96, Kar83, Kar84a, Kaw67, Kaw68, Ker91a, Ker91b, Key94, KPGR95, Kni86, Kul65, Kus72, Kus76, Kus93, KWP95, Lah94, Lai91, Lar78, LK98, LP71, LL69, Lin80, Lin90, MR18, Mer89, Mey84, Mey89a, Moi89, Mor98, Mor71, Mor88d, Nas87b, PE76, PL78, Per89, Pin77, Pit77, Por86b, Pro91, Pro95, PF96, Sch80b, Ser84b, Ser84a, Ser94, SST95, SA78, Spi81b, Spi82, Ste90b, Tze91]. Stochastic [VCKJ16, VDC22, Wal86, Wei77, Wet74b, YZZ04, ZZ18, Art83b, Art85b, BR67, BR76, Bri96, Çin93a, CK86, Doo63, Ell80, Fab80, Fle77a, Fle77b, Fle84b, Fro74, Goo91b, Gus90a, Hey93, Kar82, Kar84b, Kar87, Kaw65, Kel86a, Kus74, Lud87, Mar75a, Mor68, Mor73, Mor89c, Par67b, Paz76, Pin77, Pin87b, Pit79, Saw85, Ser80a, Ton95, Yan92]. Stochastic-Processes [Kar84a]. Stochastic-Systems [PE76]. Stock [Car71a, JRZ97, Lem70]. Stoer [Kin82, Kle72]. Stoica [Wib87]. Stokes [Dea90b, Pon90a, Sul93a, Bea90a, Bea90b, CS90, Dea89, Far96, FC99, GP79, Mey89d, PSSW06, Pon90b, Sul93b, TF96, TP78]. Stolarsky [Yan77]. Stolfi [Hug93a, Hug93b]. Stone [Koo77, Mor93, Spi80]. Stones [HSSZ98]. Stoop [Abr94]. Stopped [Sie89]. Stopping [Fri78b, Fri79a, Spi81a]. Stops [BOS73c, New72a, San87]. Storage [Cof83, Gra91, Sal64, SES75]. Stout [Sam76c]. Straeter [Dav75a]. Straffin [Saa96]. Strafin [ML85]. Straight [Mon60, Sou76]. Straight-Line [Sou76]. Strains [Cho78]. Strands [DBS^{+}07]. Strang [Cab75, Ros82, Hen86]. Strange [Hol85, Mac91b, Mac91d]. Strasser [Jow86]. Strategies [Kon62, MQ18, Mat89b, OLHB02, San87, Yan97b, Yan97a]. Strategy [BR97, Kla74b, Pac10, PL78, SS96, Saa96]. Stratification [DMT^{+}18, Eub88]. Stratified [JM78, Mar86a, She87]. Stratifying [ZKP17]. Stratonovich [Hig65]. Straub [Ruo90a, Ruo90b]. Straughan [Pro84c]. Stream [CS84, Hir78, LPL05, PLL05]. Stream-Function [Hir78]. Strecok [Bou66b, Jag72a, Sho66a]. Strength [Set63, Set62b, Tad93]. strengthened [EV91]. Stress [Cru86, Fel92, Hue91, Kel64, Mal93b, Mal93a]. Stretching [Cla64a]. Strict [LR05]. Strien [Bru95]. String [CEH12, Kru83, Mey95c, Rei59b]. Strings [Sch92, Yon06]. Strip [Cer70, Cer71]. Strogatz [Sha95a]. Strong [GP02, GST01, Mor75a, OR21, PMK96, Win93, Win96b, Win96a]. Stroock [Var85]. Strother [Tal81a]. Stroud [Ste73b]. Stroyan [Kap78]. Struble [Lev62]. Structural [Boo94, Chu89, HS15, Jen78, Koh88, Mro91, Par92b, Pit90a, Pit90b, RB23, Spi87, Spi91a, Spi91b, Tad89, Tad93, Tho77a, Yao86, Mas87, Mro91, Pao88, Sei85]. Structure [Bat21, BT19, Bel90a, BBR13, Bur83, Cor79, Deu95, EERG20, Fre86c, Giu18, GYS95, GN13, HMT11, Har72a, Jer76, Jor89c, KS95, Kri91, LB10, Lin69, Mil68, Mor75a, Mul69, Neu97a, New03, RPFM17, SCS10, Sel76, Sez90, Sha95c, Ste62, TKMP11, Wen61, Zie70, Har64a, Vog87]. Structure. [Car87]. Structured [AC78, HT12, Par10]. Structures [AN83, Aur18, EGJK04, GK95, Han75a, KR98, LLSN96, Lin91b, PL78, Spi91a, Spi91b, Dwa64, Ell90, Kee95, Lin91a, Mal87, Nic96, Rhe85, Sym92]. STT [NT78]. Stuart [Cha64a, Fri71, Hal66, Ren95, Wat63]. Stuck [SW81]. Studden [Ric67a]. Student [GH22, Mal88b, Bla87]. Students [Bla97b, Ken91b, Wax91a, Wax91b, Bla97a, Ken91a]. Studies [Col76b, CWW99, Fre83, JRZ97, RH96, Yan97b, Zie84a, Ame86b, Hei96, Ash89, Bla88b, Gol71, Sch78c, Wig85, Zie84b]. Study [Ada97b, But78, CR96, CGT^{+}97a, Cum91, EBTZ94, ETZ94, Fee00, FP21, FP04, GK95, Haw95, Hig05, JW83, Kee95, KSW96, Lau03, LF73, LR00, LH19, McM87, Mil91a, Mil91b, Nag96, Ort96b, Sau79, SS89, SN96, TS90, VBU97a, Wim88, Ada97a, Ort96a, VBU97b]. Sturm [And95, BV96, CA75, DS61, Hou74a, PA95, Ste64]. Sturmfels [Whi94]. Sturmian [Rei77a, Swa78, Laz82]. Styazhkin [Ros73]. Style [Sco92]. Subadditivity [AS80, Coh88, Coh93b, Coh93a, SW81]. Subba [Hin87]. Subcellular [But74a]. Subcommittee [LCK^{+}17]. Subgradient [Gof78]. Subgraphs [Bro66b]. Subinfinity [Kha88]. Subject [PEGK03, Sau76]. Submatrices [Bar22, BN76, SN04, Gib77]. Subroutine [Cod74, Kri86, Lon78]. Subset [Ber92c]. Subsets [Arb78, Der71, Fis88a, Zei72]. Subsidence [Gie63]. Subsonic [EMB76, Sun76, Vel21]. Subspace [Kui06, LS08, MYDF08, SS05, Xu92, Par10]. Subspaces [AMSV02, Hew76, Joh88, Ste73c]. Substitute [GM76b, GM76c]. Substructuring [BJ75]. Success [Gia63]. Successive [Eva93, Eva94, Ren94b]. Suction [Sie65, Wil78c]. Sudarshan [Gal95]. Sudden [Gia63]. Suddenly [Lin91a, Lin91b]. Suddhendu [Dur97]. Sudhist [Man61a]. Suffer [GS23a].

Sufficiency [KK98]. Sufficient [Ewi77, MW76b]. Sugihara [Giu18]. Sugiura [Mei77]. Suitably [Ewi77]. Sum [Böh90, Bre89, Car67c, Car68a, Coo93, Coo94, FSP93b, FSP^{+}93a, Gar69, Gel76, GKG^{+}86, Gla79, Gla83, Gla85b, Gro92a, GNP95, HS81, Jev95b, JG82, Jor84a, KH72, LG80, Lar80, Las07, Mer82, Mer83, Mey93a, Mey94b, NPW95, New94a, New81, OK94, Phi71a, PC90a, Rie72, RPP88, Sch61, Van98, BY98, Boe98b, PC90b, GK86a, Jor73, Jor77b, Jor83a, New82a, WJ94, WBGP91, Fis71a]. Sum-Distinct [Jev95b]. Summaries [GH13]. Summary [Wey67]. Summation [Ano89d, Ber73b, Ber74, Ber76c, Bro66a, Cha82a, CS83, Cle69, Esp95, Fet84b, Gla80, Gla81a, Gol76b, Han85b, Lam04, LB77, Lie77, Lie78, Mal70, Mal71, Maz84, RC87, Rue74a, Sar71, Sch63b, Sri73, Wah68, BPK84, Car72a, GR81, Lev77b, Nic65, RR89, Rop83, RAK88, SKGM67]. Summations [Bad60b, Car76a, Fis88c, Jor88c, Sum87, Bre77, Fis88d]. Summers [Fis88d]. Sums [BPV97b, BD93, De 79, DL66, EL85, EL86, Gla86c, Gro92b, Hen92, Jor86d, Jor93a, Kam78, Lin10, MM06, NS77, Neu97b, Ng96, Pet85, Ral65a, Rot81, Sib91, Wil92, WGLS92, BPV97a, BdD93, Fet80, JG93, Ste71]. Sun [Hil96]. Sunahara [Aok70]. Suntan [Ols03]. Super [Háj79, PAT^{+}80, Pow80]. Super-Controllability [Háj79, PAT^{+}80, Pow80]. Supercellular [But74c]. Supercomputers [And90, Kow86]. Supercomputing [Por86a]. Superconductivity [CHO92, DGP92, GP02]. Superconductors [Cha00]. Superconvergent [PG00]. Superlinear [SS05]. Superposition [Fle65]. Supersonic [Hel75, Kan62b]. Suppes [Man61d]. Support [Wal92, Wal93c, Wal93b]. Supraconservative [Vel21]. Supremum [LS71, LS72]. Sure [PE76, Sam76c]. Surface [Bél88, BT85, BL09, CF76, Gro88b, GAB^{+}88, Law78, Lud75a, Mey76a, Mor76, Mye98, NO60, OH76, Pra89b, She75, Sti87, Pat87]. Surfaces [DeR91a, DeR91b, Fel70, GWG13, Koc92, Küç92, LB78, Roc92b, VB75, Alm87, Ear81, Gre89, Roc92a]. Surfing [Sug98]. Surma [Ano82f]. Surprises [CGRG^{+}23]. Surprising [Sku04]. Survey [BS77, BD11, BGW15, Bra81, BDvD^{+}98, CS16, CS63, CG24, Cod70, Cur75, DK85, DL92, Dri91, Fre82, Hal70, HSV95, Hel78, Her73, Hig87, Hig12a, Hig12b, Hig12c, Hig12d, Hig13a, Hig13b, Hig13c, Hig14a, Hig14b, Hig14c, Hig14d, Hig15a, Hig15b, Hig15c, Hig15d, Hig15e, Hig16a, Hig16b, Hig16c, Hig16d, Hoc23d, Hoc23a, Hoc23b, Hoc23c, Hoc24a, Hoc24b, Hol81, Joy71, KM04, Kra70, Kus76, LM05, LM79, LeV03a, LeV03b, LeV03c, LeV03d, LeV04b, LeV04c, LeV04d, LeV05c, LeV05a, LeV05b, Luc10, Mir71, Miu75, Miu77b, Ove05, PR86, PWG18, Pit79, PI88, Pol73, PT07, PB74, Pow70, Rin61, San06a, San06b, San06c, San06d, San07a, San07b, San07c, San07d, San07e, San08c, San08d, San08a, San08b, San09d, San09a, San09b, San09c, San10a, San10b, San10c, San10d]. Survey [San11a, San11b, San11c, San11d, SS17a, SS17b, SS17c, SS17d, SS18a, SS18b, SS18c, SS18d, SS19a, SS19b, SS19c, SS19d, SS20a, SS20b, SS20c, SS20d, SS21a, SS21b, SS21c, SS21d, SS22a, SS22b, SS22c, SS22d, Sav97, Set77, Smi88b, TFL23, Tre99a, Tre99b, Tre99c, Tre99d, Tre99e, Tre99f, Tre00c, Tre00d, Tre00a, Tre00b, Tre01a, Tre01b, Tre01c, Tre01d, Tre02a, Tre02b, Tre02c, Tre02d, Tre02e, VBVM22, You73, BP76a, ES80, Miu76b]. Surveys [Sop59]. Survival [Pin85b]. Susanne [Pas95]. Suspected [BK83]. Suspension [LM90]. Sutin [Auc97b]. Suzko [Dei92]. Svatopluk [Ber84b, Han81a]. SVD [GGT13]. Sveltozar [Dur92b]. Sven [Kor85]. Sven-Åke [Kor85]. Sveshmikova [Ren97b]. Sveshnikov [Hoc86, Kur69]. Swamp [Whi83]. Swamy [SFH^{+}66]. Swarm [BT13]. Sweeping [GZ19]. Sweeting [Sin93a]. Swell [PB75]. Swift [Fin95]. Swimming [Bat21, Yat84b, Yat84a]. Swing [MG23]. Swinney [Hol87]. Switched [SWM^{+}07]. Switches [Jef18]. Switching [Bos73b, But61, Coh60, Man81, McM61b, McM61c, Mil59, Mil60, Mil61a, McN59]. Switzerland [LW64]. Sydney [Spi80]. Sylvan [FARE69]. Symbol [DY11]. Symbolic [Boy97, Che86, Ng94, RH76, Mor74b]. Symmetric [BG76, Bos87, Cho87, CD75, Cul78, DG68, Dru64, DU10, FOTG96, Get96, Her80b, Isr92b, Jor87b, Jor89b, KB76b, KLG69, Kla95a, KF76, Los85, Mat69a, Mor64, PR91, Saa83, Sab68, SM84, SS91, Sch67, SPR76, Wil59, ZM22]. Symmetrical [Isr92a]. Symmetries [Car70d, Olv91b, Sch93a, Sch93b, Sch88, Olv91a, Olv90, Rat87]. Symmetrized [DD02, SA78]. Symmetry [GL98, Hea91, Ric63, Ste94b, ZM20, Her97, Ant86a, Ste94a]. Symmetry-Breaking [Hea91]. Symmetry-Independent [ZM20]. Symplectic [BFM95, Iva04, Mey95b, SS16]. Symposium [Ano61f, BD64, Blu65, Cla62, Des68, Jun64, Mil61a, Moo60, Mul64a, Rei63, Man61d]. Synaptic [CSW^{+}08]. Synchronization [CLG14, Izh01, Las76, ZM20]. Synergetics [Haa93, Seg85]. Synonym [BGH^{+}04]. Synovial [ML80]. Syntactic [Ros80a, Sup77]. Synthesis [BH76a, Far85, FS12, Kha87, Kre78, Mul66b, Yau68, Sle67]. Synthetic [Che01, Che18]. Syslo [Bei85]. System [BH76a, BGF^{+}95, Bul67, Car74, Car75b, Car77, Cas82, Cha84, Coh90a, Coh90b, CBS76, Cro85, CZM96, DH09, Den78, Die62, Duf86, Fle76, GQJ96, Giv78b, GC78, GSYY22, Had65, HKS75, Hey86, Hni89, HRM96, Jev92b, Jód96, Jor90c, JH90, Kap64, Kap70b, Kha87, Kla84i, Kla84b, KW18, Kri85, Lev84, LR80, LM81, LLR^{+}81, LRJ78, LL76b, McC79, Mei68, Mor96, Mor83, New87, New76f, NJN98b, NS75, PD76, Phi87, PEGK03, Qui87, Ren92b, Ren92a, Rog83b, Ros76b, SK76a, Sem59, She09, Sle83a, Sma83, Spi71, Sta92b, Sta92c, Stu86, The91, Wes92, Wib87, YS76, Yak95, YH76, dO69, tR83, NJN98a, Bla88f, CS64, Guh66, Jag72d, Kha86, Mee96, Mor75b, Raz87, You72b]. Systèmes [Fal70a]. Systems [AE23, AG80, Arb76, Arb77, Are91a, Are91b, Arm76, AC78, BK95, BG76, Bau88, BGW15, BDD^{+}95, Ber92a, Ber84d, Ber88a, Beu92, BP94, Bon67, BEG23, Bri63b, BT20, BDE09, Bru83, BBKK22, BPJ95, But78, Cam91b, Cam91c, Cam93, CC96, CW94, Cen81, CN96, Cha78, Chi94a, Chi94b, Cle80, CP76, Coo67a, Coo91, CB09, CT95, Cul68, CBM98, Dax90, DGP^{+}16, Dic67, Don90, Dun64, DS61, Eas89, Eas93, Ede90a, Ede90b, Elc76, Ell61, Eng78, Enr78, Erm89, EERG20, FG04, FSV06, Fel78, Fis88a, Fla66, Fle65, FM18, Fou84, GPP88, Gil82, GYS95, GRL98, GT10, Gog85, Goo59, GAT75, Gra86b, Gre97, Gre59, GS23c, Gui92, Hag85b, HIC96, Hal90a, Hal90b, Han91, Har66b, Has92b, Hau89, HYS76, HNP91, HMK96, Her84b, Her84a, Her91]. Systems [Hew76, Hig86, HS95, HPS07, Hou67, JS93, Jef90, Jer95, Jos70, Jür86, Kah66, KHS96, KB76b, Kaz84b, Kaz84a, Kaz86, Kel88b, Kel88c, Kev87, Kir93b, KS90, KA95, Lai91, Lak80, Lau10, Laz91b, LM99, LmYS96, Lin69, LS22, Lio88, Lud75b, Mac91a, Mac91d, Mac91c, Mac76a, MQ18, MS76, MO78, MS96a, MSO96, Mas92, MT76, MK78, McK81, Mer89, Mey75b, Mey86a, Mey89b, Mey89a, MWL97, ML98, Mie17, Mil90a, Mil90b, MP20, MMB^{+}07b, Mor85b, Mor91a, Mor91b, Mos86, Mul66a, Neu94, Neu98, New78, PE76, PW15, PS00, Pit90a, Pit90b, Poh63, PZ81, Pro85, Rad90a, Rad90b, Raz92a, Raz92b, Rei64, Ric67a, RLT65, Riv63, Rob92b, RW95, Roo64, Rus95, Sac83a, Sag81, Sch70a, Sed96, Sel94, Ser76, Sey93a, Sey93b, Sez90, Sha89b, SES75, SWM^{+}07]. Systems [Ske78, Sle81, Sle85, Smi88b, Spi84b, Spi84a, Spi87, Spi91b, Sta75, Ste64, Ste62, Str64, Stu90, Sus88, TtK96, Tab95, Tho76a, TBBER21, Tu76, Tu92, Tze91, Urb93a, Urb93b, VG78, VBVM22, Var67, VdLBF17, Wal95, WS75b, Wes84a, Wie87, WP95, Yan90, Zak84a, ZM22, dBC^{+}08, vLMZ19, MMB^{+}07a, Bri63a, Gib97, Ham77, Stu90, Ant90, Ant86b, Aok70, Apl95, Ath87, Bak91b, Bak92, Beb79, Ber84e, Blo89, Blo91, Buc85, But74a, But74b, But74c, Cha66, Con96, Dau90, Dav91a, Dev90, Dev94b, Duf84, Dur93b, Enr84, Fab80, Fal73, Fel90, Fif86, GP88, Giu18, Gre63a, Gru93, Guc87, Gue82, Gus76, Hag83, Has92a, Hei93a, Her67a, Hig62c, Hoc94, Hol87, Jef95, Joh68, Joh73, Kal65, Kal70, Kap66, Kaz80a, Kaz82, Kaz92, Kee88b, Kir93a, Kni86, Kod96]. Systems [Koh88, Kor95b, Kov96, Kow88, Kup94, Laz91a, Leh63b, Lew91a, Lin90, Luc18, Lud91, Mac91b, Man87, Maw91, McC83, McG65b, Med78, Mei89, Mer85, Mey75b, Mey91a, Mor71, Mor73, Mor75b, Nie89a, Noh76, Old62, Oth82, Ott96, Pal95, Rat87, Rob91, Rob92a, Sac83b, Sac74, Sam76a, San82, San95, Sau79, Saw85, Sch87, Sco88, Sha95c, Sho87, Sle89b, Spi91a, Sto96, Sto62b, Str86a, Svo96, Swa96, Tab95, Ton89, Vak93, Wes84b, Wit90, Woo92, Zak84b, Zak85, Ame85, Boj93, Chu89, Cul66, Lev63a, Per89]. Systems-Analysis [Spi84a]. Systems-Rigorous [Hag85b]. Systems-The [Mor85b]. Systems. [Fro87a, O’M87b]. Systemtheorie [Hig62d]. Szabo [Flo69]. Szatzschneider [Boe89]. Szego [Sac74].

T [Art87, Bal95, Bei78, Ber91b, Ber83b, Ber85f, Bla66, Blu86, Boa85, Boa88, Bor79, Boy88, BG64, Cla62, Dav75a, Daw74, Dev90, Dur92b, Els89, Feu76, Fis93, Fle73, Fok84, Fra95a, Fri82, Gos63, Gre76, Gue89, Hag83, Han85a, Han87, Hau85, Her67b, Hin87, Hoc94, Hol87, Hun81, Isa88, Isa89a, Jag64, Jag72c, Jag80, Jar76, Joh83f, Joh83e, Kal65, Kap65, Kei65, Kel86c, Kev85, Kla92b, Klo93, Kno81, Kum82, LS66, Laz82, Lev84, Lit63, Lon87, Los83a, Lud91, McC92a, McG64, McG65a, MT61, Moi89, Nas87b, Pra86, Pre81, Rad73, Rei64, Rie72, Rog96, Ros72b, Row86, Rüd97, Rue73a, Saf96, Sal63c, Sha82a, Shi97, Sin93a, Spi74, Str62b, Sys63, Van96, Wan91, Wei69, Wen78, Wis60, Wol97, Yan94, Zie76, vdV94, Jev94]. T-Gambler [Jev94]. T. [Pin84a, Zak84a]. T.Naito [Cor92a]. Ta-Tsien [Kov96]. Table [GRS97, Gra72a]. Tableau [Hac75]. Tables [Ask75b, Ask79a, Hey82, Kel96, ZK96, Ask75b, Pap62a, Sau62, Str62a, Str62c, Str62b]. Tabor [Stu90]. Tackle [Ric90a]. Tacoma [GK12]. Tae [Ada97b]. Taha [Rub77a]. Tahir [Jód96]. Takács [Dwa70]. Takahara [Arb76]. Takahasi [Bes92]. Takayoshi [Aok70]. Takeda [Get96]. Taken [Kra96a]. Tal [Eck82, Har72b]. Tall [Win93, Win96b, Win96a]. Tamer [Bin84, Tan93b, Tan93a]. Tamhane [Spu89]. Tammer [Cra84]. Tanaka [Flo69]. Tanino [Cra86]. Taniuti [Bla66]. Tank [Gra91, SR94]. Tannehill [vL86a]. Tanner [Car93]. Tanur [Par92a]. Tapia [Goo82]. Tapir [KS73]. Taqqu [Pel97]. Tarasov [Kla84h, Kla84a]. Tarras [Sez90]. Tartar [Lan86, Nas95]. Tasks [KM12]. Taub [Jun64]. Taussky [Djo68]. Tax [Ste04]. Taxicab [Giv63]. Taylor [Art85a, Asn75, Kar83, Mey93b, Saa96, Sch95, Sop82, Wei77, vL86b, Boa70, CC78, CIS95, Stu86, Vet73]. Taylor-Vortex [Stu86]. Tchebycheff [Ric67a]. Teach [PCH98]. Teacher [GM76b, GM76c]. Teachers [Saw59]. Teaching [Pat03]. Team [KK78, Pac88b]. Teams [Kee93]. Technical [Joh84b, Joh84a]. Technique [Bos96, Bra76a, Com72, CT96, Grz89, Hir75, Pai76, Pug66a, RFS03, Rub78, Tay96, Vis01, WiOH10, Cho70]. Techniques [Ari91a, Ari91b, Ber85b, Blo76, Blo77, Car65a, Cla83a, CBS76, Din94a, Din94b, Fel76, Fla99b, Fla99c, Fla99d, Fla00a, Fla00b, Fla00c, Fla01a, Fla01b, Fla01c, Fla01d, Fla02b, Fla02c, Fla02d, Fla02e, Fla03c, Fla03d, Fla03e, Fla04b, Fla04c, Fla04d, Fla04e, Fla05a, Fla05b, Gor59, Gue83b, Hai85, Ips05, Ips06a, Ips06b, Ips06c, Ips06d, Ips07a, Ips07b, Ips07c, Ips07d, Ips08a, Ips08b, Ips09d, Ips09a, Ips09b, Ips09c, Kel84a, Kev87, Kok84, LC72, Leh63b, LH75, Lud91, Mac77, McC95, Mou67, PSSW06, Raz92a, Raz92b, RM95, SBR08, Spa62a, SBM22, Ste62, Tew77, Vie15, WP88, WS75b, Wol77, ZWGB16, Fla99a, Fla99e, Fla00d, Fla02a, Fla03a, Fla03b, Fla03f, Fla05c, Ips08c, Ips08d, Ram88, Col91, Col87b, Fle64, Foo63, How82, Man70, Seg79, Cla83b, Kel84b]. Technology [Ker92, Tul90, Fle73, Flo69, Ker92, Pfl92, Wil86b]. Ted [Coo67a]. Teicher [Sam80]. Telecommunications [Art88, Beu87]. Telegraph [Pin92c, Ren93a, Ren93c]. Teletraffic [AS80, SW81]. Telgen [Fro83]. Tells [ABGJ21]. Telser [Wet74a]. Temam [Cla78c, Far96, Liu91, Mor91a, Jol96, Mor91b]. Temme [Köl87, Olv97]. Temperament [Gol77]. Temperature [AA78, AR63, CS84, Hey63, Kla91, Tri62b, Rul98, DRWK64]. Temperature-Time [AR63]. Temperatures [Rei66]. Templeton [Beu85, Meh73]. Temporal [Erm89, GGK^{+}22]. Temps [DD74]. Ten [AD76, Lev84, Rat87, Hei93b]. Ten-Node [AD76]. Tennis [NA06]. Tension [BL09, Lew76, Mye98, Pru78]. Tensor [HKD20, Hut89, KB09, Lee84b, SG72, Vil74]. Tensors [GTH23, Gun63, Rat93]. Tenth [Fen94]. Teo [Sac83b, Sac83a]. Teodosiu [Dav75c]. Term [AB73, CS06, EC76, Gau67, GS88, Jor89a, TE96]. Terminal [Kar01, Rei78b]. Terminals [LS95, Boe96a, BGML96]. Terms [Fre86b, Jor87a, JF87, Mei68]. Ternary [BD94]. Terras [Jor87b, Jor89b]. Tessellations [DFG99]. Test [Cle59, LW59]. Testemale [McA92]. Testing [ADPR04, LMDLS23, TTY88, Kup62]. Tests [Cho78, Gho87, JMJ95, Sha97b, Jur95]. Tetrahedral [AD76, Kat78]. Tetrahedron [Hep66, Hep67]. Tetsuzo [Cra86]. Teukolsky [Isa88, Isa89a, Yan94]. Teuvo [Hun85]. Th [Dat80, Wet93, And77, Fol64, Kla76a, Ste59, PRM^{+}66]. Thaler [Old62]. Thalmann [Odg86]. Thanh [Aga93]. Thatcher [Ric74]. Theft [Hul76]. Their [Aga93, Alp85, Ask75b, Ask92a, Ask92b, Bar91a, Bar91b, Beh76, BM93, Bir69a, Cam77, Cen81, Cod78, DA79a, DA79b, DK85, Dra61a, Dri65, Eas89, Eva81, Fel70, Gra70, Haj96, Hey88, Hoc79a, Hol89a, Hor95b, Ike76, Lak80, Lan76b, Lee84a, Lud87, Luk82, May60, McM62, Mer64, MM06, Olv72, Olv91a, Olv91b, Par74, Pay67, Red83, Sha92a, Sha92b, SAH96, Sha04, Sys63, Ton95, Wah92, Was74, Whi63b, Wim65, Xu21, Zlo91, Koh86, SST95]. Thejll [Ren92b]. Them [LB78]. Theme [CJ16]. Themistocles [Neu97b]. Theodore [BR65, Dev94a, Hol70a, She77b]. Theodorescu [Kes75]. Theorem [Ahm75, And66, BD63, Bul67, Cha63, CK76, De 20, Dec65b, DM67, Dra83, Fus68a, Got60, Hei69, Hig02c, Hol70b, JP89, Jor86a, Kee93, Mac00, Maz84, Mor74b, NJN98b, Núñ96, Pon65, RB23, Sac67, San65, SC81, SS79, SK85, She66a, Van98, Wei67a, Zuo96, NJN98a, DMS23, Har78]. Theorems [AH95a, Ama77b, BH15, Bre68, Bry96, CCG^{+}02, Fra86c, Fra89, GTH23, Ike78, Jer76, Kaz67, Kem67, Kre64, Lev90b, Mac68a, PB96, Pin70, Pro91, Sie89, Wil70c, Won64b, Dut82, Coo91]. Theoretic [Boo79, BEG23, Hul76, Ike76, KKS82, Luc87, Pro85, SW93b]. Theoretical [Ano64i, Bel90a, Duf91a, FS78a, Hau85, LRJ78, Mir88, Ans80, Ham90, Hop79, Kan66, Par67a, Rei77c, Wei80b]. Theorie [Hig62a, Man61c, Lan86, BI68b, Dan72]. Theories [Coz87, Gue89, Lin86, Pat64, Tau80, Ula64, WK78, Bel90a, Jor89c, Pro84c, Sal64]. Theory [Ado89, Aeb94a, Ame86b, Ano61f, Aok70, Arb76, Arb78, Ari96, Arm90a, Arm90b, Aro69, Ars62b, Art91a, Art91b, Ask79a, Ass77, Ath76, Ath93b, Ath93a, Auc97b, Bak91a, BBN98, BT69, Bar95, BW76, Bax85, Baz73, Bei85, Bei96, BKS68, BK69, Bel90a, BI68a, Ber94, BBC11, Beu92, BR65, BR67, BR76, Bir64, Bir69a, Blu94b, BBKV96, Boo94, BBK20, Bou91b, BGF^{+}95, Bra85a, BSP95, BFM95, Bru72, BBKK22, Bry96, Bue71, Bur79, Bur81b, BS92, Car70e, Cas82, CHC95, CIPB95, Cha70a, Cha94d, CLOT99, Cha64b, CQRZ96, Che73, Chi90, Cho70, Chu82, Chu88b, Chu96, Cla62, Cla64b, Cle61, Coh76, Coh82, Coh84b, Coh84a, Col94a, Col86, CCM00, CCM01, CK18, CT76b, Con85, CL87, Cul77, CZM96, Cus96, Daa97, DC86, Dau90, DeW85, Den67, DM77, Deo64, Deu63]. Theory [DM67, Día78, Dij90, Dod73, Dol91b, Dri91, DTT98, Dri75, Dur84b, Dur84a, Dwa66, Dwa70, Dwa80, Dys66, Eas89, Edw63, EFMM13, Efr79, EKS97, Ell84, Eng17, EKZA96, Ewi77, Fal73, Fan82, FLM22, Fat93, Fej88, Fig93a, Fig93b, Fle76, FGA97, Fri80, Fro80, Gat88, GM95, Gla96a, GI97, God87, Gog85, Gol65a, Gol66b, Gol78, Goo72, Goo91b, Gos64, Gos79, Gra18, GAK17, GM85, Gra83b, Gra94a, Gra94b, Gre63a, GLO20, GM66, GK95, GK98, Gri61, Guc79, Gud84, Gun88, Gut97, Gut90, Hab83a, Hab83b, Hag85b, Hag85a, HY71, Hal90a, Hal90b, Ham90, Ham73b, Ham73c, Har60, Har64a, Har74, Har64b, Har70, Har76a, Har63, Har82, HRB96, Hau90, HHOT93, Hea91, Hel97, Her71, HK93, HR62, Hig65, Hoc66, Hoc80, Hoc92, Hoc94, Hom86, Hor83a]. Theory [How63, HRM96, Hul71, Ism83a, Ivi65, JS93, Jev93, Jev95a, Joh84a, Joh87, JH98, JMH95, Jür86, Kac62, KK78, KS95, Kan62c, Kap64, Kap67, Kap78, Kar78, KHS96, Kaz82, Kaz83, Kaz86, Kaz90b, Kaz91a, Kee95, Kee88a, Kel79a, Kel88b, KMH66, Kem67, Ken74, Kes69, Kha86, KG97, KN61b, Kno81, Koc92, KM14, Kot65, Kra24, Kri10, KG98, Küç92, Kui06, Kul69, KTSB21, Kur69, Kus76, LaS64, Lag89, Lau10, Laz82, Laz91a, Laz91b, Leh63a, Lei66, Ler63, Let69, Lev88b, Lev97, LmYS96, LW13, Liu92b, Liu94, Liu95, Luc71, Lud75a, LBL02, Lum85, Mac76b, Mac68a, Man61c, Man81, Mar66, Mar64, Mar85c, MH96, MMS98, Mat80, Mat94b, Mat94a, MQ69, May91, McA92, McG65b, McG61, McM61a, Mee96, MS96b, Mey75b, Mey75c, Mey76c, Mey79, Mey86b, Mey86a, Mey89c]. Theory [Mey89a, Mey95b, MWL97, MBK97, Mil61a, Mil90a, Mit87, Mla97, Mor97, Mor75b, Mor82, Mor83, Mor89c, Mos66a, Mos86, Mul64a, Mul69, NC76, Nas87b, Nas86, Neu94, New76e, NC09, Noe81, Noh62, Noh71, O’M86, Odg86, Ost61b, Pac86, Pac87, Pas95, Pen95, PBL98, PB96, PE96, Pic87a, PFM76, Pim93, Pin72a, PI88, PR70, Pro65a, Pro65b, Pul91, Ral62a, Ral83b, Ral83a, Rei62a, Rei87, Rhe85, Ric90b, Ric90c, Rin61, RM75, Rog94a, Ros65, Ros72a, Rox84, Rub77a, Ruo91a, Rus78, Saa96, Saa68, Sac74, Sag81, Sam80, SS79, Sch77a, Sch82b, Sch84b, Sch84a, Sch78a, Sch78b, Sei93, Ser04, Ser64, Shi71, Shu80, Shu85b, Sie69, Sie61, Sim89, Sle76, Sle81, Sle67, Slo97, SWC96, Spi82, Sta66, SF97, Ste90b, Sto96, SS96, SY69, Stu91, SH94]. Theory [Sus79, Sut95, Swa78, Swa96, Sys63, Tek89, Tew67, Tho77a, Tho60a, Tho69, TW69, TH96, Tru94, Tul68, Tur60, Uhl73a, Uhl73b, Urb93a, Urb93b, Van94c, Van94b, Var85, Var66, Vas94, Wal79, Wal93a, Wat10, Wei68, Wei79, Wei63b, Wel89, Wen78, Whi66b, Whi81, WC96, Wil70a, Wil84a, WJV79, WM15, Wim88, Win74, WA94, Wou68b, Wu75, Wyn69, You71, YH76, YL75, Zak85, ZBK96, Zie83b, Zlo83, Zwe84a, dJFO98, vH87, Auc97a, Cor16, Dut82, Gla97a, Gla98, GJ97, Ber97a, Cou97, Lie97, Aeb94b, Aga93, Ano61a, Ans64a, Ans64d, Ath94, Bei78, Ber85e, Bin84, Bjö61, Bla88b, Bou77, Bre62b, Bru73, Buc69, Cha68, Chi94a, Chi95, Chu85, Cul68, Dan91, Dat69, Day87, Des68, Deu90, Dia59, Dol91a, Eat96, Edw90, Fel72, Fis93, Fok84]. Theory [Fra86c, Gos84, Gra86b, Gre87, Hag89a, Hag80, Hag82, Har73, Har83, Har96, Her69, Hes86, Hey93, Hin91, Hir73, Hud95, Ism83b, Jd93, Jos95, Kap70b, Kar90, Kaw66, Kaz75, Kaz90a, Keh89, Kei65, Kha88, Kie94, Kus72, Lae82, Lak80, Ler64, Man89, Man63, Mat88, Maw95, McG65a, Meg88, Mer64, Mil83a, Mil90b, Moo76, Moo77b, Mor96, Mor88d, Mos93, Nel74, O’M87a, Olv86, Pac88b, Paz76, Pil87, Pit79, Ram93, Rei71a, Rei63, Rog94b, Ros63, Rox76, Roy91, Ruo91b, Sag86, Sch65a, Sch81a, Sch78c, Sei62, Ser86, Set63, She77b, Shu85a, Sle94, Smi82, Smi86, Spi87, Spi74, Spi80, Sus88, Syl96, Tho76a, Vai96, Wal63a, Wil84b, Yak95, Yan92, You72a, Zie78, Zwe84b, Bla88d, But89, Cra86, Dat78, Elu88, Fle82, Joh67, Joh84b, Kap85b, Kee61, Kel90b]. Theory [Koo77, Kre96, Lae76, Lil67, Mey82a, Pap62b, Pap63b, Ral62b, Rei77b, Rei78c, Sam79, Sch73a, Sho69, Vak93, Wou64b]. Theory-the [Nas87b]. Therapy [GAM93, SFOM99]. There [Gea81, Hil98]. Thermodynamic [DBS^{+}07, Len80]. Thermodynamics [Man90b, Sle86, Coo65, Kes85, Man90a, Owe85, vH87, Kum82]. Thermoelastic [Rog69a]. Thermoelasticity [Kno81, Kno83, Ros87]. Thermoelectric [Sta75]. Theta [And74a, And83, Car73b, Jor84b, Sal63a]. Thickness [Bar64]. Thierry [Pin84b]. Thin [BJ75, Bas86, Cla62, Cla64b, Gee78, Kon03, Mac68a, MK99, Mye98, OT12]. Thin-Layer [OT12]. Things [Lar69b]. Thinking [Efr79, OL98]. Third [Abd91, Amo65, For92, Mat69b, Nas87b, SLF66, TS90, Har64b]. Third-Order [Abd91, For92, TS90]. Thirring [Wes84b, Wes84a]. Thoft [Spi87]. Thoft-Christensen [Spi87]. Thomas [Bei80, Beu92, Goo86, Gos63, Grc18, Kie94, Mey76c, Mey95b, Min66, Nas90, Par88, Ral66a, Shu85b, Sin69, Stu94b, Zak84b, vL86b, Stu94a]. Thomason [Ros80a]. Thomée [Ode86]. Thompson [Ber83b, Goo82, Hol87, Kir90a, Sal63d, Str62c, Kir90b]. Thomson [Wol86]. Thornley [Ant91]. Thrall [Fan59, Fia73, Kla84i, Kla84b]. Three [Abr84, AB73, AL98, Blo73, Boa87, BM98, Car76a, DB70, Ell02, Gau67, Gol69b, GK81, GRK81, Gro92b, Gue61, Han98, JF71, Kla89, KR74, Knu98, Kos79b, Kuc87, Mac92, Mac98b, Mar21, MTGF78, Meg78, Meg79, Meh74d, Mor85a, Neu96, Pré95b, Quo75, Ren80b, Ric76a, Sch92, Sei97, Sei76, Sit92, Stu64, Win90, ZW94, Mor86c, Böh88c, Bre77, JG93, Kno81, Mey90, Mey91b]. Three-Coloring [Meg78, Meg79]. Three-Dimensional [KR74, Kos79b, Mac92, MTGF78, Mor85a, Quo75, Sei76, Sit92, Stu64, Win90, Mor86c, Kno81]. Three-Dimensions [ZW94]. Three-Edged [GK81, GRK81, Ren80b]. Three-Term [AB73, Gau67]. Threshold [CK76, Mul66b]. Thro [McA96a]. Throws [GL05]. Thuijsman [Fin93]. Thulasiraman [Whi83]. Thumb [Hig02c]. Tian [BKS96]. Tian-Ming [BKS96]. Tibshirani [McC92a, Siv94d, Siv94b]. Tidal [DC76]. Tikhonov [Hoc86, Nas86, Wil79]. Tim [Guc79]. Time [AA78, AW02, Ang92b, AR63, Arn92, AW19, Bad60a, Ban87, Bar65, Beb68, Ber84e, Ber84d, BC76, Blu65, BDC98, Col84, Del95, Den76, DGK^{+}12, DJLQ21, Ell80, Ell85b, Feu76, Fre75, Fri79b, Gan98b, GM81a, GM81b, Giv63, Gol65a, GST01, HIC96, HS70, Hew76, Hin87, IWMW21, JSF95, Jev92b, Kar01, Kel82, KR78, Kru83, KTSB21, Lai91, LL76a, Lev90a, Lev91, LL75, Lew83c, LL69, LH19, Mac91b, Mac91d, MM81, McC83, Mor78, Mor75b, New59, New62a, New64b, Par76, Par92b, Pea81, Ser84b, Ser84a, SK20, Sho99, Smi97, Sno76, Sül97, Szy87, Tad87, Tu76, Tze91, Van76, VB75, Wil77b, ZD95, ZN62, Bel10, Gan98a, Bou91b, Dun94, Lud77, Oco87, Con96, Fin95, Her67a, Kaz80b, McC83, Rob94a, Yak94]. Time-Dependent [Ban87, Bar65, Del95, Den76, GM81a, GM81b, Giv63, Kel82, LL76a, Smi97, Tad87, VB75, ZD95]. Time-Differencing [DJLQ21]. Time-Discrete [Ber84e, Ber84d]. Time-Discretization [Oco87]. Time-Frequency [AW02]. Time-Harmonic [Col84]. Time-Lag [Her67a]. Time-Series [AA78, Lev90a, Lev91]. Time-Stepping [Arn92]. Time-Variant [McC83]. Time-Varying [HIC96, Hew76, Con96]. Time-Volume [Ell85b]. Times [Gre67b, Har78]. Timescales [Ver15]. Timman [Col87a]. Timothy [McA96b]. Tina [McA88]. Tinbergen [Hes86]. Ting [Gun95]. Tinsley [Bal88b]. Tippe [BRMR08]. Tipping [Hep66, Hep67]. Tismenetsky [Joh87]. Tissue [Ada91]. Titchmarsh [Wal76b]. Titli [Sez90]. TLM [CZ96b, Zem96]. Toads [Mil91a, Mil91b]. Toby [Sin79b]. Tod [Lou92]. Today [For67, Ham80]. Todd [Bél88, Jun64, Leh82, Row79]. Toeplitz [Cha95d, CN96, Cha95c, Cox94, KVM78, SM14]. Tokamaks [Joh91b, Joh91a]. Tolimieri [Kel92a, Kel92b]. Tolls [Fre86c]. Tom [Bro65, Kla63a]. Tomasz [Mac91a, Mac91c]. Tome [Ost61a, BI68b]. Tomescu [Pil87]. Tomography [CIN99, Fin84a, Fir97, Kei89a, Kei89b, KMS22, Min78, MWWC96, SW11, Col96]. Tomography. [Fin84b]. Tompkins [Jun64]. Tong [Sta92b, Sta92c]. Tony [Ell92a, Ell92b]. Tool [GP92, She91, Wal97]. Toolkit [Sch94]. Tools [Ame86b, Sal64, Fla97, Car93]. Topics [Art88, Beu87, Boj97, Got78, KB59, Lai91, Ren94d, Ros72b, Tze91, Van94b, Whi66b, Ros68b, Ban87, Fos90, Hig65, Kap70b, Kap72, Lor91, Muk86, Ste93b, Van94c]. Topographically [Dre75]. Topological [FT11, vV65, Lev88a]. Topology [Dev91, WB09, Wig90, Fad62]. Topology. [Pat87]. Topping [Bjö61]. Tops [BRMR08]. Torgersen [Le 92]. Torgny [Kal93a, Kal93b]. Torn [AKP93b, AKP93a]. Tornheim [Fan59]. Torpedo [Wol60]. Torso [WiOH10]. Torsten [Wib87]. Tortoise [Emb03]. Torus [JHSB22]. Tosio [Wil70a]. Toss [DHM07]. Tossing [FB78, NW61, New77, PF62, Pul64, She63a]. Total [Fie93, Nie94b, PCLS23, WM61a, WM61b]. Toth [Van93]. Tou [Kal65, McG65a]. Tour [BS98, CKCH23, Ham62a]. Tournament [Knu86, Knu87, MP70, KL87]. Tournaments [Ano79e, Meg78, Meg79, Sch66, SAT22]. Touzot [Mea86]. Townsend [Cor79]. Toy [Som79]. Trace [BW96b, NBS12, Roe95, BW96a]. Tracer [MS75]. Tracer-Kinetic [MS75]. Traces [GZ19]. Track [Fly89a, Lew88, Fly89b]. Tracking [Bor62, JS76, Lew81b, Lew82b]. Tractable [PdHM16]. Trading [CFHP10, CJR18]. Traffic [BD11, Fre79, Lag82, Los83b, New65a, Pat64, vLMZ19, Hey86, New87, Hey89]. Trailer [Ale84, FL81, HP85, HPF85]. Trailer-Truck [Ale84, FL81, HP85, HPF85]. Trailing [GBF78]. Training [Ano75g, Car67e, HCH^{+}75]. Trait [SAT22]. Trait-Performance [SAT22]. Trajectories [HH00]. Trajectory [DMT^{+}18, Kel17]. Tranform [Bai93b]. trans [Ahl93, Bak91b, Cav90, Coo91, Dei92, Dij90, Eas91, Gui92, Kaz62, Kaz91a, Lev88b, Mal90, May90, McA92, McN89b, Mei89, Nas90, Nit90, Ock91, Pao88, Pap62b, Pul91, Ral62a, Sco88, Sim89, Tot91, Tu92, Tul90, Wil93]. Transcendental [CP89, CPRG90, EL85, EL86, Jor86d, LS02, RGJ90, Rud61, Wim67]. Transfer [Bar64, Bar67, CGT^{+}97a, Eds05, GZ19, Hoc83b, Hoc83a, May90, Pra89a, Vel71, Wil70b, CGT^{+}97b, Kev85, Pra85, Pra86, vL86a]. Transferable [Ste04]. Transform [Ang94, BS91, BLT08, Boh88b, CZ94, Che68, Che73, Daa97, Deu87, ES08, Gla81c, Gla88c, GJRW24, Hey97, Hir73, Hun76, JGK^{+}82, Kel92a, Kel92b, KN61b, Lei66, Low61, Mac67, McC60, Mcn66, Mil95b, Mil95a, Nie86, OW89a, Rao72, RY88, Sho66b, Smi71, Sta88, Sto61, Str99, Tru94, Wel89, WY70, Wit22, Gro98b, Gro98c, GL04, Ano73a, Bai93a, Fro60, Gas63, GRYL89, JBL72, Jag82, JB88, Lan67b, OW89b, Sco83, Ste67b, WS71, Wel88, Zal83]. Transformation [Abl86, Bus63, Cli76, GKLU09, Hof71, Kla62b, MCS69, Wah03, Smi87]. Transformations [Abl80, BRZS18, BCV98, Chu94b, Chu94a, GGK96, Kei64, Lee84a, Ler63, Luk82, MTGF78, OHBG00, PY98, Rus90, Van75, Ypm83, Zem68, Ber81, Kor73, Wis60, Tut94]. Transformed [Noh77]. Transforming [Na67, Na68]. Transforms [Ask75b, CW94, Cor72, Fen68, Gib73, Gol62, Har67c, HW89, Kap67, Kei89a, Kei89b, Koh86, PD97, Rei59a, Reu67, Ric76a, Smi72, Sto62a, War98, WM63, Ask75b, Mey94a, Sta89]. Transient [Rei66, ZWY14]. Transionospheric [Che18]. Transistor [Hig62b]. Transition [Ang92b, BR69, Ell92a, Ell92b, dBL75, Are93]. Transitions [FM18, Jef18]. transl [Gue92]. translation [Cuv90, Sal63b]. Translations [Ham65b]. translator [Ker92]. Transmission [Bri63a, Bri63b, BK87, CZ96b, Gla93d, Gla94c, Jar76, Lar94, PBL98, Rai59, Ran76, Wil64a, Zem96]. Transmission-Line [BK87, CZ96b, Zem96]. Transmutations [Tut94]. Transonic [Che89]. Transparent [Sei78]. Transport [Ans64a, Bar65, Bar70, CP18, FL18, Gan89, Gan90, Gla95c, HH16, KMC96, Pal96, Pri76, RBZ23, Saw85, Uhl73a, Uhl73b, VN76b, WZCX12, Whi93, WP96, Zwe84b, Zwe84a, BC96a, Reh89, Win74, Cer96]. Transportation [BI68c, BS09b, Jar76, Lee86b, TS86, Wal85, Wal86, Wal87]. transposed [BG96]. Transversal [Bru73]. Trapezoidal [For21, TW14]. Trapp [Moo77a]. Traps [IWMW21]. Traub [Nov94, Pac86, Shu87, Wil66]. Trave [Sez90]. Traveling [BDvD^{+}98, Gao18, PR91, Pat03]. Travelling [BL84, RK75]. Travelling-Wave [RK75]. Travis [Kee88b]. Treasure [Ano86l]. Treasurer [Ano75p, Ano77r, Ano78r, Ano79m, Ano80p, Ano81k, Ano82q, Ano83n, Ano84p, Ano85k, Ano87l, Ano88j, Ano89m, Ano90m, Ano91l, Ano92i, Ano93j, Ano94g, Ano95q, Ano96r]. Treat [Hor78]. Treatise [Boa65, Har61a, MT62, Wim86, Her77]. Treatment [Bal88a, Buc85, Dea89, Gos83a, Hoc92, HSSZ98, Lev65, Pre81, Rei78c, Dea90b, Slo97, Gei89]. Tree [Arm90a, Arm90b, Cah77, De 20, Chu78, Meg88]. Trees [Mor78, Mye75]. Trefethen [Cor16]. Trend [KKBG09]. Trenogin [Lae76]. Treuenfels [JS71]. Tri [Hea81, Ros80b]. Tri-Diagonal [Hea81, Ros80b]. Trials [Sab68, SM84, Ser78, Tak79, KLG69, Los85]. Triangle [Blu78, Bol91b, Bol91d, Gro93b, Kla77, Kla78a, Kla79, Kla80, McC03, New71, New72b, New92, New94b, Rei66, Sch77b, Ham77, Gro93d]. Triangles [Bur82, Deu95, Erd83, HH00, Sed91, SK85, Ste67a, Jor86a, Mil69a, Zuo96]. Triangular [DU10, Hig87, Kat76, RP96, Hor97b, Pea97]. Triangular-Block [RP96, Hor97b, Pea97]. Tricks [Boy08]. Tricomi [Gur62b, CC78]. Tridiagonal [Cho72, PH78, Rue73a]. Tridiagonalizing [CG61]. Trigger [ZWY14]. Trigiante [But89]. Trigonometric [Ano91b, AF69, Boa65, Boe95b, Böh90, Gla94b, Gla95f, Gro91c, Kim65, Lyn67, NS63, Sch65b, Ste68b, WBGP91, Jam59]. Trigonometrical [GRL87, GRLR87, HL86]. Triple [Cle59, Gla80, GR81, Gla81a, Jor73, Jor83a, KH72, Mer82, Mer83]. Triple-Diagonal [Cle59]. Tritter [Amo64, DRWK64, MS65]. Trockel [Ell85a]. Tropical [ABGJ21]. Trotter [Fis93]. Truck [Ale84, FL81, HP85, HPF85]. Trudi [Hou74c]. True [Cla78b, Cla79b]. Truemper [Kun93]. Truesdell [Sle86]. Truncated [MBOV17]. Truncation [And83, Ber75, KTSB21, Jor84b]. Truncation-Based [KTSB21]. Trust [WS13]. Trust-Region [WS13]. Truth [ES08]. Ts [Sim89]. Tsai [Wri60]. Tsai-Hwa [Wri60]. Tschauner [Hig62a]. Tschebyscheff [Kra80a]. Tsfasman [Wol93]. Tsien [Kov96]. Tsynkov [Che18]. Tsypkin [Ath87]. Tu [Gue88, Sel94]. Tube [Cel78]. Tucker [Cot94, Liu88, JP89, MW76a, TT94]. Tuckey [Loe95]. Tuckwell [Bel90a, Gus90a, Gus90b]. Tudor [Her97, Lee84b]. Tumor [RCM07, WHMN16]. Tuncer [Pra89a]. Tung [Ame91a, Mor74b, Ame91b]. Tuning [Goo91b]. Turbo [Sch94]. Turbulence [BC98, Con94, Maj91, Moi89, Bra85a, Gli91a, Lum85, Bjö61, Moi89]. Turbulent [Cor79, MQ18, May90, PL78]. Turbulent-Shear [PL78]. Turkington [Fab87]. Turn [For96, Mil98, Wol94c]. Turnbull [Ars62b]. Turner [GC69]. Turning [Olv86]. Tutorial [Buc91b, Che01, GN22, Neu98, Wal93a, vLMZ19, Buc91a]. Tutz [Mar95]. TV [UFW17]. Twelve [Ham80]. Twenty [MV03, New66, New68b, PW08, RKDD14]. Twenty-Five [MV03]. Twice [BSS14]. Twice-Ramanujan [BSS14]. Twist [Moe17]. Twists [CKCH23]. Two [AH95a, Ale95, Ano91m, Arc75, AGR83, AG83, BJ79, Bar77, Bec74, Ber90a, BR94, Bil93b, Bil94b, BV96, BGML96, Boe91, Boe96c, Bos87, BK66, Bre96, BD63, CS06, Car70b, Car71d, CM78a, CMSM79, CZ92, CS75, Cop07, De 79, Djo68, Elv75, EFA76, Eng03, Eva90, Fis88b, Ger62, Gla93d, Gla94c, GP79, GdMGR91, Gro91b, Gro91c, Gud75, Hal86, Han76, Her90, Her70, Hin89, HMO97, HMH00, IA75, KJH87, Kel89, Kla77, Kla78a, KB68, Kra96b, KS84, LS95, LR00, Lop94, Mas92, Mil66, Mor66, MCS69, NF16, New76f, New78, O’M75, O’M08e, Rai59, Ral70, Rao82, Raw83, Red80, RAH^{+}90, Rob95b, Rob95a, Ros72b, San60, Sch69a, Sch71c, Sch96, Sch76b, Sha78, SJ79, Sho66a, Siv68, Sne67, Str65a, Tau67, Tho59, Urb93a, Urb93b]. Two [Van78, Vas97, VB05, Vei67, Ver90, War69, WS75a, XRTM95, ZD78, Boe96a, New84, Ano91b, Bai74b, Bla88f, BdD94, BKS96, Bus95, Caf92, Fet80, Gas76b, Gla93b, Gro91a, Hol66, Jag79b, Jor79, Jor84c, KJ87, Lar94, LSB67, MR71, Ren90, SM79, SS70, Sle92, VBdD91]. Two-Allele [Cop07]. Two-Dimensional [Bec74, Ber90a, GP79, Mas92, New76f, VB05, ZD78, Gas76b, Hol66]. Two-Locus [Cop07]. Two-Parameter [BV96]. Two-Person [Ros72b]. Two-Phase [CS75]. Two-Point [Ale95, Eva90, GdMGR91, Gro91b, Gud75, Her70, IA75, MCS69, O’M08e, Red80, Sch76b, Tho59, Van78, WS75a, Bai74b]. Two-Term [CS06]. Two-Valued [BD63]. Type [Aga93, Bla88f, BG98, Bra96, EK87, Far60, GG73, Gro88a, Gro93d, Gro93b, HB96, HLvL83, Lev65, Los74a, Los89, New92, New94b, Par64, Per65, Sch59, Sou76, SS78, TS66, Tur68, Wal76b, War68, Sch62a, Cha00, BI70b, Hes83, Pro65c, Wim81]. Type-II [Cha00]. Types [Bus69]. Tyrell [BI70b]. Tyrrell [Kle71]. Tyuptsov [Dij90]. Tzeng [Arm90a, Arm90b]. Tzschach [Din72].

U [Ahl93, Ber91a, BR67, Don90, Fle88, Ism76, Kee95, McC95, Neu94, Ral83b, Sac83b, Sac83a, Wal79]. U.S [Por86a, HHSS75]. U.S. [YAKM20]. Uhlenbeck [Par87]. Ulf [Dwa64, Pin87a]. Ullman [Gra70]. Ulrich [Don85, Ral83a, Ruo92a, Ruo92b]. Ultrawideband [CO07]. Ul’yan [Ock91]. Umberto [Lar74]. Umbral [DF98]. Unbounded [Lea94]. Uncertain [SHL16, KS94]. Uncertainty [Bal20, Bar89, BCH14, MQ18, McA88, OSS^{+}13, Pan93a, Pan93b, PWG18, Shu87, WHZG24, ZMPK24, Sle83b, McA92]. Unconstrained [Den82, Fab80, Han86b, Hil75, Man70, Pow70]. Uncontrolled [Zar65]. Uncoupling [Mey89a]. Undamped [Jen75]. Underdetermined [CP76]. Undergraduate [Boy08, Kap68, LCK^{+}17, Ral86, SKK20, TPCC^{+}11]. Undergraduates [Sha85]. Underlying [Den67, LC72]. Undersampling [SW22]. Understand [BW11]. Understanding [LG95, MC21, Orb18, Wat82, Xu21, Gas95, Kar63]. Undiscovered [KB76a]. Unequally [Str75]. Unexpected [JM92, SR93b, SR93a]. Unfolding [SBXD06, McG93]. Unfriendly [FS62, Fri64]. Ungar [All78]. Ungleichungen [Wil67a]. Unidirectional [Mar81]. Unification [NW82]. Unified [AH95a, Bag95, Bom02, Cha70b, Gar00, Gud75, KL97b, Mor75b, Nas77b, Pic87a, Red90, Spa62b, Wit90, KL97a]. Uniform [CO07, De 00, Dre76, Ful76, Joh83b, Lud70, Mal68, Mal69, MP06, MN76, New62d, Tra66, Wil64a, Joh83d, LeV77a, Ste87]. Uniformly [Hni89, Jor90c, JH90, JGM95, Jor95, Kla94b, Kla95b, RJ96, Jor96]. Unifying [FLGM23, Kus19, Sin89, VDC22]. Uniprocessors [Kar96]. Unique [BPV97b, Car71b, Gla93a, Mey93a, Mey94b, New70a, Ng96, WJ94, WR94b, WR94a, BPV97a, New84]. Uniqueness [Boö86, Bul67, Faz96, GKR20, Jub70, Lee86c, Mac68a, Ral69a, RT97, RO66, Sac67, Sch64, Urq69, Wal76b, Wan97, Wil67b, Zah80, Ren97a, Ver98]. Unit [Blo95, Kla70a, Kol91, Kol92, Lei78a, Lei79, Men94b, Pet78, WB09]. Unitary [BS62, BG64, Mat69a, RM89, Mei77]. Units [Ber60]. Univalent [Boa87]. Univariate [Boy13, CBG75, Dub76]. Universal [BLMT18, NT78, UFW17, Ang90]. Universe [Jaf84]. University [Bor06, Hig02b, Vas94]. Unix [Kri85]. Unknowns [Mei68]. Unperturbed [PY00]. Unsolved [Cle61, Erd72, FS73, Kla92b, Kla92b]. Unstable [BLMT18, JR08]. Unsteady [Bal20, GM66, GJ19, Mar75b, Ric60, Ril75]. Unstructured [Dax90]. Unsymmetric [VG78]. Untangled [Lee97b, Lee97a]. Unthinkable [Efr79]. Untitled [Cha95e, Ips12b, Sch02b]. Up-Down [Gou68, NW67a]. Update [AC78]. Updates [DS79, Tod84]. Updating [Hag89b]. Upper [Deh89, Gre62b, Kor91, Sha64b, Tre70, Van98, JS71]. Upstream [HLvL83]. Urban [Ber95, BS09b, Kaz82, Mac76a, MN83]. Uri [Isa89b]. Urine [KMH66]. Urn [BW65, Wag63]. Ursula [Hey85]. USA [Hig00, Hig02c, Nik78]. Usami [Fri82]. Use [Gou10, GDR78, HYS76, Hue96, Hun76, KL75, Lyn83, Mil76, Par64, Par61, PC89, Ral65a, Ral69b, SS65, Tew77, VB75, Van75, Ang90, PC90b, Dra61a]. Used [Cle63, Zwe16]. Useful [Bus76a, Kot65, Tri78]. User [Bor06, Flu93a, Flu93b, Par81a, Pin85a, Sei78, SG79, Spä97]. User-Transparent [Sei78]. Users [Kri85]. Uses [Fel70, Len93c, McM62, Rap73, Len93b]. Usikov [Bry93]. Using [BW11, BDO95, BBK20, Bun83b, Bun83a, Cha95b, CT95, DM07, DEPC95, Ehr76, Els91, Gal95, GG99, GVL97, Gar75, GS22, Hir78, HF78, Kel88b, Kin94b, Kin94a, LDG95, Lev90a, Lev91, Maz84, NS77, Nie76, Olv91a, Olv91b, Orb18, Pat03, Ric88, Sch78a, Sch78b, Sed91, SK20, Shr76, SM94, ST98, SB72, Ste98, Sto76, Tab95, VLPR20, dEJL07]. Uslenghi [Mil80]. Usmani [LJW73]. Uspenskii [Ham77]. Usually [Lat23]. Utility [Kar90, Mor75a]. Utilizing [RA63, Sta75]. Uwe [Kod96].

V [Aga95a, And75a, Arb77, Aro69, ABG96, Bak91b, Bak92, Bau88, Ber85e, BR85, Bla88b, Ble81, BSW82, Bre62b, Bry94, Bry96, Buc69, But89, Can86b, Cas87b, Cha70a, Chu85, Cla64b, Con85, Dat73, Dat80, Dij90, Dol72, Dri65, Eng94, Erm89, Eva85, Fen94, Fle83, Fro86, GC69, Gol65a, GG91, Gos83a, Gue83c, Gue88, Gun88, Hai77, Ham77, Her67b, Hig62c, Hir73, Hoc80, Hol89a, How63, How83, Jag80, Jor80a, Jor90c, Kan66, Kaz91a, Kel67, Kla84h, Kla97, Kno81, Kol63, Kre96, Kus74, Lae76, Lae83b, Lag97, Lem85, Lev96, LG89, Los85, Luk75, Man70, Mar73, Maw91, Mec92, Mey77b, Mey89b, Mik89, Mil81b, Nag96, Noh71, Obe64a, Obe64b, Ock91, Olv93, Orb18, Par90, Per65, Pro83b, Pro83a, Pro65c, Ral62a, Ram93, Rei62a, Rei77c, Riv60a, RFC72, Ros67a]. V [Rox92, Rue79, RLG81, Run91, Sel94, Sib95, Sop69, Spä97, Ste90a, Sto66, Sym92, Tap86, TA90b, Whi63b, Wil60b, Wil79, Woź87, Xu96a, Yak95, Yan65, You74, Zie84b, Zit63, Lew76]. V-Splines [Lew76]. V. [Kla84a, Lae83a, Pro83a, Zie84a]. Vacillation [Jin91]. Václav [Fro87b, Kap66]. Vacuum [Sha90]. Vadim [Edw90, Koh88, Hau90]. Vaillancourt [Mey94a]. Vainberg [Lae76]. Vajda [Fro73, Gri61, Sop62, Win60]. Vakhania [Mar83b, Mar83a]. Vakhutinsky [Dau90]. Valentin [Bry96, Par80]. Valeria [Aur18]. Validation [GHW76, Szy87]. Validity [Say13]. Valuation [Cah77, Chu78, Hig02a, Kri97, Ruo92b, Ruo92a]. Value [Ale95, Ann95b, Ann95a, Ans68, AGR83, AG83, Bai74a, Bai74b, Ber84b, Ber84a, BJ76, Blu86, Blu91, Boa82a, BF82, Boy08, Bus76a, Cla67, Con66, Day89, DD02, Faz96, Fel76, Fel92, Fet81, Fix74, Fus68a, Gea85, Gei89, Gla89c, Gra67, Gro93a, Gud75, Gue69, Gun88, HJ85, Her70, Het72, Hig86, HM73, HA75, Isa89b, IA75, Joh75, Jor90b, JG90, Jub70, Kan85, Kap81, Kel79b, Kla62b, Kla64a, Kla65, Kla71a, Kla84e, LS66, Lar86, Lau74, LP75, Lev88b, Los81, Mal93a, Mal93b, Mat85, Maz84, Mey80b, Moo65, MCS69, MMH04, Mur90, Na68, NC76, O’M71, O’M88a, O’M08e, Obe64b, Pon90a, Pon90b, Pro87a, Ral66b, Ral70, Rao82, Red80, Red76b, RT97, Rog75, Roo10, Ros76b, Rus90, Sar84b, Sar84a, Sch61, Sch76b, Sco91a, Sco91b, She61, Shr76, Ste93c, SH94, TE96]. Value [Van88, Van63, Van78, Van75, WS75a, Wen68, Wil88, Wil67b, Wil73, Wil78b, Wil67c, Win67, Zay92, ZC74, AR81, Boe98a, DGH^{+}18, Ren97a, KH85, Rue72a]. Valued [BT20, BD63, CTW21, WHZG24, Sch81a]. Values [BT19, Bor79, Ike75, Rog78, Shi68]. Vanderbauwhede [Ant86a]. Vandermonde [Los75, Alp96, BNT21, Ven74, Ven75]. Vandewalle [Fie93]. Vanishing [Gla84c, Gla85d]. Vaporizing [WB76]. Varberg [Kle76]. Varga [Fix93]. Variable [Abd83, Abd85, Bal85, BL09, Buc69, Can86b, Cho70, Chr68, Gre62a, Hoc80, Loc84b, Loc84a, Mor66, Phi85, Roa78, Sch61, She75, Whi63b, Wib87, Whi66a, Whi66b]. Variables [Cha95b, DEPC95, EBTZ94, ETZ94, FS12, Han76, Kul65, Kyn76, MC91, Mac76b, MB64, Mit65, Obe64a, Pet98, PB96, Por86b, She62, She64, ST98, Sto61, Sto66, Bry96, Dra62, Fis71b, Mal80, Mot88, Ral74, Sle85, Thi63a, Whi76a]. Variables-Special [Sto66]. Variables. [Fab87]. Variance [DMGL22, HT72, Jok66a, Jow76, Meh73, PCLS23, Pri64, Ste01, SBXD06, Dwy63]. Variant [AN78, McC83]. Variants [CMSZ24]. Variate [Kel90c, Kel90d, Ste87]. Variates [Kaw63]. Variation [New66, New68b, PCLS23, She66a, Wei67a]. Variational [All92a, Ber75, Car91, Cla78c, Dam84b, Dam84a, Eva81, Eva84a, Ewi77, Gef75, Hau88, Hau90, HUS10, Kal63, LF73, MG76, Maw86, MA04, MSS96, MT96, Ono84, Ono85, Pet76a, Poh63, RM75, Tul90, Wan97, WZCX12, Ver98, All92b, Blo89, Edw90, Ell87, Eva84b, Hau85, Olk78, Ral80, Rei77a, Sha96]. Variations [CCA08, CJ16, Esl96, Her84a, MP72, Rei71a, Sag82, TL95, Ahl93, Bel63b, Bou91a, Fle63a, Fle63b, Hal66, Ham62b, Loe95]. Variations. [Her84b]. Variationsrechnung [Hes72]. Varied [Str76a]. Varieties [Som93, McN89b]. Various [Gre67b]. Varying [HIC96, Hew76, Kev87, Rei59b, Con96]. Vasek [Bla85a]. Vasil [Tot91]. Vasil’eva [Hoc86]. Vasiliev [Erm89]. Vasquez [Wet93]. Vasudevan [Mar88]. Vaulted [Spi91a, Spi91b]. Vaulting [Cor78]. Vector [And90, BDJ99, BT85, Cha87c, CLWM95, CTW21, DGK84, Fal70b, FT76, FF78, GDR78, Her91, Jef94, Kau88, Kaz84b, Kaz84a, MKR76, Mey93a, Mey94b, Mey95a, OV85, RKM76, Sch80c, Sch81c, SFS87, Tod84, TB76, VB05, WJ94, Rob94b, SFS88, Fan59, Sch81a]. Vector-Parallel [MKR76]. Vector-Valued [CTW21, Sch81a]. Vectorized [BNT76]. Vectors [LP71, SG73, Ste96, Koh98a, Gun63]. Vehicle [Mul79]. Vein [Gib87b]. Vekua [Sim89]. Velocity [Ant72, Lar86, New64a, PI88, Sim71a, Sto60, Tre82, Wil65a]. Vel’ski [GG91]. Veltkamp [Bra66, Kol65]. Vemuri [Gos83a]. Venant [Hor89]. Venit [Cur77, Los75]. Venkatarama [Bjö85]. Ventura [VN73]. Ventzel [Sei62]. Vere [Gut90]. Vere-Jones [Gut90]. Verfahren [Hig62d]. Vergnas [Whi94]. Verhulst [O’M87b]. Verification [Bos96, KK98, Mit93, Orl93b, Szy87]. Verlag [Hei96, Hig01b]. Version [Bee09, GJRW24, RB23, SP01]. Versions [BS94a, BS94b, Wag69]. versus [Bom02, Thu01]. Vertex [CHW12, DB78, Woo96]. Vertices [BGH^{+}04]. Very [Pit89, Sch93d]. Vessels [Gan98b, NN96, Gan98a]. Vetterling [Isa88, Isa89a, Yan94]. VI. [Gom67]. Via [Koc92, Küç92, Mro91, Bai76, BD94, BTKNZ00, BCV98, CESV15, CB09, ET99, FR16, FS78b, GK12, GHU22, GGT13, GO09, HSŠ16, Kre91, LS22, MO78, MMT^{+}21, Mey86a, NF16, RFP10, She09, Sie90, Smi87, VBC07, WS75a, YZZ04]. Vibrating [CMZ94, GK12, Rei59b, Sle76]. Vibration [Ban87, Ede90b, Lan64, Sto62b, Vak93, Ede90a]. Vibrations [Are89, Fre79, Hol89b, LW95, PMR76, Dav75c, Sha89b]. Vic [Sin79b]. Vich [Wel89]. Vichnevetsky [Kel83, Tre84b, Tre84a]. Victor [Chu96, Cop86, Daa97, Ebe96, Har96, Wig85]. Victoria [Wig90]. Vidar [Ode86]. Video [HH16]. Video-Based [HH16]. Vidossich [Kap85a]. Vidyasagar [Kha87]. View [Gud75, Her93a, Her93b, Hig07, Kus19, McI71, SG79]. VII. [Wey67]. Viises [Dwa66]. Vijay [Los84a, Sch85c]. Viking [Lan76a]. Villet [Jag90]. Vincent [Pin88]. Vincke [Wal88]. Viorel [Kor95b]. Viral [MXPW11, MPW23, ZWY14]. Viremia [ZWY14]. Viscoelastic [Obe77, Pip92, Riv65]. Viscoelasticity [Day87, PC61, Daf89, Day89]. Viscoplasticity [Pav94]. Viscosity [BW11, Men94a, Col89]. Viscous [Axf65, Bak91a, CL08, Riz94, Ser64, Shi71, TKG95, Gun95]. Vishik [Mie95, Mik89]. Vision [Far85, Roy93b, She09, Ste99, Tur04, Nag88, Roy93a]. Visiting [Iba73]. Visual [Hof71, Ran91b, Ros85, Sen94, Ran91a, Tre13]. Visualization [CMZ94]. Vivek [Fle90, Gut97]. Vivo [PN99]. Vlad [Con96]. Vladimir [Fir92b, Nas77b, Ock91, Par80, Fir92a]. Vladimirov [Dol72]. Vladut [Wol93]. Vleck [Wie87]. VLSI [Wol86]. VLSI-Optimal [Wol86]. Vol [Arb78, Art83a, Art88, Boe95a, But74a, But74b, But74c, CLH97, Coh93a, Cur64b, Den82, EK91a, Fir97, Fle77a, Fle77b, Gui92, Ham65b, Her77, Ker92, Kla92b, KZ93, Law72, Lin92, Lor91, McG64, Mel93, Mul64a, New94b, Noh62, Par72, Pin72b, Pin72a, Ral62b, Rei77b, Rei78c, Row79, Sal63b, Ser84a, Sik92, Sto66, Tap84a, Tu92, Tut94, Vog87, Wax91b, Wil62b, Wil86a, Win62, ADPR04, FK06, Kap65, Poh63, Ral62a, Win96a, Yan65]. vol.1 [Hig65]. Volatile [For76]. Volatility [ASPL03]. Volker [Hey85, Stu94b, Stu94a]. Volkov [Set63]. Vols [BDD^{+}95, Fif79, Mit87, Olv72, Rus95]. Volterra [Mor83, Bur83, Fus68b, Fus69, Gar75, Ole94, TeB97b, TeB97a, You76, Han85a]. Volterra/Wiener [Mor83]. Volume [Ans68, Art83b, Art85b, Ban87, Bel90a, Beu87, BPE96, Bla94, Boa87, BdDJ83, BDJ83, CHL08, Cha70a, Cul68, Dia64, Dob84, DGLZ12, Ebe96, Edw90, EK59, Ell85b, Fos90, Gol69b, Gue69, Hau90, HLS09, Jon92c, Kar60, Kel83, Kes69, Kla78b, KG79, Lew82c, Lie97, Mac69, Mac70, Mac92, Mac85, Mas87, Mel64, Mey86b, Mey89c, Muk86, Por81, Rei63, Sal62, San93, Sco94, Ser84b, Sha88, Sie69, Sit92, Sta88, Tap84b, Tho76a, Tul90, Vel21, Wax91a, Wei69, Whi63b, Whi88, Xu96c, Xu97, Sch60b, Sta85, dM94]. Volumes [Kau86]. Vor [Kni94a]. Vorganze [Hig62d]. Vorlesungen [Sch71a]. Voronoi [DFG99]. Vorst [Hig00]. Vortex [Leo78, NC09, Stu86, Caf94]. Vortical [Gun95, TKG95]. Vortices [Don23]. Vorticity [Che60, Maj91, WLP^{+}61]. Voss [New76c]. Vote [Ste96]. Votes [Ste04]. Voting [FP21, Luc87, Pac10, Hof85]. Vreugdenhil [Hol90a, Hol90b]. Vries [Miu77b, Sco91a, Sco91b, JK72, Miu75, Miu76b]. vs [HL63]. Vuorinen [TBDS92].

W [Alp85, Ano63a, Ano80d, Ans64b, Ant78, Arn89, Arn92, Ars62b, Art88, Axf65, Bar86, Bar67, Bau91, Bei85, Ber86a, Ber86c, Beu83b, Beu87, Bla70, Blu91, Boa83, Boa88, Boj93, Bos73a, Bra66, Bri96, Bru95, Bun83b, Bun83a, But92, Car73a, CL62, Cha95d, Cha68, CS64, CMR73, Cla62, Col96, Col94c, Col94d, Cot94, De 74, Deu63, Deu81c, Dev94a, Don85, Duf84, Dur93a, Eps59a, Fan82, Feu76, Fia73, Fre86a, Fre86c, Gil82, GD88, Gli91a, Gos83a, Gos84, Gou68, Gra18, GRYL89, Gus76, Ham76, Han79, Har73, Hei89, Her78a, Her65, HR62, Hul71, Ism78, Jag90, Jan70, Joh83c, Joh83d, Joh83a, Joh83b, JA80, Jor88b, Jow72, Kal63, Kap78, Kar78, Kaz92, Kee90, Kei71, Ken86b, Kes75, Kin97, Kla84b, Kle75, Kol65, Kri91, LS66, Lee82b, Lee84a, Leh63b]. W [Lem84, Les76, Los81, Luc18, Mal85, Mar66, Mar86b, Mas87, McG62b, McG65b, Meg88, Mil82a, Mil83b, Mil83c, Moo77a, Mor93, Mur94a, Mur94b, Nas77b, Nas86, NPW95, Nov94, O’M86, Olk72a, Olm77, Olv90, OG90, Ost61b, OR67, Pac88b, Par81a, Par92a, Pet95, Pim93, Pra85, RV78, Ral74, Ral83b, Ral83a, Rao94, Ren94c, Ric74, Rob94a, RB90, Ros63, Sag78, Say79, Sch89b, Sch65c, Sei86a, SM79, She91, Sho87, Shu87, Shu80, Sma76, SW81, Spi84b, Sto61, Sto62a, Str68, Szy87, Tak77, Tha82, Var85, Vog91a, Vog91b, Wal87, Wen68, Wie87, WJV79, Wil60b, WBGP91, Wim67, Wim75, Win62, Xu96b, Yan94, Zas65, ZN62, HL76]. W-A [HL76]. W. [Beu83a, Bun83a, Mer84a, Spi84a, Wes84a]. W1 [Coh75a]. Waadeland [Thr94a, Thr94b]. Wachpress [Fix78]. Wachspress [Jor90a, Var67]. Wackernagel [DK97]. Wadhwa [Dij90]. Wadsworth [Dra62]. Wagner [BW65, Ste88]. Wagon [Hig05, O’R92]. Wahba [Cle69, FSW^{+}66, Sch91a, Sch91b].

Wainstein [Zit63]. Waite [Tha82]. Waiting [San87]. Waldram [vH87]. Waldvogel [Hig05]. Walk [BGL17, BR18, Bur94c, HS70, Kem92a, Kem92b, MW61, She63b, Sku05b, WL95, New84, Sku05a, Bur94b, Law95]. Walker [Sle81]. Walks [Con71, LMSM83, LMSM84, SBH^{+}20, Sie89, Dur92a]. Wallace [Bri96]. Waller [Mei78a, Sau78]. Walling [BS67]. Wallis [Hei89]. Walsh [Bir69a, Dur87, GB89, BS70, Mor81]. Walter [Bor90, Bre94a, Bre94b, Cum96, Des68, Dre89, Ell85a, Kol76, Kow83, KZ93, Mac76a, Mar65, MNMF63, Mer84b, Sch65a, Seg06, Wes84b, Wil67a, dB65b]. Walters [Fri78a]. Waltman [Cus96, Ral70]. Wan [Mey82a]. Wang [Bla81, BKS96, Gro96h, Hab83a, Hab83b, Mey95a, Pré95b, Sha82a]. Wanner [Ale90]. Ward [Hul61a]. Warehouse [Sku05b, Sku05a]. Warga [Ces75]. Warming [Van12]. Warps [Kru83]. Warrants [VP09]. Warranty [SW93b]. Warren [MR71, Saw82]. Warwick [Ber85b]. Washline [Lew78b, LCG^{+}79]. Wasilkowski [Nov94, Shu87]. Wasow [Olv86]. Wasserman [Rat93]. Wasserstrom [Los76b]. Wassily [Man61a]. Water [Arn89, HV75, Kel88a, Pap63b, Pru00, Rog75, SR94, WPL98, YL75, Mor97, Col87a]. Waterman [Mei78a, Sau78]. Watkins [Gar93]. Wave [AN83, BS00, Bla66, Cho75, DS73, Fok84, GM81a, GM81b, Gef75, HV75, Hen76, Ike76, Kel79a, Kel86a, Kel88a, Lax75, Lud70, Lud77, Mar81, Mor88c, Pea81, PB75, RK75, Rog69b, Sah95, Sch84b, Sch84a, Sco84b, Sco84a, She87, Sug98, TWW65, Wal92, Wal93c, Wal93b, Wil64a, Wil76b, Win90, YFS76, Ben87, Dat90, Mar88, Sub90, Whe74, Wil62c]. Waveform [MBOV17]. Wavefront [LBL02]. Wavelength [Ram93]. Wavelet [CW94, Chu96, HW89, JS94, WC96]. Wavelets [Ber92a, Gar94a, KF95, Roc92b, Str89a, UB00, Wic94b, Wal97, Bre93, Fra95b, Hei93b, Roc92a, Wal93a, Wic94a]. Waves [Ame91a, Ame91b, AHP^{+}95, ADW23, Bal97, CM89, Col87a, Col84, Coo67b, EJ07, EHS17, Fej88, GAB^{+}88, HSSZ98, JK72, Kau86, KT92, Kin88, KCD^{+}97, Kri76, LM79, Lud75a, MB76, Mor97, Mor76, New76a, Quo75, Rau95, Ren97b, Ric60, Rog75, Row64, She75, Sti87, Ter93b, YL75, Zua05, Shi97, Abl77a, Ame88, Caf92, Cha78, Dat86, Dol62, Gib85, Gro88b, Kan62a, Miu76a, Ter93a, Wei90b, Mil82b]. Waves-Volumes [Kau86]. Wawrzynczyk [Var87]. Way [KK99, Kin88, Mas92]. Wayne [Kle76]. Ways [Hig02a, MV78a, MV78b, MV03, Vas97]. Weak [BK17, Dam84a, Gra86b, JK72, Bry93, Dam84b]. Weakly [Hag83, WE19, WMZ94]. Weapon [Dun64]. Weather [Ghi76, Sal63d, Som78]. Weaver [Gru88, Kel86b, Kel90b, Rog96]. Web [BGH^{+}04, DZH^{+}04, Gle15, LM05]. Weber [Kan62a, OG90, Mil89a, OGGR90]. Wechsler [Roy93a, Roy93b]. Wechsung [Ste88]. Week [FP04]. Wegert [Tre13]. Wei [Rou91, Sha82a]. Weibel [Cha87a]. Weidlich [Fre86a]. Weierstrass [Lin78b, Lin79b]. Weight [Jev92a, SG73]. Weighted [Mye75, Pro84c, Shu09, YKS78, Ant87]. Weights [For98b, SKP19, For98a]. Weinberg [Fla67]. Weinberger [Pay68]. Weingartner [Ber65a]. Weinmann [KS94]. Weinstein [Baz73]. Weispfenning [Stu94b, Stu94a]. Weiss [JW83, Law95, Wri60]. Weissblum [Gou68, Mar65, MNMF63, Sch65a, dB65b]. Well [BT20, BDvD^{+}98, Lat23, OT13]. Well-Conditioned [OT13]. Well-Posed [Lat23]. Well-Posedness [BT20]. Well-Solvable [BDvD^{+}98]. Wellner [Bic94]. Wells [Wig90]. Welsh [Gha89]. Wendell [Men94a]. Wendland [Gil82, Sch71a]. Wendroff [Par67a]. Wentzell [Mer85, Spi82]. Werner [Fal68, Gas70, Ruo92a, Ruo92b]. Werschulz [Was94a]. Westenholz [Her80a]. Weston [BdD93]. Wet [BFF^{+}99]. Wetherill [Ros68a, Sim82]. Weyl [AS85, Ano78i, Bel68, Sri84]. Whats [Hou78]. Wheeler [Bro93]. Wheeling [She63b]. Wheels [FR83]. Where [CU03]. Which [Ben70, Cor77, JvD90, JvD91, Kra80a, Smi88b, WM78, Gil65a, Kra80b, OW95b, Qui87, Urq68]. White [Ode89, Orl93a, Ser80b, Whi94]. Whiteman [Mal90]. Whitham [Miu76a]. Whitt [vLMZ19]. Whittaker [Ste81a]. Whittle [Ell85b, Per89, Ser84b, Ser84a]. Whose [Reu67, Ste59]. Wickerhauser [Chu96]. Widder [Hai77, Hir73]. Wide [DZH^{+}04]. Widths [Smi86]. Wiener [Coo67a, Mar75a, Mor83, Asm98]. Wiggins [Cra89, Pal95, Pal96]. Wiktor [O’M74]. Wilcox [Leh63b, Mar86a, Ran96b]. Wiley [Boa87, Mey72]. Wilf [Hir71, Hul61b, Lin79c, Whi80]. Wilfred [Kee88b, Sch60c]. Wilhelm [Fal68, Kaz67, Koh85]. Wilkins [Ano80d, Fet66, Jor86b, Vel71]. Wilkinson [Des68, Gra66b, Par66, Par72, CS20, Hou78]. Willard [Don85, Ral83a]. Willem [Laz91a, Laz91b]. Willems [Die92, Dis92]. Willi [Chu94b, Chu94a]. Willi-Hans [Chu94b, Chu94a]. William [Are86, Bar70, Baz73, Ber93, Ber85d, Bou66a, Cer96, Coo67a, Cyb93, Daf89, Des68, Fer79, Fis93, Fra86b, Har78, Hin91, Isa88, Isa89a, Kan68, Kap72, Kee61, Kes69, Kla84i, Koç94, Laz82, Lee86a, MR71, Man90a, Man61a, ML85, Ort96b, Pin72a, Ric67a, Ros87, Sam76c, Sie69, Ull88, War64, Wil73, Man90b]. Williams [Har74, Mla94, Por81]. Willoughby [Cho87, Par74]. Wilson [Ano71a, Kaz82, McL86b, Mor83, RB90, WS71, vdM68]. Wim [Lea94]. Wimp [Luk82]. Wind [KN61a, Kla82, Kla83a, PMK96, PB75, VK62, Win93, Win96b, Win96a].

Windowed [Lev90a, Lev91]. Windows [BGL01]. Wine [BFF^{+}99]. Wing [Ans64a, McG93]. Winners [Sch66]. Winning [Liu92a, Mar85b, Ren93d, Ren93b]. Wire [BGML96, LS95, Boe96a]. Wiring [Pom67, Ste61a]. Wirszup [Ham77]. wise [PPR11]. Within [Ros87, CS84]. without [Gol62, Nan20, O’M92, Sha04, Sku05a, Sku05b, Str62b, SKP19, Wu75]. Wittman [Jos95]. Witzgall [Kle72]. WKB [Har96, MH96]. Wohlers [Shi68]. Wojbar [Pro95]. Wojciech [Boe89]. Wold [Pap85]. Wolf [Pin92b, Wil62b]. Wolfe [YY96]. Wolfgang [Col73, Man88a, Olv86, Ruo92a, Ruo92b, Sch71a, Slo97, Wil67a, Wou64b, Pin92a]. Wolfowitz [Kem67]. Wolfram [Nit90, Wes92]. Wolfs [Noe81]. Wolk [Jag73]. Wolper [For96]. Wolsey [Bla90]. Wong [Mor73]. Wood [Arn92, Cor92b, GZ11]. Wooldridge [Gre60a, Kar60]. Worcester [Col96]. Work [BM03, Hau81, Mil76, Dan71]. Workbook [Ken91a, Ken91b, Koç94]. Workers [Sku05b, Sku05a]. Works [CGO07, Gau08]. Workshop [Por86a]. World [Cum96, Fee00, FLC96, Hig07, Mor85a, Wal92, Wal93b, Mor86c, DZH^{+}04, Len93a, Wal93c]. Wormholes [GKLU09]. Worst [Hua86]. Woyczynski [Pro95]. Wozniakowski [Nov94, Pac86, Shu87]. Wrapped [MT84]. Wray [Rus85]. Wright [Kow83, KZ93, Luc86, Par94]. Wrinkle [Wu75]. Writing [Bra94b, Bra94a]. Written [Bur78a]. Wrong [Hou78]. Wubs [Arn89]. Wuensche [Hur93a, Hur93b]. Wunsch [Col96]. Wynn [Fis76].

X [Kus93, Hig05, Wu72]. X-Ray [Wu72]. Xiang [Ath93b, Ath93a]. Xiang-Qun [Ath93b, Ath93a]. Xiaohuan [Gro96h, Pré95b]. Xie [Sav18]. xii [BC96b, Hig05]. Xin [BKS96]. Xin-Bong [BKS96]. xiv [Tre13, Dur87]. XPPAUT [GN22]. Xunjing [Sto96]. xxii [Bor06].

Y-H [Mey82a]. Y. [Hab83a]. Ya

[Ath87, Bak91b, Bak92, Ham77, How63, Kla84j, Mey94a, Pap63b, Rei62a, Rya63]. Yaglom [Lev87]. Yakhno [Erm89]. Yakov [Pan93a, Pan93b, Sha97b]. Yakubov [Tri95]. Yakubovich [Aga93, Mey77b]. Yang [Ath93b, Sav18, Ath93a]. Yanpol’skii [Kog66]. Yao [Sha95c]. Yasaimaibodi [Olk72b]. Yaspan [Sop60]. Yasutaku [Sch77a]. Year [Ano85k, Ano86l, Ano87l, Ano88j, Ano89m, Ano90m, Ano91l, Ano92i, Ano93j, Ano94g, Ano95q, Ano96r, Ber86c, Bla97b, Bla97a]. Years [Blo73, Dur93a, MV03]. Yeh [Mar75a]. Yildiz [WS71]. Yilmaz [Gol66b]. Ying [Tha96]. Ylinen [Zak85]. Yogesh [Reh89]. Yoichi [Get96]. Yonathan [Rei75]. Yong [Sto96]. Yor [Kni94b, Pit93a, Pit93b]. Yosef [Spu89]. Yoshifumi [Aok70]. Yoshihiko [Keh89]. Yoshikazu [Aok70, Cra86]. Yoshimasa [Abl86]. Yoshisada [Spi87]. Yosida [DC86]. Youcef [Cul94b, Cul94a]. Young [Fad62, Joh73, May84c, May84d, May84b, Por83, Rei71a, Saa88]. Yovits [Rei64]. Yu [Coh90a, Daa97, Erm89, Haa93, Ham77, Hey82, Jun70b, Luk75, Por85, Coh90b]. Yuan [Hab83b, Hol91b, Sam80]. Yudell [Cod78, Olv72, Tha79]. Yudin [Bla85b]. Yuri [Ock91, Vai96]. Yurii [Por89]. Yurij [Col87b]. Yushkevich [Ell81, Pin70]. Yves [Ric90b, Wic94a, Wic94b, Gar94a, Ric90c].

Z [Ath87, Ban85, Bry93, Cam77, Ham79a, Hud95, LG89, Mal77, Tad93, Thi63b]. Zabreiko [Ber85a]. Zacks [Ros72a]. Zadeh [Kap64]. Zahari [Eis86b]. Zahlten [Dev92]. Zahrt [McG93]. Zaitsev [Kla97]. Zakhariev [Dei92]. Zang [Bor90, Boy88]. Zangwill [Lem84]. Zaslavsky [Bry93]. Zauderer [Det85]. Zayed [Ann95b, Gro93a]. Zeeman [Sus79]. Zeev [Fle82]. Zegeling [Del95]. Zehnder [Din72]. Zeidler [Bor06]. Zeigler [Sam79]. Zeitouni [Din94a, Din94b]. Zelikin [Kre96]. Zemanian [Kor73, Lei66]. Zener [BI68a]. Zenisek [Wah92]. Zeon [NdA14]. Zero [And75a, Boj65, Coh75a, EBTZ94, ETZ94, Gla88d, GS63, Hog74, Hog75, wBJ77, Jor77c, Lir76, Med62, RJ89, RD65, Ste70b, Wit90, Ypm83, Ren89]. Zero-Crossings [Wit90]. Zeros [Bar76b, BK99, Boy13, Boy82, Boy83, Deu81b, FO99, Gau87, Ike76, Jin84, Jin85, Lau82, Lau83, Lau84, Lev67b, Max60, Men94b, Nev82, ND82, Rie80, RSS81, Sch81b, SFH^{+}66, Swa65, Wan84, Blo95, Jag83]. Zeta [BR94, Bus95, Coh75b, Lin78b, Lin79b]. Zeytounian [Yan91]. Zhang [Ari96, Gro96h, Pré95b]. Zharinov [Olv93]. Zheng [Hil96]. Ziebur [Rob74, Rob76]. Ziegler [Red93a, Whi94, Red93b]. Zijm [Ber86a]. Zilinskas [AKP93a, AKP93b]. Ziolkowski [Mor86a]. Zion [Bla88f]. Zisserman [Ran91a, Ran91b]. Zlatev [Eis86b]. Zlobec [Eck82]. Zohar [Tal81b]. Zolotarev [NF16]. Zone [MN65]. Zongyuan [Sav18]. Zoran [Jód96]. Zoutendijk [Him78]. Zubair [Gla93b, Mil95b]. Zubakov [Zit63]. Zubov [Dri65]. Zun [BPK84]. Zwart [Mor96]. Zweidimensionale [Alf93]. Zweig [Hur65]. Zwilliger [Lat90a, Lat90b]. Zwillinger [Hin94, Kel96]. Zyckowski [Tad89]. Zyczkowski [Tad93].

References

Alani:1978:TSA

[AA78] S. R. Alani and R. S. Alsalihi.

Time-series analysis for lowest temperature. SIAM Review, 20(3):619, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alpern:1982:BRS

[AA82] Steve Alpern and Miroslav D. Asić. Book review: Search Games (Shmuel Gal). SIAM Review, 24(2):235–236, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aguilar:2015:SPE

[AABT15] Oscar Aguilar, Moritz Allmaras, Wolfgang Bangerth, and Luis Tenorio. Statistics of parameter estimates: a concrete example. SIAM Review, 57(1):131–149, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andeen:1976:JCS

[AAJ76] G. B. Andeen, C. M. Ablow, and S. H. Johnson. Jet and column stability. SIAM Review, 18(4):792, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aalto:1971:BRP

[Aal71] S. Aalto. Book review: A Polynomial with Prescribed Properties (Edgar Cohen, Jr.). SIAM Review, 13(1): 125–126, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aarao:2007:FSS

[Aar07] Jorge Aarão. Fundamental solutions for some partial differential operators from fluid dynamics and statistical physics. SIAM Review, 49 (2):303–314, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antle:1969:PML

[AB69] Charles E. Antle and Lee J.

Bain. A property of maximum likelihood estimators of location and scale parameters. SIAM Review, 11(2): 251–253, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Amos:1973:CTT

[AB73] D. E. Amos and J. W.

Burgmeier. Computation with three-term, linear, nonhomogeneous recursion relations. SIAM Review, 15(2): 335–351, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alzer:1997:PIP

[AB97] Horst Alzer and Joel Benner. Problems: An inequality (problem 97-1). SIAM Review, 39(1):123, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aristoff:2023:BBI

[AB23] David Aristoff and Wolfgang Bangerth. A benchmark for the Bayesian inversion of coefficients in partial differential equations. SIAM Review, 65(4):1074– 1105, November 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abbott:1992:BRA

[Abb92] H. L. Abbott. Book review: Approximation for the Location of the Center of a Disk (M. L. Glasser and J. Koplowitz). SIAM Review, 34 (4):658–661, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abdelkader:1965:DDE

[Abd65] Mostafa A. Abdelkader. Duality in differential equations. SIAM Review, 7(3): 410, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abdelkader:1966:DDE

[Abd66] Mostafa A. Abdelkader. Duality in differential equations. SIAM Review, 8(4): 526–529, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abdelkader:1983:PVL

[Abd83] M. A. Abdelkader. Pendulum with variable length. SIAM Review, 25(3):402– 403, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abdelkader:1984:NDE

[Abd84]  M. A. Abdelkader.  A

nonlinear differential equation. SIAM Review, 26 (1):118, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abdelkader:1985:PVM

[Abd85] M. A. Abdelkader. Pendulum in a variable medium. SIAM Review, 27(1):80, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abdelkader:1990:GCM

[Abd90] M. A. Abdelkader. A guerrilla combat model. SIAM Review, 32(1):140, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abdelkader:1991:TOD

[Abd91]  M. A. Abdelkader.  A

third-order differential equation. SIAM Review, 33 (1):115, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abed:1988:SPS

[Abe88] Eyad H. Abed. A simple proof of stability on the center manifold for Hopf bifurcation. SIAM Review, 30 (3):487–491, September 1988.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andersen:1995:SMB

[ABG^{+}95] P. K. Andersen, O. Borgan, R. D. Gill, N. Keiding, and Ian W. McKeague. Statistical models based on counting processes. SIAM Review, 37 (3):475–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aston:1996:BRC

[ABG96] P. J. Aston, M. V. Bartuccelli, and S. A. Gourley. Book review: A Conjecture Concerning an Integral Arising in Population Modelling (M. V. Bartuccelli and S. A. Gourley). SIAM Review, 38(3): 524–525, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Allamigeon:2021:WTG

[ABGJ21] Xavier Allamigeon, Pascal Benchimol, Stéphane Gaubert, and Michael Joswig. What tropical geometry tells us about the complexity of linear programming. SIAM Review, 63(1):123– 164, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ammari:2010:MMP

[ABJK10] Habib Ammari, Emmanuel Bossy, Vincent Jugnon, and Hyeonbae Kang. Mathematical modeling in photoacoustic imaging of small absorbers. SIAM Review, 52(4): 677–695, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ablowitz:1975:NEE

[Abl75]  M. J. Ablowitz.  Non-

linear evolution equations — continuous and discrete. SIAM Review, 17(2):373– 374, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ablowitz:1977:BRN

[Abl77a] Mark J. Ablowitz. Book review: Nonlinear Waves (Sidney Leibovich and A. Richard Seebass). SIAM Review, 19 (2):368–369, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ablowitz:1977:NEE

[Abl77b] Mark J. Ablowitz. Nonlinear evolution equations, continuous and discrete. SIAM Review, 19(4):663– 684, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ablowitz:1980:BRL

[Abl80] Mark J. Ablowitz. Book review: Lie–Bäcklund Transformations in Applications (Robert L. Anderson and Nail H. Ibragimov). SIAM Review, 22(3):377, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ablowitz:1984:NPD

[Abl84a] M. J. Ablowitz. Nonlinear partial-differential equations in physical problems — Graffi, D. SIAM Review, 26 (1):125, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ablowitz:1984:BRN

[Abl84b] Mark J. Ablowitz. Book review: Nonlinear Partial Differential Equations in Physical Problems (D. Graffi). SIAM Review, 26 (1):125, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ablowitz:1986:BRB

[Abl86] Mark J. Ablowitz. Book review: Bilinear Transformation Method (Yoshimasa Matsuno). SIAM Review, 28 (1):91–92, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Allmaras:2013:EPP

[ABL^{+}13] Moritz Allmaras, Wolfgang Bangerth, Jean Marie Linhart, Javier Polanco, Fang Wang, Kainan Wang, Jennifer Webster, and Sarah Zedler. Estimating parameters in physical models through Bayesian inversion: a complete example. SIAM Review, 55(1): 149–167, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abrahams:1984:TDI

[Abr84] Julia Abrahams. Three double integrals. SIAM Review, 26(3):430, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Abraham:1994:BRE

[Abr94] Ralph H. Abraham. Book review: Encounter with Chaos (J. Peinke, J. Parisi, O. E. Rössler, and R. Stoop). SIAM Review, 36(2):288– 289, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anosov:1995:RHP

[ABS95] D. V. Anosov, A. A. Bolibruch, and Yasutaka Sibuya. The Riemann–Hilbert problem. SIAM Review, 37(4): 637–??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1976:SHF

[AC76] R. A. Askey and L. Carlitz. Series of hypergeometric functions. SIAM Review, 18 (4):768–769, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Avila:1978:UMH

[AC78] J. H. Avila and P. Concus. Update methods for highly structured systems of nonlinear equations. SIAM Review, 20(3):619, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Acker:1973:CM

[Ack73] Andrew Acker. Commuting matrices. SIAM Review, 15 (1):219, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Acker:1974:BRC

[Ack74] Andrew Acker. Book review: Commuting Matrices (Andrew Acker). SIAM Review, 16(3):391–393, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andersen:1976:EIT

[AD76] C. M. Andersen and G. C. Darden. Evaluation of integrals for a ten-node isoparametric tetrahedral finite element. SIAM Review, 18 (4):793, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1987:BRS

[AD87] Richard Askey and S. Lj.

Damjanović. Book review: On the Sign of a Set of Hypergeometric Functions (Richard Askey). SIAM Review, 29(1):129–131, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Adam:1991:AIC

[Ada91] John A. Adam. Activator inhibitor control of tissue growth. SIAM Review, 33 (3):462–466, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Adamchik:1997:IAS

[Ada97a] V. Adamchik. An integral arising in the study of a random graph. SIAM Review, 39 (1):135–136, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Adamchik:1997:SIA

[Ada97b]  Victor Adamchik.  Solu-

tions: An integral arising in the study of a random graph (problem 96-1) (Tae Chalker). SIAM Review, 39(1):135–136, March 1997.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Adomaitis:1989:KBT

[Ado89] Raymond A. Adomaitis. Kites and bifurcation theory. SIAM Review, 31(3): 478–483, September 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alon:2004:TCR

[ADPR04] Noga Alon, Seannie Dar, Michal Parnas, and Dana Ron. Testing of clustering (reprinted from SIAM Journal on Discrete Mathematics, vol 16, pg 393, 2003). SIAM Review, 46(2): 285–308, June 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/43717.

Arnold:1976:ARS

[ADR76] L. K. Arnold, L. Dodson, and L. Rosen. An average relative speed approximation. SIAM Review, 18 (3):489, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Avitabile:2023:BAW

[ADW23] Daniele Avitabile, Joshua L. Davis, and Kyle Wedgwood. Bump attractors and waves in networks of leaky integrate-and-fire neurons. SIAM Review, 65 (1):147–182, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1367246.

Ahmadi:2023:LDS

[AE23] Amir Ali Ahmadi and Bachir El Khadir. Learning dynamical systems with side information. SIAM Review, 65(1):183–223, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1388644.

Aebi:1994:SED

[Aeb94a] R. Aebi. Schrödinger equations and diffusion theory (Masao Nagasawa). SIAM Review, 36(2):288–??, June 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aebi:1994:BRS

[Aeb94b] Robert Aebi. Book review:

Schrödinger Equations and Diffusion Theory (Masao Nagasawa). SIAM Review, 36 (2):288, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Angstmann:2021:GFF

[AEH^{+}21] Christopher N. Angstmann, Austen M. Erickson, Bruce I.

Henry, Anna V. McGann, John M. Murray, and James A. Nichols. A general framework for fractional order compartment models. SIAM Review, 63(2):375–392, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1969:BRT

[AF69] Richard Askey and James Fitch. Book review: A Trigonometric Inequality (J. N. Lyness and C. B. Moler). SIAM Review, 11(1):82– 86, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1967:BRP

[AG67] George E. Andrews and I. Gerst. Book review: A Pair of Identities (L. Carlitz). SIAM Review, 9(2): 254–256, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Amos:1976:LI

[AG76] D. E. Amos and A. G. Gibbs. Limit of an integral. SIAM Review, 18(4):772, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Allgower:1980:SCM

[AG80] Eugene Allgower and Kurt Georg. Simplicial and continuation methods for approximating fixed points and solutions to systems of equations. SIAM Review, 22(1): 28–85, January 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Artz:1983:EVT

[AG83] R. E. Artz and M. A.

Golberg. On the expected value of two statistics. SIAM Review, 25(3): 405–406, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aldroubi:2001:NSR

[AG01] Akram Aldroubi and Karlheinz Gröchenig. Nonuniform sampling and reconstruction in shift-invariant spaces. SIAM Review, 43 (4):585–620, December 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38698.

Askham:2014:NPD

[AG14] Travis Askham and Leslie Greengard. Norm-preserving discretization of integral equations for elliptic PDEs with internal layers i: The one-dimensional case. SIAM Review, 56(4):625–641, ???? 2014. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Agarwal:1993:BRD

[Aga93] Ravi P. Agarwal. Book review: The Double Mellin– Barnes Type Integrals and Their Applications to Convolution Theory (Nguyen Thanh Hai and S. B. Yakubovich). SIAM Review, 35(4):652, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Agarwal:1994:BRC

[Aga94] Ravi P. Agarwal. Book review: Classical and New Inequalities in Analysis (D. S. Mitrinovic, J. E. Pecaric, and A. M. Fink). SIAM Review, 36(2):290, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Agarwal:1995:BRG

[Aga95a] Ravi P. Agarwal. Book review: Global Behavior of Nonlinear Difference Equations of Higher Order with Applications (V. L. Kocic and G. Ladas). SIAM Review, 37 (1):114–115, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Agarwal:1995:BRN

[Aga95b] Ravi P. Agarwal. Book review: Nonstandard Finite Difference Models of Differential Equations (Ronald E. Mickens). SIAM Review, 37 (3):459, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aggarwal:1974:GA

[Agg74] H. R. Aggarwal. Gravitational attraction. SIAM Review, 16(4):546, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Artz:1983:BRE

[AGR83] R. E. Artz, M. A. Golberg, and K. A. Rao. Book review: On the Expected Value of Two Statistics (K. A. Rao). SIAM Review, 25(3): 405–406, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Al-Hussaini:1995:UPT

[AH95a] A. N. Al-Hussaini. A unified proof of two theorems in statistics. SIAM Review, 37 (4):596–597, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/27305.
htm.

Axelsson:1995:ISM

[AH95b] Owe Axelsson and Martin Hanke. Iterative solution methods. SIAM Review, 37 (3):466–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Auchmuty:1998:CCI

[AH98] G. Auchmuty and J. W. He. A conjecture concerning an integral arising in population modelling. SIAM Review, 40 (3):710–714, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Adcock:2019:FNA

[AH19] Ben Adcock and Daan Huybrechs. Frames and numerical approximation. SIAM Review, 61(3):443–473, ???? 2019. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ahlfors:1961:CCA

[Ahl61] Lars V. Ahlfors. Classical and contemporary analysis. SIAM Review, 3(1): 1–9, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ahlin:1962:EBG

[Ahl62] A. C. Ahlin. On error bounds for Gaussian cubature. SIAM Review, 4(1): 25–39, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ahlberg:1982:BRS

[Ahl82] J. H. Ahlberg. Book review: Spline Functions (Larry Schumaker). SIAM Review, 24(4):490–491, October 1982.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ahlbrandt:1993:BRI

[Ahl93] Calvin D. Ahlbrandt. Book review: Introduction to the Calculus of Variations (U. Brechtken-Mandersheid; P. G. Engstrom. trans.). SIAM Review, 35(3):506– 507, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ahmed:1975:ATC

[Ahm75] N. U. Ahmed. Approximation theorem for continuous functions on L_{p}(1 ≤ p) infinity spaces. SIAM Review, 17(2):374, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anile:1995:RMN

[AHP^{+}95] A. M. Anile, J. K. Hunter, P. Pantano, G. Russo, and Jeffrey Rauch. Ray methods for nonlinear waves in fluids and plasmas. SIAM Review, 37(1):136–??, March 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aizicovici:1994:AEE

[Aiz94a] S. Aizicovici. Abstract evolution equations, periodic problems and applications (Daniel Daners and Pablo Koch Medina). SIAM Review, 36(2):

289–??, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aizicovici:1994:BRA

[Aiz94b] Sergiu Aizicovici. Book review: Abstract Evolution Equations, Periodic Problems and Applications (Daniel Daners and Pablo Koch Medina). SIAM Review, 36(2): 289–290, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alawneh:1977:SMM

[AK77] A. D. Alawneh and R. P. Kanwal. Singularity methods in mathematical physics. SIAM Review, 19(3):437– 471, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Al-Khayyal:1990:BRC

[AK90] Faiz Al-Khayyal. Book review: Constrained Global Optimization: Algorithms and Applications (P. M. Pardalos and J. B. Rosen). SIAM Review, 32(2):310– 312, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Al-Khayyal:1991:BRN

[AKP91a] Faiz Al-Khayyal and Panos Pardalos. Book review: New Computer Methods for Global Optimization (H. Ratschek and J. Rokne). SIAM Review, 33(4):684–686, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Asch:1991:FCR

[AKP^{+}91b] M. Asch, W. Kohler, G. Papanicolaou, M. Postel, and B. White. Frequency content of randomly scattered signals. SIAM Review, 33 (4):519–625, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Al-Khayyal:1993:BRG

[AKP93a] Faiz A. Al-Khayyal and Panos M. Pardalos. Book review: Global Optimization (Aimo Törn and Antanas Zilinskas). SIAM Review, 35 (3):509–510, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Al-Khayyal:1993:GOA

[AKP93b] Faiz A. Al-Khayyal and Panos M. Pardalos. Global optimization (Aimo Torn and Antanas Zilinskas). SIAM Review, 35(3):509– ??, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andersen:1998:TBC

[AL98] E. S. Andersen and M. E.

Larsen. Three binomial convolutions. SIAM Review, 40 (4):993–997, November 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Albert:2010:MGS

[Alb10] Jim Albert. Mathematics in games, sports, and gambling: The games people play. SIAM Review, 52(4): 774–775, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alexander:1984:MTT

[Ale84] J. C. Alexander. On the motion of a trailer-truck. SIAM Review, 26(4):579– 580, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alexander:1985:NCG

[Ale85] J. C. Alexander. The numerics of computing geodetic ellipsoids. SIAM Review, 27 (2):241–247, June 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alexander:1990:BRS

[Ale90] Roger Alexander. Book review: Solving Ordinary Differential Equations I: Nonstiff Problems (E. Hairer, S. P. Nørsett, and G. Wanner). SIAM Review, 32 (3):485–486, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alessandrini:1995:MEN

[Ale95] Stephen M. Alessandrini. A motivational example for the numerical solution of twopoint boundary-value problems. SIAM Review, 37(3): 423–427, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/27318.
htm.

Alessandrini:1998:MEN

[Ale98] Stephen M. Alessandrini. A motivational example for the numerical solution of the algebraic eigenvalue problem. SIAM Review, 40 (4):935–940, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32835.

Alfeld:1993:BRZ

[Alf93] Peter Alfeld. Book review:

Zweidimensionale Spline-Interpolation-Algorithmen (H. Spath). SIAM Review, 35 (3):507–509, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alhargan:1996:CMC

[Alh96] Fayez A. Alhargan. A complete method for the computations of Mathieu characteristic numbers of integer orders. SIAM Review, 38(2): 239–255, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/p239.htm.

Allen:1975:IOF

[All75] R. W. Allen. Idealized optical fibre mode conversion. SIAM Review, 17(1):167, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Allen:1976:IOF

[All76]  R. W. Allen.  Idealized

optical fiber mode conversion. SIAM Review, 18(2): 298–300, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Allegretto:1978:BRS

[All78] W. Allegretto. Book review: Solutions to Linear Partial Differential Equations Involving Arbitrary Functions (A. Ungar). SIAM Review, 20 (1):190–191, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Allen:1987:EPA

[All87] Edward J. Allen. On the error in Padé approximation. SIAM Review, 29(2): 299, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Allwrigh:1992:VPO

[All92a] D. Allwrigh. Variational problem for overlapping laser pulses. SIAM Review, 34(2): 310–312, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Allwright:1992:BRV

[All92b] David Allwright. Book review: Variational Problem for Overlapping Laser Pulses (C. E. Carroll). SIAM Review, 34(2):310–312, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Almgren:1987:BRE

[Alm87]  F. Almgren.  Book re-

view: Equilibrium Capillary Surfaces (Robert Finn). SIAM Review, 29(3):475– 476, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alpern:1985:BRG

[Alp85] Steve Alpern. Book review: Geometric Games and Their Applications (W. H. Ruckle). SIAM Review, 27 (2):257–258, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alpert:1996:CNV

[Alp96] Bradley K. Alpert. Condition number of a Vandermonde matrix. SIAM Review, 38 (2):314, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alspach:1985:BRI

[Als85] Brian Alspach. Book review: Introductory Combinatorics (Kenneth P. Bogart). SIAM Review, 27(2): 258–259, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alsalam:1986:IDH

[Als86] W. A. Alsalam. An identity for double hypergeometric series. SIAM Review, 28 (4):576–578, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alspach:1992:BRI

[Als92a] Brian Alspach. Book review: An Introduction to Combinatorics (Alan Slomson). SIAM Review, 34(4):662, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alspach:1992:ICA

[Als92b] Brian Alspach. An introduction to combinatorics (Alan Slomson). SIAM Review, 34 (4):662–??, December 1992.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alter:1976:RRE

[Alt76] R. Alter. Remarks and results on an extension of a problem of Smirnov. SIAM Review, 18(4):793, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arbib:1974:MCE

[AM74a] Michael A. Arbib and Ernest G.

Manes. Machines in a category: an expository introduction. SIAM Review, 16 (2):163–192, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Assmus:1974:CC

[AM74b] E. F. Assmus, Jr. and H. F.

Mattson, Jr. Coding and combinatorics. SIAM Review, 16(3):349–388, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Amann:1976:FPE

[Ama76]  Herbert Amann.  Fixed

point equations and nonlinear eigenvalue problems in ordered Banach spaces. SIAM Review, 18(4):620– 709, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [Ama77b].

Amann:1977:C

[Ama77a]  H. Amann.  Correction.

SIAM Review, 19(4):R7, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Amann:1977:EFP

[Ama77b] Herbert Amann. Errata: Fixed Point Theorems and Nonlinear Eigenvalue Problems in Ordered Banach Spaces. SIAM Review, 19 (4):vii, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Ama76].

Amann:1982:BRO

[Ama82] Herbert Amann. Book review: Operator Inequalities (Johann Schroeder). SIAM Review, 24(1):97, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1973:SCS

[Ame73]  W. F. Ames.  Some

computation-steeples in fluid mechanics. SIAM Review, 15 (2):524–552, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1978:ESN

[Ame78] W. F. Ames. Evolution of singularities for non-linear hyperbolic equations in multidimensional spaces. SIAM Review, 20(3):619, ???? 1978.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1981:BRS

[Ame81] W. F. Ames. Book review: Similarity, Self-Similarity, and Intermediate Asymptotics (G. I. Barenblatt). SIAM Review, 23 (4):545, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1984:BRO

[Ame84] W. F. Ames. Book review: Operator Inequalities (Johann Schroeder). SIAM Review, 26(1):144– 145, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1985:BRS

[Ame85] W. F. Ames. Book review:

Systems of Nonlinear Partial Differential Equations (J. M. Ball, ed.). SIAM Review, 27 (2):259–260, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1986:BRI

[Ame86a] W. F. Ames. Book review: Inequality Problems in Mechanics and Applications: Convex and Nonconvex Enemy Functions (P. D. Panagiotopoulos). SIAM Review, 28(3):405–406, ???? 1986.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1986:BRN

[Ame86b] W. F. Ames. Book review: Numerical Solution of Partial Differential Equations: Theory, Tools and Case Studies (D. P. Laurie. ed.). SIAM Review, 28(1):92, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1988:BRN

[Ame88] W. F. Ames. Book review: Nonlinear Diffusive Waves (P. L. Sachdev). SIAM Review, 30(3):510, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1991:BRR

[Ame91a] W. F. Ames. Book review: The Riemann Problem and Interaction of Waves in Gas Dynamics (Tung Chang and Ling Hsiao). SIAM Review, 33(1):129, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1991:RPI

[Ame91b] W. F. Ames. The Riemann problem and interaction of waves in gas dynamics (Tung Chang and Ling Hsiao). SIAM Review, 33(1): 129–??, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ames:1995:BRI

[Ame95]  Karen Ames.  Book re-

view: An Introduction to Nonlinear Partial Differential Equations (J. D. Logan). SIAM Review, 37(2): 251–252, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Amos:1964:BRR

[Amo64] D. E. Amos. Book review:

Resistance of a Cut-out Right Circular Cylinder (Alan L. Tritter). SIAM Review, 6(4): 457–459, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Amos:1965:TOD

[Amo65]  Donald E. Amos.  A

third order differential equation. SIAM Review, 7(3): 409, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Amos:1978:BRD

[Amo78a] D. E. Amos. Book review: A Definite Integral of N. Bohr (P. J. Schweitzer). SIAM Review, 20(1):188– 190, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Amos:1978:ESC

[Amo78b] Donald E. Amos. Evaluation of some cumulative distribution functions by numerical quadrature. SIAM Review, 20(4):778–800, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ahuja:1991:SRA

[AMO91] Ravindra K. Ahuja, Thomas L. Magnanti, and James B. Orlin. Some recent advances in network flows. SIAM Review, 33(2):175–219, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ahuja:1995:NF

[AMOB95] R. K. Ahuja, T. L. Magnanti, J. B. Orlin, and Michael O. Ball. Network flows. SIAM Review, 37(1): 115–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alessandrini:2004:DIE

[AMR04] Giovanni Alessandrini, Antonino Morassi, and Edi Rosset. Detecting an inclusion in an elastic body by boundary measurements. SIAM Review, 46(3):477– 498, September 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44209.

Allard:2023:RDH

[AMS^{+}23] Antoine Allard, Cristopher Moore, Samuel V. Scarpino, Benjamin M. Althouse, and Laurent Hébert-Dufresne. The role of directionality, heterogeneity, and correlations in epidemic risk and spread. SIAM Review, 65 (2):471–492, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1383811.

Absil:2002:GRQ

[AMSV02] P. A. Absil, R. Mahony, R. Sepulchre, and P. Van Dooren. A Grassmann– Rayleigh quotient iteration for computing invariant subspaces. SIAM Review, 44 (1):57–73, March 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37864.

Aiello:1978:VSB

[AN78] W. Aiello and T. V. Narayana.

A variant of Silverman’s board of directors problem. SIAM Review, 20(2): 395, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Asfar:1983:AMM

[AN83] O. R. Asfar and A. H.

Nayfeh. The application of the method of multiple scales to wave propagation in periodic structures. SIAM Review, 25(4): 455–480, October 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anderson:1966:ETS

[And66] David R. Anderson. An existence theorem for a solution of f′(x) = F (x, f(g(x))). SIAM Review, 8(3):359– 362, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1973:BRS

[And73] George E. Andrews. Book review: A Set of Nonlinear Equations (L. Carlitz). SIAM Review, 15(1): 224–225, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1974:BRP

[And74a] G. E. Andrews. Book review: A Pair of Theta Function Identities (L. Carlitz). SIAM Review, 16(4): 553–555, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1974:RRI

[And74b] G. E. Andrews. The Rogers– Ramanujan identities. SIAM Review, 16(3):390, ???? 1974.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1974:ABH

[And74c] George E. Andrews. Applications of basic hypergeometric functions. SIAM Review, 16 (4):441–484, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anderson:1975:BRB

[And75a] D. H. Anderson. Book review: Bounds for the Zero of a Polynomial (V. E. Hoggatt, Jr.). SIAM Review, 17 (4):690–691, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1975:CP

[And75b] G. E. Andrews. A combinatorial problem. SIAM Review, 17(3):566, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1976:CP

[And76] G. E. Andrews. A combinatorial problem. SIAM Review, 18(3):495–496, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1977:BRO

[And77] L. C. Andrews. Book review: An n-th Order Linear Differential Equation (M. S.

Klamkin). SIAM Review, 19 (1):155, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1978:NAD

[And78] H. C. Andrews. Numerical analysis in digital image processing. SIAM Review, 20 (3):619, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1983:TRR

[And83] G. E. Andrews. Truncation of the rogers-Ramanujan theta series. SIAM Review, 25 (3):402, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1985:PFI

[And85] G. E. Andrews. A partition function identity. SIAM Review, 27(4):575, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrews:1987:PFI

[And87] G. E. Andrews. A partition function identity. SIAM Review, 29(3):471–473, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anderson:1990:RNG

[And90] Stuart L. Anderson. Random number generators on vector supercomputers and other advanced architectures. SIAM Review, 32(2): 221–251, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrew:1995:BRN

[And95] Alan L. Andrew. Book review: Numerical Solution of Sturm–Liouville Problems (John D. Pryce). SIAM Review, 37(2):252–253, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrew:1998:CM

[And98a] A. L. Andrew. Centrosymmetric matrices. SIAM Review, 40(3):697–698, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Andrew:1998:CNC

[And98b] Alan L. Andrew. Classroom note: Centrosymmetric matrices. SIAM Review, 40 (3):697–698, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32834.

Angel:1978:IRI

[Ang78] E. S. Angel. Image restoration and image models. SIAM Review, 20(3):619, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Angus:1990:GUP

[Ang90] John E. Angus. Generalization of the “universal pivot function” and its use in constructing confidence intervals and sets. SIAM Review, 32 (4):667–670, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Angus:1992:ADA

[Ang92a] John E. Angus. An alternative derivation of asymptotic normality for sample quantiles. SIAM Review, 34(2): 304–306, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Angus:1992:SBE

[Ang92b] John E. Angus. Some bounds on the error in approximating transition probabilities in continuous-time Markov processes. SIAM Review, 34(1): 110–113, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Angus:1994:PIT

[Ang94]  John E. Angus.  The

probability integral transform and related results. SIAM Review, 36(4):652– 654, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/26453.htm.

Angus:1997:IPI

[Ang97a]  J. E. Angus.  The in-

spection paradox inequality. SIAM Review, 39(1): 95–97, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Angus:1997:CNI

[Ang97b] John E. Angus. Classroom note: The inspection paradox inequality. SIAM Review, 39(1):95–97, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29298;   http:
//www.siam.org/journals/
sirev/sirev391.htm.

Anishev:1982:RPN

[Ani82]  P. A. Anishev.  Reduc-

tion of Petri nets. (Russian). SIAM Review, 24(1):41–54, 111, January 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Annaby:1995:EBV

[Ann95a] M. H. Annaby. Existence of a boundary-value problem with prescribed eigenvalues. SIAM Review, 37(2): 237–239, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Annaby:1995:BRE

[Ann95b] Mahmoud H. Annaby. Book review: Existence of a Boundary-Value Problem with Prescribed Eigenvalues (Ahmed I. Zayed). SIAM Review, 37(2):237–239, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1960:NNa

[Ano60a] Anonymous. News and notices. SIAM Review, 2(1): 58–76, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1960:NNb

[Ano60b] Anonymous. News and notices. SIAM Review, 2(2): 164–175, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1960:NNc

[Ano60c] Anonymous. News and notices. SIAM Review, 2(3): 230–240, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1960:NNd

[Ano60d] Anonymous. News and notices. SIAM Review, 2(4): 303–317, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1961:BRS

[Ano61a] Anonymous. Book review:

Stationary Processes and Prediction Theory (Harry Furstenberg). SIAM Review, 3(1):77, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1961:NNa

[Ano61b] Anonymous. News and notices. SIAM Review, 3(1): 85–104, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1961:NNb

[Ano61c] Anonymous. News and notices. SIAM Review, 3(2): 186–209, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1961:NNc

[Ano61d] Anonymous. News and notices. SIAM Review, 3(3): 270–292, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1961:NNd

[Ano61e] Anonymous. News and notices. SIAM Review, 3(4): 346–368, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1961:RSS

[Ano61f]  Anonymous.  Report of

symposium on satellite theory. SIAM Review, 3(2): 170–177, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1962:NNa

[Ano62a] Anonymous. News and notices. SIAM Review, 4(1): 59–78, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1962:NNb

[Ano62b] Anonymous. News and notices. SIAM Review, 4(2): 166–195, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1962:NNc

[Ano62c] Anonymous. News and notices. SIAM Review, 4(3): 273–296, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1962:NNd

[Ano62d] Anonymous. News and notices. SIAM Review, 4(4): 413–430, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1962:OBR

[Ano62e] Anonymous. Other books received. SIAM Review, 4(4): 411–412, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1963:BRD

[Ano63a] Anonymous. Book review:

A Definite Integral (W. L. Bade). SIAM Review, 5(2): 157–159, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1963:NNa

[Ano63b] Anonymous. News and notices. SIAM Review, 5(1): 91–112, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1963:NNb

[Ano63c] Anonymous. News and notices. SIAM Review, 5(2): 178–202, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1963:NNc

[Ano63d] Anonymous. News and notices. SIAM Review, 5(3): 300–320, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1963:NNd

[Ano63e] Anonymous. News and notices. SIAM Review, 5(4): 385–392, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1963:OBRa

[Ano63f] Anonymous. Other books received. SIAM Review, 5 (1):88–90, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1963:OBRb

[Ano63g] Anonymous. Other books received. SIAM Review, 5(2): 175–177, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1963:OBRc

[Ano63h] Anonymous. Other books received. SIAM Review, 5(3): 298–299, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1963:OBRd

[Ano63i] Anonymous. Other books received. SIAM Review, 5 (4):384, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1964:NNa

[Ano64a] Anonymous. News and notices. SIAM Review, 6(1): 85–109, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1964:NNb

[Ano64b] Anonymous. News and notices. SIAM Review, 6(2): 193–202, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1964:NNc

[Ano64c] Anonymous. News and notices. SIAM Review, 6(3): 328–341, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1964:NNd

[Ano64d] Anonymous. News and notices. SIAM Review, 6(4): 477–511, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1964:OBRa

[Ano64e] Anonymous. Other books received. SIAM Review, 6 (1):83–84, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1964:OBRb

[Ano64f] Anonymous. Other books received. SIAM Review, 6 (2):192, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1964:OBRc

[Ano64g] Anonymous. Other books received. SIAM Review, 6(3): 325–327, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1964:OBRd

[Ano64h] Anonymous. Other books received. SIAM Review, 6(4): 474–476, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1964:TCA

[Ano64i] Anonymous. Theoretical cybernetics abstracts. SIAM Review, 6(4):463, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1965:NNa

[Ano65a] Anonymous. News and notices. SIAM Review, 7(1): 164–179, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1965:NNb

[Ano65b] Anonymous. News and notices. SIAM Review, 7(2): 310–321, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1965:NNc

[Ano65c] Anonymous. News and notices. SIAM Review, 7(3): 446–461, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1965:NNd

[Ano65d] Anonymous. News and notices. SIAM Review, 7(4): 592–623, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1965:OBRa

[Ano65e] Anonymous. Other books received. SIAM Review, 7(1): 161–163, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1965:OBRb

[Ano65f] Anonymous. Other books received. SIAM Review, 7(2): 308–309, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1965:OBRc

[Ano65g] Anonymous. Other books received. SIAM Review, 7(3): 443–445, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1965:OBRd

[Ano65h] Anonymous. Other books received. SIAM Review, 7(4): 588–591, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:Eb

[Ano66a] Anonymous. Errata. SIAM Review, 8(4):526, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:Ea

[Ano66b] Anonymous. Erratum. SIAM Review, 8(1):107, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:NNa

[Ano66c] Anonymous. News and notices. SIAM Review, 8(1): 131–143, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:NNb

[Ano66d] Anonymous. News and notices. SIAM Review, 8(2):

251–275, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:NNc

[Ano66e] Anonymous. News and notices. SIAM Review, 8(3): 417–426, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:NNd

[Ano66f] Anonymous. News and notices. SIAM Review, 8(4): 556–591, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:OBRa

[Ano66g] Anonymous. Other books received. SIAM Review, 8(1): 128–130, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:OBRb

[Ano66h] Anonymous. Other books received. SIAM Review, 8(2): 248–250, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:OBRc

[Ano66i] Anonymous. Other books received. SIAM Review, 8(3): 413–416, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1966:OBRd

[Ano66j] Anonymous. Other books received. SIAM Review, 8(4): 552–555, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1967:NNa

[Ano67a] Anonymous. News and notices. SIAM Review, 9(1): 146–164, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1967:NNb

[Ano67b] Anonymous. News and notices. SIAM Review, 9(2): 270–287, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1967:NNc

[Ano67c] Anonymous. News and notices. SIAM Review, 9(3): 619–626, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1967:NNd

[Ano67d] Anonymous. News and notices. SIAM Review, 9(4): 769–777, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1967:OBRa

[Ano67e] Anonymous. Other books received. SIAM Review, 9(1): 142–145, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1967:OBRb

[Ano67f] Anonymous. Other books received. SIAM Review, 9(2): 267–269, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1967:OBRc

[Ano67g] Anonymous. Other books received. SIAM Review, 9(3): 614–618, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1967:OBRd

[Ano67h] Anonymous. Other books received. SIAM Review, 9(4): 765–768, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1967:P

[Ano67i] Anonymous. Participants.

SIAM Review, 9(2):292, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:BRC

[Ano68a]  Anonymous.  Book re-

view: Convergence of an Iteration (Joel L. Brenner). SIAM Review, 10(3):377– 379, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:Ca

[Ano68b]  Anonymous.  Chronicle.

SIAM Review, 10(1):101– 106, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:Cb

[Ano68c]  Anonymous.  Chronicle.

SIAM Review, 10(2):244– 290, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:Cc

[Ano68d]  Anonymous.  Chronicle.

SIAM Review, 10(3):403– 405, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:E

[Ano68e] Anonymous. Erratum. SIAM Review, 10(4):500, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:NM

[Ano68f]  Anonymous.  New mem-

bers. SIAM Review, 10(4): 490–499, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:OBRa

[Ano68g] Anonymous. Other books received. SIAM Review, 10(1): 97–100, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:OBRb

[Ano68h] Anonymous. Other books received. SIAM Review, 10(2): 242–243, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:OBRc

[Ano68i] Anonymous. Other books received. SIAM Review, 10(3): 399–402, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:OBRd

[Ano68j] Anonymous. Other books received. SIAM Review, 10(4): 464–466, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:PN

[Ano68k] Anonymous. Personal notices. SIAM Review, 10 (4):500, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1968:SNM

[Ano68l] Anonymous. SIAM 1968 national meeting. SIAM Review, 10(4):467–490, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:C

[Ano69a]  Anonymous.  Chronicle.

SIAM Review, 11(2):301– 306, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:D

[Ano69b] Anonymous. Deaths. SIAM Review, 11(3):427, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:NMa

[Ano69c]  Anonymous.  New mem-

bers. SIAM Review, 11(1): 120–124, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:NMb

[Ano69d]  Anonymous.  New mem-

bers. SIAM Review, 11(3): 421–427, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:NMc

[Ano69e]  Anonymous.  New mem-

bers. SIAM Review, 11(4):

670–672, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:OBRa

[Ano69f] Anonymous. Other books received. SIAM Review, 11(1): 104–106, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:OBRb

[Ano69g] Anonymous. Other books received. SIAM Review, 11(2): 297–300, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:OBRc

[Ano69h] Anonymous. Other books received. SIAM Review, 11(3): 417–420, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:OBRd

[Ano69i] Anonymous. Other books received. SIAM Review, 11(4): 638–640, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:PN

[Ano69j] Anonymous. Pesonal notices. SIAM Review, 11 (1):125, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:SFM

[Ano69k] Anonymous. SIAM 1968 fall meeting. SIAM Review, 11 (1):107–120, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1969:SNM

[Ano69l] Anonymous. SIAM 1969 national meeting. SIAM Review, 11(4):641–669, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:Ca

[Ano70a]  Anonymous.  Chronicle.

SIAM Review, 12(1):170– 174, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:Cb

[Ano70b]  Anonymous.  Chronicle.

SIAM Review, 12(2):322– 324, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:Da

[Ano70c] Anonymous. Deaths. SIAM Review, 12(3):480, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:Db

[Ano70d] Anonymous. Deaths. SIAM Review, 12(4):602, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:NM

[Ano70e]  Anonymous.  New mem-

bers. SIAM Review, 12(4): 599–602, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:OBRa

[Ano70f] Anonymous. Other books received. SIAM Review, 12(1): 166–169, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:OBRb

[Ano70g] Anonymous. Other books received. SIAM Review, 12(2): 318–321, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:OBRc

[Ano70h] Anonymous. Other books received. SIAM Review, 12(3): 477–478, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:OBRd

[Ano70i] Anonymous. Other books received. SIAM Review, 12(4):

596–598, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1970:SRJ

[Ano70j]  Anonymous.  SIAM ref-

erees July–December 1969. SIAM Review, 12(3):479– 480, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:BRD

[Ano71a] Anonymous. Book review: A Definite Integral (A. M. Cohen and F. J. Wilson). SIAM Review, 13(3):399, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:Da

[Ano71b] Anonymous. Deaths. SIAM Review, 13(2):288, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:Db

[Ano71c] Anonymous. Deaths. SIAM Review, 13(3):423, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:Dc

[Ano71d] Anonymous. Deaths. SIAM Review, 13(4):595, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:NMa

[Ano71e]  Anonymous.  New mem-

bers. SIAM Review, 13(1): 162–167, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:NMb

[Ano71f]  Anonymous.  New mem-

bers. SIAM Review, 13(2): 285–288, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:NMc

[Ano71g]  Anonymous.  New mem-

bers. SIAM Review, 13(4): 589–595, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:OBRa

[Ano71h] Anonymous. Other books received. SIAM Review, 13(1): 111–112, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:OBRb

[Ano71i] Anonymous. Other books received. SIAM Review, 13(2): 242–245, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:OBRc

[Ano71j] Anonymous. Other books received. SIAM Review, 13(3): 414–417, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:OBRd

[Ano71k] Anonymous. Other books received. SIAM Review, 13(4): 586–588, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:SN

[Ano71l] Anonymous. Section news. SIAM Review, 13(3):418– 420, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:SFM

[Ano71m] Anonymous. SIAM 1970 fall meeting. SIAM Review, 13 (2):257–285, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:SNM

[Ano71n] Anonymous. SIAM 1970 national meeting. SIAM Review, 13(1):138–161, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:SRJa

[Ano71o]  Anonymous.  SIAM ref-

erees January–June 1970. SIAM Review, 13(1):161– 162, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1971:SRJb

[Ano71p]  Anonymous.  SIAM ref-

erees July–December 1970. SIAM Review, 13(3):421– 423, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:Da

[Ano72a] Anonymous. Deaths. SIAM Review, 14(2):384, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:Db

[Ano72b] Anonymous. Deaths. SIAM Review, 14(3):546, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:NMa

[Ano72c]  Anonymous.  New mem-

bers. SIAM Review, 14(1): 220–223, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:NMb

[Ano72d]  Anonymous.  New mem-

bers. SIAM Review, 14(2): 383–384, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:NMc

[Ano72e]  Anonymous.  New mem-

bers. SIAM Review, 14(3): 545–546, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:NMd

[Ano72f]  Anonymous.  New mem-

bers. SIAM Review, 14(4): 673–676, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:OBRa

[Ano72g] Anonymous. Other books received. SIAM Review, 14(1): 188–191, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:OBRb

[Ano72h] Anonymous. Other books received. SIAM Review, 14(2): 362–363, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:OBRc

[Ano72i] Anonymous. Other books received. SIAM Review, 14(3): 517–519, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:OBRd

[Ano72j] Anonymous. Other books received. SIAM Review, 14(4): 669–672, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:SN

[Ano72k] Anonymous. Section news.

SIAM Review, 14(1):216– 219, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:SFM

[Ano72l] Anonymous. SIAM 1971 fall meeting. SIAM Review, 14 (3):520–545, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:SNM

[Ano72m] Anonymous. SIAM 1971 national meeting. SIAM Review, 14(1):192–216, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:SRJ

[Ano72n]  Anonymous.  SIAM ref-

erees January–June 1971. SIAM Review, 14(1):219– 220, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1972:SIA

[Ano72o] Anonymous. Society for industrial and applied mathematics national officers, 1972 officers. SIAM Review, 14 (2):382, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:BRF

[Ano73a] Anonymous. Book review: A Fourier Transform (D. K. Rao). SIAM Review, 15(4): 801–802, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:Ca

[Ano73b]  Anonymous.  Chronicle.

SIAM Review, 15(1):243– 273, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:Cb

[Ano73c]  Anonymous.  Chronicle.

SIAM Review, 15(4):815– 820, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:EAP

[Ano73d] Anonymous. Erratum for April 1973 problem section. SIAM Review, 15 (4):803, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:NMa

[Ano73e]  Anonymous.  New mem-

bers. SIAM Review, 15(2): 441–443, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:NMb

[Ano73f]  Anonymous.  New mem-

bers. SIAM Review, 15(3): 683–685, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:OBRa

[Ano73g] Anonymous. Other books received. SIAM Review, 15(1): 238–242, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:OBRb

[Ano73h] Anonymous. Other books received. SIAM Review, 15(2): 407–408, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:OBRc

[Ano73i] Anonymous. Other books received. SIAM Review, 15(3): 680–682, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:OBRd

[Ano73j] Anonymous. Other books received. SIAM Review, 15(4): 809–814, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:P

[Ano73k] Anonymous. Preface. SIAM Review, 15(2):iii, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:SFM

[Ano73l] Anonymous. SIAM 1972 fall meeting. SIAM Review, 15 (2):409–440, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:SIM

[Ano73m] Anonymous. SIAM Institute for Mathematics and Society. SIAM Review, 15 (4):820, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1973:SIA

[Ano73n] Anonymous. Society for industrial and applied mathematics. SIAM Review, 15(2): 444–446, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:D

[Ano74a] Anonymous. Deaths. SIAM Review, 16(3):417, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:NMa

[Ano74b]  Anonymous.  New mem-

bers. SIAM Review, 16(2): 288–293, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:NMb

[Ano74c]  Anonymous.  New mem-

bers. SIAM Review, 16(3): 415–417, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:NMC

[Ano74d] Anonymous. New membership categories. SIAM Review, 16(3):417, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:OBRa

[Ano74e] Anonymous. Other books received. SIAM Review, 16(1): 110–112, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:OBRb

[Ano74f] Anonymous. Other books received. SIAM Review, 16(2): 276–278, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:OBRc

[Ano74g] Anonymous. Other books received. SIAM Review, 16(3): 409–411, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:OBRd

[Ano74h] Anonymous. Other books received. SIAM Review, 16(4): 572–575, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:SN

[Ano74i] Anonymous. Section news. SIAM Review, 16(3):412– 414, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:SNMa

[Ano74j] Anonymous. SIAM 1973 national meeting. SIAM Review, 16(1):113–141, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:SNMb

[Ano74k] Anonymous. SIAM 1974 national meeting. SIAM Review, 16(4):576–592, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1974:SFM

[Ano74l] Anonymous. SIAM fall meeting 1973. SIAM Review, 16 (2):279–287, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:Aa

[Ano75a]  Anonymous.  Announce-

ment. SIAM Review, 17 (1):166, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:Ab

[Ano75b]  Anonymous.  Announce-

ment. SIAM Review, 17 (3):558, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:Ac

[Ano75c]  Anonymous.  Announce-

ment. SIAM Review, 17 (4):684, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:CNA

[Ano75d]  Anonymous.  Classroom

notes in applied mathematics. SIAM Review, 17(2): 371, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:CSM

[Ano75e]  Anonymous.  Council of

SIAM meeting. SIAM Review, 17(1):194–195, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:D

[Ano75f] Anonymous. Deaths. SIAM Review, 17(1):195, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:MTN

[Ano75g]  Anonymous.  The math-

ematical training of the nonacademic mathematician. SIAM Review, 17(3):541– 557, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:NMa

[Ano75h]  Anonymous.  New mem-

bers. SIAM Review, 17(1): 186–193, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:NMb

[Ano75i]  Anonymous.  New mem-

bers. SIAM Review, 17(2): 393–394, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:NMc

[Ano75j]  Anonymous.  New mem-

bers. SIAM Review, 17(4): 720–721, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:OBRa

[Ano75k] Anonymous. Other books received. SIAM Review, 17(1): 184–185, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:OBRb

[Ano75l] Anonymous. Other books received. SIAM Review, 17(3): 582–586, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:OBRc

[Ano75m] Anonymous. Other books received. SIAM Review, 17(4): 717–719, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:SFM

[Ano75n] Anonymous. SIAM fall meetings 1974. SIAM Review, 17 (2):372–393, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:SIA

[Ano75o] Anonymous. Society for industrial and applied mathematics. SIAM Review, 17(3): 588–592, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1975:TR

[Ano75p] Anonymous. Treasurer’s report. SIAM Review, 17 (3):587, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:Aa

[Ano76a]  Anonymous.  Announce-

ment. SIAM Review, 18 (1):113, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:Ab

[Ano76b]  Anonymous.  Announce-

ment. SIAM Review, 18 (2):293, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:Ac

[Ano76c]  Anonymous.  Announce-

ment. SIAM Review, 18 (3):485, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:Ad

[Ano76d]  Anonymous.  Announce-

ment. SIAM Review, 18 (4):761, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:C

[Ano76e]  Anonymous.  Chronicle.

SIAM Review, 18(1):155– 159, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:Da

[Ano76f] Anonymous. Deaths. SIAM Review, 18(3):544, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:Db

[Ano76g] Anonymous. Deaths. SIAM Review, 18(4):841, ???? 1976.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:NMa

[Ano76h]  Anonymous.  New mem-

bers. SIAM Review, 18(3): 538–544, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:NMb

[Ano76i]  Anonymous.  New mem-

bers. SIAM Review, 18(4): 836–841, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:OBRa

[Ano76j] Anonymous. Other books received. SIAM Review, 18(1): 152–154, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:OBRb

[Ano76k] Anonymous. Other books received. SIAM Review, 18(2): 322–324, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:OBRc

[Ano76l] Anonymous. Other books received. SIAM Review, 18(3): 533–537, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:OBRd

[Ano76m] Anonymous. Other books received. SIAM Review, 18(4): 788–789, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:SN

[Ano76n] Anonymous. Section news.

SIAM Review, 18(3):538, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:SNM

[Ano76o] Anonymous. SIAM 1975 national meeting. SIAM Review, 18(2):325–349, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1976:SSF

[Ano76p] Anonymous. SIAM SIGNUM 1975 fall meeting. SIAM Review, 18(4):790–835, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:Aa

[Ano77a]  Anonymous.  Announce-

ment. SIAM Review, 19 (1):139, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:Ab

[Ano77b]  Anonymous.  Announce-

ment. SIAM Review, 19 (2):319, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:Ac

[Ano77c]  Anonymous.  Announce-

ment. SIAM Review, 19 (3):550, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:Ad

[Ano77d]  Anonymous.  Announce-

ment. SIAM Review, 19 (4):726, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:D

[Ano77e] Anonymous. Death. SIAM Review, 19(3):591, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:NJ

[Ano77f]  Anonymous.  New jour-

nals. SIAM Review, 19(1): 179–180, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:NMa

[Ano77g]  Anonymous.  New mem-

bers. SIAM Review, 19(1):

181–187, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:NMb

[Ano77h]  Anonymous.  New mem-

bers. SIAM Review, 19(3): 586–591, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:NMc

[Ano77i]  Anonymous.  New mem-

bers. SIAM Review, 19(4): 784–785, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:OBRa

[Ano77j] Anonymous. Other books received. SIAM Review, 19(1): 176–179, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:OBRb

[Ano77k] Anonymous. Other books received. SIAM Review, 19(2): 369–371, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:OBRc

[Ano77l] Anonymous. Other books received. SIAM Review, 19 (3):585, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:OBRd

[Ano77m] Anonymous. Other books received. SIAM Review, 19(4): 754–758, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:OJR

[Ano77n] Anonymous. Other journals received. SIAM Review, 19 (2):371, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:SFM

[Ano77o] Anonymous. SIAM 1976 fall meeting. SIAM Review, 19 (4):759–783, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:SNM

[Ano77p] Anonymous. SIAM 1976 national meeting. SIAM Review, 19(2):372–401, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:SIA

[Ano77q] Anonymous. Society for industrial and applied mathematics. SIAM Review, 19(4): 786–789, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1977:TR

[Ano77r] Anonymous. Treasurer’s report. SIAM Review, 19 (4):790, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:Aa

[Ano78a]  Anonymous.  Announce-

ment. SIAM Review, 20 (1):168, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:Ab

[Ano78b]  Anonymous.  Announce-

ment. SIAM Review, 20 (2):388, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:Ac

[Ano78c]  Anonymous.  Announce-

ment. SIAM Review, 20 (3):583, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:Ad

[Ano78d]  Anonymous.  Announce-

ment. SIAM Review, 20(4): 846, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:CNAa

[Ano78e]  Anonymous.  Classroom

notes in applied mathematics. SIAM Review, 20(3): 583, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:CNAb

[Ano78f]  Anonymous.  Classroom

notes in applied mathematics. SIAM Review, 20(4): 846, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:C

[Ano78g]  Anonymous.  Corrigenda.

SIAM Review, 20(4):863, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:E

[Ano78h] Anonymous. Erratum. SIAM Review, 20(1):212, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:MFJ

[Ano78i] Anonymous. In memoriam — Fritz Joachim Weyl. SIAM Review, 20(1):vi, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:NMa

[Ano78j]  Anonymous.  New mem-

bers. SIAM Review, 20(1): 209–212, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:NMb

[Ano78k]  Anonymous.  New mem-

bers. SIAM Review, 20(4): 879–886, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:OBRa

[Ano78l] Anonymous. Other books received. SIAM Review, 20(1): 205–208, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:OBRb

[Ano78m] Anonymous. Other books received. SIAM Review, 20(2): 412–416, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:OBRc

[Ano78n] Anonymous. Other books received. SIAM Review, 20(3): 615–616, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:OBRd

[Ano78o] Anonymous. Other books received. SIAM Review, 20(4): 876–878, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:SFM

[Ano78p] Anonymous. SIAM 1977 fall meeting. SIAM Review, 20 (3):617–637, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:SNM

[Ano78q] Anonymous. SIAM 1977 national meeting. SIAM Review, 20(2):417–441, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1978:TR

[Ano78r] Anonymous. Treasurer’s report. SIAM Review, 20(4): 887, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:Aa

[Ano79a]  Anonymous.  Announce-

ment. SIAM Review, 21 (1):128, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:Ab

[Ano79b]  Anonymous.  Announce-

ment. SIAM Review, 21 (2):246, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:Ac

[Ano79c]  Anonymous.  Announce-

ment. SIAM Review, 21 (3):389, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:C

[Ano79d]  Anonymous.  Chronicle.

SIAM Review, 21(2):281– 288, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:ECT

[Ano79e] Anonymous. Edge 3-coloring of tournaments. SIAM Review, 21(3):397–398, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:NMa

[Ano79f]  Anonymous.  New mem-

bers. SIAM Review, 21(1): 155–156, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:NMb

[Ano79g]  Anonymous.  New mem-

bers. SIAM Review, 21(3): 437–441, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:NMc

[Ano79h]  Anonymous.  New mem-

bers. SIAM Review, 21(4): 587–595, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:OBRa

[Ano79i] Anonymous. Other books received. SIAM Review, 21(1): 153–154, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:OBRb

[Ano79j] Anonymous. Other books received. SIAM Review, 21(2): 278–280, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:OBRc

[Ano79k] Anonymous. Other books received. SIAM Review, 21(3): 432–436, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:OBRd

[Ano79l] Anonymous. Other books received. SIAM Review, 21(4): 583–586, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1979:TR

[Ano79m] Anonymous. Treasurer’s report. SIAM Review, 21 (4):596, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:Aa

[Ano80a]  Anonymous.  Announce-

ment. SIAM Review, 22 (1):87, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:Ab

[Ano80b]  Anonymous.  Announce-

ment. SIAM Review, 22 (2):225, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:Ac

[Ano80c]  Anonymous.  Announce-

ment. SIAM Review, 22 (4):494, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:BRE

[Ano80d] Anonymous. Book review: Erratum: an Integral Inequality (J. W. Wilkins, Jr.). SIAM Review, 22(3):373– 374, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:EAOa

[Ano80e] Anonymous. Elected and appointed officials. SIAM Review, 22(3):398–399, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:EAOb

[Ano80f] Anonymous. Elected and appointed officials. SIAM Review, 22(4):535–536, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:LS

[Ano80g]  Anonymous.  Late solu-

tions. SIAM Review, 22 (1):102, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:NJ

[Ano80h]  Anonymous.  New jour-

nal. SIAM Review, 22(2): 252, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:NMa

[Ano80i]  Anonymous.  New mem-

bers. SIAM Review, 22(1): 113–117, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:NMb

[Ano80j]  Anonymous.  New mem-

bers. SIAM Review, 22(2): 256–262, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:NMc

[Ano80k]  Anonymous.  New mem-

bers. SIAM Review, 22(3): 393–397, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:OBRa

[Ano80l] Anonymous. Other books received. SIAM Review, 22(1): 109–112, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:OBRb

[Ano80m] Anonymous. Other books received. SIAM Review, 22(2): 252–255, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:OBRc

[Ano80n] Anonymous. Other books received. SIAM Review, 22(3): 388–392, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:OBRd

[Ano80o] Anonymous. Other books received. SIAM Review, 22(4): 532–534, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1980:TR

[Ano80p] Anonymous. Treasurer’s report. SIAM Review, 22 (4):537, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:A

[Ano81a]  Anonymous.  Announce-

ment. SIAM Review, 23 (3):380, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:EAO

[Ano81b] Anonymous. Elected and appointed officials. SIAM Review, 23(4):564–565, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:LS

[Ano81c]  Anonymous.  Late solu-

tions. SIAM Review, 23 (1):113, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:NMa

[Ano81d]  Anonymous.  New mem-

bers. SIAM Review, 23(1): 133–141, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:NMb

[Ano81e]  Anonymous.  New mem-

bers. SIAM Review, 23(3): 416–424, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:NMc

[Ano81f]  Anonymous.  New mem-

bers. SIAM Review, 23(4): 558–563, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:OBRa

[Ano81g] Anonymous. Other books received. SIAM Review, 23(1): 128–132, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:OBRc

[Ano81h] Anonymous. Other books received. SIAM Review, 23(3): 412–415, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:OBRd

[Ano81i] Anonymous. Other books received. SIAM Review, 23(4): 555–557, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:OBRb

[Ano81j] Anonymous. Other books recived. SIAM Review, 23(2): 275–277, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1981:TR

[Ano81k] Anonymous. Treasurer’s report. SIAM Review, 23 (4):566, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:A

[Ano82a]  Anonymous.  Announce-

ment. SIAM Review, 24 (1):72, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:AACa

[Ano82b] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 24(1): 104–108, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:AACb

[Ano82c] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 24(2): 248–253, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:AACc

[Ano82d] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 24(3): 373–377, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:AACd

[Ano82e] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 24(4): 505–510, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:BRL

[Ano82f] Anonymous. Book review:

Late Solutions (G. Fairweather and C. Surma). SIAM Review, 24(1):85, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:EAO

[Ano82g] Anonymous. Elected and appointed officials. SIAM Review, 24(1):112, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:NJ

[Ano82h]  Anonymous.  New jour-

nal. SIAM Review, 24(4): 504, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:NMa

[Ano82i]  Anonymous.  New mem-

bers. SIAM Review, 24(1): 108–111, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:NMb

[Ano82j]  Anonymous.  New mem-

bers. SIAM Review, 24(2): 253–261, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:NMc

[Ano82k]  Anonymous.  New mem-

bers. SIAM Review, 24(3): 378–380, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:NMd

[Ano82l]  Anonymous.  New mem-

bers. SIAM Review, 24(4): 511–514, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:OBRa

[Ano82m] Anonymous. Other books received. SIAM Review, 24(1): 101–103, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:OBRb

[Ano82n] Anonymous. Other books received. SIAM Review, 24(2): 244–247, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:OBRc

[Ano82o] Anonymous. Other books received. SIAM Review, 24(3): 369–371, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:OBRd

[Ano82p] Anonymous. Other books received. SIAM Review, 24(4): 502–504, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1982:TR

[Ano82q] Anonymous. Treasurer’s report. SIAM Review, 24(4): 515–516, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:AACa

[Ano83a] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 25(1): 144–148, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:AACb

[Ano83b] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 25(2): 295–300, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:AACc

[Ano83c] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 25(3): 447–451, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:AACd

[Ano83d] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 25(4): 604–609, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:EAO

[Ano83e] Anonymous. Elected and appointed officials. SIAM Review, 25(1):149, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:LS

[Ano83f]  Anonymous.  Late solu-

tions. SIAM Review, 25 (1):107, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:NMa

[Ano83g]  Anonymous.  New mem-

bers. SIAM Review, 25(2): 301–310, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:NMb

[Ano83h]  Anonymous.  New mem-

bers. SIAM Review, 25(3): 452–454, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:NMc

[Ano83i]  Anonymous.  New mem-

bers. SIAM Review, 25(4):

610–616, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:OBRa

[Ano83j] Anonymous. Other books received. SIAM Review, 25(1): 139–143, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:OBRb

[Ano83k] Anonymous. Other books received. SIAM Review, 25(2): 290–294, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:OBRc

[Ano83l] Anonymous. Other books received. SIAM Review, 25(3): 442–446, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:OBRd

[Ano83m] Anonymous. Other books received. SIAM Review, 25(4): 600–603, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1983:TR

[Ano83n] Anonymous. Treasurer’s report. SIAM Review, 25 (4):617, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:AACa

[Ano84a] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 26(1): 151–156, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:AACb

[Ano84b] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 26(2): 312–317, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:AACc

[Ano84c] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 26(3): 461–466, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:AACd

[Ano84d] Anonymous. Articles appearing in current and future issues of other SIAM journals. SIAM Review, 26(4): 624–629, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:EAO

[Ano84e] Anonymous. Elected and appointed officials. SIAM Review, 26(1):162, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:LSa

[Ano84f]  Anonymous.  Late solu-

tions. SIAM Review, 26 (2):283, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:LSb

[Ano84g]  Anonymous.  Late solu-

tions. SIAM Review, 26 (4):587, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:NMa

[Ano84h]  Anonymous.  New mem-

bers. SIAM Review, 26(1): 157–161, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:NMb

[Ano84i]  Anonymous.  New mem-

bers. SIAM Review, 26(2): 318–322, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:NMc

[Ano84j]  Anonymous.  New mem-

bers. SIAM Review, 26(3): 467–472, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:NMd

[Ano84k]  Anonymous.  New mem-

bers. SIAM Review, 26(4): 630–636, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:OBRa

[Ano84l] Anonymous. Other books received. SIAM Review, 26(1): 145–150, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:OBRb

[Ano84m] Anonymous. Other books received. SIAM Review, 26(2): 309–311, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:OBRc

[Ano84n] Anonymous. Other books received. SIAM Review, 26(3): 454–460, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:OBRd

[Ano84o] Anonymous. Other books received. SIAM Review, 26(4): 616–623, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1984:TR

[Ano84p] Anonymous. Treasurer’s report. SIAM Review, 26 (4):637, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:EAO

[Ano85a] Anonymous. Elected and appointed officials. SIAM Review, 27(1):148, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:LS

[Ano85b]  Anonymous.  Late solu-

tions. SIAM Review, 27 (1):85, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:NMa

[Ano85c]  Anonymous.  New mem-

bers. SIAM Review, 27(1): 139–147, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:NMb

[Ano85d]  Anonymous.  New mem-

bers. SIAM Review, 27(2): 328–334, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:NMc

[Ano85e]  Anonymous.  New mem-

bers. SIAM Review, 27(3): 480–484, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:NMd

[Ano85f]  Anonymous.  New mem-

bers. SIAM Review, 27(4): 618–621, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:OBRa

[Ano85g] Anonymous. Other books received. SIAM Review, 27(1): 133–138, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:OBRb

[Ano85h] Anonymous. Other books received. SIAM Review, 27(2): 315–327, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:OBRc

[Ano85i] Anonymous. Other books received. SIAM Review, 27(3): 473–479, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:OBRd

[Ano85j] Anonymous. Other books received. SIAM Review, 27(4): 610–617, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1985:RTY

[Ano85k]  Anonymous.  Report of

the Treasurer for the year ending December 31, 1984. SIAM Review, 27(4):622– 623, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:EAOa

[Ano86a] Anonymous. Elected and appointed officials. SIAM Review, 28(1):151, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:EAOb

[Ano86b] Anonymous. Elected and appointed officials. SIAM Review, 28(2):306, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:EAOc

[Ano86c] Anonymous. Elected and appointed officials. SIAM Review, 28(3):457, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:NMa

[Ano86d]  Anonymous.  New mem-

bers. SIAM Review, 28(1): 143–150, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:NMb

[Ano86e]  Anonymous.  New mem-

bers. SIAM Review, 28(2): 296–305, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:NMc

[Ano86f]  Anonymous.  New mem-

bers. SIAM Review, 28(3): 452–456, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:NMd

[Ano86g]  Anonymous.  New mem-

bers. SIAM Review, 28 (4):613–618, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:OBRc

[Ano86h] Anonymous. Other books received. SIAM Review, 28(3): 444–451, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:OBRa

[Ano86i] Anonymous. Other books recieved. SIAM Review, 28(1): 135–142, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:OBRb

[Ano86j] Anonymous. Other books recieved. SIAM Review, 28(2): 288–295, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:OBRd

[Ano86k] Anonymous. Other books reviews. SIAM Review, 28 (4):604–612, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1986:RTY

[Ano86l]  Anonymous.  Report of

the treasure for the year ending December 31, 1985. SIAM Review, 28(4):619– 620, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:BRL

[Ano87a] Anonymous. Book review:

Late Solutions (S. Lj. Damjanović). SIAM Review, 29 (1):135, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:BRP

[Ano87b] Anonymous. Book review:

Problem 85-20 (Dec. 1986, p. 574). SIAM Review, 29 (4):637, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:EAO

[Ano87c] Anonymous. Elected and appointed officials. SIAM Review, 29(1):183, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:NMa

[Ano87d]  Anonymous.  New mem-

bers. SIAM Review, 29(1): 175–182, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:NMb

[Ano87e]  Anonymous.  New mem-

bers. SIAM Review, 29(2): 349–357, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:NMc

[Ano87f]  Anonymous.  New mem-

bers. SIAM Review, 29(3): 517–523, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:NMd

[Ano87g]  Anonymous.  New mem-

bers. SIAM Review, 29(4): 689–694, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:OBRa

[Ano87h] Anonymous. Other books received. SIAM Review, 29(1): 167–174, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:OBRb

[Ano87i] Anonymous. Other books received. SIAM Review, 29(2): 337–347, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:OBRc

[Ano87j] Anonymous. Other books received. SIAM Review, 29(3): 508–516, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:OBRd

[Ano87k] Anonymous. Other books received. SIAM Review, 29(4): 680–688, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1987:RTY

[Ano87l]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1986. SIAM Review, 29(4):695– 696, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:EAO

[Ano88a] Anonymous. Elected and appointed officials. SIAM Review, 30(1):192, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:NMa

[Ano88b]  Anonymous.  New mem-

bers. SIAM Review, 30(1): 183–191, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:NMb

[Ano88c]  Anonymous.  New mem-

bers. SIAM Review, 30(2): 369–373, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:NMc

[Ano88d]  Anonymous.  New mem-

bers. SIAM Review, 30(3): 547–550, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:NMd

[Ano88e]  Anonymous.  New mem-

bers. SIAM Review, 30(4): 691–693, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:OBRa

[Ano88f] Anonymous. Other books received. SIAM Review, 30(1): 172–181, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:OBRb

[Ano88g] Anonymous. Other books received. SIAM Review, 30(2): 355–368, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:OBRc

[Ano88h] Anonymous. Other books received. SIAM Review, 30(3): 536–546, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:OBRd

[Ano88i] Anonymous. Other books received. SIAM Review, 30(4): 683–690, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1988:RTY

[Ano88j]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1987. SIAM Review, 30(4):694– 695, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:B

[Ano89a]  Anonymous.  Biography.

SIAM Review, 31(1):118– 119, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:BRP

[Ano89b] Anonymous. Book review:

A Probability Generating Function (A. A. Jagers). SIAM Review, 31(4):682– 683, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:EAO

[Ano89c] Anonymous. Elected and appointed official. SIAM Review, 31(1):190, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:MS

[Ano89d] Anonymous. A multinomial summation. SIAM Review, 31(1):122–125, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:NMa

[Ano89e]  Anonymous.  New mem-

bers. SIAM Review, 31(1): 185–189, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:NMb

[Ano89f]  Anonymous.  New mem-

bers. SIAM Review, 31(2): 360–364, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:NMc

[Ano89g]  Anonymous.  New mem-

bers. SIAM Review, 31(3): 527–532, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:NMd

[Ano89h]  Anonymous.  New mem-

bers. SIAM Review, 31(4): 721–725, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:OBRa

[Ano89i] Anonymous. Other books received. SIAM Review, 31(1): 172–184, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:OBRb

[Ano89j] Anonymous. Other books received. SIAM Review, 31(2): 350–359, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:OBRc

[Ano89k] Anonymous. Other books received. SIAM Review, 31(3): 519–526, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:OBRd

[Ano89l] Anonymous. Other books received. SIAM Review, 31(4): 711–720, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1989:RTY

[Ano89m]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1988. SIAM Review, 31(4):726– 727, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:CPL

[Ano90a] Anonymous. Cyclic pursuit with lead angle (quickie). SIAM Review, 32(4):674, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:EAO

[Ano90b] Anonymous. Elected and appointed officials. SIAM Review, 32(1):210, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:NMa

[Ano90c]  Anonymous.  New mem-

bers. SIAM Review, 32(1): 203–209, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:NMb

[Ano90d]  Anonymous.  New mem-

bers. SIAM Review, 32(2): 344–353, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:NMc

[Ano90e]  Anonymous.  New mem-

bers. SIAM Review, 32(3): 531–535, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:NMd

[Ano90f]  Anonymous.  New mem-

bers. SIAM Review, 32(4): 718–720, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:OBRa

[Ano90g] Anonymous. Other books received. SIAM Review, 32(1): 193–202, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:OBRb

[Ano90h] Anonymous. Other books received. SIAM Review, 32(2): 335–343, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:OBRc

[Ano90i] Anonymous. Other books received. SIAM Review, 32(3): 522–530, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:OBRd

[Ano90j] Anonymous. Other books received. SIAM Review, 32(4): 709–717, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:Qa

[Ano90k] Anonymous. Quickie. SIAM Review, 32(1):154, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:Qb

[Ano90l] Anonymous. Quickie. SIAM Review, 32(2):309, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1990:RTY

[Ano90m]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1989. SIAM Review, 32(4):721– 722, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:PMC

[Ano91a] Anonymous. 3 particles moving on a circle. SIAM Review, 33(1):118–119, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:BRT

[Ano91b]  Anonymous.  Book re-

view: Two Trigonometric Multiple Integrals (Quickie). SIAM Review, 33(1):127– 128, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:EAO

[Ano91c] Anonymous. Elected and appointed officials. SIAM Review, 33(1):173, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:I

[Ano91d] Anonymous. An inequality. SIAM Review, 33(1): 116–118, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:LS

[Ano91e]  Anonymous.  Late solu-

tions. SIAM Review, 33 (3):483, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:NMa

[Ano91f]  Anonymous.  New mem-

bers. SIAM Review, 33(1): 164–172, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:NMb

[Ano91g]  Anonymous.  New mem-

bers. SIAM Review, 33(2): 341–348, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:NMc

[Ano91h]  Anonymous.  New mem-

bers. SIAM Review, 33(3): 513–518, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:NMd

[Ano91i]  Anonymous.  New mem-

bers. SIAM Review, 33(4): 690–692, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:Qa

[Ano91j] Anonymous. Quickie. SIAM Review, 33(2):283, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:Qb

[Ano91k] Anonymous. Quickie. SIAM Review, 33(3):483, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:RTY

[Ano91l]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1990. SIAM Review, 33(4):693– 694, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1991:TDI

[Ano91m] Anonymous. Two definite integrals. SIAM Review, 33(1): 120–122, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1992:BRD

[Ano92a] Anonymous. Book review:

A Definite Integral Arising in Ohmic Dissipation (Quickie). SIAM Review, 34 (1):126, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1992:BRS

[Ano92b] Anonymous. Book review:

A Set of Maxima Problems (Quickie). SIAM Review, 34 (2):323–324, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1992:EAO

[Ano92c] Anonymous. Elected and appointed officials. SIAM Review, 34(1):166, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1992:E

[Ano92d] Anonymous. Errata. SIAM Review, 34(1):126, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1992:NMa

[Ano92e]  Anonymous.  New mem-

bers. SIAM Review, 34(1): 160–165, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1992:NMb

[Ano92f]  Anonymous.  New mem-

bers. SIAM Review, 34(2): 350–359, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1992:NMc

[Ano92g]  Anonymous.  New mem-

bers. SIAM Review, 34(3): 523–527, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1992:NMd

[Ano92h]  Anonymous.  New mem-

bers. SIAM Review, 34(4): 693–696, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1992:RTY

[Ano92i]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1991. SIAM Review, 34(4):697– 698, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:BRS

[Ano93a] Anonymous. Book review: A Sextuple Integral From a Cosmological Model (Quickie). SIAM Review, 35(3):504– 505, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:BRI

[Ano93b]  Anonymous.  Book re-

view: An Infinite Integral (Quickie). SIAM Review, 35 (2):307, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:BREa

[Ano93c] Anonymous. Book review:

Ellipsoid of Inertia of a Regular Simplex (Quickie). SIAM Review, 35(1):141, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:BREb

[Ano93d] Anonymous. Book review:

Errata: a Linear Integro-Differential Equation (John Melville). SIAM Review, 35 (3):505, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:EAO

[Ano93e] Anonymous. Elected and appointed officials. SIAM Review, 35(1):182, ???? 1993.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:NMa

[Ano93f]  Anonymous.  New mem-

bers. SIAM Review, 35(1): 176–181, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:NMb

[Ano93g]  Anonymous.  New mem-

bers. SIAM Review, 35(2): 349–358, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:NMc

[Ano93h]  Anonymous.  New mem-

bers. SIAM Review, 35(3): 546–550, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:NMd

[Ano93i]  Anonymous.  New mem-

bers. SIAM Review, 35(4): 687–692, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:RTY

[Ano93j]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1992. SIAM Review, 35(4):693– 694, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:SCa

[Ano93k] Anonymous. Selected collections. SIAM Review, 35(3): 543–545, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1993:SCb

[Ano93l] Anonymous. Selected collections. SIAM Review, 35(4): 685–686, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:EAO

[Ano94a] Anonymous. Elected and appointed officials. SIAM Review, 36(1):156, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:E

[Ano94b] Anonymous. Errata. SIAM Review, 36(3):499–500, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:NMa

[Ano94c]  Anonymous.  New mem-

bers. SIAM Review, 36(1): 154–155, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:NMb

[Ano94d]  Anonymous.  New mem-

bers. SIAM Review, 36(2): 328–340, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:NMc

[Ano94e]  Anonymous.  New mem-

bers. SIAM Review, 36(3): 532–536, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:NMd

[Ano94f]  Anonymous.  New mem-

bers. SIAM Review, 36 (4):687–689, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:RTY

[Ano94g]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1993. SIAM Review, 36(4):690– 691, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:SCa

[Ano94h] Anonymous. Selected collections. SIAM Review, 36(1): 150–153, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:SCb

[Ano94i] Anonymous. Selected collections. SIAM Review, 36(2): 325–327, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:SCc

[Ano94j] Anonymous. Selected collections. SIAM Review, 36(3): 530–531, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1994:SCd

[Ano94k] Anonymous. Selected collections. SIAM Review, 36 (4):683–686, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:Ca

[Ano95a]  Anonymous.  Chronicle.

SIAM Review, 37(1):142– ??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:Cb

[Ano95b]  Anonymous.  Chronicle.

SIAM Review, 37(2):288– ??, June 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:Cc

[Ano95c]  Anonymous.  Chronicle.

SIAM Review, 37(3):486– 489, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:Cd

[Ano95d]  Anonymous.  Chronicle.

SIAM Review, 37(4):644– ??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:EAO

[Ano95e] Anonymous. Elected and appointed officials. SIAM Review, 37(1):149, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:E

[Ano95f] Anonymous. Errata. SIAM Review, 37(2):250, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:LEa

[Ano95g]  Anonymous.  Later edi-

tions. SIAM Review, 37(2): 286–287, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:LEb

[Ano95h]  Anonymous.  Later edi-

tions. SIAM Review, 37 (3):484–485, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:LEc

[Ano95i]  Anonymous.  Later edi-

tions. SIAM Review, 37(4): 643, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:NMa

[Ano95j]  Anonymous.  New mem-

bers. SIAM Review, 37(1): 142–148, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:NMb

[Ano95k]  Anonymous.  New mem-

bers. SIAM Review, 37(2): 288–296, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:NMc

[Ano95l]  Anonymous.  New mem-

bers. SIAM Review, 37(3): 486–489, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:NMd

[Ano95m]  Anonymous.  New mem-

bers. SIAM Review, 37(4): 644–645, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:PSa

[Ano95n] Anonymous. Problems and solutions. SIAM Review, 37 (1):98–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:PSb

[Ano95o] Anonymous. Problems and solutions. SIAM Review, 37 (3):439–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:PSc

[Ano95p] Anonymous. Problems and solutions. SIAM Review, 37 (4):608–??, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:RTY

[Ano95q]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1994. SIAM Review, 37(4):646– 647, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:SCa

[Ano95r] Anonymous. Selected collections. SIAM Review, 37(1): 139–141, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:SCb

[Ano95s] Anonymous. Selected collections. SIAM Review, 37(2): 282–286, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:SCc

[Ano95t] Anonymous. Selected collections. SIAM Review, 37 (3):481–484, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1995:SCd

[Ano95u] Anonymous. Selected collections. SIAM Review, 37 (4):641–643, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:CPS

[Ano96a] Anonymous. Changes planned for SIAM Review. SIAM Review, 38(4):vii–viii, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:Ca

[Ano96b]  Anonymous.  Chronicle.

SIAM Review, 38(1):185– ??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:Cb

[Ano96c]  Anonymous.  Chronicle.

SIAM Review, 38(2):359– ??, June 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:Cc

[Ano96d]  Anonymous.  Chronicle.

SIAM Review, 38(3):547– ??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:Cd

[Ano96e]  Anonymous.  Chronicle.

SIAM Review, 38(4):716– ??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:EAO

[Ano96f] Anonymous. Elected and appointed officials. SIAM Review, 38(1):189, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:E

[Ano96g] Anonymous. Errata. SIAM Review, 38(1):184, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:LEa

[Ano96h]  Anonymous.  Later edi-

tions. SIAM Review, 38(2):

357–358, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:LEb

[Ano96i]  Anonymous.  Later edi-

tions. SIAM Review, 38(3): 546, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:NMa

[Ano96j]  Anonymous.  New mem-

bers. SIAM Review, 38(1): 185–188, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:NMb

[Ano96k]  Anonymous.  New mem-

bers. SIAM Review, 38(2): 359–365, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:NMc

[Ano96l]  Anonymous.  New mem-

bers. SIAM Review, 38(3): 547–551, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:NMd

[Ano96m]  Anonymous.  New mem-

bers. SIAM Review, 38(4): 716–719, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:PSa

[Ano96n] Anonymous. Problems and solutions. SIAM Review, 38 (1):146–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:PSb

[Ano96o] Anonymous. Problems and solutions. SIAM Review, 38 (2):313–??, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:PSc

[Ano96p] Anonymous. Problems and solutions. SIAM Review, 38 (3):513–??, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:PSd

[Ano96q] Anonymous. Problems and solutions. SIAM Review, 38 (4):668–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:RTY

[Ano96r]  Anonymous.  Report of

the Treasurer for the year ended December 31, 1995. SIAM Review, 38(4):720– 721, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:SCa

[Ano96s] Anonymous. Selected collections. SIAM Review, 38(1): 182–184, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:SCb

[Ano96t] Anonymous. Selected collections. SIAM Review, 38(2): 355–357, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:SCc

[Ano96u] Anonymous. Selected collections. SIAM Review, 38 (3):544–546, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1996:SCd

[Ano96v] Anonymous. Selected collections. SIAM Review, 38 (4):712–715, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:BR

[Ano97a] Anonymous. Book reviews.

SIAM Review, 39(1):142– 178, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97001.

Anonymous:1997:CPS

[Ano97b] Anonymous. Changes planned for SIAM Review. SIAM Review, 39(1):vii–viii, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:Ca

[Ano97c]  Anonymous.  Chronicle.

SIAM Review, 39(1):179– 185, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:Cb

[Ano97d]  Anonymous.  Chronicle.

SIAM Review, 39(2):375– 381, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:Cc

[Ano97e]  Anonymous.  Chronicle.

SIAM Review, 39(4):810– ??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:EP

[Ano97f] Anonymous. An extremal problem. SIAM Review, 39 (2):319–324, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:LEa

[Ano97g]  Anonymous.  Later edi-

tions. SIAM Review, 39(1): 177–178, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:LEb

[Ano97h]  Anonymous.  Later edi-

tions. SIAM Review, 39(2): 373–374, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:LEc

[Ano97i]  Anonymous.  Later edi-

tions. SIAM Review, 39(4): 808–??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:PSD

[Ano97j] Anonymous. Periodic solutions of a differential equation. SIAM Review, 39(2): 324–325, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:PS

[Ano97k] Anonymous. Problems and solutions. SIAM Review, 39 (4):761–??, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:SCa

[Ano97l] Anonymous. Selected collections. SIAM Review, 39(1): 174–176, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:SCb

[Ano97m] Anonymous. Selected collections. SIAM Review, 39(2): 369–373, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1997:SCc

[Ano97n] Anonymous. Selected collections. SIAM Review, 39(4): 805–??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1998:BR

[Ano98a] Anonymous. Book reviews. SIAM Review, 40(4):998– 1023, December 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97016.

Anonymous:1998:C

[Ano98b]  Anonymous.  Chronicle.

SIAM Review, 40(1):182– ??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1998:LE

[Ano98c]  Anonymous.  Later edi-

tions. SIAM Review, 40(1): 181–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1998:PS

[Ano98d] Anonymous. Problems and solutions. SIAM Review, 40(1):118–145, March 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1998:SC

[Ano98e] Anonymous. Selected collections. SIAM Review, 40(1): 178–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1998:S

[Ano98f]  Anonymous.  Solutions.

SIAM Review, 40(4):981– 997, December 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97015.

Anonymous:1999:BR

[Ano99a] Anonymous. Book reviews.

SIAM Review, 41(4):809– 851, December 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97040.

Anonymous:1999:GIS

[Ano99b] Anonymous. Graphical illustration of some examples related to the article “iterated random functions” by Diaconis and Freedman. SIAM Review, 41(1):77–82, ???? 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1999:Sa

[Ano99c]  Anonymous.  SIGEST.

SIAM Review, 41(1):113, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97021.

Anonymous:1999:Sb

[Ano99d] Anonymous. SIGEST. SIAM Review, 41(3):553, ???? 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:1999:SIa

[Ano99e] Anonymous. SIGEST — introduction. SIAM Review, 41(2):301, June 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97026.

Anonymous:1999:SIb

[Ano99f] Anonymous. SIGEST — introduction. SIAM Review, 41(4):775, ???? 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2000:Sa

[Ano00a] Anonymous. SIGEST. SIAM Review, 42(1):91, ???? 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2000:Sb

[Ano00b] Anonymous. SIGEST. SIAM Review, 42(2):265, ???? 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2000:Sc

[Ano00c] Anonymous. SIGEST. SIAM Review, 42(3):449, ???? 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2000:SI

[Ano00d] Anonymous. SIGEST — introduction. SIAM Review, 42(4):693, ???? 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2001:Sa

[Ano01a]  Anonymous.  SIGEST.

SIAM Review, 43(1):127, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97063.

Anonymous:2001:Sb

[Ano01b] Anonymous. SIGEST. SIAM Review, 43(2):313, ???? 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2001:Sc

[Ano01c]  Anonymous.  SIGEST.

SIAM Review, 43(3):497, September 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97073.

Anonymous:2001:Sd

[Ano01d] Anonymous. SIGEST. SIAM Review, 43(4):643, ???? 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2002:Sa

[Ano02a] Anonymous. SIGEST. SIAM Review, 44(1):93, ???? 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2002:Sb

[Ano02b] Anonymous. SIGEST. SIAM Review, 44(2):235, ???? 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2002:Sc

[Ano02c]  Anonymous.  SIGEST.

SIAM Review, 44(3):415, September 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97093.

Anonymous:2002:Sd

[Ano02d] Anonymous. SIGEST. SIAM Review, 44(4):629, ???? 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2003:Sa

[Ano03a] Anonymous. SIGEST. SIAM Review, 45(1):73, ???? 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2003:Sb

[Ano03b] Anonymous. SIGEST. SIAM Review, 45(2):289, ???? 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2003:Sc

[Ano03c] Anonymous. SIGEST. SIAM Review, 45(3):521, ???? 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2003:Sd

[Ano03d] Anonymous. SIGEST. SIAM Review, 45(4):725, ???? 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2004:Sa

[Ano04a]  Anonymous.  SIGEST.

SIAM Review, 46(1):85, March 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97123.

Anonymous:2004:Sb

[Ano04b] Anonymous. SIGEST. SIAM Review, 46(2):283, ???? 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2004:Sc

[Ano04c] Anonymous. SIGEST. SIAM Review, 46(3):475, ???? 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2004:Sd

[Ano04d] Anonymous. SIGEST. SIAM Review, 46(4):691, ???? 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2005:Sa

[Ano05a] Anonymous. SIGEST. SIAM Review, 47(1):97, ???? 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2005:Sb

[Ano05b] Anonymous. SIGEST. SIAM Review, 47(2):315, ???? 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2005:Sc

[Ano05c] Anonymous. SIGEST. SIAM Review, 47(4):746, ???? 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2005:SI

[Ano05d] Anonymous. SIGEST — introduction. SIAM Review, 47(3):517, ???? 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2006:S

[Ano06] Anonymous. SIGEST. SIAM Review, 48(4):743, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2007:I

[Ano07a]  Anonymous.  Introduc-

tion. SIAM Review, 49 (4):649, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2007:Sa

[Ano07b] Anonymous. SIGEST. SIAM Review, 49(1):89, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2007:Sb

[Ano07c] Anonymous. SIGEST. SIAM Review, 49(2):275, ???? 2007.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2008:SIa

[Ano08a] Anonymous. Sigest — introduction. SIAM Review, 50 (2):323, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2008:SIb

[Ano08b] Anonymous. Sigest — introduction. SIAM Review, 50 (3):521, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2009:SI

[Ano09] Anonymous. SIGEST — introduction. SIAM Review, 51(4):745, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2010:S

[Ano10] Anonymous. SIGEST. SIAM Review, 52(4):715, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2011:S

[Ano11] Anonymous. SIGEST. SIAM Review, 53(3):545, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2012:Sa

[Ano12a] Anonymous. SIGEST. SIAM Review, 54(1):139, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2012:Sb

[Ano12b] Anonymous. SIGEST. SIAM Review, 54(4):755, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anonymous:2018:S

[Ano18] Anonymous. SIGEST. SIAM Review, 60(2):407, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anselone:1964:BRI

[Ans64a] P. M. Anselone. Book review:

An Introduction to Transport Theory (G. M. Wing). SIAM Review, 6(4):467, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anselone:1964:BRMa

[Ans64b] P. M. Anselone. Book review: Mathematical Methods in Physics and Engineering (J. W. Dettman). SIAM Review, 6(3):323, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anselone:1964:BRMb

[Ans64c] P. M. Anselone. Book review: Methods in Analysis (Jack Indritz). SIAM Review, 6(4):474, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anselone:1964:BRS

[Ans64d] P. M. Anselone. Book review: Spectral Theory (E. R. Lorch). SIAM Review, 6(3):322, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anselone:1966:BRP

[Ans66]  P. M. Anselone.  Book

review: Partial Differential Equations (G. Hellwig). SIAM Review, 8(2): 247–248, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anselone:1968:BRB

[Ans68]  P. M. Anselone.  Book

review: Boundary Value Problems of Mathematical Physics, Volume I (Ivar Stakgold). SIAM Review, 10(1): 92–93, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anselone:1980:BRT

[Ans80] P. M. Anselone. Book review: Theoretical Numerical Analysis (Peter Linz).

SIAM Review, 22(3):376– 377, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anselone:1985:BRS

[Ans85]  P. M. Anselone.  Book

review: Spectral Approximation of Linear Operators (Françoise Chatelin). SIAM Review, 27(3):459, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antosiewicz:1963:BRO

[Ant63] H. A. Antosiewicz. Book review: Ordinary Differential Equations (G. Birkhoff and G. C. Rota). SIAM Review, 5 (2):160–161, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1972:BRA

[Ant72] Stuart S. Antman. Book review: Angular Velocity and Moment Potential for a Right Body (James G. Simmonds). SIAM Review, 14 (4):649–652, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1978:BRB

[Ant78] Stuart S. Antman. Book review: Bifurcation Problems in Nonlinear Elasticity (R. W. Dickey). SIAM Review, 20(2):402, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1982:BRI

[Ant82] Stuart S. Antman. Book review: Introduction to Partial Differential Equations and Hilbert Space Methods (Karl E. Gustafson). SIAM Review, 24(2):237–238, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1983:ICM

[Ant83a]  S. S. Antman.  An in-

troduction to continuummechanics — Gurtin, M. E. SIAM Review, 25(3): 410–411, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1983:BRI

[Ant83b] Stuart S. Antman. Book review: An Introduction to Continuum Mechanics (Morton E. Gurtin). SIAM Review, 25(3):410–411, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1986:BRL

[Ant86a] Stuart S. Antman. Book review: Local Bifurcation and Symmetry (A. Vanderbauwhede). SIAM Review, 28 (1):93–94, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antsaklis:1986:BRO

[Ant86b] Panos J. Antsaklis. Book review: Observers for Linear Systems (J. O’Reilly). SIAM Review, 28(1):94– 97, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1987:BRW

[Ant87] Stuart S. Antman. Book review: Weighted Energy Methods in Fluid Dynamics and Elasticity (Giovanni P. Galdi and Salvatore Rionero). SIAM Review, 29(4): 639–641, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1990:BRD

[Ant90] Stuart S. Antman. Book review: Dynamics of Multibody Systems (R. E. Roberson and R. Schwertassek). SIAM Review, 32(3):486– 487, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antonelli:1991:BRP

[Ant91] P. L. Antonelli. Book review: Plant and Crop Modelling (John H. M. Thornley and Ian R. Johnson). SIAM Review, 33(3):503– 504, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1992:BRM

[Ant92a] Stuart S. Antman. Book review: Mechanics: From Newton’s Laws to Deterministic Chaos (Florian Scheck). SIAM Review, 34(1):135– 137, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1992:MNL

[Ant92b] Stuart S. Antman. Mechanics: From Newton’s laws to deterministic chaos (Florian Scheck). SIAM Review, 34 (1):135–??, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antonelli:1992:ABP

[Ant92c] Peter L. Antonelli. The algorithmic beauty of plants (Przemyslaw Prusinkiewicz and Aristid Lindenmayer). SIAM Review, 34(1):142– ??, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antonelli:1992:BRA

[Ant92d] Peter L. Antonelli. Book review: The Algorithmic Beauty of Plants (Przemyslaw Prusinkiewicz and Aristid Lindenmayer). SIAM Review, 34(1):142–143, ????

1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1998:SPS

[Ant98] Stuart S. Antman. The simple pendulum is not so simple. SIAM Review, 40 (4):927–930, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32376.

Aoki:1970:BRS

[Aok70] Masanao Aoki. Book review:

Statistical Decision Theory in Adaptive Control Systems (Yoshikazu Sawaragi, Yoshifumi Sunahara and Takayoshi Nakamizo). SIAM Review, 12 (1):159–160, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aplevich:1995:BRF

[Apl95] J. D. Aplevich. Book review:

First-Order Representations of Linear Systems (Margreet Kuijper). SIAM Review, 37 (4):621–622, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Appledorn:1987:EPD

[App87] C. Robert Appledorn. The entropy of a Poisson distribution. SIAM Review, 29 (2):297, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Armitage:1963:MDT

[AR63] J. V. Armitage and R. A. Rapp. A mathematical derivation of a temperaturetime program for the passivation of Silver alloys. SIAM Review, 5(1):67–72, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1978:GBB

[AR78] Stuart S. Antman and Gerald Rosenfeld. Global behavior of buckled states of nonlinearly elastic rods. SIAM Review, 20(3):513–566, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See corrections and additions [AR80].

Antman:1980:CAG

[AR80] Stuart S. Antman and Gerald Rosenfeld. Corrections and additions: “Global behavior of buckled states of nonlinearly elastic rods” [SIAM rev. 20 (1978), no. 3, 513– 566 and MR 58 #14107]. SIAM Review, 22(2):186– 187, April 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [AR78].

Ascher:1981:RBV

[AR81] U. Ascher and R. D. Russell. Reformulation of boundary value problems into “standard” form. SIAM Review, 23 (2):238–254, April 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Antman:1995:NPE

[AR95] Stuart S. Antman and Michael Renardy. Nonlinear problems of elasticity. SIAM Review, 37(4):637– ??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arbib:1976:BRG

[Arb76] Michale A. Arbib. Book review: General Systems Theory: Mathematical Foundations (M. D. Mesarović and Y. Takahara). SIAM Review, 18(3):527–528, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arbib:1977:BRA

[Arb77] Michael A. Arbib. Book review: Applications of Fuzzy Sets to Systems Analysis (C. V. Negoita and R. A. Ralescu). SIAM Review, 19 (4):753, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arbib:1978:BRI

[Arb78] Michael A. Arbib. Book review: Introduction to the Theory of Fuzzy Subsets, Vol. 1 (A. Kaufmann). SIAM Review, 20(2):402– 403, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Archer:1975:CTD

[Arc75] D. A. Archer. Collocation in two dimensions — some numerical experiments. SIAM Review, 17(2):374, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aref:1986:BRC

[Are86]  Hassan Aref.  Book re-

view: Chaos, Fractals, and Dynamics (P. Fischer and William R. Smith, eds.). SIAM Review, 28(3):406– 407, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aref:1989:BRC

[Are89] Hassan Aref. Book review: Chaotic Vibrations: an Introduction for Applied Scientists and Engineers (Francis C. Moon). SIAM Review, 31 (2):330–331, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aref:1991:BRL

[Are91a] Hassan Aref. Book review: Lindenmayer Systems, Fractals, and Plants (Przemyslaw Prusinkiewicz and James Hanan). SIAM Review, 33 (2):284–285, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aref:1991:LSF

[Are91b] Hassan Aref. Lindenmayer systems, fractals, and plants (Przemyslaw Prusinkiewicz and James Hanan). SIAM Review, 33(2):284–??, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aref:1993:RTC

[Are93] Hassan Aref. Review: Transition to Chaos: The Orbit Diagram and the Mandelbrot Set, with Robert L. Devaney. SIAM Review, 35 (2):308–309, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/siread/v35/
i2/p308/s1;   http://www.
jstor.org/stable/2133149.

Ariyawansa:1991:APA

[Ari91a] K. A. Ariyawansa. ABS projection algorithms: Mathematical techniques for linear and nonlinear equations (Jozsef Abaffy and Emilio Spedicato). SIAM Review, 33 (2):285–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ariyawansa:1991:BRA

[Ari91b] K. A. Ariyawansa. Book review: ABS Projection Algorithms: Mathematical Techniques for Linear and Nonlinear Equations (Jozsef Abaffy and Emilio Spedicato). SIAM Review, 33(2): 285–286, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arino:1992:SSP

[Ari92] O. Arino. Some spectral properties for the asymptotic behavior of semigroups connected to population dynamics. SIAM Review, 34(3):445– 476, September 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arino:1996:BRO

[Ari96] Ovide Arino. Book review: Oscillation Theory for Functional Differential Equations (L. H. Erbe, Q. K. Kong, and B. G. Zhang). SIAM Review, 38(4):680–681, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arkin:1966:CP

[Ark66] Joseph Arkin. A combinatorial problem. SIAM Review, 8(1):107, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Armstrong:1976:DPO

[Arm76] E. S. Armstrong. Digital program for optimal regulator control of linear systems. SIAM Review, 18(4): 793–794, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Armstrong:1990:BRT

[Arm90a] Thomas E. Armstrong. Book review: A Theory of Heuristic Information in Game-Tree Search (Chun-Hung Tzeng). SIAM Review, 32 (1):155–156, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Armstrong:1990:THI

[Arm90b] Thomas E. Armstrong. A theory of heuristic information in game-tree research (Chun-Hung Tzeng). SIAM Review, 32(1):155– ??, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arms:1994:BRL

[Arm94]  Judith M. Arms.  Book

review: Lectures on Mechanics (J. E. Marsden). SIAM Review, 36(2):290– 291, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arnold:1970:PAD

[Arn70] Leslie K. Arnold. A problem in acoustic detection. SIAM Review, 12(1):143– 144, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arney:1989:BRN

[Arn89] David C. Arney. Book review: Numerical Solution of the Shallow-Water Equations (F. W. Wubs). SIAM Review, 31(4):687–688, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arney:1992:BRP

[Arn92]  David C. Arney.  Book

review: Practical Time-Stepping Schemes (W. L. Wood). SIAM Review, 34(4): 663, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aroian:1963:BRR

[Aro63] Leo A. Aroian. Book review: Reliability: Management, Methods and Mathematics (D. K. Lloyd and M. Lipow). SIAM Review, 5(2): 174–175, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aronson:1969:BRT

[Aro69] D. G. Aronson. Book review: The Theory of Difference Schemes. An Introduction (S. K. Godunov and V. S. Ryabenki). SIAM Review, 11(1):92–93, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aronson:1978:CMS

[Aro78a] D. G. Aronson. A comparison method for stability analysis of nonlinear parabolic problems. SIAM Review, 20(2): 245–264, April 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arora:1978:AFS

[Aro78b] S. K. Arora. Algorithm to find semi-linear sets representing context free languages. SIAM Review, 20 (3):619, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arsove:1960:CPL

[Ars60] Maynard G. Arsove. Continuous potentials and linear mass distributions. SIAM Review, 2(3):177–184, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arscott:1962:BRI

[Ars62a] F. M. Arscott. Book review:

Introduction to Elliptic Functions with Applications (F. Bowman). SIAM Review, 4 (4):408, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arscott:1962:BRT

[Ars62b] F. M. Arscott. Book review: The Theory of Determinants, Matrices and Invariants (H. W. Turnbull). SIAM Review, 4(3):271, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arsham:1997:CNI

[Ars97] H. Arsham. Classroom note: Initialization of the simplex algorithm: an artificial-free approach. SIAM Review, 39 (4):736–744, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30472. See er-

ratum and comments [Mac98a].

Artstein:1980:DCB

[Art80] Zvi Artstein. Discrete and continuous bang-bang and facial spaces or: Look for the extreme points. SIAM Review, 22(2):172–185, April 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arthur:1983:SMO

[Art83a] J. L. Arthur. Stochastic models in operations research, vol 1 — Heyman, D. P., Sobel, M. J. SIAM Review, 25(3):

440–441, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arthur:1983:BRS

[Art83b] Jeffrey L. Arthur. Book review: Stochastic Models in Operations Research, Volume I. (Daniel P. Heyman and Matthew J. Sobel). SIAM Review, 25(3): 440–441, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arthur:1985:BRI

[Art85a] Jeffrey L. Arthur. Book review: An Introduction to Stochastic Modeling (Howard M. Taylor and Samuel Karlin). SIAM Review, 27(4): 581–582, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arthur:1985:BRS

[Art85b] Jeffrey L. Arthur. Book review: Stochastic Models in Operations Research, Volume II. Stochastic Optimization (Daniel P. Heyman and Matthew J. Sobel). SIAM Review, 27(1): 87–88, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Artstein:1987:BRD

[Art87] Zvi Artstein. Book review: Dynamic Games and Applications in Economics. (T. Basar, ed.). SIAM Review, 29(4):641, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arthur:1988:BRM

[Art88] Jeffrey L. Arthur. Book review: Mathematical Topics in Telecommunications. Vol. 1: Optimisation Methods in Electronics and Communications (K. W. Cattermole and J. J. O’Reilly). SIAM Review, 30(1):135–136, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Artstein:1991:BRO

[Art91a]  Zvi Artstein.  Book re-

view: Optimization and Stability Theory for Economic Analysis (Brian Beavis and Ian Dobbs). SIAM Review, 33(3):502–503, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Artstein:1991:OST

[Art91b]  Zvi Artstein.  Optimiza-

tion and stability theory for economic analysis (Brian Beavis and Ian Dobbs). SIAM Review, 33(3):502– ??, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Artstein:1992:BRE

[Art92a]  Zvi Artstein.  Book re-

view: Existence and Optimality of Competitive Equilibria (Charalambos Aliprantis, Donald Brown, and Owen Burkinshaw). SIAM Review, 34(1):152–153, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Artstein:1992:EOC

[Art92b]  Zvi Artstein.  Existence

and optimality of competitive equilibria (Charalambos Aliprantis, Donald Brown, and Owen Burkinshaw). SIAM Review, 34(1): 152–??, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arthurs:1980:STS

[AS80] E. Arthurs and B. W.

Stuck. Subadditivity of teletraffic special functions. SIAM Review, 22(4):502– 503, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Al-Salam:1985:BRW

[AS85] W. A. Al-Salam. Book review: A Weyl Fractional Integral (H. M. Srivastava). SIAM Review, 27(3):448– 449, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Al-Salam:1986:BRI

[ASF86] W. A. Al-Salam and J. L. Fields. Book review: An Identity for Double Hypergeometric Series (H. M. Srivastava). SIAM Review, 28 (4):576–578, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ashley:1989:BRS

[Ash89]  Holt Ashley.  Book re-

view: Studies in Nonlinear Aeroelasticity (Earl H. Dowell and Marat Ilgamov). SIAM Review, 31(4):688– 689, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1969:PI

[Ask69] Richard Askey. A positive integral. SIAM Review, 11 (3):398, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1972:BRF

[Ask72] Richard Askey. Book review: Fourier Analysis and Approximation (Paul L. Butzer and Rolf J. Nessel). SIAM Review, 14(3): 507–508, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1973:SI

[Ask73a] R. A. Askey. A sine inequality. SIAM Review, 15 (4):788, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1973:II

[Ask73b] Richard Askey. An integral inequality. SIAM Review, 15 (2):377, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1974:OSA

[Ask74] R. Askey. One-sided approximation to special functions. SIAM Review, 16(4): 545–546, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1975:II

[Ask75a] R. Askey. Integral inequality. SIAM Review, 17(4): 691–693, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1975:BRT

[Ask75b] Richard Askey. Book review: Tables of Bessel Transforms (Fritz Oberhettinger); Fourier Expansions: a Collection of Formulas (Fritz Oberhettinger); Fourier Transforms of Distributions and Their Inverses — a Collection of Tables (Fritz Oberhettinger); Tables of Laplace Transforms (F. Oberhettinger and L. Badii). SIAM Review, 17(4): 707–712, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1976:OSA

[Ask76] R. Askey. One-sided approximation to special functions. SIAM Review, 18(1): 121–122, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1978:BRM

[Ask78] Richard Askey. Book review: Multiple Hypergeometric Functions and Applications (Harold Exton). SIAM Review, 20(4):874– 875, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1979:BRH

[Ask79a] Richard Askey. Book review: Handbook of Hypergeometric Integrals: Theory, Applications, Tables, Computer Programs (Harold Exton). SIAM Review, 21(4):578, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1979:BRS

[Ask79b] Richard Askey. Book review: Special Functions of Applied Mathematics (B. C. Carlson). SIAM Review, 21 (3):420–421, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1980:BRF

[Ask80] Richard Askey. Book review: The H-Function with Applications in Statistics and other Disciplines (A. M. Mathai and R. K. Saxena). SIAM Review, 22(3):379– 380, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1981:BRA

[Ask81] R. Askey. Book review:

Approximation with Rational Functions (D. J. Newman). SIAM Review, 23 (1):114, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1984:ECB

[Ask84]  Richard Askey.  A q-

extension of Cauchy’s beta integral. SIAM Review, 26 (2):270, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1985:BHE

[Ask85] R. Askey. Basic hypergeometric extensions of some integrals. SIAM Review, 27 (4):574, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1986:BHE

[Ask86a] R. Askey. Basic hypergeometric extensions of some integrals. SIAM Review, 28 (4):574–575, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1986:SSH

[Ask86b] Richard Askey. On the sign of a set of hypergeometric functions. SIAM Review, 28 (1):86, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1992:BRH

[Ask92a] Richard Askey. Book review: Hypergeometric Functions and Their Applications (James B. Seaborn). SIAM Review, 34(3):503, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Askey:1992:HFT

[Ask92b]  Richard Askey.  Hyper-

geometric functions and their applications (James B Seaborn). SIAM Review, 34 (3):503–??, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Asmussen:1998:PLW

[Asm98] Soren Asmussen. A probabilistic look at the Wiener– Hopf equation. SIAM Review, 40(2):189–201, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/30353.

Asner:1975:BRF

[Asn75] Bernard A. Asner. Book review: Frequency Domain Criteria for Absolute Stability (Kumpati S. Narendra and James H. Taylor). SIAM Review, 17(1):177– 178, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Asner:1978:BRO

[Asn78] B. Asner. Book review:

Ordinary and Delay Differential Equations (R. D. Driver). SIAM Review, 20 (3):608–609, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aspvall:1984:MAA

[Asp84a] B. Aspvall. Mathematics for the analysis of algorithms — Greene, D. H., Knuth, D. E. SIAM Review, 26(1):

142–143, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aspvall:1984:BRM

[Asp84b] Bengt Aspvall. Book review: Mathematics for the Analysis of Algorithms (Daniel H. Greens and Donald E. Knuth). SIAM Review, 26 (1):142–143, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Altay-Salih:2003:CNP

[ASPL03] Aslihan Altay-Salih, Mustafa C. Pinar, and Sven Leyffer. Constrained nonlinear programming for volatility estimation with GARCH models. SIAM Review, 45(3): 485–503, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40011.

Assmus:1977:BRM

[Ass77] E. F. Assmus, Jr. Book review: The Mathematical Theory of Coding (Ian F. Blake and Ronald C. Mullin). SIAM Review, 19(1):175– 176, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Astor:1970:SLE

[Ast70] Peter H. Astor. A set of linear equations. SIAM Review, 12(1):143, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Astfalk:1987:BRI

[Ast87]  Greg Astfalk.  Book re-

view: An Introduction to the Numerical Solution of Differential Equations (Douglas Quinney). SIAM Review, 29 (4):642–643, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Anderson:1976:MI

[AT76] W. Anderson, Jr. and G. Trapp. A matrix inequality. SIAM Review, 18 (2):295, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aurentz:2017:BOS

[AT17] Jared L. Aurentz and Lloyd N.

Trefethen. Block operators and spectral discretizations. SIAM Review, 59(2): 423–446, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Atherton:1976:TGT

[Ath76] R. W. Atherton. Toward a general theory of model sensitivity. SIAM Review, 18 (4):794, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Atherton:1987:BRR

[Ath87] Derek P. Atherton. Book review: Relay Control Systems (Ya. Z. Tsypkin). SIAM Review, 29(4):643– 647, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Athreya:1993:CTD

[Ath93a] Krisha B. Athreya. The construction theory of denumerable Markov processes (Xiang-qun Yang). SIAM Review, 35(3):510–??, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Athreya:1993:BRC

[Ath93b] Krishna B. Athreya. Book review: The Construction Theory of Denumerable Markov Processes (Xiang-Qun Yang). SIAM Review, 35(3):510– 513, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Athanasopoulos:1994:BRM

[Ath94] Bessy D. Athanasopoulos. Book review: Moments in Probability and Approximation Theory (George A. Anastassiou). SIAM Review, 36 (3):501–502, September 1994.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Atkin:1966:BRI

[Atk66] A. O. L. Atkin. Book review:

An Infinite Permutation (M. S. Klamkin). SIAM Review, 8 (2):234–236, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Auchmuty:1997:IGT

[Auc97a] G. Auchmuty. An inequality in gravitational theory. SIAM Review, 39(1):126– 135, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Auchmuty:1997:SIG

[Auc97b] Giles Auchmuty. Solutions:

An inequality in gravitational theory (problem 95-18) (Brian Sutin). SIAM Review, 39(1):126–135, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aurentz:2018:BRE

[Aur18] Jared L. Aurentz. Book review: Exploiting Hidden Structures in Matrix Computations: Algorithms and Applications, edited by Michele Benzi and Valeria Simoncini. SIAM Review, 60(2):470– 471, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arorat:1965:MRB

[AV65] Kasturi L. Arorat and Nguyen X. Vinh. Maximum range of ballistic missiles. SIAM Review, 7(4): 544–550, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Aharonov:2008:AQC

[AvK^{+}08] Dorit Aharonov, Wim van Dam, Julia Kempe, Zeph Landau, Seth Lloyd, and Oded Regev. Adiabatic quantum computation is equivalent to standard quantum computation. SIAM Review, 50(4):755–787, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Allen:1976:AAC

[AW76] R. C. Allen and G. M. Wing. Algorithm for accelerating convergence of Neumann series associated with certain integral equations (preliminary report). SIAM Review, 18(4):792–793, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Alm:2002:TFA

[AW02] Jeremy F. Alm and James S. Walker. Time-frequency analysis of musical instruments. SIAM Review, 44 (3):457–476, September 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38228.

Awoniyi:2019:CFC

[AW19] Samuel Awoniyi and Ira Wheaton. Case for first courses on finite Markov chain modeling to include sojourn time cycle chart. SIAM Review, 61(2):347– 360, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Axford:1965:BRS

[Axf65] Roy A. Axford. Book review: Slow Viscous Flow (W. E. Langlois). SIAM Review, 7 (1):149–150, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ying:1996:IEM

[aYT96] Lung an Ying and R. W.

Thatcher. Infinite element methods. SIAM Review, 38 (4):703–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bacon:1960:NCO

[Bac60] Ralph Hoyt Bacon. A note concerning orthogonal polynomials. SIAM Review, 2(4): 269–276, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bade:1960:MTP

[Bad60a] W. L. Bade. A minimum time path. SIAM Review, 2 (4):298, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bade:1960:SMS

[Bad60b] W. L. Bade. Some multiple summations. SIAM Review, 2(3):218, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bade:1961:DI

[Bad61] W. L. Bade. A definite integral. SIAM Review, 3 (4):329, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:1979:IEF

[BAD^{+}79] A. V. Boyd, R. Askey, D. R. Dellwo, H. E. Fettis, M. L. Glasser, O. P. Lossers, and P. T. L. M. Vanwoerkom. Integral of an elliptic function. SIAM Review, 21(3): 396–397, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bagnara:1995:UPC

[Bag95] Roberto Bagnara. A unified proof for the convergence of Jacobi and Gauss–Seidel methods. SIAM Review, 37 (1):93–97, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25266.htm.

Brinkspalink:1980:CF

[BAH^{+}80] J. Brinkspalink, J. Aczel, M. D. Humphries, G. E. Latta, S. L. Paverifontana, and G. C. Taylor. Credibility functions. SIAM Review, 22 (2):234–236, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bailey:1974:BRI

[Bai74a] P. B. Bailey. Book review: Initial Value Methods for Boundary Value Problems (Gunter H. Meyer). SIAM Review, 16(4):566– 568, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bailey:1974:BRT

[Bai74b] Paul B. Bailey. Book review: Two-point Boundary Value Problems. Shooting Methods (Sanford M. Roberts and Jerome S. Shipman). SIAM Review, 16(2): 265–266, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bailey:1976:BRI

[Bai76] Paul B. Bailey. Book review: Integral Equations via Imbedding Methods (Harriet H. Kagiwada and Robert Kalaba). SIAM Review, 18 (3):532–533, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bai:1987:BRM

[Bai87] Dov Bai. Book review: Multigrid Methods for Integral and Differential Equations (D. J. Paddon and H. Holstein, eds.). SIAM Review, 29(4): 647–648, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bailey:1993:BRC

[Bai93a] David H. Bailey. Book review: Computational Frameworks for the Fast Fourier Transform (Charles Van Loan). SIAM Review, 35(1): 142–143, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bailey:1993:CFF

[Bai93b]  David H. Bailey.  Com-

putational frameworks for the Fast Fourier Tranform (Charles Van Loan). SIAM Review, 35(1):142–??, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Baker:1966:BRI

[Bak66] Christopher T. H. Baker.

Book review: An Introduction to Numerical Linear Algebra with Exercises (L. Fox). SIAM Review, 8(1): 125–127, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Baker:1991:BRF

[Bak91a]  A. J. Baker.  Book re-

view: Finite Element Methods for Viscous Incompressible Flows: a Guide to Theory, Practice, and Algorithms (Max D. Gunzburger). SIAM Review, 33(1):129– 131, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Baker:1991:BRM

[Bak91b] George A. Baker, Jr. Book review: Mathematical Foundations of Classical Statistical Mechanics: Continuous Systems (D. Ya. Petrina, V. I. Gerasimenko, and P. V. Malyshev; trans. by P. V. Malyshev and D. V. Malyshev). SIAM Review, 33(3): 510, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Baker:1992:BRM

[Bak92] George A. Baker, Jr. Book review: Mathematical Foundations of Classical Statistical Mechanics: Continuous Systems (D. Ya. Petrina, V. I. Gerasimenko, and P. V. Malyshev). SIAM Review, 34 (2):330–331, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balinski:1970:PM

[Bal70] M. L. Balinski. On perfect matchings. SIAM Review, 12 (4):570–572, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balas:1977:C

[Bal77] E. Balas. Correction. SIAM Review, 19(1):U1, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balas:1978:FCL

[Bal78] M. J. Balas. Feedback control of linear diffusion processes. SIAM Review, 20 (3):620, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balakotaiah:1985:MVE

[Bal85] V. Balakotaiah. A multivariable extremum problem. SIAM Review, 27(3): 445, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balder:1988:BRC

[Bal88a] Erik J. Balder. Book review: Control and Optimisation: The Linear Treatment of Nonlinear Problems (J. E. Rubio). SIAM Review, 30 (4):663, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ball:1988:BRQ

[Bal88b] J. M. Ball. Book review:

Qualitative Methods in Nonlinear Mechanics (J. Tinsley Oden). SIAM Review, 30(1): 136–137, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ball:1993:BRM

[Bal93] Joseph A. Ball. Book review: Minimum Entropy H_{∞} Control (D. Mustafa and K. Glover). SIAM Review, 35 (4):652–655, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ball:1995:BRN

[Bal95] Michael O. Ball. Book review: Network Flows (R. K. Ahuja, T. L. Magnanti and J. B. Orlin). SIAM Review, 37 (1):115–116, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balk:1997:NCL

[Bal97] A. M. Balk. New conservation laws for the interaction of nonlinear waves. SIAM Review, 39(1):68– 94, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29197;   http:
//www.siam.org/journals/
sirev/sirev391.htm.

Balasuriya:2020:SSC

[Bal20] Sanjeeva Balasuriya. Stochastic sensitivity: a computable Lagrangian uncertainty measure for unsteady flows. SIAM Review, 62(4):781– 816, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Banchoff:1984:CSD

[Ban84a]  T. F. Banchoff.  Catas-

trophe — Sinha, D. K. SIAM Review, 26(2):292– 293, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Banchoff:1984:BRC

[Ban84b] Thomas F. Banchoff. Book review: Catastrophe (D. K. Sinha). SIAM Review, 26(2): 292–293, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Banchoff:1985:BRI

[Ban85] Thomas F. Banchoff. Book review: Invitation to Geometry (Z. A. Melzak). SIAM Review, 27(4):582– 583, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Banerjee:1987:BRT

[Ban87] P. K. Banerjee. Book review: Topics in Boundary Element Research. Volume 2: Time-Dependent and Vibration Problems. (C. A. Brebbia, ed.). SIAM Review, 29(1):137–138, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bang:1988:EPA

[Ban88a] S. J. Bang. On the error in Padé approximation. SIAM Review, 30(2):319– 320, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bang:1988:BRE

[Ban88b] Seung-Jin Bang. Book review: On the Error in Padé Approximation (Edward J. Allen). SIAM Review, 30(2): 319–320, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bang:1988:SI

[Ban88c] Seung-Jin Bang. An n-space integral. SIAM Review, 30 (2):309, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barrer:1959:BRQ

[Bar59] D. Y. Barrer. Book review: Queues, Inventories and Maintenance (Philip M. Morse). SIAM Review, 1(2): 186–187, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bareiss:1964:BRI

[Bar64] E. H. Bareiss. Book review: Invariant Imbedding and Radiant Transfer in Slabs of Finite Thickness (R. E. Bellman, R. E. Kalaba and M. C. Prestrud). SIAM Review, 6(2):185, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bareiss:1965:BRI

[Bar65] E. H. Bareiss. Book review: Invariant Imbedding and Time-Dependent Transport Processes (R. E. Bellman, H. H. Kagiwada, R. E. Kalaba and M. C. Prestrud). SIAM Review, 7(1): 150–151, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bareiss:1967:BRR

[Bar67] Erwin H. Bareiss. Book review: Radiative Transfer on Discrete Spaces (Rudolph W. Preisendorfer). SIAM Review, 9(1):130–131, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bareiss:1970:BRN

[Bar70] Erwin Bareiss. Book review:

Nonlinear Ordinary Differential Equations in Transport Processes (William F. Ames). SIAM Review, 12(3): 476–477, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barrucand:1971:AE

[Bar71] P. Barrucand. An asymptotic expansion. SIAM Review, 13 (4):569, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barrucand:1973:IRS

[Bar73] P. Barrucand. Integral representation for a set of polynomials. SIAM Review, 15 (1):219, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnett:1974:NLC

[Bar74] Stephen Barnett. A new look at classical algorithms for polynomials resultant and G.C.D. calculation. SIAM Review, 16(2):193–206, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barrucand:1975:CI

[Bar75] P. Barrucand. A combinatorial identity. SIAM Review, 17(1):168, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barracand:1976:CI

[Bar76a] P. Barracand. Combinatorial identity. SIAM Review, 18 (2):303–304, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barrucand:1976:ZP

[Bar76b] P. Barrucand. On the zeros of a polynomial. SIAM Review, 18(4):763, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barrucand:1977:TI

[Bar77] P. Barrucand. Two inequalities. SIAM Review, 19 (4):737, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bardavid:1978:BMS

[Bar78] T. W. Bardavid. Bettis methods — solving differential equation by linear multistep methods that know physics. SIAM Review, 20 (3):620, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnhill:1979:BRF

[Bar79] Robert E. Barnhill. Book review: The Finite Element Method for Elliptic Problems (Phillippe G. Ciarlet). SIAM Review, 21(3): 410–411, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnett:1983:MPG

[Bar83a] S. Barnett. Matrix polynomials — Gohberg, I., Lancaster, P., Rodman, L. SIAM Review, 25(4):591– 592, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnett:1983:BRM

[Bar83b] Steve Barnett. Book review:

Matrix Polynomials (I. Gohberg, P. Lancaster and L. Rodman). SIAM Review, 25 (4):591–592, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barrucand:1985:AP

[Bar85] Pierre Barrucand. An asymptotic problem. SIAM Review, 27(1):80, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barcilon:1986:BRI

[Bar86] Victor Barcilon. Book review: Inverse Problems (D. W. McLaughlin). SIAM Review, 28(1):97–98, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnard:1989:BRH

[Bar89] George A. Barnard. Book review: The History of Statistics: The Measurement of Uncertainty Before 1900 (Stephen M. Stigler). SIAM Review, 31(3):499– 502, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnard:1991:BRH

[Bar91a] George A. Barnard. Book review: A History of Probability and Statistics: And Their Applications before 1750 (Anders Hald). SIAM Review, 33 (2):286–290, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnard:1991:HPS

[Bar91b]  George A. Barnard.  A

history of probability and statistics: And their applications before 1750 (Anders Hald). SIAM Review, 33 (2):286–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barkley:1994:BRC

[Bar94]  Dwight Barkley.  Book

review: Chemical Chaos (Stephen K. Scott). SIAM Review, 36(1):116, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bardos:1995:BRM

[Bar95] Claude Bardos. Book review:

The Mathematical Theory of Dilute Gases (C. Cercignani, R. Illner, and M. Pulvierenti). SIAM Review, 37(4): 622–624, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnett:2022:HEI

[Bar22]  Alex H. Barnett.  How

exponentially ill-conditioned are contiguous submatrices of the Fourier matrix? SIAM Review, 64(1):105– 131, ???? 2022. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/20M1336837.

Bassanini:1986:ITF

[Bas86] Piero Bassanini. Interference in thin films. SIAM Review, 28(3):381–384, September 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Basar:1988:BRS

[Bas88] Tamer Basar. Book review: Social Choice and Multicriterion Decision-Making (Kenneth J. Arrow and Herve Raynaud). SIAM Review, 30 (1):137–138, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Battista:2021:FSI

[Bat21] Nicholas A. Battista. Fluidstructure interaction for the classroom: Interpolation, hearts, and swimming! SIAM Review, 63(1):181– 207, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bauer:1973:SSE

[Bau73]  F. L. Bauer.  Software

and software engineering. SIAM Review, 15(2):469– 480, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Baumeister:1988:BRS

[Bau88] Johann Baumeister. Book review: Stability of Adaptive Systems: Passivity and Averaging Analysis (B. D. O. Anderson, R. R. Bitmead, C. R. Johnson, Jr., P. Y. Kokotovic, R. L. Kosut, I. M. V. Mareels, L. Praly and B. D. Riedle). SIAM Review, 30(2): 323–324, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Baumeister:1989:BRO

[Bau89] Johann Baumeister. Book review: Optimal Control (Frank L. Lewis). SIAM Review, 31(2):331–332, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Baumeister:1991:BRA

[Bau91] Johann Baumeister. Book review: Adaptive Control and Identification: Conflict or Conflux? (J. W. Polderman). SIAM Review, 33(2): 290–291, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Baxter:1985:BRC

[Bax85]  John Baxter.  Book re-

view: Classical Potential Theory and Its Probabilistic Counterpart (J. L. Doob). SIAM Review, 27(3):460– 462, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bazley:1973:BRM

[Baz73] Norman W. Bazley. Book review: Methods of Intermediate Problems for Eigenvalues Theory and Ramifications (Alexander Weinstein and William Stenger).

SIAM Review, 15(4):805– 806, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barkauskas:1979:D

[BB79] A. E. Barkauskas and D. W. Bange. A determinant. SIAM Review, 21(1):139, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1983:GAG

[BB83a] D. Borwein and P. B. Borwein. A generalized arithmetic-geometric mean. SIAM Review, 25(3):401, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bube:1983:ODI

[BB83b] Kenneth P. Bube and Robert Burridge. The one-dimensional inverse problem of reflection seismology. SIAM Review, 25(4):497–559, October 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1984:GAG

[BB84a] D. Borwein and P. B. Borwein. A generalized arithmetic-geometric mean. SIAM Review, 26(3):433, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1984:AGM

[BB84b] J. M. Borwein and P. B.

Borwein. The arithmeticgeometric mean and fast computation of elementary functions. SIAM Review, 26 (3):351–366, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1988:CFF

[BB88] J. M. Borwein and P. B.

Borwein. On the complexity of familiar functions and numbers. SIAM Review, 30 (4):589–601, December 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bauschke:1996:PAS

[BB96] Heinz H. Bauschke and Jonathan M. Borwein. On projection algorithms for solving convex feasibility problems. SIAM Review, 38 (3):367–426, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25171.

Bertsimas:2011:TAR

[BBC11] Dimitris Bertsimas, David B.

Brown, and Constantine Caramanis. Theory and applications of robust optimization. SIAM Review, 53(3): 464–501, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i3/p464_s1.

Bossolini:2020:SOC

[BBK20] Elena Bossolini, Morten Brøns, and Kristian Uldall Kristiansen. A stiction oscillator with canards: On piecewise smooth nonuniqueness and its resolution by regularization using geometric singular perturbation theory. SIAM Review, 62(4): 869–897, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brunton:2022:MKT

[BBKK22] Steven L. Brunton, Marko Budisić, Eurika Kaiser, and J. Nathan Kutz. Modern Koopman theory for dynamical systems. SIAM Review, 64(2):229–340, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1401243.

Bobylev:1996:APN

[BBKV96] Nikolai A. Bobylev, Yuri M. Burman, Sergey K. Korovin, and Gennadi Vainikko. Approximation procedures in nonlinear oscillation theory. SIAM Review, 38(1): 181–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Broersma:1992:BRA

[BBLS92] H. J. Broersma, H. J.

Broersma, Andy Liu, and L. V. Saxton. Book review: An Affirmative Action Graph (D. J. Newman). SIAM Review, 34(4): 652–653, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balakrishnan:1998:EDT

[BBN98] N. Balakrishnan, A. P. Basu, and H. N. Nagaraja. The exponential distribution: Theory, methods and applications. SIAM Review, 40(1): 167–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Benzi:2013:DPS

[BBR13] Michele Benzi, Paola Boito, and Nader Razouk. Decay properties of spectral projectors with applications to electronic structure. SIAM Review, 55(1):3–64, ???? 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bleistein:1976:POF

[BC76] N. Bleistein and J. K. Cohen. Physical optics far field inverse scattering in time domain. SIAM Review, 18 (4):796, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bottiger:1986:NC

[BC86] Jerold R. Bottiger and David K. Cohoon. A normalization constant. SIAM Review, 28(2):233, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1996:BRE

[BC96a] J. Boersma and Robin Chapman. Book review: Exciton Transport (M. L. Glasser). SIAM Review, 38(1):152– 153, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1996:BRE

[BC96b] Jonathan M. Borwein and Rob M. Corless. Book review: The Encyclopedia of Integer Sequences, by N. J. A. Sloane and Simon Plouffe. Academic Press, San Diego, CA, 1995. $44.95. xii + 587 pp., hardcover. ISBN 0-12-558630-2. SIAM Review, 38 (2):333–337, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barenblatt:1998:NPT

[BC98] G. I. Barenblatt and A. J. Chorin. New perspectives in turbulence: Scaling laws, asymptotics, and intermittency. SIAM Review, 40(2): 265–291, June 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32004.

Bonney:1998:WMG

[BCD98a] M. Bonney, J. Coaplen, and E. Doeff. What makes a good friend? the mathematics of rock climbing. SIAM Review, 40(3):674–679, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bonney:1998:CNW

[BCD98b] Matthew Bonney, Joshua Coaplen, and Erik Doeff. Classroom note: What makes a good friend? the mathematics of rock climbing. SIAM Review, 40(3):674– 679, September 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31739.

Bakonyi:1995:SAS

[BCF95] M. Bakonyi, T. Constantinescu, and A. E. Frazho. Schur’s algorithm and several applications. SIAM Review, 37(1):121–??, March 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bush:1977:MAD

[BCG77] A. W. Bush, M. Cross, and R. D. Gibson. The monitoring of airborne dust and grit. SIAM Review, 19(1): 140–145, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bienstock:2014:CCO

[BCH14] Daniel Bienstock, Michael Chertkov, and Sean Harnett. Chance-constrained optimal power flow: Risk-aware network control under uncertainty. SIAM Review, 56(3): 461–495, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buades:2010:IDM

[BCM10] A. Buades, B. Coll, and J. M. Morel. Image denoising methods. A new nonlocal principle. SIAM Review, 52 (1):113–147, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bottou:2018:OML

[BCN18] Léon Bottou, Frank E. Curtis, and Jorge Nocedal. Optimization methods for large-scale machine learning. SIAM Review, 60(2): 223–311, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blanchard:2011:CSH

[BCT11] Jeffrey D. Blanchard, Coralia Cartis, and Jared Tanner. Compressed sensing: How sharp is the restricted isometry property? SIAM Review, 53(1):105– 125, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i1/p105_s1.

Bujosa:1998:JNF

[BCV98] A. Bujosa, R. Criado, and C. Vega. Jordan normal form via elementary transformations. SIAM Review, 40 (4):947–956, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32934.

Brimacombe:2021:CAM

[BCZ21] Chris Brimacombe, Robert M.

Corless, and Mair Zamir. Computation and applications of Mathieu functions: a historical perspective. SIAM Review, 63(4): 653–720, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brooks:1963:PTR

[BD63] Foster Brooks and N. D. Diamantides. A probability theorem for random two-valued functions, with application to autocorrelations. SIAM Review, 5(1):33–40, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Block:1964:SPM

[BD64] I. E. Block and R. F. Drenick. A symposium on the publication of mathematical literature. SIAM Review, 6(4): 431–454, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Banks:1973:II

[BD73] W. H. H. Banks and P. G. Drazin. An integral identity. SIAM Review, 15(2): 376, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1992:FPI

[BD92] J. Boersma and P. J. Dedoelder. A finite part integral. SIAM Review, 34(4): 656–658, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1993:AMS

[BD93] J. Boersma and P. J. Dedoelder. On alternating multiple sums. SIAM Review, 35 (3):497–500, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Banks:1994:SHM

[BD94] John Banks and Valentina Dragan. Smale’s horseshoe map via ternary numbers. SIAM Review, 36(2): 265–271, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25625.htm.

Bellomo:2011:MTC

[BD11] Nicola Bellomo and Christian Dogbe. On the modeling of traffic and crowds: a survey of models, speculations, and perspectives. SIAM Review, 53(3):409– 463, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i3/p409_s1.

Brockwell:1998:ITS

[BDC98] Peter J. Brockwell, Richard A.

Davis, and Collin C. Carbno. Introduction to time series and forecasting. SIAM Review, 40(1):146–147, March 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borcea:2012:FDL

[BdCPT12] L. Borcea, F. González del Cueto, G. Papanicolaou, and C. Tsogka. Filtering deterministic layer effects in imaging. SIAM Review, 54(4): 757–798, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1992:BRF

[BdD92] J. Boersma and P. J. de Doelder. Book review: A Finite Part Integral (B. Bertram and O. G. Ruehr). SIAM Review, 34(4):656– 658, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1993:BRA

[BdD93] J. Boersma and P. J. de Doelder. Book review: On Alternating Multiple Sums (Malte Henkel and R. A. Weston). SIAM Review, 35(3): 497–500, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1994:BRT

[BdD94] J. Boersma and P. J. de Doelder. Book review: Two Integrals Arising from a Reaction-Diffusion Problem (John Billingham). SIAM Review, 36(3):498– 499, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bensoussan:1995:RCI

[BDD^{+}95] A. Bensoussan, G. Da Prato, M. Delfour, S. Mitter, and D. L. Russell. Representation and control of infinite dimensional systems, vols. 1 and 2. SIAM Review, 37(3): 476–??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1983:BRV

[BdDJ83] J. Boersma, P. J. de Doelder, and J. K. M. Jansen. Book review: A Volume Problem (M. K. Lewis). SIAM Review, 25 (1):102–106, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bruckstein:2009:SSS

[BDE09] Alfred M. Bruckstein, David L.

Donoho, and Michael Elad. From sparse solutions of systems of equations to sparse modeling of signals and images. SIAM Review, 51(1): 34–81, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1983:VP

[BDJ83] J. Boersma, P. J. Dedoelder, and J. K. M. Jansen. A volume problem. SIAM Review, 25(1):102–106, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berry:1999:MVS

[BDJ99] Michael W. Berry, Zlatko Drmac, and Elizabeth R. Jessup. Matrices, vector spaces, and information retrieval. SIAM Review, 41(2): 335–362, June 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34703.

Berry:1995:ULA

[BDO95] Michael W. Berry, Susan T. Dumais, and Gavin W. O’Brien. Using linear algebra for intelligent information retrieval. SIAM Review, 37 (4):573–595, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/27260.
htm.

Braden:1992:IAL

[BDS92] B. Braden, B. Danloy, and F. Schmidt. Inequality of the AGM and the logarithmic mean. SIAM Review, 34 (4):653–654, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burkard:1998:WSS

[BDvD^{+}98] Rainer E. Burkard, Vladimir G. Deineko, René van Dal, Jack A. A. van der Veen, and Gerhard J. Woeginger. Wellsolvable special cases of the traveling salesman problem: a survey. SIAM Review, 40 (3):496–546, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29751.

Boyd:2004:FMM

[BDX04] Stephen Boyd, Persi Diaconis, and Lin Xiao. Fastest mixing Markov chain on a graph. SIAM Review, 46 (4):667–689, December 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42326.

Brady:1978:SRP

[BE78] S. W. Brady and A. Elcrat.

Some results on a postori estimation for approximate solutions of 2nd order elliptic problems. SIAM Review, 20 (3):620, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beale:1990:BRN

[Bea90a] J. Thomas Beale. Book review: Navier–Stokes Equations (Peter Constantin and Ciprian Foias). SIAM Review, 32(1):156–158, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beale:1990:NSE

[Bea90b] J. Thomas Beale. Navier– Stokes equations (Peter Constantin and Ciprian Foias). SIAM Review, 32(1):156– ??, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bebernes:1968:BRD

[Beb68] Jerrold Bebernes. Book review: Differential Equations Stability, Oscillations, Time Lags (A. Halanay). SIAM Review, 10(1):93–94, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bebernes:1979:BRS

[Beb79]  J. W. Bebernes.  Book

review: The Stability of Dynamical Systems (J. P. Lasalle). SIAM Review, 21 (3):418–420, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beckmann:1974:TDD

[Bec74] P. Beckmann. Two-dimensional discrete probability distributions. SIAM Review, 16 (4):546, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beck:1985:BRO

[Bec85] Anatole Beck. Book review: An Optimal Search (Richard Bellman). SIAM Review, 27 (3):447–448, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beck:1994:BRM

[Bec94] James V. Beck. Book review:

The Mollification Method and the Numerical Solution of Ill-Posed Problems (Diego A. Murio). SIAM Review, 36 (3):502–503, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beezer:2009:SV

[Bee09] Robert A. Beezer. Sage. version 3.4. SIAM Review, 51 (4):785–790, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brennan:2023:CIM

[BEG23] Michael C. Brennan, Mark Embree, and Serkan Gugercin. Contour integral methods for nonlinear eigenvalue problems: a systems theoretic approach. SIAM Review, 65(2):439–470, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1389303.

Behforooz:1976:PPT

[Beh76] A. Behforooz. Probabilistic partitions and their application to pattern recognition, cluster analysis, and classification. SIAM Review, 18 (4):795, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beineke:1978:BRC

[Bei78] L. W. Beineke. Book review: Chemical Applications of Graph Theory (A. T. Balaban, ed.). SIAM Review, 20(4):869–870, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beineke:1980:BRF

[Bei80] Lowell W. Beineke. Book review: The Four Color Problem: Assaults and Conquest (Thomas Saaty and Paul Kainen). SIAM Review, 22 (2):241–243, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beineke:1981:BRS

[Bei81] Lowell W. Beineke. Book review: Spectra of Graphs (D. M. Cvetković, Michael Doob and Horst Sachs). SIAM Review, 23(4):546– 548, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beineke:1985:BRG

[Bei85] Lowell W. Beineke. Book review: Graph Theory, Lagów 1981 (M. Borowiecki, J. W. Kennedy and M. M. Syslo, eds.). SIAM Review, 27(2): 260–261, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beineke:1996:BRP

[Bei96] Lowell W. Beineke. Book review: Pearls in Graph Theory: a Comprehensive Introduction (Nora Hartsfield and Gerhard Ringel). SIAM Review, 38(1):159, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bezanson:2017:JFA

[BEKS17] Jeff Bezanson, Alan Edelman, Stefan Karpinski, and Viral B. Shah. Julia: a fresh approach to numerical computing. SIAM Review, 59 (1):65–98, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bellman:1962:BRR

[Bel62] Richard Bellman. Book review: Random Processes (M. Rosenblatt). SIAM Review, 4(4):402, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bellman:1963:BRP

[Bel63a] Richard Bellman. Book review: A Primer of Real Functions (R. P. Boas). SIAM Review, 5(3):290, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bellman:1963:BRC

[Bel63b] Richard Bellman. Book review: Calculus of Variations (L. E. Elsgolc). SIAM Review, 5(2):162, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bellman:1963:OS

[Bel63c] Richard Bellman. An optimal search. SIAM Review, 5(3):274, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beltrami:1963:SAA

[Bel63d] Edward J. Beltrami. Some alternative approaches to distributions. SIAM Review, 5 (4):351–357, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bellman:1965:ADP

[Bel65] Richard Bellman. An application of dynamic programming to location-allocation problems. SIAM Review, 7 (1):126–128, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beltrami:1968:APW

[Bel68] E. J. Beltrami. Another proof of Weyl’s lemma. SIAM Review, 10(2):212– 213, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bellman:1969:CBA

[Bel69]  Richard Bellman.  On

“colonel blotto” and analogous games. SIAM Review, 11(1):66–68, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bell:1986:OPB

[Bel86] Jonathan Bell. Optimizing pacemaker battery life. SIAM Review, 28(1):73– 77, March 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Belair:1988:BRI

[Bél88] Jacques Bélair. Book review:

Intrinsic Geometry of Biological Surface Growth (Philip H. Todd). SIAM Review, 30 (1):138–139, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bell:1990:BRI

[Bel90a] Jonathan Bell. Book review: Introduction to Theoretical Neurobiology. Volume 1: Linear Cable Theory and Dendritic Structure. Volume 2: Nonlinear and Stochastic Theories (Henry C. Tuckwell). SIAM Review, 32(1): 158–160, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bell:1990:BRM

[Bel90b] Jonathan G. Bell. Book review: Mathematical Biology (J. D. Murray). SIAM Review, 32(3):487–489, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bell:2010:KFR

[Bel10] Bradley M. Bell. Kalman filtering with real-time applications [book review of MR 1724120]. SIAM Review, 52 (2):390–392, 2010. CODEN SIREAD. ISSN 0036-1445.

Benedict:1964:PI

[Ben64] Tom Benedict. A pair of identities. SIAM Review, 6 (1):60, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bennett:1970:FOL

[Ben70] Mary Katherine Bennett. A finite orthomodular lattice which does not admit a full set of states. SIAM Review, 12(2):267–271, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bender:1974:AME

[Ben74] Edward A. Bender. Asymptotic methods in enumeration. SIAM Review, 16(4): 485–515, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [Ben76b].

Bender:1976:C

[Ben76a]  E. A. Bender.  Correc-

tion. SIAM Review, 18(2): 292, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bender:1976:EAM

[Ben76b] Edward A. Bender. Errata:

Asymptotic Methods in Enumeration. SIAM Review, 18 (2):292, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Ben74].

Bender:1979:DHR

[Ben79]  Edward A. Bender.  A

driving hazard revisited. SIAM Review, 21(1):136– 138, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Benney:1987:BRW

[Ben87] David J. Benney. Book review: Wave Interactions and Fluid Flows (A. D. D. Craik). SIAM Review, 29 (4):648, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Benjamin:1995:SRR

[Ben95] Arthur T. Benjamin. Sensible rules for remembering duals — the S–O–B method. SIAM Review, 37 (1):85–87, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/26452.htm.

Bernholtz:1960:OAD

[Ber60] B. Bernholtz. Optimum allocation of discharge to units in a hydro-electric generating station. SIAM Review, 2(4): 247–258, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beranek:1962:BRQ

[Ber62] William Beranek. Book review: Quantitative Analysis for Business Decisions (H. Bierman, L. E. Fouraker, and R. K. Jaedicke). SIAM Review, 4(3):268–269, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berztiss:1964:LSF

[Ber64]  A. T. Berztiss.  Least

squares fitting of polynomials to irregularly spaced data. SIAM Review, 6(3): 203–227, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beranek:1965:BRM

[Ber65a]  William Beranek.  Book

review: Mathematical Programming and the Analysis of Capital Budgeting Problems (H. Martin Weingartner). SIAM Review, 7(2): 306–307, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berman:1965:CDE

[Ber65b] Gerald Berman. A class of difference equations with applications. SIAM Review, 7 (4):513–525, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berman:1973:RFN

[Ber73a] A. Berman. Rank factorization of nonnegative matrices. SIAM Review, 15 (3):655, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berndt:1973:DS

[Ber73b] B. C. Berndt. A double summation. SIAM Review, 15 (4):789, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berndt:1974:DS

[Ber74] B. C. Berndt. A double summation. SIAM Review, 16(4): 555–557, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berger:1975:TMS

[Ber75] A. E. Berger. Truncation method for solution of a class of variational inequalities and its relation to Ritz method. SIAM Review, 17 (2):374–375, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berliner:1976:BRC

[Ber76a] Hans Berliner. Book review: Computer Chess (Monroe Newborn). SIAM Review, 18(3):514–515, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berman:1976:GS

[Ber76b] D. N. Berman. The game of slash. SIAM Review, 18 (1):117, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berndt:1976:BFS

[Ber76c] B. C. Berndt. A Bessel function summation. SIAM Review, 18(2):296, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernfeld:1978:BRN

[Ber78] S. R. Bernfeld. Book review: Nonlinear Operators and Differential Equations in Banach Spaces (Robert H. Martin, Jr.). SIAM Review, 20 (1):202–204, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berman:1981:BRG

[Ber81] Abraham Berman. Book review: Generalized Inverses of Linear Transformations (S. L. Campbell and C. D. Meyer, Jr.). SIAM Review, 23(4):545–546, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berger:1983:BRE

[Ber83a] M. S. Berger. Book review: Les Équations De Von Kármán. (Philippe G. Ciarlet and Patrick Rabier). SIAM Review, 25(1):109, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernfeld:1983:BRI

[Ber83b] S. R. Bernfeld. Book review: Instabilities and Catastrophes in Science and Engineering (J. M. T. Thompson). SIAM Review, 25(4): 580–581, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berger:1984:SNE

[Ber84a]  M. Berger.  Solvability

of nonlinear equations and boundary-value problems — Fucik, S. SIAM Review, 26(3):446, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berger:1984:BRS

[Ber84b] Melvyn Berger. Book review: Solvability of Nonlinear Equations and Boundary Value Problems (Svatopluk Fucík). SIAM Review, 26(3):446, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berndt:1984:DI

[Ber84c] B. C. Berndt. A definite integral. SIAM Review, 26 (1):118, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernfeld:1984:NTD

[Ber84d] S. R. Bernfeld. Nonlinear time-discrete systems — Gossel, M. SIAM Review, 26 (3):443–444, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernfeld:1984:BRN

[Ber84e] Stephen R. Bernfeld. Book review: Nonlinear Time-Discrete Systems (M. Gossel). SIAM Review, 26(3): 443–444, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berger:1985:BRG

[Ber85a] Melvyn S. Berger. Book review: Geometrical Methods of Nonlinear Analysis (M.

A. Krasnosel’skii and P. P. Zabreiko). SIAM Review, 27 (2):261–262, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berk:1985:BRM

[Ber85b] Kenneth N. Berk. Book review: Multivariate Descriptive Statistical Analysis-Correspondence Analysis and Related Techniques for Large Matrices (Ludovic Lebart, Alain Morineau, and Kenneth M. Warwick). SIAM Review, 27(3):462–463, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berliner:1985:BRC

[Ber85c] Hans Berliner. Book review:

Computers in Chess-Solving Inexact Search Problems (M. M. Botvinnik). SIAM Review, 27(1):88–90, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berner:1985:BRF

[Ber85d] Andrew J. Berner. Book review: Formal Specification of Interactive Graphics Programming Languages (William R. Mallgren). SIAM Review, 27(4):583–584, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernfeld:1985:BRC

[Ber85e] Stephen R. Bernfeld. Book review: Catastrophe Theory (V. I. Arnold). SIAM Review, 27(1):90–91, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bertsekas:1985:BRN

[Ber85f] Dimitri P. Bertsekas. Book review: Network Flows and Monotropic Optimization (R. T. Rockafellar). SIAM Review, 27(4):585–589, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berzsenyi:1985:BRP

[Ber85g]  George Berzsenyi.  Book

review: Problem-Solving Through Problems (Loren C. Larson). SIAM Review, 27 (2):262–263, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berman:1986:BRN

[Ber86a] Abraham Berman. Book review: Nonnegative Matrices in Dynamic Programming (W. H. M. Zijm). SIAM Review, 28(2):245– 246, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bert:1986:FD

[Ber86b] Charles W. Bert. Falling dominoes. SIAM Review, 28 (2):219–224, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berzsenyi:1986:BRP

[Ber86c] George Berzsenyi. Book review: Problem Book for First Year Calculus (George W. Bluman). SIAM Review, 28 (2):246–247, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernfeld:1988:BRD

[Ber88a] S. R. Bernfeld. Book review: Dynamics of Hierarchical Systems: an Evolutionary Approach (J. S. Nicolis). SIAM Review, 30(2): 324–325, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernstein:1988:CME

[Ber88b] Dennis S. Bernstein. Commuting matrix exponentials. SIAM Review, 30 (1):123, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernard:1990:CDT

[Ber90a] Peter S. Bernard. Convective diffusion in two-dimensional incompressible linear flow. SIAM Review, 32(4):660– 666, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernfeld:1990:BRD

[Ber90b] S. R. Bernfeld. Book review: Differential Equations with Discontinuous Righthand Sides (A. F. Filippov). SIAM Review, 32(2): 312–315, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernstein:1990:OMM

[Ber90c] Dennis S. Bernstein. Orthogonal matrices and the matrix exponential. SIAM Review, 32(4):673, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berger:1991:BRM

[Ber91a] M. S. Berger. Book review: Mathematical Analysis of Nonlinear Dynamic Processes: an Introduction to Processes Governed by Partial Differential Equations (K.-U. Grusa). SIAM Review, 33(1):131, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berk:1991:BRF

[Ber91b] Kenneth N. Berk. Book review: Fundamentals of Mathematical Statistics (H. T. Nguyen and G. S. Rogers). SIAM Review, 33(2):291– 292, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berndt:1991:SFE

[Ber91c]  W. Berndt.  Simultane-

ous functional equations. SIAM Review, 33(4):657– 658, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berndt:1991:BRS

[Ber91d] Wolfgang Berndt. Book review: Simultaneous Functional Equations (Paul Schweitzer). SIAM Review, 33(4):657– 658, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berger:1992:RAI

[Ber92a] Marc A. Berger. Random affine iterated function systems: Curve generation and wavelets. SIAM Review, 34 (3):361–385, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berk:1992:BRR

[Ber92b] Kenneth N. Berk. Book review: Regression Analysis (Ashish Sen and Muni Srivastava). SIAM Review, 34(1): 157–158, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berk:1992:BRS

[Ber92c] Kenneth N. Berk. Book review: Subset Selection in Regression (A. J. Miller). SIAM Review, 34(2):325– 326, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beran:1993:BRS

[Ber93] Rudolf Beran. Book review: Statistical Intervals: a Guide for Practitioners (Gerald J. Hahn and William Q. Meeker). SIAM Review, 35 (2):309–310, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berger:1994:BRI

[Ber94] M. S. Berger. Book review: Infinite Dimensional Morse Theory and Multiple Solution Problems (K. C. Chang). SIAM Review, 36 (1):116–117, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beran:1995:BRC

[Ber95] R. J. Beran. Book review: Computer Intensive Statistical Methods (J. S. Urban Hjorth). SIAM Review, 37 (3):459–460, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernstein:1996:SMI

[Ber96] Dennis S. Bernstein. Some matrix integral identities. SIAM Review, 38(1):147– 148, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berliner:1997:BRB

[Ber97a] Mark Berliner. Book review:

Bayesian Theory, by Jose M. Bernardo and Adrian F. M. Smith. SIAM Review, 39(1): 142–144, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernstein:1997:SMI

[Ber97b]  D. S. Bernstein.  Some

matrix integral identities. SIAM Review, 39(4):763– 765, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2005:EI

[Ber05a] A. J. Bernoff. Education — introduction. SIAM Review, 47(3):551, ???? 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2005:Ea

[Ber05b] Andrew J. Bernoff. Education. SIAM Review, 47 (2):347, June 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97148.

Bernoff:2005:Eb

[Ber05c] Andrew J. Bernoff. Education. SIAM Review, 47 (3):551, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97153.

Bernoff:2005:Ec

[Ber05d] Andrew J. Bernoff. Education. SIAM Review, 47(4):773, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97158.

Bernoff:2006:Ea

[Ber06a] Andrew J. Bernoff. Education. SIAM Review, 48 (1):131, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2006:Eb

[Ber06b] Andrew J. Bernoff. Education. SIAM Review, 48 (2):357, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2006:Ec

[Ber06c] Andrew J. Bernoff. Education. SIAM Review, 48 (3):567, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2006:Ed

[Ber06d] Andrew J. Bernoff. Education. SIAM Review, 48 (4):769, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2007:Ea

[Ber07a] Andrew J. Bernoff. Education. SIAM Review, 49 (1):109, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2007:Eb

[Ber07b] Andrew J. Bernoff. Education. SIAM Review, 49 (2):301, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2007:Ec

[Ber07c] Andrew J. Bernoff. Education. SIAM Review, 49 (3):489, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2007:Ed

[Ber07d] Andrew J. Bernoff. Education. SIAM Review, 49 (4):671, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berzins:2007:API

[Ber07e] M. Berzins. Adaptive polynomial interpolation on evenly spaced meshes. SIAM Review, 49(4):604–627, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2008:Ea

[Ber08a] Andrew J. Bernoff. Education. SIAM Review, 50 (1):113, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2008:Eb

[Ber08b] Andrew J. Bernoff. Education. SIAM Review, 50 (4):789, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2008:EIa

[Ber08c] Andrew J. Bernoff. Education — introduction. SIAM Review, 50(2):345, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2008:EIb

[Ber08d] Andrew J. Bernoff. Education — introduction. SIAM Review, 50(3):551, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2009:Ea

[Ber09a] Andrew J. Bernoff. Education. SIAM Review, 51 (1):191, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2009:Eb

[Ber09b] Andrew J. Bernoff. Education. SIAM Review, 51 (2):415, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2009:Ec

[Ber09c] Andrew J. Bernoff. Education. SIAM Review, 51 (3):611, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2009:EI

[Ber09d] Andrew J. Bernoff. Education: Introduction. SIAM Review, 51(4):765, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2010:Ea

[Ber10a] Andrew J. Bernoff. Education. SIAM Review, 52 (1):149, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2010:Eb

[Ber10b] Andrew J. Bernoff. Education. SIAM Review, 52 (2):357, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2010:Ec

[Ber10c] Andrew J. Bernoff. Education. SIAM Review, 52 (3):543, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bernoff:2010:Ed

[Ber10d] Andrew J. Bernoff. Education. SIAM Review, 52 (4):745, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beskos:1992:BRE

[Bes92] Dimitri E. Beskos. Book review: Elastic Contact Analysis by Boundary Elements (S. Takahasi). SIAM Review, 34 (3):503–504, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beattie:2005:CPR

[BES05] Christopher A. Beattie, Mark Embree, and D. C. Sorensen. Convergence of polynomial restart Krylov methods for eigenvalue computations. SIAM Review, 47 (3):492–515, September 2005.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43307.

Bettis:1978:EER

[Bet78] D. G. Bettis. Efficient embedded Runge–Kutta method of orders one through 6. SIAM Review, 20(3):620, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beutler:1966:EFR

[Beu66] Frederick J. Beutler. Errorfree recovery of signals from irregularly spaced samples. SIAM Review, 8(3):328– 335, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beutler:1983:SSQ

[Beu83a] F. J. Beutler. The single server queue — Cohen, J. W. SIAM Review, 25(4): 589–590, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beutler:1983:BRS

[Beu83b] Frederick J. Beutler. Book review: The Single Server Queue (J. W. Cohen). SIAM Review, 25(4):589–590, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beutler:1985:BRF

[Beu85] Frederick J. Beutler. Book review: A First Course in Bulk Queues (M. L. Chaudhry and J. G. C. Templeton). SIAM Review, 27(2):263– 264, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beutler:1987:BRM

[Beu87] Frederick J. Beutler. Book review: Mathematical Topics in Telecommunications. Volume 2: Problems of Randomness in Communication Engineering (K. W. Cattermole and J. J. O’Reilly, eds.). SIAM Review, 29(3): 476–477, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beutler:1989:BRD

[Beu89] Frederick J. Beutler. Book review: Dynamic Programming: Deterministic and Stochastic Models (Dimitri P. Bertsekas). SIAM Review, 31 (1):132–133, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beutler:1992:BRC

[Beu92] Frederick J. Beutler. Book review: Computer Networks and Systems: Queueing Theory and Performance Evaluation (Thomas G. Robertazzi). SIAM Review, 34(2):

326–327, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1996:PPI

[BEV96] P. Borwein, T. Erdelyi, and Walter Van Assche. Polynomials and polynomial inequalities. SIAM Review, 38 (4):705–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boa:1982:PVI

[BF82] J. A. Boa and H. E. Fettis. Principal value of an integral. SIAM Review, 24 (1):83–85, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bandle:1996:HRC

[BF96] Catherine Bandle and Martin Flucher. Harmonic radius and concentration of energy; hyperbolic radius and Liouville’s equations ΔU = e^{U} and ΔU = U^{\frac{n+2}{n−2}} . SIAM Review, 38(2): 191–238, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/27201.htm.

Bertozzi:2016:DIM

[BF16] Andrea L. Bertozzi and Arjuna Flenner. Diffuse interface models on graphs for classification of high dimensional data. SIAM Review, 58 (2):293–328, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Broadbridge:1999:BWG

[BFF^{+}99] P. Broadbridge, G. R. Fulford, N. D. Fowkes, D. Y. C. Chan, and C. Lassig. Bubbles in wet, gummed wine labels. SIAM Review, 41(2): 363–372, June 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34702.

Bridges:1995:STE

[BFM95] Thomas Bridges, Jacques E. Furter, and Kenneth R. Meyer. Singularity theory and equivariant symplectic maps. SIAM Review, 37(1): 135–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brezina:2005:ASA

[BFM^{+}05] M. Brezina, R. Falgout, S. MacLachlan, T. Manteuffel, S. McCormick, and J. Ruge. Adaptive smoothed aggregation (αSA) multigrid. SIAM Review, 47(2): 317–346, June 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/62627.

Burrows:1964:BRP

[BG64] James W. Burrows and Wallace Givens. Book review: On a Property of a Unitary Matrix (J. L. Brenner and F. T. Smith). SIAM Review, 6(4): 460–461, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barwell:1976:CAS

[BG76] V. Barwell and A. George.

Comparison of algorithms for solving symmetric indefinite systems of equations. SIAM Review, 18(4):795, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bjorck:1977:EMA

[BG77] Åke Björck and Gene H.

Golub. Eigenproblems for matrices associated with periodic boundary conditions. SIAM Review, 19(1): 5–16, January 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2029322.

Bartuccelli:1995:CCI

[BG95a] M. V. Bartuccelli and S. A.

Gourley. A conjecture concerning an integral arising in population modelling.

SIAM Review, 37(3):440– 441, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1995:BRD

[BG95b] J. Boersma and M. L. Glasser. Book review: A Definite Integral (M. L. Glasser, J. Boersma and P. J. de Doelder). SIAM Review, 37 (4):615–618, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burnside:1996:SPT

[BG96] R. R. Burnside and P. B. Guest. A simple proof of the transposed QR algorithm. SIAM Review, 38(2): 306–308, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/p306.htm.

Bochev:1998:FEM

[BG98] Pavel B. Bochev and Max D. Gunzburger. Finite element methods of least-squares type. SIAM Review, 40 (4):789–837, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32115.

Bartel:2018:PRE

[BG18] Andreas Bartel and Michael Günther. PDAEs in refined electrical network modeling. SIAM Review, 60(1): 56–91, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:1995:LMI

[BGF^{+}95] S. Boyd, L. E. Ghaoui, E. Feron, V. Balakrishnan, and V. A. Yakubovich. Linear matrix inequalities in system and control theory. SIAM Review, 37(3):479– ??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blondel:2004:MSB

[BGH^{+}04] Vincent D. Blondel, Anahí Gajardo, Maureen Heymans, Pierre Senellart, and Paul Van Dooren. A measure of similarity between graph vertices: Applications to synonym extraction and Web searching. SIAM Review, 46 (4):647–666, December 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41596.

Bick:2023:WHO

[BGHS23] Christian Bick, Elizabeth Gross, Heather A. Harrington, and Michael T. Schaub. What are higher-order networks? SIAM Review, 65(3):686–731, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1414024.

Bojanic:1967:BRF

[BGL^{+}67] R. Bojanic, H. W. Gould, D. L. Lansing, J. M. Honer, A. Milewski, and D. E. Amos. Book review: A Finite Series (J. P. Church). SIAM Review, 9(4):749– 750, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Byrnes:2001:FCW

[BGL01] Christopher I. Byrnes, Sergei V. Gusev, and Anders Lindquist. From finite covariance windows to modeling filters: a convex optimization approach. SIAM Review, 43 (4):645–675, December 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39219.

Benson:2017:SRW

[BGL17] Austin R. Benson, David F. Gleich, and Lek-Heng Lim. The spacey random walk: a stochastic process for higherorder data. SIAM Review, 59 (2):321–345, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boehm:1996:BRE

[BGML96] W. Boehm, Carl C. Grosjean, W. Weston Meyer, and David M. Lazoff. Book review: Expected Wire Length Between Two Randomly Chosen Terminals (David M. Lazoff). SIAM Review, 38(2): 321–324, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Benner:2015:SPB

[BGW15] Peter Benner, Serkan Gugercin, and Karen Willcox. A survey of projection-based model reduction methods for parametric dynamical systems. SIAM Review, 57(4): 483–531, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnard:1976:FSS

[BH76a] R. D. Barnard and I. M.

Horowitz. Feedback system synthesis based on fixedpoint optimal-aim formulations. SIAM Review, 18(4): 794–795, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bettis:1976:OFR

[BH76b] D. G. Bettis and M. K. Horn.

Optimal M-fold Runge– Kutta methods. SIAM Review, 18(4):795, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Briggs:1990:FMA

[BH90] William L. Briggs and Van Emden Henson. The FFT as a multigrid algorithm. SIAM Review, 32(2): 252–261, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Baines:1996:MFE

[BH96] M. J. Baines and Weizhang Huang. Moving finite elements. SIAM Review, 38(3): 530–??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bank:2003:NPP

[BH03] Randolph E. Bank and Michael Holst. A new paradigm for parallel adaptive meshing algorithms. SIAM Review, 45(2):291– 323, June 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/42061.

Bindel:2015:LTN

[BH15] David Bindel and Amanda Hood. Localization theorems for nonlinear eigenvalue problems. SIAM Review, 57(4): 585–607, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bhargava:1993:BRR

[Bha93] D. K. Bhargava. Book review: A Resistance Problem (E. N. Gilbert and L. A. Shepp). SIAM Review, 35(3): 492–493, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bhaya:1998:RMP

[Bha98] Amit Bhaya. Real matrices with positive determinant are homotopic to the identity. SIAM Review, 40(2):335–340, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31623.

Butler:2009:GG

[BHKL09] Steve Butler, Mohammad T.

Hajiaghayi, Robert D. Kleinberg, and Tom Leighton. Hat guessing games. SIAM Review, 51(2):399–413, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ben-Israel:1967:BRL

[BI67] Adi Ben-Israel. Book review:

Linear Programming (Michel Simonnard). SIAM Review, 9(3):608, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ben-Israel:1968:BRG

[BI68a] Adi Ben-Israel. Book review: Geometric Programming-Theory and Application (R. J. Duffin, E. L. Peterson and C. Zener). SIAM Review, 10 (2):235–236, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ben-Israel:1968:BRM

[BI68b] Adi Ben-Israel. Book review: Methodes de Calcul Numérique-tome 2: Elements de Théorie des Matrices Carrées et Rectangles en Analyse Numérique (A. Korganoff and M. Pavel-Parvu). SIAM Review, 10 (1):95–96, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ben-Israel:1968:BRP

[BI68c] Adi Ben-Israel. Book review: Programming, Games and Transportation Networks (Claude Bergs and A. Ghouila). SIAM Review, 10 (2):238–239, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ben-Israel:1969:NPM

[BI69]  Adi Ben-Israel.  A note

on partitioned matrices and equations. SIAM Review, 11 (2):247–250, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ben-Israel:1970:BRB

[BI70a] Adi Ben-Israel. Book review: Boolean Methods in Operations Research and Related Areas (Peter L. Hammer and Sergiu Rudeanu). SIAM Review, 12(4):594– 596, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ben-Israel:1970:BRM

[BI70b] Adi Ben-Israel. Book review:

Monotone Processes of Convex and Concave Type (R. Tyrell Rockafellar). SIAM Review, 12(4):594, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ben-Israel:1970:CMC

[BI70c] Adi Ben-Israel. On conemonotonicity of complex matrices. SIAM Review, 12(1): 120–123, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brams:2018:MRS

[BI18] Steven J. Brams and Mehmet S.

Ismail. Making the rules of sports fairer. SIAM Review, 60(1):181–202, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bickel:1994:BRI

[Bic94] Peter J. Bickel. Book review: Information Bounds and Nonparametric Maximum Likelihood Estimation (Piet Groeneboom and Jon A. Wellner). SIAM Review, 36(3):503–504, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Billingham:1993:II

[Bil93a] J. Billingham. An infinite integral. SIAM Review, 35 (2):298, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Billingham:1993:TIA

[Bil93b] John Billingham. Two integrals arising from a reactiondiffusion problem. SIAM Review, 35(3):490, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Billingham:1994:II

[Bil94a] J. Billingham. An infinite integral. SIAM Review, 36 (1):115, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Billingham:1994:TIA

[Bil94b] J. Billingham. Two integrals arising from a reactiondiffusion problem. SIAM Review, 36(3):498–499, ????

1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Binmore:1984:BRD

[Bin84] K. G. Binmore. Book review: Dynamic Noncooperative Game Theory (Tamer Baser and Geert Jan Olsder). SIAM Review, 26(3): 444–445, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Birch:1964:BRI

[Bir64] John J. Birch. Book review: Introduction to the Theory of Statistics, 2nd Edition (Alexander M. Moon and Franklin A. Graybill). SIAM Review, 6(4):472– 473, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Birnbaum:1965:PEE

[Bir65]  S. Birnbaum.  Perturba-

tion of eigenvalues — with an engineering application. SIAM Review, 7(1):13–30, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Birkhoff:1969:BRT

[Bir69a] Garrett Birkhoff. Book review: The Theory of Splines and Their Applications (J. H. Ahlberg, E. N. Nilson and J. L. Walsh). SIAM Review, 11 (1):91–92, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Birkhoff:1969:MP

[Bir69b]  Garrett Birkhoff.  Math-

ematics and psychology. SIAM Review, 11(4):429– 469, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Birkhoff:1983:NFD

[Bir83] Garrett Birkhoff. Numerical fluid dynamics. SIAM Review, 25(1):1–34, January 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Birkhoff:1986:BRA

[Bir86]  Garrett Birkhoff.  Book

review: Applications of Abstract Algebra (George Mackiw). SIAM Review, 28 (2):247, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bismut:1978:IAD

[Bis78] Jean-Michel M. Bismut. An introductory approach to duality in optimal stochastic control. SIAM Review, 20 (1):62–78, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bisgard:2015:MPS

[Bis15] James Bisgard. Mountain passes and saddle points. SIAM Review, 57(2):275– 292, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barinka:1975:SPN

[BJ75] L. L. Barinka and R. L. Jennings. Substructuring procedure for numerical analysis of thin shells of revolution. SIAM Review, 17 (2):374, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1976:VFM

[BJ76] C. Blair and R. Jeroslow.

Value function of a mixed integer program I. SIAM Review, 18(4):795, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barraucand:1979:TI

[BJ79] P. Barraucand and W. B.

Jordan. Two inequalities. SIAM Review, 21(1):140– 143, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brams:2011:DCP

[BJK11] Steven J. Brams, Michael A.

Jones, and Christian Klamler. Divide-and-conquer: a proportional, minimalenvy cake-cutting algorithm. SIAM Review, 53(2):291– 307, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i2/p291_s1.

Barrenechea:2024:FEM

[BJK24] Gabriel R. Barrenechea, Volker John, and Petr Knobloch. Finite element methods respecting the discrete maximum principle for convection–diffusion equations. SIAM Review, 66(1): 3–88, February 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bjorgum:1961:BRT

[Bjö61] O. B. Björgum. Book review: Turbulence: Classic Papers on Statistical Theory (S. K. Friedlander and L. Topping, eds.). SIAM Review, 3(4): 338–339, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bjork:1985:BRN

[Bjö85] Tomas Björk. Book review: Nonlinear Filtering and Smoothing — an Introduction to Martingales, Stochastic Integrals and Estimation (Venkatarama Krishnan). SIAM Review, 27(4): 589–590, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bjorstad:1990:ILA

[Bjø90]  Petter E. Bjørstad.  An

inequality in linear algebra. SIAM Review, 32(1): 141, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brailsford:1966:TI

[BK66] A. D. Brailsford and M. S.

Klamkin. Two integrals. SIAM Review, 8(1):107, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Belinfante:1969:ILG

[BK69] J. G. Belinfante and B. Kolman. An introduction to Lie groups and Lie algebras, with applications. III: Computational methods and applications of representation theory. SIAM Review, 11(4): 510–543, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berman:1976:RCS

[BK76] David Berman and M. S.

Klamkin. A reverse card shuffle. SIAM Review, 18(3): 491–492, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burridge:1978:PSC

[BK78] R. Burridge and J. B. Keller. Peeling, slipping and cracking — some one-dimensional free-boundary problems in mechanics. SIAM Review, 20 (1):31–61, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brown:1983:SAE

[BK83] Charles C. Brown and James A. Koziol. Statistical aspects of the estimation of human risk from suspected environmental carcinogens. SIAM Review, 25 (2):151–181, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bruckstein:1987:ISD

[BK87] A. M. Bruckstein and T. Kailath. Inverse scattering for discrete transmission-line models. SIAM Review, 29(3):359– 389, September 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Batman:1988:BLM

[BK88] Don Batman and M. S. Klamkin. Bounds for the log mean radius for heat conduction in a multilagged cylindrical pipe. SIAM Review, 30 (2):298–300, June 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barbu:1995:ACN

[BK95] Viorel Barbu and Philip Korman. Analysis and control of nonlinear infinite dimensional systems. SIAM Review, 37(4):633– ??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berry:1999:RZE

[BK99] M. V. Berry and J. P.

Keating. The Riemann zeros and eigenvalue asymptotics. SIAM Review, 41(2): 236–266, June 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34749.

Benesova:2017:WLS

[BK17] Barbora Benesová and Martin Kruzík. Weak lower semicontinuity of integral functionals and applications. SIAM Review, 59(4):703– 766, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brandts:2009:NSP

[BKKŠ09] Jan Brandts, Sergey Korotov, Michal Křížek, and Jakub Šolc. On nonobtuse simplicial partitions. SIAM Review, 51(2):317– 335, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bauer:1975:MEL

[BKR75] Louis Bauer, Herbert B. Keller, and Edward L. Reiss. Multiple eigenvalues lead to secondary bifurcation. SIAM Review, 17(1):101– 122, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Belinfante:1966:ILG

[BKS66] J. G. Belinfante, B. Kolman, and H. A. Smith. An introduction to Lie groups and Lie algebras, with applications. SIAM Review, 8(1): 11–46, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Belinfante:1968:ILG

[BKS68] J. G. Belinfante, B. Kolman, and H. A. Smith. An introduction to Lie groups and Lie algebras, with applications. II: The basic methods and results of representation theory. SIAM Review, 10(2): 160–195, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Breach:1996:BRT

[BKS96] D. R. Breach, C. Krattenthaler, and Marcus Schorn. Book review: Two Combinatorial Identities (Ma Xin-Bong and Wang Tian-Ming). SIAM Review, 38(1):148– 152, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brune:2015:CSN

[BKST15] Peter R. Brune, Matthew G.

Knepley, Barry F. Smith, and Xuemin Tu. Composing scalable nonlinear algebraic solvers. SIAM Review, 57(4): 535–565, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bonomi:1984:CTS

[BL84] Ernesto Bonomi and Jean-Luc L. Lutton. The Ncity travelling salesman problem: Statistical mechanics and the Metropolis algorithm. SIAM Review, 26(4): 551–568, October 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bochev:2005:FES

[BL05] Pavel Bochev and R. B.

Lehoucq. On the finite element solution of the pure Neumann problem. SIAM Review, 47(1): 50–66, March 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/42607.

Bryan:2006:ELA

[BL06] Kurt Bryan and Tanya Leise. The $25,000,000,000 eigenvector: The linear algebra behind Google. SIAM Review, 48(3):569–581, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blanchette:2009:ECM

[BL09] François Blanchette and Yue Lei. Energy considerations for multiphase fluids with variable density and surface tension. SIAM Review, 51(2): 423–431, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryan:2010:III

[BL10] Kurt Bryan and Tanya Leise. Impedance imaging, inverse problems, and Harry Potter’s cloak. SIAM Review, 52(2): 359–377, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryan:2013:MDL

[BL13] Kurt Bryan and Tanya Leise. Making do with less: an introduction to compressed sensing. SIAM Review, 55(3): 547–566, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blanch:1964:NEC

[Bla64] G. Blanch. Numerical evaluation of continued fractions. SIAM Review, 6(4): 383–421, October 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blank:1966:BRN

[Bla66]  Albert A. Blank.  Book

review: Non-Linear Wave Propagation with Applications to Physics and Magnetohydrodynamics (A. Jeffrey and T. Taniuti). SIAM Review, 8(4):546–547, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blachman:1970:BRL

[Bla70] Nelson M. Blachman. Book review: Limit of an Integral (W. A. J. Luxemburg). SIAM Review, 12(3): 470–471, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blakley:1981:BRI

[Bla81] G. R. Blakley. Book review: Information Linkage Between Applied Mathematics and Industry (Peter C. C. Wang, Arthur L. Schoenstadt, Bert I. Russak, and Craig Comstock, eds.). SIAM Review, 23(2): 267–269, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1982:BRO

[Bla82] Charles Blair. Book review: Optimization with Disjunctive Constraints (Hanif D. Sherali and C. M. Shetty). SIAM Review, 24(3):368– 369, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blakely:1983:BRC

[Bla83] G. R. Blakely. Book review: Cryptography: a Primer (Alan G. Konheim). SIAM Review, 25(1):111– 114, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blakley:1984:BRS

[Bla84] G. R. Blakley. Book review: Secure Communications and Asymmetric Cryptosystems (Gustavus J. Simmons, ed.). SIAM Review, 26 (1):126–127, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1985:BRL

[Bla85a] Charles Blair. Book review: Linear Programming (Katta G. Murty) and Linear Programming (Vasek Chvǎtal). SIAM Review, 27 (1):91–93, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1985:BRP

[Bla85b] Charles Blair. Book review: Problem Complexity and Method Efficiency in Optimization (A. S. Nemirovsky and D. B. Yudin). SIAM Review, 27(2):264– 265, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blakley:1986:BRS

[Bla86a] G. R. Blakley. Book review: Secure Speech Communications (Henry J. Beker and Fred C. Piper). SIAM Review, 28(4):581–582, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blank:1986:BRM

[Bla86b] Albert A. Blank. Book review: Multivariable Analysis (G. Baley Price). SIAM Review, 28(2):247–248, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blachman:1987:GSD

[Bla87] N. M. Blachman. A generalization of Student’s t distribution. SIAM Review, 29 (1):123, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1988:BRP

[Bla88a] Charles Blair. Book review: Probabilistic Analysis of Algorithms: On Computing Methodologies for Computer Algorithms Performance Evaluation (Micha Hofri). SIAM Review, 30(4): 663–664, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1988:BRSa

[Bla88b] Charles Blair. Book review: Studies in Complexity Theory (Ronald V. Book, Ker-I. Ko, Stephen R. Mahaney and Kenneth Mcaloon, eds.). SIAM Review, 30(1): 139–140, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1988:BRSb

[Bla88c] Charles Blair. Book review: The Simplex Method: a Probabilistic Analysis (Karl Heinz Borgwardt). SIAM Review, 30(3):510–511, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1988:BRT

[Bla88d] Charles Blair. Book review: Theory of Linear and Integer Programming (Alexander Schrijver). SIAM Review, 30 (2):325–326, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blakley:1988:BRP

[Bla88e] G. R. Blakley. Book review: Primality and Cryptography (Evangelos Kranakis). SIAM Review, 30(3):511– 512, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blakley:1988:BRT

[Bla88f] G. R. Blakley. Book review: Two Issues in Public Key Cryptography: RSA Bit Security and a New Knapsack Type System (Ben-Zion Chor). SIAM Review, 30(4): 664–665, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1989:BRG

[Bla89a] Charles Blair. Book review:

Geometric Algorithms and Combinatorial Optimization (Martin Grotschel, Laszló Lovász, and Alexander Schrijver). SIAM Review, 31(2): 332–333, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1989:BRL

[Bla89b] Charles Blair. Book review: Linear Complementarily, Linear and Nonlinear Programming (K. G. Murty). SIAM Review, 31 (2):333–334, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blair:1990:BRI

[Bla90] Charles Blair. Book review: Integer and Combinatorial Optimization (George L. Nemhauser and Laurence A. Wolsey). SIAM Review, 32(2):315, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blaine:1991:PI

[Bla91] Larry G. Blaine. A probability integral. SIAM Review, 33(2):276, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blanchard:1993:BRI

[Bla93] Paul Blanchard. Book review: Iteration of Rational Functions (Alan F. Beardon). SIAM Review, 35(2): 310–312, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blachman:1994:MV

[Bla94] N. M. Blachman. A maximum volume. SIAM Review, 36(2):278–281, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Black:1997:PCO

[Bla97a]  K. Black.  Putting con-

straints in optimization for first-year calculus students. SIAM Review, 39(2):310– 312, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Black:1997:CNP

[Bla97b] Kelly Black. Classroom note:

Putting constraints in optimization for first-year calculus students. SIAM Review, 39(2):310–312, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29481.

Bleistein:1981:BRB

[Ble81] Norman Bleistein. Book review: The Boundary-Layer Method in Diffraction Problems (V. M. Basic and N. Y. Kirpicnikova). SIAM Review, 23(4):539, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Biondini:2018:UBM

[BLMT18] Gino Biondini, Sitai Li, Dionyssios Mantzavinos, and Stefano Trillo. Universal behavior of modulationally unstable media. SIAM Review, 60(4):888–908, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Block:1973:SFT

[Blo73] I. Edward Block. SIAM — its first three years. SIAM Review, 15(2):v–ix, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blomqvist:1976:BRS

[Blo76] Nils Blomqvist. Book review: Statistical Techniques in Simulation, Part I (Jack P. C. Kleijnen). SIAM Review, 18 (2):314–316, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blomqvist:1977:BRS

[Blo77] Nils Blomqvist. Book review: Statistical Techniques in Simulation, Part II (Jack P. C. Kleijnen). SIAM Review, 19 (4):752–753, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bloch:1989:BRV

[Blo89] Anthony M. Bloch. Book review: Variational and Hamiltonian Control Systems (P. E. Crouch and A. J. van der Schaft). SIAM Review, 31(2): 334–336, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bloch:1991:BRN

[Blo91] Anthony M. Bloch. Book review: Nonlinear Dynamical Control Systems (H. Nijmeijer and A. J. van der Schaft). SIAM Review, 33 (2):292–294, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blom:1994:OBS

[Blo94] G. Blom. Overlapping binary sequences. SIAM Review, 36 (4):657, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blondel:1995:BRZ

[Blo95] Vincent Blondel. Book review: Zeros in the Unit Disk (M. Meneguetle). SIAM Review, 37(4):619, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Broersma:1992:AAG

[BLS92] H. J. Broersma, A. Liu, and L. V. Saxton. An affirmativeaction graph. SIAM Review, 34(4):652–653, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bauschke:2008:HTO

[BLT08] Heinz H. Bauschke, Yves Lucet, and Michael Trienis. How to transform one convex function continuously into another. SIAM Review, 50(1):115–132, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blum:1965:BRP

[Blu65] J. R. Blum. Book review: Proceedings of the Symposium on Time Series Analysis (Murray Rosenblatt, ed.). SIAM Review, 7(1): 154–159, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blundon:1978:BRT

[Blu78] W. J. Blundon. Book review: A Triangle Inequality (I. J. Schoenberg). SIAM Review, 20(2):399–400, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bluman:1986:BRG

[Blu86] George W. Bluman. Book review: Group Invariance in Engineering Boundary Value Problems (R. Seshadri and T. Y. Na). SIAM Review, 28 (2):248–249, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bluman:1991:BRN

[Blu91] George Bluman. Book review: Nonlinear Boundary Value Problems in Science and Engineering (C. Rogers and W. F. Ames). SIAM Review, 33(2):294– 296, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bluman:1994:BRD

[Blu94a] George Bluman. Book review: Differential Equations and Group Methods (James M. Hill). SIAM Review, 36 (1):117, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bluman:1994:BRI

[Blu94b] George Bluman. Book review: Infinite Dimensional Morse Theory and Multiple Solution Problems (K. C. Chang). SIAM Review, 36 (1):117, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bertozzi:1993:MRG

[BM93] Andrea Bertozzi and James McKenna. Multidimensional residues, generating functions, and their application to queueing networks. SIAM Review, 35(2): 239–268, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1995:GSP

[BM95] J. Boersma and J. Molenaar. Geometry of the shoulder of a packaging machine. SIAM Review, 37(3):406– 422, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/27080.htm.

Boros:1998:ITP

[BM98] George Boros and Victor H. Moll. An integral with three parameters. SIAM Review, 40 (4):972–980, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31740.

Bakker:2003:HDP

[BM03] Eric Bakker and A. J. Meir. How do pulsed amperometric ion sensors work? A simple PDE model. SIAM Review, 45(2):327–344, June 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37712.

Breach:1976:BRC

[BMMO76] D. R. Breach, D. McCarthy, D. Monk, and P. E. O’Neil. Book review: A Combinatorial Identity (P. Barrucand). SIAM Review, 18(2): 303–304, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berman:1976:MS

[BN76] A. Berman and M. Neumann. Monotone submatrices. SIAM Review, 18 (3):490, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1985:I

[BNR85] J. L. Brenner, W. A. Newcomb, and O. G. Ruehr. An inequality. SIAM Review, 27 (4):574, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barinka:1976:PSV

[BNT76] L. L. Barinka, K. W.

Neves, and P. G. Tuttle. Performance of some vectorized mathematical software. SIAM Review, 18 (4):794, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Babuska:2010:SCM

[BNT10] Ivo Babuška, Fabio Nobile, and Raúl Tempone. A stochastic collocation method for elliptic partial differential equations with random input data. SIAM Review, 52(2):317– 355, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brubeck:2021:VA

[BNT21] Pablo D. Brubeck, Yuji Nakatsukasa, and Lloyd N. Trefethen. Vandermonde with Arnoldi. SIAM Review, 63(2):405–415, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1965:BRT

[Boa65] R. P. Boas, Jr. Book review: A Treatise on Trigonometric Series (N. K. Bary). SIAM Review, 7(3):429– 431, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1970:GTS

[Boa70] R. P. Boas, Jr. Generalized Taylor series, quadrature formulas, and a formula by Kronecker. SIAM Review, 12(1): 116–119, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1976:BRA

[Boa76]  R. P. Boas.  Book re-

view: Applied and Computational Complex Analysis (Peter Henrici). SIAM Review, 18(3):506–510, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1978:BRA

[Boa78]  R. P. Boas.  Book re-

view: Applied and Computational Complex Analysis (Peter Henrici). SIAM Review, 20(4):871, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//www.jstor.org/stable/
2029767.

Boas:1981:BRI

[Boa81] R. P. Boas. Book review:

An Introduction to Classical Complex Analysis (Robert B. Burckel). SIAM Review, 23 (3):410, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boa:1982:BRP

[Boa82a] James A. Boa. Book review:

Principal Value of an Integral (H. E. Fettis). SIAM Review, 24(1):83–85, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1982:BRA

[Boa82b] R. P. Boas. Book review: Aspects of Contemporary Complex Analysis (D. A. Brannan and J. G. Clunie, eds.). SIAM Review, 24(3):369, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1983:BRS

[Boa83] R. P. Boas. Book review: A Second Course on Real Functions (A. C. M. Van Rooij and W. H. Schikhof). SIAM Review, 25(3):410, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1984:BRI

[Boa84] R. P. Boas. Book review: Intermediate Real Analysis (Emanuel Fischer). SIAM Review, 26(1):129, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1985:BRP

[Boa85]  R. P. Boas.  Book re-

view: Primer of Modern Analysis (Kennan T. Smith). SIAM Review, 27(1):93– 94, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1987:BRA

[Boa87] R. P. Boas. Book review: Applied and Computational Complex Analysis. Volume Three: Discrete Fourier Analysis-Cauchy Integrals-Construction of Conformal Maps-Univalent Functions (Peter Henrici. John Wiley). SIAM Review, 29 (1):138, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//www.jstor.org/stable/
pdf/2030950.pdf.

Boas:1988:BRF

[Boa88]  R. P. Boas.  Book re-

view: Fourier Analysis (T. W. Körner). SIAM Review, 30(4):665–666, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bobisud:1966:SFO

[Bob66] Larry Bobisud. On the single first-order partial differential equation with a small parameter. SIAM Review, 8 (4):479–493, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boehme:1968:CI

[Boe68] T. K. Boehme. The convolution integral. SIAM Review, 10(4):407–416, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1976:BRL

[Boe76] J. Boersma. Book review:

Limit of an Integral (M. L. Glasser). SIAM Review, 18 (4):770–772, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boehm:1988:BRP

[Boe88a] Wolfgang Boehm. Book review: Projective Geometry and Its Applications to Computer Graphics (Michael A. Penna and Richard R. Patterson). SIAM Review, 30(1): 140–141, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1988:EPD

[Boe88b] J. Boersma. The entropy of a Poisson distribution. SIAM Review, 30(2):315– 317, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boehm:1989:BRI

[Boe89] Walter Boehm. Book review: An Inequality (Wojciech Szatzschneider). SIAM Review, 31(4):677–678, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1991:TIA

[Boe91]  J. Boersma.  Two inte-

grals arising from a cloud model. SIAM Review, 33(3): 474–477, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boehm:1993:OL

[Boe93] W. Boehm. An operator limit. SIAM Review, 35 (4):642, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boehm:1994:OL

[Boe94] W. Boehm. An operator limit. SIAM Review, 36(2): 277–278, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boehm:1995:OLV

[Boe95a] W. Boehm. An operator limit (vol 36, pg 659, 1994). SIAM Review, 37(2):250, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1995:BRN

[Boe95b] J. Boersma. Book review:

A Nonharmonic Trigonometric Series (M. L. Glasser). SIAM Review, 37(3):443– 448, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boehm:1996:EWL

[Boe96a]  W. Boehm.  Expected

wire length between two randomly chosen terminals. SIAM Review, 38(2):321– 324, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1996:BRF

[Boe96b] J. Boersma. Book review:

A Function Arising in One-Dimensional Percolation (M. L. Glasser). SIAM Review, 38 (4):671–675, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1996:TCI

[Boe96c] J. Boersma. Two combinatorial identities. SIAM Review, 38(1):152–153, ???? 1996.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1998:PIA

[Boe98a] J. Boersma. A parametric integral arising from a mixed boundary value problem for the Laplacian. SIAM Review, 40(4):984– 986, November 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1998:SBS

[Boe98b] J. Boersma. A sum over binary sequences. SIAM Review, 40(2):372–374, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bogdanoff:1973:H

[Bog73] D. A. Bogdanoff. A hypervolume. SIAM Review, 15 (3):655, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boggs:1975:CBI

[Bog75] P. T. Boggs. Convergence of Ben–Israel iteration for nonlinear least-squares problems. SIAM Review, 17 (2):375, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bohm:1988:CI

[Boh88a] W. Bohm. 3 convolution integrals. SIAM Review, 30(4): 661–662, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bohm:1988:AAT

[Boh88b] W. Bohm. The attenuated Abel transform. SIAM Review, 30(1):125–128, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bohm:1988:BRT

[Böh88c] Walter Böhm. Book review:

Three Convolution Integrals (F. Kuchuk). SIAM Review, 30(4):661–662, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bohm:1990:TS

[Böh90] W. Böhm. A trigonometric sum. SIAM Review, 32 (1):141, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bojanic:1965:BRA

[Boj65] R. Bojanic. Book review: An Asymptotic Result on a Zero of a Function (R. L. Graham and L. Shepp). SIAM Review, 7(2):280–283, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bojadziev:1993:BRS

[Boj93] George Bojadziev. Book review: Systems of Nonlinear Partial Differential Equations, Applications to Biology and Engineering (Anthony W. Leung). SIAM Review, 35(1):143–145, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bojanov:1997:BRS

[Boj97] Borislav Bojanov. Book review: Selected Topics in Approximation and Computation, by Marek A. Kowalski, Krzysztof A. Sikorski, and Frank Stenger. SIAM Review, 39(2):333–334, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boltyanski:1991:BRG

[Bol91a] Vladimir G. Boltyanski. Book review: Games, Set, & Math: Enigmas and Conundrums (Ian Stewart). SIAM Review, 33(2):296– 297, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boltyanski:1991:BRH

[Bol91b] Vladimir G. Boltyanski. Book review: How Does One Cut a Triangle? (Alexander Soifer). SIAM Review, 33(2): 297–299, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boltyanski:1991:GSM

[Bol91c] Vladimir G. Boltyanski.

Games, set, and math: Enigmas and conundrums (Ian Stewart). SIAM Review, 33 (2):296–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boltyanski:1991:HDO

[Bol91d] Vladimir G. Boltyanski. How does one cut a triangle? (Alexander Soifer). SIAM Review, 33(2):297–??, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bomze:2002:RVD

[Bom02] Immanuel M. Bomze. Regularity versus degeneracy in dynamics, games, and optimization: a unified approach to different aspects. SIAM Review, 44(3):394– 414, September 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37565.

Bongiorno:1967:RLS

[Bon67] J. J. Bongiorno, Jr. On the response of linear systems to inputs with limited amplitudes and slopes. SIAM Review, 9(3):554– 563, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Book:1979:BRA

[Boo79] Ronald V. Book. Book review: Algebraic and Automata-Theoretic Properties of Formal Languages (Seymour Ginsburg). SIAM Review, 21 (4):573–576, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Booning:1986:UP

[Boö86]  DankMar Boöning.  A

uniqueness problem. SIAM Review, 28(3):397–398, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Book:1994:RPN

[Boo94] Ronald V. Book. Relativizations of the P =? NP and other problems: Developments in structural complexity theory. SIAM Review, 36 (2):157–175, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/23994.htm.

Bordelon:1962:BLE

[Bor62] Derrill J. Bordelon. Best linear estimate of missile location based on diverse simultaneous tracking data. SIAM Review, 4(2):91–93, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bordelon:1973:BRI

[Bor73] D. J. Bordelon. Book review: Inequalities for Special Functions (D. K. Ross). SIAM Review, 15(3):665– 670, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boroson:1979:BRE

[Bor79] D. M. Boroson. Book review:

Expected Values for Random Regions of a Circle (T. D. Rogers). SIAM Review, 21 (4):562–564, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1980:PP

[Bor80] P. B. Borwein. A polynomial problem. SIAM Review, 22(4):501, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1990:BRG

[Bor90] Jonathan Borwein. Book review: Generalized Concavity (Mordecai Avriel, Walter E. Diewert, Siegfried Schaible, and Israel Zang). SIAM Review, 32(4):689– 690, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:2006:FRO

[Bor06] Jonathan M. Borwein. Featured review: Oxford Users’ Guide to Mathematics. Edited by Eberhard Zeidler. Oxford University Press, Oxford, 2004. $59.50. xxii + 1285 pp., softcover. ISBN 0-19-850763-1. SIAM Review, 48 (3):585–594, September 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/doi/abs/
                                        10.1137/SIREAD000048000003000585000001;
                                         http://www.jstor.org/
                                        stable/20453842.

Borwein:2009:PCM

[Bor09] Jonathan M. Borwein. The Princeton companion to mathematics [book review of 2467561]. SIAM Review, 51(4):790–794, 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bosch:1973:BRD

[Bos73a] A. J. Bosch. Book review: A Determinant Identity (K. W. Schmidt). SIAM Review, 15 (4):792–793, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bosch:1973:BRS

[Bos73b] A. J. Bosch. Book review: On a Switching Circuit (Herbert A. Cohen). SIAM Review, 15 (4):789–790, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryant:1973:BRE

[BOS73c] Peter Bryant, Patrick E.

O’Neil, and F. W. Steutel. Book review: Expected Number of Stops for an Elevator (D. J. Newman). SIAM Review, 15(4): 793–796, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bosch:1987:NFS

[Bos87] A. J. Bosch. Note on the factorization of a square matrix into two Hermitian or symmetric matrices. SIAM Review, 29(3):463– 468, September 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bos:1995:BRC

[Bos95] Len Bos. Book review: Computer Aided Geometric Design (Josef Hoschek and Dieter Lasser). SIAM Review, 37(3):460–461, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bosley:1996:TNV

[Bos96] D. L. Bosley. A technique for the numerical verification of asymptotic expansions. SIAM Review, 38(1):

128–135, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/p128.htm.

Bosch:1999:IPC

[Bos99]  Robert A. Bosch.  In-

teger programming and Conway’s game of Life. SIAM Review, 41(3):594– 604, September 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33825.

Bottema:1972:EP

[Bot72] O. Bottema. Equilibrium points. SIAM Review, 14 (3):493, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bottema:1974:SP

[Bot74] O. Bottema. A stability problem. SIAM Review, 16 (4):545, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1965:BRR

[Bou65a] C. J. Bouwkamp. Book review: A Resistance Problem, by Ron L. Graham. SIAM Review, 7(2):286– 290, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1965:BRG

[Bou65b] C. J. Bouwkamp. Book review: Gravitational Attraction (Morton L. Slater). SIAM Review, 7(4):562– 564, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1966:BRR

[Bou66a] C. J. Bouwkamp. Book review: Resistance of a Ladder Network (William D. Fryer). SIAM Review, 8(1): 111–112, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1966:BRS

[Bou66b] C. J. Bouwkamp. Book review: Solution to an Integral Equation (A. J. Strecok). SIAM Review, 8(3): 393–395, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1967:IE

[Bou67a] C. J. Bouwkamp. An integral equation. SIAM Review, 9(4):748, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1967:NEP

[Bou67b] C. J. Bouwkamp. Nonlinear eigenvalue problem. SIAM Review, 9(1):120, ???? 1967.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1968:NEP

[Bou68] C. J. Bouwkamp. A nonlinear eigenvalue problem. SIAM Review, 10(1):114– 115, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1971:DI

[Bou71] C. J. Bouwkamp. A double integral. SIAM Review, 13(4):570, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1976:BRG

[Bou76a] C. J. Bouwkamp. Book review: Gravitational Attraction (H. R. Aggarwal). SIAM Review, 18(1):123– 125, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1976:GA

[Bou76b] C. J. Bouwkamp. Gravitational attraction. SIAM Review, 18(1):123–126, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwer:1977:BRA

[Bou77]  Izak Z. Bouwer.  Book

review: Algebraic Graph Theory (Norman Biggs).

SIAM Review, 19(3):570– 571, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouwkamp:1986:BRC

[Bou86] C. J. Bouwkamp. Book review: A Conjectured Definite Integral (A. H. Nuttall). SIAM Review, 28 (4):568–569, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouchitte:1991:BRD

[Bou91a] Guy Bouchitte. Book review:

Direct Methods in the Calculus of Variations (B. Dacorogna). SIAM Review, 33 (2):299–300, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boulware:1991:BRA

[Bou91b] David G. Boulware. Book review: Aspects of Quantum Field Theory in Curved Space-Time (S. A. Fulling). SIAM Review, 33(3):508, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boute:2021:PRM

[Bou21] Raymond Boute. A precise and reliable multivariable chain rule. SIAM Review, 63 (4):854–864, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:1970:BRI

[Boy70]  David W. Boyd.  Book

review: An Integral Inequality (D. J. Newman). SIAM Review, 12(4):590– 591, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:1978:IEF

[Boy78] A. V. Boyd. An integral of an elliptic function. SIAM Review, 20(3):593, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:1980:DI

[Boy80] A. V. Boyd. A determinant and an identity. SIAM Review, 22(1):97, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boynton:1982:ZD

[Boy82] Nancy J. Boynton. Zeros of a determinant. SIAM Review, 24(2):224, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boynton:1983:ZD

[Boy83] N. J. Boynton. Zeros of a determinant. SIAM Review, 25 (2):271–273, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:1988:BRS

[Boy88] John P. Boyd. Book review:

Spectral Methods in Fluid Dynamics (C. Canuto, M. Y. Hussaini, A. Quarteroni, and T. A. Zang). SIAM Review, 30(4):666–668, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:1990:BRI

[Boy90a] John P. Boyd. Book review: An Introduction to the Numerical Analysis of Spectral Methods (Bertrand Mercier). SIAM Review, 32 (3):489–490, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:1990:INA

[Boy90b]  John P. Boyd.  An in-

troduction to the numerical analysis of spectral methods (Bertrand Mercier). SIAM Review, 32(3):489– ??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyle:1997:BRI

[Boy97] Mike Boyle. Book review: An Introduction to Symbolic Dynamics and Coding, by Douglas Lind and Brian Marcus. SIAM Review, 39(2):

334–335, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:2005:HLB

[Boy05]  John P. Boyd.  Hyper-

asymptotics and the linear boundary layer problem: Why asymptotic series diverge. SIAM Review, 47 (3):553–575, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44436.

Boyd:2008:BFC

[Boy08] John P. Boyd. The Blasius function: Computations before computers, the value of tricks, undergraduate projects, and open research problems. SIAM Review, 50 (4):791–804, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boyd:2013:FZU

[Boy13]  John P. Boyd.  Finding

the zeros of a univariate equation: Proxy rootfinders, Chebyshev interpolation, and the companion matrix. SIAM Review, 55(2): 375–396, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balas:1976:SPS

[BP76a] Egon Balas and Manfred W.

Padberg. Set partitioning: a survey. SIAM Review, 18(4): 710–760, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnhill:1976:ISA

[BP76b] R. E. Barnhill and C. C.

Poepplemeier. Interpolation scheme to arbitrary data for computer-aided geometric design. SIAM Review, 18 (4):795, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blows:1994:BLC

[BP94] T. R. Blows and L. M. Perko.

Bifurcation of limit cycles from centers and separatrix cycles of planar analytic systems. SIAM Review, 36 (3):341–376, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/363-1.
htm.

Brown:1995:MRC

[BP95] Philip J. Brown and S. Panchapakesan. Measurement, regression, and calibration. SIAM Review, 37(4):635– ??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bini:1996:PMC

[BPE96] Dario Bini, Victor Pan, and Wayne Eberly. Polynomial and matrix computations volume 1: Fundamental algorithms. SIAM Review, 38(1): 161–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butkovskiy:1995:CDP

[BPJ95] A. G. Butkovskiy, L. M. Pustyl’nikov, and Alan Jeffrey. Characteristics of distributed parameter systems. SIAM Review, 37(3):468– ??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Breach:1984:BRB

[BPK84] D. R. Breach, H. Prodinger, and D. E. Knuth. Book review: A Binomial Summation (Gengzhe Chang and Zun Shan). SIAM Review, 26 (1):122–124, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boehm:1995:GCG

[BPP95] W. Boehm, H. Prautzsch, and Jorg Peters. Geometric concepts for geometric design. SIAM Review, 37(3): 473–??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bienfait:1997:SUS

[BPV97a] J. M. Bienfait, P. Prea, and Y. Vaxes. Sequences with unique sums and differences. SIAM Review, 39(1): 138–140, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bienfait:1997:SSU

[BPV97b] Jean-Marie Bienfait, Pascal Préa, and Yann Vaxes. Solutions: Sequences with unique sums and differences (problem 96-3) (Kim Chuen Ng). SIAM Review, 39(1): 138–140, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bharucha-Reid:1962:BRS

[BR62] A. T. Bharucha-Reid. Book review: Statistical Inference for Markov Processes (P. Billingsley). SIAM Review, 4 (3):262–263, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bharucha-Reid:1965:BRT

[BR65] A. T. Bharucha-Reid. Book review: The Theory of Branching Processes (Theodore E. Harris). SIAM Review, 7 (2):302–303, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bharucha-Reid:1967:BRS

[BR67] A. T. Bharucha-Reid. Book review: Stochastic Processes: Basic Theory and Applications (N. U. Prabhu). SIAM Review, 9(1):142, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bharucha-Reid:1969:BRM

[BR69] A. T. Bharucha-Reid. Book review: Markov Chains with Stationary Transition Probabilities (Kai Lai Chung). SIAM Review, 11(3):416, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balinski:1974:AP

[BR74] M. L. Balinski and Andrew Russakoff. On the assignment polytope. SIAM Review, 16 (4):516–525, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bharucha-Reid:1976:BRS

[BR76] A. T. Bharucha-Reid. Book review: Stochastic Differential Equations: Theory and Applications (Ludwig Arnold). SIAM Review, 18 (1):149, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bharucha-Reid:1979:BRI

[BR79] A. T. Bharucha-Reid. Book review: Innovation Processes (Y. A. Rozanov). SIAM Review, 21(4):579, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bharucha-Reid:1985:BRR

[BR85] A. T. Bharucha-Reid. Book review: Random Linear Operators (A. V. Skorohod). SIAM Review, 27(3):463– 464, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bertram:1991:FPI

[BR91] B. Bertram and O. G. Ruehr.

A finite part integral. SIAM Review, 33(4):656, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bertram:1994:TIE

[BR94] Barbara S. Bertram and Otto G. Ruehr. Two integrals evaluated by zeta functions. SIAM Review, 36 (1):108, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balinski:1997:SMG

[BR97] Michel Balinski and Guillaume Ratier. Of stable marriages and graphs, and strategy and polytopes.

SIAM Review, 39(4):575– 604, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29451.

Bou-Rabee:2018:SSR

[BR18] Nawaf Bou-Rabee. SPEC-TRWM: Spectral random walk method for the numerical solution of stochastic partial differential equations. SIAM Review, 60(2): 386–406, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braver:1963:BRO

[Bra63]  Fred Braver.  Book re-

view: Ordinary Differential Equations (L. S. Pontryagin). SIAM Review, 5(3): 289, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brauer:1965:BRO

[Bra65]  Fred Brauer.  Book re-

view: Ordinary Differential Equations (Philip Hartman). SIAM Review, 7(4): 576, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brands:1966:BRP

[Bra66] J. J. M. Brands. Book review: A Property of Real Solutions to Bessel’s Equation (G. W. Veltkamp). SIAM Review, 8 (2):236–237, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braddock:1976:EGR

[Bra76a] R. D. Braddock. Extensions to grid refinement technique. SIAM Review, 18 (4):796, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bradley:1976:MNA

[Bra76b] J. H. S. Bradley. Maximizing number of acceptable eigenfunctions. SIAM Review, 18 (4):796, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braun:1981:MRV

[Bra81] Martin Braun. Mathematical remarks on the Van Allen radiation belt: a survey of old and new results. SIAM Review, 23(1): 61–93, January 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brauer:1982:BRM

[Bra82] Fred Brauer. Book review:

Management and Analysis of Biological Populations (Bean-San Goh). SIAM Review, 24 (1):89–90, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bradshaw:1985:BRM

[Bra85a] Peter Bradshaw. Book review: The Mathematical Theory of Turbulence (M. M. Stanisić). SIAM Review, 27 (4):590–592, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brauer:1985:BRP

[Bra85b] Fred Brauer. Book review: Population Biology (Simon A. Levin, ed.). SIAM Review, 27(3):464–465, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braunfeld:1991:BRM

[Bra91a] Peter Braunfeld. Book review: The Mathematics of Great Amateurs (Julian Lowell Coolidge). SIAM Review, 33(4):663–664, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braunfeld:1991:MGA

[Bra91b] Peter Braunfeld. The mathematics of great amateurs (Julian Lowell Coolidge). SIAM Review, 33(4):663– ??, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braden:1994:HWM

[Bra94a] B. Braden. Handbook of writing for the mathematical sciences (Nicholas J Higham). SIAM Review, 36 (2):291–??, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braden:1994:BRH

[Bra94b]  Bart Braden.  Book re-

view: Handbook of Writing for the Mathematical Sciences (Nicholas J. Higham). SIAM Review, 36(2):291– 292, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brakhage:1996:BRC

[Bra96] Helmut Brakhage. Book review: Conjugate Gradient Type Methods for Ill-Posed Problems (Martin Hanke). SIAM Review, 38(4):682– 683, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braverman:2017:IIC

[Bra17]  Mark Braverman.  Inter-

active information complexity. SIAM Review, 59(4): 803–846, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1961:EMM

[Bre61] J. L. Brenner. Expanded matrices from matrices with complex elements. SIAM Review, 3(2):165–166, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1962:BRLa

[Bre62a] Joel Brenner. Book review: Lectures on Linear Algebra (I. M. Gel’fand). SIAM Review, 4(1):55, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1962:BRLb

[Bre62b] Joel Brenner. Book review: Linear Algebra and Group Theory (V. I. Smirnov; R. A. Silverman, ed.). SIAM Review, 4(2):160, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1963:CR

[Bre63] Joel Brenner. On commutative rotations. SIAM Review, 5(2):156, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1964:BRI

[Bre64a] J. L. Brenner. Book review: Introduction to Linear Algebra (Frank M. Stewart. D. Van Nostrand). SIAM Review, 6(2):191, ???? 1964.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1964:PCI

[Bre64b] Joel L. Brenner. A pair of combinatorial identities. SIAM Review, 6(2):177, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1967:CI

[Bre67] Joel L. Brenner. Convergence of an iteration. SIAM Review, 9(3):590, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1968:BRS

[Bre68] J. L. Brenner. Book review: Stability Theorems for Linear Motions. With an Introduction to Liapunov’s Direct Method (Siegfried H. Lehnigk). SIAM Review, 10 (4):459–460, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Breach:1970:BRC

[Bre70] D. R. Breach. Book review: A Combinatorial Problem (Melda Hayes). SIAM Review, 12(2):292–297, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1971:I

[Bre71] Joel Brenner. An inequality. SIAM Review, 13(2): 247, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Breach:1974:BRD

[Bre74] D. R. Breach. Book review: Determination of a Sequence (Michael A. Narcowich). SIAM Review, 16 (1):97, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Breach:1977:BRT

[Bre77] D. R. Breach. Book review: Three Multiple Summations (L. Carlitz). SIAM Review, 19(3):567–568, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bressoud:1981:BRR

[Bre81] D. M. Bressoud. Book review: The Rogers-Ramanujan Identities (G. E. Andrews). SIAM Review, 23(1):101– 104, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Breach:1989:BRD

[Bre89] D. R. Breach. Book review: Double Sum of Bessel Functions in Computer Music (A. Ronveaux, J. P. Palamin, and P. Palamin). SIAM Review, 31(4):683–684, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brewster:1993:BRI

[Bre93] M. E. Brewster. Book review: An Introduction to Wavelets (Charles K. Chui). SIAM Review, 35(2):312– 313, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bressan:1994:BRM

[Bre94a] Alberto Bressan. Book review: Multivalued Differential Equations (Klaus Deimling and Walter de Gruyter). SIAM Review, 36(1):117– 119, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bressan:1994:MDE

[Bre94b]  Alberto Bressan.  Multi-

valued differential equations (Klaus Deimling and Walter de Gruyter). SIAM Review, 36(1):117–??, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Breach:1996:TCI

[Bre96] D. R. Breach. Two combinatorial identities. SIAM Review, 38(1):148–149, ???? 1996. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Bou-Rabee:2020:SBM

[BRHC20] Nawaf Bou-Rabee and Miranda C. Holmes-Cerfon. Sticky Brownian motion and its numerical solution. SIAM Review, 62(1):164– 195, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bridgland:1963:ESL

[Bri63a] T. F. Bridgland, Jr. Errata: Stability of Linear Signal Transmission Systems. SIAM Review, 5(3):273, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Bri63b].

Bridgland:1963:SLS

[Bri63b] T. F. Bridgland, Jr. Stability of linear signal transmission systems. SIAM Review, 5(1):7–32, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [Bri63a].

Brick:1964:BRO

[Bri64]  Donald B. Brick.  Book

review: Optical Character Recognition (George L. Fischer, Jr., Donald K. Pollock, Bernard Radack and Mary Elizabeth Stevens, eds.). SIAM Review, 6(1): 78–79, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brillinger:1965:PLP

[Bri65]  David R. Brillinger.  A

property of low-pass filters. SIAM Review, 7(1): 65–67, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brige:1996:BRS

[Bri96] John R. Brige. Book review:

Stochastic Programming (Peter Kall and Stein W. Wallace.). SIAM Review, 38(4): 681–682, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bou-Rabee:2008:DIH

[BRMR08] Nawaf M. Bou-Rabee, Jerrold E. Marsden, and Louis A. Romero. Dissipation-induced heteroclinic orbits in Tippe tops. SIAM Review, 50(2): 325–344, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Reprinted from SIAM Journal on Applied Dynamical Systems, vol 3, pg 352-377, 2004.

Brock:1959:OSP

[Bro59a] Paul Brock. Optimum sorting procedure. SIAM Review, 1(1):70–71, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brownell:1959:BRO

[Bro59b] F. H. Brownell. Book review: Ordinary Difference-Differential Equations (Edmond Pinney). SIAM Review, 1(2):182–184, ???? 1959. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brock:1960:BIA

[Bro60a] Paul Brock. On a binomial identity arising from a sorting problem. SIAM Review, 2(1):40, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brock:1960:OSP

[Bro60b] Paul Brock. Optimum sorting procedure. SIAM Review, 2(2):155–156, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bronwell:1961:DWN

[Bro61] Arthur Bronwell. Do we need a forum for research philosophy. SIAM Review, 3(3): 256–259, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brown:1963:DI

[Bro63] J. L. Brown, Jr. A definite integral. SIAM Review, 5(3):274, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brock:1964:PS

[Bro64] Paul Brock. On a permutation sequence. SIAM Review, 6(3):310, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brown:1965:BRP

[Bro65] J. L. Brown, Jr. Book review:

A Pair of Identities (Tom Benedict). SIAM Review, 7 (3):420–422, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brook:1966:MS

[Bro66a] Paul Brook. A multinomial summation. SIAM Review, 8(2):233, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brown:1966:RNS

[Bro66b] David P. Brown. On the rank and nullity of subgraphs. SIAM Review, 8(2): 220–223, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brockett:1976:BRB

[Bro76a] R. W. Brockett. Book review:

Bilinear Control Processes (Ronald R. Mohler). SIAM Review, 18(4):774, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brownell:1976:BRS

[Bro76b] Frank Brownell. Book review: Spectra of Partial Differential Operators (M. Schechter). SIAM Review, 18 (1):140–141, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brogan:1977:BRA

[Bro77] William L. Brogan. Book review: Applied Optimal Estimation (Arthur Gels, ed.). SIAM Review, 19(1):172– 175, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Broyden:1982:MP

[Bro82] C. G. Broyden. A matrix problem. SIAM Review, 24 (2):223, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brownell:1983:BRL

[Bro83a] F. H. Brownell. Book review: The Logic of Quantum Mechanics (E. G. Beltrametti and Gianni Cassinelli). SIAM Review, 25(3):429– 431, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Broyden:1983:BRM

[Bro83b] C. G. Broyden. Book review: A Matrix Problem (C.

G. Broyden). SIAM Review, 25(3):405, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brownell:1990:BRS

[Bro90] F. H. Brownell. Book review:

Solvable Models in Quantum Mechanics (S. Albeverio, F. Gesztesy, R. Hoegh-Krohn, and H. Holden). SIAM Review, 32(1):160– 167, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brown:1993:BRJ

[Bro93] Bryson Brown. Book review: A Journey into Gravity and Spacetime (John A. Wheeler). SIAM Review, 35 (2):314–315, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burton:1988:CSS

[BRP88] R. M. Burton, Jr., K. Ravishankar, and E. D. Palmes. Convergence to steady state in a diffusional sampler. SIAM Review, 30(3):482– 486, September 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brugia:1965:NMP

[Bru65] Odoardo Brugia. A noniterative method for the partial fraction expansion of a rational function with high order poles. SIAM Review, 7(3): 381–387, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brualdi:1972:BRF

[Bru72] R. A. Brualdi. Book review: On the Foundations of Combinatorial Theory: Combinatorial Geometries (Henry H. Crapo and Gian Carlo Rota). SIAM Review, 14(1): 184–185, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brualdi:1973:BRT

[Bru73] R. A. Brualdi. Book review: Transversal Theory (L. Mirsky). SIAM Review, 15 (2):393–395, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brumfield:1983:BRC

[Bru83] Jeffrey A. Brumfield. Book review: Computer Systems Performance Modeling (Charles H. Sauer and K. Mani Chandy). SIAM Review, 25(1):124–125, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brumfield:1986:BRC

[Bru86] Jeffrey A. Brumfield. Book review: Computer Performance Modeling Handbook (Stephen S. Lavenberg, ed.). SIAM Review, 28(2):249– 250, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brucks:1995:BRO

[Bru95] K. M. Brucks. Book review: One-dimensional Dynamics (W. de Melo and S. van Strien). SIAM Review, 37 (3):461–462, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryant:1990:BRE

[Bry90a] Peter J. Bryant. Book review: From Equilibrium to Chaos: Practical Bifurcation and Stability Analysis (Rudiger Seydel). SIAM Review, 32 (3):490–491, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryant:1990:ECP

[Bry90b] Peter J. Bryant. From equilibrium to chaos: Practical bifurcation and stability analysis (Rudiger Seydel). SIAM Review, 32(3): 490–??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryant:1991:BRC

[Bry91a] Peter J. Bryant. Book review:

Chaotic Dynamics: an Introduction (Gregory Baker and Jerry Gollub). SIAM Review, 33(4):666–668, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryant:1991:CDI

[Bry91b] Peter J. Bryant. Chaotic dynamics: An introduction (Gregory Baker and Jerry Gollub). SIAM Review, 33 (4):666–??, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryant:1993:BRW

[Bry93]  Peter J. Bryant.  Book

review: Weak Chaos and Quasi-Regular Patterns (G. M. Zaslavsky, R. Z. Sagdeev, D. A. Usikov, and A. A. Chernikov). SIAM Review, 35(1):145–146, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryant:1994:BRN

[Bry94]  Peter J. Bryant.  Book

review: Nonlinearities in Action (A. V. Gaponov-Grekhov and M. I. Rabinovich). SIAM Review, 36 (2):292–293, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bryc:1996:BRL

[Bry96] Wlodzimierz Bryc. Book review: Limit Theorems of Probability Theory: Sequences of Independent Random Variables (Valentin V. Petrov). SIAM Review, 38 (3):527, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brezinski:2018:SST

[BRZS18] Claude Brezinski, Michela Redivo-Zaglia, and Yousef Saad. Shanks sequence transformations and Anderson acceleration. SIAM Review, 60 (3):646–669, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1962:PUM

[BS62] J. L. Brenner and F. T.

Smith. On a property of a unitary matrix. SIAM Review, 4(4):395, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berman:1963:ARL

[BS63] Gerald Berman and R. G.

Stanton. The asymptotes of the root locus. SIAM Review, 5(3):209–218, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boas:1966:IIR

[BS66] R. P. Boas, Jr. and Lowell Schoenfeld. Indefinite integration by residues. SIAM Review, 8(2):173–183, ????

1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barrodale:1967:BRL

[BS67] Ian Barrodale and R. R. Singleton. Book review: A Least Absolute Fit (Derald Walling). SIAM Review, 9 (1):122–124, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Byrnes:1970:IWF

[BS70] J. S. Byrnes and D. A. Swick. Instant Walsh functions. SIAM Review, 12 (1):131, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barnett:1977:RAC

[BS77] S. Barnett and D. D. Siljak. Routh’s algorithm: a centennial survey. SIAM Review, 19 (3):472–489, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brink-Spalink:1980:BRC

[BS80] Jan Brink-Spalink. Book review: Credibility Functions (Y. P. Sabharwal and J. Kumar). SIAM Review, 22(2): 234–236, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bosch:1986:BRI

[BS86] A. J. Bosch and F. W. Steutel. Book review: An Identity (M. S. Klamkin and O. G. Ruehr). SIAM Review, 28 (2):243–244, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bailey:1991:FFT

[BS91] David H. Bailey and Paul N.

Swarztrauber. The fractional Fourier transform and applications. SIAM Review, 33 (3):389–404, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butzer:1992:STN

[BS92] P. L. Butzer and R. L.

Stens. Sampling theory for not necessarily bandlimited functions: a historical overview. SIAM Review, 34 (1):40–53, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bullett:1993:RSP

[BS93] Shaun Bullett and Jaroslav Stark. Renormalizing the simple pendulum. SIAM Review, 35(4):631–640, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Babuska:1994:PHP

[BS94a] I. Babuska and M. Suri. The P and H-P versions of the finite-element method, basic principles and properties. SIAM Review, 36 (4):578–632, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Babuska:1994:VFE

[BS94b] Ivo Babuška and Manil Suri. The p and h–p versions of the finite element method, basic principles and properties. SIAM Review, 36(4):578– 632, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/23906.htm.

Bonnans:1998:OPP

[BS98] J. Frédéric Bonnans and Alexander Shapiro. Optimization problems with perturbations: a guided tour. SIAM Review, 40(2):228– 264, June 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/30264.

Babuska:2000:PEF

[BS00] Ivo M. Babuška and Stefan A. Sauter. Is the Pollution effect of the FEM avoidable for the Helmholtz equation considering high wave numbers? SIAM Review, 42 (3):451–484, September 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36779. Reprinted

from SIAM Journal Numerical Analysis, vol 32, pg 2392– 2423, 1997.

Borzi:2009:MMP

[BS09a] Alfio Borzì and Volker Schulz. Multigrid methods for PDE optimization. SIAM Review, 51(2):361– 395, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buttazzo:2009:MTM

[BS09b] Giuseppe Buttazzo and Filippo Santambrogio. A mass transportation model for the optimal planning of an urban region. SIAM Review, 51(3): 593–610, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bosch:1986:I

[BSD^{+}86] A. J. Bosch, F. W. Steutel, S. L. Damjanovic, A. A. Jagers, W. B. Jordan, and G. Parker. An identity. SIAM Review, 28(2):243– 244, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1982:SII

[BSFW82] J. M. Borwein, G. P. H. Styan, L. V. Foster, and H. Wolkowicz. Some inequalities involving statistical expressions. SIAM Review, 24 (3):340–342, July 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brenner:1995:MTF

[BSP95] Susanne C. Brenner, L. Ridgway Scott, and Joseph E. Pasciak. The mathematical theory of finite element methods. SIAM Review, 37(3): 472–??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Braga:2005:CPB

[BSS05] Gastão A. Braga, Rémy Sanchis, and Tiago A. Schieber. Critical percolation on a Bethe lattice revisited. SIAM Review, 47(2):349–365, June 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43368.

Batson:2014:TRS

[BSS14] Joshua Batson, Daniel A. Spielman, and Nikhil Srivastava. Twice-Ramanujan sparsifiers. SIAM Review, 56 (2):315–334, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1982:SSI

[BSW82] J. M. Borwein, G. P. H.

Styan, and H. Wolkowicz. Solutions: Some inequalities involving statistical expressions (L. V. Foster). SIAM Review, 24(3):340–342, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balinski:1969:DTL

[BT69] M. L. Balinski and A. W.

Tucker. Duality theory of linear programs: a constructive approach with applications. SIAM Review, 11(3): 347–377, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Blackmore:1985:SIM

[BT85] Denis Blackmore and Lu Ting.

Surface integral of its mean curvature vector. SIAM Review, 27(4):569–572, December 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berrut:2004:BLI

[BT04] Jean-Paul P. Berrut and Lloyd N. Trefethen. Barycentric Lagrange interpolation. SIAM Review, 46(3):501– 517, September 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/41771.

Betcke:2005:RMP

[BT05] Timo Betcke and Lloyd N. Trefethen. Reviving the method of particular solutions. SIAM Review, 47 (3):469–491, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43733.

Bernoff:2013:NAM

[BT13] Andrew J. Bernoff and Chad M. Topaz. Nonlocal aggregation models: a primer of swarm equilibria. SIAM Review, 55(4): 709–747, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beckermann:2019:BSV

[BT19] Bernhard Beckermann and Alex Townsend. Bounds on the singular values of matrices with displacement structure. SIAM Review, 61(2): 319–344, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Brogliato:2020:DSC

[BT20] Bernard Brogliato and Aneel Tanwani. Dynamical systems coupled with monotone set-valued operators:

Formalisms, applications, well-posedness, and stability. SIAM Review, 62(1): 3–129, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ben-Tal:2000:FMD

[BTKNZ00] A. Ben-Tal, M. Kovara, A. Nemirovski, and J. Zowe. Free material design via semidefinite programming: The multiload case with contact conditions. SIAM Review, 42(4):695–715, December 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37208.

Buck:1969:BRF

[Buc69]  R. C. Buck.  Book re-

view: Functions of a Complex Variable, Constructive Theory (V. I. Smirnov and N. S. Lebedev). SIAM Review, 11(2):290–291, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buckmaster:1985:BRA

[Buc85]  John Buckmaster.  Book

review: Asymptotic Treatment of Chemically Reacting Systems (Ashwani K. Kapila). SIAM Review, 27 (1):94–95, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buck:1991:BRM

[Buc91a] R. Creighton Buck. Book review: Modern Cryptology: a Tutorial (Gilles Brassard). SIAM Review, 33(3): 487–489, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buck:1991:MCT

[Buc91b] R. Creighton Buck. Modern cryptology: a tutorial (Gilles Brassard). SIAM Review, 33 (3):487–??, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bueckner:1966:BRP

[Bue66] Hans F. Bueckner. Book review: Positive Solutions of Operator Equations (M. A. Krasnoselskii). SIAM Review, 8(1):122–123, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bueckner:1971:BRC

[Bue71] Hans F. Bueckner. Book review: Crack Problems in the Classical Theory of Elasticity (I. N. Sneddon and M. Lowengrub). SIAM Review, 13(2):241–242, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bullock:1967:UTD

[Bul67]  Robert M. Bullock.  A

uniqueness theorem for a delay-differential system. SIAM Review, 9(4):737–740, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buneman:1983:CSU

[Bun83a] O. Buneman. Computer simulation using particles — Hockney, R. W., Eastwood, J. W. SIAM Review, 25(3): 425–426, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buneman:1983:BRC

[Bun83b] Oscar Buneman. Book review: Computer Simulation Using Particles (R. W. Hockney and J. W. Eastwood). SIAM Review, 25(3): 425–426, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burt:1963:ER

[Bur63] Oscar R. Burt. Economic replacement. SIAM Review, 5 (3):203–208, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [Bur64].

Burt:1964:EER

[Bur64] Oscar R. Burt. Errata: Economic Replacement. SIAM Review, 6(1):59, ???? 1964.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Bur63].

Burlakoff:1978:MSL

[Bur78a] M. Burlakoff. Mathematical science library written in a high-order language. SIAM Review, 20(3):620, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burrows:1978:FDF

[Bur78b] J. W. Burrows. Failure detector false alarm rates. SIAM Review, 20(3):620– 621, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burr:1979:BRT

[Bur79] Stefan A. Burr. Book review: The Theory of Partitions (George E. Andrews). SIAM Review, 21(3):430– 432, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burnside:1981:BRN

[Bur81a] R. R. Burnside. Book review: A Nonoscillation Result (D. K. Ross). SIAM Review, 23 (2):257–258, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burr:1981:BRL

[Bur81b] Stefan A. Burr. Book review:

Location on Networks: Theory and Algorithms (Gabriel Y. Handler and Pitu B. Mirchandani). SIAM Review, 23 (1):123–124, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burnashev:1982:BTH

[Bur82] M. Burnashev. “bad” triangles in Hilbert space. SIAM Review, 24(4):477– 478, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burton:1983:SSV

[Bur83] T. A. Burton. Structure of solutions of Volterra equations. SIAM Review, 25(3): 343–364, July 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burton:1989:AC

[Bur89] T. A. Burton. Author’s comment. SIAM Review, 31(3): 517–519, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burton:1993:BRC

[Bur93] T. A. Burton. Book review: A Course on Integral Equations (A. C. Pipkin). SIAM Review, 35 (4):655–656, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burckel:1994:BRH

[Bur94a] R. B. Burckel. Book review: A History of Complex Dynamics from Schröder to Fatou and Julia (Daniel S. Alexander). SIAM Review, 36(4):663–664, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burdzy:1994:BRS

[Bur94b]  Krzysztof Burdzy.  Book

review: The Self-Avoiding Walk (Neal Madras and Gordon Slade). SIAM Review, 36(1):119, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burdzy:1994:SAW

[Bur94c] Krzysztof Burdzy. The selfavoiding walk (Neal Madras and Gordon Slade). SIAM Review, 36(1):119–??, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buschman:1963:IIG

[Bus63] R. G. Buschman. An inversion integral for a general Legendre transformation. SIAM Review, 5(3): 232–233, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buschman:1964:CEG

[Bus64] R. G. Buschman. Convolution equations with generalized Laguerre polynomial kernels. SIAM Review, 6(2): 166–167, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bushaw:1969:SLP

[Bus69] D. Bushaw. Stabilities of Lyapunov and Poisson types. SIAM Review, 11(2):214– 225, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bushaw:1970:BRD

[Bus70] D. Bushaw. Book review:

Discontinuous and Optimal Control (Irmgard Flügge– Lotz). SIAM Review, 12(2): 317–318, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bushard:1976:CUB

[Bus76a] L. B. Bushard. Computationally useful bounds for singular nonlinear 2nd order boundary-value problems. SIAM Review, 18 (4):796, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bushard:1976:EEM

[Bus76b] L. B. Bushard. On the extreme eigenvalues of an n × n matrix. SIAM Review, 18(4): 762–763, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buschman:1986:BRM

[Bus86] R. G. Buschman. Book review: A Moment Generating Function (E. Q. George and C. C. Rousseau). SIAM Review, 28(4):575– 576, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buschman:1995:BRT

[Bus95] R. G. Buschman. Book review: Two Integrals Evaluated by Zeta Functions (Barbra S. Bertram and Otto G. Ruehr). SIAM Review, 37(1): 110–112, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butts:1961:BRM

[But61] Layton E. Butts. Book review: A Minimum Switching Network (Raphael Miller). SIAM Review, 3(3):263– 266, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butz:1974:BRFa

[But74a] Edward Butz. Book review:

Foundations of Mathematical Biology, Vol. 1: Subcellular Systems (Robert Rosen, ed.). SIAM Review, 16(2): 274–276, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butz:1974:BRFb

[But74b] Edward Butz. Book review:

Foundations of Mathematical Biology. Vol. II: Cellular Systems (Robert Rosen, ed.). SIAM Review, 16(3): 401–403, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butz:1974:BRFc

[But74c] Edward G. Butz. Book review: Foundations of Mathematical Biology. Vol. III Supercellular Systems (Robert Rosen, ed.). SIAM Review, 16(4):558–561, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butz:1978:BRB

[But78] Edward Butz. Book review: Biochemical Systems Analysis. A Study of Function and Design in Molecular Biology (Michael A. Savageau). SIAM Review, 20(3): 613–615, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butz:1979:BRI

[But79] Edward Butz. Book review: An Introduction to Mathematical Modelling (Edward A. Bender). SIAM Review, 21 (3):402–404, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butcher:1989:BRT

[But89] J. C. Butcher. Book review: Theory of Difference Equations: Numerical Methods and Applications (V. Lakshmikantham and D. Trigiante). SIAM Review, 31(4): 689–690, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Butcher:1992:BRD

[But92] J. C. Butcher. Book review: Difference Equations from Differential Equations (W. J. Lick). SIAM Review, 34(1):140, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Buzbee:1978:LSS

[Buz78] B. L. Buzbee. Large-scale scientific computing at Los Alamos. SIAM Review, 20 (3):621, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Binding:1996:ETP

[BV96] Paul Binding and Hans Volkmer. Eigencurves for twoparameter Sturm–Liouville equations. SIAM Review, 38 (1):27–48, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/26508.htm.

Barclay:1978:DPM

[BvdD78] H. Barclay and P. van den Driesschei. Deterministic population models and stability. SIAM Review, 20(2): 389–393, April 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Besseling:1995:MMI

[BvdGH95] J. F. Besseling, E. van der Giessen, and K. S. Havner. Mathematical modelling of inelastic deformation. SIAM Review, 37(3):467–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bouten:2009:DIQ

[BvJ09] Luc Bouten, Ramon van Handel, and Matthew R. James. A discrete invitation to quantum filtering and feedback control. SIAM Review, 51(2):239– 316, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Berterman:1965:BRU

[BW65] John Berterman and Alan Weinstein. Book review: An Urn Problem (Daniel H. Wagner). SIAM Review, 7 (1):136–138, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Bordelon:1973:BRE

[BW73] D. J. Bordelon and A. S. Williamson. Book review: Equilibrium Points (O. Bottema). SIAM Review, 15(3): 663–665, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Barker:1976:ITO

[BW76] W. H. Barker and D. H. Wagner. Information theory and optimal detection search. SIAM Review, 18 (4):794, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burnside:1996:CTI

[BW96a] R. R. Burnside and W. C. Waterhouse. A conditional trace inequality. SIAM Review, 38(2):325–326, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burnside:1996:BRC

[BW96b] R. R. Burnside and William C. Waterhouse. Book review:

A Conditional Trace Inequality (Paul A. Roediger). SIAM Review, 38(2): 324–326, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Beck:2011:UGI

[BW11] Margaret Beck and C. Eugene Wayne. Using global invariant manifolds to understand metastability in the Burgers equation with small viscosity. SIAM Review, 53 (1):129–153, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i1/p129_s1.

Berlekamp:1996:MGC

[BWYY96] Elwyn Berlekamp, David Wolfe, Zhiping You, and James A. Yorke. Mathematical go chilling gets the last point. SIAM Review, 38(3): 543–??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Balinski:1978:JMA

[BY78] M. L. Balinski and H. P.

Young. The Jefferson method of apportionment. SIAM Review, 20(2):278– 284, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1998:ASK

[BY98] J. Boersma and S. B. Yakubovich. The asymptotic sum of a Kapteyn series. SIAM Review, 40(4):986– 990, November 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Byrd:1965:BRA

[Byr65] Paul F. Byrd. Book review: An Approximation of a Definite Integral (David Rothman). SIAM Review, 7(1): 131–132, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Burq:2005:BBM

[BZ05] Nicolas Burq and Maciej Zworski. Bouncing ball modes and quantum chaos. SIAM Review, 47(1):43– 49, March 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/42924.

Comstock:1975:SRS

[CA75] C. Comstock and D. A. Archer. Some results with singular Sturm–Liouville problems. SIAM Review, 17 (2):377, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cabay:1974:BRS

[Cab74] Stan Cabay. Book review:

Spline Analysis (Martin H. Schultz). SIAM Review, 16 (1):105–106, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cabay:1975:BRA

[Cab75]  Stan Cabay.  Book re-

view: An Analysis of the Finite Element Method (Gilbert Strang and George J. Fix). SIAM Review, 17(3):577– 578, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Caflisch:1992:BRT

[Caf92] Russel Caflisch. Book review: Two Phase Flows and Waves (Daniel D. Joseph and David G Schaeffer, eds.). SIAM Review, 34(1):138– 140, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Caflisch:1994:BRV

[Caf94] Russel E. Caflisch. Book review: Vortex Dynamics (P. G. Saffman). SIAM Review, 36(2):293–297, June 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cahit:1977:CMV

[Cah77] I. Cahit. A conjectured minimum valuation tree.

SIAM Review, 19(3):564– 565, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cortez:2023:ICC

[CAiJS23] Mark Jayson Cortez, Alan Eric Akil, Kre imir Josi, and Alexander J. Stewart. Incorporating computational challenges into a multidisciplinary course on stochastic processes. SIAM Review, 65(4):1152–1170, November 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Campbell:1977:BRM

[Cam77] L. L. Campbell. Book review: On Measures of Information and Their Characterizations (J. Aczél and Z. Daróczy). SIAM Review, 19 (3):577–579, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Campbell:1991:BRI

[Cam91a] L. L. Campbell. Book review: An Introduction to Digital Signal Processing (John H. Karl). SIAM Review, 33(1): 131–132, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Campbell:1991:BRN

[Cam91b] Stephen L. Campbell. Book review: The Numerical Solution of Differential-Algebraic Systems by Runge– Kutta Methods (Ernst Hairer, Christian Lubich, and Michel Roche). SIAM Review, 33(3): 504–506, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Campbell:1991:NSD

[Cam91c] Stephen L. Campbell. The numerical solution of differentialalgebraic systems by Runge– Kutta methods (Ernst Hairer, Christian Lubich, and Michel Roche). SIAM Review, 33(3): 504–??, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Campbell:1993:BRM

[Cam93] Stephen L. Campbell. Book review: Methods of Solving Singular Systems of Ordinary Differential Equations (Y. Boyarintsev). SIAM Review, 35(4):656–658, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cantor:1986:MEP

[Can86a]  D. G. Cantor.  A mul-

tivariable extremum problem. SIAM Review, 28(3): 398–399, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cantor:1986:BRM

[Can86b] David G. Cantor. Book review: A multi-Variable Extremum Problem (V. Balakotaiah). SIAM Review, 28(3): 398–399, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Canfield:1991:BRC

[Can91a] E. Rodney Canfield. Book review: Combinatorial Search (Martin Aigner). SIAM Review, 33(1):132–133, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Canfield:1991:CSM

[Can91b] E. Rodney Canfield. Combinatorial search (Martin Aigner). SIAM Review, 33 (1):132–??, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1964:BIA

[Car64] L. Carlitz. A binomial identity arising from a sorting problem. SIAM Review, 6 (1):20–30, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carter:1965:BRP

[Car65a] D. S. Carter. Book review: Perturbation Techniques in Mathematics, Physics and Engineering (Richard Bellman). SIAM Review, 7(3): 433, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carter:1965:BRA

[Car65b] Harvey P. Carter. Book review: Automatic Data Processing (Frederick P. Brooks, Jr., and Kenneth E. Iverson). SIAM Review, 7(1): 148, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1966:BRR

[Car66a] L. Carlitz. Book review:

A Recursion Relation (E. L. Pugh). SIAM Review, 8(4): 539–540, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1966:PI

[Car66b] L. Carlitz. A pair of identities. SIAM Review, 8(2): 233–234, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carr:1966:BRI

[Car66c] John W. Carr III. Book review: Introduction to AL-GOL (R. Baumann, M. Feliciano, F. L. Bauer and K. Samelson). SIAM Review, 8(3):406–407, July 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //www.jstor.org/stable/
                                        2028240.

Carlitz:1967:BI

[Car67a] L. Carlitz. A binomial identity. SIAM Review, 9(2): 229–231, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1967:BRC

[Car67b] L. Carlitz. Book review: A Combinatorial Problem (Joseph Arkin). SIAM Review, 9(1):124, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1967:DS

[Car67c]  L. Carlitz.  A double

sum. SIAM Review, 9(1): 121, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1967:QI

[Car67d] L. Carlitz. A quadruple integral. SIAM Review, 9 (4):749, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carrier:1967:ITA

[Car67e] George F. Carrier. IV. training in applied mathematics research. SIAM Review, 9(2):

347–365, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1968:DS

[Car68a] L. Carlitz. A double sum.

SIAM Review, 10(1):117– 119, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1968:ISS

[Car68b]  L. Carlitz.  An infinite

set of simultaneous equations. SIAM Review, 10 (2):224, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1969:ISS

[Car69]  L. Carlitz.  An infinite

set of simultaneous equations. SIAM Review, 11(4): 624–625, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1970:BRP

[Car70a] L. Carlitz. Book review:

A Positive Integral (Richard Askey). SIAM Review, 12(4): 588–590, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1970:TLE

[Car70b] L. Carlitz. Two Laurent expansions. SIAM Review, 12 (2):288, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlson:1970:IMA

[Car70c] B. C. Carlson. An inequality of mixed arithmetic and geometric means. SIAM Review, 12(2):287–288, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlson:1970:HSS

[Car70d] Bille Chandler Carlson. Hidden symmetries of special functions. SIAM Review, 12(3):332–345, July 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2028552.

Carrier:1970:SPT

[Car70e] G. F. Carrier. Singular perturbation theory and geophysics. SIAM Review, 12(2): 175–193, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1971:BRS

[Car71a] L. Carlitz. Book review: A Stock Market Investment Inequality (M. J. Lempel). SIAM Review, 13(3):404– 406, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1971:BRU

[Car71b] L. Carlitz. Book review: A Unique Solution (D. J. Newman). SIAM Review, 13(4): 574–575, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1971:PI

[Car71c] L. Carlitz. A positive integral. SIAM Review, 13 (2):248, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1971:TLE

[Car71d] L. Carlitz. Two Laurent expansions. SIAM Review, 13 (2):255–256, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1972:BRI

[Car72a] L. Carlitz. Book review:

An Infinite Summation (S. Saran). SIAM Review, 14(3): 502–504, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1972:SNE

[Car72b] L. Carlitz. A set of nonlinear equations. SIAM Review, 14(1):165, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1973:BRB

[Car73a] L. Carlitz. Book review: Biorthogonal Conditions for a Class of Polynomials (W. A. Chai). SIAM Review, 15 (3):670–672, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1973:PTF

[Car73b] L. Carlitz. A pair of theta function identities. SIAM Review, 15(4):788, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1974:SDD

[Car74] L. Carlitz. A system of difference-differential equations. SIAM Review, 16 (1):94, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1975:PSS

[Car75a] L. Carlitz. Permutations, sequences and special functions. SIAM Review, 17(2): 298–322, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1975:SDD

[Car75b] L. Carlitz. A system of difference-differential equations. SIAM Review, 17(1): 171–173, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1976:TMS

[Car76a] L. Carlitz. Three multiple summations. SIAM Review, 18(3):489–490, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlson:1976:BRS

[Car76b] B. C. Carlson. Book review: A Singular Integral Equation (M. L. Glasser). SIAM Review, 18(3):493– 495, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1977:SSO

[Car77] L. Carlitz. A system of second order differential equations. SIAM Review, 19 (4):736, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1980:BRH

[Car80] L. Carlitz. Book review:

A Hyperbolic Series Identity (Chih-Bong Ling). SIAM Review, 22(2):231, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlson:1981:BRN

[Car81] David Carlson. Book review: Nonnegative Matrices in the Mathematical Sciences (Abraham Berman and Robert J. Plemmons). SIAM Review, 23(3):409– 410, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carter:1987:BRE

[Car87] E. M. Carter. Book review: Elementary Multivariate Analysis for the Behavioral Sciences-Applications of Basic Structure. (Clifford E. Lunneborg and Robert D. Abbott). SIAM Review, 29 (1):139, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carroll:1991:VPO

[Car91] C. E. Carroll. Variational problem for overlapping laser pulses. SIAM Review, 33(2): 275–276, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlin:1993:BRT

[Car93] Bradley P. Carlin. Book review: Tools for Statistical Inference: Observed Data and Data Augmentation Methods (Martin A. Tanner). SIAM Review, 35(1): 146–147, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carter:2004:M

[Car04] J. D. Carter. Mathematica 5.0. SIAM Review, 46(3): 564–568, ???? 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carter:2008:FRM

[Car08]  John D. Carter.  Fea-

tured review: Mathematica 6. SIAM Review, 50(1): 149–152, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cassity:1972:BRR

[Cas72a] C. R. Cassity. Book review:

Root of an Octic Equation (H. B. Coonce). SIAM Review, 14(2):372, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Casti:1972:FFI

[Cas72b]  J. L. Casti.  A family

of Fredholm integral equations. SIAM Review, 14(1): 164–165, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cassity:1975:SSO

[Cas75] C. R. Cassity. Solutions of sixth order Runge–Kutta equations. SIAM Review, 17 (2):375, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Casti:1980:LQC

[Cas80] J. Casti. The linear-quadratic control problem: Some recent results and outstanding problems. SIAM Review, 22(4): 459–485, October 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Casti:1982:RDF

[Cas82] John L. Casti. Recent developments and future perspectives in nonlinear system theory. SIAM Review, 24 (3):301–331, July 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Case:1987:OCP

[Cas87a] J. H. Case. An optimal control problem. SIAM Review, 29(1):124–127, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Case:1987:BRO

[Cas87b] James H. Case. Book review: An Optimal Control Problem (J. C. Moorhouse and J. V. Baxley). SIAM Review, 29(1):124–127, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Casti:1988:AC

[Cas88] John L. Casti. Author’s comment. SIAM Review, 30(4):

680–683, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Castro:1994:BRS

[Cas94a] A. Castro. Book review:

Solvability and Bifurcations of Nonlinear Equations (P. Drábek). SIAM Review, 36 (1):119–122, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Castro:1994:SBN

[Cas94b]  A. Castro.  Solvability

and bifurcations of nonlinear equations (P Drabek). SIAM Review, 36(1):119– ??, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cassels:1995:BRC

[Cas95] J. W. S. Cassels. Book review: Catalan’s Conjecture (P. Ribenboim). SIAM Review, 37(1):116–117, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Catinas:2021:HMS

[Cat21] Emil Catinas. How many steps still left to x *? SIAM Review, 63(3):585– 624, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cavendish:1974:NME

[Cav74] J. C. Cavendish. On the norm of a matrix exponential. SIAM Review, 16 (1):94, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cavaretta:1990:BRL

[Cav90] Alfred S. Cavaretta, Jr. Book review: Lectures on Complex Approximation (Dieter Gaier; Renate McLaughlin, trans.). SIAM Review, 32(1): 167–168, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1976:SCI

[CB76] J. K. Cohen and N. Bleistein. Solutions of a class of inverse problems in acoustics and electromagnetics. SIAM Review, 18(4):798, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cortes:2009:NCG

[CB09] Jorge Cortés and Francesco Bullo. Nonsmooth coordination and geometric optimization via distributed dynamical systems. SIAM Review, 51(1):163–189, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cavendish:1975:MAL

[CBG75] J. C. Cavendish, G. Birkhoff, and W. J. Gordon. Multivariate approximation by locally blended univariate interplants. SIAM Review, 17 (2):375, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cushman:1998:GAC

[CBM98] Richard H. Cushman, Larry M.

Bates, and Richard Montgomery. Global aspects of classical integrable systems. SIAM Review, 40(1): 164–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Conn:1976:MTP

[CBS76] A. R. Conn, R. H. Bartels, and J. W. Sinclair. Minimization techniques for piecewise differentiable functions — L_{1} solution to an overdetermined linear-system. SIAM Review, 18(4):799, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Curry:1998:I

[CBW98] R. L. Curry, B. Bertram, and J. E. Wilkins. An inequality. SIAM Review, 40(3): 714–718, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Charalambous:1975:EMS

[CC75] C. Charalambous and A. R. Conn. Efficient method to solve minimax problem directly. SIAM Review, 17 (2):376, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cooper:1976:EAN

[CC76] L. Cooper and M. W. Cooper. Exact algorithm for nonlinear integer programming. SIAM Review, 18 (4):799, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1978:SNT

[CC78] Y. F. Chang and G. Corliss. Solving a nonhomogeneous Tricomi equation with Taylor series. SIAM Review, 20 (3):621, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chui:1990:ECD

[CC90] Charles K. Chui and Guanrong Chen. Estimation, control, and the discrete Kalman filter (Donald E Catlin). SIAM Review, 32(3):493– ??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cantrell:1996:MPP

[CC96] R. S. Cantrell and C. Cosner. Models for predatorprey systems at multiple scales. SIAM Review, 38(2): 256–286, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/27277.htm.

Chan:2023:CGR

[CC23] Eunice Y. S. Chan and Robert M. Corless. Chaos game representation. SIAM Review, 65(1):261–290, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://epubs.siam.org/
doi/10.1137/20M1386438.

Castillo:2008:SAC

[CCA08] Enrique Castillo, Antonio J.

Conejo, and Ernesto Aranda. Sensitivity analysis in calculus of variations. some applications. SIAM Review, 50(2): 294–312, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Castillo:2006:SOD

[CCCM06] Enrique Castillo, Antonio J.

Conejo, Carmen Castillo, and Roberto Mínguez. Solving ordinary differential equations with range conditions. applications. SIAM Review, 48(2): 307–317, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coffman:2002:PPT

[CCG^{+}02] E. G. Coffman, Jr., C. Courcoubetis, M. R. Garey, D. S. Johnson, P. W. Shor, R. R. Weber, and M. Yannakakis. Perfect packing theorems and the average-case behavior of optimal and online bin packing. SIAM Review, 44(1):95–108, March 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39542.

Colton:2000:RDI

[CCM00] David Colton, Joe Coyle, and Peter Monk. Recent developments in inverse acoustic scattering theory. SIAM Review, 42(3):369– 414, September 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v42/i3/p369_s1.

Colton:2001:RDI

[CCM01] David Colton, Joe Coyle, and Peter Monk. Recent developments in inverse acoustic scattering theory. SIAM Review, 42(3): 369–414, September 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36773.

Carrier:1962:AMW

[CCR^{+}62] George F. Carrier, R. Courant, Paul Rosenbloom, C. N. Yang, and H. J. Greenberg. Applied mathematics what is needed in research and education. SIAM Review, 4(4): 297–320, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cullum:1975:BLA

[CD75] J. Cullum and W. E. Donath.

Block Lanczos algorithm for computing Q algebraically largest eigenvalues and a corresponding eigenspace of a large, sparse, real symmetric matrix. SIAM Review, 17(2):377, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carey:1993:GAQ

[CD93] G. F. Carey and H. T. Dinh.

A Grid anomaly in quadrilateral elements. SIAM Review, 35(1):130–134, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chappell:1994:ODN

[CD94] David Chappell and Karen Dury. On the optimal depletion of a nonrenewable natural resource under conditions of increasing marginal extraction costs. SIAM Review, 36(1):102–106, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/25107.
htm.

Carmona:2003:PHS

[CD03] René Carmona and Valdo Durrleman. Pricing and hedging spread options. SIAM Review, 45(4):627– 685, December 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/42479.

Carmon:2020:FOM

[CD20] Yair Carmon and John C. Duchi. First-order methods for nonconvex quadratic minimization. SIAM Review, 62 (2):395–436, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:2001:ADB

[CDS01] Scott Shaobing B. Chen, David L. Donoho, and Michael A. Saunders. Atomic decomposition by basis pursuit. SIAM Review, 43 (1):129–159, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37906.

Chan:1989:FAI

[CE89] Tony F. Chan and Howard C.

Elman. Fourier analysis of iterative methods for elliptic problems. SIAM Review, 31 (1):20–49, March 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cox:2012:OCH

[CEH12] Steven J. Cox, Mark Embree, and Jeffrey M. Hokanson. One can hear the composition of a string: Experiments with an inverse eigenvalue problem. SIAM Review, 54(1): 157–178, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v54/i1/p157_s1.

Celmins:1978:CRO

[Cel78] A. Celmins. Critical review of one-dimensional tube flow equations. SIAM Review, 20 (3):621, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Censor:1981:RAM

[Cen81] Yair Censor. Row-action methods for huge and sparse systems and their applications. SIAM Review, 23(4): 444–466, October 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cerimele:1970:CMC

[Cer70] B. J. Cerimele. Conformal mapping of a cross slit strip. SIAM Review, 12 (1):145, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cerimele:1971:CMC

[Cer71] B. J. Cerimele. Conformal mapping of a cross slit strip. SIAM Review, 13 (2):250, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cercignani:1996:BRT

[Cer96] Carlo Cercignani. Book review: Transport Simulation in Microelectronics (Alfred Kersch and William J. Morokoff). SIAM Review, 38(2): 337–341, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cesari:1975:BRO

[Ces75]  Lamberto Cesari.  Book

review: Optimal Control of Differential and Functional Equations (J. Warga). SIAM Review, 17(3):579– 580, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Candes:2015:PRM

[CESV15] Emmanuel J. Candès, Yonina C. Eldar, Thomas Strohmer, and Vladislav Voroninski. Phase retrieval via matrix completion. SIAM Review, 57(2): 225–251, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Concus:1976:HCS

[CF76] P. Concus and R. Finn.

Height of a capillary surface. SIAM Review, 18(4): 798–799, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cantoni:1980:SDM

[CF80] Vittorio Cantoni and Amalia Ercoli Finzi. Slowest descent to the Moon. SIAM Review, 22 (4):495–499, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chu:1995:ROR

[CFG95] Moody T. Chu, Robert E.

Funderlic, and Gene H. Golub. A rank-one reduction formula and its applications to matrix factorizations. SIAM Review, 37 (4):512–530, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/27264.
htm.

Carmona:2010:MDE

[CFHP10] René Carmona, Max Fehr, Juri Hinz, and Arnaud Porchet. Market design for emission trading schemes. SIAM Review, 52(3):403– 452, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1980:HSI

[CFL^{+}80] L. Carlitz, H. E. Fettis, O. P. Lossers, W. A. Alsalam, B. C. Berndt, R. C. Carson, Z. Drezner, G. Gripenberg, A. A. Jagers, F. Milinazzo, J. Piepers, O. G. Ruehr, P. Wagner, and H. T. Wang. A hyperbolic series identity. SIAM Review, 22 (2):231, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Causey:1961:LAT

[CG61] R. L. Causey and R. T. Gregory. On Lanczos’ algorithm for tridiagonalizing matrices. SIAM Review, 3 (4):322–328, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Conway:1966:SP

[CG66] J. H. Conway and R. K. Guy. Stability of polyhedra. SIAM Review, 8(3):381, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chan:1976:ALM

[CG76] T. F. Chan and R. Glowinski. Augmented Lagrangian methods for solving mildly nonlinear elliptic equations and comparisons with other methods. SIAM Review, 18 (4):796–797, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chung:2024:CML

[CG24] Julianne Chung and Silvia Gazzola. Computational methods for large-scale inverse problems: a survey on hybrid projection methods. SIAM Review, 66(2): 205–284, May 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chan:2007:MMC

[CGO07] Raymond H. Chan, Chen Greif, and Dianne P. O’Leary, editors. Milestones in Matrix Computation: the Selected Works of Gene H. Golub with Commentaries. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920681-3. xi + 565 + 3 pp. LCCN QA188 .G67 2007.

URL http://www.loc.gov/
catdir/enhancements/fy0737/
2007276086-d.html.

Chen:2021:SCL

[CGP21] Yongxin Chen, Tryphon T.

Georgiou, and Michele Pavon.

Stochastic control liaisons: Richard Sinkhorn meets Gaspard Monge on a Schrödinger bridge. SIAM Review, 63(2): 249–313, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clark:2023:SCB

[CGRG^{+}23] William A. Clark, Mario W. Gomes, Arnaldo Rodriguez-Gonzalez, Leo C. Stein, and Steven H. Strogatz. Surprises in a classic boundarylayer problem. SIAM Review, 65(1):291–315, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1436087.

Chui:1997:CSI

[CGT^{+}97a] W. K. Chui, J. Glimm, F. M. Tangerman, A. P. Jardine, J. S. Madsen, T. M. Donnellan, and R. Leek. Case study from industry: Process modeling in resin transfer molding as a method to enhance product quality. SIAM Review, 39 (4):714–727, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30854.

Chui:1997:PMR

[CGT^{+}97b] W. K. Chui, J. Glimm, F. M. Tangerman, J. S. Jardine, J. S. Madsen, T. M. Donnellan, and R. Leek. Processing modeling in resin transfer molding as a method to enhance product quality. SIAM Review, 39(4): 714–727, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cartwright:1961:NLD

[CH61] Dorwin Cartwright and Frank Harary. The number of lines in a digraph of each connectedness category. SIAM Review, 3(4): 309–314, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cobb:1966:CSS

[CH66] E. B. Cobb and Bernard Harris. The characterization of the solution sets for generalized reduced moment problems and its application to numerical integration. SIAM Review, 8(1): 86–99, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cholewinski:1968:CAG

[CH68] Frank M. Cholewinski and Deborah Tepper Haimo. Classical analysis and the generalized heat equation. SIAM Review, 10(1):67– 80, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chartres:1963:GPT

[Cha63] B. A. Chartres. A geometrical proof of a theorem due to D. Slepian. SIAM Review, 5 (4):335–341, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Charnes:1964:BRA

[Cha64a]  A. Charnes.  Book re-

view: Applied Dynamic Programming (Richard E. Bellman and Stuart E. Dreyfus). SIAM Review, 6(2): 184–185, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Charnes:1964:BRI

[Cha64b]  A. Charnes.  Book re-

view: Introduction to the Theory of Games (Ewald Burger). SIAM Review, 6 (3):324–325, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1966:BRR

[Cha66] A. Chang. Book review: Random Processes in Nonlinear Control Systems (A. A. Pervozvanskii). SIAM Review, 8 (3):400–401, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chandra:1967:BRO

[Cha67a] Jagdish Chandra. Book review: Ordinary Differential Equations (Fred Brauer and John Nohel). SIAM Review, 9 (2):264–265, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chawla:1967:CPS

[Cha67b] M. M. Chawla. On the Chebyshev polynomials of the second kind. SIAM Review, 9(4):729–733, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chazan:1968:BRI

[Cha68]  Dan Chazan.  Book re-

view: Introduction to Approximation Theory (E. W. Cheney). SIAM Review, 10 (3):393–394, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chandrasekaran:1969:PI

[Cha69] R. Chandrasekaran. A probability inequality. SIAM Review, 11(1):74, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chandra:1970:BRD

[Cha70a] Jagdish Chandra. Book review: Differential and Integral Inequalities: Theory and Applications, Volume 1 (V. Lakshmikantham and S. Leela). SIAM Review, 12(1): 164–165, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chatterjea:1970:UPC

[Cha70b] S. K. Chatterjea. On the unified presentation of classical orthogonal polynomials. SIAM Review, 12(1): 124–126, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chat:1972:BCC

[Cha72] W. A. Chat. Biorthogonal conditions for a class of polynomials. SIAM Review, 14 (3):494–495, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chandra:1978:BRQ

[Cha78] Jagdish Chandra. Book review: Quasilinear Hyperbolic Systems and Waves (A. Jeffrey). SIAM Review, 20(4): 867–868, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chatelin:1981:SAL

[Cha81a] Françoise Chatelin. The spectral approximation of linear operators with applications to the computation of eigenelements of differential and integral operators. SIAM Review, 23(4): 495–522, October 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chatfield:1981:BRP

[Cha81b] J. A. Chatfield. Book review: Product Integration with Applications to Differential Equations (John D. Dollard and Charles N. Friedman). SIAM Review, 23(4): 548–551, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1982:BCS

[Cha82a]  Gengzhe Chang.  A bi-

nomial coefficient summation. SIAM Review, 24(4): 478, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chao:1982:BRC

[Cha82b] C. Y. Chao. Book review:

Circulant Matrices (Philip J. Davis). SIAM Review, 24(3): 356–357, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1983:LPM

[Cha83a] Gengzhe Chang. Limit of a power of a matrix. SIAM Review, 25(1):97, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chao:1983:BRE

[Cha83b] C. Y. Chao. Book review: Elements of Algebra and Algebraic Computing (John D. Lipson). SIAM Review, 25 (3):426, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1984:SOD

[Cha84] Gengzhe Chang. A system of ordinary differential equations. SIAM Review, 26 (3):429, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1985:LBP

[Cha85] G. Chang. A limit of Bernstein polynomials. SIAM Review, 27(1):80, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chao:1986:BRG

[Cha86]  C. Y. Chao.  Book re-

view: Graphs and Algorithms (M. Gondran and M. Minoux). SIAM Review, 28 (4):583–584, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chahine:1987:BRF

[Cha87a] Moustafa T. Chahine. Book review: Foundations of Radiation Hydrodynamics (Dimitri Mihalas and Barbara Weibel Mihalas). SIAM Review, 29(4):648–650, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chawla:1987:BPP

[Cha87b] M. M. Chawla. Bounded polynomials with a prescribed behaviour at the origin. SIAM Review, 29(4): 631–632, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chawla:1987:CVP

[Cha87c] M. M. Chawla. A column vector problem from numerical integration. SIAM Review, 29 (3):469–470, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chafee:1994:BRN

[Cha94a]  N. Chafee.  Book re-

view: Nonlinear Ordinary Differential Equations (R. Grimshaw). SIAM Review, 36(4):664–??, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chafes:1994:BRN

[Cha94b]  Nathaniel Chafes.  Book

review: Nonlinear Ordinary Differential Equations (R. Grimshaw). SIAM Review, 36(4):664–665, December 1994. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1994:BRL

[Cha94c] T. C. Chang. Book review: Large Sample Methods in Statistics: An Introduction with Applications (Pranab K. Sen and Julio M. Singer). SIAM Review, 36 (4):666, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1994:BRT

[Cha94d] T. C. Chang. Book review: The Theory of Linear Models (Bent Jorgensen). SIAM Review, 36(4):665– 666, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chamberland:1995:PIG

[Cha95a] Marc Chamberland. A problem from integral geometry. SIAM Review, 37(4): 610, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chan:1995:BRM

[Cha95b] Wai Chan. Book review: Modelling Covariances and Latent Variables Using EQS (G. Dunn, B. Everitt, and A. Pickles). SIAM Review, 37 (3):462–464, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1995:TME

[Cha95c] J. Chapman. A Toeplitz matrix eigenproblem. SIAM Review, 37(2):245–249, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1995:BRT

[Cha95d] R. J. Chapman. Book review: A Toeplitz Matrix Eigenproblem (R. W. Cox). SIAM Review, 37(2):245– 249, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1995:U

[Cha95e]  R. J. Chapman.  Unti-

tled. SIAM Review, 37(3): 445–448, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1996:PIG

[Cha96a]  R. Chapman.  A prob-

lem from integral geometry. SIAM Review, 38(4): 678–679, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1996:SBF

[Cha96b] R. Chapman. A series of Bessel functions. SIAM Review, 38(2):329–332, ???? 1996. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1996:BRP

[Cha96c] Robin Chapman. Book review: A Problem from Integral Geometry (Marc Chamberland). SIAM Review, 38 (4):678–679, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1997:HPI

[Cha97]  R. Chapman.  A Her-

mite polynomial identity. SIAM Review, 39(3):521– 522, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:2000:HMT

[Cha00] S. J. Chapman. A hierarchy of models for Type-II superconductors. SIAM Review, 42 (4):555–598, December 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37191.

Castrigiano:1995:CT

[CHC95] Domenico P. L. Castrigiano, Sandra A. Hayes, and David Chillingworth. Catastrophe theory. SIAM Review, 37(1): 118–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheng:1960:VI

[Che60] Sin-I Cheng. Vorticity interaction. SIAM Review, 2 (2):151, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheney:1965:BRA

[Che65] E. W. Cheney. Book review:

Approximation von Funktionen and ihre numerische Behandlung (Gunter Meinardus). SIAM Review, 7(4): 585, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:1967:DGS

[Che67a]  Wai-Kai Chen.  On di-

rected graph solutions of linear algebraic equations. SIAM Review, 9(4):692– 707, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheney:1967:BRB

[Che67b] E. W. Cheney. Book review: Book Review: Chebyshev Methods in Numerical Approximation (Martin Avery Snyder). SIAM Review, 9(2):264, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:1968:ILT

[Che68]  Wai-Kai Chen.  Inverse

of a Laplace transform.

SIAM Review, 10(2):224– 225, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:1973:ILT

[Che73] Wai-Kai Chen. The inverse of a Laplace transform in electrical filter theory. SIAM Review, 15(4): 790–792, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chernoff:1975:BRC

[Che75] Herman Chernoff. Book review: Cluster Analysis for Applications (Michael R. Anderberg). SIAM Review, 17 (3):580–582, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheung:1976:RDS

[Che76] T. Y. Cheung. Recent development in solution of partialdifferential equations by linear programming. SIAM Review, 18(4):797, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheung:1978:RDN

[Che78] To-Yat Y. Cheung. Recent developments in the numerical solution of partial differential equations by linear programming. SIAM Review, 20(1):139–167, ???? 1978.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cherry:1986:BRC

[Che86] Guy W. Cherry. Book review:

Computer Algebra-Symbolic and Algebraic Computation. 2nd Edition (B. Buchberger, G. E. Collins and R. Loos, ed.). SIAM Review, 28(3): 407–408, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheng:1989:BRT

[Che89] H. K. Cheng. Book review: Transonic Aerodynamics (J. D. Cole and L. P. Cook). SIAM Review, 31(2): 336–338, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:1991:BRN

[Che91] Guanrong Chen. Book review: Numerical Solution of Optimal Control Problems with State Constraints by Sequential Quadratic Programming in Function Spaces (K. C. P. Machielsen). SIAM Review, 33(1):133–135, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:1992:BRL

[Che92a] Guanrong Chen. Book review: Linear Controller Design: Limits of Performance (Stephen P. Boyd and Craig H. Barratt). SIAM Review, 34(4):663–665, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheng:1992:ADD

[Che92b] H. K. Cheng. Airfoil design and data (Richard Eppler). SIAM Review, 34(2): 331–??, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheng:1992:BRA

[Che92c] H. K. Cheng. Book review: Airfoil Design and Data (Richard Eppler). SIAM Review, 34(2):331–333, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:1996:PSD

[Che96a] H. W. Chen. Periodic solutions of a differential equation. SIAM Review, 38(2): 327–328, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:1996:BRP

[Che96b] Hongwei Chen. Book review: Periodic Solutions of a Differential Equation (Philip Korman). SIAM Review, 38 (2):326–328, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:1998:PFR

[Che98] Hongwei W. Chen. The Poisson formula revisited. SIAM Review, 40(2):353– 355, June 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31873.

Cheney:2001:MTS

[Che01]  Margaret Cheney.  A

mathematical tutorial on synthetic aperture radar. SIAM Review, 43(2):301– 312, June 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/36885.

Cheney:2018:BRT

[Che18]  Margaret Cheney.  Book

review: Transionospheric Synthetic Aperture Imaging, by Mikhail Gilman, Erick Smith, and Semyon Tsynkov. SIAM Review, 60(2):471– 473, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Childs:1971:BRI

[Chi71] Donald R. Childs. Book review: An Infinite Integral (M. L. Glasser). SIAM Review, 13 (1):130–131, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Childress:1979:BRM

[Chi79] Stephen Childress. Book review: Magnetic Field Generation in Electrically Conducting Fluids (H. K. Moffatt). SIAM Review, 21(3): 407–409, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Childress:1983:BRM

[Chi83] Stephen Childress. Book review: Mathematical Models in Molecular and Cellular Biology (Lee A. Segel, ed.). SIAM Review, 25(1): 138–139, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chicone:1990:BRI

[Chi90] Carmen Chicone. Book review: Introduction to the Algebraic Theory of Invariants of Differential Equations (K. S. Sibirsky). SIAM Review, 32(3):491–493, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chillingworth:1994:BRF

[Chi94a] D. R. J. Chillingworth. Book review: Fundamentals of Dynamical Systems and Bifurcation Theory (M. Medved). SIAM Review, 36(1):122– 123, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chillingworth:1994:FDS

[Chi94b] D. R. J. Chillingworth. Fundamentals of dynamical systems and bifurcation (M Medved). SIAM Review, 36 (1):122–??, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chillingworth:1995:BRC

[Chi95] David Chillingworth. Book review: Catastrophe Theory (Domenico P. L. Cast-Rigiano and Sandra A. Hayes). SIAM Review, 37(1): 118–119, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Calhoun:2008:LRG

[CHL08] Donna A. Calhoun, Christiane Helzel, and Randall J. LeVeque. Logically rectangular grids and finite volume methods for PDEs in circular and spherical domains. SIAM Review, 50(4): 723–752, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:1966:BRI

[Cho66] T. Y. Chow. Book review: Intermediate Differential Equations (Earl D. Rainville). SIAM Review, 8 (3):409–411, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:1969:CHM

[Cho69] T. S. Chow. A class of Hessenberg matrices with known eigenvalues and inverses. SIAM Review, 11(3): 391–395, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:1970:BRF

[Cho70] T. Y. Chow. Book review: Functions of a Complex Variable: Theory and Technique (George F. Carrier, Max Krook and Carl E. Pearson). SIAM Review, 12 (1):162–163, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:1972:ERT

[Cho72] T. S. Chow. Eigenvalues of a reflected tridiagonal matrix. SIAM Review, 14(3): 495, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:1975:PMS

[Cho75] Pao Liu Chow. Perturbation methods in stochastic wave propagation. SIAM Review, 17(1):57–81, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Choi:1976:APL

[Cho76a] K. Choi. Asymptotic properties of least-squares estimators for discrete compound distributions. SIAM Review, 18(4):797, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:1976:IPP

[Cho76b] P. L. Chow. Inverse problem in propagation of noise. SIAM Review, 18(4):797– 798, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chong:1978:DRD

[Cho78]  K. P. Chong.  Deriva-

tion of relaxation data from creep tests, for small strains. SIAM Review, 20(3):621– 622, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:1987:BRL

[Cho87] Joe H. Chow. Book review: Lanczos Algorithms for Large Symmetric Eigenvalue Computations (Jane K. Cullum and Ralph A. Willoughby). SIAM Review, 29(2):307–309, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1992:MMS

[CHO92] S. J. Chapman, S. D.

Howison, and J. R. Ockendon. Macroscopic models for superconductivity.

SIAM Review, 34(4):529– 560, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Choudhury:1994:BRD

[Cho94] S. Roy Choudhury. Book review: Dynamics and Bifurcations (Jack K. Hale and Huseyin Kocak). SIAM Review, 36(2):297–299, June 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Christy:1968:DVS

[Chr68] R. F. Christy. The dynamics of variable stars. SIAM Review, 10(3):291– 302, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:2015:MFC

[CHT15] S. Jonathan Chapman, David P. Hewett, and Lloyd N. Trefethen. Mathematics of the Faraday cage. SIAM Review, 57(3):398–417, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Colbrook:2021:CSM

[CHT21] Matthew Colbrook, Andrew Horning, and Alex Townsend. Computing spectral measures of self-adjoint operators. SIAM Review, 63 (3):489–524, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Church:1966:FS

[Chu66] J. P. Church. A finite series. SIAM Review, 8(3): 381–382, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Church:1967:FS

[Chu67] J. P. Church. A finite series. SIAM Review, 9(1): 120, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chung:1978:BRC

[Chu78] F. R. K. Chung. Book review: A Conjectured Minimum Valuation Tree (I. Cahit). SIAM Review, 20(3): 601–604, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chui:1981:BRG

[Chu81] Charles K. Chui. Book review: The General Problem of Approximation and Spline Functions (B. N. Sahney and A. S. B. Holland). SIAM Review, 23(2):262, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chui:1982:BRA

[Chu82] Charles K. Chui. Book review: Approximation Theory and Methods (M. J. D. Powell). SIAM Review, 24(4): 493–494, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chui:1985:BRK

[Chu85]  Charles K. Chui.  Book

review: Kalman Filtering Theory (A. V. Balakrishnan). SIAM Review, 27 (2):265, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chu:1988:CRI

[Chu88a] Moody T. Chu. On the continuous realization of iterative processes. SIAM Review, 30(3):375–387, September 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chui:1988:BRD

[Chu88b] Charles K. Chui. Book review: Digital Filters. Theory and Applications (N. K. Boss). SIAM Review, 30 (1):141, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chui:1989:BRS

[Chu89] Charles K. Chui. Book review: Systems Analysis by Graphs and Matroids: Structural Solvability and Controllability (Kazuo Murota).

SIAM Review, 31(3):502– 503, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chui:1990:BRE

[Chu90] Charles K. Chui. Book review: Estimation, Control, and the Discrete Kalman Filter (Donald E. Calin). SIAM Review, 32(3):493– 494, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chunqing:1992:BRR

[Chu92] Lu Chunqing. Book review:

Reaction-Diffusion Equations (K. J. Brown and A. A. Lacey, eds.). SIAM Review, 34(1):131–??, March 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Churchill:1994:IPT

[Chu94a]  R. Churchill.  Invertible

point transformations and nonlinear differential equations (Willi-Hans Steeb). SIAM Review, 36(4):666– ??, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Churchill:1994:BRI

[Chu94b] Richard Churchill. Book review: Invertible Point Transformations and Nonlinear Differential Equations (Willi-Hans Steeb). SIAM Review, 36(4):666, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chui:1996:BRA

[Chu96] Charles K. Chui. Book review: Adapted Wavelet Analysis from Theory to Software (Mladen Victor Wickerhauser). SIAM Review, 38 (1):160, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chu:1998:IEP

[Chu98]  Moody T. Chu.  In-

verse eigenvalue problems. SIAM Review, 40(1):1–39, March 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/30398.

Csorba:2012:ANG

[CHW12] Péter Csorba, Cor A. J. Hurkens, and Gerhard J. Woeginger. The Alcuin number of a graph and its connections to the vertex cover number. SIAM Review, 54 (1):141–154, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v54/i1/p141_s1.

Ciarlet:1992:ARR

[Cia92a] P. G. Ciarlet. Author’s reply to the review of Introduction to Numerical Linear Algebra and Optimization. SIAM Review, 34(1): 159–??, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ciarlet:1992:BRA

[Cia92b] Philippe G. Ciarlet. Book review: Author’s reply to the review of Introduction to Numerical Linear Algebra and Optimisation (P. G. Ciarlet). SIAM Review, 34(1): 159, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cinlar:1993:BRS

[Çin93a] Erhan Çinlar. Book review: Stochastic Flows and Stochastic Differential Equations (Hiroshi Kunita). SIAM Review, 35(3):513–514, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cinlar:1993:SFS

[Cin93b] Erhan Cinlar. Stochastic flows and stochastic differential equations (Hiroshi Kunita). SIAM Review, 35(3): 513–??, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheney:1999:EIT

[CIN99] Margaret Cheney, David Isaacson, and Jonathan C. Newell. Electrical impedance tomography. SIAM Review, 41(1):85–101, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33361.

Cercignani:1995:MTD

[CIPB95] C. Cercignani, R. Illner, M. Pulvierenti, and Claude Bardos. The mathematical theory of dilute gases. SIAM Review, 37(4):622– ??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chossat:1995:CTP

[CIS95] Pascal Chossat, Gerard Iooss, and Juergen Scheurle. The Couette–Taylor problem. SIAM Review, 37(2): 279–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corless:2016:VTE

[CJ16] Robert M. Corless and Julia E. Jankowski. Variations on a theme of Euler. SIAM Review, 58(4): 775–792, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cartea:2018:ATS

[CJR18] Álvaro Cartea, Sebastian Jaimungal, and Jason Ricci. Algorithmic trading, stochastic control, and mutually exciting processes. SIAM Review, 60(3):673–703, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cooke:1976:PTT

[CK76] K. L. Cooke and J. Kaplan.

Periodicity threshold theorem for epidemics and population growth. SIAM Review, 18(4):799, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corduneanu:1986:BRS

[CK86] C. Corduneanu and A. Korzeniowski. Book review: Stochastic Functional Differential Equations (S-E. A. Mohammed). SIAM Review, 28 (3):411–413, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Colton:2018:LBI

[CK18] David Colton and Rainer Kress. Looking back on inverse scattering theory. SIAM Review, 60(4):779– 807, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Condat:2023:PSA

[CKCH23] Laurent Condat, Daichi Kitahara, Andrés Contreras, and Akira Hirabayashi. Proximal splitting algorithms for convex optimization: a tour of recent advances, with new twists. SIAM Review, 65 (2):375–435, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1379344.

Cassity:1962:BRS

[CL62] C. R. Cassity and N. A. Lindberger. Book review: On a Set of Linear Equations (B. W. Rosen). SIAM Review, 4 (4):398–401, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:1971:FD

[CL71] T. S. Chow and N. A. Lindberger. Factorization of a determinant. SIAM Review, 13 (2):246, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1972:BRP

[CL72] C. L. Chang and R. C. T. Lee. Book review: Problem-Solving Methods in Artificial Intelligence (Nils J. Nilsson). SIAM Review, 14(3): 513–514, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ciment:1975:HOC

[CL75] M. Ciment and S. H. Leventhal. Higher-order compact implicit schemes for hyperbolic equations. SIAM Review, 17(2):376–377, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cressie:1987:RST

[CL87] Noel Cressie and G. M.

Laslett. Random set theory and problems of modeling. SIAM Review, 29 (4):557–574, December 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chadwick:2008:HFO

[CL08] Richard S. Chadwick and Zhijie Liao. High-frequency oscillations of a sphere in a viscous fluid near a rigid plane. SIAM Review, 50(2): 313–322, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chamoin:2023:IRP

[CL23] Ludovic Chamoin and Frédéric Legoll. An introductory review on a posteriori error estimation in finite element computations. SIAM Review, 65(4):963–1028, November 2023. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Clark:1962:BRP

[Cla62] R. A. Clark. Book review: Proceedings of the I.U.T.A.M. Symposium of the Theory of Thin Elastic Shells (W. T. Koiter, ed.). SIAM Review, 4(2): 159, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clark:1964:BRB

[Cla64a] R. A. Clark. Book review: The Bending and Stretching of Plates (E. H. Mansfield). SIAM Review, 6(4): 468, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clark:1964:BRT

[Cla64b] R. A. Clark. Book review: The Theory of Thin Shells (V. V. Novozhilov). SIAM Review, 6(2):189– 190, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clark:1967:ADE

[Cla67] Colin Clark. The asymptotic distribution of eigenvalues and eigenfunctions for elliptic boundary value problems. SIAM Review, 9(4): 627–646, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clark:1978:MRN

[Cla78a] C. W. Clark. Models for renewable natural-resource management. SIAM Review, 20(3):622, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clarke:1978:OCT

[Cla78b] F. H. Clarke. Optimal control and true Hamiltonian. SIAM Review, 20(3):622, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clarks:1978:BRC

[Cla78c] Frank H. Clarks. Book review: Convex Analysis and Variational Problems (Ivar Ekeland and Roger Temam). SIAM Review, 20(1):192– 194, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clark:1979:MME

[Cla79a] Colin W. Clark. Mathematical models in the economics of renewable resources. SIAM Review, 21(1): 81–99, January 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clarke:1979:OCT

[Cla79b] Frank H. Clarke. Optimal control and the true Hamiltonian. SIAM Review, 21(2): 157–166, April 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clarke:1983:TAP

[Cla83a] T. L. Clarke. Techniques and applications of path integration — Schulman, L. S. SIAM Review, 25(3): 433–434, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clarks:1983:BRT

[Cla83b] Thomas L. Clarks. Book review: Techniques and Applications of Path Integration (Lawrence S. Schulman). SIAM Review, 25(3): 433–434, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clement:1959:CTD

[Cle59] Paul A. Clement. A class of triple-diagonal matrices for test purposes. SIAM Review, 1(1):50–52, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clemence:1961:UPS

[Cle61] G. M. Clemence. Unsolved problems in satellite theory. SIAM Review, 3(2):

156–161, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clement:1963:CBF

[Cle63] Preston R. Clement. On completeness of basis functions used for signal analysis. SIAM Review, 5(2): 131–139, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cleveland:1969:BRF

[Cle69] William S. Cleveland. Book review: A Family of Summation Formulas (Grace Wahba). SIAM Review, 11 (3):409–410, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clements:1980:RMS

[Cle80]  R. R. Clements.  Relia-

bility of multichannel systems. SIAM Review, 22(1): 88–95, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:2014:SPI

[CLG14] Ge Chen, Zhixin Liu, and Lei Guo. The smallest possible interaction radius for synchronization of self-propelled particles. SIAM Review, 56 (3):499–521, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ciarlet:1997:HNA

[CLH97] P. G. Ciarlet, J. L. Lions, and Weimin Han. Handbook of numerical analysis. vol. IV. finite element methods (part 2) and numerical methods for solids (part 2). SIAM Review, 39(4):796– ??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cline:1964:NGI

[Cli64]  Randall E. Cline.  Note

on the generalized inverse of the product of matrices. SIAM Review, 6(1): 57–58, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cline:1976:TQP

[Cli76] A. K. Cline. Transformation of a quadrative programming problem into solvable form. SIAM Review, 18 (4):798, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cline:1978:DIR

[Cli78] R. E. Cline. Drazin inverse for rectangular matrices. SIAM Review, 20(3): 622, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1999:TCR

[CLOT99] S. J. Chapman, J. M. H.

Lawry, J. R. Ockendon, and R. H. Tew. On the theory of complex rays. SIAM Review, 41(3):417– 509, September 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35205.

Chow:1995:NFB

[CLWM95] Shee-Nee Chow, Chengzhi Li, Duo Wang, and Kenneth R. Meyer. Normal forms and bifurcation of planar vector fields. SIAM Review, 37(2): 269–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Childs:1976:CNO

[CM76] S. B. Childs and M. J.

Maron. Constrained nonlinear optimization — direct approach. SIAM Review, 18 (4):797, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cavendish:1978:TRR

[CM78a] J. C. Cavendish and W. W.

Meyer. Two recurrence relations for Hermite basis polynomials. SIAM Review, 20 (1):182, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Childs:1978:NSC

[CM78b] B. Childs and M. J. Maron. Numerical solution of a class of control problems. SIAM Review, 20(3):621, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cameron:1983:CMS

[CM83] R. Cameron and S. McKee. A conjectured monotonic sequence. SIAM Review, 25(2): 267–268, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Choi:1989:ASA

[CM89] Y. S. Choi and Andrew Majda. Amplification of smallamplitude high-frequency waves in a reactive mixture. SIAM Review, 31(3):401– 427, September 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Criminale:1991:EBN

[CM91] W. O. Criminale and T. F. Mar. The electric Bell as a nonlinear oscillator. SIAM Review, 33(4):644– 649, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ciment:1973:BRC

[CMR73] Melvyn Ciment, Cleve Moler, and Otto G. Ruehr. Book review: A Conjectured Positive Definite Matrix (K. W. Schmidt). SIAM Review, 15 (1):222–223, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:2023:HS

[CMR23] Amenda N. Chow, Kirsten A.

Morris, and Gina F. Rabbah. Hysteresis and stability. SIAM Review, 65(4): 1171–1184, November 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cavendish:1979:TRR

[CMSM79] J. C. Cavendish, W. W.

Meyer, A. Sharma, and K. K. Mather. Two recurrence relations for Hermite basis polynomials. SIAM Review, 21 (1):143–144, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:2024:NOS

[CMSZ24] Shixiang Chen, Shiqian Ma, Anthony Man-Cho So, and Tong Zhang. Nonsmooth optimization over the Stiefel manifold and beyond: Proximal gradient method and recent variants. SIAM Review, 66(2):319–352, May 2024. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chen:1994:VSE

[CMZ94] Goong Chen, Philip J. Morris, and Jian-Xin X. Zhou. Visualization of special eigenmode shapes of a vibrating elliptical membrane. SIAM Review, 36(3):453– 469, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/23332.htm.

Chan:1996:CGM

[CN96] Raymond H. Chan and Michael K. Ng. Conjugate gradient methods for Toeplitz systems. SIAM Review, 38 (3):427–482, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27647.

Cooley:2005:IID

[CN05] Bryan Cooley and Paul K. Newton. Iterated impact dynamics of N-beads on a ring. SIAM Review, 47(2): 273–300, June 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/42658.

Cropp:2015:PIE

[CN15] R. A. Cropp and J. Norbury. Population interactions in ecology: a rule-based approach to modeling ecosystems in a mass-conserving framework. SIAM Review, 57 (3):437–465, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cropp:2019:OMR

[CN19] Roger Cropp and John Norbury. Obligate mutualism in a resource-based framework. SIAM Review, 61(3): 596–622, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cartwright:2007:UAA

[CO07] Natalie A. Cartwright and Kurt E. Oughstun. Uniform asymptotics applied to ultrawideband pulse propagation. SIAM Review, 49(4): 628–648, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cochran:1984:BRL

[Coc84] James A. Cochran. Book review: Linear Integral Operators (K. Jörgens). SIAM Review, 26(1):128–129, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cochran:1986:BRI

[Coc86] James A. Cochran. Book review: Introduction to Integral Equations with Applications (Abdul J. Jerri). SIAM Review, 28(3):408– 409, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cody:1970:SPR

[Cod70] W. J. Cody. A survey of practical rational and polynomial approximation of functions. SIAM Review, 12(3): 400–423, July 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Reprinted in Studies in Applied Mathematics, 6, SIAM, 1970, pp. 86–109.

Cody:1974:CNS

[Cod74] W. J. Cody. The construction of numerical subroutine libraries. SIAM Review, 16 (1):36–46, January 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cody:1978:BRM

[Cod78]  W. J. Cody.  Book re-

view: Mathematical Functions and Their Approximations (Yudell L. Luke). SIAM Review, 20(1):194– 195, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coffman:1983:ICM

[Cof83] E. G. Coffman, Jr. An introduction to combinatorial models of dynamic storage allocation. SIAM Review, 25 (3):311–325, July 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1960:SC

[Coh60] Herbert A. Cohen. On a switching circuit. SIAM Review, 2(4):297, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1969:PPP

[Coh69] Edgar Cohen, Jr. A polynomial with proscribed properties. SIAM Review, 11 (4):620, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1975:ZDB

[Coh75a] E. A. Cohen. Zero distribution and behavior of orthogonal polynomials in Sobolev space w1,2[-1,1]. SIAM Review, 17(2):377, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohn:1975:BRR

[Coh75b] Harvey Cohn. Book review:

Riemann’s Zeta Function (H. M. Edwards). SIAM Review, 17(4):697–699, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1976:BRM

[Coh76] Donald S. Cohen. Book review: The Mathematical Theory of Diffusion and Reaction in Permeable Catalysts (Rutherford Aris). SIAM Review, 18 (4):783, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1981:LS

[Coh81] A. M. Cohen. Late solutions. SIAM Review, 23 (4):532, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohn:1982:BRA

[Coh82] Harvey Cohn. Book review: Analytic Number Theory: an Introduction (Richard Bellman). SIAM Review, 24 (1):99, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohn:1984:ITN

[Coh84a] H. Cohn. Introduction to the theory of numbers — Shapiro, H. N. SIAM Review, 26(3):452–454, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohn:1984:BRI

[Coh84b] Harvey Cohn. Book review: Introduction to the Theory of Numbers (Harold N. Shapiro). SIAM Review, 26(3):452–454, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1988:SGP

[Coh88] Joel E. Cohen. Subadditivity, generalized products of random matrices and operations research. SIAM Review, 30(1):69–86, March 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See erratum [Coh93b].

Cohen:1990:BRP

[Coh90a] Joel E. Cohen. Book review:

Population System Control (Jian Song and Jingyuan Yu). SIAM Review, 32(3): 494–500, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1990:PSC

[Coh90b] Joel E. Cohen. Population system control (Jian Song and Jingyuan Yu). SIAM Review, 32(3):494– ??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1993:SGP

[Coh93a] J. E. Cohen. Subadditivity, generalized products of random matrices and operations research (vol 30, pg 69, 1988). SIAM Review, 35 (1):124, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1993:ESG

[Coh93b]  Joel E. Cohen.  Erra-

tum: “Subadditivity, generalized products of random matrices and operations research” [SIAM rev. 30 (1988), no. 1, 69–86 and MR 89g:60300]. SIAM Review, 35 (1):124, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Coh88].

Coleman:1973:BRS

[Col73] Courtney Coleman. Book review: Stability of Motion (Wolfgang Hahn). SIAM Review, 15(1):228–229, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cole:1976:BRP

[Col76a] Julian D. Cole. Book review: Perturbation Methods (Ali Hasan Nayfeh). SIAM Review, 18(1):139– 140, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Colvin:1976:WNM

[Col76b] B. H. Colvin. What nonacademic mathematicians do — examples and case studies. SIAM Review, 18(4):

798, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cole:1977:BRP

[Col77] Julian D. Cole. Book review: Perturbation Methods in Fluid Mechanics (Milton Van Dyke). SIAM Review, 19(3):572, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Colton:1984:ISP

[Col84]  David Colton.  The in-

verse scattering problem for time-harmonic acoustic waves. SIAM Review, 26(3): 323–350, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Colonius:1986:BRO

[Col86] F. Colonius. Book review:

Oscillation Theory of Optimal Processes (George M. Smirnov). SIAM Review, 28 (3):410–411, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cole:1987:BRW

[Col87a] Susan Cole. Book review:

Water Waves and Ship Hydrodynamics: an Introduction (R. Timman, A. J. Hermans and G. C. Hsiao). SIAM Review, 29(1):139– 140, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coleman:1987:BRN

[Col87b] Tom Coleman. Book review: Numerical Optimization Techniques (Yurij G. Evtushenko). SIAM Review, 29 (2):309–310, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Colonius:1989:BRV

[Col89] Fritz Colonius. Book review: Viscosity Solutions and Optimal Control (Robert J. Elliott). SIAM Review, 31 (3):505, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cole:1991:BRP

[Col91] Susan Cole. Book review: Partial Differential Equations: Analytical Solution Techniques (J. Kevorkian). SIAM Review, 33(1):135– 136, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coleman:1992:BRN

[Col92] William P. Coleman. Book review: Nonlinear Estimation (Gavin J. S. Ross). SIAM Review, 34(1):156– 157, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cole:1994:DPT

[Col94a] Julian D. Cole. The development of perturbation theory at Galcit. SIAM Review, 36 (3):425–430, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/25790.
htm.

Colombeau:1994:BRM

[Col94b]  J.-F. Colombeau.  Book

review: Multiplication of Distributions and Applications to Partial Differential Equations (M. Oberguggenberger). SIAM Review, 36 (2):299–301, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Colton:1994:BRI

[Col94c] David Colton. Book review: Inverse Problems in the Mathematical Sciences (Charles W. Groetsch). SIAM Review, 36(1):123– 124, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Colton:1994:IPM

[Col94d] David Colton. Inverse problems in the mathematical sciences (Charles W Groetsch). SIAM Review, 36(1):123– ??, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Collins:1996:BRO

[Col96] Michael D. Collins. Book review: Ocean Acoustic Tomography (W. Munk, P. Worcester, and C. Wunsch). SIAM Review, 38(4): 683–684, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Comstock:1972:PLP

[Com72]  Craig Comstock.  The

Poincaré–Lighthill perturbation technique and its generalizations. SIAM Review, 14 (3):433–446, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Conte:1966:NSL

[Con66]  S. D. Conte.  The nu-

merical solution of linear boundary value problems. SIAM Review, 8(3):309– 321, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Conolly:1971:RRW

[Con71] B. W. Conolly. On randomized random walks. SIAM Review, 13(1):81–99, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Conley:1985:BRG

[Con85] Charles Conley. Book review: Geometrical Methods in the Theory of Ordinary Differential Equations (V. I. Arnold). SIAM Review, 27 (1):95–96, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Constantin:1994:GST

[Con94]  Peter Constantin.  Geo-

metric statistics in turbulence. SIAM Review, 36(1): 73–98, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/24714.htm.

Constantinescu:1996:BRT

[Con96] T. Constantinescu. Book review: Time-Varying Discrete Linear Systems (Aristide Halanay and Vlad Ionescu ). SIAM Review, 38(4):684– 685, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Constanda:1997:PCF

[Con97] Christian Constanda. Problems: Composition formulae for boundary operators (problem 97-3). SIAM Review, 39(1):124–125, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Constanda:1998:CFB

[Con98] C. Constanda. Composition formulae for boundary operators. SIAM Review, 40(1): 128–132, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cooper:1964:HML

[Coo64]  Leon Cooper.  Heuris-

tic methods for locationallocation problems. SIAM Review, 6(1):37–53, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cooper:1965:BRM

[Coo65] J. L. B. Cooper. Book review: Mathematical Foundations of Thermodynamics (R. Giles). SIAM Review, 7(3): 431–433, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cook:1966:BRP

[Coo66]  J. M. Cook.  Book re-

view: Perturbation of Spectra in Hubert Space (K. O. Friedrichs). SIAM Review, 8 (4):545–546, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cook:1967:BRD

[Coo67a] J. M. Cook. Book review: Differential Space, Quantum Systems, and Prediction (Norbert Wiener, Armand Siegel, Bayard’ Rankin and William Ted Martin). SIAM Review, 9(3):601– 602, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cooper:1967:POD

[Coo67b] Henry F. Cooper, Jr. Propagation of one-dimensional waves in inhomogeneous elastic media. SIAM Review, 9 (4):671–679, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coonce:1971:ROE

[Coo71] H. B. Coonce. Root of an octic equation. SIAM Review, 13(2):248, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cook:1973:BRQ

[Coo73]  J. M. Cook.  Book re-

view: Quantum Mechanics in Hilbert Space (Eduard Prugovecki). SIAM Review, 15 (2):403–404, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cook:1989:BRM

[Coo89] L. Pamela Cook. Book review: Mathematics for Dynamic Modeling (Edward Beltrami). SIAM Review, 31 (3):505–506, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coomes:1991:BRM

[Coo91] Brian Coomes. Book review: Multiphase Averaging for Classical Systems: With Applications to Adiabatic Theorems (P. Lochak and C. Meunier; trans. by H. S. Dumas). SIAM Review, 33 (3):495–496, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cooper:1993:ISE

[Coo93] Robert B. Cooper. An infinite sum of Erlang loss functions. SIAM Review, 35 (1):137, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cooper:1994:ISE

[Coo94] R. B. Cooper. An infinite sum of Erlang loss functions. SIAM Review, 36(1): 112–113, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Copson:1971:HES

[Cop71] E. T. Copson. Hadamard’s elementary solution and Frobenius’s method. SIAM Review, 13(2):222–230, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coppersmith:1986:BRH

[Cop86] Don Coppersmith. Book review: How to Multiply Matrices Faster (Victor Pan). SIAM Review, 28(2):250– 252, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Copeland:2007:BAT

[Cop07] Arthur H. Copeland, Jr.

Buchberger’s algorithm and the two-locus, two-allele model. SIAM Review, 49(3): 421–433, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corrington:1968:NES

[Cor68] Murlan S. Corrington. Numerical evaluation of some integrals. SIAM Review, 10 (2):224, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cornille:1972:CHT

[Cor72]  P. Cornille.  Computa-

tion of Hankel transforms. SIAM Review, 14(2):278– 285, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corliss:1977:WRD

[Cor77] George Corliss. Which root does the bisection algorithm find? SIAM Review, 19(2): 325–327, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corliss:1978:NSO

[Cor78] G. Corliss. Numerical solution of ODEs in complex plane pole vaulting. SIAM Review, 20(3):623, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corrsin:1979:BRS

[Cor79] Stanley Corrsin. Book review: The Structure of Turbulent Shear Flow (A. A. Townsend). SIAM Review, 21 (4):581–583, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corduneanu:1992:BRF

[Cor92a] C. Corduneanu. Book review: Functional Differential Equations with Infinite Delay (Y. Hino, S. Murakami, and T.Naito). SIAM Review, 34 (3):504–506, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corsten:1992:BRS

[Cor92b] L. C. A. Corsten. Book review: Statistical Methods:

The Geometric Approach (David J. Saville and Graham R. Wood). SIAM Review, 34 (3):506–508, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corduneanu:1995:BRE

[Cor95] C. Corduneanu. Book review: Evolutionary Integral Equations and Applications (J. Pruss). SIAM Review, 37 (1):117–118, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corless:2016:BRT

[Cor16] Robert M. Corless. Book review of: L. N. Trefethen, Approximation theory and approximation practice. SIAM Review, 58(1):159–163, 2016. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Costa:1993:MDI

[Cos93] Peter J. Costa. Matrix differentiation identities. SIAM Review, 35(4):642, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cockayne:2019:BPN

[COSG19] Jon Cockayne, Chris J.

Oates, T. J. Sullivan, and Mark Girolami. Bayesian probabilistic numerical methods. SIAM Review, 61(4):

756–789, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cottle:1973:BRP

[Cot73] R. W. Cottle. Book review: A Parametric Linear Complementarily Problem (G. Maier). SIAM Review, 15(2): 381–384, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cottle:1976:SCL

[Cot76]  R. W. Cottle.  Special

class of large quadratic programs. SIAM Review, 18 (4):799, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cottle:1994:BRL

[Cot94] R. W. Cottle. Book review: Linear Programs and Related Problems (Evar D. Nering and Albert W. Tucker). SIAM Review, 36(4):666– 668, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coulton:1997:BRI

[Cou97] David L. Coulton. Book review: An Introduction to Linear and Nonlinear Scattering Theory, by G. F. Roach. SIAM Review, 39(2): 336–337, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chapman:1996:BRS

[COW96] Robin Chapman, Norbert Ortner, and Peter Wagner. Book review: A Series of Bessel Functions (Norbert Ortner and Peter Wagner). SIAM Review, 38(2): 328–332, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coxeter:1965:BRP

[Cox65] H. S. M. Coxeter. Book review: Packing and Covering (C. A. Rogers). SIAM Review, 7(2):303–305, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cox:1994:TME

[Cox94] R. W. Cox. A Toeplitz matrix eigenproblem. SIAM Review, 36(2):278, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cozzens:1985:BRM

[Coz85] Margaret Cozzens. Book review: Mathematical Models in the Social and Behavioral Sciences (Anatol Rapoport). SIAM Review, 27 (2):266–267, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cozzens:1987:BRM

[Coz87] Margaret B. Cozzens. Book review: Mathematics in the Social and Life Sciences: Theories, Models and Methods (M. A. Ball). SIAM Review, 29(2):310– 311, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cline:1976:SUL

[CP76] R. E. Cline and R. J. Plemmons. L_{2}-solutions to underdetermined linear systems. SIAM Review, 18(1): 92–106, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Concus:1978:NSB

[CP78] P. Concus and W. Proskurowski. Numerical solution of Buckley– Leverett equation by random choice method. SIAM Review, 20(3):622, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohoon:1989:NRT

[CP89] D. K. Cohoon and R. M. Purcell. On the number of roots of a transcendental equation. SIAM Review, 31 (4):675, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cuturi:2018:SRO

[CP18] Marco Cuturi and Gabriel Peyré. Semidual regularized optimal transport. SIAM Review, 60(4):941– 965, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohoon:1990:NRT

[CPRG90] D. K. Cohoon, R. M. Purcell, Michael Renardy, and C. C. Grosjean. On the number of roots of a transcendental equation. SIAM Review, 32 (4):682–683, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Calvetti:2018:BMK

[CPSV18] D. Calvetti, F. Pitolli, E. Somersalo, and B. Vantaggi. Bayes meets Krylov: Statistically inspired preconditioners for CGLS. SIAM Review, 60(2):429–461, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chandrasekaran:2012:CGI

[CPW12] Venkat Chandrasekaran, Pablo A.

Parrilo, and Alan S. Willsky. Convex graph invariants. SIAM Review, 54(3): 513–541, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chaudhry:1996:DIP

[CQRZ96] M. Aslam Chaudhry, Asghar Qadir, M. Rafique, and S. M. Zubair. A definite integral in probability theory. SIAM Review, 38(3): 514–515, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cooper:1967:SPA

[CR67] J. N. Cooper and R. A. Restrepo. Some problems of attack and defense. SIAM Review, 9(4): 680–691, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cope:1978:GRR

[CR78] J. E. Cope and B. W. Rust. Generalization of ridge regression methods to interval estimation. SIAM Review, 20 (3):622, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clark:1982:I

[CR82] Jo Ann Clark and O. G. Ruehr. An identity. SIAM Review, 24(2):224, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clark:1983:I

[CR83] J. A. Clark and O. G. Ruehr. An identity. SIAM Review, 25(2):273–274, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chalker:1996:IAS

[CR96] Tae Chalker and Joshua Radcliff. An integral arising in the study of a random graph. SIAM Review, 38 (1):146, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Crawford:1974:MP

[Cra74] C. R. Crawford. A matrix problem. SIAM Review, 16 (3):391, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Crawford:1975:MP

[Cra75] C. R. Crawford. Matrix problem. SIAM Review, 17(4): 695–696, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Craven:1984:BRN

[Cra84] B. D. Craven. Book review: Non-Linear Parametric Optimization (B. Bank, J. Guddat, D. Klatte, B. Kummer and K. Tammer). SIAM Review, 26(4):594– 595, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Craven:1986:BRT

[Cra86] B. D. Craven. Book review: Theory of Multiobjective Optimization (Yoshikazu Sawaragi, Hirotaka Nakayama and Tetsuzo Tanino). SIAM Review, 28(4):584–586, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Crawford:1989:BRG

[Cra89] John David Crawford. Book review: Global Bifurcations and Chaos: Analytical Methods (S. Wiggins). SIAM Review, 31(4):690– 692, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cressie:1997:BRI

[Cre97]  Noel Cressie.  Book re-

view: Introduction to Disjunctive Kriging and Non-Linear Geostatistics, by J. Rivoirard. SIAM Review, 39 (2):337–340, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Crow:1973:BRI

[Cro73]  Steven C. Crow.  Book

review: Introduction to Mathematical Fluid Dynamics (Richard E. Meyer). SIAM Review, 15(2):401– 403, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cronin:1977:SMB

[Cro77] Jane Cronin. Some mathematics of biological oscillations. SIAM Review, 19(1): 100–138, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Crow:1985:BRS

[Cro85] John A. Crow. Book review:

A System of Ordinary Differential Equations (Gengzhe Chang). SIAM Review, 27 (3):452–453, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cruse:1986:BRB

[Cru86] T. A. Cruse. Book review:

The Boundary Integral Equation Method in Axisymmetric Stress Analysis Problems (A. A. Bakr). SIAM Review, 28 (4):586, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cheney:1963:SMR

[CS63] E. W. Cheney and T. H.

Southard. A survey of methods for rational approximation, with particular reference to a new method based on a formula of Darboux. SIAM Review, 5(3): 219–231, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chew:1964:BRL

[CS64] Victor Chew and M. J. Synge. Book review: A Least-Squares Solution of a Linear System (Franklin W. Diederich). SIAM Review, 6 (2):182–183, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlitz:1970:BRS

[CS70] L. Carlitz and R. Scoville. Book review: Sequences of Absolute Differences (John M. Hammersley). SIAM Review, 12(2): 297–300, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Childress:1975:AIT

[CS75] W. S. Childress and E. A. Spiegel. Archimedean instabilities in two-phase flows. SIAM Review, 17(1):136– 165, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chang:1983:BS

[CS83] Gengzhe Chang and Zun Shan. A binomial summation. SIAM Review, 25 (1):97, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chiappetta:1984:TDW

[CS84] L. M. Chiappetta and D. R.

Sobel. The temperature distribution within a hemisphere exposed to a hot gas stream. SIAM Review, 26(4): 575–577, October 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cuvelier:1990:SNM

[CS90] C. Cuvelier and R. M. S. M.

Schulkes. Some numerical methods for the computation of capillary free boundaries governed by the Navier– Stokes equations. SIAM Review, 32(3):355–423, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corrlieb:1997:GPR

[CS97] David Corrlieb and Ch Wang Shu. On the Gibbs phenomenon and its resolution. SIAM Review, 39(4):644– ??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cain:2006:TTA

[CS06] John W. Cain and David G.

Schaeffer. Two-term asymptotic approximation of a cardiac restitution curve. SIAM Review, 48(3):537– 546, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Calamoneri:2016:PCG

[CS16] Tiziana Calamoneri and Blerina Sinaimeri. Pairwise compatibility graphs: a survey. SIAM Review, 58(3): 445–460, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chan:2017:FRM

[CS17] Patrick Chan and Ronnie Sircar. Fracking, renewables, and mean field games. SIAM Review, 59(3):588– 615, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Corless:2020:REI

[CS20] Robert M. Corless and Leili Rafiee Sevyeri. The Runge example for interpolation and Wilkinson’s examples for rootfinding. SIAM Review, 62(1):231– 243, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clauset:2009:PLD

[CSN09] Aaron Clauset, Cosma Rohilla Shalizi, and M. E. J. Newman. Power-law distributions in empirical data. SIAM Review, 51(4):661– 703, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Catlla:2008:SMS

[CSW^{+}08] Anne J. Catllá, David G.

Schaeffer, Thomas P. Witelski, Eric E. Monson, and Anna L. Lin. On spiking models for synaptic activity and impulsive differential equations. SIAM Review, 50 (3):553–569, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chow:1975:SSN

[CT75] P. L. Chow and W. C.

Tam. Solutions to some nonlinear diffusion equations in population-growth models. SIAM Review, 17(2): 376, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chvatal:1976:DOG

[CT76a] V. Chvatal and C. Thomassen.

Distances in orientations of graphs. SIAM Review, 18 (4):798, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Comstock:1976:GTL

[CT76b] C. Comstock and J. G.

Taylor. General theory for Lanchester combat models. SIAM Review, 18(4): 798, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cox:1995:FSH

[CT95] Earl Cox and Rod Taber. The fuzzy systems handbook: a practitioner’s guide to building, using, and maintaining fuzzy systems. SIAM Review, 37(2):281–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cordes:1996:TPO

[CT96] H. O. Cordes and Michael Taylor. The technique of pseudodifferential operators. SIAM Review, 38(3):540– ??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Clason:2021:CRD

[CTW21] Christian Clason, Carla Tameling, and Benedikt Wirth. Convex relaxation of discrete vector-valued optimization problems. SIAM Review, 63(4):783–821, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cox:2003:WBH

[CU03] Steven J. Cox and Paul X. Uhlig. Where best to hold a drum fast. SIAM Review, 45(1):75–92, March 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41427.

Culver:1966:BRS

[Cul66] Walter J. Culver. Book review: Systems and Simulation (Dimitris N. Chorafas). SIAM Review, 8(3):407– 409, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Culver:1968:BRC

[Cul68] Walter J. Culver. Book review: Control Systems Functions and Programming Approaches — Volume A: Theory (Dimitris N. Chorafas). SIAM Review, 10(4):462– 464, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cullum:1977:BRO

[Cul77]  Jane Cullum.  Book re-

view: Optimization Theory; The Finite Dimensional Case (Magnus R. Hestenes). SIAM Review, 19(4):750– 752, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cullum:1978:SCL

[Cul78]  J. Cullum.  Simultane-

ous computation of largest and smallest eigenvalues and corresponding eigenvectors of large, symmetric matrices. SIAM Review, 20(3): 623, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cullum:1994:NML

[Cul94a] J. Cullum. Numerical methods for large eigenvalue problems (Youcef Saad). SIAM Review, 36(2):301–??, June 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cullum:1994:BRN

[Cul94b] Jane Cullum. Book review: Numerical Methods for Large Eigenvalue Problems (Youcef Saad). SIAM Review, 36(2): 301–303, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cumberbatch:1991:BRM

[Cum91] Ellis Cumberbatch. Book review: Mathematical Modelling: a Case Study Approach (R. R. Clements). SIAM Review, 33(1):136, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cumberbatch:1992:BRC

[Cum92a] Ellis Cumberbatch. Book review: A Concrete Approach to Mathematical Modelling (Michael Mesterton-Gibbons). SIAM Review, 34 (1):140–141, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cumberbatch:1992:CAM

[Cum92b] Ellis Cumberbatch. A concrete approach to mathematical modelling (Michael Mesterton-Gibbons). SIAM Review, 34(1):140–??, March 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cumberbatch:1996:BRI

[Cum96] Ellis Cumberbatch. Book review: Industrial Mathematics: a Course in Solving Real-World Problems (Avner Friedman and Walter Littman). SIAM Review, 38 (2):341–342, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Curtis:1964:BRR

[Cur64a] C. W. Curtis. Book review: Representations of Groups (Hermann Boerner). SIAM Review, 6(1):73–74, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Curtiss:1964:BRN

[Cur64b] C. F. Curtiss. Book review: Non-Equilibrium Statistical Mechanics, Vol. 1 (I. Prigogine). SIAM Review, 6(3):320, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Curtain:1975:SID

[Cur75] Ruth Curtain. A survey of infinite-dimensional filtering. SIAM Review, 17(3): 395–411, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Curtis:1977:BRM

[Cur77] A. R. Curtis. Book review: A Matrix Stability Problem (S. Venit). SIAM Review, 19(2): 331–332, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cushing:1996:BRT

[Cus96] J. M. Cushing. Book review: The Theory of the Chemostat: Dynamics of Microbial Competition (Hal L. Smith and Paul Waltman). SIAM Review, 38(1): 160–161, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cuvelier:1990:BRM

[Cuv90] C. Cuvelier. Book review: Méthodes des Éléments Finis pour les Fluides (O. Pironneau); English translation: Finite Element Methods for Fluids (O. Pironneau). SIAM Review, 32(2): 315–316, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Curtis:1977:MSP

[CV77] A. R. Curtis and S. Venit.

Matrix stability problem. SIAM Review, 19(2):331– 332, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Carlson:1991:IAL

[CV91] B. C. Carlson and M. Vuorinen. Inequality of the AGM and the logarithmic mean. SIAM Review, 33 (4):655, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Cohen:1970:DI

[CW70] A. M. Cohen and F. J.

Wilson. A definite integral. SIAM Review, 12(2): 288–289, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Casey:1994:SCE

[CW94] Stephen D. Casey and David F. Walnut. Systems of convolution equations, deconvolution, Shannon sampling, and the wavelet and Gabor transforms. SIAM Review, 36(4):537–577, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/24616.htm.

Czyzyk:1999:OCS

[CWW99] Joseph Czyzyk, Timothy Wisniewski, and Stephen J. Wright. Optimization case studies in the NEOS guide. SIAM Review, 41(1):148– 163, March 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33487.

Cybenko:1993:BRC

[Cyb93] George Cybenko. Book review: Computation and Cognition (C. William Gear, ed.). SIAM Review, 35(1): 147–148, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Coleman:1989:BRD

[CZ89] Thomas F. Coleman and Earl Zmijewski. Book review: Direct Methods for Sparse Matrices (I. S. Duff, A. M. Erisman, and J. K Reid). SIAM Review, 31(3): 503–504, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chaudhry:1992:TIA

[CZ92] M. A. Chaudhry and S. M. Zubair. Two integrals arising in generalized inverse Gaussian model and heat conduction problems. SIAM Review, 34(3):498, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chaudhry:1994:LTI

[CZ94] M. Aslam Chaudhry and S. M. Zubair. Laplace transform involving the Macdonald function. SIAM Review, 36(3):489, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chaudhry:1996:SII

[CZ96a] M. Aslam Chaudhry and S. M. Zubair. A semiinfinite integral involving the Macdonald function. SIAM Review, 38(4):669– 670, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Christopolous:1996:TLM

[CZ96b] Christos Christopolous and A. H. Zemanian. The transmission-line modeling method TLM. SIAM Review, 38(4):711–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Chaudhry:1997:SII

[CZ97] M. A. Chaudhry and S. M.

Zubair. A semi-infinite integral involving the Macdonald function. SIAM Review, 39 (4):786–789, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Curtain:1996:IID

[CZM96] R. F. Curtain, H. J. Zwart, and K. A. Morris. An introduction to infinitedimensional linear system theory. SIAM Review, 38(3): 536–??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Diaz:1978:CGM

[DA78] J. C. Diaz and D. Archer. Collocation-Galerkin method for 1st order hyperbolic equations. SIAM Review, 20 (3):623, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1979:NAA

[DA79a] P. J. Davis and J. A. Anderson. Non-analytic aspects of mathematics and their implication for research and education. SIAM Review, 21(1): 112–127, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1979:NAM

[DA79b] Philip J. Davis and James A. Anderson. Nonanalytic aspects of mathematics and their implication for research and education. SIAM Review, 21(1):112–127, ???? 1979. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Daalhuis:1997:BRB

[Daa97] Adri Olde Daalhuis. Book review: Borel–Laplace Transform and Asymptotic Theory: Introduction to Resurgent Analysis, by Boris Yu Sternin and Victor E. Shatalov. SIAM Review, 39(1): 159–161, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dafermos:1989:BRM

[Daf89] C. M. Dafermos. Book review: Mathematical Problems in Viscoelasticity (Michael Renardy, William J. Hrusa, and John A. Nobel). SIAM Review, 31(2):338–339, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Daley:1976:AD

[Dal76] D. J. Daley. Average distance. SIAM Review, 18(3): 497–500, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Damlamian:1984:WVM

[Dam84a] A. Damlamian. Weak and variational methods for moving boundary problems — Elliott, C. M., Ockendon, J. R. SIAM Review, 26(1): 137–138, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Damlamian:1984:BRW

[Dam84b] Alain Damlamian. Book review: Weak and Variational Methods for Moving Boundary Problems (C. M. Elliott and J. R. Ockendon). SIAM Review, 26(1): 137–138, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Damjanovia:1985:HFI

[Dam85a] S. Lj. Damjanovia. A hypergeometric function integral. SIAM Review, 27(4): 574, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Damjanovic:1985:BFI

[Dam85b] S. Damjanovic. A Bessel function integral. SIAM Review, 27(3):445, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Damjanovic:1986:BFI

[Dam86a] S. Damjanovic. A Bessel function integral. SIAM Review, 28(3):399–400, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Damjanovic:1986:IHF

[Dam86b] S. Lj. Damjanović. An identity for a _{4}F_{3} hypergeometric function. SIAM Review, 28 (1):86–87, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Damjanovic:1987:IHF

[Dam87] S. Lj. Damjanović. An identity for a _{4}F_{3} hypergeometric function. SIAM Review, 29(1):131–132, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Damjanovic:1988:I

[Dam88] S. L. Damjanovic. An identity. SIAM Review, 30(2): 318–319, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Daniel:1966:BRS

[Dan66] Cuthbert Daniel. Book review: Statistics and Experimental Design in Engineering and the Physical Sciences (Norman L. Johnson and Fred C. Leone). SIAM Review, 8(3):402– 404, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Daniel:1971:BRN

[Dan71] James W. Daniel. Book review: Numerical Methods That Work (Forman S. Acton). SIAM Review, 13(4): 581–582, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Daniel:1972:BRO

[Dan72] James W. Daniel. Book review: Optimisation Théorie et Algorithmes (Jean Céa). SIAM Review, 14(3):509– 511, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dancer:1991:BRE

[Dan91] E. N. Dancer. Book review: Elementary Stability and Bifurcation Theory (Gerard Loos and Daniel D. Joseph). SIAM Review, 33(1):136–137, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Danby:1993:BRO

[Dan93] J. M. Anthony Danby. Book review: Ordinary Differential Equations (G. F. Carrier and E. E. Pearson). SIAM Review, 35(4):658– 659, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Darling:1997:CNC

[Dar97] R. W. R. Darling. Classroom note: Converting matrix Riccati equations to second-order linear ODE. SIAM Review, 39(3):508– 510, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29875.

Datko:1967:BRI

[Dat67] R. F. Datko. Book review: An Introduction to Optimal Control (George Leitmann). SIAM Review, 9(4): 764–765, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Datko:1969:BRF

[Dat69] R. Datko. Book review:

Foundations of Optimal Control Theory (E. Bruce Lee and Lawrence Markus). SIAM Review, 11(1):93–95, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Datko:1973:BRM

[Dat73] Richard F. Datko. Book review: Mathematical Methods of Optimal Control (V. G. Boltyanskii). SIAM Review, 15(1):231–232, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Datko:1978:BRT

[Dat78] Richard F. Datko. Book review: Theory of Functional Differential Equations (Jack Hale). SIAM Review, 20(3): 610–612, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Datko:1980:BRC

[Dat80] Richard F. Datko. Book review: Convexity and Optimization in Banach Spaces (V. Barbu and Th. Precipanu). SIAM Review, 22(3): 375–376, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Datta:1982:SLE

[Dat82] B. N. Datta. On the solution of a Lyapunov equation. SIAM Review, 24(3): 339, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Datta:1986:BRE

[Dat86] Subhendu K. Datta. Book review: The Excitation and Propagation of Elastic Waves (J. A. Hudson). SIAM Review, 28(3):413– 414, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Datta:1990:BRW

[Dat90] Subhendu K. Datta. Book review: Wave Propagation in Solids and Fluids (J. L. Davis). SIAM Review, 32(3): 500–??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dauer:1990:BRI

[Dau90]  Jerald P. Dauer.  Book

review: Iterative Aggregation Theory. Mathematical Methods of Coordinating Detailed and Aggregate Problems in Large Control Systems (Lev M. Dudkin, Hya Rabinovich, and Ilya Vakhutinsky). SIAM Review, 32(1):168–172, March 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davidon:1975:BRD

[Dav75a] William C. Davidon. Book review: Davidon’s Cubic Interpolation (S. K. Park and T. A. Straeter). SIAM Review, 17(1):170–171, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davidson:1975:DCI

[Dav75b] W. C. Davidson. Davidon’s cubic interpolation. SIAM Review, 17(1):170– 171, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1975:BRN

[Dav75c] Paul Wm. Davis. Book review: Nonlinear and Random Vibrations (Florea Dinca and Christian Teodosiu). SIAM Review, 17(3):578– 579, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1977:BRA

[Dav77] Martin Davis. Book review: Algorithms and Recursive Functions (A. I. Mal’cev). SIAM Review, 19 (3):573, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1989:BRP

[Dav89] Philip J. Davis. Book review: Perturbation Bounds for Matrix Eigenvalues (Rajendra Bhatia). SIAM Review, 31(1):133, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1991:BRN

[Dav91a] Jon H. Davis. Book review: Nonnegative Matrices in Dynamic Systems (Abraham Berman, Michael Neumann, and Donald J. Stern). SIAM Review, 33(2): 300–301, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1991:BRA

[Dav91b] Paul Davis. Book review: Applied Mathematics: a Contemporary Approach (J. David Logan). SIAM Review, 33(1):137–139, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1994:BRI

[Dav94a] A. M. J. Davis. Book review: An Infinite Integral (J. Billingham). SIAM Review, 36(1):115, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1994:MKD

[Dav94b] P. Davis. Murray Klamkin — dedication. SIAM Review, 36(1):U1, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dawson:1974:BRR

[Daw74] Donald A. Dawson. Book review: Random Integral Equations (A. T. Bharucha-Reid). SIAM Review, 16(2): 266–268, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dax:1990:CLS

[Dax90] Achiya Dax. The convergence of linear stationary iterative processes for solving singular unstructured systems of linear equations. SIAM Review, 32(4):611– 635, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dax:1997:EPF

[Dax97a] A. Dax. An elementary proof of Farkas’ lemma. SIAM Review, 39(3):503– 507, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dax:1997:CNE

[Dax97b]  Achiya Dax.  Classroom

note: An elementary proof of Farkas’ lemma. SIAM Review, 39(3):503–507, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

                                        http://epubs.siam.org/
                                        sam-bin/dbq/article/29550.

Daykin:1971:FD

[Day71] D. E. Daykin. Falling dominoes. SIAM Review, 13 (4):569, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Day:1987:BRL

[Day87]  W. A. Day.  Book re-

view: Lectures on Viscoelasticity Theory (A. C. Pipkin). SIAM Review, 29(3): 477–478, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Day:1989:BRB

[Day89] W. A. Day. Book review: Boundary Value Problems in Linear Viscoelasticity (J. M.

Golden and G. A. C. Graham). SIAM Review, 31 (2):339, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deBruijn:1964:AS

[dB64] N. G. de Bruijn. An asymptotic series. SIAM Review, 6(2):177, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deBruijn:1965:AS

[dB65a] N. G. de Bruijn. An asymptotic series. SIAM Review, 7(4):564–565, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deBruijn:1965:BRA

[dB65b] N. G. de Bruijn. Book review: Asymptotic Distribution of Lattice Points in a Random Rectangle (Walter Weissblum). SIAM Review, 7 (2):274–276, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dhir:1970:DIC

[DB70] S. K. Dhir and J. S. Brock.

A definite integral containing three parameters. SIAM Review, 12(3):468, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deBoor:1974:BES

[dB74] Carl de Boor. Bounding the error in spline interpolation. SIAM Review, 16(4): 531–544, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dube:1978:SIV

[DB78] R. P. Dube and J. Brown. Smooth interpolation with vertex functions. SIAM Review, 20(3):623, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

diBernardo:2008:BND

[dBC^{+}08] Mario di Bernardo, Chris J. Budd, Alan R. Champneys, Piotr Kowalczyk, Arne B. Nordmark, Gerard Olivar Tost, and Petri T. Piiroinen. Bifurcations in nonsmooth dynamical systems. SIAM Review, 50(4): 629–701, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deBoor:1995:BS

[dBHRR95] C. de Boor, K. Hollig, S. Riemenschneider, and Amos Ron. Box splines. SIAM Review, 37(2):275– ??, June 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deBoer:1975:ETS

[dBL75] P. C. T. de Boer and G. S. S.

Ludford. Existence of a transition solution. SIAM Review, 17(1):169, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dirks:2007:TAI

[DBS^{+}07] Robert M. Dirks, Justin S.

Bois, Joseph M. Schaeffer, Erik Winfree, and Niles A. Pierce. Thermodynamic analysis of interacting nucleic acid strands. SIAM Review, 49(1):65–88, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Debnath:1976:NMM

[DC76] L. Debnath and A. K. Chatterjee. Nonlinear mathematical model of junctions of tidal rivers. SIAM Review, 18 (4):800, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dales:1986:BRO

[DC86] H. G. Dales and D. G.

Crighton. Book review: Operational Calculus — a Theory of Hyperfunctions (K. Yosida). SIAM Review, 28 (1):98–99, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Derriennic:1986:LBP

[DCL^{+}86] M. M. Derriennic, G. Chang, E. Lee, R. Rice, and H. P. Yang. A limit of Bernstein polynomials. SIAM Review, 28(1):89–90, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Doleans-Dads:1974:BRM

[DD74] C. Doléans-Dads. Book review: Martingales à Temps Discret (Jacques Neveu). SIAM Review, 16(3):398– 399, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeSchutter:2002:DSV

[DD02] Bart De Schutter and Bart De Moor. The QR decomposition and the singular value decomposition in the symmetrized max-plus algebra revisited. SIAM Review, 44 (3):417–454, September 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40396.

Davison:2004:HPP

[DD04] Matt Davison and Andrea Doeschl. A hyperbolic PDE with parabolic behavior. SIAM Review, 46 (1):115–127, March 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40900.

Dolev:2003:NC

[DDN03] Danny Dolev, Cynthia Dwork, and Moni Naor. Nonmalleable cryptography. SIAM Review, 45(4):727–784, December 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/42985.

DeSimone:2022:SAI

[DdSG^{+}22] Valentina De Simone, Daniela di Serafino, Jacek Gondzio, Spyridon Pougkakiotis, and Marco Viola. Sparse approximations with interior point methods. SIAM Review, 64(4):954–988, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1401103.

DeBoer:1974:BRI

[De 74] F. J. De Boer. Book review: An Integral Identity (W. H. H. Banks and P. G. Drazin). SIAM Review, 16 (2):259–260, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeDoelder:1979:TIS

[De 79] P. J. De Doelder. Two infinite sums. SIAM Review, 21 (3):395–396, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeKleine:1984:NAS

[De 84] H. Arthur De Kleine. A note on the asymptotic stability of periodic solutions of autonomous differential equations. SIAM Review, 26(3): 417–421, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeYoung:2000:ERD

[De 00] Gary W. De Young. Exploring reflection: Designing light reflectors for uniform illumination. SIAM Review, 42 (4):727–735, December 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35885.

DeLeenheer:2020:EPM

[De 20] Patrick De Leenheer. An elementary proof of a matrix tree theorem for directed graphs. SIAM Review, 62(3): 716–726, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deakin:1989:SFE

[Dea89] Michael A. B. Deakin. Stokes flow: an elementary treatment. SIAM Review, 31 (4):667–670, December 1989.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See addendum [Dea90b].

Deakin:1990:A

[Dea90a] M. A. B. Deakin. Addition. SIAM Review, 32(4): 671, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deakin:1990:ASF

[Dea90b] Michael A. B. Deakin. Addendum: “Stokes flow: an elementary treatment” [SIAM rev. 31 (1989), no. 4, 667– 670 and MR 90k:76043]. SIAM Review, 32(4):671, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Dea89].

Deboor:1976:BAS

[Deb76]  C. Deboor.  L_{∞} bound

for L_{2}-approximation by splines. SIAM Review, 18 (4):800, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Decell:1965:AFG

[Dec65a] Henry P. Decell, Jr. An alternate form of the generalized inverse of an arbitrary complex matri. SIAM Review, 7 (3):356–358, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Decell:1965:ACH

[Dec65b] Henry P. Decell, Jr. An application of the Cayley– Hamilton theorem to generalized matrix inversion. SIAM Review, 7(4):526– 528, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Defigueiredo:1976:S

[Def76] R. J. P. Defigueiredo. L_{m}- G splines. SIAM Review, 18 (4):800, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dehesa:1989:UBF

[Deh89]  J. S. Dehesa.  Upper

bounds for frequency moments. SIAM Review, 31(3): 494–495, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deift:1992:BRD

[Dei92] Percy Deift. Book review: Direct and Inverse Problems: Potentials in Quantum Scattering (Boris Zakhariev and Allina Suzko; G. Pontecorvo, trans.). SIAM Review, 34(1): 141–142, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

dAspremont:2007:DFS

[dEJL07] Alexandre d’Aspremont, Laurent El Ghaoui, Michael I.

Jordan, and Gert R. G. Lanckriet. A direct formulation for sparse PCA using semidefinite programming. SIAM Review, 49(3): 434–448, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Delillo:1995:BRM

[Del95] Thomas K. Delillo. Book review: Moving-Grid Methods for Time-Dependent Partial Differential Equations (P. A. Zegeling). SIAM Review, 37 (1):120–121, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Delgado:1997:LCM

[Del97a] M. Delgado. The Lagrange– Charpit method. SIAM Review, 39(2):298–304, ???? 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Delgado:1997:CNL

[Del97b] Manuel Delgado. Classroom note: The Lagrange–Charpit method. SIAM Review, 39(2):298–304, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29353.

DeMar:1965:BRP

[DeM65] Richard DeMar. Book review: Polynomial Expansions of Analytic Functions, Second Edition (Ralph P. Boas, Jr. and R. C. Buck). SIAM Review, 7(2):301– 302, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Demetrius:1974:IIP

[Dem74] Lloyd Demetrius. Invariant ideals of positive operators. SIAM Review, 16(4): 428–440, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Demmel:1986:BRM

[Dem86] James W. Demmel. Book review: Matrix Computations (Gene H. Golub and Charles F. van Loan). SIAM Review, 28(2):252–255, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Demmel:1990:BRM

[Dem90] James W. Demmel. Book review: Matrix Computations; Second Edition (Gene Golub and Charles F. Van Loan). SIAM Review, 32 (4):690–691, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Denardo:1967:CMT

[Den67] Eric V. Denardo. Contraction mappings in the theory underlying dynamic programming. SIAM Review, 9 (2):165–177, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Denman:1968:IM

[Den68] Harry Denman. The inverse of a matrix. SIAM Review, 10(3):376, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Denman:1969:IM

[Den69] Harry Denman. The inverse of a matrix. SIAM Review, 11 (4):630–632, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dendy:1976:ASA

[Den76] J. E. Dendy. Analysis of some alternating direction methods for some nonlinear timedependent problems. SIAM Review, 18(4):801, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Denning:1978:BRD

[Den78] Peter J. Denning. Book review: Decomposability. Queueing and Computer System Applications (P. J. Courtois). SIAM Review, 20(4): 871–873, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dennis:1982:BRP

[Den82] John E. Dennis, Jr. Book review: Practical Methods of Optimization, Vol. 1: Unconstrained Optimization (R. Fletcher). SIAM Review, 24 (1):97–98, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deng:1990:EMO

[Den90] Lih-Yuan Deng. Elements of maximum order in a matrix group. SIAM Review, 32(3):479, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2017:Ea

[Den17a] Darinka Dentcheva. Education. SIAM Review, 59 (1):197, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2017:Eb

[Den17b] Darinka Dentcheva. Education. SIAM Review, 59(2): 391–392, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2017:Ec

[Den17c] Darinka Dentcheva. Education. SIAM Review, 59(3): 647–648, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2017:Ed

[Den17d] Darinka Dentcheva. Education. SIAM Review, 59 (4):847, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2018:Ea

[Den18a] Darinka Dentcheva. Education. SIAM Review, 60(1): 179–180, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2018:Eb

[Den18b] Darinka Dentcheva. Education. SIAM Review, 60(2): 427–428, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2018:Ec

[Den18c] Darinka Dentcheva. Education. SIAM Review, 60(3): 705–706, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2018:Ed

[Den18d] Darinka Dentcheva. Education. SIAM Review, 60 (4):967, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2019:Ea

[Den19a] Darinka Dentcheva. Education. SIAM Review, 61(1): 157–158, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2019:Eb

[Den19b] Darinka Dentcheva. Education. SIAM Review, 61(2): 345–346, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2019:Ec

[Den19c] Darinka Dentcheva. Education. SIAM Review, 61(3): 547–548, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2019:Ed

[Den19d] Darinka Dentcheva. Education. SIAM Review, 61(4): 837–838, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2020:Ea

[Den20a] Darinka Dentcheva. Education. SIAM Review, 62(1): 229–230, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2020:Eb

[Den20b] Darinka Dentcheva. Education. SIAM Review, 62(2): 437–438, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2020:Ec

[Den20c] Darinka Dentcheva. Education. SIAM Review, 62 (3):683, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2020:Ed

[Den20d] Darinka Dentcheva. Education. SIAM Review, 62(4): 899–900, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2021:Ea

[Den21a] Darinka Dentcheva. Education. SIAM Review, 63(1): 165–166, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2021:Eb

[Den21b] Darinka Dentcheva. Education. SIAM Review, 63 (2):393, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2021:Ec

[Den21c] Darinka Dentcheva. Education. SIAM Review, 63 (3):583, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2021:Ed

[Den21d] Darinka Dentcheva. Education. SIAM Review, 63 (4):823, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dentcheva:2022:Ea

[Den22a] Darinka Dentcheva. Education. SIAM Review, 64 (1):179, ???? 2022. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/22N975408.

Dentcheva:2022:Eb

[Den22b] Darinka Dentcheva. Education. SIAM Review, 64 (2):467, ???? 2022. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/22N975457.

Dentcheva:2022:Ec

[Den22c] Darinka Dentcheva. Education. SIAM Review, 64 (3):711–712, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975536.

Dentcheva:2022:Ed

[Den22d] Darinka Dentcheva. Education. SIAM Review, 64 (4):1029, ???? 2022. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/22N975585.

Deo:1964:GT

[Deo64] Narsingh Deo. Graph theory. SIAM Review, 6 (3):310, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeLillo:2001:SCM

[DEP01] T. K. DeLillo, A. R. Elcrat, and J. A. Pfaltzgraff. Schwarz–Christoffel mapping of the annulus. SIAM Review, 43(3):469– 477, September 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37528.

Dunn:1995:MCL

[DEPC95] G. Dunn, B. Everitt, A. Pickles, and Wai Chan. Modelling covariances and latent variables using EQS. SIAM Review, 37(3):462– ??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Derby:1971:SSF

[Der71] Royce Derby. Special subsets of a finite group. SIAM Review, 13(2):246, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Derriennic:1986:BRL

[Der86] M. M. Derriennic. Book review: A Limit of Bernstein Polynomials (G. Chang). SIAM Review, 28(1):89– 90, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeRose:1991:BRC

[DeR91a] Tony DeRose. Book review: Curves and Surfaces for Computer Aided Geometric Design: a Practical Guide (Gerald Farin). SIAM Review, 33(2):301– 302, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeRose:1991:CSC

[DeR91b] Tony DeRose. Curves and surfaces for computer aided geometric design: a practical guide (Gerald Farin). SIAM Review, 33(2):301– ??, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Descloux:1968:BRP

[Des68]  A. Descloux.  Book re-

view: Proceedings of the Symposium on Congestion Theory (Walter L. Smith and William E. Wilkinson, eds.). SIAM Review, 10(3): 388–391, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeSanti:1988:MPA

[DeS88] Albert J. DeSanti. A model for predicting aircraft altitude loss in a pull-up from a dive. SIAM Review, 30 (4):625–628, December 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dettman:1985:BRP

[Det85] John W. Dettman. Book review: Partial Differential Equations of Applied Mathematics (Erich Zauderer). SIAM Review, 27(1): 96–97, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deutsch:1963:BRP

[Deu63] Ralph Deutsch. Book review: Principles of the Statistical Theory of Communication (W. W. Harman). SIAM Review, 5(4): 381–382, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deutsch:1981:DPM

[Deu81a] E. Deutsch. Determinant of a partitioned matrix. SIAM Review, 23(2):259, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deutsch:1981:LZP

[Deu81b] E. Deutsch. Location of the zeros of a polynomial. SIAM Review, 23(1):100, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deutsch:1981:BRD

[Deu81c]  G. Deutsch.  Book re-

view: Determinant of a Partitioned Matrix (M. S. Klamkin, A. Sharma and P. W. Smith). SIAM Review, 23(2):259, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deuflhard:1985:RPE

[Deu85]  P. Deuflhard.  Recent

progress in extrapolation methods for ordinary differential equations. SIAM Review, 27(4):505–535, December 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deutsch:1987:AAT

[Deu87] Moshe Deutsch. The attenuated Abel transform. SIAM Review, 29(1):121, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deutsch:1990:BRN

[Deu90] Frank Deutsch. Book review: Nonlinear Approximation Theory (Dietrich Braess). SIAM Review, 32 (1):172, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deutsch:1995:RLS

[Deu95]  Sid Deutsch.  A reso-

nant line structure consisting of rational right triangles. SIAM Review, 37(1): 90–92, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/27020.htm.

Devaney:1990:BRP

[Dev90] Robert L. Devaney. Book review: Practical Numerical Algorithms for Chaotic Systems (T. S. Parker and L. O. Chua). SIAM Review, 32 (3):501–503, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Devaney:1991:BRM

[Dev91] Robert L. Devaney. Book review: Measure, Topology, and Fractal Geometry (Gerald A. Edgar). SIAM Review, 33(4):668– 669, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Devaney:1992:BRF

[Dev92] Robert L. Devaney. Book review: Fractals: An Animated Discussion with Edward Lorenz and Benoît Mandelbrot (H.-O. Peitgen, H. Jurgens, D. Saupe, and C. Zahlten). SIAM Review, 34 (2):333–335, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Devaney:1994:BRC

[Dev94a] R. L. Devaney. Book review: Complex Dynamics (Lennart Carleson and Theodore W. Gamelin). SIAM Review, 36 (3):504–??, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Devaney:1994:BRR

[Dev94b] Robert L. Devaney. Book review: Rational Iteration: Complex Analytic Dynamical Systems (Norbert Steinmetz). SIAM Review, 36 (3):504–505, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeWitt:1985:BRM

[DeW85] Bryce DeWitt. Book review: The Mathematical Theory of Black Holes (S. Chandrasekhar). SIAM Review, 27 (1):97–99, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DiBucchianico:1998:PAA

[DF98] A. Di Bucchianico and Philip Feinsilver. Probabilistic and analytical aspects of the umbral calculus. SIAM Review, 40(1):147–??, March 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Diaconis:1999:IRF

[DF99] Persi Diaconis and David Freedman. Iterated random functions. SIAM Review, 41(1):45–76, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33844.

Du:1999:CVT

[DFG99] Qiang Du, Vance Faber, and Max Gunzburger. Centroidal Voronoi tessellations: Applications and algorithms. SIAM Review, 41(4):637– 676, December 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35283.

Djokovic:1968:BRE

[DG68] D. Z. Djoković and W. B.

Gragg. Book review: Elementary Symmetric Polynomials (Paul J. Schweitzer). SIAM Review, 10(3):379– 380, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeMeyer:1996:BRN

[DG96] H. E. De Meyer and C. C. Grosjean. Book review: Nonsymmetric Cyclic Pursuit on a Sphere (M. S. Klamkin). SIAM Review, 38 (1):153–155, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Drexler:1998:CB

[DG98] Michael Drexler and Martin J. Gander. Circular billiard. SIAM Review, 40(2):315–323, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31087.

Dubeau:2014:FPN

[DG14] François Dubeau and Calvin Gnang. Fixed point and Newton’s methods for solving a nonlinear equation: From linear to high-order convergence. SIAM Review, 56(4): 691–708, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duersch:2020:RPR

[DG20] Jed A. Duersch and Ming Gu. Randomized projection for rank-revealing matrix factorizations and low-rank approximations. SIAM Review, 62(3):661–682, ???? 2020. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dongarra:2018:SVD

[DGH^{+}18] Jack Dongarra, Mark Gates, Azzam Haidar, Jakub Kurzak, Piotr Luszczek, Stanimire Tomov, and Ichitaro Yamazaki. The singular value decomposition: anatomy of optimizing an algorithm for extreme scale. SIAM Review, 60(4):808–865, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dongarra:1984:ILA

[DGK84] J. J. Dongarra, F. G. Gustavson, and A. Karp. Implementing linear algebra algorithms for dense matrices on a vector pipeline machine. SIAM Review, 26(1): 91–112, January 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Desroches:2012:MMO

[DGK^{+}12] Mathieu Desroches, John Guckenheimer, Bernd Krauskopf, Christian Kuehn, Hinke M. Osinga, and Martin Wechselberger. Mixed-mode oscillations with multiple time scales. SIAM Review, 54(2): 211–288, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Druskin:2016:NOP

[DGK16] Vladimir Druskin, Stefan Güttel, and Leonid Knizhnerman. Near-optimal perfectly matched layers for indefinite Helmholtz problems. SIAM Review, 58(1): 90–116, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Du:2012:AAN

[DGLZ12] Qiang Du, Max Gunzburger, R. B. Lehoucq, and Kun Zhou. Analysis and approximation of nonlocal diffusion problems with volume constraints. SIAM Review, 54 (4):667–696, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Du:1992:AAG

[DGP92] Qiang Du, Max D. Gunzburger, and Janet S. Peterson. Analysis and approximation of the Ginzburg– Landau model of superconductivity. SIAM Review, 34 (1):54–81, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Desroches:2016:CFN

[DGP^{+}16] Mathieu Desroches, Antoni Guillamon, Enrique Ponce, Rafael Prohens, Serafim Rodrigues, and Antonio E. Teruel. Canards, folded nodes, and mixed-mode oscillations in piecewise-linear slow-fast systems. SIAM Review, 58(4):653–691, ???? 2016. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dawande:2005:DCS

[DGS05] Milind W. Dawande, H. Neil Geismar, and Suresh P. Sethi. Dominance of cyclic solutions and challenges in the scheduling of robotic cells. SIAM Review, 47 (4):709–721, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44138.

Dekking:2009:PAI

[DH09] Michel Dekking and André Hensbergen. A problem with the assessment of an Iris identification system. SIAM Review, 51(2):417– 422, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dunning:2017:JML

[DHL17] Iain Dunning, Joey Huchette, and Miles Lubin. JuMP: a modeling language for mathematical optimization. SIAM Review, 59(2):295– 320, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Diaconis:2007:DBC

[DHM07] Persi Diaconis, Susan Holmes, and Richard Montgomery. Dynamical bias in the coin toss. SIAM Review, 49(2): 211–235, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Diaz:1959:BRD

[Dia59] J. B. Diaz. Book review: Differential Equations: Geometric Theory (Solomon Lefschetz). SIAM Review, 1(2): 179–181, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Diaz:1964:BRM

[Dia64]  J. B. Diaz.  Book re-

view: Methods of Mathematical Physics, Volume II (R. Courant and D. Hilbert). SIAM Review, 6(4):463– 466, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Diaz:1978:MPE

[Día78] J. B. Díaz. Mathematical prolegomena to every theory of homogeneous heat engines. SIAM Review, 20(2): 265–277, April 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dickson:1967:BRC

[Dic67] L. E. Dickson. Book review: On Certain Systems Matrices (A. A. Mullin). SIAM Review, 9(4):750– 751, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dickson:1975:AP

[Dic75] L. J. Dickson. Accident probability. SIAM Review, 17 (3):565, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Diederich:1962:LSS

[Die62] Franklin W. Diederich. A least-squares solution of a linear system. SIAM Review, 4(2):150, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dieci:1992:BRR

[Die92]  Luca Dieci.  Book re-

view: The Riccati Equation (S. Bittanti, A. J. Laub, and J. C. Willems, eds.). SIAM Review, 34(4):665– ??, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dijkstra:1990:BRL

[Dij90] Hendrik A. Dijkstra. Book review: Low-Gravity Fluid Mechanics: Mathematical Theory of Capillary Phenomena (A. D. Myshkis, V. G. Babskii, N. D. Kopachevskii, L. A. Slobozhanin, and A. D. Tyuptsov; R. S. Wadhwa, trans.). SIAM Review, 32 (4):691–693, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dimock:1991:BRF

[Dim91a]  J. Dimock.  Book re-

view: Functional Analysis: an Introduction for Physicists (Nino Boccara). SIAM Review, 33(3):509– 510, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dimock:1991:FAI

[Dim91b] J. Dimock. Functional analysis: An introduction for physicists (Nino Boccara). SIAM Review, 33(3):509– ??, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dinkel:1972:BRN

[Din72]  John Dinkel.  Book re-

view: Numerical Methods of Mathematical Optimization with ALGOL and FORTRAN Programs (Hans P. Kunzi, H.

G. Tzschach and C. A. Zehnder). SIAM Review, 14(2): 359–361, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dinwoodie:1994:BRL

[Din94a]  I. H. Dinwoodie.  Book

review: Large Deviations Techniques and Applications (Amir Dembo and Ofer Zeitouni). SIAM Review, 36 (2):303–304, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dinwoodie:1994:LDT

[Din94b] I. H. Dinwoodie. Large deviations techniques and applications (Amir Dembo and Ofer Zeitouni). SIAM Review, 36 (2):303–??, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Disci:1992:BRR

[Dis92] Luca Disci. Book review:

The Riccati Equation (S. Bittanti, A. J. Laub, and J C. Willems). SIAM Review, 34 (4):665–666, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dixon:1991:BRP

[Dix91] Philip M. Dixon. Book review: Parameter Estimation in Ecology: The Link between Data and Models (O.

Richter and D. Söndgerath). SIAM Review, 33(4):687– 688, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deJager:1998:TSP

[dJFO98] E. M. de Jager, Jiang Furu, and Robert E. O’Malley, Jr. The theory of singular perturbations. SIAM Review, 40 (1):168–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Du:2021:MBP

[DJLQ21] Qiang Du, Lili Ju, Xiao Li, and Zhonghua Qiao. Maximum bound principles for a class of semilinear parabolic equations and exponential time-differencing schemes. SIAM Review, 63 (2):317–359, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Djokovic:1968:BRT

[Djo68] D. Z. Djoković. Book review: On Two Matrices (Olga Taussky). SIAM Review, 10(3):379, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Disney:1985:QNS

[DK85] Ralph L. Disney and Dieter König. Queueing networks: a survey of their random processes. SIAM Review, 27 (3):335–403, September 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dassios:1989:KIL

[DK89] George Dassios and R. E.

Kleinman. On Kelvin inversion and low-frequency scattering. SIAM Review, 31 (4):565–585, December 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeOliveira:1997:BRM

[DK97] Victor De Oliveira and Benjamin Kedem. Book review: Multivariate Geostatistics, by Hans Wackernagel. SIAM Review, 39(2): 340–341, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duncan:1996:RPE

[DKS96] J. Duncan, D. Khavinson, and H. Shapiro. Rectangular parallelepipeds in ellipsoids. SIAM Review, 38 (4):655–657, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29301.

Dodwell:2019:MMC

[DKST19] T. J. Dodwell, C. Ketelsen, R. Scheichl, and A. L. Teckentrup. Multilevel Markov Chain Monte Carlo. SIAM Review, 61(3): 509–545, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Datta:2009:OPM

[DKW^{+}09] Kaushik Datta, Shoaib Kamil, Samuel Williams, Leonid Oliker, John Shalf, and Katherine Yelick. Optimization and performance modeling of stencil computations on modern microprocessors. SIAM Review, 51(1): 129–159, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Decell:1966:NIR

[DL66] Henry P. Decell, Jr. and Robert N. Lea. Numerical integration and Riemann– Stieltjes sums. SIAM Review, 8(2):196–200, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Diprima:1976:ESF

[DL76] R. C. Diprima and N. Liron. Effect of secondary flow on decaying oscillations of a sphere and a disk. SIAM Review, 18(4):801, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dontchev:1992:DMD

[DL92] Asen Dontchev and Frank Lempio. Difference methods for differential inclusions: a survey. SIAM Review, 34(2): 263–294, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Davis:1995:MMO

[DL95] M. H. A. Davis and Suzanne Lenhart. Markov models and optimization. SIAM Review, 37(2):262–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deutsch:1967:AHB

[DM67] F. R. Deutsch and P. H.

Maserick. Applications of the Hahn–Banach theorem in approximation theory. SIAM Review, 9(3): 516–530, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dennis:1977:QNM

[DM77] J. E. Dennis, Jr. and Jorge J.

Moré. Quasi-Newton methods, motivation and theory. SIAM Review, 19(1): 46–89, January 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deMeyer:1994:BRM

[dM94] H. de Meyer. Book review:

A Maximum Volume (Bill Sands). SIAM Review, 36(2): 278–281, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Driscoll:2007:SRG

[DM07] Tobin A. Driscoll and Kara L. Maki. Searching for rare growth factors using multicanonical Monte Carlo methods. SIAM Review, 49(4): 673–692, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Devriendt:2022:VCD

[DMGL22] Karel Devriendt, Samuel Martin-Gutierrez, and Renaud Lambiotte. Variance and covariance of distributions on graphs. SIAM Review, 64(2):343–359, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://epubs.siam.org/
doi/10.1137/20M1361328.

Durmus:2022:PMC

[DMP22] Alain Durmus, Éric Moulines, and Marcelo Pereyra. A proximal Markov chain Monte Carlo method for Bayesian inference in imaging inverse problems: When Langevin meets Moreau. SIAM Review, 64(4):991–1028, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://epubs.siam.org/
doi/10.1137/22M1522917.

Leenheer:2023:CPB

[DMS23] Patrick De Leenheer, John Musgrove, and Tyler Schimleck. A comprehensive proof of Bertrand’s Theorem. SIAM Review, 65 (2):563–588, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1436658.

Dinner:2018:TSS

[DMT^{+}18] Aaron R. Dinner, Jonathan C.

Mattingly, Jeremy O. B. Tempkin, Brian Van Koten, and Jonathan Weare. Trajectory stratification of stochastic dynamics. SIAM Review, 60(4):909–938, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deMelo:1995:ODD

[dMvSB95] W. de Melo, S. van Strien, and K. M. Brucks. Onedimensional dynamics. SIAM Review, 37(3):461–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dent:1959:POD

[DN59] Benjamin A. Dent and Albert Newhouse. Polynomials orthogonal over discrete domains. SIAM Review, 1 (1):55–59, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

deOliveira:1969:CSM

[dO69] G. N. de Oliveira. On a certain system of matrix equations. SIAM Review, 11(3): 378–381, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Daalhuis:1998:ANS

[DO98] A. B. Olde Daalhuis and F. W. J. Olver. On the asymptotic and numerical solution of linear ordinary differential equations. SIAM Review, 40(3):463– 495, September 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31534.

Doberkat:1984:HV

[Dob84] E. E. Doberkat. A hypervolume. SIAM Review, 26 (4):580, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Doberkat:1987:BRF

[Dob87] Ernst-Erich Doberkat. Book review: Fundamentals of the Average Case Analysis of Particular Algorithms (Rainer Kemp). SIAM Review, 29(2):311–314, ????

1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dodge:1973:BRF

[Dod73] Philip G. Dodge, Jr. Book review: Foundations o f the Theory o f Plasticity (L. M. Kachanov). SIAM Review, 15 (2):400–401, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dolph:1962:BRE

[Dol62] C. L. Dolph. Book review: Electromagnetic Waves (R. E. Langer, ed.). SIAM Review, 4(4):409–411, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dolph:1972:BRE

[Dol72] C. L. Dolph. Book review:

Equations of Mathematical Physics (V. S. Vladimirov). SIAM Review, 14(1):186– 187, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dold:1991:BRM

[Dol91a] J. W. Dold. Book review:

Mathematical Problems from Combustion Theory (Jerrold Bebernes and David Eberly). SIAM Review, 33(3):484– 486, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dold:1991:MPC

[Dol91b] J. W. Dold. Mathematical problems from combustion theory (Jerrold Bebernes and David Eberly). SIAM Review, 33(3):484– ??, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dold:1993:NIP

[Dol93] J. W. Dold. A note on integral properties of periodic orbits. SIAM Review, 35(1): 125–129, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dongarra:1985:BRN

[Don85] Jack Dongarra. Book review: A New Approach to Scientific Computation (Ulrich W. Kulisch and Willard L. Miranker, eds.). SIAM Review, 27(2):267–268, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dontchev:1990:BRO

[Don90] Asen L. Dontchev. Book review: Optimization and Identification of Systems Governed by Evolution Equations on Banach Space (N. U. Ahmed). SIAM Review, 32(3):503, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Donati:2023:ICP

[Don23] Martin Donati. Improbability of collisions of pointvortices in bounded planar domains. SIAM Review, 65(1):227–257, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1523340.

Doob:1963:BRS

[Doo63]  J. L. Doob.  Book re-

view: Stochastic Processes (E. Parzen). SIAM Review, 5 (2):172–173, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dorr:1970:DSD

[Dor70] Fred W. Dorr. The direct solution of the discrete Poisson equation on a rectangle. SIAM Review, 12(2): 248–263, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dorr:1973:DSD

[Dor73] F. W. Dorr. The direct solution of the discrete Poisson equation in O(n^{2}) operations. SIAM Review, 15(??): 412–415, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dorr:1974:BRD

[Dor74] Fred W. Dorr. Book review: Dynamic programming and Partial Differential Equations (Edward Angel and Richard Bellman). SIAM Review, 16(1):104– 105, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dorr:1975:DSD

[Dor75] Fred W. Dorr. The direct solution of the discrete Poisson equation in O(N^{2}) operations. SIAM Review, 17(3): 412–415, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dorfman:1989:BRB

[Dor89] J. R. Dorfman. Book review: The Boltzmann Equation and Its Applications (Carlo Cercignani). SIAM Review, 31 (2):339–340, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dantzig:1976:API

[DP76] G. B. Dantzig and S. C. Parikh. Assessing physical impact on economy of a changing energy picture. SIAM Review, 18(4):799– 800, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dorr:1973:AMP

[DPS73] F. W. Dorr, S. V. Parter, and L. F. Shampine. Applications of the maximum principle to singular perturbation problems. SIAM Review, 15 (1):43–88, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Draper:1961:BRG

[Dra61a] Norman R. Draper. Book review: Great Ideas of Modern Mathematics: Their Nature and Use (Jagjit Singh). SIAM Review, 3(3):267– 268, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Draper:1961:BRQ

[Dra61b] Norman R. Draper. Book review: Quantitative Methods in Pharmacology (H. de Jonge). SIAM Review, 3(4): 339–340, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Draper:1962:BRI

[Dra62] Norman R. Draper. Book review: Introduction to Probability and Random Variables (G. P. Wadsworth and J. G. Bryan). SIAM Review, 4 (2):158, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Drager:1983:STR

[Dra83] Lance D. Drager. A simple theorem on Riemann integration, based on classroom experience. SIAM Review, 25 (2):255–259, April 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Drenick:1960:RPS

[Dre60] R. F. Drenick. On the role of professional societies in stimulating and guiding of researchi. SIAM Review, 2(4): 241–246, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Drew:1975:TFC

[Dre75] D. A. Drew. Topographically forced component of atmospheric motion. SIAM Review, 17(2):377, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Drew:1976:ELF

[Dre76] D. A. Drew. Effect of lift force on stability of uniform fluidization. SIAM Review, 18(4):801, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Drew:1989:BRL

[Dre89] Donald Drew. Book review: Learning the Art of Mathematical Modelling (Cross and A. O. Moscardini); Concepts of Mathematical Modeling (Walter J. Meyer). SIAM Review, 31(2): 340–341, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Driver:1963:BRD

[Dri63] Rodney D. Driver. Book review: Differential Equations (L. E. El’sgol’ts). SIAM Review, 5(1):79–80, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Driver:1965:BRM

[Dri65] Rodney D. Driver. Book review: Methods of A. M. Lyapunov and Their Application (V. I. Zubov). SIAM Review, 7(4):570–571, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Driver:1968:RAC

[Dri68]  Rodney D. Driver.  On

Ryabov’s asymptotic characterization of the solutions of quasi-linear differential equations with small delays. SIAM Review, 10(3): 329–341, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Driver:1975:BRI

[Dri75] R. D. Driver. Book review: Introduction to the Theory and Application of Differential Equations with Deviating Arguments (L. E. El’sgol’ts and S. B. Norkin). SIAM Review, 17(1):178– 179, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Driver:1981:BRI

[Dri81] R. D. Driver. Book review: Integrodifferential Equations and Delay Models in Population Dynamics (J. M. Gushing). SIAM Review, 23(3): 411, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Driessen:1991:SCP

[Dri91] Theo S. H. Driessen. A survey of consistency properties in cooperative game theory. SIAM Review, 33(1): 43–59, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Driscoll:1997:EID

[Dri97] Tobin A. Driscoll. Eigenmodes of isospectral drums. SIAM Review, 39(1):1–17, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28506;   http:
//www.siam.org/journals/
sirev/sirev391.htm.

Drucker:1964:BRL

[Dru64] D. C. Drucker. Book review: Limit Analysis of Rotationally Symmetric Plates and Shells (Philip G. Dodge). SIAM Review, 6(2):188, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Deutsch:1964:BRS

[DRWK64] E. Deutsch, Thomas Rogge, J. Ernest Wilkins, Jr., and M. S. Klamkin. Book review: A Steady-State Temperature (Alan L. Tritter). SIAM Review, 6(2):178– 180, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dunn:1961:FSL

[DS61] Eldon L. Dunn and F. Max Stein. Families of Sturm– Liouville systems. SIAM Review, 3(1):54–65, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Devaney:1973:PWR

[DS73] A. J. Devaney and George C.

Sherman. Plane-wave representations for scalar wave fields. SIAM Review, 15(4): 765–786, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dennis:1979:LCS

[DS79] J. E. Dennis, Jr. and R. B. Schnabel. Least change secant updates for quasi-Newton methods. SIAM Review, 21(4):443–459, October 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dorman:2004:GBP

[DSL04] Karin S. Dorman, Janet S. Sinsheimer, and Kenneth Lange. In the garden of branching processes. SIAM Review, 46(2):202–229, June 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/41784.

Dantzig:1998:LPI

[DTG98] George B. Dantzig, Mukund N. Thapa, and Saul I. Gass. Linear programming 1: Introduction. SIAM Review, 40 (1):149–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Driscoll:1998:PTM

[DTT98] Tobin A. Driscoll, Kim-Chuan C. Toh, and Lloyd N. Trefethen. From potential theory to matrix iterations in six steps. SIAM Review, 40 (3):547–578, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30558.

Deconinck:2014:MFS

[DTV14] Bernard Deconinck, Thomas Trogdon, and Vishal Vasan. The method of Fokas for solving linear partial differential equations. SIAM Review, 56 (1):159–186, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duff:2010:BTF

[DU10] Iain S. Duff and Bora Uçar.

On the block triangular form of symmetric matrices. SIAM Review, 52(3): 455–470, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dube:1976:UBF

[Dub76] R. P. Dube. Univariate blending functions and alternatives. SIAM Review, 18 (4):801, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dubeau:1997:HCC

[Dub97a] F. Dubeau. Horizontal circular curves and cubics. SIAM Review, 39(1):110– 117, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dubeau:1997:CNH

[Dub97b] François Dubeau. Classroom note: Horizontal circular curves and cubics. SIAM Review, 39(1):110– 117, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29320;   http:
//www.siam.org/journals/
sirev/sirev391.htm.

Duffin:1969:ACS

[Duf69] R. J. Duffin. Algorithms for classical stability problems. SIAM Review, 11(2): 196–213, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duffin:1970:LGP

[Duf70]  R. J. Duffin.  Lineariz-

ing geometric programs. SIAM Review, 12(2):211– 227, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duff:1984:BRC

[Duf84]  Iain S. Duff.  Book re-

view: Computer Solution of Large Sparse Positive Definite Systems (Alan George and Joseph W. Liu). SIAM Review, 26(2):289–291, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duffy:1985:BRK

[Duf85] Joseph Duffy. Book review: Kinematics (Joseph Stiles Beggs). SIAM Review, 27(1):99–100, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duffy:1986:BRC

[Duf86] Joseph Duffy. Book review:

Computer Aided Analysis and Optimization of Mechanical System Dynamics (E. J. Haug, ed.). SIAM Review, 28 (1):99–100, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duffy:1991:BRI

[Duf91a] Joseph Duffy. Book review: Introduction to Theoretical Kinematics (J. M. Mc-Carthy). SIAM Review, 33 (2):302, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duffy:1991:BRK

[Duf91b] Joseph Duffy. Book review: Kinematic Geometry of Mechanisms (K. H. Hunt). SIAM Review, 33 (4):678–679, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dumey:1965:BRS

[Dum65] Arnold Dumey. Book review: A Sharp Autocorrelation Function (Henry D. Friedman). SIAM Review, 7 (2):283–284, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dumas:1996:BRI

[Dum96] H. Scott Dumas. Book review: Intermediate Classical Dynamics with Applications to Beam Physics (Leo Michelotti). SIAM Review, 38(4): 685–687, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duncan:1964:HPM

[Dun64] R. L. Duncan. Hit probabilities for multiple weapon systems. SIAM Review, 6(2): 111–114, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dunham:1968:ECC

[Dun68] C. B. Dunham. Existence and continuity of the Chebyshev operator. SIAM Review, 10(4):444–446, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duncan:1994:BRN

[Dun94] T. E. Duncan. Book review: Numerical Methods for Stochastic Control Problems in Continuous Time (Harold J. Kushner and Paul G. Dupuis). SIAM Review, 36 (4):668–669, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dupac:1993:BRA

[Dup93] Vaclav Dupac. Book review: Adaptive Algorithms and Stochastic Approximations (A. Benveniste, M. Métivier, and P. Priouret). SIAM Review, 35(2):313– 314, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1984:PTM

[Dur84a] R. T. Durrett. Percolation theory for mathematicians — Kesten, H. SIAM Review, 26 (3):446–448, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1984:BRP

[Dur84b] Richard T. Durrett. Book review: Percolation Theory for Mathematicians (Harry Kesten). SIAM Review, 26 (3):446–448, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1987:BRE

[Dur87]  Rick Durrett.  Book re-

view: École d’Été de Probabilités de Saint-Flour XIV (R. Carmona, H. Kesten and J. B. Walsh; P. L. Hennequin, ed.). SIAM Review, 29(3):478, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1990:BRP

[Dur90a] Richard T. Durrett. Book review: Percolation (Geoffrey Grimmett). SIAM Review, 32(4):693, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1990:PGG

[Dur90b] Richard T. Durrett. Percolation (Geoffrey Grimmett). SIAM Review, 32(4):693– ??, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1992:BRI

[Dur92a] Richard T. Durrett. Book review: Intersections of Random Walks (Gregory F. Lawler). SIAM Review, 34 (4):666, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1992:BRP

[Dur92b]  Rick Durrett.  Book re-

view: Probability Metrics and the Stability of Stochastic Models (Sveltozar T. Rachev). SIAM Review, 34 (1):147, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Duren:1993:BRI

[Dur93a] Peter Duren. Book review:

Inequalities: Fifty Years on from Hardy, Littlewood and Pólya (W. Norrie Everitt, ed.). SIAM Review, 35(2): 315–316, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1993:BRM

[Dur93b]  Rick Durrett.  Book re-

view: From Markov Chains to Nonequilibrium Particle Systems (Mu Fa Chen). SIAM Review, 35(3):514, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1997:BRA

[Dur97]  Rick Durrett.  Book re-

view: Applied Stochastic Processes: a Biostatistics and Population Oriented Approach, by Suddhendu Biswas. SIAM Review, 39 (1):144, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Durrett:1999:SSM

[Dur99] Rick Durrett. Stochastic spatial models. SIAM Review, 41 (4):677–718, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35470.

Dutton:1982:FTC

[Dut82] John A. Dutton. Fundamental theorems of climate theory—some proved, some conjectured. SIAM Review, 24(1):1–33, January 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Daley:1976:BRA

[DV76] D. J. Daley and R. J. Vaughan. Book review: An Average Distance (H. J. Oser). SIAM Review, 18(3): 497–501, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dongarra:1995:SLL

[DW95] Jack J. Dongarra and David W. Walker. Software libraries for linear algebra computations on high performance computers. SIAM Review, 37(2): 151–180, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/25458.htm.

Dwass:1964:BRP

[Dwa64] Meyer Dwass. Book review: Probabilities on Algebraic Structures (Ulf Grenander). SIAM Review, 6(3): 323–324, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dwass:1966:BRM

[Dwa66] Meyer Dwass. Book review: Mathematical Theory of Probability and Statistics (Richard von Viises). SIAM Review, 8(1):117, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dwass:1970:BRC

[Dwa70] Meyer Dwass. Book review: Combinatorial Methods in the Theory of Stochastic Processes (Lajos Takács). SIAM Review, 12(1):158– 159, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dwass:1972:BRM

[Dwa72] Meyer Dwass. Book review:

Markov Chains (David Freedman). SIAM Review, 14(1): 180–181, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dwass:1980:BRA

[Dwa80] Meyer Dwass. Book review: The Asymptotic Theory of Extreme Order Statistics (Janos Galambos). SIAM Review, 22(3):379, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dwass:1985:BRI

[Dwa85] Meyer Dwass. Book review: Introductory Applied Probability (G. P. Beaumont). SIAM Review, 27 (1):100, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dwyer:1963:BRA

[Dwy63] Paul S. Dwyer. Book review: The Analysis of Variance (Henry Scheffé). SIAM Review, 5(1):84–86, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dyakonov:1998:OSE

[DX98] Eugene G. D’yakonov and Jinchao Xu. Optimization in solving elliptic problems. SIAM Review, 40(1): 177–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Demanet:2011:DSC

[DY11] Laurent Demanet and Lexing Ying. Discrete symbol calculus. SIAM Review, 53 (1):71–104, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i1/p71_s1.

Dyson:1966:AGT

[Dys66] Freeman J. Dyson. Applications of group theory in particle physics. SIAM Review, 8(1):1–10, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/

2028169. Von Neumann Lecture given at Cornell University, Ithaca, on September 2, 1965, for the Society of Industrial and Applied Mathematics.

Ding:2004:LAH

[DZH^{+}04] Chris H. Q. Ding, Hongyuan Zha, Xiaofeng He, Parry Husbands, and Horst D. Simon. Link analysis: Hubs and authorities on the World Wide Web. SIAM Review, 46(2):256–268, June 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38921.

Earls:1981:BRQ

[Ear81] Clifford J. Earls. Book review: Quasiconformal Mappings and Riemann Surfaces (Samuil L. Krushkal). SIAM Review, 23(1):124, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Easton:1989:BRL

[Eas89] Robert W. Easton. Book review: Linear Dynamical Systems. A Revised Edition of “Dynamical Systems and Their Applications: Linear Theory” (John L. Casti). SIAM Review, 31(2):341– 342, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Easton:1991:BRO

[Eas91] Robert W. Easton. Book review: Ordinary Differential Equations: An Introduction to Nonlinear Analysis (Herbert Amann; Gerhard Metzen, trans.). SIAM Review, 33(4):669–670, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Easton:1993:BRI

[Eas93] Robert W. Easton. Book review: Introduction to Hamiltonian Dynamical Systems and the N-Body Problem (K. R. Meyer and G. R. Hall). SIAM Review, 35(4): 659, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eaton:1996:BRP

[Eat96]  Morris L. Eaton.  Book

review: Parametric Statistical Theory (Johann Pfanzagl). SIAM Review, 38(3): 527–529, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1988:BRE

[EB88] Ronald J. Evans and J. Boersma. Book review: The Entropy of a Poisson Distribution (C. Robert Appledorn). SIAM Review, 30(2): 314–317, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ebanks:1982:BRF

[Eba82] B. R. Ebanks. Book review: A Functional Equation (K. R. Parthasarathy). SIAM Review, 24(4):478, October 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eberlein:1975:MNC

[Ebe75] P. J. Eberlein. Minimum number of comparisons required to find median of Nelements. SIAM Review, 17 (2):377–378, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eberly:1996:BRP

[Ebe96] Wayne Eberly. Book review: Polynomial and Matrix Computations Volume 1: Fundamental Algorithms (Dario Bini and Victor Pan). SIAM Review, 38(1):161– 165, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

El-Bakry:1994:SII

[EBTZ94] A. S. El-Bakry, R. A. Tapia, and Y. Zhang. A study of indicators for identifying zero variables in interior-point methods. SIAM Review, 36 (1):45–72, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/22277.htm.

Ewing:1976:DST

[EC76] R. E. Ewing and J. R. Cannon. Determination of a source term in a linear parabolic partial-differential equation. SIAM Review, 18 (4):802, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eckhaus:1972:BLL

[Eck72] Wiktor Eckhaus. Boundary layers in linear elliptic singular perturbation problems. SIAM Review, 14(2): 225–270, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eckhaus:1977:FAS

[Eck77] Wiktor Eckhaus. Formal approximations and singular perturbations. SIAM Review, 19(4):593–633, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ecker:1980:GPM

[Eck80]  J. G. Ecker.  Geomet-

ric programming: Methods, computations and applications. SIAM Review, 22(3):

338–362, July 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ecker:1982:BRO

[Eck82] J. G. Ecker. Book review:

Optimality in Nonlinear Programming (Adi Ben-Israel, Aharon Ben–Tal and Sanjo Zlobec). SIAM Review, 24 (3):362, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eckhaus:1994:FCM

[Eck94] Wiktor Eckhaus. Fundamental concepts of matching. SIAM Review, 36(3):431– 439, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25791.htm.

Eden:1990:BRV

[Ede90a] Alp Eden. Book review: Vibration and Coupling of Continuous Systems: Asymptotic Methods (Jacqueline Sanchez Hubert and Enrique Sanchez Palencia). SIAM Review, 32 (4):693–694, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eden:1990:VCC

[Ede90b] Alp Eden. Vibration and coupling of continuous systems: Asymptotic methods (Jacqueline Sanchez Hubert and Enrique Sanchez Palencia). SIAM Review, 32(4): 693–??, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Edelman:1997:MPD

[Ede97] Alan Edelman. The mathematics of the Pentium division bug. SIAM Review, 39(1):54–67, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29395;   http:
                                        //www-math.mit.edu/~edelman/
                                        homepage/papers/pentiumbug.
                                        pdf;   http://www.siam.
                                        org/journals/sirev/sirev391.
                                        htm.

Editors:2011:Sd

[Edi11] The Editors. SIGEST. SIAM Review, 53(4):745, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2012:Sa

[Edi12a]  The Editors.  SIGEST.

SIAM Review, 54(1):139, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v54/i1/p139_s1.

Editors:2012:Sb

[Edi12b] The Editors. SIGEST. SIAM Review, 54(2):325, ???? 2012.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2012:Sc

[Edi12c] The Editors. SIGEST. SIAM Review, 54(3):543, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2014:Sb

[Edi14a] The Editors. SIGEST. SIAM Review, 56(2):313, ???? 2014. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2014:Sc

[Edi14b] The Editors. SIGEST. SIAM Review, 56(3):497, ???? 2014. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2014:Sd

[Edi14c] The Editors. SIGEST. SIAM Review, 56(4):643, ???? 2014. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2015:Sa

[Edi15a] The Editors. SIGEST. SIAM Review, 57(1):93, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2015:Sb

[Edi15b] The Editors. SIGEST. SIAM Review, 57(2):223, ???? 2015.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2015:Sc

[Edi15c] The Editors. SIGEST. SIAM Review, 57(3):419, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2015:Sd

[Edi15d] The Editors. SIGEST. SIAM Review, 57(4):583, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2016:Sc

[Edi16a] The Editors. SIGEST. SIAM Review, 58(3):515, ???? 2016. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2016:Sd

[Edi16b] The Editors. SIGEST. SIAM Review, 58(4):737, ???? 2016. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2017:Sa

[Edi17a] The Editors. SIGEST. SIAM Review, 59(1):151, ???? 2017. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2017:Sb

[Edi17b] The Editors. SIGEST. SIAM Review, 59(2):367, ???? 2017.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2017:Sc

[Edi17c] The Editors. SIGEST. SIAM Review, 59(3):617, ???? 2017. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2017:Sd

[Edi17d] The Editors. SIGEST. SIAM Review, 59(4):801, ???? 2017. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2018:Sa

[Edi18a] The Editors. SIGEST. SIAM Review, 60(1):137, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2018:Sb

[Edi18b] The Editors. SIGEST. SIAM Review, 60(3):671, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2018:Sc

[Edi18c] The Editors. SIGEST. SIAM Review, 60(4):939, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2019:Sa

[Edi19a] The Editors. SIGEST. SIAM Review, 61(1):119, ???? 2019.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2019:Sb

[Edi19b] The Editors. SIGEST. SIAM Review, 61(2):317, ???? 2019. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2019:Sc

[Edi19c] The Editors. SIGEST. SIAM Review, 61(3):507, ???? 2019. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2019:Sd

[Edi19d] The Editors. SIGEST. SIAM Review, 61(4):791, ???? 2019. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2020:Sa

[Edi20a] The Editors. SIGEST. SIAM Review, 62(1):197, ???? 2020. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2020:Sb

[Edi20b] The Editors. SIGEST. SIAM Review, 62(2):393, ???? 2020. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2020:Sc

[Edi20c] The Editors. SIGEST. SIAM Review, 62(3):659, ???? 2020.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2020:Sd

[Edi20d] The Editors. SIGEST. SIAM Review, 62(4):867, ???? 2020. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2021:Sa

[Edi21a] The Editors. SIGEST. SIAM Review, 63(1):121, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2021:Sb

[Edi21b] The Editors. SIGEST. SIAM Review, 63(2):373, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2021:Sc

[Edi21c] The Editors. SIGEST. SIAM Review, 63(3):557, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2021:Sd

[Edi21d] The Editors. SIGEST. SIAM Review, 63(4):781, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2022:Sa

[Edi22a] The Editors. SIGEST. SIAM Review, 64(1):151, ???? 2022.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975391.

Editors:2022:Sb

[Edi22b] The Editors. SIGEST. SIAM Review, 64(2):423, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21N975448.

Editors:2022:Sc

[Edi22c] The Editors. SIGEST. SIAM Review, 64(3):687, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21N975527.

Editors:2022:Sd

[Edi22d] The Editors. SIGEST. SIAM Review, 64(4):989, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975573.

Editors:2023:Sa

[Edi23a] The Editors. SIGEST. SIAM Review, 65(1):225, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N975636.

Editors:2023:Sb

[Edi23b] The Editors. SIGEST. SIAM Review, 65(2):493, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N975697.

Editors:2023:Sc

[Edi23c] The Editors. SIGEST. SIAM Review, 65(3):829, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N975740.

Editors:2023:S

[Edi23d]  The Editors.  SIGEST.

SIAM Review, 65(4):1107, November 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2024:Sa

[Edi24a]  The Editors.  SIGEST.

SIAM Review, 66(1):123, February 2024. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2024:Sb

[Edi24b] The Editors. SIGEST. SIAM Review, 66(2):317, May 2024. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Edstrom:2005:FSS

[Eds05] Per Edström. A fast and stable solution method for the radiative transfer problem. SIAM Review, 47(3): 447–468, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43871.

Edwards:1963:BRF

[Edw63] R. H. Edwards. Book review: Flight Mechanics — 1. Theory of Flight Paths (A. Miele). SIAM Review, 5(3): 297–298, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Edward:1990:BRV

[Edw90] J. Haug Edward. Book review: Variational Principles of Continuum Mechanics with Engineering Applications. Volume 2: Introduction to Optimal Design Theory (Vadim Komkov). SIAM Review, 32(4):696– 697, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Edwards:2000:AEM

[Edw00] D. A. Edwards. An alternative example of the method of multiple scales. SIAM Review, 42(2):317– 332, June 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34845.

Estrada:2020:MNI

[EERG20] Ernesto Estrada, Gissell Estrada-Rodriguez, and Heiko Gimperlein. Metaplex networks: Influence of the exo–endo structure of complex systems on diffusion. SIAM Review, 62(3):617– 645, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1963:MDF

[EF63] Carl Evans and John Fleck.

A monotonic distribution function. SIAM Review, 5 (2):156–157, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1965:BRM

[EF65] Carl Evans and John Fleck.

Book review: A Monotonic Distribution Function (Carl Evans and John Fleck). SIAM Review, 7(2):277– 278, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elveback:1976:SSM

[EFA76] Lila R. Elveback, John P.

Fox, and Eugene Ackerman. Stochastic simulation models for two immunization problems. SIAM Review, 18(1): 52–61, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Edwards:2013:AMT

[EFMM13] David A. Edwards, Joseph D. Fehribach, Richard O. Moore, and Colin J. McKinstrie. An application of matrix theory to the evolution of coupled modes. SIAM Review, 55(4): 764–787, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eden:1996:EAD

[EFN^{+}96] A. Eden, C. Foias, B. Nicolaenko, R. Temam, and Michael S. Jolly. Exponential attractors for dissipative evolution equations. SIAM Review, 38(1):171– ??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Efron:1979:CTS

[Efr79] Bradley Efron. Computers and the theory of statistics: Thinking the unthinkable. SIAM Review, 21(4): 460–480, October 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Efron:1988:CIM

[Efr88] Bradley Efron. Computerintensive methods in statistical regression. SIAM Review, 30(3):421–449, September 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Epureanu:1998:FBA

[EG98] Bogdan I. Epureanu and Henry S. Greenside. Fractal basins of attraction associated with a damped Newton’s method. SIAM Review, 40(1):102–109, March 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31003.

Elmroth:2004:RBA

[EGJK04] Erik Elmroth, Fred Gustavson, Isak Jonsson, and Bo Kågström. Recursive blocked algorithms and hybrid data structures for dense matrix library software. SIAM Review, 46 (1):3–45, March 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/42869.

El-Hodiri:1978:SPR

[EH78] Mohamed El-Hodiri. A simple proof of the Ramsey savings equation. SIAM Review, 20(1):169–170, ???? 1978.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elaydi:1998:C

[EH98] Saber N. Elaydi and William A. Harris, Jr. On the computation of A_{n}. SIAM Review, 40 (4):965–971, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31923.

Estrada:2010:NPR

[EH10] Ernesto Estrada and Desmond J. Higham. Network properties revealed through matrix functions. SIAM Review, 52 (4):696–714, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Estrada:2016:CAS

[EH16] Ernesto Estrada and Naomichi Hatano. Communicability angle and the spatial efficiency of networks. SIAM Review, 58(4):692– 715, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ehrlich:1976:SEU

[Ehr76] L. W. Ehrlich. Some experience using matrix splitting and conjugate gradient. SIAM Review, 18(4): 801, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

El:2017:DDD

[EHS17] G. A. El, M. A. Hoefer, and M. Shearer. Dispersive and diffusive–dispersive shock waves for nonconvex conservation laws. SIAM Review, 59(1):3–61, ???? 2017. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eicker:1967:BRM

[Eic67] F. Eicker. Book review:

Mathematical Foundations of the Calculus of Probability (Jacques Neveu). SIAM Review, 9(1):136–139, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eichinger:1984:I

[Eic84] B. E. Eichinger. An integral on SO(3). SIAM Review, 26(1):117, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eisemann:1962:OLC

[Eis62] Kurt Eisemann. The optimum location of a center. SIAM Review, 4(4): 394–395, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eisele:1986:BRE

[Eis86a] Theodore Eisele. Book review: Entropy, Large Deviations, and Statistical Mechanics (Richard S. Ellis). SIAM Review, 28(4):586– 587, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eisenstat:1986:BRD

[Eis86b] Stanley C. Eisenstat. Book review: Direct Methods for Sparse Matrices (Ole østerby and Zahari Zlatev). SIAM Review, 28(3):414, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Efthimiou:2007:DW

[EJ07] C. J. Efthimiou and M. D. Johnson. Domino waves. SIAM Review, 49(1):111– 120, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eisenstein:1959:DV

[EK59] Maurice Eisenstein and M. S. Klamkin. N-dimensional volume. SIAM Review, 1(1): 69–70, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elkies:1986:II

[EK86] N. Elkies and M. S. Klamkin. An isoperimetric inequality. SIAM Review, 28(1):

88–89, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Estrada:1987:CTS

[EK87] R. Estrada and R. P. Kanwal. The Carleman type singular integral equations. SIAM Review, 29(2):263– 290, June 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Edelstein-Keshet:1991:BRM

[EK91a] Leah Edelstein-Keshet. Book review: Mathematical Methods and Models in the Biological Sciences Vol. 1 (Martin Eisen). SIAM Review, 33 (1):139–141, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Edelstein-Keshet:1991:MMM

[EK91b] Leah Edelstein-Keshet. Mathematical methods and models in the biological sciences (Martin Eisen). SIAM Review, 33(1):139–??, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Engelberg:2013:IRC

[EK13] Shlomo Engelberg and Osnat Keren. An introduction to robust codes over finite fields. SIAM Review, 55(4): 751–763, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ekstrom:1960:IIC

[Eks60] Joel L. Ekstrom. On infinite integrals containing products of Bessel functions. SIAM Review, 2(1): 23–26, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Egorov:1997:STE

[EKS97] Yuri Egorov, Vladimir Kondratiev, and Martin Schechter. On spectral theory of elliptic operators. SIAM Review, 39 (4):802–??, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Estrada:1996:AAD

[EKW96] R. Estrada, R. P. Kanwal, and Jet Wimp. Asymptotic analysis: a distributional approach. SIAM Review, 38(4): 706–??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erbe:1996:OTF

[EKZA96] L. H. Erbe, Q. K. Kong, B. G. Zhang, and Ovide Arino. Oscillation theory for functional differential equations. SIAM Review, 38(4): 680–??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elishakoff:1985:SPR

[EL85] I. Elishakoff and D. Livshits.

Sums of powers of roots of a transcendental equation. SIAM Review, 27(3): 446, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elishakoff:1986:SPR

[EL86] I. Elishakoff and D. Livshits.

Sums of powers of roots of a transcendental equation. SIAM Review, 28(3): 400–402, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elcrat:1976:CIO

[Elc76] A. Elcrat. Coercivity inequality for 2nd order elliptic systems with applications to nonlinear systems. SIAM Review, 18(4): 801–802, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elhodiri:1978:SPR

[Elh78] M. Elhodiri. Simple proof of Ramsey savings equation. SIAM Review, 20(1): 169–170, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elkies:1986:BRI

[Elk86] Noam Elkies. Book review:

An Isoperimetric Inequality (M. S. Klamkin). SIAM Review, 28(1):88–89, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elliott:1961:NSL

[Ell61] David L. Elliott. A note on systems of linear equations. SIAM Review, 3(1): 66–69, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elliott:1980:BRS

[Ell80] R. J. Elliott. Book review: Stochastic Optimal Control: The Discrete Time Case (Dimitri P. Bertsekas and Steven F. Shreve). SIAM Review, 22(2):237– 238, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elliott:1981:BRC

[Ell81] Robert J. Elliott. Book review: Controlled Markov Processes (E. B. Dynkin and A. A. Yushkevich). SIAM Review, 23(2):269– 270, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ellerman:1984:ATM

[Ell84] David P. Ellerman. Arbitrage theory — a mathematical introduction. SIAM Review, 26 (2):241–261, April 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ellickson:1985:BRM

[Ell85a]  Bryan Ellickson.  Book

review: Market Demand — an Analysis of Large Economies with Non-Convex Preferences (Walter Trockel). SIAM Review, 27(4):592– 593, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elliott:1985:BRO

[Ell85b] Robert J. Elliott. Book review: Optimization Over Time-Volume II: Dynamic Programming and Stochastic Control (Peter Whittle). SIAM Review, 27(1): 100–101, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elliott:1987:BRV

[Ell87] C. M. Elliott. Book review: Variational and Quasivariational Inequalities Applications to Free-Boundary Problems (Claudio Baiocchi And António Capelo). SIAM Review, 29(2):314– 315, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elliott:1988:BRF

[Ell88] C. M. Elliott. Book review: Free and Moving Boundary Problems (John Crank). SIAM Review, 30(4):668– 669, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ellingwood:1990:BRR

[Ell90] Bruce Ellingwood. Book review: Reliability of Randomly Excited Hysteretic Structures (B. F. Spencer, Jr.). SIAM Review, 32(1): 172–173, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ellickson:1992:BRE

[Ell92a] Bryan Ellickson. Book review: The Econometric Analysis of Transition Data (Tony Lancaster). SIAM Review, 34 (1):151–152, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ellickson:1992:EAT

[Ell92b] Bryan Ellickson. The econometric analysis of transition data (Tony Lancaster). SIAM Review, 34(1):151– ??, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ellis:1994:BRC

[Ell94] Steven P. Ellis. Book review:

A Course on Point Processes (R.-D. Reiss). SIAM Review, 36(1):124–125, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ellis:2002:FLT

[Ell02] Steven P. Ellis. Fitting a line to three or four points on a plane. SIAM Review, 44 (4):616–628, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37677.

Englar:1966:BRP

[ELRS66] Thomas S. Englar, Jr., E. S.

Langford, Judy Richman, and Sidney Spital. Book review: On a Permutation Sequence (Paul Brook). SIAM Review, 8(1):108, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elsworth:1989:BRB

[Els89] Derek Elsworth. Book review: Boundary Element Analysis in Computational Fracture Mechanics (T. A. Cruse). SIAM Review, 31 (4):692, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elsworth:1991:BRS

[Els91] Derek Elsworth. Book review: Slope Analysis Using Boundary Elements (Y. S. Jiang). SIAM Review, 33(1): 141–142, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eluhu:1988:BRT

[Elu88] C. Calmelet Eluhu. Book review: Theory of Hydromagnetic Stability (Bhimsen K. Shivamoggi). SIAM Review, 30(2):326–327, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elveback:1975:SSM

[Elv75] L. R. Elveback. Stochastic simulation models for two immunization problems. SIAM Review, 17(2):378, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eiseman:1976:SFC

[EMB76] P. R. Eiseman, H. Mcdonald, and W. R. Briley. Subsonic flows in a curved Duct. SIAM Review, 18(4):801, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Embree:2003:THR

[Emb03] Mark Embree. The tortoise and the hare restart GMRES. SIAM Review, 45(2):259–266, June 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39961.

Engeln-Mullges:1998:BRN

[EMUP98] Gisela Engeln-Mullges, Frank Uhlig, and John D. Pryce. Book reviews: Numerical Algorithms with C and Numerical Algorithms with Fortran. SIAM Review, 40(1): 171–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Engel:1978:BRE

[Eng78] Joseph H. Engel. Book review: Elementary Systems Mathematics (Robert E. Machol). SIAM Review, 20(4): 873, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Engl:1994:BRR

[Eng94] Heinz W. Engl. Book review: Regularization Methods for Ill-Posed Problems (V. A. Morozow). SIAM Review, 36 (3):505–506, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Engelberg:2003:SST

[Eng03] Shlomo Engelberg. Spread spectrum from two perspectives. SIAM Review, 45(3):

574–587, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40692.

Engelberg:2017:ENT

[Eng17]  Shlomo Engelberg.  Ele-

mentary number theory and Rader’s FFT. SIAM Review, 59(3):671–678, ???? 2017. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Enright:1978:ESL

[Enr78] W. H. Enright. Efficient solution of linear constant coefficient systems of ODEs. SIAM Review, 20(3):623, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Enright:1984:BRS

[Enr84] W. Enright. Book review: Stable Recursions with Applications to the Numerical Solution of Stiff Systems (J. R. Cash). SIAM Review, 26(3):450, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Enright:1987:BRC

[Enr87] W. H. Enright. Book review: Computer Science. A Mathematical Introduction (Art Lew). SIAM Review, 29(4): 650–651, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Enright:1989:BRN

[Enr89] W. H. Enright. Book review: The Numerical Analysis of Ordinary Differential Equations: Runge–Kutta and General Linear Methods (J. C. Butcher). SIAM Review, 31(4):693, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Epperson:1986:AIO

[Epp86] James F. Epperson. Analysis and implementation of an oxygen absorption model. SIAM Review, 28(3):389– 393, September 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Epstein:1959:BRI

[Eps59a] Benjamin Epstein. Book review: Introduction to Statistical Analysis (W. J. Dixon and F. J. Massey, Jr.). SIAM Review, 1(1): 75–77, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Epstein:1959:BRN

[Eps59b] Benjamin Epstein. Book review: Nonparametric Methods in Statistics (D. A. S. Fraser). SIAM Review, 1(2): 174–175, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erlebach:1979:ABS

[ER79] L. Erlebach and O. Ruehr. Asymptotic behavior of a sequence. SIAM Review, 21 (1):140, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erdos:1972:SUG

[Erd72] P. Erdős. A solved and unsolved graph coloring problem. SIAM Review, 14(2): 364, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erdos:1977:BRA

[Erd77] P. Erdős. Book review: An Arithmetic Conjecture (D. J. Newman). SIAM Review, 19 (1):154–155, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erdos:1983:BRB

[Erd83] P. Erdős. Book review: “Bad” Triangles in Hilbert Space (M. Burnashev). SIAM Review, 25(4):574–575, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erdos:1992:OMA

[Erd92a] P. Erdős. Optimum multiple access coding. SIAM Review, 34(2):309–310, ???? 1992.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erdos:1992:BRO

[Erd92b]  Paul Erdős.  Book re-

view: Optimum Multiple Access Coding (D.B. Jevtić). SIAM Review, 34(2):309– 310, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erdos:1992:EP

[Erd92c] Paul Erdős. An extremal problem. SIAM Review, 34 (2):308, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erdos:1994:EP

[Erd94] P. Erdős. An extremal problem. SIAM Review, 36(3): 493–495, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eberle:1995:NSE

[ERHO95] Albrecht Eberle, Arthur Rizzi, Ernst Heinrich Hirschel, and Paul D. Orkwis. Numerical solutions of the Euler equations for steady flow problems. SIAM Review, 37 (2):271–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ericksen:1966:NEM

[Eri66] Wilhelm S. Ericksen. Note on the equations of motion of a rigid body. SIAM Review, 8 (2):206–210, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ermentrout:1989:BRA

[Erm89] G. B. Ermentrout. Book review: Autowave Processes in Kinetic Systems: Spatial and Temporal Self-Organization in Physics, Chemistry, Biology, and Medicine (V. A. Vasiliev, Yu. M. Romanovskii, D. S. Chernavskii, and V. G. Yakhno). SIAM Review, 31(1):133– 136, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elman:2014:ICL

[ERS14] Howard C. Elman, Alison Ramage, and David J. Silvester. IFISS: a computational laboratory for investigating incompressible flow problems. SIAM Review, 56 (2):261–273, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eiseman:1980:CLF

[ES80] P. R. Eiseman and A. P. Stone. Conservation laws of fluid dynamics — a survey. SIAM Review, 22(1):

12–27, January 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Epstein:2008:BTA

[ES08] Charles L. Epstein and John Schotland. The bad truth about Laplace’s transform. SIAM Review, 50(3):504– 520, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Escary:1970:BRP

[Esc70] J. D. Escary. Book review: A Probability Inequality (R. Chandrasekaran). SIAM Review, 12(2):300– 302, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eslami:1993:OSB

[Esl93] Mansour Eslami. Optimization of a stability bound. SIAM Review, 35(4):625– 630, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eslami:1996:ESB

[Esl96] Mansour Eslami. On the extension of a stability bound: The multiple largeparameter variations case. SIAM Review, 38(1):120– 127, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/p120.htm.

Espelid:1995:FPS

[Esp95] T. O. Espelid. On floatingpoint summation. SIAM Review, 37(4):603–607, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/27303.htm.

Efroymson:1979:MEP

[ESS79] G. Efroymson, A. Steger, and S. Steinberg. A matrix eigenvalue problem. SIAM Review, 21(1):139, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Embree:1999:GFM

[ET99] Mark Embree and Lloyd N. Trefethen. Green’s functions for multiply connected domains via conformal mapping. SIAM Review, 41 (4):745–761, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34927.

Elbakry:1994:SII

[ETZ94] A. S. Elbakry, R. A. Tapia, and Y. Zhang. A study of indicators for identifying zero variables in interiorpoint methods. SIAM Review, 36(1):45–72, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eubank:1988:OGS

[Eub88] R. L. Eubank. Optimal grouping, spacing, stratification, and piecewise constant approximation. SIAM Review, 30(3):404–420, September 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Eijkhout:1991:RSC

[EV91] Victor Eijkhout and Panayot Vassilevski. The role of the strengthened Cauchy– Buniakowskiĭ–Schwarz inequality in multilevel methods. SIAM Review, 33(3): 405–419, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Elmachtoub:2010:RPE

[EV10] Adam N. Elmachtoub and Charles F. Van Loan. From random polygon to ellipse: An eigenanalysis. SIAM Review, 52(1):151–170, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1981:BRI

[Eva81] Lawrence C. Evans. Book review: An Introduction to Variational Inequalities and Their Applications (D. Kinderlehrer and G. Stampacchia). SIAM Review, 23 (4):539–543, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1984:VPF

[Eva84a] L. C. Evans. Variational principles and free-boundary problems — Friedman, A. SIAM Review, 26(2):307– 308, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1984:BRV

[Eva84b] Lawrence C. Evans. Book review: Variational Principles and Free Boundary Problems. (Avner Friedman). SIAM Review, 26(2):307– 308, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1985:BRH

[Eva85] Lawrence C. Evans. Book review: Hamilton–Jacobi Equations in Hilbert Spaces (V. Barbu and G. DaPrato). SIAM Review, 27(1):101– 102, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1988:EPD

[Eva88]  R. J. Evans.  The en-

tropy of a Poisson distribution. SIAM Review, 30(2): 314–315, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1990:TPB

[Eva90] Richard B. Evans. A twopoint boundary problem for Airy functions. SIAM Review, 32(3):478–479, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1993:IRS

[Eva93] Richard B. Evans. An integral relation for successive eigenvalues. SIAM Review, 35(3):489, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Evans:1994:IRS

[Eva94] R. B. Evans. An integral relation for successive eigenvalues. SIAM Review, 36 (3):497, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Estabrook:1975:GAS

[EW75] F. B. Estabrook and H. D.

Wahlquist. The geometric approach to sets of ordinary differential equations and Hamiltonian dynamics. SIAM Review, 17(2): 201–220, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ehrhardt:2019:NMP

[EW19] Beate Ehrhardt and Patrick J. Wolfe. Network modularity in the presence of covariates. SIAM Review, 61(2): 261–276, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ewing:1967:LSA

[Ewi67]  George M. Ewing.  On

a least square approximation. SIAM Review, 9(2): 251–252, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ewing:1977:SCG

[Ewi77] George M. Ewing. Sufficient conditions for global minima of suitably convex functionals from variational and control theory. SIAM Review, 19 (2):202–220, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fabian:1980:BRS

[Fab80] Väclav Fabian. Book review: Stochastic Approximation Methods for Constrained and Unconstrained Systems (Harold L. Kushner and Dean S. Clark). SIAM Review, 22 (3):382–384, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fabian:1987:BRI

[Fab87] Václav Fabian. Book review: Instrumental Variables. (Roger J. Bowden and Darrell A. Turkington). SIAM Review, 29(1):140– 142, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fadell:1962:BRT

[Fad62] Edward Fadell. Book review:

Topology (J. G. Hocking and G. S. Young). SIAM Review, 4(1):54, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fairweather:1971:EEC

[Fai71] Graeme Fairweather. On the eigenvalues and eigenvectors of a class of Hessenberg matrices. SIAM Review, 13(2): 220–221, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Falk:1968:BRN

[Fal68] James E. Falk. Book review: Nonlinear Programming (Hans Paul Künzi and Wilhelm Krelle in collaboration with Werner Oettli). SIAM Review, 10(1): 96–97, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Falb:1970:BRC

[Fal70a] Peter L. Falb. Book review: Contrôle Optimal de Systèmes Gouvernés des Équations aux Dérivées Partielles (J. L. Lions). SIAM Review, 12(2):308–309, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Falk:1970:BRO

[Fal70b] James E. Falk. Book review: Optimization by Vector Space Methods (David G. Luenberger). SIAM Review, 12 (2):315–316, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Falk:1973:BRO

[Fal73] James E. Falk. Book review: Optimization Theory for Large Systems (Leon S. Lasdon). SIAM Review, 15 (2):405–406, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Falk:1993:BRM

[Fal93a]  Richard S. Falk.  Book

review: Mixed and Hybrid Finite Element Methods (Franco Brezzi and Michel Fortin). SIAM Review, 35 (3):514–517, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Falk:1993:MFE

[Fal93b] Richard S. Falk. Mixed and finite element methods (Franco Brezzi and Michel Fortin). SIAM Review, 35(3): 515–??, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fan:1959:BRV

[Fan59] Ky Fan. Book review: Vector Spaces and Matrices (Robert M. Thrall and Leonard Tornheim). SIAM Review, 1(1): 72–73, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fan:1982:BRI

[Fan82] Ky Fan. Book review: Inequalities: Theory of Majorization and Its Applications (Albert W. Marshall and Ingram Olkin). SIAM Review, 24(1):86– 87, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farnell:1960:CRB

[Far60] A. B. Farnell. Characteristic root bounds of Gerschgorin type. SIAM Review, 2(1): 36–38, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farkas:1965:BRD

[Far65] I. Farkas. Book review: On a Definite Integral (Robert Klein). SIAM Review, 7(1): 138–139, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farkas:1969:BRN

[Far69] I. Farkas. Book review: Numerical Evaluation of Some Integrals (Murlan S. Corrington). SIAM Review, 11(4): 625–628, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farges:1984:ND

[Far84] Jacques Farges. A nonnegative determinant. SIAM Review, 26(3):431–432, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farin:1985:BRG

[Far85] Gerald Farin. Book review: Geometry of a Spatial Form. Analysis, Synthesis, Concept Formulation and Space Vision for CAD (Peter C. Gasson). SIAM Review, 27(4): 593–594, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farin:1993:BRG

[Far93a] G. Farin. Book review: Geometric Modelling for Numerically Controlled Machining (K. Marciniak). SIAM Review, 35(3):517–518, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farin:1993:GMN

[Far93b] G. Farin. Geometric modelling for numerically controlled machining (K Marciniak). SIAM Review, 35(3):517– ??, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farebrother:1994:BRP

[Far94] Richard William Farebrother.

Book review: Pitman’s Measure of Closeness — a Comparison of Statistical Estimators (Jerome P. Keating, Robert L. Mason, and Pranab K. Sen). SIAM Review, 36(3):507, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farwig:1996:BRN

[Far96] R. Farwig. Book review:

Navier–Stokes Equations and Nonlinear Functional Analysis (Roger Temam). SIAM Review, 38(4):687–689, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Foschini:1969:BRR

[FARE69] G. L. Foschini, Carlene Arthur, Cecil Rousseau, and S. H. Eisman. Book review: Rank and Eigenvalues of a Matrix (Sylvan Katz and M. S. Klamkin). SIAM Review, 11(4):628–630, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fateman:1976:MAP

[Fat76] R. J. Fateman. Macsyma arbitrary precision floating point arithmetic package — philosophy and an overview of its implementation. SIAM Review, 18(4):802, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fateman:1993:BRA

[Fat93] Richard J. Fateman. Book review: Automatic Differentiation of Algorithms: Theory, Implementation, and Application (Andreas Griewank and George F. Corliss, eds.). SIAM Review, 35(4):659– 660, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fazio:1996:EUB

[Faz96]  Riccardo Fazio.  Exis-

tence and uniqueness for a boundary value problem. SIAM Review, 38(1):146– 147, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fazio:1998:SAN

[Faz98] Riccardo Fazio. A similarity approach to the numerical solution of free boundary problems. SIAM Review, 40 (3):616–635, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28505.

Foley:1978:BRC

[FB78] R. D. Foley and J. C.

Butcher. Book review: A Coin Tossing Problem (Danny Newman). SIAM Review, 20(3):597–598, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Funderlic:1976:WMR

[FC76] R. E. Funderlic and R. E.

Cline. What makes rank of a matrix change. SIAM Review, 18(4):804, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fernandez-Cara:1999:ANC

[FC99]  Enrique  Fernández-Cara.

On the approximate and null controllability of the Navier–Stokes equations. SIAM Review, 41(2):269– 277, June 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33458.

Feeman:2000:EAW

[Fee00] Timothy G. Feeman. Equal area world maps: a case study. SIAM Review, 42 (1):109–114, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35899.

Feinberg:1996:BRM

[Fei96] A. Feinberg. Book review: Markov Decision Processes: Discrete Stochastic Dynamic Programming (Martin L. Puterman). SIAM Review, 38(4):689, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fejer:1988:BRP

[Fej88]  J. A. Fejer.  Book re-

view: The Propagation of Radio Waves: The Theory of Radio Waves of Low Power in the Ionosphere and Magnetosphere (K. G. Budden). SIAM Review, 30 (2):327, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Felsen:1970:RDS

[Fel70] Leopold B. Felsen. Rays, dispersion surfaces and their uses for radiation and diffraction problems. SIAM Review, 12(3):424–448, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Felsen:1972:BRA

[Fel72] L. B. Felsen. Book review: Applied Diffraction Theory (F. H. Northover). SIAM Review, 14(2):354– 356, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Felippa:1976:CDT

[Fel76] C. A. Felippa. Constraint definition techniques in solution of discrete boundary-value problems. SIAM Review, 18 (4):802–803, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Felippa:1978:LIP

[Fel78] C. A. Felippa. Local iterative procedures for improving penalty function solutions of finite-element systems. SIAM Review, 20(3):624, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feliachi:1990:BRC

[Fel90] Ali Feliachi. Book review:

Control of Partially-Known Dynamical Systems (A. A. Bahnasawi and M. S. Mahmoud). SIAM Review, 32 (4):694–695, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fellman:1992:MVS

[Fel92] J. O. Fellman. Maximum value of a shear stress. SIAM Review, 34(2):307, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fenyo:1968:RCC

[Fen68] I. Fenyö. Remark on certain convolution transforms. SIAM Review, 10(4):449– 450, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fenstad:1994:BRH

[Fen94] Jens Erik Fenstad. Book review: Hilbert’s Tenth Problem (Y. V. Matiyasevich). SIAM Review, 36(4):669– 671, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ferry:1979:BRN

[Fer79] John M. Ferry. Book review: Numerical Methods for Partial Differential Equations (William F. Ames). SIAM Review, 21(3):410, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fernandez:1997:LLM

[Fer97a] L. A. Fernandez. On the limits of the Lagrange multiplier rule. SIAM Review, 39(2):

292–297, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fernandez:1997:CNL

[Fer97b] Luis A. Fernández. Classroom note: On the limits of the Lagrange multiplier rule. SIAM Review, 39(2):292–297, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29352.

Fettis:1965:BRI

[Fet65] H. E. Fettis. Book review: On Integrals Related to Bessel Functions (Robert E. Shafer). SIAM Review, 7(3): 412–415, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1966:BRE

[Fet66] Henry E. Fettis. Book review:

An Expansion for the Absorptivity of a Gas (J. Ernest Wilkins, Jr.). SIAM Review, 8(2):237–238, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1968:BRD

[Fet68a]  Henry E. Fettis.  Book

review: A Definite Integral (M. Lawrence Glasser). SIAM Review, 10(1):116– 117, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1968:BRA

[Fet68b] Henry E. Fettis. Book review: An Approximation of a Definite Integral (David Rothman). SIAM Review, 10(1): 110–112, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1968:GBF

[Fet68c] Henry E. Fettis. Generalization of a Bessel function integral. SIAM Review, 10(2): 214–215, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1972:BRE

[Fet72a] H. E. Fettis. Book review: An Elliptic Integral (M. L. Glasser). SIAM Review, 14 (3):497–498, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1972:RR

[Fet72b] H. E. Fettis. A recurrence relation. SIAM Review, 14 (1):165, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1980:BRT

[Fet80] H. E. Fettis. Book review: Two Infinite Sums (P. J. De Doelder). SIAM Review, 22 (3):365–368, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1981:PVI

[Fet81] H. E. Fettis. Principal value of an integral. SIAM Review, 23(1):101, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1984:PI

[Fet84a] Henry E. Fettis. A polynomial identity. SIAM Review, 26(2):271–272, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1984:SBF

[Fet84b] Henry E. Fettis. A summation of Bessel functions. SIAM Review, 26(3):430– 431, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feuerverger:1976:BRS

[Feu76] Andrey Feuerverger. Book review: The Statistical Analysis of Time Series (T. W. Anderson). SIAM Review, 18 (2):313–314, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fyfe:1961:NSH

[FEY61] I. M. Fyfe, R. C. Eng, and D. M. Young. On the numerical solution of the hydrodynamic equations. SIAM Review, 3(4):298– 308, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feldstein:1978:OCV

[FF78] A. Feldstein and R. M. Firestone. Order of convergence for vector recurrences. SIAM Review, 20(3): 623–624, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farkas:1995:PM

[FF95] Miklos Farkas and H. I. Freedman. Periodic motions. SIAM Review, 37(4): 626–??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feldstein:1976:CFI

[FG76] A. Feldstein and R. Goodman. Conjecture on a finite and an infinite product. SIAM Review, 18(3): 490–491, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fairweather:2004:AAB

[FG04] Graeme Fairweather and Ian Gladwell. Algorithms for almost block diagonal linear systems. SIAM Review, 46(1):49–58, March 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40506.

Fennell:2019:MDP

[FG19] Peter G. Fennell and James P.

Gleeson. Multistate dynamical processes on networks: Analysis through degreebased approximation frameworks. SIAM Review, 61(1): 92–118, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fonseca:1997:DTA

[FGA97] Irene Fonseca, Wilfrid Gangbo, and Antonio Ambrosetti. Degree theory in analysis and applications. SIAM Review, 39(3):528–529, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feldstein:1981:BRC

[FGB81] A. Feldstein, R. Goodman, and J. Bustoz. Book review: Conjecture on a Finite and an Infinite Product (A. Feldstein and R. Goodman). SIAM Review, 23(1): 104–105, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ford:1992:ODE

[FGD^{+}92] W. F. Ford, C. C. Grosjean, H. Demeyer, W. B. Jordan, A. C. King, A. D. Osborne, and J. E. Wilkins. A 3rdorder differential-equation. SIAM Review, 34(1):121– 122, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feng:2013:RDN

[FGN13] Xiaobing Feng, Roland Glowinski, and Michael Neilan. Recent developments in numerical methods for fully nonlinear second order partial differential equations. SIAM Review, 55(2):205– 267, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fawzi:2022:LSC

[FGP^{+}22] Hamza Fawzi, Joao Gouveia, Pablo A. Parrilo, James Saunderson, and Rekha R. Thomas. Lifting for simplicity: Concise descriptions of convex sets. SIAM Review, 64(4):866–918, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1324417.

Froyland:2016:OME

[FGTW16] Gary Froyland, Cecilia González-Tokman, and Thomas M.

Watson. Optimal mixing enhancement by local perturbation. SIAM Review, 58(3): 494–513, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Forsgren:2002:IMN

[FGW02] Anders Forsgren, Philip E.

Gill, and Margaret H. Wright. Interior methods for nonlinear optimization. SIAM Review, 44 (4):525–597, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41494.

Fox:1970:DIS

[FH70] L. Fox and Linda Hayes.

On the definite integration of singular integrands. SIAM Review, 12(3):449– 457, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frost:1976:EPP

[FH76] P. A. Frost and E. Y.

Harper. An extended Padé procedure for constructing global approximations from asymptotic expansions: an explication with examples. SIAM Review, 18(1):62– 91, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Falk:1995:LSN

[FHRO95] Michael Falk, Jurg Husler, Rolf-Dieter Reiss, and George L. O’Brien. Laws of small numbers: Extremes and rare events. SIAM Review, 37(4): 634–??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fiala:1973:BRL

[Fia73]  F. Fiala.  Book review:

Linear Optimization (W. Allen Spivey and Robert M. Thrall). SIAM Review, 15(4): 807–809, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ficken:1960:LNE

[Fic60] F. A. Ficken. Linear and nonlinear equations; local and global properties. SIAM Review, 2(1):11–14, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fierro:1993:BRT

[Fie93] Ricardo D. Fierro. Book review: The Total Least Squares Problem: Computational Aspects and Analysis (S. Van Huffel and J. Vandewalle). SIAM Review, 35 (4):660–662, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fife:1979:BRS

[Fif79] Paul C. Fife. Book review: Stability of Fluid Motions Vols. I and II (Daniel D. Joseph). SIAM Review, 21 (3):415–418, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fife:1986:BRG

[Fif86] Paul C. Fife. Book review:

Global Solutions of Reaction-Diffusion Systems (Franz Rothe). SIAM Review, 28 (1):100–102, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Figotin:1993:BRS

[Fig93a]  Alex Figotin.  Book re-

view: Spectral Theory of Random Schrödinger Operators (René Carmona and Jean Lacroix). SIAM Review, 35(2):316–318, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Figotin:1993:STR

[Fig93b] Alex Figotin. Spectral theory of random schrödinger operators (Rene Carmona and Jean Lacroix). SIAM Review, 35(2):316–??, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fink:1971:CCI

[Fin71] James P. Fink. Complete continuity of the inverse of the Laplacian and Green’s function. SIAM Review, 13 (1):47–54, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fink:1972:APF

[Fin72] A. M. Fink. Almost periodic functions invented for specific purposes. SIAM Review, 14 (4):572–581, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fink:1976:BRD

[Fin76a] A. M. Fink. Book review: n-Dimensional Simple Harmonic Motion (I. J. Schoenberg). SIAM Review, 18(4): 772–773, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fink:1976:BRN

[Fin76b]  A. M. Fink.  Book re-

view: Non-Linear Differential Equations of Higher Order (R. Reissig, G. Sansone and R. Conti). SIAM Review, 18(1):143–144, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Finch:1984:CTS

[Fin84a] D. V. Finch. Computed tomography — Shepp, L.

A. SIAM Review, 26(3): 450–452, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Finch:1984:BRC

[Fin84b] David V. Finch. Book review: Computed Tomography. (Lawrence A. Shepp, ed). SIAM Review, 26(3): 450–452, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fink:1993:BRO

[Fin93] A. M. Fink. Book review:

Optimality and Equilibria in Stochastic Games (F. Thuijsman). SIAM Review, 35 (3):518–519, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Findley:1995:BRT

[Fin95] David F. Findley. Book review: Time Series: Forecasting, Simulation, Applications (Gareth Janacek and Louise Swift). SIAM Review, 37(2): 253–257, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Firey:1992:GEC

[Fir92a]  William J. Firey.  Geo-

metric etudes in combinatorial mathematics (Vladimir G Boltyanski and Alexander Soifer). SIAM Review, 34(3):

508–??, September 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Firey:1992:BRG

[Fir92b] Wm. J. Firey. Book review: Geometric Etudes in Combinatorial Mathematics (Vladimir G. Boltyanski and Alexander Soifer). SIAM Review, 34(3):508, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Firey:1997:BRG

[Fir97] Wm. J. Firey. Book review: Geometric Tomography Encyclopedia of Mathematics and its Applications. Vol. 58, by Richard J. Gardner. SIAM Review, 39(1): 144–145, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fisher:1971:BRS

[Fis71a] Michael E. Fisher. Book review: Sum of Inverse Powers of Cosines (L. A. Gardner, Jr.). SIAM Review, 13 (1):116–119, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fisher:1971:BRI

[Fis71b] Stephen Fisher. Book review: Introduction to Complex Variables (Norman Levinson and Raymond Redheffer). SIAM Review, 13(3): 409–411, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fishburn:1974:SCF

[Fis74] Peter C. Fishburn. Social choice functions. SIAM Review, 16(1):63–90, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fisher:1976:BRI

[Fis76] Walter D. Fisher. Book review: An Introduction to Applied Econometric Analysis (R. F. Wynn and K. Holden). SIAM Review, 18 (2):307, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fishburn:1988:COP

[Fis88a] Peter C. Fishburn. Combinatorial optimization problems for systems of subsets. SIAM Review, 30 (4):578–588, December 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fisher:1988:PPT

[Fis88b] Dawn Fisher. Pressure potential in two dimensions. SIAM Review, 30(2):305– 307, June 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fisher:1988:DS

[Fis88c] M. E. Fisher. Double summations. SIAM Review, 30 (2):320–322, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fisher:1988:BRD

[Fis88d] Michael. E. Fisher. Book review: Double Summations (Danny Summers). SIAM Review, 30(2):320– 322, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fishburn:1993:BRC

[Fis93] Peter C. Fishburn. Book review: Combinatorics and Partially Ordered Sets: Dimension Theory (William T. Trotter). SIAM Review, 35 (3):519–520, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fix:1974:BRA

[Fix74]  George Fix.  Book re-

view: Approximation of Elliptic Boundary-Value Problems (Jean-Pierre Aubin). SIAM Review, 16(3):400– 401, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fix:1975:FES

[Fix75] G. J. Fix. Finite elements and scientific computing. SIAM Review, 17(2):378, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fix:1976:HFE

[Fix76]  George M. Fix.  Hy-

brid finite element methods. SIAM Review, 18(3): 460–484, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fix:1978:BRR

[Fix78] George J. Fix. Book review: A Rational Finite Element Basis (Eugene L. Wachpress). SIAM Review, 20(1): 195–196, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fix:1993:BRS

[Fix93] George J. Fix. Book review:

Scientific Computations on Mathematical Problems and Conjectures (Richard S. Varga). SIAM Review, 35 (2):318–320, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Filip:2019:SRF

[FJT19] Silviu Filip, Aurya Javeed, and Lloyd N. Trefethen. Smooth random functions, random ODEs, and Gaussian processes. SIAM Review, 61 (1):185–205, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flatto:1962:RDI

[FK62] Leopold Flatto and Alan G. Konheim. The random division of an interval and the random covering of a circle. SIAM Review, 4(3): 211–222, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feige:2006:PAM

[FK06] U. Feige and R. Krauthgamer. A polylogarithmic approximation of the minimum bisection (reprinted from SIAM journal on computing, vol 31, 2002). SIAM Review, 48(1): 99–130, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Finkel:2006:CAM

[FKL^{+}06] Daniel E. Finkel, Christopher Kuster, Matthew Lasater, Rachel Levy, Jill P. Reese, and Ilse C. F. Ipsen. Communicating applied mathematics: Four examples. SIAM Review, 48(2):359– 389, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fields:1968:HS

[FL68] Jerry L. Fields and Yudell L. Luke. A hypergeometric series. SIAM Review, 10(4):

452, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fields:1970:HS

[FL70] Jerry L. Fields and Yudell L.

Luke. A hypergeometric series. SIAM Review, 12(4): 586–588, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fossum:1981:MMT

[FL81] Timothy V. Fossum and Gilbert N. Lewis. A mathematical model for trailertruck jackknifing. SIAM Review, 23(1):95–99, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fisher:1982:MP

[FL82] N. I. Fisher and A. J. Lee.

A maximum problem. SIAM Review, 24(1):76, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fields:1984:IRH

[FL84] Jerry L. Fields and Yudell L.

Luke. Inequality for the roots of a hypergeometric polynomial. SIAM Review, 26(2): 269–270, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feppon:2018:DON

[FL18] Florian Feppon and Pierre F. J. Lermusiaux. Dynamically orthogonal numerical schemes for efficient stochastic advection and Lagrangian transport. SIAM Review, 60 (3):595–625, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flanigan:1966:BRC

[Fla66] Larry K. Flanigan. Book review: Computer Software. Programming Systems for Digital Computers (Ivan Flores). SIAM Review, 8(4): 549–550, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flanigan:1967:BRP

[Fla67] L. K. Flanigan. Book review: PL/I Programming Primer (Gerald M. Weinberg). SIAM Review, 9(3):613, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flajolet:1997:BRA

[Fla97] Philippe Flajolet. Book review: Analysis of Algorithms — Computational Methods and Mathematical Tools, by Micha Hofri. SIAM Review, 39(2):341–345, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flaherty:1999:PTIa

[Fla99a] J. Flaherty. Problems and techniques — introduction. SIAM Review, 41(2):267– 268, ???? 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flaherty:1999:PTa

[Fla99b] Joe Flaherty. Problems and techniques. SIAM Review, 41(1):83–84, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97020.

Flaherty:1999:PTb

[Fla99c] Joe Flaherty. Problems and techniques. SIAM Review, 41(2):267–268, June 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97025.

Flaherty:1999:PTc

[Fla99d] Joe Flaherty. Problems and techniques. SIAM Review, 41(3):511, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97030.

Flaherty:1999:PTIb

[Fla99e]  Joe Flaherty.  Problems

and techniques — introduction. SIAM Review, 41(4):719, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97036.

Flaherty:2000:PTa

[Fla00a]  Joe Flaherty.  Problems

and techniques. SIAM Review, 42(1):41, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97042.

Flaherty:2000:PTb

[Fla00b]  Joe Flaherty.  Problems

and techniques. SIAM Review, 42(2):231, June 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97046.

Flaherty:2000:PTc

[Fla00c] Joe Flaherty. Problems and techniques. SIAM Review, 42(3):415, September 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97052.

Flaherty:2000:PTI

[Fla00d]  Joe Flaherty.  Problems

and techniques — introduction. SIAM Review, 42(4):655, December 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97057.

Flaherty:2001:PTa

[Fla01a]  Joe Flaherty.  Problems

and techniques. SIAM Review, 43(1):87, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97062.

Flaherty:2001:PTb

[Fla01b]  Joe Flaherty.  Problems

and techniques. SIAM Review, 43(2):287, June 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97067.

Flaherty:2001:PTc

[Fla01c] Joe Flaherty. Problems and techniques. SIAM Review, 43(3):467, September 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97072.

Flaherty:2001:PTd

[Fla01d] Joe Flaherty. Problems and techniques. SIAM Review, 43(4):621, December 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97077.

Flaherty:2002:PTI

[Fla02a]  J. Flaherty.  Problems

and techniques — introduction. SIAM Review, 44 (4):599, ???? 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flaherty:2002:PTa

[Fla02b]  Joe Flaherty.  Problems

and techniques. SIAM Review, 44(1):55, March 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97082.

Flaherty:2002:PTb

[Fla02c]  Joe Flaherty.  Problems

and techniques. SIAM Review, 44(2):225, June 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97087.

Flaherty:2002:PTc

[Fla02d] Joe Flaherty. Problems and techniques. SIAM Review, 44(3):371, September 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97092.

Flaherty:2002:PTd

[Fla02e] Joe Flaherty. Problems and techniques. SIAM Review, 44(4):599, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97097.

Flaherty:2003:PTIa

[Fla03a]  J. Flaherty.  Problems

and techniques — introduction. SIAM Review, 45 (1):51, ???? 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flaherty:2003:PTIb

[Fla03b]  J. Flaherty.  Problems

and techniques — introduction. SIAM Review, 45 (2):257, ???? 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flaherty:2003:PTa

[Fla03c]  Joe Flaherty.  Problems

and techniques. SIAM Review, 45(1):51, March 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97102.

Flaherty:2003:PTb

[Fla03d]  Joe Flaherty.  Problems

and techniques. SIAM Review, 45(2):257, June 2003.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97107.

Flaherty:2003:PTc

[Fla03e] Joe Flaherty. Problems and techniques. SIAM Review, 45(4):687, December 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97117.

Flaherty:2003:PTIc

[Fla03f]  Joe Flaherty.  Problems

and techniques — introduction. SIAM Review, 45 (3):483, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97112.

Flaherty:2004:I

[Fla04a] J. Flaherty. Introduction. SIAM Review, 46(4):635– 636, ???? 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flaherty:2004:PTa

[Fla04b]  Joe Flaherty.  Problems

and techniques. SIAM Review, 46(1):47, March 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97122.

Flaherty:2004:PTb

[Fla04c] Joe Flaherty. Problems and techniques. SIAM Review, 46(2):187–188, June 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97126.

Flaherty:2004:PTc

[Fla04d] Joe Flaherty. Problems and techniques. SIAM Review, 46 (3):415–416, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97131.

Flaherty:2004:PTd

[Fla04e] Joe Flaherty. Problems and techniques. SIAM Review, 46 (4):635–636, December 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97136.

Flaherty:2005:PTa

[Fla05a] Joe Flaherty. Problems and techniques. SIAM Review, 47(1):41–42, March 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97141.

Flaherty:2005:PTb

[Fla05b] Joe Flaherty. Problems and techniques. SIAM Review, 47(2):245–246, June 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97146.

Flaherty:2005:PTI

[Fla05c]  Joe Flaherty.  Problems

and techniques — introduction. SIAM Review, 47(3): 445–446, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97151.

Friedman:1996:IMC

[FLC96] Avner Friedman, Walter Littman, and Ellis Cumberbatch. Industrial mathematics: a course in solving real-world problems. SIAM Review, 38(2):341–??, June 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleishman:1962:BRI

[Fle62] B. A. Fleishman. Book review: Introduction to Partial Differential Equations (Donald Greenspan). SIAM Review, 4(3):260–261, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1963:BRC

[Fle63a] W. Fleming. Book review:

The Calculus of Variations (N. I. Akhiezer). SIAM Review, 5(2):161–162, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1963:BRI

[Fle63b] W. H. Fleming. Book review: An Introduction to the Calculus of Variations (L. A. Pars). SIAM Review, 5 (4):382, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1964:BRM

[Fle64] Wendell Fleming. Book review: Mathematical Optimization Techniques (Richard Bellman, ed.). SIAM Review, 6(1):71–73, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleishman:1965:FOC

[Fle65] B. A. Fleishman. Forced oscillations and convex superposition in piecewise-linear systems. SIAM Review, 7(2): 205–222, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1969:OCP

[Fle69] Wendell H. Fleming. Optimal continuous-parameter stochastic control. SIAM Review, 11(4):470–509, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1972:BRD

[Fle72] Wendell H. Fleming. Book review: Differential Games (A. Friedman). SIAM Review, 14 (4):661, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleishman:1973:BRM

[Fle73] B. A. Fleishman. Book review: Modern Mathematical Methods in Technology (S. Fenyo and T. Frey). SIAM Review, 15(2):404– 405, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleishman:1976:BRN

[Fle76] B. A. Fleishman. Book review: Nonlinear System Theory: a Functional Analysis Approach (Jack M. Holtzman). SIAM Review, 18(2): 308–310, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1977:BRSa

[Fle77a] Wendell H. Fleming. Book review: Stochastic Differential Equations and Applications, Vol. 1 (Avner Friedman). SIAM Review, 19(2):

366–367, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1977:BRSb

[Fle77b] Wendell H. Fleming. Book review: Stochastic Differential Equations and Applications, Vol. 2 (Avner Friedman). SIAM Review, 19(2): 367–368, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1979:BRP

[Fle79] Wendell H. Fleming. Book review: Probability Methods for Approximations in Stochastic Control and Elliptic Equations (Harold J. Kushner). SIAM Review, 21 (1):149–150, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1982:BRT

[Fle82] Wendell H. Fleming. Book review: Theory and Applications of Stochastic Differential Equations (Zeev Schuss). SIAM Review, 24(3):357– 358, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1983:BRC

[Fle83] Wendell H. Fleming. Book review: Controlled Diffusion Processes (N. V. Krylov).

SIAM Review, 25(2):287– 288, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1984:SCA

[Fle84a] W. H. Fleming. Stochastic calculus and applications — Elliott, R. J. SIAM Review, 26(2):303–304, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1984:BRS

[Fle84b] Wendell H. Fleming. Book review: Stochastic Calculus and Applications (Robert J. Elliott). SIAM Review, 26 (2):303–304, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1985:BRG

[Fle85] Wendell H. Fleming. Book review: Generalized Solutions of Hamilton–Jacobi Equations (P.-L. Lions). SIAM Review, 27(2):268–269, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1988:BRS

[Fle88] W. H. Fleming. Book review: A Stochastic Maximum Principle for Optimal Control of Diffusions (U. G. Haussmann). SIAM Review, 30(3):512, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1989:BRP

[Fle89] W. H. Fleming. Book review:

Perturbation Methods in Optimal Control (Alain Bensoussan). SIAM Review, 31 (4):693–694, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fleming:1990:BRO

[Fle90] Wendell H. Fleming. Book review: Optimal Control of Diffusion Processes (Vivek S. Borkar). SIAM Review, 32 (3):503–504, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fountoulakis:2023:FBA

[FLGM23] Kimon Fountoulakis, Meng Liu, David F. Gleich, and Michael W. Mahoney. Flowbased algorithms for improving clusters: a unifying framework, software, and performance. SIAM Review, 65(1):59–143, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1333055.

Farina:2022:LLI

[FLM22] Leandro Farina, Guillaume Lang, and P. A. Martin. Love–Lieb integral equations:

Applications, theory, approximations, and computations. SIAM Review, 64 (4):831–865, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1371038.

Fosdick:2018:CRG

[FLNU18] Bailey K. Fosdick, Daniel B. Larremore, Joel Nishimura, and Johan Ugander. Configuring random graph models with fixed degree sequences. SIAM Review, 60 (2):315–355, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flores:1969:BRR

[Flo69]  Ivan Flores.  Book re-

view: Residue Arithmetic and Its Application to Computer Technology (Nicholas S. Szabo and Richard I. Tanaka). SIAM Review, 11 (1):103–104, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flor:1977:CI

[Flo77] Peter Flor. Conjectured inequalities. SIAM Review, 19 (3):563, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1981:IS

[FLR81a] H. E. Fettis, Y. L. Luke, and O. G. Ruehr. An infinite series. SIAM Review, 23(2): 258–259, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1981:BRI

[FLR^{+}81b] H. E. Fettis, Y. L. Luke, O. G. Ruehr, A. A. Jagers, O. P. Lossers, and O. G. Ruehr. Book review: An Infinite Series (M. L. Glasser and A. J. Jerri). SIAM Review, 23 (2):258–259, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fluellen:1978:CMM

[Flu78]  A. H. Fluellen.  Conroy

method of multidimensional integration for multiplescattering problems. SIAM Review, 20(3):624, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flury:1990:ARS

[Flu90] Bernard D. Flury. Acceptancerejection sampling made easy. SIAM Review, 32(3):474– 476, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fluky:1993:BRU

[Flu93a] Bernard Fluky. Book review:

A User’s Guide to Principal Components (Edward Jackson). SIAM Review, 35(1): 148–149, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flury:1993:UGP

[Flu93b] Bernard Flury. A user’s guide to principal components (Edward Jackson). SIAM Review, 35(1):148– ??, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flynn:1976:SSP

[Fly76] J. Flynn. Steady-state policies in dynamic programming — general case. SIAM Review, 18(4):803, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flynn:1989:SRC

[Fly89a] L. E. Flynn. Shape of a rollercoaster track. SIAM Review, 31(2):320–324, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flynn:1989:BRS

[Fly89b] Lawrence E. Flynn. Book review: Shape of a Roller-Coaster Track (John S. Lew). SIAM Review, 31(2):320– 324, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fillmore:1968:LRS

[FM68] Jay P. Fillmore and Morris L.

Marx. Linear recursive sequences. SIAM Review, 10 (3):342–353, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Forgoston:2018:PNI

[FM18] Eric Forgoston and Richard O.

Moore. A primer on noiseinduced transitions in applied dynamical systems. SIAM Review, 60(4):969– 1009, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flatto:1973:IDP

[FN73] L. Flatto and D. J. Newman. An inverse drum problem. SIAM Review, 15(4): 788, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fabijonas:1999:RAE

[FO99] Bruce R. Fabijonas and F. W. J. Olver. On the reversion of an asymptotic expansion and the zeros of the Airy functions. SIAM Review, 41 (4):762–773, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34953.

Fokas:1984:BRA

[Fok84]  A. S. Fokas.  Book re-

view: Asymptotic Methods in Nonlinear Wave Theory (A. Jeffrey and T. Kawahara). SIAM Review, 26(1): 141–142, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Foley:1964:TPM

[Fol64] James F. Foley. N th power of a matrix. SIAM Review, 6(4):456, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Foley:1973:SCM

[Fol73] J. F. Foley. A special class of matrices. SIAM Review, 15(4):787, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Foots:1963:BRO

[Foo63]  J. R. Foots.  Book re-

view: Optimization Techniques (G. Leitmann, ed.). SIAM Review, 5(3):291– 292, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Forsythe:1959:RNR

[For59] George E. Forsythe. Reprint of a note on rounding-off errors. SIAM Review, 1(1):

66–67, January 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Forsythe:1967:TCM

[For67] George E. Forsythe. Today’s computational methods of linear algebra. SIAM Review, 9(3):489–515, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Reprinted in Studies in Numerical Analysis 1, SIAM, Philadelphia, 1968.

Foregger:1975:BRM

[For75] Thomas Foregger. Book review: A Matrix Problem (C. R. Crawford). SIAM Review, 17(4):695–696, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fordlivene:1976:MMK

[For76] C. Fordlivene. Mathematical model of kinetics of coal carbonization to Coke and volatile by-products. SIAM Review, 18(4):803, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ford:1992:BRT

[For92] William. F. Ford. Book review: A Third-Order Differential Equation (M. A. Abdelkader). SIAM Review, 34 (1):121–122, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ford:1996:CTP

[For96] William F. Ford. Clarification of: “Turn performance of aircraft” [SIAM Rev. 36 (1994), no. 3, 470–473] by J. S. Wolper. SIAM Review, 38 (2):309–312, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/p309.htm.  See

[Wol94c].

Fornberg:1998:CWF

[For98a] B. Fornberg. Calculation of weights in finite difference formulas. SIAM Review, 40 (3):685–691, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fornberg:1998:CNC

[For98b] Bengt Fornberg. Classroom note: Calculation of weights in finite difference formulas. SIAM Review, 40(3): 685–691, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32250.

Fornberg:2021:IAT

[For21] Bengt Fornberg. Improving the accuracy of the trapezoidal rule. SIAM Review, 63 (1):167–180, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fosdick:1968:MCM

[Fos68] Lloyd D. Fosdick. The Monte Carlo method in quantum statistics. SIAM Review, 10 (3):315–328, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Foster:1981:SII

[Fos81] L. V. Foster. Some inequalities involving statistical expressions. SIAM Review, 23 (2):256, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fossum:1990:BRT

[Fos90] A. F. Fossum. Book review: Topics in Boundary Element Research, Volume 4: Applications in Geomechanics (C. A. Brebbia, ed.). SIAM Review, 32(1): 173–174, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fukushima:1996:DFS

[FOTG96] Masatoshi Fukushima, Yoichi Oshima, Masayoshi Takeda, and Ronald Getoor. Dirichlet forms and symmetric Markov processes. SIAM Review, 38 (1):166–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fourer:1984:SMS

[Fou84] Robert Fourer. Staircase matrices and systems. SIAM Review, 26(1):1–70, January 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fox:1967:ESO

[Fox67] Bennett Fox. Existence of stationary optimal policies for some Markov renewal programs. SIAM Review, 9(3): 573–576, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fox:1971:BRM

[Fox71]  Bennett Fox.  Book re-

view: Markovian Decision Processes (H. Mine and S. Osaki). SIAM Review, 13(4): 585–586, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ford:1996:ACN

[FP96] William F. Ford and James A. Pennline. Accelerated convergence in Newton’s method. SIAM Review, 38(4):658– 659, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/29297.

Ferris:1997:EEA

[FP97a] M. C. Ferris and J. S. Pang. Engineering and economic applications of complementarity problems. SIAM Review, 39(4):669– 713, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28596.

Ford:1997:I

[FP97b] W. F. Ford and J. A.

Pennline. An integral. SIAM Review, 39(3):524– 527, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fitt:2004:SCM

[FP04] Alistair D. Fitt and Colin P.

Please. On the separation of coconuts: a modeling week study. SIAM Review, 46(1):128–139, March 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42106.

Feng:2021:PHG

[FP21] Michelle Feng and Mason A.

Porter. Persistent homology of geospatial data: a case study with voting. SIAM Review, 63(1):67–99, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ferziger:1997:CMF

[FPM97] Joel H. Ferziger, Milovan Peric, and K. W. Morton. Computational methods for fluid dynamics. SIAM Review, 39(4):801– ??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fraser:1963:BRL

[FR63] W. Fraser and Pierre Robert. Book review: On a Least Square Approximation (D. J. Newman). SIAM Review, 5 (4):374–375, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fox:1966:CMD

[FR66] David W. Fox and Werner C. Rheinboldt. Computational methods for determining lower bounds for eigenvalues of operators in Hilbert spaces. SIAM Review, 8(4): 427–462, October 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freedman:1983:DPR

[FR83] H. I. Freedman and S. D. Riemenschneider. Determining the path of the rear wheels of a bus. SIAM Review, 25(4):561–567, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franz:2011:CNM

[FR11] Sebastian Franz and Hans-Görg Roos. The capriciousness of numerical methods for singular perturbations. SIAM Review, 53(1): 157–173, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i1/p157_s1.

Fercoq:2016:OHD

[FR16] Olivier Fercoq and Peter Richtárik. Optimization in high dimensions via accelerated, parallel, and proximal coordinate descent. SIAM Review, 58(4):739– 771, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franck:1965:OSPa

[Fra65a] Wallace Franck. An optimal search problem. SIAM Review, 7(4):503–512, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franck:1965:OSPb

[Fra65b] Wallace Franck. An optimal search problem. SIAM Review, 7(4):562, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [Fra66a].

Franklin:1965:NSS

[Fra65c] Joel N. Franklin. Numerical simulation of stationary and non-stationary Gaussian random processes. SIAM Review, 7(1):68–80, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franck:1966:EOS

[Fra66a]  Wallace Franck.  Errata:

“An Optimal Search Problem”. SIAM Review, 8(4): 524, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Fra65b].

Frank:1966:BRF

[Fra66b] Evelyn Frank. Book review: Factorization of a Set of Polynomials (N. Mullineux). SIAM Review, 8(1):115– 116, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frankel:1972:BRB

[Fra72] A. Frankel. Book review: Bigradients (Alston S. Householder). SIAM Review, 14(3): 495–496, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franke:1975:CAR

[Fra75] R. Franke. Convergence of an algorithm for rational Chebyshev approximation. SIAM Review, 17(2):378, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fransen:1983:FPI

[Fra83] A. Fransén. A factorial polynomial inequality. SIAM Review, 25(1):97, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franck:1986:BRO

[Fra86a] Wallace Franck. Book review:

An Optimal Search (Richard Bellman). SIAM Review, 28 (2):233–234, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franke:1986:BRE

[Fra86b] Richard Franke. Book review: The Elements of Graphing Data (William S. Cleveland). SIAM Review, 28 (3):414–416, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franklin:1986:BRF

[Fra86c] Joel Franklin. Book review:

Fixed Point Theorems with Applications to Economics and Game Theory (Kim C. Border). SIAM Review, 28 (2):255, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franklin:1988:BRC

[Fra88] Wm. Randolph Franklin. Book review: Computational Geometry: an Introduction (Franco P. Preparata and Michael Ian Shamos). SIAM Review, 30(4):669– 670, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Franklin:1989:BRA

[Fra89] Wm. Randolph Franklin. Book review: Art Gallery Theorems and Algorithms (Joseph O’Rourke). SIAM Review, 31(2):342–343, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frazho:1995:BRS

[Fra95a] A. E. Frazho. Book review: Schur’s Algorithm and Several Applications (M. Bakonyi and T. Constantinescu). SIAM Review, 37(1): 121–122, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frazier:1995:BRF

[Fra95b] Michael Frazier. Book review: A Friendly Guide to Wavelets (Gerald Kaiser). SIAM Review, 37(4):624– 626, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frankowska:2023:Ea

[Fra23a] Hélène Frankowska. Education. SIAM Review, 65 (1):259–260, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N975648.

Frankowska:2023:Eb

[Fra23b] Hélène Frankowska. Education. SIAM Review, 65 (2):537, ???? 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/23N975703.

Frankowska:2023:Ec

[Fra23c] Hélène Frankowska. Education. SIAM Review, 65 (3):867, ???? 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/23N975752.

Frankowska:2023:E

[Fra23d] Hèléne Frankowska. Education. SIAM Review, 65(4): 1135, November 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frankowska:2024:Ea

[Fra24a] Hélène Frankowska. Education. SIAM Review, 66(1): 147, February 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frankowska:2024:Eb

[Fra24b] Hélène Frankowska. Education. SIAM Review, 66 (2):353, May 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freedman:1969:EEN

[Fre69] H. I. Freedman. An explicit estimate on the norm of the inverse of a matrix. SIAM Review, 11(2): 254–256, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freedman:1975:PAM

[Fre75] M. I. Freedman. Perturbation analysis for minimum time problem in optimal control. SIAM Review, 17(2): 378, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freedman:1978:BRC

[Fre78] H. I. Freedman. Book review: Calculator Calculus (George McCarty). SIAM Review, 20 (1):196–197, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freedman:1979:BRM

[Fre79] H. I. Freedman. Book review: Mathematical Models.

Mechanical Vibrations, Population Dynamics, and Traffic Flow (Richard Haberman). SIAM Review, 21(3): 413–414, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frederickson:1980:BRD

[Fre80] P. O. Frederickson. Book review: A Differential Equation (J. D. Love). SIAM Review, 22(4):503–504, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fredricksen:1982:SFL

[Fre82] Harold Fredricksen. A survey of full length nonlinear shift register cycle algorithms. SIAM Review, 24(2):195–221, April 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2029360.

Freedman:1983:BRC

[Fre83] H. I. Freedman. Book review:

Case Studies in Mathematical Modelling (R. Bradley, R. D. Girson and M. Cross, eds.). SIAM Review, 25(3): 409–410, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freedman:1986:BRC

[Fre86a] H. I. Freedman. Book review: Concepts and Models of a Quantitative Sociology-The Dynamics of Interacting Populations (W. Weidlich and G. Haag). SIAM Review, 28(2): 255–256, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frenzen:1986:SPT

[Fre86b] C. L. Frenzen. Sign pattern of terms of a Maclaurin series. SIAM Review, 28(4): 565, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freund:1986:BRC

[Fre86c] Robert M. Freund. Book review: Complementary Pivoting on a Pseudomanifold Structure with Applications in the Decision Sciences (F. J. Gould and J. W. Tolls). SIAM Review, 28(3):416– 417, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freudenstein:1990:BRR

[Fre90] Ferdinand Freudenstein. Book review: Rational Kinematics (Jorge Angeles). SIAM Review, 32(4):695–696, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freedman:1995:BRP

[Fre95]  H. I. Freedman.  Book

review: Periodic Motions (Miklós Farkas). SIAM Review, 37(4):626, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freitas:2014:OBP

[Fre14] Pedro Freitas. Optimal ball placement in rugby conversions. SIAM Review, 56(4): 673–690, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fritz:1959:BRN

[Fri59] W. Barkley Fritz. Book review: Numerical Analysis (Kaiser S. Kunz). SIAM Review, 1(2):175–176, ???? 1959. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Friedman:1963:SAF

[Fri63]  Henry D. Friedman.  A

sharp autocorrelation function. SIAM Review, 5(3): 275, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Friedman:1964:BRU

[Fri64] Henry D. Friedman. Book review: An Unfriendly Seating Arrangement (Dave Freedman). SIAM Review, 6(2): 180–182, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Frisch:1971:BRS

[Fri71] H. L. Frisch. Book review: The Statistical Mechanics of Simple Liquids (Stuart A. Rice and Peter Gray). SIAM Review, 13(4):586, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Friedlander:1978:BRR

[Fri78a] Susan Friedlander. Book review: Rheometry (K. Walters). SIAM Review, 20(1): 197–198, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Friedman:1978:OSP

[Fri78b] A. Friedman. Optimal stopping problems in stochastic control. SIAM Review, 20 (3):624, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Friedman:1979:OSP

[Fri79a] Avner Friedman. Optimal stopping problems in stochastic control. SIAM Review, 21 (1):71–80, January 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Friedman:1979:TDF

[Fri79b] Avner Friedman. Time dependent free boundary problems. SIAM Review, 21(2):

213–221, April 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Friedrichs:1980:NHS

[Fri80] K. O. Friedrichs. Von Neumann’s Hilbert space theory and partial differential equations. SIAM Review, 22(4): 486–493, October 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Friedlander:1982:BRF

[Fri82] Susan Friedlander. Book review: Free Oscillations of the Earth (E. R. Lapwood and T. Usami). SIAM Review, 24(4): 496–497, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Friedlander:1986:BRI

[Fri86] Susan Friedlander. Book review: Imaging the Earth’s Interior (Jon F. Claerbou). SIAM Review, 28(2):256– 257, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Froese:1960:BRI

[Fro60] Charlotte Froese. Book review: Introduction to the Laplace Transform (Dio L. Holl and Clair G. Maple). SIAM Review, 2(3):223, ???? 1960. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Fromovitz:1973:BRP

[Fro73] Stan Fromovitz. Book review: Probabilistic Programming (S. Vajda). SIAM Review, 15(2):401, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fromovitz:1974:BRS

[Fro74] Stan Fromovitz. Book review: Stochastic Programming: Methods and Applications (Jati K. Sengupta). SIAM Review, 16(3):399– 400, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fromovitz:1980:BRN

[Fro80] Stan Fromovitz. Book review: Nonlinear Programming: Theory and Algorithms (Mokhtar S. Bazaraa and C. M. Shetty). SIAM Review, 22(4):523–524, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fromovitz:1983:BRR

[Fro83] Stan Fromovitz. Book review: Redundancy and Linear Programs (J. Telgen). SIAM Review, 25(1):117– 118, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fromovitz:1985:BRM

[Fro85] Stan Fromovitz. Book review: Mathematical Methods in Management (Geoffrey Gregory). SIAM Review, 27 (3):465, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fromovitz:1986:BRM

[Fro86] Stan Fromovitz. Book review: Methods for Solving Incorrectly Posed Problems (V. A. Morozov). SIAM Review, 28(1):102, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fromovitz:1987:BRA

[Fro87a] Stan Fromovitz. Book review: Algorithms and Approximations for Queueing Systems. (M. H. van Hoorn). SIAM Review, 29(1):142– 143, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fromovitz:1987:BRI

[Fro87b] Stan Fromovitz. Book review: Introduction to Probability and Mathematical Statistics (Václav Fabian and James Hannan). SIAM Review, 29(3):478–479, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fryer:1964:RLN

[Fry64] William D. Fryer. Resistance of a ladder network. SIAM Review, 6(3):310, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Freedman:1962:USA

[FS62] Dave Freedman and Larry Shepp. An unfriendly seating arrangement. SIAM Review, 4(2):150, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1968:BRI

[FS68] Henry E. Fettis and Anthony J. Strecok. Book review: An Integral Recurrence Relation (M. B. Mc-Neil). SIAM Review, 10(2): 227–229, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Faudree:1973:BRS

[FS73] R. J. Faudree and R. H. Schelp. Book review: A Solved and Unsolved Graph Coloring Problem (P. Erdős). SIAM Review, 15(2):377– 380, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fox:1976:SSO

[FS76] D. W. Fox and V. G. Sigillito. Steady-state oscillations in a buoyant fluid. SIAM Review, 18(4):803, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feldman:1978:TCB

[FS78a] M. W. Feldman and L. C.

Sforza. Theoretical connections between biological and cultural evolution. SIAM Review, 20(3):623, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fordlivene:1978:OCI

[FS78b] C. Fordlivene and S. T.

Seo. Optimal control of infestations in harvested fruits via dynamic programming. SIAM Review, 20 (3):624, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flatto:1989:PN

[FS89] Leopold Flatto and Larry Shepp. Problem of the Nile. SIAM Review, 31(2): 318, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Faber:1996:FDP

[FS96] T. E. Faber and P. G.

Saffman. Fluid dynamics for physicists. SIAM Review, 38 (4):700–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fokas:2012:SOS

[FS12] A. S. Fokas and E. A. Spence. Synthesis, as opposed to separation, of variables. SIAM Review, 54(2): 291–324, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flajolet:1993:FSP

[FSP^{+}93a] P. Flajolet, B. Salvy, H. Prodinger, M. L. Larsen, and J. L. Lavoie. A finite sum of products of binomial coefficients. SIAM Review, 35(4): 645–647, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flajolet:1993:BRF

[FSP93b] P. Flajolet, B. Salvy, and Helmut Prodinger. Book review: A Finite Sum of Products of Binomial Coefficients (C. C. Grosjean). SIAM Review, 35 (4):645–647, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Flatto:1990:PN

[FSRJ90] L. Flatto, L. Shepp, M. Renardy, and A. A. Jagers. Problem of the Nile. SIAM Review, 32(2):302–304, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Falk:2006:RAN

[FSV06] James E. Falk, Nozer D.

Singpurwalla, and Yefim Y. Vladimirsky. Reliability allocation for networks and systems. SIAM Review, 48(1): 43–65, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Farrell:1966:BRL

[FSW^{+}66] J. L. Farrell, J. C. Stuelpnagel, R. H. Wessner, J. R. Velman, and J. E. Brook. Book review: A Least Squares Estimate of Satellite Attitude (Grace Wahba). SIAM Review, 8(3):384– 386, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fu:1965:NMG

[FSY65] Yumin Fu, Mandyam D. Srinath, and Andrew T. Yen. A note on multivariate generating functions and applications to discrete stochastic processes. SIAM Review, 7 (1):31–41, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Feldstein:1976:ABV

[FT76] A. Feldstein and J. F. Traub.

Asymptotic behavior of vector recurrences with applications. SIAM Review, 18 (4):802, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Finn:2011:TOR

[FT11] Matthew D. Finn and Jean-Luc Thiffeault. Topological optimization of rod-stirring devices. SIAM Review, 53 (4):723–743, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v53/i4/p723_s1.

Fahrmeir:1995:MSM

[FTM95] Ludwig Fahrmeir, Gerhard Tutz, and Brian D. Marx. Multivariate statistical modelling based on generalized linear models. SIAM Review, 37(4):633–??, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fullerton:1976:NBU

[Ful76] L. W. Fullerton. Nearly best uniform rational approximations. SIAM Review, 18(4): 803–804, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fullerton:1984:IP

[Ful84] L. W. Fullerton. An iteration problem. SIAM Review, 26 (2):279–281, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fusaro:1968:CKT

[Fus68a] B. A. Fusaro. The Cauchy– Kowalewski theorem and a singular initial value problem. SIAM Review, 10(4): 417–421, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fusaro:1968:VIE

[Fus68b] B. A. Fusaro. A Volterra integral equation. SIAM Review, 10(3):377, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fusaro:1969:VIE

[Fus69] B. A. Fusaro. A Volterra integral equation. SIAM Review, 11(4):632, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fulkerson:1962:ASM

[FW62] D. R. Fulkerson and P. Wolfe.

An algorithm for scaling matrices. SIAM Review, 4(2): 142–146, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fraser:1966:RCC

[FW66] W. Fraser and M. W. Wilson. Remarks on the Clenshaw–Curtis quadrature scheme. SIAM Review, 8(3): 322–327, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fettis:1982:BRN

[FW82] H. E. Fettis and G. Wanner. Book review: A Nonlinear Second Order Differential Equation (E. Y. Rodin). SIAM Review, 24(4): 480–481, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Furstenberg:1963:BRE

[FZ63] Harry Furstenberg and F. Zetto. Book review: An Expectation Problem (P. L. van Albada). SIAM Review, 5(3): 283–287, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1963:BRI

[GA63] R. C. Gilbert and U. S. Army. Book review: Introduction to Hilbert Space (S. K. Berberian). SIAM Review, 5(2): 162–163, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grossman:1988:DIS

[GAB^{+}88] N. Grossman, M. Ashbaugh, W. Bohm, D. R. Breach, J. Case, C. Qun, C. Georghiou, M. L. Glasser, J. A. Grzesik, E. Harrell, R. N. Hill, W. B. Jordan, H. K. Kuiken, K. W.

Lau, O. P. Lossers, M. Renardy, R. Richberg, O. G. Ruehr, P. Smith, and J. E. Wilkins. A definite integral from surface gravity waves. SIAM Review, 30(4): 660–661, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ghigliazza:2005:SSR

[GAHK05] R. M. Ghigliazza, R. Altendorfer, P. Holmes, and D. Koditschek. A simply stabilized running model. SIAM Review, 47(3):519– 549, September 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/62659.

Granados:2017:PAI

[GAK17] Albert Granados, Lluís Alsedà, and Maciej Krupa. The period adding and incrementing bifurcations: From rotation theory to applications. SIAM Review, 59(2): 225–292, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gallopoulos:1995:BRP

[Gal95] E. Gallopoulos. Book review:

Parallel Computing Using the Prefix Problem (S. Lakshmivarahan and Sudarshan K. Dhall). SIAM Review, 37(4): 626–628, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gupta:1993:MMD

[GAM93] Sunetra Gupta, Roy M. Anderson, and Robert M. May. Mathematical models and the design of public health policy: HIV and antiviral therapy. SIAM Review, 35 (1):1–16, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ganapol:1989:PTP

[Gan89] Barry D. Ganapol. Photon transport in plant canopies. SIAM Review, 31(2):319– 320, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ganapol:1990:PTP

[Gan90] Barry D. Ganapol. Photon transport in plant canopies. SIAM Review, 32(2):305– 307, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gander:1998:NOI

[Gan98a] M. J. Gander. Note on the optimal intercept time of vessels to a nonzero range. SIAM Review, 40(3):673, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gander:1998:CNN

[Gan98b] Martin J. Gander. Classroom note: Note on the optimal intercept time of vessels to a nonzero range. SIAM Review, 40(3):673, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31924.

Gao:2018:MPT

[Gao18] Zhihan Gao. On the metric s–t path traveling salesman problem. SIAM Review, 60 (2):409–426, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gardner:1960:SSD

[Gar60] G. H. F. Gardner. Steadystate diffusion-convection. SIAM Review, 2(1):39–40, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gardner:1961:BRS

[Gar61] G. H. F. Gardner. Book review: Steady-State Diffusion-Convection (G. H. F. Gardner). SIAM Review, 3(1): 74–76, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Garfinkel:1967:BRM

[Gar67] Boris Garfinkel. Book review:

Mathematical Introduction to Celestial Mechanics (Harry Pollard). SIAM Review, 9(1): 139–140, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gardner:1969:SIP

[Gar69] L. A. Gardner, Jr. Sum of inverse powers of cosines. SIAM Review, 11(4):621, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Garey:1975:SPV

[Gar75] L. Garey. Starting procedures for Volterra integral equations using generalized quadrature rules. SIAM Review, 17(2):378, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Garabedian:1989:CMF

[Gar89] P. R. Garabedian. Computational models in fusion research. SIAM Review, 31 (4):542–559, December 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gardner:1991:BRS

[Gar91] Carl L. Gardner. Book review: Semiconductor Equations (P. A. Markowich, C. A. Ringhofer, and C. Schmeiser). SIAM Review, 33 (4):675–677, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gardiner:1993:BRF

[Gar93] Judith D. Gardiner. Book review: Fundamentals of Matrix Computations (David S. Watkins). SIAM Review, 35 (3):520–521, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gardner:1994:WAA

[Gar94a]  W. Gardner.  Wavelets:

Algorithms and applications (Yves Meyer): Author’s comment. SIAM Review, 36(3): 528–??, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gardner:1994:AC

[Gar94b] William A. Gardner. Authors comment. SIAM Review, 36 (3):528–530, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Garcia:2000:OSF

[Gar00] Antonio G. Garcia. Orthogonal sampling formulas: a unified approach. SIAM Review, 42(3):499– 512, September 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/36349.

Gaskell:1960:PMG

[Gas60] R. E. Gaskell. Professional mathematicians in government and industry. SIAM Review, 2(1):3–10, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gaskell:1963:BRF

[Gas63] R. E. Gaskell. Book review: Fundamentals of the Laplace Transform (C. J. Savant and E. C. Levy). SIAM Review, 5 (3):289–290, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gass:1970:BRS

[Gas70] Saul I. Gass. Book review: Sensitivitätsanalysen und Parametrische Programmierung (Werner Dinklebach). SIAM Review, 12(1): 165–166, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gasper:1971:BRP

[Gas71] George Gasper. Book review: A Positive Integral (Richard Askey). SIAM Review, 13(3): 396–397, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gasper:1976:DDP

[Gas76a] G. Gasper. 2-dimensional discrete probability distributions. SIAM Review, 18(1):

126–129, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gasper:1976:BRT

[Gas76b] George Gasper. Book review:

Two-Dimensional Discrete Probability Distributions (P. Beckmann). SIAM Review, 18(1):126–129, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gasper:1985:EHF

[Gas85a]  G. Gasper.  Evaluation

of a hypergeometric function. SIAM Review, 27 (2):255, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gasper:1985:BRE

[Gas85b] George Gasper. Book review: Evaluation of a Hypergeometric Function (Otto G. Ruehr). SIAM Review, 27(2):255, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gass:1995:BRU

[Gas95] Frederick Gass. Book review: Understanding the Infinite (Shaughan Lavine). SIAM Review, 37(3):464– 466, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gates:1964:NSD

[Gat64] Leslie D. Gates, Jr. Numerical solution of differential equations by repeated quadratures. SIAM Review, 6 (2):134–147, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gottlieb:1975:FOA

[GAT75] D. Gottlieb, S. S. Abarbanel, and E. Turkel. Fourth order accuracy finite-difference methods for hyperbolic systems. SIAM Review, 17 (2):379, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gatica:1988:BRO

[Gat88] Juan A. Gatica. Book review: Ordinary Differential Equations: Introduction to the Theory of Ordinary Differential Equations in the Real Domain (Jaroslav Kurzweil). SIAM Review, 30 (3):512–513, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gautschi:1962:DFS

[Gau62]  Walter Gautschi.  Diffu-

sion functions for small argument. SIAM Review, 4(3): 227–229, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gautschi:1967:CAT

[Gau67] Walter Gautschi. Computational aspects of threeterm recurrence relations. SIAM Review, 9(1):24–82, January 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/?SIR/9/24/
1.

Gautschi:1987:CIH

[Gau87] Walter Gautschi. A conjectured inequality for Hermite interpolation at the zeros of Jacobi polynomials. SIAM Review, 29(2): 297–298, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gautschi:2008:LEH

[Gau08] Walter Gautschi. Leonhard Euler: His life, the man, and his works. SIAM Review, 50 (1):3–33, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/?SIR/50/3/
1.

Gaviano:1982:LS

[Gav82] M. Gaviano. Limit of a sequence. SIAM Review, 24(3): 339–340, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gay:1976:IBM

[Gay76] D. M. Gay. Implementing Brown’s method. SIAM Review, 18(4):804, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gay:1978:SCP

[Gay78] D. M. Gay. Some convergence properties of Broyden method. SIAM Review, 20 (3):624, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenspan:1989:BRN

[GB89] D. Greenspan and B. F. Bowman. Book review: Numerical Algorithms (J. L. Mohamed and J. Walsh, eds.). SIAM Review, 31(2):343– 344, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibbs:1990:NES

[GBC^{+}90] A. Gibbs, J. P. Boyd, S. P. Chicatelli, J. A. Crow, C. Dorman, M. Gupta, A. Genz, C. C. Grosjean, H. Guggenheimer, M. Haag, S. Haber, K. Imre, H. Weitzner, W. B. Jordan, J. F. Kinkel, D. A. Kurtze, J. K. Larsen, K. W. Lau, P. D. Mccray, A. J. Oconnor, M. Renardy, R. S. Sezginer, E. L. Smith, H. A. Steinberg, R. Terrell, P. Wagner, N. Waller, J. E. Wilkins, G. Wilson, and J. C. Xu. Numerical evaluation of a slowly convergent series. SIAM Review, 32(3):481– 483, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1994:DI

[GBdD94] M. L. Glasser, J. Boersma, and P. J. de Doelder. A definite integral. SIAM Review, 36(4):656, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goodman:1978:ITD

[GBF78] R. Goodman, J. Bustoz, and A. Feldstein. Improved trailing digits estimates applied to optimal computer arithmetic. SIAM Review, 20 (3):625, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ghosh:2008:MER

[GBS08] Arpita Ghosh, Stephen Boyd, and Amin Saberi. Minimizing effective resistance of a graph. SIAM Review, 50 (1):37–66, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gobel:1969:BRN

[GC69] F. Göbel and L. Carlitz.

Book review: A Network Inequality (J. C. Turner and V.

Conway). SIAM Review, 11 (3):402–406, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Givens:1978:SOD

[GC78] C. Givens and L. Carlitz. System of 2nd order differential equations. SIAM Review, 20 (4):859, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Givens:1979:BRC

[GC79] C. Givens and D. G. Cantor. Book review: A Conjecture on Determinants (H. L. Langhaar and R. E. Miller). SIAM Review, 21(1): 144–146, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1978:BP

[GD78] M. L. Glasser and S. G. Davison. A bundling problem. SIAM Review, 20(1): 178–180, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1988:BRI

[GD88] M. L. Glasser and Lj. Damjanovr. Book review: An Identity (John W. Moon). SIAM Review, 30(2):317– 319, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1995:BRM

[GdM95] Carl C. Grosjean and Hans E.

de Meyer. Book review: A Minimization Problem (Jeffrey Shallit). SIAM Review, 37(3):451–458, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1991:BRTb

[GdMGR91] C. C. Grosjean, H. de Meyer, Alan Gibbs, and Michael Renardy. Book review: A Two-Point Boundary Problem for Airy Functions (Richard B. Evans). SIAM Review, 33(3): 477–479, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenbaum:1978:AIM

[GDR78] A. Greenbaum, P. Dubois, and G. Rodrigue. Approximating inverse of a matrix for use in iterative algorithms on vector processors. SIAM Review, 20(3): 625–626, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Geismar:2005:DCS

[GDS05] H. Neil Geismar, Milind W.

Dawande, and Suresh P. Sethi. Dominance of cyclic solutions and challenges in the scheduling of robotic cells. SIAM Review, 47(4): 709–721, ???? 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Groenewoud:1963:SCC

[GE63] C. Groenewoud and L. Eusanio. The smallest covering cone or sphere. SIAM Review, 5(2):156, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gear:1974:BRA

[Gea74] Bill Gear. Book review: Analysis of Discretization Methods for Ordinary Differential Equations (Hans J. Stetter). SIAM Review, 16 (4):561–562, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gear:1980:BRN

[Gea80]  C. W. Gear.  Book re-

view: Numerical Solution of Differential Equations (Isaac Fried). SIAM Review, 22(4): 512–513, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gear:1981:NSO

[Gea81] C. W. Gear. Numerical solution of ordinary differential equations: Is there anything left to do? SIAM Review, 23 (1):10–24, January 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gear:1985:BRD

[Gea85]  C. W. Gear.  Book re-

view: Discretization Methods for Stable Initial Value Problems (Eckart Gekeler). SIAM Review, 27(2):269, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Geer:1978:EFA

[Gee78] J. Geer. Electrostatic field about a thin disk. SIAM Review, 20(3):625, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Geffen:1975:VFN

[Gef75] N. Geffen. Variational formulation for nonlinear-wave calculations. SIAM Review, 17 (2):378–379, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Geist:1989:BRF

[Gei89]  Bruce Geist.  Book re-

view: Functional Analysis and Boundary-Value Problems: an Introductory Treatment (B. Dayanand Reddy). SIAM Review, 31(1):136– 137, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Geller:1976:IS

[Gel76] Murray Geller. An infinite sum. SIAM Review, 18 (1):117, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Geldenhuys:1979:DAI

[Gel79] G. Geldenhuys. Design of an asymmetric ideal cascade for isotope enrichment. SIAM Review, 21(3):390– 394, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gentleman:1976:BRS

[Gen76] W. Morven Gentleman. Book review: Solving Least Squares Problems (Charles L. Lawson and Richard J. Hanson). SIAM Review, 18(3): 518–520, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gentle:1986:BRS

[Gen86] James E. Gentle. Book review: Statistical Computation (J. H. Maindonald). SIAM Review, 28(2):257– 259, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Genz:1996:BRL

[Gen96] Alan Genz. Book review: Lattice Methods for Multiple Integration (I. H. Sloan and S. Joe). SIAM Review, 38 (1):165–166, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Geoffrion:1967:IPI

[Geo67] Arthur M. Geoffrion. Integer programming by implicit enumeration and Balas’ method. SIAM Review, 9(2): 178–190, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Geoffrion:1971:DNP

[Geo71] A. M. Geoffrion. Duality in nonlinear programming: a simplified applicationsoriented development. SIAM Review, 13(1):1–37, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

George:1976:MPA

[Geo76] J. H. George. Mathematical problems arising in insitu retorting of oil shale. SIAM Review, 18(4):804, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Georghion:1984:LDE

[Geo84a] C. Georghion. A linear differential equation. SIAM Review, 26(2):282–283, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Georghiou:1984:LDE

[Geo84b] C. Georghiou. A lineardifferential equation. SIAM Review, 26(2):282–283, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

George:1995:PLD

[Geo95] E. O. George. A property of the logistic distribution. SIAM Review, 37(4): 608–609, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gerst:1962:BGF

[Ger62] Irving Gerst. The bivariate generating function and two problems in discrete stochastic processes. SIAM Review, 4(2):105–114, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gerlach:1994:ACN

[Ger94] Jürgen Gerlach. Accelerated convergence in Newton’s method. SIAM Review, 36 (2):272–276, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/25467.htm.

Gerlach:1996:BRM

[Ger96] Jürgen Gerlach. Book review: Mathematical Analysis in Engineering (Chiang C. Mei). SIAM Review, 38 (2):342, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Getoor:1996:BRD

[Get96] Ronald Getoor. Book review: Dirichlet Forms and Symmetric Markov Processes (Masatoshi Fukushima, Yoichi Oshima, and Masayoshi Takeda). SIAM Review, 38 (1):166–167, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Geveci:1982:SSD

[Gev82] T. Geveci. The significance of the stability of difference schemes in different l^{p}- spaces. SIAM Review, 24(4): 413–426, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goodman:1976:EGD

[GF76] R. Goodman and A. Feldstein. Effect of guard digits and normalization options on floating point multiplication. SIAM Review, 18 (4):806, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golberg:1978:PMC

[GF78] M. A. Golberg and J. A.

Fromme. Projection methods for a class of operator equations with applications to some singular integral equations. SIAM Review, 20(3):625, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1986:FSS

[GF86] M. L. Glasser and H. E. Fettis. A Fourier sine series. SIAM Review, 28(2): 233, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gunderson:1968:ECS

[GG68] R. W. Gunderson and J. H. George. On the equivalence of certain stability properties. SIAM Review, 10(1): 81–84, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldstein:1973:HTB

[GG73] A. J. Goldstein and R. L. Graham. A Hadamard-type bound on the coefficients of a determinant of polynomials. SIAM Review, 15(3): 657–658, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Garcia:1980:RBS

[GG80] C. B. García and F. J. Gould. Relations between several path following algorithms and local and global Newton methods. SIAM Review, 22 (3):263–274, July 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1988:DI

[GG88] M. L. Glasser and N. Grossman. A definite integral. SIAM Review, 30(1):130– 132, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1989:SI

[GG89] M. L. Glasser and C. C.

Grosjean. An N-space integral. SIAM Review, 31 (2):329, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1991:BRP

[GG91] Andrew V. Goldberg and Dan Gusfield. Book review: Potokovije Algoritmij (Flow Algorithms) (G. M. Adel’son-Vel’ski, E. A. Dinits, and A. V. Karzanov). SIAM Review, 33(2):306–314, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibbons:1995:SMG

[GG95] Robert D. Gibbons and Edward J. Gilroy. Statistical methods for groundwater monitoring. SIAM Review, 37(4):628–??, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gander:1999:DNM

[GG99] Walter Gander and Dominik Gruntz. Derivation of numerical methods using computer algebra. SIAM Review, 41 (3):577–593, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35093.

Goldberg:1969:BRS

[GGG69] Michael Goldberg, R. K. Guy, and R. K. Guy. Book review: Stability oF Polyhedra (J. H. Conway and R. K. Guy). SIAM Review, 11 (1):78–82, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1996:HAC

[GGK96] R. B. Guenther, J. A. Gottsch, and D. B. Kramer. The Herglotz algorithm for constructing canonical transformations. SIAM Review, 38 (2):287–293, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/p287.htm.

Gauvin:2022:RRM

[GGK^{+}22] Laetitia Gauvin, Mathieu Génois, Márton Karsai, Mikko Kivelä, Taro Takaguchi, Eugenio Valdano, and Christian L. Vestergaard. Randomized reference models for temporal networks. SIAM Review, 64(4):763–830, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1242252.

Glasser:1988:NI

[GGR^{+}88] M. L. Glasser, J. Gevirtz, O. G. Ruehr, C. C. Grosjean, and W. B. Jordan. A nonnegative integral. SIAM Review, 30(3):506–507, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gonnet:2013:RPA

[GGT13] Pedro Gonnet, Stefan Güttel, and Lloyd N. Trefethen. Robust Padé approximation via SVD. SIAM Review, 55(1): 101–117, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldstein:1986:BRI

[GH86] J. A. Goldstein and Kosaki H. Book review: An Integral Identity (M. L. Glasser). SIAM Review, 28(4):570, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gravesen:2001:GSC

[GH01] Jens Gravesen and Christian Henriksen. The geometry of the scroll compressor. SIAM Review, 43 (1):113–126, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36212.

Grindrod:2013:MID

[GH13] Peter Grindrod and Desmond J. Higham. A matrix iteration for dynamic network summaries. SIAM Review, 55(1): 118–128, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guasoni:2022:MRC

[GH22] Paolo Guasoni and Yu-Jui Huang. Minimizing the repayment cost of Federal student loans. SIAM Review, 64(3):689–709, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1505840.

Ghandehari:1989:BRC

[Gha89] Mostafa Ghandehari. Book review: Codes and Cryptography (Dominic Welsh). SIAM Review, 31(4):694– 695, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ghandehari:1994:BRF

[Gha94] Mostafa Ghandehari. Book review: Foundations of Coding (J. Adamek). SIAM Review, 36(1):125–126, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gatica:1997:BFE

[GHC97] Gabriel N. Gatica, George C.

Hsiao, and Carsten Carstensen. Boundary-field equation methods for a class on nonlinear problems. SIAM Review, 39 (3):529–??, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ghil:1976:ARI

[Ghi76] M. Ghil. Analytical results on initialization in numerical weather prediction. SIAM Review, 18(4):804– 805, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ghosh:1987:BRS

[Gho87] B. K. Ghosh. Book review: Sequential Analysis: Tests and Confidence Intervals (David Siegmund). SIAM Review, 29(2):315– 318, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gaudioso:2022:DPN

[GHU22] Manlio Gaudioso and Jean-Baptiste Hiriart-Urruty. Deforming ||.||_{1} into ||.||_{∞} via polyhedral norms: a pedestrian approach. SIAM Review, 64(3):713–727, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1391481.

Golub:1976:CVO

[GHW76] G. H. Golub, M. Heath, and G. Wahba. Cross validation and optimum ridge regression. SIAM Review, 18 (4):806, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glassey:1997:CPK

[GI97] Robert T. Glassey and Reinhard Illner. The Cauchy problem in kinetic theory. SIAM Review, 39(3):532– ??, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Giammo:1963:PSS

[Gia63] T. P. Giammo. On the probability of success in a sudden death search with intermittent moves confined to a finite area. SIAM Review, 5 (1):41–51, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibes:1973:BRL

[Gib73] A. G. Gibes. Book review: Laplace Transforms of Airy Functions (P. Smith). SIAM Review, 15(4):796– 798, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibson:1977:BRM

[Gib77] P. M. Gibson. Book review: Monotone Submatrices (A. Berman). SIAM Review, 19(3):568–569, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibbon:1985:BRN

[Gib85] J. D. Gibbon. Book review:

Nonlinear Waves (Lokenath Debnath, ed.). SIAM Review, 27(3):465–466, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibson:1987:CMD

[Gib87a]  P. M. Gibson.  A con-

jectured minimum degree determinantal polynomial. SIAM Review, 29(4):634– 635, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibson:1987:BRC

[Gib87b] Peter M. Gibson. Book review: A Conjectured Minimum Degree Determinantal Polynomial (P. R. Vein).

SIAM Review, 29(4):634– 635, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibes:1990:BRN

[Gib90] Alan Gibes. Book review: Numerical Evaluation of a Slowly Convergent Series (R. E. Shafer). SIAM Review, 32 (3):481–483, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibbon:1995:BRI

[Gib95] J. D. Gibbon. Book review: Introduction to Multidimensional Integrable Equations (B. G. Konopelchenko). SIAM Review, 37(2):257– 258, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gibson:1997:BRM

[Gib97] Peter M. Gibson. Book review: Matrices of Sign-Solvable Linear Systems, by Richard A. Brualdi and Bryan L. Shader. SIAM Review, 39(1):145–146, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Giese:1963:IEO

[Gie63] J. H. Giese. An integral equation occurring in a problem of subsidence. SIAM Review, 5(1):1–6, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1962:GIC

[Gil62a]  E. N. Gilbert.  Games

of identification or convergence. SIAM Review, 4(1): 16–24, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilliland:1962:NNC

[Gil62b] Dennis C. Gilliland. On the non-negativity of the coefficients of a certain series expansion. SIAM Review, 4(4): 395–396, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1965:LSW

[Gil65a] E. N. Gilbert. Latin squares which contain no repeated digrams. SIAM Review, 7(2): 189–198, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1965:BRD

[Gil65b] R. C. Gilbert. Book review: Differentialoperatoren der Mathematischen Physik (Gunter Hellwig). SIAM Review, 7(3):436, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gillies:1969:EI

[Gil69] A. W. Gillies. An elliptic integral. SIAM Review, 11 (4):620, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gillies:1970:CMG

[Gil70] A. W. Gillies. On the classification of matrix generalized inverses. SIAM Review, 12 (4):573–576, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1974:DM

[Gil74] E. N. Gilbert. Distortion in maps. SIAM Review, 16 (1):47–62, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1978:BRP

[Gil78] R. P. Gilbert. Book review: Partial Differential Equations in the Complex Domain (D. L. Colton). SIAM Review, 20(1):198– 200, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1982:BRE

[Gil82] R. P. Gilbert. Book review: Elliptic Systems in the Plane (W. L. Wendland). SIAM Review, 24(2):

241–244, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1990:BRM

[Gil90a] Robert P. Gilbert. Book review: Methods of Complex Analysis in Partial Differential Equations with Applications (Manfred Kracht and Erwin Kreyszig). SIAM Review, 32(1):174–175, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1990:MCA

[Gil90b] Robert P. Gilbert. Methods of complex analysis in partial differential equations with applications (Manfred Kracht and Erwin Kreyszig). SIAM Review, 32(1):174– ??, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilroy:1995:BRS

[Gil95] Edward J. Gilroy. Book review: Statistical Methods for Groundwater Monitoring (Robert D. Gibbons). SIAM Review, 37(4):628– 629, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ginzburg:1985:BRD

[Gin85] Lev Ginzburg. Book review: Demography Through Problems (Nathan Keyfitz and John A. Beekman). SIAM Review, 27(2):269– 270, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Giusti:2018:BRN

[Giu18]  Chad Giusti.  Book re-

view: Niche Hierarchy: Structure, Organization and Assembly in Natural Systems, by George Sugihara. SIAM Review, 60(2):473– 475, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Giveen:1963:TPT

[Giv63]  Samuel M. Giveen.  A

taxicab problem with timedependent arrival rates. SIAM Review, 5(2):119– 127, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Givens:1978:BRM

[Giv78a] C. Givens. Book review: A Matrix Convergence Problem (G. K. Kristiansen). SIAM Review, 20(4):857– 858, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Givens:1978:BRS

[Giv78b] C. Givens. Book review: A System of Second Order Differential Equations (L. Carlitz). SIAM Review, 20(4): 859, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1980:IS

[GJ80] M. L. Glasser and A. J. Jerri.

An infinite series. SIAM Review, 22(2):230, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosifan:1990:ABD

[GJ90] C. C. Grosifan and W. B.

Jordan. Asymptotic behavior of a definite integral. SIAM Review, 32(3): 480–481, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1997:IEG

[GJ97] C. C. Grosjean and W. B.

Jordan. An integral from electron gas theory. SIAM Review, 39(2):325–327, ???? 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greengard:2019:NMP

[GJ19] Leslie Greengard and Shidong Jiang. A new mixed potential representation for unsteady, incompressible flow. SIAM Review, 61(4):733– 755, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greengard:2024:NVA

[GJRW24] Leslie F. Greengard, Shidong Jiang, Manas Rachh, and Jun Wang. A new version of the adaptive fast Gauss transform for discrete and continuous sources. SIAM Review, 66(2):287–315, May 2024. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Givens:1978:MCP

[GK78] C. Givens and G. K. Kristiansen. Matrix convergence problem. SIAM Review, 20 (4):857, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1981:BRD

[GK81] M. Goldberg and J. D. E. Konhauser. Book review: Designing a Three-Edged Reamer (B. C. Rennie). SIAM Review, 23 (4):525, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Georghiou:1986:BRI

[GK86a] C. Georghiou and I. P. E. Kinnmark. Book review: An Infinite Sum (M. L. Glasser). SIAM Review, 28 (2):242, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldstein:1986:II

[GK86b] J. A. Goldstein and H. Kosaki.

An integral identity. SIAM Review, 28(4):570, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1991:ALC

[GK91] M. L. Glasser and J. Koplowitz. Approximation for the location of the center of a disk. SIAM Review, 33 (4):657, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grenander:1995:GPT

[GK95] U. Grenander and Daniel Keenan. General pattern theory: a mathematical study of regular structures. SIAM Review, 37(2): 258–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grenander:1998:EPT

[GK98] U. Grenander and Daniel Keenan. Elements of pattern theory. SIAM Review, 40(1): 155–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gander:2012:CFT

[GK12] Martin J. Gander and Felix Kwok. Chladni figures and the Tacoma Bridge: Motivating PDE eigenvalue problems via vibrating plates. SIAM Review, 54(3):573– 596, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Girard:2015:PNO

[GK15] Anouck R. Girard and Pierre T. Kabamba. Proportional navigation: Optimal homing and optimal evasion. SIAM Review, 57(4): 611–624, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Georghiou:1986:IS

[GKG^{+}86] C. Georghiou, I. P. E. Kinnmark, E. P. Gray, W. B. Jordan, K. W. Lau, and O. Ruehr. An infinite sum. SIAM Review, 28 (2):242, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenleaf:2009:CDE

[GKLU09] Allan Greenleaf, Yaroslav Kurylev, Matti Lassas, and Gunther Uhlmann. Cloaking devices, electromagnetic wormholes, and transformation optics. SIAM Review, 51(1):3–33, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Green:1975:SIT

[GKP75] M. W. Green, A. J. Korsak, and M. C. Pease. Simultaneous iteration towards all roots of a complex polynomial. SIAM Review, 17 (3):567, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grohs:2020:PRU

[GKR20] Philipp Grohs, Sarah Koppensteiner, and Martin Rathmair. Phase retrieval: Uniqueness and stability. SIAM Review, 62(2):301– 350, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

George:1989:EMD

[GL89] Alan George and Joseph W. H. Liu. The evolution of the minimum degree ordering algorithm. SIAM Review, 31 (1):1–19, March 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Govinder:1998:EDE

[GL98] K. S. Govinder and P. G. L.

Leach. An elementary demonstration of the existence of sℓ(3, R) symmetry for all second-order linear ordinary differential equations. SIAM Review, 40 (4):945–946, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31683.

Greengard:2004:ANF

[GL04] Leslie Greengard and June-Yub Y. Lee. Accelerating the Nonuniform Fast Fourier Transform. SIAM Review, 46 (3):443–454, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43200.

Gablonsky:2005:MBF

[GL05] Joerg M. Gablonsky and Andrew S. I. D. Lang. Modeling basketball free throws. SIAM Review, 47(4):775– 798, December 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/33955.

Glasser:1967:DI

[Gla67] M. Lawrence Glasser. A definite integral. SIAM Review, 9(1):121, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1970:II

[Gla70] M. L. Glasser. An infinite integral. SIAM Review, 12 (1):144, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1971:EL

[Gla71] M. L. Glasser. An elliptic lntegral. SIAM Review, 13 (3):389, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1975:LI

[Gla75a] M. L. Glasser. Limit of an integral. SIAM Review, 17 (4):686, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See erratum [Gla76].

Glasser:1975:SIE

[Gla75b] M. L. Glasser. A singular integral equation. SIAM Review, 17(3):565, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1976:ELI

[Gla76] M. L. Glasser. Erratum:

Limit of an Integral. SIAM Review, 18(2):306, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Gla75a].

Glasser:1977:CII

[Gla77a] M. L. Glasser. A conjectured increasing infinite series. SIAM Review, 19(1): 148, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1977:DI

[Gla77b] M. L. Glasser. A definite integral. SIAM Review, 19 (2):329, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1978:BRW

[Gla78a] M. L. Glasser. Book review: When is the Modified Bessel Function Equal to Its Derivative? (I. Nasell). SIAM Review, 20 (4):862, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1978:DIa

[Gla78b] M. L. Glasser. A definite integral. SIAM Review, 20(3): 595–596, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1978:DIb

[Gla78c] M. L. Glasser. A definite integral. SIAM Review, 20 (4):855, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1979:SBF

[Gla79] M. L. Glasser. A sum of Bessel functions. SIAM Review, 21(4):559, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1980:ITS

[Gla80] M. L. Glasser. An infinite triple summation. SIAM Review, 22(3):363–364, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1981:ITS

[Gla81a] M. L. Glasser. An infinite triple summation. SIAM Review, 23(3):393–394, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1981:PSE

[Gla81b] M. L. Glasser. Power series of an elliptic function. SIAM Review, 23(3):391– 392, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1981:PIL

[Gla81c] M. L. Glasser. A property of an inverse Laplace transform. SIAM Review, 23 (2):256, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1982:BFI

[Gla82a] M. L. Glasser. Bessel function integrals. SIAM Review, 24(2):224, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1982:PSE

[Gla82b] M. L. Glasser. Power series of an elliptic function. SIAM Review, 24(4): 479, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1983:SLP

[Gla83] M. L. Glasser. A sum of Legendre polynomials. SIAM Review, 25(4):571, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1984:ADI

[Gla84a] M. L. Glasser. Another definite integral. SIAM Review, 26(2):271, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1984:RS

[Gla84b] M. L. Glasser. A rectangular source. SIAM Review, 26 (2):276–278, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1984:VDI

[Gla84c] M. L. Glasser. A vanishing definite integral. SIAM Review, 26(1):118, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1985:BRA

[Gla85a] M. L. Glasser. Book review:

Another Definite Integral (M. L. Glasser). SIAM Review, 27(2):254, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1985:IS

[Gla85b] M. L. Glasser. An infinite sum. SIAM Review, 27 (2):250, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1985:II

[Gla85c] M. L. Glasser. An integral identity. SIAM Review, 27 (4):573, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1985:VDI

[Gla85d] M. L. Glasser. A vanishing definite integral. SIAM Review, 27(1):85, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glass:1986:BRM

[Gla86a] Leon Glass. Book review: Modeling Dynamic Phenomena in Molecular and Cellular Biology (Lee A. Segel). SIAM Review, 28(2):259– 260, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1986:BRH

[Gla86b] M. L. Glasser. Book review: A Hypergeometric Function Integral (S. Lj. D Iuanović). SIAM Review, 28(4):572– 573, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1986:BRI

[Gla86c] M. L. Glasser. Book review: Infinite Sums of Bessel Functions (D. M. Petković). SIAM Review, 28(3):402– 403, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1987:BRN

[Gla87a] M. L. Glasser. Book review: A Normalization Constant (Jerold R. Bottiger and David K. Cohoon). SIAM Review, 29(2):302– 303, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1987:DI

[Gla87b] M. L. Glasser. A definite integral. SIAM Review, 29(1): 122–123, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1987:DIB

[Gla87c] M. L. Glasser. A definite integral of Bessel functions. SIAM Review, 29 (4):632, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1988:BRN

[Gla88a] M. L. Glasser. Book review: A Nonnegative Integral (J. Gevirtz and O. G. Ruehr). SIAM Review, 30(3): 506–507, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1988:I

[Gla88b] M. L. Glasser. An identity. SIAM Review, 30(2): 317–318, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1988:LT

[Gla88c] M. L. Glasser. A Laplace transform. SIAM Review, 30 (3):503, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1988:ZLM

[Gla88d] M. L. Glasser. Zero of least modulus. SIAM Review, 30 (1):124, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1989:ABD

[Gla89a] M. L. Glasser. Asymptotic behavior of a definite integral. SIAM Review, 31 (3):495, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1989:BRS

[Gla89b] M. L. Glasser. Book review: An n-Space Integral (Seung-Jin Bang). SIAM Review, 31 (2):328–329, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1989:PVQ

[Gla89c] M. L. Glasser. On the principal value of a quadruple integral. SIAM Review, 31(1): 119–120, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1991:ECH

[Gla91] M. L. Glasser. Exponentials of certain Hilbert space operators. SIAM Review, 33 (3):472, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gladwell:1993:URR

[Gla93a] G. M. L. Gladwell. A unique real root. SIAM Review, 35 (4):643, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1993:BRT

[Gla93b] M. L. Glasser. Book review: Two Integrals Arising in Generalized Inverse Gaussian Model and Heat Conduction Problems (M. A. Chaudry AND S. M. Zubair). SIAM Review, 35 (3):503–504, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1993:IIF

[Gla93c] M. L. Glasser. An infinite integral of four Bessel functions. SIAM Review, 35 (2):299, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1993:TDI

[Gla93d] M. L. Glasser. Two definite integrals arising in light transmission through a crystal. SIAM Review, 35(1): 136, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glass:1994:BRI

[Gla94a] Edward Glass. Book review:

An Introduction to Mathematical Cosmology (J. N. Islam). SIAM Review, 36 (3):507–508, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1994:NTS

[Gla94b] M. L. Glasser. A nonharmonic trigonometric series. SIAM Review, 36(3): 489, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1994:TDI

[Gla94c] M. L. Glasser. Two definite integrals arising in light transmission through a Crystal. SIAM Review, 36(1): 111–112, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gladwell:1995:BRN

[Gla95a] I. Gladwell. Book review: Numerical Solution of Ordinary Differential Equations (L. F. Shampine). SIAM Review, 37(1):122– 123, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1995:BRD

[Gla95b] M. L. Glasser. Book review: A Definite Integral (Norbert Ortner). SIAM Review, 37 (3):450–451, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1995:ET

[Gla95c] M. L. Glasser. Exciton transport. SIAM Review, 37(1): 98–99, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1995:FAO

[Gla95d] M. L. Glasser. A function arising in one-dimensional percolation. SIAM Review, 37(4):608, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1995:IAC

[Gla95e]  M. L. Glasser.  An in-

tegral arising in computing the energy of crystals. SIAM Review, 37(2):235– 236, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1995:NTS

[Gla95f] M. L. Glasser. A nonharmonic trigonometric series. SIAM Review, 37(3): 443–445, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1996:IEG

[Gla96a] M. L. Glasser. An integral from electron gas theory. SIAM Review, 38(2):

313, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1996:PI

[Gla96b] M. L. Glasser. Parlon integrals. SIAM Review, 38 (4):669, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1997:DIP

[Gla97a] M. L. Glasser. A definite integral in probability theory. SIAM Review, 39(3): 522–523, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1997:SIS

[Gla97b] M. L. Glasser. Solutions: An integral from a Stefan problem (problem 96-4) (Glenn Ledder). SIAM Review, 39 (1):141, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1998:DIP

[Gla98] M. L. Glasser. A definite integral in probability theory. SIAM Review, 40 (4):983–984, November 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gleason:1981:FHP

[Gle81] A. Gleason. Factorization of a homogeneous polynomial. SIAM Review, 23 (2):255, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gleason:1983:BRF

[Gle83] A. M. Gleason. Book review:

Factorization of a Homogeneous Polynomial (A. M. Gleason). SIAM Review, 25 (2):268–269, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gleason:1985:FHP

[Gle85] A. M. Gleason. Factorization of a homogeneous polynomial. SIAM Review, 27 (1):80–81, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gleich:2015:PBW

[Gle15] David F. Gleich. PageRank beyond the Web. SIAM Review, 57(3):321–363, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glimm:1991:BRP

[Gli91a] James Glimm. Book review:

The Physics of Fluid Turbulence (W. D. McComb). SIAM Review, 33(4):673– 674, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glimm:1991:NSP

[Gli91b] James Glimm. Nonlinear and stochastic phenomena: The grand challenge for partial differential equations. SIAM Review, 33(4):626– 643, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Georghiou:1988:I

[GLL^{+}88] C. Georghiou, C. A. Libis, O. P. Lossers, R. E. Shafer, and J. A. Wilson. An identity. SIAM Review, 30(2): 319, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenbaum:2020:FOP

[GLO20] Anne Greenbaum, Ren-Cang Li, and Michael L. Overton. First-order perturbation theory for eigenvalues and eigenvectors. SIAM Review, 62(2): 463–482, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Getty:2018:LLQ

[GLY^{+}18] Daniel Getty, Hao Li, Masayuki Yano, Charles Gao, and A. E. Hosoi. Luck and the law: Quantifying chance in fantasy sports and other contests. SIAM Review, 60 (4):869–887, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenberg:1966:AEC

[GM66] M. D. Greenberg and John P.

Moran. Approximate evaluation of certain functions arising in unsteady airfoil theory. SIAM Review, 8(1): 103–105, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gallie:1976:CMP

[GM76a] T. M. Gallie and M. N.

Morrow. Calculation of Moore–Penrose pseudoinverse. SIAM Review, 18 (4):804, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golberg:1976:OSS

[GM76b] M. A. Golberg and J. Moore.

Optimal size of a substitute teacher pool. SIAM Review, 18(1):114–116, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golbergt:1976:OSS

[GM76c] Michael A. Golbergt and John Moore. The optimal size of a substitute teacher pool. SIAM Review, 18(1): 114–116, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Garcia:1981:ASW

[GM81a] C. Garcia and A. A. Minzoni. An asymptotic solution for the wave equation in a timedependent domain. SIAM Review, 23(1):1–9, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

GarciaR:1981:ASW

[GM81b] Catherine Garcia R. and A. A. Minzoni. An asymptotic solution for the wave equation in a time-dependent domain. SIAM Review, 23 (1):1–9, January 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Graves-Morris:1985:BRP

[GM85] P. R. Graves-Morris. Book review: Padé Approximants for Operators: Theory and Applications (Annie Cuyt). SIAM Review, 27(3):467, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goguen:1987:BRF

[GM87] J. A. Goguen and J. Meseguer. Book review: Fundamentals of Algebraic Specification 1: Equations and Initial Semantics (H. Ehrig and B. Mahr). SIAM Review, 29(2):

318–322, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Graves-Morris:1992:DI

[GM92] P. R. Graves-Morris. A determinant identity. SIAM Review, 34(4):649–650, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ghoussoub:1995:DPM

[GM95] N. Ghoussoub and Jean Mawhin. Duality and perturbation methods in critical point theory. SIAM Review, 37(2):265–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grotzes:1996:MLI

[GM96] George Grotzes and Kenneth R. Meyer. Math into LATEX: An introduction to LATEX and AMS-LATEX. SIAM Review, 38(3):535– ??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Graybill:1966:NCG

[GMP66] F. A. Graybill, C. D. Meyer, and R. J. Painter. Note on the computation of the generalized inverse of a matrix. SIAM Review, 8(4): 522–524, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gebremedhin:2005:WCY

[GMP05] Assefaw Hadish Gebremedhin, Fredrik Manne, and Alex Pothen. What color is your Jacobian? graph coloring for computing derivatives. SIAM Review, 47 (4):629–705, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44471.

Gill:2005:SSA

[GMS05] Philip E. Gill, Walter Murray, and Michael A. Saunders. SNOPT: An SQP algorithm for large-scale constrained optimization. SIAM Review, 47(1):99– 131, March 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44609. Reprinted

from SIAM Journal Optimization, vol 12, pg 979– 1006, 2002.

Gill:1991:ICM

[GMSW91] P. E. Gill, W. Murray, M. A. Saunders, and M. H. Wright. Inertia-controlling methods for general quadratic programming. SIAM Review, 33 (1):1–36, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1978:WMB

[GN78] M. L. Glasser and I. Nasell.

When is modified Bessel function equal to its derivative. SIAM Review, 20(4): 862, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grebenkov:2013:GSL

[GN13] D. S. Grebenkov and B.-T.

Nguyen. Geometrical structure of Laplacian eigenfunctions. SIAM Review, 55(4): 601–667, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gandy:2022:APF

[GN22] Demi L. Gandy and Martin R. Nelson. Analyzing pattern formation in the Gray– Scott model: an XPPAUT tutorial. SIAM Review, 64 (3):728–747, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1402868.

Grosjean:1995:ISD

[GNP95] C. C. Grosjean, I. Nemes, and P. Paule. An infinite sum from a diffusion problem. SIAM Review, 37(1): 105–107, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guo:2020:CLS

[GNZ20] Ling Guo, Akil Narayan, and Tao Zhou. Constructing least-squares polynomial approximations. SIAM Review, 62(2):483–508, ???? 2020. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gray:1967:LFD

[GO67] H. L. Gray and P. L. Odell. On least favorable density functions. SIAM Review, 9 (4):715–720, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golub:1989:SHC

[GO89] Gene H. Golub and Dianne P. O’Leary. Some history of the conjugate gradient and Lanczos algorithms: 1948–1976. SIAM Review, 31(1):50–102, March 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        2030847.

Gottwald:2009:BDI

[GO09] Georg A. Gottwald and Marcel Oliver. Boltzmann’s Dilemma: An introduction to statistical mechanics via the Kac ring. SIAM Review, 51 (3):613–635, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gobel:1970:BRC

[Göb70] F. Göbel. Book review:

A Combinatorial Problem (Melda Hayes). SIAM Review, 12(1):151–153, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Godsil:1987:BRG

[God87] Chris Godsil. Book review:

Graphical Evolution. an Introduction to the Theory of Random Graphs (Edgar M. Palmer). SIAM Review, 29 (4):651–652, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goffin:1978:SOR

[Gof78] J. L. Goffin. Subgradient optimization and relaxation method. SIAM Review, 20 (3):625, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goguen:1985:BRF

[Gog85] J. A. Goguen. Book review: Fuzzy Sets and Systems: Theory and Applications (Didier Dubois and Henri Prade). SIAM Review, 27(2):270–274, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goka:1970:OBS

[Gok70] T. Goka. An operator on binary sequences. SIAM Review, 12(2):264–266, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golomb:1961:PCS

[Gol61] Solomon W. Golomb. Permutations by cutting and shuffling. SIAM Review, 3(4): 293–297, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldenberg:1962:EIL

[Gol62] H. Goldenberg. The evaluation of inverse Laplace transforms without the aid of contour integration. SIAM Review, 4(2):94–104, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1965:BRT

[Gol65a] Joshua N. Goldberg. Book review: The Theory of Space Time and Gravitation, Second Edition (V. Fock). SIAM Review, 7(1): 153–154, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1965:BRC

[Gol65b]  R. Goldberg.  Book re-

view: Counterexamples in Analysis (Bernard R. Gelbaum and John M. H. Olmsted). SIAM Review, 7(2): 307–308, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1966:BDP

[Gol66a] J. L. Goldberg. Bounds on the derivatives of positive functions. SIAM Review, 8 (3):343–345, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1966:BRI

[Gol66b] J. N. Goldberg. Book review:

Introduction to the Theory of Relativity and the Principles of Modern Physics (Huseyin Yilmaz). SIAM Review, 8(1): 123–125, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1969:BRM

[Gol69a] Michael Goldberg. Book review: Moving Furniture Through a Hallway (Leo Moser). SIAM Review, 11 (1):75–78, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1969:TVP

[Gol69b] Michael Goldberg. Three volume problems. SIAM Review, 11(3):396–397, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1971:BRS

[Gol71] Richard R. Goldberg. Book review: Studies in Hebrew Astronomy and Mathematics (Solomon Gandz). SIAM Review, 13(4):584, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1973:BRF

[Gol73a] Richard R. Goldberg. Book review: Fourier Series and Integrals (H. Dym and H. P. McKean). SIAM Review, 15 (3):679–680, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golub:1973:SMM

[Gol73b]  Gene H. Golub.  Some

modified matrix eigenvalue problems. SIAM Review, 15(2):318–334, April 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

                                        //www.jstor.org/stable/

2028604. Cited in Åke Björck’s bibliography on least squares, which is available at

                                        ftp://math.liu.se/pub/references.

Reprinted in [CGO07].

Golberg:1975:ML

[Gol75a]  M. Golberg.  A matrix

limit. SIAM Review, 17 (3):567, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1975:ML

[Gol75b]  M. Goldberg.  Matrix

limit. SIAM Review, 17 (3):567, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldstein:1975:FDP

[Gol75c] M. J. Goldstein. Further decomposition of pseudoinverse and eigensystem of a Hermitian persymmetric matrix. SIAM Review, 17(2): 379, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golberg:1976:SNN

[Gol76a] M. A. Golberg. Some new numerical methods for solving convolution integral equations. SIAM Review, 18 (4):805, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldstein:1976:FCF

[Gol76b] M. J. Goldstein. Further comparison of floating point summation methods. SIAM Review, 18(4):805, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golland:1976:MSI

[Gol76c] R. W. Golland. Method for solving ill-conditioned linear equation problem. SIAM Review, 18(4):805– 806, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldstein:1977:OT

[Gol77] A. A. Goldstein. Optimal temperament. SIAM Review, 19(3):554–562, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golubitsky:1978:ICT

[Gol78] Martin Golubitsky. An introduction to catastrophe theory and its applications. SIAM Review, 20(2):352– 387, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1981:MP

[Gol81] M. Goldberg. A minimization problem. SIAM Review, 23(4):524, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldstein:1993:BRA

[Gol93] Jerome A. Goldstein. Book review: Attractors for Semigroups and Evolution Equations (Olga Ladyzhenskaya).

SIAM Review, 35(4):664– 666, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gomory:1967:VPD

[Gom67] R. E. Gomory. VI. postdoctoral experience. SIAM Review, 9(2):388–410, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gonzaga:1992:PFM

[Gon92] Clovis C. Gonzaga. Pathfollowing methods for linear programming. SIAM Review, 34(2):167–224, June 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gonzalez:2006:OPI

[Gon06] Luis González. Orthogonal projections of the identity: Spectral analysis and applications to approximate inverse preconditioning. SIAM Review, 48(1):66–75, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1959:SLR

[Goo59] R. A. Good. Systems of linear relations. SIAM Review, 1(1):1–31, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1965:FM

[Goo65a]  R. A. Good.  f-finger

morra. SIAM Review, 7(1): 81–87, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goodwin:1965:IEN

[Goo65b]  Bruce E. Goodwin.  In-

tegral equations with nonlinear eigenvalue parameters. SIAM Review, 7(3): 388–394, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gootkind:1968:BRI

[Goo68] David Gootkind. Book review: An Integral Equation (C. J. Bouwkamp). SIAM Review, 10(4):454– 455, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1972:BRM

[Goo72] I. J. Good. Book review: Multitype Branching Processes. Theory and Applications (Charles J. Mode). SIAM Review, 14(3):512– 513, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1982:BRN

[Goo82] I. J. Good. Book review: Nonparametric Probability Density Estimation (Richard A. Tapia and James R. Thompson). SIAM Review, 24(3):364–365, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1984:BRF

[Goo84] I. J. Good. Book review:

The Fractal Geometry of Nature, by Benoît B. Mandelbrot. SIAM Review, 26(1): 131–132, January 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/siread/v26/
i1/p131/s1;   http://www.
jstor.org/stable/2029696.

Good:1986:BRS

[Goo86]  I. J. Good.  Book re-

view: Statistical Inference Based on Ranks (Thomas P. Hettmansperger). SIAM Review, 28(4):587–588, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1990:BRF

[Goo90a] I. J. Good. Book review:

Fractals Everywhere (Michael Barnsley). SIAM Review, 32 (2):317–318, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1990:FEM

[Goo90b] I. J. Good. Fractals everywhere (Michael Barnsley). SIAM Review, 32(2): 317–??, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Good:1991:BRS

[Goo91a] I. J. Good. Book review: The Science of Fractal Images (Heinz-Otto Peitgen and Dietmar Saupe, eds.). SIAM Review, 33(3):496– 499, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goodwin:1991:BRS

[Goo91b] G. C. Goodwin. Book review: Stochastic Optimal Control Theory with Application in Self-Tuning Control (K. J. Hunt). SIAM Review, 33(1): 142–144, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gorn:1959:BRP

[Gor59]  Saul Gorn.  Book re-

view: Principles and Techniques of Applied Mathematics (Bernard Friedman). SIAM Review, 1(1):77–78, ???? 1959. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gorn:1963:CIS

[Gor63]  Saul Gorn.  The com-

puter and information sciences: a new basic discipline. SIAM Review, 5(2): 150–155, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gordon:1975:RNA

[Gor75] W. J. Gordon. Role of numerical analysis in automotive design and production. SIAM Review, 17(2):379, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goss:1963:BRP

[Gos63]  Robert N. Goss.  Book

review: Plastic Flow and Fracture in Solids (T. Y. Thomas). SIAM Review, 5 (3):296–297, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goss:1964:BRS

[Gos64]  R. N. Goss.  Book re-

view: Some Basic Problems of the Mathematical Theory of Elasticity (N. I. Muskhelishvili). SIAM Review, 6 (3):320–321, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goss:1979:BRI

[Gos79] R. N. Goss. Book review: Information Theory with Applications (Silviu Guiasu). SIAM Review, 21(4):579– 580, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goss:1983:BRD

[Gos83a] R. N. Goss. Book review: Digital Computer Treatment of Partial Differential Equations (V. Vemuri and W. Karplus). SIAM Review, 25 (1):131–132, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goss:1983:BRN

[Gos83b] R. N. Goss. Book review: Numerical Solution of Partial Differential Equations in Science and Engineering (Leon Lapidus and George F. Pinder). SIAM Review, 25(4): 581–582, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goss:1984:BRA

[Gos84] R. N. Goss. Book review: Algebraic Automata Theory (W. M. L. Holcombe). SIAM Review, 26(2):302, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gott:1960:TD

[Got60] Euyen Gott. A theorem on determinants. SIAM Review, 2(4):288–291, ???? 1960.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gottlieb:1978:TNS

[Got78] D. Gottlieb. Topics in numerical spectral methods. SIAM Review, 20(3):625, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gould:1968:BRP

[Gou68] H. W. Gould. Book review: “Up-Down” Permutations (D. J. Newman and W. Weissblum). SIAM Review, 10(2):225–226, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gould:2010:MHG

[Gou10] Ronald Gould. Mathletics: How gamblers, managers, and sports enthusiasts use mathematics in baseball, basketball, and football. SIAM Review, 52(4): 776–777, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goulet:2016:MSP

[Gou16] D. Goulet. Modeling, simulating, and parameter fitting of biochemical kinetic experiments. SIAM Review, 58(2): 331–353, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glowinski:1979:NMF

[GP79] R. Glowinski and O. Pironneau. Numerical methods for the first biharmonic equation and for the twodimensional Stokes problem. SIAM Review, 21(2): 167–212, April 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Georghiou:1988:BRA

[GP88] C. Georghiou and A. N. Philippou. Book review: Asymptotic Reliability of Consecutive k-out-of-n: F Systems (A. N. Philippou). SIAM Review, 30(1):133– 134, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1992:MTD

[GP92] M. Goldberg and D. L. Powers. A machine-tool design. SIAM Review, 34(1): 114–117, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Giorgi:2002:BSD

[GP02] T. Giorgi and D. Phillips. The breakdown of superconductivity due to strong fields for the Ginzburg–Landau model. SIAM Review, 44(2): 237–256, June 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/39951.

Georghiou:1988:ARC

[GPP88] C. Georghiou, A. N. Philippou, and S. Papastauridis. Asymptotic reliability of consecutive K-out-of-N-F systems. SIAM Review, 30(1): 133–134, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gallivan:1990:PAD

[GPS90] K. A. Gallivan, R. J. Plemmons, and A. H. Sameh. Parallel algorithms for dense linear algebra computations. SIAM Review, 32(1):54– 135, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gajic:1996:LME

[GQJ96] Zoran Gajic, Muhammad Tahir Javed Qureshi, and Lucas Jodar. Lyapunov matrix equation in system stability and control. SIAM Review, 38(4):691– ??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasser:1981:BRI

[GR81] M. L. Glasser and O. G.

Ruehr. Book review: An Infinite Triple Summation (M. L. Glasser). SIAM Review, 23 (3):393–394, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

George:1985:MGF

[GR85] E. O. George and C. C. Rousseau. A moment generating function. SIAM Review, 27(4):575, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gevirtz:1987:NI

[GR87] J. Gevirtz and O. G. Ruehr. A nonnegative integral. SIAM Review, 29 (3):470, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gelbrich:1995:BRN

[GR95] Matthias Gelbrich and Werner Römisch. Book review: Numerical Solution of Stochastic Differential Equations (Peter E. Kloeden and Eckhard Platen). SIAM Review, 37 (2):272–275, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Graham:1963:RP

[Gra63] Ron L. Graham. A resistance problem. SIAM Review, 5 (4):370, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gram:1966:BRF

[Gra66a]  Chr. Gram.  Book re-

view: Funktional Analysis and Numerische Mathematik (Lothar Collatz). SIAM Review, 8(3):396–397, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grau:1966:BRR

[Gra66b] A. A. Grau. Book review:

Rounding Errors in Algebraic Processes (J. H. Wilkinson). SIAM Review, 8(3): 397–398, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gragg:1967:BRN

[Gra67] W. B. Gragg. Book review: Numerical Solution of Initial Value Problems (F. Ceschino and J. Kuntzmann). SIAM Review, 9(3): 609–610, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grau:1970:BRF

[Gra70] Albert A. Grau. Book review:

Formal Languages and Their Relation to Automata (John E. Hopcroft and Jeffrey D. Ullman). SIAM Review, 12 (2):316–317, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gragg:1972:PTR

[Gra72a] W. B. Gragg. The Padé table and its relation to certain algorithms of numerical analysis. SIAM Review, 14 (1):1–62, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Graham:1972:BRP

[Gra72b] R. L. Graham. Book review: Principles of Combinatorics (C. Berg). SIAM Review, 14 (2):344–346, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gray:1979:BRD

[Gra79] Ernest P. Gray. Book review: A Definite Integral (M. L. Glasser). SIAM Review, 21 (4):568–569, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grassman:1983:BRM

[Gra83a] W. K. Grassman. Book review: Matrix-Geometric Solutions in Stochastic Models (Marcel F. Neuts). SIAM Review, 25(1):127–128, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gravesen:1983:CTC

[Gra83b] Jens Gravesen. Catastrophe theory and caustics.

SIAM Review, 25(2):239– 247, April 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Graham:1986:BRR

[Gra86a] James H. Graham. Book review: Robotics and Artificial Intelligence (M. Brady, L. A. Gerhardt and H. F. Davidson, ed.). SIAM Review, 28 (1):102–104, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grassman:1986:BRA

[Gra86b] Winfried K. Grassman. Book review: Approximation and Weak Convergence Methods for Random Processes with Applications to Stochastic Systems Theory (Harold J. Kushner). SIAM Review, 28 (1):104, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gray:1991:STD

[Gra91] L. J. Gray. Storage tank design. SIAM Review, 33(2): 271–274, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gray:1993:BRE

[Gra93a] J. Brian Gray. Book review:

Exploratory and Multivariate Data Analysis (Michel Jambu). SIAM Review, 35 (3):521–524, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gray:1993:EMD

[Gra93b] J. Brian Gray. Exploratory and multivariate data analysis (Michel Jambu). SIAM Review, 35(3):521–??, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gray:1994:BRG

[Gra94a] Jeremy J. Gray. Book review: Galois’ Dream-Group Theory and Differential Equations (Michio Kuga). SIAM Review, 36(1):126, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gray:1994:GDG

[Gra94b] Jeremy J. Gray. Galois’ dream-group theory and differential equations (Michio Kuga). SIAM Review, 36(1): 126–??, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grandine:2000:AC

[Gra00] Thomas A. Grandine. Applications of contouring. SIAM Review, 42(2):297– 316, June 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/36403.

Grabovsky:2018:BRE

[Gra18] Yury Grabovsky. Book review: Extending the Theory of Composites to Other Areas of Science, edited and authored by Graeme W. Milton. SIAM Review, 60(2): 475–481, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grcar:2011:JNA

[Grc11] Joseph F. Grcar. John von Neumann’s analysis of Gaussian elimination and the origins of modern numerical analysis. SIAM Review, 53 (4):607–682, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i4/p607_s1.

Grcar:2018:BRE

[Grc18]  Joseph F. Grcar.  Book

review: ENIAC in Action: Making and Remaking the Modern Computer, by Thomas Haigh, Mark Priestly, and Crispin Rope. SIAM Review, 60(2):465– 470, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greville:1959:PRS

[Gre59] T. N. E. Greville. The pseudoinverse of a rectangular or singular matrix and its application to the solution of systems of linear equations. SIAM Review, 1(1): 38–43, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Green:1960:BRH

[Gre60a] T. F. Green. Book review: Handbook of Automation, Computation, and Control (E. M. Grabbe, S. Ramo, and D. E. Wooldridge). SIAM Review, 2(1):46–48, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gregory:1960:DDM

[Gre60b] Robert T. Gregory. Defective and derogatory matrices. SIAM Review, 2(2): 134–139, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greville:1960:SAP

[Gre60c] T. N. E. Greville. Some applications of the pseudoinverse of a matrix. SIAM Review, 2(1):15–22, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenspan:1962:BRD

[Gre62a] Donald Greenspan. Book review: Discrete Variable Methods in Ordinary Differential Equations (P. Henrici). SIAM Review, 4(3):261– 262, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grey:1962:UB

[Gre62b] Louis D. Grey. On an upper bound. SIAM Review, 4(1):48, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Green:1963:BRM

[Gre63a] Thomas F. Green. Book review: The Mathematical Theory of Linear Systems (B. M. Brown). SIAM Review, 5 (2):163–164, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grey:1963:BRA

[Gre63b] Louis D. Grey. Book review:

Another Satellite Communications Problem (Isidore Silberman). SIAM Review, 5 (3):287–288, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenstein:1965:BRI

[Gre65] David S. Greenstein. Book review: Interpolation and Approximation (Philip J. Davis). SIAM Review, 7(1): 151–152, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greville:1966:NGI

[Gre66] T. N. E. Greville. Note on the generalized inverse of a matrix product. SIAM Review, 8(4):518–521, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See erratum [Gre67c].

Greenberg:1967:P

[Gre67a]  H. J. Greenberg.  Pref-

ace. SIAM Review, 9(2): 291, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenberg:1967:BSQ

[Gre67b] Irwin Greenberg. The behavior of a simple queue at various times and epochs. SIAM Review, 9(2):234– 248, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greville:1967:EGI

[Gre67c] T. N. E. Greville. Erratum: Generalized Inverse of a Matrix Product. SIAM Review, 9(2):249, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Gre66].

Greville:1972:BRB

[Gre72] T. N. E. Greville. Book review: Bigradients (Alston S. Householder). SIAM Review, 14(1):171–172, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grenander:1973:CPS

[Gre73] Ulf Grenander. Computational probability and statistics. SIAM Review, 15(1): 134–192, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenstadt:1976:BRI

[Gre76]  John Greenstadt.  Book

review: Iterative Methods for Nonlinear Optimization Problems (S. L. S. Jacoby, J. S. Kowalik, and J. T. Pizzo). SIAM Review, 18(2): 318–320, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenwell:1987:BRI

[Gre87] Raymond N. Greenwell.

Book review: Introduction to Probability Theory (Kiyosi Itô). SIAM Review, 29(2): 322–323, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenspan:1988:BRN

[Gre88] Donald Greenspan. Book review: Numerical Solutions of the N-Body Problem (Andrzej Marciniak). SIAM Review, 30(1):142, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Greenwell:1989:BRE

[Gre89] Raymond N. Greenwell. Book review: Empirical Model-Building and Response Surfaces (George E. P. Box and Norman R. Draper). SIAM Review, 31(1):137– 139, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Green:1997:BRD

[Gre97] Leon W. Green. Book review: Dynamical Systems of Algebraic Origin, by Klaus Schmidt. SIAM Review, 39(1):146–148, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Griesmer:1961:BRI

[Gri61] James H. Griesmer. Book review: An Introduction to Linear Programming and the Theory of Games (S. Vajda). SIAM Review, 3(1): 77–78, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grinold:1972:SAL

[Gri72] Richard C. Grinold. Steepest ascent for large scale linear programs. SIAM Review, 14 (3):447–464, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grimson:1977:PHP

[Gri77] R. C. Grimson. The pig house problem or the detection and measurement of pairwise associations. SIAM Review, 19 (4):727–731, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Griewank:1985:SNE

[Gri85] A. Griewank. On solving nonlinear equations with simple singularities or nearly singular solutions. SIAM Review, 27(4):537–563, December 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Griewank:1988:BRN

[Gri88] Andreas Griewank. Book review: Numerical Derivatives and Nonlinear Analysis (Harriet Kagiwada, Robert Kalaba, Nima Rosakhoo, and Karl Spingarn). SIAM Review, 30(2):327–329, June 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Goldberg:1981:DTE

[GRK81] M. Goldberg, B. C. Rennie, and J. D. E. Konhauser. Designing a threeedged reamer. SIAM Review, 23(4):525, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Graham:1987:BRC

[GRL87] S. W. Graham, O. Ruehr, and O. P. Lossers. Book review: Conjectured Trigonometrical Identities (M. Henkel and J. Lacki). SIAM Review, 29(1):132–135, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Godlewski:1998:NAH

[GRL98] Edwige Godlewski, Pierre-Arnaud Raviart, and Randall J. LeVeque. Numerical approximation of hyperbolic systems of conservation laws. SIAM Review, 40(1): 160–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Graham:1987:CTI

[GRLR87] S. W. Graham, O. Ruehr, O. P. Lossers, and M. Renardy. Conjectured trigonometrical identities. SIAM Review, 29(1):132–135, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Groetsch:1987:BRC

[Gro87] C. W. Groetsch. Book review: Computational Methods for Integral Equations (L.

M. Delves and J. L. Mohamed). SIAM Review, 29 (2):323–324, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1988:GHT

[Gro88a] C. C. Grosjean. A generalized hypergeometric-type identity. SIAM Review, 30 (3):502, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grossman:1988:BRD

[Gro88b] Nathaniel Grossman. Book review: A Definite Integral from Surface Gravity Waves (M. Stiassnie). SIAM Review, 30(4):660–661, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1990:BRA

[Gro90] C. C. Grosjean. Book review: Asymptotic Behavior of a Definite Integral (M. L. Glasser). SIAM Review, 32 (3):480–481, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1991:BRTa

[Gro91a] C. C. Grosjean. Book review: Two Definite Integrals (I. Herron). SIAM Review, 33 (1):120–122, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1991:TPB

[Gro91b] C. C. Grosjean. A two-point boundary problem for Airy functions. SIAM Review, 33 (3):477–479, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1991:TTM

[Gro91c]  C. C. Grosjean.  Two

trigonometric multiple integrals. SIAM Review, 33 (1):114, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1992:FSP

[Gro92a] C. C. Grosjean. A finite sum of products of binomial coefficients. SIAM Review, 34 (4):649, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1992:TIS

[Gro92b] C. C. Grosjean. Three infinite sums. SIAM Review, 34 (2):307–308, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1993:BREb

[Gro93a] C. C. Grosjean. Book review: Existence of a Boundary-Value Problem with Prescribed Eigenvalues (Ahmed I. Zayed). SIAM Review, 35 (3):493–497, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1993:ETT

[Gro93b] C. C. Grosjean. Expected type of triangle. SIAM Review, 35(1):139–141, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1993:SIC

[Gro93c] C. C. Grosjean. A sextuple integral from a cosmological model. SIAM Review, 35(3):488, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1993:BREa

[Gro93d] Carl C. Grosjean. Book review: Expected Type of Triangle (D. J. Newman). SIAM Review, 35(1):139– 141, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosiean:1994:BRM

[Gro94a] C. C. Grosiean. Book review: Maximum Gravitational Attraction (K. M. Seymour). SIAM Review, 36 (4):660–661, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1994:E

[Gro94b]  Carl C. Grosjean.  Er-

rata. SIAM Review, 36 (1):115, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1996:EGA

[Gro96a]  C. C. Grosjean.  Ex-

treme gravitational attraction. SIAM Review, 38(3): 515–520, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1996:IBE

[Gro96b] C. C. Grosjean. Integrals of Bessel and elliptic functions from heat conduction. SIAM Review, 38(3):523– 524, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1996:NCP

[Gro96c] C. C. Grosjean. Nonsymmetric cyclic pursuit on a sphere. SIAM Review, 38(1): 156–158, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1996:OSP

[Gro96d] C. C. Grosjean. Optimal sensor placement. SIAM Review, 38(4):670–671, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1996:BRE

[Gro96e] Carl C. Grosjean. Book review: Extreme Gravitational Attraction (M. S. Klamkin). SIAM Review, 38(3):515– 520, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1996:BRIa

[Gro96f] Carl C. Grosjean. Book review: Infinite Series of Binomial Coefficients (L. A. Skinner). SIAM Review, 38(1): 155–158, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1996:BRIb

[Gro96g] Carl C. Grosjean. Book review: Integrals of Bessel and Elliptic Functions from Heat Conduction (Ping Hui). SIAM Review, 38(3):523– 524, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1996:BRO

[Gro96h] Carl C. Grosjean. Book review: Optimal Sensor Placement (Hong Zhang and Xiaohuan Wang). SIAM Review, 38(4):670–671, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1998:FIO

[Gro98a] C. C. Grosjean. A family of integrals occurring in quantum mechanics. SIAM Review, 40(2):374–381, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1998:LTD

[Gro98b] C. C. Grosjean. A Laplace transform from a diffusion problem. SIAM Review, 40 (3):723–726, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1998:MT

[Gro98c] C. C. Grosjean. A Mellin transform. SIAM Review, 40 (3):726–729, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1998:PFD

[Gro98d] C. C. Grosjean. A partition function difference related to chiral discrimination. SIAM Review, 40(2):390, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1998:PI

[Gro98e] C. C. Grosjean. Polaron integrals. SIAM Review, 40(1): 118–126, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gradshteyn:1997:TIS

[GRS97] I. S. Gradshteyn, I. M. Ryzik, and Joseph J. Shirron. Table of integrals, series, and products. SIAM Review, 39(4): 803–??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gruber:1974:BRL

[Gru74] B. Gruber. Book review: Lie Groups and Lie Algebras with Applications and Computational Methods (Johan G. F. Belinfante and Bernard Kolman). SIAM Review, 16(1): 102–104, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gruendler:1983:LDE

[Gru83] J. Gruendler. A linear differential equation. SIAM Review, 25(2):268, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gruber:1988:BRL

[Gru88] Bruno Gruber. Book review: Lie Groups and Algebras with Applications to Physics, Geometry, and Mechanics (D. H. Sattinger and O. L. Weaver). SIAM Review, 30(2):329–330, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gruendler:1993:BRN

[Gru93] Joseph Gruendler. Book review: Nonlinear Systems (P. G. Drazin). SIAM Review, 35 (4):662–663, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1989:BRI

[GRYL89] C. C. Grosjean, O. G. Ruehr, W. R. Young, and J. S. Lew. Book review: An Inverse Laplace Transform (O. G. Ruehr and W. R. Young). SIAM Review, 31 (4):685–686, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grzesik:1983:BRA

[Grz83] J. Grzesik. Book review: Asymptotic Behavior of an nfold Integral (A. M. Odlyzko, L. A. Shepp and D. Slepian). SIAM Review, 25(4):576– 577, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grzesik:1989:NRT

[Grz89] J. A. Grzesik. Von Neumann’s rejection technique reexamined. SIAM Review, 31(3):486–489, September 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Graham:1963:ARZ

[GS63] R. L. Graham and L. Shepp.

An asymptotic result on a zero of a function. SIAM Review, 5(3):274–275, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gustafson:1967:INP

[GS67] Karl Gustafson and Vincent Sigillito. Inequalities for nonlocal parabolic and higher order elliptic equations. SIAM Review, 9(3): 531–541, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gilbert:1984:RP

[GS84] E. N. Gilbert and L. A.

Shepp. A resistance problem. SIAM Review, 26(3): 429, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gerasoulis:1988:CMT

[GS88] A. Gerasoulis and R. P.

Srivastav. Conjectured maximum term of a sequence. SIAM Review, 30 (1):124, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gottlieb:1997:GPR

[GS97] David Gottlieb and Chi-Wang W. Shu. On the Gibbs phenomenon and its resolution. SIAM Review, 39 (4):644–668, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30139.

Gneiting:2004:SMS

[GS04] Tilmann Gneiting and Martin Schlather. Stochastic models that separate fractal dimension and the Hurst effect. SIAM Review, 46(2): 269–282, June 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/39438.

Gouveia:2022:GRR

[GS22] Bernardo Gouveia and Howard A. Stone. Generating resonant and repeated root solutions to ordinary differential equations using perturbation methods. SIAM Review, 64(2):485–499, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1395922.

Galkowski:2023:DHB

[GS23a] J. Galkowski and E. A. Spence. Does the Helmholtz boundary element method suffer from the pollution effect? SIAM Review, 65 (3):806–828, ???? 2023.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1474199.

Ginn:2023:CMM

[GS23b] Timothy Ginn and Lynn Schreyer. Compartment models with memory. SIAM Review, 65(3):774–805, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1437160.

Guida:2023:RMN

[GS23c] Margherita Guida and Carlo Sbordone. The reflection method for the numerical solution of linear systems. SIAM Review, 65(4):1137– 1151, November 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gray:1979:DI

[GSL79] E. P. Gray, R. E. Shafer, and O. P. Lossers. Definite integral. SIAM Review, 21(4): 568–569, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gottlieb:2001:SSP

[GST01] Sigal Gottlieb, Chi-Wang W.

Shu, and Eitan Tadmor. Strong stability-preserving high-order time discretization methods. SIAM Review, 43(1):89–112, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36757.

Glowinski:2022:BOC

[GSYY22] Roland Glowinski, Yongcun Song, Xiaoming Yuan, and Hangrui Yue. Bilinear optimal control of an advection–reaction–diffusion system. SIAM Review, 64 (2):392–421, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1389778.

Goebel:2010:PSH

[GT10] Rafal Goebel and Andrew R. Teel. Preasymptotic stability and homogeneous approximations of hybrid dynamical systems. SIAM Review, 52 (1):87–109, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gautier:2023:NPF

[GTH23] Antoine Gautier, Francesco Tudisco, and Matthias Hein. Nonlinear Perron–Frobenius theorems for nonnegative tensors. SIAM Review, 65 (2):495–536, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1557489.

Guanrong:1992:BRL

[Gua92] Chen Guanrong. Book review: Linear Controller Design: Limits of Performance (Stephen P Boyd and Craig H. Barratt). SIAM Review, 34(4):663–??, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guckheimer:1979:BRC

[Guc79] John Guckheimer. Book review: Catastrophe Theory and Its Applications (Tim Poston and Ian Stewart). SIAM Review, 21(4):572– 573, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guckenheimer:1987:BRP

[Guc87] John Guckenheimer. Book review: Probabilistic Properties of Deterministic Systems (Andrzej Lasota and Michael C. Mackey). SIAM Review, 29(2):324–326, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guderley:1975:UVS

[Gud75] Karl G. Guderley. A unified view of some methods for stiff two-point boundary value problems. SIAM Review, 17 (3):416–442, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gudder:1977:CM

[Gud77] Stanley P. Gudder. Convexity and mixtures. SIAM Review, 19(2):221–240, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See erratum [Gud78b].

Gudder:1978:C

[Gud78a]  S. P. Gudder.  Correc-

tion. SIAM Review, 20(4): 837, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gudder:1978:ECM

[Gud78b] Stanley P. Gudder. Erratum: “Convexity and mixtures” (SIAM Rev. 19 (1977), no. 2, 221–240). SIAM Review, 20 (4):837, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Gud77].

Gudder:1984:ECM

[Gud84] Stanley P. Gudder. An extension of classical measure theory. SIAM Review, 26(1): 71–89, January 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1961:PCR

[Gue61] William C. Guenther. On the probability of capturing a randomly selected point in three dimensions. SIAM Review, 3(3):247– 251, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1969:BRB

[Gue69] Ronald B. Guenther. Book review: Boundary Value Problems of Mathematical Physics, Volume 2 (Ivar Stakgold). SIAM Review, 11 (2):289, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1972:BRL

[Gue72] Ronald B. G. Guenther.

Book review: Linear Integral Equations (Ram P. Kanwal). SIAM Review, 14 (4):669, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1979:BRF

[Gue79] Ronald B. Guenther. Book review: Functional Analysis in Modern Applied Mathematics (Ruth G. Curtain and A. J. Pritchard). SIAM Review, 21(3):409– 410, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1982:BRS

[Gue82] Ronald B. Guenther. Book review: State Models of Dynamic Systems (N. H. Mc-Clamroch). SIAM Review, 24 (2):234–235, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1983:AMC

[Gue83a] R. B. Guenther. Applying mathematics — a course in mathematical modeling — Burghes, D. N., Huntley, I., Mcdonald, J. SIAM Review, 25(4):588–589, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1983:BRAb

[Gue83b] R. B. Guenther. Book review: Applied Mathematics: Principles, Techniques and Applications (James A. Cochran). SIAM Review, 25 (4):593–594, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1983:BRQ

[Gue83c] R. B. Guenther. Book review: Qualitative Analysis of Physical Problems (M. Gitterman and V. Halpern). SIAM Review, 25(1):132, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1983:BRAa

[Gue83d] Ronald B. Guenther. Book review: Applying Mathematics: a Course in Mathematical Modelling (D. N. Burghes, I. Huntley and J. McDonald). SIAM Review, 25(4): 588–589, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1984:BRD

[Gue84] R. B. Guenther. Book review:

Differential Equation Models (Martin Braun, Courtney S. Coleman and Donald A. Drew, eds.). SIAM Review, 26(2):297–298, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1985:BRI

[Gue85] Ronald B. Guenther. Book review: Introduction to Partial Differential Equations with Applications (Mark A. Pinsky). SIAM Review, 27 (4):594, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1988:BRI

[Gue88] Ronald B. Guenther. Book review: Introductory Optimization Dynamics-Optimal Control with Economics and Management Science Applications (Pierre N. V. Tu). SIAM Review, 30(1):142– 143, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1989:BRD

[Gue89] Ronald B. Guenther. Book review: Differential Geometry, Gauge Theories, and Gravity (M. Gockeler and T. Schucker). SIAM Review, 31 (2):344, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1992:BRA

[Gue92] Ronald B. Guenther. Book review: Augustin-Louis Cauchy: a Biography (Bruno Belhoste; Frank Ragland, transl.). SIAM Review, 34(3):508– 509, September 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guenther:1995:BRP

[Gue95] Ronald B. Guenther. Book review: Partial Differential Equations in Classical Mathematical Physics (I. Rubinstein and L. Rubinstein). SIAM Review, 37(1):123– 124, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guha:1966:BRB

[Guh66] D. K. Guha. Book review: Backorder Costs in a Periodic Review Inventory System (G. Hadley). SIAM Review, 8(4):532– 535, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guha:1968:BRG

[Guh68] D. K. Guha. Book review:

A Generating Function (P. I. Richards). SIAM Review, 10 (4):455–456, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guizhang:1992:BRI

[Gui92] Tu Guizhang. Book review: Integrable Systems of Classical Mechanics and Lie Algebras, Vol. I (A. M. Perelomov; A. G. Reyman, trans.). SIAM Review, 34(1): 137–??, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gundersen:1963:BRV

[Gun63] Roy M. Gundersen. Book review: Vectors, Tensors and the Basic Equations of Fluid Mechanics (R. Aris). SIAM Review, 5(2):165, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gunzburger:1988:BRF

[Gun88] Max Gunzburger. Book review: Finite Element Solution of Boundary Value Problems-Theory and Computation (O. Axelsson and V. A. Barker). SIAM Review, 30 (1):143–144, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gunzburger:1995:BRV

[Gun95] Max D. Gunzburger. Book review: Viscous Vortical Flows (L. Ting and R. Klein). SIAM Review, 37(1):124– 127, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gupta:1975:NF

[Gup75] M. M. Gupta. A nonnegative form. SIAM Review, 17(1):169, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gupta:1982:BRS

[Gup82] Shanti S. Gupta. Book review: Selecting and Ordering Populations: a New Statistical Methodology (Jean Dickinson Gibbons, Ingram Olkin and Milton Sobel). SIAM Review, 24(3):350– 351, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gupta:1989:BRN

[Gup89] Murli M. Gupta. Book review: The Numerical Solution of Ordinary and Partial Differential Equations (Granville Sewell). SIAM Review, 31(4):695–696, ???? 1989. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Gupta:1991:BRN

[Gup91a] Murli M. Gupta. Book review: Numerical Methods and Software (David Kahaner, Cleve Moler, and Stephen Nash). SIAM Review, 33(1):144–147, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gupta:1991:NMS

[Gup91b]  Murli M. Gupta.  Nu-

merical methods and software (David Kahaner, Cleve Moler, and Stephen Nash). SIAM Review, 33(1):144– ??, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gupta:1996:BRA

[Gup96]  Arjun K. Gupta.  Book

review: Applied Discriminant Analysis (Carl J. Huberty). SIAM Review, 38(1): 167–168, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gupta:1998:PSL

[Gup98a] R. C. Gupta. On particular solutions of linear difference equations with constant coefficients. SIAM Review, 40 (3):680–684, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gupta:1998:CNP

[Gup98b] Ramesh C. Gupta. Classroom note: On particular solutions of linear difference equations with constant coefficients. SIAM Review, 40 (3):680–684, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32019.

Gurel:1962:BRB

[Gur62a]  Okan Gurel.  Book re-

view: Boundary and Eigenvalue Problems in Mathematical Physics (H. Sagan). SIAM Review, 4(4):407– 408, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gurel:1962:BRD

[Gur62b] Okan Gurel. Book review: Differential Equations. (F. G. Tricomi). SIAM Review, 4 (3):269–270, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gurley:1963:BRI

[Gur63] J. G. Gurley. Book review: Inertial Guidance (G. R. Pitman, Jr., ed.). SIAM Review, 5(1):81–82, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gurland:1966:BRM

[Gur66] John Gurland. Book review:

Multidimensional Gaussian Distributions (Kenneth S. Miller). SIAM Review, 8 (4):551, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gustafson:1976:BRD

[Gus76]  G. B. Gustafson.  Book

review: Differential Equations, Linear Algebra and Dynamical Systems (Morris W. Hirsch and Stephen Smale). SIAM Review, 18(1): 137–138, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gustafson:1990:BRS

[Gus90a] Karl Gustafson. Book review: Stochastic Processes in the Neurosciences (Henry C. Tuckwell). SIAM Review, 32 (2):318–322, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gustafson:1990:SPN

[Gus90b] Karl Gustafson. Stochastic processes in the neurosciences (Henry C Tuckwell). SIAM Review, 32(2): 318–??, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guttmann:1974:ISM

[Gut74] A. J. Guttmann. An integral from statistical mechanics. SIAM Review, 16 (3):389, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guttorp:1982:BRL

[Gut82] P. Guttorp. Book review: A Learning Model (Y. P. Sabharwal and B. R. Gupta). SIAM Review, 24(4):479– 480, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guttorp:1988:BRS

[Gut88] Peter Guttorp. Book review: Series of Irregular Observations: Forecasting and Model Building (Robert Azencott and Didier Dacunha-Castelle). SIAM Review, 30 (3):513–514, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Guttorp:1990:BRI

[Gut90] Peter Guttorp. Book review: An Introduction to the Theory of Point Processes (D. J. Daley and D. Vere-Jones). SIAM Review, 32(1): 175–176, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gut:1997:BRP

[Gut97] Allan Gut. Book review:

Probability Theory: An Advanced Course, by Vivek S. Borkar. SIAM Review, 39(1): 149–150, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Grosjean:1997:PLP

[GV97] C. C. Grosjean and J. Van der Jeugt. Problems: a Laguerre polynomial identity (problem 97-5). SIAM Review, 39(1):126, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ganzha:1997:NSP

[GVL97] Victor G. Ganzha, Evgenii V.

Vorozhtsov, and Biyue Liu. Numerical solutions for partial differential equations: Problem solving using mathematics. SIAM Review, 39 (4):800–??, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Golub:1976:ICE

[GW76] G. H. Golub and J. H.

Wilkinson. Ill-conditioned eigensystems and the computation of the Jordan canonical form. SIAM Review, 18(4):578–619, October 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gulliver:1995:CHF

[GW95] R. Gulliver and N. B. Willms. A conjectured heatflow problem. SIAM Review, 37(1):100–104, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gander:2012:ERG

[GW12] Martin J. Gander and Gerhard Wanner. From Euler, Ritz, and Galerkin to modern computing. SIAM Review, 54 (4):627–666, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gordon-Wright:2013:GFS

[GWG13] Rachael Gordon-Wright and Pierre A. Gremaud. Granular free surfaces. SIAM Review, 55(1):168–184, ???? 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gulliver:1995:BRC

[GWK95] R. Gulliver, N. B. Willms, and B. Kawohl. Book review: A Conjectured Heat Flow Problem (M. S. Klamkin). SIAM Review, 37(1):100– 105, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Groote:2016:PAG

[GWZ16] Jan Friso Groote, Freek Wiedijk, and Hans Zantema. A probabilistic analysis of the game of the goose. SIAM Review, 58(1): 143–155, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Glasserman:1995:MSD

[GYS95] P. Glasserman, D. D. Yao, and Alexander Shapiro. Monotone structure in discreteevent systems. SIAM Review, 37(3):474–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gourley:2011:MMC

[GZ11] Stephen A. Gourley and Xingfu Zou. A mathematical model for the control and eradication of a wood boring beetle infestation. SIAM Review, 53(2): 321–345, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i2/p321_s1.

Gander:2019:CIS

[GZ19] Martin J. Gander and Hui Zhang. A class of iterative solvers for the Helmholtz equation: Factorizations, sweeping preconditioners, source transfer, single layer potentials, polarized traces, and optimized Schwarz methods. SIAM Review, 61(1): 3–76, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1971:BRS

[HA71] Alston S. Householder and Harry Appelgate. Book review: Some Algebraic Identities (Stan Kaplan). SIAM Review, 13(1):132– 136, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hsiao:1975:CME

[HA75] G. C. Hsiao and J. F. Ahner. Constructive method for exterior boundary-value problems of elasticity. SIAM Review, 17(2):381–382, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haake:1993:BRF

[Haa93] F. Haake. Book review: Foundations of Synergetics II, Complex Patterns (A. S. Mikhailov and A. Yu. Loskutov). SIAM Review, 35 (3):524–525, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haber:1970:NEM

[Hab70] Seymour Haber. Numerical evaluation of multiple integrals. SIAM Review, 12 (4):481–526, October 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2028488.

Haber:1983:ANT

[Hab83a] S. Haber. Applications of number theory of numerical analysis — Hua, L. K., Wang, Y. SIAM Review, 25(3): 414–416, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haber:1983:BRA

[Hab83b] Seymour Haber. Book review: Applications of Number Theory of Numerical Analysis (Loo Keng Hua and Yuan Wang). SIAM Review, 25(3):414–416, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hachtel:1975:AST

[Hac75]  G. D. Hachtel.  Astap

— sparse tableau program for solution and optimization of differential equations for engineering applications. SIAM Review, 17(2): 379–380, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hadley:1965:BCP

[Had65] G. Hadley. Backorder costs in a periodic review inventory system. SIAM Review, 7(4): 560–562, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hager:1980:BRM

[Hag80] William W. Hager. Book review: Mathematical Programming and Control Theory (B. D. Craven). SIAM Review, 22(3):377–378, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hager:1982:BRF

[Hag82] William W. Hager. Book review: Functional Analysis and Linear Control Theory (J. R. Leigh). SIAM Review, 24(3):349, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hager:1983:BRA

[Hag83] William W. Hager. Book review: Analysis of Global Expansion Methods: Weakly Asymptotically Diagonal Systems (L. M. Delves and T. L. Freeman). SIAM Review, 25 (1):123–124, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hagedorn:1985:BRSb

[Hag85a] George A. Hagedorn. Book review: Scattering Theory by the Enss Method (Peter A. Perry) (B. Simon, ed.). SIAM Review, 27(2): 274–275, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hagedorn:1985:BRSa

[Hag85b] George A. Hagedorn. Book review: Scattering Theory for Many-Body Quantum Mechanical Systems-Rigorous Results (Israel Michael Sigal). SIAM Review, 27 (1):103, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hager:1985:BRO

[Hag85c] William W. Hager. Book review: Optimal Processes on Manifolds (Roel Nottrot). SIAM Review, 27 (1):104, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hager:1986:BRC

[Hag86] William W. Hager. Book review: Computational Functional Analysis (Ramon E. Moore). SIAM Review, 28 (2):260, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hagedorn:1989:BRQ

[Hag89a] George A. Hagedorn. Book review: Quantum Scattering and Spectral Theory (D. B. Pearson). SIAM Review, 31 (3):506, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hager:1989:UIM

[Hag89b] William W. Hager. Updating the inverse of a matrix. SIAM Review, 31(2): 221–239, June 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haimo:1977:BRH

[Hai77] Deborah Tepper Haimo. Book review: The Heat Equation (D. V. Widder). SIAM Review, 19(2):364– 365, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haigh:1978:PP

[Hai78] John Haigh. On a Poisson process. SIAM Review, 20 (2):394, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haight:1985:BRM

[Hai85] Frank A. Haight. Book review: Mathematical Techniques of Applied Probability (Jeffrey J. Hunter). SIAM Review, 27(4):595– 596, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hajek:1979:SC

[Háj79] O. Hájek. Super-controllability.

SIAM Review, 21(2):256, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hajnal:1996:BRD

[Haj96]  Peter Hajnal.  Book re-

view: Davenport–Schinzel Sequences and Their Geometric Applications (Micha Sharir and Pankaj K. Agarwal). SIAM Review, 38(4): 689–691, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Halkin:1966:BRD

[Hal66] Huber Halkin. Book review: Dynamic Programming and the Calculus of Variations (Stuart E. Dreyfus). SIAM Review, 8(4): 547–549, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Halton:1970:RPS

[Hal70] John H. Halton. A retrospective and prospective survey of the Monte Carlo method. SIAM Review, 12 (1):1–63, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Halton:1986:TAP

[Hal86] John H. Halton. Two algebraic problems. SIAM Review, 28(2):231–232, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hallam:1990:BRT

[Hal90a] Thomas G. Hallam. Book review: The Theory of Evolution and Dynamical Systems: Mathematical Aspects of Selection (Josef Hofbauer and Karl Sigmund). SIAM Review, 32(2):322– 323, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hallam:1990:TED

[Hal90b] Thomas G. Hallam. The theory of evolution and dynamical systems: Mathematical aspects of selection (Josef Hofbauer and Karl Sigmund). SIAM Review, 32 (2):322–??, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hall:1995:BRC

[Hal95] Glen Richard Hall. Book review: Codes, Puzzles, and Conspiracy (Dennis Shasha). SIAM Review, 37(1):127– 128, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1961:BRN

[Ham61] Preston C. Hammer. Book review: Numerical Methods for Science and Engineering (R. G. Stanton). SIAM Review, 3(4):342, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1962:BRA

[Ham62a] P. C. Hammer. Book review: Adaptive Control Processes: a Guided Tour (R. Bellman). SIAM Review, 4 (2):163, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1962:BRC

[Ham62b] P. C. Hammer. Book review:

The Calculus of Variations (A. R. Forsyth). SIAM Review, 4(1):57–58, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1965:BRQ

[Ham65a] Preston C. Hammer. Book review: Qualitative Methods in Mathematical Analysis (L. È. Èl’sgol’c). SIAM Review, 7(4):570, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1965:BRT

[Ham65b] Preston C. Hammer. Book review: Translations of Mathematical Monographs, Vol. I. (A. D. Aleksandrov, A. N. Kolmogorov, and M. A. Lavrent’ev eds.). SIAM Review, 7(1):143– 148, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammersley:1969:SAD

[Ham69] John M. Hammersley. Sequences of absolute differences. SIAM Review, 11(1): 73–74, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1970:BRM

[Ham70] Preston C. Hammer. Book review: The Mathematical Sciences: a Collection of Essays (COSRIMS). SIAM Review, 12(3):474– 476, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1971:BRN

[Ham71] Preston C. Hammer. Book review: The Nature and Growth of Modern Mathematics (Edna E. Kramer). SIAM Review, 13(2):238– 240, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1973:BRN

[Ham73a] Preston C. Hammer. Book review: The Nonmathematical Foundations of Mathematics (R. J. Diamond). SIAM Review, 15(2):392– 393, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hamming:1973:BRNa

[Ham73b] Richard W. Hamming. Book review: Numerical Analysis and Computation: Theory and Practice (E. K. Blum). SIAM Review, 15 (4):807, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hamming:1973:BRNb

[Ham73c] Richard W. Hamming. Book review: Numerical Analysis and Computation: Theory and Practice (E. K. Blum). SIAM Review, 15 (4):807, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammersley:1974:SSS

[Ham74] J. M. Hammersley. Some speculations on a sense of nicely calculated chances. SIAM Review, 16(2):237– 255, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1976:BRF

[Ham76] Preston C. Hammer. Book review: The Future of Applied Mathematics (W. Freiberger, ed.). SIAM Review, 18(2):310–311, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1977:BRP

[Ham77] Preston C. Hammer. Book review: Popular Lectures in Mathematics (Izaak Wirszup, ed.); Inversions(I. Ya. Bakel’Man); Number Systems (S. V. Fomin); The Application of Mechanics to Geometry (B. Yu. Kogan); What is Distance? (Yu. A. Shreider); The Monte Carlo Method (I. M. Sobol’); Pascal’s Triangle (V. A. Uspenskii). SIAM Review, 19(1):170– 172, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1979:BRMa

[Ham79a] Preston C. Hammer. Book review: Mathematical Ideas, Modelling and Applications (Z. A. Melzak). SIAM Review, 21(1):148–149, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1979:BRMb

[Ham79b] Preston C. Hammer. Book review: Mathematical Methods in Science (George Polya). SIAM Review, 21 (3):424, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hammer:1980:BRM

[Ham80] Preston C. Hammer. Book review: Mathematics Today: Twelve Informal Essays (Lynn Arthur Steen, ed.). SIAM Review, 22(2): 250–252, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hamacher:1985:BRA

[Ham85] Horst W. Hamacher. Book review: Algebraische Methoden in der Ganzzahligen Optimierung (Joachim Piehler). SIAM Review, 27(3):467– 468, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hamacher:1987:BRO

[Ham87] Horst W. Hamacher. Book review: Optimisation Combinatoire — Méthodes Mathématiques et Algorithmiques: Programmation Discrète; Optimisation Combinatoire Méthodes Mathématiques et Algorithmiques: Graphes et Programmation Linéaire (Michel Sakarovitch). SIAM Review, 29(1):143–144, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hamacher:1988:BRM

[Ham88] Horst W. Hamacher. Book review: Minimal Cost Flow in Processing Networks, A Primal Approach (J. Koene). SIAM Review, 30(2):330– 331, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hamacher:1990:BRT

[Ham90] Horst W. Hamacher. Book review: Theoretical and Computational Aspects of Simulated Annealing (P. J. M. van Laarhoven); Simulated Annealing: Theory and Applications (P. J. M. van Laarhoven and E. H. L. Aarts). SIAM Review, 32 (3):504–506, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1964:BRN

[Han64] Eldon Hansen. Book review: On the Nonnegativity of the Coefficients of a Certain Series Expansion (Dennis C. Gilliland). SIAM Review, 6 (4):461–462, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1967:BRI

[Han67] Eldon Hansen. Book review: Interval Analysis (Ramon E. Moore). SIAM Review, 9(3): 610–612, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Handelman:1975:SMS

[Han75a] George H. Handelman. A simple model of stability of structures. SIAM Review, 17 (4):593–604, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hannsgen:1975:BRO

[Han75b] Kenneth B. Hannsgen. Book review: Ordinary Differential Equations (Philip Hartman). SIAM Review, 17(4): 704–705, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hanson:1976:PPP

[Han76] R. J. Hanson. Poorly posed problems of data fitting and integral equations of 1st kind in one and two variables. SIAM Review, 18 (4):807, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hannsgen:1979:BRN

[Han79] Kenneth B. Hannsgen. Book review: Nonlinear Ordinary Differential Equations (D. W. Jordan and P. Smith). SIAM Review, 21(2):264, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hannsgen:1981:BRN

[Han81a] Kenneth B. Hannsgen. Book review: Nonlinear Differential Equations (Svatopluk Fucik and Alois Kufner). SIAM Review, 23(3):402– 403, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1981:BRM

[Han81b] Eldon Hansen. Book review: Methods and Applications of Interval Analysis (Ramon E. Moore). SIAM Review, 23(1): 121–123, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1983:NOC

[Han83a]  E. R. Hansen.  Numer-

ical optimization of computer models — Schwefel, H. P. SIAM Review, 25(3): 431–433, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1983:BRN

[Han83b] Eldon R. Hansen. Book review: Numerical Optimization of Computer Models (Hans-Paul Schwefel). SIAM Review, 25(3):431– 433, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Handelman:1984:P

[Han84] George Handelman. Preface. SIAM Review, 26(4): 589–590, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hannsgen:1985:BRV

[Han85a] Kenneth B. Hannsgen. Book review: Volterra Integral and Differential Equations (T. A. Burton). SIAM Review, 27 (1):104–105, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1985:BRS

[Han85b] Eldon Hansen. Book review: A Summation of Bessel Functions (Henry E. Fettis). SIAM Review, 27(3): 455–457, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1985:SIB

[Han85c] Eldon Hansen. Series involving Bessel functions. SIAM Review, 27(4):576, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Handelman:1986:SRA

[Han86a] George H. Handelman. Some reminiscences about Richard C. DiPrima:1927-1984. SIAM Review, 28(3):307–313, September 1986. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1986:BRN

[Han86b] Eldon Hansen. Book review: Numerical Methods for Unconstrained Optimization and Nonlinear Equations (J. E. Dennis, Jr., and Robert B. Schnabel). SIAM Review, 28 (3):417–419, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hannsgen:1987:BRS

[Han87] Kenneth B. Hannsgen. Book review: Stability and Periodic Solutions of Ordinary and Functional Differential Equations (T. A. Burton). SIAM Review, 29(4):652– 654, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1989:MER

[Han89] Vagn Lundsgaard Hansen. Minima for elastics on rigid cylinders. SIAM Review, 31 (2):310–317, June 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Handscomb:1990:ICT

[Han90] D. C. Handscomb. Impossibility of a conic through six points. SIAM Review, 32(2): 300–301, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hannsgen:1991:BRR

[Han91] Kenneth B. Hannsgen. Book review: Retarded Dynamical Systems: Stability and Characteristic Functions (G. Stépán). SIAM Review, 33 (1):147, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1992:ADI

[Han92] Per Christian Hansen. Analysis of discrete ill-posed problems by means of the Lcurve. SIAM Review, 34 (4):561–580, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hanke:1994:CIL

[Han94a] M. Hanke. Convergence of iterations for linear equations (Olavi Nevanlinna). SIAM Review, 36(3):508– ??, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hanks:1994:BRC

[Han94b] Martin Hanks. Book review:

Convergence of Iterations for Linear Equations (Olavi Nevanlinna). SIAM Review, 36(3):508–509, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hanke:1995:BRI

[Han95] Martin Hanke. Book review: Iterative Solution Methods (Owe Axeisson). SIAM Review, 37(3):466– 467, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1998:TBC

[Han98] E. R. Hansen. Three binomial convolutions. SIAM Review, 40(4):991–993, November 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harary:1960:SHI

[Har60] Frank Harary. Some historical and intuitive aspects of graph theory. SIAM Review, 2(2):123–131, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1961:BRT

[Har61a] W. A. Harris, Jr. Book review: A Treatise on the Calculus of Finite Differences (George Boole). SIAM Review, 3(4):341, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1961:BRF

[Har61b] W. A. Harris, Jr. Book review: Finite Difference Equations (H. Levy and F. Lessman). SIAM Review, 3(4):

340–341, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harary:1962:DAM

[Har62] Frank Harary. The determinant of the adjacency matrix of a graph. SIAM Review, 4 (3):202–210, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hartley:1963:BRC

[Har63]  H. O. Hartley.  Book

review: Contributions to the Theory of Estimation from Grouped and Partially Grouped Samples (G. Kulldorf). SIAM Review, 5(4): 379–380, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harary:1964:BRA

[Har64a] Frank Harary. Book review:

Applications of Graph Theory to Group Structure (Claude Flament). SIAM Review, 6 (2):185–186, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1964:BRP

[Har64b] Bernard Harris. Book review: Probability Theory, third edition (Michel Loeve). SIAM Review, 6(3): 318–319, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1966:BRE

[Har66a] Bernard Harris. Book review: The Estimation of Probabilities: an Essay on Modern Bayesian Methods (I. J. Good). SIAM Review, 8(1): 118–119, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harvey:1966:MCL

[Har66b] C. A. Harvey. Minimax control of linear systems with amplitude bounded disturbances. SIAM Review, 8(3): 370–373, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harary:1967:GM

[Har67a] Frank Harary. Graphs and matrices. SIAM Review, 9 (1):83–90, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harkin:1967:MSI

[Har67b] J. B. Harkin. Mean Stieltjes integrals of order p. SIAM Review, 9(1):107– 114, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harrington:1967:PMT

[Har67c] W. J. Harrington. A property of Mellin transforms. SIAM Review, 9(3):542– 547, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1967:BRL

[Har67d] Bernard Harris. Book review: The Logic of Decision (Richard C. Jeffrey). SIAM Review, 9(4):761– 763, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1970:BRA

[Har70] William A. Harris, Jr. Book review: Asymptotic Methods in the Theory of Linear Differential Equations (S. F. Feshchenko, N. I. Shkil’ and L. D. Nikolenko). SIAM Review, 12(1):163–164, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1972:BRS

[Har72a] B. Harris. Book review: The Structure of Inference (D. A. S. Fraser). SIAM Review, 14 (3):511–512, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hartmanis:1972:BRL

[Har72b] Juris Hartmanis. Book review: Logic, Automata, and Algorithms (Mark A. Aisermann, Leonid A. Gusev, Lev I. Rozonoer, Irena M. Smirnova and Aleksey A. Tal’). SIAM Review, 14(2): 343–344, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harary:1973:BRA

[Har73] Frank Harary. Book review: Applied Graph Theory (W. K. Chen). SIAM Review, 15 (1):230, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harary:1974:BRG

[Har74] Frank Harary. Book review: Graph Theory in Modern Engineering (Ernest J. Henley and R. A. Williams). SIAM Review, 16(2):269– 270, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hartmanis:1975:BRF

[Har75] J. Hartmanis. Book review: Formal Languages (Arto Salomaa). SIAM Review, 17(1): 179–180, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1976:BRP

[Har76a] Bernard Harris. Book review: Probability Theory: a Historical Sketch (L. E.

Maistrov). SIAM Review, 18 (4):786–788, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1976:BRI

[Har76b] W. A. Harris, Jr. Book review: Introduction To Singular Perturbations (Robert E. O’Malley, Jr.). SIAM Review, 18(3):517–518, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harvey:1976:IPQ

[Har76c] C. A. Harvey. Independent parameters in quadratic performance indexes. SIAM Review, 18(4):807, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:1978:BRL

[Har78] Bernard Harris. Book review: The Life and Times of the Central Limit Theorem (William J. Adams). SIAM Review, 20(3):609– 610, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hartmanis:1982:BRC

[Har82] Juris Hartmanis. Book review: Computers and Intractability: a Guide to the Theory of NP-Completeness (Michael R. Garey and David S. Johnson). SIAM Review, 24(1):90–91, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harary:1983:BRE

[Har83] Frank Harary. Book review: Examples and Counterexamples in Graph Theory (Michael Capobianco and John C. Molluzzo). SIAM Review, 25(1):136– 138, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harter:1993:BRO

[Har93] H. Leon Harter. Book review: Order Statistics and Inference: Estimation Methods (N. Balakrishnan and A. Clifford Cohen). SIAM Review, 35(1):149–150, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harrell:1996:BRC

[Har96] Evans M. Harrell II. Book review: The Complex WKB Method for Nonlinear Equations 1: Linear Theory (Victor P. Maslov). SIAM Review, 38(1):168–171, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hastings:1986:BRC

[Has86] Alan Hastings. Book review: Continuous and Discrete Dynamics Near Manifolds of Equilibria (Bernd Aulbach). SIAM Review, 28 (1):105, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hastings:1992:BRD

[Has92a] S. P. Hastings. Book review: Differential Equations and Dynamical Systems (Lawrence Perko). SIAM Review, 34(1):129–131, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hastings:1992:DED

[Has92b]  S. P. Hastings.  Differ-

ential equations and dynamical systems (Lawrence Perko). SIAM Review, 34(1): 129–??, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haussmann:1975:PSE

[Hau75] U. G. Haussmann. A power series expansion. SIAM Review, 17(1):168, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haussmann:1976:PSE

[Hau76] U. G. Haussmann. Powerseries expansion. SIAM Review, 18(2):301–302, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haussmann:1978:NAS

[Hau78] U. G. Haussmann. Numerical approximation of stochastic optimal controls. SIAM Review, 20(3):626, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haussmann:1981:SEO

[Hau81] U. G. Haussmann. Some examples of optimal stochastic controls OR: The stochastic maximum principle at work. SIAM Review, 23(3): 292–307, July 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haug:1985:BRV

[Hau85] Edward J. Haug. Book review: Variational Methods in Theoretical Mechanics (J. T. Oden and J. N. Reddy). SIAM Review, 27 (1):105–106, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haug:1986:BRM

[Hau86] Edward J. Haug. Book review: Methods of Functional Analysis for Application in Solid Mechanics (Jayme Mason). SIAM Review, 28 (4):588–589, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haug:1988:BRA

[Hau88] Edward J. Haug. Book review: Applied Functional Analysis and Variational Methods in Engineering (J. N. Reddy). SIAM Review, 30 (1):144–145, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haussmann:1989:BRL

[Hau89] U. G. Haussmann. Book review: Linear Systems and Optimal Control (C. K. Chui and G. Chen). SIAM Review, 31(4):696–698, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haug:1990:VPC

[Hau90] Edward J. Haug. Variational principles of continuum mechanics with engineering applications. volume 2: Introduction to optimal design theory (Vadim Komkov). SIAM Review, 32 (4):696–??, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Havner:1995:BRM

[Hav95] K. S. Havner. Book review:

Mathematical Modelling of Inelastic Deformation (J. F. Besseling and E. van der Giessen). SIAM Review, 37 (3):467–468, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hawkins:1995:ASP

[Haw95] William G. Hawkins. Analytic study of the post attenuation correction algorithm. SIAM Review, 37(3): 439–440, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hayes:1968:CP

[Hay68] Melda Hayes. A combinatorial problem. SIAM Review, 10(4):451, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hjorth:1995:CIS

[HB95] J. S. Urban Hjorth and R. J. Beran. Computer intensive statistical methods. SIAM Review, 37(3):459– ??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hanke:1996:CGT

[HB96] Martin Hanke and Helmut Brakhage. Conjugate gradient type methods for ill-posed problems. SIAM Review, 38 (4):682–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hemmer:1966:BRP

[HBO66] P. C. Hemmer, A. J. Bosch, and W. Oettli. Book review: On a Probability of Overlap (Murray S. Klamkin). SIAM Review, 8(1):112– 115, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes-Cerfon:2016:ERS

[HC16] Miranda C. Holmes-Cerfon.

Enumerating rigid sphere packings. SIAM Review, 58 (2):229–244, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Handelman:1975:MTN

[HCH^{+}75] G. H. Handelman, B. H.

Colvin, E. W. Hart, A. J. Hoffman, M. S. Klamkin, and J. J. Seidel. Mathematical training of nonacademic mathematician. SIAM Review, 17(3):541–557, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hull:1962:RNG

[HD62] T. E. Hull and A. R. Dobell. Random number generators. SIAM Review, 4(3): 230–254, July 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://

dspace.library.uvic.ca:
8443/bitstream/handle/

                                        1828/3142/Random_Number_
                                        Generators.pdf?sequence=
                                        3.

Hanau:1991:OSB

[HE91] P. T. L. M. Hanau and M. Eslami. Optimization of a stability bound. SIAM Review, 33(2):277–281, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hearon:1974:BRS

[Hea74] John Z. Hearon. Book review: A Special Class of Matrices (J. F. Foley). SIAM Review, 16(4):548–549, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heath:1976:SAI

[Hea76] M. T. Heath. Some aspects of interval estimation for illposed problems. SIAM Review, 18(4):808, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hearon:1981:BRE

[Hea81] J. Z. Hearon. Book review: Eigenvalues of a Tri-Diagonal Matrix (D. K. Ross). SIAM Review, 23(1): 112–113, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Healey:1991:BRS

[Hea91] Timothy J. Healey. Book review: Singularity Theory, Rod Theory and Symmetry-Breaking Loads (John F. Pierce). SIAM Review, 33 (2):314–315, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heidel:1969:SPA

[Hei69] J. W. Heidel. A short proof of Atkinson’s oscillation theorem. SIAM Review, 11(3): 389–390, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heinrich:1989:BRC

[Hei89] Katherine Heinrich. Book review: Combinatorial Designs (W. D. Wallis). SIAM Review, 31(4):698–699, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heij:1993:BRI

[Hei93a] Christiaan Heij. Book review:

Implicit Linear Systems (J. D. Aplevich). SIAM Review, 35(1):150–151, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heil:1993:BRT

[Hei93b]  Christopher Heil.  Book

review: Ten Lectures on Wavelets, by Ingrid Daubechies.

SIAM Review, 35(4):666– 669, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2132406.

Heijmans:1995:MMM

[Hei95] Henk J. A. M. Heijmans. Mathematical morphology: a modern approach in image processing based on algebra and geometry. SIAM Review, 37(1):1–36, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/23546.htm.

Heitjan:1996:BRO

[Hei96] Daniel F. Heitjan. Book review: Observational Studies (Paul R. Rosenbaum. Springer-Verlag). SIAM Review, 38(3):529, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heinrich:1997:BRC

[Hei97] Katherine Heinrich. Book review: The CRC Handbook of Combinatorial Designs, by Charles J. Colbourn and Jeffrey H. Dinitz. SIAM Review, 39(2):345–346, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Helfand:1975:AMM

[Hel75] H. M. Helfand. Application of method of multiple scales to problems arising in supersonic aerodynamics. SIAM Review, 17(2):380, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heller:1978:SPA

[Hel78] Don Heller. A survey of parallel algorithms in numerical linear algebra. SIAM Review, 20(4):740–777, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Helffer:1997:BRI

[Hel97] Bernard Helffer. Book review: Introduction to Spectral Theory, with Applications to Schrödinger Operators, by P. D. Hislop and I. M. Sigal. SIAM Review, 39(1): 150–152, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hennart:1975:EFE

[Hen75]  J. P. Hennart.  Extrap-

olated finite-element procedures over Cartesian product domains. SIAM Review, 17 (2):380–381, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hendricks:1976:CRW

[Hen76] F. Hendricks. Computations of resonant wave interactions in Blasius flow. SIAM Review, 18(4):808, ???? 1976.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Henrici:1979:FFM

[Hen79] Peter Henrici. Fast Fourier methods in computational complex analysis. SIAM Review, 21(4):481–527, October 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL https://www.
                                        jstor.org/stable/pdf/10.
                                        2307/2030105.pdf.

Henrici:1982:DIP

[Hen82]  Peter Henrici.  A defi-

nite integral: Problem 82-17. SIAM Review, 24(4): 477, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //www.jstor.org/stable/
                                        2029543.

Henrici:1986:BRI

[Hen86] Peter Henrici. Book review: Introduction to Applied Mathematics by Gilbert Strang. SIAM Review, 28 (4):590–592, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //www.jstor.org/stable/
                                        2031128.

Henkel:1992:AMS

[Hen92] Malts Henkel. On alternating multiple sums. SIAM Review, 34(3):496, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heppes:1966:DTT

[Hep66] A. Heppes. A double tipping tetrahedron. SIAM Review, 8(3):381, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heppes:1967:BRD

[Hep67] A. Heppes. Book review: A Double Tipping Tetrahedron (A. Heppes). SIAM Review, 9 (3):599–600, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermes:1965:BRP

[Her65] H. Hermes. Book review:

Periodic Solutions of Perturbed Second-Order Autonomous Equations (W. S. Loud). SIAM Review, 7 (2):300, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermes:1967:BRT

[Her67a]  H. Hermes.  Book re-

view: Time-Lag Control Systems (M. Namik Ogustoreli). SIAM Review, 9(1): 131–132, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermes:1967:BRN

[Her67b] Henry Hermes. Book review: Nonlinear Differential Equations (T. V. Davies and Eleanor M. James). SIAM Review, 9(2):262– 264, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermes:1969:BRO

[Her69] Henry Hermes. Book review: Optimal Control Theory (R. Pallu de la Barrièrs). SIAM Review, 11(3):417, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hertling:1970:BRN

[Her70] Jörg Hertling. Book review: Numerical Methods for Two-Point Boundary Value Problems (Herbert B. Keller). SIAM Review, 12(2):313– 315, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermes:1971:BRQ

[Her71] Henry Hermes. Book review: The Qualitative Theory of Ordinary Differential Equations (Fred Brauer and John A. Nohel). SIAM Review, 13 (1):107–108, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermes:1973:SRR

[Her73] Henry Hermes. A survey of recent results in differential equations. SIAM Review, 15 (2):453–468, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermann:1977:BRT

[Her77] Robert Hermann. Book review: Treatise on Analysis, Vol. IV (J. Dieudonné). SIAM Review, 19(4):750, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermann:1978:BRS

[Her78a] Robert Hermann. Book review: Singular and Degenerate Cauchy Problems (R. W. Carroll and R. E. Showalter). SIAM Review, 20(3): 612–613, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hersh:1978:HOA

[Her78b] R. Hersh. Higher-order approximations for C_{0} semigroups in Banach-spaces. SIAM Review, 20(3):626, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermann:1980:BRDa

[Her80a]  Robert Hermann.  Book

review: Differential Forms in Mathematical Physics (C. Von Westenholz). SIAM Review, 22(1):106–107, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermann:1980:BRDb

[Her80b] Robert Hermann. Book review: Differential Geometry, Lie Groups, and Symmetric Spaces (Sigurdur Helgason). SIAM Review, 22(4): 524–526, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Herold:1981:IMP

[Her81] Edward W. Herold. Inflation mathematics for professionals. SIAM Review, 23(3): 381–384, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermann:1983:GLG

[Her83a]  R. Hermann.  Global

Lorentzian geometry — Beem, J. K., Ehrlich, P. E. SIAM Review, 25(3): 417–418, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermann:1983:BRG

[Her83b] Robert Hermann. Book review: Global Lorentzian Geometry (John K. Beem and Paul E. Ehrlich). SIAM Review, 25(3):417–418, ????

1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermann:1984:EDS

[Her84a] R. Hermann. Exterior differential systems and the calculus of variations — Griffiths, P. A. SIAM Review, 26(2): 304–306, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermann:1984:BRE

[Her84b] Robert Hermann. Book review: Exterior Differential Systems and the Calculus of Variations. (Phillip A. Griffiths). SIAM Review, 26(2): 304–306, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Herron:1987:OSE

[Her87] Isom H. Herron. The Orr– Sommerfeld equation on infinite intervals. SIAM Review, 29(4):597–620, December 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hersch:1988:BRI

[Her88] Joseph Hersch. Book review:

Isoperimetric Monotonicity: Some Properties and Conjectures (Connections between Isoperimetric Inequalities). SIAM Review, 30(4):551– 577, December 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Herron:1990:TDI

[Her90] I. Herron. Two definite integrals. SIAM Review, 32 (1):140, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermes:1991:NHO

[Her91] Henry Hermes. Nilpotent and high-order approximations of vector field systems. SIAM Review, 33(2): 238–264, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hersh:1992:NDK

[Her92a] R. Hersh. Nicholas D. Kazarinoff — in memoriam. SIAM Review, 34(1):127, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hersh:1992:MND

[Her92b] Reuben Hersh. In memoriam: Nicholas D. kazarinoff. SIAM Review, 34 (1):127, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hersh:1993:BRE

[Her93a] Reuben Hersh. Book review: Ethnomathematics: a Multicultural View of Mathematical Ideas (Marcia Ascher). SIAM Review, 35(1): 151–153, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hersh:1993:EMV

[Her93b] Reuben Hersh. Ethnomathematics: a multicultural view of mathematical ideas (Marcia Ascher). SIAM Review, 35(1):151–??, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hermann:1997:BRI

[Her97] Robert Hermann. Book review: Introduction to Mechanics and Symmetry, by Jerrold E. Marsden and Tudor S. Ratiu. SIAM Review, 39(1):152, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hestene:1972:BRM

[Hes72]  M. R. Hestene.  Book

review: Mehrdimensionale Variationsrechnung (R. Klotzler). SIAM Review, 14(2): 351–352, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hess:1983:BRN

[Hes83]  Peter Hess.  Book re-

view: Nonlinear Mappings of Monotone Type (Dan Pascali and Silvu-Sburlan). SIAM Review, 25(1):116, ???? 1983.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hess:1986:BRT

[Hes86] James D. Hess. Book review: Tinbergen Lectures on Organization Theory (Martin J. Beckmann). SIAM Review, 28(2):260–261, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hethcote:1972:GEP

[Het72] Herbert W. Hethcote. Geometric existence proofs for nonlinear boundary value problems. SIAM Review, 14 (1):121–128, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hethcote:2000:MID

[Het00] Herbert W. Hethcote. The mathematics of infectious diseases. SIAM Review, 42 (4):599–653, December 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37190.

Hewer:1976:SST

[Hew76] G. A. Hewer. Stabilization of subspaces of time-varying linear systems. SIAM Review, 18(4):808, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heyda:1963:TDO

[Hey63]  James F. Heyda.  Tem-

perature distribution in an orbiting sphere with alternate heating and cooling. SIAM Review, 5(2): 113–118, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heyman:1982:BRQ

[Hey82] Daniel P. Heyman. Book review: Queueing Tables and Graphs (Frederick S. Hillier and Oliver S. Yu with D. M. Avis, L. D. Fosset, F. D. Lo, M. I. Reiman). SIAM Review, 24(3):362–363, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heyman:1985:BRQ

[Hey85] Daniel P. Heyman. Book review: Queues and Point Processes (Peter Franken, Dieter König, Ursula Arndt and Volker Schmidt). SIAM Review, 27(2):275–276, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heyman:1986:BRM

[Hey86] Daniel P. Heyman. Book review: The M/M/Infinity Service System with Ranked Servers in Heavy Traffic (G. F. Newell). SIAM Review, 28 (3):419–420, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heyman:1988:BRP

[Hey88] Daniel P. Heyman. Book review: Point Processes and Their Statistical Inference (Alan F. Karr). SIAM Review, 30(3):514–515, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heyman:1989:BRT

[Hey89] Daniel P. Heyman. Book review: Traffic Processes in Queueing Networks: a Markov Renewal Approach (Ralph L. Disney and Peter C. Kiessler). SIAM Review, 31(3):506–507, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heyman:1993:BRS

[Hey93] Daniel P. Heyman. Book review: Stochastic Models in Queueing Theory (J. Medhi). SIAM Review, 35 (2):320–321, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heywood:1997:BRH

[Hey97] Phillp Heywood. Book review: The Hilbert Transform of Schwartz Distributions and Applications, by J. N. Pandey. SIAM Review, 39 (2):346–347, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoover:1978:SMQ

[HF78] W. E. Hoover and J. S.

Frame. Some multidimensional quadrature formulas using partial derivatives. SIAM Review, 20 (3):627, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hersh:2005:BRM

[HF05] Reuben Hersh and Roger Frye. Book review: Mathematics by Experiment: Plausible Reasoning in the 21st Century. SIAM Review, 47 (4):832–833, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
20453726.

Holmes:2006:DLL

[HFKG06] Philip Holmes, Robert J.

Full, Dan Koditschek, and John Guckenheimer. The dynamics of legged locomotion: Models, analyses, and challenges. SIAM Review, 48(2): 207–304, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harris:2001:MEA

[HFS01] William A. Harris, Jr., Jay P.

Fillmore, and Donald R. Smith. Matrix exponentials — another approach.

SIAM Review, 43(4):694– 706, December 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/36240.

Haigh:1979:PP

[HG79] John Haigh and G. Gripenberg. On a Poisson process. SIAM Review, 21(4): 560–561, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Huberty:1996:ADA

[HG96] Carl J. Huberty and Arjun K. Gupta. Applied discriminant analysis. SIAM Review, 38 (1):167–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Henrici:1975:CCC

[HH75] Peter Henrici and Christian Hoffmann. Chess champ’s chances: an exercise in asymptotics. SIAM Review, 17(3):559–564, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Halbeisen:2000:PBT

[HH00] Lorenz Halbeisen and Norbert Hungerbühler. On periodic billiard trajectories in Obtuse triangles. SIAM Review, 42(4):657– 670, December 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35572.

Hadjighasem:2016:GTB

[HH16] Alireza Hadjighasem and George Haller. Geodesic transport barriers in Jupiter’s atmosphere: a video-based analysis. SIAM Review, 58 (1):69–89, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2019:DLI

[HH19] Catherine F. Higham and Desmond J. Higham. Deep learning: an introduction for applied mathematicians. SIAM Review, 61(4):860– 891, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Head:1993:ETD

[HHOT93] A. K. Head, S. D. Howison, J. R. Ockendon, and S. P. Tighe. An equilibrium theory of dislocation continua. SIAM Review, 35 (4):580–609, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hall:1975:SEU

[HHSS75] H. H. Hall, E. O. Heady, A. Stoecker, and V. A.

Sposito. Spatial equilibrium in U.S. agriculture: a quadratic programming analysis. SIAM Review, 17(2): 323–338, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Halanay:1996:TVD

[HIC96] Aristide Halanay, Vlad Ionescu, and T. Constantinescu. Timevarying discrete linear systems. SIAM Review, 38(4): 684–??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higgins:1962:BRE

[Hig62a] T. J. Higgins. Book review: Einführung in die Theorie der Abtastsysteme (J. Tschauner and R. Oldenbourg). SIAM Review, 4(1): 56–57, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higgins:1962:BRT

[Hig62b] T. J. Higgins. Book review: Transistor Logic Circuits (R. B. Hurley). SIAM Review, 4 (1):55–56, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higgins:1962:BRI

[Hig62c] Thomas J. Higgins. Book review: Introduction to the Statistical Dynamics of Automatic Control Systems (V. V. Solodovnikov). SIAM Review, 4(1):52–53, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higgins:1962:BRS

[Hig62d] Thomas J. Higgins. Book review: Systemtheorie für regellose Vorganze: Statische Verfahren für Nachrichtenfind Regelungstechnik (H. S. Schlitt). SIAM Review, 4(2): 161–162, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higgins:1965:BRT

[Hig65] Thomas J. Higgins. Book review: Topics in the Theory of Random Noise, vol.1 (R. L. Stratonovich). SIAM Review, 7(4):574–576, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higdon:1986:IBV

[Hig86] Robert L. Higdon. Initialboundary value problems for linear hyperbolic systems. SIAM Review, 28(2): 177–217, June 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:1987:SCN

[Hig87] Nicholas J. Higham. A survey of condition number estimation for triangular matrices. SIAM Review, 29 (4):575–596, December 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:1992:BRO

[Hig92] Desmond J. Higham. Book review: Order Stars (A. Iserles and S. P. Nørsett). SIAM Review, 34(4):666– 668, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:1999:FBR

[Hig99] Nicholas J. Higham. Featured book review: Selected books on numerical linear algebra. SIAM Review, 41 (3):607–610, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2000:RJJ

[Hig00] Nicholas J. Higham. Review of “Jack J. Dongarra, Iain S. Duff, Danny C. Sorensen, and Henk A. van der Vorst, Numerical Linear Algebra for High-Performance Computers, Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1998”. SIAM Review, 42(3): 529, September 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2001:AIN

[Hig01a] Desmond J. Higham. An algorithmic introduction to numerical simulation of stochastic differential equations. SIAM Review, 43(3):525– 546, September 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37830.

Higham:2001:RMA

[Hig01b] Nicholas J. Higham. Review of “Michael Alley, The Craft of Editing: A Guide for Managers, Scientists and Engineers. Springer-Verlag, 2000”. SIAM Review, 43 (1):202, March 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2002:NWI

[Hig02a] Desmond J. Higham. Nine ways to implement the binomial method for option valuation in MATLAB. SIAM Review, 44(4):661–677, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/39326.

Higham:2002:RAN

[Hig02b] Nicholas J. Higham. Review of “Arnold Neumaier. Introduction to Numerical Analysis. Cambridge University Press, 2001”. SIAM Review, 44(3):492–493, September 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2002:RMO

[Hig02c] Nicholas J. Higham. Review of “Michael Overton, Numerical Computing with IEEE Floating Point Arithmetic: Including One Theorem, One Rule of Thumb, and One Hundred and One Exercises. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2001”. SIAM Review, 44(2): 287–288, June 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2003:OMP

[Hig03] Nicholas J. Higham. Jorthogonal matrices: Properties and generation. SIAM Review, 45(3):504–519, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/41493.

Higham:2005:BRS

[Hig05] Nicholas J. Higham. Book review: The SIAM 100-Digit Challenge: A Study in High-Accuracy Numerical Computing. By Folkmar Bornemann, Dirk Laurie, Stan Wagon, and Jörg Waldvogel. SIAM, Philadelphia, 2004. $57.00. xii + 306 pp., hardcover. ISBN 0-89871-561-X. SIAM Review, 47(2):382–383, June 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97149.

Higham:2006:RAJ

[Hig06] Nicholas J. Higham. Review of “Alan J. Laub. Matrix Analysis for Scientists and Engineers”. SIAM Review, 48(1):170–171, March 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2007:MPV

[Hig07] Desmond J. Higham. A matrix perturbation view of the small world phenomenon. SIAM Review, 49 (1):91–108, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2008:MSC

[Hig08] Desmond J. Higham. Modeling and simulating chemical reactions. SIAM Review, 50 (2):347–368, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2009:SSM

[Hig09] Nicholas J. Higham. The scaling and squaring method for the matrix exponential revisited. SIAM Review, 51(4): 747–764, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2012:SRa

[Hig12a] Desmond J. Higham. Survey and review. SIAM Review, 54(1):1, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v54/i1/p1_s1.

Higham:2012:SRb

[Hig12b] Desmond J. Higham. Survey and review. SIAM Review, 54(2):209, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2012:SRc

[Hig12c] Desmond J. Higham. Survey and review. SIAM Review, 54(3):419, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2012:SRd

[Hig12d] Desmond J. Higham. Survey and review. SIAM Review, 54(4):625, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2013:SRa

[Hig13a] Desmond J. Higham. Survey and review. SIAM Review, 55(1):1, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2013:SRb

[Hig13b] Desmond J. Higham. Survey and review. SIAM Review, 55(3):413, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2013:SRc

[Hig13c] Desmond J. Higham. Survey and review. SIAM Review, 55(4):599, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2014:SRa

[Hig14a] Desmond J. Higham. Survey and review. SIAM Review, 56(1):1, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2014:SRb

[Hig14b] Desmond J. Higham. Survey and review. SIAM Review, 56(2):211, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2014:SRc

[Hig14c] Desmond J. Higham. Survey and review. SIAM Review, 56(3):383, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2014:SRd

[Hig14d] Desmond J. Higham. Survey and review. SIAM Review, 56(4):575, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2015:SRa

[Hig15a] Desmond J. Higham. Survey and review. SIAM Review, 57(1):1, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2015:SRb

[Hig15b] Desmond J. Higham. Survey and review. SIAM Review, 57(2):165, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2015:SRc

[Hig15c] Desmond J. Higham. Survey and review. SIAM Review, 57(2):165, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2015:SRd

[Hig15d] Desmond J. Higham. Survey and review. SIAM Review, 57(3):319, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2015:SRe

[Hig15e] Desmond J. Higham. Survey and review. SIAM Review, 57(4):481, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2016:SRa

[Hig16a] Desmond J. Higham. Survey and review. SIAM Review, 58(1):1, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2016:SRb

[Hig16b] Desmond J. Higham. Survey and review. SIAM Review, 58(2):175, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2016:SRc

[Hig16c] Desmond J. Higham. Survey and review. SIAM Review, 58(3):375, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2016:SRd

[Hig16d] Desmond J. Higham. Survey and review. SIAM Review, 58(4):601, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hillstrom:1975:CMP

[Hil75] K. E. Hillstrom. Comments on modularization of a package of algorithms for unconstrained nonlinear optimization problem. SIAM Review, 17(2):381, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hill:1976:BRF

[Hil76]  R. N. Hill.  Book re-

view: Fourier Coefficients of a Function Involving Elliptic Integrals of the First Kind (E. O. Schulz-DuBois). SIAM Review, 18(1):130– 131, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hill:1996:BRI

[Hil96] Mary C. Hill. Book review: Inverse Problems in Groundwater Modeling (Ne-Zheng Sun). SIAM Review, 38(2):342–347, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hill:1998:TSP

[Hil98] R. N. Hill. Is there a simple proof? SIAM Review, 40 (2):385–388, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hillen:2007:CSP

[Hil07] Thomas Hillen. A classification of spikes and plateaus. SIAM Review, 49(1):35– 51, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Himmelblau:1978:BRM

[Him78] D. Himmelblau. Book review: Mathematical Programming Methods (G. Zoutendijk). SIAM Review, 20 (3):605–606, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hinich:1987:BRI

[Hin87] Melvin J. Hinich. Book review: An Introduction to Bispectral Analysis and Bilinear Time Series Models (T. Subba Rao and M. M. Gabr). SIAM Review, 29(2): 326–329, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hinton:1989:BRN

[Hin89] Don Hinton. Book review:

Non-Oscillation Domains of Differential Equations with Two Parameters (Angelo B. Mingarelli and S. Gotskalk Halvorsen). SIAM Review, 31 (4):699–700, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hinich:1991:BRS

[Hin91] Melvin J. Hinich. Book review: Statistical Spectral Analysis: a Nonprobabilistic Theory (William A. Gardner). SIAM Review, 33 (4):677–678, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hinton:1994:BRH

[Hin94]  Don Hinton.  Book re-

view: Handbook of Differential Equations, Second Edition (Daniel Zwillinger). SIAM Review, 36(1):126– 127, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hirschman:1971:BRF

[Hir71] I. I. Hirschman, Jr. Book review: Finite Sections of Some Classical Inequalities (Herbert S. Wilf). SIAM Review, 13(3):412–413, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hirschman:1973:BRI

[Hir73] I. I. Hirschman, Jr. Book review: An Introduction to Transform Theory (D. V. Widder). SIAM Review, 15 (2):396–397, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hirsh:1975:AFO

[Hir75] R. S. Hirsh. Application of a fourth order differencing technique to fluid mechanics problems. SIAM Review, 17 (2):381, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hirsh:1978:SSF

[Hir78] J. E. Hirsh. Solution for stream-function in an irregular domain using a fast direct method. SIAM Review, 20 (3):626–627, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Histand:1977:BRM

[His77] Michael B. Histand. Book review: Mathematical Biofluiddynamics (James Lighthill). SIAM Review, 19(3):576– 577, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hagstrom:1985:MV

[HJ85] T. M. Hagstrom and W. B.

Jordan. A minimum value.

SIAM Review, 27(2):256, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hulsbergen:1990:EEC

[HJ90] W. Hulsbergen and A. A. Jagers. Eigenvalues and eigenvectors of a combinatorial matrix. SIAM Review, 32 (1):148–149, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hall:2000:PMD

[HJ00] Rachel W. Hall and Kresimir Josic. Planetary motion and the duality of force laws. SIAM Review, 42 (1):115–124, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34600.

Holloway:1969:Q

[HK69] H. Holloway and M. S. Klamkin. A quartic. SIAM Review, 11(1):74–75, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holloway:1970:Q

[HK70] H. Holloway and M. S. Klamkin. A quartic. SIAM Review, 12(3):471–473, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hanan:1972:RPQ

[HK72] Maurice Hanan and Jerome M.

Kurtzberg. A review of the placement and quadratic assignment problems. SIAM Review, 14(2):324–342, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hajmirzaahmad:1992:SOO

[HK92a] M. Hajmirzaahmad and A. M. Krall. Singular 2ndorder operators — the maximal and minimal operators, and self-adjoint operators in between. SIAM Review, 34 (4):614–634, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hajmirzaahmad:1992:SSO

[HK92b] Mojdeh Hajmirzaahmad and Allan M. Krall. Singular second-order operators: The maximal and minimal operators, and selfadjoint operators in between. SIAM Review, 34(4):614– 634, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hettich:1993:SIP

[HK93] R. Hettich and K. O. Kortanek. Semi-infinite programming: Theory, methods, and applications. SIAM Review, 35(3):380–429, September 1993. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Holloway:1994:Q

[HK94] H. Holloway and M. S. Klamkin. A quartic. SIAM Review, 36(3):491–493, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heck:1995:IM

[HK95] Andre Heck and Wolfram Koepf. Introduction to Maple. SIAM Review, 37(1): 132–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hong:2020:GCP

[HKD20] David Hong, Tamara G. Kolda, and Jed A. Duersch. Generalized canonical polyadic tensor decomposition. SIAM Review, 62(1): 133–163, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heidel:1975:PIM

[HKS75] J. W. Heidel, A. J. Kretchmar, and S. M. Serbin. Parameter identification in a metabolic system. SIAM Review, 17(2):380, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hale:1963:DEL

[HL63] Jack K. Hale and Joseph P.

LaSalle. Differential equations: Linearity vs. nonlinearity. SIAM Review, 5(3): 249–272, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hangelbroek:1976:DHS

[HL76] R. J. Hangelbroek and C. G.

Lekkerkerker. Decompositions of a Hilbert space and factorization of a W-A determinant. SIAM Review, 18 (4):806, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Henkel:1986:CTI

[HL86] M. Henkel and J. Lacki.

Conjectured trigonometrical identities. SIAM Review, 28 (1):87–88, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoschek:1995:CAG

[HLB95] Josef Hoschek, Dieter Lasser, and Len Bos. Computer aided geometric design. SIAM Review, 37(3): 460–??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hartline:2003:CCM

[HLH03] Jeffrey R. Hartline and Ran Libeskind-Hadas. The computational complexity of motion planning. SIAM Review, 45(3):543–557, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/39517.

Henrion:2009:AVI

[HLS09] D. Henrion, J. B. Lasserre, and C. Savorgnan. Approximate volume and integration for basic semialgebraic sets. SIAM Review, 51(4): 722–743, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Harten:1983:UDG

[HLvL83] Amiram Harten, Peter D. Lax, and Bram van Leer. On upstream differencing and Godunov-type schemes for hyperbolic conservation laws. SIAM Review, 25(1): 35–61, January 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hsiao:1973:SBV

[HM73] George Hsiao and R. C. Mac-Camy. Solution of boundary value problems by integral equations of the first kind. SIAM Review, 15(4): 687–705, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hicks:1978:APO

[HM78] D. L. Hicks and M. M.

Madsen. Accuracy property for one-dimensional Lagrangean hydrodynamic difference schemes. SIAM Review, 20(3):626, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Highfill:2001:AOC

[HM01] Jannett Highfill and Michael McAsey. An application of optimal control to the economics of recycling. SIAM Review, 43(4):679–693, December 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/36269.

Hubert:2000:TPM

[HMH00] Lawrence Hubert, Jacqueline Meulman, and Willem Heiser. Two purposes for matrix factorization: a historical appraisal. SIAM Review, 42(1):68–82, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34048.

Helmke:1996:ODS

[HMK96] Uwe Helmke, John B. Moore, and Yuji Kodama. Optimization and dynamical systems. SIAM Review, 38(3):

531–??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Howison:1997:CCT

[HMO97] S. D. Howison, J. D. Morgan, and J. R. Ockendon. A class of codimensiontwo free boundary problems. SIAM Review, 39(2): 221–253, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28062.

Halko:2011:FSR

[HMT11] N. Halko, P. G. Martinsson, and J. A. Tropp. Finding structure with randomness: Probabilistic algorithms for constructing approximate matrix decompositions. SIAM Review, 53(2): 217–288, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i2/p217_s1.

Hnizdo:1989:RSU

[Hni89] V. Hnizdo. Radiation from a system of uniformly circling charges. SIAM Review, 31 (1):121–122, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hnizdo:1992:C

[Hni92] V. Hnizdo. Correction. SIAM Review, 34(1):126, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heath:1991:PAS

[HNP91] Michael T. Heath, Esmond Ng, and Barry W. Peyton. Parallel algorithms for sparse linear systems. SIAM Review, 33(3):420– 460, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1996:IPM

[HO96] M. H. Holmes and Robert E.

O’Malley, Jr. Introduction to perturbation methods. SIAM Review, 38(3): 538–??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1960:FCP

[Hoc60] Harry Hochstadt. On the factorization of certain polynomials. SIAM Review, 2(2): 140–147, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1965:BRP

[Hoc65]  Harry Hochstadt.  Book

review: Periodic Differential Equations (F. M. Arscott). SIAM Review, 7(3):

439–440, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1966:ECI

[Hoc66] Harry Hochstadt. On the eigenvalues of a class of integral equations arising in laser theory. SIAM Review, 8(1): 62–65, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1967:MCF

[Hoc67]  Harry Hochstadt.  The

mean convergence of Fourier– Bessel series. SIAM Review, 9 (2):211–218, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1970:PCS

[Hoc70] Harry Hochstadt. On perturbations of complete sets. SIAM Review, 12(1):103– 106, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1979:BRD

[Hoc79a]  Harry Hochstadt.  Book

review: Differential Equations and Their Applications, 2nd edition (Martin Braun). SIAM Review, 21 (2):264–266, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1979:BRM

[Hoc79b] Harry Hochstadt. Book review: Mathematics Applied to Continuum Mechanics (Lee A. Segel). SIAM Review, 21(3):414–415, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1980:BRC

[Hoc80] Harry Hochstadt. Book review: Complex Analysis: an Introduction to the Theory of Analytic Functions of One Complex Variable; 3rd ed. (Lars V. Ahlfors). SIAM Review, 22(3):378, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1982:BRA

[Hoc82]  Harry Hochstadt.  Book

review: Arithmetic Applied Mathematics (Donald Greenspan). SIAM Review, 24(1):100, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1983:GFT

[Hoc83a] H. Hochstadt. Green functions and transfer functions handbook — Butkovskiy, A. G. SIAM Review, 25(3): 435–436, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1983:BRG

[Hoc83b] Harry Hochstadt. Book review: Green’S Functions and Transfer Functions Handbook (Anatoliy G. Butkovskiy). SIAM Review, 25(3):435– 436, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1986:BRD

[Hoc86] Harry Hochstadt. Book review: Differential Equations (A. N. Tikhonov, A. B. Vasil’eva and A. G. Sveshnikov). SIAM Review, 28(1): 105–106, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochstadt:1992:BRI

[Hoc92]  Harry Hochstadt.  Book

review: Integral Equations, A Practical Treatment, From Spectral Theory to Applications (David Porter and David S. G. Stirling). SIAM Review, 34(3): 509–511, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hockett:1994:BRQ

[Hoc94] Kevin Hockett. Book review: Qualitative Theory of Dynamical Systems (L. Dingjun and T. Libang). SIAM Review, 36(4):671– 672, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochbruck:2023:SRa

[Hoc23a]  Marlis Hochbruck.  Sur-

vey and review. SIAM Review, 65(1):1, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N975612.

Hochbruck:2023:SRb

[Hoc23b]  Marlis Hochbruck.  Sur-

vey and review. SIAM Review, 65(2):329, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N975673.

Hochbruck:2023:SRc

[Hoc23c]  Marlis Hochbruck.  Sur-

vey and review. SIAM Review, 65(3):599, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N975727.

Hochbruck:2023:SR

[Hoc23d] Marlis Hochbruck. Survey and review. SIAM Review, 65 (4):917, November 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochbruck:2024:SRa

[Hoc24a] Marlis Hochbruck. Survey and review. SIAM Review, 66(1):1, February 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hochbruck:2024:SRb

[Hoc24b] Marlis Hochbruck. Survey and review. SIAM Review, 66(2):203, May 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoffman:1971:VIA

[Hof71] William C. Hoffman. Visual illusions of angle as an application of Lie transformation groups. SIAM Review, 13(2): 169–184, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoffman:1985:BRA

[Hof85] Dale T. Hoffman. Book review: Approval Voting (Steven J. Brams and Peter C. Fishburn). SIAM Review, 27(2):276–278, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoffmann:1995:BRC

[Hof95] Christoph M. Hoffmann. Book review: Computational Geometry in C (J. O’Rourke). SIAM Review, 37 (1):128–129, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hogan:1973:PSM

[Hog73] William W. Hogan. Pointto-set maps in mathematical programming. SIAM Review, 15(3):591–603, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoggatt:1974:BZP

[Hog74] V. E. Hoggatt, Jr. Bounds for the zero of a polynomial. SIAM Review, 16(2): 258–259, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoggatt:1975:BZP

[Hog75] V. E. Hoggatt. Bounds for zero of a polynomial. SIAM Review, 17(4):690– 691, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holt:1966:BRT

[Hol66] Maurice Holt. Book review:

Two-Dimensional Problems in Hydrodynamics and Aerodynamics (L. I. Sedov). SIAM Review, 8(3):404– 405, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1970:BRI

[Hol70a] Richard Holmes. Book review: An Introduction to the Approximation of Functions (Theodore J. Rivlin). SIAM Review, 12(2):311– 312, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holtzman:1970:EIF

[Hol70b] J. M. Holtzman. Explicit ε and δ for the implicit function theorem. SIAM Review, 12 (2):284–286, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holt:1975:BRI

[Hol75] A. R. Holt. Book review: An Integral Inequality (R. Askey). SIAM Review, 17(4): 691–693, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hollingsworth:1978:P

[Hol78] B. J. Hollingsworth. Problem 76-17. SIAM Review, 20 (4):856, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1979:MFS

[Hol79] Richard B. Holmes. Mathematical foundations of signal processing. SIAM Review, 21 (3):361–388, July 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holland:1981:SDA

[Hol81] A. S. B. Holland. A survey of degree of approximation of continuous functions. SIAM Review, 23(3): 344–379, July 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1985:BRL

[Hol85] Philip Holmes. Book review: The Lorenz Equations: Bifurcations, Chaos, and Strange Attractors (Colin Sparrow). SIAM Review, 27 (1):106–110, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1986:BRP

[Hol86a] Mark Holmes. Book review: Problems in Perturbation (Ali Hasan Nayfeh). SIAM Review, 28(3):420– 421, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1986:BRC

[Hol86b] Philip Holmes. Book review: Chaos (Hao Bai-Lin). SIAM Review, 28(1):106– 109, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1987:BRI

[Hol87] Philip Holmes. Book review: An Introduction to Chaotic Dynamical Systems (Robert L. Devaney); Nonlinear Dynamics and Chaos: Geometrical Methods for Engineers and Scientists (J. M. T. Thompson and H. B. Stewart) (with the Assistance of R. Ghaffari and C. Franciosi and a Contribution by H. L. Swinney). SIAM Review, 29 (4):654–658, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1989:BRMa

[Hol89a] Mark H. Holmes. Book review: Mechanoreceptors, Their Functional Organization (G. N. Akoev, N. P. Alekseev, and B. V. Krylov). SIAM Review, 31(2):344– 346, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1989:BRMb

[Hol89b] Mark H. Holmes. Book review: The Mechanics of Vibrations of Cylindrical Shells (Stefan Markus). SIAM Review, 31(2):346– 347, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holz:1990:BRC

[Hol90a] K.-Peter Holz. Book review:

Computational Hydraulics: an Introduction (Cornelis B. Vreugdenhil). SIAM Review, 32(4):697–698, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holz:1990:CHI

[Hol90b] K.-Peter Holz. Computational hydraulics: An introduction (Cornelis B Vreugdenhil). SIAM Review, 32(4): 697–??, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hollman:1991:EMO

[Hol91a] H. O. L. Hollman. Elements of maximum order in a matrix group. SIAM Review, 33 (3):479–481, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hollmann:1991:BRE

[Hol91b] Henk D. L. Hollmann. Book review: Elements of Maximum Order in a Matrix Group (Lih-Yuan Deng). SIAM Review, 33(3):479– 481, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmes:1995:BRE

[Hol95] Philip Holmes. Book review:

The Essence of Chaos (E. N.

Lorenz). SIAM Review, 37 (1):129–131, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Homles:1986:BRS

[Hom86] F. A. Homles. Book review: Singular-Perturbation Theory: an Introduction with Applications (Donald R. Smith). SIAM Review, 28(3): 421–423, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoppensteadt:1975:ASP

[Hop75] Frank Hoppensteadt. Analysis of some problems having matched asymptotic expansion solutions. SIAM Review, 17(1):123–135, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoppensteadt:1979:BRT

[Hop79] F. C. Hoppensteadt. Book review: Theoretical Immunology (George I. Bell, Alan S. Perelson and George H. Pimbley, Jr., eds.). SIAM Review, 21(4):577–578, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoppensteadt:1983:BRM

[Hop83] Frank Hoppensteadt. Book review: Mathematical Modelling in Epidemiology (J. C. Frauenthal). SIAM Review, 25(1):131, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoppensteadt:1992:SPM

[Hop92] F. C. Hoppensteadt. Signal processing by model neural networks. SIAM Review, 34 (3):426–444, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Horn:1978:ERK

[Hor78] M. K. Horn. Extensions of Runge–Kutta algorithm to treat rejected steps and data output efficiently. SIAM Review, 20(3):627, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Horgan:1983:BRI

[Hor83a] C. O. Horgan. Book review:

An Introduction to the Theory of Elasticity (R. J. Atkin and N. Fox). SIAM Review, 25(3):437–438, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Horgan:1983:BRC

[Hor83b] C. O. Horgan. Book review:

Continuum Mechanics (A. J. M. Spencer). SIAM Review, 25(3):436–437, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Horgan:1989:BRS

[Hor89] Cornelius O. Horgan. Book review: Saint-Venant’s Problem (Dorin lesan). SIAM Review, 31(1):139–141, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Horn:1993:BRF

[Hor93] Paul S. Horn. Book review: A First Course in Order Statistics (Barry C. Arnold, N. Balakrishnan, and H. N. Nagaraja). SIAM Review, 35(3): 525, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Horgan:1995:APS

[Hor95a] C. O. Horgan. Anti-plane shear deformations in linear and nonlinear solid mechanics. SIAM Review, 37(1): 53–81, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/26520.htm.

Horgan:1995:KIT

[Hor95b] C. O. Horgan. Korn’s inequalities and their applications in continuum mechanics. SIAM Review, 37(4):491– 511, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Horgan:1997:BRQ

[Hor97a] Cornelius O. Horgan. Book review: Qualitative Estimates for Partial Differential Equations: An Introduction, by James N. Flavin and Salvatore Rionero. SIAM Review, 39(2): 347–350, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Horn:1997:DTB

[Hor97b] R. A. Horn. The determinant of a triangular-block matrix. SIAM Review, 39(3): 517–520, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1965:KSR

[Hou65] A. S. Householder. The Kantorovich and some related inequalities. SIAM Review, 7 (4):463–473, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1967:BRS

[Hou67] A. S. Householder. Book review: Solution of Equations and Systems of Equations (A. M. Ostrowski). SIAM Review, 9(3):608– 609, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1968:BPR

[Hou68a]  A. S. Householder.  Bi-

gradients and the problem of Routh and Hurwitz. SIAM Review, 10(1): 56–66, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Hou68b].

Householder:1968:EBP

[Hou68b] A. S. Householder. Erratum: Bigradients and the Problem of Routh and Hurwitz. SIAM Review, 10(4):438, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Hou68a].

Householder:1970:BEL

[Hou70a]  A. S. Householder.  Be-

zoutiants, elimination and localization. SIAM Review, 12 (1):73–78, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1970:BRD

[Hou70b] A. S. Householder. Book review: A Determinant (J. Prasad). SIAM Review, 12 (2):302–304, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1971:BRD

[Hou71a] A. S. Householder. Book review: A Determinant (J. Prasad). SIAM Review, 13 (3):395, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1971:B

[Hou71b] Altson S. Householder. Brigadients. SIAM Review, 13(1): 114–115, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1974:BES

[Hou74a] A. S. Householder. Bigradients and the Euclid–Sturm algorithm. SIAM Review, 16 (2):207–213, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1974:G

[Hou74b] A. S. Householder. The gatlinburgs. SIAM Review, 16(3):340–343, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1974:STH

[Hou74c] A. S. Householder. Schröder and Trudi: a historical excursion. SIAM Review, 16(3): 344–348, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hough:1978:WWW

[Hou78] D. Hough. Whats wrong with Wilkinson polynomial. SIAM Review, 20(3):627, ???? 1978.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hou:1998:SPL

[Hou98a] S. H. Hou. A simple proof of the Leverrier–Faddeev characteristic polynomial algorithm. SIAM Review, 40(3): 706–709, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hou:1998:CNS

[Hou98b] Shui-Hung Hou. Classroom note: a simple proof of the Leverrier–Faddeev characteristic polynomial algorithm. SIAM Review, 40 (3):706–709, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32076.

Howe:1963:BRE

[How63] Wm. G. Howe. Book review: An Elementary Introduction to the Theory of Probability (B. V. Gnedenko and A. Ya. Khinchin). SIAM Review, 5 (4):383, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Howes:1982:BRI

[How82] F. A. Howes. Book review: Introduction to Perturbation Techniques (Ali Hasan Nayfeh). SIAM Review, 24 (3):355–356, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Howes:1983:BRR

[How83] F. A. Howes. Book review: Random Differential Inequalities (G. S. Ladde and V. Lakshmikantham). SIAM Review, 25(1):122– 123, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Howison:1994:BRR

[How94] R. Clarissa Howison. Book review: Russian–English, English–Russian Dictionary on Probability, Statistics, and Combinatorics (K. A. Borovkov). SIAM Review, 36 (4):672, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hansen:1976:EMR

[HP76] E. Hansen and M. Patrick.

Estimating multiplicity of a root. SIAM Review, 18 (4):807, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Halphen:1985:MTT

[HP85] B. Halphen and S. L. Paverifontana. On the motion of a trailer-truck. SIAM Review, 27(4):578–579, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Halphen:1985:BRM

[HPF85] B. Halphen and S. L. Paveri-Fontana. Book review: On the Motion of a Trailer-Truck (J. C. Alexander). SIAM Review, 27(4):578– 579, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Homescu:2007:EER

[HPS07] Chris Homescu, Linda R. Petzold, and Radu Serban. Error estimation for reducedorder models of dynamical systems. SIAM Review, 49 (2):277–299, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hideout:1962:BRS

[HR62] V. C. Hideout and V. Rajaraman. Book review: Statistical Theory of Communication (Y. W. Lee). SIAM Review, 4 (4):402–404, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hornor:1985:BRD

[HR85] W. E. Hornor and C. C. Rousseau. Book review: A Definite Integral (B. C. Berndt). SIAM Review, 27 (1):84, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hartsfield:1996:PGT

[HRB96] Nora Hartsfield, Gerhard Ringel, and Lowell W. Beineke. Pearls in graph theory: a comprehensive introduction. SIAM Review, 38 (1):159–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoyland:1996:SRT

[HRM96] Arnljot Hoyland, Marvin Rausand, and William Q. Meeker. System reliability theory: Models and statistics methods. SIAM Review, 38 (1):175–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Humpherys:2012:FLK

[HRW12] Jeffrey Humpherys, Preston Redd, and Jeremy West. A fresh look at the Kalman filter. SIAM Review, 54(4): 801–823, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hardin:1970:MTA

[HS70] J. C. Hardin and A. L.

Sweet. Moments of the time to absorption in the random walk between a reflecting and an absorbing barrier. SIAM Review, 12(1):

140–142, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Householder:1971:NFP

[HS71] A. S. Householder and G. W. Stewart. The numerical factorization of a polynomial. SIAM Review, 13(1): 38–46, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hindmarsh:1972:BRF

[HS72] A. C. Hindmarsh and L. A. Shepp. Book review: A Functional Equation (L. A. Shepp). SIAM Review, 14(1): 169–170, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Henderson:1981:DIS

[HS81] H. V. Henderson and S. R. Searle. On deriving the inverse of a sum of matrices. SIAM Review, 23(1): 53–60, January 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hua:1988:CMM

[HS88] Y. Hua and T. K. Sarkar. Conditions for maximizing and minimizing diagonal elements of a Hermitian matrix. SIAM Review, 30(4): 650–651, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holmgren:1995:FCD

[HS95] R. A. Holmgren and James T.

Sandefur. A first course in discrete dynamical systems. SIAM Review, 37(1): 137–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Heijmans:1996:MIO

[HS96] Henk J. A. M. Heijmans and Jean Serra. Morphological image operators. SIAM Review, 38(1):178– ??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hoffman:2003:SMS

[HS03] Kathleen Hoffman and Fadil Santosa. A simple model of sheet metal assembly. SIAM Review, 45(3):558– 573, September 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38230.

Holcman:2014:NEP

[HS14] D. Holcman and Z. Schuss.

The narrow escape problem. SIAM Review, 56(2): 213–257, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Howison:2015:RNP

[HS15] Sam Howison and Daniel Schwarz. Risk-neutral pricing of financial instruments in emission markets: a structural approach. SIAM Review, 57(1):95–127, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Higham:2016:RDS

[HSŠ16] Nicholas J. Higham, Nataša Strabić, and Vedran Šego. Restoring definiteness via shrinking, with an application to correlation matrices with a fixed block. SIAM Review, 58(2):245– 263, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Howle:1998:LTK

[HSSZ98] Laurens Howle, David G. Schaeffer, Michael Shearer, and Pei Zhong. Lithotripsy: The treatment of kidney stones with shock waves. SIAM Review, 40(2):356– 371, June 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32263.

Hartl:1995:SMP

[HSV95] Richard F. Hartl, Suresh P. Sethi, and Raymond G. Vickson. A survey of the maximum principles for optimal control problems with state constraints. SIAM Review, 37 (2):181–218, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/23508.htm.

Hauer:1972:MVF

[HT72] E. Hauer and J. G. C. Templeton. Mean and variance of a function of a random permutation. SIAM Review, 14 (4):645, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hanson:1978:ADD

[HT78] F. B. Hanson and H. C.

Tuckwell. Analysis of a differential-delay equation, singular at both endpoints. SIAM Review, 20 (3):626, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Howard:1981:BRE

[HT81] L. N. Howard and M. Tortorella. Book review: An Extremal Sphere (D. Singmaster). SIAM Review, 23(3): 394–395, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Holtz:2012:SMC

[HT12] Olga Holtz and Mikhail Tyaglov. Structured matrices, continued fractions, and root localization of polynomials. SIAM Review, 54(3): 421–509, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hu:1968:RAN

[Hu68] T. C. Hu. Recent advances in network flows. SIAM Review, 10(3):354–359, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hiriart-Urruty:2007:PCO

[HU07] Jean-Baptiste Hiriart-Urruty. Potpourri of conjectures and open questions in nonlinear analysis and optimization. SIAM Review, 49(2): 255–273, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Huang:1986:WPA

[Hua86] J. S. Huang. Worst polynomial approximation to f(x) = x. SIAM Review, 28(3):397, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Huang:1996:BRM

[Hua96] Weizhang Huang. Book review: Moving Finite Elements (M. J. Baines). SIAM Review, 38(3):530– 531, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hubbell:1983:RS

[Hub83] J. H. Hubbell. A rectangular source. SIAM Review, 25 (2):267, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hubner:1988:BRM

[Hüb88] Gerhard Hübner. Book review: Markovian Control Problems: Functional Equations and Algorithms (A. Federgruen). SIAM Review, 30(4):670–671, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hubner:1991:BRA

[Hüb91] Gerhard Hübner. Book review: Adaptive Markov Control Processes (D. Hernandez-Lerma). SIAM Review, 33 (4):686–687, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hudson:1995:BRO

[Hud95] W. N. Hudson. Book review: Operator-Limit Distributions in Probability Theory (Z. J. Jurek and J. D. Mason). SIAM Review, 37(1): 131–132, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Huestis:1991:SCO

[Hue91] Stephen P. Huestis. Some comments on the optimization problem for shear stress magnitude. SIAM Review, 33(1):106–108, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Huestis:1992:IFO

[Hue92a] Stephen P. Huestis. Interpolation formulas for oversampled band-limited functions. SIAM Review, 34 (3):477–481, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Huestis:1992:SNS

[Hue92b] Stephen P. Huestis. Smallest nonnegative solutions to linear inverse problems. SIAM Review, 34(4):642– 647, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Huestis:1996:ULP

[Hue96] Stephen P. Huestis. The use of linear programming in the construction of extremal solutions to linear inverse problems. SIAM Review, 38(3): 496–506, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28895.

Hughes:1993:BRO

[Hug93a]  John F. Hughes.  Book

review: Oriented Projective Geometry: a Framework for Geometric Computations (Jorge Stolfi). SIAM Review, 35(2):321–323, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hughes:1993:OPG

[Hug93b]  John F. Hughes.  Ori-

ented projective geometry: a framework for geometric computations (Jorge Stolfi). SIAM Review, 35(2):321– ??, June 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hui:1995:IBE

[Hui95] Ping Hui. Integrals of Bessel and elliptic functions from heat conduction. SIAM Review, 37(3):440, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hui:1996:IHC

[Hui96] Ping Hui. Integrals from heat conduction. SIAM Review, 38 (4):668–669, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hui:1997:IHC

[Hui97] P. Hui. Integrals from heat conduction. SIAM Review, 39 (4):775–779, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hull:1961:BRD

[Hul61a] T. E. Hull. Book review: Digital Computers and Nuclear Reactor Calculations (Ward C. Sangren). SIAM Review, 3(1):78–80, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hull:1961:BRM

[Hul61b]  T. E. Hull.  Book re-

view: Mathematical Methods for Digital Computers (Anthony Ralston and Herbert S. Wilf, eds.). SIAM Review, 3 (2):181–182, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hull:1967:SOM

[Hul67]  T. E. Hull.  A search

for optimum methods for the numerical integration of ordinary differential equations. SIAM Review, 9(4): 647–654, October 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hull:1971:BRC

[Hul71]  T. E. Hull.  Book re-

view: Computation and Theory in Ordinary Differential Equations (James W. Daniel and Ramon E.

Moore). SIAM Review, 13 (3):413–414, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hulme:1976:GTM

[Hul76] B. L. Hulme. Graph theoretic models of theft problems. SIAM Review, 18 (4):809, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hulsbergen:1989:EEC

[Hul89] W. Hulsbergen. Eigenvalues and eigenvectors of a combinatorial matrix. SIAM Review, 31(1):120–121, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hunter:1976:IES

[Hun76] C. Hunter. Integral equation solvable by use of a Stieltjes transform. SIAM Review, 18 (4):809–810, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hunt:1981:BRI

[Hun81] B. R. Hunt. Book review:

Image Reconstruction from Projections: Implementation and Applications (G. T. Herman, Editor). SIAM Review, 23(3):399–400, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hunt:1984:BRP

[Hun84] B. R. Hunt. Book review: Pattern Models (Narenda Ahuja and Bruce J. Schachter). SIAM Review, 26(4):592– 593, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hunt:1985:BRS

[Hun85] B. R. Hunt. Book review: Self-Organization and Associative Memory (Teuvo Kohonen). SIAM Review, 27(2): 278–279, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hurd:1965:BRS

[Hur65] R. A. Hurd. Book review: A Set of Linear Equations (M. J. Pascual and J. E. Zweig). SIAM Review, 7(3): 424–428, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hurley:1993:BRG

[Hur93a]  Mike Hurley.  Book re-

view: The Global Dynamics of Cellular Automata (Andrew Wuensche and Mike Lesser). SIAM Review, 35(4): 669–670, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hurley:1993:GDC

[Hur93b] Mike Hurley. The global dynamics of cellular automata (Andrew Wuensche and Mike Lesser). SIAM Review, 35(4): 669–??, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hiriart-Urruty:2010:VAC

[HUS10] J.-B. Hiriart-Urruty and A. Seeger. A variational approach to copositive matrices. SIAM Review, 52(4): 593–629, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hutchinson:1989:BRA

[Hut89] John W. Hutchinson. Book review: Applied Elasticity: Matrix and Tensor Analysis of Elastic Continua (J. D. Renton). SIAM Review, 31 (1):141–142, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hutter:1991:BRE

[Hut91]  K. Hutter.  Book re-

view: Electrodynamics of Continua (A. Cemal Eringen and Gerard A. Maugin). SIAM Review, 33(2): 315–320, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haussling:1975:NSL

[HV75] H. J. Haussling and R. T. Vaneseltine. Numerical solution of linear and nonlinear water wave problems. SIAM Review, 17(2):380, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Horgan:1976:EDE

[HW76] C. O. Horgan and L. T. Wheeler. Exponential decay estimates for 2nd order quasilinear elliptic equations. SIAM Review, 18(4): 808–809, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Haaland:1977:DCA

[HW77] C. M. Haaland and E. P. Wigner. Defense of cities by anti-ballistic missiles. SIAM Review, 19(2):279– 296, April 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2029503.

Heil:1989:CDW

[HW89] Christopher E. Heil and David F. Walnut. Continuous and discrete wavelet transforms. SIAM Review, 31 (4):628–666, December 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Halanay:1971:SNR

[HY71] A. Halanay and James A.

Yorke. Some new results and problems in the theory of differential-delay equations. SIAM Review, 13(1): 55–80, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hyde:1986:FPI

[Hyd86] Michael F. Hyde. Financial planning in an inflationary economy. SIAM Review, 28 (4):551–564, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hyers:1973:BRF

[Hye73] D. H. Hyers. Book review: The Functions of Mathematical Physics (Harry Hochstadt). SIAM Review, 15(1):235–236, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Hayes:1976:UAS

[HYS76] L. Hayes, D. M. Young, and E. Schleicher. Use of accelerated SSOR method to solve large linear systems. SIAM Review, 18(4):808, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ito:1975:CMT

[IA75] T. Ito and J. H. Ahlberg.

Collocation method for twopoint boundary-value problems. SIAM Review, 17 (2):382, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ibaraki:1973:AOS

[Iba73] Toshihide Ibaraki. Algorithms for obtaining shortest paths visiting specified nodes. SIAM Review, 15(2): 309–317, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ierley:1996:CEO

[Ier96] Glenn R. Ierley. A combinatorial expression from oceanography. SIAM Review, 38(4):668, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ingle:1990:PE

[IGWL90] R. Ingle, A. Genz, P. Wollan, and T. W. Lockhart. Probabilistic ethics. SIAM Review, 32(4):686–688, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ikebe:1972:GMN

[Ike72]  Yasuhiko Ikebe.  The

Galerkin method for the numerical solution of Fredholm integral equations of the second kind. SIAM Review, 14 (3):465–491, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ikebe:1975:NCC

[Ike75] Y. Ikebe. Numerical computation of characteristic values of mathieus equation. SIAM Review, 17(2):382, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ikebe:1976:CZB

[Ike76] Y. Ikebe. Computing zeros of Bessel and regular Coulomb wave functions and of their derivatives by matrix theoretic approach — practical accuracy criteria. SIAM Review, 18(4):810, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ikebe:1978:PTM

[Ike78] Y. Ikebe. Perturbation theorems for matrix eigenvalues. SIAM Review, 20(3): 627, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Iserlas:1997:FCN

[IL97] Arish Iserlas and Raytcho Lazarov. A first course in the numerical analysis of differential equations. SIAM Review, 39(3):541– ??, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ikonic:1988:ESD

[IM88] Z. Ikonic and V. Milanovic. Exact solutions of a differential equation. SIAM Review, 30(2):308, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ikonic:1989:ESD

[IM89] Z. Ikonic and V. Milanovic. Exact solutions of a differential equation. SIAM Review, 31(2):324–325, June 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Innis:1969:EBS

[Inn69]  George S. Innis.  Exis-

tence of binary sequences with prescribed properties. SIAM Review, 11(4):621– 622, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Inselberg:1978:CDE

[Ins78]  A. Inselberg.  Cochlear

dynamics: The evolution of a mathematical model. SIAM Review, 20(2):301– 351, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:1997:CEI

[Ips97] Ilse C. F. Ipsen. Computing an eigenvector with inverse iteration. SIAM Review, 39(2):254–291, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30077.

Ipsen:2005:PT

[Ips05] Ilse Ipsen. Problems and techniques. SIAM Review, 47(4):707, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97156.

Ipsen:2006:PTa

[Ips06a] Ilse Ipsen. Problems and techniques. SIAM Review, 48 (1):41–42, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2006:PTb

[Ips06b] Ilse Ipsen. Problems and techniques. SIAM Review, 48 (2):305, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2006:PTc

[Ips06c] Ilse Ipsen. Problems and techniques. SIAM Review, 48 (3):485, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2006:PTd

[Ips06d] Ilse Ipsen. Problems and techniques. SIAM Review, 48 (4):679–680, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:PTa

[Ips07a] Ilse Ipsen. Problems and techniques. SIAM Review, 49 (1):33–34, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:PTb

[Ips07b] Ilse Ipsen. Problems and techniques. SIAM Review, 49 (2):209, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:PTc

[Ips07c] Ilse Ipsen. Problems and techniques. SIAM Review, 49 (3):419–420, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2007:PTd

[Ips07d] Ilse Ipsen. Problems and techniques. SIAM Review, 49 (4):593, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:PTa

[Ips08a] Ilse Ipsen. Problems and techniques. SIAM Review, 50(1):35, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:PTb

[Ips08b] Ilse Ipsen. Problems and techniques. SIAM Review, 50 (4):721, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:PTIa

[Ips08c] Ilse Ipsen. Problems and techniques — introduction. SIAM Review, 50(2):273, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2008:PTIb

[Ips08d] Ilse Ipsen. Problems and techniques — introduction. SIAM Review, 50(3):483, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2009:PTa

[Ips09a] Ilse Ipsen. Problems and techniques. SIAM Review, 51 (1):127, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2009:PTb

[Ips09b] Ilse Ipsen. Problems and techniques. SIAM Review, 51 (2):337, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2009:PTc

[Ips09c] Ilse Ipsen. Problems and techniques. SIAM Review, 51 (3):543, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2009:PTI

[Ips09d] Ilse Ipsen. Problems and techniques: Introduction. SIAM Review, 51(4):705, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2010:ERPa

[Ips10a] Ilse Ipsen. Expository research papers. SIAM Review, 52(1):55, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2010:ERPb

[Ips10b] Ilse Ipsen. Expository research papers. SIAM Review, 52(2):267, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2010:ERPc

[Ips10c] Ilse Ipsen. Expository research papers. SIAM Review, 52(3):453, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2010:ERPd

[Ips10d] Ilse Ipsen. Expository research papers. SIAM Review, 52(4):675, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ipsen:2011:ERPa

[Ips11a] Ilse Ipsen. Expository research papers. SIAM Review, 53(1):69, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i1/p69_s1.

Ipsen:2011:ERPb

[Ips11b] Ilse Ipsen. Expository research papers. SIAM Review, 53(2):289, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i2/p289_s1.

Ipsen:2011:ERPc

[Ips11c] Ilse Ipsen. Expository research papers. SIAM Review, 53(3):503, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v53/i3/p503_s1.

Ipsen:2011:ERPd

[Ips11d] Ilse Ipsen. Expository research papers. SIAM Review, 53(4):721, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v53/i4/p721_s1.

Ipsen:2012:RS

[Ips12a] Ilse Ipsen. Research spotlights. SIAM Review, 54 (1):119, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v54/i1/p119_s1.

Ipsen:2012:U

[Ips12b] Ilse Ipsen. Untitled. SIAM Review, 54(1):119–120, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ierley:1997:CEO

[IR97] G. R. Ierley and O. G. Ruehr. A combinatorial expression from oceanography. SIAM Review, 39(4): 772–774, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Isaacson:1988:BRN

[Isa88] Eugene Isaacson. Book review: Numerical Recipes: The Art of Scientific Computing (William H. Press, Brian P. Flannery, Saul A. Teukolsky, and William T. Vetterling). SIAM Review, 30 (2):331–332, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Isaacson:1989:BRNa

[Isa89a] Eugene Isaacson. Book review: Numerical Recipes in C: The Art of Scientific Computing (William H. Press, Brian P. Flannery, Saul A. Teukolsky, and William T. Vetterling); Numerical Recipes: Example Book (C) (William T. Vetterling, Saul A. Teukolsky, William H. Press, and Brian P. Flannery). SIAM Review, 31 (1):142, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Isaacson:1989:BRNb

[Isa89b] Eugene Isaacson. Book review: Numerical Solution of Boundary Value Problems for Ordinary Differential Equations (Uri M. Ascher, Robert M. M. Mattheij, and Robert D. Russell). SIAM Review, 31 (4):700–701, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Isbell:1968:P

[Isb68] John R. Isbell. On playing M_{3}. SIAM Review, 10(4): 442–443, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ismail:1976:BRP

[Ism76] M. E. H. Ismail. Book review: A Power Series Expansion (U. G. Haussmann). SIAM Review, 18(2): 301–303, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ismail:1978:BRG

[Ism78] M. E. H. Ismail. Book review: A Gaussian Hypergeometric Function Identity (P. W. Karlsson). SIAM Review, 20(1):187–188, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ismail:1983:SFQ

[Ism83a] M. Ismail. Special functions in queuing theory — Srivastava, H. M., Kashyap, R. K. SIAM Review, 25(4): 582–585, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ismail:1983:BRS

[Ism83b] Mourad Ismail. Book review: Special Functions in Queuing Theory (H. M. Srivastava and R. K. Kashya). SIAM Review, 25(4):582– 585, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ismail:1985:BRH

[Ism85] Mourad E. H. Ismail. Book review: q-Hypergeometric Functions and Applications (H. Exton). SIAM Review, 27 (2):279–281, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ismail:1996:IAC

[Ism96a] M. E. H. Ismail. An integral arising in computing the energy of crystals. SIAM Review, 38(2):324, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ismail:1996:BRI

[Ism96b] Mourad E. H. Ismail. Book review: An Integral Arising in Computing the Energy of Crystals (M. L. Glassar). SIAM Review, 38(2): 324, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Israel:1992:EMO

[Isr92a] R. B. Israel. Even minus odd involutions in the symmetrical group. SIAM Review, 34 (2):315–317, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Israel:1992:BRE

[Isr92b] Robert B. Israel. Book review: Even Minus Odd Involutions in the Symmetric Group (Frank Schmidt and Rodica Simion). SIAM Review, 34(2):315–317, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ivancevic:2004:SRG

[Iva04] V. Ivancevic. Symplectic rotational geometry in human biomechanics. SIAM Review, 46(3):455–474, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

                                        http://epubs.siam.org/
                                        sam-bin/dbq/article/41313.

Iviarchani:1965:BRM

[Ivi65] E. W. Iviarchani. Book review: Mathematical Theory of Optics (R. K. Luneburg). SIAM Review, 7(3): 433–435, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Iyaniwura:2021:OMF

[IWMW21] Sarafa A. Iyaniwura, Tony Wong, Colin B. Macdonald, and Michael J. Ward. Optimization of the mean first passage time in neardisk and elliptical domains in 2-D with small absorbing traps. SIAM Review, 63(3): 525–555, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Izhikevich:2001:SEB

[Izh01] Eugene M. Izhikevich. Synchronization of elliptic bursters. SIAM Review, 43(2):315– 344, June 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38206.

Jordan:1980:BRD

[JA80] William B. Jordan and E. L. Allgower. Book review: A Determinant (A. E. Barkauskas and D. W. Bange). SIAM Review, 22 (1):100–101, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jackson:1969:CEI

[Jac69] L. W. Jackson. A comparison of ellipsoidal and interval arithmetic error bounds, numerical solutions of nonlinear problems (notice). SIAM Review, 11(??):??, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jackson:1976:BRN

[Jac76] L. W. Jackson. Book review: Numerical Methods (Germund Dahlquist and Åke Björck). SIAM Review, 18 (4):774–776, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jackson:1997:BRP

[Jac97] Ken Jackson. Book review: Parallel and Sequential Methods for Ordinary Differential Equations, by Kevin Burrage. SIAM Review, 39(2): 349–350, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jacobsen:2007:FP

[Jac07] Jon Jacobsen. As flat as possible. SIAM Review, 49(3): 491–507, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jaffe:1984:OUR

[Jaf84] Arthur Jaffe. Ordering the universe — the role of mathematics. SIAM Review, 26(4): 473–500, October 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagerman:1964:BRI

[Jag64] David L. Jagerman. Book review: Introduction to Nonlinear Differential and Integral Equations (Harold T.

Davis). SIAM Review, 6(1): 76–77, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1971:BRM

[Jag71] A. A. Jagers. Book review:

A Minimization for Markov Chains (Paul J. Schweitzer). SIAM Review, 13(4):573– 574, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1972:BRA

[Jag72a] A. A. Jagers. Book review:

An Antiderivative (Anthony J. Strecok). SIAM Review, 14 (1):172–173, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1972:BRI

[Jag72b] A. A. Jagers. Book review: An Inequality (Joel Brenner). SIAM Review, 14 (2):371–372, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1972:BRFa

[Jag72c] A. A. Jagers. Book review:

Factorization of a Determinant (T. S. Chow and N. A. Lindberger). SIAM Review, 14(2):367–368, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1972:BRN

[Jag72d] A. A. Jagers. Book review: Noncomplete Residue System (Robert Spira). SIAM Review, 14(1):173–174, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1973:BRB

[Jag73] A. A. Jagers. Book review: A Binomial Expansion Identity (Barry Wolk). SIAM Review, 15(4):798–799, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1974:BRS

[Jag74a] A. A. Jagers. Book review: A Sine Inequality (R. A. Askey). SIAM Review, 16 (4):550–553, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1974:BRI

[Jag74b] A. A. Jagers. Book review: An Integral Inequality (Richard Askey). SIAM Review, 16(2):262–263, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1976:BRI

[Jag76a] A. A. Jagers. Book review: Inequalities for the Gamma Function (D. K. Ross). SIAM Review, 18(3):

496–497, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1976:MPM

[Jag76b] A. A. Jagers. A monotonicity property for moments. SIAM Review, 18(4):762, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1979:BRE

[Jag79a] A. A. Jagers. Book review:

Evaluation of a Determinant (D. Slepian). SIAM Review, 21(3):400–401, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1979:BRT

[Jag79b] A. A. Jagers. Book review:

Two Equal Determinants (R. Shantaram). SIAM Review, 21(3):401, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1980:BRC

[Jag80] A. A. Jagers. Book review: A Combinatorial Identity (T. V. Narayana and M. Özsoyoglu). SIAM Review, 22(3):369, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1982:BRP

[Jag82] A. A. Jagers. Book review: A Property of an Inverse Laplace Transform (M. L. Glasser). SIAM Review, 24 (2):228–229, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1983:BRZ

[Jag83] A. A. Jagers. Book review: Zeros of a Determinant (Nancy J. Boynton). SIAM Review, 25(2): 271–273, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1987:BRO

[Jag87a] A. A. Jagers. Book review: An OC Curve Inequality (P. A. Roediger and J. G. Mardo). SIAM Review, 29 (4):636–637, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1987:BRR

[Jag87b] A. A. Jagers. Book review: On the Roots of a Bessel Function Equation (Anthony D. Rawlins). SIAM Review, 29(2):301–302, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1987:BNH

[Jag87c] A. A. Jagers. Bounds for nonlinear heat conduction. SIAM Review, 29(4):631, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1988:BRD

[Jag88a] A. A. Jagers. Book review: A Definite Integral of Bessel Functions (M. L. Glasser). SIAM Review, 30 (4):659–660, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1988:BRB

[Jag88b] A. A. Jagers. Book review: Bounds for Nonlinear Heat Conduction (A. A. Jagers). SIAM Review, 30(4): 655–657, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1988:PGF

[Jag88c] A. A. Jagers. A probability generating function. SIAM Review, 30(4):651– 652, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1990:BRC

[Jag90] A. A. Jagers. Book review: A Conjectured Matrix Inequality (W. H. Steeb and C. M. Villet). SIAM Review, 32(3): 483–484, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1992:BRI

[Jag92] A. A. Jagers. Book review: An Inequality for a Family of Means (H.-J. Seiffert). SIAM Review, 34(3):502, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1993:CMP

[Jag93] A. A. Jagers. A constrained minimum problem. SIAM Review, 35(2):298, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1994:CMP

[Jag94]  A. A. Jagers.  A con-

strained minimum problem. SIAM Review, 36(2):281– 283, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

James:1959:BRT

[Jam59] R. D. James. Book review: Trigonometric Series (R. L. Jeffery). SIAM Review, 1 (1):74–75, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jansen:1970:BRE

[Jan70] J. K. M. Jansen. Book review: An Elliptical Integral (A. W. Gillies). SIAM Review, 12(4):591–592, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jarvis:1976:BRC

[Jar76] John J. Jarvis. Book review:

Communication, Transmission, and Transportation Networks (Howard Frank and Ivan T. Frish). SIAM Review, 18(1):142, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1988:BRA

[JB88] W. B. Jordan and Walter Böhm. Book review: The Attenuated Abel Transform (Moshe Deutsch). SIAM Review, 30(1):125–128, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1990:CMI

[JBG^{+}90] A. A. Jagers, W. Boehm, C. C. Grosjean, R. Mathias, A. Y. Ozban, M. Renardy, R. S. Sezginer, W. So, B. A. Watson, and J. C. Xu. A conjectured matrix inequality. SIAM Review, 32(3): 483–484, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1972:BRFb

[JBL72] A. A. Jagers, T. J. Baker, and G. S. S. Ludford. Book review: A Fourier Sine Transform (S. H. Smith). SIAM Review, 14(2):368– 371, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1988:BPP

[JC88a] W. B. Jordan and J. Case. Bounded polynomials with a prescribed behavior at the origin. SIAM Review, 30(4): 657–659, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1988:MBF

[JC88b] W. B. Jordan and J. A. Cochran. Monotonicity of Bessel functions. SIAM Review, 30(3):504–506, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jones:2012:MGB

[JC12] Gareth Wyn Jones and S. Jonathan Chapman. Modeling growth in biological materials. SIAM Review, 54(1): 52–118, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v54/i1/p52_s1.

Joag-dev:1993:BRF

[Jd93] Kumar Joag-dev. Book review: Functional Equations in Probability Theory (B. Ramachandran and Ka-Sing Lau). SIAM Review, 35(2): 325–326, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jean:1978:GE

[Jea78] R. V. Jean. Growth and entropy. SIAM Review, 20 (3):627, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jeffries:1990:CRN

[Jef90] Clark Jeffries. Code recognition with neural network dynamical systems. SIAM Review, 32(4):636–651, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jeffries:1992:NCL

[Jef92] Clark Jeffries. A new conservation law for classical electrodynamics. SIAM Review, 34(3):386–405, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See comments [Rob94b] and response [Jef94].

Jeffries:1994:PVC

[Jef94] Clark Jeffries. Poynting’s vector: Comments on a recent paper by Clark Jeffries: Response to a commentary by F. N. H. Robinson. SIAM Review, 36(4):638– 641, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.

                                        siam.org/25583.htm. See

[Jef92, Rob94b].

Jeferey:1995:BRC

[Jef95]  Alan Jeferey.  Book re-

view: Characteristics of Distributed Parameter Systems (A. G. Butkovskiy and L. M. Pustyl’nikov). SIAM Review, 37(3):468–469, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jeffrey:2018:GDQ

[Jef18]  Mike R. Jeffrey.  The

ghosts of departed quantities in switches and transitions. SIAM Review, 60(1): 116–136, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jeltsch:1987:BRS

[Jel87] Rolf Jeltsch. Book review: Stiff Computation (Richard C. Aiken, ed.). SIAM Review, 29(2):329–330, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jensen:1975:SSM

[Jen75] P. S. Jensen. Stiffly stable methods for undamped second-order equations of motion. SIAM Review, 17 (2):382, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jensen:1978:LNP

[Jen78] S. Jensen. Large numerical problems in structural mechanics. SIAM Review, 20(3): 627–628, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jeroslow:1976:SSB

[Jer76] R. G. Jeroslow. Some structure and basis theorems for integral monoids. SIAM Review, 18(4):810, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jerome:1981:BRA

[Jer81] Joe Jerome. Book review:

Applied Semigroups and Evolution Equations (Aldo Belleni-Morante). SIAM Review, 23(3):408–409, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jerri:1982:DI

[Jer82] A. J. Jerri. A definite integral. SIAM Review, 24 (1):77, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jerome:1991:BRE

[Jer91a] Joseph W. Jerome. Book review: Electro-Diffusion of Ions (Isaak Rubinstein). SIAM Review, 33(4):682– 684, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jerome:1991:EDI

[Jer91b] Joseph W. Jerome. Electrodiffusion of ions (Isaak Rubinstein). SIAM Review, 33 (4):682–??, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jerome:1995:APD

[Jer95] Joseph W. Jerome. The approximation problem for drift-diffusion systems. SIAM Review, 37(4):552–572, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25888.htm.

Jespersen:1977:EEN

[Jes77a] D. C. Jespersen. Extreme eigenvalues of an nxn matrix. SIAM Review, 19(4): 742–743, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jespersen:1977:BRE

[Jes77b] Dennis C. Jespersen. Book review: On the Extreme Eigenvalues of an n × n Matrix (L. B. Bushard). SIAM Review, 19(4):742– 743, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jespersen:1993:FCE

[Jes93a] D. Jespersen. A family of cubic equations. SIAM Review, 35(3):503–504, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jespersen:1993:BRF

[Jes93b] Dennis Jespersen. Book review: A Family of Cubic Equations (Gerhard Metzen and Cecil Rousseau). SIAM Review, 35(3):503, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jetter:1991:BRA

[Jet91]  Kurt Jetter.  Book re-

view: Approximation by Spline Functions (Günther Nürnberger). SIAM Review, 33(3):501–502, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jevtic:1991:OMA

[Jev91] D. B. Jevtić. Optimum multiple access coding. SIAM Review, 33(1):114–115, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jevtic:1992:ODC

[Jev92a] D. B. Jevtić. Optimum disjoint constant-weight codebooks for multiple-access OR channels. SIAM Review, 34 (3):497–498, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jevtic:1992:OTE

[Jev92b] D. B. Jevtić. Optimum time to empty a random access FHMA system. SIAM Review, 34(4):648–649, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jevtic:1993:NTP

[Jev93]  D. B. Jevtić.  A num-

ber theory problem from multiple-access adder channels. SIAM Review, 35(3): 489–490, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jevtic:1994:GFR

[Jev94] D. B. Jevtić. A T-gambler FHMA ruin. SIAM Review, 36(4):655–656, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jevtic:1995:NTP

[Jev95a]  D. B. Jevtić.  A num-

ber theory problem from multiple-access adder channels. SIAM Review, 37(2): 240–241, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jevtic:1995:SDS

[Jev95b] D. B. Jevtić. Sum-distinct sets. SIAM Review, 37(1): 99–100, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1971:BRD

[JF71] W. B. Jordan and Henry E.

Fettis. Book review: A Definite Integral Containing Three Parameters (S. K. Dhir and J. S. Brook). SIAM Review, 13(3):401– 404, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordon:1987:SPT

[JF87] W. B. Jordon and C. L.

Frenzen. Sign pattern of terms of a Maclaurin series. SIAM Review, 29(4): 634–635, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1970:BRD

[JG70] William B. Jordan and M. L.

Glasser. Book review: A Definite Integral (B. F. Logan, C. L. Mallows and L. A. Shepp). SIAM Review, 12 (1):153–154, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1982:LS

[JG82] W. B. Jordan and J. Grzesik. Limit of a sum. SIAM Review, 24(2):226– 228, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1990:PVQ

[JG90] W. B. Jordan and M. L. Glasser. On the principal value of a quadruple integral. SIAM Review, 32(1): 143–145, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1993:BRT

[JG93] W. B. Jordan and C. C. Grosjean. Book review: Three Infinite Sums (C. C. Grosjean). SIAM Review, 35(2): 304–306, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1988:DIB

[JGA^{+}88] A. A. Jagers, N. Grossman, A. Adams, O. G. Ruehr, and O. P. Lossers. A definite integral of Bessel functions. SIAM Review, 30(4): 659–660, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1987:BRF

[JGF87] A. A. Jagers, M. L. Glasser, and H. E. Fettis. Book review: A Fourier Sine Series (M. L. Glasser and H. E. Fettis). SIAM Review, 29(2): 303–305, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1982:PIL

[JGK^{+}82] A. A. Jagers, N. Grossman, I. I. Kotlarski, N. Nasir, and R. Piessens. A property of an inverse Laplace transform. SIAM Review, 24(2): 228–229, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1995:BRP

[JGM95] W. B. Jordan, C. C. Grosjean, and Ernst W. Mayer. Book review: A Possible Characterization of Uniformly Accelerated Motion (M. S. Klamkin). SIAM Review, 37(2):249–250, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1990:RSU

[JH90] W. B. Jordan and V. Hnizdo.

Radiation from a system of uniformly circling charges. SIAM Review, 32(1):149– 154, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1992:DIA

[JH92] A. A. Jagers and E. M. J. Hinessen. A definite integral arising in Ohmic dissipation. SIAM Review, 34 (1):120, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jurdevic:1998:GCT

[JH98] Velimir Jurdevic and Robert Hermann. Geometric control theory. SIAM Review, 40(1): 154–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jeffrey:2011:GGS

[JH11] M. R. Jeffrey and S. J. Hogan. The geometry of generic sliding bifurcations. SIAM Review, 53(3):505– 525, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v53/i3/p505_s1.

Jafarpour:2022:FEN

[JHSB22] Saber Jafarpour, Elizabeth Y. Huang, Kevin D. Smith, and Francesco Bullo. Flow and elastic networks on the n-torus: Geometry, analysis, and computation. SIAM Review, 64 (1):59–104, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/18M1242056.

Jinghuang:1984:ZDI

[Jin84] Tian Jinghuang. Zeros of a definite integral. SIAM Review, 26(1):118, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jinghuang:1985:ZDI

[Jin85] Tian Jinghuang. Zeros of a definite integral. SIAM Review, 27(1):82, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jin:1991:BRV

[Jin91] Fei-Fei Jin. Book review:

Vacillation and Predictability Properties of Low-Order Atmospheric Spectral Models (H. E. deSwart). SIAM Review, 33(2):320–321, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1959:BP

[JK59] Mary Johnson and M. S.

Klamkin. The ballot problem. SIAM Review, 1(1): 68–69, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jeffrey:1972:WND

[JK72] A. Jeffrey and T. Kakutani. Weak nonlinear dispersive waves: a discussion centered around the Korteweg–de Vries equation. SIAM Review, 14(4):582– 643, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1988:BNH

[JK88] A. A. Jagers and T. Kauffman. Bounds for nonlinear heat conduction. SIAM Review, 30(4):655–657, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

John:2017:DCM

[JLM^{+}17] Volker John, Alexander Linke, Christian Merdon, Michael Neilan, and Leo G. Rebholz. On the divergence constraint in mixed finite element methods for incompressible flows. SIAM Review, 59(3):492–544, ???? 2017. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnston:1978:CED

[JM78] R. L. Johnston and R. Mathon. Computation of electric dipole fields in stratified conducting media. SIAM Review, 20(3):628, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Janssen:1992:URC

[JM92] A. J. E. M. Janssen and J. B. M. Melissen. An unexpected result in classical electrostatics. SIAM Review, 34 (4):648, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jia:2015:ODE

[JMFB15] Peng Jia, Anahita MirTabatabaei, Noah E. Friedkin, and Francesco Bullo. Opinion dynamics and the evolution of social power in influence networks. SIAM Review, 57 (3):367–397, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jurek:1995:OLD

[JMH95] Z. J. Jurek, J. D. Mason, and W. N. Hudson. Operatorlimit distributions in probability theory. SIAM Review, 37(1):131–??, March 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Janssen:1995:NSR

[JMJ95] A. Janssen, D. M. Mason, and Jana Jureckova. Nonstandard rank tests. SIAM Review, 37(2):261–??, June 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Joppich:1996:MMP

[JMX96] W. Joppich, S. Mijalkovic, and Jinchao Xu. Multigrid methods for process simulation. SIAM Review, 38(3): 543–??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jodar:1996:BRL

[Jód96]  Lucas Jódar.  Book re-

view: Lyapunov Matrix Equation in System Stability and Control (Zoran Gajic and Muhammad Tahir Javed Qureshi). SIAM Review, 38 (4):691, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1967:BRT

[Joh67] Millard W. Johnson, Jr. Book review: Theory of Elasticity of an Anisotropic Elastic Body (S. G. Lekhnitskii). SIAM Review, 9(1): 136, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnston:1968:BRC

[Joh68] R. L. Johnston. Book review: Computer Solution of Linear Algebraic Systems (George E. Forsythe and Cleve B. Moler). SIAM Review, 10(3): 384–385, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnston:1973:BRI

[Joh73] R. L. Johnston. Book review: Iterative Solution of Large Linear Systems (David M. Young). SIAM Review, 15 (1):236–237, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1975:IEF

[Joh75] H. L. Johnson. Integralequation formulation of a mixed boundary-value problem on a sphere. SIAM Review, 17(2):382–383, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1981:BRA

[Joh81] Charles R. Johnson. Book review: The Algebra of Econometrics (D. S. G. Pollock). SIAM Review, 23(3):406– 407, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1983:EMC

[Joh83a] C. Johnson. Economics for mathematicians — Cassels, J. W. S. SIAM Review, 25 (4):596–597, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1983:UNM

[Joh83b] C. Johnson. Uniform numerical methods for problems with initial and boundary layers — Doolan, E. D., Miller, J. J. H., Schilders, W. H. A. SIAM Review, 25(4): 597–598, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1983:BRE

[Joh83c] Charles Johnson. Book review: Economics For Mathematicians (J. W. S. Cassels). SIAM Review, 25(4): 596–597, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1983:BRU

[Joh83d] Claes Johnson. Book review: Uniform Numerical Methods for Problems with Initial and Boundary Layers (E. P. Doolan, J. J. H. Miller and W. H. A. Schilders). SIAM Review, 25 (4):597–598, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1983:CAH

[Joh83e] D. B. Johnson. Combinatorial algorithms — Hu, T. C. SIAM Review, 25 (4):599, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1983:BRC

[Joh83f] Donald B. Johnson. Book review: Combinatorial Algorithms (T. C. Hu). SIAM Review, 25(4):599, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1983:DI

[Joh83g] H. L. Johnson. A definite integral. SIAM Review, 25 (4):573, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1984:TTC

[Joh84a] C. R. Johnson. Theory of technical change and economic invariance — application of Lie groups — Sato, R. SIAM Review, 26(4): 609–610, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1984:BRT

[Joh84b] Charles R. Johnson. Book review: Theory of Technical Change and Economic Invariance: Application of Lie Groups (Ryuzo Sato). SIAM Review, 26(4):609– 610, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1987:BRT

[Joh87] Charles R. Johnson. Book review: The Theory of Matrices. Second Edition (with Applications) (Peter Lancaster and Miron Tismenetsky). SIAM Review, 29(3): 479–480, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1988:BRI

[Joh88] Charles R. Johnson. Book review: Invariant Subspaces of Matrices with Applications (I. Gohberg, P. Lancaster, and L. Rodman). SIAM Review, 30(4):671–672, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1991:BRN

[Joh91a] John L. Johnson. Book review: Numerical Simulation and Optimal Control in Plasma Physics: With Applications to Tokamaks (Jacques Blum). SIAM Review, 33(1): 147–150, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1991:NSO

[Joh91b]  John L. Johnson.  Nu-

merical simulation and optimal control in plasma physics: With applications to tokamaks (Jacques Blum). SIAM Review, 33(1):147– ??, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Joksch:1966:RVL

[Jok66a] H. C. Joksch. The reduction of the variance by least squares of polynomial approximations. SIAM Review, 8(2):211–219, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Joksch:1966:RED

[Jok66b] Hans C. Joksch. Random errors of derivatives obtained from least squares approximations to empirical functions. SIAM Review, 8(1): 47–56, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jolly:1996:BRE

[Jol96] Michael S. Jolly. Book review: Exponential Attractors for Dissipative Evolution Equations (A. Eden, C. Foias, B. Nicolaenko, and R. Témam). SIAM Review, 38 (1):171–172, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jones:1971:BCE

[Jon71] Michael Jones. Balancing chemical equations. SIAM Review, 13(4):571, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jones:1972:ABI

[Jon72]  D. S. Jones.  Asymp-

totic behavior of integrals. SIAM Review, 14(2):286– 317, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jones:1992:BRG

[Jon92a] Alan K. Jones. Book review: Geometric & Solid Modeling: an Introduction (Christoph M. Hoffman). SIAM Review, 34(2):327–329, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jones:1992:GSM

[Jon92b] Alan K. Jones. Geometric and solid modeling: An introduction (Christoph M Hoffman). SIAM Review, 34 (2):327–??, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jones:1992:BRS

[Jon92c]  D. S. Jones.  Book re-

view: The Scientific Letters and Papers of James Clerk Maxwell Volume 1: 1846–1862. (P. M. Harman ed.). SIAM Review, 34 (3):511–513, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1972:BRA

[Jor72] W. B. Jordan. Book review:

An Asymptotic Expansion (P. Barrucand). SIAM Review, 14(3):501–502, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1973:BRT

[Jor73] W. B. Jordan. Book review: A Triple Sum (P. G. Kirmser and K. K. Hu). SIAM Review, 15(4):802– 803, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1977:BRA

[Jor77a] W. B. Jordan. Book review: An Average Relative Speed Approximation (L. K. Arnold , L. Dodson and L. Rosen). SIAM Review, 19(3): 565–567, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1977:BRI

[Jor77b] W. B. Jordan. Book review: An Infinite Sum (Murray Geller). SIAM Review, 19 (1):150–152, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1977:ZMM

[Jor77c] W. B. Jordan. Zero of maximum multiplicity. SIAM Review, 19(4):743–744, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1978:BRD

[Jor78a] W. B. Jordan. Book review: A Double Integral from Demagnetizing Fields (R. I. Joseph). SIAM Review, 20 (1):184–186, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1978:BRP

[Jor78b] W. B. Jordan. Book review: A Property of the First Erlang Function (Ilia Kaufman). SIAM Review, 20(3): 599–601, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1979:BRT

[Jor79] W. B. Jordan. Book review: Two Inequalities (P. Barraucand). SIAM Review, 21(1): 140–143, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1980:BRC

[Jor80a] W. B. Jordan. Book review: A Conjectured Property of Legendre Functions (A. K. Raina and V. Singh). SIAM Review, 22(3):369– 372, July 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1980:FSC

[Jor80b] W. B. Jordan. Fourier series for a combination of Jacobian elliptic functions. SIAM Review, 22(2):232– 233, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1980:BRF

[Jor80c] William B. Jordan. Book review: Fourier Series for a Combination of Jacobian Elliptic Functions (N. R. Pereira). SIAM Review, 22 (2):232–234, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1981:BRM

[Jor81] W. B. Jordan. Book review:

A Matrix Stability Problem (K. Sourisseau and M. F. Doherty). SIAM Review, 23(1): 107–112, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1982:BRP

[Jor82] W. B. Jordan. Book review: Power Series of an Elliptic Function (M. L. Glasser). SIAM Review, 24 (4):479, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1983:BRT

[Jor83a] W. B. Jordan. Book review: A Triple Sum (P. E. Merilees). SIAM Review, 25(2): 269–271, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1983:GFE

[Jor83b] W. B. Jordan. Gamma function expansions. SIAM Review, 25(2):268, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1984:BRS

[Jor84a] W. B. Jordan. Book review: A Sum of Legendre Polynomials (M. L. Glasser). SIAM Review, 26(4):586– 587, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1984:BRTa

[Jor84b] W. B. Jordan. Book review: Truncation of the Rogers– Ramanujan Theta Series (G. E. Andrews). SIAM Review, 26(3):433–434, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1984:BRTb

[Jor84c] W. B. Jordan. Book review: Two Legendre Polynomial Identities (A. D. Rawlins). SIAM Review, 26(4):

582–584, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1986:BRE

[Jor86a] W. B. Jordan. Book review:

Extension of Routh’s Theorem to Spherical Triangles (A. Sharma and Murray S. Klamkin). SIAM Review, 28 (3):403–404, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1986:BRR

[Jor86b] W. B. Jordan. Book review: Radius of Convergence of a Power Series (J. E. Wilkins). SIAM Review, 28 (4):570–572, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1986:RCP

[Jor86c] W. B. Jordan. Radius of convergence of a power series. SIAM Review, 28 (4):571–572, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1986:SPR

[Jor86d] W. B. Jordan. Sums of powers of roots of a transcendental equation. SIAM Review, 28(3):400–402, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordon:1987:BRS

[Jor87a] W. B. Jordon. Book review: Sign Pattern of Terms of a Maclaurin Series (C. L. Frenzen). SIAM Review, 29 (4):634, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jorgensen:1987:BRH

[Jor87b] Pallee E. T. Jorgensen. Book review: Harmonic Analysis on Symmetric Spaces and Applications I (Aunxey Terras). SIAM Review, 29(2): 330–332, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1988:BRB

[Jor88a] W. B. Jordan. Book review: Bounded Polynomials with a Prescribed Behavior at the Origin (M. M. Chawla). SIAM Review, 30 (4):657–659, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1988:BRM

[Jor88b] W. B. Jordan. Book review: Monotonicity of Bessel Functions (S. W. Rienstra). SIAM Review, 30(3):504– 506, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1988:DS

[Jor88c] W. B. Jordan. Double summations. SIAM Review, 30 (2):322, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1989:BRC

[Jor89a] W. B. Jordan. Book review: Conjectured Maximum Term of a Sequence (A. Gerasoulis and R. P. Srivastav). SIAM Review, 31(1): 127–129, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jorgensen:1989:BRH

[Jor89b] Palls Jorgensen. Book review: Harmonic Analysis on Symmetric Spaces and Applications II (Audrey Terras). SIAM Review, 31(4): 701, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jorgensen:1989:BRG

[Jor89c] Palls E. Jorgensen. Book review: Group Structure of Gauge Theories (L. O’Raifeartaigh). SIAM Review, 31(1):142– 144, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1990:BRE

[Jor90a] W. B. Jordan. Book review: An Elliptic Function Identity (E. L. Wachspress). SIAM Review, 32(1): 142–143, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1990:BRP

[Jor90b] W. B. Jordan. Book review: On the Principal Value of a Quadruple Integral (M. L. Glasser). SIAM Review, 32 (1):143–145, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1990:BRR

[Jor90c] W. B. Jordan. Book review: Radiation from a System of Uniformly Circling Charges (V. Hnizdo). SIAM Review, 32(1):149–153, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1993:IS

[Jor93a] W. B. Jordan. 3 infinite sums. SIAM Review, 35(2): 304–306, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1993:BRL

[Jor93b] W. B. Jordan. Book review: Limit of a Sequence (M. Gaviano). SIAM Review, 35(3): 490–492, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1995:PCU

[Jor95] W. B. Jordan. A possible characterization of uniformly accelerated motion. SIAM Review, 37(2):249– 250, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1996:CUA

[Jor96] W. B. Jordan. A characterization of uniformly accelerated motion. SIAM Review, 38(3):525–526, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Joshi:1970:NSR

[Jos70] V. N. Joshi. A note on the solution of rectangular linear systems by iteration. SIAM Review, 12(3): 463–466, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Joseph:1976:DID

[Jos76] R. I. Joseph. A double integral from demagnetizing fields. SIAM Review, 18 (4):762, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Joseph:1995:BRI

[Jos95] R. I. Joseph. Book review: An Integral Occurring in Coherence Theory (R. C. Wittman and B. K. Alpert). SIAM Review, 37 (4):610–615, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jowett:1972:BRS

[Jow72] David Jowett. Book review: Statistical Design and Analysis of Experiments (Peter W. M. John). SIAM Review, 14 (1):177–178, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jowett:1976:BRA

[Jow76] David Jowett. Book review: Analysis of Variance in Complex Experimental Designs (Harold R. Lindman). SIAM Review, 18(1):134– 137, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jowett:1986:BRI

[Jow86] David Jowett. Book review: Infinitely Divisible Statistical Experiments (Arnold Janssen, Hartmut Milbrodt, and Helmut Strasser). SIAM Review, 28(3):423–424, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Joyce:1971:SEP

[Joy71] D. C. Joyce. Survey of extrapolation processes in numerical analysis. SIAM Review, 13(4):435–490, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jaffray:1989:DPK

[JP89] J. Y. Jaffray and J. Ch.

Pomerol. A direct proof of the Kuhn–Tucker necessary optimality theorem for convex and affine inequalities. SIAM Review, 31 (4):671–674, December 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1987:RBF

[JR87] A. A. Jagers and A. D.

Rawlins. On the roots of a Bessel-function equation. SIAM Review, 29(2): 301–302, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jensen:1996:SA

[JR96] Jens Ledet Jensen and N. Reid. Saddlepoint approximations. SIAM Review, 38 (4):696–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Josic:2008:USN

[JR08] Krešimir Josić and Robert Rosenbaum. Unstable solutions of nonautonomous linear differential equations. SIAM Review, 50(3):570– 584, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:2022:MMO

[JRA22] Joseph D. Johnson, Adam M. Redlich, and Daniel M. Abrams. A mathematical model for the origin of name brands and generics. SIAM Review, 64 (3):625–639, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1360888.

Jamison:1971:BRP

[JRR71] J. E. Jamison, C. C. Rousseau, and J. W. Riese. Book review: A Positive Functional (Richard Kellner). SIAM Review, 13(3): 399–401, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1997:CSI

[JRZ97] M. P. Johnson, C. Rennick, and E. Zak. Case studies from industry: Skiving addition to the cutting stock problem in the paper industry. SIAM Review, 39 (3):472–483, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31004.

Jones:1968:EOC

[JS68] G. Stephen Jones and Aaron Strauss. An example of optimal control. SIAM Review, 10(1):25–55, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jaffey:1971:BRU

[JS71] Arthur H. Jaffey and Anthony J. Strecok. Book review: Upper Bound for the Abscissa of the χ^{2}- Distribution (Peter Treuenfels). SIAM Review, 13(1): 136–137, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1976:BEB

[JS76] O. G. Johnson and J. L.

Stiles. Best ellipsoidal bounds for tracking problem sensitivity. SIAM Review, 18 (4):810, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jacquez:1993:QTC

[JS93] John A. Jacquez and Carl P.

Simon. Qualitative theory of compartmental systems. SIAM Review, 35(1): 43–79, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jawerth:1994:OWB

[JS94] Björn Jawerth and Wim Sweldens. An overview of wavelet based multiresolution analyses. SIAM Review, 36 (3):377–412, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/24402.
htm.

Ju:2020:DAF

[JS20] Caleb Ju and Edgar Solomonik. Derivation and analysis of fast bilinear algorithms for convolution. SIAM Review, 62(4):743–777, ???? 2020. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Janacek:1995:TSF

[JSF95] Gareth Janacek, Louise Swift, and David F. Findley. Time series: Forecasting, simulation, applications. SIAM Review, 37(2): 253–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jones:1982:BRE

[JSR82a] P. W. Jones, P. Smith, and D. E. Raeside. Book review: Estimation of the Median Effective Dose in Bioassay (D. E. Raeside). SIAM Review, 24(3):342–344, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jones:1982:EME

[JSR82b] P. W. Jones, P. Smith, and D. E. Raeside. Estimation of the median effective dose in bioassay. SIAM Review, 24 (3):342–345, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Johnson:1976:CDE

[JU76] G. W. Johnson and R. G.

Underwood. Controllability of differential equations with phase constraints. SIAM Review, 18(4):810, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Juberg:1970:USC

[Jub70]  R. K. Juberg.  Unique-

ness of solutions for certain boundary value problems. SIAM Review, 12(3): 461–462, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Juncosa:1964:BRP

[Jun64] Mario Juncosa. Book review: Proceedings of the 15th Symposium in Applied Mathematics of the A.M.S. Experimental Arithmetic, High Speed Computing and Mathematics (N. C. Metropolis, A. H. Taub, John Todd C. B. Tompkins). SIAM Review, 6(4):468–471, October 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

                                        //links.jstor.org/sici?
                                        sici=0036-1445%28196410%
                                        296%3A4%3C468%3APOT1SI%
                                        3E2.0.CO%3B2-T.

Juncosa:1965:BRM

[Jun65] M. L. Juncosa. Book review: Monte Carlo Methods (J. M. Hammersley and D. C. Handscomb). SIAM Review, 7(3): 435–436, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Juncosa:1970:BRR

[Jun70a] Mario Juncosa. Book review: Random Number Generators (Birger Jansson). SIAM Review, 12(2):310– 311, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        link.aip.org/link/?SIR/
                                        12/310/1.

Juncosa:1970:BRM

[Jun70b] Mario Juncosa. Book review: The Monte Carlo Method (Yu. A. Shreider). SIAM Review, 12(2):309– 310, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jurgen:1986:RDT

[Jür86]  Moser Jürgen.  Recent

developments in the theory of Hamiltonian systems.

SIAM Review, 28(4):459– 485, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jureckova:1995:BRN

[Jur95] Jana Jurećková. Book review: Non-Standard Rank Tests (A. Janssen and D. M. Mason). SIAM Review, 37 (2):261–262, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1990:CFW

[JvD90] A. A. Jagers and E. A. van Doorn. Convexity of functions which are generalizations of the Erlang loss function and the Erlang delay function. SIAM Review, 32 (2):301–302, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jagers:1991:CFW

[JvD91] A. A. Jagers and E. A. van Doorn. Convexity of functions which are generalizations of the Erlang loss function and the Erlang delay function. SIAM Review, 33 (2):281–282, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jamison-Waldner:1983:BRG

[JW83] Robert E. Jamison-Waldner.

Book review: Geometry and Convexity: a Study in Mathematical Methods (Paul J. Kelly and Max L. Weiss ). SIAM Review, 25(2): 278–279, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1990:EFI

[JW90] W. B. Jordan and E. L. Wachspress. An elliptic function identity. SIAM Review, 32(1):142–143, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kleinman:1976:ORP

[KA76] R. E. Kleinman and T. S. Angell. Optimization of radiated power in an angular sector. SIAM Review, 18 (4):812, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kim:1989:EAA

[KA89] Young J. Kim and Joan L. Aron. On the equality of average age and average expectation of remaining life in a stationary population. SIAM Review, 31(1): 110–113, March 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kuijper:1995:FOR

[KA95] Margreet Kuijper and J. D. Aplevich. First-order representations of linear systems. SIAM Review, 37(4): 621–??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kabamba:1993:NNA

[Kab93]  Pierre T. Kabamba.  A

note on navigation accuracy. SIAM Review, 35 (3):481–487, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kac:1962:PTR

[Kac62] M. Kac. Probability theory:

Its role and its impact. SIAM Review, 4(1):1–11, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kahan:1966:CEB

[Kah66] W. M. Kahan. A computable error-bound for systems of ordinary differential equations (abstract). SIAM Review, 8 (??):568–569, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kahan:1976:PIP

[Kah76] W. M. Kahan. Pitfalls in ill-posed problems. SIAM Review, 18(4):810–811, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaijtz:1959:SP

[Kai59] William H. Kaijtz. A sorting problem. SIAM Review, 1(2):173, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kalaba:1963:BRD

[Kal63] Robert Kalaba. Book review: Dynamic Optimization and Control — A Variational Approach (W. Kipiniak). SIAM Review, 5(1):81, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kalman:1965:BRO

[Kal65] R. E. Kalman. Book review: Optimum Design of Digital Control Systems (Julius T. Tou). SIAM Review, 7(1): 140–141, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kalman:1970:BRA

[Kal70] R. E. Kalman. Book review: Absolute Stability of Regulator Systems (M. A. Aizerman and F. R. Gantmacher). SIAM Review, 12 (1):161–162, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kallenberg:1993:BRL

[Kal93a] Olav Kallenberg. Book review: Lectures on the Coupling Method (Torgny Lindvall). SIAM Review, 35(3): 525–527, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kallenberg:1993:LCM

[Kal93b] Olav Kallenberg. Lectures on the coupling method (Torgny Lindvall). SIAM Review, 35 (3):525–??, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kammler:1976:NEE

[Kam76] D. W. Kammler. Numerical evaluation of exp(at) when a is a companion matrix. SIAM Review, 18(4):811, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kammler:1978:ACM

[Kam78] D. W. Kammler. L_{1} approximation of completely monotonic functions by sums of exponentials. SIAM Review, 20 (3):628, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kanwal:1962:BRG

[Kan62a] R. P. Kanwal. Book review: General Relativity and Gravitational Waves (J. Weber). SIAM Review, 4(1): 58, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kanwal:1962:BRS

[Kan62b] R. P. Kanwal. Book review: Supersonic Aerodynamics (E. R. C. Miles). SIAM Review, 4(3):272, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kanwal:1962:BRD

[Kan62c] R. P. Kanwal. Book review: The Dynamical Theory of Sound (H. Lamb). SIAM Review, 4(1):58, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kanwal:1966:BRT

[Kan66] R. P. Kanwal. Book review: Theoretical Hydromechanics (N. E. Kochin, I. A. Kibel, and N. V. Roze). SIAM Review, 8(1):120– 121, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kanwal:1968:BRS

[Kan68]  Ram P. Kanwal.  Book

review: Stability of Parallel Flows (Robert Betchov and William O. Criminale, Jr.). SIAM Review, 10(3): 396, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaneko:1978:BRN

[Kan78] Ikuyo Kaneko. Book review: Nonlinear Programming (Mordecai Avriel). SIAM Review, 20(1):204– 205, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kantrowitz:1985:MV

[Kan85] M. Kantrowitz. A minimum value. SIAM Review, 27(2): 255–256, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kanwal:1988:DGS

[Kan88] R. P. Kanwal. A differential geometric solution of a problem in electrostatics. SIAM Review, 30(2):301, June 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1964:BRL

[Kap64] Wilfred Kaplan. Book review: Linear System Theory, the State Space Approach (Lotfi A. Zadeh and Charles A. Desoer). SIAM Review, 6 (3):319–320, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1965:BRG

[Kap65] Wilfred Kaplan. Book review: Generalized Functions and Direct Operational Methods, vol 1 (T. P. G. Liverman). SIAM Review, 7 (2):299, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1966:BRD

[Kap66] Wilfred Kaplan. Book review: Dynamics of Linear Systems (Václav Dolezal). SIAM Review, 8(2):246– 247, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1967:BRF

[Kap67] Wilfred Kaplan. Book review: Fourier Transforms and the Theory of Distributions (J. Arsac). SIAM Review, 9(3):606–607, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1968:BRA

[Kap68] Wilfred Kaplan. Book review: Applications of Undergraduate Mathematics in Engineering (Ben Noble). SIAM Review, 10(3):383– 384, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1970:SAI

[Kap70a] Stan Kaplan. Some algebraic identities. SIAM Review, 12 (1):144–145, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1970:BRT

[Kap70b] Wilfred Kaplan. Book review: Topics in Mathematical System Theory (Rudolf E. Kalman, Peter L. Falb and Michael A. Arbib). SIAM Review, 12(1):157– 158, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1971:BRI

[Kap71] Wilfred Kaplan. Book review: Introduction to Applied Mathematics (Joseph Genin and John S. Maybes). SIAM Review, 13(4):584– 585, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1972:BRT

[Kap72] Wilfred Kaplan. Book review: Topics in Ordinary Differential Equations: a Potpourri (William D. Lakin and David A. Sanchez). SIAM Review, 14(3):508– 509, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1976:BRM

[Kap76] Wilfred Kaplan. Book review: Mathematics Applied to Deterministic Problems in the Natural Sciences (C. C. Lin and L. A. Segel). SIAM Review, 18(1):146– 148, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1977:BRO

[Kap77] Wilfred Kaplan. Book review: Ordinary Differential Equations in the Complex Domain (Einar Hills). SIAM Review, 19(4):749, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1978:BRI

[Kap78] Wilfred Kaplan. Book review: Introduction to the Theory of Infinitesimals (J. D. Stroyan and W. A. J. Luxemburg). SIAM Review, 20 (2):403, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaplan:1981:BRG

[Kap81] Wilfred Kaplan. Book review: Green’s Functions and Boundary Value Problems (Ivar Stackgold). SIAM Review, 23(1):117–118, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaper:1985:BRO

[Kap85a] Hans G. Kaper. Book review: Ordinary Differential Equations in R^{n}-Problems and Methods (L. C. Piccinini, G. Stampacchia and G. Vidossich). SIAM Review, 27(4): 596–598, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kapila:1985:BRT

[Kap85b] A. K. Kapila. Book review:

Theory of Laminar Flames (J. D. Buckmaster and G. S. S. Ludford). SIAM Review, 27(2):281–283, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kapila:1990:BRD

[Kap90a] A. K. Kapila. Book review:

Dynamics of Curved Fronts (Pierre Pelce, ed.). SIAM Review, 32(3):507, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kapila:1990:BRM

[Kap90b] A. K. Kapila. Book review: Mathematics in Industrial Problems Mathematics in Industrial Problems, Part 2. (Avner Friedman). SIAM Review, 32(3):506– 507, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kapila:1990:DCF

[Kap90c]  A. K. Kapila.  Dynam-

ics of curved fronts (Pierre Pelce). SIAM Review, 32(3): 507–??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kapila:1990:MIP

[Kap90d] A. K. Kapila. Mathematics in industrial problems and mathematics in industrial problems, part 2 (Avner Friedman). SIAM Review, 32 (3):506–??, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karp:1960:BRH

[Kar60] Richard M. Karp. Book review: Handbook of Automation, Computation, and Control, Volume 2, Computers and Data Processing (E. M. Grabbe, S. Ramo, and D. E, Wooldridge, eds.). SIAM Review, 2(4): 299–302, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karp:1963:BRU

[Kar63]  R. M. Karp.  Book re-

view: Understanding Digital Computers (P. Siegel). SIAM Review, 5(2):169– 170, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karreman:1964:BRL

[Kar64] H. F. Karreman. Book review: Linear Programming (G. Hadley). SIAM Review, 6 (2):191–192, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karp:1976:PMC

[Kar76] R. M. Karp. Probabilistic models of combinatorial optimization methods. SIAM Review, 18(4):811, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karlson:1977:GHF

[Kar77] P. W. Karlson. A Gaussian hypergeometric function identity. SIAM Review, 19 (1):146, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karr:1978:BRM

[Kar78] Alan F. Karr. Book review:

Markov Chains: Theory and Applications (Dean L. Isaacson and Richard W. Madsen). SIAM Review, 20(3): 606–607, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karr:1982:BRS

[Kar82] Alan F. Karr. Book review: Stochastic Processes and Integration (M. M. Rao).

SIAM Review, 24(2):238– 239, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karr:1983:BRS

[Kar83] Alan F. Karr. Book review: A Second Course in Stochastic Processes (Samuel Karlin and Howard M. Taylor). SIAM Review, 25(2): 280–281, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karr:1984:SPR

[Kar84a] A. F. Karr. Stochasticprocesses — Ross, S. M. SIAM Review, 26(3):448– 449, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karr:1984:BRS

[Kar84b] Alan F. Karr. Book review: Stochastic Processes (Sheldon M. Ross). SIAM Review, 26(3):448–449, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karwan:1987:BRS

[Kar87] Mark H. Karwan. Book review: Stochastic Programming with Multiple Objective Functions (I. M. Stancu-Minasian). SIAM Review, 29 (3):480–481, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karni:1990:BRN

[Kar90] Edi Karni. Book review:

Nonlinear Preference and Utility Theory (Peter C. Fishburn). SIAM Review, 32(1): 176–178, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karp:1996:BRU

[Kar96] Alan H. Karp. Bit reversal on uniprocessors. SIAM Review, 38(1):1–26, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25698.htm.

Karger:2001:RFP

[Kar01] David R. Karger. A randomized fully polynomial time approximation Scheme for the all-terminal network reliability problem. SIAM Review, 43(3):499– 522, September 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38714.

Kardos:2022:HCP

[Kar22] Frantisek Kardos. Hamiltonicity of cubic planar graphs with bounded face sizes. SIAM Review, 64 (2):425–465, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1476915.

Katsanis:1965:BRC

[Kat65] Theodore Katsanis. Book review: On Commutative Rotations (Joel Brenner). SIAM Review, 7(2):276, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Katz:1968:REM

[Kat68] Sylvan Katz. Rank and eigenvalues of a matrix. SIAM Review, 10(2):225, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Katz:1976:CTF

[Kat76] I. N. Katz. Conforming triangular finite elements of arbitrary polynomial order. SIAM Review, 18(4):811– 812, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Katz:1978:HCC

[Kat78] I. N. Katz. Hierarchical, complete, conforming tetrahedral elements of arbitrary polynomial order. SIAM Review, 20(3):628, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaufman:1977:PFE

[Kau77]  Ilia Kaufman.  A prop-

erty of the first Erlang function. SIAM Review, 19(3): 564, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaup:1986:BRA

[Kau86] David J. Kaup. Book review: Advances in Nonlinear Waves-Volumes I/II (Lokenath Debnath, ed.). SIAM Review, 28(2):261– 262, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kautsky:1988:BRC

[Kau88] J. Kautsky. Book review:

A Column Vector Problem From Numerical Integration (M. M. Chawla). SIAM Review, 30(4):653– 655, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kawata:1963:BRI

[Kaw63] T. Kawata. Book review:

An Introduction to Infinitely Many Variates (E. A. Robinson). SIAM Review, 5(1): 86–87, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kawata:1965:BRS

[Kaw65] Tatsuo Kawata. Book review: Stochastic Process in Mathematical Physics and Engineering (Richard Bellman, eds.). SIAM Review, 7(4): 586–588, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kawata:1966:BRP

[Kaw66] Tatsuo Kawata. Book review: Probability Theory (K. Krickeberg). SIAM Review, 8 (3):411–413, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kawata:1967:BRF

[Kaw67] Tatsuo Kawata. Book review: A First Course in Stochastic Processes (Samuel Karlin). SIAM Review, 9(2): 265–266, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kawata:1968:BRS

[Kaw68] Tatstuo Kawata. Book review: Stationary and Related Stochastic Processes (Harald Cramér and M. R. Leadbetter). SIAM Review, 10(1): 94–95, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kawohl:1995:CHF

[Kaw95]  B. Kawohl.  A conjec-

tured heat-flow problem. SIAM Review, 37(1):104– 105, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1962:BRA

[Kaz62] N. D. Kazarinoff. Book review: Asymptotic Estimates and Entire Functions (M. A. Evgrafov; A. Shields, trans.). SIAM Review, 4(3): 270–271, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1967:BRF

[Kaz67] Nicholas D. Kazarinoff. Book review: Formulas and Theorems for the Special Functions of Mathematical Physics (Wilhelm Magnus, Fritz Oberhettinger and Raj Pal Soni). SIAM Review, 9 (4):755, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1975:BRF

[Kaz75] Nicholas D. Kazarinoff. Book review: Foundations of Modern Potential Theory (N. S. Landkof). SIAM Review, 17 (1):180, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1980:BRM

[Kaz80a] N. D. Kazarinoff. Book review: Mathematical Aspects of Reacting and Diffusing Systems (Paul C. Fife). SIAM Review, 22(4):520, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1980:BRT

[Kaz80b] Nicholas D. Kazarinoff. Book review: Time Lags in Biological Models (Norman Mac-Donald). SIAM Review, 22 (2):245–246, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1982:BRC

[Kaz82] N. D. Kazarinoff. Book review: Catastrophe Theory and Bifurcations: Applications to Urban and Regional Systems (A. G. Wilson). SIAM Review, 24(4): 497–498, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1983:BRQ

[Kaz83] N. D. Kazarinoff. Book review: Quadratic Forms Theory and Differential Equations (John Gregory). SIAM Review, 25(2):288–289, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1984:NOD

[Kaz84a] N. D. Kazarinoff. Nonlinear oscillations, dynamical systems, and bifurcations of vector fields — Guckenheimer, J., Holmes, P. SIAM Review, 26(4):600–601, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1984:BRN

[Kaz84b] Nicholas D. Kazarinoff. Book review: Nonlinear Oscillations, Dynamical Systems, and Bifurcations of Vector Fields (John Guckenheimer and Philip Holmes). SIAM Review, 26(4):600– 601, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1986:BRM

[Kaz86] Nicholas D. Kazarinoff. Book review: Methods in the Qualitative Theory of Dynamical Systems in Astrophysics and Gas Dynamics (Oleg I. Bogoyavlensky). SIAM Review, 28(4):592– 593, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1989:BRP

[Kaz89] Nicholas D. Kazarinoff. Book review: A Primer of Diffusion Problems (Richard Ghez). SIAM Review, 31 (3):508, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1990:BRE

[Kaz90a] Nicholas D. Kazarinoff. Book review: Elements of Differentiable Dynamics and Bifurcation Theory (David Ruelle). SIAM Review, 32 (4):698, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1990:EDD

[Kaz90b] Nicholas D. Kazarinoff. Elements of differentiable dynamics and bifurcation theory (David Ruelle). SIAM Review, 32(4):698–??, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1991:BRH

[Kaz91a] Nicholas D. Kazarinoff. Book review: Huygens and Barrow, Newton and Hooke: Pioneers in Mathematical Analysis and Catastrophe Theory from Evolvents to Quasicrystals (V. I. Arnol’d; trans. by Eric J. F. Primrose). SIAM Review, 33(3):493– 495, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1991:BRJ

[Kaz91b] Nicholas D. Kazarinoff. Book review: Joseph Liouville 1809-1882: Master of Pure and Applied Mathematics (Jesper Lützen). SIAM Review, 33(2):321–323, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazarinoff:1992:BRI

[Kaz92] Nicholas D. Kazarinoff. Book review: IBM’s 360 and Early 370 Systems (Emerson W. Pugh, Lyle R. Johnson, and John H. Palmer). SIAM Review, 34(1):158, March 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Koopmans:1959:STE

[KB59] Tjalling C. Koopmans and Augustus F. Bausch. Selected topics in economics involving mathematical reasoning. SIAM Review, 1(2): 79–148, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klinger:1968:MTF

[KB68] Allen Klinger and Thomas A.

Brown. Monotonicity of two functions. SIAM Review, 10 (4):452, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaufman:1976:EUO

[KB76a] G. M. Kaufman and E. Barouch. Estimating undiscovered oil and gas. SIAM Review, 18 (4):812, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaufman:1976:ACD

[KB76b] L. Kaufman and J. R. Bunch. Algorithm for computing a decomposition for symmetric matrices for solving linear systems and calculating inertia. SIAM Review, 18 (4):812, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolda:2009:TDA

[KB09] Tamara G. Kolda and Brett W. Bader. Tensor decompositions and applications. SIAM Review, 51(3): 455–500, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Komen:1997:DMO

[KCD^{+}97] G. J. Komen, L. Cavaleri, M. Donelan, K. Hasselmann, S. Hasselmann, P. E. A. M. Janssen, and Sen-Huei Chen. Dynamics and modelling of ocean waves. SIAM Review, 39(4):790– ??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knuth:1980:ABS

[KCI^{+}80] D. E. Knuth, J. Case, M. E. H. Ismail, R. Piessens, J. R. Pounder, N. M. Temme, J. H. Vanlint, D. J. Daley, Z. Drezner, T. O. Espelid, G. Grippenberg, W. B. Jordan, W. J. Rey, P. Sermer, M. Klawe, W. Pulleyblank, and W. Vollmerhaus. Asymptotic behavior of a sequence. SIAM Review, 22(1): 101–102, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krueger:1976:IAG

[KD76] R. J. Krueger and J. P.

Dauer. Iterative approach to goal programming. SIAM Review, 18(4):813, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krueger:1978:MOM

[KD78] R. J. Krueger and J. P.

Dauer. Multiobjective optimization model. SIAM Review, 20(3):629, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klapper:2010:MDM

[KD10] Isaac Klapper and Jack Dockery. Mathematical description of microbial biofilms. SIAM Review, 52 (2):221–265, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1988:OMR

[KDJM88] M. S. Klamkin, F. Dubeau, W. B. Jordan, and W. W. Meyer. An optimum multistage rocket design. SIAM Review, 30(3):508–509, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kellogg:1998:BRc

[Ke98] R. B. Kellogg and editor. Book reviews. SIAM Review, 40(3):732–760, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

                                        http://epubs.siam.org/
                                        sam-bin/dbq/article/97013.

Kearsley:1980:BRN

[Kea80] D. A. Kearsley. Book review: A Nonlinear Differential Equation (D. K. Ross). SIAM Review, 22(3):364– 365, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keating:1993:BRQ

[Kea93a] J. P. Keating. Book review: Quantum Signatures of Chaos (Fritz Haake). SIAM Review, 35(3):527– 529, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keating:1993:QSC

[Kea93b] J. P. Keating. Quantum signatures of chaos (Fritz Haake). SIAM Review, 35(3): 527–??, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keane:1995:EEF

[Kea95]  Joe Keane.  Expansions

for elementary functions. SIAM Review, 37(4):609– 610, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keane:1996:EEF

[Kea96a] J. Keane. Expansions for elementary functions. SIAM Review, 38(4):675–678, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keane:1996:BRE

[Kea96b]  Joe Keane.  Book re-

view: Expansions for Elementary Functions (Joe Keane). SIAM Review, 38 (4):675–678, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keeping:1961:BRT

[Kee61] E. S. Keeping. Book review: Theory of Probability (William Burnside). SIAM Review, 3(2):184, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keer:1987:BRC

[Kee87] Leon M. Keer. Book review: Contact Mechanics (K. L. Johnson). SIAM Review, 29 (2):332–333, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keener:1988:BRP

[Kee88a] James P. Keener. Book review: Perturbation Methods, Bifurcation Theory and Computer Algebra (Richard H. Rand and Dieter Armbruster). SIAM Review, 30 (3):516–517, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keesling:1988:BRP

[Kee88b] James Keesling. Book review: Positive Feedback in Natural Systems (Donald L. DeAngelis, Wilfred M. Post, and Curtis C. Travis). SIAM Review, 30(2):332–335, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keesling:1990:BRN

[Kee90] James Keesling. Book review: Natural Resource Economics: Notes and Problems (Jon M. Conrad and Colin W. Clark). SIAM Review, 32 (2):323–325, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keener:1993:PFT

[Kee93]  James P. Keener.  The

Perron–Frobenius theorem and the ranking of football teams. SIAM Review, 35 (1):80–93, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keenan:1995:BRG

[Kee95] Daniel Keenan. Book review:

General Pattern Theory: a Mathematical Study of Regular Structures (U. Grenander). SIAM Review, 37(2): 258–261, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kehoe:1989:BRM

[Keh89] Timothy J. Kehoe. Book review: Microeconomic Theory (Yoshihiko Otani and Mohamed El-Hodiri). SIAM Review, 31(4):701–703, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keilson:1964:BRN

[Kei64] Julian Keilson. Book review: Nonlinear Transformations of Random Processes (R. Deutsch). SIAM Review, 6(1):77–78, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keilson:1965:BRE

[Kei65] Julian Keilson. Book review: Elements of Queuing Theory (T. L. Saaty). SIAM Review, 7(1):151, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keilson:1971:BRS

[Kei71] Julian Keilson. Book review: The Single Server Queue (J. W. Cohen). SIAM Review, 13 (1):106–107, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keinert:1989:IKP

[Kei89a] F. Keinert. Inversion of kappa-plane transforms and applications in computer tomography. SIAM Review, 31 (2):273–298, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keinert:1989:IPT

[Kei89b] Fritz Keinert. Inversion of kplane transforms and applications in computer tomography. SIAM Review, 31(2): 273–298, June 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelley:1959:MCM

[Kel59] James E. Kelley, Jr. Modern computing methods. SIAM Review, 1(2):185–186, ????

1959. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelisky:1960:BRI

[Kel60a] Richard P. Kelisky. Book review: An Introduction to the Calculus of Finite Differences and Difference Equations (Kenneth S. Miller). SIAM Review, 2(3):223– 224, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1960:SBI

[Kel60b] Herbert B. Keller. Special block iterations with applications to Laplace and biharmonic difference equations. SIAM Review, 2(4): 277–287, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1964:PSD

[Kel64] Herbert B. Keller. Propagation of stress discontinuities in inhomogeneous elastic media. SIAM Review, 6 (4):356–382, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1967:BRE

[Kel67] Joseph B. Keller. Book review: Electromagnetic Diffraction and Propagation Problems (V. A. Fock).

SIAM Review, 9(1):132– 133, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kellner:1970:PF

[Kel70] Richard Kellner. A positive functional. SIAM Review, 12 (3):468, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1975:CBL

[Kel75a] H. B. Keller. Computation of boundary-layer flows. SIAM Review, 17(2):383, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelman:1975:BRM

[Kel75b] Robert B. Kelman. Book review: Mathematical Physiology: Blood Flow and Electrically Active Cells (H. Melvin Lieberstein). SIAM Review, 17(1):181–182, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1979:PPT

[Kel79a] Joseph B. Keller. Progress and prospects in the theory of linear wave propagation. SIAM Review, 21(2): 229–245, April 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelman:1979:LSF

[Kel79b] Robert B. Kelman. Least squares Fourier series solutions to boundary value problems. SIAM Review, 21(3): 329–338, July 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1980:PFU

[Kel80a] Joseph B. Keller. Plate failure under pressure. SIAM Review, 22(2):227–228, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1980:SBC

[Kel80b] Joseph B. Keller. Some bubble and contact problems. SIAM Review, 22(4): 442–458, October 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1982:TDQ

[Kel82] Joseph B. Keller. Timedependent queues. SIAM Review, 24(4):401–412, October 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kellogg:1983:BRC

[Kel83] R. B. Kellogg. Book review:

Computer Methods for Partial Differential Equations, Volume L. Elliptic Equations and the Finite-Element Method (Robert Vichnevetsky). SIAM Review, 25(3): 434–435, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1984:TAP

[Kel84a] J. B. Keller. Techniques and applications of path integration — Schulman, L. S. SIAM Review, 26(1): 125, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1984:BRT

[Kel84b] Joseph B. Keller. Book review: Techniques and Applications of Path Integration (Lawrence S. Schulman). SIAM Review, 26 (1):125, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1984:PSR

[Kel84c] Joseph B. Keller. Probability of a shutout in racquetball. SIAM Review, 26(2): 267–268, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kellogg:1984:BRN

[Kel84d] R. B. Kellogg. Book review: Nonlinear Equivalence, Reduction of PDEs to ODEs and Fast Convergent Numerical Methods (E. E.

Rosinger). SIAM Review, 26(3):439, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1985:SM

[Kel85] Joseph B. Keller. Semiclassical mechanics. SIAM Review, 27(4):485–504, December 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1986:BRS

[Kel86a] Joseph B. Keller. Book review: Stochastic Wave Propagation (Kazimierz Sobczyk). SIAM Review, 28(4):593– 594, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelman:1986:BRA

[Kel86b] Robert B. Kelman. Book review: Applications of Discrete and Continuous Fourier Analysis (H. Joseph Weaver). SIAM Review, 28 (1):109, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelton:1986:BRE

[Kel86c] W. David Kelton. Book review: Elements of Simulation (Byron J. T. Morgan). SIAM Review, 28(2): 262–265, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kellogg:1988:BRS

[Kel88a] R. B. Kellogg. Book review: The Shallow Water Wave Equations: Formulation, Analysis and Application (I. Kinnmark). SIAM Review, 30(3):517–518, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelly:1988:DNM

[Kel88b] S. Graham Kelly. Development of normal mode theory for multi-degree-of-freedom systems using an energy inner product. SIAM Review, 30 (2):302–304, June 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelly:1988:NFM

[Kel88c]  S. Graham Kelly.  Nat-

ural frequency and mode shape perturbations for linear multidegree of freedom systems. SIAM Review, 30 (4):634–638, December 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelsey:1989:TDI

[Kel89] Ralph Kelsey. Two determinant inequalities. SIAM Review, 31(4):675, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelly:1990:NPC

[Kel90a] S. Graham Kelly. Nonlinear phenomena in a column of liquid in a rotating manometer. SIAM Review, 32 (4):652–659, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelman:1990:BRT

[Kel90b] Robert B. Kelman. Book review: Theory of Discrete and Continuous Fourier Analysis (H. Joseph Weaver). SIAM Review, 32(3):507– 508, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelton:1990:BRP

[Kel90c] W. David Kelton. Book review: Principles of Random Variate Generation (John Dagpunar). SIAM Review, 32 (3):508–511, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelton:1990:PRV

[Kel90d] W. David Kelton. Principles of random variate generation (John Dagpunar). SIAM Review, 32(3):508– ??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelman:1992:ADF

[Kel92a] Robert B. Kelman. Algorithms for discrete Fourier transform and convolution (Richard Tolimieri, Myoung An, and Chao Lu). SIAM Review, 34(4):668– ??, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelman:1992:BRA

[Kel92b] Robert B. Kelman. Book review: Algorithms for Discrete Fourier Transform and Convolution (Richard Tolimieri, Myoung An, and Chao Lu). SIAM Review, 34(4): 668–669, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keller:1995:HMS

[Kel95]  Joseph B. Keller.  How

many shuffles to mix a deck? SIAM Review, 37 (1):88–89, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/26629.htm.

Kellogg:1996:BRC

[Kel96] Bruce Kellogg. Book review: CRC Standard Mathematical Tables and Formulae (Daniel Zwillinger, ed.). SIAM Review, 38(4):691– 692, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kellogg:1997:BRa

[Kel97a] R. B. Kellogg. Book reviews. SIAM Review, 39(2): 333–374, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97003.

Kellogg:1997:BRb

[Kel97b] R. B. Kellogg. Book reviews. SIAM Review, 39 (3):528–569, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97005.

Kellogg:1997:BRc

[Kel97c] R. B. Kellogg. Book reviews. SIAM Review, 39(4): 790–809, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kellogg:1997:BRd

[Kel97d] R. B. Kellogg. Book reviews. SIAM Review, 39 (4):790–809, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97007.

Kellogg:1998:BRa

[Kel98a] R. B. Kellogg. Book reviews. SIAM Review, 40 (1):146–181, March 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97009.

Kellogg:1998:BRb

[Kel98b] R. B. Kellogg. Book reviews. SIAM Review, 40(2): 391–431, June 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97011.

Kellogg:1999:BRa

[Kel99a] R. Bruce Kellogg. Book reviews. SIAM Review, 41 (1):165–195, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97023.

Kellogg:1999:BRb

[Kel99b] R. Bruce Kellogg. Book reviews. SIAM Review, 41 (2):373, June 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97028.

Kellogg:1999:BRc

[Kel99c] R. Bruce Kellogg. Book reviews. SIAM Review, 41 (3):605–633, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97033.

Kellogg:1999:BRI

[Kel99d] R. Bruce Kellogg. Book reviews introduction. SIAM Review, 41(4):807, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/97039.

Kellogg:2000:BR

[Kel00] R. Bruce Kellogg. Book reviews. SIAM Review, 42 (1):125–157, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97045.

Kelly:2017:ITO

[Kel17] Matthew Kelly. An introduction to trajectory optimization: How to do your own direct collocation. SIAM Review, 59(4):849– 904, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelley:2022:NMM

[Kel22]  C. T. Kelley.  Newton’s

method in mixed precision. SIAM Review, 64(1):191– 211, ???? 2022. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.
                                        1137/20M1342902.

Kemperman:1967:BRC

[Kem67] J. H. B. Kemperman. Book review: Coding Theorems of Information Theory. Second edition (J. Wolfowitz). SIAM Review, 9(3):605– 606, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kemperman:1985:BRP

[Kem85] J. H. B. Kemperman. Book review: Probability (A. N. Shiryayev). SIAM Review, 27(1):110, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kemperman:1992:BRR

[Kem92a] J. H. B. Kemperman. Book review: Random Walk in Random and Non-Random Environments (Pál Revesz). SIAM Review, 34(2):335, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kemperman:1992:RWR

[Kem92b] J. H. B. Kemperman. Random walk in random and non-random environments (Pal Revesz). SIAM Review, 34(2):335–??, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kennedy:1974:BRM

[Ken74] Kenneth W. Kennedy. Book review: A Mathematical Theory of Global Program Optimization (Marvin Schaefer). SIAM Review, 16(4): 565–566, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kenkel:1986:BRC

[Ken86a] N. C. Kenkel. Book review: Canonical Analysis — a Review with Application in Ecology (R. Gittins). SIAM Review, 28(2): 265–266, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kenkel:1986:BRE

[Ken86b] N. C. Kenkel. Book review: Environmental and Natural Resource Mathematics (Robert W. McKelvey, ed.). SIAM Review, 28 (4):594–595, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kenkel:1991:BRM

[Ken91a] N. C. Kenkel. Book review: Mathematical Modeling in Ecology: a Workbook for Students (Clark Jeffries). SIAM Review, 33(1): 150–151, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kenkel:1991:MME

[Ken91b] N. C. Kenkel. Mathematical modeling in ecology: a workbook for students (Clark Jeffries). SIAM Review, 33(1): 150–??, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kenkel:1993:BRM

[Ken93] N. C. Kenkel. Book review: Mathematical Ecology of Plant Species Competition: a Class of Deterministic Models for Binary Mixtures of Plant Genotypes (Anthony G. Pakes and R. A. Maller). SIAM Review, 35 (2):323–324, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kern:1991:BRS

[Ker91a] W. Kern. Book review: Simulated Annealing and Boltzmann Machines: a Stochastic Approach to Combinatorial Optimization and Neural Computing (Emile Aarts and Jan Korst). SIAM Review, 33(2):323, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kern:1991:SAB

[Ker91b] W. Kern. Simulated annealing and Boltzmann machines: a stochastic approach to combinatorial optimization and neural computing (Emile Aarts and Jan Korst). SIAM Review, 33(2): 323–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kershaw:1992:BRM

[Ker92] D. Kershaw. Book review:

Mathematical Analysis and Numerical Methods for Science and Technology Vol. 4. Integral Equations and Numerical Methods for Science and Technology (Robert Dautray and Jacques-Louis Lions) (John C. Amson translator). SIAM Review, 34(1): 132–134, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kesten:1965:BRA

[Kes65] H. Kesten. Book review:

Applications of Characteristic Functions (E. Lukacs and R. A. Laha). SIAM Review, 7(2):305, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kesten:1969:BRI

[Kes69] H. Kesten. Book review:

An Introduction to Probability Theory and Its Applications, Volume I, (William Feller). SIAM Review, 11 (1):96, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kesten:1975:BRC

[Kes75] Harry Kesten. Book review: Concentration Functions (W. Hengartner and R. Theodorescu). SIAM Review, 17(3): 576–577, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kestin:1985:BRF

[Kes85] Joseph Kestin. Book review: A First Course in the Mathematical Foundations of Thermodynamics (David R. Owen). SIAM Review, 27(1): 110–113, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kevorkian:1985:BRP

[Kev85] J. Kevorkian. Book review: Perturbation Methods in Heat Transfer (A. Aziz and T. Y. Na). SIAM Review, 27(4): 598–600, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kevorkian:1987:PTO

[Kev87] J. Kevorkian. Perturbation techniques for oscillatory systems with slowly varying coefficients. SIAM Review, 29 (3):391–461, September 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keyfitz:1973:MP

[Key73] Nathan Keyfitz. Mathematics and population. SIAM Review, 15(2):370–375, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Key:1994:BRI

[Key94] Eric S. Key. Book review: An Introduction to Probability and Stochastic Processes (M. A. Berger). SIAM Review, 36(1):128–129, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Konheim:1962:BRP

[KF62] Alan G. Konheim and Leopold Flatto. Book review: A Parking Problem (M. S. Klamkin, D. J. Newman and L. Shepp). SIAM Review, 4 (3):257–258, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krohn:1976:CDS

[KF76] B. J. Krohn and K. Fox.

Cholesky decomposition of symmetric idempotent matrices. SIAM Review, 18 (4):813, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaiser:1995:FGW

[KF95] Gerald Kaiser and Michael Frazier. A friendly guide to wavelets. SIAM Review, 37 (4):624–??, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krauthgamer:2006:PAM

[KF06] Robert Krauthgamer and Uriel Feige. A polylogarithmic approximation of the minimum Bisection. SIAM Review, 48(1):99–130, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1979:VIP

[KG79] M. S. Klamkin and L. Gerber. A volume inequality for a pair of associated simplexes. SIAM Review, 21(4): 569–570, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1989:BRI

[KG89] M. S. Klamkin and Carl C. Grosjean. Book review: An Inverse Potential Problem (M. S. Klamkin). SIAM Review, 31(4):678–682, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Konopelchenko:1995:IMI

[KG95] B. G. Konopelchenko and J. D. Gibbon. Introduction to multidimensional integrable equations. SIAM Review, 37 (2):257–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirsch:1997:IMT

[KG97] Andreas Kirsch and Charles Groetsch. An introduction to the mathematical theory of inverse problems. SIAM Review, 39(4):794– ??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krotov:1998:GMO

[KG98] Vadim F. Krotov and Mostafa Ghandehari. Global methods in optimal control theory. SIAM Review, 40(1): 152–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirmser:1972:TS

[KH72] P. G. Kirmser and K. K. Hu.

A triple sum. SIAM Review, 14(4):646–647, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kantrowitz:1985:BRM

[KH85] Mark Kantrowitz and T. M.

Hagstrom. Book review:

A Minimum Value (M. S. Klamkin). SIAM Review, 27 (2):255–256, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kearfott:1998:RGS

[KH98] R. Baker Kearfott and Eldon Hansen. Rigorous global search: Continuous problems. SIAM Review, 40(1): 153–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Khayyam:1971:BRL

[Kha71] Oah Tajmah Khayyam. Book review: Limit of a Probability Density Function (D. J. Newman). SIAM Review, 13 (4):572–573, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Khargonekar:1986:BRS

[Kha86] Pramod Khargonekar. Book review: System Theory: a Hilbert Space Approach (Avraham Feintuch and Richard Saeks). SIAM Review, 28(4):595–597, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Khargonekar:1987:BRC

[Kha87] Pramod P. Khargonekar. Book review: Control System Synthesis: a Factorization Approach (M. Vidyasagar).

SIAM Review, 29(4):658– 660, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Khargonekar:1988:BRC

[Kha88] Pramod P. Khargonekar.

Book review: A Course in H-Subinfinity Control Theory (Bruce A. Francis). SIAM Review, 30(2):335– 336, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krishnaiah:1963:NBC

[KHS63] P. R. Krishnaiah, Peter Hagis, Jr., and Leon Steinberg. A note on the bivariate chi distribution. SIAM Review, 5(2):140– 144, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Katok:1996:IMT

[KHS96] Anatole Katok, Boris Hasselblatt, and Richard Swanson. Introduction to the modern theory of dynamical systems. SIAM Review, 38(1): 180–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kieffer:1994:BRE

[Kie94]  John Kieffer.  Book re-

view: Elements of Information Theory (Thomas M.

Cover and Joy A Thomas). SIAM Review, 36(3):509– 511, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kijima:1993:NCR

[Kij93] Masaaki Kijima. Numerical calculation of ruin probabilities for skip-free Markov chains. SIAM Review, 35 (4):621–624, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2018:RSa

[Kil18a] Misha E. Kilmer. Research spotlights. SIAM Review, 60 (1):93, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2018:RSb

[Kil18b] Misha E. Kilmer. Research spotlights. SIAM Review, 60 (2):313, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2018:RSc

[Kil18c] Misha E. Kilmer. Research spotlights. SIAM Review, 60 (3):593, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2018:RSd

[Kil18d] Misha E. Kilmer. Research spotlights. SIAM Review, 60 (4):867–868, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2019:RSa

[Kil19a] Misha E. Kilmer. Research spotlights. SIAM Review, 61 (1):77, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2019:RSb

[Kil19b] Misha E. Kilmer. Research spotlights. SIAM Review, 61 (2):259, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2019:RSc

[Kil19c] Misha E. Kilmer. Research spotlights. SIAM Review, 61 (3):441, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2019:RSd

[Kil19d] Misha E. Kilmer. Research spotlights. SIAM Review, 61 (4):731, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2020:RSa

[Kil20a] Misha E. Kilmer. Research spotlights. SIAM Review, 62 (1):131, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2020:RSb

[Kil20b] Misha E. Kilmer. Research spotlights. SIAM Review, 62 (2):351, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2020:RSc

[Kil20c] Misha E. Kilmer. Research spotlights. SIAM Review, 62 (3):615, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2020:RSd

[Kil20d] Misha E. Kilmer. Research spotlights. SIAM Review, 62 (4):779–780, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2021:RSa

[Kil21a] Misha E. Kilmer. Research spotlights. SIAM Review, 63 (1):65, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2021:RSb

[Kil21b] Misha E. Kilmer. Research spotlights. SIAM Review, 63 (2):315, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2021:RSc

[Kil21c] Misha E. Kilmer. Research spotlights. SIAM Review, 63 (3):487, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2021:RSd

[Kil21d] Misha E. Kilmer. Research spotlights. SIAM Review, 63 (4):721, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kilmer:2022:RSa

[Kil22a] Misha E. Kilmer. Research spotlights. SIAM Review, 64(1):57–58, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N97538X.

Kilmer:2022:RSb

[Kil22b] Misha E. Kilmer. Research spotlights. SIAM Review, 64(2):341–342, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975433.

Kilmer:2022:RSc

[Kil22c] Misha E. Kilmer. Research spotlights. SIAM Review, 64(3):623–624, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975512.

Kilmer:2022:RSd

[Kil22d]  Misha E. Kilmer.  Re-

search spotlights. SIAM Review, 64(4):919, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975561.

Kimme:1965:DPS

[Kim65] Ernest G. Kimme. Distributions of phases of sinusoidal components of ergodic trigonometric series. SIAM Review, 7(1):88–99, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

King:1976:BRP

[Kin76] J. Thomas King. Book review: Polynomial Approximation (Robert P. Feinerman and Donald J. Newman). SIAM Review, 18(1): 142–143, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

King:1982:BRI

[Kin82] J. Thomas King. Book review: Introduction to Numerical Analysis (J. Stoer and R. Bulirsch). SIAM Review, 24(1):96, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kingston:1988:OWW

[Kin88] J. G. Kingston. One-way waves. SIAM Review, 30 (4):645–649, December 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kincaid:1994:NCU

[Kin94a] D. R. Kincaid. Numerical computation using C (Robert Glassey). SIAM Review, 36 (4):672–??, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kincaid:1994:BRN

[Kin94b] David R. Kincaid. Book review: Numerical Computation Using C (Robert Glassey). SIAM Review, 36 (4):672–673, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

King:1994:BRM

[Kin94c]  J. Thomas King.  Book

review: Multigrid Methods (James H. Bramble).

SIAM Review, 36(3):512– 514, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kincaid:1997:BRA

[Kin97] David R. Kincaid. Book review: Afternotes on Numerical Analysis, by G. W. Stewart. SIAM Review, 39 (1):153, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirchgaessner:1975:BNH

[Kir75a] Klaus Kirchgaessner. Bifurcation in nonlinear hydrodynamic stability. SIAM Review, 17(4):652–683, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirchgassner:1975:BNH

[Kir75b] Klaus Kirchgässner. Bifurcation in nonlinear hydrodynamic stability. SIAM Review, 17(4):652–683, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirmser:1976:BRN

[Kir76] P. G. Kirmser. Book review: A Nonlinear Integral Equation (O. G. Ruehr). SIAM Review, 18(4):769– 770, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirkpatrick:1990:BRC

[Kir90a] T. R. Kirkpatrick. Book review: Classical Equilibrium Statistical Mechanics (Colin J. Thompson). SIAM Review, 32(3):511, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirkpatrick:1990:CES

[Kir90b] T. R. Kirkpatrick. Classical equilibrium statistical mechanics (Colin J Thompson). SIAM Review, 32(3): 511–??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirchgassner:1993:BRI

[Kir93a] Klaus Kirchgässner. Book review: Instabilities and Fronts in Extended Systems (Pierre Collet and Jean-Pierre Eckmann). SIAM Review, 35(1): 154–155, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirchgassner:1993:IFE

[Kir93b] Klaus Kirchgassner. Instabilities and fronts in extended systems (Pierre Collet and Jean-Pierre Eckmann). SIAM Review, 35(1): 154–??, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirby:2010:FAI

[Kir10] Robert C. Kirby. From functional analysis to iterative methods. SIAM Review, 52 (2):269–293, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirkinis:2012:RGP

[Kir12] Eleftherios Kirkinis. The renormalization group: a perturbation method for the graduate curriculum. SIAM Review, 54(2):374– 388, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karrila:1987:BRT

[KJ87] Seppo Karrila and A. A. Jagers. Book review: Two Algebraic Problems (John H. Halton). SIAM Review, 29 (2):299–301, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karrila:1987:TAP

[KJH87] S. Karrila, A. A. Jagers, and J. H. Halton. Two algebraic problems. SIAM Review, 29 (2):299–301, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kittappa:1975:ASP

[KK75] R. Kittappa and R. E. Kleinman. Acoustic scattering by penetrable homogeneous objects. SIAM Review, 17 (2):383, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kagiwada:1978:TTE

[KK78] H. Kagiwada and R. Kalaba. Team theory and extension of Sobolev–Bellman– Krein formula for Fredholm resolvents. SIAM Review, 20 (3):628, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelly:1998:NVS

[KK98] Terrence K. Kelly and Michael Kupferschmid. Numerical verification of secondorder sufficiency conditions for nonlinear programming. SIAM Review, 40(2):310– 314, June 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/30205.

Karypis:1999:PMS

[KK99] George Karypis and Vipin Kumar. Parallel multilevel series k-way partitioning Scheme for irregular graphs. SIAM Review, 41(2):278–300, June 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33413.

Kim:2009:TF

[KKBG09] Seung-Jean Kim, Kwangmoo Koh, Stephen Boyd, and Dimitry Gorinevsky. ℓ_{1} trend filtering. SIAM Review, 51 (2):339–360, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kracht:1982:IOF

[KKS82] M. Kracht, E. Kreyszig, and G. Schröder. Integral operators in the function theoretic method for partial differential equations. SIAM Review, 24(1):52–71, January 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krasovskii:1995:CUL

[KKY95] A. N. Krasovskii, N. N. Krasovskii, and Jiongmin Yong. Control under lack of information. SIAM Review, 37(4):640–??, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaper:1975:URC

[KL75] H. G. Kaper and G. K. Leaf. Use of rational coordinate functions for interpolation and approximation over a convex quadrilateral. SIAM Review, 17(2):383, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kailath:1976:NAS

[KL76] T. Kailath and L. Ljung. New approach to solution of linear integral equations with nondisplacement kernels. SIAM Review, 18(4):811, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1980:DP

[KL80] M. S. Klamkin and A. Liu. A dice problem. SIAM Review, 22(1):99, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knuth:1987:BRR

[KL87] D. E. Knuth and O. P.

Lossers. Book review: A Random Knockout Tournament (D. E. Knuth). SIAM Review, 29(1):127– 129, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karro:1997:UEA

[KL97a] J. Karro and C. K. Li. A unified elementary approach to canonical forms of matrices. SIAM Review, 39(2): 305–309, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Karro:1997:CNU

[KL97b] John Karro and Chi-Kwong Li. Classroom note: a unified elementary approach to canonical forms of matrices. SIAM Review, 39(2): 305–309, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/29486.

Kaloshin:2008:GAD

[KL08] Vadim Kaloshin and Mark Levi. Geometry of Arnold diffusion. SIAM Review, 50(4): 702–720, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1959:SEA

[Kla59] M. S. Klamkin. The smallest escape asteroid. SIAM Review, 1(2):172, ???? 1959. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1960:CGP

[Kla60] M. S. Klamkin. A center of gravity perturbation. SIAM Review, 2(1):41, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1961:CR

[Kla61] Murray S. Klamkin. On cooking a roast. SIAM Review, 3 (2):167–169, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1962:PHF

[Kla62a] M. S. Klamkin. A property of harmonic functions. SIAM Review, 4(4):396, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1962:TCB

[Kla62b] Murray S. Klamkin. On the transformation of a class of boundary value problems into initial value problems for ordinary differential equations. SIAM Review, 4(1): 43–47, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1963:BRO

[Kla63a]  M. S. Klamkin.  Book

review: One-Dimensional Steady-State Ablation (Tom Munson). SIAM Review, 5 (1):76–77, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1963:IP

[Kla63b] M. S. Klamkin. An infinite permutation. SIAM Review, 5(3):275–276, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1964:BVP

[Kla64a] M. S. Klamkin. A boundary value problem. SIAM Review, 6(2):177–178, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1964:PCS

[Kla64b] M. S. Klamkin. A physical characterization of a sphere. SIAM Review, 6 (1):61, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1964:PO

[Kla64c] Murray S. Klamkin. On a probability of overlap. SIAM Review, 6(3):311, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1965:BVP

[Kla65] M. S. Klamkin. A boundary value problem. SIAM Review, 7(4):566–567, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1966:DRD

[Kla66]  Murray S. Klamkin.  A

duality relation in differential equations and some associated functional equations. SIAM Review, 8(4): 494–500, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1970:CUR

[Kla70a]  M. S. Klamkin.  Con-

ductors of unit resistance. SIAM Review, 12(2):289– 290, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1970:DP

[Kla70b] M. S. Klamkin. On the diagonals of a polygon. SIAM Review, 12(4):581, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1971:EV

[Kla71a] M. S. Klamkin. An expected value. SIAM Review, 13 (2):248, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1971:I

[Kla71b] M. S. Klamkin. An inequality. SIAM Review, 13 (4):572, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1971:DP

[Kla71c] M. S. Klamkin. On the diagonals of a polygon. SIAM Review, 13(3):388, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1973:PDE

[Kla73] M. S. Klamkin. A polynomial Diophantine equation. SIAM Review, 15(1):220, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1974:MAP

[Kla74a] M. S. Klamkin. Moving axes and the principle of the complementary function. SIAM Review, 16(3):295– 302, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1974:OS

[Kla74b] M. S. Klamkin. An optimal strategy. SIAM Review, 16 (4):547, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1976:OLD

[Kla76a] M. S. Klamkin. An n-th order linear differential equation. SIAM Review, 18(1): 118, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1976:ELF

[Kla76b] M. S. Klamkin. On extreme length flight paths. SIAM Review, 18(3):486– 488, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1977:TPT

[Kla77] M. S. Klamkin. A two point triangle inequality. SIAM Review, 19(2):329, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1978:TPT

[Kla78a] M. S. Klamkin. A two point triangle inequality. SIAM Review, 20(2):400– 401, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1978:VIP

[Kla78b] M. S. Klamkin. A volume inequality for a pair of associated simplexes. SIAM Review, 20(4):856, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1979:TI

[Kla79] M. S. Klamkin. A triangle inequality. SIAM Review, 21 (4):559, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1980:BRT

[Kla80] M. S. Klamkin. Book review:

A Triangle Inequality (M. S. Klamkin). SIAM Review, 22 (4):509–511, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1982:FIW

[Kla82] M. S. Klamkin. Flight in an irrotational wind field. II. SIAM Review, 24(3): 340, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1983:BRF

[Kla83a] M. S. Klamkin. Book review: Flight in an Irrotational Wind Field. II (M. S. Klamkin). SIAM Review, 25 (3):406–407, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1983:II

[Kla83b] M. S. Klamkin. An isoperimetric inequality. SIAM Review, 25(1):98, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:CBC

[Kla84a]  M. S. Klamkin.  Calcu-

lus — basic concepts for high schools — Tarasov, L. V. SIAM Review, 26(4): 611–612, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:DSM

[Kla84b] M. S. Klamkin. Discrete and system models — Lucas, W. F., Roberts, F. S., Thrall, R. M. SIAM Review, 26(4):

606–608, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:FLA

[Kla84c] M. S. Klamkin. Fundamentals of linear algebra and analytical geometry — Bugrov, Y. S., Nokolsky, S. M. SIAM Review, 26(4):595, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:II

[Kla84d] M. S. Klamkin. An isoperimetric inequality. SIAM Review, 26(2):275–276, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:MV

[Kla84e] M. S. Klamkin. A minimum value. SIAM Review, 26(2):273, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:PSN

[Kla84f] M. S. Klamkin. A problem seminar — Newman, D. J. SIAM Review, 26(2): 286–289, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:BRP

[Kla84g] Murray S. Klamkin. Book review: A Problem Seminar (Donald J. Newman). SIAM Review, 26(2):286– 289, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:BRC

[Kla84h] Murray S. Klamkin. Book review: Calculus: Basic Concepts for High Schools (L. V. Tarasov). SIAM Review, 26 (4):611–612, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:BRD

[Kla84i] Murray S. Klamkin. Book review: Discrete and System Models (William F. Lucas, Fred S. Roberts and Robert M. Thrall, eds.). SIAM Review, 26(4):606– 608, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1984:BRF

[Kla84j] Murray S. Klamkin. Book review: Fundamentals of Linear Algebra and Analytical Geometry (Ya. S. Bugrov and S. M. Nokolsky). SIAM Review, 26(4):595, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1985:IS

[Kla85] M. S. Klamkin. Inequality for a simplex. SIAM Review, 27(4):576, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1986:BRI

[Kla86a]  M. S. Klamkin.  Book

review: Inequality for a Simplex (M. S. Klamkin). SIAM Review, 28(4):579– 580, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1986:IS

[Kla86b] M. S. Klamkin. Inequality for a simplex. SIAM Review, 28 (4):578–580, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1987:OMR

[Kla87] M. S. Klamkin. An optimum multistage rocket design. SIAM Review, 29(3): 470–471, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1988:BRO

[Kla88a] M. S. Klamkin. Book review: An Optimum Multistage Rocket Design (M. S. Klamkin). SIAM Review, 30 (3):508–509, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1988:CPM

[Kla88b] M. S. Klamkin. Characterizations of parabolic motion. SIAM Review, 30(1): 125, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1988:IPP

[Kla88c] M. S. Klamkin. An inverse potential problem. SIAM Review, 30(3):504, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1989:TPM

[Kla89] Murray S. Klamkin. Three particles moving on a circle. SIAM Review, 31(2): 320, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1990:BRP

[Kla90a] M. S. Klamkin. Book review: Polynomials (E. J. Barbeau). SIAM Review, 32 (4):699–700, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1990:CPL

[Kla90b]  M. S. Klamkin.  Cyclic

pursuit with lead angle (quickie). SIAM Review, 32 (4):673, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1990:II

[Kla90c] M. S. Klamkin. An integral inequality. SIAM Review, 32 (1):141, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1991:PMT

[Kla91] M. S. Klamkin. Point of minimum temperature. SIAM Review, 33(3):474, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1992:EGA

[Kla92a]  M. S. Klamkin.  Ex-

treme gravitational attraction. SIAM Review, 34(1): 120–121, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1992:BRU

[Kla92b] Murray S. Klamkin. Book review: Unsolved Problems in Intuitive Mathematics, Vol. II, Unsolved Problems in Geometry (H. T. Croft, K. J. Falconer, and R. K Guy). SIAM Review, 34(2): 335–338, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1994:CHF

[Kla94a] M. S. Klamkin. A conjectured heat flow problem. SIAM Review, 36(1):

107, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1994:PCU

[Kla94b] M. S. Klamkin. A possible characterization of uniformly accelerated motion. SIAM Review, 36(2):278, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1995:NSC

[Kla95a]  M. S. Klamkin.  Non-

symmetric cyclic pursuit on a sphere. SIAM Review, 37 (1):99, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1995:CUA

[Kla95b] Murray S. Klamkin. A characterization of uniformly accelerated motion. SIAM Review, 37(3):441–442, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kocic:1995:GBN

[KLA95c] V. L. Kocic, G. Ladas, and Ravi P. Agarwal. Global behavior of nonlinear difference equations. SIAM Review, 37 (1):114–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1997:BRH

[Kla97] Murray S. Klamkin. Book review: Handbook of Exact Solutions for Ordinary Differential Equations by A. D. Polyanin and V. F. Zaitsev. SIAM Review, 39 (2):350, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klein:1963:DI

[Kle63] Robert Klein. On a definite integral. SIAM Review, 5(1):73, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klee:1971:BRC

[Kle71] Victor Klee. Book review:

Convex Analysis (R. Tyrrell Rockafellar). SIAM Review, 13(2):233–238, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klee:1972:BRC

[Kle72] Victor Klee. Book review:

Convexity and Optimization in Finite Dimension (Josef Stoer and Christoph Witzgall). SIAM Review, 14(2): 357–359, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kleitman:1975:BRP

[Kle75] Daniel Kleitman. Book review: A Probabilistic Inequality (F. W. Steutel). SIAM Review, 17(4):693– 694, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klee:1976:BRC

[Kle76] Victor Klee. Book review: Convex Functions (A. Wayne Roberts and Dale E. Varberg). SIAM Review, 18(1): 133–134, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klebanov:1979:FE

[Kle79] L. B. Klebanov. A functional equation. SIAM Review, 21 (2):256, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klein:1983:CSD

[Kle83] Morton Klein. Complementary slackness and dynamic programming. SIAM Review, 25(2):249–253, April 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kimeldorf:1969:BRS

[KLG69] George S. Kimeldorf, Harry Lass, and Peter Gottlieb. Book review: A Sequence of Symmetric Bernoulli Trials (Y. P. Sabharwal). SIAM Review, 11(3):406–409, ???? 1969. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Klip:1978:EDB

[Kli78] D. A. Klip. Empirical data for base-30 integer GCD algorithm relative to classical Euclidean algorithm. SIAM Review, 20(3):628, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klosek:1993:BRM

[Klo93] Malgorzta M. Klosek. Book review: Markov Processes: an Introduction for Physical Scientists (Daniel T. Gillespie). SIAM Review, 35(2): 324–325, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolda:2003:ODS

[KLT03] Tamara G. Kolda, Robert Michael Lewis, and Virginia Torczon. Optimization by direct search: New perspectives on some classical and modern methods. SIAM Review, 45 (3):385–482, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolda:2004:ODS

[KLT04] Tamara G. Kolda, Robert Michael Lewis, and Virginia Torczon. Optimization by direct search: New perspectives on some classical and modern methods. SIAM Review, 45 (3):385–482, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42889.

Klamkin:1980:ICN

[KM80] M. S. Klamkin and A. Meir. An identity for complex numbers. SIAM Review, 22 (3):364, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

King:1985:AFD

[KM85] A. C. King and M. E. Mortimer. Approximating factorials: a difference equation approach. SIAM Review, 27 (4):565–567, December 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kulisch:1986:ADC

[KM86] Ulrich W. Kulisch and Willard L. Miranker. The arithmetic of the digital computer: a new approach. SIAM Review, 28(1):1–40, March 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krasnoselskii:1996:ANO

[KM96] Alexander M. Krasnosel’skii and Jean Mawhin. Asymptotics of nonlinearities and operator equations. SIAM Review, 38(3):533–??, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Koblitz:2004:SPK

[KM04] Neal Koblitz and Alfred J.

Menezes. A survey of public-key cryptosystems. SIAM Review, 46(4):599– 634, December 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43919.

Kimura:2012:GDM

[KM12] Margot Kimura and Jeff Moehlis. Group decision making models for sequential tasks. SIAM Review, 54(1): 121–138, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v54/i1/p121_s1.

Kollar:2014:GKS

[KM14] Richard Kollár and Peter D. Miller. Graphical Krein signature theory and Evans–Krein functions. SIAM Review, 56(1): 73–123, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kersch:1996:TSM

[KMC96] Alfred Kersch, William J.

Morokoff, and Carlo Cercignani. Transport simulation in microelectronics. SIAM Review, 38(2): 337–??, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kelman:1966:NSL

[KMH66] Robert B. Kelman, Donald J. Marsh, and Henry C. Howard. Nonmonotonicity of solutions of linear differential equations occurring in the theory of urine formation. SIAM Review, 8(4): 463–478, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1990:PMC

[KMH^{+}90] M. S. Klamkin, W. W. Meyer, F. Hezhi, S. G. Kelly, and O. P. Lossers. 3 particles moving on a circle. SIAM Review, 32(2): 307–309, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kittenberger:2022:COT

[KMS22] Axel Kittenberger, Leonidas Mindrinos, and Otmar Scherzer. Computed origami tomography. SIAM Review, 64 (2):469–484, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1378247.

Klamkin:1961:FIW

[KN61a] M. S. Klamkin and D. J.

Newman. Flight in an irrotational wind field. SIAM Review, 3(1):72, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1961:PAT

[KN61b] M. S. Klamkin and D. J.

Newman. The philosophy and applications of transform theory. SIAM Review, 3(1): 10–36, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1967:ERS

[KN67] Murray S. Klamkin and Donald J. Newman. Extended reducibility of some differential operators. SIAM Review, 9 (3):577–580, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Keeler:1998:CGC

[KN98] Stephen P. Keeler and Yves Nievergelt. Computing geodetic coordinates. SIAM Review, 40(2):300– 309, June 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32392.

Knight:1984:BMD

[Kni84a] F. Knight. Brownian motion and diffusion — Freedman, D. SIAM Review, 26 (2):302–303, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knight:1984:BRB

[Kni84b] Frank Knight. Book review: Brownian Motion and Diffusion (David Freedman). SIAM Review, 26(2):302– 303, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knight:1986:BRB

[Kni86] Frank B. Knight. Book review: Brownian Motion and Stochastic Flow Systems (J. Michael Harrison). SIAM Review, 28(3):424– 426, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knight:1994:SAB

[Kni94a] F. Knight. Some aspects of Brownian motion, part 1 (Marc Vor). SIAM Review, 36(3):511–??, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knight:1994:BRS

[Kni94b] Frank Knight. Book review: Some Aspects of Brownian Motion, Part 1 (Marc Yor). SIAM Review, 36(3): 511–512, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knops:1981:BRT

[Kno81] R. J. Knops. Book review:

Three-Dimensional Problems of the Mathematical Theory of Elasticity and Thermoelasticity (V. D. Kupradze, T. G. Gegelia, M. O. Basheleishvili, and T. V. Burchuladze). SIAM Review, 23(4):543– 545, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knops:1983:BRD

[Kno83] R. J. Knops. Book review: Dynamic Coupled Thermoelasticity (Ranjit S. Dhaliwal and Avtar Singh). SIAM Review, 25(1):126– 127, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knoerr:1998:DPM

[Kno98] Alan P. Knoerr. A dynamical proof of the method of Lagrange. SIAM Review, 40 (4):941–944, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32833.

Klamkin:1960:PP

[KNS60] M. S. Klamkin, D. J. Newman, and L. Shepp. A parking problem. SIAM Review, 2(3):219, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knuth:1977:BRR

[Knu77a] D. E. Knuth. Book review: A Reverse Card Shuffle (David Berman and M. S. Klamkin). SIAM Review, 19 (4):739–740, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knuth:1977:RCS

[Knu77b] D. E. Knuth. Reverse card shuffle. SIAM Review, 19(4): 739–741, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knuth:1980:BRA

[Knu80] Donald K. Knuth. Book review: Asymptotic Behavior of a Sequence ( L. Erlebach and O. Ruehr). SIAM Review, 22(1):101– 102, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knuth:1986:RKT

[Knu86]  D. E. Knuth.  A ran-

dom knockout tournament.

SIAM Review, 28(1):85– 86, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knuth:1987:RKT

[Knu87]  D. E. Knuth.  A ran-

dom knockout tournament. SIAM Review, 29(1):127– 129, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Knuth:1998:TBC

[Knu98] D. E. Knuth. Three binomial convolutions. SIAM Review, 40(4):990–991, November 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Koblitz:2010:BRD

[Kob10] Neal Koblitz. Book review:

Decrypted Secrets: Methods and Maxims of Cryptology. Fourth Edition. By Friedrich L. Bauer. SIAM Review, 52 (4):777–779, December 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kocak:1989:BRD

[Koç89] Hüseyin Koçak. Book review:

Deterministic Chaos: an Introduction; Second revised Revised Edition (Heinz Georg Schuster). SIAM Review, 31 (1):144–145, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kocak:1992:NMS

[Koc92] Huseyin Kocak. Natural minimal surfaces: Via theory and computation (with David Hoffman). SIAM Review, 34 (4):691–??, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kocak:1994:BRD

[Koç94] Hüseyin Koçak. Book review: Differential Equations Laboratory Workbook: a Collection of Experiments, Explorations and Modeling Projects for the Computer (Robert L. Borrelli, Courtney Coleman, and William E. Boyce). SIAM Review, 36(1): 129–130, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kodama:1996:BRO

[Kod96] Yuji Kodama. Book review: Optimization and Dynamical Systems (Uwe Helmke and John B. Moore). SIAM Review, 38(3):531– 533, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Koepf:1995:BRI

[Koe95] Wolfram Koepf. Book review: Introduction to Maple (André Heck). SIAM Review, 37(1):

132–135, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kogbetliantz:1966:BRH

[Kog66] Ervand G. Kogbetliantz.

Book review: Handbook for Computing Elementary Functions (L. A. Lyusternik, O. A. Chervonenkis, and A. R. Yanpol’skii). SIAM Review, 8(3):398–399, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kohn:1972:IRR

[Koh72]  W. Kohn.  Improvement

of Rayleigh–Ritz eigenfunctions. SIAM Review, 14(3): 399–419, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Koh:1985:BRC

[Koh85]  Eusebio L. Koh.  Book

review: The Convolution Product (Wilhelm Kecs). SIAM Review, 27(2):283– 285, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Koh:1986:BRI

[Koh86] Eusebio L. Koh. Book review:

Integral Transforms and their Applications — Second Edition (B. Davies). SIAM Review, 28(2):266, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kohn:1988:BRD

[Koh88]  Robert V. Kohn.  Book

review: Design Sensitivity Analysis of Structural Systems (Edward J. Haug, Kyung K. Choi, and Vadim Komkov). SIAM Review, 30 (3):518–519, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kohaupt:1998:BEP

[Koh98a] L. Kohaupt. Basis of eigenvectors and principal vectors associated with Gauss–Seidel matrix of A = tridiag[−12 − 1]. SIAM Review, 40(4): 959–964, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33138.

Kohaupt:1998:CCF

[Koh98b]  L. Kohaupt.  Calcula-

tion of cam-form errors. SIAM Review, 40(2):292– 299, June 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/32251.

Kokotovic:1984:ASP

[Kok84] Petar V. Kokotović. Applications of singular perturbation techniques to control problems. SIAM Review, 26(4): 501–550, October 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolodnea:1963:BRE

[Kol63] I. Kolodnea. Book review:

Elements of Functional Analysis (L. A. Liusternik and V. J. Sobolev). SIAM Review, 5 (1):80–81, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolodner:1965:BRP

[Kol65] I. I. Kolodner. Book review:

On a Periodic Solution of a Differential Equation (G. W. Veltkamp). SIAM Review, 7 (2):290–293, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolodner:1976:BRI

[Kol76] I. I. Kolodner. Book review: An Integral Inequality (J. Walter). SIAM Review, 18 (4):766–767, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolbig:1983:BRP

[Köl83] K. S. Kölbig. Book review: Polylogarithms and Associated Functions (Leonard Lewin). SIAM Review, 25(1): 133–136, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolbig:1987:BRC

[Köl87] K. S. Kölbig. Book review: Calculation of Special Functions, the Gamma Function, the Exponential Integrals and Error-Like Functions (C. G. van der Laan and N. M. Temme). SIAM Review, 29 (4):660–661, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolodner:1991:CSC

[Kol91] Ignace I. Kolodner. Convexity for the self-circumference of a unit disk. SIAM Review, 33(2):277, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolodner:1992:CSC

[Kol92] Ignace I. Kolodner. Convexity for the self circumference of a unit disk. SIAM Review, 34(2):317–323, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Konheim:1962:NBS

[Kon62] Alan G. Konheim. The number of bisecting strategies for the game Γ(N). SIAM Review, 4(4):379– 383, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kondic:2003:IGD

[Kon03] L. Kondic. Instabilities in gravity driven flow of thin fluid films. SIAM Review, 45(1):95–115, March 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40135.

Koopman:1977:BRT

[Koo77] Bernard O. Koopman. Book review: Theory of Optimal Search (Lawrence D. Stone). SIAM Review, 19(2): 361–364, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Korevaar:1965:BRD

[Kor65] Jacob Korevaar. Book review: Differential Equations of Mathematical Physics (N. S. Koshlyakov, M. M. Smirnov, and E. B. Gliner). SIAM Review, 7(3):440– 441, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Korevaar:1968:BRM

[Kor68] Jacob Korevaar. Book review: Mathematics for the Physical Sciences (Laurent Schwartz). SIAM Review, 10 (2):233–234, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Korevaar:1973:BRG

[Kor73] Jacob Korevaar. Book review: Generalized Integral Transformations (A. H. Zemanian). SIAM Review, 15 (1):232–234, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kortanek:1976:CPD

[Kor76] K. O. Kortanek. Constructing a perfect duality for semiinfinite programming and generalized linear programming. SIAM Review, 18(4): 812–813, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kortanek:1985:BRL

[Kor85]  K. O. Kortanek.  Book

review: Linear Optimization and Approximation (Klaus Glashoff and Sven-Åke Gustafson). SIAM Review, 27(2):285–286, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Korman:1991:UB

[Kor91] Philip Korman. An upper bound. SIAM Review, 33 (2):276, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Korman:1995:BRI

[Kor95a] Philip Korman. Book review: An Introduction to Partial Differential Equations (M. Renardv and R. C. Rogers). SIAM Review, 37 (3):469–470, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Korman:1995:BRA

[Kor95b] Philip Korman. Book review: Analysis and Control of Nonlinear Infinite Dimensional Systems (Viorel Barbu). SIAM Review, 37 (4):633, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Korman:1995:PSD

[Kor95c] Philip Korman. Periodic solutions of a differential equation. SIAM Review, 37(2): 236, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kosko:1979:DPJ

[Kos79a] E. Kosko. 3-dimensional pipe joining. SIAM Review, 21(4): 542–549, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kosko:1979:TDP

[Kos79b] Eryk Kosko. Three-dimensional pipe joining. SIAM Review, 21(4):542–549, October 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kotz:1965:SIC

[Kot65] Samuel Kotz. Some inequalities for convex functions useful in information theory. SIAM Review, 7(3): 395–402, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kovalyov:1996:BRG

[Kov96] Mikhael Kovalyov. Book review: Global Classical Solutions for Quasilinear Hyperbolic Systems (Li Ta-Tsien). SIAM Review, 38(2):347– 348, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kowalik:1983:BRP

[Kow83] Janusz S. Kowalik. Book review: Practical Optimization (Philip E. Gill, Walter Murray and Margaret H. Wright). SIAM Review, 25 (2):283–284, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kowalik:1986:BRS

[Kow86] Janusz Kowalik. Book review: Supercomputers and Parallel Computation (D. J. Paddon, ed.). SIAM Review, 28(1):109–111, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kowalik:1988:BRA

[Kow88] Janusz Kowalik. Book review: ACTORS: a Model of Concurrent Computation in Distributed Systems (Gul Agha). SIAM Review, 30 (1):146, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kaas-Petersen:1987:CML

[KP87] Chr. Kaas-Petersen. Continuation methods as the link between perturbation analysis and asymptotic analysis. SIAM Review, 29(1): 115–120, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kruzik:2006:RDM

[KP06] Martin Kruzík and Andreas Prohl. Recent developments in the modeling, analysis, and numerics of ferromagnetism. SIAM Review, 48(3): 439–483, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kloeden:1995:NSS

[KPGR95] Peter E. Kloeden, Eckhard Platen, Matthias Gelbrich, and Werner Romisch. Numerical solution of stochastic differential equations. SIAM Review, 37(2):272– ??, June 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kloeden:1996:NSS

[KPSP96] P. E. Kloeden, E. Platen, H. Schurz, and Philip Protter. Numerical solution of SDE through computer experiments. SIAM Review, 38 (1):177–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kleinman:1974:BIE

[KR74] R. E. Kleinman and G. F. Roach. Boundary integral equations for the threedimensional Helmholtz equation. SIAM Review, 16(2): 214–236, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirk:1978:MES

[KR78] B. L. Kirk and B. W. Rust. Maximum entropy spectral analysis of biological population time series. SIAM Review, 20(3):628, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1985:I

[KR85] M. S. Klamkin and O. G. Ruehr. An identity. SIAM Review, 27(2):250, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kamenjarzh:1998:LAS

[KR98] Jacov A. Kamenjarzh and B. Daya Reddy. Limit analysis of solids and structures. SIAM Review, 40(1): 173–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krall:1966:MCE

[Kra66] Allan M. Krall. On the michailov criterion for exponential polynomials. SIAM Review, 8(2):184–187, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kranzer:1967:BRL

[Kra67] Herbert C. Kranzer. Book review: Linearized Analysis of One-Dimensional Magnetohydrodynamic Flows (Roy M. Gundersen). SIAM Review, 9 (1):140–142, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kramer:1968:BRC

[Kra68] Henry P. Kramer. Book review: A Counting Circuit Problem (John H. Rowland). SIAM Review, 10(1): 112–113, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krall:1970:RLM

[Kra70]  Allan M. Krall.  The

root locus method: a survey. SIAM Review, 12(1): 64–72, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krall:1980:TSP

[Kra80a] A. M. Krall. Tschebyscheff sets of polynomials which satisfy an ordinary differentialequation. SIAM Review, 22 (4):436–441, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krall:1980:CSP

[Kra80b] Allan M. Krall. Chebyshev sets of polynomials which satisfy an ordinary differential equation. SIAM Review, 22 (4):436–441, October 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kraft:1996:BRH

[Kra96a] Roger L. Kraft. Book review: Hyperbolicity & Sensitive Chaotic Dynamics at Homoclinic Bifurcations (J. Palis and F. Taken). SIAM Review, 38(2):348–349, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krattenthaler:1996:TCI

[Kra96b] C. Krattenthaler. Two combinatorial identities. SIAM Review, 38(1):149–150, ???? 1996. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Krantz:2024:PMB

[Kra24]  Steven G. Krantz.  The

Poincaré metric and the Bergman theory. SIAM Review, 66(2):355–367, May 2024. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kreith:1964:CTC

[Kre64]  Kurt Kreith.  Compari-

son theorems for constrained rods. SIAM Review, 6(1): 31–36, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kreith:1966:MAC

[Kre66] Kurt Kreith. Mathematics and arms control. SIAM Review, 8(3):363–369, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krener:1978:SOC

[Kre78]  A. J. Krener.  Singular

optimal control — analysis, synthesis and applications. SIAM Review, 20 (3):629, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kreith:1991:EHG

[Kre91]  Kurt Kreith.  Establish-

ing hyperbolic Green’s functions via Leibniz’s rule.

SIAM Review, 33(1):101– 105, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krener:1996:BRT

[Kre96] Arthur J. Krener. Book review: Theory of Chattering Control (M. I. Zelikin and V. F. Borisov). SIAM Review, 38(1):172–173, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krishnan:1967:NBC

[Kri67] Marakatha Krishnan. The noncentral bivariate chi distribution. SIAM Review, 9 (4):708–714, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kriegsmann:1976:AMG

[Kri76] G. A. Kriegsmann. Application of method of geometrical optics to scattering of plane electromagnetic waves off cylindrically confined cold plasmas. SIAM Review, 18 (4):813, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kristiansen:1977:MCP

[Kri77] G. K. Kristiansen. A matrix convergence problem. SIAM Review, 19(3):564, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kristiansen:1982:BRP

[Kri82] G. K. Kristiansen. Book review: A Polynomial Problem (P. B. Borwein). SIAM Review, 24(1):77–78, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krishnamoorthy:1985:BRU

[Kri85] M. S. Krishnamoorthy. Book review: The Unix System Guidebook: an Introductory Guide for Serious Users (Peter P. Silvester). SIAM Review, 27(1):113– 114, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krishnamoorthy:1986:BRF

[Kri86] Mukkai S. Krishnamoorthy.

Book review: FORTRAN-Scientific Subroutine Library (by Peerless Engineering Service). SIAM Review, 28(1): 111–112, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krishnamoorthy:1991:BRS

[Kri91] Mukkai S. Krishnamoorthy.

Book review: The Stable Marriage Problem: Structure and Algorithms (Dan Gusfield and Robert W. Irving). SIAM Review, 33(2):

323–324, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krishnarao:1997:PIV

[Kri97] P. Krishnarao. Problems: An integral for valuation of spread options (problem 97-2). SIAM Review, 39(1): 123, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krishnamoorthy:2010:ECT

[Kri10] Mukkai Krishnamoorthy. Elements of computation theory. SIAM Review, 52(4): 779, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krolak:1966:PKC

[Kro66] P. Krolak. A property of the Krolak–Cooper extension of Fibonaccian search. SIAM Review, 8(4):510– 517, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kruskal:1983:OSC

[Kru83]  Joseph B. Kruskal.  An

overview of sequence comparison: Time warps, string edits, and macromolecules. SIAM Review, 25(2):201– 237, April 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kuttler:1968:LBS

[KS68] J. R. Kuttler and V. G.

Sigillito. Lower bounds for Stekloff and free membrane eigenvalues. SIAM Review, 10(3):368–370, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Konheim:1973:BRE

[KS73] A. G. Konheim and R. T.

Shield. Book review: An Extremum Problem (R. Tapir). SIAM Review, 15(2):386– 387, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kupperman:1977:RPD

[KS77] Robert H. Kupperman and Harvey A. Smith. The role of population defense in mutual deterrence. SIAM Review, 19 (2):297–318, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kuttler:1984:ELT

[KS84] J. R. Kuttler and V. G.

Sigillito. Eigenvalues of the Laplacian in two dimensions. SIAM Review, 26(2): 163–193, April 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kirchgraber:1990:CBS

[KS90] U. Kirchgraber and D. Stoffer. Chaotic behaviour in simple dynamical systems. SIAM Review, 32(3):424– 452, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kenney:1994:BRU

[KS94] Charles Kenney and Roy Smith. Book review: Uncertain Models and Robust Control (A. Weinmann). SIAM Review, 36(1):127– 128, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kailath:1995:DST

[KS95] Thomas Kailath and Ali H. Sayed. Displacement structure — theory and applications. SIAM Review, 37 (3):297–386, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/24648.
htm.

Knupp:1995:FGG

[KSM95] P. Knupp, S. Steinberg, and C. Wayne Mastin. Fundamentals of grid generation. SIAM Review, 37(2): 264–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kearsley:1980:NLD

[KSS80a] D. A. Kearsley, W. N.

Selander, and W. Squire. A non-linear differentialequation. SIAM Review, 22 (3):364–365, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Klamkin:1980:DPM

[KSS80b] M. S. Klamkin, A. Sharma, and P. W. Smith. Determinant of a partitioned matrix. SIAM Review, 22(2): 230, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Koch:1996:CAP

[KSW96] Hans Koch, Alain Schenkel, and Peter Wittwer. Computerassisted proofs in analysis and programming in logic: a case study. SIAM Review, 38 (4):565–604, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28418.

Keener:1992:DSW

[KT92] James P. Keener and John J.

Tyson. The dynamics of scroll waves in excitable media. SIAM Review, 34(1): 1–39, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kuntz:2021:SDC

[KTSB21] Juan Kuntz, Philipp Thomas, Guy-Bart Stan, and Mauricio Barahona. Stationary distributions of continuous-time Markov chains: a review of theory and truncation-based approximations. SIAM Review, 63(1):3–64, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kuchuk:1987:TCI

[Kuc87] Fikri Kuchuk. Three convolution integrals. SIAM Review, 29(4):632–633, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kucak:1992:NMS

[Küç92] Huseyin Küçak. Natural minimal surfaces: Via theory and computation with David Hoffman. SIAM Review, 34 (4):691–692, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kuijlaars:2006:CAK

[Kui06] Arno B. J. Kuijlaars. Convergence analysis of Krylov subspace iterations with methods from potential theory. SIAM Review, 48(1): 3–40, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kullback:1960:BRI

[Kul60] Solomon Kullback. Book review: An Introduction to Combinatorial Analysis (John Riordan). SIAM Review, 2(1):54, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kullback:1965:BRP

[Kul65] Solomon Kullback. Book review: Probability, Random Variables, and Stochastic Processes (Athanasios Papoulis). SIAM Review, 7(3): 441–443, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kullback:1969:BRI

[Kul69] S. Kullback. Book review:

Information Theory and Reliable Communication (R. G. Gallager). SIAM Review, 11 (4):633–634, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kumar:1982:BRT

[Kum82] Pradeep Kumar. Book review: Thermodynamics and Statistical Mechanics (Peter T. Landsberg). SIAM Review, 24(3):363–364, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kumei:1988:BRA

[Kum88]  Sukeyuki Kumei.  Book

review: Applications of Lie Groups to Differential Equations (Peter J. Olver). SIAM Review, 30(2):336– 337, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kung:1993:BRM

[Kun93] Joseph P. S. Kung. Book review: Matroid Decomposition (K. Truemper). SIAM Review, 35(3):529–531, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kupperman:1962:BRT

[Kup62] Morton Kupperman. Book review: Testing Statistical Hypotheses (E. L. Lehmann). SIAM Review, 4(4):407, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kupperman:1975:BRL

[Kup75] Morton Kupperman. Book review: Linear Statistical Inference and Its Applications 2nd Edition (C. Radhakrishna Rao). SIAM Review, 17 (4):705–707, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kupershmidt:1994:BRS

[Kup94] B. A. Kupershmidt. Book review: Soliton Equations and Hamiltonian Systems (L. A. Dickey). SIAM Review, 36(1):130–131, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kurtz:1969:BRP

[Kur69] Thomas G. Kurtz. Book review: Problems in Probability Theory, Mathematical Statistics and Theory of Random Functions (A. A. Sveshnikov, ed.). SIAM Review, 11(2):290, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kushner:1972:BRI

[Kus72] Harold J. Kushner. Book review: Introduction to Stochastic Control Theory (Karl J. Astrom). SIAM Review, 14(1):183, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kushner:1974:BRS

[Kus74] Harold Kushner. Book review: Stochastic Differential Equations (I. I. Gihman and A. V. Skorohod). SIAM Review, 16(2):271– 272, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kushner:1976:SSA

[Kus76] Harold J. Kushner. A survey of some applications of probability and stochastic control theory to finite difference methods for degenerate elliptic and parabolic equations. SIAM Review, 18(4): 545–577, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kushner:1977:C

[Kus77]  H. J. Kushner.  Correc-

tion. SIAM Review, 19 (1):U1, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kushner:1993:BRS

[Kus93] Harold Kushner. Book review: Stability of Stochastic Differential Equations with Respect to Semi-martingales (X. Mao). SIAM Review, 35 (4):670–671, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kushel:2019:UMS

[Kus19] Olga Y. Kushel. Unifying matrix stability concepts with a view to applications. SIAM Review, 61(4): 643–729, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kutz:2006:MLS

[Kut06] J. Nathan Kutz. Mode-locked soliton lasers. SIAM Review, 48(4):629–678, January 2006.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kan:1993:BRP

[Kv93] A. H. G. Rinnooy Kan and A. van Vliet. Book review: Probabilistic Analysis of Packing and Partitioning Algorithms (E. G. Coffman, Jr. and George S. Lueker). SIAM Review, 35(1):153–154, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kailath:1978:ITO

[KVM78] T. Kailath, A. Vieira, and M. Morf. Inverses of Toeplitz operators, innovations, and orthogonal polynomials. SIAM Review, 20 (1):106–119, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kallina:1971:LPR

[KW71] Carl Kallina and A. C.

Williams. Linear programming in reflexive spaces. SIAM Review, 13(3):350– 376, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kahaner:1976:ADA

[KW76] D. K. Kahaner and M. B.

Wells. Algorithm for Ndimensional adaptive quadrature. SIAM Review, 18 (4):811, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kruk:1999:PP

[KW99] Serge Kruk and Henry Wolkowicz. Pseudolinear programming. SIAM Review, 41 (4):795–805, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33525.

Kolokolnikov:2018:PFR

[KW18] Theodore Kolokolnikov and Juncheng Wei. Pattern formation in a reaction-diffusion system with space-dependent feed rate. SIAM Review, 60 (3):626–645, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kwapisz:1998:CNP

[Kwa98]  M. Kwapisz.  Classroom

note: The power of a matrix. SIAM Review, 40(3): 703–705, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32074.

Kall:1996:SP

[KWB96] Peter Kall, Stein W. Wallace, and John R. Birge. Stochastic programming. SIAM Review, 38(4):681–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kwapien:1995:RSS

[KWP95] Stanislaw Kwapien, Wojbar A. Woyczynski, and Philip Protter. Random series and stochastic integrals: Single and multiple. SIAM Review, 37(1): 135–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kyner:1976:OVO

[Kyn76] W. T. Kyner. Oscillator variables in orbital mechanics. SIAM Review, 18(4): 813, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kowalik:1993:BRN

[KZ93] Janusz S. Kowalik and Karel Zikan. Book review: Numerical Linear Algebra and Optimization: Vol. 1 (Philip E. Gill, Walter Murray, and Margaret H. Wright). SIAM Review, 35(1):155– 158, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leventhal:1975:NSM

[LA75] S. H. Leventhal and A. K.

Aziz. Numerical solution of mixed hyperbolic-elliptic equations. SIAM Review, 17 (2):384, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Logan:1995:INP

[LA95] J. D. Logan and Karen Ames. An introduction to nonlinear partial differential equations. SIAM Review, 37(2): 251–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laetsch:1976:BRT

[Lae76] Theodore Laetsch. Book review: Theory of Branching of Solutions of Non-Linear Equations (M. M. Vainberg and V. A. Trenogin). SIAM Review, 18(1):145– 146, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laetsch:1982:BRD

[Lae82] Theodore Laetsch. Book review: Differential Equations-Introduction and Qualitative Theory (Jane Cronin). SIAM Review, 24(3):358– 360, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laetsch:1983:AFA

[Lae83a] T. Laetsch. Applied functional analysis — Balakrishnan, A. V. SIAM Review, 25 (3):439–440, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laetsch:1983:BRA

[Lae83b] Theodore Laetsch. Book review: Applied Functional Analysis (A. V. Balakrishnan). SIAM Review, 25(3): 439–440, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laffin:1993:APS

[Laf93] Steven Laffin. All purpose stamp design. SIAM Review, 35(2):299, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lagarias:1982:MTL

[Lag82] J. C. Lagarias. Malevolent traffic lights. SIAM Review, 24(4):477, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lagarias:1989:BRC

[Lag89] Jeffrey C. Lagarias. Book review: A Course in Number Theory and Cryptography (Neil Koblitz). SIAM Review, 31(3):508–510, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lagnese:1997:BRE

[Lag97] John E. Lagnese. Book review: Exact Controllability and Stabilization: The Multiplier Method, by V. Komornik. SIAM Review, 39 (2):351–352, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laha:1994:BRI

[Lah94] R. G. Laha. Book review: Identifiability in Stochastic Models (B. L. S. Prakasa Rao). SIAM Review, 36(2): 304–307, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laing:1978:BRA

[Lai78] Richard Laing. Book review: Adaptive Information Processing (Jeffrey R. Sampson). SIAM Review, 20(2): 404–405, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lai:1991:BRO

[Lai91] Tze Leung Lai. Book review: Optimization of Stochastic Systems: Topics in Discrete-Time Dynamics (Masanao Aoki). SIAM Review, 33(4): 671–672, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lakshmikantham:1980:BRD

[Lak80] V. Lakshmikantham. Book review: Dynamical Systems and Their Applications: Linear Theory (John L. Casti). SIAM Review, 22(4):521– 522, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lamborn:1969:EDD

[Lam69] B. N. A. Lamborn. An expression for the Dirac delta function. I. SIAM Review, 11(4):603, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lamborn:1970:EDD

[Lam70] B. N. A. Lamborn. An expression for the Dirac delta function. II. SIAM Review, 12(4):567–569, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lambert:1985:CNO

[Lam85] J. P. Lambert. Constructing nodes for optimal quadrature. SIAM Review, 27 (2):249, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lambert:1986:BRN

[Lam86] Robert J. Lambert. Book review: Numerical Solution of Differential Equations-Second Edition (M. K. Jain). SIAM Review, 28(1):112– 113, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lampret:2004:IHI

[Lam04] Vito Lampret. An invitation to Hermite’s integration and summation: a comparison between Hermite’s and Simpson’s rules. SIAM Review, 46(2):311–328, June 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41630.

Lancaster:1964:BLR

[Lan64] P. Lancaster. Bounds for latent roots in damped vibration problems. SIAM Review, 6(2):121–125, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lansing:1967:AE

[Lan67a] Donald L. Lansing. An asymptotic expansion. SIAM Review, 9(2):251, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lansing:1967:BRA

[Lan67b] Donald L. Lansing. Book review: Asymptotic Expansion of a Laplace Transform (M. B. McNeil). SIAM Review, 9 (3):593–599, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lancaster:1970:ESL

[Lan70] Peter Lancaster. Explicit solutions of linear matrix equations. SIAM Review, 12(4): 544–566, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lanzano:1976:VIF

[Lan76a] B. C. Lanzano. Viking interpretive flight dynamics simulation program. SIAM Review, 18(4):813–814, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lanzano:1976:RFA

[Lan76b] P. Lanzano. Recursion formulas for Appell hypergeometric series of 1st kind and their application to an astrophysical problem. SIAM Review, 18(4):814, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lancaster:1981:BRM

[Lan81]  P. Lancaster.  Book re-

view: Minimal Factorization of Matrix and Operator Functions (H. Bart, I. Gohberg, and M. A. Kaashoek). SIAM Review, 23(4):551– 555, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lanchon:1986:BRM

[Lan86]  Hélène Lanchon.  Book

review: Les Methodes de L’homogénéisation: Théorie et Applications en Physique (D. J. Berg-Man, J. L. Lions, G. C. Papanicolaou, F. Murat, L. Tartar and E. Sanchez-Palencia). SIAM Review, 28(3):426– 427, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lardner:1969:RBB

[Lar69a] Thomas J. Lardner. Relations between _{0}F_{3} and Bessel functions. SIAM Review, 11 (1):69–72, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Larkin:1969:PST

[Lar69b] P. A. Larkin. The possible shapes of things to come. SIAM Review, 11 (1):1–6, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Larson:1974:BRN

[Lar74] Roland E. Larson. Book review: Nonserial Dynamic Programming (Umberto Bertelè and Francesco Brioschi). SIAM Review, 16 (4):571–572, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Larson:1978:BRD

[Lar78] Roland E. Larson. Book review: Dynamic Programming and Stochastic Control (Dimitri P. Bertsekas). SIAM Review, 20(2):403– 404, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Larsen:1980:SBF

[Lar80] J. Larsen. A sum of Bessel functions. SIAM Review, 22 (4):508–509, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lardner:1986:MVP

[Lar86] Thomas J. Lardner. Minimum value of a projectile velocity. SIAM Review, 28(3): 385–388, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Larsen:1994:BRT

[Lar94] Jesper Larsen. Book review: Two Definite Integrals Arising in Light Transmission Through a Crystal (M. L. Glasser). SIAM Review, 36 (1):111–112, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

LaSalle:1964:RAL

[LaS64] J. P. LaSalle. Recent advances in Liapunov stability theory. SIAM Review, 6 (1):1–11, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lassez:1976:CCS

[Las76] J. L. Lassez. Circular codes and synchronization. SIAM Review, 18(4):814, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lasserre:2007:SSA

[Las07] Jean B. Lasserre. A sum of squares approximation of nonnegative polynomials. SIAM Review, 49(4): 651–669, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Latta:1969:BRP

[Lat69] Gordon E. Latta. Book review: Perturbation Methods in Applied Mathematics (Julian D. Cole). SIAM Review, 11(3):412, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Latta:1990:BRH

[Lat90a] Gordon Latta. Book review: Handbook of Differential Equations (Daniel Zwilliger). SIAM Review, 32 (3):511–512, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Latta:1990:HDE

[Lat90b] Gordon Latta. Handbook of differential equations (Daniel Zwilliger). SIAM Review, 32 (3):511–??, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Latz:2023:BIP

[Lat23] Jonas Latz. Bayesian inverse problems are usually well-posed. SIAM Review, 65(3):831–865, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1556435.

Lau:1974:BRN

[Lau74] Brian R. Lau. Book review:

Non-Homogeneous Boundary Value Problems and Applications (J. L. Lions and E. Magenes). SIAM Review, 16(1): 101–102, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lau:1982:CNR

[Lau82] T. C. Y. Lau. A conjecture on the number of real zeros. SIAM Review, 24 (3):340, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lau:1983:CNR

[Lau83] T. C. Y. Lau. A conjecture on the number of real zeros. SIAM Review, 25 (2):267, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lau:1984:CNR

[Lau84] T. C. Y. Lau. A conjecture on the number of real zeros. SIAM Review, 26(2): 273–275, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lau:1998:COG

[Lau98] K. W. Lau. Connection of oscillation and growth rates of entire functions. SIAM Review, 40(3):730– 731, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laumakis:2003:DKC

[Lau03]  Paul J. Laumakis.  De-

signing a 401(k): a case study. SIAM Review, 45 (4):806–813, December 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39299.

Laub:2010:LST

[Lau10] Alan J. Laub. Linear systems theory. SIAM Review, 52(4): 779–781, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lawson:1965:BRS

[Law65] C. L. Lawson. Book review:

The Smallest Covering Cone or Sphere (C. Groenewod and L. Eusanio). SIAM Review, 7 (3):415–416, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lawson:1966:BRS

[Law66] J. D. Lawson. Book review: The Shape of Miler’s Lamp (Roland Silver). SIAM Review, 8(4):530–532, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lawson:1972:BRA

[Law72] Charles L. Lawson. Book review: The Approximation of Functions, Vol. II (John R. Rice). SIAM Review, 14(1): 187–188, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lawson:1978:ESS

[Law78] C. L. Lawson. Estimating slopes for C_{1} surface interpolation. SIAM Review, 20 (3):629, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lawler:1995:BRA

[Law95] Gregory F. Lawler. Book review: Aspects and Applications of the Random Walk (G. H. Weiss). SIAM Review, 37(3):470–471, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Law:2018:BRD

[Law18] K. J. H. Law. Book review: Data Assimilation: Methods, Algorithms, and Applications, by Mark Asch, Marc Bocquet, and Maëlle Nodet. SIAM Review, 60(2): 481–483, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lax:1967:IMC

[Lax67] P. D. Lax. II. the mathematics curriculum. SIAM Review, 9(2):312–325, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lax:1969:NPD

[Lax69] Peter D. Lax. Nonlinear partial differential equations and computing. SIAM Review, 11 (1):7–19, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lax:1975:NWM

[Lax75] P. D. Lax. Nonlinear-wave motion. SIAM Review, 17 (2):383, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lax:1976:APS

[Lax76] Peter D. Lax. Almost periodic solutions of the KdV equation. SIAM Review, 18 (3):351–375, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lax:1989:FAM

[Lax89] Peter D. Lax. The flowering of applied mathematics in America. reprinted from A century of mathematics, Part II, 455–466, Amer. Math. Soc., Providence, RI, 1989 [see MR 90a:01065]. SIAM Review, 31 (4):533–541, December 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Layton:2024:BRa

[Lay24a] Anita T. Layton. Book reviews. SIAM Review, 66(1): 193–201, February 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Layton:2024:BRb

[Lay24b] Anita T. Layton. Book reviews. SIAM Review, 66(2):

391–399, May 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lazer:1982:BRS

[Laz82] Alan Lazer. Book review: Sturmian Theory for Ordinary Differential Equations (William T. Reid). SIAM Review, 24(4):485– 489, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lazer:1991:BRC

[Laz91a] Alan C. Lazer. Book review: Critical Point Theory and Hamiltonian Systems (Jean Mawhin and Michel Willem). SIAM Review, 33 (2):324–327, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lazer:1991:CPT

[Laz91b] Alan C. Lazer. Critical point theory and Hamiltonian systems (Jean Mawhin and Michel Willem). SIAM Review, 33(2):324–??, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levy:1977:BFS

[LB77] P. Levy and B. C. Berndt. Bessel function summation. SIAM Review, 19(2):333– 334, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Little:1978:SHF

[LB78] F. F. Little and R. E. Barnhill. Surfaces and how to find them. SIAM Review, 20(3): 629–630, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lefrancois:2010:IFS

[LB10] Emmanuel Lefrançois and Jean-Paul Boufflet. An introduction to fluid-structure interaction: Application to the piston problem. SIAM Review, 52(4):747–767, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1986:NMD

[LBD^{+}86] G. N. Lewis, W. Bohm, A. A. Diwan, A. A. Jagers, Z. J. Kabala, and D. F. Lockhart. On a nonlinear matrix differential-equation. SIAM Review, 28(2):241– 242, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luke:2002:OWR

[LBL02] D. Russell Luke, James V.

Burke, and Richard G. Lyon. Optical wavefront reconstruction: Theory and numerical methods. SIAM Review, 44(2):169–224, June 2002. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic). URL

                                        http://epubs.siam.org/
                                        sam-bin/dbq/article/39075.

Leadbetter:1965:CCN

[LC65] M. R. Leadbetter and J. D. Cryer. Curve crossings by normal processes and reliability implications. SIAM Review, 7(2):241–250, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lagerstrom:1972:BCU

[LC72] P. A. Lagerstrom and R. G. Casten. Basic concepts underlying singular perturbation techniques. SIAM Review, 14(1):63–120, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lempel:1976:NCB

[LCE76] A. Lempel, M. Cohn, and W. L. Eastman. New class of balanced binary sequences with optimal autocorrelation properties. SIAM Review, 18 (4):815, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1979:PBW

[LCG^{+}79] J. S. Lew, F. H. Clarke, J. Grzesik, D. C. Jespersen, W. B. Jordan, and W. W. Meyer. Paul Bunyan’s washline. SIAM Review, 21(4):

565–566, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levy:2017:SES

[LCK^{+}17] Rachel Levy, Edmond Chow, Byong Kwon, Katherine Socha, Maeve McCarthy, and Peter R. Turner. SIAM Education Subcommittee report on undergraduate degree programs in applied mathematics. SIAM Review, 59(1): 199–204, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lakshmivarahan:1995:PCU

[LDG95] S. Lakshmivarahan, Sudarshan K. Dhall, and E. Gallopoulos. Parallel computing using the prefix problem. SIAM Review, 37(4): 626–??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

LeCam:1992:BRC

[Le 92] Lucien Le Cam. Book review: Comparison of Statistical Experiments (Erik Torgersen). SIAM Review, 34 (4):669–671, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leavens:1994:BRP

[Lea94] Gary T. Leavens. Book review: Programs, Recursion and Unbounded Choice (Wim H. Hesselink). SIAM Review, 36(1):131–133, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lebovitz:1982:PEP

[Leb82] N. R. Lebovitz. Perturbation expansions on perturbed domains. SIAM Review, 24(4): 381–400, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leddder:1996:ISP

[Led96] Glenn Leddder. An integral from a Stefan problem. SIAM Review, 38(1):147, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1977:BRC

[Lee77]  K. K. Lee.  Book re-

view: Classical Electrodynamics 2nd Ed. (John David Jackson). SIAM Review, 19 (1):169–170, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1982:BRF

[Lee82a] K. K. Lee. Book review: A First Course in Differential Geometry (Chuan-Chih Hsiung). SIAM Review, 24(3): 365–366, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1982:BRI

[Lee82b] K. K. Lee. Book review:

Introduction to Global Analysis (Donald W. Kahn). SIAM Review, 24(1):93– 95, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1984:BRB

[Lee84a] K. K. Lee. Book review:

Bäcklund Transformations and Their Applications (C. Rogers and W. F. Shadwick). SIAM Review, 26(1): 138–141, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1984:BRM

[Lee84b]  K. K. Lee.  Book re-

view: Manifolds, Tensor Analysis, and Applications (Ralph Abraham, Jerrold E. Marsden and Tudor Ratiu). SIAM Review, 26(4):590– 592, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1986:BRA

[Lee86a] K. K. Lee. Book review:

Applied Differential Geometry (William L. Burke). SIAM Review, 28(3):427– 428, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1986:CRM

[Lee86b] T. S. Lee. A complete russels method for the transportation problem. SIAM Review, 28(4):547–549, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1986:UMP

[Lee86c] Tze-San Lee. On the uniqueness of a minimum problem. SIAM Review, 28(3): 395, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1997:HCU

[Lee97a] J. Lee. Hoffman’s circle untangled. SIAM Review, 39 (1):98–105, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1997:CNH

[Lee97b] Jon Lee. Classroom note: Hoffman’s circle untangled. SIAM Review, 39(1):98– 105, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/29299;   http:
                                        //www.siam.org/journals/
                                        sirev/sirev391.htm.

Lefschetz:1965:RAS

[Lef65] Solomon Lefschetz. Recent advances in the stability of nonlinear controls. SIAM Review, 7(1):1–12, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lehman:1960:RNI

[Leh60]  Alfred Lehman.  A re-

sistor network inequality. SIAM Review, 2(2):152– 153, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lehman:1963:BRE

[Leh63a] A. Lehman. Book review:

Economic Applications of the Theory of Graphs (G. Avondo-Bodino). SIAM Review, 5(2):171–172, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lehman:1963:BRR

[Leh63b] A. Lehman. Book review:

Redundancy Techniques for Computing Systems (R. H. Wilcox and W. C Mann, eds.). SIAM Review, 5(2): 166, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lehman:1965:PNB

[Leh65] Alfred Lehman. Postulates for a normed Boolean algebra. SIAM Review, 7(2): 253–273, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lehmer:1982:BRE

[Leh82] D. H. Lehmer. Book review: Error-Free Computation: Why It is Needed and Methods for Doing It (Robert Todd Gregory). SIAM Review, 24(1):99– 100, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leibowitz:1966:BRD

[Lei66] Gerald M. Leibowitz. Book review: Distribution Theory and Transform Analysis (A. H. Zemanian). SIAM Review, 8(2):245–246, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leipnik:1976:ESM

[Lei76] R. B. Leipnik. Explicit solution of matrix Riccati differential equation in noncommuting case. SIAM Review, 18(4):814–815, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leipala:1978:ADU

[Lei78a]  Timo Leipälä.  Average

distance in a unit cube. SIAM Review, 20(2):394– 395, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leipnik:1978:DSQ

[Lei78b] R. B. Leipnik. Dirichlet series for quadratic difference equations. SIAM Review, 20 (3):629, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leipala:1979:ADU

[Lei79]  Timo Leipälä.  Average

distance in a unit cube. SIAM Review, 21(2):260– 263, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lempel:1970:SMI

[Lem70] M. J. Lempel. A stock market investment inequality. SIAM Review, 12(3):469, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lemoine:1983:APH

[Lem83a] A. Lemoine. Applied probability — Haight, F. SIAM Review, 25(3):418–419, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lemoine:1983:BRA

[Lem83b] Austin Lemoine. Book review: Applied Probability (Frank Haight). SIAM Review, 25(3):418–419, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lemke:1984:BRP

[Lem84] C. E. Lemke. Book review: Pathways to Solutions, Fixed Points, and Equilibria (C. B. Garcia and W. J. Zangwill). SIAM Review, 26(3): 445–446, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lemke:1985:BRI

[Lem85] C. E. Lemke. Book review: Introduction to Sensitivity and Stability Analysis in Nonlinear Programming (Anthony V. Fiacco). SIAM Review, 27(1):114– 115, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lemarechal:1988:BRM

[Lem88] Claude Lemarechal. Book review: Methods of Descent for Nondifferentiable Optimization (Krzysztof C. Kiwiel). SIAM Review, 30(1): 146–147, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lenard:1980:BRT

[Len80]  Andrew Lenard.  Book

review: Thermodynamic Formalism (David Ruelle). SIAM Review, 22(2):248– 250, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lengyel:1993:CIW

[Len93a] Tamás Lengyel. A combinatorial identity and the World Series. SIAM Review, 35(2): 294–297, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lenstra:1993:BRU

[Len93b] Arjen K. Lenstra. Book review: Uses of Randomness in Algorithms and Protocols (Joe Kilian). SIAM Review, 35(2):326–328, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lenstra:1993:URA

[Len93c] Arjen K. Lenstra. Uses of randomness in algorithms and protocols (Joe Kilian). SIAM Review, 35(2): 326–??, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lenhart:1995:BRM

[Len95] Suzanne Lenhart. Book review: Markov Models and Optimization (M. H. A. Davis). SIAM Review, 37(2): 262–263, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leonard:1978:CFM

[Leo78]  A. Leonard.  Computing

fluid motions with vortex elements. SIAM Review, 20 (3):629, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leonard:1996:ME

[Leo96] I. E. Leonard. The matrix exponential. SIAM Review, 38 (3):507–512, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28648.

Leray:1963:FTR

[Ler63] J. Leray. The functional transformations required by the theory of partial differential equations. SIAM Review, 5(4):321–334, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lerin:1964:BRD

[Ler64] J. J. Lerin. Book review: Differential Equations: Geometric Theory (Second Edition) (S. Lefschetz). SIAM Review, 6(1):79–80, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leslie:1976:BRI

[Les76]  A. S. Leslie.  Book re-

view: Idealized Optical Fibre Mode Conversion (R. W. Allen). SIAM Review, 18(2): 298–300, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Letcher:1969:BLT

[Let69] John S. Letcher, Jr. On boundary layers in the theory of heat conduction in solids. SIAM Review, 11(1): 20–29, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levin:1962:BRN

[Lev62] J. J. Levin. Book review:

Nonlinear Differential Equations (R. A. Struble). SIAM Review, 4(4):404, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levine:1963:BRS

[Lev63a] Leon Levine. Book review:

Systems Engineering Mathematics (J. E. Alexander and J. M. Bailey). SIAM Review, 5(2):164–165, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levy:1963:BRI

[Lev63b] B. R. Levy. Book review:

An Introduction to Phase-Integral Methods (J. Heading). SIAM Review, 5(3): 290, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levinson:1965:STI

[Lev65]  Norman Levinson.  Sim-

plified treatment of integrals of Cauchy type, the Hilbert problem and singular integral equations. appendix: Poincaré–Bertrand formula. SIAM Review, 7(4): 474–502, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levin:1967:BRO

[Lev67a] Jacob J. Levin. Book review: Ordinary Differential Equations (I. G. Petrovski). SIAM Review, 9(2): 266–267, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levit:1967:ZHP

[Lev67b]  R. J. Levit.  The ze-

ros of the Hahn polynomials. SIAM Review, 9(2): 191–203, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

LeVeque:1977:BRU

[LeV77a] William J. LeVeque. Book review: Uniform Distribution of Sequences (L. Kuipers and H. Niederreiter). SIAM Review, 19(1): 168–169, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levy:1977:BRB

[Lev77b]  P. Levy.  Book review:

A Bessel Function Summation (B. C. Berndt). SIAM Review, 19(2):333– 335, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levinger:1984:BRS

[Lev84] B. W. Levinger. Book review: A System of Probabilistic Inequalities (T. Ten Raa). SIAM Review, 26(4): 584–585, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levitin:1987:BRP

[Lev87] Lev B. Levitin. Book review:

Probability and Information. (A. M. Yaglom and I. M. Yaglom). SIAM Review, 29(1): 144–146, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levi:1988:SIP

[Lev88a] Mark Levi. Stability of the inverted pendulum — a topological explanation. SIAM Review, 30(4):639– 644, December 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levine:1988:BRP

[Lev88b] Howard A. Levine. Book review: Partial Differential Equations: an Introduction to a General Theory of Linear Boundary Value Problems (Aleksei A. Dezin; Ralph P. Boas, trans.). SIAM Review, 30(4):672– 673, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lever:1989:CSF

[Lev89] K. V. Lever. The complexity of the standard form of an integer. SIAM Review, 31 (3):493–494, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lever:1990:EWE

[Lev90a] Ken Lever. Exponentially windowed estimates of a periodically sampled time-series, and its first and second derivatives using quadratic regression. SIAM Review, 32 (4):672–673, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levine:1990:RCE

[Lev90b] Howard A. Levine. The role of critical exponents in blowup theorems. SIAM Review, 32(2):262–288, June 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lever:1991:EWE

[Lev91] Ken Lever. Exponentially windowed estimates of a periodically sampled time-series, and its first and second derivatives using quadratic regression. SIAM Review, 33 (4):658–661, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

LeVeque:1993:BRA

[LeV93] R. J. LeVeque. Book review: Adaptive Finite Element Solution Algorithm for the Euler Equations (R. A. Shapiro). SIAM Review, 35 (3):531–532, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levine:1996:BRB

[Lev96] Howard Levine. Book review: Blow-Up in Quasilinear Parabolic Equations (A. A. Samarskii, V. A. Galaktionov, S. P. Kurdyumov, and A. P. Mikhailov). SIAM Review, 38(4):692– 694, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levitin:1997:PTM

[Lev97] Evgenij S. Levitin. Perturbation theory in mathematical programming and its applications. SIAM Review, 39(4): 799–??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

LeVeque:2003:SR

[LeV03a] Randy LeVeque. Survey and review. SIAM Review, 45 (4):625–626, December 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97116.

LeVeque:2003:SRIa

[LeV03b]  Randy LeVeque.  Survey

and review — introduction. SIAM Review, 45 (1):1, March 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97101.

LeVeque:2003:SRIb

[LeV03c]  Randy LeVeque.  Survey

and review — introduction. SIAM Review, 45 (2):165, June 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97106.

LeVeque:2003:SRIc

[LeV03d]  Randy LeVeque.  Survey

and review — introduction. SIAM Review, 45 (3):383, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97111.

LeVeque:2004:I

[LeV04a]  R. LeVeque.  Introduc-

tion. SIAM Review, 46 (4):597, ???? 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

LeVeque:2004:SRa

[LeV04b]  Randy LeVeque.  Survey

and review. SIAM Review, 46(1):1, March 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97121.

LeVeque:2004:SRb

[LeV04c]  Randy LeVeque.  Survey

and review. SIAM Review, 46(3):375, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97130.

LeVeque:2004:SRc

[LeV04d]  Randy LeVeque.  Survey

and review. SIAM Review, 46(4):597, December 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97135.

LeVeque:2005:SRa

[LeV05a]  Randy LeVeque.  Survey

and review. SIAM Review, 47(1):1, March 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97140.

LeVeque:2005:SRb

[LeV05b]  Randy LeVeque.  Survey

and review. SIAM Review, 47(2):195, June 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97145.

LeVeque:2005:SRI

[LeV05c]  Randy LeVeque.  Survey

and review — introduction. SIAM Review, 47 (3):405, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97150.

Lewis:1969:OH

[Lew69] J. A. Lewis. Ohmic heating. SIAM Review, 11(4): 620, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1971:OH

[Lew71] J. A. Lewis. Ohmic heating. SIAM Review, 13(1):

123–125, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1973:ACC

[Lew73] James T. Lewis. Approximation with convex constraints. SIAM Review, 15 (1):193–217, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1976:BSB

[Lew76] J. W. Lewis. B-spline bases for splines under tension, Vsplines, and fractional order splines. SIAM Review, 18 (4):815, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1978:AID

[Lew78a] J. S. Lew. Approximate invariance of disc averages. SIAM Review, 20(1):181– 182, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1978:PBW

[Lew78b] J. S. Lew. Paul Bunyan’s washline. SIAM Review, 20 (4):855, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1979:PFD

[Lew79] J. A. Lewis. Post-furnace drawdown of glass fiber. SIAM Review, 21(2):252– 255, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1981:OR

[Lew81a] J. S. Lew. An optimal raster. SIAM Review, 23 (2):255, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1981:OT

[Lew81b] J. S. Lew. Optimal tracking. SIAM Review, 23(1): 100–101, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1982:OR

[Lew82a] J. S. Lew. An optimal raster. SIAM Review, 24(2): 224–226, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1982:OT

[Lew82b] J. S. Lew. Optimal tracking. SIAM Review, 24(1): 81–83, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1982:VP

[Lew82c] M. K. Lewis. A volume problem. SIAM Review, 24 (1):77, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1983:BRM

[Lew83a] G. N. Lewis. Book review:

A Minimization Problem (M. Goldberg). SIAM Review, 25 (1):98–99, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1983:ND

[Lew83b] G. N. Lewis. A nonnegative determinant. SIAM Review, 25(3):401, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1983:ATN

[Lew83c] Stephen D. Lewis. Adjustment time in the neoclassical growth model: an optimal control solution. SIAM Review, 25(3):395–400, July 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1986:BRN

[Lew86] Gilbert N. Lewis. Book review: On a Nolnlinear Matrix Differential Equation (P. Schweitzer). SIAM Review, 28(2):241–242, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1988:SRC

[Lew88] John S. Lew. Shape of a roller-coaster track. SIAM Review, 30(2):308, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1989:ESD

[Lew89] J. S. Lew. Exact-solutions of a differential-equation. SIAM Review, 31(2):324– 325, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1991:BRS

[Lew91a] F. L. Lewis. Book review: Singular Control Systems (L. Dai). SIAM Review, 33(1): 151–152, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1991:BRG

[Lew91b] Gilbert N. Lewis. Book review: Guerrilla Combat Model (M. A. Abdelkader). SIAM Review, 33(1):122– 124, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lansing:1966:BRD

[LF66] D. L. Lansing and H. E. Fettis. Book review: A Double Integral (N. Mullineux and J.

R. Reed). SIAM Review, 8 (3):391–392, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1973:VSN

[LF73] E. H. Lee and G. E.

Forsythe. Variational study of nonlinear spline curves. SIAM Review, 15(1):120– 133, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1978:ADC

[LFK78] John S. Lew, James C.

Frauenthal, and Nathan Keyfitz. On the average distances in a circular disc. SIAM Review, 20(3):584– 592, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1980:BRI

[LFR80] O. P. Lossers, H. E. Fettis, and D. K. Ross. Book review: An Identity (J. D. Love). SIAM Review, 22(3): 372–373, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1974:IEC

[LG74] J. S. Lew and R. Ghez. An integral equation for crystal growth. SIAM Review, 16 (2):258, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1975:IEC

[LG75] J. S. Lew and R. Ghez. An integral equation for crystal growth. SIAM Review, 17(4): 687–690, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Larsen:1980:BRS

[LG80] Jesper Larsen and Gustave Gripenberg. Book review: A Sum of Bessel Functions (M. L. Glasser). SIAM Review, 22 (4):508–509, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lew:1989:BRE

[LG89] John S. Lew and C. C. Grosjean. Book review: Exact Solutions of a Differential Equation (Z. Ikonić and V. Milanović). SIAM Review, 31 (2):324–325, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lavine:1995:UI

[LG95] Shaughan Lavine and Frederick Gass. Understanding the infinite. SIAM Review, 37(3): 464–??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lord:1975:PML

[LH75] W. Lord and J. H. Hwang.

Predicting magnetic leakage fields by finite-element techniques. SIAM Review, 17 (2):384, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lorenz:1995:EC

[LH95] E. N. Lorenz and Philip Holmes. The essence of chaos. SIAM Review, 37(1): 129–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luo:2019:FFT

[LH19] Guo Luo and Thomas Y.

Hou. Formation of finitetime singularities in the 3D axisymmetric Euler equations: a numerics guided study. SIAM Review, 61(4): 793–835, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Leon:1997:ACE

[LHFE97] Steven Leon, Eugene Herman, Richard Faulkenberry, and Richard Elderkin. AT-LAST: Computer exercises for linear algebra. SIAM Review, 39(3):530–??, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liebetrau:1977:IS

[Lie77] A. M. Liebetrau. An infinite summation. SIAM Review, 19(4):736–737, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liebetrau:1978:IS

[Lie78] A. M. Liebetrau. An infinite summation. SIAM Review, 20(4):859–861, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liess:1995:BRL

[Lie95] Otto Liess. Book review: Linear Partial Differential Operators in Gevrey Spaces (Luigi Rodino). SIAM Review, 37(2):263–264, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lieberman:1997:BRL

[Lie97] Gary M. Lieberman. Book review: Linear and Quasilinear Parabolic Problems. Volume 1: Abstract Linear Theory, by Herbert Amann. SIAM Review, 39(1):153– 154, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lighthill:1976:FHJ

[Lig76]  James Lighthill.  Flagel-

lar hydrodynamics — John von Neumann Lecture, 1975.

SIAM Review, 18(2):161– 230, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liggett:1986:BRB

[Lig86] James A. Liggett. Book review: Boundary Integral Equation Analyses of Singular, Potential, and Biharmonic Problems (Derek B. Ingham and Mark A. Kelmanson). SIAM Review, 28 (1):113–115, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lilliefors:1967:BRT

[Lil67] Hubert W. Lilliefors. Book review: Theory of Probability (Bernard Harris). SIAM Review, 9(4):763–764, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lilliefors:1968:BRE

[Lil68] Hubert W. Lilliefors. Book review: Elements of Nonparametric Statistics (Gottfried E. Noether). SIAM Review, 10(3):394–396, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lim:2020:HLG

[Lim20] Lek-Heng Lim. Hodge Laplacians on graphs. SIAM Review, 62(3):685–715, ???? 2020. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Lima:2023:HPM

[Lim23] Rafael Lima. Hawkes processes modeling, inference, and control: an overview. SIAM Review, 65(2):331– 374, ???? 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.
                                        1137/21M1396927.

Lindgren:1959:SFF

[Lin59]  B. W. Lindgren.  On

Samulon’s formula for frequency response from step response. SIAM Review, 1 (1):47–49, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lin:1967:OAM

[Lin67] C. C. Lin. I. Objectives of applied mathematics education. SIAM Review, 9(2): 293–311, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lin:1969:DSG

[Lin69] C. C. Lin. Dynamics of selfgravitating systems: Structure of galaxies. SIAM Review, 11(2):127–151, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lin:1976:AAA

[Lin76] C. C. Lin. Academic activities in applied mathematics. SIAM Review, 18(4): 815, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lin:1978:EAM

[Lin78a]  C. C. Lin.  Education

of applied mathematicians. SIAM Review, 20(4):838– 845, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ling:1978:EWZ

[Lin78b] Chin-Bing Ling. Evaluation of Weierstrass zeta functions. SIAM Review, 20 (1):183, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ling:1979:HSI

[Lin79a] Chih-Bing Ling. A hyperbolic series identity. SIAM Review, 21(2):257, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ling:1979:EWZ

[Lin79b] Chih-Bing B. Ling. Evaluation of Weierstrass zeta functions. SIAM Review, 21(1): 146–147, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ling:1979:BRS

[Lin79c] Robert F. Ling. Book review: Statistical Methods for Digital Computers (Kurt Enslein, Anthony Ralston and Herbert S. Wilf, eds.). SIAM Review, 21(2):276– 277, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lindquist:1980:BRL

[Lin80] Anders Lindquist. Book review: Linear Estimation and Stochastic Control (M. H. A. Davis). SIAM Review, 22(2): 239–241, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ling:1986:AAI

[Lin86] Frederick F. Ling. Asymptotic analyses in isothermal fluid film lubrication theories. SIAM Review, 28 (3):343–366, September 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lindquist:1990:BRL

[Lin90]  Anders Lindquist.  Book

review: Linear Stochastic Systems (Peter E. Caines). SIAM Review, 32(2):325– 328, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lindberg:1991:BRD

[Lin91a] Herbert E. Lindberg. Book review: Dynamic Stability of Suddenly Loaded Structures (George J. Simitses). SIAM Review, 33(2):327– 328, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lindberg:1991:DSS

[Lin91b] Herbert E. Lindberg. Dynamic stability of suddenly loaded structures (George J Simitses). SIAM Review, 33 (2):327–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ling:1992:BRA

[Lin92] Robert F. Ling. Book review:

Applied Multivariate Data Analysis, Vol. I: Regression and Experimental Design (J. D. Jobson). SIAM Review, 34 (4):671–673, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Linton:2010:LSH

[Lin10]  C. M. Linton.  Lattice

sums for the Helmholtz equation. SIAM Review, 52(4): 630–674, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lions:1982:EPS

[Lio82] P. L. Lions. On the existence of positive solutions of semilinear elliptic equations. SIAM Review, 24(4): 441–467, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lions:1988:ECS

[Lio88] J. L. Lions. Exact controllability, stabilization and perturbations for distributed systems. SIAM Review, 30 (1):1–68, March 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liron:1976:ZMM

[Lir76] N. Liron. A zero of maximum multiplicity. SIAM Review, 18(4):763, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Littauer:1963:BRM

[Lit63] Sebastian B. Littauer. Book review: Mathematical Methods of Operations Research (T. L. Saaty). SIAM Review, 5(2):170–171, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liu:1983:IP

[Liu83] P. C. Liu. An iteration problem. SIAM Review, 25 (2):268, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liu:1988:BRA

[Liu88] Andy Liu. Book review:

Applied Combinatorics. Second Edition (Alan Tucker). SIAM Review, 30(2):337– 339, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liu:1991:BRI

[Liu91] Xiaosong Liu. Book review: Integral Manifolds and Inertial Manifolds for Dissipative Partial Differential Equations (P. Constantin, C. Foias, B. Nicolaenko, and R. Temam). SIAM Review, 33(3):489, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liu:1992:WGL

[Liu92a] Andy Liu. On winning in the game of lotto. SIAM Review, 34(1):120, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liu:1992:MMS

[Liu92b] Joseph W. H. Liu. The multifrontal method for sparse matrix solution: Theory and practice. SIAM Review, 34 (1):82–109, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liu:1994:LPT

[Liu94] Yunkang Liu. A limit problem from the theory of delay equations. SIAM Review, 36(3):490, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liu:1995:LPT

[Liu95] Yunkang Liu. A limit problem from the theory of delay equations. SIAM Review, 37 (3):448–450, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lockhart:1990:PE

[LIW90] Ted W. Lockhart, Robert Ingle, and Peter Wollan. Probabilistic ethics. SIAM Review, 32(4):686–688, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liz:1998:NME

[Liz98a] E. Liz. A note on the matrix exponential. SIAM Review, 40(3):700–702, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liz:1998:CNN

[Liz98b] Eduardo Liz. Classroom note: a note on the matrix exponential. SIAM Review, 40(3):700–702, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/32075.

Lossers:1983:BRD

[LJ83] O. P. Lossers and A. J.

Jerri. Book review: A Definite Integral (A. J. Jerri). SIAM Review, 25(1):101– 102, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1973:BRC

[LJW73] O. P. Lossers, A. A. Jagers, and A. N. Willson, Jr. Book review: Conjectured Monotonicity of a Matrix (R. A. Usmani). SIAM Review, 15 (2):387–390, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luther:1965:SFO

[LK65] H. A. Luther and H. P.

Konen. Some fifth-order classical Runge–Kutta formulas. SIAM Review, 7(4): 551–558, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lawler:1998:ISP

[LK98] Gregory F. Lawler and Eric S. Key. Introduction to stochastic processes. SIAM Review, 40(1):159– ??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Li:2007:DNC

[LKGK07] Ju Li, Panayotis G. Kevrekidis, C. William Gear, and Ioannis G. Kevrekidis. Deciding the nature of the coarse equation through microscopic simulations: The babybathwater Scheme. SIAM Review, 49(3):469–487, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liggett:1969:SGP

[LL69] Thomas M. Liggett and Steven A. Lippman. Stochastic games with perfect information and time average payoff. SIAM Review, 11(4): 604–607, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levine:1975:OAS

[LL75] L. E. Levine and E. S. Lubot. Optimal approximate solutions and time scales. SIAM Review, 17(2):384, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lavita:1976:TDN

[LL76a] J. Lavita and W. Loehr. Time-dependent, non-Paretian, policy dictated moves. SIAM Review, 18(4):814, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lynn:1976:OLS

[LL76b] R. Y. S. Lynn and M. L. Levitan. Overdetermined linearsystem. SIAM Review, 18 (4):816, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laffin:1994:APS

[LL94] Steven Laffin and A. Liu.

All purpose stamp design. SIAM Review, 36(2):283– 285, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Liberti:2014:EDG

[LLMM14] Leo Liberti, Carlile Lavor, Nelson Maculan, and Antonio Mucherino. Euclidean distance geometry and applications. SIAM Review, 56 (1):3–69, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lockhart:1981:NLS

[LLR^{+}81] D. F. Lockhart, G. N. Lewis, O. G. Ruehr, W. W. Meyer, H. E. Fettis, D. Powers, and R. A. Struble. A non-linear system of differential equations. SIAM Review, 23(4): 525–527, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lagnese:1996:MAC

[LLSN96] J. E. Lagnese, G. Leugeuring, E. J. P. G. Schmidt, and Serge Nicaise. Modeling, analysis and control of dynamic elastic multi-link structures. SIAM Review, 38 (3):537–??, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Langhaar:1978:CD

[LM78] H. L. Langhaar and R. E. Miller. A conjecture on determinants. SIAM Review, 20(1):182, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

LeBlond:1979:OWS

[LM79] Paul H. LeBlond and Lawrence A. Mysak. Ocean waves: a survey of some recent results. SIAM Review, 21(3): 289–328, July 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lockhart:1981:BRN

[LM81] Deborah Frank Lockhart and W. Weston Meyer. Book review: A Nonlinear System of Differential Equations Problem (G. N. Lewis and O. G. Ruehr). SIAM Review, 23(4): 525–527, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lazer:1990:LAP

[LM90] A. C. Lazer and P. J.

Mckenna. Large-amplitude periodic oscillations in suspension bridges: Some new connections with nonlinear analysis. SIAM Review, 32 (4):537–578, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lewis:1999:CCS

[LM99] Andrew D. Lewis and Richard M.

Murray. Configuration controllability of simple mechanical control systems. SIAM Review, 41(3):555– 574, September 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35106.

Li:2000:ECS

[LM00] Chi-Kwong K. Li and Roy Mathias. Extremal characterizations of the Schur complement and resulting inequalities. SIAM Review, 42(2):233–246, June 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33729.

Langville:2005:SEM

[LM05] Amy N. Langville and Carl D. Meyer. A survey of eigenvector methods for Web information retrieval. SIAM Review, 47 (1):135–161, March 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42478.

Llorente:2023:MLC

[LMDLS23] F. Llorente, L. Martino, D. Delgado, and J. López-Santiago. Marginal likelihood computation for model selection and hypothesis testing: an extensive review. SIAM Review, 65 (1):3–58, ???? 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/20M1310849.

Langhaar:1979:CD

[LMGC79] H. L. Langhaar, R. E. Miller, C. Gzven, and G. D. Cantor. Conjecture on determinants. SIAM Review, 21(1): 144–146, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lu:2021:DDL

[LMMK21] Lu Lu, Xuhui Meng, Zhiping Mao, and George Em Karniadakis. DeepXDE: a deep learning library for solving differential equations. SIAM Review, 63(1):208– 228, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Logan:1968:DI

[LMS68] B. F. Logan, C. L. Mallows, and L. A. Shepp. A definite integral. SIAM Review, 10(4):451, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lagarias:1983:IWG

[LMSM83] J. C. Lagarias, J. E. Mazo, L. A. Shepp, and B. McKay. An inequality for walks in a graph. SIAM Review, 25 (3):403, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lagarias:1984:IWG

[LMSM84] J. C. Lagarias, J. E. Mazo, L. A. Shepp, and B. McKay. An inequality for walks in a graph. SIAM Review, 26(4): 580–582, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Li:1996:OCT

[LmYS96] Xunjing Li, Jiong min Yong, and Srdjan Stojanovic. Optimal control theory for infinite dimensional systems. SIAM Review, 38(2):352– ??, June 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lockhart:1984:PVL

[Loc84a] D. F. Lockhart. Pendulum with variable length. SIAM Review, 26(3):436– 437, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lockhart:1984:BRP

[Loc84b] Deborah Frank Lockhart. Book review: Pendulum with Variable Length (M. A. Abdelkader). SIAM Review, 26 (3):436–437, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lockhart:1989:PE

[Loc89] Ted W. Lockhart. Probabilistic ethics. SIAM Review, 31 (4):676–677, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Loeser:1992:BRG

[Loe92a] François Loeser. Book review: Generalized Hypergeometric Functions (Bernard Dwork). SIAM Review, 34 (4):673–675, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Loeser:1992:GHF

[Loe92b] François Loeser. Generalized hypergeometric functions (Bernard Dwork).

SIAM Review, 34(4):673– ??, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Loeb:1995:BRN

[Loe95] Peter A. Loeb. Book review:

Nonstandard Methods in the Calculus of Variations (Curtis Tuckey). SIAM Review, 37 (3):474–475, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Logan:1997:BRM

[Log97] J. David Logan. Book review: Modelling Mathematical Methods and Scientific Computation, by Nicola Bellomo and Luigi Preziosi. SIAM Review, 39(1):154– 156, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Logan:1998:SSH

[Log98] J. David Logan. Similarity solution to a heat exchange problem. SIAM Review, 40 (4):918–921, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32391.

Lonseth:1977:SAI

[Lon77] Arvid T. Lonseth. Sources and applications of integral equations. SIAM Review, 19 (2):241–278, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Longley:1978:SAC

[Lon78] J. W. Longley. Subroutine for avoiding catastrophic roundoff in solution of large matrices. SIAM Review, 20 (3):630, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Loncaric:1987:BRR

[Lon87] Josip Loncarić. Book review: Robot Hands and the Mechanics of Manipulation (Matthew T. Mason and J. Kenneth Salisbury, Jr.). SIAM Review, 29(3): 481–483, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Loper:1994:RTD

[Lop94] D. E. Loper. On the ratio of two double integrals. SIAM Review, 36(2):277, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lopez:2008:AEM

[Lóp08] José L. López. Asymptotic expansions of Mellin convolution integrals. SIAM Review, 50(2):275–293, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lord:1991:BRT

[Lor91] W. Lord. Book review: Topics in Boundary Element Research. Vol. 6: Electromagnetic Applications (C. A. Brebbia, ed.). SIAM Review, 33(3):510–511, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1973:BRL

[Los73] O. P. Lossers. Book review: Limit of an Integral (N. Mullineux and J. R. Reed). SIAM Review, 15(2): 390–391, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1974:BRH

[Los74a] O. P. Lossers. Book review:

A Hadamard-Type Bound on the Coefficients of a Determinant of Polynomials (A. J. Goldstein and R. L. Graham). SIAM Review, 16(3): 394–395, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1974:BRP

[Los74b] O. P. Lossers. Book review: A Polynomial Diophantine Equation (M. S. Klamkin). SIAM Review, 16 (1):99–100, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1975:BRG

[Los75] O. P. Lossers. Book review: A Generalization of the Vandermonde Determinant (S. Venit). SIAM Review, 17(4): 694–695, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1976:BRS

[Los76a] O. P. Lossers. Book review: A Stability Problem (O. Bottema). SIAM Review, 18(1): 118–119, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1976:BRE

[Los76b] O. P. Lossers. Book review: An Eigenvalue Problem (E. Wasserstrom). SIAM Review, 18(3):502–505, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1976:NIE

[Los76c] O. P. Lossers. Nonlinear integral equation. SIAM Review, 18(4):770, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1981:BRE

[Los81] O. P. Lossers. Book review: Extreme Value of an Integral (W. W. Meyer). SIAM Review, 23(3):392– 393, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1983:BRMa

[Los83a] O. P. Lossers. Book review: A Maximum, Minimum Problem (T. Sekiguchi). SIAM Review, 25(1):106, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1983:BRMb

[Los83b] O. P. Lossers. Book review:

Malevolent Traffic Lights (J. C. Lagarias). SIAM Review, 25(4):571–572, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1983:DI

[Los83c] O. P. Lossers. A definite integral. SIAM Review, 25(1): 101–102, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1984:BRC

[Los84a] O. P. Lossers. Book review:

A Covering Problem (J. Vijay ). SIAM Review, 26 (1):124, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1984:GFE

[Los84b] O. P. Lossers. Gamma function expansions. SIAM Review, 26(2):278–279, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1985:BRP

[Los85] O. P. Lossers. Book review: Patterns in a Sequence of Symmetric Bernoulli Trials (Y. P. Sabharwal and V. K. Malhotra). SIAM Review, 27 (3):453–454, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1986:BRS

[Los86a] O. P. Lossers. Book review: Series Involving Bessel Functions (Eldon Hansen). SIAM Review, 28(4):578– 579, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1986:SIB

[Los86b] O. P. Lossers. Series involving Bessel-functions. SIAM Review, 28(4):579, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1989:BRG

[Los89]  O. P. Lossers.  Book

review: A Generalized Hypergeometric-Type Identity (C. C. Grosjean). SIAM Review, 31(3):496–497, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lossers:1995:BRO

[Los95] O. P. Lossers. Book review: Overlapping Binary Sequences (G. Blom). SIAM Review, 37(4):619–620, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lavoie:1976:FDS

[LOT76a] J. L. Lavoie, T. J. Osler, and R. Tremblay. Fractional derivatives and special functions. SIAM Review, 18(2): 240–268, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lovoie:1976:FDS

[LOT76b] J. L. Lovoie, T. J. Osler, and R. Tremblay. Fractional derivatives and special functions. SIAM Review, 18(2): 240–268, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Loud:1970:SEG

[Lou70] Warren S. Loud. Some examples of generalized Green’s functions and generalized Green’s matrices. SIAM Review, 12(2):194–210, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Louko:1992:BRI

[Lou92] Jorma Louko. Book review: An Introduction to General Relativity (L. P. Hughston and K. P. Tod). SIAM Review, 34(4):675, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Love:1979:DE

[Lov79a] J. D. Love. A differential equation. SIAM Review, 21 (4):560, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Love:1979:I

[Lov79b]  J. D. Love.  An iden-

tity. SIAM Review, 21(3): 396, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lowndes:1961:NIT

[Low61] J. S. Lowndes. Note on an integral transform. SIAM Review, 3(2):162–164, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levy:1971:FPV

[LP71] J. Levy and M. H. Pearl. Fixed probability vectors of some stochastic matrices. SIAM Review, 13(2): 246–247, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lee:1975:NMM

[LP75] D. Lee and S. Preiser. Nonlinear multistep methods for solving initial value problems in ordinary differential equations. SIAM Review, 17(2): 383–384, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lutscher:2005:EDP

[LPL05] Frithjof Lutscher, Elizaveta Pachepsky, and Mark A. Lewis. The effect of dispersal patterns on stream populations. SIAM Review, 47 (4):749–772, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/63615.

Lewis:1980:NSD

[LR80] G. N. Lewis and O. G. Ruehr.

A nonlinear system of differential equations. SIAM Review, 22(4):501, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Li:2000:NSF

[LR00] Jie Li and Yuriko Renardy.

Numerical study of flows of two immiscible liquids at low Reynolds number. SIAM Review, 42(3):417– 439, September 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/35460.

Lancaster:2005:CFH

[LR05] Peter Lancaster and Leiba Rodman. Canonical forms for Hermitian matrix pairs under strict equivalence and congruence. SIAM Review, 47 (3):407–443, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44556.

Lu:1978:SLS

[LRJ78] P. Lu, P. E. Rubbert, and F. T. Johnson. Solving a large system of non-linear equations arising from a theoretical model of an aerodynamics problem. SIAM Review, 20(3):630, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lancaster:1996:ARE

[LRM96] Peter Lancaster, Leiba Rodman, and Volker Mehrmann. Algebraic Riccati equations. SIAM Review, 38(4):694– ??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lugue:1976:PDL

[LRV76] B. S. Lugue, B. W. Rust, and W. Vanwinkle. Problem of determining length of prediction error filter in maximum entropy spectral analysis. SIAM Review, 18(4): 815–816, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lindamood:1963:MCO

[LS63] George E. Lindamood and George Shapiro. Magnitude comparison and overflow detection in modular arithmetic computers. SIAM Review, 5 (4):342–350, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lansing:1966:BRL

[LS66] D. L. Lansing and F. W.

Steutel. Book review: Limiting Value of a Solution of an Integral Equation (W. T. Moody). SIAM Review, 8(4): 535–539, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Logan:1971:SI

[LS71] B. F. Logan and L. A.

Shepp. Supremum of an integral. SIAM Review, 13 (3):389, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Logan:1972:SI

[LS72] B. F. Logan and L. A.

Shepp. Supremum of an integral. SIAM Review, 14(3): 500–501, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ladany:1978:MPJ

[LS78] Shaul P. Ladany and Jagbir Singh. On maximizing the probability of jumping over a ditch. SIAM Review, 20(1): 171–177, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luxemburg:1984:BRF

[LS84] W. A. J. Luxemburg and Robert E. Shafer. Book review: A Factorial Polynomial Inequality (A. Fransén). SIAM Review, 26(1):118– 121, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levin:1985:PGS

[LS85] Simon A. Levin and Lee A. Segel. Pattern generation in space and aspect. SIAM Review, 27(1):45–67, March 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lazoff:1995:EWL

[LS95] David M. Lazoff and Alan T. Sherman. Expected wire length between two randomly chosen terminals. SIAM Review, 37(2):235, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Langer:1996:LAK

[LS96] Joel Langer and David A.

Singer. Lagrangian aspects of the Kirchhoff elastic rod. SIAM Review, 38 (4):605–618, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25329.

Luck:2002:EST

[LS02] Rogelio Luck and James W.

Stevens. Explicit solutions for transcendental equations. SIAM Review, 44(2): 227–233, June 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38611.

Liesen:2008:OSR

[LS08] J. Liesen and Z. Strakoš.

On optimal short recurrences for generating orthogonal Krylov subspace bases. SIAM Review, 50(3):485– 503, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Linderman:2022:DRD

[LS22] George C. Linderman and Stefan Steinerberger. Dimensionality reduction via dynamical systems: The case of t-SNE. SIAM Review, 64(1):153–178, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1446769.

Luke:1967:BRT

[LSB67] Yudell Luke, A. J. Strecok, and J. L. Brown, Jr. Book review: Two Integrals (A. D. Brailsford and M. S. Klamkin). SIAM Review, 9 (1):124–127, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lynn:1998:BAN

[LSS98] Nicholas Lynn, Nozer Singpurwalla, and Adrian Smith. Bayesian assessment of network reliability. SIAM Review, 40(2):202–227, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

                                        http://epubs.siam.org/
                                        sam-bin/dbq/article/30678.

Lin:2016:ASD

[LSY16] Lin Lin, Yousef Saad, and Chao Yang. Approximating spectral densities of large matrices. SIAM Review, 58(1): 34–65, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Laurie:1990:LS

[LT90] D. P. Laurie and B. Tibken.

Late solutions. SIAM Review, 32(4):674, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levy:1998:SFD

[LT98] Doron Levy and Eitan Tadmor. From semidiscrete to fully discrete: Stability of Runge–Kutta schemes by the energy method. SIAM Review, 40(1):40– 73, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31625.

Luss:2013:CGA

[LT13] Ronny Luss and Marc Teboulle. Conditional gradient algorithms for rankone matrix approximations with a sparsity constraint. SIAM Review, 55(1):65– 98, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lu:1992:BRR

[Lu92] Chunqing Lu. Book review: Reaction-Diffusion Equations. (K. J. Brown and A. A. Lacey, eds.). SIAM Review, 34(1):131– 132, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lucas:1971:SRD

[Luc71] W. F. Lucas. Some recent developments in n-person game theory. SIAM Review, 13(4): 491–523, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luchins:1986:BRD

[Luc86] Edith H. Luchins. Book review: Discrete Mathematics (Kenneth A. Ross and Charles R. B. Wright) and Applied Combinatorics (Fred S. Roberts). SIAM Review, 28(2):267–269, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lucas:1987:BRG

[Luc87] William F. Lucas. Book review: Game Theoretic Analysis of Voting in Committees (Bezalel Peleg). SIAM Review, 29(4):661–663, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lucet:2010:WSY

[Luc10] Yves Lucet. What shape is your conjugate? A survey of computational convex analysis and its applications. SIAM Review, 52(3): 505–542, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luchtenburg:2018:BRD

[Luc18] Dirk M. Luchtenburg. Book review: Dynamic Mode Decomposition: Data-Driven Modeling of Complex Systems, by J. Nathan Kutz, Steven L. Brunton, Bingni W. Brunton, and Joshua L. Proctor. SIAM Review, 60 (2):483, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ludwig:1970:UAE

[Lud70] Donald Ludwig. Uniform asymptotic expansions for wave propagation and diffraction problems. SIAM Review, 12(3):325–331, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ludwig:1975:GTS

[Lud75a]  Donald Ludwig.  Geo-

metrical theory for surface waves. SIAM Review, 17 (1):1–15, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ludwig:1975:PDS

[Lud75b] Donald Ludwig. Persistence of dynamical systems under random perturbations. SIAM Review, 17(4):605– 640, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ludwig:1977:BRW

[Lud77] Donald A. Ludwig. Book review: The Wave Equation on a Curved Space-Time (F. G. Friedlander). SIAM Review, 19(3):582–583, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ludwig:1982:BRG

[Lud82] Donald Ludwig. Book review: The Genetics of Altruism (S. A. Boorman and P. R. Leavitt). SIAM Review, 24 (3):366–368, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ludwig:1987:BRS

[Lud87] Donald Ludwig. Book review: Stochastic Processes in Demography and Their Computer Implementation (C. J. Mode). SIAM Review, 29(1): 146–149, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ludwig:1991:BRE

[Lud91] Donald Ludwig. Book review: Estimation Techniques for Distributed Parameter Systems (H. T. Banks and K. Kunisch). SIAM Review, 33 (3):486–487, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luke:1968:BRA

[Luk68] Yudell L. Luke. Book review: An Asymptotic Expansion (Donald L. Lansing). SIAM Review, 10(2): 229–232, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lukacs:1975:BRC

[Luk75] Eugene Lukacs. Book review: Characterization Problems in Mathematical Statistics (A. M. Kagan, Yu. V. Linnik and C. Radhakrishna Rao). SIAM Review, 17(4): 712–714, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luke:1982:BRS

[Luk82] Yudell L. Luke. Book review: Sequence Transformations and Their Applications (Jet Wimp). SIAM Review, 24(4):489–490, October 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lumley:1985:BRD

[Lum85] J. Lumley. Book review:

Developments in the Theory of Turbulence (D. C. Leslie). SIAM Review, 27(3): 468–469, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lusk:1987:BRC

[Lus87] Ewing Lusk. Book review: The Computer Modelling of Mathematical Reasoning (Alan Bundy). SIAM Review, 29(3):483–484, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luther:1966:FEF

[Lut66] H. A. Luther. Further explicit fifth-order Runge–Kutta formulas. SIAM Review, 8(3): 374–380, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luxemburg:1969:LI

[Lux69] W. A. Luxemburg. Limit of an integral. SIAM Review, 11(1):74, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lux:1979:I

[Lux79] I. Lux. An inequality. SIAM Review, 21(1):139, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Luxemburg:1984:FPI

[Lux84] W. A. J. Luxemburg. A factorial polynomial inequality. SIAM Review, 26(1): 118–121, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lomont:1959:DEC

[LW59] J. S. Lomont and R. A.

Willoughby. Dominant eigenvectors of a class of test matrices. SIAM Review, 1(1): 64–65, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

LaSalle:1964:BRS

[LW64] J. P. LaSalle and W. M. Wonham. Book review: The Second International Congress of IFAC on Automatic Control (August 27–September 4, 1963, Basel, Switzerland). SIAM Review, 6(3): 296–309, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lih:1981:PCP

[LW81] Ko-Wei Lih and E. T. H.

Wang. Positivity of the coefficients of a polynomial. SIAM Review, 23(3):391, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Levi:1995:SIL

[LW95] Mark Levi and Warren Weckesser. Stabilization of the inverted linearized pendulum by high frequency vibrations. SIAM Review, 37 (2):219–223, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/26260.htm.

Li:2004:LRS

[LW04] Jing-Rebecca R. Li and Jacob White. Low-rank solution of Lyapunov equations. SIAM Review, 46 (4):693–713, December 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44338. Reprinted

from SIAM Journal on Matrix Analysis and Applications, vol 24, pg 260–280, 2002.

Lieberman:2013:GOI

[LW13] Chad Lieberman and Karen Willcox. Goal-oriented inference: Approach, linear theory, and application to advection diffusion. SIAM Review, 55(3):493–519, ???? 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lim:2017:FRI

[LW17] Lek-Heng Lim and Jonathan Weare. Fast randomized iteration: Diffusion Monte Carlo through the lens of numerical linear algebra. SIAM Review, 59(3):547–587, ???? 2017. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lyness:1967:TI

[Lyn67] J. N. Lyness. A trigonometric inequality. SIAM Review, 9(2):250, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lyness:1983:WUA

[Lyn83] J. N. Lyness. When not to use an automatic quadrature routine. SIAM Review, 25(1): 63–87, January 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mickens:1995:NFD

[MA95] Ronald E. Mickens and Ravi P. Agarwal. Nonstandard finite difference models of differential equations. SIAM Review, 37(3):459– ??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moreau:2004:NVM

[MA04] Luc Moreau and Dirk Aeyels.

A novel variational method for deriving Lagrangian and Hamiltonian models of inductor-capacitor circuits. SIAM Review, 46(1): 59–84, March 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/40902.

Machol:1961:BRL

[Mac61] Robert E. Machol. Book review: Linear Graphs and Electrical Networks (S. Seshu and M. B. Reed). SIAM Review, 3(4):343–345, ????

1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mackenzie:1967:EFT

[Mac67]  J. K. Mackenzie.  Eval-

uation of a Fourier transform. SIAM Review, 9(2): 219–222, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mackie:1968:SCE

[Mac68a] A. G. Mackie. Some comments on existence and uniqueness theorems in applied mathematics with an application to thin airfoil theory. SIAM Review, 10(2): 196–207, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacLaren:1968:BRL

[Mac68b] M. Donald MacLaren. Book review: List Processing (J. M. Foster). SIAM Review, 10(3):396, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacLaren:1969:BRA

[Mac69] M. Donald MacLaren. Book review: The Art of Computer Programming — Volume 1: Fundamental Algorithms (Donald E. Knuth). SIAM Review, 11(1):89– 91, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacLaren:1970:BRA

[Mac70] M. Donald MacLaren. Book review: The Art of Computer Programming. Volume 2: Seminumerical Algorithms (Donald E. Knuth). SIAM Review, 12(2):306– 308, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Machol:1976:BRU

[Mac76a] Robert E. Machol. Book review: Urban Systems Models (Walter Helly). SIAM Review, 18(3):516–517, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Macki:1976:ACV

[Mac76b] Jack W. Macki. Applications of change of variables in the qualitative theory of second order linear ordinary differential equations. SIAM Review, 18(2):269–274, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mackie:1977:EMI

[Mac77] A. G. Mackie. An example from mechanics illustrating classical and modern mathematical techniques. SIAM Review, 19(3):517– 535, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Macdonald:1985:HV

[Mac85] I. G. Macdonald. A hyper volume. SIAM Review, 27(4): 579–580, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacKay:1991:BRCb

[Mac91a] R. S. MacKay. Book review: Chaos in Systems with Noise (Tomasz Kapitaniak). SIAM Review, 33(4):666, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacKay:1991:BRCa

[Mac91b] R. S. MacKay. Book review: Chaotic Evolution and Strange Attractors: The Statistical Analysis of Time Series for Deterministic Nonlinear Systems (David Ruelle). SIAM Review, 33(2): 328–329, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacKay:1991:CSN

[Mac91c] R. S. MacKay. Chaos in systems with noise (Tomasz Kapitaniak). SIAM Review, 33(4):666–??, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacKay:1991:CES

[Mac91d] R. S. MacKay. Chaotic evolution and strange attractors: The statistical analysis of time series for deterministic nonlinear systems (David Ruelle). SIAM Review, 33 (2):328–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacSithigh:1992:BRM

[Mac92] Gearoid P. Mac Sithigh.

Book review: Mathematical Elasticity: Volume One: Three-Dimensional Elasticity (Phillipe G. Ciarlet). SIAM Review, 34(3):513– ??, September 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Macki:1998:NEC

[Mac98a] Jack Macki. Notice: Erratum and comments on “Initialization of the Simplex Algorithm: An Artificial-Free Approach” [SIAM Review, 39, (1997), no. 4, 736–744]. SIAM Review, 40(4):1024, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97017.  See

[Ars97].

Macki:1998:TNE

[Mac98b] Jack Macki. Three notes on the exponential of a matrix and applications. SIAM Review, 40(3):699, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97014.

MacCluer:2000:MPA

[Mac00]  C. R. MacCluer.  The

many proofs and applications of Perron’s theorem. SIAM Review, 42(3):487– 498, September 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35944.

Magnon:1986:BRS

[Mag86] Anne Magnon. Book review: Semi-Riemannian Geometrywith Applications to Relativity (Barrett O’Neill). SIAM Review, 28(2):269– 270, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mahrous:1975:QPR

[Mah75]  M. A. Mahrous.  Qual-

itative properties of reformulated Hodgkin–Huxley partial-differential equation. SIAM Review, 17(2):385, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mahaffy:1994:BRD

[Mah94] Joseph M. Mahaffy. Book review: The Dynamics of Cellular Motility (M. Murase). SIAM Review, 36(2):307– 308, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Maier:1972:PLC

[Mai72] G. Maier. A parametric linear complementarily problem. SIAM Review, 14(2): 364–365, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Majda:1991:VTA

[Maj91] Andrew J. Majda. Vorticity, turbulence, and acoustics in fluid flow. SIAM Review, 33 (3):349–388, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Makhoul:1986:CLM

[Mak86] John Makhoul. Conjectured location of maximum amplitude in an all-pass sequence. SIAM Review, 28(3): 395–397, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallows:1962:PS

[Mal62] C. L. Mallows. Patience sorting. SIAM Review, 4(2):

148–149, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallows:1963:PS

[Mal63] C. L. Mallows. Patience sorting. SIAM Review, 5(4): 375–376, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallows:1968:EEU

[Mal68] C. L. Mallows. Extrema of expectations of uniform order-statistics. SIAM Review, 10(1):109, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallows:1969:EEU

[Mal69] C. L. Mallows. Extrema of expectations of uniform order-statistics. SIAM Review, 11(3):410–411, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Malmuth:1970:SS

[Mal70] Norman D. Malmuth. Summation of a series. SIAM Review, 12(1):146, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Malmuth:1971:SS

[Mal71] Norman D. Malmuth. Summation of a series. SIAM Review, 13(2):250–253, ????

1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallik:1977:BRS

[Mal77] Ashim Mallik. Book review: Sequential Statistical Procedure (Z. Govindarajulu). SIAM Review, 19(1): 167–168, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallows:1978:MI

[Mal78] C. L. Mallows. Minimizing an integral. SIAM Review, 20(1):183, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Malik:1980:BRA

[Mal80] Henrick John Malik. Book review: The Algebra of Random Variables (M. D. Springer). SIAM Review, 22 (4):522, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Malliaris:1981:MMF

[Mal81] A. G. Malliaris. Martingale methods in financial decision making. SIAM Review, 23(4): 434–443, October 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Malliaris:1983:ICF

[Mal83] A. G. Malliaris. Itô’s calculus in financial decision making. SIAM Review, 25(4): 481–496, October 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Malkus:1985:BRF

[Mal85] David S. Malkus. Book review: The Finite Element Method — a Basic Introduction for Engineers (K. C. Rockey, H. R. Evans, D. W. Griffiths, and D. A. Nethercot) and Finite Elements: an Introduction for Engineers (R. K. Livesley). SIAM Review, 27(2): 286–288, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Malkus:1987:BRM

[Mal87] David S. Malkus. Book review: Matrix and Finite Element Displacement Analysis of Structures (D. J. Dawe). SIAM Review, 29(1): 149–151, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Malkus:1988:BRF

[Mal88a] David S. Malkus. Book review: The Finite Element Method and Its Applications (Masatake Mori). SIAM Review, 30(2):339– 340, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallows:1988:BRG

[Mal88b] C. L. Mallows. Book review:

A Generalization of Student’s t Distribution (N. M. Blachman). SIAM Review, 30(1): 132–133, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Malkus:1990:BRF

[Mal90] David S. Malkus. Book review: Finite Element Methods (H. R. Schwarz; J. R. Whiteman, ed.;Caroline M. Whiteman, trans.). SIAM Review, 32(3):512–513, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallet:1993:BRM

[Mal93a] Russell L. Mallet. Book review: Maximum Value of a Shear Stress (J. O. Fellman). SIAM Review, 35(2): 299–304, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallett:1993:MVS

[Mal93b] R. L. Mallett. Maximum value of a shear stress. SIAM Review, 35(2):299– 304, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mandelbrot:1961:BRC

[Man61a] Benoît Mandelbrot. Book review: Contributions to Probability and Statistics: Essays in Honor of Harold Hotelling (Ingram Olkin, Sudhist G. Ghurye, Wassily Hoeffding, William G. Madow, and Henry B. Mann, eds.). SIAM Review, 3(1): 80, January 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/siread/v3/
i1/p80/s1;   http://www.
jstor.org/stable/2027260.

Mandelbrot:1961:BRI

[Man61b] Benoît Mandelbrot. Book review: Information and Decision Processes (Robert E. Machol, ed.). SIAM Review, 3(2):182–183, April 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

link.aip.org/link/siread/
v3/i2/p182/s1;   http:/
/www.jstor.org/stable/
2027402.

Mandelbrot:1961:BRT

[Man61c] Benoît Mandelbrot. Book review: La Theorie Physique au Sens de Boltzmann (René Dugas). (French) [The Physical Theory According to Boltzmann (René Dugas)]. SIAM Review, 3(2):183– 184, April 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/siread/v3/
i2/p183/s1;   http://www.
jstor.org/stable/2027403.

Mandelbrot:1961:BRM

[Man61d] Benoît Mandelbrot. Book review: Mathematical Methods in the Social Sciences, 1959, Proceedings of the First Stanford Symposium (Kenneth J. Arrow, Samuel Karlin, and Patrick Suppes, eds.). SIAM Review, 3(1): 80–81, January 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/siread/v3/
i1/p80/s2;   http://www.
jstor.org/stable/2027261.

Mann:1963:BRS

[Man63] H. B. Mann. Book review: Statistical Independence in Probability, Analysis and Number Theory (M. Kac). SIAM Review, 5(3): 292–293, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mangasarian:1968:CRM

[Man68] O. L. Mangasarian. Characterizations of real matrices of monotone kind. SIAM Review, 10(4):439– 441, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mangasarian:1970:BRN

[Man70] O. L. Mangasarian. Book review: Nonlinear Programming: Sequential Unconstrained Minimization Techniques (Anthony V. Fiacco and Garth P. McCormick). SIAM Review, 12(4):593– 594, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mansfield:1976:ABF

[Man76] L. Mansfield. Approximation of boundary in finite-element solution of 4th order problems. SIAM Review, 18(4): 816–817, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Manes:1981:BRF

[Man81] Ernest G. Manes. Book review: Fuzzy Switching and Automata: Theory and Applications (Abraham Kandel and Samuel C. Lee). SIAM Review, 23(2):271– 273, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mangel:1982:AMN

[Man82] Marc Mangel. Applied mathematicians and Naval operators. SIAM Review, 24(3):

289–300, July 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Manitius:1985:BRO

[Man85] Andrzej Manitius. Book review: Optimization and Nonsmooth Analysis (Frank H. Clarke). SIAM Review, 27 (2):288–291, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mangel:1987:BRS

[Man87] Marc Mangel. Book review:

Singular Perturbation Analysis of Discrete Control Systems (D. S. Naidu and A. K. Rao). SIAM Review, 29(4): 664–665, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mandel:1988:BRM

[Man88a] Jan Mandel. Book review:

Multi-Grid Methods and Applications (Wolfgang Hackbusch). SIAM Review, 30(3): 519–520, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mangel:1988:BRS

[Man88b] Marc Mangel. Book review: Singular Perturbation Methods in Control: Analysis and Design (Petar Kokotovic, Hassan K. Khalil, and John O’Reilly). SIAM Review, 30 (4):674–675, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Manacher:1989:BRA

[Man89] Glenn K. Manacher. Book review: Algorithmic Graph Theory (Alan Gibbons). SIAM Review, 31(1):145– 147, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Man:1990:BRC

[Man90a] Chi-Sing Man. Book review: A Commentary on Thermodynamics (William Alan Day). SIAM Review, 32(1): 178–180, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Man:1990:CTW

[Man90b]  Chi-Sing Man.  A com-

mentary on thermodynamics (William Alan Day). SIAM Review, 32(1):178– ??, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mangel:1990:BRI

[Man90c] Marc Mangel. Book review: An Introduction to Mathematical Physiology and Biology (J. Mazumdar). SIAM Review, 32(3):513–514, ???? 1990. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Mangel:1990:IMP

[Man90d] Marc Mangel. An introduction to mathematical physiology and biology (J Mazumdar). SIAM Review, 32(3): 513–??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Manning:2009:CPR

[Man09] Robert S. Manning. Conjugate points revisited and Neumann–Neumann problems. SIAM Review, 51(1): 193–212, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marimont:1960:AGB

[Mar60] Rosalind B. Marimont. Applications of graphs and Boolean matrices to computer programming. SIAM Review, 2(4):259–268, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcus:1964:BRA

[Mar64] Marven Marcus. Book review: Applications of the Theory of Matrices (F. R. Gantmacher). SIAM Review, 6(2):190–191, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Martin:1965:BRS

[Mar65] John L. Martin. Book review: Spatial Navigation (Walter Weissblum). SIAM Review, 7 (1):132–134, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marchand:1966:BRE

[Mar66] Erich W. Marchand. Book review: Electromagnetic Theory and Geometrical Optics (Morris Kline and Irvin W. Kay). SIAM Review, 8(1): 119–120, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcus:1973:BRM

[Mar73] Marvin Marcus. Book review: Multiparameter Eigenvalue Problems: Matrices and Compact Operators (F. V. Atkinson). SIAM Review, 15(3):678–679, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcus:1975:BRS

[Mar75a] Michael Marcus. Book review: Stochastic Processes and the Wiener Integral (James J. Yeh). SIAM Review, 17(1):182–183, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Margolis:1975:FEM

[Mar75b]  S. G. Margolis.  Finite-

element methods for unsteady compressible gas dynamics. SIAM Review, 17 (2):385, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marsden:1976:NF

[Mar76a] M. Marsden. Nonnegative form. SIAM Review, 18(4): 763–765, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marsden:1976:BRN

[Mar76b] Martin Marsden. Book review: A Nonnegative Form (M. M. Gupta). SIAM Review, 18(4):763–766, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcuvitz:1981:BRU

[Mar81] N. Marcuvitz. Book review:

Unidirectional Wave Motion (H. Levine). SIAM Review, 23(1):125–126, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcus:1983:PDL

[Mar83a]  M. Marcus.  Probabil-

ity distributions on linear spaces — Vakhania, N. N. SIAM Review, 25(4):594– 596, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcus:1983:BRP

[Mar83b] Michael Marcus. Book review: Probability Distributions on Linear Spaces (N. N. Vakhania). SIAM Review, 25 (4):594–596, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marchand:1985:BRO

[Mar85a] Erich W. Marchand. Book review: Optique de Fourier en Microscopie Électronique (A. Lannes and J. Ph. Pérez). SIAM Review, 27(4):600– 601, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcus:1985:PWG

[Mar85b] David J. Marcus. Probability of winning a game of racquetball. SIAM Review, 27 (3):443–444, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marsden:1985:BRB

[Mar85c] Jerry Marsden. Book review: Bifurcation Theory, Mechanics and Physics (C. P. Bruter, A. Aragnol and A. Lichnerowicz, eds.). SIAM Review, 27(1):116, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcuvitz:1986:BRS

[Mar86a] N. Marcuvitz. Book review: Sound Propagation in Stratified Fluids (Calvin H. Wilcox). SIAM Review, 28 (2):270–271, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marsland:1986:BRC

[Mar86b] T. A. Marsland. Book review: Chess Skill in Man and Machine-Second Edition (Peter W. Frey, ed.). SIAM Review, 28(1):115– 116, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Martin:1987:BRS

[Mar87] Robert H. Martin, Jr. Book review: Semigroups of Linear Operators and Applications (Jerome A. Goldstein). SIAM Review, 29(1): 151–153, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcuvitz:1988:BRW

[Mar88] Nathan Marcuvitz. Book review: Wave Propagation: an Invariant Imbedding Approach (R. Bellman and R. Vasudevan). SIAM Review, 30(1):147–148, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcotte:1989:SSB

[Mar89] Patrice Marcotte. To serve or not to serve — a badminton dilemma. SIAM Review, 31 (2):303–305, June 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcus:1993:CS

[Mar93] Marvin Marcus. A convex set. SIAM Review, 35 (1):136, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marcus:1994:CS

[Mar94]  M. Marcus.  A convex

set. SIAM Review, 36(1): 111, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marx:1995:BRM

[Mar95] Brian D. Marx. Book review: Multivariate Statistical Modelling Based on Generalized Linear Models (Ludwig Fahrmeir and Gerhard Tutz). SIAM Review, 37(4): 633–634, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Maritz:2021:RTD

[Mar21] Milton F. Maritz. Rotations in three dimensions.

SIAM Review, 63(2):395– 404, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Masur:1987:BRS

[Mas87] E. F. Masur. Book review:

Structural Optimization. Volume 1: Optimality Criteria (M. Save and W. Prager, eds.). SIAM Review, 29(1): 153–154, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Massell:1992:QWP

[Mas92] Paul B. Massell. A quick way to prove the existence of centers in twodimensional almost linear systems. SIAM Review, 34 (3):492–495, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mastin:1995:BRF

[Mas95] C. Wayne Mastin. Book review: Fundamentals of Grid Generation (P. Knupp and S. Steinberg). SIAM Review, 37 (2):264–265, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mathis:1969:CSU

[Mat69a] Robert F. Mathis. Completion of a symmetric unitary matrix. SIAM Review, 11(2):

261–263, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mathsen:1969:NST

[Mat69b] Ronald M. Mathsen. A note on solutions of third order ordinary differential equations. SIAM Review, 11(2): 283–284, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Matthews:1970:ID

[Mat70] James W. Matthews. Interpolation with derivatives. SIAM Review, 12(1):127– 128, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Matkowsky:1975:BLP

[Mat75] Bernard J. Matkowsky. On boundary layer problems exhibiting resonance. SIAM Review, 17(1):82–100, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [Mat76b].

Matkowsky:1976:C

[Mat76a] B. J. Matkowsky. Correction. SIAM Review, 18(1): 112, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Matkowsky:1976:EBL

[Mat76b] Bernard J. Matkowsky. Errata: On Boundary Layer Problems Exhibiting Resonance. SIAM Review, 18 (1):112, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Mat75].

Mattson:1980:BRT

[Mat80] H. F. Mattson, Jr. Book review: The Theory of Error-Correcting Codes (F. J. MacWilliams and N. J. A. Sloane). SIAM Review, 22 (4):513–519, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mattheij:1985:DSA

[Mat85] R. M. M. Mattheij. Decoupling and stability of algorithms for boundary value problems. SIAM Review, 27 (1):1–44, March 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mattson:1988:BRF

[Mat88] H. F. Mattson, Jr. Book review: A First Course in Coding Theory (Raymond Hill). SIAM Review, 30(1): 148–150, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mathis:1989:BRC

[Mat89a] Frank Mathis. Book review: Characterizations of Parabolic Motion (M. S. Klamkin). SIAM Review, 31 (1):129–131, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mathis:1989:EFR

[Mat89b] Frank Mathis. The effect of fatigue on running strategies. SIAM Review, 31(2): 306–309, June 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mathis:1991:GBP

[Mat91] Frank H. Mathis. A generalized birthday problem. SIAM Review, 33(2):265– 270, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mattson:1994:ECC

[Mat94a] H. F. Mattson. Error correcting codes-theory and applications (Alain Poli and Llorenc Huguet). SIAM Review, 36 (2):309–??, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mattson:1994:BRE

[Mat94b] H. F. Mattson, Jr. Book review: Error Correcting Codes-Theory and Applications (Alai Poli and Llorenc Huguet). SIAM Review, 36 (2):309–313, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1977:MI

[MATL77] M. H. Moore, W. Anderson, G. Trapp, and E. H. Lieb. Matrix inequality. SIAM Review, 19(2):329– 330, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mawhin:1986:BRC

[Maw86] Jean Mawhin. Book review: Convexity Methods in Variational Calculus (Peter Smith). SIAM Review, 28(3): 428–430, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mawhin:1991:BRS

[Maw91] Jean Mawhin. Book review:

Stability Analysis of Nonlinear Systems (V. Lakshmikantham, S. Leela, and A. A. Martynyuk). SIAM Review, 33(1):152–154, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mawhin:1993:BRN

[Maw93a] Jean Mawhin. Book review: Non-Linear Phenomena in Science and Engineering (George Rowlands). SIAM Review, 35(1):158– 159, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mawhin:1993:NLP

[Maw93b] Jean Mawhin. Non-linear phenomena in science and engineering (George Rowlands). SIAM Review, 35(1): 158–??, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mawhin:1995:BRD

[Maw95] Jean Mawhin. Book review: Duality and Perturbation Methods in Critical Point Theory (N. Ghoussoub). SIAM Review, 37(2): 265–266, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mawhin:1996:BRA

[Maw96] Jean Mawhin. Book review: Asymptotics of Nonlinearities and Operator Equations (Alexander M. Krasnosel’skii). SIAM Review, 38 (3):533–535, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Maxfield:1960:ISF

[Max60] John E. Maxfield. An iterative scheme for finding the real zeros of certain polynomials. SIAM Review, 2(2): 148–150, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mayer:1960:RTA

[May60] Arthur Mayer. Rotations and their algebra. SIAM Review, 2(2):77–122, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mayberry:1984:C

[May84a] J. P. Mayberry. Correction. SIAM Review, 26(4): 615, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mayberry:1984:FRB

[May84b] J. P. Mayberry. Fair representation — Balinski, M. L., Young, H. P. SIAM Review, 26(1):135–137, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mayberry:1984:BRE

[May84c] John P. Mayberry. Book review: Erratum: Fair Representation (Michel L. Balinski and H. Peyton Young). SIAM Review, 26(4):615, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mayberry:1984:BRF

[May84d] John P. Mayberry. Book review: Fair Representation (Michel L. Balinski and H. Peyton Young). SIAM Review, 26(1):135– 137, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mayle:1990:BRH

[May90] R. E. Mayle. Book review: Heat Transfer in Turbulent Mixed Convection (B. S. Petukhov and A. F. Polyakov; R. Hainsworth and G. Rybchinskaya, trans.; B. E. Launder, ed., English edition). SIAM Review, 32(2): 328–329, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Maybee:1991:BRP

[May91] John S. Maybee. Book review: Pearls in Graph Theory: a Comprehensive Introduction (N. Hartsfield and G. Ringel). SIAM Review, 33 (4):664–665, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mazaika:1984:LSU

[Maz84] Paul K. Mazaika. A lattice summation using the mean value theorem for harmonic functions. SIAM Review, 26 (1):113–115, January 1984.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marsaglia:1964:CMG

[MB64] G. Marsaglia and T. A.

Bray. A convenient method for generating normal variables. SIAM Review, 6(3): 260–264, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2027592.

Maslowe:1976:AEE

[MB76] S. A. Maslowe and D. J. Benney. Amplitude evolution equation for nonlinear waves in parallel shear flows. SIAM Review, 18(4):817, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moorhouse:1986:OCP

[MB86] J. C. Moorhouse and J. V.

Baxley. An optimal control problem. SIAM Review, 28 (1):85, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McLachlan:1983:BRF

[MBCG83a] B. G. McLachlan, G. Beaupre, A. B. Cox, and L. Gore. Book review: Falling Dominoes (D. E. Daykin). SIAM Review, 25 (3):403–404, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mclachlan:1983:FD

[MBCG83b] B. G. Mclachlan, G. Beaupre, A. B. Cox, and L. Gore. Falling dominoes. SIAM Review, 25(3):403–404, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mika:1997:SPE

[MBK97] J. R. Mika, J. Banasiak, and Hans G. Kaper. Singularly perturbed evolution equations with applications to kinetic theory. SIAM Review, 39(3):533– ??, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:2008:MCS

[MBK08] R. O. Moore, G. Biondini, and W. L. Kath. A method to compute statistics of large, noise-induced perturbations of nonlinear Schrödinger solitons. SIAM Review, 50(3): 523–549, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Metivier:2017:FWI

[MBOV17] L. Métivier, R. Brossier, S. Operto, and J. Virieux. Full waveform inversion and the truncated Newton method. SIAM Review, 59 (1):153–195, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mickens:1998:MD

[MBR98a] R. E. Mickens, D. N. Brewley, and M. L. Russell. A model of dieting. SIAM Review, 40 (3):667–672, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mickens:1998:CNM

[MBR98b] Ronald E. Mickens, Denise N.

Brewley, and Matasha L. Russell. Classroom note: a model of dieting. SIAM Review, 40(3):667–672, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/31874.

More:1976:NCN

[MC76] J. J. More and M. Cosnard. Numerical comparison of nonlinear equation solvers. SIAM Review, 18 (4):818, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacCluer:1991:CIP

[MC91] C. R. MacCluer and Yossi Chait. Choosing an inner product that separates variables. SIAM Review, 33 (3):467–471, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Menon:2021:ATU

[MC21] Shreya Menon and John W. Cain. Approaches toward understanding delayinduced stability and instability. SIAM Review, 63(3): 625–637, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McAvaney:1979:MNM

[McA79] K. L. McAvaney. A maximum number of matrices. SIAM Review, 21(1):139– 140, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McAllister:1987:ANR

[McA87] M. Luisa N. McAllister. An algorithm for networks of rules. SIAM Review, 29(2): 291–296, June 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McAllister:1988:BRF

[McA88] Marialuisa N. McAllister. Book review: Fuzzy Sets, Uncertainty, and Information (George J. Klir and Tina A. Folger). SIAM Review, 30(3): 520–522, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McAllister:1992:BRP

[McA92] M. Luisa N. McAllister. Book review: Possibility Theory: an Approach to Computerized Processing of Uncertainty (Didier Dubois and Henri Prade, with the collaboration of Henri Farreny, Roger Martin-Clouaire, and Claudette Testemale; E. F. Handing, trans.). SIAM Review, 34(1):147– 148, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McArthur:1994:NMB

[McA94a] K. M. McArthur. Numerical methods based on sinc and analytic functions (Frank Stenger). SIAM Review, 36 (4):673–??, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McArthur:1994:RNM

[McA94b]  K. M. McArthur.  Re-

view: Numerical Methods Based on Sinc and Analytic Functions (Frank Stenger). SIAM Review, 36(4):673– 674, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//links.jstor.org/sici?
sici=0036-1445%28199412%
2936%3A4%3C673%3ANMBOSA%
3E2.0.CO%3B2-J.

McArthur:1994:BRN

[McA94c] Kelly M. McArthur. Book review: Numerical Methods Based on Sine and Analytic Functions, by Frank Stenger. SIAM Review, 36 (4):673–674, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McAllister:1996:BRFa

[McA96a] Marialuisa N. McAllister. Book review: Fuzzy Logic: a Practical Approach (M. Mc-Neill and E. Thro). SIAM Review, 38(1):173–174, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McAllister:1996:BRFb

[McA96b] Marialuisa N. McAllister. Book review: Fuzzy Logic with Engineering Applications (Timothy Ross). SIAM Review, 38(1):174–175, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCully:1960:LT

[McC60] Joseph McCully. The Laguerre transform. SIAM Review, 2(3):185–191, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCormick:1971:BRN

[McC71] Garth P. McCormick. Book review: Nonlinear Programming (Olvi L. Mangasarian). SIAM Review, 13(2): 240–241, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mccormick:1978:SOR

[Mcc78] S. F. Mccormick. Simultaneous optimization of Rayleigh quotient for computing a few extremal solutions of Ax = ΛBx. SIAM Review, 20(3): 630–631, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McClamroch:1979:BRM

[McC79] N. H. McClamroch. Book review: Multivariate System Control (A. J. Fossard). SIAM Review, 21(3): 406–407, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McClamroch:1983:BRT

[McC83] N. Harris McClamroch.

Book review: Time-Variant Discrete-Time Systems (Gunter Ludyk). SIAM Review, 25 (1):116–117, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McClamroch:1985:BRG

[McC85] N. Harris McClamroch. Book review: A Guide to Simulation (Paul Bratley, Bennett L. Fox, and Linus E. Schrage). SIAM Review, 27 (2):292–293, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCoy:1986:BRC

[McC86] T. L. McCoy. Book review: Constructing Nodes for Optimal Quadrature (J. P. Lambert). SIAM Review, 28(2): 238–240, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCabe:1991:BRC

[McC91a] J. H. McCabe. Book review: Continued Fractions in Statistical Applications (K. O. Bowman and L. R. Shenton). SIAM Review, 33(2): 329–330, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCall:1991:BRM

[McC91b] Brian P. McCall. Book review: Multi-Armed Bandit Allocation Indices (J. C. Gittins). SIAM Review, 33(1): 154–155, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCaffrey:1992:BRG

[McC92a] Daniel F. McCaffrey. Book review: Generalized Additive Models (T. J. Hastie and R. J. Tibshirani). SIAM Review, 34(4):675–678, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCartin:1992:AAD

[McC92b] Brian J. McCartin. An alternative analysis of Duffing’s equation. SIAM Review, 34 (3):482–491, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCabe:1993:BRH

[McC93a] J. H. McCabe. Book review: History of Continued Fractions and Padé Approximants (Claude Brezinski). SIAM Review, 35(1): 159–160, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCabe:1993:HCF

[McC93b]  J. H. McCabe.  History

of continued fractions and Padé approximants (Claude Brezinski). SIAM Review, 35 (1):159–??, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCormick:1995:BRM

[McC95] Steve McCormick. Book review: Mathematical Computational Techniques for Multilevel Adaptive Methods (U. Rüde). SIAM Review, 37(2): 266–269, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McCartin:2003:EET

[McC03] Brian J. McCartin. Eigenstructure of the equilateral triangle, Part I: The Dirichlet problem. SIAM Review, 45(2):267–287, June 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38720.

McGregor:1961:BRM

[McG61]  J. L. McGregor.  Book

review: Modern Probability Theory and Its Applications (Emanuel Parzen). SIAM Review, 3(2):184– 185, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGee:1962:BM

[McG62a] W. C. McGee. A binary multiplication. SIAM Review, 4 (3):256, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGee:1962:BRP

[McG62b] W. C. McGee. Book review: Programming and Coding for Automatic Digital Computers (G. W. Evans II and C. L. Perky). SIAM Review, 4(2): 157–158, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGrath:1964:BRP

[McG64] R. J. McGrath. Book review: Progress In Control Engineering, Vol. I. (R. H. Macmillan, T. J. Higgins, P. Naslin, eds.). SIAM Review, 6(1):80–81, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGrath:1965:BRM

[McG65a] R. J. McGrath. Book review:

Modern Control Theory (J. T. Tou). SIAM Review, 7(4): 582–583, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGrath:1965:BRO

[McG65b] Richard J. McGrath. Book review: Optimization Theory and the Design of Feedback Control Systems (C. W. Merriam III). SIAM Review, 7 (2):296–297, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGee:1967:BRA

[McG67a] William C. McGee. Book review: Advances in Programming & Non-Numerical Computation (L. Fox). SIAM Review, 9(1):133–136, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGee:1967:BRL

[McG67b] William C. McGee. Book review: Logic and Algorithms: With Applications to the Computer and Information Sciences (Robert R. Korfhage). SIAM Review, 9(3): 612–613, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGarvey:1982:AGS

[McG82] James F. McGarvey. Approximating the general solution of a differential equation. SIAM Review, 24(3): 333–337, July 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McGrath:1993:BRP

[McG93] Joseph F. McGrath. Book review: A Primer on Integral Equations of the First Kind: The Problem of Deconvolution and Unfolding (G. Milton Wing with the assistance of John D. Zahrt). SIAM Review, 35(2):328– 330, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McIlroy:1971:BRV

[McI71] M. D. McIlroy. Book review: A View of Programming Languages (B. A. Galler and A. J. Perlis). SIAM Review, 13 (4):583–584, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McKelvey:1981:BRM

[McK81] Robert McKelvey. Book review: Models for Public Systems Analysis (Edward J. Beltrami). SIAM Review, 23 (1):118–120, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McKeague:1995:BRS

[McK95] Ian W. McKeague. Book review: Statistical Models Based on Counting Processes (P. K. Andersen, Ø. Borgan; R. D. Gill, and N. Keiding). SIAM Review, 37(3): 475–476, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McLeish:1979:BRP

[McL79]  D. McLeish.  Book re-

view: A Probabilistic Inequality (L. A. Shepp and A. M. Odlyzko). SIAM Review, 21 (4):564–565, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McLaughlin:1986:AMR

[McL86a] Joyce R. McLaughlin. Analytical methods for recovering coefficients in differential equations from spectral data. SIAM Review, 28(1):

53–72, March 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McLeod:1986:BRM

[McL86b] Robin McLeod. Book review: The Mathematical Description of Shape and Form (E. A. Lord and C. B. Wilson). SIAM Review, 28 (4):597–599, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McMillan:1961:EAT

[McM61a] Brockway McMillan. An elementary approach to the theory of information. SIAM Review, 3(3):211–229, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McMorris:1961:AMS

[McM61b]  A. H. McMorris.  An-

other minimum switching network. SIAM Review, 3 (1):70, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McMorris:1961:BRM

[McM61c] A. H. McMorris. Book review: A Minimum Switching Network (Raphael Miller). SIAM Review, 3(1):72–74, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McMillan:1962:MTU

[McM62] Brockway McMillan. Mathematicians and their uses. SIAM Review, 4(2):79–90, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McMorris:1987:BRI

[McM87] F. R. McMorris. Book review: Interval Orders and Interval Graphs — a Study of Partially Ordered Sets (Peter C. Fishburn). SIAM Review, 29(3):484–486, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McNaughton:1959:BRS

[McN59] Robert McNaughton. Book review: Switching Circuits and Logical Design (Samuel H. Caldvvell). SIAM Review, 1(2):181–182, ???? 1959. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mcneil:1966:AEL

[Mcn66] M. B. Mcneil. Asymptotic expansion of a Laplace transform. SIAM Review, 8(2): 234, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McNeil:1967:IRR

[McN67] M. B. McNeil. An integral recurrence relation. SIAM Review, 9(2):250, ???? 1967.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McNaughton:1989:BRI

[McN89a] Robert McNaughton. Book review: Introduction to Formal Languages (G. E. Revesz). SIAM Review, 31 (2):348, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McNaughton:1989:BRV

[McN89b] Robert McNaughton. Book review: Varieties of Formal Languages (J. E. Pin; A. Howie, trans.). SIAM Review, 31(2):347–348, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mufti:1969:SIC

[MCS69] I. H. Mufti, C. K. Chow, and F. T. Stock. Solution of ill-conditioned linear twopoint boundary value problems by the Riccati transformation. SIAM Review, 11(4): 616–619, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Michelotti:1996:ICD

[MD96] Leo Michelotti and H. Scott Dumas. Intermediate classical dynamics with applications to beam physics. SIAM Review, 38(4):685– ??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meade:1986:BRF

[Mea86] Kevin P. Meade. Book review: The Finite Element Method Displayed (Gouri Dhatt and Gilbert Touzot). SIAM Review, 28(3):430– 431, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meade:1988:BRD

[Mea88a] Kevin P. Meade. Book review: Difference Equations (Ronald E. Mickens). SIAM Review, 30(4):675, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meade:1988:BRF

[Mea88b]  Kevin P. Meade.  Book

review: Finite Difference Methods in the Dynamics of Continuous Media (Julian L. Davis). SIAM Review, 30 (1):150, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meade:1998:OMP

[Mea98] Douglas B. Meade. ODE models for the parachute problem. SIAM Review, 40(2):327–332, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31624.

Mecke:1992:BRF

[Mec92] J. Mecke. Book review: Factorization Calculus and Geometric Probability (R. V. Ambartzumian). SIAM Review, 34(3):514, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Medick:1962:ZIF

[Med62] M. A. Medick. On a zero of an implicit function. SIAM Review, 4(1):49, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meditch:1978:BRS

[Med78] James S. Meditch. Book review: State-Space Methods for Control Systems (F. Csaki). SIAM Review, 20(4): 864–865, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meeker:1996:BRS

[Mee96] William Q. Meeker. Book review: System Reliability Theory: Models and Statistics Methods (Arnljot Høyland and Marvin Rausand). SIAM Review, 38(1): 175–177, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Megiddo:1978:ETC

[Meg78] N. Megiddo. Edge threecoloring of tournaments. SIAM Review, 20(3):593, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Megiddo:1979:ETC

[Meg79] N. Megiddo. Edge threecoloring of tournaments. SIAM Review, 21(3):397– 398, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Megiddo:1988:BRT

[Meg88] Nimrod Megiddo. Book review: Tree Network and Planar Rectilinear Location Theory (A. J. W. Kolen). SIAM Review, 30(2):340– 341, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mehran:1973:BRM

[Meh73] Farhad Mehran. Book review: Mean and Variance of a Function of a Random Permutation (E. Hauer and J. G. C. Templeton). SIAM Review, 15(4):800– 801, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mehta:1974:MF

[Meh74a] M. L. Mehta. Monotonicity of a function. SIAM Review, 16(1):91–92, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mehta:1974:NC

[Meh74b] M. L. Mehta. A normalization constant. SIAM Review, 16(1):92–93, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mehta:1974:PCO

[Meh74c] M. L. Mehta. Positivity conditions for orthogonal polynomials. SIAM Review, 16 (2):257–258, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mehta:1974:TMI

[Meh74d] M. L. Mehta. Three multiple integrals. SIAM Review, 16 (2):256–257, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mehrmann:1996:BRA

[Meh96] Volker Mehrmann. Book review: Algebraic Riccati Equations (Peter Lancaster and Leiba Rodman). SIAM Review, 38(4):694–695, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meicler:1968:CSI

[Mei68] Marcel Meicler. Chebyshev solution of an inconsistent system of n + 1 linear equations in n unknowns in terms of its least squares solution. SIAM Review, 10(3): 373–375, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meisters:1977:BRU

[Mei77] Gary H. Meisters. Book review: Unitary Representations and Harmonic Analysis (Mitsuo Sugiura). SIAM Review, 19(3):579– 582, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meir:1978:BRP

[Mei78a] A. Meir. Book review: Percentiles for the Gamma Distribution (R. A. Waller and M. S. Waterman). SIAM Review, 20(1):186–187, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meir:1978:II

[Mei78b] A. Meir. An integral inequality. SIAM Review, 20 (4):855, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meir:1983:BRC

[Mei83a]  A. Meir.  Book review:

A Covering Problem (D. J. Newman). SIAM Review, 25 (1):99–101, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meir:1983:CP

[Mei83b] A. Meir. A covering problem. SIAM Review, 25(1): 99–100, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meisters:1989:BRG

[Mei89] Gray H. Meisters. Book review: Global Stability of Dynamical Systems (Michael Shub with the collaboration of Albert Fathi and Remi Langevin; Joseph Christy, trans.). SIAM Review, 31(1): 147–151, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meinsma:2019:DSA

[Mei19] Gjerrit Meinsma. Dimensional and scaling analysis. SIAM Review, 61(1): 159–184, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Melzak:1964:EVP

[Mel64] Z. A. Melzak. Estimating the volume from parallel crosssection areas. SIAM Review, 6(2):131–133, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Melville:1991:LID

[Mel91] John Melville. A linear integro-differential equation. SIAM Review, 33(4):655– 656, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Melville:1992:LIE

[Mel92a] J. Melville. A linear integrodifferential equation. SIAM Review, 34(4):654– 656, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Melville:1992:BRL

[Mel92b] John Melville. Book review: A Linear Integro-Differential Equation (John Melville). SIAM Review, 34 (4):654–656, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Melville:1993:LIE

[Mel93] J. Melville. A linear integrodifferential equation (vol 34, pg 654, 1992). SIAM Review, 35(3):505, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Melman:1997:CNG

[Mel97] A. Melman. Classroom note: Geometry and convergence of Euler’s and Halley’s methods. SIAM Review, 39(4): 728–735, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30114.

Melman:2012:MGD

[Mel12] Aaron Melman. Modified Gershgorin disks for companion matrices. SIAM Review, 54(2):355–373, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Menaldi:1994:BRC

[Men94a] Jose Luis Menaldi. Book review: Controlled Markov Processes and Viscosity Solutions (Wendell H. Fleeting and H. Mete Soner). SIAM Review, 36(1):133– 134, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meneguette:1994:ZUD

[Men94b] M. Meneguette. Zeros in the unit disk. SIAM Review, 36 (4):656–657, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Merkes:1964:BRA

[Mer64] E. P. Merkes. Book review:

The Application of Continued Fractions and Their Generalizations to Problems in Approximation Theory (A. B. Khovanskii). SIAM Review, 6 (2):188–189, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Merrill:1981:BRM

[Mer81] Stephen J. Merrill. Book review: Mathematical Models in Cell Biology and Cancer Chemotherapy (Martin Eisen). SIAM Review, 23(3): 401–402, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Merilees:1982:TS

[Mer82] P. E. Merilees. A triple sum. SIAM Review, 24 (2):223, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Merilees:1983:TS

[Mer83] P. E. Merilees. A triple sum. SIAM Review, 25(2): 269–271, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Merz:1984:CCI

[Mer84a] G. Merz. Complex contour integral representation of cardinal spline functions — Schempp, W. SIAM Review, 26(1):130–131, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Merz:1984:BRC

[Mer84b] Gerhard Merz. Book review: Complex Contour Integral Representation of Cardinal Spline (Walter Schempp). SIAM Review, 26(1):130– 131, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Merrill:1985:BRR

[Mer85] Stephen Merrill. Book review: Random Perturbations of Dynamical Systems (M. I. Freidlin and A. D. Wentzell). SIAM Review, 27 (2):295–296, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Merrill:1989:SDS

[Mer89] Stephen J. Merrill. Stochastic differential systems: Analysis and filtering. SIAM Review, 31(1):147, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1969:RP

[Mey69a] Carl D. Meyer, Jr. On ranks of pseudoinverses. SIAM Review, 11(3):382– 385, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1969:CSM

[Mey69b] Carl D. Meyer, Jr. On the construction of solutions to the matrix equations AX = A and Y A = A. SIAM Review, 11(4): 612–615, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1972:BRO

[Mey72] Kenneth R. Meyer. Book review: Ordinary Differential Equations (Jack K. Hale. John Wiley). SIAM Review, 14(2):348–350, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1975:FSF

[Mey75a] G. G. L. Meyer. Finite solvability of a family of approximating problems. SIAM Review, 17(2):385, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1975:BRQ

[Mey75b] Kenneth R. Meyer. Book review: Qualitative Theory of Second-Order Dynamic Systems (A. A. Andronov, E. A. Leontovich, I. I. Gordon and A. G. Maier); Theory of Bifurcations of Dynamic Systems on a Plane (A. A. Andronov, E. A. Leontovich, I. I. Gordon and A. G. Maier). SIAM Review, 17 (1):183–184, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1975:RGG

[Mey75c] Carl D. Meyer, Jr. The role of the group generalized inverse in the theory of finite Markov chains. SIAM Review, 17(3): 443–464, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1976:LOD

[Mey76a] G. H. Meyer. Locally onedimensional methods for multidimensional parabolic freesurface problems. SIAM Review, 18(4):818, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1976:IPP

[Mey76b] R. R. Meyer. Integer programming — possibilities and impossibilities. SIAM Review, 18(4):818, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1976:BRG

[Mey76c] Carl D. Meyer, Jr. Book review: Generalized Inverses (Theory and Applications) (Adi Ben–Israel and Thomas N. E. Greville).

SIAM Review, 18(2):320– 322, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1977:ODP

[Mey77a] G. H. Meyer. One-dimensional parabolic free boundary problems. SIAM Review, 19 (1):17–34, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1977:BRL

[Mey77b] Kenneth R. Meyer. Book review: Linear Differential Equations with Periodic Coefficients (V. A. Yakubovich and V. M. Starzhinskii). SIAM Review, 19(1):166– 167, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1978:BRF

[Mey78] W. W. Meyer. Book review:

A First Order Nonlinear Differential Equation (I. Rubinstein). SIAM Review, 20(4): 858–859, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1979:BRI

[Mey79] R. R. Meyer. Book review: Integer and Mixed Programming; Theory and Applications (Arnold Kaufmann and Arnaud Henry-Labordère). SIAM Review, 21(3):411–412, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1980:EA

[Mey80a] R. E. Meyer. Exponential asymptotics. SIAM Review, 22(2):213–224, April 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1980:EVI

[Mey80b] W. W. Meyer. Extreme value of an integral. SIAM Review, 22(3):363, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1982:BRT

[Mey82a] Kenneth R. Meyer. Book review: Theory and Applications of Hopf Bifurcation (D. D. Hassard, N. D. Kazarinoff and Y-H Wan). SIAM Review, 24(4):498– 499, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1982:BRP

[Mey82b] R. E. Meyer. Book review: Perturbation Methods in Applied Mathematics (J. Kevorkian and J. D. Cole). SIAM Review, 24 (4):501, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1983:BRI

[Mey83] K. R. Meyer. Book review:

An Introduction to Nonlinear Oscillations (Ronald Elbert Mickens). SIAM Review, 25 (2):279–280, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1984:BRR

[Mey84] K. R. Meyer. Book review: Regular and Stochastic Motion (A. J. Lichtenberg and M. A. Lieberman). SIAM Review, 26(2): 298–299, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1986:CED

[Mey86a]  K. R. Meyer.  Counter-

examples in dynamical systems via normal form theory. SIAM Review, 28(1): 41–51, March 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1986:BRS

[Mey86b] Kenneth R. Meyer. Book review: Singularities and Groups in Bifurcation Theory, Volume I (Martin Golubitsky and David G. Schaeffer). SIAM Review, 28(2): 271–272, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1988:BRH

[Mey88] Kenneth R. Meyer. Book review: The Hamiltonian Hopf Bifurcation (Jan-Cees van der Meer). SIAM Review, 30(3):522–523, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1989:SCU

[Mey89a] Carl D. Meyer. Stochastic complementation, uncoupling Markov chains, and the theory of nearly reducible systems. SIAM Review, 31 (2):240–272, June 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1989:BRD

[Mey89b] Kenneth R. Meyer. Book review: Dynamical Systems III (V. I. Arnold, ed.). SIAM Review, 31(3):510, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1989:BRS

[Mey89c] Kenneth R. Meyer. Book review: Singularities and Groups in Bifurcation Theory. Volume II (Martin Golubitsky, Ian Stewart, and David G. Schaeffer). SIAM Review, 31(4):703–704, ???? 1989. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1989:SES

[Mey89d] R. E. Meyer. A simple explanation of the Stokes phenomenon. SIAM Review, 31 (3):435–445, September 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1990:BRT

[Mey90] W. Weston Meyer. Book review: Three Particles Moving on a Circle (Murray S. Klamkin). SIAM Review, 32 (2):307–309, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1991:BRI

[Mey91a] Kenneth R. Meyer. Book review: An Introduction to Dynamical Systems (D. K. Arrowsmith and C. M. Place). SIAM Review, 33(4): 664, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1991:BRT

[Mey91b] W. Weston Meyer. Book review: Three Particles Moving on a Circle (Murray S. Klamkin). SIAM Review, 33 (1):118–119, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1993:UVS

[Mey93a] Rudolf X. Meyer. A unique vector sum. SIAM Review, 35(3):488, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1993:BRP

[Mey93b] Y. Meyer. Book review: Pseudodifferential Operators and Nonlinear Partial Differential Equations (M. E. Taylor). SIAM Review, 35 (4):671–672, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1994:BRA

[Mey94a] Gunter H. Meyer. Book review: Applied Integral Transforms (M. Ya. Antimirov, A. A. Kolyshin, and Remi Vaillancourt). SIAM Review, 36 (4):674–675, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1994:UVS

[Mey94b] R. X. Meyer. A unique vector sum. SIAM Review, 36 (3):495–496, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1995:BRN

[Mey95a] Kenneth R. Meyer. Book review: Normal Forms and Bifurcation of Planar Vector Fields (Shee-Nee Chow, Chengzhi Li, and Duo Wang). SIAM Review, 37(2): 269–270, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1995:BRS

[Mey95b] Kenneth R. Meyer. Book review: Singularity Theory and Equivariant Symplectic Maps (Thomas J. Bridges and Jacques E. Furter). SIAM Review, 37(1):135, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1995:MS

[Mey95c] Rudolf X. Meyer. Motion of a string. SIAM Review, 37 (3):441, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1996:BRM

[Mey96] Kenneth R. Meyer. Book review: Math into L^{A}TEX: an Introduction to L^{A}TEX and AmS-L^{A}TEX (George Gräztzer). SIAM Review, 38 (3):535, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Matkowsky:1976:VAB

[MG76] B. J. Matkowsky and J. Grasman. Variational approach to boundary-layer problems and resonance for ordinary and partial-differential equations. SIAM Review, 18 (4):817, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mcleish:1979:PI

[MG79] D. Mcleish and G. Gripenberg. Probabilistic inequality. SIAM Review, 21(4): 564–565, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Myers:1995:BRI

[MG95] Christopher W. Myers and Frank H. Guenther. Book review: 4An Introduction to the Modeling of Neural Networks (P. Peretto). SIAM Review, 37(4):629–632, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mei:1996:MAE

[MG96] Chiang C. Mei and Jurgen Gerlach. Mathematical analysis in engineering. SIAM Review, 38(2): 342–??, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murphy:2023:PSM

[MG23] Brigid Murphy and Paul Glendinning. Piecewise smooth models of pumping a child’s swing. SIAM Review, 65(3):887–902, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1268574.

Maslov:1996:CWM

[MH96] Victor P. Maslov and Evans M.

Harrell II. The complex WKB method for nonlinear equations 1: Linear theory. SIAM Review, 38(1): 168–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McNamara:2001:OMB

[MHC01] J. M. McNamara, A. I. Houston, and E. J. Collins. Optimality models in behavioral biology. SIAM Review, 43 (3):413–466, September 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38526.

Mickens:1988:ESP

[Mic88] Ronald E. Mickens. Exact solutions to a population model: The logistic equation with advection. SIAM Review, 30(4):629– 633, December 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Michael:1992:CRS

[Mic92a] Anthony N. Michael. Code recognition and set selection with neural networks (Clark Jeffries). SIAM Review, 34 (4):678–??, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Michel:1992:BRC

[Mic92b] Anthony N. Michel. Book review: Code Recognition and Set Selection with Neural Networks (Clark Jeff fries). SIAM Review, 34(4):678– 679, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Michalowski:1995:PC

[Mic95a] B. Michalowski. Path of a car. SIAM Review, 37(2): 241–244, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Michalowski:1995:BRP

[Mic95b] Brian Michalowski. Book review: Path of a Car (R. C. Moore). SIAM Review, 37 (2):241–244, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mielke:1995:BRA

[Mie95] Alexander Mielke. Book review: Asymptotic Behaviour of Solutions of Evolutionary Equations (M. I. Vishik). SIAM Review, 37(3):471– 472, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mierczynski:2017:ILC

[Mie17] Janusz Mierczyński. Instability in linear cooperative systems of ordinary differential equations. SIAM Review, 59 (3):649–670, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mikelic:1989:BRM

[Mik89] Andro Mikelić. Book review: Mathematical Problems of Statistical Hydromechanics (M. J. Vishik and A. V. Fursikov). SIAM Review, 31(4): 704–706, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1959:MSN

[Mil59] Raphael Miller. A minimum switching network. SIAM Review, 1(2):171– 172, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1960:MSN

[Mil60] Raphael Miller. A minimum switching network. SIAM Review, 2(3):219– 222, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1961:BRP

[Mil61a] Raymond E. Miller. Book review: Proceedings of an International Symposium on the Theory of Switching, Parts I and II. SIAM Review, 3(1):81–84, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mills:1961:SSP

[Mil61b] Harlan D. Mills. Smoothing in servo processes. SIAM Review, 3(2):131–139, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Millar:1966:ABT

[Mil66] R. F. Millar. On the asymptotic behavior of two classes of integrals. SIAM Review, 8 (2):188–195, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Millham:1968:SEP

[Mil68] C. B. Millham. On the structure of equilibrium points in bimatrix games. SIAM Review, 10(4):447–448, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miles:1969:BRP

[Mil69a] Roger E. Miles. Book review: Probability Distribution of a Network of Triangles (Mary Beth Stearns). SIAM Review, 11(3):399–402, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1969:CGP

[Mil69b] K. S. Miller. Complex Gaussian processes. SIAM Review, 11(4):544–567, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1973:CLL

[Mil73] Kenneth S. Miller. Complex linear least squares. SIAM Review, 15(4):706– 726, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1976:UMR

[Mil76] O. W. Miller. Use of multiple regression in developing work measurement standards — proposed application. SIAM Review, 18(4): 818, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Millar:1980:BRE

[Mil80] Robert F. Millar. Book review: Electromagnetic Scattering (Piergiorgio L. E. Uslenghi, ed.). SIAM Review, 22(3):385–387, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mills:1981:BRD

[Mil81a] W. H. Mills. Book review: Difference Sets in Elementary Abelian Groups (Paul Camion). SIAM Review, 23 (2):270–271, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mills:1981:BRO

[Mil81b] W. H. Mills. Book review: Orthogonal Designs: Quadratic Forms and Hadamard Matrices (Anthony V. Geramita and Jennifer Seberry). SIAM Review, 23(1):116–117, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miles:1982:BRH

[Mil82a] John Miles. Book review: Hydrodynamic Stability (P. G. Drazin and W. H. Reid). SIAM Review, 24(3):351– 354, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Millar:1982:BRW

[Mil82b] Robert F. Millar. Book review: Waves in Layered Media (L. M. Brekhovskikh). SIAM Review, 24(2):239– 241, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1982:BRM

[Mil82c] Webb Miller. Book review:

Matrix Computations and Mathematical Software (John R. Rice). SIAM Review, 24 (2):232–233, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miles:1983:BRE

[Mil83a] John W. Miles. Book review: Elements of Soliton Theory (G. L. Lamb, Jr.). SIAM Review, 25(1): 118–121, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mills:1983:AMO

[Mil83b] W. H. Mills. Algorithmic methods in optimal control — Gruver, W. A., Sach, E. SIAM Review, 25(4): 579–580, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mills:1983:BRA

[Mil83c] Wendell H. Mills. Book review: Algorithmic Methods in Optimal Control (W. A. Gruver and E. Sach). SIAM Review, 25(4):579– 580, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miles:1989:CWI

[Mil89a] J. W. Miles. A complement of Weber’s integral. SIAM Review, 31(4):675– 676, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milton:1989:BRB

[Mil89b] John G. Milton. Book review: Biograph: a Graphical Computer Simulation Package with Exercises to Accompany Lee A. Segel’s Modeling Dynamic Phenomena in Molecular Cellular Biology (Garrett M. Odell and Lee A. Segel). SIAM Review, 31(1): 151–153, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milton:1990:BDS

[Mil90a] John G. Milton. Biological delay systems: Linear stability theory (N MacDonald). SIAM Review, 32(3): 514–??, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milton:1990:BRB

[Mil90b] John G. Milton. Book review: Biological Delay Systems: Linear Stability Theory (N. MacDonald). SIAM Review, 32(3):514–516, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milton:1991:BRD

[Mil91a] John G. Milton. Book review: Depth Perception in Frogs and Toads: a Study in Neural Computing (Donald House). SIAM Review, 33 (1):155–156, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milton:1991:DPF

[Mil91b] John G. Milton. Depth perception in frogs and toads: a study in neural computing (Donald House). SIAM Review, 33(1):155– ??, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1992:BRP

[Mil92] A. R. Miller. Book review:

A Probability Integral (Larry G. Blaine). SIAM Review, 34 (2):312–315, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miles:1993:EBR

[Mil93] John Miles. The electric Bell revisited. SIAM Review, 35 (2):289–293, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1994:BRI

[Mil94a] A. R. Miller. Book review: An Infinite Integral of Four Bessel Functions (M. L. Glasser). SIAM Review, 36 (2):285–287, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milton:1994:BRM

[Mil94b] John Milton. Book review: Mathematics in Medicine and the Life Sciences (F. C. Hoppensteadt and C. S. Peskin). SIAM Review, 36(1): 134–135, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1995:LTI

[Mil95a] A. R. Miller. Laplace transform involving the Macdonald function. SIAM Review, 37(3):442–443, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1995:BRL

[Mil95b] Allen R. Miller. Book review: Laplace Transform Involving the Macdonald Function (M. Aslam Chaudhry and S. M. Zubair). SIAM Review, 37 (3):442–443, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milgram:1998:TPA

[Mil98] Judah Milgram. Turn performance of aircraft, revisited. SIAM Review, 40 (4):957–958, December 1998.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33180.

Minty:1966:BRF

[Min66] George J. Minty. Book review: Finite Graphs and Networks (Robert G. Busacker and Thomas L. Saaty). SIAM Review, 8(2):243– 245, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Minieka:1970:CP

[Min70] Edward Minieka. The mcenter problem. SIAM Review, 12(1):138–139, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Minerbo:1978:MME

[Min78] G. N. Minerbo. Ment — maximum entropy algorithm for computerized axial tomography. SIAM Review, 20 (3):631, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miranker:1971:SPN

[Mir71] W. L. Miranker. A survey of parallelism in numerical analysis. SIAM Review, 13(4): 524–547, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mironer:1985:BRF

[Mir85] Alan Mironer. Book review: A First Course in Fluid Dynamics (A. R. Paterson). SIAM Review, 27(2): 294–295, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mironer:1988:BRI

[Mir88] Alan Mironer. Book review: An Informal Introduction to Theoretical Fluid Mechanics (James Lighthill). SIAM Review, 30(3):523– 524, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miranker:1997:IEC

[Mir97] Willard L. Miranker. Interference effects in computation. SIAM Review, 39 (4):630–643, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29626.

Mitchell:1965:BRH

[Mit65] Josephine Mitchell. Book review: Harmonic Analysis of Functions of Several Complex Variables in the Classical Domains (L. K. Hua). SIAM Review, 7(1): 148–149, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mitra:1987:BRG

[Mit87] Nilotpal Mitra. Book review:

Group Theory in Physics, Vols. I and II (J. F. Cornwell). SIAM Review, 29(1): 154–155, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mityagin:1991:ILA

[Mit91a] B. Mityagin. An inequality in linear algebra. SIAM Review, 33(1):125–127, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mityagin:1991:BRI

[Mit91b] Boris Mityagin. Book review:

An Inequality in Linear Algebra (Petter E. Bjørstad). SIAM Review, 33(1):125– 127, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mitchell:1993:BRV

[Mit93] John C. Mitchell. Book review: Verification of Sequential and Concurrent Programs (Krzysztof R. Apt and Ernst-Rudiger Olderog). SIAM Review, 35(2):330– ??, June 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mittelmann:1994:BRP

[Mit94] H. Mittelmann. Book review: PLTMG: a Software Package for Solving Elliptic Partial Differential Equations (Randolph E. Bank). SIAM Review, 36(1):135– 136, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miura:1975:KVE

[Miu75] R. M. Miura. Korteweg–de Vries equation — survey of results. SIAM Review, 17(2): 385–386, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miura:1976:BRL

[Miu76a] Robert M. Miura. Book review: Linear and Nonlinear Waves (G. B. Whitham). SIAM Review, 18(4):776– 781, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miura:1976:KVE

[Miu76b]  Robert M. Miura.  The

Korteweg–de Vries equation: a survey of results. SIAM Review, 18(3):412– 459, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [Miu77b].

Miura:1977:C

[Miu77a]  R. M. Miura.  Correc-

tion. SIAM Review, 19 (4):R6, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miura:1977:EKV

[Miu77b]  Robert M. Miura.  Er-

rata: The Korteweg–de Vries Equations: a Survey of Results. SIAM Review, 19 (4):vi, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Miu76b].

Miller:1961:MI

[MJ61] Arthur E. Miller and Bradley M.

Johnston. A multiple integral. SIAM Review, 3(1): 70–71, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Matula:1978:DNR

[MK78] D. W. Matula and P. Kornerup. Distribution of numbers representable in finite precision number systems. SIAM Review, 20 (3):630, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morton:1997:NSC

[MK97] K. W. Morton and R. Bruce Kellogg. Numerical solution of convection-diffusion problems. SIAM Review, 39(3): 535–??, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mahadevan:1999:PFT

[MK99] L. Mahadevan and Joseph B. Keller. Periodic folding of thin sheets. SIAM Review, 41(1):115–131, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33916. Reprinted

from SIAM Journal on Applied Mathematics, vol 55, pg 1609–1624, 1995.

Madsen:1976:MMD

[MKR76] N. K. Madsen, J. I. Karush, and G. H. Rodrigue. Matrix multiplication by diagonals on vector-parallel processors. SIAM Review, 18 (4):816, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mccoy:1986:CNO

[MKS86] T. L. Mccoy, Z. K. Kabala, and R. E. Shafer. Constructing nodes for optimal quadrature. SIAM Review, 28(2): 238–240, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mow:1980:RDS

[ML80] Van C. Mow and W. Michael Lai. Recent developments in synovial joint biomechanics. SIAM Review, 22(3): 275–317, July 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McKelvey:1985:BRP

[ML85] Robert McKelvey and Roland Lamberson. Book review: Political and Related Models (Steven J. Brams, William F. Lucas and Philip D. Strafin, Jr., eds.). SIAM Review, 27 (2):293–294, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mickens:1998:OPD

[ML98] Ronald E. Mickens and Mark Levi. Oscillations in planar dynamic systems. SIAM Review, 40(1):162– ??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mladineo:1994:BRM

[Mla94] Regina Hunter Mladineo.

Book review: Model Building in Mathematical Programming (H. P. Williams). SIAM Review, 36(2):313– 315, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mladenov:1997:BRG

[Mla97] Ivailo M. Mladenov. Book review: Group Theory and Physics, by S. Sternberg. SIAM Review, 39(2):352– 354, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Maxfield:1981:BT

[MM81] Margaret W. Maxfield and Naomi J. McCarty. The best time to die. SIAM Review, 23 (3):385–389, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Manas:1989:PCI

[MM89] G. J. Manas and D. H. Meyer. On a problem of coin identification. SIAM Review, 31(1): 114–117, March 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mathis:1995:NPA

[MM95] Frank H. Mathis and Lenora Jane Mathis. A nonlinear programming algorithm for hospital management. SIAM Review, 37(2):230–234, June 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

                                        http://epubs.siam.org/
                                        23719.htm.

McCluskey:1998:SIB

[MM98] C. Connell McCluskey and James S. Muldowney. Stability implications of Bendixson’s criterion. SIAM Review, 40(4):931–934, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32844.

Miller:2006:DSC

[MM06] Allen R. Miller and Ira S.

Moskowitz. Difference of sums containing products of binomial coefficients and their logarithms. SIAM Review, 48(2):318–325, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mitchell:2010:ASR

[MM10] S. L. Mitchell and T. G. Myers. Application of standard and refined heat balance integral methods to one-dimensional Stefan problems. SIAM Review, 52(1): 57–86, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moennigmann:2007:HAE

[MMB^{+}07a] Martin Moennigmann, Wolfgang Marquardt, Christian H. Bischof, Thomas Beelitz, Bruno Lang, and Paul Willems. A hybrid approach for efficient robust design of dynamic systems. SIAM Review, 49(2): 236–254, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Monnigmann:2007:HAE

[MMB^{+}07b] Martin Mönnigmann, Wolfgang Marquardt, Christian H. Bischof, Thomas Beelitz, Bruno Lang, and Paul Willems. A hybrid approach for efficient robust design of dynamic systems. SIAM Review, 49(2): 236–254, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Muller:2004:SVD

[MMH04] Neil Muller, Lourenço Magaia, and B. M. Herbst. Singular value decomposition, eigenfaces, and 3D reconstructions. SIAM Review, 46 (3):518–545, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38751.

Mattheij:1998:ODE

[MMS98] R. M. M. Mattheij, J. Molenaar, and Lawrence F. Shampine. Ordinary differential equations in theory and practice. SIAM Review, 40 (1):175–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mengaldo:2021:IRI

[MMT^{+}21] Gianmarco Mengaldo, David Moxey, Michael Turner, Rodrigo Costa Moura, Ayad Jassim, Mark Taylor, Joaquim Peiró, and Spencer Sherwin. Industry-relevant implicit large-eddy simulation of a high-performance road car via spectral/hp element methods. SIAM Review, 63 (4):723–755, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Martin:1965:NZE

[MN65] E. J. Martin, Jr. and N. S.

Nahman. On the near-zone evaluation of circular-loop field integrals. SIAM Review, 7(4):539–543, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morgan:1976:UAS

[MN76] A. Morgan and K. S. Narendra. Uniform asymptotic stability of linear nonautonomous equation X = −U(T)U(T)′X. SIAM Review, 18(4):818, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McKelvey:1983:BRU

[MN83] Robert McKelvey and Hien Nguyen. Book review: Urban Operations Research (Richard C. Larson and Amedeo R. Odoni). SIAM Review, 25(1):129–131, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meany:1971:BRI

[MNC71] R. K. Meany, S. A. Nelson, and B. C. Carlson. Book review: An Inequality of Mixed Arithmetic and Geometric Means (B. C. Carlson). SIAM Review, 13(2): 253–255, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mckinney:1963:BRS

[MNMF63] A. W. Mckinney, D. J. Newman, F. J. Macwilliams, and Henry D. Friedman. Book review: A Set o f Convolution Equations (Walter Weissblum). SIAM Review, 5 (3):276–283, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morin:2002:CAF

[MNS02] Pedro Morin, Ricardo H. Nochetto, and Kunibert G. Siebert. Convergence of adaptive finite element methods. SIAM Review, 44(4): 631–658, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40909.

Macki:1993:MMH

[MNZ93] Jack W. Macki, Paolo Nistri, and Pietro Zecca. Mathematical models for hysteresis. SIAM Review, 35(1): 94–123, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murley:1966:LS

[MO66] P. C. Murley and R. R.

O’Brien. Late solution. SIAM Review, 8(1):107, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morris:1968:CGI

[MO68] Gerald L. Morris and Patrick L.

Odell. A characterization for generalized inverses of matrices. SIAM Review, 10(2): 208–211, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Manitius:1978:FCL

[MO78] A. Manitius and A. W.

Olbrot. Feedback control of linear retarded systems via finite spectrum assignment. SIAM Review, 20 (3):630, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mock:1985:BRA

[Moc85] Michael S. Mock. Book review: Analysis and Simulation of Semiconductor Devices (Siegfried Selberherr). SIAM Review, 27(3):469– 470, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mode:1992:BRR

[Mod92] Charles J. Mode. Book review: Random Number Generators and Simulation, by Istvan Deak. SIAM Review, 34(2):338–341, June 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moeckel:2017:FPT

[Moe17] Richard Moeckel. The Foucault pendulum (with a twist). SIAM Review, 59(2): 369–389, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mohanty:1966:SCM

[Moh66]  S. G. Mohanty.  Some

convolutions with multinomial coefficients and related probability distributions. SIAM Review, 8(4): 501–509, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moin:1989:BRT

[Moi89]  Parviz Moin.  Book re-

view: Turbulence in Fluids: Stochastic and Numerical Modelling (M. Lesieur); Turbulence and Random Processes in Fluid Mechanics (M. T. Landahl and E. Mollo-Christensen). SIAM Review, 31(3):510–512, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Monson:1960:BRA

[Mon60] Philip R. Monson. Book review: Analysis of Straight Line Data (Forman S. Acton). SIAM Review, 2(1): 48–49, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Montroll:1967:ISC

[Mon67] E. W. Montroll. III. the science curriculum. SIAM Review, 9(2):326–346, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mond:1970:MCM

[Mon70] Bertram Mond. On the monotonicity of complex matrices. SIAM Review, 12(4): 577–579, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Monegato:1982:SPR

[Mon82] Giovanni Monegato. Stieltjes polynomials and related quadrature rules. SIAM Review, 24(2):137–158, April 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1960:BRN

[Moo60] Ramon E. Moore. Book review: On Numerical Approximation (Proceedings of a Symposium) (R. E. Langer, ed.). SIAM Review, 2(1):

49–50, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1963:BRN

[Moo63] R. Moore. Book review: Numerical Analysis (N. Macon). SIAM Review, 5(4): 383–384, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moody:1965:LVS

[Moo65] W. T. Moody. Limiting value of a solution of an integral equation. SIAM Review, 7 (4):562, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1969:BRA

[Moo69] R. H. Moore. Book review: Analysis of Numerical Methods (Eugene Isaacson and Herbert Bishop Keller). SIAM Review, 11(1):100– 101, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moon:1974:BRH

[Moo74a] J. W. Moon. Book review: A Handbook of Integer Sequences (N. J. A. Sloane). SIAM Review, 16 (4):572, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moon:1974:BRG

[Moo74b] J. W. Moon. Book review: Graphical Enumeration (Frank Harary and Edgar M. Palmer). SIAM Review, 16 (2):264–265, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moors:1974:CMI

[Moo74c] Michael H. Moors. Constrained minimization of an integral functional. SIAM Review, 16(4):545, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moon:1975:BRP

[Moo75] J. W. Moon. Book review: Probabilistic Methods in Combinatorics (Paul Erdős and Joel Spencer). SIAM Review, 17(4):714– 715, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1976:BRC

[Moo76] R. H. Moore. Book review:

Collectively Compact Operator Approximation Theory (Phillip M. Ansélone). SIAM Review, 18(2):311– 312, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1977:BRM

[Moo77a] M. H. Moore. Book review: A Matrix Inequality (W. Anderson, Jr. and G. Trapp). SIAM Review, 19(2): 329–331, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1977:BRL

[Moo77b] Robert T. Moore. Book review: Linear Analysis and Representation Theory (Steven A. Gaal). SIAM Review, 19(2):358–361, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1978:BSF

[Moo78] R. E. Moore. Bounding sets in function spaces with applications to nonlinear operator equations. SIAM Review, 20 (3):492–512, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moore:1985:BRI

[Moo85] R. E. Moore. Book review: Introduction to Interval Computations (Götz Alefeld and Jürgen Herzberger). SIAM Review, 27(2):296– 297, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moon:1987:I

[Moo87] John W. Moon. An identity. SIAM Review, 29(2): 298–299, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moors:1994:PC

[Moo94] R. C. Moors. Path of a car. SIAM Review, 36(2): 277, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morgan:1964:NCS

[Mor64] Samuel P. Morgan. Note on complex symmetric kernels and the Rayleigh–Ritz procedure. SIAM Review, 6 (3):265–268, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morrison:1966:CMM

[Mor66] J. A. Morrison. Comparison of the modified method of averaging and the two variable expansion procedure. SIAM Review, 8(1): 66–85, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1968:BRS

[Mor68] R. E. Mortensen. Book review: Stochastic Stability and Control (Harold J. Kushner). SIAM Review, 10(4):

460–462, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morton:1969:BSL

[Mor69] Thomas E. Morton. Bounds on the solution of the lagged optimal inventory equation with no demand backlogging and proportional costs. SIAM Review, 11(4):572– 596, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1971:BRO

[Mor71] R. E. Mortensen. Book review: Optimization of Stochastic Systems (Masanao Aoki). SIAM Review, 13(4): 579–581, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1973:BRS

[Mor73] R. E. Mortensen. Book review: Stochastic Processes in Information and Dynamical Systems (Eugene Wong). SIAM Review, 15(2):397– 400, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

More:1974:CCN

[Mor74a] Jorge J. Moré. Coercivity conditions in nonlinear complementarity problems. SIAM Review, 16(1):

1–16, January 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morgan:1974:BRS

[Mor74b] Charles G. Morgan. Book review: Symbolic Logic and Mechanical Theorem Proving (Chin-Liang Chang and Richard Char-Tung Lee). SIAM Review, 16(3):403– 407, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morrison:1975:SUF

[Mor75a] J. A. Morrison. Structure of a utility function under strong risk invariance. SIAM Review, 17(2):386, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1975:BRS

[Mor75b] Richard E. Mortensen. Book review: System Theory — a Unified State-Space Approach to Continuous and Discrete Time Systems (Louis Padulo and Michael A. Arbib). SIAM Review, 17(4): 699–703, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morrison:1976:HFA

[Mor76]  J. A. Morrison.  High-

frequency approximations for elastic surface waves propagating along cylinders of general cross section. SIAM Review, 18(4):818–819, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morrison:1978:CLA

[Mor78] J. A. Morrison. Combinatorial lemma and its application to concentrating trees of discrete-time queues. SIAM Review, 20(3):631, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morettin:1981:WSA

[Mor81] Pedro A. Morettin. Walsh spectral analysis. SIAM Review, 23(3):279–291, July 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1982:BRI

[Mor82] Richard E. Mortensen. Book review: Introduction to Control Theory, Including Optimal Control (David Burghes and Alexander Graham). SIAM Review, 24(1):87– 89, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1983:BRN

[Mor83] R. E. Mortensen. Book review: Nonlinear System Theory: The Volterra/Wiener Approach (Wilson J. Rugh).

SIAM Review, 25(3):412– 414, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morrison:1984:DI

[Mor84] J. A. Morrison. A definite integral. SIAM Review, 26 (2):271, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morley:1985:SPW

[Mor85a]  Tom Morley.  A simple

proof that the world is threedimensional. SIAM Review, 27(1):69–71, March 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [Mor86c].

Mortensen:1985:BRI

[Mor85b] Richard E. Mortensen. Book review: Identification of Continuous Dynamical Systems-The Poisson Moment Functional (PMF) Approach (Dines Chandra Saha and Ganti Prasada Rao). SIAM Review, 27(1):116–118, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morgan:1986:BRD

[Mor86a] Thomas R. Morgan. Book review: Deconvolution (Anton Ziolkowski). SIAM Review, 28(2):272–273, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morley:1986:C

[Mor86b] T. Morley. Correction. SIAM Review, 28(2):229, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morley:1986:ESP

[Mor86c] Tom Morley. Errata: “A simple proof that the world is three-dimensional” [SIAM rev. 27 (1985), no. 1, 69–71 and MR 86f:35111]. SIAM Review, 28(2):229, June 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Mor85a].

Morgan:1988:SCI

[Mor88a] Thomas L. Morgan. Six combinatorial identities. SIAM Review, 30(2):308–309, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morgenthaler:1988:BRM

[Mor88b] S. Morgenthaler. Book review: Mathematical Nonparametric Statistics (Edward B. Manoukian). SIAM Review, 30(3):524–525, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morley:1988:RWP

[Mor88c] Tom Morley. Radial wave propagation and distortionless signaling. SIAM Review, 30(1):115–118, March 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1988:BRO

[Mor88d] Richard E. Mortensen. Book review: Optimal Estimation: With An Introduction to Stochastic Control Theory (Frank L. Lewis). SIAM Review, 30(2):341– 343, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morgan:1989:CI

[Mor89a] T. L. Morgan. 6 combinatorial identities. SIAM Review, 31(2):326–328, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morgan:1989:SCI

[Mor89b] Thomas L. Morgan. Six combinatorial identities. SIAM Review, 31(2):325–328, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1989:BRS

[Mor89c] R. E. Mortensen. Book review: Stochastic Optimal Control: Theory and Application (Robert F. Stengel). SIAM Review, 31(1): 153–154, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1991:BRI

[Mor91a] Richard E. Mortensen. Book review: Infinite-Dimensional Dynamical Systems in Mechanics and Physics (Roger Temam). SIAM Review, 33 (1):156–158, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mortensen:1991:IDD

[Mor91b] Richard E. Mortensen. Infinitedimensional dynamical systems in mechanics and physics (Roger Temam). SIAM Review, 33(1):156– ??, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morgan:1992:CPG

[Mor92] Frank Morgan. Calculus, planets, and General Relativity. SIAM Review, 34(2): 295–299, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morris:1993:BRL

[Mor93] Walter Morris. Book review: The Linear Complementarily Problem (R. W. Cottle, J.-S. Pang, and R. E. Stone). SIAM Review, 35 (4):673–674, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morris:1996:BRI

[Mor96] K. A. Morris. Book review:

An Introduction to Infinite-Dimensional Linear System Theory (R. F. Curtain and H. J. Zwart). SIAM Review, 38 (3):536–537, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morlet:1997:BRW

[Mor97] Anne C. Morlet. Book review: Water Waves: Relating Modern Theory to Advanced Engineering Practice, by Matiur Rahman. SIAM Review, 39(1):156– 157, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morokoff:1998:GQR

[Mor98] William J. Morokoff. Generating quasi-random paths for stochastic processes. SIAM Review, 40(4):765–788, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31795.

Moser:1966:TQM

[Mos66a] J. Moser. On the theory of quasiperiodic motions. SIAM Review, 8(2): 145–172, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moser:1966:MFT

[Mos66b] Leo Moser. Moving furniture through a hallway. SIAM Review, 8(3):381, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moser:1986:RDT

[Mos86]  Jürgen Moser.  Recent

developments in the theory of Hamiltonian systems. SIAM Review, 28(4):459– 485, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moser:1992:CPD

[Mos92] Louise E. Moser. Calculus and probability determine the performance of a disk drive. SIAM Review, 34 (4):635–641, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moskowitz:1993:BRI

[Mos93] Ira S. Moskowitz. Book review: Information Theory (Robert B. Ash). SIAM Review, 35(4):674–675, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1997:FNM

[MOSK97] J. J. H. Miller, E. O’Riordan, G. I. Shishkin, and R. Bruce Kellogg. Fitted numerical methods for singular perturbation problems. SIAM Review, 39(3):535– ??, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Motteler:1988:BRC

[Mot88] Zane C. Motteler. Book review: Complex Variables (Stephen D. Fisher). SIAM Review, 30(2):343–344, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moursund:1967:ANM

[Mou67] David G. Moursund. Application of numerical maximization and minimization techniques to Chebyshev approximation. SIAM Review, 9(1):100–106, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McKean:1962:BRP

[MP62] Henry McKean and Jan Petersen. Book review: A Pie Problem (Larry Shepp). SIAM Review, 4(1):49–51, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moon:1970:GTM

[MP70] J. W. Moon and N. J.

Pullman. On generalized tournament matrices.

SIAM Review, 12(3):384– 399, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miele:1972:NSS

[MP72] Angelo Miele and Robert E. Pritchard. Numerical solutions in the simplest problem of the calculus of variations. SIAM Review, 14(3): 385–398, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moses:1993:PEF

[MP93] Harry E. Moses and Reese T. Prosser. Propagation of an electromagnetic field through a planar slab. SIAM Review, 35(4):610–620, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Menon:2006:DSS

[MP06] Govind Menon and Robert L. Pego. Dynamical scaling in Smoluchowski’s coagulation equations: Uniform convergence. SIAM Review, 48 (4):745–768, January 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mironchenko:2020:ISS

[MP20] Andrii Mironchenko and Christophe Prieur. Inputto-state stability of infinitedimensional systems: Recent results and open questions. SIAM Review, 62(3): 529–614, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Madhukar:2002:EMC

[MPRW02] A. Madhukar, T. Parent, I. G. Rosen, and C. Wang. An elementary model for control of a semiconductor etching process. SIAM Review, 44 (4):678–695, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38229.

Medvinsky:2002:SCP

[MPT^{+}02] Alexander B. Medvinsky, Sergei V. Petrovskii, Irene A. Tikhonova, Horst Malchow, and Bai-Lian Li. Spatiotemporal complexity of plankton and fish dynamics. SIAM Review, 44(3): 311–370, September 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40444.

Miao:2023:EIN

[MPW23] Hongyu Miao, Alan S. Perelson, and Hulin Wu. Erratum: On identifiability of nonlinear ODE models and applications in viral dynamics. SIAM Review, 65(3):732, ???? 2023.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1568958.  See

[MXPW11].

Maybe:1969:QPM

[MQ69] John Maybe and James Quirk. Qualitative problems in matrix theory. SIAM Review, 11(1):30–51, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Micchelli:1995:MAG

[MQ95] Charles A. Micchelli and Ewald Quak. Mathematical aspects of geometric modeling. SIAM Review, 37(4): 636–??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Majda:2018:SRO

[MQ18] Andrew J. Majda and Di Qi. Strategies for reduced-order models for predicting the statistical responses and uncertainty quantification in complex turbulent dynamical systems. SIAM Review, 60(3): 491–549, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mullineux:1965:DI

[MR65a] N. Mullineux and J. R. Reed. A double integral. SIAM Review, 7(3):

410–411, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murdeshwar:1965:BRP

[MR65b] M. G. Murdeshwar and V. K. Rohatgi. Book review: A Pair of Combinatorial Identities (Joel L. Brenner). SIAM Review, 7(4): 566, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mahoney:1971:BRT

[MR71] J. J. Mahoney and L. Rubenfeld. Book review: Two Singular Integrals (William E. Warren). SIAM Review, 13 (1):121–123, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mullineux:1972:LI

[MR72] N. Mullineux and J. R. Reed.

Limit of an integral. SIAM Review, 14(2):366, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mcallister:1976:IIS

[MR76] G. T. Mcallister and S. M.

Rohde. Increasing integrals of solutions of some distributed control problems. SIAM Review, 18(4):817, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1978:ONL

[MR78] W. W. Meyer and I. Rubinstein. 1st order nonlinear differential-equation. SIAM Review, 20(4):858, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Metzen:1992:FCE

[MR92] Gerhard Metzen and Cecil Rousseau. A family of cubic equations. SIAM Review, 34(3):497, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Melman:2000:EMC

[MR00] A. Melman and G. Rabinowitz. An efficient method for a class of continuous nonlinear knapsack problems. SIAM Review, 42(3): 440–448, September 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33017.

Masuda:2018:GAN

[MR18] Naoki Masuda and Luis E. C. Rocha. A Gillespie algorithm for non-Markovian stochastic processes. SIAM Review, 60 (1):95–115, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mroz:1991:BRS

[Mro91] Zenon Mroz. Book review:

Structural Design Via Optimality Criteria: The Prager Approach to Structural Optimization (George I. N. Rozvany). SIAM Review, 33(1): 158–159, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morgan:1965:BRR

[MS65] Samuel P. Morgan and W. R.

Smythe. Book review: Resistance of a Cut-Out Right Circular Cylinder (Alan L. Tritter). SIAM Review, 7(3): 411–412, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mazo:1972:ACN

[MS72] J. E. Mazo and L. A. Shepp.

Asymptotic conjecture for a nonlinear differential equation. SIAM Review, 14(3): 494, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Markham:1975:AML

[MS75] J. Markham and D. L. Snyder. Algorithm for maximumlikelihood estimation for tracer-kinetic data. SIAM Review, 17(2):385, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mallory:1976:CAS

[MS76] S. R. Mallory and D. F. Stubbs. Comparison of algorithms for solution of dense linear systems. SIAM Review, 18(4):816, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mathis:1989:CPM

[MS89] F. Mathis and S. Sivakumar. Characterizations of parabolic motion. SIAM Review, 31(1):129–131, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

MacBain:1991:SIR

[MS91] John MacBain and Bruce Secrest. Source identification in remote sensing problems. SIAM Review, 33(1): 109–113, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marchuk:1996:AEA

[MS96a] Guri I. Marchuk and Thomas P. Svobodny. Adjoint equations and analysis of complex systems. SIAM Review, 38(2): 353–??, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Melrose:1996:GST

[MS96b] Richard B. Melrose and John Sylvester. Geometric scattering theory. SIAM Review, 38 (4):702–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moiola:2014:HER

[MS14] Andrea Moiola and Euan A.

Spence. Is the Helmholtz equation really sign-indefinite? SIAM Review, 56(2):274– 312, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marek:1996:CBD

[MSO96] Milos Marek, Igor Schreiber, and Edward Ott. Chaotic behavior of deterministic dissipative systems. SIAM Review, 38(4):695–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morel:1996:VMI

[MSS96] Jean-Michel Morel, Sergio Solimini, and Jayant Shah. Variational methods in image segmentation. SIAM Review, 38(1):179–??, March 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milne-Thomson:1961:BRP

[MT61] L. M. Milne-Thomson. Book review: Photoelasticity, Principles and Methods (H. T. Jessop and F. C. Harris). SIAM Review, 3(3):

268, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milne-Thomson:1962:BRT

[MT62] L. M. Milne-Thomson. Book review: A Treatise on Hydrodynamics (A. B. Basset). SIAM Review, 4(4): 402, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Milne-Thomson:1964:BRS

[MT64] L. M. Milne-Thomson. Book review: Some Problems of the Mechanics of Deformable Media (I. I. Gol’denblat). SIAM Review, 6(3):317– 318, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mastin:1976:GCC

[MT76] C. W. Mastin and J. F. Thompson. Generation of curvilinear coordinate systems and control of coordinate line spacing. SIAM Review, 18(4):817, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mathis:1984:ELM

[MT84] Frank H. Mathis and Danny W. Turner. Estimating the length of material wrapped around a cylindrical core.

SIAM Review, 26(2):263– 266, April 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Morris:1996:VCA

[MT96] K. A. Morris and K. J. Taylor. A variational calculus approach to the modelling of flexible manipulators. SIAM Review, 38(2): 294–305, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/p294.htm.

Motsch:2014:HDE

[MT14] Sebastien Motsch and Eitan Tadmor. Heterophilious dynamics enhances consensus. SIAM Review, 56(4): 577–621, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mastin:1978:ATD

[MTGF78] C. W. Mastin, J. F. Thompson, A. Ghosh, and T. H. Fu. Application of threedimensional harmonic transformations in solution of fluid-flow problems. SIAM Review, 20(3):630, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

McNeill:1996:FLP

[MTM96] M. McNeill, E. Thro, and Marialuisa N. McAllister.

Fuzzy logic: a practical approach. SIAM Review, 38(1): 173–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mukherjee:1986:BRT

[Muk86] Subrata Mukherjee. Book review: Topics in Boundary Element Research. Volume 1, Basic Principles and Applications (C. A. Brebbia, ed.). SIAM Review, 28(3): 431–432, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mullin:1964:BRP

[Mul64a] Albert A. Mullin. Book review: Proceedings of the Symposium on Mathematical Theory of Automata, Vol. 12 (Jerome Fox, ed.). SIAM Review, 6(4):466– 467, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mullineux:1964:FSP

[Mul64b] N. Mullineux. Factorization of a set of polynomials. SIAM Review, 6(4):455, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mullin:1966:CSM

[Mul66a] A. A. Mullin. On certain systems matrices. SIAM Review, 8(4):525, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mullin:1966:BRT

[Mul66b] Albert A. Mullin. Book review: Threshold Logic: a Synthesis Approach (Michael L. Dertouzos). SIAM Review, 8(3):405–406, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mullin:1967:BRA

[Mul67] Albert A. Mullin. Book review: Applied Boolean Algebra (Franz E. Hohn). SIAM Review, 9(3):613– 614, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mullin:1969:BRA

[Mul69] Albert A. Mullin. Book review: Algebraic Structure Theory of Sequential Machines (J. Hartmanis and R. E. Stearns). SIAM Review, 11(1):101–103, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mullin:1975:AGC

[Mul75] A. A. Mullin. Applications of geometry to crystallography. SIAM Review, 17 (2):386, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Muller:1976:ASO

[Mul76] G. M. Muller. Averaging spline operators and digital filtering. SIAM Review, 18 (4):819, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Muller:1979:IBA

[Mul79] E. R. Muller. Ice breaking with an air cushion vehicle. SIAM Review, 21(1): 129–135, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Munson:1961:ODS

[Mun61] Tom Munson. One-dimensional steady-state ablation. SIAM Review, 3(3):263, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murty:1971:ACP

[Mur71] Katta G. Murty. Adjacency on convex polyhedra. SIAM Review, 13(3): 377–386, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murray:1990:MVI

[Mur90] K. S. Murray. Minimum value of an integral. SIAM Review, 32(2):302, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murray:1992:SMP

[Mur92] K. S. Murray. A set of maxima problems. SIAM Review, 34(2):308, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murray:1993:EIR

[Mur93] K. S. Murray. Ellipsoid of inertia of a regular simplex. SIAM Review, 35 (1):136, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murdock:1994:BRP

[Mur94a] James Murdock. Book review: Perturbation Methods for Engineers and Scientists (Alan W. Bush). SIAM Review, 36(1):136– 137, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murdock:1994:PME

[Mur94b] James Murdock. Perturbation methods for engineers and scientists (Alan W Bush). SIAM Review, 36(1): 136–??, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Murio:1997:BRE

[Mur97]  Diego A. Murio.  Book

review: Elliptic Marching Methods for Domain Decomposition, by Patrick J. Roache. SIAM Review, 39 (2):354–355, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mandelbrot:1968:FBM

[MV68] Benoît B. Mandelbrot and John W. Van Ness. Fractional Brownian motions, fractional noises and applications. SIAM Review, 10(4): 422–437, October 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/siread/v10/
i4/p422/s1;   http://www.
jstor.org/stable/2027184.

Moler:1978:DWC

[MV78a] C. Moler and C. Vanloan. 19 dubious ways to compute exponential of a matrix. SIAM Review, 20(4): 801–836, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moler:1978:NDW

[MV78b] Cleve Moler and Charles Van Loan. Nineteen dubious ways to compute the exponential of a matrix. SIAM Review, 20(4):801–836, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miller:1993:ERG

[MV93] Allen R. Miller and Emanuel Vegh. Exact result for the grazing angle of specular reflection from a sphere. SIAM Review, 35(3):472– 480, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moler:2003:NDW

[MV03] Cleve Moler and Charles Van Loan. Nineteen dubious ways to compute the exponential of a matrix, twenty-five years later. SIAM Review, 45(1): 3–49, March 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/41801.

Missov:2015:MIM

[MV15] Trifon I. Missov and James W.

Vaupel. Mortality implications of mortality plateaus. SIAM Review, 57(1):61– 70, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mitchell:1961:RWH

[MW61] R. H. G. Mitchell and R. F.

Wheeling. A random walk in a hyper-cube. SIAM Review, 3(3):262, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miranker:1966:FOC

[MW66] W. L. Miranker and B. Weiss. The Feynman operator calculus. SIAM Review, 8(2): 224–232, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mancini:1976:GKT

[MW76a] L. J. Mancini and D. J. Wilde. Generating Kuhn– Tucker regions in nonlinear programming with interval arithmetic. SIAM Review, 18 (4):816, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mcneill:1976:SCG

[MW76b] B. Mcneill and D. J. Wilde. Sufficient conditions for global minimum of a signomial by construction of lower bounding functions. SIAM Review, 18(4):817– 818, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Mishra:1995:AA

[MW95] B. Mishra and Franz Winkler. Algorithmic algebra. SIAM Review, 37(3):479– ??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Moreira:1997:GSR

[MW97] H. N. Moreira and Y. Q.

Wang. Global stability in an s-¿i-¿r-¿i model. SIAM Review, 39(3):496–502, ???? 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Myers:2006:SNA

[MW06] Amy N. Myers and Herbert S.

Wilf. Some new aspects of the Coupon Collector’s Problem. SIAM Review, 48(3): 549–565, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Michel:1997:QTD

[MWL97] Anthony N. Michel, Kaining Wang, and V. Lakshmikantham. Qualitative theory of dynamical systems: The role of stability preserving mappings. SIAM Review, 39(3): 540–??, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Munk:1996:OAT

[MWWC96] W. Munk, P. Worcester, C. Wunsch, and Michael D. Collins. Ocean acoustic tomography. SIAM Review, 38 (4):683–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Miao:2011:INO

[MXPW11] Hongyu Miao, Xiaohua Xia, Alan S. Perelson, and Hulin Wu. On identifiability of nonlinear ODE models and applications in viral dynamics. SIAM Review, 53(1):3–39, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sirev/
resource/1/siread/v53/

i1/p3_s1. See erratum [MPW23].

Moreira:1997:CNG

[MY97] Helmar Nunes Moreira and Wang Yuquan. Classroom note: Global stability in an S → I → R → I model. SIAM Review, 39 (3):496–502, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29587.

Ma:2008:ESA

[MYDF08] Yi Ma, Allen Y. Yang, Harm Derksen, and Robert Fossum. Estimation of subspace arrangements with applications in modeling and segmenting mixed data. SIAM Review, 50(3):413–458, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Myers:1975:STW

[Mye75] B. R. Myers. On spanning trees, weighted compositions, Fibonacci numbers, and resistor networks. SIAM Review, 17(3):465– 474, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Myers:1998:TFH

[Mye98] T. G. Myers. Thin films with high surface tension. SIAM Review, 40(3):441– 462, September 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29284.

Na:1967:TBC

[Na67]  Tsung Yen Na.  Trans-

forming boundary conditions to initial conditions for ordinary differential equations. SIAM Review, 9(2): 204–210, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Na:1968:FET

[Na68] Tsung Yen Na. Further extension on transforming from boundary value to initial value problems. SIAM Review, 10(1):85–87, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newton:2006:MCT

[NA06] Paul K. Newton and Kamran Aslam. Monte Carlo tennis. SIAM Review, 48(4): 722–742, January 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Naatz:2001:NQC

[Naa01] Michael Naatz. A note on a question of C. D. Savage. SIAM Journal on Discrete Mathematics, 14 (1):116–120, February 2001. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36919.  See

[Sav97].

Nagy:1988:BRR

[Nag88] George Nagy. Book review: Robot Vision (Berthold Klaus Paul Horn). SIAM Review, 30(1):150–152, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nagaraja:1996:BRI

[Nag96] H. N. Nagaraja. Book review: The Inverse Gaussian Distribution: a Case Study in Exponential Families (V. Seshadri). SIAM Review, 38 (2):350, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nannicini:2020:IQC

[Nan20] Giacomo Nannicini. An introduction to quantum computing, without the physics. SIAM Review, 62(4):936– 981, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Narcowich:1973:DS

[Nar73] Michael A. Narcowich. Determination of a sequence. SIAM Review, 15(1):220, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nashed:1965:BRP

[Nas65] M. Z. Nashed. Book review:

The Problem of the Minimum of a Quadratic Functional (S. G. Mikhlin). SIAM Review, 7(4):568–569, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Naasell:1977:WMB

[Nås77a] I. Nåsell. When is the modified Bessel function equal to its derivative? SIAM Review, 19(4):737, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nashed:1977:BRA

[Nas77b] M. Z. Nashed. Book review: An Algorithmic Approach to Nonlinear Analysis and Optimization (Edward J. Beltrami); The Approximate Minimization of Functionals (James W. Daniel); Approximate Methods in Optimization Problems (Vladimir F. Demyanov and Aleksandr M. Rumnov); Computational Methods in Optimization: a Unified Approach (E. Polak); Necessary Conditions for an Extremum (N. Pshenichnyi). SIAM Review, 19(2):341– 358, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nashed:1986:BRT

[Nas86] M. Z. Nashed. Book review: The Theory of Tikhonov Regularization for Fredholm Equations of the First Kind (C. W. Groetsch). SIAM Review, 28(1):116–118, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nash:1987:BRA

[Nas87a] John C. Nash. Book review: Actuarial Mathematics (Harry H. Panjer). SIAM Review, 29(4):665– 667, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nash:1987:BRR

[Nas87b] John C. Nash. Book review: Risk Theory-the Stochastic Basis of Insurance. Third Edition. (R. E. Beard, T. Pentikainen and E. Pesonen). SIAM Review, 29(1): 155–156, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nash:1990:BRI

[Nas90] John C. Nash. Book review: Introduction to Numerical Linear Algebra and Optimisation (Philippe G. Ciarlet with Bernadette Miara and Jean-Marie Thomas; A. Buttigieg, trans.). SIAM Review, 32(4):700–702, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nash:1992:RRC

[Nas92] John Nash. The Reviewer’s response to Ciarlet’s reply. SIAM Review, 34(1): 159, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nason:1995:BRM

[Nas95] Guy Nason. Book review:

Model-Free Curve Estimation (M. E. Tartar and M. D. Lock). SIAM Review, 37(2): 270–271, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nauss:1986:BRL

[Nau86] Robert M. Nauss. Book review: Linear and Nonlinear Programming — an Introduction to Linear Methods in Mathematical Programming (Roger Hartley). SIAM Review, 28(3):432– 433, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nazareth:1976:RBB

[Naz76] L. Nazareth. Relationship between BFGS algorithm and conjugate gradient algorithm. SIAM Review, 18 (4):819, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nazareth:1980:SRA

[Naz80] Larry Nazareth. Some recent approaches to solving large residual nonlinear least squares problems. SIAM Review, 22(1):1–11, January 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nazareth:1986:CGM

[Naz86]  J. L. Nazareth.  Conju-

gate gradient methods less dependent on conjugacy. SIAM Review, 28(4):501– 511, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nazareth:1994:NCP

[Naz94] J. L. Nazareth. The Newton and Cauchy perspectives on computational nonlinear optimization. SIAM Review, 36 (2):215–225, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/24004.htm.

Nayfeh:1996:AND

[NBR96] Ali H. Nayfeh, Balakumar Balachandran, and R. H. Rand. Applied nonlinear dynamics: Analytical, computational and experimental methods. SIAM Review, 38 (2):350–??, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ngo:2012:TRO

[NBS12] T. T. Ngo, M. Bellalij, and Y. Saad. The trace ratio optimization problem. SIAM Review, 54(3):545– 569, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nachman:1976:ASN

[NC76] A. Nachman and A. Callegari. Analytical solution of a nonlinear, singular boundary-value problem in theory of pseudoplastic fluids. SIAM Review, 18 (4):819, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newton:2009:VLT

[NC09] Paul K. Newton and George Chamoun. Vortex lattice theory: a particle interaction perspective. SIAM Review, 51(3):501–542, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nevai:1982:LZP

[ND82] P. G. Nevai and E. Deutsch. Location of the zeros of a polynomial. SIAM Review, 24(1):80, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neto:2014:ZAC

[NdA14] Antônio Francisco Neto and Petrus H. R. dos Anjos. Zeon algebra and combinatorial identities. SIAM Review, 56(2):353–370, ???? 2014. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nelson:1976:CEI

[NE76] P. Nelson and I. T. Elder. Calculation of eigenfunctions by integration to blow-up. SIAM Review, 18 (4):820, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neidinger:2010:IAD

[Nei10] Richard D. Neidinger. Introduction to automatic differentiation and MAT-LAB object-oriented programming. SIAM Review, 52 (3):545–563, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neidinger:2019:MPI

[Nei19] Richard D. Neidinger. Multivariate polynomial interpolation in Newton forms. SIAM Review, 61(2):361– 381, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nelson:1974:BRM

[Nel74] Paul Nelson, Jr. Book review: Mathematical Optimization and Economic Theory (Michael D. Intriligator). SIAM Review, 16(1): 107–108, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nerode:1970:BRM

[Ner70]  Anil Nerode.  Book re-

view: Mathematical Foundations of Network Analysis (Paul Slepian). SIAM Review, 12(3):476, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neto:2020:MAO

[Net20] Antônio Francisco Neto. Matrix analysis and omega calculus. SIAM Review, 62(1): 264–280, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neuberger:1994:BRS

[Neu94] J. W. Neuberger. Book review: Semigroup Theory with Applications to Systems and Control (N U. Ahmed). SIAM Review, 36(1):137–139, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neuman:1996:TIA

[Neu96]  Edward Neuman.  Three

inequalities for the arithmetic, identric, and geometric means. SIAM Review, 38 (2):315, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neumaier:1997:MMP

[Neu97a] Arnold Neumaier. Molecular modeling of proteins and mathematical prediction of protein structure. SIAM Review, 39(3):407– 460, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/27806.

Neuman:1997:BRF

[Neu97b] Frantisek Neuman. Book review: Finite Sums Decompositions in Mathematical Analysis, by Themistocles M. Rassias and Jaromir Simsa. SIAM Review, 39(1): 157–159, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neumaier:1998:SIC

[Neu98]  Arnold Neumaier.  Solv-

ing ill-conditioned and singular linear systems: a tutorial on regularization. SIAM Review, 40(3):636– 666, September 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32190.

Nevai:1982:BRL

[Nev82]  P. G. Nevai.  Book re-

view: Location of the Zeros of a Polynomial (E. Deutsch). SIAM Review, 24 (1):80, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1959:MTP

[New59] D. J. Newman. A minimum time path. SIAM Review, 1 (2):172–173, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1960:PLD

[New60] D. J. Newman. A parking lot design. SIAM Review, 2 (4):297–298, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1962:MTP

[New62a] D. J. Newman. A minimum time path. SIAM Review, 4 (2):148, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1962:LSA

[New62b] D. J. Newman. On a least square approximation. SIAM Review, 4(1):48, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1962:PA

[New62c] D. J. Newman. On a polynomial approximation. SIAM Review, 4(1):48, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1962:UA

[New62d] D. J. Newman. A uniform approximation. SIAM Review, 4(4):396, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1963:BME

[New63] D. J. Newman. On bounds for matrix eigenvalues. SIAM Review, 5(1):73, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1964:EVD

[New64a] D. J. Newman. Escape velocity with drag. SIAM Review, 6(1):60–61, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1964:MTP

[New64b] D. J. Newman. A minimum time path. SIAM Review, 6 (3):311–313, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1964:PLD

[New64c] D. J. Newman. A parking lot design. SIAM Review, 6 (1):62–66, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newell:1965:AMQ

[New65a] G. F. Newell. Approximation methods for queues with application to the fixed-cycle traffic light. SIAM Review, 7 (2):223–240, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1965:PBM

[New65b] D. J. Newman. Problem 63-2: On bounds for matrix eigenvalues. SIAM Review, 7(1):138, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Includes solutions by A. S. Householder, Marvin D. Marcus, J. H. Van Lint, and others.

Newman:1966:VGT

[New66] D. J. Newman. A variation of the game of “twenty questions”. SIAM Review, 8 (4):526, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1967:DCR

[New67a] D. J. Newman. A doublecone radar search. SIAM Review, 9(2):250, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1967:II

[New67b] D. J. Newman. An integral inequality. SIAM Review, 9 (3):590, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1968:II

[New68a] D. J. Newman. An integral inequality. SIAM Review, 10 (4):453–454, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1968:VGT

[New68b] D. J. Newman. A variation of the game of “Twenty Questions”. SIAM Review, 10(1): 113–114, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1969:II

[New69] D. J. Newman. An integral inequality. SIAM Review, 11 (3):398, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1970:US

[New70a] D. J. Newman. A unique solution. SIAM Review, 12 (4):580, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newton:1970:IPP

[New70b] R. G. Newton. Inverse problems in physics. SIAM Review, 12(3):346–356, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1971:TC

[New71] D. J. Newman. Triangle on a checkerboard. SIAM Review, 13(1):115, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1972:ENS

[New72a] D. J. Newman. Expected number of stops for an elevator. SIAM Review, 14 (4):644, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1972:TC

[New72b] D. J. Newman. Triangle on a checkerboard. SIAM Review, 14(4):647, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newton:1974:BRC

[New74] R. G. Newton. Book review: Causality and Dispersion Relations (H. M. Nussenzveig). SIAM Review, 16(4):570–571, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newell:1976:LWS

[New76a] A. C. Newell. Long waves short waves — solvable model. SIAM Review, 18 (4):820, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newhall:1976:CNR

[New76b]  X. X. Newhall.  Con-

cise numerical representation of planetary ephemerides. SIAM Review, 18(4):820,

????  1976.  CODEN

SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newhouse:1976:BRS

[New76c] A. Newhouse. Book review: The Spectral Radius of A Matrix (D. A. Voss). SIAM Review, 18(2):304– 306, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1976:AC

[New76d] D. J. Newman. An arithmetic conjecture. SIAM Review, 18 (1):118, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newton:1976:BRT

[New76e] Roger G. Newton. Book review: The Theory of Relativity (R. K. Pathria). SIAM Review, 18(3):526– 527, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newton:1976:CPP

[New76f] T. A. Newton. Classification procedure for phase portrait of two-dimensional homogeneous quadratic differential system. SIAM Review, 18(4):820, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1977:CTP

[New77] Danny Newman. A coin tossing problem. SIAM Review, 19(3):563, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newton:1978:TDH

[New78]  Tyre A. Newton.  Two

dimensional homogeneous quadratic differential systems. SIAM Review, 20(1): 120–138, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1981:LS

[New81] D. J. Newman. Limit of a sum. SIAM Review, 23 (2):255, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1982:BRL

[New82a] D. J. Newman. Book review: Limit of a Sum (D. J. Newman). SIAM Review, 24(2): 226–228, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1982:CP

[New82b] D. J. Newman. A covering problem. SIAM Review, 24(1):77, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1984:RWN

[New84] David Newman. In a random walk the number of “unique experiences” is two on the average. SIAM Review, 26(4): 573–574, October 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newell:1987:BRA

[New87] Gordon F. Newell. Book review: Author’s Comments on Daniel P. Heyman’s Review of: The M/M/Infinity Service System with Ranked Servers in Heavy Traffic (G. F. Newell). SIAM Review, 29 (3):507–508, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newell:1990:BRA

[New90] G. F. Newell. Book review: Asymptotic Behavior of a Melting Interface (O. G. Ruehr). SIAM Review, 32(1): 145–148, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newberger:1991:II

[New91a] Barry Newberger. An infinite integral. SIAM Review, 33(3):473, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1991:AAG

[New91b] D. J. Newman. An affirmative action graph. SIAM Review, 33(4):655, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1992:ETT

[New92] D. J. Newman. Expected type of triangle. SIAM Review, 34(1):120, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newell:1994:BRI

[New94a] G. F. Newell. Book review: An Infinite Sum of Erlang Loss Functions (Robert B. Cooper). SIAM Review, 36 (1):112–114, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1994:ETT

[New94b] D. J. Newman. Expected type of triangle (vol 35, pg 139, 1993). SIAM Review, 36(1):115, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:2003:SFC

[New03] M. E. J. Newman. The structure and function of complex networks. SIAM Review, 45(2):167–256, June 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42480.

Nakatsukasa:2016:CFM

[NF16] Yuji Nakatsukasa and Roland W.

Freund. Computing fundamental matrix decompositions accurately via the matrix sign function in two iterations: The power of Zolotarev’s functions. SIAM Review, 58(3):461–493, ???? 2016. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ng:1994:SCO

[Ng94] Kevin Y. K. Ng. Symbolic computations in operations research. SIAM Review, 36 (4):642–648, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/26153.
htm.

Ng:1996:SUS

[Ng96] Kam Chuen Ng. Sequences with unique sums and differences. SIAM Review, 38 (1):147, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ning:2015:LMB

[NGTB15] Lipeng Ning, Tryphon T.

Georgiou, Allen Tannenbaum, and Stephen P. Boyd. Linear models based on noisy data and the Frisch scheme.

SIAM Review, 57(2):167– 197, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nguyenmanhhung:1976:OGI

[Ngu76] Nguyenmanhhung. Optimal growth, investment decision and depletion of a replenishable resource. SIAM Review, 18(4):809, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neves:1978:DSS

[NH78] K. W. Neves and F. A. Hale. Design of software for solving delay differential equations. SIAM Review, 20 (3):631, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nicholson:1965:BRD

[Nic65] W. L. Nicholson. Book review: On a Double Summation (Perry Scheinok). SIAM Review, 7(2):294– 295, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nicolas:1994:BRE

[Nic94] Jean-Louis Nicolas. Book review: An Extremal Problem (Paul Erdős). SIAM Review, 36(3):493–495, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nicaise:1996:BRM

[Nic96] Serge Nicaise. Book review: Modeling, Analysis and Control of Dynamic Elastic Multi-Link Structures (J. E. Lagnese, G. Leugeuring, and E. J. P. G. Schmidt. Birkhäuser). SIAM Review, 38(3):537–538, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nievergelt:1966:CA

[Nie66] J. Nievergelt. On a control automaton. SIAM Review, 8 (3):382–383, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nielson:1976:ICD

[Nie76] G. M. Nielson. Interactive curve design using V - splines. SIAM Review, 18(4): 820–821, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nievergelt:1983:CEE

[Nie83] Yves Nievergelt. The concept of elasticity in economics. SIAM Review, 25(2): 261–265, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nievergelt:1984:NRO

[Nie84] Yves Nievergelt. Navigation on Riemannian oceans. SIAM Review, 26(3):423– 427, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nievergelt:1986:EIR

[Nie86] Yves Nievergelt. Elementary inversion of Radon’s transform. SIAM Review, 28(1): 79–84, March 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nielsen:1989:BRS

[Nie89a] Hans B. Nielsen. Book review: Sensitivity Analysis in Linear Systems (Assem Deif). SIAM Review, 31(4): 706–707, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nievergelt:1989:MHM

[Nie89b] Yves Nievergelt. The mathematics of a home-made parabolic reflector. SIAM Review, 31(3):490–492, September 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nield:1994:OEF

[Nie94a] D. A. Nield. Odd-even factorization results for eigenvalue problems. SIAM Review, 36 (4):649–651, December 1994.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nievergelt:1994:TLS

[Nie94b] Yves Nievergelt. Total least squares: State-of-the-art regression in numerical analysis. SIAM Review, 36(2): 258–264, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25606.htm.

Nievergelt:1996:IPH

[Nie96]  Yves Nievergelt.  Inter-

sections of planes and helices, or lines and sinusoids. SIAM Review, 38(1): 136–145, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/28040.htm.

Nievergelt:1998:SIP

[Nie98] Yves Nievergelt. Solution to an inverse problem in diffusion. SIAM Review, 40(1):74–80, March 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30757.

Nikolai:1978:USF

[Nik78] P. J. Nikolai. USA Standard Fortran implementation of simultaneous iteration for real diagonizable matrices. SIAM Review, 20(3):

631–632, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nitecki:1985:BRI

[Nit85]  Zbigniew Nitecki.  Book

review: Iterates of Maps on an Interval (Chris Preston). SIAM Review, 27(3): 470–471, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nitecki:1990:BRN

[Nit90]  Zbigniew Nitecki.  Book

review: Non-Linear Oscillations. Second edition (Peter Hagedorn; Wolfram Stadler, trans. and ed.). SIAM Review, 32(3):516– 517, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nord:1998:GPS

[NJN98a] G. Nord, D. Jabon, and J. Nord. The global positioning system and the implicit function theorem. SIAM Review, 40(3):692– 696, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nord:1998:CNG

[NJN98b] Gail Nord, David Jabon, and John Nord. Classroom note: The global positioning system and the implicit function theorem. SIAM Review, 40 (3):693–697, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32734.

Neuringer:1967:EP

[NN67] J. Neuringer and D. J. Newman. An extremal problem. SIAM Review, 9(4): 748, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nguyen:1996:OIC

[NN96] B. U. Nguyen and D. Nguyen.

Optimal intercept course of vessels to a nonzero range. SIAM Review, 38(4):647– 649, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28650.

Nestor:1960:NMR

[NO60] O. H. Nestor and H. N. Olsen.

Numerical methods for reducing line and surface probe data. SIAM Review, 2(3): 200–207, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Narayana:1979:CI

[NÖ79] T. V. Narayana and M. Özsoyoglu.

A combinatorial identity. SIAM Review, 21(3):396,

????  1979.  CODEN

SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Noether:1972:BRN

[Noe72] Gottfried E. Noether. Book review: Nonparametric Statistical Inference (Jean Dickinson Gibbons). SIAM Review, 14(2):346–348, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Noether:1981:BRI

[Noe81] Gottfried E. Noether. Book review: Introduction to the Theory of Nonparametric Statistics (Ronald H. Randles and Douglas A. Wolfs). SIAM Review, 23(2):260– 261, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Noether:1982:BRN

[Noe82] Gottfried E. Noether. Book review: Nonparametric Sequential Selection Procedures (H. Büringer, H. Martin, and K.-H. Schriever). SIAM Review, 24(1):95–96, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nohel:1962:BRC

[Noh62]  J. A. Nohel.  Book re-

view: Contributions to the Theory of Nonlinear Oscillations, Vol. 5 (L. Cesari, J. LaSalle, and S. Lefschetz, eds.). SIAM Review, 4(3): 264–266, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nohel:1967:BRN

[Noh67] John A. Nohel. Book review: Nonlinear Differential Equations (G. Sansone and R. Conti). SIAM Review, 9 (2):258–260, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nohel:1971:BRN

[Noh71] John A. Nohel. Book review: Nonlocal Problems of the Theory of Oscillations (V. A. Pliss). SIAM Review, 13(1):100–102, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nohel:1976:BRI

[Noh76] John A. Nohel. Book review:

Integral Equations and Stability of Feedback Systems (C. Corduneanu). SIAM Review, 18(3):520–526, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nohel:1977:BRE

[Noh77] John A. Nohel. Book review:

Equations with Transformed Argument (D. Przeworska-Rolewicz). SIAM Review, 19 (1):165–166, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Norman:1974:MLP

[Nor74] M. Frank Norman. Markovian learning processes. SIAM Review, 16(2):143– 162, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Norris:1989:LI

[Nor89] Andrew N. Norris. A Lagrangian inequality. SIAM Review, 31(2):318–319, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Norris:1990:BRA

[Nor90] Andrew N. Norris. Book review: Acoustics of Solids (A. I. Seltzer). SIAM Review, 32 (1):180, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

North:1996:I

[Nor96]  Roy North.  An inte-

gral. SIAM Review, 38 (3):515, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Notz:1994:BRO

[Not94] William I. Notz. Book review: Optimum Experimental Designs (A. C. Atkinson and A. N. Donev). SIAM Review, 36(2):315–316, June 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Novak:1994:BRI

[Nov94] Erich Novak. Book review:

Information-Based Complexity (by Joseph F. Traub, G. W. Wasilkowski, and H. Wozniakowski). SIAM Review, 36 (3):514–515, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nemes:1995:BRI

[NPW95] István Nemes, Peter Paule, and James A. Wilson. Book review: An Infinite Sum from a Diffusion Problem (A. W. Overhauser and Y J. Kim). SIAM Review, 37(1): 105–107, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newell:1990:ABM

[NR90] G. F. Newell and O. G.

Ruehr. Asymptotic behavior of a melting interface. SIAM Review, 32(1): 145–148, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nijholt:2019:CMC

[NRS19] Eddie Nijholt, Bob Rink, and Jan Sanders. Center manifolds of coupled cell networks. SIAM Review, 61(1): 121–155, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1963:CTP

[NS63] D. J. Newman and L. A. Shepp. On the coefficients of a trigonometric polynomial. SIAM Review, 5(4): 371, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1966:LP

[NS66] D. J. Newman and L. A. Shepp. The limit of a product. SIAM Review, 8(1): 106, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Norris:1975:LSI

[NS75] D. O. Norris and L. E. Snyder. Least-squares identification of parameters of a linearsystem. SIAM Review, 17(2): 386–387, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Neuman:1977:ESC

[NS77] C. P. Neuman and D. I. Schonbach. Evaluation of sums of convolved powers using Bernoulli numbers. SIAM Review, 19(1): 90–99, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newbury:2002:IGL

[NS02] Peter R. Newbury and Raymond J. Spiteri. Inverting gravitational lenses. SIAM Review, 44(1):111– 130, March 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/38093.

Nielsen:2024:SFG

[NS24] Bjørn Fredrik Nielsen and Zdeněk Strakoš. A simple formula for the generalized spectrum of second order self-adjoint differential operators. SIAM Review, 66(1): 125–146, February 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nanya:1978:UMS

[NT78] T. Nanya and Y. Tohma.

Universal multicode STT state assignments for asynchronous sequential machines. SIAM Review, 20 (3):631, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nunez:1996:DAM

[Núñ96] Manuel Núñez. The decay of axisymmetric magnetic fields: a review of Cowling’s theorem. SIAM Review, 38 (4):553–564, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26157.

Nuttall:1985:CDI

[Nut85] A. H. Nuttall. A conjectured definite integral. SIAM Review, 27(4):573, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1961:CTP

[NW61] D. J. Newman and Walter Weissblum. A coin tossing problem. SIAM Review, 3 (4):329, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1967:P

[NW67a] D. J. Newman and W. Weissblum. “up-down” permutations. SIAM Review, 9 (1):121, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1967:ECR

[NW67b] D. J. Newman and W. E. Weissblum. Expectations in certain reliability problems. SIAM Review, 9(4): 744–747, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1970:LPD

[NW70] D. J. Newman and W. Weissblum. Limit of a probability density function. SIAM Review, 12(2):288, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Newman:1981:BRC

[NW81] D. J. Newman and D. V.

Widder. Book review: A Conjectured Limit (H. P. Robinson). SIAM Review, 23 (2):256–257, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Nguyen:1982:RUL

[NW82] V. V. Nguyen and E. F.

Wood. Review and unification of linear identifiability concepts. SIAM Review, 24 (1):34–51, January 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

NunesMoreira:1997:CNG

[NW97] Helmar Nunes Moreira and Yuquan Wang. Classroom note: Global stability in an S ⟶ I ⟶ R ⟶ I model. SIAM Review, 39 (3):496–502, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29587.

Nicolaides:1997:MCI

[NWF97] Roy Nicolaides, Noel Walkington, and Patrick Fitzpatrick. Maple: a comprehensive introduction. SIAM Review, 39(4):792–??, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Norrts:1990:CIA

[NY90] A. N. Norrts and Y. Yang. Conjectured identities from an acoustic scattering problem. SIAM Review, 32(3): 477, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OBrien:1995:BRL

[O’B95] George L. O’Brien. Book review: Laws of Small Numbers: Extremes and Rare Events (Michael Falk, Jurg Husler, and Rolf-Dieter Reiss). SIAM Review, 37(4): 634–635, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OBrien:2011:LSS

[O’B11]  S. B. G. O’Brien.  Lin

& Segel’s standing gradient problem revisited: a lesson in mathematical modeling and asymptotics. SIAM Review, 53(4):775–796, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sirev/
resource/1/siread/v53/
i4/p775_s1.

Oberhettinger:1964:BRC

[Obe64a] F. Oberhettinger. Book review: Complex Variables and Applications, Second Edition (Ruel V. Churchill). SIAM Review, 6(3):321–322, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Oberhettinger:1964:BRF

[Obe64b] F. Oberhettinger. Book review: Fourier Series and Boundary Value Problems (Ruel V. Churchill). SIAM Review, 6(3):322, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Oberhettinger:1977:BRC

[Obe77] F. Oberhettinger. Book review: Continuum Mechanics of Viscoelastic Liquids (R. R. Huilgol). SIAM Review, 19 (4):747–749, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ockendon:1991:BRN

[Ock91] J. R. Ockendon. Book review: Nonequilibrium Condensation in High-Speed Gas Flows (Yuri A. Ryzhov, Ul’yan G. Pirumov, and Vladimir N. Gorbunov; V. A. Khokhrayakov, trans.).

SIAM Review, 33(1):159– 160, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ockendon:1992:BRM

[Ock92a] J. R. Ockendon. Book review: Mathematics in Industrial Problems:Part 4 (Avner Friedman). SIAM Review, 34 (4):679–680, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ockendon:1992:MIP

[Ock92b] J. R. Ockendon. Mathematics in industrial problems: Part 4 (Avner Friedman). SIAM Review, 34(4): 679–??, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ocone:1987:BRC

[Oco87] Daniel Ocone. Book review: Controlled Markov Processes: Time-Discretization (N. M. Van Dijk). SIAM Review, 29 (2):333–334, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Odeh:1982:OMM

[Ode82] Aziz S. Odeh. An overview of mathematical modeling of the behavior of hydrocarbon reservoirs. SIAM Review, 24 (3):263–273, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Oden:1985:EEL

[Ode85] N. Oden. Eigenvalues and eigenvectors of a large matrix. SIAM Review, 27(2): 250, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Oden:1986:BRG

[Ode86] Tinsley J. Oden. Book review: Galerkin Finite Element Methods for Parabolic Problems (Vidar Thomée). SIAM Review, 28(2):273– 275, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Oden:1989:BRI

[Ode89] J. Tinsley Oden. Book review: An Introduction to the Finite Element Method with Applications to Nonlinear Problems (R. E. White). SIAM Review, 31(3):512, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Odgers:1986:BRC

[Odg86] Christopher R. Odgers. Book review: Computer Animation: Theory and Practice (Nadia Magnenat-Thalmann and Daniel Thalmann). SIAM Review, 28(4):600– 601, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1990:BRC

[OG90] Norbert Ortner and M. L. Glasser. Book review: A Complement of Weber’s Integral (J. W. Miles). SIAM Review, 32(4):683–684, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1990:CWI

[OGGR90] N. Ortner, M. L. Glasser, C. C. Grosjean, and M. Renardy. A complement of Weber integral. SIAM Review, 32(4):683–684, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olmstead:1976:DSI

[OH76] W. E. Olmstead and R. A. Handelsman. Diffusion in a semi-infinite region with nonlinear surface dissipation. SIAM Review, 18(2): 275–291, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

ORourke:1995:CGC

[OH95] J. O’Rourke and Christoph M. Hoffmann. Computational geometry in C. SIAM Review, 37(1):128–??, March 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Orum:2000:EHA

[OHBG00] J. C. Orum, R. T. Hudspeth, W. Black, and R. B. Guenther. Extension of the Herglotz algorithm to nonautonomous canonical transformations. SIAM Review, 42(1):83–90, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34762.

Overhauser:1994:ISD

[OK94] A. W. Overhauser and Y. J.

Kim. An infinite sum from a diffusion problem. SIAM Review, 36(1):107, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1998:TAO

[OL98] Robert E. O’Malley, Jr. and J. David Logan. Thinking about ordinary differential equations. SIAM Review, 40 (1):163–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Oldenburger:1962:BRA

[Old62] Rufus Oldenburger. Book review: Analysis and Design of Nonlinear Feedback Control Systems (G. J. Thaler and M. P. Pastel). SIAM Review, 4 (4):405, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Oleszkiewicz:1968:BRP

[Ole68]  M. Oleszkiewicz.  Book

review: On a Probability of Overlap (Murray S. Klamkin). SIAM Review, 10 (1):112, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olek:1994:ASM

[Ole94] Shmuel Olek. An accurate solution to the multispecies Lotka–Volterra equations. SIAM Review, 36 (3):480–488, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

                                        //epubs.siam.org/24342.
                                        htm.

Oliker:2002:EOS

[OLHB02] Leonid Oliker, Xiaoye Li, Parry Husbands, and Rupak Biswas. Effects of ordering strategies and programming paradigms on sparse matrix computations. SIAM Review, 44(3):373– 393, September 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/38207.

Olkin:1959:BRS

[Olk59] Ingram Olkin. Book review: Some Aspects Of Multivariate Analysis (S. N. Roy). SIAM Review, 1 (2):184, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olkin:1972:BRD

[Olk72a] I. Olkin. Book review: A Determinant Identity (K. W. Schmidt). SIAM Review, 14 (3):499–500, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olkin:1972:BRI

[Olk72b] Ingram Olkin. Book review: An Infinite Integral (Hedayat Yasaimaibodi). SIAM Review, 14(4):654– 656, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olkin:1977:BRM

[Olk77] I. Olkin. Book review: A Monotonicity Property for Moments (A. A. Jagers). SIAM Review, 19(4):741– 742, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olkin:1978:BRV

[Olk78] Ingram Olkin. Book review: Variational Methods in Statistics (Jagdish S. Rustagi). SIAM Review, 20 (1):201–202, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olmstead:1976:ELS

[Olm76] W. E. Olmstead. Extent of leftward solution branch in bifurcation problems. SIAM Review, 18(4):821, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olmstead:1977:BRS

[Olm77] W. E. Olmstead. Book review: Similarity Methods for Differential Equations (G. W. Bluman and J. D. Cole). SIAM Review, 19(1): 164–165, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olsder:2003:BRM

[Ols03] Geert Jan Olsder. Bicycle routing for maximum suntan. SIAM Review, 45(2): 345–358, June 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/39612.

Olver:1970:WSD

[Olv70] F. W. J. Olver. Why steepest descents? SIAM Review, 12 (2):228–247, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1972:BRS

[Olv72] Frank W. J. Olver. Book review: The Special Functions and Their Approximations (Vols. I & II Yudell L. Luke). SIAM Review, 14(1): 181–183, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1978:NAE

[Olv78]  F. W. J. Olver.  New

approach to error arithmetic. SIAM Review, 20 (3):632, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1980:AAE

[Olv80] F. W. J. Olver. Asymptotic approximations and error bounds. SIAM Review, 22 (2):188–203, April 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1986:BRL

[Olv86] Frank W. J. Olver. Book review: Linear Turning Point Theory (Wolfgang Wasow). SIAM Review, 28(3):433– 435, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1990:BRS

[Olv90] Peter J. Olver. Book review: Symmetries and Differential Equations (G. W. Bluman and S. Kumei). SIAM Review, 32(3):517– 519, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1991:BRD

[Olv91a] Peter J. Olver. Book review: Differential Equations: Their Solution Using Symmetries (Hans Stephani) (Malcolm MacCallum, ed.). SIAM Review, 33(2):330– 332, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1991:DET

[Olv91b] Peter J. Olver. Differential equations: Their solution using symmetries (Hans Stephani). SIAM Review, 33 (2):330–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1993:BRG

[Olv93] Peter J. Olver. Book review: Geometrical Aspects of Partial Differential Equations (V. V. Zharinov). SIAM Review, 35(4):675– 676, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1997:BRS

[Olv97] Frank W. J. Olver. Book review: Special Functions: An Introduction to the Classical Functions of Mathematical Physics, by Nico M.

Temme. SIAM Review, 39 (2):355–356, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1971:BLM

[O’M71] R. E. O’Malley. Boundary layer methods for nonlinear initial value problems. SIAM Review, 13(4): 425–434, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1974:BRM

[O’M74] R. E. O’Malley, Jr. Book review: Matched Asymptotic Expansions and Singular Perturbations (Wiktor Eckhaus). SIAM Review, 16 (4):564–565, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1975:TMS

[O’M75] R. E. O’Malley, Jr. On two methods of solution for a singularly perturbed linear state regulator problem. SIAM Review, 17(1):16–37, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1986:BRN

[O’M86] R. E. O’Malley, Jr. Book review: Nonlinear Singular Perturbation Phenomena: Theory and Application (K. W. Chang and F.

A. Howes). SIAM Review, 28 (1):118–120, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1987:BRF

[O’M87a] R. E. O’Malley, Jr. Book review: A First Look at Perturbation Theory (James G. Simmonds and James E. Mann, Jr.). SIAM Review, 29(4):667, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1987:BRA

[O’M87b] R. E. O’Malley, Jr. Book review: Averaging Methods in Nonlinear Dynamical Systems. (J. A. Sanders and F. Verhulst). SIAM Review, 29(1):157–158, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1988:NSP

[O’M88a] R. E. O’Malley, Jr. On nonlinear singularly perturbed initial value problems. SIAM Review, 30(2): 193–212, June 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1988:BRA

[O’M88b] Robert E. O’Malley, Jr. Book review: Asymptotic Methods for Relaxation Oscillations and Applications (Johan Grasman). SIAM Review, 30 (2):344–345, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1992:BRS

[O’M92] R. E. O’Malley, Jr. Book review: Singular Perturbations I. Spaces and Singular Perturbations on Manifolds without Boundaries (L. S. Frank). SIAM Review, 34 (1):142, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:1994:SPE

[O’M94] Robert E. O’Malley, Jr. On singular perturbations, especially matching. SIAM Review, 36(3):413–414, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
25788.htm.

OMalley:1996:BRI

[O’M96] Robert E. O’Malley, Jr. Book review: Introduction to Perturbation Methods (M. H. Holmes). SIAM Review, 38 (3):538, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2000:BRa

[O’M00a] Bob O’Malley. Book reviews. SIAM Review, 42(2):

333–365, June 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97049.

OMalley:2000:BRb

[O’M00b] Bob O’Malley. Book reviews. SIAM Review, 42 (3):513–552, September 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97055.

OMalley:2000:BRc

[O’M00c] Bob O’Malley. Book reviews. SIAM Review, 42 (4):737–768, December 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97060.

OMalley:2001:BRa

[O’M01a] Bob O’Malley. Book reviews. SIAM Review, 43 (1):179–232, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97065.

OMalley:2001:BRb

[O’M01b] Bob O’Malley. Book reviews. SIAM Review, 43(2): 369–409, June 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97070.

OMalley:2001:BRc

[O’M01c] Bob O’Malley. Book reviews. SIAM Review, 43 (3):547–581, September 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97075.

OMalley:2001:BRd

[O’M01d] Bob O’Malley. Book reviews. SIAM Review, 43(4): 707, December 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97080.

OMalley:2002:BRa

[O’M02a] Bob O’Malley. Book reviews. SIAM Review, 44 (1):131–165, March 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97085.

OMalley:2002:BRb

[O’M02b] Bob O’Malley. Book reviews. SIAM Review, 44 (2):279, June 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97090.

OMalley:2002:BRc

[O’M02c] Bob O’Malley. Book reviews. SIAM Review, 44 (3):477–522, September 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97095.

OMalley:2002:BRd

[O’M02d] Bob O’Malley. Book reviews. SIAM Review, 44 (4):697–747, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97100.

OMalley:2003:BRa

[O’M03a] Bob O’Malley. Book reviews. SIAM Review, 45 (1):125–164, March 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97105.

OMalley:2003:BRb

[O’M03b] Bob O’Malley. Book reviews. SIAM Review, 45(2): 359–381, June 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97110.

OMalley:2003:BRc

[O’M03c] Bob O’Malley. Book reviews. SIAM Review, 45 (3):591–624, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97115.

OMalley:2003:BRd

[O’M03d] Bob O’Malley. Book reviews. SIAM Review, 45 (4):815–845, December 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97120.

OMalley:2004:BRa

[O’M04a] Bob O’Malley. Book reviews. SIAM Review, 46 (1):143–185, March 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97125.

OMalley:2004:BRb

[O’M04b] Bob O’Malley. Book reviews. SIAM Review, 46(2): 347–374, June 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97129.

OMalley:2004:BRc

[O’M04c] Bob O’Malley. Book reviews. SIAM Review, 46 (3):547–595, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97134.

OMalley:2004:BRd

[O’M04d] Bob O’Malley. Book reviews. SIAM Review, 46 (4):740–771, December 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97139.

OMalley:2005:BRa

[O’M05a] Bob O’Malley. Book reviews. SIAM Review, 47 (1):163–193, March 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97144.

OMalley:2005:BRb

[O’M05b] Bob O’Malley. Book reviews. SIAM Review, 47 (3):577–626, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97154.

OMalley:2005:BRc

[O’M05c] Bob O’Malley. Book reviews. SIAM Review, 47 (4):799–858, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97159.

OMalley:2005:BRR

[O’M05d] Bob O’Malley. Book reviews: Recent books on the mathematics of finance. SIAM Review, 47(2):367– 403, June 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97149.

OMalley:2006:BRa

[O’M06a] Bob O’Malley. Book reviews. SIAM Review, 48(1): 149–204, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2006:BRb

[O’M06b] Bob O’Malley. Book reviews. SIAM Review, 48(2): 393–436, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2006:BRc

[O’M06c] Bob O’Malley. Book reviews. SIAM Review, 48(3): 585–626, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2006:BRd

[O’M06d] Bob O’Malley. Book reviews. SIAM Review, 48(4): 785–821, January 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2007:BRa

[O’M07a] Bob O’Malley. Book reviews. SIAM Review, 49(1): 123–176, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2007:BRb

[O’M07b] Bob O’Malley. Book reviews. SIAM Review, 49(2): 327–367, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2007:I

[O’M07c] Bob O’Malley. Introduction. SIAM Review, 49(4): 693, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2007:BRc

[O’M07d] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 49 (3):511–542, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2007:BRd

[O’M07e] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 49 (4):695–732, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2008:BRa

[O’M08a] Bob O’Malley. Book reviews. SIAM Review, 50(1): 149–195, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2008:BRb

[O’M08b] Bob O’Malley. Book reviews. SIAM Review, 50(2): 371–410, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2008:BRc

[O’M08c] Bob O’Malley. Book reviews. SIAM Review, 50(3): 587–626, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2008:BRd

[O’M08d] Bob O’Malley. Book reviews. SIAM Review, 50(4): 807–819, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2008:SPL

[O’M08e] Robert E. O’Malley, Jr. Singularly perturbed linear twopoint boundary value problems. SIAM Review, 50(3): 459–482, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2009:BRa

[O’M09a] Bob O’Malley. Book reviews. SIAM Review, 51(1): 215–236, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2009:BRb

[O’M09b] Bob O’Malley. Book reviews. SIAM Review, 51(2): 435–452, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2009:BRc

[O’M09c] Bob O’Malley. Book reviews. SIAM Review, 51(3): 639–658, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2009:BRd

[O’M09d] Bob O’Malley. Book reviews. SIAM Review, 51(4): 785–807, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2010:BRa

[O’M10a] Bob O’Malley. Book reviews. SIAM Review, 52(1): 173–217, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2010:BRb

[O’M10b] Bob O’Malley. Book reviews. SIAM Review, 52(2): 381–400, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2010:BRc

[O’M10c] Bob O’Malley. Book reviews. SIAM Review, 52(3): 567–590, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2010:BRd

[O’M10d] Bob O’Malley. Book reviews. SIAM Review, 52(4): 771–791, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2010:LNA

[O’M10e] Robert E. O’Malley, Jr. Lecture notes on applied analysis. SIAM Review, 52(4): 782, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2010:ODE

[O’M10f] Robert E. O’Malley, Jr. Ordinary differential equations: Applications, models, and computing. SIAM Review, 52 (4):781–782, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2011:BRa

[O’M11a] Bob O’Malley. Book reviews. SIAM Review, 53(1): 177–213, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i1/p177_s1.

OMalley:2011:BRb

[O’M11b] Bob O’Malley. Book reviews. SIAM Review, 53(2): 375–405, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i2/p375_s1.

OMalley:2011:BRc

[O’M11c] Bob O’Malley. Book reviews. SIAM Review, 53(3): 577–603, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i3/p577_s1.

OMalley:2011:BRd

[O’M11d] Bob O’Malley. Book reviews. SIAM Review, 53(4): 799–823, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i4/p799_s1.

OMalley:2012:BRa

[O’M12a] Bob O’Malley. Book reviews. SIAM Review, 54(1):

181–208, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v54/i1/p181_s1.

OMalley:2012:BRb

[O’M12b] Bob O’Malley. Book reviews. SIAM Review, 54(2): 391–417, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2012:BRc

[O’M12c] Bob O’Malley. Book reviews. SIAM Review, 54(3): 609–623, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2012:BRd

[O’M12d] Bob O’Malley. Book reviews. SIAM Review, 54(4): 827–847, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2013:BRa

[O’M13a] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 55 (1):187–202, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2013:BRb

[O’M13b] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 55 (2):399–412, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2013:BRc

[O’M13c] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 55 (3):569–597, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2013:BRd

[O’M13d] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 55 (4):791–809, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2014:BRa

[O’M14a] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 56 (1):189–209, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2014:BRb

[O’M14b] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 56 (2):373–382, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2014:BRc

[O’M14c] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 56 (3):549–574, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

OMalley:2014:BRd

[O’M14d] Robert E. O’Malley, Jr. Book reviews. SIAM Review, 56 (4):711–726, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Onofri:1984:NVP

[Ono84] E. Onofri. A nonlinear variational problem. SIAM Review, 26(4):579, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Onofri:1985:NVP

[Ono85] E. Onofri. A nonlinear variational problem. SIAM Review, 27(4):576–578, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ockendon:2017:HMS

[OO17] H. Ockendon and J. R. Ockendon. How to mitigate sloshing. SIAM Review, 59 (4):905–911, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Orlowski:1988:BRM

[OP88] M. Orlowski and M. Pachter.

Book review: Maximal Stability (David Singmaster). SIAM Review, 30(2):311– 314, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Overhauser:1967:BRM

[OR67] A. W. Overhauser and Paul I. Richards. Book review: The Minimum of a Double Integral (A. W. Overhauser). SIAM Review, 9(1): 127–129, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

ORourke:1992:BRM

[O’R92] Joseph O’Rourke. Book review: Mathematics in Action (Stan Wagon). SIAM Review, 34(3):515–516, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

ORourke:2010:GPC

[O’R10] Joseph O’Rourke. Games, puzzles, and computation. SIAM Review, 52(4):782– 785, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Owen:2021:SLL

[OR21] Art B. Owen and Daniel Rudolf. A strong law of large numbers for scrambled net integration. SIAM Review, 63 (2):360–372, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Orban:2018:BRI

[Orb18] Dominique Orban. Book review: Introduction to Computation and Programming Using Python. Second Edition, with Application to Understanding Data, by John V. Guttag. SIAM Review, 60(2): 483–485, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Orkwis:1995:BRN

[Ork95] Paul D. Orkwis. Book review: Numerical Solutions of the Euler Equations for Steady Flow Problems (Albrecht Eberle, Arthur Rizzi, and Ernst Heinrich Hirschel). SIAM Review, 37(2):271– 272, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Orlik:1993:BRM

[Orl93a] Peter Orlik. Book review:

Matroid Applications (Neil White, ed.). SIAM Review, 35(2):331, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Orlik:1993:BRV

[Orl93b]  Peter Orlik.  Book re-

view: Verification of Sequential and Concurrent Programs (Krzysztof R. Apt and Ernst-Rudiger Olderog). SIAM Review, 35(2):330– 331, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortega:1974:BRS

[Ort74] James M. Ortega. Book review: Solution of Equations in Euclidean and Banach Spaces (A. M. Ostrowski). SIAM Review, 16 (4):564, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1994:DI

[Ort94] Norbert Ortner. A definite integral. SIAM Review, 36 (3):490, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1995:DI

[Ort95] N. Ortner. A definite integral. SIAM Review, 37(3): 450–451, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1996:ASP

[Ort96a] N. Ortner. Analytic study of the post attenuation correction algorithm. SIAM Review, 38(3):520–522, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1996:BRA

[Ort96b] Norbert Ortner. Book review: Analytic Study of the Post Attenuation Correction Algorithm (William G. Hawkins and Stephen Glick). SIAM Review, 38(3): 520–522, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Osborne:1985:BRN

[Osb85] A. D. Osborne. Book review: A Nonlinear Differential Equation (M. A. Abdelkader). SIAM Review, 27 (1):83, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Oser:1975:AD

[Ose75] H. J. Oser. An average distance. SIAM Review, 17 (3):566, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Odlyzko:1982:ABF

[OSS82] A. M. Odlyzko, L. A. Shepp, and D. Slepian. Asymptotic behavior of an n-fold integral. SIAM Review, 24(4): 478, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Owhadi:2013:OUQ

[OSS^{+}13] H. Owhadi, C. Scovel, T. J.

Sullivan, M. McKerns, and M. Ortiz. Optimal uncertainty quantification. SIAM Review, 55(2):271–345, ???? 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Owhadi:2015:BBI

[OSS15] Houman Owhadi, Clint Scovel, and Tim Sullivan. On the brittleness of Bayesian inference. SIAM Review, 57(4): 566–582, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ostrowski:1961:BRM

[Ost61a] A. M. Ostrowski. Book review: Méthodes de Calcul Numérique. Tome 1, Algèbre Non Linéaire (A. Korganoff). SIAM Review, 3 (3):269, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ostrowski:1961:BRT

[Ost61b] A. M. Ostrowski. Book review: The Theory of Equations with an Introduction to the Theory of Binary Algebraic Forms (W. S. Burnside and A. W. Panton). SIAM Review, 3(3):268, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Oster:1991:LDI

[Ost91] C. A. Oster. Limit of a definite integral. SIAM Review, 33(1):115–116, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olshevsky:2011:CSD

[OT11] Alex Olshevsky and John N.

Tsitsiklis. Convergence speed in distributed consensus and averaging. SIAM Review, 53 (4):747–772, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i4/p747_s1.

Ockendon:2012:TLS

[OT12] J. R. Ockendon and R. H.

Tew. Thin-layer solutions of the Helmholtz and related equations. SIAM Review, 54 (1):3–51, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v54/i1/p3_s1.

Olver:2013:FWC

[OT13] Sheehan Olver and Alex Townsend. A fast and wellconditioned spectral method. SIAM Review, 55(3):462– 489, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Othmer:1982:BRS

[Oth82] Hans G. Othmer. Book review: Self-Organization in Nonequilibrium Systems (G.

Nicolis and I. Prigogine). SIAM Review, 24(4):483– 485, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ott:1996:BRC

[Ott96] Edward Ott. Book review: Chaotic Behavior of Deterministic Dissipative Systems (Milos Marek and Igor Schreiber). SIAM Review, 38 (4):695–696, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortega:1985:SPD

[OV85] James M. Ortega and Robert G. Voigt. Solution of partial differential equations on vector and parallel computers. SIAM Review, 27(2): 149–240, June 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Overhauser:1966:MDI

[Ove66] A. W. Overhauser. The minimum of a double integral. SIAM Review, 8(2): 233, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Overton:2005:SR

[Ove05] Michael Overton. Survey and review. SIAM Review, 47(4):627, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97155.

Ortner:1988:RDI

[OW88] Norbert Ortner and Peter Wagner. Reducing a double integral. SIAM Review, 30 (3):503–504, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1989:LT

[OW89a] N. Ortner and P. Wagner. A Laplace transform. SIAM Review, 31(3):497– 498, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1989:BRL

[OW89b] Norbert Ortner and Peter Wagner. Book review: A Laplace Transform (M. L. Glasser). SIAM Review, 31 (3):497–498, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1995:SBF

[OW95a] Norbert Ortner and Peter Wagner. A series of Bessel functions. SIAM Review, 37 (2):236–237, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ortner:1995:SCW

[OW95b] Norbert Ortner and Peter Wagner. Series, the convergence of which should be interpreted in the sense of L. Schwartz’s distributions. SIAM Review, 37(3):428– 435, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/27304.htm.

Owen:1985:BRA

[Owe85] David R. Owen. Book review: Author’s Comments on Joseph Kestin’s Review of: a First Course in the Mathematical Foundations of Thermodynamics (David R. Owen). SIAM Review, 27(2): 313–314, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Owen:2006:HSA

[Owe06] Art B. Owen. Halton sequences avoid the origin. SIAM Review, 48(3):487– 503, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Owhadi:2017:MRC

[Owh17] Houman Owhadi. Multigrid with rough coefficients and multiresolution operator decomposition from hierarchical information games. SIAM Review, 59(1):99– 149, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pryce:1995:NSS

[PA95] John D. Pryce and Alan L.

Andrew. Numerical solution of Sturm–Liouville problems. SIAM Review, 37(2): 252–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Petzold:2001:GEC

[PAB^{+}01] L. R. Petzold, U. Ascher, H. T. Banks, J. Crowley, W. Gander, L. Greengard, M. Heath, A. Lumsdaine, C. Moler, T. Oden, R. Schnabel, K. Stewart, and A. Trefethen. Graduate education in computational science and engineering. SIAM Review, 43(1):163–177, ???? 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Packel:1986:BRG

[Pac86] Edward W. Packel. Book review: A General Theory of Optimal Algorithms (J. F. Traub and H. Wozniakowski). SIAM Review, 28 (3):435–437, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Packel:1987:BRG

[Pac87] Edward W. Packel. Book review: Game Theory in the Social Sciences Concepts and Solutions (Martin Shubik). SIAM Review, 29(1): 158–160, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pachter:1988:BCB

[Pac88a] M. Pachter. Birthday candles blowout problem. SIAM Review, 30(4):650, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Packel:1988:BRT

[Pac88b] Edward W. Packel. Book review: Team Theory (K. H. Kim and F. W. Roush). SIAM Review, 30(4):676– 677, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Packel:1988:DLP

[Pac88c] Edward W. Packel. Do linear problems have linear optimal algorithms? SIAM Review, 30(3):388–403, September 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Packel:2010:BRM

[Pac10] Edward W. Packel. Book review: Mathematics and Politics: Strategy, Voting, Power, and Proof. Second Edition. SIAM Review, 52 (4):785–786, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Paige:1976:CTE

[Pai76] C. C. Paige. Computational technique for estimation and control. SIAM Review, 18 (4):821, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Palmer:1984:DEC

[Pal84a] K. Palmer. Differential equations — classical to controlled — Lukes, D. L. SIAM Review, 26(2):291– 292, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Palmer:1984:BRD

[Pal84b]  Kenneth Palmer.  Book

review: Differential Equations: Classical to Controlled (Dahlard L. Lukes). SIAM Review, 26(2):291– 292, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Palmore:1994:BRC

[Pal94] Julian Palmore. Book review: Chaos, Dynamics and Fractals: an Algorithmic Approach to Deterministic Chaos (J. L. McCauley).

SIAM Review, 36(3):515– 517, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Palmer:1995:BRN

[Pal95] Kenneth J. Palmer. Book review: Normally Hyperbolic Invariant Manifolds in Dynamical Systems (Stephen Wiggins). SIAM Review, 37 (3):472, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Palmer:1996:BRG

[Pal96] Kenneth J. Palmer. Book review: Global Dynamics, Phase Space Transport, Orbits Homoclinic to Resonances, and Applications (Stephen Wiggins). SIAM Review, 38(3):538– 539, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pan:1984:HCW

[Pan84] Victor Pan. How can we speed up matrix multiplication? SIAM Review, 26 (3):393–415, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pan:1986:ACD

[Pan86] Victor Pan. Author’s comments on Don Coppersmith’s review. SIAM Review, 28(3):

443–444, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pan:1992:CCM

[Pan92]  Victor Pan.  Complex-

ity of computations with matrices and polynomials. SIAM Review, 34(2):225– 262, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Panagiotopoulos:1993:BRC

[Pan93a] P. D. Panagiotopoulos. Book review: Convex Models of Uncertainty in Applied Mechanics (Yakov Ben-Haim and Isaac Elishakoff). SIAM Review, 35(2):331–332, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Panagiotopoulos:1993:CMU

[Pan93b] P. D. Panagiotopoulos. Convex models of uncertainty in applied mechanics (Yakov Ben-Haim and Isaac Elishakoff). SIAM Review, 35 (2):331–??, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Panchapakesan:1995:BRM

[Pan95] S. Panchapakesan. Book review: Measurement, Regression, and Calibration (Philip J. Brown). SIAM Review, 37 (4):635–636, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pan:1997:SPE

[Pan97]  Victor Y. Pan.  Solv-

ing a polynomial equation: Some history and recent progress. SIAM Review, 39(2):187–220, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28855.

Pao:1988:BRS

[Pao88] Yoh-Han Pao. Book review: Structural Methods in Pattern Recognition (Laurent Miclet; J. Howlett, trans.). SIAM Review, 30(1):152– 153, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pao:1990:BRF

[Pao90a] Yoh-Han Pao. Book review: Fundamentals of Pattern Recognition (Monique Pavel). SIAM Review, 32 (4):702, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pao:1990:FPR

[Pao90b] Yoh-Han Pao. Fundamentals of pattern recognition (Monique Pavel). SIAM Review, 32(4):702–??, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pao:1993:BRS

[Pao93] Yoh-Han Pao. Book review: Statistical Pattern Recognition. Second edition (Keinosuke Fukunaga). SIAM Review, 35(1):161, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Papadopoulos:1962:BRTa

[Pap62a] Michael Papadopoulos. Book review: Tables of Indefinite Integrals (G. Petit Bois). SIAM Review, 4 (3):272, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Papadopoulos:1962:BRTb

[Pap62b] Michael Papadopoulos. Book review: Theory of Functions as Applied to Engineering Problems (R. Rothe, F. Ollendorff and K. Pohlhasen, eds.; A, Herzenberg, trans.). SIAM Review, 4(4):405, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Papadopoulos:1963:BRC

[Pap63a]  M. Papadopoulos.  Book

review: Classical Electrodynamics (J. D. Jackson).

SIAM Review, 5(1):82–83, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Papadopoulos:1963:BRT

[Pap63b]  M. Papadopoulos.  Book

review: Theory of Ground Water Movement (P. Ya. Polubarinova-Kochina). SIAM Review, 5(3):297, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Papoulis:1985:LAW

[Pap85] A. Papoulis. Levinson’s algorithm, Wold’s decomposition, and spectral estimation. SIAM Review, 27 (3):405–441, September 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pappas:1986:DGS

[Pap86] Richard C. Pappas. Differentialgeometric solution of a problem in electrostatics. SIAM Review, 28(2):225–227, June 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parter:1961:ULG

[Par61] S. Parter. The use of linear graphs in Gauss elimination. SIAM Review, 3(2): 119–130, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1964:DUM

[Par64] Beresford Parlett. The development and use of methods of LR type. SIAM Review, 6 (3):275–295, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1966:BRA

[Par66] B. N. Parlett. Book review: The Algebraic Eigenvalue Problem (J. H. Wilkinson). SIAM Review, 8(4): 543–545, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parter:1967:BRT

[Par67a] Seymour S. Parter. Book review: Theoretical Numerical Analysis (Burton Wendroff). SIAM Review, 9(4): 758–759, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parzen:1967:BRS

[Par67b] Emanuel Parzen. Book review: Stochastic Processes (M. Girault). SIAM Review, 9(3):602, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Partek:1969:BRC

[Par69] Seymour V. Partek. Book review: A Course in Numerical Analysis (H. Melvin Lieberstein). SIAM Review, 11(2): 294–295, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1971:AAR

[Par71] Beresford N. Parlett. Analysis of algorithms for reflections in bisectors. SIAM Review, 13(2):197–208, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://www.math.utah.
edu/pub/bibnet/subjects/
acc-stab-num-alg-2ed.bib;
 https://www.math.utah.
edu/pub/bibnet/subjects/
acc-stab-num-alg.bib;
https://www.math.utah.
edu/pub/tex/bib/gvl.bib.

Partlett:1972:BRH

[Par72]  B. Partlett.  Book re-

view: Handbook for Automatic Computation, Vol. II, Linear Algebra (J. H. Wilkinson and C. Reinsch). SIAM Review, 14(4):658– 661, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1974:BRL

[Par74] B. N. Parlett. Book review: Large Sparse Sets of Linear Equations (J. K. Reid, ed.); Sparse Matrices and Their Applications (Donald J. Rose and Ralph A. Willoughby, ed.). SIAM Review, 16(3): 396–398, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parzen:1976:BRS

[Par76] Emanuel Parzen. Book review: The Spectral Analysis of Time Series (L. H. Koopmans). SIAM Review, 18(3): 515–516, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1978:PNA

[Par78] Beresford N. Parlett. Progress in numerical analysis. SIAM Review, 20(3):443–456, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

ftp://ftp.math.utah.edu/
pub/bibnet/subjects/acc-
stab-num-alg.bib.

Park:1980:BRR

[Par80]  C. J. Park.  Book re-

view: Random Allocations (Valentin F. Kolchin, Boris A. Sevast’yanov and Vladimir P. Chistyakov). SIAM Review, 22(1):104, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1981:BRL

[Par81a] B. N. Parlett. Book review:

LINPACK Users’ Guide (J. J. Dongarra, J. R. Bunch, C.

B. Moler and G. W. Stewart). SIAM Review, 23(1): 126–128, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parthasarathy:1981:FE

[Par81b] K. R. Parthasarathy. A functional equation. SIAM Review, 23(1):100, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parter:1986:IME

[Par86] Seymour V. Parter. Iterative methods for elliptic problems and the discovery of “q”. SIAM Review, 28(2): 153–175, June 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parker:1987:BRI

[Par87] Thomas Parker. Book review: Instantons and Four-Manifolds (Daniel S. Freed and Karen K. Uhlenbeck). SIAM Review, 29(1):160– 162, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pardoux:1988:BRM

[Par88]  Etienne Pardoux.  Book

review: Markov Processes: Characterization and Convergence (Stewart N. Ethier and Thomas G. Kurtz).

SIAM Review, 30(3):525– 526, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parker:1990:BRN

[Par90] D. F. Parker. Book review:

Nonlinear Evolution Equations (J. K. Engelbrecht, V. E. Fridman, and E. N. Pelinovski). SIAM Review, 32 (3):519–520, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parzen:1992:BRS

[Par92a] Emanuel Parzen. Book review: A Statistical Model: Frederick Mosteller’s Contributions to Statistics, Science, and Public Policy (S. E. Fienberg, D. C. Hoaglin, W. H. Kruskal, and J. M. Tanur, eds.). SIAM Review, 34 (1):147, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parzen:1992:BRF

[Par92b] Emanuel Parzen. Book review: Forecasting, Structural Time Series and the Kalman Filter (Andrew C. Harvey). SIAM Review, 34(1):155–156, March 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pardalos:1994:BRO

[Par94] Panos M. Pardalos. Book review: Optimization Software Guide (Jorge J. Moré and Stephen J. Wright). SIAM Review, 36(3):518, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:2010:BRM

[Par10] Beresford Parlett. Book review: The matrix eigenvalue problem: GR and Krylov subspace methods [book review of MR2383888]; Numerical methods for general and structured eigenvalue problems [book review of MR2164298]. SIAM Review, 52(4):771–774, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pasciak:1995:BRM

[Pas95] Joseph E. Pasciak. Book review: The Mathematical Theory of Finite Element Methods (Susanne C. Brenner and L. Ridgway Scott). SIAM Review, 37(3):472– 473, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Patterson:1964:BRM

[Pat64] Richard L. Patterson. Book review: Mathematical Theories of Traffic Flow (Frank A.

Haight). SIAM Review, 6(2): 187–188, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Powers:1980:SC

[PAT^{+}80] D. Powers, W. N. Anderson, G. E. Trapp, A. E. Eckberg, W. H. Gustafson, J. Z. Hearon, R. B. Kellogg, A. Khadr, O. P. Lossers, S. L. Paverifontana, P. Vandendriessche, and J. C. Willems. Super-controllability. SIAM Review, 22(2):230–231, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Patterson:1987:BRS

[Pat87] Richard R. Patterson. Book review: Surface Topology. (P. A. Firby and C. F. Gardiner). SIAM Review, 29(3): 486–487, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pataki:2003:TIP

[Pat03]  Gábor Pataki.  Teaching

integer programming formulations using the traveling salesman problem. SIAM Review, 45(1):116– 123, March 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/36855.

Pataki:2019:CBS

[Pat19] Gábor Pataki. Characterizing bad semidefinite programs: Normal forms and short proofs. SIAM Review, 61(4):839–859, ???? 2019. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pavel:1994:BRF

[Pav94]  N. H. Pavel.  Book re-

view: Functional and Numerical Methods in Viscoplasticity (I. R. Ionescu and M. Sofonea). SIAM Review, 36 (4):675–676, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Payne:1967:IIT

[Pay67] L. E. Payne. Isoperimetric inequalities and their applications. SIAM Review, 9(3): 453–488, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Payne:1968:BRM

[Pay68] Lawrence E. Payne. Book review: Maximum Principles in Differential Equations (Murray H. Protter and Hans F. Weinberger). SIAM Review, 10(3):386– 387, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Paz:1976:BRS

[Paz76]  A. Paz.  Book review:

Stochastic Automata Constructive Theory (A. A. Lorenz). SIAM Review, 18 (3):510–511, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Poole:1974:SM

[PB74] George Poole and Thomas Boullion. A survey on Mmatrices. SIAM Review, 16 (4):419–427, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Phillips:1975:WBP

[PB75] O. M. Phillips and M. L.

Banner. Wave breaking in presence of wind drift and swell. SIAM Review, 17 (2):387, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Petrov:1996:LTP

[PB96] Valentin V. Petrov and Wlodzimierz Bryc. Limit theorems of probability theory: Sequences of independent random variables. SIAM Review, 38(3):527– ??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peres:1998:TLM

[PBL98] Pedro L. D. Peres, Ivanil S. Bonatti, and Amauri Lopes. Transmission line modeling: a circuit theory approach. SIAM Review, 40(2):347–352, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31604.

Parks:1961:ANI

[PC61] J. R. Parks and L. Cooper. The application of numerical integration to a problem in viscoelasticity. SIAM Review, 3(4):315–321, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Piegorsch:1989:EUM

[PC89] Walter W. Piegorsch and George Casella. The early use of matrix diagonal increments in statistical problems. SIAM Review, 31 (3):428–434, September 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See erratum [PC90b].

Piegorsch:1990:ISM

[PC90a] W. W. Piegorsch and G. Casella. Inverting a sum of matrices. SIAM Review, 32(3): 470, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Piegorsch:1990:EIS

[PC90b] Walter W. Piegorsch and George Casella. Erratum. inverting a sum of matrices: “The early use of matrix diagonal increments in statistical problems” [SIAM rev. 31 (1989), no. 3, 428– 434 and MR 90g:65044]. SIAM Review, 32(3):470, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [PC89].

Pruss:1995:EIE

[PC95] J. Pruss and C. Corduneanu.

Evolutionary integral equations and applications. SIAM Review, 37(1):117–??, March 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Powell:1998:GTM

[PCH98] James A. Powell, James S.

Cangelosi, and Ann Marie Harris. Games to teach mathematical modelling. SIAM Review, 40(1):87–95, March 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/31002.

Pragliola:2023:BTV

[PCLS23] Monica Pragliola, Luca Calatroni, Alessandro Lanza, and Fiorella Sgallari. On and beyond total variation regularization in imaging: The role of space variance. SIAM Review, 65 (3):601–685, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1410683.

Perelson:1976:ISN

[PD76] A. S. Perelson and C. Delisi. An infinite system of nonlinear differential equations. SIAM Review, 18(2):296– 298, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pennell:1989:CPA

[PD89] S. Pennell and J. Deignan. Computing the projected area of a cone. SIAM Review, 31(2):299–302, June 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Poularikas:1997:TAH

[PD97] Alexander D. Poularikas and Lokenath Debnath. The transforms and applications handbook. SIAM Review, 39 (4):791–??, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Postek:2016:CTC

[PdHM16] Krzysztof Postek, Dick den Hertog, and Bertrand Melenberg. Computationally tractable counterparts of distributionally robust constraints on risk measures. SIAM Review, 58(4):603– 650, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pridor:1978:SSF

[PDS78] A. Pridor, R. C. Diprima, and J. T. Stuart. Stability of spiral flow. SIAM Review, 20(3):633, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parthasarathy:1976:ASS

[PE76] A. Parthasarathy and R. M.

Evaniwanowski. Almost-sure stability of linear stochasticsystems. SIAM Review, 18 (4):821, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pfanzagl:1996:PST

[PE96] Johann Pfanzagl and Morris L. Eaton. Parametric statistical theory. SIAM Review, 38(3):527–??, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pearson:1965:NSD

[Pea65] Carl E. Pearson. A note on the second derivative of a P. R. function. SIAM Review, 7 (2):251–252, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pearson:1981:DTS

[Pea81] Carl E. Pearson. Dual time scales in a wave problem governed by coupled nonlinear equations. SIAM Review, 23 (4):425–433, October 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pearson:1982:OMS

[Pea82] Carl E. Pearson. Optimal mapping from a sphere onto a plane. SIAM Review, 24(4): 469–475, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pearce:1997:DTB

[Pea97] C. E. M. Pearce. The determinant of a triangular-block matrix. SIAM Review, 39(3): 517–520, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peck:1962:PCF

[Pec62] J. E. L. Peck. Polynomial curve fitting with constraint. SIAM Review, 4(2): 135–141, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pieroux:2003:HBS

[PEGK03] Didier Pieroux, Thomas Erneux, Athanasios Gavrielides, and Vassilios Kovanis. Hopf bifurcation subject to a large delay in a laser system. SIAM Review, 45(3): 523–540, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42355.

Peizhuang:1983:BRP

[Pei83] Wang Peizhuang. Book review: Pattern Recognition with Fuzzy Objective Function Algorithms (James C. Bezdek). SIAM Review, 25 (3):442, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peligrad:1997:BRS

[Pel97] Magda Peligrad. Book review: Stable Non-Gaussian Random Processes, by Gennady Samorodnitsky and Murad S. Taqqu. SIAM Review, 39(2):356–358, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pentikainen:1988:AC

[Pen88] Teivo Pentikäinen. Author’s comment. SIAM Review, 30 (3):533–536, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Penrice:1995:AEP

[Pen95] Stephen G. Penrice. Applying elementary probability theory to the NBA draft lottery. SIAM Review, 37 (4):598–602, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/27302.
htm.

Perry:1965:BRI

[Per65] John M. Perry. Book review: Integration of Equations of Parabolic Type by the Method of Nets (V. K. Saul’yev). SIAM Review, 7 (4):576–577, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Perry:1968:BRN

[Per68] John M. Perry. Book review: Numerical Solution of Partial Differential Equations (G. D. Smith). SIAM Review, 10(3): 396–397, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Perry:1976:BRD

[Per76] John M. Perry. Book review: Discrete Numerical Methods in Physics and Engineering (Donald GreenSpan). SIAM Review, 18(1):148– 149, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pereira:1979:FSC

[Per79] N. R. Pereira. Fourier series for a combination of Jacobian elliptic functions. SIAM Review, 21(2):257, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Percus:1989:BRS

[Per89] J. K. Percus. Book review: Systems in Stochastic Equilibrium (Peter Whittle). SIAM Review, 31(1): 154–155, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Perry:1993:BRL

[Per93a]  Nigel Perry.  Book re-

view: Lazy Functional Languages: Abstract Interpretation and Compilation (Geoffrey Burn). SIAM Review, 35 (2):332–334, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Perry:1993:LFL

[Per93b] Nigel Perry. Lazy functional languages: Abstract interpretation and compilation (Geoffrey Burn). SIAM Review, 35(2):332–??, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Petty:1968:BRQ

[Pet68] James S. Petty. Book review: A Quadruple Integral (L. Carlitz). SIAM Review, 10(4):457–458, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peterson:1973:RCQ

[Pet73] David W. Peterson. A review of constraint qualifications in finite-dimensional spaces. SIAM Review, 15(3): 639–654, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peterson:1976:DVP

[Pet76a] E. L. Peterson. Dual variational principles for predicting economic equilibria. SIAM Review, 18(4): 822, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peterson:1976:GP

[Pet76b] Elmor L. Peterson. Geometric programming. SIAM Review, 18(1):1–51, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peters:1978:ODH

[Pet78] David A. Peters. Optimum design of a hydraulic cushion unit. SIAM Review, 20(4): 847–849, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Petkovic:1985:ISB

[Pet85] D. M. Petković. Infinite sums of Bessel functions. SIAM Review, 27(3):446, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peterson:1991:SNM

[Pet91]  Ivars Peterson.  Search-

ing for new mathematics. SIAM Review, 33(1): 37–42, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peters:1995:BRG

[Pet95]  Jörg Peters.  Book re-

view: Geometric Concepts for Geometric Design (W. Boehm and H. Prautzsch). SIAM Review, 37(3):473– 474, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peters:1997:OSD

[Pet97a] D. A. Peters. Optimum spring-damper design for mass impact. SIAM Review, 39(1):118–122, ???? 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peters:1997:CNO

[Pet97b] David A. Peters. Classroom note: Optimum springdamper design for mass impact. SIAM Review, 39 (1):118–122, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28154;   http:
//www.siam.org/journals/
sirev/sirev391.htm.

Peterson:1998:EGE

[Pet98] T. E. Peterson. Eliminating Gibb’s effect from separation of variables solutions. SIAM Review, 40(2): 324–326, June 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31622.

Persico:1962:ACT

[PF62] Donald J. Persico and Henry C. Friedman. Another coin tossing problem. SIAM Review, 4(3):255, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Puterman:1996:MDP

[PF96] Martin L. Puterman and Eugene A. Feinberg. Markov decision processes: Discrete stochastic dynamic programming. SIAM Review, 38(4): 689–??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pavon:2013:GME

[PF13] Michele Pavon and Augusto Ferrante. On the geometry of maximum entropy problems. SIAM Review, 55(3): 415–439, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pflederer:1992:BRN

[Pfl92] Steven J. Pflederer. Book review: Numerical Control: Making a New Technology (J. Francis Reintjes). SIAM Review, 34(3): 517, September 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Piele:1976:ACM

[PFM76] D. T. Piele, M. W. Firebaugh, and R. J. Manulik. Applications of conformal mapping to potential theory through computer graphics. SIAM Review, 18 (4):822, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pierce:2000:ARS

[PG00] Niles A. Pierce and Michael B.

Giles. Adjoint recovery of superconvergent functionals from PDE approximations. SIAM Review, 42(2): 247–264, June 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34942.

Pinault:1988:SCF

[PGH^{+}88] S. C. Pinault, R. W. Gatterdam, J. A. Heminger, S. L. Paverifontana, and W. B. Jordan. Safe care following. SIAM Review, 30(3): 507–508, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Poole:1975:ARB

[PGS75] W. G. Poole, N. E. Gibbs, and P. K. Stockmeyer. Algorithm for reducing bandwidth and profile of a sparse matrix. SIAM Review, 17 (2):387, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Patrick:1976:FIR

[PH76] M. Patrick and E. Hansen. Family of iterative rootfinding methods. SIAM Review, 18(4):821, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1978:NPG

[PH78] B. N. Parlett and W. Hoffman. New proof of global convergence for tridiagonal QL algorithm. SIAM Review, 20(3):632, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Previte:2013:PDC

[PH13] Joseph P. Previte and Kathleen A. Hoffman. Period doubling cascades in a predatorprey model with a scavenger. SIAM Review, 55(3): 523–546, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Phipps:1968:CFR

[Phi68] Thomas E. Phipps, Jr.

A continued fraction representation of eigenvalues. SIAM Review, 10(2):223– 224, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Phipps:1971:DSI

[Phi71a] T. E. Phipps, Jr. A doublesum identity. SIAM Review, 13(3):389, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Phipps:1971:CFR

[Phi71b] Thomas E. Phipps, Jr.

A continued fraction representation of eigenvalues. SIAM Review, 13(3):390– 395, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Philippou:1985:MRV

[Phi85] A. N. Philippou. Modes of a random variable. SIAM Review, 27(1):79, ???? 1985.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Philippou:1987:ARC

[Phi87] A. N. Philippou. Asymptotic reliability of consecutive k-out-of- n: F system. SIAM Review, 29(1): 123–124, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Platkowski:1988:DVM

[PI88] Tadeusz Płatkowski and Reinhard Illner. Discrete velocity models of the Boltzmann equation: a survey on the mathematical ASPECTS of the theory. SIAM Review, 30(2):213–255, June 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Picard:1987:BRU

[Pic87a] Rainer H. Picard. Book review: A Unified Theory of Nonlinear Operator and Evolution Equations with Applications. A New Approach to Nonlinear Partial Differential Equations (Mieczyslaw Altman). SIAM Review, 29 (4):668–670, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pickands:1987:BRS

[Pic87b] James Pickands III. Book review: Stationary Sequences and Random Fields (Murray Rosenblatt). SIAM Review, 29(4):670–671, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Picard:1990:BRN

[Pic90] Rainer H. Picard. Book review: Nonlinear Evolution Equations and Applications (Gheorghe Morosanu). SIAM Review, 32(1):180– 182, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Picard:1992:BRS

[Pic92] Rainer Picard. Book review: Singular Electromagnetic Fields and Sources (J. Van Bladel). SIAM Review, 34(3):517–518, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Piless:1987:BRP

[Pil87] Vera Piless. Book review:

Problems In Combinatorics and Graph Theory (Ioan Tomescu). SIAM Review, 29 (1):164, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pimbley:1987:BRS

[Pim87] Joseph M. Pimbley. Book review: The Stationary Semiconductor Device Equations. (Peter A. Markowich).

SIAM Review, 29(4):671– 673, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pimbley:1993:BRQ

[Pim93]  Joe Pimbley.  Book re-

view: Quantum Theory of the Optical and Electronic Properties of Semiconductors (H. Haug and Stephan W. Koch). SIAM Review, 35 (4):676–677, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinsky:1969:BRC

[Pin69] Mark Pinsky. Book review: Convergence of Probability Measures (Patrick Billingsley). SIAM Review, 11(3): 414–415, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinsky:1970:BRM

[Pin70] Mark Pinsky. Book review: Markov Processes, Theorems and Problems (Evgenii B. Dynkin and Aleksandr A. Yushkevich). SIAM Review, 12(2):305–306, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinsky:1972:BRI

[Pin72a] Mark A. Pinsky. Book review: An Introduction to Probability Theory and Its Applications, Vol. 2 (William Feller). SIAM Review, 14(4): 662–663, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinsky:1972:BRB

[Pin72b] Mark A. Pinsky. Book review: Brownian Motion and Diffusion, Vol. II (David Freedman). SIAM Review, 14 (4):667–668, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinsky:1977:BRS

[Pin77] Mark A. Pinsky. Book review: Stochastic Calculus and Stochastic Models (E. J. McShane). SIAM Review, 19 (3):583–584, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinsky:1984:NAM

[Pin84a]  M. A. Pinsky.  Nonlin-

ear analysis on manifolds — Monge–Ampere equations — Aubin, T. SIAM Review, 26 (4):593–594, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinsky:1984:BRN

[Pin84b]  Mark A. Pinsky.  Book

review: Nonlinear Analysis on Manifolds. Monge– Ampere Equations (Thierry Aubin). SIAM Review, 26(4):

593–594, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinkham:1985:BRAa

[Pin85a]  Roger Pinkham.  Book

review: Abstract Algebra — a Computational Approach (Charles C. Sims) and A CLASSLIB User’s Manual (Charles C. Sims). SIAM Review, 27(3):471– 472, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinkham:1985:BRAb

[Pin85b] Roger S. Pinkham. Book review: Analysis of Survival Data (D. R. Cox and D. Oakes). SIAM Review, 27 (4):602, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinkham:1987:BRM

[Pin87a] Roger Pinkham. Book review: Mathematical Experiments on the Computer (Ulf Grenander). SIAM Review, 29(1):162–163, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinsky:1987:BRS

[Pin87b]  Mark Pinsky.  Book re-

view: Stochastic Differential Equations. An Introduction with Applications (B. Oksendal). SIAM Review, 29(1): 163–164, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinault:1988:BRS

[Pin88] Steven C. Pinault. Book review: Safe Car Following (Vincent Salmon). SIAM Review, 30(3):507–508, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinkham:1992:ANR

[Pin92a] Roger Pinkham. Applied nonparametric regression (Wolfgang Hardle). SIAM Review, 34(2):341–??, June 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinkham:1992:BRA

[Pin92b] Roger Pinkham. Book review: Applied Nonparametric Regression (Wolf gang Hardle). SIAM Review, 34(2): 341–342, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pinsky:1992:AST

[Pin92c] Mark A. Pinsky. Asymptotic solution of the telegraph equation. SIAM Review, 34 (2):308–309, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pipkin:1981:BRE

[Pip81] A. C. Pipkin. Book review: Elongational Flows: Aspects of the Behavior of Model Elasticoviscous Fluids (C. J. S. Petrie). SIAM Review, 23 (1):114–115, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pipkin:1986:BRO

[Pip86]  Allen C. Pipkin.  Book

review: Orienting Polymers: Proceedings, Minneapolis 1983 (J. L. Ericksen, ed.). SIAM Review, 28 (1):120, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pipkin:1992:BRF

[Pip92] Allen C. Pipkin. Book review: Fluid of Dynamics of Viscoelastic Liquids (D. D. Joseph). SIAM Review, 34 (1):138, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pittenger:1974:BRA

[Pit74] A. O. Pittenger. Book review: Approximating Countable Markov Chains (David Freedman). SIAM Review, 16 (4):562–564, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pittenger:1977:BRI

[Pit77] A. O. Pittenger. Book review:

Introduction to Stochastic Processes (Erhan Çinlar). SIAM Review, 19(3):573– 575, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pittenger:1979:BRS

[Pit79] Arthur O. Pittenger. Book review: Stochastic Processes: a Survey of the Mathematical Theory (John Lamperti). SIAM Review, 21(3):421– 422, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pittenger:1981:BRM

[Pit81] A. O. Pittenger. Book review:

Markov Chain Models-Rarity and Exponentiality (Julian Keilson). SIAM Review, 23 (2):261–262, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pitman:1989:BRV

[Pit89] E. Bruce Pitman. Book review: Very Slow Flow of Solids: Basics of Modeling in Geodynamics and Glaciology (L. A. Lliboutry). SIAM Review, 31(4):707– 708, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pitman:1990:ASS

[Pit90a] E. Bruce Pitman. Analysis of skeletal structural systems in the elastic and elastic-plastic range (Adam Borkowski). SIAM Review, 32(4):702– ??, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pitman:1990:BRA

[Pit90b] E. Bruce Pitman. Book review: Analysis of Skeletal Structural Systems in the Elastic and Elastic-Plastic Range (Adam Borkowski). SIAM Review, 32(4):702– 703, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pitman:1991:BRN

[Pit91] E. Bruce Pitman. Book review: Numerical Methods for Conservation Laws (Randall J. LeVeque). SIAM Review, 33(3):506–507, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pitman:1993:BRC

[Pit93a]  Jim Pitman.  Book re-

view: Continuous Martingales and Brownian Motion (Daniel Revuz and Mare Yor). SIAM Review, 35(3): 532–533, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pitman:1993:CMB

[Pit93b] Jim Pitman. Continuous martingales and Brownian motion (Daniel Revuz and Marc Yor). SIAM Review, 35 (3):532–??, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Powers:1975:NME

[PJ75] D. L. Powers and R. Jeltsch.

Norm of a matrix exponential. SIAM Review, 17(1): 174–176, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Payne:1978:LSE

[PL78] F. R. Payne and L. A. Lemmerman. Large-scale eigenstructures of turbulent-shear flows — dual strategy for prediction and measurement of stochastic fields. SIAM Review, 20(3):632, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Plakhov:2015:PMR

[Pla15] Alexander Plakhov. Problems of minimal resistance and the Kakeya problem. SIAM Review, 57(3):421– 434, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pearson:1996:ADC

[PLC96] J. M. Pearson, N. G. Lloyd, and C. J. Christopher. Algorithmic derivation of centre conditions. SIAM Review, 38 (4):619–636, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28357.

Plemmons:1988:BRM

[Ple88] Robert J. Plemmons. Book review: Matrix Analysis (Roger A. Horn and Charles R. Johnson). SIAM Review, 30(1):153–154, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pachepsky:2005:EDP

[PLL05] Elizaveta Pachepsky, Frithjof Lutscher, and Mark A. Lewis. The effect of dispersal patterns on stream populations. SIAM Review, 47(4):749– 772, ???? 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Plant:1987:MSA

[PM87] Richard E. Plant and Marc Mangel. Modeling and simulation in agricultural pest management. SIAM Review, 29(2):235–261, June 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peretto:1995:IMN

[PMG95] P. Peretto, Christopher W.

Myers, and Frank H. Guenther. An introduction to the modeling of neural networks. SIAM Review, 37(4): 629–??, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Philipsen:1996:MSC

[PMK96] C. Philipsen, S. Markvorsen, and W. Kliem. Modelling the stem curve of a Palm in a strong wind. SIAM Review, 38(3):483– 484, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/26382.

Putnick:1975:SSR

[PMR75] L. J. Putnick, B. J. Matkowsky, and E. L. Reiss. Secondary states of rectangular plates. SIAM Review, 17 (2):387, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Putnick:1976:SBN

[PMR76] L. J. Putnick, B. J. Matkowsky, and E. L. Reiss. Secondary bifurcations of nonlinear plate vibrations. SIAM Review, 18(4):823, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Perelson:1999:MAH

[PN99] Alan S. Perelson and Patrick W. Nelson. Mathematical analysis of HIV-1 dynamics in vivo. SIAM Review, 41(1): 3–44, March 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/33510.

Proskurowski:1976:FMI

[PO76] A. Proskurowski and M. Overton. Finding maximal incidence matrix of a large graph. SIAM Review, 18 (4):822, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pohle:1963:BRD

[Poh63] Frederick V. Pohle. Book review: Dynamics; Particles, Rigid Bodies, and Systems (vol. I). Systems, Variational Methods, and Relativity (vol. II) (R. L. Halfman). SIAM Review, 5(2): 166–169, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pohl:1978:IDK

[Poh78]  I. Pohl.  Improvements

to Dinitz–Karzanov network flow algorithm. SIAM Review, 20(3):633, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Polaneczky:1959:BRA

[Pol59] Aloysius J. Polaneczky. Book review: Analysis for Production Management (E. H. Bowman and R. B. Fetter). SIAM Review, 1(2): 176, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pollak:1965:BRF

[Pol65] H. O. Pollak. Book review:

The Fourier Integral and Its Applications (Athanasios Papoulis). SIAM Review, 7(1): 159–161, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pollard:1969:BRA

[Pol69] Harry Pollard. Book review: Analytical and Numerical Methods of Celestial Mechanics (G. A. Chebotarev). SIAM Review, 11 (1):98, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Polak:1973:HSC

[Pol73]  E. Polak.  An histori-

cal survey of computational methods in optimal control. SIAM Review, 15(2): 553–584, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pollard:1976:BRC

[Pol76] Harry Pollard. Book review: Celestial Mechanics, Part I (Shlomo Sternberg). SIAM Review, 18(1):132, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Polak:1987:MFN

[Pol87] E. Polak. On the mathematical foundations of nondifferentiable optimization in engineering design. SIAM Review, 29(1):21–89, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pomentale:1967:MBW

[Pom67] T. Pomentale. On minimization of backboard wiring functions. SIAM Review, 9 (3):564–568, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ponstein:1965:EMM

[Pon65]  J. Ponstein.  An exten-

sion of the min-max theorem. SIAM Review, 7(2): 181–188, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ponstein:1967:SKC

[Pon67] J. Ponstein. Seven kinds of convexity. SIAM Review, 9 (1):115–119, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ponstein:1983:ASM

[Pon83] J. Ponstein. Applying some modern developments to choosing your own Lagrange multipliers. SIAM Review, 25 (2):183–199, April 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ponce:1990:BRI

[Pon90a] Gustavo Ponce. Book review:

Initial-Boundary Value Problems and the Navier–Stokes Equations (Heinz-Otto Kreiss and Jens Lorenz). SIAM Review, 32(4):703–704, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ponce:1990:IBV

[Pon90b]  Gustavo Ponce.  Initial-

boundary value problems and the Navier–Stokes equations (Heinz-Otto Kreiss and Jens Lorenz). SIAM Review, 32 (4):703–??, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Porter:1964:NSL

[Por64]  R. W. Porter.  Numeri-

cal solution for local emission coefficients in axisymmetric self-absorbed sources. SIAM Review, 6(3):228– 242, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Porsching:1972:BRL

[Por72] T. A. Porsching. Book review: Lectures on the Numerical Solution of Linear, Singular, and Nonlinear Differential Equations (D. Greenspan). SIAM Review, 14(1):179–180, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Porsching:1980:BRN

[Por80]  T. A. Porsching.  Book

review: Numerical Solution of Partial Differential Equations: Finite Difference Methods (G. D. Smith). SIAM Review, 22 (3):376, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Port:1981:BRD

[Por81] Sidney C. Port. Book review: Diffusions, Markov Processes, & Martin gales: Volume 1: Foundations (David. Williams). SIAM Review, 23 (1):120, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Porsching:1983:BRA

[Por83]  T. A. Porsching.  Book

review: Applied Iterative Methods (Louis A. Hageman and David M. Young).

SIAM Review, 25(2):281– 283, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Porsching:1985:BRM

[Por85] T. A. Porsching. Book review: The Method of Differential Approximation (Yu. I. Shokin). SIAM Review, 27 (2):297–298, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Porsching:1986:BRP

[Por86a] Thomas A. Porsching. Book review: Progress and Supercomputing in Computational Fluid Dynamics: Proceedings of U.S.–Israel Workshop, 1984 (Earll M. Murman and Saul S. Abarbanel, eds.). SIAM Review, 28(4): 601, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Port:1986:BRP

[Por86b]  Sidney C. Port.  Book

review: Probability, Random Variables, and Stochastic Processes-Second Edition (Athanasios Papoulis). SIAM Review, 28(2):275– 276, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Porsching:1988:BRN

[Por88] Thomas A. Porsching. Book review: Numerical Methods in Fluid Dynamics-Como, 1983 (F. Brezzi, ed.). SIAM Review, 30(2):345– 346, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Port:1989:BRI

[Por89]  Sidny Port.  Book re-

view: Introduction to Random Processes (Yurii A. Rozanov). SIAM Review, 31 (1):155, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Port:1992:BRC

[Por92] Daniel Port. Book review: A Course in Simulation (Sheldon M. Ross). SIAM Review, 34(2):342–343, June 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Powell:1970:SNM

[Pow70] M. J. D. Powell. A survey of numerical methods for unconstrained optimization. SIAM Review, 12(1): 79–97, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Powers:1975:BRN

[Pow75] David L. Powers. Book review: On the Norm of a Matrix Exponential (J. C. Cavendish). SIAM Review, 17(1):174–176, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Powers:1980:BRS

[Pow80] David Powers. Book review: Super-Controllability (O. Hájek). SIAM Review, 22 (2):230–231, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Powell:1986:CPA

[Pow86] M. J. D. Powell. Convergence properties of algorithms for nonlinear optimization. SIAM Review, 28 (4):487–500, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Powers:1993:PC

[Pow93] David L. Powers. Probabilities for a cuboctahedral die. SIAM Review, 35(1): 135–136, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Powers:1994:PC

[Pow94] David L. Powers. Probabilities for a cuboctahedral die. SIAM Review, 36(1): 109–110, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pagh:2011:LPW

[PPR11] Anna Pagh, Rasmus Pagh, and Milan Ruzić. Linear probing with 5-wise independence. SIAM Review, 53(3): 547–558, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i3/p547_s1.

Pringle:1970:EGI

[PR70] R. M. Pringle and A. A. Rayner. Expressions for generalized inverses of a bordered matrix with application to the theory of constrained linear models. SIAM Review, 12(1):107– 115, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pardalos:1986:MGC

[PR86] P. M. Pardalos and J. B. Rosen. Methods for global concave minimization: a bibliographic survey. SIAM Review, 28(3):367–379, September 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pincheon:1990:OP

[PR90] E. Pincheon and C. C. Rousseau. An optimization problem. SIAM Review, 32 (2):302, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Padberg:1991:BCA

[PR91] Manfred Padberg and Giovanni Rinaldi. A branch-and-cut algorithm for the resolution of large-scale symmetric traveling salesman problems. SIAM Review, 33(1): 60–100, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prager:1968:BRF

[Pra68] William Prager. Book review: Foundations of the Non-Linear Mechanics of Continua (L. I. Sedov). SIAM Review, 10(3):385– 386, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prasad:1969:D

[Pra69]  J. Prasad.  A determi-

nant. SIAM Review, 11 (1):74, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prakash:1985:BRN

[Pra85] Chander Prakash. Book review: Numerical Methods in Heat Transfer (R. W. Lewis, K. Morgan, and B. A. Schreifler, eds.). SIAM Review, 27 (1):118–119, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prakash:1986:BRN

[Pra86] Chander Prakash. Book review: Numerical Heat Transfer (T. M. Shin). SIAM Review, 28(2):276–277, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prakash:1989:BRP

[Pra89a] C. Prakash. Book review: Physical and Computational Aspects of Convective Heat Transfer; Solutions Manual; and Computer Programs (Tuncer Cebeci and Peter Bradshaw). SIAM Review, 31 (4):708–709, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prautzsch:1989:BRC

[Pra89b] Hartmut Prautzsch. Book review: Curve and Surface Fitting: an Introduction (Peter Lancaster and Kestutis Salkauskas). SIAM Review, 31(1):155–157, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prenter:1981:BRN

[Pre81] P. M. Prenter. Book review: The Numerical Treatment of Integral Equations (C. T. H. Baker). SIAM Review, 23(2): 266–267, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prea:1995:OSP

[Pre95a]  P. Prea.  Optimal sen-

sor placement in 3 dimensions. SIAM Review, 37(1): 112–113, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prea:1995:BRO

[Pré95b]  Pascal Préa.  Book re-

view: Optimal Sensor Placement in Three Dimensions (Hong Zhang and Xiaohuan Wang). SIAM Review, 37(1): 112–113, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Price:1964:MPM

[Pri64]  Charles M. Price.  The

matrix pseudoinverse and minimal variance estimates. SIAM Review, 6(2):115– 120, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pridor:1976:MMS

[Pri76] A. Pridor. Method of moments for solving transport problems. SIAM Review, 18 (4):822, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pritchard:1993:MMR

[Pri93] W. G. Pritchard. Mathematical models of running.

SIAM Review, 35(3):359– 379, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ponting:1966:BRT

[PRM^{+}66] F. W. Ponting, R. F. Rinehart, V. O. Mowery, J. Astin, T. Krishnan, and T. J. Rao. Book review: Nth Power of a Matrix (James F. Foley). SIAM Review, 8(2): 238–241, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Proschan:1961:PLS

[Pro61] Frank Proschan. Precision of least squares polynomial estimates. SIAM Review, 3(3): 230–236, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prosser:1965:BRG

[Pro65a] Reese T. Prosser. Book review: Group Theory, the Application to Quantum Mechanics (Paul Meijer and Edmond Bauer). SIAM Review, 7(2):300–301, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prosser:1965:BRI

[Pro65b] Reese T. Prosser. Book review: Introduction to Field Theory and Dispersion Relations (R. Hagedorn). SIAM Review, 7(3): 438–439, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Protter:1965:BRE

[Pro65c] Murray H. Protter. Book review: Equations of the Mixed Type (A. V. Bitsadze). SIAM Review, 7(2):298, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prosser:1983:SCA

[Pro83a] R. Prosser. Semi-classical approximation in quantum mechanics — Maslov, V. P, Fedoriuk, M. V. SIAM Review, 25(3):426–429, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prosser:1983:BRS

[Pro83b] Reese Prosser. Book review: Semi-Classical Approximation in Quantum Mechanics (V. P. Maslov and M. V. Fedoriuk). SIAM Review, 25(3):426–429, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prosser:1984:GQA

[Pro84a] R. T. Prosser. Geometric quantization in action — Hurt, N. E. SIAM Review, 26 (4):602–603, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prosser:1984:BRG

[Pro84b] Reese T. Prosser. Book review: Geometric Quantization in Action (Norman E. Hurt). SIAM Review, 26(4): 602–603, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Protter:1984:BRI

[Pro84c] M. H. Protter. Book review: Instability, Nonexistence and Weighted Energy Methods in Fluid Dynamics and Related Theories (B. Straughan). SIAM Review, 26(2):285–286, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Protter:1985:BRF

[Pro85] M. H. Protter. Book review: First Order Elliptic Systems: a Function Theoretic Approach (Robert P. Gilbert and James L. Buchanan). SIAM Review, 27 (1):119–121, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Protter:1987:BRB

[Pro87a] M. H. Protter. Book review: The Boundary Value Problems of Mathematical Physics (O. A. Ladyzhenskaya). SIAM Review, 29(2):

334–335, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Protter:1987:COH

[Pro87b] M. H. Protter. Can one hear the shape of a drum — revisited. SIAM Review, 29 (2):185–197, June 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Protter:1991:BRL

[Pro91] Philip Protter. Book review:

Limit Theorems for Stochastic Processes (J. Jacod and A . N. Shiryaev). SIAM Review, 33(2):332–333, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Protter:1995:BRR

[Pro95] Philip Protter. Book review:

Random Series and Stochastic Integrals: Single and Multiple (Stanislaw Kwapien and Wojbar A. Woyczynski). SIAM Review, 37(1): 135–136, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Protter:1996:BRN

[Pro96] Philip Protter. Book review: Numerical Solution of SDE through Computer Experiments (P. E. Kloeden, E. Platen, and H. Schurz). SIAM Review, 38 (1):177–178, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pruess:1975:ALF

[Pru75] Steven Pruess. The approximation of linear functionals and h^{2}-extrapolation. SIAM Review, 17(4):641– 651, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pruess:1978:AES

[Pru78] S. Pruess. Alternatives to exponential spline in tension. SIAM Review, 20(3): 633, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prusa:2000:HWR

[Pru00] Joseph M. Prusa. Hydrodynamics of a water rocket. SIAM Review, 42(4):719– 726, December 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/34822.

Pollak:1964:AE

[PS64] H. O. Pollak and L. Shepp. An asymptotic expansion. SIAM Review, 6(1):60, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Park:1974:DCI

[PS74] S. K. Park and T. A.

Straeter. Davidon’s cubic interpolation. SIAM Review, 16(1):93–94, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pierre:2000:BRD

[PS00] Michel Pierre and Didier Schmitt. Blowup in reactiondiffusion systems with dissipation of mass. SIAM Review, 42(1):93–106, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/35973.

Persson:2004:SMG

[PS04] Per-Olof O. Persson and Gilbert Strang. A simple mesh generator in MAT-LAB. SIAM Review, 46(2): 329–345, June 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/42912.

Pachter:2007:MP

[PS07] Lior Pachter and Bernd Sturmfels. The mathematics of phylogenomics. SIAM Review, 49(1):3–31, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pawlowski:2006:GTN

[PSSW06] Roger P. Pawlowski, John N. Shadid, Joseph P. Simonis, and Homer F. Walker. Globalization techniques for Newton–Krylov methods and applications to the fully coupled solution of the Navier– Stokes equations. SIAM Review, 48(4):700–721, January 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peletier:2012:CRL

[PSV12] Mark A. Peletier, Giuseppe Savaré, and Marco Veneroni. Chemical reactions as Γ-limit of diffusion. SIAM Review, 54 (2):327–352, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Parlett:1981:BRM

[PT81] B. Parlett and R. Tolimieri. Book review: A Matrix Eigenvalue Problem (G. Efroymson, A. Steger and S. Steinberg). SIAM Review, 23 (1):105, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Polik:2007:SL

[PT07] Imre Pólik and Tamás Terlaky. A survey of the Slemma. SIAM Review, 49(3): 371–418, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Palis:1996:HSC

[PTK96] J. Palis, F. Takens, and Roger L. Kraft. Hyperbolicity and sensitive chaotic dynamics at homoclinic bifurcations. SIAM Review, 38 (2):348–??, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Platte:2011:IFS

[PTK11] Rodrigo B. Platte, Lloyd N.

Trefethen, and Arno B. J. Kuijlaars. Impossibility of fast stable approximation of analytic functions from equispaced samples. SIAM Review, 53(2):308– 318, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i2/p308_s1.

Pugh:1966:GTA

[Pug66a] E. L. Pugh. A gradient technique of adaptive Monte Carlo. SIAM Review, 8(3): 346–355, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pugh:1966:RR

[Pug66b] E. L. Pugh. A recursion relation. SIAM Review, 8 (1):106, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pugliese:1986:BRD

[Pug86]  Andrea Pugliese.  Book

review: Deterministic Aspects of Mathematical Demography (John Impagliazzo). SIAM Review, 28(2): 277–278, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pullen:1964:BRA

[Pul64] J. Pullen. Book review: Another Coin Tossing Problem Donald (J. Persico and Henry C. Friedman). SIAM Review, 6(3):313–314, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pullin:1991:BRG

[Pul91]  Jorge Pullin.  Book re-

view: General Relativity. An Introduction to the Theory of the Gravitational Field (Hans Stephani; Martin Pollock and John Stewart, trans.; John Stewart, ed.). SIAM Review, 33(4):674– 675, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pereira:2023:AGS

[PvySF23] Rodrigo M. Pereira, Natacha Nguyen van yen, Kai Schneider, and Marie Farge. Are adaptive Galerkin schemes dissipative? SIAM Review, 65(4):1109–1134, November 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peters:1979:III

[PW79] G. Peters and J. H. Wilkinson. Inverse iteration, illconditioned equations, and Newton’s method. SIAM Review, 21(3):339–360, July 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pemantle:2008:TCE

[PW08] Robin Pemantle and Mark C.

Wilson. Twenty combinatorial examples of asymptotics derived from multivariate generating functions. SIAM Review, 50(2):199– 272, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pestana:2015:NPI

[PW15] Jennifer Pestana and Andrew J. Wathen. Natural preconditioning and iterative methods for saddle point systems. SIAM Review, 57(1): 71–91, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Peherstorfer:2018:SMM

[PWG18] Benjamin Peherstorfer, Karen Willcox, and Max Gunzburger. Survey of multifidelity methods in uncertainty propagation, inference, and optimization. SIAM Review, 60(3):550–591, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Petkovsek:1997:AB

[PWZK97] Marko Petkovsek, Herbert S. Wilf, Doron Zeilberger, and Wolfram Koepf. A = B. SIAM Review, 39(3):538– ??, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pakdemirli:1998:STP

[PY98] Mehmet Pakdemirli and Muhammet Yurusoy. Similarity transformations for partial differential equations. SIAM Review, 40(1):96– 101, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31001.

Palacian:2000:RPP

[PY00] Jesús Palacián and Patricia Yanguas. Reduction of polynomial planar Hamiltonians with quadratic unperturbed part. SIAM Review, 42 (4):671–691, December 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36232.

Pascual:1964:SLE

[PZ64] M. J. Pascual and J. E.

Zwetg. A set of linear equations. SIAM Review, 6 (1):61, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Pritchard:1981:SSI

[PZ81] A. J. Pritchard and J. Zabczyk.

Stability and stabilizability of infinite-dimensional systems. SIAM Review, 23(1): 25–52, January 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Quak:1995:BRM

[Qua95] Ewald Quak. Book review:

Mathematical Aspects of Geometric Modeling (Charles A. Micchelli). SIAM Review, 37 (4):636–637, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Quinn:1987:DID

[Qui87]  Dennis W. Quinn.  De-

termining the input for a dynamic system which is not completely controllable. SIAM Review, 29(4):621– 623, December 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Quinn:1992:CGP

[Qui92a] J. Quinn. A center of gravity perturbation. SIAM Review, 34(4):650–652, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Quinn:1992:BRC

[Qui92b] John Quinn. Book review: A Center of Gravity Perturbation (M. S. Klamkin). SIAM Review, 34(4):650– 652, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Quon:1975:OSC

[Quo75] C. Quon. Optimum simplification in computation of three-dimensional annulus waves. SIAM Review, 17 (2):387, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1963:BRP

[RA63] L. B. Rall and U. S. Army. Book review: Programming and Utilizing Digital Computers (R. S. Ledley). SIAM Review, 5(3): 295–296, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rabinowitz:1968:ALP

[Rab68a] Philip Rabinowitz. Applications of linear programming to numerical analysis. SIAM Review, 10(2): 121–159, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rabinowitz:1968:OSP

[Rab68b] Stanley Rabinowitz. Optimal search paths. SIAM Review, 10(3):376, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rabinowitz:1994:BRQ

[Rab94] Stanley Rabinowitz. Book review: A Quartic (H. Holloway and M. S. Klamkin). SIAM Review, 36(3):491– 493, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rafter:2002:MCM

[RAB02] John A. Rafter, Martha L.

Abell, and James P. Braselton. Multiple comparison methods for means. SIAM Review, 44(2):259– 278, June 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35723.

Radwin:1973:BRP

[Rad73] Mark S. Radwin. Book review: PL/1 for Scientific Programmers (C. T. Fike). SIAM Review, 15(1): 237–238, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rademacher:1990:PRI

[Rad90a] F. J. Rademacher. Probabilistic reasoning in intelligent systems: Networks of plausible inference (Judea Pearl). SIAM Review, 32(4): 704–??, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Radermacher:1990:BRP

[Rad90b] F. J. Radermacher. Book review: Probabilistic Reasoning in Intelligent Systems: Networks of Plausible Inference (Judea Pearl). SIAM Review, 32(4):704–707, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Raeside:1981:EME

[Rae81] D. E. Raeside. Estimation of the median effective dose in bioassay. SIAM Review, 23(3):391, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rousseau:1972:BRP

[RAG72] C. C. Rousseau, R. Askey, and G. Gasper. Book review: A Positive Integral (L. Carlitz). SIAM Review, 14(2): 372–378, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rahman:1987:IHF

[Rah87a] M. Rahman. An identity for a 4f3 hypergeometric function. SIAM Review, 29(1): 131–132, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rahman:1987:SSH

[Rah87b] M. Rahman. On the sign of a set of hypergeometric functions. SIAM Review, 29(1): 129–131, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1990:TDI

[RAH^{+}90] M. Renardy, M. D. Asic, D. Hewgill, P. Vandendriessche, A. A. Jagers, E. Loucks, G. Taubin, and P. J. Thomas. Two determinant inequalities. SIAM Review, 32(4): 681–682, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Raisbeck:1959:MLT

[Rai59] Gordon Raisbeck. Minimumloss two conductor transmission lines. SIAM Review, 1 (1):71, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1988:BRM

[RAK88] O. G. Ruehr, George E.

Andrews, and Louis W.

Kolitsch. Book review: A Multinomial Summation (Donald Richards and Stamatis Cambanis). SIAM Review, 30(1):128–130, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1962:BRE

[Ral62a] L. B. Rall. Book review: Elements of the Theory of Functions and Functional Analysis vol. I (A. N. Kolmogorov and S. V. Fomin); Metric and Normed Spaces vol. II (L. F. Boron, trans.); Measure — The Lebesgue Integral Hilbert Space (H. Kamel and H. Komm). SIAM Review, 4(3):266, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1962:BRT

[Ral62b] L. B. Rall. Book review: Theory of Linear Operators in Hilbert Space. Vol. I (N. I. Akhiezer and I. M. Glazman). SIAM Review, 4(2): 159, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1965:NIS

[Ral65a] L. B. Rall. Numerical integration and the solution of integral equations by the use of Riemann sums. SIAM Review, 7(1):55–64, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ralston:1965:RSN

[Ral65b] Anthony Ralston. Relative stability in the numerical solution of ordinary differential equations. SIAM Review, 7 (1):114–125, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1966:BRN

[Ral66a]  L. B. Rall.  Book re-

view: Nonlinear Mathematics (Thomas L. Saaty and Joseph Bram). SIAM Review, 8(1):117–118, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1966:BRQ

[Ral66b] L. B. Rall. Book review:

Quasilinearization and Nonlinear Boundary-Value Problems (R. E. Bellman and R. E. Kalaba). SIAM Review, 8 (3):401–402, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1969:USQ

[Ral69a] L. B. Rall. On the uniqueness of solutions of quadratic equations. SIAM Review, 11 (3):386–388, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1969:SLI

[Ral69b] L. B. Rall. Solution of linear integral equations by the use of invariant functions. SIAM Review, 11(3): 307–316, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1970:BRN

[Ral70]  L. B. Rall.  Book re-

view: Nonlinear Two Point Boundary Value Problems (Paul B. Bailey, Lawrence F. Shampine and Paul E. Waltman). SIAM Review, 12(1): 160–161, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1974:BRI

[Ral74]  L. B. Rall.  Book re-

view: Iterative Solution of Nonlinear Equations in Spneod Variables (J. M. Ortega and W. C. Rheinboldt). SIAM Review, 16(2):272– 274, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1980:BRV

[Ral80]  L. B. Rall.  Book re-

view: Variational Methods in Mathematics, Science and Engineering (Karel Rektorys). SIAM Review, 22(4): 528–532, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1983:BRCb

[Ral83a] L. B. Rall. Book review:

Computer Arithmetic in Theory and Practice (Ulrich W. Kulisch and Willard L. Miranker). SIAM Review, 25 (4):585–588, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rall:1983:BRCa

[Ral83b] L. B. Rall. Book review of “Computer Arithmetic In Theory and Practice”, by U. W. Kulisch and W. L. Miranker. SIAM Review, 25 (??):585–588, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ralston:1986:BRC

[Ral86] Anthony Ralston. Book review: The Carnegie-Mellon Curriculum for Undergraduate Computer Science (Mary Shaw, ed.). SIAM Review, 28 (1):120–123, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ralston:1993:BRM

[Ral93a] Anthony Ralston. Book review: Microcomputer Algorithms, Action from Algebra (John P. Killingbeck). SIAM Review, 35(3):533– 534, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ralston:1993:MAA

[Ral93b] Anthony Ralston. Microcomputer algorithms, action from algebra (John P Killingbeck). SIAM Review, 35(3): 533–??, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ramshaw:1978:NMM

[Ram78] J. D. Ramshaw. Numerical methods for multidimensional fluid-dynamics problems. SIAM Review, 20 (3):633, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ramaswami:1988:NME

[Ram88]  V. Ramaswami.  Nonlin-

ear matrix equations in applied probability — solution techniques and open problems. SIAM Review, 30(2): 256–263, June 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ramsey:1989:BBS

[Ram89] L. Thomas Ramsey. Best bounds for the stability of DARMA filters with constant coefficients. SIAM Review, 31 (3):365–400, September 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ramkrishna:1990:BRP

[Ram90a] D. Ramkrishna. Book review:

Process Modeling (Morton M. Denn). SIAM Review, 32(2): 329–330, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ramm:1990:AC

[Ram90b] A. G. Ramm. Author’s comment. SIAM Review, 32(1): 192–193, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ramm:1993:BRS

[Ram93] A. G. Ramm. Book review:

Short-Wavelength Diffraction Theory (V. M. Babic and V. S. Buldyrev). SIAM Review, 35(1):161–164, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ramer:1997:PI

[Ram97] A. Ramer. A projection inequality. SIAM Review, 39 (3):516–517, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rang:1976:TBI

[Ran76] E. R. Rang. Transmission of bands of infrared radiation. SIAM Review, 18 (4):823, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rand:1989:BRM

[Ran89] Richard H. Rand. Book review: Manifolds and Mechanics (Arthur Jones, Alistair Gray, and Robert Hutton). SIAM Review, 31(1): 157–158, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rangarajan:1991:BRV

[Ran91a] Anand Rangarajan. Book review: Visual Reconstruction (Andrew Blake and Andrew Zisserman). SIAM Review, 33(2):333–334, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rangarajan:1991:VRA

[Ran91b]  Anand Rangarajan.  Vi-

sual reconstruction (Andrew Blake and Andrew Zisserman). SIAM Review, 33 (2):333–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rand:1996:BRA

[Ran96a]  R. H. Rand.  Book re-

view: Applied Nonlinear Dynamics: Analytical, Computational and Experimental Methods (Ali H. Nayfeh and Balakumar Balachandran). SIAM Review, 38(2): 350–352, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rand:1996:BRP

[Ran96b]  R. H. Rand.  Book re-

view: Perturbation Methods in the Computer Age (David C. Wilcox). SIAM Review, 38 (3):539–540, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rao:1972:FT

[Rao72] D. K. Rao. A Fourier transform. SIAM Review, 14 (4):645, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rao:1982:EVT

[Rao82] K. A. Rao. On the expected value of two statistics. SIAM Review, 24(3):340, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Raoult:1994:BRE

[Rao94] Annie Raoult. Book review: Elliptic Differential Equations (W. Hackbusch). SIAM Review, 36(3):518– 519, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rapoport:1973:UMO

[Rap73]  Anatol Rapoport.  Uses

of mathematics outside the physical sciences. SIAM Review, 15(2):481–502, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ratiu:1987:BRS

[Rat87] Tudor Ratiu. Book review: Symmetries for Dynamical and Hamiltonian Systems (by H. M. M. ten Eikelder). SIAM Review, 29(1):165– 166, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ratiu:1992:BRC

[Rat92a] Tudor Ratiu. Book review: Convexity Methods in Hamiltonian Mechanics (Ivar Ekeland). SIAM Review, 34(2): 343–345, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ratiu:1992:CMH

[Rat92b]  Tudor Ratiu.  Convex-

ity methods in Hamiltonian mechanics (Ivar Ekeland). SIAM Review, 34 (2):343–??, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ratiu:1993:BRT

[Rat93]  Tudor Ratiu.  Book re-

view: Tensors and Manifolds (Robert H. Wasserman). SIAM Review, 35(2): 334–335, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rauch:1995:BRR

[Rau95] Jeffrey Rauch. Book review: Ray Methods for Nonlinear Waves in Fluids and Plasmas (A. M. Anile, J K Hunter, P. Pantano, and G. Russo). SIAM Review, 37(1): 136–137, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rawlings:1970:CSN

[Raw70] Robert R. Rawlings, Jr.

Concerning the solution of a nonlinear operator equation. SIAM Review, 12(1): 129–130, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rawlins:1983:TLP

[Raw83] A. D. Rawlins. Two Legendre polynomial identities. SIAM Review, 25(4):569, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rawlins:1986:RBF

[Raw86] Anthony D. Rawlins. On the roots of Bessel function equation. SIAM Review, 28(2): 232–233, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Raymon:1969:ACF

[Ray69] L. Raymon. Approximation of continuous functions on curves. SIAM Review, 11(2): 264–267, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Raz:1987:BRS

[Raz87] Tzvi Raz. Book review: System Reliability; Evaluation & Prediction in Engineering (Alain Pages and Michel Gondran). SIAM Review, 29 (4):673–674, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Raz:1992:ACS

[Raz92a] Tzvi Raz. The art of computer systems performance analysis: Techniques for experimental design, measurement, simulation, and modeling (Raj Jain). SIAM Review, 34(3):518–??, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Raz:1992:BRA

[Raz92b] Tzvi Raz. Book review: The Art of Computer Systems Performance Analysis: Techniques for Experimental Design, Measurement, Simulation, and Modeling (Raj Jain). SIAM Review, 34(3):

518–519, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1990:BRF

[RB90] D. G. Rogers and Peter Borwein. Book review: A Functional Equation (David W. Wilson). SIAM Review, 32 (4):684–686, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rahman:2023:ODV

[RB23] Aminur Rahman and D. Blackmore. The one-dimensional version of Peixoto’s structural stability theorem: a calculus-based proof. SIAM Review, 65(3):869–886, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://epubs.siam.org/
doi/10.1137/21M1426572.

Rogers:1990:FE

[RBB^{+}90] D. G. Rogers, P. Borwein, D. R. Breach, J. L. Buhite, A. G. Gibbs, C. C. Grosjean, H. Guggenheimer, R. Ingle, W. B. Jordan, D. A. Kurtze, G. N. Lewis, L. J. Mathis, F. Mathis, D. C. Nairn, M. Renardy, H. J. Seiffert, J. Vankan, and H. Wiener. A functional equation. SIAM Review, 32 (4):684–686, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rozier:2007:ROK

[RBC^{+}07] D. Rozier, F. Birol, E. Cosme, P. Brasseur, J. M. Brankart, and J. Verron. A reducedorder Kalman filter for data assimilation in physical oceanography. SIAM Review, 49(3):449–465, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruiz-Balet:2023:NOC

[RBZ23] Domènec Ruiz-Balet and Enrique Zuazua. Neural ODE control for classification, approximation, and transport. SIAM Review, 65(3):735–773, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1411433.

Richards:1987:MS

[RC87] Donald Richards and Stamatis Cambanis. A multinomial summation. SIAM Review, 29(1):121–122, March 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ribenboim:1995:CC

[RC95] P. Ribenboim and J. W. S. Cassels. Catalan’s conjecture. SIAM Review, 37(1): 116–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rainey:1971:BRS

[RCC71] J. W. Rainey, L. Carlitz, and R. C. Cavanagh. Book review: A Set of Linear Equations (Peter H. Astor). SIAM Review, 13(1):126– 130, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Raeside:1978:MAF

[RCC78] D. E. Raeside, W. K.

Chu, and P. A. N. Chandraratna. Medical application of Fourier analysis. SIAM Review, 20(4):850– 854, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roose:2007:MMA

[RCM07] Tiina Roose, S. Jonathan Chapman, and Philip K. Maini. Mathematical models of avascular tumor growth. SIAM Review, 49(2):179– 208, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rothman:1965:BRZ

[RD65] David Rothman and Kenneth S. Dams. Book review: On a Zero of an Implicit Function (M. A. Medick). SIAM Review, 7(1):

129–131, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reddy:1976:MHF

[Red76a] J. N. Reddy. Mixed-hybrid finite-element approximations of biharmonic equation. SIAM Review, 18 (4):823, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Redheffer:1976:DIC

[Red76b] Ray Redheffer. Differential inequalities and characteristic value problems. SIAM Review, 18(2):231– 239, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reddien:1980:PMT

[Red80] G. W. Reddien. Projection methods for two-point boundary value problems. SIAM Review, 22(2):156– 171, April 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Redheffer:1983:BRM

[Red83] R. M. Redheffer. Book review: Maximum Principles and Their Applications (René Sperb). SIAM Review, 25(1): 128–129, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reddy:1989:BRH

[Red89] J. N. Reddy. Book review:

Hamilton’s Principle in Continuum Mechanics (A. Bedford). SIAM Review, 31(3): 512–513, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reddy:1990:BRM

[Red90] B. D. Reddy. Book review: Maximum and Minimum Principles: a Unified Approach, with Applications. (M. J. Sewell). SIAM Review, 32(1):182–184, March 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reddy:1993:BRM

[Red93a] J. N. Reddy. Book review: Mechanics of Solids and Fluids (Franz Ziegler). SIAM Review, 35(1):164– 165, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reddy:1993:MSF

[Red93b] J. N. Reddy. Mechanics of solids and fluids (Franz Ziegler). SIAM Review, 35 (1):164–??, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ramani:2019:CFB

[REG19] Arjun S. Ramani, Nicole Eikmeier, and David F. Gleich. Coin-flipping, ball-dropping, and grass-hopping for generating random graphs from matrices of edge probabilities. SIAM Review, 61(3): 549–595, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rehm:1989:BRB

[Reh89] Ronald G. Rehm. Book review: Buoyancy-Induced Flows and Transport (Benjamin Gebhart, Yogesh Jaluria, Roop L. Mahajan, and Bahgat Sammakia). SIAM Review, 31(1):158– 159, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1959:FT

[Rei59a] Walter P. Reid. Finite transforms. SIAM Review, 1(1): 44–46, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1959:VSV

[Rei59b] Walter P. Reid. Vibrating string of varying length. SIAM Review, 1(2):158– 159, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1962:BRE

[Rei62a] H. E. Reinhardt. Book review: An Elementary Introduction to the Theory of Probability (B. V. Gnedenko and A. Ya. Khinchin). SIAM Review, 4(2):162, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1962:BRP

[Rei62b] H. E. Reinhardt. Book review: Probability and Experimental Errors in Science (L. G. Parratt). SIAM Review, 4 (2):163–164, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1963:BRP

[Rei63]  H. E. Reinhardt.  Book

review: Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability. Volume II: Probability Theory (J. Neyman ed.). SIAM Review, 5 (1):86, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reitman:1964:BRS

[Rei64] Walter R. Reitman. Book review: Self-Organizing Systems, 1962 (Marshall C. Yovits, George T. Jacobi, Gordon D. Goldstein, eds.). SIAM Review, 6(1):81–83,

????  1964.  CODEN

SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1966:TTR

[Rei66] Walter P. Reid. Transient temperatures in a 30^{∘} right triangle. SIAM Review, 8(3): 356–358, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1969:BRP

[Rei69]  H. E. Reinhardt.  Book

review: Probabilistic Approaches to Design (Edward B. Haugen). SIAM Review, 11(2):292–294, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1971:BRL

[Rei71a]  W. T. Reid.  Book re-

view: Lectures on the Calculus of Variations and Optimal Control Theory (L. C. Young). SIAM Review, 13 (2):231–233, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1971:IMM

[Rei71b]  William T. Reid.  Inte-

gral means and mollifiers. SIAM Review, 13(3):333– 349, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1971:BRC

[Rei71c] H. E. Reinhardt. Book review: The Chi-Squared Distribution (H. O. Lancaster). SIAM Review, 13(1):102– 103, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reiss:1971:MAE

[Rei71d] Edward L. Reiss. On multivariable asymptotic expansions. SIAM Review, 13(2): 189–196, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1975:BRN

[Rei75]  H. E. Reinhardt.  Book

review: Nonlinear Parameter Estimation (Yonathan Bard). SIAM Review, 17(4): 703–704, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1977:BRV

[Rei77a] Williams T. Reid. Book review: Variational Analysis: Critical Extremals and Sturmian Extensions (Marston Morse). SIAM Review, 19(1): 159–164, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1977:BRTa

[Rei77b]  H. E. Reinhardt.  Book

review: Theory of Probability, Vol. 1 (Bruno de Finetti). SIAM Review, 19 (2):339–341, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1977:BRTb

[Rei77c] Howard Reinhardt. Book review: Theoretical Statistics (D. R. Cox and D. V. Hinkley). SIAM Review, 19(4): 746–747, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1978:HNM

[Rei78a] William T. Reid. A historical note on the maximum principle. SIAM Review, 20 (3):580–582, July 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reifler:1978:OOC

[Rei78b]  F. Reifler.  Optimal or-

thogonal control law for maximizing terminal missile speed. SIAM Review, 20 (3):633, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1978:BRT

[Rei78c] Howard E. Reinhardt. Book review: Theory of Probability: a Critical Introductory Treatment, Vol. 2 (Bruno de Finetti). SIAM Review, 20 (1):200–201, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1979:BRM

[Rei79] Howard Reinhardt. Book review: Multivariate Statistical Inference (Narayan C. Giri). SIAM Review, 21(3): 423–424, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1983:BRR

[Rei83a] H. E. Reinhardt. Book review: Robust Statistics (Peter J. Huber). SIAM Review, 25(1):129, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1983:IDA

[Rei83b]  Howard Reinhardt.  Ion-

ization in diprotic acids. SIAM Review, 25(1):89– 91, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1987:BRS

[Rei87] Howard E. Reinhardt. Book review: Statistical Decision Theory and Bayesian Analysis. Second Edition (James O. Berger). SIAM Review, 29 (3):487–489, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1988:BRP

[Rei88] Howard E. Reinhardt. Book review: Probability and Statistical Inference (J. G. Kalbfleisch). SIAM Review, 30(2):347, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reinhardt:1989:BRI

[Rei89] Howard E. Reinhardt. Book review: Introduction to Statistical Inference (Jack Carl Kiefer; Gary Lorden, ed.). SIAM Review, 31(1):159– 160, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1996:BRS

[Rei96]  N. Reid.  Book review:

Saddlepoint Approximations (Jens Ledet Jensen). SIAM Review, 38(4):696–697, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1997:SEP

[Rei97a] R. M. Reid. Some eigenvalue properties of persymmetric matrices. SIAM Review, 39 (2):313–316, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reid:1997:CNS

[Rei97b] Russell M. Reid. Classroom note: Some eigenvalue properties of persymmetric matrices. SIAM Review, 39(2):313–316, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29480.

Rennie:1980:BRR

[Ren80a] B. C. Rennie. Book review: Resistances in an n-Dimensional Cube (D. Singmaster). SIAM Review, 22(4):504–508, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rennie:1980:DTE

[Ren80b]  B. C. Rennie.  Design-

ing a three-edged reamer. SIAM Review, 22(4):501– 502, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1984:LPM

[Ren84a] M. Renardy. Limit of a power of a matrix. SIAM Review, 26 (1):121–122, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1984:BRL

[Ren84b] Michael Renardy. Book review: Limit of a Power of a Matrix (Gengzhe Chang). SIAM Review, 26(1):121– 122, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1989:BRZ

[Ren89] Michael Renardy. Book review: Zero of Least Modulus (M. L. Glasser). SIAM Review, 31(1):126–127, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1990:BRT

[Ren90] Michael Renardy. Book review: Two Determinant Inequalities (Ralph Kelsey). SIAM Review, 32(4):681– 682, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1992:LSP

[Ren92a]  M. Renardy.  Linear-

system with positive solutions. SIAM Review, 34(3): 500–502, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1992:BRL

[Ren92b] Michael Renardy. Book review: Linear System with Positive Solutions (Peter Thejll). SIAM Review, 34(3): 500–502, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1993:AST

[Ren93a] M. Renardy. Asymptotic solution of the telegraph equation. SIAM Review, 35(2): 306–307, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1993:WGL

[Ren93b]  M. Renardy.  On win-

ning in the game of lotto. SIAM Review, 35(1):137– 139, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1993:BRA

[Ren93c]  Michael Renardy.  Book

review: Asymptotic Solution of the Telegraph Equation (Mark A. Pinsky). SIAM Review, 35(2):306– 307, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1993:BRW

[Ren93d] Michael Renardy. Book review: On Winning in the Game of Lotto (Andy Liu). SIAM Review, 35(1):137– 139, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1994:BRC

[Ren94a] Michael Renardy. Book review: A Convex Set (Marvin Marcus). SIAM Review, 36(1):111, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1994:BRI

[Ren94b] Michael Renardy. Book review: An Integral Relation for Successive Eigenvalues (Richard B. Evans). SIAM Review, 36(3):497, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1994:BRO

[Ren94c] Michael Renardy. Book review: An Operator Limit (W. Boehm). SIAM Review, 36 (4):659, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1994:BRM

[Ren94d]  Michael Renardy.  Book

review: Mathematical Topics in Fluid Mechanics (J. F. Rodrigues and A. Sequeira). SIAM Review, 36(1): 139–140, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1995:BRN

[Ren95]  Michael Renardy.  Book

review: Nonlinear Problems of Elasticity (Stuart S. Antman). SIAM Review, 37 (4):637, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1997:EUB

[Ren97a] M. Renardy. Existence and uniqueness for a boundary value problem. SIAM Review, 39(1):136–138, ???? 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1997:BRN

[Ren97b]  Michael Renardy.  Book

review: Nonlinear Waves in Elastic Media, by A. Kulikovskii and E. Sveshmikova. SIAM Review, 39 (2):358, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1998:I

[Ren98] M. Renardy. An inequality. SIAM Review, 40(3): 729–730, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reudink:1967:CTW

[Reu67] D. O. Reudink. Convolution transforms whose inversions have the same kernel. SIAM Review, 9(4): 721–725, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reza:1962:BDP

[Rez62] F. M. Reza. A bound for the derivative of positive real functions. SIAM Review, 4 (1):40–42, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rohatgi:1972:BRC

[RFC72] V. K. Rohatgi, J. S. Frame, and L. Carlitz. Book review: A Combinatorial Identity (V. K. Rohatgi). SIAM Review, 14(1):166–169, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Recht:2010:GMR

[RFP10] Benjamin Recht, Maryam Fazel, and Pablo A. Parrilo. Guaranteed minimum-rank solutions of linear matrix equations via nuclear norm minimization. SIAM Review, 52(3):471–501, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Remacle:2003:ADG

[RFS03] Jean-François Remacle, Joseph E. Flaherty, and Mark S. Shephard. An adaptive discontinuous Galerkin technique with an orthogonal basis applied to compressible flow problems. SIAM Review, 45(1):53–72, March 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38305.

Renardy:1990:NRT

[RGJ90] M. Renardy, C. C. Grosjean, and W. B. Jordan. On the number of roots of a transcendental equation. SIAM Review, 32(4): 682–683, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Russo:1976:SON

[RH76] P. M. Russo and J. N. Hanson. Symbolic optimization by Newton’s method. SIAM Review, 18(a4):824, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenbaum:1996:OS

[RH96] Paul R. Rosenbaum and Daniel F. Heitjan. Observational studies. SIAM Review, 38(3):529–??, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rheinboldt:1960:BRA

[Rhe60] Werner C. Rheinboldt. Book review: Asymptotic Methods in Analysis (N. G. de Bruijn). SIAM Review, 2(3): 227–229, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rheinboldt:1970:BRC

[Rhe70] Werner C. Rheinboldt. Book review: Computational Solution of Nonlinear Operator Equations (L. B. Rall). SIAM Review, 12(2):312– 313, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rheinboldt:1985:BRC

[Rhe85] Werner C. Rheinboldt. Book review: Computational Methods in Bifurcation Theory and Dissipative Structures (M. Kubícek and M. Marek). SIAM Review, 27(2):298– 299, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rheinboldt:1988:BRL

[Rhe88] Werner C. Rheinboldt. Book review: Lectures on Numerical Methods in Bifurcation Problems (Herbert B. Kelley). SIAM Review, 30(4): 677, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richards:1960:NSU

[Ric60] Paul I. Richards. Numerical stability and unsteady shock waves. SIAM Review, 2(3): 208–216, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richards:1961:PRC

[Ric61] Paul B. Richards. Plemelj relations for a class of Cauchy integrals. SIAM Review, 3 (3):252–255, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richards:1963:SNM

[Ric63] Paul I. Richards. Symmetry in numerical matrix calculations. SIAM Review, 5(3): 234–248, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rice:1967:BRT

[Ric67a] John R. Rice. Book review: Tchebycheff Systems: With Applications in Analysis and Statistics (Samuel Karlin and William J. Studden). SIAM Review, 9(2): 257–258, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richards:1967:GF

[Ric67b] P. I. Richards. A generating function. SIAM Review, 9 (4):748–749, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richards:1967:PCA

[Ric67c] Paul I. Richards. Precedence constraints and Arrow diagrams. SIAM Review, 9(3):

548–553, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rice:1974:BRC

[Ric74] John R. Rice. Book review: Complexity of Computer Computations (Raymond E. Miller and James W. Thatcher, eds.). SIAM Review, 16(3):407–409, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rice:1976:TIL

[Ric76a]  S. A. Rice.  Three in-

verse Laplace transforms. SIAM Review, 18(1):117– 118, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richter:1976:NSE

[Ric76b] G. R. Richter. Numerical solution of elliptic differential equations as integral equations of 1st kind. SIAM Review, 18(4):823, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richards:1985:I

[Ric85a] D. Richards. An integral on SO(3). SIAM Review, 27 (1):81–82, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richards:1985:BRI

[Ric85b] Donald Richards. Book review: An Integral on SO(3) (B. E. Eichinger). SIAM Review, 27(1):81–82, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richard:1988:BRB

[Ric88] Jean-François Richard. Book review: Bayesian Full Information Analysis of Simultaneous Equation Models Using Integration by Monte Carlo (Luc Bauwens). SIAM Review, 30(4):677– 679, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richardson:1989:IAP

[Ric89] S. Richardson. An identity arising in a problem of conformal mapping. SIAM Review, 31(3):484– 485, September 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richardson:1990:HTS

[Ric90a] S. Richardson. How Not to tackle some singular perturbation problems. SIAM Review, 32(3):471–473, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richter:1990:BRF

[Ric90b] Donald K. Richter. Book review: Foundations of the Theory of General Equilibrium (Yves Balasko). SIAM Review, 32(2):330–331, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richter:1990:FTG

[Ric90c] Donald K. Richter. Foundations of the theory of general equilibrium (Yves Balasko). SIAM Review, 32 (2):330–??, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richberg:1998:PFD

[Ric98] R. Richberg. A partition function difference related to chiral discrimination. SIAM Review, 40(2):389, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Riese:1972:BRD

[Rie72] J. W. Riese. Book review: A Double-Sum Identity (T. E. Phipps, Jr.). SIAM Review, 14(3):498–499, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Riemenschneider:1980:ZP

[Rie80] S. Riemenschneider. Zeros of a polynomial. SIAM Review, 22(3):363, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rienstra:1987:MBF

[Rie87] S. W. Rienstra. Monotonicity of Bessel functions. SIAM Review, 29(3):469, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Riley:1968:SI

[Ril68] James D. Riley. A set of inequalities. SIAM Review, 10(3):377, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Riley:1975:ULB

[Ril75] N. Riley. Unsteady laminar boundary layers. SIAM Review, 17(2):274–297, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rindler:1961:SRT

[Rin61] W. Rindler. Survey of Relativity theory. SIAM Review, 3(2):105–118, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rine:1974:DM

[Rin74] D. Rine. A discrete minimization. SIAM Review, 16 (3):391, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rivlin:1959:SI

[Riv59] T. J. Rivlin. Smooth interpolation. SIAM Review, 1 (1):60–63, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rivlin:1960:BRA

[Riv60a] T. J. Rivlin. Book review: Approximate Methods of Higher Analysis (L. V. Kantorovich and V. I. Krylov). SIAM Review, 2 (4):299, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rivlin:1960:NSI

[Riv60b] T. J. Rivlin. A note on smooth interpolation. SIAM Review, 2(1):27–30, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rivlin:1963:OSL

[Riv63] T. J. Rivlin. Overdetermined systems of linear equations. SIAM Review, 5(1): 52–66, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rivlin:1965:NVS

[Riv65] R. S. Rivlin. Nonlinear viscoelastic solids. SIAM Review, 7(3):323–340, ????

1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rizzo:1994:BRB

[Riz94] Frank Rizzo. Book review:

Boundary Integral and Singularity Methods for Linearized Viscous Flow (C. Pozrikidis). SIAM Review, 36(1):140–141, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1989:ZLM

[RJ89] M. Renardy and W. B. Jordan. Zero of least modulus. SIAM Review, 31(1): 126–127, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1990:BRP

[RJ90] Michael Renardy and A. A.

Jagers. Book review: Problem of the Nile (Leopold Flatto and Larry Shepp). SIAM Review, 32(2):302– 304, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1996:BRC

[RJ96] Michael Renardy and W. B.

Jordan. Book review: A Characterization of Uniformly Accelerated Motion (Murray S. Klamkin). SIAM Review, 38(3):525–526, ???? 1996. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Rinzel:1975:SIT

[RK75] J. Rinzel and J. B. Keller. Spatial instability of travellingwave solutions of a nerveconduction equation. SIAM Review, 17(2):387–388, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruhe:1978:NCM

[RK78] A. Ruhe and B. Kagstrom. Numerical computation of matrix functions. SIAM Review, 20(3):634, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rokicki:2014:DRC

[RKDD14] Tomas Rokicki, Herbert Kociemba, Morley Davidson, and John Dethridge. The diameter of the Rubik’s cube group is twenty. SIAM Review, 56(4):645– 670, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rodrigue:1976:HFB

[RKM76] G. Rodrigue, J. Karush, and N. Madsen. Having fun with band matrices on a vector processor. SIAM Review, 18 (4):823, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reza:1962:BRR

[RL62] Fazlollah Reza and Alfred Lehman. Book review: A Resistor Network Inequality (Alfred Lehman). SIAM Review, 4(2):150–155, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1981:DI

[RL81] O. G. Ruehr and O. P.

Lossers. A determinant and an identity. SIAM Review, 23 (1):105–107, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rodino:1995:LPD

[RL95] Luigi Rodino and Otto Liess. Linear partial differential operators in Gevrey spaces. SIAM Review, 37 (2):263–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1981:BRD

[RLG81] O. G. Ruehr, O. P. Lossers, and C. Givens. Book review: A Determinant and an Identity (A. V. Boyd). SIAM Review, 23(1):105– 107, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roanes-Lozano:2006:ACA

[RLGBRMC06] Eugenio Roanes-Lozano, Alfonso González-Bermejo, Eugenio Roanes-Macías, and Justo Cabezas. An application of computer algebra to pharmacokinetics: The Bateman equation. SIAM Review, 48(1):133–146, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richards:1965:NIL

[RLT65] P. I. Richards, W. D. Lanning, and M. D. Torrey. Numerical integration of large, highly-damped, nonlinear systems. SIAM Review, 7(3): 376–380, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1975:IPV

[RM75] E. H. Rogers and J. F. Menick. Intermediate problems and variational theory for nonlinear spectra. SIAM Review, 17(2):388, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roediger:1986:OCI

[RM86] P. A. Roediger and J. G. Mardo. An OC curve inequality. SIAM Review, 28 (4):568, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Regalia:1989:KPU

[RM89] Phillip A. Regalia and Sanjit K. Mitra. Kronecker products, unitary matrices and signal processing applications. SIAM Review, 31 (4):586–613, December 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rude:1995:MCT

[RM95] U. Rude and Steve Mc-Cormick. Mathematical computational techniques for multilevel adaptive methods. SIAM Review, 37(2): 266–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ross:1996:FLE

[RM96] Timothy Ross and Marialuisa N. McAllister. Fuzzy logic with engineering applications. SIAM Review, 38 (1):174–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rifkin:2003:AMM

[RMT^{+}03] Ryan Rifkin, Sayan Mukherjee, Pablo Tamayo, Sridhar Ramaswamy, Chen-Hsiang Yeang, Michael Angelo, Michael Reich, Tomaso Poggio, Eric S. Lander, Todd R. Golub, and Jill P. Mesirov. An analytical method for multiclass molecular cancer classification. SIAM Review, 45(4):706– 723, December 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/41198.

Richmond:1966:BRU

[RO66] Judy Richmond and W. Oettli. Book review: On the Uniqueness of a Probability Density Function (Perry Scheinok). SIAM Review, 8 (1):108–110, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roache:1978:DSV

[Roa78] P. J. Roache. Direct solution of variable coefficient elliptic equations in 3 dimensions. SIAM Review, 20(3): 633–634, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1962:BRC

[Rob62a] Enders A. Robinson. Book review: Characteristic Functions (E. Lukacs). SIAM Review, 4(3):264, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1962:BRP

[Rob62b] Enders A. Robinson. Book review: The Passage Problem for a Stationary Markcov Chain (J. H. B. Kemperman). SIAM Review, 4(3): 263–264, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1974:BRI

[Rob74] D. W. Robinson. Book review: Integral Representation for the Moore–Penrose Generalized Inverse of a Matrix (Allen D. Ziebur). SIAM Review, 16(1):94–96, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1976:BRI

[Rob76] D. W. Robinson. Book review: Integral Representation for the Moore–Penrose Generalized Inverse of a Matrix (A. D. Ziebur). SIAM Review, 18(3):492, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1980:CL

[Rob80] H. P. Robinson. A. Conjectured limit. SIAM Review, 22(2):229, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robbin:1981:BRF

[Rob81] Joel W. Robbin. Book review: Foundations of Mechanics (Ralph Abraham and Jerrold E. Marsden). SIAM Review, 23(1):120– 121, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robbin:1985:BRI

[Rob85] Joel W. Robbin. Book review: Introduction to Dynamics (Ian Percival and Derek Richards). SIAM Review, 27(1):121–123, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1991:BRC

[Rob91] Clark Robinson. Book review: Chaotic Dynamics of Nonlinear Systems (S. Neil Rasband). SIAM Review, 33 (2):334–335, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1992:BRC

[Rob92a]  Clark Robinson.  Book

review: Chaotic Behavior of Deterministic Dissipative Systems (Milos Marek and Igor Schreiber). SIAM Review, 34(4):680–681, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1992:CBD

[Rob92b] Clark Robinson. Chaotic behavior of deterministic dissipative systems (Milos Marek and Igor Schreiber). SIAM Review, 34(4):680– ??, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robertson:1994:BRC

[Rob94a] James B. Robertson. Book review: Continuous-Time Markov Chains (W. J. Anderson). SIAM Review, 36 (2):316–317, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1994:PVC

[Rob94b] F. N. H. Robinson. Poynting’s vector. comments on a recent paper by C. Jeffries: “A new conservation law for classical electrodynamics” [SIAM Rev. 34 (1992), no. 3, 386–405 and MR 93i:78005]. With a reply by Jeffries. SIAM Review, 36(4):633– 637, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25039.htm. See

[Jef92, Jef94].

Roberts:1995:RTD

[Rob95a] P. H. Roberts. On the ratio of two double integrals. SIAM Review, 37(2): 244–245, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roberts:1995:BRR

[Rob95b]  Paul H. Roberts.  Book

review: On the Ratio of Two Double Integrals (D. E. Loper). SIAM Review, 37(2): 244–245, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roberts:1997:I

[Rob97] A. J. Roberts. An integral. SIAM Review, 39(3): 524–527, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rochberg:1992:BRW

[Roc92a] Richard Rochberg. Book review: Wavelets and Singular Integrals on Curves and Surfaces (Guy David). SIAM Review, 34(4):681– 682, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rochberg:1992:WSI

[Roc92b] Richard Rochberg. Wavelets and singular integrals on curves and surfaces (Guy David). SIAM Review, 34(4): 681–??, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rockafellar:1993:LMO

[Roc93] R. Tyrrell Rockafellar. Lagrange multipliers and optimality. SIAM Review, 35(2): 183–238, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rodin:1981:NSO

[Rod81] E. Y. Rodin. A nonlinear second order differential equation. SIAM Review, 23 (4):524, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roediger:1995:CTI

[Roe95] Paul A. Roediger. A conditional trace inequality. SIAM Review, 37(2):236, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1960:BRA

[Rog60] Jack O. Rogers. Book review: Analyses of Industrial Operations (Edward H. Bowman and Robert B. Fetter). SIAM Review, 2(1): 55–56, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogge:1969:CTP

[Rog69a] Thomas Rogge. A coupled thermoelastic problem. SIAM Review, 11(2):277– 282, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogge:1969:MSW

[Rog69b] Thomas R. Rogge. A method of solving wave propagation problems in bounded anisotropic media. SIAM Review, 11(4):608–611, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1973:FEG

[Rog73] C. Rogers. A functional equation from gas dynamics. SIAM Review, 15(3): 655–657, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1975:IVP

[Rog75]  J. C. W. Rogers.  Ini-

tial value problem for water waves. SIAM Review, 17 (2):388, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1978:EVR

[Rog78] T. D. Rogers. Expected values for random regions of a circle. SIAM Review, 20 (3):594, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1983:LRP

[Rog83a] Joseph W. Rogers. Locations of roots of polynomials. SIAM Review, 25(3): 327–342, July 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1983:ISD

[Rog83b] T. D. Rogers. An infinite system of differential equations. SIAM Review, 25 (4):570, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1994:MPC

[Rog94a]  R. C. Rogers.  Mathe-

matical problems of classical nonlinear electromagnetic theory (Frederick Bloom). SIAM Review, 36(2):317– ??, June 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1994:BRM

[Rog94b] Robert C. Rogers. Book review: Mathematical Problems of Classical Nonlinear Electromagnetic Theory (Fredrick Bloom). SIAM Review, 36 (2):217–219, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rogers:1996:BRM

[Rog96] Robert C. Rogers. Book review: Mathematical Methods for Geo-Electromagnetic Induction (J. T. Weaver). SIAM Review, 38(4):697– 698, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rohde:1966:SRG

[Roh66] Charles A. Rohde. Some results on generalized inverses. SIAM Review, 8(2): 201–205, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rohatgi:1970:CI

[Roh70] V. K. Rohatgi. A combinatorial identity. SIAM Review, 12(4):580, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rohatgi:1971:CI

[Roh71] V. K. Rohatgi. A combinatorial identity. SIAM Review, 13(3):388, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ron:1995:BRB

[Ron95] Amos Ron. Book review: Box Splines (C. de Boor, K. Hollig, and S. Riemenschneider). SIAM Review, 37(2): 275–276, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roos:1964:NGD

[Roo64]  B. W. Roos.  Note on

generalized dynamical systems. SIAM Review, 6(3): 269–274, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roos:2010:NAM

[Roo10] Hans-Goerg Roos. Numerical approximation methods for elliptic boundary value problems: Finite and boundary elements. SIAM Review, 52 (4):786–787, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roppert:1983:BRB

[Rop83] J. Roppert. Book review:

A Binomial Coefficient Summation (Gengzhe Chang). SIAM Review, 25(4):575– 576, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosen:1961:SLE

[Ros61] B. W. Rosen. A set of linear equations. SIAM Review, 3(2):180, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenblatt:1963:BRS

[Ros63] Judah I. Rosenblatt. Book review: Statistical Theory (B. W. Lindgren). SIAM Review, 5(3):293–295, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenblatt:1965:BRM

[Ros65] Judah Rosenblatt. Book review: Mathematical Theory of Reliability. (Richard E.

Barlow and Frank Porschan). SIAM Review, 7(4):580– 582, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosser:1967:VRP

[Ros67a] J. B. Rosser. V. the research program in applied mathematics. SIAM Review, 9(2): 366–387, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosser:1967:RKA

[Ros67b]  J. Barkley Rosser.  A

Runge–Kutta for all seasons. SIAM Review, 9(3): 417–452, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenblatt:1968:BRS

[Ros68a] Judah Rosenblatt. Book review: Sequential Methods in Statistics (G. Barrie Wetherill). SIAM Review, 10(2): 236–237, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenblum:1968:BRC

[Ros68b] Marvin Rosenblum. Book review: Commutation Properties of Hilbert Space Operators and Related Topics (C. R. Putnam). SIAM Review, 10(3):391, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenberg:1969:RDG

[Ros69] Milton Rosenberg. Range decomposition and generalized inverse of nonnegative Hermitian matrices. SIAM Review, 11(4):568–571, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosen:1971:BRP

[Ros71]  Saul Rosen.  Book re-

view: Programming Languages: History and Fundamentals (Jean E. Sammet). SIAM Review, 13(1): 108–109, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenblatt:1972:BRT

[Ros72a] Judah Rosenblatt. Book review: The Theory of Statistical Inference (Shelemyahu Zacks). SIAM Review, 14(2): 361–362, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenthal:1972:BRS

[Ros72b] Robert W. Rosenthal. Book review: Some Topics in Two-Person Games (T. Parthasarathy and T. E. S. Raghavan). SIAM Review, 14 (2):356–357, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ross:1972:ISF

[Ros72c] D. K. Ross. Inequalities for special functions. SIAM Review, 14(3):494, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenbloom:1973:BRH

[Ros73] P. G. Rosenbloom. Book review: History of Mathematical Logic from Leibniz to Peano (N. I. Styazhkin). SIAM Review, 15(2):395– 396, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ross:1975:IGF

[Ros75] D. K. Ross. Inequalities for the gamma function. SIAM Review, 17(3):566, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rose:1976:ACN

[Ros76a] D. J. Rose. Aspects of combinatorial numerical analysis. SIAM Review, 18(4): 823–824, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosen:1976:AIG

[Ros76b] J. B. Rosen. Applications of an interactive graphical system to boundary-value problems and optimization. SIAM Review, 18(4):824, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenblatt:1978:BRH

[Ros78] Judah Rosenblatt. Book review: Handbook of Statistical Distributions (Jagdish K. Patel, C. H. Kapadia and D. B. Owen, eds.). SIAM Review, 20(4):868–869, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ross:1979:NLD

[Ros79] D. K. Ross. A non-linear differential equation. SIAM Review, 21(3):395, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenfeld:1980:BRS

[Ros80a]  Azriel Rosenfeld.  Book

review: Syntactic Pattern Recognition: an Introduction (Rafael C. Gonzalez and Michael G. Thomason). SIAM Review, 22(1):103– 104, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ross:1980:ETD

[Ros80b] D. K. Ross. Eigenvalues of a Tri-diagonal matrix. SIAM Review, 22(1):98, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ross:1980:NR

[Ros80c] D. K. Ross. A nonoscillation result. SIAM Review, 22(2):229, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rose:1982:BRL

[Ros82] Nicholas J. Rose. Book review: Linear Algebra and Its Applications (Gilbert Strang). SIAM Review, 24 (4):499–501, October 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenfeld:1985:BRM

[Ros85] Azriel Rosenfeld. Book review: The Measurement of Visual Motion (Ellen Catherine Hildreth). SIAM Review, 27(2):299–300, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rostamain:1987:BRH

[Ros87] Rouben Rostamain. Book review: Heat Conduction Within Linear Thermoelasticity (William Alan Day). SIAM Review, 29(2):335– 336, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rosenthal:1995:CRM

[Ros95] Jeffrey S. Rosenthal. Convergence rates for Markov chains. SIAM Review, 37 (3):387–405, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/26674.
htm.

Rossi:2011:Ea

[Ros11a]  Louis F. Rossi.  Educa-

tion. SIAM Review, 53 (1):155, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i1/p155_s1.

Rossi:2011:Eb

[Ros11b]  Louis F. Rossi.  Educa-

tion. SIAM Review, 53 (2):347, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i2/p347_s1.

Rossi:2011:Ec

[Ros11c]  Louis F. Rossi.  Educa-

tion. SIAM Review, 53 (3):559, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i3/p559_s1.

Rossi:2011:Ed

[Ros11d]  Louis F. Rossi.  Educa-

tion. SIAM Review, 53 (4):773, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v53/i4/p773_s1.

Rossi:2012:Ea

[Ros12a]  Louis F. Rossi.  Educa-

tion. SIAM Review, 54 (1):155, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v54/i1/p155_s1.

Rossi:2012:Eb

[Ros12b]  Louis F. Rossi.  Educa-

tion. SIAM Review, 54(2): 353, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2012:Ec

[Ros12c]  Louis F. Rossi.  Educa-

tion. SIAM Review, 54(3): 571, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2012:Ed

[Ros12d]  Louis F. Rossi.  Educa-

tion. SIAM Review, 54(4): 799, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2013:Ea

[Ros13a]  Louis F. Rossi.  Educa-

tion. SIAM Review, 55(1): 147, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2013:Eb

[Ros13b]  Louis F. Rossi.  Educa-

tion. SIAM Review, 55(3): 521, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2013:Ec

[Ros13c]  Louis F. Rossi.  Educa-

tion. SIAM Review, 55(4): 749–750, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2014:Ea

[Ros14a]  Louis F. Rossi.  Educa-

tion. SIAM Review, 56(1): 157–158, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2014:Eb

[Ros14b]  Louis F. Rossi.  Educa-

tion. SIAM Review, 56(2): 335, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2014:Ec

[Ros14c]  Louis F. Rossi.  Educa-

tion. SIAM Review, 56(3): 523, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2014:Ed

[Ros14d]  Louis F. Rossi.  Educa-

tion. SIAM Review, 56(4): 671–672, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2015:Ea

[Ros15a]  Louis F. Rossi.  Educa-

tion. SIAM Review, 57(1): 129, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2015:Eb

[Ros15b]  Louis F. Rossi.  Educa-

tion. SIAM Review, 57(2): 253–254, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2015:Ec

[Ros15c]  Louis F. Rossi.  Educa-

tion. SIAM Review, 57(3): 435–436, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2015:Ed

[Ros15d]  Louis F. Rossi.  Educa-

tion. SIAM Review, 57(4): 609, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2016:Ea

[Ros16a]  Louis F. Rossi.  Educa-

tion. SIAM Review, 58(1): 141, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2016:Eb

[Ros16b]  Louis F. Rossi.  Educa-

tion. SIAM Review, 58(2): 329, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2016:Ec

[Ros16c]  Louis F. Rossi.  Educa-

tion. SIAM Review, 58(3): 553–554, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rossi:2016:Ed

[Ros16d]  Louis F. Rossi.  Educa-

tion. SIAM Review, 58(4): 773, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rothman:1962:ADI

[Rot62] David Rothman. An approximation of a definite integral. SIAM Review, 4 (3):255, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rothblum:1981:ESM

[Rot81] Uriel G. Rothblum. Expansions of sums of matrix powers. SIAM Review, 23(2): 143–164, April 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rousseau:1972:BRI

[Rou72] C. C. Rousseau. Book review: An Inequality (M. S. Klamkin). SIAM Review, 14 (4):656–657, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rousseau:1973:BRN

[Rou73]  C. C. Rousseau.  Book

review: A Nonlinear Differential Equation (O. G. Ruehr). SIAM Review, 15(2): 384–386, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rouvray:1988:BRM

[Rou88] Dennis H. Rouvray. Book review: Mathematical Concepts in Organic Chemistry (Ivan Gutman and Oskar E. Polansky). SIAM Review, 30(2): 348–350, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rousseau:1991:BRM

[Rou91] C. C. Rousseau. Book review: A Multidimensional Integral (Gaoyuan Wei). SIAM Review, 33(3):481– 483, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roussel:1997:ACM

[Rou97a] M. R. Roussel. An analytic center manifold for a simple epidemiological model. SIAM Review, 39(1):106– 109, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roussel:1997:CNA

[Rou97b] Marc R. Roussel. Classroom note: An analytic center manifold for a simple epidemiological model. SIAM Review, 39(1):106– 109, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29303;   http:
//www.siam.org/journals/
sirev/sirev391.htm.

Row:1964:BRI

[Row64] Ronald V. Row. Book review: An Introduction to Waves, Rays and Radiation in Plasma Media (Julius J. Brandstatter). SIAM Review, 6(1):75–76, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rowland:1966:CCP

[Row66] John H. Rowland. A counting circuit problem. SIAM Review, 8(4):526, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rowland:1976:BRF

[Row76] John H. Rowland. Book review: Floating-Point Computation (Pat H. Sterbenz). SIAM Review, 18(1):138– 139, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rowland:1979:BRB

[Row79] John H. Rowland. Book review: Basic Numerical Mathematics, Vol. 2: Numerical Algebra (John Todd). SIAM Review, 21(3):404, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rowland:1986:BRC

[Row86] John H. Rowland. Book review: Computational Mathematics — an Introduction to Numerical Approximation (T. R. F. Nonweilerz). SIAM Review, 28(3): 437–438, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roxin:1976:BRO

[Rox76] Emilio Roxin. Book review:

Optimal Control Theory (L. D. Berkovitz). SIAM Review, 18(2):316, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roxin:1977:BRP

[Rox77] Emilio O. Roxin. Book review: Pursuit Games (Otomar Hajek). SIAM Review, 19(3):584–585, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roxin:1984:BRO

[Rox84] E. O. Roxin. Book review: Optimization Theory and Applications (Lamberto Cesari). SIAM Review, 26 (3):441–443, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roxin:1990:BRO

[Rox90a] Emilio O. Roxin. Book review: Optimal Periodic Control (Fritz Colonius). SIAM Review, 32(2):331– 332, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roxin:1990:OPC

[Rox90b] Emilio O. Roxin. Optimal periodic control (Fritz Colonius). SIAM Review, 32(2): 331–??, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roxin:1992:BRS

[Rox92] Emilio O. Roxin. Book review: Stability of Functional Differential Equations (V. B. Kolmanovskii and V. R. Nosov). SIAM Review, 34 (4):682–684, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roy:1991:BRI

[Roy91] Dilip N. Ghosh Roy. Book review: Inverse Problems in Quantum Scattering Theory (K. Chadan and P. C. Sabatier). SIAM Review, 33 (2):302–306, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roysam:1993:BRC

[Roy93a] Badrinath Roysam. Book review: Computational Vision (Harry Wechsler). SIAM Review, 35(1):165–166, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roysam:1993:CVH

[Roy93b] Badrinath Roysam. Computational vision (Harry Wechsler). SIAM Review, 35(1): 165–??, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roughan:1996:DTB

[RP96] Matthew Roughan and Kenneth Pope. The determinant of a triangular-block matrix. SIAM Review, 38(3): 513–514, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rombach:2017:CPS

[RPFM17] Puck Rombach, Mason A.

Porter, James H. Fowler, and Peter J. Mucha. Coreperiphery structure in networks (revisited). SIAM Review, 59(3):619–646, ???? 2017. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ronveaux:1988:DSB

[RPP88] A. Ronveaux, J. P. Palamin, and P. Palamin. Double sum of Bessel functions in computer music. SIAM Review, 30(4):652, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Richmond:1989:BRM

[RR89] Bruce Richmond and Cecil Rousseau. Book review: A Multinomial Summation (Donald Richards and Stamatis Cambanis). SIAM Review, 31(1):122– 125, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rousseau:1995:PS

[RR95] Cecil C. Rousseau and Otto G. Ruehr. Problems and solutions. SIAM Review, 37 (2):235–250, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rorres:1997:FCC

[RR97a] C. Rorres and D. G. Romano. Finding the center of a circular starting line in an ancient Greek stadium. SIAM Review, 39(4): 745–754, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rorres:1997:CNF

[RR97b] Chris Rorres and David Gilman Romano. Classroom note: Finding the center of a circular starting line in an ancient Greek stadium. SIAM Review, 39(4):745–754, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/30572.

Rousseau:1998:PSa

[RR98a] C. Rousseau and O. Ruehr.

Problems and solutions. SIAM Review, 40(1):118– 145, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97008.

Rousseau:1998:PSb

[RR98b] Cecil C. Rousseau and Otto G. Ruehr. Problems and solutions. SIAM Review, 40(2):372–390, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97010.

Rousseau:1997:PSa

[RRE97a] Cecil C. Rousseau, Otto G.

Ruehr, and Editors. Problems and solutions. SIAM Review, 39(1):123–141, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/97000.

Rousseau:1997:PSb

[RRE97b] Cecil C. Rousseau, Otto G.

Ruehr, and Editors. Problems and solutions. SIAM Review, 39(2):317–332, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/97002.

Rousseau:1997:PSc

[RRE97c] Cecil C. Rousseau, Otto G. Ruehr, and Editors. Problems and solutions. SIAM Review, 39(3):514–527, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

                                        http://epubs.siam.org/
                                        sam-bin/dbq/article/97004.

Rousseau:1997:PSd

[RRE97d] Cecil C. Rousseau, Otto G. Ruehr, and Editors. Problems and solutions. SIAM Review, 39(4):761–789, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97006.

Rousseau:1998:PSc

[RRe98] Cecil C. Rousseau, Otto G. Ruehr, and editors. Problems and solutions. SIAM Review, 40(3):710–731, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

                                        http://epubs.siam.org/
                                        sam-bin/dbq/article/97012.

Rubinstein:1995:PDE

[RRG95] I. Rubinstein, L. Rubinstein, and Ronald B. Guenther. Partial differential equations in classical mathematical physics. SIAM Review, 37 (1):123–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1995:IPD

[RRK95] M. Renardy, R. C. Rogers, and Philip Korman. An introduction to partial differential equations. SIAM Review, 37(3):469–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rubinstein:2007:DSD

[RRW07] Aviad Rubinstein, Jacob Rubinstein, and Gershon Wolansky. Determining sets for the discrete Laplacian. SIAM Review, 49(2):315– 324, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Russell:1971:EEI

[RS71] R. D. Russell and L. F.

Shampine. Existence of eigenvalues of integral equations. SIAM Review, 13(2): 209–219, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rubenfeld:1975:SMA

[RS75] L. A. Rubenfeld and W. L.

Siegmann. Some mathematical aspects of double-diffusive convection. SIAM Review, 17(2):388, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Raina:1979:CPL

[RS79] A. K. Raina and V. Singh. A conjectured property of Legendre functions. SIAM Review, 21(3):396, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Reidys:2002:CL

[RS02] Christian M. Reidys and Peter F. Stadler. Combinatorial landscapes. SIAM Review, 44 (1):3–54, March 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/39595.

Rosenstock:1965:BRD

[RSS65] Herbert B. Rosenstock, Perry Scheinok, and Sidney Spital. Book review: A Definite Integral (P. J. Short). SIAM Review, 7(4):565, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Riemenschneider:1981:ZP

[RSS81] S. Riemenschneider, A. Sharma, and P. W. Smith. Zeros of a polynomial. SIAM Review, 23(3):390, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Roos:1997:NMS

[RSTK97] H-G Roos, M. Stynes, L. Tobiska, and R. Bruce Kellogg.

Numerical methods for singularly perturbed differential equations. SIAM Review, 39 (3):535–??, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1973:BRR

[RSTW73] O. G. Ruehr, J. Swetits, A. Thyagaraja, and J. S. White. Book review: A Recurrence Relation (H. E. Fettis). SIAM Review, 15(1): 225–227, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Renardy:1997:SEU

[RT97] Michael Renardy and Robert E.

Terrell. Solutions: Existence and uniqueness for a boundary value problem (problem 96-2) (Ricardo Fazio). SIAM Review, 39(1):136– 138, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rubenfeld:1969:BRG

[Rub69]  Lester Rubenfeld.  Book

review: Generalized Functions and Partial Differential Equations (Georgi E. Shilov). SIAM Review, 11(4): 635–638, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rubin:1977:BRI

[Rub77a]  David S. Rubin.  Book

review: Integer Programming: Theory, Applications, and Computations (Hamdy A. Taha). SIAM Review, 19 (2):338–339, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rubinstein:1977:FON

[Rub77b] I. Rubinstein. A first order nonlinear differential equation. SIAM Review, 19(4): 736, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rubenfeld:1978:DET

[Rub78] L. A. Rubenfeld. On a derivative-expansion technique and some comments on multiple scaling in the asymptotic approximation of solutions of certain differential equations. SIAM Review, 20(1):79–105, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rudin:1961:BRT

[Rud61] Walter Rudin. Book review: Transcendental and Algebraic Numbers (A. D. Gelfond). SIAM Review, 3 (3):267, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rude:1997:BRS

[Rüd97]  Ulrich Rüde.  Book re-

view: The Splitting Extrapolation Method, by C. B. Liem and T. Lü and T. M. Shih. SIAM Review, 39(1): 161–162, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1972:BREa

[Rue72a] O. G. Ruehr. Book review:

An Expected Value (M. S. Klamkin). SIAM Review, 14 (2):378–381, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1972:NDE

[Rue72b] O. G. Ruehr. A nonlinear differential equation. SIAM Review, 14(2):365, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1972:BRA

[Rue72c] Otto G. Ruehr. Book review:

An Asymptotic Expansion (P. Barrucand). SIAM Review, 14(4):647–649, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1972:BREb

[Rue72d] Otto G. Ruehr. Book review: Expansion of an Integral (Robert E. Shafer).

SIAM Review, 14(4):652– 654, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1973:BRE

[Rue73a] O. G. Ruehr. Book review: Eigenvalues of a Reflected Tridiagonal Matrix (T. S. Chow). SIAM Review, 15(3): 672–673, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1973:BRF

[Rue73b] Otto G. Ruehr. Book review: A Family of Fredholm Integral Equations (J. L. Casti). SIAM Review, 15 (1):220–221, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1973:NDE

[Rue73c] Otto G. Ruehr. A nonlinear differential equation. SIAM Review, 15(2):377, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1974:BRS

[Rue74a] Otto G. Ruehr. Book review: A Summation Identity (H. M. Srivastava). SIAM Review, 16(2):260–261, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1974:NDE

[Rue74b] Otto G. Ruehr. A nonlinear differential equation. SIAM Review, 16(2):261– 262, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1975:NIE

[Rue75] O. G. Ruehr. A nonlinear integral equation. SIAM Review, 17(4):686, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1976:SHF

[Rue76a] O. G. Ruehr. Series of hypergeometric functions. SIAM Review, 18(4):767– 769, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1976:BRS

[Rue76b] Otto G. Ruehr. Book review: A Series of Hypergeometric Functions (H. M. Srivastava). SIAM Review, 18 (4):767–769, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1978:BRC

[Rue78] O. G. Ruehr. Book review:

Conjectured Inequalities (Peter Flor). SIAM Review, 20 (3):598–599, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1979:BRI

[Rue79] Otto G. Ruehr. Book review: An Integral of an Elliptic Function (A. V. Boyd). SIAM Review, 21(3):396– 397, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1981:BRI

[Rue81a] O. G. Ruehr. Book review: An Identity for Complex Numbers (M. S. Klamkin and A. Meir). SIAM Review, 23(3):395–396, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1981:IHF

[Rue81b]  O. G. Ruehr.  Identity

for hypergeometric functions. SIAM Review, 23 (3):392, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1982:II

[Rue82a] O. G. Ruehr. Inequality or identity. SIAM Review, 24 (4):481–482, October 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1982:BRI

[Rue82b] Otto G. Ruehr. Book review: Inequality or Identity (E. A. van Doorn). SIAM Review, 24(4):481–482, October 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1984:EHF

[Rue84] Otto G. Ruehr. Evaluation of a hypergeometric function. SIAM Review, 26(2): 272, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1989:ABM

[Rue89] O. G. Ruehr. Asymptotic behavior of a melting interface. SIAM Review, 31 (1):120, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1998:PHF

[Rue98] O. G. Ruehr. A property of harmonic functions. SIAM Review, 40(4):981– 983, November 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rukhin:1993:BEE

[Ruk93a] Andrew Rukhin. Bayesian estimation and experimental design in linear regression models (Jurgen Pilz). SIAM Review, 35(1):166– ??, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rukhin:1993:BRB

[Ruk93b] Andrew Rukhin. Book review: Bayesian Estimation and Experimental Design in Linear Regression Models (Jürgen Pilz). SIAM Review, 35(1):166–167, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rulla:1998:ATR

[Rul98] J. Rulla. Average temperature in a reaction-diffusion process. SIAM Review, 40(2): 382–385, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rump:1999:ICM

[Rum99] Siegfried M. Rump. Illconditioned matrices are componentwise near to singularity. SIAM Review, 41 (1):102–112, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32321.

Rundle:1991:BRP

[Run91] John B. Rundle. Book review: Principles of Earthquake Source Mechanics (B. V. Kostrov and Shamita Das). SIAM Review, 33(2): 335–337, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruohonen:1990:BRN

[Ruo90a] Matti Ruohonen. Book review: Non-Life Insurance Mathematics (Erwin Straub). SIAM Review, 32(1):184– 185, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruohonen:1990:NLI

[Ruo90b] Matti Ruohonen. Non-life insurance mathematics (Erwin Straub). SIAM Review, 32 (1):184–??, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruohonen:1991:ART

[Ruo91a] Matti Ruohonen. Aspects of risk theory (Jan Grandell). SIAM Review, 33(4):688– ??, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruohonen:1991:BRA

[Ruo91b] Matti Ruohonen. Book review: Aspects of Risk Theory (Jan Grandell). SIAM Review, 33(4):688–689, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruohonen:1992:BRE

[Ruo92a] Matti Ruohonen. Book review: Engineering Risks: Evaluation and Valuation (Ulrich Hauptmanns and Wolfgang Werner). SIAM Review, 34(2):345–346, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruohonen:1992:ERE

[Ruo92b]  Matti Ruohonen.  En-

gineering risks: Evaluation and valuation (Ulrich Hauptmanns and Wolfgang Werner). SIAM Review, 34 (2):345–??, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Russell:1960:BRE

[Rus60] J. P. Russell. Book review: Elementary Matrix Algebra (Franz E. Hohn). SIAM Review, 2(2):160– 161, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rust:1976:CIE

[Rus76] B. W. Rust. Constrained interval estimation for ill-posed problems. SIAM Review, 18 (4):824–825, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Russell:1978:CST

[Rus78]  David L. Russell.  Con-

trollability and stabilizability theory for linear partial differential equations: Recent progress and open questions. SIAM Review, 20(4):

639–739, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rust:1985:BRC

[Rus85] Bert W. Rust. Book review: Conjugate Duality and the Exponential Fourier Spectrum (Wray Britton). SIAM Review, 27(4):602– 605, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Russell:1990:BRC

[Rus90] R. D. Russell. Book review: Continuous Decoupling Transformations for Linear Boundary Value Problems (P. M. van Loon). SIAM Review, 32(4):707– 708, December 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Russel:1995:BRR

[Rus95] D. L. Russel. Book review:

Representation and Control of Infinite Dimensional Systems, Vols. 1 and 2 (A. Bensonssan, G. Da Prato, M. Delfour, and S. Mitter). SIAM Review, 37(3): 476–477, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rago:1978:BRM

[RV78] Paul J. Rago and George M. Van Dyne. Book review: Mathematical Bioeconomics: The Optimal Management of Renewable Resources (Colin W. Clark). SIAM Review, 20(4):865–867, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Riman:2021:DBO

[RVBE21] Nour Riman, Jonathan D. Victor, Sebastian D. Boie, and Bard Ermentrout. The dynamics of bilateral olfactory search and navigation. SIAM Review, 63(1): 100–120, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rice:1964:NSE

[RW64] John R. Rice and John S. White. Norms for smoothing and estimation. SIAM Review, 6(3):243–256, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruhe:1980:ASN

[RW80] Axel Ruhe and Per Åke Wedin. Algorithms for separable nonlinear least squares problems. SIAM Review, 22 (3):318–337, July 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Redner:1984:MDM

[RW84] Richard A. Redner and Homer F. Walker. Mixture densities, maximum likelihood and the EM algorithm. SIAM Review, 26(2): 195–239, April 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Robinson:1995:DSS

[RW95] Clark Robinson and Russell Walker. Dynamical systems: Stability, dynamics, and chaos. SIAM Review, 37 (4):639–??, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rude:2018:REC

[RWMD18] Ulrich Rüde, Karen Willcox, Lois Curfman McInnes, and Hans De Sterck. Research and education in computational science and engineering. SIAM Review, 60(3): 707–754, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ruehr:1988:ILT

[RY88] O. G. Ruehr and W. R.

Young. An inverse Laplace transform. SIAM Review, 30 (4):652, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ryan:1963:BRC

[Rya63]  R. Ryan.  Book review:

Convex Functions and Orlicz Spaces (M. A. Krasnoesl’skii and Ya. B. Rutickii). SIAM Review, 5(3):290–291, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saaty:1975:OGN

[SA75] Thomas L. Saaty and Joyce M. Alexander. Optimization and the geometry of numbers: Packing and covering. SIAM Review, 17(3): 475–519, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sibul:1978:SSN

[SA78] L. H. Sibul and G. Adomian. Symmetrized solutions for non-linear stochastic differential equations. SIAM Review, 20(3):634, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saaty:1968:BRT

[Saa68]  T. L. Saaty.  Book re-

view: The Theory of Max-Min (John M. Danskin). SIAM Review, 10(3):387– 388, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saari:1973:BRL

[Saa73] Donald G. Saari. Book review: Lectures on Celestial Mechanics (C. L. Siegel and J. K. Moser). SIAM Review, 15(3):675–676, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saad:1983:BRS

[Saa83] Youcef Saad. Book review:

The Symmetric Eigenvalue Problem (Beresford N. Parlett). SIAM Review, 25(2): 286–287, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saari:1988:BRF

[Saa88] Donald G. Saari. Book review: Fair Allocation (H. Peyton Young). SIAM Review, 30(2):350–352, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saari:1995:CEA

[Saa95] Donald G. Saari. A chaotic exploration of aggregation paradoxes. SIAM Review, 37 (1):37–52, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25849.htm.

Saari:1996:BRG

[Saa96] Donald G. Saari. Book review: Game Theory and Strategy (Philip D. Straffin); Mathematics and Politics (Alan D. Taylor). SIAM Review, 38(4):698–700, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sabharwal:1968:SSB

[Sab68] Y. P. Sabharwal. A sequence of symmetric Bernoulli trials. SIAM Review, 10(1): 108–109, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shi:2019:DSN

[SAB19] Guodong Shi, Claudio Altafini, and John S. Baras. Dynamics over signed networks. SIAM Review, 61(2):229– 257, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sackman:1967:UTN

[Sac67]  Jerome L. Sackman.  A

uniqueness theorem for a nonlinear, steadily creeping body. SIAM Review, 9(4): 741–743, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sacker:1974:BRS

[Sac74] Robert J. Sacker. Book review: Stability Theory of Dynamical Systems (N. P. Bhatia and G. P. Szegő).

SIAM Review, 16(1):106– 107, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sachs:1983:OCD

[Sac83a] E. W. Sachs. Optimal control of distributed parameter systems — Ahmed, N. U., Teo, K. L. SIAM Review, 25(3): 441–442, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sachs:1983:BRO

[Sac83b] Ekkehard W. Sachs. Book review: Optimal Control of Distributed Parameter Systems (N. U. Ahmed and K. L. Teo). SIAM Review, 25 (3):441–442, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saenz:1963:BRE

[Sae63] A. W. Saenz. Book review: On the Eigenvalues of a Matrix (Carl E. Sealander). SIAM Review, 5(1): 74–75, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saeks:1970:CHS

[Sae70] R. Saeks. Causality in Hilbert space. SIAM Review, 12(3): 357–383, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saeks:1973:FAH

[Sae73a] R. Saeks. Fourier analysis in Hilbert spaces. SIAM Review, 15(3):604–638, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saeks:1973:SHS

[Sae73b] R. Saeks. State in Hilbert space. SIAM Review, 15(2): 283–308, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saffman:1996:BRF

[Saf96] P. G. Saffman. Book review: Fluid Dynamics for Physicists (T. E. Faber). SIAM Review, 38(4):700– 702, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sagan:1977:BRS

[Sag77] Hans Sagan. Book review: Singular Optimal Control Problems (D. J. Bell and D. H. Jacobson). SIAM Review, 19(3):575–576, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sagan:1978:BRO

[Sag78]  Hans Sagan.  Book re-

view: Optimization (Lucien W. Neustadt). SIAM Review, 20(2):405, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sagan:1981:BRM

[Sag81]  Hans Sagan.  Book re-

view: Mathematics of Finite-Dimensional Control Systems: Theory and Design (David L. Russell). SIAM Review, 23(1):115– 116, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sagan:1982:BRH

[Sag82]  Hans Sagan.  Book re-

view: A History of the Calculus of Variations from the 17th Through the 19th Century (Herman H. Goldstine). SIAM Review, 24(4): 494–495, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sagan:1983:BRI

[Sag83]  Hans Sagan.  Book re-

view: An Introduction to Applied Optimal Control (Greg Knowles). SIAM Review, 25 (2):284–286, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sagan:1986:BRB

[Sag86] Hans Sagan. Book review:

Boundary Methods: an Algebraic Theory (Ismael Herrera). SIAM Review, 28(2): 278–279, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sahay:1995:SHR

[Sah95] Pratap N. Sahay. Spherical harmonics representation of an inhomogeneous plane wave. SIAM Review, 37 (3):436–438, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

                                        //epubs.siam.org/27320.
                                        htm.

Sharir:1996:DSS

[SAH96] Micha Sharir, Pankaj K. Agarwal, and Peter Hajnal. Davenport–Schinzel sequences and their geometric applications. SIAM Review, 38(4):689–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saigal:1991:BRN

[Sai91a] Romesh Saigal. Book review: Numerical Continuation Methods, An Introduction (Eugene Allgower and Kurt Georg). SIAM Review, 33(4):675, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saigal:1991:NCM

[Sai91b] Romesh Saigal. Numerical continuation methods, an introduction (Eugene Allgower and Kurt Georg). SIAM Review, 33(4):675– ??, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saltzer:1962:BRP

[Sal62] Charles Saltzer. Book review: Progress in Operations Research, Volume I (R. L. Ackoff, ed.). SIAM Review, 4(1):52, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saltzer:1963:BRB

[Sal63a] Charles Saltzer. Book review: A Brief Introduction to Theta Functions (R. Bellman). SIAM Review, 5(2): 162, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saltzer:1963:BRI

[Sal63b] Charles Saltzer. Book review: Information Retrieval and Machine Translation. (Vol. III, Part 2, of Advances in Documentation and Library Science.) (A. Kent ed.). SIAM Review, 5(1):88, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saltzer:1963:BRM

[Sal63c] Charles Saltzer. Book review: Modern Computing Methods — National Physical Laboratory Notes on Applied Science No. 16. (E. T. Goodwin). SIAM Review, 5(2): 172, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saltzer:1963:BRN

[Sal63d] Charles Saltzer. Book review: Numerical Weather Analysis and Prediction (P. D. Thompson). SIAM Review, 5(1):87–88, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Salton:1964:BRI

[Sal64] Gerard Salton. Book review: Information Storage and Retrieval: Tools, Elements, Theories (Joseph Becker and Robert M. Hayes). SIAM Review, 6(1):70–71, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Salton:1966:BRI

[Sal66] Gerard Salton. Book review: Information Science (Manfred Kochen). SIAM Review, 8(2):242–243, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Salinetti:1980:A

[Sal80] G. Salinetti. Addition. SIAM Review, 22(1):86, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sallee:1985:BRI

[Sal85] G. T. Sallee. Book review:

An Introduction to Convex Polytopes (Arne Brøndsted). SIAM Review, 27(1):123– 124, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Salmon:1987:SCF

[Sal87] Vincent Salmon. Safe car following. SIAM Review, 29 (3):470, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Samuelson:1973:MSP

[Sam73] Paul A. Samuelson. Mathematics of speculative price. SIAM Review, 15(1):1–42, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sampson:1976:BRAa

[Sam76a] Jeffrey R. Sampson. Book review: Adaptation in Natural and Artificial Systems (John H. Holland). SIAM Review, 18(3):529–530, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sampson:1976:BRAb

[Sam76b] Jeffrey R. Sampson. Book review: Artificial Intelligence (Earl M. Hunt). SIAM Review, 18(4):784–786, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Samuels:1976:BRA

[Sam76c] S. M. Samuels. Book review: Almost Sure Convergence (William F. Stout). SIAM Review, 18(2):317– 318, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sampson:1979:BRT

[Sam79] Jeffrey R. Sampson. Book review: Theory of Modelling and Simulation (Bernard P. Zeigler). SIAM Review, 21 (1):150–153, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Samuels:1980:BRP

[Sam80] Stephen M. Samuels. Book review: Probability Theory: Independence, Interchangeability, Martingales (Yuan Shih Chow and Henry Teicher). SIAM Review, 22(4): 522–523, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Samarov:1986:BRN

[Sam86] Alexander Samarov. Book review: Nonparametric Functional Estimation (B. L. S. Prakasa Rao). SIAM Review, 28(1):123–125, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sangren:1960:ITD

[San60] Ward C. Sangren. Interfaces in two dimensions. SIAM Review, 2(3):192– 199, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sandberg:1965:NAC

[San65] I. W. Sandberg. A note on the application of the contraction-mapping fixed point theorem to a class of nonlinear functional equations. SIAM Review, 7(2): 199–204, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santalo:1972:MI

[San72] L. A. Santalo. A multidimensional integral. SIAM Review, 14(1):164, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sangiovannivincentelli:1976:BGN

[San76] A. Sangiovannivincentelli.

Bipartite graphs and nonsymmetric permutations on sparse matrices. SIAM Review, 18(4):825–826, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanchez:1977:PH

[Sán77] David A. Sánchez. Populations and harvesting. SIAM Review, 19(3):551– 553, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanchez:1982:BRS

[San82] David A. Sanchez. Book review: Stability of Linear Systems (C. J. Harris and J. F. Miles). SIAM Review, 24 (2):231, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanchez:1983:MMP

[San83a] D. A. Sanchez. Mathematical models of population biology — Hoppensteadt, F. F. SIAM Review, 25(3): 421–422, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanchez:1983:BRM

[Sán83b] David A. Sánchez. Book review: Mathematical Models of Population Biology (Frank F. Hoppensteadt). SIAM Review, 25(3):421– 422, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saniee:1987:NOW

[San87] Iraj Saniee. A note on optimal waiting strategies at bus stops. SIAM Review, 29 (4):625–629, December 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sands:1993:MV

[San93] Bill Sands. A maximum volume. SIAM Review, 35 (2):298, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sandefur:1995:BRF

[San95] James T. Sandefur. Book review: A First Course in Discrete Dynamical Systems (R. A. Holmgren). SIAM Review, 37(1):137–138, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2006:SRa

[San06a] Fadil Santosa. Survey and review. SIAM Review, 48 (1):1, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2006:SRb

[San06b] Fadil Santosa. Survey and review. SIAM Review, 48 (2):205, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2006:SRc

[San06c] Fadil Santosa. Survey and review. SIAM Review, 48 (3):437, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2006:SRd

[San06d] Fadil Santosa. Survey and review. SIAM Review, 48 (4):627, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2007:SRa

[San07a] Fadil Santosa. Survey and review. SIAM Review, 49 (1):1, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2007:SRb

[San07b] Fadil Santosa. Survey and review. SIAM Review, 49 (2):177, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2007:SRc

[San07c] Fadil Santosa. Survey and review. SIAM Review, 49 (3):ix, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2007:SRd

[San07d] Fadil Santosa. Survey and review. SIAM Review, 49 (3):369, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2007:SRe

[San07e] Fadil Santosa. Survey and review. SIAM Review, 49 (4):543, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2008:SRa

[San08a] Fadil Santosa. Survey and review. SIAM Review, 50 (1):1, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2008:SRb

[San08b] Fadil Santosa. Survey and review. SIAM Review, 50 (4):627, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2008:SRIa

[San08c] Fadil Santosa. Survey and review — introduction. SIAM Review, 50(2):197, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2008:SRIb

[San08d] Fadil Santosa. Survey and review — introduction. SIAM Review, 50(3):411, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2009:SRa

[San09a] Fadil Santosa. Survey and review. SIAM Review, 51 (1):1, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2009:SRb

[San09b] Fadil Santosa. Survey and review. SIAM Review, 51 (2):237, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2009:SRc

[San09c] Fadil Santosa. Survey and review. SIAM Review, 51 (3):453, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2009:SRI

[San09d] Fadil Santosa. Survey and review: Introduction. SIAM Review, 51(4):659, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2010:SRa

[San10a] Fadil Santosa. Survey and review. SIAM Review, 52 (1):1, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2010:SRb

[San10b] Fadil Santosa. Survey and review. SIAM Review, 52 (2):219, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2010:SRc

[San10c] Fadil Santosa. Survey and review. SIAM Review, 52 (3):401, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2010:SRd

[San10d] Fadil Santosa. Survey and review. SIAM Review, 52 (4):591, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Santosa:2011:SRa

[San11a] Fadil Santosa. Survey and review. SIAM Review, 53 (1):1, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i1/p1_s1.

Santosa:2011:SRb

[San11b] Fadil Santosa. Survey and review. SIAM Review, 53 (2):215, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v53/i2/p215_s1.

Santosa:2011:SRc

[San11c] Fadil Santosa. Survey and review. SIAM Review, 53 (3):407, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v53/i3/p407_s1.

Santosa:2011:SRd

[San11d] Fadil Santosa. Survey and review. SIAM Review, 53 (4):605, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v53/i4/p605_s1.

Sander:2014:RS

[San14] Evelyn Sander. Research spotlights. SIAM Review, 56 (4):623, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2015:RSa

[San15a] Evelyn Sander. Research spotlights. SIAM Review, 57 (1):59, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2015:RSb

[San15b] Evelyn Sander. Research spotlights. SIAM Review, 57 (2):199, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2015:RSc

[San15c] Evelyn Sander. Research spotlights. SIAM Review, 57 (3):365, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2015:RSd

[San15d] Evelyn Sander. Research spotlights. SIAM Review, 57 (4):533, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2016:RSa

[San16a] Evelyn Sander. Research spotlights. SIAM Review, 58 (1):67, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2016:RSb

[San16b] Evelyn Sander. Research spotlights. SIAM Review, 58 (2):227, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2016:RSc

[San16c] Evelyn Sander. Research spotlights. SIAM Review, 58 (3):443, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2016:RSd

[San16d] Evelyn Sander. Research spotlights. SIAM Review, 58 (4):651, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2017:RSa

[San17a] Evelyn Sander. Research spotlights. SIAM Review, 59 (1):63, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2017:RSb

[San17b] Evelyn Sander. Research spotlights. SIAM Review, 59 (2):293, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2017:RSc

[San17c] Evelyn Sander. Research spotlights. SIAM Review, 59 (3):545, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sander:2017:RSd

[San17d] Evelyn Sander. Research spotlights. SIAM Review, 59 (4):767, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saran:1971:IS

[Sar71] S. Saran. An infinite summation. SIAM Review, 13 (4):570, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sarason:1984:HBV

[Sar84a]  L. Sarason.  Hyperbolic

boundary-value problems — Sakamoto, R. SIAM Review, 26(4):608–609, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sarason:1984:BRH

[Sar84b]  Leonard Sarason.  Book

review: Hyperbolic Boundary Value Problems (Reiko Sakamoto). SIAM Review, 26 (4):608–609, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sarlet:2010:AM

[Sar10] Willy Sarlet. Analytical mechanics. SIAM Review, 52(4): 787–790, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sattinger:1978:BRH

[Sat78] D. H. Sattinger. Book review: The Hopf Bifurcation and Its Applications (J. E. Marsden and M. Mc-Cracken). SIAM Review, 20 (2):410–412, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strang:2022:NHQ

[SAT22] Alexander Strang, Karen C. Abbott, and Peter J. Thomas. The network HHD: Quantifying cyclic competition in trait-performance models of tournaments. SIAM Review, 64(2):360–391, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1321012.

Saunders:1962:BRT

[Sau62] Sam C. Saunders. Book review: Tables of the Hypergeometric Probability Distribution (G. J. Liberman and D. B. Owen). SIAM Review, 4(1):54, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saunders:1976:NOS

[Sau76] M. Saunders. Nonlinear optimization subject to linearprogramming constraints. SIAM Review, 18(4):825– 826, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saunders:1978:BRP

[Sau78] I. W. Saunders. Book review: Percentiles for the Gamma Distribution (R. A. Waller and M. S. Waterman). SIAM Review, 20 (4):856, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saunders:1979:BRM

[Sau79] Sam Saunders. Book review: Mathematical Models for the Study of Reliability of Systems (A. Kaufman, D. Grouchko and R. Groun). SIAM Review, 21 (1):148, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Savage:1997:SCG

[Sav97]  Carla Savage.  A sur-

vey of combinatorial Gray codes. SIAM Review, 39 (4):605–629, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29527.  See

note [Naa01].

Savas:2018:BRC

[Sav18] Berkant Savas. Book review:

Computational and Statistical Methods for Analysing Big Data with Applications, by Shen Liu, James Mc-Gree, Zongyuan Ge, and Yang Xie. SIAM Review, 60 (2):485–487, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sawyer:1959:PAM

[Saw59] W. Warwick Sawyer. Pressures on American mathematics teachers. SIAM Review, 1(1):32–37, ???? 1959. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sawyer:1982:BRM

[Saw82] Stanley Sawyer. Book review: Mathematical Population Genetics (Warren J. Ewens). SIAM Review, 24(4): 491–492, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sawyer:1985:BRS

[Saw85] Stanley Sawyer. Book review: Stochastic Transport in Discrete Biological Systems (Eckhart Frehland). SIAM Review, 27(1):124– 125, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saylor:1979:BRF

[Say79] Paul E. Saylor. Book review: Fundamentals of Numerical Reservoir Simulation (Donald W. Peaceman). SIAM Review, 21(3):404– 406, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sayas:2013:VJN

[Say13] Francisco-Javier Sayas. The validity of Johnson–Nédélec’s BEM–FEM coupling on polygonal interfaces. SIAM Review, 55(1):131–146, ???? 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stallings:1972:CPM

[SB72] W. T. Stallings and T. L.

Boullion. Computation of pseudoinverse matrices using residue arithmetic. SIAM Review, 14(1):152–163, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanchez:1998:CCL

[SB98a] Angel Sánchez and A. R.

Bishop. Collective coordinates and length-scale competition in spatially inhomogeneous soliton-bearing equations. SIAM Review, 40 (3):579–615, September 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31741.

Sauren:1998:DQM

[SB98b] M. Sauren and H. M. Bücker. On deriving the quasiminimal residual method. SIAM Review, 40(4):922– 926, December 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/31936.

Smith:1998:DDP

[SBGP98] Barry F. Smith, Petter E. Bjorstad, William D. Gropp, and Joseph E. Pasciak. Domain decomposition: Parallel multilevel methods for elliptic partial differential equations. SIAM Review, 40(1): 169–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schaub:2020:RWS

[SBH^{+}20] Michael T. Schaub, Austin R. Benson, Paul Horn, Gabor Lippner, and Ali Jadbabaie. Random walks on simplicial complexes and the normalized Hodge 1-Laplacian. SIAM Review, 62(2):353– 391, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spantini:2022:CTN

[SBM22] Alessio Spantini, Ricardo Baptista, and Youssef Marzouk. Coupling techniques for nonlinear ensemble filtering. SIAM Review, 64 (4):921–953, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1312204.

Schenk:2008:LSD

[SBR08] Olaf Schenk, Matthias Bollhöfer, and Rudolf A. Römer. On large-scale diagonalization techniques for the Anderson model of localization. SIAM Review, 50(1):91– 112, ???? 2008. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sun:2006:FMM

[SBXD06] Jun Sun, Stephen Boyd, Lin Xiao, and Persi Diaconis. The fastest mixing Markov process on a graph and a connection to a maximum variance unfolding problem. SIAM Review, 48(4):681–699, January 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stanton:1970:NSF

[SC70] R. G. Stanton and D. D.

Cowan. Note on a “square” functional equation. SIAM Review, 12(2):277–279, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sarlet:1981:GNT

[SC81] Willy Sarlet and Frans Cantrijn. Generalizations of Noether’s theorem in classical mechanics. SIAM Review, 23(4):467–494, October 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scarf:1959:BRG

[Sca59] Herbert E. Scarf. Book review: Games and Decisions (R. D. Luce and H. Raiffa). SIAM Review, 1(2): 178–179, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schoenfeld:1959:IT

[Sch59] Lowell Schoenfeld. On integrals of the type ∫_{0}^{∞} f(x) log |\frac{x+1x−1}{|}dx. SIAM Review, 1(2):154– 157, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schniewind:1960:BRA

[Sch60a] Maria Schniewind. Book review: Automation and Computing (A. D. Booth). SIAM Review, 2(1):52–54, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schoenberg:1960:BRD

[Sch60b] I. J. Schoenberg. Book review: N-Dimensional Volume (Maurice Eisenstein and M. S. Klamkin). SIAM Review, 2(1):41–45, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schuster:1960:BRO

[Sch60c] Seymour Schuster. Book review: Ordinary Differential Equations (Wilfred Kaplan). SIAM Review, 2(1): 50–52, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scheuer:1961:CDS

[Sch61] Ernest M. Scheuer. The cumulative distribution of the sum or difference of a normal random variable and the absolute value of a normal random variable. SIAM Review, 3(2):152–155, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schoenfeld:1962:IT

[Sch62a] L. Schoenfeld. On integrals of the type ∫_{0}^{∞} f(x) log |\frac{x+1}{x−1} | dx. SIAM Review, 4(1):154– 157, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). A remark on a previous paper, SIAM Review 4, 147 (1962).

Schoenfeld:1962:RPP

[Sch62b] Lowell Schoenfeld. A remark on a previous paper. SIAM Review, 4(2):147, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1963:BRP

[Sch63a]  Martin Schechter.  Book

review: Partial Differential Equations. An Introduction (B. Epstein). SIAM Review, 5(1):78–79, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scheinok:1963:DS

[Sch63b] Perky Scheinok. On a double summation. SIAM Review, 5(4):371, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schoenfeld:1963:ECD

[Sch63c] Lowell Schoenfeld. The evaluation of certain definite integrals. SIAM Review, 5(4): 358–369, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scheinok:1964:UPD

[Sch64] Perry Scheinok. On the uniqueness of a probability density function. SIAM Review, 6(3):310, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scheinok:1965:BRR

[Sch65a] Perry A. Scheinok. Book review: Reliability Theory (Walter Weissblum). SIAM Review, 7(3):416–417, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schoenfeld:1965:BRC

[Sch65b] Lowell Schoenfeld. Book review: On The Coefficients of a Trigonometric Polynomial (D. J. Newman). SIAM Review, 7(2): 293–294, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schreiber:1965:BRB

[Sch65c] Alvin L. Schreiber. Book review: A Binary Multiplication (W. C. McGee). SIAM Review, 7(1):134– 136, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schwartz:1966:PWP

[Sch66] B. L. Schwartz. Possible winners in partially completed tournaments. SIAM Review, 8(3):302–308, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1967:ESP

[Sch67] Paul J. Schweitzer. Elementary symmetric polynomials. SIAM Review, 9(3):

590–591, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1968:BM

[Sch68] K. W. Schmidt. Binary matrices. SIAM Review, 10 (4):452, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1969:BMT

[Sch69a] K. W. Schmidt. Bounds for maximal two element determinants. SIAM Review, 11 (3):398, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1969:IBM

[Sch69b] K. W. Schmidt. Irreducible binary matrices. SIAM Review, 11(2):257–260, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schipper:1970:SNM

[Sch70a] J. F. Schipper. Some nuances in the matrix analysis of dynamical systems with constraints imposed. SIAM Review, 12(2):280– 283, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1970:MMC

[Sch70b] Paul J. Schweitzer. A minimization for Markov chains. SIAM Review, 12(3):469– 470, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1971:BRV

[Sch71a] M. Schechter. Book review: Vorlesungen über Partielle und Pfaffsche Differentialgleichungen (Wolfgang Haack and Wolfgang Wendland). SIAM Review, 13(3): 411–412, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1971:BM

[Sch71b] K. W. Schmidt. Binary matrices. SIAM Review, 13(1): 119–121, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1971:BMT

[Sch71c] K. W. Schmidt. Bounds for maximal two element determinants. SIAM Review, 13 (2):248–249, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1971:DI

[Sch71d] K. W. Schmidt. A determinant identity. SIAM Review, 13(3):389, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1971:MMC

[Sch71e]  Paul J. Schweitzer.  A

minimization for Markov chains. SIAM Review, 13 (1):114, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1972:BOM

[Sch72a] K. W. Schmidt. A bound for a 4^{k}-order maximal (0, 1) determinant. SIAM Review, 14(3):495, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1972:BMD

[Sch72b] K. W. Schmidt. Bounds for maximal (1, −1) determinants of order 1(mod4). SIAM Review, 14(4):644– 645, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1972:CPD

[Sch72c] K. W. Schmidt. A conjectured positive definite matrix. SIAM Review, 14(1): 165, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1973:BRT

[Sch73a]  Martin Schechter.  Book

review: Theory of Partial Differential Equations (H. Melvin Lieberstein). SIAM Review, 15(4):809, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1973:BOM

[Sch73b] K. W. Schmidt. A bound for a 4^{k}-order maximal (0, 1) determinant. SIAM Review, 15 (3):673–674, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1973:BMD

[Sch73c] K. W. Schmidt. Bounds for maximal (1, −1) determinants of order 1(mod4). SIAM Review, 15(4):799– 800, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1973:HD

[Sch73d]  K. W. Schmidt.  On

Hadamard determinants. SIAM Review, 15(4):788, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1973:HM

[Sch73e]  K. W. Schmidt.  On

Hadamard matrices. SIAM Review, 15(2):377, ???? 1973.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1973:MHD

[Sch73f] K. W. Schmidt. On the minors of a Hadamard determinant. SIAM Review, 15 (1):220, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1974:MHD

[Sch74] K. W. Schmidt. On the minors of a Hadamard determinant. SIAM Review, 16 (1):98–99, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scheid:1975:NFG

[Sch75a] F. J. Scheid. Nonlinear feature of golf course rating with application to handicapping. SIAM Review, 17(2): 388–389, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schoenberg:1975:DSH

[Sch75b]  I. J. Schoenberg.  n-

dimensional simple harmonic motion. SIAM Review, 17 (4):687, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schelin:1976:NEA

[Sch76a] C. W. Schelin. Numerical evaluation of argument principle. SIAM Review, 18 (4):826, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schreiber:1976:NMS

[Sch76b]  R. Schreiber.  Numerical

methods for singular twopoint boundary-value problems. SIAM Review, 18 (4):826, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schubert:1976:BRC

[Sch76c] C. F. Schubert. Book review: Constrained Minimization of an Integral Functional (Michael H. Moors). SIAM Review, 18(1):119– 120, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schaar:1977:BRG

[Sch77a] Richard J. Schaar. Book review: Global Theory of a Second Order Linear Ordinary Differential Equation with a Polynomial Coefficient (Yasutaku Sibuya). SIAM Review, 19(2):336– 338, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schoenberg:1977:TI

[Sch77b] I. J. Schoenberg. A triangle inequality. SIAM Review, 19(2):329, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1977:DIB

[Sch77c] P. J. Schweitzer. A definite integral of N. Bohr. SIAM Review, 19(1):147, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schroder:1978:MGU

[Sch78a] J. Schroder. M-matrices and generalizations using an operator theory approach. SIAM Review, 20(2):213– 244, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schroder:1978:MMG

[Sch78b] Johann Schröder. M-matrices and generalizations using an operator theory approach. SIAM Review, 20(2):213– 244, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schumaker:1978:BRS

[Sch78c] Larry L. Schumaker. Book review: Studies in Spline Functions and Approximation Theory (Samuel Karlin, Charles A. Micchelli, Allan Pinkus and I. J. Schoenberg). SIAM Review, 20(2): 406–407, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1978:SAM

[Sch78d] P. J. Schweitzer. Spectral analysis of a matrix. SIAM Review, 20(3):594, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schild:1980:BRM

[Sch80a] Albert Schild. Book review:

Mathematics Applied to Continuum Mechanics (Lee A. Segel). SIAM Review, 22(3): 384–385, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schuss:1980:SPM

[Sch80b] Zeev Schuss. Singular perturbation methods in stochastic differential equations of mathematical physics. SIAM Review, 22(2):119–155, April 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1980:NVF

[Sch80c] P. J. Schweitzer. A nonlinear vector finite difference Scheme. SIAM Review, 22 (4):503, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmitendorf:1981:BRV

[Sch81a] W. E. Schmitendorf. Book review: Vector-Valued Optimization Problems in Control Theory (M. E. Salukvadze). SIAM Review, 23(2): 273–274, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schrader:1981:DZF

[Sch81b] K. Schrader. Distinct zeros of a function. SIAM Review, 23(3):391, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1981:BRN

[Sch81c] P. J. Schweitzer. Book review: A Nonlinear Vector Finite Difference Scheme (P. J. Schweitzer). SIAM Review, 23(4):528–532, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1982:BRA

[Sch82a] Martin Schechter. Book review: Applied Functional Analysis (Jean-Pierre Aubin). SIAM Review, 24(1):101, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1982:BRS

[Sch82b] Martin Schechter. Book review: Spectral Theory of Ordinary Differential Operators (Erich Muller-Pfeiffer). SIAM Review, 24(4):495– 496, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1983:PO

[Sch83] P. Schweitzer. Properties of an operator. SIAM Review, 25(4):570–571, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1984:STW

[Sch84a] M. Schechter. Spectral theory and wave operators for the Schrödinger-equation — Berthier, A. M. SIAM Review, 26(2):299, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1984:BRS

[Sch84b]  Martin Schechter.  Book

review: Spectral Theory and Wave Operators for the Schrödinger Equation (A. M. Berthier). SIAM Review, 26 (2):299, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schumaker:1985:BRC

[Sch85a] Larry L. Schumaker. Book review: Computational Methods for Parsimonious Data Fitting (Marjan Ribaric).

SIAM Review, 27(4):605– 606, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1985:NMD

[Sch85b] P. Schweitzer. On a nonlinear matrix differential equation. SIAM Review, 27(2): 250, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schwenk:1985:BRC

[Sch85c] A. J. Schwenk. Book review: A Covering Problem (J. Vijay). SIAM Review, 27(2): 250–252, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1986:BRC

[Sch86a] Martin Schechter. Book review: A Course in Functional Analysis (John B. Conway). SIAM Review, 28 (1):126, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1986:BRE

[Sch86b]  Martin Schechter.  Book

review: Elliptic Problems in Nonsmooth Domains (P. Grisvard). SIAM Review, 28 (1):125, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schinzinger:1987:BRQ

[Sch87] Roland Schinzinger. Book review: Qualitative Reasoning about Physical Systems (Daniel G. Bobrow, eds.). SIAM Review, 29(3):489– 492, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schwarz:1988:SDE

[Sch88] Fritz Schwarz. Symmetries of differential equations: From sophus Lie to computer algebra. SIAM Review, 30 (3):450–481, September 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schechter:1989:BRS

[Sch89a] Martin Schechter. Book review: Semi-Classical Analysis for the Schrödinger Operator and Applications (Bernard Helffer). SIAM Review, 31(4):709, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schiffman:1989:BRF

[Sch89b] Robert L. Schiffman. Book review: The Finite Element Method in the Deformation and Consolidation of Porous Media (Roland W. Lewis and Bernard A. Schrefler). SIAM Review, 31(1): 160–161, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1990:SFE

[Sch90]  Paul Schweitzer.  Simul-

taneous functional equations. SIAM Review, 32 (4):672, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schumaker:1991:BRS

[Sch91a] Larry L. Schumaker. Book review: Spline Models for Observational Data (Grace Wahba). SIAM Review, 33 (3):502, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schumaker:1991:SMO

[Sch91b] Larry L. Schumaker. Spline models for observational data (Grace Wahba). SIAM Review, 33(3):502–??, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmid:1992:SKQ

[Sch92]  Rudolf Schmid.  Strings,

knots, and quantum groups: a glimpse at three 1990 Fields Medalists. SIAM Review, 34 (3):406–425, September 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schattschneider:1993:BRC

[Sch93a] Doris Schattschneider. Book review: Crystalline Symmetries: an Informal Mathematical Introduction (Marjorie Senechal). SIAM Review, 35(2):335–336, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schattschneider:1993:CSI

[Sch93b] Doris Schattschneider. Crystalline symmetries: An informal mathematical introduction (Marjorie Senechal). SIAM Review, 35(2):335– ??, June 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1993:BRC

[Sch93c] F. Schmidt. Book review:

Circular Binary Sequences (Mark Stamp). SIAM Review, 35(3):500–502, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schreiber:1993:BRV

[Sch93d] Robert Schreiber. Book review: Very Large Scale Computation in the 21st Century (Jill P Mesirov, ed.). SIAM Review, 35(3):534– 535, September 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1994:BRS

[Sch94] D. S. Schmidt. Book review: Scientific Programmer’s Toolkit — Turbo Pascal Edition (M. H. Beilby, R. D. Harding, and M. R. Manning). SIAM Review, 36(2): 319–320, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scheurle:1995:BRC

[Sch95] Juergen Scheurle. Book review: The Couette–Taylor Problem (Pascal Chossat and Gerard boss). SIAM Review, 37(2):279–280, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schorn:1996:TCI

[Sch96] M. Schorn. Two combinatorial identities. SIAM Review, 38(1):150–151, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:1999:EIb

[Sch99a] B. Schnabel. Education — introduction. SIAM Review, 41(4):793, ???? 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:1999:Ea

[Sch99b] Bobby Schnabel. Education. SIAM Review, 41 (1):133–134, March 1999.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97022.

Schnabel:1999:Eb

[Sch99c] Bobby Schnabel. Education. SIAM Review, 41 (3):575, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97032.

Schnabel:1999:Ec

[Sch99d] Bobby Schnabel. Education. SIAM Review, 41(4): 793, December 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97038.

Schnabel:1999:EIa

[Sch99e] Bobby Schnabel. Education — introduction. SIAM Review, 41(2):333, June 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97027.

Schnabel:2000:EIa

[Sch00a] B. Schnabel. Education — introduction. SIAM Review, 42(3):485, ???? 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:2000:EIb

[Sch00b] B. Schnabel. Education — introduction. SIAM Review, 42(4):717, ???? 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:2000:Ea

[Sch00c] Bobby Schnabel. Education. SIAM Review, 42 (1):107, March 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97044.

Schnabel:2000:Eb

[Sch00d] Bobby Schnabel. Education. SIAM Review, 42 (2):295, June 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97048.

Schnabel:2000:Ec

[Sch00e] Bobby Schnabel. Education. SIAM Review, 42 (3):485, September 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97054.

Schnabel:2000:Ed

[Sch00f] Bobby Schnabel. Education. SIAM Review, 42(4):

717, December 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97059.

Schabel:2001:E

[Sch01a]  Bobby Schabel.  Educa-

tion. SIAM Review, 43 (2):345, June 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97069.

Schnabel:2001:Eb

[Sch01b] B. Schnabel. Education.

SIAM Review, 43(2):345, ???? 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:2001:Ea

[Sch01c] Bobby Schnabel. Education. SIAM Review, 43 (1):161, March 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97064.

Schnabel:2001:Ec

[Sch01d] Bobby Schnabel. Education. SIAM Review, 43 (3):523, September 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97074.

Schnabel:2001:Ed

[Sch01e] Bobby Schnabel. Education. SIAM Review, 43(4): 677, December 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97079.

Schnabel:2002:EI

[Sch02a] B. Schnabel. Education — introduction. SIAM Review, 44(4):659, ???? 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:2002:U

[Sch02b] B. Schnabel. Untitled. SIAM Review, 44(3):455, ???? 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:2002:Ea

[Sch02c] Bobby Schnabel. Education. SIAM Review, 44 (1):109, March 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97084.

Schnabel:2002:Eb

[Sch02d] Bobby Schnabel. Education. SIAM Review, 44 (2):257, June 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97089.

Schnabel:2002:Ec

[Sch02e] Bobby Schnabel. Education. SIAM Review, 44 (3):455, September 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97094.

Schnabel:2002:Ed

[Sch02f] Bobby Schnabel. Education. SIAM Review, 44(4): 659, December 2002. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97099.

Schnabel:2003:EIa

[Sch03a] B. Schnabel. Education — introduction. SIAM Review, 45(1):93, ???? 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:2003:EIb

[Sch03b] B. Schnabel. Education — introduction. SIAM Review, 45(2):325, ???? 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:2003:EIc

[Sch03c] B. Schnabel. Education — introduction. SIAM Review, 45(3):541, ???? 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schnabel:2003:Ea

[Sch03d] Bobby Schnabel. Education. SIAM Review, 45 (1):93, March 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97104.

Schnabel:2003:Eb

[Sch03e] Bobby Schnabel. Education. SIAM Review, 45 (2):325, June 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97109.

Schnabel:2003:Ec

[Sch03f] Bobby Schnabel. Education. SIAM Review, 45 (3):541, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97114.

Schnabel:2003:Ed

[Sch03g] Bobby Schnabel. Education. SIAM Review, 45(4):

785, December 2003. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97119.

Schafer:2004:LCP

[Sch04a] Uwe Schäfer. A linear complementarity problem with a P -matrix. SIAM Review, 46(2):189–201, June 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42008.

Schnabel:2004:Ea

[Sch04b] Bobby Schnabel. Education. SIAM Review, 46 (1):113, March 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97124.

Schnabel:2004:Eb

[Sch04c] Bobby Schnabel. Education. SIAM Review, 46 (2):309, June 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97128.

Schnabel:2004:Ec

[Sch04d] Bobby Schnabel. Education. SIAM Review, 46 (3):499, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97133.

Schnabel:2004:Ed

[Sch04e] Bobby Schnabel. Education. SIAM Review, 46(4): 715, December 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97138.

Schnabel:2005:E

[Sch05] Bobby Schnabel. Education. SIAM Review, 47 (1):133, March 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97143.

Schulz:2018:BRb

[Sch18a] Volker Schulz. Book reviews. SIAM Review, 60(2): 465–487, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2018:BRc

[Sch18b] Volker Schulz. Book reviews. SIAM Review, 60(3): 757–776, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2018:BRa

[Sch18c] Volker H. Schulz. Book reviews. SIAM Review, 60(1): 205–220, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2018:BRd

[Sch18d] Volker H. Schulz. Book reviews. SIAM Review, 60(4): 1013–1027, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2019:BRa

[Sch19a] Volker H. Schulz. Book reviews. SIAM Review, 61(1): 209–225, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2019:BRb

[Sch19b] Volker H. Schulz. Book reviews. SIAM Review, 61(2): 385–400, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2019:BRc

[Sch19c] Volker H. Schulz. Book reviews. SIAM Review, 61(3): 625–640, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2019:BRd

[Sch19d] Volker H. Schulz. Book reviews. SIAM Review, 61(4): 895–906, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2020:BRa

[Sch20a] Volker H. Schulz. Book reviews. SIAM Review, 62(1): 283–297, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2020:BRb

[Sch20b] Volker H. Schulz. Book reviews. SIAM Review, 62(2): 511–526, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2020:BRc

[Sch20c] Volker H. Schulz. Book reviews. SIAM Review, 62(3): 729–739, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2020:BRd

[Sch20d] Volker H. Schulz. Book reviews. SIAM Review, 62(4): 985–994, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2021:BRa

[Sch21a] Volker H. Schulz. Book reviews. SIAM Review, 63(1): 231–245, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2021:BRb

[Sch21b] Volker H. Schulz. Book reviews. SIAM Review, 63(2): 419–431, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2021:BRc

[Sch21c] Volker H. Schulz. Book reviews. SIAM Review, 63(3): 641–650, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2021:BRd

[Sch21d] Volker H. Schulz. Book reviews. SIAM Review, 63(4): 867–875, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz:2022:BRa

[Sch22a] Volker H. Schulz. Book reviews. SIAM Review, 64 (1):215–225, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N97541X.

Schulz:2022:BRb

[Sch22b] Volker H. Schulz. Book reviews. SIAM Review, 64 (2):503–513, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975469.

Schulz:2022:BRc

[Sch22c] Volker H. Schulz. Book reviews. SIAM Review, 64 (3):751–759, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975548.

Schulz:2022:BRd

[Sch22d] Volker H. Schulz. Book reviews. SIAM Review, 64 (4):1083–1095, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975597.

Schulz:2023:BRa

[Sch23a] Volker H. Schulz. Book reviews. SIAM Review, 65 (1):319–328, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N97565X.

Schulz:2023:BRb

[Sch23b] Volker H. Schulz. Book reviews. SIAM Review, 65 (2):591–598, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N975715.

Schulz:2023:BRc

[Sch23c] Volker H. Schulz. Book reviews. SIAM Review, 65 (3):905–915, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23N975764.

Schulz:2023:BR

[Sch23d] Volker H. Schulz. Book reviews. SIAM Review, 65(4): 1187–1197, November 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1959:ESS

[Sco59] E. J. Scott. An eigenfunction series solution of a certain hyperbolic partial differential equation. SIAM Review, 1 (2):160–166, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1983:BRS

[Sco83]  Alwyn Scott.  Book re-

view: Solitons and the Inverse Scattering Transform (Mark J. Ablowitz and Harvey Segur). SIAM Review, 25 (2):289–290, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1984:SNW

[Sco84a]  A. Scott.  Solitons and

nonlinear wave equations — Dodd, R. K., Eilbeck, J. C., Gibbon, J. D., Morris, H. C. SIAM Review, 26(2): 293–294, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1984:BRS

[Sco84b] Alwyn Scott. Book review: Solitons and Nonlinear Wave Equations (R. K. Dodd, J. C. Eilbeck, J. D. Gibbon and H. C. Morris). SIAM Review, 26 (2):293–294, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1985:BRS

[Sco85]  Alwyn Scott.  Book re-

view: Solitons (P. G. Drazin). SIAM Review, 27(1):125, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1988:BRS

[Sco88]  Alwyn Scott.  Book re-

view: Solitons in Molecular Systems (A. S. Davydov; Eugene S. Kryachko, trans.). SIAM Review, 30(1): 155–156, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1990:BRSa

[Sco90a] Alwyn Scott. Book review:

Solitons: an Introduction (P. G. Drazin and R. S. Johnson). SIAM Review, 32(1): 185–186, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1990:BRSb

[Sco90b] Alwyn Scott. Book review:

Solitons: An Introduction (P. G. Drazin and R. S. Johnson). SIAM Review, 32(1): 185–??, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1991:BRAa

[Sco91a]  Alwyn Scott.  Book re-

view: Asymptotic Analysis of a Class of Perturbed Korteweg–de Vries Initial Value Problems (F. DeKerf). SIAM Review, 33(2): 337–338, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1991:BRAb

[Sco91b]  Alwyn Scott.  Book re-

view: Asymptotic Analysis of a Class of Perturbed Korteweg–de Vries Initial Value Problems (F. DeKerf). SIAM Review, 33(2):

337–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1992:BRM

[Sco92] Peter D. Scott. Book review: Minimalist Mobile Robotics. A Colony-Style Architecture for an Artificial Creature (Jonathan H. Connell). SIAM Review, 34(2): 329–330, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1994:BRH

[Sco94] L. Ridgway Scott. Book review: Handbook of Numerical Analysis, Volume II. Finite Element Methods (Part I) (P. G. Ciarlet and J. L. Lions, eds.). SIAM Review, 36 (2):320–321, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saad:2010:NME

[SCS10] Yousef Saad, James R. Chelikowsky, and Suzanne M. Shontz. Numerical methods for electronic structure calculations of materials. SIAM Review, 52(1): 3–54, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schulz-DuBois:1974:FCF

[SD74] E. O. Schulz-DuBois. Fourier coefficients of a function involving elliptic integrals of the first kind. SIAM Review, 16(4):547, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simons:2006:SCS

[SDW06] Frederik J. Simons, F. A.

Dahlen, and Mark A. Wieczorek. Spatiospectral concentration on a sphere. SIAM Review, 48(3):504– 536, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sealander:1961:EM

[Sea61]  Carl E. Sealander.  On

the eigenvalues of a matrix. SIAM Review, 3(3): 262, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sebastian:1970:BRM

[Seb70] James D. Sebastian. Book review: Moving Furniture Through a Hallway (Leo Moser). SIAM Review, 12 (4):582–586, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sebastian:1975:CNF

[Seb75] J. D. Sebastian. Computation of Neumann functions of complex argument and real order. SIAM Review, 17 (2):389, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sedney:1988:BRN

[Sed88]  Raymond Sedney.  Book

review: Nonlinear Conical Flow (B. M. Bulakh). SIAM Review, 30(2):352– 353, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sedaghat:1991:MSA

[Sed91] Hassan Sedaghat. Measuring the speed and altitude of an aircraft using similar triangles. SIAM Review, 33 (4):650–654, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sedaghat:1996:GPF

[Sed96] Hassan Sedaghat. Geometric properties of factorable planar systems of differential equations. SIAM Review, 38 (4):660–665, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29302.

Sedaghat:2002:COC

[Sed02]  Hassan Sedaghat.  Con-

vergence, oscillations, and chaos in a discrete model of combat. SIAM Review, 44(1):74–92, March 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38299.

Segel:1972:SS

[Seg72] Lee A. Segel. Simplification and scaling. SIAM Review, 14(4):547–571, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Segel:1979:BRM

[Seg79] Lee A. Segel. Book review: Mathematical Modelling Techniques (R. Aris). SIAM Review, 21(4):571– 572, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Segel:1980:BRA

[Seg80] Lee A. Segel. Book review:

Applied Mathematics: an Intellectual Orientation (Francis J. Murray). SIAM Review, 22(2):246–247, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Segel:1985:BRS

[Seg85] Lee A. Segel. Book review: Synergetics of the Brain (E. Basar, H. Flohr, H. Haken and A. J. Mandell, eds.). SIAM Review, 27(2):

300–301, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Segel:1988:BRM

[Seg88] Lee A. Segel. Book review: Mathematical Models in Biology (Leah Edelstein-Keshet). SIAM Review, 30 (4):679–680, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Segura:2006:ROP

[Seg06] Javier Segura. Review of Orthogonal Polynomials: Computation and Approximation by Walter Gautschi. SIAM Review, 48(2):431– 433, June 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //www.jstor.org/stable/
                                        20453824.

Seidman:1962:BRL

[Sei62] Thomas Seidman. Book review: Lectures on Game Theory (E. S. Ventzel). SIAM Review, 4(4):406, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seifert:1965:SRR

[Sei65] George Seifert. Some recent results for almost periodic solutions of differential equations. SIAM Review, 7(4):

529–538, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seitelman:1976:ENM

[Sei76] L. H. Seitelman. Efficient numerical method for solution of three-dimensional incompressible potential flow problems. SIAM Review, 18 (4):826, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seitelman:1978:NUT

[Sei78]  L. H. Seitelman.  New

user-transparent end condition for cubic spline data fitting. SIAM Review, 20 (3):634, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seidman:1985:BRS

[Sei85] Stephen B. Seidman. Book review: Structural Models in Anthropology (Per Hags and Frank Harary). SIAM Review, 27(2):301– 303, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seiffert:1986:BRI

[Sei86a] H.-J. Seiffert. Book review:

An Inequality (J. L. Brenner, W. A. Newcomb and O. G. Rueh). SIAM Review, 28 (4):573–574, December 1986.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seiffert:1986:I

[Sei86b] H. J. Seiffert. An inequality. SIAM Review, 28(4): 574, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seiffert:1987:C

[Sei87]  H. J. Seiffert.  Correc-

tion. SIAM Review, 29(4): 637, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seiffert:1991:IFM

[Sei91] H.-J. Seiffert. An inequality for a family of means. SIAM Review, 33(3):473, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seifert:1993:BRO

[Sei93] George Seifert. Book review: Oscillation Theory of Delay Differential Equations (I. Györi and G. Ladas); Oscillation Theory for Neutral Differential Equations with Delay (D. D. Bainov and D. P. Mishev). SIAM Review, 35 (4):677–678, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seiffert:1997:TIA

[Sei97] H. J. Seiffert. Three inequalities for the arithmetic, identric, and geometric means. SIAM Review, 39(2):330– 332, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sekiguchi:1982:MMP

[Sek82] T. Sekiguchi. A maximum, minimum problem. SIAM Review, 24(1):77, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sellers:1976:BRM

[Sel76] Peter H. Sellers. Book review: Mathematical Structure in Human Affairs (R. H. Atkin). SIAM Review, 18(2): 312–313, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Selgrade:1994:BRD

[Sel94] James F. Selgrade. Book review: Dynamical Systems — an Introduction with Applications in Economics and Biology (Pierre N. V. Tu). SIAM Review, 36(2): 321–322, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Semarne:1959:NDM

[Sem59] H. Manvel Semarne. New direct method of solution of a system of simultaneous linear equations. SIAM Review, 1 (1):53–54, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Senn:1989:AMD

[Sen89] Peter Senn. An average maximum distance. SIAM Review, 31(2):319, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Senechal:1994:BRV

[Sen94] Marjorie Senechal. Book review: The Visual Mind (M. Emmer, ed.). SIAM Review, 36(4):676–677, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serrin:1964:BRM

[Ser64] James Serrin. Book review: The Mathematical Theory of Viscous Incompressible Flow (O. A. Ladyzhenskaya). SIAM Review, 6(3):315–316, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serbin:1976:SMP

[Ser76] S. M. Serbin. Solution of matrix pencil systems. SIAM Review, 18(4):826, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serfling:1978:SER

[Ser78] R. J. Serfling. Some elementary results on Poisson approximation in a sequence of Bernoulli trials. SIAM Review, 20(3):567– 579, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serfling:1980:BRS

[Ser80a] R. J. Serfling. Book review: Stochastic Convergence; 2nd Edition (Eugene Lukacs). SIAM Review, 22 (2):243–245, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serfozo:1980:BRF

[Ser80b] Richard Serfozo. Book review: Finite Dynamic Programming: an Approach to Finite Markov Decision Processes (D. J. White). SIAM Review, 22(4):527– 528, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serfozo:1984:OTV

[Ser84a] R. Serfozo. Optimization over time, vol 1, dynamic programming and stochastic control — Whittle, P. SIAM Review, 26(2):299– 302, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serfozo:1984:BRO

[Ser84b] Richard Serfozo. Book review: Optimization Over Time Volume 1: Dynamic Programming and Stochastic Control (Peter Whittle). SIAM Review, 26(2): 299–302, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serfozo:1986:BRA

[Ser86] Richard F. Serfozo. Book review: Asymptotic Methods in Queueing Theory (A. A. Borovkov). SIAM Review, 28 (2):279–283, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serfozo:1994:BRS

[Ser94] Richard Serfozo. Book review: Stationary Stochastic Models (A. Brands, P. Franken, and B. Lisek). SIAM Review, 36(1):141– 143, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serra:1996:BRM

[Ser96] Jean Serra. Book review: Morphological Image Operators (Henk J. A. M. Heijmans). SIAM Review, 38(1): 178–179, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Serrano:2004:TIS

[Ser04] Roberto Serrano. The theory of implementation of social choice rules. SIAM Review, 46(3):377–414, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/43594.

Sherman:1975:DMS

[SES75] A. H. Sherman, S. C. Eisenstat, and M. H. Schultz. Direct methods for solution of sparse systems of linear equations with limited core storage. SIAM Review, 17(2): 389, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seth:1962:BRM

[Set62a] B. R. Seth. Book review:

Mechanics (J. P. Den Hartog). SIAM Review, 4(3): 272, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seth:1962:BRS

[Set62b] B. R. Seth. Book review:

Strength of Materials (J. P. Den Hartog). SIAM Review, 4(3):272, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seth:1963:BRS

[Set63] B. R. Seth. Book review: Statistical Strength Theory (S. D. Volkov). SIAM Review, 5(1):84, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sethi:1977:DOC

[Set77]  Suresh P. Sethi.  Dy-

namic optimal control models in advertising: a survey. SIAM Review, 19(4): 685–725, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sethian:1999:FMM

[Set99] J. A. Sethian. Fast marching methods. SIAM Review, 41(2):199–235, June 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34705.

Seydel:1993:AMD

[Sey93a] Rudiger Seydel. The art of modeling dynamic systems: Forecasting for chaos, randomness, and determinism (Foster Morrison). SIAM Review, 35(1):167–??, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seydel:1993:BRA

[Sey93b] Rüdiger Seydel. Book review: The Art of Modeling Dynamic Systems: Forecasting for Chaos, Randomness, and Determinism (Foster Morrison). SIAM Review, 35(1): 167–168, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seymour:1993:MGA

[Sey93c]  K. M. Seymour.  Max-

imum gravitational attraction. SIAM Review, 35(4): 642, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sezer:1990:BRL

[Sez90] M. Erol Sezer. Book review:

Large Scale Systems: Decentralization, Structure Constraints and Fixed Modes (L. Trave, A. Titli, and A. Tarras). SIAM Review, 32 (4):708–709, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steele:1997:PTC

[SF97] J. Michael Steele and Alan Frieze. Probability theory and combinatorial optimization. SIAM Review, 39(4): 793–??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strecok:1966:BRZ

[SFH^{+}66] A. J. Strecok, H. E. Fettis, W. W. Hooker, L. Magagna, and E. P. Merkes. Book review: On the Zeros of a Set of Polynomials (M. N. S. Swamy). SIAM Review, 8(3): 388–391, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepard:1999:ODR

[SFOM99] David M. Shepard, Michael C. Ferris, Gustavo H. Olivera, and T. Rockwell Mackie. Optimizing the delivery of radiation therapy to cancer patients. SIAM Review, 41 (4):721–744, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34203.

Smith:1987:EMV

[SFS87] David A. Smith, William F. Ford, and Avram Sidi. Extrapolation methods for vector sequences. SIAM Review, 29(2):199–233, June 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See correction [SFS88].

Smith:1988:CEM

[SFS88] David A. Smith, William F. Ford, and Avram Sidi. Correction to: “Extrapolation methods for vector sequences”. SIAM Review, 30 (4):623–624, December 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [SFS87].

Scholten:1972:DTF

[SG72] William B. Scholten and Richard A. Gaggioli. Differentiation of tensor fields. SIAM Review, 14(2):271– 277, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scholten:1973:NCH

[SG73] William B. Scholten and Richard A. Gaggioli. Normals to coordinate hypersurfaces as weight minusone covariant base vectors. SIAM Review, 15(2):275– 282, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shampine:1976:QRS

[SG76a] L. F. Shampine and M. K.

Gordon. Qn — root solving code. SIAM Review, 18 (4):826, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slepian:1976:DOC

[SG76b] D. Slepian and E. N. Gilbert.

Doubly orthogonal concentrated polynomials. SIAM Review, 18(4):828, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shampine:1979:UVS

[SG79] L. F. Shampine and C. W. Gear. A User’s view of solving stiff ordinary differential equations. SIAM Review, 21 (1):1–17, January 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sabharwal:1981:LM

[SG81] Y. P. Sabharwal and B. R. Gupta. A learning model. SIAM Review, 23(4):523– 524, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shawagfeh:1982:PSE

[SG82] N. T. Shawagfeh and M. L. Glasser. Power series of an elliptic function. SIAM Review, 24(3):345–346, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shampine:1995:NSO

[SG95] L. F. Shampine and I. Gladwell. Numerical solution of ordinary differential equations. SIAM Review, 37(1): 122–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sidi:1998:FMP

[SGD98] A. Sidi, C. C. Grosjean, and E. De Meyer. A family of matrix problems. SIAM Review, 40(3):718–723, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Samarskii:1996:BQP

[SGK^{+}96] A. A. Samarskii, V. A.

Galaktionov, S. P. Kurdyumov, A. P. Mikhailov, and Howard Levine. Blow-up in quasilinear parabolic equations. SIAM Review, 38(4): 692–??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stuart:1994:MPN

[SH94] A. M. Stuart and A. R.

Humphries. Model problems in numerical stability theory for initial value problems. SIAM Review, 36(2): 226–257, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/24014.htm.

Shasha:1995:CPC

[SH95] Dennis Shasha and Glen Richard Hall. Codes, puzzles, and conspiracy. SIAM Review, 37 (1):127–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sun:1996:IPG

[SH96] Ne-Zheng Sun and Mary C.

Hill. Inverse problems in groundwater modeling.

SIAM Review, 38(2):342– ??, June 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1963:IRBa

[Sha63a] Robert E. Shafer. On an integral related to Bessel functions. SIAM Review, 5(1): 73–74, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1963:IRBb

[Sha63b] Robert E. Shafer. On integrals related to Bessel functions. SIAM Review, 5(4): 370, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shanks:1964:DRE

[Sha64a] Daniel Shanks. A duality relation in exterior ballistics. SIAM Review, 6(1): 54–56, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sharpe:1964:UBR

[Sha64b] C. B. Sharpe. Upper bounds for resistance and reactance integrals of analytic functions. SIAM Review, 6(2): 126–130, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shanks:1965:BRI

[Sha65] Daniel Shanks. Book review:

An Infinite Permutation (M. S. Klamkin). SIAM Review, 7 (2):284–286, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1970:GSH

[Sha70] Robert E. Shafer. A generalization of spherical harmonics. SIAM Review, 12(3): 467–468, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1971:EI

[Sha71a] Robert E. Shafer. Expansion of an integral. SIAM Review, 13(4):571–572, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1971:FS

[Sha71b] Robert E. Shafer. Fourier series. SIAM Review, 13 (1):115, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1971:GSHa

[Sha71c] Robert E. Shafer. A generalization of spherical harmonics. SIAM Review, 13(1): 113–114, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1971:GSHb

[Sha71d] Robert E. Shafer. A generalization of spherical harmonics. SIAM Review, 13(3): 387–388, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1972:FS

[Sha72a] Robert E. Shafer. Fourier series. SIAM Review, 14(1): 174–176, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sharma:1972:SPN

[Sha72b] A. Sharma. Some poised and nonpoised problems of interpolation. SIAM Review, 14 (1):129–151, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1973:GSH

[Sha73a] Robert E. Shafer. A generalization of spherical harmonics. SIAM Review, 15(3): 658–663, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1973:II

[Sha73b] Robert E. Shafer. An integral inequality. SIAM Review, 15 (1):219, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1974:BRI

[Sha74] Robert E. Shafer. Book review: An Integral Equality (Robert E. Shafer). SIAM Review, 16(1):96–97, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shampine:1975:SNS

[Sha75]  L. F. Shampine.  Solv-

ing non-stiff ordinary differential equations — state of art. SIAM Review, 17(2): 389, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shantaram:1978:TED

[Sha78] R. Shantaram. Two equal determinants. SIAM Review, 20(3):594, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1982:BRP

[Sha82a] R. E. Shafer. Book review: Positivity of the Coefficients of a Polynomial (Ko-Wei Lih and E. T. H. Wang). SIAM Review, 24(3): 344–345, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shawagfeh:1982:BRP

[Sha82b] N. T. Shawagfeh. Book review: Power Series of an Elliptic Function (M. L.

Glasser). SIAM Review, 24 (3):345–346, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1983:BRI

[Sha83] Robert E. Shafer. Book review: An Identity (Jo Ann Clark and O. G. Ruehr). SIAM Review, 25(2):273– 274, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shanno:1985:BRC

[Sha85] David F. Shanno. Book review: Combinatorial Optimization for Undergraduates (L. R. Foulds). SIAM Review, 27(4):606–607, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shah:1988:VIC

[Sha88] G. A. Shah. Volume of intersection of a cone with a sphere. SIAM Review, 30(2): 310–311, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1989:NES

[Sha89a] R. E. Shafer. Numerical evaluation of a slowly convergent series. SIAM Review, 31(3):495, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shaw:1989:BRV

[Sha89b] Steven Shaw. Book review:

Vibrations in Mechanical Systems: Analytical Methods and Applications (Maurice Roseau). SIAM Review, 31 (3):513–514, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shaffer:1990:MRP

[Sha90] Dorothy Browne Shaffer. The maximum range of a projectile in a vacuum revisited. SIAM Review, 32(2): 289–293, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shanmugam:1992:BRP

[Sha92a] Ramalingam Shanmugam.

Book review: Point Processes and Their Statistical Inference (Alan F. Karr). SIAM Review, 34(4):684, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shanmugam:1992:PPT

[Sha92b] Ramalingam Shanmugam.

Point processes and their statistical inference (Alan F Karr). SIAM Review, 34(4): 684–??, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shallit:1994:MP

[Sha94] Jeffrey Shallit. A minimization problem. SIAM Review, 36(3):490–491, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shafer:1995:BRN

[Sha95a] Douglas S. Shafer. Book review: Nonlinear Dynamics and Chaos: With Applications to Physics, Biology, Chemistry, and Engineering (Steven H. Strogatz). SIAM Review, 37(2): 280–281, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shallit:1995:MP

[Sha95b] J. Shallit. A minimization problem. SIAM Review, 37 (3):451–458, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shapiro:1995:BRM

[Sha95c] Alexander Shapiro. Book review: Monotone Structure in Discrete-Event Systems (P. Glasserman and D. D. Yao). SIAM Review, 37 (3):474, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shah:1996:BRV

[Sha96]  Jayant Shah.  Book re-

view: Variational Methods in Image Segmentation (Jean-Michel Morel and Sergio Solimini). SIAM Review, 38(1): 179–180, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shampine:1997:BRN

[Sha97a] Lawrence F. Shampine. Book review: Numerical Methods for Differential Equations: A Computational Approach, by John R. Dormand. SIAM Review, 39(2): 358–360, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shanmugam:1997:BRA

[Sha97b] Ramalingam Shanmugam.

Book review: Asymptotic Efficiency of Nonparametric Tests, by Yakov Nikitin. SIAM Review, 39(1):162– 163, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sharygin:1998:MMP

[Sha98] I. F. Sharygin. A maximum, minimum problem. SIAM Review, 40(3):710, ???? 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shaw:2004:RHF

[Sha04]  William T. Shaw.  Re-

covering holomorphic functions from their real or imaginary parts without the Cauchy–Riemann equations. SIAM Review, 46(4):717– 728, December 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/43215.

Shepp:1960:PP

[She60a] Larry Shepp. A pie problem. SIAM Review, 2(4): 297, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sherman:1960:FBP

[She60b]  Chris Sherman.  A free

boundary problem. SIAM Review, 2(2):154–155, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1961:EVP

[She61] Lawrence Shepp. The expected value of a product. SIAM Review, 3(4): 329, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1962:NFN

[She62] Larry Shepp. Normal functions of normal random variables. SIAM Review, 4(3): 255–256, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1963:BRC

[She63a] L. A. Shepp. Book review:

A Coin Tossing Problem (D. J. Newman). SIAM Review, 5(4):372, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1963:BRR

[She63b] Larry Shepp. Book review:

A Random Walk in a Hyper-Cube (R. H. G. Mitchell and R. F. Wheeling). SIAM Review, 5(1):75–76, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1964:NFN

[She64] Larry Shepp. Normal functions of normal random variables. SIAM Review, 6(4): 459–460, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1965:D

[She65] L. A. Shepp. A determinant. SIAM Review, 7(4): 559–560, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1966:VTK

[She66a] L. Shepp. A variation of a theorem of Kakutani. SIAM Review, 8(2):233, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1966:BRD

[She66b] L. A. Shepp. Book review: A Determinant (L. A. Shepp). SIAM Review, 8(4): 529–530, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepherd:1967:MMD

[She67] William L. Shepherd. Modular measurements of a distance. SIAM Review, 9(2): 223–228, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1971:FE

[She71] L. A. Shepp. A functional equation. SIAM Review, 13 (1):114, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shere:1972:RAP

[She72] K. D. Shere. A resource allocation problem. SIAM Review, 14(3):493, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shen:1975:RMS

[She75] M. C. Shen. Ray method for surface waves on fluid of variable depth. SIAM Review, 17 (1):38–56, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sheskin:1976:MSM

[She76]  T. J. Sheskin.  Merging

states in a Markov-chain with known steady-state probabilities. SIAM Review, 18 (4):827, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sheep:1977:BC

[She77a] L. A. Sheep. A boundedness condition. SIAM Review, 19 (2):328, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shen:1977:BRM

[She77b] Ruth R. Shen. Book review:

Macroeconomic Theory (Paul Burrows and Theodore Hitiris). SIAM Review, 19(3): 571–572, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shepp:1978:BC

[She78] L. A. Shepp. A boundedness condition. SIAM Review, 20 (2):398–399, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sheriff:1987:BRS

[She87] R. E. Sheriff. Book review:

Seismic Wave Propagation in Stratified Media (B. L. N. Kennett). SIAM Review, 29 (3):492–493, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sherman:1991:BRM

[She91] Arthur Sherman. Book review: Mathematical Modelling. A Tool for Problem Solving in Engineering, Physical, Biological and Social Sciences (D. N. P. Murthy, N. W. Page, and E. Y. Rodin). SIAM Review, 33 (4):681–682, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shen:2009:LSH

[She09] Jianhong (Jackie) Shen. Least-squares halftoning via human vision system and Markov gradient descent (LS-MGD): Algorithm and analysis. SIAM Review, 51(3): 567–589, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shinbrot:1968:BRD

[Shi68] Marvin Shinbrot. Book review: Distributions and the Boundary Values of Analytic Functions (E. J. Beltrami and M. R. Wohlers). SIAM Review, 10(1):95, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shinbrot:1969:BRM

[Shi69] Marvin Shinbrot. Book review: Mathematical Methods (Jacob Korevaar). SIAM Review, 11(2):291–292, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shinbrot:1971:BRM

[Shi71]  Marvin Shinbrot.  Book

review: The Mathematical Theory of Viscous Incompressible Flow (O. A. Ladyzhenskaya). SIAM Review, 13(1):103–106, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shippy:1993:BRB

[Shi93] David J. Shippy. Book review: The Boundary Integral Approach to Static and Dynamic Contact Problems (H. Antes and P. D. Panagiotopoulos). SIAM Review, 35(4):678–680, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shields:1994:BRP

[Shi94] Paul C. Shields. Book review: Probability, Random Processes, and Ergodic Properties (Robert M. Gray). SIAM Review, 36(1):143– 145, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shirron:1997:BRH

[Shi97] Joseph J. Shirron. Book review: Handbook of Radiation and Scattering of Waves, by Adrianus T. de Hoop. SIAM Review, 39(1): 163–164, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shitov:2017:NRM

[Shi17] Yaroslav Shitov. The nonnegative rank of a matrix: Hard problems, easy solutions. SIAM Review, 59(4): 794–800, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schweitzer:1979:SAM

[SHL^{+}79] P. J. Schweitzer, J. Z. Hearon, C. E. Langenhop, W. B. Jordan, and E. L. Wachspress. Spectral analysis of a matrix. SIAM Review, 21(3):398–400, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schick:2016:NGM

[SHL16] M. Schick, V. Heuveline, and O. P. Le Ma. A Newton–Galerkin method for fluid flow exhibiting uncertain periodic dynamics. SIAM Review, 58(1):119– 140, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Short:1964:DI

[Sho64] P. J. Short. A definite integral. SIAM Review, 6 (2):177, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Short:1966:BRR

[Sho66a] P. J. Short. Book review: A Ratio of Two Definite Multiple Integrals (A. J. Strecok). SIAM Review, 8(3): 393, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Short:1966:ILT

[Sho66b] P. J. Short. An incomplete Laplace transform. SIAM Review, 8(4):525–526, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shoenfield:1969:BRT

[Sho69]  J. R. Shoenfield.  Book

review: Theory of Recursive Functions and Effective Computability (Hartley Rogers, Jr.). SIAM Review, 11(3):415–416, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shor:1978:CI

[Sho78] Peter Shor. A combinatorial identity. SIAM Review, 20(2):394, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Showalter:1987:BRA

[Sho87]  R. E. Showalter.  Book

review: Approximation of Nonlinear Evolution Systems (Joseph W. Jerome). SIAM Review, 29(3):493– 495, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shor:1999:PTA

[Sho99] Peter W. Shor. Polynomialtime algorithms for prime factorization and discrete logarithms on a quantum computer. SIAM Review, 41(2):303–332, June 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34701.

Shrager:1976:IMA

[Shr76] R. I. Shrager. Implementing marquardts algorithm using singular value decomposition. SIAM Review, 18 (4):827, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shulman:1962:PSP

[Shu62] Harold Shulman. The pole spotting problem. SIAM Review, 4(1):12–15, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shubik:1980:BRO

[Shu80] Martin Shubik. Book review: Oligopoly and the Theory of Games (James W. Friedman). SIAM Review, 22 (1):107–109, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shubik:1982:BRA

[Shu82] Martin Shubik. Book review: Axiomatic Models of Bargaining (Alvin E. Roth). SIAM Review, 24(4):501– 502, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shubik:1985:BRGa

[Shu85a] Martin Shubik. Book review:

Game Theory (Guillermo Owen). SIAM Review, 27(2): 303–305, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shubik:1985:BRGb

[Shu85b] Martin Shubik. Book review:

Games, Theory and Applications (L. C. Thomas). SIAM Review, 27(4):608, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shub:1987:BRI

[Shu87] Michael Shub. Book review: Information, Uncertainty, Complexity (J. F. Traub, G. W. Wasilkowski and H. Woźniakowski). SIAM Review, 29(3):495–497, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shubik:1990:BRS

[Shu90a] Martin Shubik. Book review: Stability and Perfection of Nash Equilibria (Eric van Damme). SIAM Review, 32(2):332–333, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shubik:1990:SPN

[Shu90b] Martin Shubik. Stability and perfection of Nash equilibria (Eric van Damme). SIAM Review, 32(2):332– ??, June 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shu:2009:HOW

[Shu09] Chi-Wang Shu. High order weighted essentially nonoscillatory schemes for convection dominated problems. SIAM Review, 51(1):82– 126, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

SWCE:2001:GEC

[SIA01] SIAM Working Group on CSE Education. Graduate education in computational science and engineering. SIAM Review, 43 (1):163–177, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37974.

Sibuya:1991:ISB

[Sib91] Masaaki Sibuya. An identity for sums of binomial coefficients. SIAM Review, 33 (1):116, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sibuya:1995:BRR

[Sib95] Yasutaka Sibuya. Book review: The Riemann–Hilbert Problem (D. V. Anosov and A. A. Bolibruch). SIAM Review, 37(4):637–639, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Siekmann:1961:CIO

[Sie61] Julius Siekmann. Concerning an integral occurring in aerofoil theory. SIAM Review, 3 (3):243–246, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Siekmann:1965:NCS

[Sie65] Julius Siekmann. Note on the calculation of the suction force. SIAM Review, 7(3): 349–355, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Siegmund:1969:BRI

[Sie69]  David Siegmund.  Book

review: An Introduction to Probability Theory and Its Applications, Volume 2 (William Feller). SIAM Review, 11(2):295–297, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Siegmund:1989:BRS

[Sie89]  David Siegmund.  Book

review: Stopped Random Walks: Limit Theorems and Applications (Allan Gut). SIAM Review, 31(4):709– 710, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Siegmund:1990:BRP

[Sie90] David Siegmund. Book review: Probability Approximations via the Poisson Clumping Heuristic (D. Aldous). SIAM Review, 32(1): 186–187, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sigillito:1967:PBS

[Sig67] Vincent G. Sigillito. Pointwise bounds for solutions of semilinear parabolic equations. SIAM Review, 9(3): 581–585, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sikorski:1992:BRH

[Sik92] K. Sikorski. Book review:

Handbook of Numerical Analysis. Vol. 1: Finite Difference Methods (Part 1) and Solution of Equations in R^{n} (Part 1) (P. G. Ciarlet and J L. Lions, eds.). SIAM Review, 34(1):143–146, March 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Silberman:1961:ASC

[Sil61]  Isidore Silberman.  An-

other satellite communications problem. SIAM Review, 3(2):179, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Silver:1965:SML

[Sil65] Roland Silver. The shape of Milner’s lamp. SIAM Review, 7(4):560, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Silvestri:1966:CPC

[Sil66]  A. Silvestri.  Connect-

ing points in a cubic lattice. SIAM Review, 8(4): 525, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Silvestri:1967:CPC

[Sil67]  A. Silvestri.  Connect-

ing points in a cubic lattice. SIAM Review, 9(4): 751–752, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Silber:1978:BRG

[Sil78] R. Silber. Book review: The Game of Slash (D. N. Berman). SIAM Review, 20 (1):183–184, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simmonds:1970:LFI

[Sim70] James G. Simmonds. A Legendre function inequality. SIAM Review, 12(4): 581, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simmonds:1971:AVM

[Sim71a] James G. Simmonds. Angular velocity and moment potential for a rigid body. SIAM Review, 13(4):571, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simmonds:1971:LFI

[Sim71b] James G. Simmonds. A Legendre function inequality. SIAM Review, 13(4): 577–578, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simmons:1975:RPR

[Sim75] G. J. Simmons. The regiment problem revisited. SIAM Review, 17(1):167, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simmons:1976:BRR

[Sim76] G. J. Simmons. Book review: The Regiment Problem Revisited (G. J. Simmons). SIAM Review, 18(2): 300–301, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simons:1982:BRS

[Sim82] Gordon Simons. Book review: Sequential Methods in Statistics (G. Barrie Wetherill). SIAM Review, 24(3): 360–361, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simmonds:1989:BRS

[Sim89] James G. Simmonds. Book review: Shell Theory: General Methods of Construction (I. N. Vekua; Ts. Gabeskiria, trans.). SIAM Review, 31(1): 161–164, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simo:1993:BRM

[Sim93a] J. C. Simo. Book review: Mixed and Hybrid Finite Element Methods (Franco Brezzi and Michel Fortin). SIAM Review, 35(2):336– 339, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simo:1993:MHF

[Sim93b] J. C. Simo. Mixed and hybrid finite element methods (Franco Brezzi and Michel Fortin). SIAM Review, 35 (2):336–??, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simmonds:1996:AFI

[Sim96]  J. G. Simmonds.  Ana-

lytic functions, ideal fluid flow, and Bernoulli’s equation. SIAM Review, 38 (4):666–667, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29300.

Simoncini:2016:CML

[Sim16a] V. Simoncini. Computational methods for linear matrix equations. SIAM Review, 58 (3):377–441, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simpson:2016:BCB

[Sim16b] D. J. W. Simpson. Bordercollision bifurcations in R^{N}. SIAM Review, 58(2):177– 226, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Singer:1959:BRD

[Sin59] T. Singer. Book review:

Digital Computer Programming (D. D. McCracken). SIAM Review, 1(2):176– 177, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sinden:1961:SC

[Sin61] Frank W. Sinden. Satellite communications. SIAM Review, 3(2):179, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Singleton:1969:BRC

[Sin69] Robert Singleton. Book review: A Continued Fraction Representation of Eigenvalues (Thomas E. Phipps, Jr.). SIAM Review, 11(4): 622–624, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sinha:1976:NCA

[Sin76] A. S. C. Sinha. New concept in attitude control of AMCD spacecraft. SIAM Review, 18 (4):827, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Singmaster:1979:RDC

[Sin79a]  D. Singmaster.  Resis-

tances in an n-dimensional cube. SIAM Review, 21 (4):559, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sinha:1979:BRO

[Sin79b]  S. K. Sinha.  Book re-

view: Outliers in Statistical Data (Vic Barnett and Toby Lewis). SIAM Review, 21(4): 576–577, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Singmaster:1980:ES

[Sin80] D. Singmaster. An extremal sphere. SIAM Review, 22 (3):364, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Singmaster:1986:MS

[Sin86] David Singmaster. Maximum stability. SIAM Review, 28 (4):565, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Singpurwalla:1988:FIR

[Sin88] Nozer D. Singpurwalla. Foundational issues in reliability and risk analysis. SIAM Review, 30(2):264–282, June 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Singpurwalla:1989:UPS

[Sin89] Nozer D. Singpurwalla. A unifying perspective on statistical modeling. SIAM Review, 31(4):560–564, December 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Singpurwalla:1993:BRS

[Sin93a] Nozer D. Singpurwalla. Book review: Statistical Analysis of Reliability Data (M. J. Crowder, A. C. Kimber, R. L. Smith, and T. J. Sweeting). SIAM Review, 35 (3):535–538, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sinha:1993:BRM

[Sin93b] Bikas Kumar Sinha. Book review: Minimax Solutions in Sampling from Finite Populations (Siegfried Gabler). SIAM Review, 35(2):339– 340, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sinha:1993:MSS

[Sin93c] Bikas Kumar Sinha. Minimax solutions in sampling from finite populations (Siegfried Gabler). SIAM Review, 35(2):339–??, June 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sithigh:1992:BRM

[Sit92] Gearoid P. Mac Sithigh. Book review: Mathematical Elasticity Volume One: Three-Dimensional Elasticity (Phillipe G. Ciarlet). SIAM Review, 34(3):513– 514, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sivazlian:1968:TDI

[Siv68] B. D. Sivazlian. Two double integrals. SIAM Review, 10(3):377, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sivazlian:1969:GDM

[Siv69] B. D. Sivazlian. The generalized Dirichlet’s multiple integral. SIAM Review, 11(2): 285–288, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sivasubramanian:1972:IE

[Siv72] A. Sivasubramanian. An integral equation. SIAM Review, 14(4):645, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sivaganesan:1994:BRP

[Siv94a] S. Sivaganesan. Book review: Predictive Inference: an Introduction (Seymour Geisser). SIAM Review, 36 (3):519–520, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sivaganesan:1994:IBB

[Siv94b] S. Sivaganesan. An introduction to the bootstrap (Bradley Efron and Robert J Tibshirani). SIAM Review, 36(4):677–??, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sivaganesan:1994:PII

[Siv94c]  S. Sivaganesan.  Predic-

tive inference: an introduction (Seymour Geisser). SIAM Review, 36(3):519– ??, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sivaganesan:1994:BRI

[Siv94d] Siva Sivaganesan. Book review: An Introduction to the Bootstrap (Bradley Efron and Robert J. Tibshirani). SIAM Review, 36(4):677– 678, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shantaram:1979:TED

[SJ79] R. Shantaram and A. A. Jagers. Two equal determinants. SIAM Review, 21 (3):401, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sloan:1996:LMM

[SJG96] I. H. Sloan, S. Joe, and Alan Genz. Lattice methods for multiple integration. SIAM Review, 38(1): 165–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sameh:1976:SPL

[SK76a] A. H. Sameh and D. J. Kuck. Stable parallel linear-system solvers. SIAM Review, 18 (4):825, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Skeel:1976:IMD

[SK76b] R. Skeel and A. Kong. Imitating multi-derivative multistep formulas. SIAM Review, 18(4):827, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Snyder:1978:CSN

[SK78] W. V. Snyder and F. T. Krogh. Comparison of software for numerical quadrature. SIAM Review, 20 (3):634, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sabharwal:1979:CF

[SK79] Y. P. Sabharwal and J. Kumar. Credibility functions. SIAM Review, 21 (2):257, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sekiguchi:1980:BRM

[SK80] T. Sekiguchi and N. Kimura.

Book review: A Matrix Eigenvalue Problem (G. Efroymson, A. Steger and S. Steinberg). SIAM Review, 22 (1):99–100, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sharma:1985:ERT

[SK85] A. Sharma and Murray S.

Klamkin. Extension of Routh’s theorem to spherical triangles. SIAM Review, 27 (3):446–447, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shang:2020:TCF

[SK20] Xiaocheng Shang and Martin Kröger. Time correlation functions of equilibrium and nonequilibrium Langevin dynamics: Derivations and numerics using random numbers. SIAM Review, 62(4): 901–935, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Skeel:1978:SSL

[Ske78] R. D. Skeel. Scaling systems of linear equations. SIAM Review, 20(3):634, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Skeel:1995:BRN

[Ske95] Robert D. Skeel. Book review: Numerical Hamiltonian Problems (J. M. Sanz-Serna and M. P. Calvo). SIAM Review, 37(2):277– 279, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sekiguchi:1980:MEP

[SKG^{+}80] T. Sekiguchi, N. Kimura, C. Givens, O. Ruehr, M. H. Gutknecht, W. B. Jordan, O. P. Lossers, P. J. Nikolai, and K. Wang. Matrix eigenvalue problem. SIAM Review, 22(1):99–100, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steutel:1967:BRM

[SKGM67] F. W. Steutel, S. S. Kapur, H. W. Gould, and A. Makowski. Book review: A Multinomial Summation (Paul Brock). SIAM Review, 9(2):252–253, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Skinner:1995:ISB

[Ski95]  L. A. Skinner.  Infinite

series of binomial coefficients. SIAM Review, 37 (1):99, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stepien:2020:MCU

[SKK20] Tracy L. Stepien, Eric J.

Kostelich, and Yang Kuang. Mathematics + cancer: an undergraduate “bridge” course in applied mathematics. SIAM Review, 62(1):244– 263, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Surace:2019:HAC

[SKP19] Simone Carlo Surace, Anna Kutschireiter, and Jean-Pascal Pfister. How to avoid the curse of dimensionality: Scalability of particle filters with and without importance weights. SIAM Review, 61 (1):79–91, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Skufca:2004:ASM

[Sku04] Joseph D. Skufca. Analysis still matters: a surprising instance of failure of Runge–Kutta–Felberg ODE solvers. SIAM Review, 46 (4):729–737, December 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42911.

Skufca:2005:KWC

[Sku05a] J. D. Skufca. k workers in a circular warehouse: A random walk on a circle, without passing. SIAM Review, 47(2): 301–314, ???? 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Skufca:2005:WCW

[Sku05b] Joseph D. Skufca. k workers in a circular warehouse: a random walk on a circle, without passing. SIAM Review, 47(2): 301–314, June 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/43201.

Slater:1964:GA

[Sla64] Morton L. Slater. Gravitational attraction. SIAM Review, 6(2):177, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slepian:1962:BRB

[Sle62] David Slepian. Book review:

On a Binomial Identity Arising from a Sorting Problem (Paul Brook). SIAM Review, 4(4):396–398, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slepian:1967:BRN

[Sle67]  Paul Slepian.  Book re-

view: Network Theory: Analysis and Synthesis (Shlomo Karni). SIAM Review, 9(2): 260–262, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sleator:1974:BRA

[Sle74] F. B. Sleator. Book review: Applied Mathematics: an Introduction (Harry Pollard). SIAM Review, 16(2): 268–269, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slemrod:1976:ABV

[Sle76] M. Slemrod. Asymptotic behavior of a vibrating membrane with nonlinear boundary damping — application of theory of nonlinear semigroups. SIAM Review, 18(4): 827–828, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slepian:1978:ED

[Sle78] D. Slepian. Evaluation of a determinant. SIAM Review, 20(3):594, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slemrod:1979:BRM

[Sle79a] Marshall Slemrod. Book review: Modern Methods in Partial Differential Equations: an Introduction (Martin Schechter). SIAM Review, 21(4):577, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slepian:1979:ED

[Sle79b] D. Slepian. Evaluation of a determinant. SIAM Review, 21(3):400–401, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slemrod:1981:BRD

[Sle81] M. Slemrod. Book review: Dynamical Systems and Evolution Equations; Theory and Applications (J. A. Walker). SIAM Review, 23(3):398– 399, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slemrod:1983:BRP

[Sle83a] M. Slemrod. Book review: Power System Stability (M. A. Pai). SIAM Review, 25 (3):419–420, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slepian:1983:SCF

[Sle83b] David Slepian. Some comments on Fourier analysis, uncertainty and modeling. SIAM Review, 25(3): 379–393, July 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slepian:1984:MNB

[Sle84a] P. Slepian. The mathematics of networks — Burr, S. A. SIAM Review, 26(4): 595–596, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slepian:1984:BRM

[Sle84b]  Paul Slepian.  Book re-

view: The Mathematics of Networks (Stefan A. Burr). SIAM Review, 26(4):595– 596, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slemrod:1985:BRC

[Sle85] Marshall Slemrod. Book review: Compressible Fluid Flow and Systems of Conservation Laws in Several Space Variables (A. Majda). SIAM Review, 27(3): 472–473, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slemrod:1986:BRR

[Sle86] Marshall Slemrod. Book review: Rational Thermodynamics, Second Edition (C. Truesdell). SIAM Review, 28 (1):126–127, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slemrod:1987:BRH

[Sle87] Marshall Slemrod. Book review: Homogenization and Effective Moduli of Materials and Media (J. L. Ericksen, D. Kinderlehrer, R. Kohn and J.-L. Lions, eds.). SIAM Review, 29(3):497– 498, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sleeman:1989:BRA

[Sle89a] B. D. Sleeman. Book review: Applied Analysis (Allan M. Krall). SIAM Review, 31(3): 514–515, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Slemrod:1989:BRS

[Sle89b] Marshall Slemrod. Book review: Sensors and Controls in the Analysis of Distributed Systems (A. El Jai and A. J. Pritchard). SIAM Review, 31(4):710, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sleeman:1992:BRT

[Sle92] B. D. Sleeman. Book review:

Two Parameter Eigenvalue Problems in Ordinary Differential Equations (M. Faierman). SIAM Review, 34 (4):684–687, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sleeman:1994:BRI

[Sle94] B. D. Sleeman. Book review:

Inverse Acoustic and Electromagnetic Scattering Theory (D. Colton and R. Kress). SIAM Review, 36(3):520– 523, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spital:1966:BRT

[SLF66] Sidney Spital, D. L. Lansing, and H. E. Fettis. Book review: A Third Order Differential Equation (Donald E. Amos). SIAM Review, 8(3): 386–388, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sloan:1993:BRR

[Slo93] Ian H. Sloan. Book review:

Random Number Generation and Quasi-Monte Carlo Methods (H. Niederreiter). SIAM Review, 35(4):680– 681, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sloan:1997:BRI

[Slo97] Ian H. Sloan. Book review: Integral Equations: Theory and Numerical Treatment, by Wolfgang Hackbusch. SIAM Review, 39(2):360, June 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spital:1965:LS

[SM65] Sidney Spital and J. K. Mackenzie. Late solutions. SIAM Review, 7 (2):274, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanderson:1978:RSF

[SM78] J. G. Sanderson and G. N. Minerbo. Reconstruction of a source from a few (2 or 3) projections. SIAM Review, 20(3):634, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sharma:1979:BRT

[SM79] A. Sharma and K. K. Mathur. Book review: Two Recurrence Relations for Hermite Basis Polynomials (J. C. Cavendish and W. W. Meyer). SIAM Review, 21(1): 143–144, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sabharwal:1984:PSS

[SM84] Y. P. Sabharwal and V. K.

Malhotra. Patterns in a sequence of symmetric Bernoulli trials. SIAM Review, 26(3):430, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spanier:1994:QRM

[SM94] Jerome Spanier and Earl H.

Maize. Quasi-random methods for estimating integrals using relatively small samples. SIAM Review, 36(1): 18–44, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/22673.htm.

Strang:2014:FDM

[SM14] Gilbert Strang and Shev MacNamara. Functions of difference matrices are Toeplitz plus Hankel. SIAM Review, 56(3):525–546, ???? 2014. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strang:2022:LCE

[SM22] Gilbert Strang and Cleve Moler. LU and CR elimination. SIAM Review, 64 (1):181–190, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1358694.

Small:1976:BRS

[Sma76] R. D. Small. Book review: Simultaneous Iteration Towards All Roots of a Complex Polynomial (M. W. Green, A. J. Korsak and M. C. Pease). SIAM Review, 18(3): 501–502, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Small:1983:PGC

[Sma83] Robert D. Small. Population growth in a closed system. SIAM Review, 25(1): 93–95, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smale:1990:SRF

[Sma90] Steve Smale. Some remarks on the foundations of numerical analysis. SIAM Review, 32(2):211–220, June 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1971:FST

[Smi71] S. H. Smith. A Fourier sine transform. SIAM Review, 13 (2):247, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1972:LTA

[Smi72] P. Smith. Laplace transforms of Airy functions. SIAM Review, 14(4):644, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smillie:1974:BRC

[Smi74] K. W. Smillie. Book review: Computer Approaches to Mathematical Problems (J. Nievergelt, J. C. Farrar and E. M. Reingold). SIAM Review, 16(3):409, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1975:MMS

[Smi75] Donald R. Smith. The multivariable method in singular perturbation analysis. SIAM Review, 17(2): 221–273, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1980:SLI

[Smi80a] P. W. Smith. A set of linearly independent polynomials. SIAM Review, 22 (2):229, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1980:BRP

[Smi80b] Philip W. Smith. Book review: A Practical Guide to Splines (Carl de Boor).

SIAM Review, 22(4):520– 521, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1981:SLI

[Smi81] P. W. Smith. A set of linearly independent polynomials. SIAM Review, 23 (3):390, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1982:BRC

[Smi82] Carl H. Smith. Book review: Computability: an Introduction to Recursive Function Theory (N. J. Cutland). SIAM Review, 24 (1):98, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smillie:1983:BRC

[Smi83] K. W. Smillie. Book review: Charles Babbage, Pioneer of the Computer, by Anthony Hyman. SIAM Review, 25(3):422–423, July 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2029408.

Smith:1986:BRW

[Smi86] Philip W. Smith. Book review: n-Widths in Approximation Theory (Allan Pinkus). SIAM Review, 28 (2):283–284, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1987:DOR

[Smi87] Donald R. Smith. Decoupling and order reduction via the Riccati transformation. SIAM Review, 29(1): 91–113, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1988:C

[Smi88a] D. A. Smith. Correction.

SIAM Review, 30(4):623– 624, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1988:SOD

[Smi88b] Hal L. Smith. Systems of ordinary differential equations which generate an order preserving flow. A survey of results. SIAM Review, 30(1): 87–113, March 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1988:EPD

[Smi88c] P. Smith. The entropy of a Poisson distribution. SIAM Review, 30(2):317, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1997:CNT

[Smi97] S. H. Smith. Classroom note: Time-dependent Poiseuille flow. SIAM Review, 39(3): 511–513, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30113.

Smoller:1991:ALP

[Smo91a] Joel A. Smoller. The analysis of linear partial differential operators, I (Lars Hormander). SIAM Review, 33 (2):338–??, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smoller:1991:BRA

[Smo91b] Joel A. Smoller. Book review: The Analysis of Linear Partial Differential Operators, I (Lars Hörmander). SIAM Review, 33(2):338, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seshadri:1996:IGD

[SN96] V. Seshadri and H. N. Nagaraja. The inverse Gaussian distribution: a case study in exponential families. SIAM Review, 38(2): 350–??, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strang:2004:IRS

[SN04] Gilbert Strang and Tri Nguyen. The interplay of ranks of submatrices. SIAM Review, 46(4):637– 646, December 2004. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43438.

Sneddon:1967:EII

[Sne67]  Ian N. Sneddon.  The

evaluation of an integral involving the product of two Gegenbauer polynomials. SIAM Review, 9(3): 569–572, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Snow:1976:BRF

[Sno76] Donald R. Snow. Book review: Functional Analysis and Time Optimal Control (Henry Hermes and Joseph P. LaSalle). SIAM Review, 18 (2):307–308, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scheinok:1966:BRL

[SNS66] Perry Scheinok, D. J. Newman, and L. A. Shepp. Book review: The Limit of a Product (D. J. Newman and L. A. Shepp). SIAM Review, 8(4): 540–542, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Snyder:1997:SCN

[Sny97a] J. Snyder. Stability considerations for numerical methods. SIAM Review, 39(4): 755–760, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Snyder:1997:CNS

[Sny97b] Johnny Snyder. Classroom note: Stability considerations for numerical methods. SIAM Review, 39(4): 755–760, December 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30590.

Shepp:1978:PI

[SO78] L. A. Shepp and A. M. Odlyzko. A probabilistic inequality. SIAM Review, 20 (4):855, October 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sofo:1996:RRP

[Sof96] Anthony Sofo. Real roots of polynomials. SIAM Review, 38(3):515, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sofo:1997:RRP

[Sof97] A. Sofo. Real roots of polynomials. SIAM Review, 39 (4):765–772, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sollfrey:1967:BFI

[Sol67] William Sollfrey. On a Bessel function integral. SIAM Review, 9(3):586–589, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Somerville:1978:CSL

[Som78] R. C. J. Somerville. Current status of large-scale numerical weather prediction. SIAM Review, 20(3):635, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sommerer:1979:TAC

[Som79] John C. Sommerer. The toy airplane catastrophe. SIAM Review, 21(2):247– 251, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sommese:1993:BRI

[Som93] Andrew J. Sommese. Book review: Ideals, Varieties, and Algorithms — an Introduction to Computational Algebraic Geometry and Commutative Algebra (D. Cox, J. Little, and D. O’Shea).

SIAM Review, 35(4):682, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Soper:1959:BRS

[Sop59] Willis A. Soper, Jr. Book review: Surveys in Mechanics (G. K. Batchelor and R. M. Davies). SIAM Review, 1 (1):73–74, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sopka:1960:BRO

[Sop60] John J. Sopka. Book review: Operations Research-Methods and Problems (M. Sasieni, A. Yaspan, L. Friedman). SIAM Review, 2(2): 159–160, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sopka:1962:BRM

[Sop62] John J. Sopka. Book review: Mathematical Programming (S. Vajda). SIAM Review, 4 (2):160–161, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sopka:1965:BRO

[Sop65] John J. Sopka. Book review: Operational Calculus and Generalized Functions (Arthur Erdélyi). SIAM Review, 7(3):437–438, ???? 1965. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Sopka:1969:BRF

[Sop69] John J. Sopka. Book review: Functional Analysis in Normed Spaces (L. V. Kantorovich and G. P. Akilov). SIAM Review, 11(3): 412–413, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sopka:1972:BRP

[Sop72] John J. Sopka. Book review:

Principles of Functional Analysis (Martin Schechter). SIAM Review, 14(4):668– 669, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sopka:1979:BRI

[Sop79] John J. Sopka. Book review: Introductory Functional Analysis with Applications (Erwin Kreyszig). SIAM Review, 21(3):412– 413, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sopka:1982:BRI

[Sop82] John J. Sopka. Book review: Introduction To Functional Analysis; 2nd edition (Angus E. Taylor and David C. Lay). SIAM Review, 24 (1):89, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Soukup:1976:ASL

[Sou76] J. Soukup. Algorithm for a straight-line representation of planar graphs and its application to PERT-type networks. SIAM Review, 18 (4):828, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sourisseau:1980:MSP

[Sou80] K. Sourisseau. A matrix stability problem. SIAM Review, 22(1):97–98, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stvincent:1982:FE

[SP82] M. Stvincent and K. R. Parthasarathy. A functional equation. SIAM Review, 24 (1):78–80, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sun:2001:MVF

[SP01] Xiaobai B. Sun and Nikos P. Pitsianis. A matrix version of the fast multipole method. SIAM Review, 43(2):289–300, June 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37083.

Spang:1962:RMT

[Spa62a] H. A. Spang III. A review of minimization techniques for nonlinear functions. SIAM Review, 4(4): 343–365, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spanier:1962:UAM

[Spa62b] J. Spanier. A unified approach to Monte Carlo methods and an application to a multigroup calculation of absorption rates. SIAM Review, 4(2):115–134, April 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2028365.

Spanier:1977:EAM

[Spa77a] J. Spanier. Education in applied mathematics — Claremont Mathematics Clinic. SIAM Review, 19(3):536– 550, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spanier:1977:CMC

[Spa77b] Jerome Spanier. The Claremont Mathematics Clinic. SIAM Review, 19(3):536– 549, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spath:1997:BRH

[Spä97] Helmut Späth. Book review: Handbook on Splines for the User, by Eugene V. Shikin and Alexander I. Plis. SIAM Review, 39(2): 360–362, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sloane:1996:EIS

[SPBC96] N. J. A. Sloane, Simon Plouffe, J. M. Borwein, and R. M. Corless. The encyclopedia of integer sequences. SIAM Review, 38 (2):333–??, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spira:1971:NRS

[Spi71] Robert Spira. Noncomplete residue system. SIAM Review, 13(1):115, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spitzer:1974:BRF

[Spi74] F. Spitzer. Book review: Fourier Analysis in Probability Theory (T. Kawata). SIAM Review, 16(4):566, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spinetto:1976:FLP

[Spi76] R. D. Spinetto. A facility location problem. SIAM Review, 18(2):294–295, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spitzer:1980:BRB

[Spi80] Frank Spitzer. Book review:

Brownian Motion and Classical Potential Theory (Sydney C. Port and Charles J. Stone). SIAM Review, 22(2): 247–248, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spitzer:1981:BRO

[Spi81a] Frank Spitzer. Book review:

Optimal Stopping Rules (A. N. Shiryayev). SIAM Review, 23(1):117, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spitzer:1981:BRR

[Spi81b] Frank Spitzer. Book review:

Reversibility and Stochastic Networks (F. P. Kelly). SIAM Review, 23(3):400– 401, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spitzer:1982:BRC

[Spi82] Frank Spitzer. Book review:

A Course in the Theory of Stochastic Processes (A. D. Wentzell). SIAM Review, 24 (3):361–362, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spillers:1984:CES

[Spi84a] W. R. Spillers. Civil engineering systems-analysis and design — Smith, A. A., Hinton, E., Lewis, R. W. SIAM Review, 26(4): 610–611, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spillers:1984:BRC

[Spi84b] William R. Spillers. Book review: Civil Engineering Systems Analysis and Design (Alan A. Smith, Ernest Hinton and Roland W. Lewis). SIAM Review, 26(4):610– 611, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spinetto:1986:OED

[Spi86] Richard D. Spinetto. An optimal error distribution problem. SIAM Review, 28(2): 231, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spillers:1987:BRA

[Spi87] W. R. Spillers. Book review: Application of Structural Systems Reliability Theory (Palls Thoft-Christensen and Yoshisada Murotsu). SIAM Review, 29(4):674– 675, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spillers:1991:BRI

[Spi91a] W. R. Spillers. Book review:

An Introduction to the History of Structural Mechanics, Part I: Statics and Resistance of Solids, Part II: Vaulted Structures and Elastic Systems (Edoardo Benvenuto). SIAM Review, 33 (4):679–680, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spillers:1991:IHS

[Spi91b] W. R. Spillers. An introduction to the history of structural mechanics Part I: Statics and resistance of solids, Part II: Vaulted structures and elastic systems (Edoardo Benvenuto). SIAM Review, 33(4):679– ??, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spohn:1985:BRN

[Spo85] Herbert Spohn. Book review: Non-Equilibrium Entropy and Irreversibility (G. Lindblad). SIAM Review, 27 (1):125–127, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spohn:1988:BRI

[Spo88] Herbert Spohn. Book review: Introduction to Modern Statistical Mechanics (David Chandler). SIAM Review, 30 (4):680, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1976:ACE

[SPR76] F. Stenger, W. Petrick, and Z. Rotsides. Algorithm for computing electromagnetic scattered field from an axially-symmetric body with an impedance boundary condition. SIAM Review, 18(4): 828–829, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Spurrier:1989:BRM

[Spu89] John D. Spurrier. Book review: Multiple Comparison Procedures (Yosef Hochberg and Ajit C. Tamhane). SIAM Review, 31(2):349– 350, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Squire:1961:FP

[Squ61] William Squire. Flame propagation. SIAM Review, 3(2): 178–179, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Squire:1967:GFQ

[Squ67] W. Squire. Gauss–Fourier quadrature. SIAM Review, 9(4):749, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Simmons:1970:PB

[SR70] G. J. Simmons and D. Rawlinson. Palindromic biquadrates. SIAM Review, 12 (2):287, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Siegmann:1975:DBN

[SR75] W. L. Siegmann and L. A.

Rubenfeld. Dynamic behavior near onset of doublediffusive convection. SIAM Review, 17(2):389, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shor:1979:BRC

[SR79] Peter Shor and Otto G.

Ruehr. Book review: A Combinatorial Identity (Peter Shor). SIAM Review, 21 (2):258–260, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schuur:1993:URC

[SR93a] P. C. Schuur and R. Richberg. An unexpected result in classical electrostatics. SIAM Review, 35(4):

643–645, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schuur:1993:BRU

[SR93b] P. C. Schuur and Rolf Richberg. Book review: An Unexpected Result in Classical Electrostatics (A. J. E. M. Janssen and J. B. M. Melissen). SIAM Review, 35(4): 643–645, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seliktar:1994:FHW

[SR94] Miriam Seliktar and Chris Rorres. The flow of hot water from a distant hotwater tank. SIAM Review, 36 (3):474–479, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/25826.
htm.

Shklyar:1998:CSO

[SR98] Alexander Ya Shklyar and Michael Renardy. Complete second order linear differential equations in Hilbert spaces. SIAM Review, 40(1): 175–??, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Srivastava:1973:SI

[Sri73] H. M. Srivastava. A summation identity. SIAM Review, 15(2):376, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Srivastava:1975:SHF

[Sri75] H. M. Srivastava. A series of hypergeometric functions. SIAM Review, 17(4): 685–686, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Srivastava:1984:WFI

[Sri84] H. M. Srivastava. A Weyl fractional integral. SIAM Review, 26(2):272–273, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Srivastava:1985:IDH

[Sri85] H. M. Srivastava. An identity for double hypergeometric series. SIAM Review, 27 (4):575–576, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shampine:1999:SID

[SRK99] Lawrence F. Shampine, Mark W. Reichelt, and Jacek A. Kierzenka. Solving index-1 DAEs in MATLAB and Simulink. SIAM Review, 41(3):538–552, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/33425.

Smith:1961:NEC

[SS61] E. H. Smith and D. E. Stone. A note on the expected coverage of one circle by another, the case of offset aim. SIAM Review, 3 (1):51–53, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sheldon:1962:ROC

[SS62] J. W. Sheldon and D. P. Squier. Remarks on the order of convergence of discrete analogs for secondorder elliptic partial differential equations. SIAM Review, 4(4):366–378, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saaty:1965:NPM

[SS65] T. L. Saaty and G. Suzuki. A nonlinear programming model in optimum communication satellite use. SIAM Review, 7(3):403–408, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sivazlian:1970:BRT

[SS70] B. D. Sivazlian and Kasum Soni. Book review: Two Double Integrals (B. D. Sivazlian). SIAM Review, 12(1): 146–150, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Starner:1978:EID

[SS78] J. W. Starner and W. L.

Shepherd. Efficient inverse distribution type normal random number generator. SIAM Review, 20 (3):635, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saunders:1979:AGS

[SS79] B. David Saunders and Hans Schneider. Applications of the Gordan–Stiemke theorem in combinatorial matrix theory. SIAM Review, 21(4): 528–541, October 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Scott:1988:GS

[SS88] Dan M. Scott and Gary J.

Sherman. The great shootout. SIAM Review, 30(3):492– 497, September 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Segel:1989:QSS

[SS89] Lee A. Segel and Marshall Slemrod. The quasisteady-state assumption: a case study in perturbation. SIAM Review, 31(3):446– 477, September 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schmidt:1991:EMO

[SS91] Frank Schmidt and Rodica Simion. Even minus odd involutions in the symmetric group. SIAM Review, 33(2): 276–277, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strogatz:1995:NDC

[SS95] Steven H. Strogatz and Douglas S. Shafer. Nonlinear dynamics and chaos: With applications to physics biology, chemistry, and engineering. SIAM Review, 37(2): 280–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Straffin:1996:GTS

[SS96] Philip D. Straffin and Donald G. Saari. Game theory and strategy. SIAM Review, 38(4):698–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shirvani:1998:SLD

[SS98] Mazi Shirvani and Joseph W.-H. H. So. Solutions of linear differential algebraic equations. SIAM Review, 40(2):344–346, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31369.

Simoncini:2005:OSC

[SS05] Valeria Simoncini and Daniel B.

Szyld. On the occurrence of superlinear convergence of exact and inexact Krylov subspace methods. SIAM Review, 47(2): 247–272, June 2005. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/42443.

Sanz-Serna:2016:SRK

[SS16] J. M. Sanz-Serna. Symplectic Runge–Kutta schemes for adjoint equations, automatic differentiation, optimal control, and more. SIAM Review, 58(1):3–33, ???? 2016. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2017:SRa

[SS17a] J. M. Sanz-Serna. Survey and review. SIAM Review, 59(1):1, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2017:SRb

[SS17b] J. M. Sanz-Serna. Survey and review. SIAM Review, 59(2):223, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2017:SRc

[SS17c] J. M. Sanz-Serna. Survey and review. SIAM Review, 59(3):467, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2017:SRd

[SS17d] J. M. Sanz-Serna. Survey and review. SIAM Review, 59(4):701, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2018:SRa

[SS18a] J. M. Sanz-Serna. Survey and review. SIAM Review, 60(1):1, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2018:SRb

[SS18b] J. M. Sanz-Serna. Survey and review. SIAM Review, 60(2):221, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2018:SRc

[SS18c] J. M. Sanz-Serna. Survey and review. SIAM Review, 60(3):489, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2018:SRd

[SS18d] J. M. Sanz-Serna. Survey and review. SIAM Review, 60(4):777, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2019:SRa

[SS19a] J. M. Sanz-Serna. Survey and review. SIAM Review, 61(1):1, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2019:SRb

[SS19b] J. M. Sanz-Serna. Survey and review. SIAM Review, 61(2):227, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2019:SRc

[SS19c] J. M. Sanz-Serna. Survey and review. SIAM Review, 61(3):401, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2019:SRd

[SS19d] J. M. Sanz-Serna. Survey and review. SIAM Review, 61(4):641, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2020:SRa

[SS20a] J. M. Sanz-Serna. Survey and review. SIAM Review, 62(1):1, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2020:SRb

[SS20b] J. M. Sanz-Serna. Survey and review. SIAM Review, 62(2):299, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2020:SRc

[SS20c] J. M. Sanz-Serna. Survey and review. SIAM Review, 62(3):527, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2020:SRd

[SS20d] J. M. Sanz-Serna. Survey and review. SIAM Review, 62(4):741, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2021:SRa

[SS21a] J. M. Sanz-Serna. Survey and review. SIAM Review, 63(1):1, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2021:SRb

[SS21b] J. M. Sanz-Serna. Survey and review. SIAM Review, 63(2):247, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2021:SRc

[SS21c] J. M. Sanz-Serna. Survey and review. SIAM Review, 63(3):433, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2021:SRd

[SS21d] J. M. Sanz-Serna. Survey and review. SIAM Review, 63(4):651, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:2022:SRa

[SS22a]  J. M. Sanz-Serna.  Sur-

vey and review. SIAM Review, 64(1):1, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975378.

Sanz-Serna:2022:SRb

[SS22b]  J. M. Sanz-Serna.  Sur-

vey and review. SIAM Review, 64(2):227, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N975421.

Sanz-Serna:2022:SRc

[SS22c]  J. M. Sanz-Serna.  Sur-

vey and review. SIAM Review, 64(3):515, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22N975500.

Sanz-Serna:2022:SRd

[SS22d]  J. M. Sanz-Serna.  Sur-

vey and review. SIAM Review, 64(4):761, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22N97555X.

Silva:2024:DSG

[SS24] Hannelore De Silva and Karl Sigmund. Dynamics of signaling games. SIAM Review, 66(2):368–387, May 2024. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanz-Serna:1995:NHP

[SSCS95] J. M. Sanz-Serna, M. P. Calvo, and Robert D. Skeel. Numerical Hamiltonian problems. SIAM Review, 37(2): 277–??, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shaked:1995:SOT

[SST95] Moshe Shaked, J. George Shanthikumar, and Y. L.

Tong. Stochastic orders and their applications. SIAM Review, 37(3):477–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Schrijver:1999:RLP

[SSW99] Alexander Schrijver, Paul Seymour, and Peter Winkler. The ring loading problem. SIAM Review, 41 (4):777–791, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35647.

StVincent:1982:BRF

[St.82] M. St. Vincent. Book review:

A Functional Equation (K. R. Parthasarathy). SIAM Review, 24(1):78–80, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stone:1991:ESP

[ST91a] Richard E. Stone and Craig A.

Tovey. Erratum: “The simplex and projective scaling algorithms as iteratively reweighted least squares methods”. SIAM Review, 33(3):461, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [ST91b].

Stone:1991:SPS

[ST91b] Richard E. Stone and Craig A.

Tovey. The simplex and projective scaling algorithms as iteratively reweighted least squares methods. SIAM Review, 33(2):220–237, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See erratum [ST91a].

Shapiro:1994:EPB

[ST94] Harold S. Shapiro and Max Tegmark. An elementary proof that the biharmonic Green function of an eccentric ellipse changes sign. SIAM Review, 36(1):99– 101, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/24859.htm.

Squire:1998:UCV

[ST98] William Squire and George Trapp. Using complex variables to estimate derivatives of real functions. SIAM Review, 40(1):110– 112, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31241.

Stakgold:1966:BRT

[Sta66] Ivar Stakgold. Book review: The Theory of Electromagnetism (D. S. Jones). SIAM Review, 8(3):399– 400, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stakgold:1971:BSN

[Sta71] Ivar Stakgold. Branching of solutions of nonlinear equations. SIAM Review, 13(3): 289–332, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [Sta72].

Stakgold:1972:EBS

[Sta72]  Ivar Stakgold.  Errata:

Branching of Solutions of Nonlinear Equations. SIAM Review, 14(3):492, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Sta71].

Starr:1975:ACS

[Sta75]  J. B. Starr.  Analysis

of control systems utilizing thermoelectric coolers. SIAM Review, 17(2):389– 390, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stallman:1976:IBD

[Sta76]  F. W. Stallman.  Inte-

gral bounds derived from inversion of integral equations. SIAM Review, 18 (4):828, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stanley:1985:BRH

[Sta85] Richard Stanley. Book review: A Hyper-Volume (E. E. Doberkat). SIAM Review, 27 (4):579–580, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stark:1988:BRI

[Sta88]  Henry Stark.  Book re-

view: Image Processing-Continuous to Discrete, Volume 1: Geometric, Transform, and Statistical Methods (Edward R. Dougherty and Charles R. Giardina). SIAM Review, 30(3):526– 527, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stark:1989:BRT

[Sta89]  Henry Stark.  Book re-

view: Transforms for Engineers: a Guide to Signal Processing (I. G. Beauchamp). SIAM Review, 31(3):515– 516, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stark:1990:BRM

[Sta90] Henry Stark. Book review: Morphological Methods in Image and Signal Processing (Charles R. Giardina and Edward R. Dougherty). SIAM Review, 32(2):333, June 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stamp:1992:CBS

[Sta92a] Mark Stamp. Circular binary sequences. SIAM Review, 34 (3):496–497, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stark:1992:BRN

[Sta92b] Jaroslav Stark. Book review:

Non-linear Series: a Dynamical System Approach (Howell Tong). SIAM Review, 34 (1):149–151, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stark:1992:NLS

[Sta92c] Jaroslav Stark. Non-linear series: a dynamical system approach (Howell Tong). SIAM Review, 34(1):149– ??, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stein:1959:OFW

[Ste59] F. Max Stein. Orthogonal functions whose k th derivatives are also orthogonal. SIAM Review, 1(2): 167–170, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steinberg:1961:BWP

[Ste61a] Leon Steinberg. The backboard wiring problem: a placement algorithm. SIAM Review, 3(1):37–50, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stelson:1961:RCL

[Ste61b]  Hugh Stelson.  On the

reducibility of certain linear differential operators. SIAM Review, 3(2):144– 151, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steward:1962:ATA

[Ste62] Donald V. Steward. On an approach to techniques for the analysis of the structure of large systems of equations. SIAM Review, 4(4): 321–342, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stein:1964:SLS

[Ste64]  F. Max Stein.  Sturm–

Liouville systems that generate families. SIAM Review, 6(1):12–19, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stearns:1967:PDN

[Ste67a] Mary Beth Stearns. Probability distribution of a network of triangles. SIAM Review, 9(3):591, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stein:1967:BRI

[Ste67b] S. Stein. Book review: An Incomplete Laplace Transform (P. J. Short). SIAM Review, 9(4):752–754, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1968:BRN

[Ste68a] Frank Stenger. Book review: Numerical Integration (Philip J. Davis and Philip Rabinowitz). SIAM Review, 10(2):239–240, April 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//links.jstor.org/sici?
sici=0036-1445%28196804%
2910%3A2%3C239%3ANI%3E2.
0.CO%3B2-Y.

Steutel:1968:BRT

[Ste68b] F. W. Steutel. Book review:

A Trigonometric Inequality (J. N. Lyness). SIAM Review, 10(2):226–227, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steutel:1970:BRS

[Ste70a] F. W. Steutel. Book review:

A Set of Inequalities (James D. Riley). SIAM Review, 12 (2):290–292, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1970:CSA

[Ste70b] G. W. Stewart. On the convergence of Sebastião E. Silva’s method for finding a zero of a polynomial. SIAM Review, 12(3):458– 460, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1971:BRI

[Ste71] Frank Stenger. Book review: Integrals and Sums (P. C. Chakravarti). SIAM Review, 13(4):582–583, October 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//links.jstor.org/sici?
sici=0036-1445%28197110%
2913%3A4%3C582%3AIAS%3E2.
0.CO%3B2-1.

Stenger:1972:BRQ

[Ste72] Frank Stenger. Book review: Quadrature Formulae (A. Ghizetti and A. Ossicini). SIAM Review, 14(4):662, October 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//links.jstor.org/sici?
sici=0036-1445%28197210%
2914%3A4%3C662%3AQF%3E2.
0.CO%3B2-J.

Stein:1973:MNN

[Ste73a] S. Stein. A minimum number of normals. SIAM Review, 15(3):658, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1973:BRA

[Ste73b] Frank Stenger. Book review: Approximate Calculation of Multiple Integrals (A. H. Stroud). SIAM Review, 15(1):234–235, January 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//links.jstor.org/sici?
sici=0036-1445%28197301%
2915%3A1%3C234%3AACOMI%
3E2.0.CO%3B2-7.

Stewart:1973:EPB

[Ste73c] G. W. Stewart. Error and perturbation bounds for subspaces associated with certain eigenvalue problems. SIAM Review, 15(4):727– 764, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steutel:1974:PI

[Ste74] F. W. Steutel. A probabilistic inequality. SIAM Review, 16 (3):390–391, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steutel:1975:PI

[Ste75] F. W. Steutel. Probabilistic inequality. SIAM Review, 17 (4):693–694, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1976:DCR

[Ste76] G. W. Stewart. Detecting and correcting rank degeneracy. SIAM Review, 18 (4):829, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1977:PPI

[Ste77] G. W. Stewart. On the perturbation of pseudo-inverses, projections, and linear least squares problems. SIAM Review, 19(4):634–662, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1981:NMB

[Ste81a]  Frank Stenger.  Numeri-

cal methods based on Whittaker cardinal, or sinc functions. SIAM Review, 23(2): 165–224, April 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1981:BRN

[Ste81b] William Stenger. Book review: Numerical and Quantitative Analysis (G. Fichera).

SIAM Review, 23(3):411– 412, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewartson:1981:DP

[Ste81c] Keith Stewartson. D’alembert’s paradox. SIAM Review, 23 (3):308–343, July 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steutel:1985:PPB

[Ste85] F. W. Steutel. Poisson processes and a Bessel function integral. SIAM Review, 27 (1):73–77, March 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steele:1987:BRN

[Ste87] J. Michael Steele. Book review: Non-Uniform Random Variate Generation (Luc Devroye). SIAM Review, 29(4): 675–676, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1988:BRC

[Ste88] Frank Stenger. Book review: Computational Complexity (K. Wagner and G. Wechsung). SIAM Review, 30(2):353–354, June 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//links.jstor.org/sici?

sici=0036-1445%28198806%
2930%3A2%3C353%3ACC%3E2.
0.CO%3B2-B.

Stenger:1990:BRR

[Ste90a] Frank Stenger. Book review: Rational Approximation of Real Functions (P. P. Petrushev and V. I. Popov). SIAM Review, 32(1):187–188, March 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//links.jstor.org/sici?
sici=0036-1445%28199003%
2932%3A1%3C187%3ARAORF%
3E2.0.CO%3B2-R.

Stewart:1990:SPT

[Ste90b] G. W. Stewart. Stochastic perturbation theory. SIAM Review, 32(4):579–610, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stenger:1993:BRS

[Ste93a] Frank Stenger. Book review: Sinc Methods for Quadrature and Differential Equations (J. Lund and K. L. Bowers). SIAM Review, 35 (4):682–683, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//links.jstor.org/sici?
sici=0036-1445%28199312%
2935%3A4%3C682%3ASMFQAD%
3E2.0.CO%3B2-F.

Stewart:1993:BRT

[Ste93b] G. W. Stewart. Book review:

Topics in Matrix Analysis (Roger A. Horn and Charles R. Johnson). SIAM Review, 35(1):168–169, March 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1993:EHS

[Ste93c] G. W. Stewart. On the early history of the singular value decomposition. SIAM Review, 35(4):551– 566, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1994:BRS

[Ste94a] Ian Stewart. Book review:

Symmetry in Chaos — a Search for Pattern in Mathematics, Art, and Nature (Michael Field and Martin Golubitsky). SIAM Review, 36(1):145–148, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1994:SCS

[Ste94b] Ian Stewart. Symmetry in chaos — a search for pattern in mathematics, art, and nature (Michael Field and Martin Golubitsky). SIAM Review, 36(1):145– ??, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stensholt:1996:CPV

[Ste96]  Eivind Stensholt.  Circle

pictograms for vote vectors. SIAM Review, 38(1): 96–119, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/27363.htm.

Steinberg:1997:ME

[Ste97a] S. Steinberg. Matrix exponentials. SIAM Review, 39 (2):327–330, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stewart:1997:BRA

[Ste97b] G. W. Stewart. Book review: Accuracy and Stability of Numerical Algorithms, by Nicholas J. Higham. SIAM Review, 39(1):164– 166, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stevenson:1998:ICS

[Ste98] D. E. Stevenson. Introducing computational science methods using parallax. SIAM Review, 40(1): 81–86, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/30713.

Stewart:1999:RPE

[Ste99] Charles V. Stewart. Robust parameter estimation in computer vision. SIAM Review, 41(3):513–537, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/34580.

Stewart:2000:RBD

[Ste00] David E. Stewart. Rigidbody dynamics with friction and impact. SIAM Review, 42(1):3–39, March 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/36011.

Steinbach:2001:MRM

[Ste01] Marc C. Steinbach. Markowitz revisited: Mean-variance models in financial portfolio analysis. SIAM Review, 43(1):31–85, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37665.

Stensholt:2004:STV

[Ste04]  Eivind Stensholt.  Single

transferable votes with tax cuts. SIAM Review, 46(3): 417–442, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39265.

Stiassnie:1987:DIS

[Sti87] M. Stiassnie. A definite integral from surface gravity waves. SIAM Review, 29 (4):632, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stoller:1960:FPD

[Sto60] David S. Stoller. Fitting position data to minimize velocity errors. SIAM Review, 2 (1):31–35, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stone:1961:BRC

[Sto61] W. M. Stone. Book review: Complex Variables and the Laplace Transform for Engineers (W. R. LePage). SIAM Review, 3(4): 342–343, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stone:1962:BRI

[Sto62a] W. M. Stone. Book review: Introduction to Laplace Transforms for Radio and Electronic Engineers (W. D. Day). SIAM Review, 4(3): 271, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stone:1962:BRS

[Sto62b] W. M. Stone. Book review: Shock and Vibration in Linear Systems (Paul A. Crafton). SIAM Review, 4 (4):409, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stone:1966:BRC

[Sto66] William M. Stone. Book review: Complex Variables-Special Functions. Vol. III-2 (V. I. Smirnov). SIAM Review, 8(1):122, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stockton:1973:CC

[Sto73]  Fred G. Stockton.  On

the characterization of a curve. SIAM Review, 15 (1):218, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stoutemyer:1976:ASI

[Sto76] D. R. Stoutemyer. Analytical solution of integral equations, using computer algebra. SIAM Review, 18(4): 829, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stone:1981:BRS

[Sto81] Lawrence D. Stone. Book review: Search and Screening: General Principles with Historical Applications (B. O. Koopman). SIAM Review, 23 (4):533–539, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stone:1991:C

[Sto91]  R. E. Stone.  Correc-

tion. SIAM Review, 33 (3):461, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stojanovic:1996:BRO

[Sto96] Srdjan Stojanovic. Book review: Optimal Control Theory for Infinite Dimensional Systems (Xunjing Li and Jiongmin Yong). SIAM Review, 38(2):352–353, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stockie:2011:MAD

[Sto11] John M. Stockie. The mathematics of atmospheric dispersion modeling. SIAM Review, 53(2):349–372, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sirev/
resource/1/siread/v53/
i2/p349_s1.

Struble:1962:BRTa

[Str62a] George Struble. Book review: Tables for Bivariate Osculatory Interpolation over a Cartesian Grid (Herbert E. Salzer and G. M. Kimbro). SIAM Review, 4(2): 164, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Struble:1962:BRTc

[Str62b] George Struble. Book review:

Tables of Coefficients for Obtaining the Second Derivative without Differences (H. E. Salzer and P. T. Roberson). SIAM Review, 4(2): 165, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Struble:1962:BRTb

[Str62c] George Struble. Book review:

Tables of Osculatory Integration Coefficients (H. Salzer, D. C. Schoultz and E. P. Thompson). SIAM Review, 4(2):164, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Struble:1964:NDO

[Str64] Raimond A. Struble. A note on damped oscillations in nonlinear systems. SIAM Review, 6(3):257–259, July 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strecok:1965:RTD

[Str65a] A. J. Strecok. A ratio of two definite multiple integrals. SIAM Review, 7(4):

559, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strecok:1965:SIE

[Str65b] A. J. Strecok. Solution to an integral equation. SIAM Review, 7(4):559, October 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strecok:1968:BRG

[Str68] Anthony J. Strecok. Book review: Gauss–Fourier Quadrature (W. Squire). SIAM Review, 10(4):456–457, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strecok:1970:BRS

[Str70] Anthony J. Strecok. Book review: A Set of Inequalities (James D. Riley). SIAM Review, 12(1):150–151, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strecok:1971:A

[Str71] Anthony J. Strecok. An antiderivative. SIAM Review, 13(1):115, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Streit:1975:OFP

[Str75] R. Streit. Optimum field patterns for unequally spaced linear antenna arrays. SIAM Review, 17(2):390, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strand:1976:SAB

[Str76a] O. N. Strand. Some aspects of behavior of regularized solutions as amount of smoothing is varied. SIAM Review, 18(4):829, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stroud:1976:BRM

[Str76b] A. H. Stroud. Book review: Methods of Numerical Integration (Philip J. Davis and Philip Rabinowitz). SIAM Review, 18 (3):528–529, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Struble:1977:AAA

[Str77] Raimond A. Struble. Analytical and algebraic aspects of the operational calculus. SIAM Review, 19(3): 403–436, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stroud:1982:BRN

[Str82] A. H. Stroud. Book review: Numerical Quadrature and Cubature (H. Engels). SIAM Review, 24(2): 233–234, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strang:1986:BRO

[Str86a] Gilbert Strang. Book review: Optimal Shape Design for Elliptic Systems (Olivier Pironneau). SIAM Review, 28(1): 127–128, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strauss:1986:GCM

[Str86b] David Strauss. On a general class of models for interaction. SIAM Review, 28 (4):513–527, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stroud:1986:BRM

[Str86c] A. H. Stroud. Book review: Methods of Numerical Integration-Second Edition (Philip J. Davis and Philip Rabinowitz). SIAM Review, 28(1):128, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strang:1988:FEE

[Str88] Gilbert Strang. A framework for equilibrium equations. SIAM Review, 30(2): 283–297, June 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strang:1989:WDE

[Str89a] Gilbert Strang. Wavelets and dilation equations. A brief introduction. SIAM Review, 31 (4):614–627, December 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strawderman:1989:BRS

[Str89b] William E. Strawderman.

Book review: Statistical Analysis with Missing Data (Roderick J. A. Little and Donald B. Rubin). SIAM Review, 31(2):348–349, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strang:1999:DCT

[Str99] Gilbert Strang. The discrete cosine transform. SIAM Review, 41(1):135–147, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/33674.

Stuelpnagel:1964:PTD

[Stu64] John Stuelpnagel. On the parametrization of the threedimensional rotation group. SIAM Review, 6(4):422– 430, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stuart:1986:TVF

[Stu86] J. T. Stuart. Taylor-vortex flow: a dynamical system.

SIAM Review, 28(3):315– 342, September 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stuart:1989:NID

[Stu89]  Andrew Stuart.  Nonlin-

ear instability in dissipative finite difference schemes. SIAM Review, 31(2):191– 220, June 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stuart:1990:BRH

[Stu90] J. T. Stuart. Book review: Hamiltonian Systems: Chaos and Quantization (Alfredo M. Ozorio de Almeida). Chaos and Integrability in Nonlinear Dynamics: an Introduction (Michael Tabor). An Introduction to Chaotic Dynamical Systems (Robert L. Devaney). SIAM Review, 32(1):188–190, March 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stuart:1991:BRM

[Stu91] Jeffrey L. Stuart. Book review: Matrix Theory and Applications (Charles R. Johnson, ed.). SIAM Review, 33 (3):511–512, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sturmfels:1994:GBC

[Stu94a]  B. Sturmfels.  Groebner

bases — a computational approach to commutative algebra (Thomas Becker and Volker Weispfenning). SIAM Review, 36(2):322–??, June 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sturmfels:1994:BRG

[Stu94b] Bernd Sturmfels. Book review: Gröbner Bases — a Computational Approach to Commutative Algebra (Thomas Becker and Volker Weispfenning). SIAM Review, 36(2):323, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Subhendu:1990:BRW

[Sub90] Datta K. Subhendu. Book review: Wave Propagation in Solids and Fluids (J. L. Davis). SIAM Review, 32(3): 500–501, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sugai:1976:GP

[Sug76] Iwao Sugai. Geometric probability. SIAM Review, 18 (1):118, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sugai:1977:GP

[Sug77] Iwao Sugai. Geometric probability. SIAM Review, 19(1): 152–154, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sugimoto:1998:HRW

[Sug98]  Takeshi Sugimoto.  How

to ride a wave: Mechanics of surfing. SIAM Review, 40(2):341–343, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31621.

Sulsky:1993:BRM

[Sul93a] Deborah Sulsky. Book review: Multigrid and Defect Correction for the Steady Navier–Stokes Equations Applications to Aerodynamics (Barry Koren). SIAM Review, 35(2):340–341, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sulsky:1993:MDC

[Sul93b]  Deborah Sulsky.  Multi-

grid and defect correction for the steady Navier–Stokes equations applications to aerodynamics (Barry Koren). SIAM Review, 35(2): 340–??, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Suli:1997:BRT

[Sül97] Endre Süli. Book review:

Time Dependent Problems and Difference Methods, by Bertil Gustafsson and Heinz-Otto Kreiss and Joseph Oliger. SIAM Review, 39(2): 362–366, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Summers:1987:DS

[Sum87] Danny Summers. Double summations. SIAM Review, 29(2):299, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sundstrom:1976:FNB

[Sun76] A. Sundstrom. Formulation of numerical boundary conditions for subsonic fluid-flow problems at high Reynolds numbers. SIAM Review, 18 (4):830, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Suppes:1977:BRS

[Sup77] Patrick Suppes. Book review:

Syntactic Methods in Pattern Recognition (K. S. Fu). SIAM Review, 19(4):746, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sussmann:1979:BRC

[Sus79] Héctor J. Sussmann. Book review: Catastrophe Theory, Selected Papers 1972-1977 (E. C. Zeeman). SIAM Review, 21(2):268–276, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sussmann:1988:BRN

[Sus88] H. J. Sussmann. Book review: Nonlinear Systems Theory (John L. Casti). SIAM Review, 30(1):156– 165, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sutin:1995:IGT

[Sut95] Brian Sutin. An inequality in gravitational theory. SIAM Review, 37(4):609, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Steeb:1989:CMI

[SV89] W. H. Steeb and C. M. Villet. A conjectured matrix inequality. SIAM Review, 31 (3):495, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sanchez-Velasco:1998:SDR

[SV98] Eduardo Sánchez-Velasco. A simple derivation of a result in electrostatics. SIAM Review, 40(4):915–917, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/32454.

Sleijpen:2000:JDI

[SV00] Gerard L. G. Sleijpen and Henk A. Van der Vorst. A Jacobi–Davidson iteration method for linear eigenvalue problems. SIAM Review, 42(2):267–293, June 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36308.

Svobodny:1996:BRA

[Svo96] Thomas P. Svobodny. Book review: Adjoint Equations and Analysis of Complex Systems (Guri I. Marchuk). SIAM Review, 38(2):353– 355, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Salinetti:1979:CSC

[SW79] Gabriella Salinetti and Roger J.-B. B. Wets. On the convergence of sequences of convex sets in finite dimensions. SIAM Review, 21(1): 18–33, January 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Salinetti:1980:ACC

[SW80] G. Salinetti and R. Wets. Addendum: “On the convergence of convex sets in finite dimensions” [SIAM rev. 21 (1979), no. 1, 18–33 and MR 80h:52007]. SIAM Review, 22 (1):86, January 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1981:BRS

[SW81] D. R. Smith and W. Whitt. Book review: Subadditivity of Teletraffic Special Functions (E. Arthurs and B. W. Stuck). SIAM Review, 23(4): 527–528, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sigman:1993:RRP

[SW93a] Karl Sigman and Ronald W. Wolff. A review of regenerative processes. SIAM Review, 35(2):269–288, June 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Singpurwalla:1993:WPS

[SW93b] Nozer D. Singpurwalla and Simon Wilson. The warranty problem: Its statistical and game theoretic aspects. SIAM Review, 35(1): 17–42, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Seo:2011:MRE

[SW11] Jin Keun Seo and Eung Je Woo. Magnetic Resonance Electrical Impedance Tomography (MREIT). SIAM Review, 53(1):40–68, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sirev/
resource/1/siread/v53/
i1/p40_s1.

Snieder:2022:WRH

[SW22] Roel Snieder and Michael B.

Wakin. When randomness helps in undersampling. SIAM Review, 64 (4):1062–1080, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1441006.

Swamy:1965:ZSP

[Swa65] M. N. S. Swamy. On the zeros of a set of polynomials. SIAM Review, 7(3): 409–410, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Swarztrauber:1977:MCR

[Swa77] Paul N. Swarztrauber. The methods of cyclic reduction, Fourier analysis and the FACR algorithm for the discrete solution of Poisson’s equation on a rectangle. SIAM Review, 19(3): 490–501, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Swanson:1978:DPS

[Swa78]  C. A. Swanson.  A di-

chotomy of PDE Sturmian theory. SIAM Review, 20(2): 285–300, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Swanson:1996:BRI

[Swa96] Richard Swanson. Book review: Introduction to the Modern Theory of Dynamical Systems (Anatole Katok and Boris Hasselblatt). SIAM Review, 38(1):180– 181, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Smith:1996:TCD

[SWC96] Hal L. Smith, Paul Waltman, and J. M. Cushing. The theory of the chemostat: Dynamics of microbial competition. SIAM Review, 38(1): 160–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shampine:1976:SNO

[SWD76] L. F. Shampine, H. A. Watts, and S. M. Davenport. Solving nonstiff ordinary differential equations — the state of the art. SIAM Review, 18 (3):376–411, July 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Swift:1960:CMI

[Swi60] George Swift. A comment on matrix inversion by partition. SIAM Review, 2(2): 132–133, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shorten:2007:SCS

[SWM^{+}07] Robert Shorten, Fabian Wirth, Oliver Mason, Kai Wulff, and Christopher King. Stability criteria for switched and hybrid systems. SIAM Review, 49(4):545–592, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shaidurov:1996:MMF

[SX96] V. V. Shaidurov and Jinchao Xu. Multigrid methods for finite elements. SIAM Review, 38(3):542– ??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shen:2019:NCE

[SXY19] Jie Shen, Jie Xu, and Jiang Yang. A new class of efficient and robust energy stable schemes for gradient flows. SIAM Review, 61(3): 474–506, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Strauss:1969:FTD

[SY69] Aaron Strauss and James A. Yorke. On the fundamental theory of differential equations. SIAM Review, 11(2): 236–246, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Stolarsky:1975:PI

[SY75] K. B. Stolarsky and L. J. Yang. A polynomial inequality. SIAM Review, 17 (4):686, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Sylvester:1996:BRG

[Syl96] John Sylvester. Book review: Geometric Scattering Theory (Richard B. Melrose). SIAM Review, 38(4): 702–703, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Symonds:1992:BRI

[Sym92] P. S. Symonds. Book review: Introduction to Optimization of Structures (N. V. Banichuk). SIAM Review, 34 (4):687–689, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Synowiec:1983:BRI

[Syn83] John Synowiec. Book review: Introduction to Partial Differential Equations and Hilbert Space Methods (Karl E. Gustafson). SIAM Review, 25(1):110–111, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Synowiec:1993:BRI

[Syn93a] John A. Synowiec. Book review: Introduction to Hilbert Spaces with Applications (Lokenath Debnath and Piotr Mikusiński). SIAM Review, 35(2):341–342, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Synowiec:1993:IHS

[Syn93b] John A. Synowiec. Introduction to Hilbert spaces with applications (Lokenath Debnath and Piotr Mikusinski). SIAM Review, 35(2): 341–??, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Syski:1963:BRE

[Sys63]  R. Syski.  Book review:

Elements of the Theory of Markov Processes and Their Applications (A. T. Bharucha-Reid). SIAM Review, 5(2):173–174, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Saigal:1996:LPM

[SZ96] Romesh Saigal and Yin Zhang. Linear programming: a modern integrated analysis. SIAM Review, 38(4): 711–??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Shirron:1999:SMI

[SZ99] J. J. Shirron and D. Zwillinger. Standard math interactive. SIAM Review, 41(3): 628–630, ???? 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Szatzschneider:1988:I

[Sza88] Wojciech Szatzschneider. An inequality. SIAM Review, 30 (3):502–503, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Szatzschneider:1991:I

[Sza91] Wojciech Szatzschneider. An inequality. SIAM Review, 33 (1):116–118, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Szymanski:1987:BRV

[Szy87] Boleslaw K. Szymanski. Book review: Verification and Validation of Real Time Software (W. J. Quirk, eds.). SIAM Review, 29(4):676– 678, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

TaAsan:1990:BRM

[Ta’90a] Shlomo Ta’Asan. Book review: Multigrid Solution of the Steady Euler Equations (S. P. Spekreijse). SIAM Review, 32(1):190– 191, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Terrell:1990:BRC

[TA90b] Robert E. Terrell and H. L.

Abbott. Book review: The Complexity of the Standard Form of an Integer (K. V. Lever). SIAM Review, 32(4): 674–681, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taber:1995:BRF

[Tab95] Rod Taber. Book review:

The Fuzzy Systems Handbook: a Practitioner’s Guide to Building, Using, and Maintaining Fuzzy Systems (Earl Cox). SIAM Review, 37 (2):281–282, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taber:1997:BRF

[Tab97] Rod Taber. Book review:

Fuzzy Logic for Business and Industry, by Earl D. Cox. SIAM Review, 39(1):

166–168, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tadmor:1987:SAF

[Tad87]  Eitan Tadmor.  Stability

analysis of finite difference, pseudospectral and Fourier– Galerkin approximations for time-dependent problems. SIAM Review, 29(4):525– 555, December 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tadjbakhsh:1989:BRO

[Tad89] Iradj G. Tadjbakhsh. Book review: Optimal Structural Design under Stability Constraints (A. Gajewski and M. Zyckowski). SIAM Review, 31(3):516, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tadjbakhsh:1993:BRS

[Tad93]  Iradj Tadjbakhsh.  Book

review: Strength of Structural Elements (Z. Brzoska, Z. Kaczkowski, J. Lipka, Z. Olesiak, and M. Zyczkowski). SIAM Review, 35 (2):342–343, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Takacs:1977:BRE

[Tak77] Lajos Takács. Book review: Einführung in die Bedienungstheorie (B. W. Gnedenko and I. N. Kowalenko). SIAM Review, 19(2):336, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Takacs:1979:FPB

[Tak79] Lajos Takács. Fluctuation problems for Bernoulli trials. SIAM Review, 21(2): 222–228, April 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Takacs:1983:BRL

[Tak83] L. F. Takacs. Book review:

Lattice Path Counting and Applications (Sri Gopal Mohanty). SIAM Review, 25(4): 592–593, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Talcott:1981:BRC

[Tal81a] Carolyn Talcott. Book review: A Computational Logic (Robert S. Boyer and J. Strother Moore). SIAM Review, 23(2):264–266, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Talcott:1981:BRL

[Tal81b] Carolyn Talcott. Book review: Lectures on the Logic of Computer Programming (Zohar Manna). SIAM Review, 23(3):397–398, ???? 1981.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tannenbaum:1993:OCR

[Tan93a] Allen Tannenbaum. H^{∞}- optimal control and related minimax design problems (Tamer Baser and Pierre Bernhard). SIAM Review, 35 (3):538–??, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tannenbaum:1993:BRO

[Tan93b] Allen Tannenbaum. Book review: H^{∞}-Optimal Control and Related Minimax Design Problems (Tamer Basar and Pierre Bernhard). SIAM Review, 35(3):538– 540, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tapia:1972:EP

[Tap72] R. Tapia. An extremum problem. SIAM Review, 14 (2):365, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tapia:1976:QNM

[Tap76] R. Tapia. Quasi-Newton methods for constrained optimization. SIAM Review, 18 (4):830, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tapia:1984:PMO

[Tap84a] R. Tapia. Practical methods of optimization, vol 2, constrained optimization — Fletcher, R. SIAM Review, 26(1):143–144, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tapia:1984:BRP

[Tap84b] Richard Tapia. Book review:

Practical Methods of Optimization, Volume 2: Constrained Optimization (R. Fletcher). SIAM Review, 26 (1):143–144, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tapia:1986:BRE

[Tap86] Richard Tapia. Book review: Engineering Optimization: Methods and Applications (G. V. Reklaitis, A. Ravindran and K. M. Ragsdell). SIAM Review, 28(2): 284–285, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tarjan:1978:CCA

[Tar78] Robert Endre Tarjan. Complexity of combinatorial algorithms. SIAM Review, 20(3): 457–491, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tarjan:1981:BRG

[Tar81] Robert E. Tarjan. Book review: Graphs and Networks (Bernard Carré). SIAM Review, 23(3):397, ???? 1981. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taussky:1967:TM

[Tau67] Olga Taussky. On two matrices. SIAM Review, 9 (3):590, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taub:1980:BRR

[Tau80]  A. H. Taub.  Book re-

view: Relativistic Theories of Materials (Aldo Bressan). SIAM Review, 22(3): 380–382, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taylor:1968:BRB

[Tay68] H. Taylor. Book review: Bayesian Decision Problems and Markov Chains (J. J. Martin). SIAM Review, 10 (3):392–393, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taylor:1973:MK

[Tay73] Geoffrey I. Taylor. Memories of von Kármán. SIAM Review, 15(2):447–452, ???? 1973. CODEN SIREAD.

ISSN 0036-1445 (print), 1095-7200 (electronic).

Taylor:1986:BRP

[Tay86] Howard M. Taylor. Book review: Probabilistic Models of Cumulative Damage (J. L. Bogdanoff and F. Kozin). SIAM Review, 28(3): 438–440, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taylor:1989:BRS

[Tay89] Michael E. Taylor. Book review: Scattering by Obstacles ( A. G. Ramm). SIAM Review, 31(1):164– 168, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taylor:1996:BRT

[Tay96] Michael Taylor. Book review:

The Technique of Pseudodifferential Operators (H. O. Cordes). SIAM Review, 38 (3):540–542, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taylor:1997:BRF

[Tay97] Michael E. Taylor. Book review: Fourier Integral Operators, by J. J. Duistermaat. SIAM Review, 39(1): 169–171, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tuttle:1976:ISE

[TB76] P. G. Tuttle and R. L. Brown. Implementation of selected eigenvalue algorithms on a vector computer. SIAM Review, 18(4):831, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Torres:2021:WHW

[TBBER21] Leo Torres, Ann S. Blevins, Danielle Bassett, and Tina Eliassi-Rad. The why, how, and when of representations for complex systems. SIAM Review, 63(3):435– 485, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Todd:1992:BRI

[TBDS92] John Todd, Bill Braden, Bernard Danloy, and Frank Schmidt. Book review: Inequality of the AGM and the Logarithmic Mean (B. C. Carlson and M. Vuorinen). SIAM Review, 34(4): 653–654, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Turner:1968:NI

[TC68] J. C. Turner and V. Conway. A network inequality. SIAM Review, 10(1): 107–108, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tian:2020:ACS

[TD20] Xiaochuan Tian and Qiang Du. Asymptotically compatible schemes for robust discretization of parametrized problems with applications to nonlocal models. SIAM Review, 62(1):199–227, ???? 2020. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tapia:2018:ISI

[TDS18] Richard A. Tapia, J. E.

Dennis, Jr., and Jan P. Schäfermeyer. Inverse, shifted inverse, and Rayleigh quotient iteration as Newton’s method. SIAM Review, 60(1):3–55, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Troutt:1996:PVS

[TE96] Marvin D. Troutt and Hussein H. Elsaid. The potential value of Saaty’s eigenvector scaling method for short-term forecasting of currency exchange rates. SIAM Review, 38(4):650– 654, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/25234.

TeBeest:1997:NAS

[TeB97a] K. G. TeBeest. Numerical and analytical solutions of Volterra’s population model. SIAM Review, 39(3):484– 493, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

TeBeest:1997:CNN

[TeB97b] Kevin G. TeBeest. Classroom note: Numerical and analytical solutions of Volterra’s population model. SIAM Review, 39(3):484– 493, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/29485.

Tekalp:1989:BRI

[Tek89] A. Murat Tekalp. Book review: Image Recovery: Theory and Application (Henry Stark, ed.). SIAM Review, 31 (1):168–170, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tenorio:2001:SRI

[Ten01] Luis Tenorio. Statistical regularization of inverse problems. SIAM Review, 43(2): 347–366, June 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/35823.

Terras:1982:NEH

[Ter82]  Audrey Terras.  Non–

Euclidean harmonic analysis. SIAM Review, 24(2): 159–193, April 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Terrell:1992:BRE

[Ter92] R. E. Terrell. Book review: Exponentials of Certain Hilbert Space Operators (M. L. Glasser). SIAM Review, 34(3):498–500, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Terman:1993:BRP

[Ter93a] David Terman. Book review:

Patterns and Waves (Peter Grindrod). SIAM Review, 35 (4):683–684, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Terman:1993:PWP

[Ter93b] David Terman. Patterns and waves (Peter Grindrod). SIAM Review, 35(4):683– ??, December 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Terrell:1996:ME

[Ter96] Robert E. Terrell. Matrix exponentials. SIAM Review, 38 (2):313–314, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tewarson:1966:PFI

[Tew66] R. P. Tewarson. On the product form of inverses of sparse matrices. SIAM Review, 8(3): 336–342, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tewarson:1967:PFI

[Tew67] R. P. Tewarson. The product form of inverses of sparse matrices and graph theory. SIAM Review, 9(1): 91–99, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tewarson:1969:SRG

[Tew69] R. P. Tewarson. On some representations of generalized inverses. SIAM Review, 11(2): 272–276, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tewarson:1970:CSM

[Tew70] R. P. Tewarson. Computations with sparse matrices. SIAM Review, 12(4): 527–543, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tewarson:1977:USD

[Tew77]  R. P. Tewarson.  Use

of smoothing and damping techniques in the solution of nonlinear equations. SIAM Review, 19(1): 35–45, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thrall:1965:SRC

[TF65] R. M. Thrall and J. E.

Falk. Some results concerning the kernel of a game. SIAM Review, 7(3): 359–375, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Temam:1996:NSE

[TF96] Roger Temam and R. Farwig.

Navier–Stokes equations and nonlinear functional analysis. SIAM Review, 38(4): 687–??, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tillquist:2023:GLL

[TFL23] Richard C. Tillquist, Rafael M.

Frongillo, and Manuel E. Lladser. Getting the lay of the land in discrete space: a survey of metric dimension and its applications. SIAM Review, 65(4):919– 962, November 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thornton:1996:CTI

[TH96] Barry S. Thornton and W. T. Hung. Catastrophe theory implications for rightsizing when planning interim solutions for progressing from a partial mainframe to clientserver distributed databases: 3D previewing of possible problems. SIAM Review, 38 (3):487–495, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29468.

Thacher:1979:BRA

[Tha79] Henry C. Thacher, Jr. Book review: Algorithms for the Computation of Mathematical Functions (Yudell L. Luke). SIAM Review, 21(3): 429–430, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thacher:1982:BRS

[Tha82] Henry C. Thacher, Jr. Book review: Software Manual for the Elementary Functions (W. J. Cody, Jr. and W. Waite). SIAM Review, 24 (1):91–93, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thatcher:1996:BRI

[Tha96]  R. W. Thatcher.  Book

review: Infinite Element Methods (Lung-an Ying). SIAM Review, 38(4):703– 705, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thejll:1991:LSP

[The91] Peter Thejll. Linear system with positive solutions. SIAM Review, 33(3):472– 473, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:1999:GIS

[The99a] The Editors. Graphical illustration of some examples related to the article “Iterated Random Functions” by Diaconis and Freedman. SIAM Review, 41(1): 77–82, March 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/34475.

Editors:1999:Sa

[The99b]  The Editors.  SIGEST.

SIAM Review, 41(3):553, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97031.

Editors:1999:Sb

[The99c]  The Editors.  SIGEST.

SIAM Review, 41(4):775, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97037.

Editors:2000:Sa

[The00a]  The Editors.  SIGEST.

SIAM Review, 42(1):91, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97043.

Editors:2000:Sb

[The00b] The Editors. SIGEST. SIAM Review, 42(2):265, June 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97047.

Editors:2000:Sc

[The00c]  The Editors.  SIGEST.

SIAM Review, 42(3):449, September 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97053.

Editors:2000:Sd

[The00d]  The Editors.  SIGEST.

SIAM Review, 42(4):693, December 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97058.

Editors:2001:Sa

[The01a] The Editors. SIGEST. SIAM Review, 43(2):313, June 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97068.

Editors:2001:Sb

[The01b]  The Editors.  SIGEST.

SIAM Review, 43(4):643, December 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97078.

Editors:2002:Sa

[The02a]  The Editors.  SIGEST.

SIAM Review, 44(1):93, March 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97083.

Editors:2002:Sb

[The02b] The Editors. SIGEST. SIAM Review, 44(2):235, June 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97088.

Editors:2002:Sc

[The02c]  The Editors.  SIGEST.

SIAM Review, 44(4):629, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97098.

Editors:2003:Sa

[The03a]  The Editors.  SIGEST.

SIAM Review, 45(1):73, March 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97103.

Editors:2003:Sb

[The03b] The Editors. SIGEST. SIAM Review, 45(2):289, June 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97108.

Editors:2003:Sc

[The03c]  The Editors.  SIGEST.

SIAM Review, 45(3):521, September 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97113.

Editors:2003:Sd

[The03d]  The Editors.  SIGEST.

SIAM Review, 45(4):725, December 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97118.

Editors:2004:Sa

[The04a] The Editors. SIGEST. SIAM Review, 46(2):283, June 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97127.

Editors:2004:Sb

[The04b]  The Editors.  SIGEST.

SIAM Review, 46(3):475, September 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97132.

Editors:2004:Sc

[The04c]  The Editors.  SIGEST.

SIAM Review, 46(4):691, December 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97137.

Editors:2005:Sa

[The05a]  The Editors.  SIGEST.

SIAM Review, 47(1):97, March 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97142.

Editors:2005:Sb

[The05b] The Editors. SIGEST. SIAM Review, 47(2):315, June 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97147.

Editors:2005:Sc

[The05c]  The Editors.  SIGEST.

SIAM Review, 47(3):517, September 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97152.

Editors:2005:Sd

[The05d]  The Editors.  SIGEST.

SIAM Review, 47(4):746, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97157.

Editors:2006:Sa

[The06a]  The Editors.  SIGEST.

SIAM Review, 48(1):97– 98, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2006:Sb

[The06b] The Editors. SIGEST. SIAM Review, 48(2):327, ???? 2006.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2006:Sc

[The06c] The Editors. SIGEST. SIAM Review, 48(3):547, ???? 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2007:Sa

[The07a] The Editors. SIGEST. SIAM Review, 49(3):467, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2007:Sb

[The07b] The Editors. SIGEST. SIAM Review, 49(4):649, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2008:Sa

[The08a] The Editors. SIGEST. SIAM Review, 50(1):89, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2008:Sb

[The08b] The Editors. SIGEST. SIAM Review, 50(2):323, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2008:Sc

[The08c] The Editors. SIGEST. SIAM Review, 50(3):521, ???? 2008.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2008:Sd

[The08d] The Editors. SIGEST. SIAM Review, 50(4):753, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2009:Sa

[The09a] The Editors. SIGEST. SIAM Review, 51(1):161, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2009:Sb

[The09b] The Editors. SIGEST. SIAM Review, 51(2):397, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2009:Sc

[The09c] The Editors. SIGEST. SIAM Review, 51(3):591, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2009:Sd

[The09d] The Editors. SIGEST. SIAM Review, 51(4):745, ???? 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2010:Sa

[The10a] The Editors. SIGEST. SIAM Review, 52(1):111, ???? 2010.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2010:Sb

[The10b] The Editors. SIGEST. SIAM Review, 52(2):315, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2010:Sc

[The10c] The Editors. SIGEST. SIAM Review, 52(3):503, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2010:Sd

[The10d] The Editors. SIGEST. SIAM Review, 52(4):715, ???? 2010. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2011:Sa

[The11a]  The Editors.  SIGEST.

SIAM Review, 53(1):127, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i1/p127_s1.

Editors:2011:Sb

[The11b]  The Editors.  SIGEST.

SIAM Review, 53(2):319, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.

siam.org/sirev/resource/
1/siread/v53/i2/p319_s1.

Editors:2011:Sc

[The11c]  The Editors.  SIGEST.

SIAM Review, 53(3):545, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i3/p545_s1.

Editors:2011:Se

[The11d]  The Editors.  SIGEST.

SIAM Review, 53(4):745, ???? 2011. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i4/p745_s1.

Editors:2012:Sd

[The12] The Editors. SIGEST. SIAM Review, 54(4):755, ???? 2012. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2013:Sa

[The13a] The Editors. SIGEST. SIAM Review, 55(1):129, ???? 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2013:Sb

[The13b] The Editors. SIGEST. SIAM Review, 55(3):491, ???? 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2013:Sc

[The13c] The Editors. SIGEST. SIAM Review, 55(4):707, ???? 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2014:Sa

[The14] The Editors. SIGEST. SIAM Review, 56(1):125, ???? 2014. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2016:Sa

[The16a] The Editors. SIGEST. SIAM Review, 58(1):117, ???? 2016. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Editors:2016:Sb

[The16b] The Editors. SIGEST. SIAM Review, 58(2):290, ???? 2016. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thickstun:1963:BRE

[Thi63a] W. R. Thickstun. Book review: Elements of Complex Variables (Louis L. Pennisi). SIAM Review, 5(4): 379, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thickstun:1963:BRI

[Thi63b] W. R. Thickstun. Book review: Introduction to Complex Analysis (Z. Nehari). SIAM Review, 5 (2):161, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thisted:1984:SSC

[Thi84a]  R. A. Thisted.  Statis-

tical software — a comparative review — Francis, I. SIAM Review, 26(2): 294–297, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thisted:1984:BRS

[Thi84b] Ronald A. Thisted. Book review: Statistical Software: a Comparative Review (Ivor Francis). SIAM Review, 26 (2):294–297, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thompson:1959:BRN

[Tho59] G. T. Thompson. Book review: The Numerical Solution of Two-Point Boundary Problems in Ordinary Differential Equations (L. Fox). SIAM Review, 1(2): 177–178, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thomasian:1960:BRI

[Tho60a] Aram J. Thomasian. Book review: Information Theory and Statistics (Solomon Kullback). SIAM Review, 2(2): 157–158, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thomsen:1960:RP

[Tho60b] D. L. Thomsen, Jr. Report of the President for 1959. SIAM Review, 2(1):1–2, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thomee:1969:STP

[Tho69] Vidar Thomée. Stability theory for partial difference operators. SIAM Review, 11(2): 152–195, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thomas:1974:BRR

[Tho74] L. B. Thomas. Book review: Rank Factorization of Nonnegative Matrices (A. Berman). SIAM Review, 16 (3):393–394, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thompson:1975:BRM

[Tho75] M. E. Thompson. Book review: Markov Processes and Learning Models (M. Frank Norman). SIAM Review, 17 (3):575–576, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thomas:1976:BRQ

[Tho76a] Marlin U. Thomas. Book review: Queueing Systems. Volume 1: Theory (Leonard Kleinrock). SIAM Review, 18 (3):512–514, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thompson:1976:OSC

[Tho76b] S. Thompson. Optimally stable corrector methods. SIAM Review, 18(4):830, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thom:1977:SSC

[Tho77a]  René Thom.  Structural

stability, catastrophe theory, and applied mathematics. SIAM Review, 19(2): 189–201, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thomas:1977:MDC

[Tho77b] L. E. Thomas. A model for drug concentration. SIAM Review, 19(4):732–735, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thomas:1987:BRN

[Tho87] J. W. Thomas. Book review: Numerical Methods for Scientific and Engineering Computation (M. K. Jain, S. R. K. Iyengar and R. K. Jain, eds). SIAM Review, 29(1): 166–167, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thrall:1979:BRD

[Thr79] Robert M. Thrall. Book review: Discrete Mathematical Models (Fred S. Roberts). SIAM Review, 21(3):424– 425, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thron:1994:BRC

[Thr94a] W. J. Thron. Book review:

Continued Fractions with Applications (Lisa Lorentzen and Haakon Waadeland). SIAM Review, 36(1):148– 149, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thron:1994:CFA

[Thr94b] W. J. Thron. Continued fractions with applications (Lisa Lorentzen and Haakon Waadeland). SIAM Review, 36(1):148–??, March 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thunberg:2001:PVC

[Thu01]  Hans Thunberg.  Peri-

odicity versus chaos in one-dimensional dynamics. SIAM Review, 43(1):3–30, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/37664.

Thyagaraja:1973:EA

[Thy73] A. Thyagaraja. Ensemble averages. SIAM Review, 15 (4):787, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thyagaraja:1974:BRE

[Thy74] A. Thyagaraja. Book review: Ensemble Averages (A. Thyagaraja). SIAM Review, 16 (4):548, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tijs:1981:NEN

[Tij81] Steve H. Tijs. Nash equilibria for noncooperative nperson games in normal form. SIAM Review, 23(2): 225–237, April 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ting:1976:EBD

[Tin76] A. S. C. Ting. Error bounds on discretized quasi-Newton methods. SIAM Review, 18 (4):830, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tismenetsky:1983:KPM

[Tis83a]  M. Tismenetsky.  Kro-

necker products and matrix calculus — Graham, A. SIAM Review, 25(3): 420–421, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tismenetsky:1983:BRK

[Tis83b] Miron Tismenetsky. Book review: Kronecker Products and Matrix Calculus (Alexander Graham). SIAM Review, 25(3):420–421, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tien:1982:MSA

[TK82] James M. Tien and Angelica Kamiyama. On manpower scheduling algorithms. SIAM Review, 24(3):275– 287, July 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tanabe:1997:FAM

[TK97] Hiroki Tanabe and R. Bruce Kellogg. Functional analytic methods for partial differential equations. SIAM Review, 39(4):798– ??, December 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ting:1995:VVF

[TKG95] L. Ting, R. Klein, and Max D. Gunzburger. Viscous vortical flows. SIAM Review, 37(1):124–??, March 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Traud:2011:CCS

[TKMP11] Amanda L. Traud, Eric D. Kelsic, Peter J. Mucha, and Mason A. Porter. Comparing community structure to characteristics in online collegiate social networks. SIAM Review, 53(3):526– 543, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i3/p526_s1.

Tuckey:1995:NMC

[TL95] Curtis Tuckey and Peter A. Loeb. Nonstandard methods in the calculus of variations. SIAM Review, 37(3): 474–??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Teo:2002:MRF

[TL02] Chung-Piaw P. Teo and Siew Meng Leong. Managing risk in a four-digit number game. SIAM Review, 44 (4):601–615, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36698.

Terrell:1990:CSF

[TLM90] R. E. Terrell, K. V. Lever, and G. Mccaughan. The complexity of the standard form of an integer. SIAM Review, 32(4):674–681, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tartar:1995:MFC

[TLN95] M. E. Tartar, M. D. Lock, and Guy Nason. Modelfree curve estimation. SIAM Review, 37(2):270–??, June 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tisseur:2001:QEP

[TM01] Françoise Tisseur and Karl Meerbergen. The quadratic eigenvalue problem. SIAM Review, 43(2):235–286, June 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/38198.

Todd:1974:JNN

[Tod74] John Todd. John von Neumann and the national accounting machine. SIAM Review, 16(4):526–530, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Todd:1975:NAN

[Tod75] John Todd. Numerical analysis at the National Bureau of Standards. SIAM Review, 17 (2):361–370, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Todd:1984:QNU

[Tod84] Michael J. Todd. Quasi-Newton updates in abstract vector spaces. SIAM Review, 26(3):367–378, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tokarsky:1994:IPS

[Tok94] G. Tokarsky. An impossible pool shot? SIAM Review, 36 (1):107–108, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tokarksy:1995:IPS

[Tok95] G. Tokarksy. An impossible pool shot? SIAM Review, 37 (1):107–109, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tong:1988:SMI

[Ton88] Y. L. Tong. Some majorization inequalities in multivariate statistical analysis. SIAM Review, 30(4):602– 622, December 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tongue:1989:BRC

[Ton89] B. H. Tongue. Book review: Cell-to-Cell Mapping: a Method of Global Analysis for Nonlinear Systems (C. S. Hsu). SIAM Review, 31(1): 170–171, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tong:1995:BRS

[Ton95] Y. L. Tong. Book review:

Stochastic Orders and Their Applications (Moshe Shaked and J. George Shanthikumar). SIAM Review, 37(3): 477–479, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tornheim:1962:DLP

[Tor62]  Leonard Tornheim.  De-

termination of large parallelepipeds. SIAM Review, 4 (3):223–226, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Totik:1991:BRA

[Tot91] Vilmos Totik. Book review: The Averaged Moduli of Smoothness: With Applications in Numerical Methods and Approximation (Blagovest Sendov and Vasil A. Popov; trans. ed., G. M. Phillips). SIAM Review, 33 (3):499–501, September 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thomas:1978:EDS

[TP78] J. P. Thomas and S. L. Passman. Exact dalembert solutions of Navier–Stokes equations. SIAM Review, 20 (3):635, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Turner:2011:UCS

[TPCC^{+}11] Peter Turner, Linda Petzold, Co-Chairs, Angela Shiflet, Ignatios Vakalis, Kirk Jordan, and Samuel St. John. Undergraduate computational science and engineering education. SIAM Review, 53(3): 561–574, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sirev/resource/
                                        1/siread/v53/i3/p561_s1.

The authors are members of the SIAM Working Group on CSE Undergraduate Education.

tenRaa:1983:SPI

[tR83]  T. ten Raa.  A system

of probabilistic inequalities. SIAM Review, 25(4):569– 570, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Traub:1966:APU

[Tra66] J. F. Traub. Associated polynomials and uniform methods for the solution of linear problems. SIAM Review, 8 (3):277–301, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trench:1965:BRA

[Tre65] William I. Trench. Book review: An Asymptotic Expansions (H. O. Pollak and L. Shepp). SIAM Review, 7(3): 417–420, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Treuenfels:1970:UBA

[Tre70]  Peter Treuenfels.  Upper

bound for the abscissa of the χ^{2}-distribution. SIAM Review, 12(1):145–146, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trebels:1973:CI

[Tre73] W. Trebels. Conditions for an inequality. SIAM Review, 15(4):789, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Treadway:1978:CNM

[Tre78] A. H. Treadway. Comparison of numerical methods for multicomponent radionuclide ion migration in adsorbing media. SIAM Review, 20 (3):635, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:1982:GVF

[Tre82] Lloyd N. Trefethen. Group velocity in finite difference schemes. SIAM Review, 24 (2):113–136, April 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:1984:FAN

[Tre84a] L. N. Trefethen. Fourier analysis of numerical approximations of hyperbolic equations — Vichnevetsky, R., Bowles, J. B. SIAM Review, 26(3): 439–441, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:1984:BRF

[Tre84b] Lloyd N. Trefethen. Book review: Fourier Analysis of Numerical Approximations of Hyperbolic Equations (R. Vichnevetsky and J. B. Bowles). SIAM Review, 26 (3):439–441, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:1997:PLO

[Tre97] Lloyd N. Trefethen. Pseudospectra of linear operators. SIAM Review, 39(3): 383–406, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29528.

Trefethen:1999:SRIa

[Tre99a] N. Trefethen. Survey and review — introduction. SIAM Review, 41(2):197, ???? 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:1999:SRIb

[Tre99b] N. Trefethen. Survey and review — introduction. SIAM Review, 41(4):635, ???? 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:1999:SRa

[Tre99c]  Nick Trefethen.  Survey

and review. SIAM Review, 41(1):1–2, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97019.

Trefethen:1999:SRb

[Tre99d]  Nick Trefethen.  Survey

and review. SIAM Review, 41(2):197, June 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97024.

Trefethen:1999:SRc

[Tre99e]  Nick Trefethen.  Survey

and review. SIAM Review, 41(3):415, September 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97029.

Trefethen:1999:SRd

[Tre99f]  Nick Trefethen.  Survey

and review. SIAM Review, 41(4):635, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97035.

Trefethen:2000:SRa

[Tre00a]  Nick Trefethen.  Survey

and review. SIAM Review, 42(1):1, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97041.

Trefethen:2000:SRb

[Tre00b]  Nick Trefethen.  Survey

and review. SIAM Review, 42(2):159, June 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97050.

Trefethen:2000:SRIa

[Tre00c]  Nick Trefethen.  Survey

and review — introduction. SIAM Review, 42 (3):367, September 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97051.

Trefethen:2000:SRIb

[Tre00d]  Nick Trefethen.  Survey

and review — introduction. SIAM Review, 42(4): 553, December 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/97056.

Trefethen:2001:SRa

[Tre01a]  Nick Trefethen.  Survey

and review. SIAM Review, 43(1):1, March 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97061.

Trefethen:2001:SRb

[Tre01b]  Nick Trefethen.  Survey

and review. SIAM Review, 43(2):233, June 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97066.

Trefethen:2001:SRc

[Tre01c]  Nick Trefethen.  Survey

and review. SIAM Review, 43(3):411, September 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97071.

Trefethen:2001:SRd

[Tre01d]  Nick Trefethen.  Survey

and review. SIAM Review, 43(4):583, December 2001. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97076.

Trefethen:2002:SRI

[Tre02a] N. Trefethen. Survey and review — introduction. SIAM Review, 44(4):523, ???? 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2002:SRa

[Tre02b]  Nick Trefethen.  Survey

and review. SIAM Review, 44(1):1, March 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97081.

Trefethen:2002:SRb

[Tre02c]  Nick Trefethen.  Survey

and review. SIAM Review, 44(2):167, June 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/97086.

Trefethen:2002:SRc

[Tre02d]  Nick Trefethen.  Survey

and review. SIAM Review, 44(3):309, September 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97091.

Trefethen:2002:SRd

[Tre02e]  Nick Trefethen.  Survey

and review. SIAM Review, 44(4):523, December 2002. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/97096.

Trefethen:2008:GQB

[Tre08]  Lloyd N. Trefethen.  Is

Gauss quadrature better than Clenshaw–Curtis? SIAM Review, 50(1):67–87, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://link.aip.org/link/
?SIR/50/67/1.

Trefethen:2013:BRV

[Tre13] Lloyd N. Trefethen. Book review: Visual complex functions: an introduction with phase portraits, by Elias Wegert. Birkhäuser/ Springer Basel AG, Basel, 2012. $59.95. xiv + 360 pp., softcover. ISBN 978-3-0348-0179-9. SIAM Review, 55 (4):791–797, 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2015:PCA

[Tre15] Lloyd N. Trefethen. The Princeton companion to applied mathematics [book review of MR3380576]. SIAM Review, 57(3):469–473, 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2017:CAI

[Tre17] Lloyd N. Trefethen. Cubature, approximation, and isotropy in the hypercube. SIAM Review, 59(3):469– 491, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2020:EPE

[Tre20] Lloyd N. Trefethen. Eight perspectives on the exponentially ill-conditioned equation εy′′ − xy′ + y = 0. SIAM Review, 62(2):439– 462, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2022:EQF

[Tre22] Lloyd N. Trefethen. Exactness of quadrature formulas. SIAM Review, 64 (1):132–150, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1389522.

Tritter:1962:RCR

[Tri62a] Alan L. Tritter. Resistance of a cut-out right circular cylinder. SIAM Review, 4 (2):150, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tritter:1962:SST

[Tri62b] Alan L. Tritter. A steadystate temperature. SIAM Review, 4(2):148, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trigiante:1978:SCR

[Tri78] D. Trigiante. Some comparison result useful in numerical analysis. SIAM Review, 20(3):635, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tripp:1986:LS

[Tri86] C. E. Tripp. Late solution. SIAM Review, 28(3): 404, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Triebel:1995:BRC

[Tri95] Hans Triebel. Book review:

Completeness of Root Functions of Regular Differential Operators (S. Yakubov). SIAM Review, 37(1):138,

????  1995.  CODEN

SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trombetta:1971:GRP

[Tro71] Michael L. Trombetta. A Gambler’s ruin problem. SIAM Review, 13(4):569– 570, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trummer:1994:BRC

[Tru94] Manfred R. Trummer. Book review: The Cauchy Transform, Potential Theory and Conformal Mapping (Steven R. Bell). SIAM Review, 36 (3):523–524, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trench:1965:BRD

[TS65] William F. Trench and Alan Schumitzky. Book review: A Definite Integral (J. L. Brown, Jr.). SIAM Review, 7 (2):278–280, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trench:1966:IHT

[TS66] William F. Trench and Perry A. Scheinok. On the inversion of a Hilbert type matrix. SIAM Review, 8 (1):57–61, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tze-San:1986:CRM

[TS86]  Lee Tze-San.  A com-

plete Russel’s method for the transportation problem. SIAM Review, 28(4):547– 549, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tuck:1990:NAS

[TS90] E. O. Tuck and L. W.

Schwartz. A numerical and asymptotic study of some third-order ordinary differential equations relevant to draining and coating flows. SIAM Review, 32(3):453– 469, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taylor:1996:MP

[TS96] Alan D. Taylor and Donald G. Saari. Mathematics and politics. SIAM Review, 38(4):698–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Taylor:1976:BRA

[TT76] I. D. S. Taylor and J. G. C.

Templeton. Book review: Accident Probability (L. J. Dickson). SIAM Review, 18 (3):493, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tapia:1994:EKK

[TT94] R. A. Tapia and M. W. Trosset. An extension of the Karush–Kuhn–Tucker necessity conditions to infinite programming. SIAM Review, 36 (1):1–17, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/22073.htm.

Ta-tsien:1996:GCS

[TtK96] Li Ta-tsien and Mikhael Kovalyov. Global classical solutions for quasilinear hyperbolic systems. SIAM Review, 38(2):347–??, June 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Torres-Torriti:2005:SPL

[TTM05] Miguel Torres-Torriti and Hannah Michalska. A software package for Lie algebraic computations. SIAM Review, 47(4):722–745, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/41042.

Turner:1988:CBA

[TTY88] D. W. Turner, F. E. Tidmore, and D. M. Young. A calculus based approach to the blood testing problem. SIAM Review, 30(1): 119–122, March 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tu:1976:RTF

[Tu76] K. W. Tu. Real-time flow network computational problem in power generating systems simulation. SIAM Review, 18 (4):830–831, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tu:1992:BRI

[Tu92] Guizhang Tu. Book review: Integrable Systems of Classical Mechanics and Lie Algebras. Vol. I (A. M. Perelomov; A. G. Reyman, trans.). SIAM Review, 34(1): 137–138, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tucker:1973:PGA

[Tuc73] Alan Tucker. Perfect graphs and an application to optimizing municipal services. SIAM Review, 15(3):585– 590, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tudor:1990:DMH

[Tud90] David Tudor. A deterministic model for herpes infections in human and animal populations. SIAM Review, 32(1): 136–139, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tudor:1992:MEP

[Tud92]  David Tudor.  Modeling

the effect of public health campaigns on the spread of aids. SIAM Review, 34(2): 300–303, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tukey:1967:DS

[Tuk67] John W. Tukey. Discussion of session 1. SIAM Review, 9 (??):308–310, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tukey:1987:CP

[Tuk87] John W. Tukey. Configural polysampling. SIAM Review, 29(1):1–20, March 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tulcea:1968:BRE

[Tul68] A. Ionescu Tulcea. Book review: Ergodic Theory and Information (Patrick Billingsley). SIAM Review, 10(4): 459, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tulovsky:1990:BRM

[Tul90] Vladimir Tulovsky. Book review: Mathematical Analysis and Numerical Methods for Science and Technology, Volume 2: Functional and Variational Methods (Robert Daut ray and Jacques-Louis Lions; Ian N. Sneddon, trans.). SIAM Review, 32(1): 191–192, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tuminaro:2012:RSa

[Tum12a] Ray Tuminaro. Research spotlights. SIAM Review, 54 (2):289, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tuminaro:2012:RSb

[Tum12b] Ray Tuminaro. Research spotlights. SIAM Review, 54 (3):511, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tuminaro:2012:RSc

[Tum12c] Ray Tuminaro. Research spotlights. SIAM Review, 54 (4):697, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tuminaro:2013:RSa

[Tum13a] Ray Tuminaro. Research spotlights. SIAM Review, 55 (1):99, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tuminaro:2013:RSb

[Tum13b] Ray Tuminaro. Research spotlights. SIAM Review, 55 (3):441, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tuminaro:2013:RSc

[Tum13c] Ray Tuminaro. Research spotlights. SIAM Review, 55 (4):669, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tuminaro:2014:RS

[Tum14] Ray Tuminaro. Research spotlights. SIAM Review, 56 (1):71, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Turanski:1960:BRT

[Tur60] W. J. Turanski. Book review: The Theory of Groups (Marshall Hall, Jr.). SIAM Review, 2(2):161–163, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Turski:1968:IEC

[Tur68] A. J. Turski. An integral equation of convolution type. SIAM Review, 10 (1):108, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Turkel:1978:HOM

[Tur78] E. Turkel. High-order methods for large-scale problems. SIAM Review, 20(3):635– 636, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Turski:2004:GFA

[Tur04]  Jacek Turski.  Geomet-

ric Fourier analysis of the conformal camera for active vision. SIAM Review, 46(2):230–255, June 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40096.

Tutschke:1994:BRT

[Tut94] Wolfgang Tutschke. Book review: Transformations, Transmutations, and Kernel Functions, Vol. 1 (H. Begehr and R. P. Gilbert). SIAM Review, 36(3):524– 525, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Thompson:1969:FRB

[TW69] Gerald L. Thompson and Roman L. Weil, Jr. Further relations between game theory and eigensystems.

SIAM Review, 11(4):597– 602, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Trefethen:2014:ECT

[TW14] Lloyd N. Trefethen and J. A. C. Weideman. The exponentially convergent trapezoidal rule. SIAM Review, 56 (3):385–458, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tierney:1965:SPO

[TWW65] Martin S. Tierney, Paul Waltman, and G. Milton Wing. Some problems in the optimal design of shields and reflectors in particle and wave physics. SIAM Review, 7(1): 100–113, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Tze:1991:OSS

[Tze91]  Leung Lai Tze.  Opti-

mization of stochastic systems: Topics in discrete-time dynamics (Masanao Aoki). SIAM Review, 33(4):671– ??, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ucar:2007:RHM

[UA07] Bora Uçar and Cevdet Aykanat. Revisiting hypergraph models for sparse matrix partitioning. SIAM Review, 49(4):595–603, ???? 2007. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Unser:2000:FSW

[UB00] Michael Unser and Thierry Blu. Fractional splines and wavelets. SIAM Review, 42(1):43–67, March 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34943.

Udwadia:1989:SRM

[Udw89] Firdaus E. Udwadia. Some results on maximum entropy distributions for parameters known to Lie in finite intervals. SIAM Review, 31(1): 103–109, March 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Unser:2017:SUS

[UFW17] Michael Unser, Julien Fageot, and John Paul Ward. Splines are universal solutions of linear inverse problems with generalized TV regularization. SIAM Review, 59(4): 769–793, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Uhlenbeck:1973:BRMa

[Uhl73a] G. E. Uhlenbeck. Book review: Mathematical Theory of Transport Processes in Gases (J. H. Ferziger and H. G. Kaper). SIAM Review, 15 (4):804–805, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Uhlenbeck:1973:BRMb

[Uhl73b] G. E. Uhlenbeck. Book review: Mathematical Theory of Transport Processes in Gases (J. H. Ferziger and H. G. Kaper). SIAM Review, 15 (4):804–805, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ulam:1964:CAI

[Ula64]  S. Ulam.  Combinato-

rial analysis in infinite sets and some physical theories. SIAM Review, 6(4): 343–355, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ulam:1967:EAM

[Ula67] Stanisław M. Ulam. An education in applied math. SIAM Review, 9(??):343– 344, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Proceedings of May 24–27, 1966 SIAM Conference (Aspen, Colorado).

Ulkucu:1976:RPM

[Ulk76] A. Ulkucu. Random polyhedra in mathematical programming. SIAM Review, 18 (4):831, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ullman:1988:BRM

[Ull88] Jeffrey D. Ullman. Book review: Mathematics of Information Processing (Michael Anshel and William Gewirtz, eds.). SIAM Review, 30(1): 165–168, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Underwood:1976:GPM

[Und76] R. G. Underwood. Gradient projection method for solving continuous games. SIAM Review, 18(4):831, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Underwood:1978:NCN

[Und78] R. G. Underwood. Null controllability of neutral differential equations. SIAM Review, 20(3):636, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ungar:1977:SLP

[Ung77] A. Ungar. Solutions to linear partial differential equations involving arbitrary functions. SIAM Review, 19(1):147– 148, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Unguru:1985:BRA

[Ung85]  Sabetai Unguru.  Book

review: Astronomy and History-Selected Essays (O. Neugebauer). SIAM Review, 27(2):305–309, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Urbas:1993:BRR

[Urb93a] John Urbas. Book review: Regularity Theory for Quasilinear Elliptic Systems and Monge–Ampere Equations in Two Dimensions (Friedmar Schulz). SIAM Review, 35 (1):169–172, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Urbas:1993:RTQ

[Urb93b]  John Urbas.  Regularity

theory for quasilinear elliptic systems and Monge– Ampere equations in two dimensions (Friedmar Schulz). SIAM Review, 35(1):169– ??, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Uribe:1992:BRP

[Uri92] Alejandro Uribe. Book review: Pseudo-Differential Operators (S. R. Simanca). SIAM Review, 34(1):128– 129, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Urquhart:1968:CGI

[Urq68] N. S. Urquhart. Computation of generalized inverse matrices which satisfy specified conditions. SIAM Review, 10 (2):216–218, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Urquhart:1969:NLU

[Urq69] N. S. Urquhart. The nature of the lack of uniqueness of generalized inverse matrices. SIAM Review, 11(2): 268–271, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Usmani:1972:CMM

[Usm72]  R. A. Usmani.  Conjec-

tured monotonicity of a matrix. SIAM Review, 14(2): 365–366, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Utz:1979:MMP

[Utz79] W. R. Utz. A min-max problem. SIAM Review, 21 (4):559, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Uhler:2013:PEO

[UW13] Caroline Uhler and Stephen J. Wright. Packing ellipsoids with overlap. SIAM Review, 55(4):671–706, ???? 2013. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanAlbada:1961:EP

[vA61] P. L. van Albada. An expectation problem. SIAM Review, 3(1):71–72, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vainikko:1996:BRA

[Vai96] Gennadi Vainikko. Book review: Approximation Procedures in Nonlinear Oscillation Theory (Nikolai A. Bobylev, Yuri M. Burman, and Sergey K. Korovin). SIAM Review, 38(1):181– 182, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vakakis:1993:BRT

[Vak93] Alexander F. Vakakis. Book review: Theory of Vibration: I. An Introduction and II: Discrete and Continuous Systems (A. A. Shabana). SIAM Review, 35 (3):540–541, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Valentine:1967:MPF

[Val67] Charles W. Valentine. A method for partial fraction decomposition. SIAM Review, 9(2):232–233, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanYzeren:1963:BRE

[Van63] J. Van Yzeren. Book review: The Expected Value of a Product (L. A. Shepp). SIAM Review, 5(4):373– 374, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanLint:1966:BRA

[Van66] J. H. Van Lint. Book review: An Asymptotic Expansion (H. O. Pollak and L. Shepp). SIAM Review, 8(3): 383–384, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanNess:1968:BRP

[Van68] John W. Van Ness. Book review: Probability (John Lamperti). SIAM Review, 10(2): 237–238, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vantuyl:1975:CLL

[Van75] A. H. Vantuyl. Calculation of limiting lines of solutions of analytic initial value problems by use of power series and nonlinear sequence transformations. SIAM Review, 17 (2):390–391, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanNess:1976:BRM

[Van76] John W. Van Ness. Book review: Multiple Time Series (E. J. Hannan). SIAM Review, 18(1):132–133, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vandevender:1978:SII

[Van78] W. H. Vandevender. Stability of an invariant imbedding algorithm for numerical solution of two-point boundaryvalue problems. SIAM Review, 20(3):636, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanAssche:1986:BRA

[Van86] Walter Van Assche. Book review: An Asymptotic Problem (Pierre Barrucand). SIAM Review, 28(2):234– 238, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanLeer:1988:BRN

[Van88] Bram Van Leer. Book review: Numerical Methods in Fluid Dynamics: Initial and Initial Boundary-Value Problems (Gary A. Sod). SIAM Review, 30(1): 168–169, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vance:1993:BRK

[Van93] Pamela H. Vance. Book review: Knapsack Problems: Algorithms and Computer Implementations (S. Martello and P. Toth). SIAM Review, 35(4):684–685, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanDyke:1994:NCR

[Van94a] Milton Van Dyke. Nineteenth-Century roots of the boundarylayer idea. SIAM Review, 36 (3):415–424, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/25789.
htm.

Vanderbauwhede:1994:TBT

[Van94b] A. Vanderbauwhede. Topics in bifurcation theory and applications (Gerard Iooss and Moritz Adelmeyer). SIAM Review, 36(2):323– ??, June 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vanderbauwhede:1994:BRT

[Van94c] André Vanderbauwhede. Book review: Topics in Bifurcation Theory and Applications (Gerard Ioss and Moritz Adelmeyer). SIAM Review, 36(2):323–324, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vandyke:1994:CRB

[Van94d] M. Vandyke. 19th-Century roots of the boundary-layer idea. SIAM Review, 36(3): 415–424, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanAssche:1996:BRP

[Van96] Walter Van Assche. Book review: Polynomials and Polynomial Inequalities (P. Borwein and T. Erdélyi). SIAM Review, 38(4):705– 706, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanPutten:1998:RSU

[Van98] Maurice H. P. M. Van Putten. A Riemann sum upper bound in the Riemann– Lebesgue theorem. SIAM Review, 40(2):333–334, June 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

                                        http://epubs.siam.org/
                                        sam-bin/dbq/article/31189.

Vanderbei:2012:LW

[Van12] Robert J. Vanderbei. Local warming. SIAM Review, 54 (3):597–606, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Varga:1966:BRT

[Var66]  R. S. Varga.  Book re-

view: The Theory of Matrices in Numerical Analysis (Alston S. Householder). SIAM Review, 8(1):127– 128, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Varga:1967:BRI

[Var67] Richard S. Varga. Book review: Iterative Solution of Elliptic Systems and Applications to the Neutron Diffusion Equations of Reactor Physics (Eugene L. Wachspress). SIAM Review, 9(4): 756–758, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Varah:1976:MNS

[Var76] J. M. Varah. Methods for numerical solution of linear discrete ill-posed problems. SIAM Review, 18 (4):832, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Varah:1979:PES

[Var79]  J. M. Varah.  A prac-

tical examination of some numerical methods for linear discrete ill-posed problems. SIAM Review, 21(1): 100–111, January 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Varma:1980:CFI

[Var80] A. Varma. A closed form integration. SIAM Review, 22(1):97, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Varadhan:1985:BRI

[Var85] Srinivasa S. R. Varadhan. Book review: An Introduction to the Theory of Large Deviations (D. W. Stroock). SIAM Review, 27(4):608– 610, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Varadarajan:1987:BRG

[Var87] V. S. Varadarajan. Book review: Group Representations and Special Functions (by Antoni Wawrzynczyk). SIAM Review, 29(2):336– 337, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vasilieva:1994:DSP

[Vas94] A. B. Vasilieva. On the development of singular perturbation theory at Moscow state University and elsewhere. SIAM Review, 36 (3):440–452, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/25729.
htm.

Vassilevski:1997:TWS

[Vas97] Panayot S. Vassilevski. On two ways of stabilizing the hierarchical basis multilevel methods. SIAM Review, 39(1):18–53, March 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28221;   http:
//www.siam.org/journals/
sirev/sirev391.htm.

Vandervorst:1975:UFE

[VB75] M. J. Vandervorst and A. E.

Berger. Use of finite-element method in a marker and Cell program for solving timedependent fluid-flow problems with free surfaces. SIAM Review, 17(2):390, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vandenberghe:1996:SP

[VB96] Lieven Vandenberghe and Stephen Boyd. Semidefinite programming. SIAM Review, 38(1):49–95, March 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/27182.
htm.

Vastenhouw:2005:TDD

[VB05] Brendan Vastenhouw and Rob H. Bisseling. A twodimensional data distribution method for parallel sparse matrix-vector multiplication. SIAM Review, 47(1):67–95, March 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40901.

Vandenberghe:2007:GCB

[VBC07] Lieven Vandenberghe, Stephen Boyd, and Katherine Comanor. Generalized Chebyshev bounds via semidefinite programming. SIAM Review, 49(1):52–64, ???? 2007. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Verlinde:1991:BRT

[VBdD91] Johannes Verlinde, J. Boersma, and P. J. de Doelder. Book review: Two Integrals Arising from a Cloud Model (Johannes Verlinde). SIAM Review, 33(3):474–477, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Veldt:2022:HCG

[VBK22] Nate Veldt, Austin R. Benson, and Jon Kleinberg. Hypergraph cuts with general splitting functions. SIAM Review, 64(3):650–685, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

https://epubs.siam.org/
doi/10.1137/20M1321048.

Villacampa:1997:CNS

[VBU97a] Y. Villacampa, A. Balaguer, and J. L. Usó. Classroom note: a study of a semi-infinite integral. SIAM Review, 39(3):494– 495, September 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28678.

Villacampa:1997:SSI

[VBU97b] Y. Villacampa, A. Balaguer, and J. L. Uso. A study of a semi-infinite integral. SIAM Review, 39(3): 494–495, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanderBoor:2022:SLB

[VBVM22] Mark Van der Boor, Sem C.

Borst, Johan S. H. Van Leeuwaarden, and Debankur Mukherjee. Scalable load balancing in networked systems: a survey of recent advances. SIAM Review, 64(3):554–622, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1323746.

Vellekoop:2006:NFA

[VC06] M. H. Vellekoop and J. M. C. Clark. A nonlinear filtering approach to changepoint detection problems: Direct and differential-geometric methods. SIAM Review, 48(2): 329–356, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vanassche:1986:AP

[VCG^{+}86] W. Vanassche, J. A. Crow, T. A. Grandine, Z. J. Kabala, H. J. Seiffert, and W. B. Jordan. An asymptotic problem. SIAM Review, 28(2): 234–238, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Veliz-Cuba:2016:SME

[VCKJ16] Alan Veliz-Cuba, Zachary P. Kilpatrick, and Kresimir Josić. Stochastic models of evidence accumulation in changing environments. SIAM Review, 58(2): 264–289, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Veinott:1968:IEP

[VD68] Arthur F. Veinott, Jr. and George B. Dantzig. Integral extreme points. SIAM Review, 10(3):371–372, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanDoorn:1981:II

[vD81] E. A. van Doorn. Inequality or identity. SIAM Review, 23(4):524, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vono:2022:HDG

[VDC22] Maxime Vono, Nicolas Dobigeon, and Pierre Chainais. High-dimensional Gaussian sampling: A review and a unifying approach based on a stochastic proximal point algorithm. SIAM Review, 64(1):3–56, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1371026.

Vomel:2017:SLK

[VdLBF17] Christof Vömel, Flavio de Lorenzi, Samuel Beer, and Erwin Fuchs. The secret life of keys: On the calculation of mechanical lock systems. SIAM Review, 59(2):393– 422, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanderMeiden:1968:BRA

[vdM68] W. van der Meiden. Book review: On the Average of a Convex Function (Howell K. Wilson). SIAM Review, 10 (1):115–116, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanderVorst:1994:BRP

[vdV94] Henk van der Vorst. Book review: Parallel Numerical Algorithms (T. L. Freeman and C. Phillips). SIAM Review, 36(4):678–679, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL http://www.
                                        siam.org/journals/sirev/
                                        sirev364.htm.

Vein:1967:ETS

[Vei67] P. R. Vein. On existence of two sequences. SIAM Review, 9(2):250–251, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vein:1986:CMD

[Vei86] P. R. Vein. A conjectured minimum degree determinantal polynomial. SIAM Review, 28(4):567, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Veltkamp:1963:PSD

[Vel63] G. W. Veltkamp. On a periodic solution of a differential equation. SIAM Review, 5 (4):370–371, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Veltkamp:1964:PRS

[Vel64] G. W. Veltkamp. A property of real solutions to Bessel’s equation. SIAM Review, 6 (4):455–456, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Veltkamp:1971:BRH

[Vel71] G. W. Veltkamp. Book review: A Heat Transfer Problem (J. Ernest Wilkins, Jr.). SIAM Review, 13(4): 575–577, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Veldman:2021:SFV

[Vel21] Arthur E. P. Veldman.

Supraconservative finite-volume methods for the Euler equations of subsonic compressible flow. SIAM Review, 63 (4):756–779, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ventura:1972:CMD

[Ven72] J. R. Ventura, Jr. Construction of a maximal (0, 1) determinant. SIAM Review, 14 (2):366, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Venit:1974:GVD

[Ven74] S. Venit. A generalization of the Vandermonde determinant. SIAM Review, 16 (3):391, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Venit:1975:GVD

[Ven75]  S. Venit.  Generalization

of Vandermonde determinant. SIAM Review, 17(4): 694–695, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Venit:1976:MSP

[Ven76] S. Venit. A matrix stability problem. SIAM Review, 18(2):295, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Verlinde:1990:TIA

[Ver90] Johannes Verlinde. Two integrals arising from a cloud model. SIAM Review, 32 (3):478, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Verriest:1998:EUV

[Ver98] E. Verriest. Existence and uniqueness for a variational problem. SIAM Review, 40 (1):132–143, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Verzani:2009:CSH

[Ver09]  John Verzani.  Compu-

tational statistics handbook with MATLAB. Second edition [book review of MR2404036]. SIAM Review, 51(3):654– 656, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Verhulst:2015:PPT

[Ver15] Ferdinand Verhulst. Profits and pitfalls of timescales in asymptotics. SIAM Review, 57(2):255–274, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Veselic:1995:FCM

[Ves95]  K. Veselić.  Finite cate-

nary and the method of Lagrange. SIAM Review, 37 (2):224–229, June 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/23998.htm.

Vetter:1973:MCO

[Vet73] William J. Vetter. Matrix calculus operations and Taylor expansions. SIAM Review, 15(2):352–369, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanLoan:1978:SUP

[VG78] C. Van Loan and G. H. Golub. Solving unsymmetric positive definite linear systems. SIAM Review, 20 (3):636, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

VanderJeugt:1996:HPI

[VG96] J. Van der Jeugt and C. C. Grosjean. A Hermite polynomial identity. SIAM Review, 38(3):514, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vatsa:1978:MGP

[VH78] V. N. Vatsa and J. E. Harris. Minimization of gridpoint requirements for solution of boundary-layer equations. SIAM Review, 20 (3):636, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanHemmen:1987:BRT

[vH87] J. L. van Hemmen. Book review: The Theory of Thermodynamics (J. R. Waldram). SIAM Review, 29(3): 498–499, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vial:1985:BRN

[Via85] Jean-Phillippe Vial. Book review: Nonlinear Programming (Garth P. McCormick). SIAM Review, 27(1):128– 130, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vielma:2015:MIL

[Vie15] Juan Pablo Vielma. Mixed integer linear programming formulation techniques. SIAM Review, 57(1):3–57, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vijay:1983:CP

[Vij83] J. Vijay. A covering problem. SIAM Review, 25(1): 97–98, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vilms:1974:DTF

[Vil74]  Jaak Vilms.  Differentia-

tion of tensor fields. II. SIAM Review, 16(3):303– 308, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Viswanath:2001:LPT

[Vis01] Divakar Viswanath. The Lindstedt–Poincaré technique as an algorithm for computing periodic orbits. SIAM Review, 43(3):478– 495, September 2001. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/37529.

VanTuyl:1983:BRD

[VJ83] Andrew H. Van Tuyl and Harry L. Johnson. Book review: A Definite Integral (Peter Henrici). SIAM Review, 25(4):572–573, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Veltkamp:1962:BRF

[VK62] G. W. Veltkamp and M. S. Klamkin. Book review: Flight in An Irrotational Wind Field (M. S. Klamkin). SIAM Review, 4(2):155– 156, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanKan:1979:BRI

[vK79] J. van Kan. Book review: An Integral Inequality (A. Meir). SIAM Review, 21(4): 567–568, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Verma:1984:HRM

[VK84] Ghasi R. Verma and Joseph B. Keller. Hanging rope of minimum elongation. SIAM Review, 26(4):569–571, October 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vankan:1979:II

[VKK79] J. Vankan, G. K. Kristiansen, and M. K. Kwong. Integral inequality. SIAM Review, 21 (4):567–568, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanLint:1960:ASP

[vL60] J. H. van Lint. Another sorting problem. SIAM Review, 2(3):217, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanLint:1961:ASP

[vL61] J. H. van Lint. Another sorting problem. SIAM Review, 3 (4):336–337, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanLeer:1986:BRCb

[vL86a] Bram van Leer. Book review: Computational Fluid Mechanics and Heat Transfer (Dale A. Anderson, John C. Tannehill and Richard H. Pletcher). SIAM Review, 28 (4):602–603, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanLeer:1986:BRCa

[vL86b] Bram van Leer. Book review: Computational Methods for Fluid Flow (Roger Peyret and Thomas D. Taylor). SIAM Review, 28(3): 440–442, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanLeeuwaarden:2019:ESM

[vLMZ19] Johan S. H. van Leeuwaarden, Britt W. J. Mathijsen, and Bert Zwart. Economies-of-scale in manyserver queueing systems: Tutorial and partial review of the QED Halfin– Whitt heavy-traffic regime. SIAM Review, 61(3):403– 440, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Volkening:2020:FEU

[VLPR20] Alexandria Volkening, Daniel F. Linder, Mason A. Porter, and Grzegorz A. Rempala. Forecasting elections using compartmental models of infection. SIAM Review, 62(4): 837–865, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vishik:1995:ABS

[VM95] M. I. Vishik and Alexander Mielke. Asymptotic behaviour of solutions of evolutionary equations. SIAM Review, 37(3):471–??, September 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ventura:1973:BRC

[VN73] J. R. Ventura, Jr. and P. J.

Nikolai. Book review: Construction of a Maximal (0, 1) Determinant (J. R. Ventura, Jr.). SIAM Review, 15(2): 391, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vagners:1976:OLH

[VN76a] J. Vagners and R. D.

Neal. Optimal laser heating of a magnetically confined plasma. SIAM Review, 18(4): 831–832, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Victory:1976:MPS

[VN76b] H. D. Victory and P. Nelson.

Measures on phase space as solutions of one-dimensional neutron-transport equation. SIAM Review, 18(4):832, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vogt:1987:BRA

[Vog87] Andrew Vogt. Book review: An Axiomatic Basis for Quantum Mechanics: Vol. 1, Derivation of Hilbert Space Structure (Günther Ludwig). SIAM Review, 29 (3):499–501, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vogt:1991:BRI

[Vog91a] Andrew Vogt. Book review: An Introduction to Hilbert Space and Quantum Logic (David W. Cohen). SIAM Review, 33(1): 160–161, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vogt:1991:IHS

[Vog91b] Andrew Vogt. An introduction to Hilbert space and quantum logic (David W Cohen). SIAM Review, 33(1): 160–??, March 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vogt:1993:BRQ

[Vog93] Andrew Vogt. Book review: Quantum Mechanics, Algebras and Distributions (D. A. Dubin and M. A. Hennings). SIAM Review, 35(2): 343–345, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Voss:1975:SRM

[Vos75] D. A. Voss. The spectral radius of a matrix. SIAM Review, 17(1):169– 170, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Voss:1976:SRM

[Vos76] D. A. Voss. Spectral radius of a matrix. SIAM Review, 18 (2):304–306, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vanderbei:2009:PAP

[VP09] Robert J. Vanderbei and Mustafa Ç. Pınar. Pricing American perpetual warrants by linear programming. SIAM Review, 51(4): 767–782, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vasko:1995:OCC

[VS95] Francis J. Vasko and Kenneth L. Stott. Optimizing continuous Caster product dimensions: An example of a nonlinear design problem in the steel industry. SIAM Review, 37(1): 82–84, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/26136.htm.

Viswanath:2010:CSL

[VŞ10] Divakar Viswanath and Sönmez Şahutoğlu. Complex singularities and the Lorenz attractor. SIAM Review, 52 (2):294–314, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vanderbei:2014:RAF

[VSM14] Robert J. Vanderbei, Gordon Scharf, and Daniel Marlow. A regression approach to fairer grading. SIAM Review, 56 (2):337–352, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanVleck:1965:BRF

[vV65] F. S. van Vleck. Book review: Fixed Points and Topological Degree in Nonlinear Analysis (Jane Cronin). SIAM Review, 7(1):141–143, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanWoerkom:1990:OSB

[vW90] P. Th. L. M. van Woerkom. Optimization of a stability bound. SIAM Review, 32 (2):300, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

vanWoerkom:1991:BRO

[vWHE91] P. Th. L. M. van Woerkom, P. Th. L. M. Hanau, and Mansour Eslami. Book review: Optimization of a Stability Bound (P. Th. L. M. van Woerkom). SIAM Review, 33(2):277–281, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wouk:1963:BRI

[WA63] A. Wouk and U. S. Army.

Book review: An Introduction to Electronic Analogue Computers (M. G. Hartley). SIAM Review, 5(3): 296, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wittmann:1994:IOC

[WA94] R. C. Wittmann and B. K.

Alpert. An integral occurring in coherence theory. SIAM Review, 36(4): 655, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wachspress:1989:EFI

[Wac89] E. L. Wachspress. An elliptic function identity. SIAM Review, 31(1):119, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wagner:1963:UP

[Wag63] Daniel H. Wagner. An urn problem. SIAM Review, 5 (1):73, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wagner:1969:NFV

[Wag69] Daniel H. Wagner. Nonlinear functional versions of the Neyman–Pearson lemma.

SIAM Review, 11(1):52– 65, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wagner:1992:BRL

[Wag92] Peter Wagner. Book review: Limit of a Definite Integral (C. A. Oster). SIAM Review, 34(1):122–123, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wahba:1965:LSE

[Wah65]  Grace Wahba.  A least

squares estimate of satellite attitude. SIAM Review, 7 (3):409, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wahba:1968:FSF

[Wah68] Grace Wahba. A family of summation formulae. SIAM Review, 10(1):109, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wahlbin:1992:BRN

[Wah92] Lars B. Wahlbin. Book review: Nonlinear Elliptic and Evolution Problems and Their Finite Element Approximations (A. Ženišek). SIAM Review, 34(2):346– 347, June 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wahlberg:2003:ORF

[Wah03] Bo Wahlberg. Orthogonal rational functions: a transformation analysis. SIAM Review, 45(4):689–705, December 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37663.

Wake:1986:NDE

[Wak86]  G. C. Wake.  A non-

linear difference equation. SIAM Review, 28(4):566– 567, December 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walsh:1963:BRA

[Wal63a] J. L. Walsh. Book review:

Analytic Function Theory (E. Hills). SIAM Review, 5(4): 377–378, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waltman:1963:OSN

[Wal63b] P. Waltman. Oscillation of solutions of a nonlinear equation. SIAM Review, 5(2): 128–130, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walling:1966:LAF

[Wal66] Derald Walling. A least absolute fit. SIAM Review, 8 (1):107, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walter:1970:FSA

[Wal70]  Gilbert Walter.  Fourier

series and analytic representation of distributions. SIAM Review, 12(2):272– 276, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walter:1975:II

[Wal75a] J. Walter. An integral inequality. SIAM Review, 17 (4):685, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waltman:1975:BRM

[Wal75b] Paul Waltman. Book review: Models in Ecology (J. Maynard Smith). SIAM Review, 17(4):716–717, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walter:1976:NC

[Wal76a] Wolfgang Walter. A note on contraction. SIAM Review, 18(1):107–111, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walton:1976:QUD

[Wal76b] J. R. Walton. Question of uniqueness for dual integral equations of Titchmarsh type. SIAM Review, 18 (4):832, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wallis:1979:BRG

[Wal79] W. D. Wallis. Book review:

Graph Theory with Applications (J. A. Bondy and U. S. R. Murty). SIAM Review, 21(3):429, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wallace:1985:IMT

[Wal85] S. W. Wallace. Incidence matrix for a transportation problem. SIAM Review, 27 (1):79, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wallace:1986:STP

[Wal86] S. W. Wallace. A stochastic transportation problem. SIAM Review, 28(3):397, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wallace:1987:BRI

[Wal87] S. W. Wallace. Book review: Incidence Matrix for a Transportation Problem (S. W. Wallace). SIAM Review, 29(4):633, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wallis:1988:BRP

[Wal88] W. D. Wallis. Book review: Preference Modelling (Marc Roubens and Philippe Vincke). SIAM Review, 30 (3):527–528, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walter:1992:WLP

[Wal92] Johann Walter. World lines of photons and distributional solutions with small support of the homogeneous wave equation. SIAM Review, 34 (4):649, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walnut:1993:BRW

[Wal93a] David F. Walnut. Book review: Wavelets: a Tutorial in Theory and Applications (Charles K Chui, ed.). SIAM Review, 35(2): 345–348, June 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walter:1993:WLP

[Wal93b] J. Walter. World lines of photons and distributed solutions with small support of the homogeneous wave equation. SIAM Review, 35(4): 647–649, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walter:1993:BRW

[Wal93c] Johann Walter. Book review:

World Lines of Photons and Distributional Solutions with Small Support of the Homogeneous Wave Equation (Johann Walter). SIAM Review, 35(4):647–649, ???? 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walker:1995:BRD

[Wal95] Russell Walker. Book review: Dynamical Systems: Stability, Dynamics, and Chaos (Clark Robinson). SIAM Review, 37(4):639– 640, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walnut:1997:BRW

[Wal97] David Walnut. Book review: Wavelets: An Analysis Tool, by M. Holschneider. SIAM Review, 39(1): 172–174, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walkington:2023:NMC

[Wal23] Noel J. Walkington. Nesterov’s method for convex optimization. SIAM Review, 65(2):539–562, ???? 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1390037.

Wang:1979:DIC

[Wan79] Chung Lie Wang. On development of inverses of the Cauchy and Hölder inequalities. SIAM Review, 21(4): 550–557, October 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wang:1983:PRF

[Wan83] P. H. Wang. Pattern recognition with fuzzy objective function algorithms — Bezdek, J. C. SIAM Review, 25(3):442, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wanner:1984:CNR

[Wan84] G. Wanner. A conjecture on the number of real zeros. SIAM Review, 26(2): 273–275, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wan:1991:BRH

[Wan91] Yieh-Hei Wan. Book review: The Hyperbolic Map and Applications to the Linear Quadratic Regulator (B. J. Daiuto, T. T. Hartley, and S. P. Chicatelli). SIAM Review, 33(2):338, June 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wang:1997:PEU

[Wan97] Yongji Wang. Problems: Existence and uniqueness for a variational problem (problem 97-4). SIAM Review, 39(1): 125–126, March 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Warren:1964:BRF

[War64] Joseph E. Warren. Book review: Flame Propagation (William Squire). SIAM Review, 6(1):66–69, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Warga:1968:RCC

[War68] J. Warga. The reduction of certain control problems to an “ordinary differential” type. SIAM Review, 10(2): 219–222, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Warren:1969:TSI

[War69] William E. Warren. Two singular integrals. SIAM Review, 11(3):398, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ward:1970:MSB

[War70] L. E. Ward, Sr. Motion of a sliding bead. SIAM Review, 12(3):469, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ward:1971:MSB

[War71] L. E. Ward, Sr. Motion of a sliding bead. SIAM Review, 13(3):407–408, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ward:1975:CSA

[War75] R. C. Ward. Combination shift QZ algorithm. SIAM Review, 17(2):391, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ward:1976:NCM

[War76] R. C. Ward. Numerical computation of matrix exponential. SIAM Review, 18(4): 832–833, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Warchall:1992:BRS

[War92] Henry Warchall. Book review: Solitons in Physics, Mathematics, and Nonlinear Optics (P. J. Olver and D. H. Sattinger, eds.). SIAM Review, 34(4):689– 691, December 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ware:1998:FAF

[War98]  Antony F. Ware.  Fast

approximate Fourier transforms for irregularly spaced data. SIAM Review, 40 (4):838–856, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31533.

Wasserstrom:1973:NSC

[Was73] E. Wasserstrom. Numerical solutions by the continuation method. SIAM Review, 15 (1):89–119, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wasow:1974:BRA

[Was74]  Wolfgang Wasow.  Book

review: Asymptotic Expansions: Their Derivation and Interpretation (R. B. Dingle). SIAM Review, 16(4): 568–570, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wasserstrom:1975:EP

[Was75] E. Wasserstrom. An eigenvalue problem. SIAM Review, 17(3):567–568, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wasilkowski:1994:BRC

[Was94a] Grzegorz W. Wasilkowski.

Book review: The Computational Complexity of Differential and Integral Equations: an Information-Based Approach (Arthur G. Werschulz). SIAM Review, 36 (3):525–526, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wasow:1994:BRS

[Was94b] Wolfgang Wasow. Book review: Singular Perturbation Methods for Ordinary Differential Equations (Robert E. O’Malley, Jr.). SIAM Review, 36(1):149, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waterman:1963:BRI

[Wat63] D. Waterman. Book review: An Introduction to Fourier Analysis (R. D. Stuart). SIAM Review, 5(2): 160, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waterman:1965:BRR

[Wat65] D. Waterman. Book review: Real Analysis (H. L. Royden). SIAM Review, 7(3): 436–437, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:1982:UA

[Wat82]  David S. Watkins.  Un-

derstanding the QR algorithm. SIAM Review, 24(4): 427–440, October 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:1984:IF

[Wat84] David S. Watkins. Isospectral flows. SIAM Review, 26 (3):379–391, July 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:1986:BRE

[Wat86a] David S. Watkins. Book review: Eigenvalues and Eigenvectors of a Large Matrix (N. Oden). SIAM Review, 28(2): 240–241, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watson:1986:BRG

[Wat86b] G. S. Watson. Book review: Generalized Linear Models (P. Mccullagh and J. A. Nelder). SIAM Review, 28 (1):128–130, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watson:1986:NLA

[Wat86c] Layne T. Watson. Numerical linear algebra aspects of globally convergent homotopy methods. SIAM Review, 28 (4):529–545, December 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watson:1990:BRA

[Wat90] John G. Watson. Book review: An Average Maximum Distance (Peter Senn). SIAM Review, 32(2):304– 305, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waterhouse:1991:OMM

[Wat91a] W. C. Waterhouse. Orthogonal matrices and the matrix exponential. SIAM Review, 33(4):661–662, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waterhouse:1991:BRO

[Wat91b] William C. Waterhouse. Book review: Orthogonal Matrices and the Matrix Exponential (Dennis S. Bernstein). SIAM Review, 33(4): 661–662, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:1993:SPE

[Wat93] David S. Watkins. Some perspectives on the eigenvalue problem. SIAM Review, 35 (3):430–471, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waterhouse:1996:SBP

[Wat96] William C. Waterhouse. Some boundary problems with even or odd solutions. SIAM Review, 38 (4):645–646, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28649.

Watkins:2005:PEP

[Wat05] David S. Watkins. Product eigenvalue problems. SIAM Review, 47(1):3–40, March 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/44311.

Watkins:2008:AR

[Wat08] David S. Watkins. The QR algorithm revisited. SIAM Review, 50(1):133–145, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2010:BRM

[Wat10] David S. Watkins. Book review: Matrix Mathematics: Theory, Facts, and Formulas. Second Edition. SIAM Review, 52(4):790– 791, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2015:BRa

[Wat15a] David S. Watkins. Book reviews. SIAM Review, 57(1):

153–164, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2015:BRb

[Wat15b] David S. Watkins. Book reviews. SIAM Review, 57(2): 295–318, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2015:BRc

[Wat15c] David S. Watkins. Book reviews. SIAM Review, 57(3): 469–480, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2015:BRd

[Wat15d] David S. Watkins. Book reviews. SIAM Review, 57(4): 627–643, ???? 2015. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2016:BRa

[Wat16a] David S. Watkins. Book reviews. SIAM Review, 58(1): 159–173, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2016:BRb

[Wat16b] David S. Watkins. Book reviews. SIAM Review, 58(2): 357–374, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2016:BRc

[Wat16c] David S. Watkins. Book reviews. SIAM Review, 58(3): 577–599, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2016:BRd

[Wat16d] David S. Watkins. Book reviews. SIAM Review, 58(4): 795–807, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2017:BRa

[Wat17a] David S. Watkins. Book reviews. SIAM Review, 59(1): 207–222, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2017:BRb

[Wat17b] David S. Watkins. Book reviews. SIAM Review, 59(2): 449–465, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2017:BRc

[Wat17c] David S. Watkins. Book reviews. SIAM Review, 59(3): 681–699, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:2017:BRd

[Wat17d] David S. Watkins. Book reviews. SIAM Review, 59(4): 915–928, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wathen:2022:SCP

[Wat22] Andy Wathen. Some comments on preconditioning for normal equations and least squares. SIAM Review, 64 (3):640–649, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1387948.

Waxler:1991:BRC

[Wax91a] Roger Waxler. Book review: A Course in Mathematics for Students of Physics. Volume 2 (Paul Bamberg and Shlomo Sternberg). SIAM Review, 33 (4):670–671, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waxler:1991:CMS

[Wax91b] Roger Waxler. A course in mathematics for students of physics, vol. 2 (Paul Bamberg and Shlomo Sternberg). SIAM Review, 33(4): 670–??, December 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waymire:1982:PLM

[Way82] E. Waymire. A probabilistic look at mixing and cooling processes. SIAM Review, 24(1):73–75, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1976:VDC

[WB76] S. S. Wilson and A. C. Buckingham. Vaporizing droplet in a convective field. SIAM Review, 18(4):833, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wadbro:2009:MTO

[WB09] Eddie Wadbro and Martin Berggren. Megapixel topology optimization on a graphics processing unit. SIAM Review, 51(4):707– 721, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1991:BRT

[WBGP91] James A. Wilson, W. Bohm, C. C. Grosjean, and D. Pickett. Book review: A Trigonometric Sum (W. Böhm). SIAM Review, 33(1):124, ???? 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1977:BRZ

[wBJ77] w. B. Jordan. Book review:

A Zero of Maximum Multiplicity (N. Liron and L. A. Rubenfeld). SIAM Review, 19(4):743–744, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wermuth:1989:CME

[WBLQ89] E. M. E. Wermuth, J. Brenner, F. S. Leite, and J. F. Queiro. Commuting matrix exponentials. SIAM Review, 31(1):125–126, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wong:1977:ADD

[WC77] C. K. Wong and Kai-Ching C. Chu. Average distances in L_{p} disks. SIAM Review, 19(2):320– 324, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watkins:1986:EEL

[WC86] D. S. Watkins and C. R. Crawford. Eigenvalues and eigenvectors of a large matrix. SIAM Review, 28(2): 240–241, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wickerhauser:1996:AWA

[WC96] Mladen Victor Wickerhauser and Charles K. Chui. Adapted wavelet analysis from theory to software. SIAM Review, 38 (1):160–??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wedin:1976:ICN

[WE76] P. A. Wedin and B. Ekenberg. Ill-conditioned nonlinear parameter-estimation problems. SIAM Review, 18 (4):833, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:2019:APR

[WE19] Dan Wilson and Bard Ermentrout. Augmented phase reduction of (not so) weakly perturbed coupled oscillators. SIAM Review, 61(2): 277–315, ???? 2019. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Webb:1985:BRS

[Web85] Glenn F. Webb. Book review:

Semigroups of Linear Operators and Applications to Partial Differential Equations (A. Pazy). SIAM Review, 27 (2):309–311, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weissblum:1960:SCE

[Wei60a] Walter Weissblum. A set of convolution equations. SIAM Review, 2(3):217– 218, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weissblum:1960:SP

[Wei60b] Walter Weissblum. A sorting problem. SIAM Review, 2(4):298, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weissblum:1962:SN

[Wei62] Walter Weissblum. Spatial navigation. SIAM Review, 4 (3):255, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weissblum:1963:ADL

[Wei63a] Walter Weissblum. Asymptotic distribution of lattice points in a random rectangle. SIAM Review, 5(2): 156, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weissblum:1963:RT

[Wei63b] Walter Weissblum. Reliability theory. SIAM Review, 5 (4):372, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weissblum:1967:BRV

[Wei67a] Walter Weissblum. Book review: A Variation of a Theorem of Kakutani (L. Shepp). SIAM Review, 9(2):

253–254, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weissblum:1967:BRS

[Wei67b] Walter Weissblum. Book review: Sorting Problem (Walter Weissblum). SIAM Review, 9(1):121–122, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weil:1968:GTE

[Wei68] Roman L. Weil, Jr. Game theory and eigensystems. SIAM Review, 10(3):360– 367, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weiss:1969:BRP

[Wei69] George H. Weiss. Book review: Probabilistic Methods in Applied Mathematics, Volume 1 (A. T. Bharucha-Reid, ed.). SIAM Review, 11(1): 96–98, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weiss:1977:BRF

[Wei77] George H. Weiss. Book review: A First Course in Stochastic Processes, 2nd sd. (Samuel Karlin and Howard M. Taylor). SIAM Review, 19 (1):158–159, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weiss:1979:BRT

[Wei79] George H. Weiss. Book review: The Theory of Gambling and Statistical Logic (Richard A. Epstein). SIAM Review, 21(2):267–268, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weinberg:1980:EPM

[Wei80a] Alvin M. Weinberg. Energy policy and mathematics. SIAM Review, 22(2): 204–212, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weinstein:1980:BRT

[Wei80b] Alan Weinstein. Book review: Theoretical Kinematics (O. Bottema and B. Roth). SIAM Review, 22(4): 519–520, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weinstein:1984:EMG

[Wei84a] A. Weinstein. The elements of mechanics — Gallavotti, G. SIAM Review, 26(4): 598, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weinstein:1984:BRE

[Wei84b] Alan Weinstein. Book review:

The Elements of Mechanics (Giovanni Gallavotti). SIAM Review, 26(4):598, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weiser:1987:BRE

[Wei87] Alan Weiser. Book review:

Elliptic Problem Solvers II (Garrett Birkhoff and Arthur Schoenstadt, eds.). SIAM Review, 29(3):502–503, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weinstein:1988:BRM

[Wei88] Alan Weinstein. Book review: Manifolds and Mechanics (Arthur Jones, Alistair Gray, and Robert Hutton). SIAM Review, 30 (3):530, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wei:1990:MI

[Wei90a] Gaoyuan Wei. A multidimensional integral. SIAM Review, 32(3):479, ???? 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weitzner:1990:BRI

[Wei90b] Harold Weitzner. Book review: Introduction to Nonlinear Fluid-Plasma Waves (Bhimsen K. Shivamoggi). SIAM Review, 32(3):520– 521, September 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weiser:1991:BRM

[Wei91] Alan Weiser. Book review: Multilevel Adaptive Methods for Partial Differential Equations (Stephen F. Mc-Cormick). SIAM Review, 33 (4):680–681, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Welland:1988:BRH

[Wel88] Grant V. Welland. Book review: The Hartley Transform (Ronald N. Bracewell). SIAM Review, 30(1):170, ???? 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Welland:1989:BRT

[Wel89] Grant V. Welland. Book review: Z Transform Theory and Applications (Robert Vich). SIAM Review, 31(1): 171–172, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wendroff:1961:SCF

[Wen61] Burton Wendroff. The structure of certain finite difference schemes. SIAM Review, 3(3):237–242, ???? 1961. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wendroff:1968:BRD

[Wen68] Burton Wendroff. Book review: Difference Methods for Initial-Value Problems (Robert D. Richtmyer and K. W. Morton). SIAM Review, 10(3):381–383, ???? 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wendroff:1978:BRI

[Wen78]  Burton Wendroff.  Book

review: An Introduction to the Mathematical Theory of Finite Elements (J. T. Oden and J. N. Reddy). SIAM Review, 20(2):407– 408, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wertz:1963:BRM

[Wer63] H. J. Wertz. Book review:

Matrix Methods for Engineering (Louis A. Pipes). SIAM Review, 5(4):380– 381, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wermuth:1989:BRC

[Wer89] Edgar M. E. Wermuth. Book review: Commuting Matrix Exponentials (Dennis S. Bernstein). SIAM Review, 31 (1):125–126, ???? 1989. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Westmann:1965:SPD

[Wes65] R. A. Westmann. Simultaneous pairs of dual integral equations. SIAM Review, 7 (3):341–348, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Westwater:1984:QML

[Wes84a] J. Westwater. Quantum mechanics of large systems — Thirring, W. SIAM Review, 26(4):612–615, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Westwater:1984:BRQ

[Wes84b] John Westwater. Book review: Quantum Mechanics of Large Systems (Walter Thirring). SIAM Review, 26 (4):612–615, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wester:1992:BRM

[Wes92] Michael Wester. Book review: Mathematics: a System for Doing Mathematics by Computer, Second Edition (Stephen Wolfram). SIAM Review, 34(3):519– 522, September 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wets:1974:BRF

[Wet74a]  Roger Wets.  Book re-

view: Functional Analysis in Mathematical Economics. Optimization over Infinite Horizons (Lester G. Telser and Robert L. Graves). SIAM Review, 16(2):270– 271, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wets:1974:SPF

[Wet74b] Roger J.-B. Wets. Stochastic programs with fixed recourse: The equivalent deterministic program. SIAM Review, 16 (3):309–339, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wets:1975:BRO

[Wet75] Roger J. B. Wets. Book review: Optimization Methods for Resource Allocation (Richard Cottle and Jacob Krarup, eds.). SIAM Review, 17(4):715–716, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wetterling:1993:BRP

[Wet93] Wolfgang Wetterling. Book review: Parametric Optimization: Singularities, Pathfollowing and Jumps (J. Guddat and F. Guerra Vasquez with H. Th. Jongen). SIAM Review, 35 (1):172, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weyl:1962:PCG

[Wey62]  F. Joachim Weyl.  The

perplexing case of geometry. SIAM Review, 4(3): 197–201, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weyl:1967:VSC

[Wey67]  F. J. Weyl.  VII. sum-

mary of the conference. SIAM Review, 9(2):411– 415, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wijnberg:1976:BFS

[WG76] L. Wijnberg and M. L. Glasser. A Bessel function series. SIAM Review, 18 (2):296, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wijnberg:1977:BFSa

[WG77a] L. Wijnberg and M. L. Glasser. A Bessel function series. SIAM Review, 19 (1):148, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wijnberg:1977:BFSb

[WG77b] L. Wijnberg and M. L.

Glasser. A Bessel function series. SIAM Review, 19(2): 332–333, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Woodcock:1993:DI

[WG93] C. F. Woodcock and P. R.

Gravesmorris. A determinant identity. SIAM Review, 35 (4):649–651, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watson:1990:AMD

[WGG^{+}90] J. G. Watson, M. X.

Goemans, J. A. Grzesik, D. A. Kurtze, T. S. Lewis, O. P. Lossers, I. Mareels, F. Mathis, D. Ullman, and J. R. Varma. An average maximum distance. SIAM Review, 32(2):304– 305, ???? 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wagner:1992:LDI

[WGG^{+}92] P. Wagner, A. Gibbs, N. L.

Glasser, C. C. Grosjean, W. B. Jordan, N. Ortner, D. Pickett, M. Renardy, P. Svobodny, and J. A. Wilson. Limit of a definite integral. SIAM Review, 34(1): 122–123, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1992:ISB

[WGLS92] J. A. Wilson, C. C. Grosjean, O. P. Lossers, and M. Stamp. An identity for sums of binomial coefficients. SIAM Review, 34(1): 123–126, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Woodcock:1993:BRD

[WGM93] C. F. Woodcock and P. R. Graves-Morris. Book review: A Determinant Identity (P. R. Graves-Morris). SIAM Review, 35(4):649– 651, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wheeler:1974:BRW

[Whe74] Lewis T. Wheeler. Book review: Wave Propagation in Elastic Solids (J. D. Achenbach). SIAM Review, 16 (4):558, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wheeler:1976:BRB

[Whe76] Roger F. Wheeler. Book review: Boolean Functions and Equations (Sergiu Rudeanu). SIAM Review, 18(2):316– 317, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1963:BRC

[Whi63a] A. M. White. Book review: Complex Numbers and Conformal Mapping (A. I. Markusuevich). SIAM Review, 5(1):79, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1963:BRF

[Whi63b] Alvin M. White. Book review: Functions of a Complex Variable: and Some of Their Applications, Volume II (B. A. Fuchs and V. I. Levin). SIAM Review, 5(4): 378–379, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1966:BRF

[Whi66a] A. M. White. Book review: Functions of a Complex Variable (Gino Moretti). SIAM Review, 8(1):121– 122, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1966:BRS

[Whi66b] Alvin M. White. Book review: Selected Topics in the Classical Theory of Functions of a Complex Variable (Maurice Heins). SIAM Review, 8(4):552, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1967:BRA

[Whi67] Alvin M. White. Book review: Analytic Functions (M. A. Evgrafov). SIAM Review, 9(4):755–756, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1969:BRI

[Whi69] Lee J. White. Book review: Introduction to Combinatorial Mathematics (C. L. Liu). SIAM Review, 11(4): 634–635, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1976:BRC

[Whi76a]  Alvin M. White.  Book

review: Complex Variables (George Polya and Gordon Latta). SIAM Review, 18(3): 531–532, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1976:BRL

[Whi76b] Lee J. White. Book review: Linear Numerical Analysis (Noel Gastinel). SIAM Review, 18(1):149– 152, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Whitehead:1980:BRC

[Whi80] Earl Glen Whitehead, Jr.

Book review: Combinatorial Algorithms for Computers and Calculators; 2nd Edition (Albert Nijenhuis and Herbert S. Wilf). SIAM Review, 22(2):238–239, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Whitehead:1981:BRC

[Whi81] Earl Glen Whitehead, Jr.

Book review: Combinatorial Complexes: a Mathematical Theory of Algorithms (Peter H. Sellers). SIAM Review, 23 (2):263–264, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Whitehead:1983:BRG

[Whi83] Earl Glen Whitehead, Jr.

Book review: Graphs, Networks and Algorithms (M. N. S. Swamp and K. Thulasiraman). SIAM Review, 25 (1):115, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Whitehead:1984:CEG

[Whi84a] E. G. Whitehead. Combinatorial enumeration — Goulden, I. P., Jackson, D. M. SIAM Review, 26(4): 599, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Whitehead:1984:BRC

[Whi84b] Earl Glen Whitehead, Jr. Book review: Combinatorial Enumeration (Ian P. Goulden and David M. Jackson). SIAM Review, 26 (4):599, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Whitehead:1988:BRE

[Whi88] Earl Glen Whitehead, Jr. Book review: Enumerative Combinations Volume I (Richard P. Stanley). SIAM Review, 30(1):170– 171, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1992:BRS

[Whi92a] Donald B. White. Book review: Sampling of Populations: Methods and Applications (Paul Levy and Stanley Lemeshow). SIAM Review, 34(2):347–349, ???? 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

White:1992:SPM

[Whi92b] Donald B. White. Sampling of populations: Methods and applications (Paul Levy and Stanley Lemeshow). SIAM Review, 34(2):347–??, June 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Whitaker:1993:BRI

[Whi93] Stephen Whitaker. Book review: Interfacial Transport Phenomena (John C. Slattery). SIAM Review, 35(1): 172–174, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Whittle:1994:BRO

[Whi94] Geoff Whittle. Book review: Oriented Matroids (A. Bjorner, M. Las Vergnas, B. Sturmfels, N. White, and G. Ziegler). SIAM Review, 36 (4):679–680, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

West:2016:EMT

[WHMN16] Jeffrey West, Zaki Hasnain, Paul Macklin, and Paul K. Newton. An evolutionary model of tumor cell kinetics and the emergence of molecular heterogeneity driving Gompertzian growth. SIAM Review, 58(4):716– 736, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walz:2024:EUQ

[WHZG24] Eva-Maria Walz, Alexander Henzi, Johanna Ziegel, and Tilmann Gneiting. Easy uncertainty quantification (EasyUQ): Generating predictive distributions from single-valued model output. SIAM Review, 66(1):91– 122, February 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wiberg:1987:BRI

[Wib87] Donald M. Wiberg. Book review: Instrumental Variable Methods for System Identification (Torsten Söderström and Petre Gheorghe Stoica). SIAM Review, 29(3): 503–505, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wickerhauser:1994:BRWa

[Wic94a] M. Victor Wickerhauser. Book review: Wavelets: Algorithms and Applications (Yves Meyer). SIAM Review, 36(3):526–528, ???? 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wickerhauser:1994:BRWb

[Wic94b] Mladen Victor Wickerhauser. Book review: Wavelets: Algorithms and Applications by Yves Meyer. SIAM Review, 36(3):526–528, September 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Book review.

Wiegand:1987:BRL

[Wie87] Sylvia Wiegand. Book review: Linear Systems Over Commutative Rings (James W. Brewer, John W. Bunce and F. S. van Vleck). SIAM Review, 29(4):678– 680, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wigner:1967:RMP

[Wig67] Eugene P. Wigner. Random matrices in physics. SIAM Review, 9(1):1–23, January 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://link.
aip.org/link/siread/v9/
i1/p1/s1;   http://www.
jstor.org/stable/2027409.

Wightman:1985:BRS

[Wig85] A. S. Wightman. Book review: Studies in Applied Mathematics (Victor Guillemin, ed.). SIAM Review, 27(1):127–128, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wightman:1990:BRR

[Wig90] Arthur S. Wightman. Book review: Riemann, Topology, and Physics, by Michael Monastyrsky, Freeman J. Dyson, James King, Victoria King, R. O. Wells, Jr. SIAM Review, 32(1): 192, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        2030418.

Wightman:1991:BRF

[Wig91] Arthur S. Wightman. Book review: Finite Quantum Electrodynamics (G. Scharf). SIAM Review, 33(3):508– 509, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1959:CHI

[Wil59] J. Ernest Wilkins, Jr. Conduction of heat in an insulated spherical Shell with arbitrary spherically symmetric sources. SIAM Review, 1(2): 149–153, ???? 1959. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1960:BRI

[Wil60a] R. A. Willoughby. Book review: Introduction to Matrix Analysis (Richard Bellman). SIAM Review, 2(3): 225–226, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1960:BRM

[Wil60b] R. A. Willoughby. Book review: Matrices (W. V. Parker and J. C. Eaves). SIAM Review, 2(3):226, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Williams:1961:BRE

[Wil61] J. W. Williams. Book review: The Electrical Double Layer Around a Spherical Colloid Particle (A. L. Loeb). SIAM Review, 3(4): 341–342, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1962:BRM

[Wil62a] C. H. Wilcox. Book review:

Modern Geometrical Optics (M. Herzberger). SIAM Review, 4(1):54–55, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1962:BRP

[Wil62b] C. H. Wilcox. Book review:

Progress in Optics. Vol. I (E. Wolf). SIAM Review, 4(1):58, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1962:BRW

[Wil62c] C. H. Wilcox. Book review: Wave Propagation in a Random Medium (Lev A. Chernov). SIAM Review, 4 (1):55, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1964:EWP

[Wil64a] Calvin H. Wilcox. Electric wave propagation on nonuniform coupled transmission lines. SIAM Review, 6(2): 148–165, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1964:EAG

[Wil64b] J. Ernest Wilkins, Jr. An expansion for the absorptivity of a gas. SIAM Review, 6(4):456, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1964:BRE

[Wil64c] R. A. Willoughby. Book review: Error Propagation for Difference Methods (Peter Henrici). SIAM Review, 6 (3):316–317, July 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1965:BRE

[Wil65a] J. Ernest Wilkins, Jr. Book review: Escape Velocity with Drag (D. J. Newman). SIAM Review, 7(3): 423–424, ???? 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1965:BRE

[Wil65b] Ralph A. Willoughby. Book review: Elements of Numerical Analysis (Peter Henrici). SIAM Review, 7(4):571– 572, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//www.jstor.org/stable/
2027792.

Wilf:1966:BRI

[Wil66] Herbert S. Wilf. Book review: Iterative Methods for the Solution of Equations (J. F. Traub). SIAM Review, 8 (4):550–551, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1967:BRD

[Wil67a] Calvin H. Wilcox. Book review: Differential- and Integral- Ungleichungen (Wolfgang Walter). SIAM Review, 9(3):603–605, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willett:1967:USB

[Wil67b] D. Willett. Uniqueness of solutions of boundary value problems. SIAM Review, 9 (4):726–728, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1967:SBV

[Wil67c] Howell K. Wilson. On a singular boundary value problem. SIAM Review, 9(4): 734–736, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1967:ACF

[Wil67d] Howell K. Wilson. On the average of a convex function. SIAM Review, 9(1): 120–121, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1969:BRE

[Wil69] J. Ernest Wilkins, Jr. Book review: An Extremal Problem (J. Neuringer and D. J. Newman). SIAM Review, 11 (1):86–88, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1970:BRP

[Wil70a] Calvin H. Wilcox. Book review: Perturbation Theory for Linear Operators (Tosio Kato). SIAM Review, 12(1): 155–157, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1970:HTP

[Wil70b] J. Ernest Wilkins, Jr. A heat transfer problem. SIAM Review, 12(4):580–581, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Williams:1970:CTL

[Wil70c] A. C. Williams. Complementarity theorems for linear programming. SIAM Review, 12 (1):135–137, ???? 1970. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkinson:1971:MEA

[Wil71] J. H. Wilkinson. Modern error analysis. SIAM Review, 13(4):548–568, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). The 1970 von Neumann lecture.

Willoughby:1972:BRN

[Wil72] R. A. Willoughby. Book review: Numerical Solution of Ordinary Differential Equations (Leon Lapidus and John H. Seinfeld). SIAM Review, 14(2):352–354, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1973:BRN

[Wil73] Ralph A. Willoughby. Book review: Numerical Initial Value Problems in Ordinary Differential Equations (C. William Gear). SIAM Review, 15(3):676–678, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilde:1976:MPE

[Wil76a] D. J. Wilde. Mathematical properties of engineering design optimization problems. SIAM Review, 18 (4):833, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1976:WPH

[Wil76b]  L. O. Wilson.  Wave

propagation in hexagonal 622 or 6 mm piezoelectric fibers. SIAM Review, 18 (4):833, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1977:II

[Wil77a] J. E. Wilkins, Jr. An integral inequality. SIAM Review, 19 (2):328, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1977:BRF

[Wil77b] Ralph A. Willoughby. Book review: Fourier Analysis of Time Series: an Introduction (Peter Bloomfield). SIAM Review, 19(4):745– 746, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1977:BRG

[Wil77c] D. J. Wilson. Book review: The Game of Slash (D. N. Berman). SIAM Review, 19 (1):149–150, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1978:II

[Wil78a] J. E. Wilkins, Jr. An integral inequality. SIAM Review, 20(2):396–398, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See erratum [Wil80c].

Willoughby:1978:BRC

[Wil78b] Ralph A. Willoughby. Book review: Construction of Integration Formulas for Initial Value Problems (P. J. Van Der Houwen). SIAM Review, 20(2):408–409, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1978:FBS

[Wil78c] L. O. Wilson. Flow between a stationary and a rotating disk with suction. SIAM Review, 20(3):637, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1979:BRS

[Wil79] Ralph A. Willoughby. Book review: Solutions of Ill-Posed Problems (A. N. Tikhonov and V. Y. Arsenin). SIAM Review, 21(2): 266–267, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1980:Ca

[Wil80a]  J. E. Wilkins.  Correc-

tion. SIAM Review, 22(1):

102, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1980:Cb

[Wil80b]  J. E. Wilkins.  Correc-

tion. SIAM Review, 22(3): 373–374, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1980:EII

[Wil80c] J. E. Wilkins, Jr. Erratum: An Integral Inequality. SIAM Review, 22(1):102, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Wil78a].

Willoughby:1981:BRC

[Wil81] Ralph A. Willoughby. Book review: Compact Numerical Methods for Computers: Linear Algebra and Function Minimization (J. C. Nash). SIAM Review, 23(1): 124–125, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1982:IHF

[Wil82a]  J. A. Wilson.  Identity

for hypergeometric functions. SIAM Review, 24(3):346– 347, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1982:BRI

[Wil82b] James A. Wilson. Book review: Identity for Hypergeometric Functions (O. G. Ruehr). SIAM Review, 24(3): 346–347, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1984:IEM

[Wil84a] C. H. Wilcox. Integral equation method in scatteringtheory — Colton, D., Kress, R. SIAM Review, 26(4): 597–598, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1984:BRI

[Wil84b] Calvin H. Wilcox. Book review: Integral Equation Method in Scattering Theory (David Colton and Rainer Kress). SIAM Review, 26(4): 597–598, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1985:BRA

[Wil85a] Calvin H. Wilcox. Book review: The Analysis of Linear Partial Differential Operators I, II (Lars Hörmander). SIAM Review, 27(2):311– 313, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1985:RCP

[Wil85b] J. E. Wilkins. Radius of convergence of a power Serieì. SIAM Review, 27(4): 573–574, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1985:BRE

[Wil85c] J. A. Wilson. Book review: A q-Extension of Cauchy’s Beta Integral (Richard Askey). SIAM Review, 27(2):252– 253, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1986:BRA

[Wil86a] Calvin H. Wilcox. Book review: The Analysis of Linear Partial Differential Operators. Vol. III: Pseudo-Differential Operators (Lars Hörmander) and The Analysis of Linear Partial Differential Operators. Vol. IV: Fourier Integral Operators (Lars Hörmander). SIAM Review, 28(2):285–287, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1986:BRS

[Wil86b] Ralph A. Willoughby. Book review: Sparse Matrix Technology (Sergio Pissanetsky). SIAM Review, 28(1):130– 131, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1988:BRI

[Wil88] Calvin H. Wilcox. Book review: Initial Boundary Value Problems in Mathematical Physics (R. Leis). SIAM Review, 30(2):354– 355, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1989:FE

[Wil89] David W. Wilson. A functional equation. SIAM Review, 31(4):676, ???? 1989. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Williams:1991:BRC

[Wil91a] William O. Williams. Book review: Continua with Microstructure (Gianfranco Capriz). SIAM Review, 33 (1):161–163, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Williams:1991:CMG

[Wil91b] William O. Williams. Continua with microstructure (Gianfranco Capiz). SIAM Review, 33(1):161–??, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1992:BRI

[Wil92] James A. Wilson. Book review: An Identity for Sums of Binomial Coefficients (Masaaki Sibuya). SIAM Review, 34(1):123– 126, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1993:BRN

[Wil93] Ralph A. Willoughby. Book review: Numerical Mathematics (Gunther Hammerline and Karl-Heinz Hoffman; Larry Schumaker, trans.). SIAM Review, 35(1):174– 175, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1996:II

[Wil96] J. E. Wilkins, Jr. An integral inequality. SIAM Review, 38 (2):316–319, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1998:I

[Wil98a] J. E. Wilkins. An inequality. SIAM Review, 40(1): 126–128, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1998:LPI

[Wil98b]  J. A. Wilson.  A La-

guerre polynomial identity.

SIAM Review, 40(1):143– 145, ???? 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Williams:2021:CCF

[Wil21] R. Ryan Williams. From circuit complexity to faster allpairs shortest paths. SIAM Review, 63(3):559–582, ???? 2021. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wild:2023:RSa

[Wil23a] Stefan M. Wild. Research spotlights. SIAM Review, 65 (1):145, ???? 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/23N975624.

Wild:2023:RSb

[Wil23b] Stefan M. Wild. Research spotlights. SIAM Review, 65 (2):437, ???? 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/23N975685.

Wild:2023:RSc

[Wil23c] Stefan M. Wild. Research spotlights. SIAM Review, 65 (3):733, ???? 2023. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:/

/epubs.siam.org/doi/10.
1137/23N975739.

Wild:2023:RS

[Wil23d] Stefan M. Wild. Research spotlights. SIAM Review, 65(4):1029, November 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wild:2024:RSa

[Wil24a] Stefan M. Wild. Research spotlights. SIAM Review, 66 (1):89, February 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wild:2024:RSb

[Wil24b] Stefan M. Wild. Research spotlights. SIAM Review, 66(2):285, May 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willms:2024:RFO

[Wil24c] Allan R. Willms. Resonantly forced ODEs and repeated roots. SIAM Review, 66(1): 149–160, February 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1965:BRS

[Wim65] Jet Wimp. Book review: Special Functions and Their Applications (N. N. Lebedev). SIAM Review, 7(4): 577–580, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1967:BRT

[Wim67]  Jet Wimp.  Book re-

view: Transcendental Functions Satisfying Nonhomogeneous Linear Differential Equations (A. W. Babister). SIAM Review, 9(4): 759–761, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1972:BRO

[Wim72] Jet Wimp. Book review:

Obtaining Generating Functions (Elna B. McBride). SIAM Review, 14(4):663– 667, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1975:BRA

[Wim75]  Jet Wimp.  Book re-

view: Asymptotics and Special Functions (F. W. J. Olver). SIAM Review, 17(3): 569–575, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1980:BRB

[Wim80] Jet Wimp. Book review:

Bessel Polynomials (Emil Grosswald). SIAM Review, 22(1):104–106, ???? 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1981:BRP

[Wim81]  Jet Wimp.  Book re-

view: Padé-Type Approximation and General Orthogonal Polynomials (Claude Brezinski). SIAM Review, 23(3): 403–406, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1986:BRT

[Wim86] Jet Wimp. Book review: A Treatise on Generating Functions (H. M. Srivastava and H. L. Manocha). SIAM Review, 28(1):131– 135, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1988:BRP

[Wim88] Jet Wimp. Book review: Pi and the AGM: a Study in Analytic Number Theory and Computational Complexity (Jonathan M. Borwein and Peter B. Borwein). SIAM Review, 30(3): 530–533, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1991:BHS

[Wim91a]  Jet Wimp.  Basic hy-

pergeometric series (George Gasper and Mizan Rahman). SIAM Review, 33(3):489– ??, September 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1991:BRB

[Wim91b] Jet Wimp. Book review:

Basic Hypergeometric Series (George Gasper and Mizan Rahman). SIAM Review, 33 (3):489–493, ???? 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1996:BRA

[Wim96] Jet Wimp. Book review:

Asymptotic Analysis: a Distributional Approach (R. Estrada and R. P. Kanwal). SIAM Review, 38(4): 706–711, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Winston:1960:BRR

[Win60] Clement Winston. Book review: Readings In Linear Programming (S. Vajda). SIAM Review, 2(1): 56–57, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Winston:1962:BRM

[Win62] Clement Winston. Book review: Management Models and Industrial Applications of Linear Programming, Vol. 1 (A. Charnes and W. W. Cooper). SIAM Review, 4(3): 267–268, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wing:1967:CFI

[Win67] G. M. Wing. On certain Fredholm integral equations reducible to initial value problems. SIAM Review, 9(4): 655–670, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wing:1974:BRI

[Win74] G. Milton Wing. Book review: Invariant Imbedding Theory of Nuclear Transport (J. O. Mingle). SIAM Review, 16(4):568, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wingate:1975:EOC

[Win75] J. W. Wingate. Existence of optimal controls. SIAM Review, 17(2):392, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Winfree:1990:SPL

[Win90]  A. T. Winfree.  Stable

particle-like solutions to the nonlinear wave equations of three-dimensional excitable media. SIAM Review, 32 (1):1–53, March 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Winter:1993:SCT

[Win93] Donald F. Winter. On the stem curve of a tall Palm in a strong wind. SIAM Review, 35(4):567–579, December 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See erratum [Win96b].

Winkler:1995:BRA

[Win95] Franz Winkler. Book review: Algorithmic Algebra (B. Mishra). SIAM Review, 37(3):479, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Winter:1996:SCT

[Win96a] D. F. Winter. On the stem curve of a tall palm in a strong wind (vol 35, pg 567, 1993). SIAM Review, 38(3): 485–486, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Winter:1996:ERS

[Win96b] Donald F. Winter. Erratum and reformulation: “On the stem curve of a tall palm in a strong wind” [SIAM Rev. 35 (1993), no. 4, 567–579, MR 94g:73028]. SIAM Review, 38 (3):485–486, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29892.  See

[Win93].

Washio:2010:PMT

[WiOH10] Takumi Washio, Jun ichi Okada, and Toshiaki Hisada. A parallel multilevel technique for solving the bidomain equation on a human heart with Purkinje fibers and a torso model. SIAM Review, 52(4):717– 743, ???? 2010. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wisner:1960:BRI

[Wis60] Robert J. Wisner. Book review: Introduction to Matrices and Linear Transformations (Daniel T. Finkbeiner). SIAM Review, 2(3):226– 227, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Witzgall:1975:ASP

[Wit75] C. Witzgall. Algorithm selection for parameter estimation in logit models. SIAM Review, 17(2):392, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wittke:1990:BRU

[Wit90] P. H. Wittke. Book review: A Unified Approach to Zero-Crossings and Nonuniform Sampling: Of Single and Multidimensional Signals and Systems (Farokh A. Marvasti). SIAM Review, 32 (3):521–522, September 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Withers:2022:GDT

[Wit22] L. P. Withers. A generalized dual transform: Linear algebra and geometry of (pseudo)Inverting a matrix. SIAM Review, 64 (4):1031–1061, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1270410.

Wixom:1976:BFI

[Wix76] J. A. Wixom. Blendingfunction interpolation over non-rectangular domains. SIAM Review, 18(4):834, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Warnock:1975:MSI

[WJ75] R. L. Warnock and P. W.

Johnson. Marginally singular integral operators with divided-difference kernels. SIAM Review, 17(2):391, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waterhouse:1994:BRU

[WJ94] William C. Waterhouse and A. A. Jagers. Book review: A Unique Vector Sum (Rudolf K. Meyer). SIAM Review, 36 (3):495–496, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Williams:1979:BRE

[WJV79] B. K. Williams, L. A. Joyce, and G. M. Van Dyne. Book review: Ecosystem Modeling in Theory and Practice: an Introduction with Case Histories (Charles A. S. Hall and John W. Day, Jr., eds.). SIAM Review, 21(3): 425–429, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Whittemore:1978:QTC

[WK78] Alice Whittemore and Joseph B. Keller. Quantitative theories of carcinogenesis. SIAM Review, 20(1):1–30, January 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wang:1975:BRS

[WKW75] P. T. R. Wang, D. S. Kerr, and L. J. White. Bandwidth reduction of sparse matrices by row and column permutations. SIAM Review, 17 (2):391, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weiss:1995:AAR

[WL95] G. H. Weiss and Gregory F. Lawler. Aspects and applications of the random walk. SIAM Review, 37(3):470– ??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1961:BRV

[WLP^{+}61] J. Ernest Wilkins, Jr., Yudell L. Luke, Jan Petersson, P. J. de Doelder, J. H. van Lint, and M. S. Klamkin. Book review: Vorticity Interaction (Sin-I Cheng). SIAM Review, 3(4): 329–336, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilkins:1961:CTF

[WM61a] J. Ernest Wilkins, Jr. and Harold Mechanic. Calculation of total field integrals for air scattered dose rates. SIAM Review, 3(2): 140–143, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See errata [WM61b].

Wilkins:1961:ECT

[WM61b] J. Ernest Wilkins, Jr. and Harold Mechanic. Errata: Calculation of Total Field Integrals for Air Scattered Dose Rates. SIAM Review, 3(3): 260–261, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [WM61a].

Weiss:1963:PMT

[WM63] L. Weiss and R. N. Mc-Donough. Prony’s method, Z-transforms, and Padé approximation. SIAM Review, 5(2):145–149, ???? 1963. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Willoughby:1978:FWP

[WM78] R. A. Willoughby and C. A. Micchelli. Functions which preserve class of Stieltjes matrices. SIAM Review, 20(3): 636–637, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1985:BRR

[WM85] J. A. Wilson and J. A. MacBain. Book review: A Resistance Problem (E. N. Gilbert and L. A. Shepp). SIAM Review, 27(3):449– 452, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:2015:EPR

[WM15] Dan Wilson and Jeff Moehlis. Extending phase reduction to excitable media: Theory and applications. SIAM Review, 57(2):201–222, ???? 2015. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wollkind:1994:WNS

[WMZ94] David J. Wollkind, Valipuram S. Manoranjan, and Limin Zhang. Weakly nonlinear stability analyses of prototype reaction-diffusion model equations. SIAM Review, 36(2):176–214, June 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wojdylo:2006:CCL

[Woj06] John Wojdylo. Computing the coefficients in Laplace’s method. SIAM Review, 48 (1):76–96, ???? 2006. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolk:1960:THP

[Wol60] E. S. Wolk. Torpedo hit probabilities. SIAM Review, 2(4): 292–296, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolfe:1965:CSA

[Wol65] Philip Wolfe. The composite simplex algorithm. SIAM Review, 7(1):42–54, ???? 1965. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolfe:1969:CCA

[Wol69] Philip Wolfe. Convergence conditions for ascent methods. SIAM Review, 11(2):

226–235, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolfe:1971:CCA

[Wol71] Philip Wolfe. Convergence conditions for ascent methods. II: Some corrections. SIAM Review, 13(2):185– 188, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolfs:1972:BRD

[Wol72a] Peter Wolfs. Book review: A Double Integral (C. J. Bouwkamp). SIAM Review, 14(3):505–506, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolk:1972:BEI

[Wol72b] Barry Wolk. A binomial expansion identity. SIAM Review, 14(4):644, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolfe:1975:SPM

[Wol75] P. Wolfe. Some progress in mathematical programming. SIAM Review, 17 (2):392, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wollkind:1977:SPT

[Wol77] David J. Wollkind. Singular perturbation techniques: a comparison of the method of matched asymptotic expansions with that of multiple scales. SIAM Review, 19 (3):502–516, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolkowicz:1981:CMO

[Wol81]  H. Wolkowicz.  A con-

strained matrix optimization problem. SIAM Review, 23 (1):101, ???? 1981. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolkowicz:1982:CMO

[Wol82]  H. Wolkowicz.  A con-

strained matrix optimization problem. SIAM Review, 24 (1):83, ???? 1982. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolf:1986:BRC

[Wol86]  Wayne Wolf.  Book re-

view: Complexity Issues in VLSI-Optimal Layouts for the Shuffle-Exchange Graph and Other Networks (Frank Thomson Leighton). SIAM Review, 28(2):287–288, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolper:1988:CNR

[Wol88] James S. Wolper. (celestial) navigation on a Riemannian Earth. SIAM Review, 30 (3):498–501, September 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolper:1993:BRA

[Wol93] James S. Wolper. Book review: Algebraic-Geometric Codes (M. A. Tsfasman and S. G. Vlad̆uţ). SIAM Review, 35(3):541–543, September 1993. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolkowicz:1994:BRM

[Wol94a] H. Wolkowicz. Book review: Matrix Differentiation Identities (Peter J. Costa and Stanley Rabinowitz). SIAM Review, 36(4):657– 659, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolkowicz:1994:PS

[Wol94b] H. Wolkowicz. Problem solution 93-17. SIAM Review, 36 (4):657–659, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wolper:1994:TPA

[Wol94c] James S. Wolper. Turn performance of aircraft.

SIAM Review, 36(3):470– 473, September 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25697.htm. See

clarification [For96].

Wolfe:1997:BRA

[Wol97]  Peter Wolfe.  Book re-

view: Applied Continuum Mechanics, by T. J. Chung. SIAM Review, 39(2):366– 367, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wong:1964:LSP

[Won64a] Eugene Wong. A linear search problem. SIAM Review, 6(2):168–174, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wong:1964:NBT

[Won64b] James S. W. Wong. A note on boundedness theorems of certain second order differential equations. SIAM Review, 6 (2):175–176, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wong:1968:NSO

[Won68] James S. W. Wong. A note on second order nonlinear oscillation. SIAM Review, 10 (1):88–91, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wong:1975:GEF

[Won75] James S. W. Wong. On the generalized Emden–Fowler equation. SIAM Review, 17 (2):339–360, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wong:1980:EBA

[Won80] R. Wong. Error bounds for asymptotic expansions of integrals. SIAM Review, 22(4): 401–435, October 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wong:1983:COA

[Won83a] R. T. Wong. Combinatorial optimization — algorithms and complexity — Papadimitriou, C. H. , Steiglitz, K. SIAM Review, 25 (3):424–425, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wong:1983:BRC

[Won83b] Richard T. Wong. Book review: Combinatorial Optimization: Algorithms and Complexity (Christos H. Papadimitriou and Kenneth Steiglitz). SIAM Review, 25 (3):424–425, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Woods:1992:BRR

[Woo92]  John W. Woods.  Book

review: Random Signals and Systems (Richard E. Mortensen). SIAM Review, 34(1):148–149, March 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wood:1996:VLE

[Woo96] T. E. Wood. Vertex latitudes on ellipsoid geodesics. SIAM Review, 38(4):637– 644, December 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28344.

Wouk:1962:RMD

[Wou62] Arthur Wouk. Recursive methods, dynamic programming, and invariant imbedding. SIAM Review, 4(4): 384–393, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wouk:1964:BRS

[Wou64a] Arthur Wouk. Book review:

Stability of Motion (N. N. Krasovskii). SIAM Review, 6 (4):471–472, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wouk:1964:BRT

[Wou64b] Arthur Wouk. Book review: Theory and Application of Liapunov’s Direct Method (Wolfgang Hahn). SIAM Review, 6(2):186– 187, ???? 1964. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wouk:1966:NSR

[Wou66]  A. Wouk.  A note on

square roots of positive operators. SIAM Review, 8(1): 100–102, ???? 1966. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wouk:1967:BRC

[Wou67] Arthur Wouk. Book review: Constructive Real Analysis (Allen A. Goldstein). SIAM Review, 9(4):759, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wouk:1968:BRA

[Wou68a] Arthur Wouk. Book review: Approximate Methods for Solution of Differential and Integral Equations (S. G. Mikhlin and K. L. Smolitskiy). SIAM Review, 10(3): 391–392, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wouk:1968:BRI

[Wou68b] Arthur Wouk. Book review: Introduction to the Theory of Differential Equations with Deviating Arguments (L. E.

El’sgol’ts). SIAM Review, 10 (2):240–241, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wozniakowski:1987:BRN

[Woź87] H. Woźniakowski. Book review: Nondiscrete Induction and Iterative Processes (F.-A. Potra and V. Ptak). SIAM Review, 29(3):505– 506, ???? 1987. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Walker:1978:IPO

[WP78] H. F. Walker and B. C.

Peters. Iterative procedure for obtaining maximumlikelihood estimates of parameters for a mixture of normal distributions. SIAM Review, 20(3):636, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wang:1988:BRF

[WP88] Pei-Zhuang Wang and Xian-Tu Peng. Book review: Fuzzy Mathematical Techniques with Applications (Abraham Kandel). SIAM Review, 30 (3):528–530, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wiggins:1995:NHI

[WP95] Stephen Wiggins and Kenneth J. Palmer. Normally hyperbolic invariant manifolds in dynamical systems. SIAM Review, 37(3):472– ??, September 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wiggins:1996:GDP

[WP96] Stephen Wiggins and Kenneth J. Palmer. Global dynamics, phase space transport, orbits homoclinic to resonances, and applications. SIAM Review, 38(3):538– ??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weaver:1998:ERN

[WPL98] James W. Weaver and Freda Porter-Locklear. Estimating the rate of natural bioattenuation of ground water contaminants by a mass conservation approach. SIAM Review, 40(1):113– 117, March 1998. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/31240.

Wu:2001:FNA

[WPTZ01] Zhijun Wu, George N. Phillips, Jr., Richard Tapia, and Yin Zhang. A fast Newton algorithm for entropy maximization in phase determination. SIAM Review, 43 (4):623–642, December 2001.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37173.

Wilson:1994:URR

[WR94a] J. A. Wilson and M. Renardy. A unique real root. SIAM Review, 36(4):661– 662, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1994:BRU

[WR94b] James A. Wilson and Michael Renardy. Book review: A Unique Real Root (G. M. Gladwell). SIAM Review, 36 (4):661–662, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weaver:1996:MMG

[WR96a] J. T. Weaver and Robert C.

Rogers. Mathematical methods for geo-electromagnetic induction. SIAM Review, 38 (4):697–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilcox:1996:PMC

[WR96b] David C. Wilcox and R. H.

Rand. Perturbation methods in the computer age. SIAM Review, 38(3):539– ??, September 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wright:1960:BRP

[Wri60] Stephen E. Wright. Book review: Programming Business Computers (Daniel Mc-Cracken, Harold Weiss, and Tsai-Hwa Lee). SIAM Review, 2(2):158–159, ???? 1960. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wright:1999:P

[Wri99] Margaret H. Wright. Preface. SIAM Review, 41(1): ix–x, March 1999. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/97018.

Wright:2014:RSa

[Wri14a] Stephen J. Wright. Research spotlights. SIAM Review, 56 (2):259, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wright:2014:RSb

[Wri14b] Stephen J. Wright. Research spotlights. SIAM Review, 56 (3):459, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watson:1971:BRL

[WS71] G. N. Watson and Robert E. Shafer. Book review: A Laplace Transform (Lewis R. Wilson and A. Yildiz). SIAM Review, 13(3):397– 399, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Watts:1975:OCC

[WS75a] H. A. Watts and M. R.

Scott. Orthog — computer code for two-point boundaryvalue problems via orthonormalization. SIAM Review, 17 (2):392, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Weinstein:1975:CTO

[WS75b] Mills B. Weinstein and Donald R. Smith. Comparison techniques for overdamped systems. SIAM Review, 17 (3):520–540, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wild:2013:GCR

[WS13] Stefan M. Wild and Christine Shoemaker. Global convergence of radial basis function trust-region algorithms for derivative-free optimization. SIAM Review, 55(2): 349–371, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wu:2023:GQF

[WS23] Kailiang Wu and Chi-Wang Shu. Geometric quasilinearization framework for analysis and design of boundpreserving schemes. SIAM Review, 65(4):1031–1073, November 2023. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wu:1972:SMP

[Wu72] T. T. Wu. Some mathematical problems in fiber Xray crystallography. SIAM Review, 14(3):420–432, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wu:1975:TPL

[Wu75] C. H. Wu. Theory of plane linear wrinkle elasticity without body force. SIAM Review, 17(2):392, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waller:1977:PGD

[WW77] R. A. Waller and M. S. Waterman. Percentiles for the gamma distribution. SIAM Review, 19(1):146, ???? 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Waller:1978:PGD

[WW78] R. A. Waller and M. S. Waterman. Percentiles for Gamma distribution. SIAM Review, 20(4):856, October 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wilson:1970:LT

[WY70] Lewis R. Wilson and A. Yildiz.

A Laplace transform. SIAM Review, 12(2):287, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wyner:1969:CIT

[Wyn69] A. D. Wyner. On cading and information theory. SIAM Review, 11(3): 317–346, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimmer:1972:SME

[WZ72] Harald Wimmer and Allen D.

Ziebur. Solving the matrix equation ∑_{ρ=1}^{r} f_{ρ}(A)Xg_{ρ}(B) = C. SIAM Review, 14(2): 318–323, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wei:2012:VMM

[WZCX12] Guo-Wei Wei, Qiong Zheng, Zhan Chen, and Kelin Xia. Variational multiscale models for charge transport. SIAM Review, 54(4):699– 754, ???? 2012. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xanthopoulos:1987:BRR

[Xan87] Basilis C. Xanthopoulos. Book review: Rotating Fields in General Relativity (Jamal Nazrul Islam). SIAM Review, 29(3):506–507, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xu:2011:FAM

[XC11] Zhenli Xu and Wei Cai. Fast analytical methods for macroscopic electrostatic models in biomolecular simulations. SIAM Review, 53 (4):683–720, ???? 2011. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/sirev/resource/
1/siread/v53/i4/p683_s1.

Xin:2000:FPH

[Xin00] Jack Xin. Front propagation in heterogeneous media. SIAM Review, 42(2): 161–230, June 2000. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/36429.

Xu:1994:BRM

[XL94] Yongzhi Xu and Wei Lin. Book review: Methods of Singular Integral Equations (A. Dzhuraev). SIAM Review, 36 (2):324–325, June 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xin-Rong:1995:TCI

[XRTM95] Ma Xin-Rong and Wang Tian-Ming. Two combinatorial identities. SIAM Review, 37(1):98, ???? 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xu:1992:IMS

[Xu92] Jinchao C. Xu. Iterative methods by space decomposition and subspace correction. SIAM Review, 34 (4):581–613, December 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xu:1996:BRMa

[Xu96a] Jinchao Xu. Book review:

Multigrid Methods for Finite Elements (V. V. Shaidurov). SIAM Review, 38(3):542– 543, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xu:1996:BRMb

[Xu96b] Jinchao Xu. Book review:

Multigrid Methods for Process Simulation (W. Joppich and S. Mijalkovic). SIAM Review, 38(3):543, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xu:1996:VIC

[Xu96c] Liqun Xu. The volume of the intersection of a cube and a ball in N-space. SIAM Review, 38(4):669, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xu:1997:VIC

[Xu97] L. Q. Xu. The volume of the intersection of a cube and a ball in N-space. SIAM Review, 39(4):779– 786, ???? 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xu:2021:UGE

[Xu21] Mengjia Xu. Understanding graph embedding methods and their applications. SIAM Review, 63(4):825– 853, ???? 2021. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Xu:1998:SND

[XZ98] Jinchao C. Xu and Jun Zou. Some nonoverlapping domain decomposition methods. SIAM Review, 40(4): 857–914, December 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30680.

Yu:2018:PIB

[YA18] Sanghyeon Yu and Habib Ammari. Plasmonic interaction between nanospheres.

SIAM Review, 60(2):356– 385, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yakowitz:1994:BRT

[Yak94] Sidney Yakowitz. Book review: Time Series Analysis of Higher Order Crossings (B. Kedem). SIAM Review, 36(4):680–682, December 1994. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yakubovich:1995:BRL

[Yak95] V. A. Yakubovich. Book review: Linear Matrix Inequalities in System and Control Theory (S. Boyd, L. E. Ghaoui, E. Feron, and V. Balakrishnan). SIAM Review, 37(3):479–481, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yang:2020:WUP

[YAKM20] Vicky Chuqiao Yang, Daniel M.

Abrams, Georgia Kernell, and Adilson E. Motter. Why are U.S. parties so polarized? A “satisficing” dynamical model. SIAM Review, 62 (3):646–657, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yanowitch:1965:BRI

[Yan65] Michael Yanowitch. Book review: Integral Equations and Partial Differential Equations, vol. IV (V. I. Smirnov). SIAM Review, 7(4):572– 574, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yanowitch:1976:BRI

[Yan76] Michael Yanowitch. Book review: Integral Equations (Harry Hochstadt). SIAM Review, 18(4):782, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yang:1977:BRP

[Yan77]  L. J. Yang.  Book re-

view: A Polynomial Inequality (K. B. Stolarsky and L. J. Yang). SIAM Review, 19 (4):737–739, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yanushevsky:1990:PFI

[Yan90] R. T. Yanushevsky. Parametric feedback improving control systems efficiency. SIAM Review, 32(2):294– 299, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yang:1991:BRA

[Yan91] Huijun Yang. Book review: Asymptotic Modeling of Atmospheric Flows (R. Zeytounian). SIAM Review, 33 (4):672–673, December 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yang:1992:BRS

[Yan92] Grace L. Yang. Book review:

Stochastic Processes in Epidemic Theory (J.-P. Gabriel, C. Léfvre and P. Picard, eds.). SIAM Review, 34(1): 153–155, March 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yanik:1994:BRN

[Yan94] Elizabeth Greenwell Yanik.

Book review: Numerical Recipes in FORTRAN-The Art of Scientific Computing 2nd Ed. (W. H. Press, W. T. Vetterling, S. A. Teukolsky and B. P. Flannery). SIAM Review, 36 (1):149, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yan:1997:ODS

[Yan97a] D. Yan. Optimal and dominating strategies for determining continuous caster product dimensions. SIAM Review, 39(3):461–471, ???? 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yan:1997:CSI

[Yan97b]  Dicky Yan.  Case stud-

ies from industry: Optimal and dominating strategies for determining continuous Caster product dimensions. SIAM Review, 39 (3):461–471, September 1997. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30988.

Yao:1986:BRP

[Yao86] J. T. P. Yao. Book review: Probabilistic Methods in Structural Engineering (G. Augusti, A. Baratta and F. Casciati). SIAM Review, 28 (3):442–443, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yasaimaibodi:1971:II

[Yas71] Hedayat Yasaimaibodi. An infinite integral. SIAM Review, 13(4):572, ???? 1971. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yates:1984:MSF

[Yat84a] G. T. Yates. Mechanics of swimming and flying — Childress, S. SIAM Review, 26 (4):596–597, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yates:1984:BRM

[Yat84b] George T. Yates. Book review: Mechanics of Swimming and Flying (Stephen Childress). SIAM Review, 26 (4):596–597, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yau:1968:BRL

[Yau68] S. S. Yau. Book review:

Linear Sequential Circuits Analysis, Synthesis, and Applications (Arthur Gill). SIAM Review, 10(3):397– 399, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yau:1969:BRC

[Yau69] S. S. Yau. Book review:

Computation: Finite and Infinite Machines (Marvin L. Minsky). SIAM Review, 11 (1):99–100, ???? 1969. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ye:1994:BRI

[Ye94]  Yinyu Ye.  Book re-

view: Interior-Point Polynomial Algorithms in Convex Programming (Y. Nesterov and A. Nemirovskii). SIAM Review, 36(4):682– 683, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yuen:1978:DNL

[YF78] H. C. Yuen and W. E. Ferguson. 2-dimensional nonlinear Schrödinger equation with hyperbolic spatial operator. SIAM Review, 20 (3):637, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yuen:1976:MEB

[YFS76] H. C. Yuen, W. E. Ferguson, and P. G. Saffman. Model equation for breaking and capillary wave generation by steep gravity wave. SIAM Review, 18(4):834– 835, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yousif:1976:RGO

[YH76] S. M. Yousif and W. W. Happ. Representation of generalized operator in linearsystem theory. SIAM Review, 18(4):834, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yoh-Han:1993:SPR

[YH93] Pao Yoh-Han. Statistical pattern recognition (Keinosuke Fukunaga). SIAM Review, 35 (1):161–??, March 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yuan:1996:PI

[YKF96] Bo Yuan, George J. Klir, and Jiri Fridrich. A projection inequality. SIAM Review, 38 (2):315, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yakowitz:1978:WMC

[YKS78] S. Yakowitz, J. Krimmel, and F. Szidarovszky. Weighted Monte–Carlo integration. SIAM Review, 20 (3):637, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yuen:1975:NDW

[YL75] H. C. Yuen and B. M.

Lake. Nonlinear deep-water waves — theory and experiment. SIAM Review, 17 (2):393, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yasar:2003:ECS

[YL03] Osman Yasar and Rubin H.

Landau. Elements of computational science and engineering education. SIAM Review, 45(4):787–805, December 2003. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/40807.

Yohe:1979:INA

[Yoh79] J. M. Yohe. Implementing nonstandard arithmetics. SIAM Review, 21(1):34– 56, ???? 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yong:1995:BRC

[Yon95] Jiongmin Yong. Book review: Control under Lack of Information (A. N. Krasovskii and N. N. Krasovskii). SIAM Review, 37(4):640–641, ???? 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yong:2006:SCR

[Yon06] Darryl Yong. Strings, chains, and ropes. SIAM Review, 48(4):771–781, January 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yorke:1975:DSD

[Yor75] J. A. Yorke. Dynamics of spread of diseases. SIAM Review, 17(2):392, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yos:1960:SSP

[Yos60]  Jerry Yos.  Steady-state

plasma arc. SIAM Review, 2 (2):153–154, ???? 1960. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yousif:1971:BRA

[You71] Salah M. Yousif. Book review: Algebraic Theory of Automata (Abraham Ginsburg). SIAM Review, 13(1): 109–110, ???? 1971. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Younger:1972:BRG

[You72a] D. H. Younger. Book review: Graph Theory (Frank Harary). SIAM Review, 14 (2):350–351, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yousif:1972:BRS

[You72b] Salah M. Yousif. Book review: System Identification (Andrew P. Sage and James L. Melsa). SIAM Review, 14 (3):514–517, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Young:1973:SMN

[You73] David M. Young. A survey of modern numerical analysis. SIAM Review, 15(2): 503–523, April 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yousif:1974:BRS

[You74] Salah M. Yousif. Book review: Statistics of Directional Data (K. V. Mardia). SIAM Review, 16(1):

108–109, ???? 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Young:1976:CLV

[You76] D. F. Young. Control of linear Volterra–Stieltjes equations. SIAM Review, 18 (4):834, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ypma:1983:FMZ

[Ypm83] T. J. Ypma. Finding a multiple zero by transformations and Newton-like methods. SIAM Review, 25(3): 365–378, July 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ypma:1995:HDN

[Ypm95]  Tjalling J. Ypma.  His-

torical development of the Newton–Raphson method. SIAM Review, 37(4):531– 551, December 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/23425.
htm;   http://link.aip.
org/link/?SIR/37/531/1.

Yakowitz:1976:NRA

[YS76] S. Yakowitz and E. Schuster. Nonparametric regression with application to system identification. SIAM Review, 18(4):834, ???? 1976.

CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yakubov:1995:CRF

[YT95] S. Yakubov and Hans Triebel.

Completeness of root functions of regular differential operators. SIAM Review, 37 (1):138–??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

You:1996:BRM

[YY96] Zhiping You and James A.

Yorke. Book review: Mathematical Go Chilling Gets the Last Point (Elwyn Berlekamp and David Wolfe). SIAM Review, 38(3):543–544, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yao:2004:SLQ

[YZZ04] David D. Yao, Shuzhong Zhang, and Xun Yu Zhou. Stochastic linear-quadratic control via primal-dual semidefinite programming. SIAM Review, 46(1):87–111, March 2004. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/
sam-bin/dbq/article/43420.

Zadoff:1967:BRS

[Zad67] Solomon A. Zadoff. Book review: A Sharp Autocorrelation Function (Henry D. Friedman). SIAM Review, 9 (3):591–593, ???? 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zahler:1980:BRQ

[Zah80] Raphael Zahler. Book review: Questions of Uniqueness and Resolution in Reconstruction from Projections (Myron Bernard Katz). SIAM Review, 22(3):387– 388, ???? 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zak:1984:LSK

[Zak84a] S. H. Zak. Linear systems — Kailath, T. SIAM Review, 26 (1):132–135, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zak:1984:BRL

[Zak84b] Stanislaw H. Zak. Book review: Linear Systems (Thomas Kailath). SIAM Review, 26(1):132–135, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zak:1985:BRA

[Zak85] Stanislaw H. Zak. Book review: Algebraic Theory for Multivariable Linear Systems (H. Blomberg and R. Ylinen). SIAM Review, 27(1): 130–131, ???? 1985. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zalcman:1983:BRR

[Zal83] Lawrence Zalcman. Book review: The Radon Transform (Sigurdur Helgason). SIAM Review, 25(2):275– 278, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zaroodny:1965:BRF

[Zar65] Serge J. Zaroodny. Book review: The Flight of Uncontrolled Rockets (F. R. Gantmacher and L. M. Levin). SIAM Review, 7(4):583– 585, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zaremba:1968:MBM

[Zar68] S. K. Zaremba. The mathematical basis of Monte Carlo and quasi-Monte Carlo methods. SIAM Review, 10(3):303–314, July 1968. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2027655.

Zassenhaus:1965:BRG

[Zas65] Hans Zassenhaus. Book review: Generators and Relations for Discrete Groups, second edition (H. S. M. Coxeter and W. O. J. Moser). SIAM Review, 7 (4):588, October 1965. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zayed:1992:EBV

[Zay92] Ahmed I. Zayed. Existence of a boundary-value problem with prescribed eigenvalues. SIAM Review, 34(1): 118–119, ???? 1992. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zelikin:1996:TCC

[ZBK96] M. I. Zelikin, V. F. Borisov, and Arthur J. Krener. Theory of chattering control. SIAM Review, 38(1):172– ??, March 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zischka:1974:NIB

[ZC74] K. A. Zischka and P. S. Chow. On nonlinear initial boundary value problems of heat conduction and diffusion. SIAM Review, 16(1): 17–35, January 1974. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zabusky:1978:CDI

[ZD78] N. J. Zabusky and G. S. Deem. Contour dynamics for inviscid incompressible twodimensional flows. SIAM Review, 20(3):637, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zegeling:1995:MGM

[ZD95] P. A. Zegeling and Thomas K.

DeLillo. Moving-grid methods for time-dependent partial differential equations. SIAM Review, 37(1):120– ??, March 1995. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zeiberger:1972:BRS

[Zei72] D. Zeiberger. Book review:

Special Subsets of a Finite Group (Royce Derby). SIAM Review, 14(2):367, ???? 1972. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zelmer:1977:IPM

[Zel77] G. Zelmer. On interpolating polynomials of minimal degree. SIAM Review, 19 (1):1–4, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zemanian:1968:IGF

[Zem68] A. H. Zemanian. An introduction to generalized functions and the generalized Laplace and Legendre transformations. SIAM Review, 10 (1):1–24, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zemanian:1996:BRT

[Zem96] A. H. Zemanian. Book review: The Transmission-Line Modeling Method TLM (Christos Christopolous). SIAM Review, 38(4):711, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:1988:CE

[Zha88] Weijian Zhang. A conditional expectation. SIAM Review, 30(1):124, ???? 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:1996:BRL

[Zha96]  Yin Zhang.  Book re-

view: Linear Programming: a Modern Integrated Analysis (Romesh Saigal). SIAM Review, 38(4):711–712, ???? 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:1997:BRC

[Zha97] Zhimin Zhang. Book review: Conservative Finite-Difference Methods on General Grids, by Mikhail Shashkov. SIAM Review, 39 (2):367–??, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:2013:DRL

[Zha13] Qinghai Zhang. On donating regions: Lagrangian flux through a fixed curve. SIAM Review, 55(3):443– 461, ???? 2013. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ziebur:1970:DSA

[Zie70] Allen D. Ziebur. On determining the structure of A by analyzing e^{At}. SIAM Review, 12(1):98–102, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ziebur:1973:IRM

[Zie73] Allen D. Ziebur. Integral representation for the Moore– Penrose generalized inverse of a matrix. SIAM Review, 15 (1):218, ???? 1973. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zierler:1976:BRE

[Zie76]  Neal Zierler.  Book re-

view: Error Coding for Arithmetic Processors (T. R. N. Rao). SIAM Review, 18 (1):144, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zierler:1978:BRI

[Zie78]  Neal Zierler.  Book re-

view: An Introduction to Algebraic and Combinatorial Coding Theory (Ian F. Blake and Ronald C. Mullin).

SIAM Review, 20(3):607– 608, ???? 1978. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ziebur:1979:NDL

[Zie79] Allen D. Ziebur. New directions in linear differential equations. SIAM Review, 21 (1):57–70, January 1979. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zierler:1983:ECF

[Zie83a]  N. Zierler.  Elementary

computability, formal languages, and automata — Mcnaughton, R. SIAM Review, 25(3):438, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zierler:1983:BRR

[Zie83b]  Neal Zierler.  Book re-

view: A Relational Theory of Computing (John G. Sanderson). SIAM Review, 25(1): 121–122, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zierler:1983:BRE

[Zie83c] Neal Zierler. Book review: Elementary Computability, Formal Languages, and Automata (Robert McNaughton). SIAM Review, 25(3):438, ???? 1983. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zierler:1984:SCS

[Zie84a] N. Zierler. Studies in computer science — Pollack, S. V. SIAM Review, 26(4): 608, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zierler:1984:BRS

[Zie84b] Neal Zierler. Book review:

Studies in Computer Science (S. V. Pollack, ed.). SIAM Review, 26(4):608, ???? 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zimmermann:1990:BRD

[Zim90] Uwe Zimmermann. Book review: Discrete Optimization (R. Gary Parker and Ronald L. Rardin). SIAM Review, 32 (2):333–334, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zitron:1963:BRE

[Zit63] N. Zitron. Book review:

Extraction of Signals from Noise (L. A. Wainstein and V. D. Zubakov). SIAM Review, 5(1):83, ???? 1963. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zitron:1976:CMO

[Zit76] N. R. Zitron. Continuous model of optimal-cost routes in a circular city. SIAM Review, 18(4):835, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zwillinger:1996:CSM

[ZK96] Daniel Zwillinger and Bruce Kellogg. CRC Standard Mathematical Tables and Formulae. SIAM Review, 38 (4):691–??, December 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ziegelmeier:2017:SHD

[ZKP17] Lori Ziegelmeier, Michael Kirby, and Chris Peterson. Stratifying high-dimensional data based on proximity to the convex hull boundary. SIAM Review, 59(2): 346–365, ???? 2017. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zlobec:1970:EFM

[Zlo70] Sanjo Zlobec. An explicit form of the Moore–Penrose inverse of an arbitrary complex matrix. SIAM Review, 12(1):132–134, ???? 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zlobec:1977:BRN

[Zlo77] S. Zlobec. Book review: Nonlinear Programming (Bela Martos). SIAM Review, 19 (1):156–158, ???? 1977. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zlobec:1983:BRA

[Zlo83] S. Zlobec. Book review:

Approaches to the Theory of Optimization (J. Ponstein). SIAM Review, 25(3): 416–417, ???? 1983. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zlobec:1985:BRC

[Zlo85] S. Zlobec. Book review: Constrained Optimization and Lagrange Multiplier Rules (Dimitri P. Bertsekas). SIAM Review, 27(1):131–132, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zlobec:1991:BRI

[Zlo91] S. Zlobec. Book review:

2-Inverses and Their Statistical Application (Albert J. Getson and Francis C. Hsuan). SIAM Review, 33 (2):338–340, June 1991. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:2020:SIS

[ZM20] Yuanzhao Zhang and Adilson E. Motter. Symmetryindependent stability analysis of synchronization patterns. SIAM Review, 62(4): 817–836, ???? 2020. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zou:2022:DGM

[ZM22] Qinmeng Zou and Frédéric Magoulès. Delayed gradient methods for symmetric and positive definite linear systems. SIAM Review, 64(3):517–553, ???? 2022. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1321140.

Zou:2024:NCL

[ZMPK24] Zongren Zou, Xuhui Meng, Apostolos F. Psaros, and George E. Karniadakis. NeuralUQ: a comprehensive library for uncertainty quantification in neural differential equations and operators. SIAM Review, 66(1): 161–190, February 2024. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zetto:1962:BRM

[ZN62] F. Zetto and D. J. Newman. Book review: A Minimum Time Path (W. L. Bade). SIAM Review, 4(3): 258–259, ???? 1962. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zuazua:2005:POC

[Zua05]  Enrique Zuazua.  Prop-

agation, observation, and control of waves approximated by finite difference methods. SIAM Review, 47(2):197–243, June 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43286.

Zuo:1996:BRE

[Zuo96] Quan-Ru Zuo. Book review:

Extensions of Routh’s Theorem to Spherical Triangles (A. Sharma and Murray S. Klamkin). SIAM Review, 38 (2):319–320, ???? 1996. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:1993:OSP

[ZW93] Hong Zhang and Xiaohuan Wang. Optimal sensor placement. SIAM Review, 35 (4):641, ???? 1993. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:1994:OSP

[ZW94] Hong Zhang and Xiaohuan Wang. Optimal sensor placement in three-dimensions. SIAM Review, 36(1):108– 109, ???? 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zweifel:1984:SML

[Zwe84a] P. F. Zweifel. Spectral methods in linear transport theory — Kaper, H. G., Lekkerkerker, C. G., Hejtmanek, J. SIAM Review, 26(4): 604–606, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zweifel:1984:BRS

[Zwe84b]  Paul F. Zweifel.  Book

review: Spectral Methods in Linear Transport Theory (H. G. Kaper, C. G. Lekkerkerker and J. Hejtmanek). SIAM Review, 26 (4):604–606, ???? 1984. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zweck:2016:AMU

[Zwe16] John Zweck. Analysis of methods used to reconstruct the flight path of Malaysia Airlines Flight 370. SIAM Review, 58(3):555– 574, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:2016:HSA

[ZWGB16] Guannan Zhang, Clayton G. Webster, Max Gunzburger, and John Burkardt. Hyperspherical sparse approximation techniques for highdimensional discontinuity detection. SIAM Review, 58(3): 517–551, ???? 2016. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:2014:VBM

[ZWY14] Wenjing Zhang, Lindi M.

Wahl, and Pei Yu. Viral blips may not need a trigger: How transient viremia can arise in deterministic in-host models. SIAM Review, 56(1): 127–155, ???? 2014. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zha:2009:SPA

[ZZ09] Hongyuan Zha and Zhenyue Zhang. Spectral properties of the alignment matrices in manifold learning. SIAM Review, 51(3):545– 566, ???? 2009. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Zhang:2018:SON

[ZZ18] Haisen Zhang and Xu Zhang.

Second-order necessary conditions for stochastic optimal control problems. SIAM Review, 60(1):139–178, ???? 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).